A Complete Bibliography of Publications in Applied Mathematics and Computation: 2015–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

03 June 2025 Version 1.77

Title word cross-reference

(1 + 1) [HGH19a]. (1 + 2) [HGH19a]. (2 + 1) [FMC15, MMIA19, OW18, ZH15a, ZW15b]. (2R + 1) [SdR19]. (3 + 1) [DT16, CL15b, LL18a, VK19, WTXZ16]. (4, 6) [SZ19c]. (d, 1) [DSZY19]. (n + 1) [YJSW16]. (n, m) [MBW19]. (n − 1) [HHZ18]. (P, Q) [DYW15, CZ16, KS18, MG17, MAK15a, MAK15b, MAK15c, MAK16, MNK19]. (q, h) [WDCJ19]. (r, r − 1, r − 1) [ZZCY16]. (s, m) [DLY17b]. (x, u, v) [SMG19]. 0 [Cih15]. 1 [BAV18, Cih15, FK17, KNK15, LXPD15, OGC19, VG18]. 1 + α [BMO15]. 1 + \sqrt{2} [RA15]. 1, 2 [WX18]. {1, α, 0} [CCY19b]. 1/2 [Fuk15d]. 1/coshn(αx + βt) [VDI17]. 2 [ADSS16, ABGL17, AV15, AO15c, BZ17, BFR18, BP17b, BS19b, CSV19, CLR16b, Che15e, CW17, CJ16, CVA18, DHX^{+}19, DDD19, DA19, Dob19, LYDH18, LZL15b, LMMZ18, LLdM15, LLdM16, LA15, LT18b, MN17, MS15d, NKP16, NSM16, NB19b, ODP17, OA15, ÖK19, QY15, QFTT15, SZL^{+}18, TX18, TSA15, WY17, WDG19b, XJW17, YFL16, ZWF15, ZLDZ16]. 2 + 1 [ACE19b]. 2 + 1/2 [JL15]. 2 1/2 [LL16c]. {2, 3} [SKP16]. {2, 3˜} [PS15a]. {2, 4} [SKP16]. {2, 4˜} [PS15a]. 2 − (v, k, 1) [LLL17a, LLL19b]. 21/2 [Zha16c]. 2^{m}p^{n} [YFH19]. 2N [HEA15a, HEA15b]. 2 × 2 [EEST16, HSSH18]. 3 [AV15, ATX19, BKW18, CWG15, CSV19, CDL18, DL16, Ers16b, GLW18, GT15b, JOÖ16, JDL15, Jin17, KCWK15, Kan16, LTY16, LMYW17, LZLG19, LTL16, LMMZ18, LDL15, MGK^{+}16, MKN15b, NKP16, ÖAD15, ÖÖAD15, Qia16a, RZH^{+}19, RSC16, SA19, SuSR15, SKCS15, SH19, SDWG19, WL16c, WW18, Yan18a, YLL18, YZ17b, YB15, YL19c, ZWaY18]. 3D [GCM^{+}18]. 3 × 3 [XLW15]. 4 [ACRZ18, LYW17, QY15, RS16a, RGH^{+}15, SuSR15, SuSB15, WL18a, WD19, YCH19, ZT16a, ZH19b, ZLSX18]. 4 − (v, k, 3) [DL18a]. 5 [GLQ^{+}18, GWX17, Han16]. 6 [LWZ15, RS15b]. 7 [Xu19d]. [−1, 1] [ZMQ16]. [1, 2] [CMSZ18]. ^{+} [MK15b]. ^{p} [DL17b]. ^{rα} [CVMGPPGC15]. ^{th} [ZS16b]. _{2} [ATK^{+}16, KSKW^{+}19]. _{3} [ATK^{+}16]. _{3}F_{2} [SS15b]. _{4}^{+} [HJ18]. _{n} [GL18b]. _{x} [MNYL18]. A [RSKB15, Wan17c, HM17]. A − ϕ [KCZK15]. A_{α} [CLWM19, LCM19]. AB [GL15b]. α [FPUK15, LT15, SPKC16, YwFW^{+}19, YKLS19]. AX = B [ZWLZ15]. AXA^{∗} = B [Liu15f]. AXB + CY D = E [KM15b]. AXB = C [LZZ^{+}19, TTLX17]. AXB = D [YZ15c]. B [BM15b, Han15a, KBS17, VK19, HM17, RDLG16]. ḡ [Jan15]. β [DÇO15, GR17a]. C [GLS16, GLS19, LLL19a, WTD^{+}19]. C[a, b] [KM19]. C^{∗} [MD15b]. C^{0} [Pöt19]. C^{1} [GK16b, NKP16, SA19]. C^{2} [LLz18, SN15c, Zhu18]. A [RS17a]. C^{0} [ALP18]. H_{2} [GR19b]. H_{∞} [GSPH19, LPJ18]. L_{2} [LDS^{+}18]. L_{∞} [LDS^{+}18]. CXC^{∗} ≥ D [Liu15f]. D [CGAL19, LDL15, ZZX18, BGMZ17, GFE19]. Δ [WA15a]. Δ_{k} [TQ18]. E [MK15a]. e_{j} [Bir15]. Ed(2)(Z) [SZ18a]. ℓ_{2} [LSCK15]. ℓ_{∞} [LSCK15]. ℓ_{ρ}^{p} [ZZ16b]. ϵ [GK16a, HTZY17, RR15, SM18]. η [FPUK15, HWZ17, Liu18c, RWH15]. F [Dan15a, CYZ17, CDK15]. F_{c,2n} [DÖA15]. F_{p} + uF_{p} + vF_{p} + uvF_{p} [ZK17]. F_{s,2n} [DÖA15]. G [CLT18, Dan15a, RWY19, WY19, LHG18, WM18a]. G^{3} [FKV16]. G_{b} [LZCW19]. γ [YL18]. GM(1, 1) [WLFX15]. H [LGK15, WS16, ZGL15, ZVL19, MNNA15]. H^{1} [HZLY17]. H^{2} [JKK15]. H_{1} [BS19b, ZZ15a, CH18b]. H_{2}/H_{∞} [GSZ15]. H_{∞} [CHL15, CLCW17, KPP17, LCY16, LSZY17, LSLZ18, LZ18c, LSZ^{+}19b, LXTF18, LTZ^{+}17, MFJZ15, MC15a, MJL16, QG16, RWW18, SMG19, SMZ^{+}17, SLS^{+}18, WWZD16, WRX^{+}19, XXW^{+}19, XZHX17, YKJY17, ZLDZ16, ZZL16a, ZSS16, ZML^{+}17, ZJYD17, ZPM16, ZSL^{+}18]. H_{∞}/l_{2} [HXW^{+}19b]. hp [Gia15a, Dol15]. I_{2} [ZML^{+}17]. I_{∞} [ZML^{+}17]. J [ZY19]. j ≥ 1 [Bir15]. K [Her19, ZSA^{+}15, AM18d, BGMO18, BS19c, CHL17, CLL17, GZhL^{+}19, HLX17, HL18, HHT19, KS19a, LLW17, LZ18e, QGH19, RP18, RR18a, STS15, SRT18, SJL17, WGL18, YEMRV17, ZY18a, ZWL18, dFY15]. κ_{3} [LW18b]. kp [ZZLY15]. L [Bab16, BB15a, DL17b, PS16, ZY18a]. L(Z) = L[V (X + Z)] [Zha19c]. L = (1/2)(−u^{′2} + u^{2}) + ϵ^{i}G_{i}(u, u′, u′′) [JM18]. L^{2} [SST18, YB15, ZL16a]. L^{∞} [LZ17f]. L^{p} [APTS15, TY16]. L^{p}(R_{+}) [KAH15]. L_{1} [CCQS16, ZZC16, LX16b, ZWY15b]. L_{2} [BS19b, SZPH16, WR19]. L_{4n} [DÖA15]. L_{∞} [SZPH16, TF18, HXW^{+}19b, LX16b]. L_{p} [AK15a, YSD^{+}17]. L_{q}(q = 1/2, 1, 2) [CC18a]. λ [AS19b, RM18b, SGG15a]. Λ^{2} [BLS15]. λ_{1} + λ_{2}u + λ_{3}v + λ_{4}uv [ZK17]. λ_{3} [LW18b]. λμ [HQJ^{+}15]. ⌊(n − 3)/4⌋ [CL15a]. Lip(ω(t), p) [Zha15d]. M [Gua19, HM15c, APK15, DPL^{+}18b, Rom19, SA18b]. m^{th} [LKM17]. H_{∞} [Kim19, YCX^{+}19]. R^{2} [Vol15]. R^{3} [Hua15, Qia16b, Val15, ZZ15d]. S^{3} [AMN19]. Z^{3}Δ [PBMP15]. Z_{p} [KKD15]. C^{(∞)} [PBMP15]. H_{∞} [CYSY16, SYZP16, SS15c, TPL16, XSSW19]. N [KKD17, LXPD15, LDL15, ARB19, CZ19, DLY^{+}17a, HKM18, KLS15, LZH16, PP12, PP15b, San16, SV17b, XYDL16, ZYH15, ZWK16]. ω [SXS15, dSH15]. p [AS19b, BM19b, BC16, CHRD16, CCC19, FFT15, HLZ15, Jia15, KKD15, LJ15a, Lv18, ÖAD15, ÖÖAD15, XHW18, XHW15]. p(·) [GW16]. p(x) [All18b]. p^{5} [CC18c]. P_{3} [ZTW19]. P_{4} [BKKBS19, DVJ18, Yue16]. P_{4n} [DÖA15]. P_{k} [Tu17]. ϕ [GS11, JK15]. ψ [KCWK15]. PSL(2, q) [DL18a]. Q [JFT17, YZYgG18, Acu15, AK16a, AFK15, ANT^{+}16, AHG15, BM15b, BA17, KIAG16, KTA15, KAG15, KM15c, Koc15, Sha15d, SIA15, TZ19b, Uça16]. Q_{1} [AGZ15]. q ∈ (1, 2) [Zha15f]. R [And18, BMRW19, Ili18, JXW15, JZ15a, LT19, MWL^{+}18, MC15b]. R^{2} [ZL17b]. R^{k} [HTM18]. R^{p} × R^{q} [AOS19]. rd [WA15a]. RL_{β}C_{α} [JCW18]. S [DFP15, HPR16, SXS15, Yor15, ZLDZ17, dSH15]. σ [Rét19]. ∑_{i=1}^{s} A_{i}XB_{i} + ∑_{j=1}^{t} C_{j}Y D_{j} = E [YZ15b]. Sz(q) [LLL19b]. T [BLM^{+}16, KCWK15, ZLDZ17, ZWD15, Har15]. θ [ARB19, HH18b, MDZ17, TRP15, ZCXX18]. U(a, b, z) [GST15a]. uu_{xx} = −2 [Boy17]. φ [ZS15d]. W [Kyr15, Miy19, WMS17, LR19a]. X + M(X) = C [LH19b]. x_{n} = x_{n−1}y_{n−2}/(ay_{n−2} + by_{n−1}) [SDIS15]. X ± A^{∗}X^{q}A = Q [LWS15]. y′′′ = f(x, y) [FT17]. y^{(4)} = f(x, y) [FT16]. y_{n} = y_{n−1}x_{n−2}/(cx_{n−2} + dx_{n−1}) [SDIS15]. Z [KCSD^{+}18, Pos17]. Z_{t} [WDB18].

-adaptive [Dol15, Gia15a]. -adic [KKD15]. -algebras [MD15b]. -algorithm [GK16a]. -analogue [CZ16, MAK15a, MAK15b, MAK15c, MAK16]. -anti-Hermitian [Liu18c]. -Appell [KM15c]. -arboricity [HHZ18]. -ary [HM17]. -asymptotically [JK15, DFP15, HPR16, SXS15, dSH15]. -axis [Han16]. -B-spline [HQJ^{+}15]. -basis [Han15a]. -Baskakov-Beta [MG17]. -Bell [MC15b]. -Bernstein [BA17]. -Boussinesq [VK19]. -Brownian [WY19, RWY19]. -Cayley [HKM18]. -circulant [CZ19, JXW15, JZ15a]. -class [Zha15d]. -colorable [ZLSX18]. -coloring [CYZ17]. -compact [WR19]. -complex [ZSA^{+}15]. -connected [Dob19, SJL17, TX18, ZWL18]. -connection [LLW17]. -connectivity [GLW18, LTY16, LMYW17, LW18b, LZLG19, ZH19b]. -constacyclic [ZK17].

-continuous [BM15b]. -convergence [LJ15a, ZZX18]. -convex [DLY17b, MNNA15]. -critical [YL18]. -cycles [San16, Xu19d]. -cyclic [WTD^{+}19, SdR19]. -cytokine-host [HJ18]. -D [DHX^{+}19, WL18a, Che15e, FK17, LZL15b, LDL15, WD19, ZT16a]. -dependent [SMG19]. -designs [YFL16]. -determinantal [GLS16]. -difference [AK16a]. -Dimensional [YJSW16, ARB19, BGMZ17, CL15b, DPL^{+}18b, FMC15, HGH19a, KLS15, LL18a, MMIA19, OW18, SV17b, VK19, WTXZ16, YCH19, ZH15a, ZW15b]. -distance [LR19a]. -divergence [CDK15]. -domination [CMSZ18]. -Drazin [CLT18]. -dynamic [WA15a]. -eigenvalue [YZYgG18]. -eigenvalues [LLL19a, WDB18]. -error [LZ17f]. -extendible [DVJ18]. -extra [WM18a]. -Favard [KTA15]. -finite [Su19]. -fold [APK15]. -fractional [AS19b, Koc15, STS15, TZ19b, WDCJ19]. -fullerenes [SZ19c]. -gain [LX16b, ZZC16]. -Galerkin [CH18b, HZLY17]. -generalized [SGG15a]. -good [LHG18, LT19]. -good-neighbor [WY17, WX18]. -graphs [MBW19]. -Hamiltonian [ZY19]. -Hamiltonian/skew-Hamiltonian [ZY19]. -Hermicity [HWZ17, RWH15]. -index [Rét19, YKLS19]. -infinity [LGK15]. -integers [Acu15, AFK15, BM15b, KS18, KAG15, SIA15]. -integral [Jan15]. -integro-difference [ANT^{+}16]. -inverses [PS15a, SKP16]. -labeling [GZhL^{+}19]. -Laplacian [FFT15, HLZ15, All18b, CHRD16, CCC19, GW16, Jia15, Lv18, XHW18]. -lattices [AHG15]. -Lupas [Sha15d]. -main [JFT17]. -matrices [WS16, ZVL19]. -matrix [Gua19, ZGL15]. -meshes [BLM^{+}16]. -method [MDZ17, TRP15, ZCXX18]. -metric [RDLG16, YEMRV17, LZCW19]. -Milstein [ARB19]. -minimization [ZWY15b]. -modes [CHL17].

-mono-components [HTZY17]. -monotone [FPUK15]. -nonexpansive [SPKC16]. -nonnesting [Yan18a]. -norm [TY16, YSD^{+}17]. -number [KCSD^{+}18]. -numerical [GLS19]. -optimal [GR19b]. -order [BMO15, SMM18, ZZCY16]. -orthogonal [AHG15, SST18]. -out-of- [LXPD15]. -outer [DYW15]. -partiteness [HLX17, HHT19]. -path [GFE19, LZ18e, RZH^{+}19]. -paths [San16]. -periodic [dSH15]. -permutation [Yor15]. -person [DLY^{+}17a]. -piecewise [WA15a]. -planar [OGC19]. -point [ACRZ18, HM17, RS15b, RS16a, SuSR15, SuSB15]. -policy [ZWD15]. -polytopes [Cih15]. -positive [SXS15]. -proper [CLL17]. -QG [GR17a]. -quadrilateral [HL18]. -quasiconformal [SA18b]. -rainbow [BS19c, SRT18]. -rank [ZYH15]. -reducible [Yue16]. -refinable [Rom19]. -regular [GWX17, Jin17, LYW17]. -restricted [BGMO18, BM19b]. -Ritz [ZZLY15]. -section [PP12, PP15b]. -series [BB15a]. -setting [Pöt19]. -space [GT15b]. -spaces [Bab16]. -sparse [DVJ18, Yue16]. -species [XYDL16]. -spectral [CLWM19, LCM19]. -splines [PS16]. -splitting [HM15c]. -stabilization [CGAL19]. -stable [RSKB15, RS17a, Wan17c]. -starlike [Uça16]. -step [RP18, RR18a]. -Stirling [BMRW19]. -strongly [GS11, ZS15d]. -subtree [YwFW^{+}19]. -summability [KBS17]. -super [MK15a]. -symmetries [RM18b]. -Szász [MNK19]. -tidy [BKKBS19]. -total [DSZY19]. -transform [Pos17]. -tree [QGH19, ZY18a]. -trees [ATX19]. -triangle [HL18, WGL18]. -tridiagonal [dFY15]. -type [VK19]. -uniform [GLQ^{+}18, SM18]. -valent [AS19b]. -values [Har15]. -vector [AM18d]. -vertex [KS19a]. -vertex-connectedness [HTM18]. -Weighted [BLS15, Kyr15, WMS17]. -Wright [BC16].

/passive [SS15c].

1 [Amm15, KGGAV^{+}18, Boy15, FQMR19, LXT17, SJ19b, WGLZ16, YLW16, ZWD15]. 101 [EL19a]. 149 [Sab18]. 155876 [SPGP18]. 191408 [SPGP18].

2-extra [WY17]. 2009 [LLL16a]. 2014 [TT16, ZS16b]. 219 [Pan15, PP15b, Tom15]. 232 [AZ19, EH15]. 233 [Rho19]. 236 [ZZ15c]. 246 [PK18a]. 261 [Pon17]. 264 [MAK15a]. 265 [Bag19, DN16]. 266 [MAK16, VK17a]. 268 [YLWF16]. 291 [WL18a].

305 [LLZD18a]. 333 [SN18a]. 334 [FMB^{+}19]. 346 [ZL19b]. 362 [PPS25, RRR20].

4th [CN16a].

5- [YL18].

60 [SPGP18]. 601 [DN16].

9 [ZS16b].

Abaffian [AP18b]. abatement [ZWY19]. ABC [Ers16a, DU18, LCW^{+}18, SWG^{+}17, WZ19]. ABC-flow [DU18]. ABC-minimal [WZ19]. Abel [BS15c, De 17, SAC18]. Abel-type [BS15c]. abelian [ML17, ACKS16]. ability [LZ17d, ŚKK^{+}18]. above [LCOS19]. ABS [AP18b]. ABS-type [AP18b]. Absolute [Bor15a, ZZC16, EHS17, GLL15a, KM17b, MYH15, MKN^{+}15a, Zha15b, ZG15a]. absolutely [WZWJ18]. Absorbing [RMD18]. Absorbing/emitting [RMD18]. absorption [ZJZ15, ZMH18]. abstract [HO15, HPR16, MD18c, WLZW15, dSH15]. Accelerated [DFSC15, Pet15b, Bor19, KDS16, LM19, LMMR15, PV16, UHLM15, ZM18b].

Accelerating [LCZ19, XY18]. acceleration [HP15c, PPS25, Pon17, PP18a, RRR19, RRR20, SSG15b, VS19]. accelerators [SLTH15]. acceptance [XYZK15]. access [HHR17]. according [GN19a]. account [SWSS18]. accounts [FHM16]. accretion [GZG^{+}15]. Accretive [TÖGE18, CS15b, GS11, KT16, ZS15d]. accumulation [WLFX15]. accuracy [AKL19, AAC16, AP15, JSM16, LW15b, MDH15, MK16, NRY16, PZ17, RAV15, YCZ19, ZZM18]. Accurate [DP15b, DP17, DBJC17, JGBY16, Lam15, SS18c, YDWZ19, FK17, FG18, Gra18, GR15b, KSZ15, LYDH18, MSL15, PVL19, SM19, XH19b, ZYY17, ZCU15]. Achieving [XY17]. acknowledgments [AM19]. acoustic [AAC16, CDP19, LLX18a, LYDH18, Mis15, PFH^{+}15, QA19, RL19, Wan16b]. acoustically [KDS16]. across [How15a]. acting [ZL19a, ZL19b]. action [Ace19a, Cha19, Kri19b, Lu16]. Activation [SRJ^{+}15, CCH15, NLC19, TLX19]. activations [PRC^{+}19]. active [ABD16, BFJ^{+}17, BL15, CK17, DG15a, GMPP18, MGPL15]. actuated [DMA^{+}19]. actuator [FM16, KLKU18, KRA^{+}18b, MFJZ15, MJL16, QG16, QKGW18, SKKL17, SJWK18, TFLZ19, WWHZ18, ZPZ^{+}19, ZW19b]. actuators [LLZ17]. Acyclic [CR19b, Yue16]. acyclically [ZLSX18]. Adamchik [You15]. Adams [ZJJ18, WNMC15, YR15]. adaptation [OA18, RAV15, RDG18]. adapted [FKR15, GVA16, ATV16]. Adaptive [ARS15, BCNN19, BS15a, Bor19, CLPY18, DR19, Gia15a, HR18, LY15d, LABK15, LK18, zMnZyZ15, ODP17, SGY^{+}15, TDL^{+}17, XD19, XCDZ19, YSWZ17, YWH15, ZDPZ15, ZW19b, AHK^{+}19, BB15b, BBGR15, BDH16, CC18a, DZD^{+}16, Dol15, FY19, GHL19, GSD16, GYZC15, GHZ19, JM15, KC19, KLKU18, KKS^{+}16, KMP15, LS17a, LLCR15, LLC18, LZjy18, LSX19, LMZ^{+}19, LGL19c, MRQZ19, NB15, QFTT15, Raz15, SGKR19, TVK^{+}16, TFLZ19, WZF^{+}19, WWHZ18, XZLZ17, XZLL17, YBZ18, YNMW16, ZWF15, ZZL16a, ZLM^{+}18, ZWY^{+}15a, ZYLY17, ZSXW18, ZWO^{+}19, ZQG16]. adaptive-feedback [LLCR15]. adaptive-gridding [GYZC15]. Adaptively [BJM16]. adaptivity [Cha19, KMP15]. adding [TSA15]. addition [Far18b]. additional [LJHA18, Sah15, SSB15a]. additive [BCO15, DJ15, HMD19, LZH^{+}18, LT18a, SM15b, SSZ^{+}18, SZ15, TM19, Vis15, XCCL18, YYP18, ZGL15]. Adequate [DGL^{+}15]. ADI [BTWY15, CDL19, DL18b, LZ17a, QXW19, WW18]. adic [KKD15]. adjacency [DGF17, TS18, XLYZ19]. Adjacent [Khl18]. Adjoint [DKL15, QFTT15, Lia15b, Liu15g, YMeFG19]. Adjoint-based [DKL15, Lia15b]. adjointness [TBG16]. adjusted [YJYH15]. adjustment [ASV17, LSX19, YLH^{+}19]. ADM [KS16b]. Admissibility [GHQ17, FLL15, PLL16]. admission [SJ19a]. admit [CN19a]. admitted [ASM16, JM18]. admitting [LLL19b]. ADMM [Gon18]. ADP [CXW^{+}19]. advanced [ABLZ15, CDJ19, ZS15b]. Advances [DGRR16, SGGD^{+}18, Far16, Zho15c]. advantageous [CFdCdS18]. advection [ADG16, BKW18, BKLW15, CL19a, CGW16, DBK15, DN15c, FM18, GPHAPRW16, HGH19a, HP17, HV19b, KD16, LL15e, LRL18, QG19, VD16, WA18, Zha18b, ZBH^{+}15, ZZL16b]. advection-diffusion [ADG16, DBK15, KD16]. advection-diffusion-reaction [BKW18, LRL18, WA18]. advection-dispersion [Zha18b, ZBH^{+}15]. advection-dominated [CGW16]. adversary [DLCV17]. adverse [XP17]. advertising [MZ15, WWZ19]. advice [GOH18]. AE [HP15b, Pop17]. AE-solution [HP15b, Pop17]. Aedes [ZGH16]. aegypti [ZGH16]. aerated [KKM15]. aerodynamics [ALDHA^{+}17]. aeroelastic [Ber15, Ili19b, Ili19a, SH15d]. aerospace [TT16]. affect [Kur19, Tan17b]. affine [BC16, Fio15, Zha15i, ZLZW15, ZSZW16, ZYZ^{+}16a]. Africa [DCK^{+}18]. AFT [PW19b]. aftershocks [CVJ19]. against [SJWK18, SWA^{+}19]. age [AA16, CMW17, FLMC15, FG15, KLZ19, LKZ18, LYZ19, MDZ15, PYZS18, TRP15, TMPG17, Ter15, WWK19, XGZ17, YW19a, ZG18, ZYZ15]. age-dependent [KLZ19, LKZ18, MDZ15, PYZS18, TRP15, TMPG17]. age-structure [ZYZ15]. age-structured [AA16, CMW17, FG15, LYZ19, WWK19, XGZ17, YW19a, ZG18]. agent [CLZW19, FW18a, GSD16, HWY15a, HW17b, HCY^{+}18, LL15b, LZjy18, LS19c, LSL19b, LGHW19, MM15b, Ma17b, MGCAPP16, SM17, SSCX18, SMG19, SY18a, SWZL19, TFLZ19, WS19a, WXXY19, XH16, YJMH18, ZLM^{+}18, ZJ15b, ZYLY17, ZJYD17, ZSL^{+}18]. agents [AJ15, hCWW^{+}15, Lu19, WYW16]. agglomeration [LLMS16]. aggregate [AGM16, TDA17]. aggregation [LMF17, MPLS19]. AGM [ANGC17]. Agros2D [NKKS16]. ahead [Lah18]. AHP [KWGRP16]. AHSS [ZM16]. aimed [RM16]. AINV [RBB19]. air [DMSS18, FCW^{+}17, FVMO15, GSDH15, RSLL^{+}19, SWSS18, ZHWW15]. air-filled [GSDH15]. airfoil [SH15d]. Airy [BCD^{+}19]. Aitken [PC16, YKZ15]. Al [ATK^{+}16]. ALE [JOÖ16, GS17a]. ALE-SUPG [GS17a]. algebra [SSA19, VD15]. Algebraic [ABGL17, BN15, FFN18, ZJJ15, AAH19, BM15a, BBF^{+}18, CD15b, DLQ17, FMB^{+}18, FMB^{+}19, FNS19, Gua19, HW17a, KDS16, QTF^{+}16, QYF^{+}18, SR19, WSZ18]. algebras [ASM16, AADMM16, BCSG18, FFN18, HS17, MD15b, RMSD16]. Algorithm [BKZA18, CYCW15, EH15, EL19a, LYW18, AT15a, ATV18, ALMCDH^{+}18, ASA19, AADLF15, AIG17, ACC15, Arà16, ACC16, AF15, Bas19, Bog18, Boi15, BFJ^{+}17, BPR18, CHL17, CC18a, CJ15, CCX15, CKW15, CB17, CLT^{+}15, CPS18, De 17, DP15b, DCS15, DNC^{+}17, DBJC17, DZY^{+}18, EH14, EL19b, EL19c, FWLC15, FLCSV18, fGlHyL^{+}15, wG15a, Gar16b, GK16a, Gol15, Gra18, GOA19, HSAR19, HMF15, Her19, HP15c, HGM15a, HM18a, HCY15, HTZY17, HM18b, Hua18, Ili19a, IKHS16, JPTL15, JLL15b, JLX15b, KSK^{+}15, KTD15, KSS19, KLZT15, LHJ15, LQZ^{+}18, LLz18, LY18a, LCZ19, LH19a, LABK15, Lin15, LS15b, LF15, LWWW15, Liu15b, LWSM16, LG19, LDL15, MZP15, MH19a, MBHS18, MFF15, MN15, MPLS19, Naa15, NSR^{+}19, OW18, PKHD18, PZ16a, QTF^{+}16, RCTZ18, SSK^{+}19, Sch16, SNAR15, SBG^{+}17, SRH15, SY18a, SKBR18, SDK15, SKCS15, SZRC15, TD18]. algorithm [TSJ16, TTD18, Tsy19, Tu17, UHN^{+}18, WW16a, WZL15, WXW18, WEH18, WYL^{+}19b, WLZ19b, WF15c, Wu17, WLFH17, Xu15b, XLW16, YWS18, YLX18, YZ19, YZMM15, Zad16, ZBKS16, ZN15, ZYY17, ZLZ^{+}19, ZTW19, Zha15i, ZCL18, ZSZW16, ZYZ^{+}16a, ZXDP18, Zhu15, ZWY15b, ZK18, ATV16, GGAVGGRC15]. Algorithmic [Shi15b]. Algorithms [BKS18, Alz15, ASZD18, ABB^{+}18, BLP17, CLY17, DZW15, DG18b, lDqCWzH18, FPKF15, GdS18, GMPP18, GJP18, Guo16, HW17a, HS15c, HM15b, cJzS15, KZZ15, KM18b, LYW17, LM18, LH19b, LWS18, LP15c, MMG15, MM18b, MMB15, PG15b, PS15c, PL15d, QW19, QLZ15, RZH^{+}19, SA16a, SIA^{+}17, SD17, WDLS18, WL17, WSZ18, XS15, YJW16, ZM18b, ZS19b, ZLZW15]. alighting [YYX^{+}19]. Allee [MMS17a, PZ16b]. Allen [LL19a, LG17, VD15]. allocation [DCK^{+}18, GKDHR19, KPLC15, LJ17, SSS15, WHC18, YZ15a]. allowable [JYVS15]. Almost [APS16, LZ15b, LDH18, XZCC15, XDH19, APL15, Bor15a, CZP15, DZ16, LZ19c, MZM15, MO16, RN19, WA15a, ZZ15b]. alone [FFH^{+}18]. along [KHN19, SBH17, TÖGE18, WLQ^{+}19, ZCPM17]. Alsaedi [PK18a]. alternately [Gua19]. Alternating [KM15b, MN17, SPW15, GLL15b, HW19, HH17, LYXL15, LDL15, MNNP19, WY17, YLCF17, ZM16]. alternative [Liu17a, MYHS18, PMC16]. aluminium [NDZ15]. always [SSGG17]. ambient [SNS^{+}16]. Ambrosetti [GS15b]. AMC [EH15]. American [LLZD18a, ASV17, CW17, LD17, LW15a, MBHS18, RPB15]. American-style [LLZD18a, LD17]. AMF [GPHAPRW16]. AMF-W-methods [GPHAPRW16]. Amin [WL18a]. among [LSX19, LP16a, Par18b, SdA18]. Ampère [MG15]. amplification [DZ17a]. amplifier [SRH15]. Amplitude [NC16]. Amplitude-frequency [NC16]. Analog [SMM18, TD18]. analogue [CZ16, MAK15a, MAK15b, MAK15c, MAK16]. analogues [TCY15, WYL19a]. analyses [LCL17a, LX19b, LL19b, NMNTBXVD15]. Analysis [ABD16, BMDS17, BM18b, Chy16, DU18, DLCT18, EFPT18, FQMR19, GAB17, HMT15a, JCW18, JT16, KKD17, KSS19, KG17, LW15b, LD19a, LZL^{+}19b, LZ17c, LSZF18, LYXL15, LZ15d, LDD15, Liu15d, MMS18, MGN18, MPHG15, NSM16, QA19, SPG17, SSM17, STK15, TVK^{+}16, TLHC17, Vid18, Vil16, WWK19, Wei17, YWL18, YMMH15, YLL^{+}16, ZS16b, ZPT^{+}15, AR16a, ARB19, AM15b, ADSS16, AB15b, ACDT18, Amm15, AHY19, AB18, AAH19, AADF18, Bag19, BMM19, Ber15, BMT19, BF15, BRV18, Boy15, BCD^{+}19, BE17, CSZ15, CWFL15, CKW15, Che15e, CD16a, CS18, CS15a, CN15b, CN16a, CNK^{+}18, CLP^{+}18, CDFP17, CGMT16, Cui15, CnTVsL18, DDY15, DWRZ17, DN15b, DN16, DA17, DW17, DRS16, DV15, DHS19, DL15a, DNC^{+}17, DFJ16, DOZ18, DBJC17, DMSS18, EAY17, Elm15, EYE18, EGB15, ES17b, EHVV15, FLL15, FM17b, FMRRT15]. analysis [GS15a, GHQ17, GhDwZS17, GSDH15, GK15, GZW15, GZLL17, GYC19, GSV18b, GSVK17, HN14, HL15a, HSS15, HTZY17, Hua18, Ino16, JDL15, JSW^{+}18, JPZ^{+}19, KCSD^{+}18, KLZ19, KM15a, KLW15, Kim15b, Kim15a, KSL19, KSN15, KZ19, KVS15, KKM15, Kri19b, Kun16, KKG18, Lah18, LLC^{+}15, LAVME18, LK15, LL15d, LCW15, LJ15a, LC16a, LX16b, LWL^{+}18, LQC^{+}18, LJYH19, LXD19, LZ19a, LHHS16, LdCC15, LLZ^{+}16, LZG16, LYZL18, LXZ19b, LXZL19, LCW19b, LFC^{+}19, LLX^{+}19, LWZZ18, LZH^{+}18, LJHW19, LJ15b, LCLZ19, fLLPY16, MHN15, MWWM15, MWLL17, MA15b, MS17b, MPW15, MSM15, MPPP19, MW18, zMnZyZ15, MLG18, MWY15, MKMC15, MR16a, MB17, MFMK17, MK15b, NR16, NM19, Nas19, NS15b, NLC19, NKKS16, NS17b, NN19, Oka18b, PAM19, PTA16, PZJ17, PL15c, PD15, PBB18, Pon18, PRMC19, PRC^{+}19, PP18b, Rad15, RZSS15, RzS15, RWY19, RSAF16]. analysis [RL15, RHJ18, RM19b, RC17b, SJD^{+}18, SFLM19, SM15b, SRJ^{+}15, SA16b, SSS17, SBG^{+}17, SY15a, SSA18, SAES19, SLW16a, SW17b, SW17a, SY18b, SY19, SRS17, Ste19, SRBA15, SZT17, SW17c, SC17, SCM19, TLX19, TXG15, TXL19, TMM18, TB15, TZD^{+}19, VLCT18, VVD15, Wan17a, WXY15, WPL15, WA15a, WSZ16, WMZ16, WK16b, WYCZ17, WL18a, WCH19, WSC19, WWZ19, WSXR19, WCB19, WWC^{+}18, WZY^{+}19, XZ19, XW19, Xu15a, XCL18, Xu19b, XS19b, YWL15, YLHS17, YZ18b, YWYS18, YYM19, YLL15, Yil16, YmZHC15, YZ16, YCLS19, ZZZ15a, ZP16, ZT16a, ZLHH15, ZLW19, ZLDZ17, ZL15a, ZLJ15, ZLWW15, ZZLY15, ZZC16, ZZZ16, ZHJ^{+}17, ZSZ17, ZYC18, ZG18, ZW18, ZWBC18, ZC19b, ZWY19, ZHX^{+}17, ZGZ16, ZYZW16, dDSI19, dFGAN18, uRAD19]. Analytic [ABN^{+}17, LZ19f, LZWZ19, RLH15, RC17b, ZYJ15, AZAM15, CNQRR18, DCT19, ERFE16, FÖB19, GGL15, GA15, HZ17b, KMP19, KMT16, MF17, MFOP17, Mil16, RS15a, SGG15b, VK19, ZSY15]. Analytical [HG18, KZ15, KR16, LRF^{+}18, MDH16, SSB18, SG18, WWZ19, YRA15, ZBR15, ZLLA19, ZS17, Ber15, CM18, EYE18, GHN^{+}16, HN14, KZA19, KD18, KJ15b, Kor18, LHHS16, MWH^{+}18, NN19, RSO^{+}18, SHA15a, VDV15]. analyticity [ARSV15]. analyzer [CMW15]. Analyzing [FSW16, LLMW18, LLMS16, RG15, SA16b, SLL^{+}16, WQ15]. anchor [STK15]. anew [ZLDZ17]. ANFIS [SJ19a]. Angkor [PG15a]. angle [BCK15, HV19a, MZP15, MRF18]. angle-ply [MRF18]. Anisotropic [Dol15, BE19, BMM19, cCpSkZ15, GB19, LRL18, NM19, Por18, SS18c, SZL^{+}18, SMR15, Sin17, YW19b, ZQG16]. annealing [Mat15, Pal15, PTO19, RRTR15]. annuli [FW18b]. annulus [GA15, LLdM16, Mah15, NKNA16]. anomalous [PeO18]. Anopheles [SWA^{+}19, MTMB17]. ansätz [ND15]. Ant [DCA15, GWC^{+}16, SZRC15, WW15]. Anti [GZhL^{+}19, Jin17, AM15b, BPC15, DMR15, DZ17b, HGM15b, Liu18c, SAM^{+}16]. Anti- [GZhL^{+}19]. anti-bisymmetric [HGM15b]. Anti-Ramsey [Jin17]. anti-reflective [DMR15]. anti-symmetric [AM15b]. anti-synchronization [BPC15, DZ17b, SAM^{+}16]. antibiotics [WLG17]. anticipated [DWS19]. anticipation [RHJ18]. antiferromagnetic [KSD^{+}15]. antigen [PRMC19]. antigen-driven [PRMC19]. Antimagic [LMS19, YHY^{+}18]. antimaximum [CCC19]. antiperiodic [ALP15]. antiretroviral [MK15b, XGZ17]. antisymmetric [MRF18]. any [GG19b, LZS15]. AOR [HXLC15]. aperiodically [LCLZ19]. APFs [BL15]. Apostol [BN15, HASA15, Luo15]. apparent [BKZA18]. Appell [KM15c]. Appl [MAK15a, MAK16, Sab18, Tom15, VK17a, WL18a]. applicability [AM15e, AM17]. Application [AKAA15, AQ15, CLR15, Fio16, HSAR19, LGO17, MH16b, PZJ17, SPS18, Sow18, USCAAA15, ZKV^{+}15, ATV19, Alf17, ÁL15a, AV18, AY16b, Ata16, BKW18, BS15b, Bot15, CH18a, CP15a, CCD18, CWL19, CCL^{+}16, EFPT18, HRS15, HQJ^{+}15, HYL^{+}19b, JGBY16, KS16b, LLL16a, LW18c, LABK15, LX19d, MPR^{+}17, NKS18, PSP17, PCA16, QZL18b, RDLG16, RSLL^{+}19, SRSZ15, SZ19a, SRH15, SGFCRM^{+}18, Suk19, SD19c, TD18, TB15, Vol15, WQK15, WDG19a, WLZ15, WE19, WWMX16, WZ18b, XS19a, XCCL18, ZZ17a, ZMA^{+}18, ZLD18, ZL19a, ZL19b, dSH17, CDP19]. Applications [AK18b, Awa15, EHdLS16, AED15, AHG15, ALQ15, AHP19, AL15b, AVC^{+}19, ACSS15, BB15a, BLK19, BMRW19, BM19d, Bou18, CDK15, CC15, CJ17b, CKSV19, CT17b, DP15a, DOR18, DD19, Dok18, DBJC17, EAAAS15, EJF16, EYE18, ES16, GT15a, GLL15a, GHL19, Gav15, GE16, GDF^{+}17, HWZ17, HZL18, HL18, JS15, JDS18, KBS17, KM19, KOR19, KKT19, KZ19, KSP16, Kun16, LLO16, LH17b, LW19c, LZCW19, LAAZH19, LWWW15, Liu16b, LZZT18, LB19, LHX15, LYY15, MZ15, MS19b, MSS15, MNNP19, Mor17, NA19, NKKS16, OB18, PCW15, PPV15, Pon18, Pro16, QZ15a, RZD^{+}19, Reh18, SSK^{+}19, SA19, SWY18, SY18c, SI15, TBSR16, TZH15, VRC15, VDV15, Wan15e, WA15a, WGL18, WZ18a, XH19b, YC15, YZ17a, YP18, ZA16, ZWLZ15, ZWD15, Zha16d, Zha18a, ZCPM17]. Applied [AZ19, Bag19, DN16, DSDL19, EL19a, FMB^{+}19, LLZD18a, Pan15, PK18a, PP15b, Pon17, Rho19, SN18a, YLWF16, ZZ15c, ZL19b, AHLR19, BBG19, BST18, FCW^{+}17, Flo18, GMPP18, HP17, KC19, LYXL15, Mis16, PPS25, RRR19, RRR20, WZZ18, Ano19w]. Applying [CCNQ19, MH15a]. apprehend [NSN^{+}18]. approach [ABMBR15, ABGS15, AM15a, ARB16, ARS15, BD18, BS19a, Bor19, BPZB18, BS15c, CCTV15, CCGP16, CLR15, CPP15, DP15a, DKN16, DK15c, DLjZ18, DBK15, Du17a, Elm15, EGH^{+}19, ET15, FMB^{+}18, FMB^{+}19, Fio16, FY18, FGPR18, GW15c, GWSC16, GGMPC15, GDS16, GIS15, GMPP18, GCM^{+}18, GWXZ18, GG15c, HN14, HAH19, HST17, HXS^{+}16, HXM18, HDG15, HI15b, Ili19b, IS16, JNGB18, KBR15, KKKD18, KLW15, KR15, KK15a, KJ15b, KCL19, KEE17, KVS15, KKG18, KAS16, Lac15, LLP17, LZL15a, LGO17, LW15a, LZZL15, LL15b, LSZ^{+}19b, LTQF19, LW19b, LGC19b, LP15b, LHZM19, LWY^{+}19, LMMR15, LZ19g, MZM16, MRC^{+}18, MP19a, MH17c, MWH^{+}18, NC16, NJM^{+}18, NY15, NN19, OLZ19, OA18, PS18a, PWFY18, PMC15, PMC16, PZLZ15, Pon17, RRRR16, RXQ19, RRV15, RM16, SWUE19, SH15b, SJ19b, SRJ^{+}15, SSA19, SYZP16, SY17a, SS19b, Shi15b, SSG15b]. approach [SPS16, TZTG17, TT18, TH19, UKE15, UMF19, UGY16, Van15, WZYS15, WDZ17, WJZF19, WX16, WMSP19, XJW17, XYG^{+}18, YWL15, YMS16, YWS^{+}19, YAIN17, YMeFG19, Yüz18, ZZLZ17, ZLDZ17, ZHJ^{+}17, ZHWW17, ZCAN19, ZCW^{+}19b, ZDZ15, ZWL16, ZHX^{+}17, ddGL15]. approaches [AKL19, AELH15, BR16a, CLXW17, JT17, NJTP19, PL15d, Sah19b, WR15, ZPRZ18]. Approximate [AS15, BT15a, CGM16, EM15, GZK16, HKL15, JM18, KCK16, LLB16, SHA15a, Sin18, Wan15c, YL17, YW16, Bag19, BS17, BKLW15, CCZ15, Deh16, EGB15, Fan15a, GGL15, LW18c, Mah18b, Reu15, RM19a, SHEBS19]. approximately [KRS18]. approximates [Kor18]. Approximating [BM15b, CRCVMS18, CB17, GZK16, Mah18b, TST16, VLCT18]. Approximation [Ade19, AIK15, AM16, BZ17, CST18, CCH15, DPD18, DMH19, GA15, GO15a, GM15b, GAS17, Hus19, KT16, MG17, MSH15, PN15b, RZH^{+}19, Tha15, TTP19, AQ18, Agr15, AD16, AK15a, ARGF15, ABB^{+}18, BA17, BH17, Bor15b, CY15a, CC18b, CVA18, DHQ19, DA17, Du15, EB16, FP18a, FA18, FY18, Fuk15c, Fuk15d, Fuk15b, Gat18a, GSV17, GDS16, GIT15, HZ17a, HGM15b, HLJ18, IZ16, IAW18, Jeo16, JCW^{+}15, KBS17, KMP19, KS18, KMT16, Kri19a, KK17, KS19c, LS17a, LCW15, LH17c, MS19a, Mah17, MRE18, Mal15, MRS18, MP19b, Mil16, MS16b, MMS17b, MDG18, MO15b, MAK15a, MAK15c, MNK19, Mus18, Oka18a, OA15, PZJ17, Pis18, PIP19, PV19, Prz15, QM15a, RV16, Rub19, Seg15, SN18c, Suk19, SH18b, TMS15, UDA15, WLLL15, Yua15, ZYH15, Zha15d, ZS16c]. approximation [ZZY18, ZLD18, ZZS^{+}19, ZZC19a, ZEW19, ZH16, Zil15]. Approximations [PKB15, Rem18, ALL18a, AL18c, Bea19, DQH18, KH17, LLC19, LSM15, NH17, NMM16, Nis15, RG18a, San16, SE18, TTD17, TF18]. Approximative [Mig15]. apps [DLCV17]. aquifer [LCW19b]. arbitrage [Cim18]. arbitrage-free [Cim18]. arbitrarily [JA17, Neu16]. arbitrary [ABC^{+}16, DEdlT17, GO15a, Gar16a, HAC18, JOÖ16, PMC16, SST18, XZW^{+}15, ZXDP18, dlRGB16]. arboricity [HHZ18, Xu19d]. arc [BGMO18, CC18c, FHK^{+}19, LLL18a, Val18, YFH19]. arc-transitive [FHK^{+}19, LLL18a, YFH19]. arched [ANGC17]. architecture [GWK15, KC18c, Xu19e, ZZMS19]. area [Chi15, DD19, YYX^{+}19]. areas [BCT15]. argument [IT15, Mus18, PKB15, SS15b]. arguments [Can15, CÖ15, Gao17a, GST15a, JLX15a, LLW19, YSL19, ZS15b, ZYZW16]. arisen [AAMC19]. arises [Cov19, ID19]. arising [AM15a, AAA16, CM17a, CLP^{+}18, DBV18, DWTT17, HN14, HRS15, KM17a, LWZ19a, LZL19a, LBB18, MG15, MT15, NN19, ÖK19, SH15a, Sah19b, WLZ19a, Zha18e]. arithmetic [FTF16, GFWH17, GJP18, HHZ15, MMM18, RS16b, SG19a]. ARMA [lZA^{+}19]. Arnold [Ers16a]. Arnoldi [DMR15, LG19, YZ19]. arrangement [YSZ19]. array [IMTJ19, MSdC^{+}18]. Arrhenius [MP18]. arrival [Upa16]. arrivals [CSA18]. Arrow [YJK19]. arrowhead [SKK19]. Arsham [wG15a]. arterial [MSM15, PPS25, PP18a, RRR19, RRR20, SCZ^{+}17]. artery [BG18, JV18, MGN18, Pon17, SSG15b]. article [Ano19w]. Artificial [EM16, fGlHyL^{+}15, wG15a, JNGB18, KSP16, KPLC15, MHZC15, PAE^{+}17, ZCW^{+}19b]. artificial-free [wG15a]. ary [HM17]. Asad [PK18a]. ascent [Yan18a]. Asian [AM15d, CXL15, CDL15]. Aspects [Mah19, Alf17, BPR18, HP15c, KCK19, NBAGS19, Sch16, TQC15]. Aspiration [LSLL19, Pla15, Du19]. Aspiration-based [Pla15]. aspirations [LJR19]. assembly [AV15]. Assessing [Bhu15, SG19b, XSS^{+}17]. Assessment [Ery15, RDG18, Zei16, BBGR15, JV18, MTCV^{+}16, NKL15, PK15, PSP17, SG17]. asset [CW17]. assigning [JLL16]. assignment [LY16b, RKQ19, XY15b]. assimilation [GCM^{+}18]. associate [BR16a]. associated [Agr15, AIIK15, AHM16, CE15, Che15a, Che15b, Che16a, DDSL15, DSDL19, DD16b, GR15b, HHZ15, KKD15, LZ15a, LLJ19, MYH15, Mor17, Ras15, SZ18a, SGG15a, SKA16, USCA15, ZW15b]. association [AGC19, KPLC15, SA18b]. associative [RMB15]. assortativity [APPE17]. Asymmetric [Wan19b, AL18a, AK18a, EEST16, GG19a, HHL^{+}19, ILPE17, LW15c, LMW^{+}19, TXM15, WDLS18, ZH15a]. Asymmetrical [ZAZ19]. asymmetry [Ued19]. asymptotes [WJGM19]. Asymptotic [Aca15, AZ18, ACSS15, BHRW16, CE15, Dan15b, HL15b, LC15c, LPHH15, MH19a, MPPP19, MNR19, PYZS18, PBMP15, PD15, RN18, Reh17, TCS15, WL16b, WDCJ19, YC15, ZDK15, ABLZ15, AK15d, BMR17, BE17, CP15b, Che15a, Che16b, Che16a, CP17, CC17, EO15, FLMC15, FLS18, Kan19, KH17, LL15a, Lia15a, LP15b, LSM15, MDZ15, MHM15, MZM15, NY15, Prz15, SL15b, Tan15d, XHW15, Xu17a, XDH19, YPZ18, YCZ19, YB15, Zha17b, ZG15b, ZXC19]. Asymptotical [WCS16, ZSL15, ZS15b, OL15, YW19a]. Asymptotically [Mig16, MMZ19, DFP15, GN19a, GG19b, HPR16, JK15, SB16, SXS15, ZZZ15b, dSH15, Bor19]. Asymptotics [MMMMB17, ZWK16, CDK17b, FVMO15, PR15b, YWL19]. asynchronism [SFSR18]. Asynchronous [TFLZ19, XXW^{+}19, XSSW19, ZLDZ16, CCYT19, CYSY16, GHQ17, LTZ^{+}17, SSCX18, WXX18]. Atkinson [CZ18a]. atmosphere [CDP19]. atmospheric [DMSS18, FWLC15, WF15c]. atom [DDdF16, Dim17, GFWH17, Pal17, Pal19]. atom-bond [DDdF16, Dim17, GFWH17, Pal17, Pal19]. attachment [MB15b]. attacks [LXTF18]. Attenuation [ANU16, KKG18, Wan19a]. attitude [CST^{+}15, HWY15b]. Attiya [SGG15b]. attracting [ZZ16b]. attraction [ADMV19a, CN15c, CN15d, NCS16, SA16d, SLLC19]. attractive [CVA18, IJ15, LZCL18, YZH^{+}19, YL19b]. attractivity [HQ15, MS16a]. attractor [DK15a, Fan18, YLL18]. Attractors [Kan16, MX17, SB15, SSPCPG^{+}16, WL19a, ZZ16b, ZPT^{+}15, ZTW16]. attribute [MS15a]. attributes [CHL17]. augmentative [GGM15]. Augmented [KBK15, SGDX18, CWLA18, GHW19, LM19, LJHW19, Pal17, Pal19, WH18, WLZ19b]. August [DN16]. autapse [GJX^{+}19, Uzu17]. authentication [CZ15b]. authors [RSGFV16]. autoclaved [KKM15]. Autoencoders [GS19]. automata [Ort15a, Sch16, SZO16, RLH15]. Automatic [ABMBR15, OFwCW19, KC19]. automaton [CM17b, HZCY17, SdR19]. automorphic [APL15, CZP15]. Automorphism [HH17, GS16, GS18a, LLL19b]. automorphisms [FHK^{+}19, LLL17a]. autonomous [DS16a, GRV15, HGG16, HJ15, HS15a, Kan16, LC15c, Liu15d, LR15, WL18a, WD19, WL19a, YLL18, ZSS15, ZT16a, Zha15g, ZZ16b, ZTW16]. auxiliary [MS15a, SYZP16, ZA16]. Availability [Xu15a, HSKHA15]. available [WLZ15]. Average [Jur17, LZLW17, AY16b, Bog16, Du17b, ES15, Gao18, LW17a, NT19, PS16, ZLLZ16]. averaged [FFH^{+}15, KMB16, RXQ19]. averaging [Fio16, GLL15a, GNW15, LZS15]. averaging-hyperbolic [GLL15a]. aversion [GLL15a, SBLZ17, WRZ19]. AVF [JSLW17]. avian [SZJ19, XSS^{+}17]. aware [BBF^{+}18, QZJY15]. Awareness [GRS^{+}15, WGSX19]. axi [BBGR15, LXL17]. axi-symmetric [BBGR15, LXL17]. axial [BKT16, KG17]. axis [ÁL15a, Han16]. axis-symmetric [ÁL15a]. Axisymmetric [ERFE16, ZLL16, FF18, Muk16, Zha17c]. axonal [CM16].

B [RG19, SR15d, Bej17, BLM^{+}16, BJM16, DBK15, ET15, GR19a, HH19, HMT15a, HQJ^{+}15, IAW18, KZ16, KIZ18, KD16, LX15, LQZ^{+}18, MD17, SR16b, SuSR16, TSJ16, Tho17, UKE15, XzXqT^{+}15, YAIN17, ZZG18]. B-basis [SR15d]. B-spline [RG19, GR19a, DBK15, ET15, HH19, HMT15a, IAW18, KZ16, KD16, LX15, LQZ^{+}18, MD17, SR16b, TSJ16, Tho17, UKE15, XzXqT^{+}15, YAIN17, ZZG18]. B-splines [Bej17, BLM^{+}16, BJM16, SuSR16]. background [HD16b, LZX15]. Bäcklund [LXZ19c, WTSQ15]. backordering [MMB15]. Backstepping [MTB18]. backtracking [LPW15, SH16]. Backward [DJ15, GR15a, LRH15, OLZ19, SSGG17, Su19, WY19, WL19a, WZ18b]. backwards [TNPL15]. bacteria [DY17, Vid18, WLG17]. Bacterial [AADLF15, BS15a]. bacteriocin [DY17]. bacteriophage [Vid18]. baffled [CKKF16]. Bagley [KCK16, ZV17]. Balaban [KKST16, KKST17]. balance [AZAM15, BST18, CB19, DD19, LMF17, RMM19]. Balanced [DHQ19, DQH18, DB16, GWK16, JYF19, KLZ19, LCQ15, LX17, LJD17, LZ19c, MO16, QLLY18, RCTZ18, RN18, Tha18a, WLQ^{+}19, YFLZ18, hY18]. Balancedness [Par18b]. balking [YLW16]. Ball [AG15c, AG15d, HVR16, How15a, Mor17]. Baltic [ZCW^{+}19b]. BAM [CLPY18, GZLL17, LJ15a, MRC^{+}18, MPS15, SAJ18, SAM^{+}16, TXG15, ZWBC18]. BAM-type [MRC^{+}18]. Bampton [KSL19]. Banach [ALP15, AG15c, AM15g, AK15c, CLL16, CB17, CEHVT15, JK15, KT16, LL15c, MSHG16, Nak15, RMSD16, SSA19, WIF15]. band [Dan18, GLA^{+}17]. band-limited [Dan18]. band-parameter [Dan18]. banded [RM15]. bandwidth [XS19a]. bang [ASS16]. bang-bang [ASS16]. bankers [EG17]. Banzhaf [AMCMFJ15]. Bargaining [ZLC^{+}15]. bargraphs [CMS19]. Barnes [DKKM15a, Kim15c]. Barnes-type [DKKM15a]. barrier [Alz15]. barriers [RDM19]. barycentric [LHZM19, MZM16]. Barzilai [WSdSY15]. base [LW18a]. baseball [ALMCDH^{+}18].

Based [CDK17a, Acu15, AK16a, AIK15, AFK15, ATV16, ALMCDH^{+}18, AD15, APMM18, AB18, ARK15, AADMM16, BM19a, BMDS17, BPC15, BM15b, BS17, BCNN19, BB18b, BKW18, BKZA18, BRV18, BPZB18, CFR18, CCC^{+}17, CD15a, CL19b, CLHC15, CC18a, CDS15, CCX15, CKW15, CS16b, CXL16a, CLPY18, CYGL18, CL18, CH18c, CLSS19, CL19c, CZM^{+}19, CN15b, CNK^{+}18, CM17b, DCW15, DHX^{+}19, DLBQ19, DGNA15, DDD19, DA17, DLJT16, DSM15, DESS15, DHY^{+}18, DDM16, DCA15, DS16b, DKL15, DLH15, Du17a, DQW^{+}19, DN15c, Els17, EGH^{+}19, EFL18, FXC17, FDPP18, FPKF15, FTF16, FGPR18, FFL16, GI16, GFE19, GWSC16, GMOGCCC15, GOH18, GDZ18, GPRCTRY19, GNR^{+}19, GM17b, GG18, GL15a, GIS18, GZLL17, GFD16, HAC18, Han16, HYL19a, HCS19, How15b, How16, HGM15a, HWBL19, HXW^{+}19a, HZYW15, HI15a, HM15b, HYW^{+}15, HM15c, HZCY17, HLFZ19, ID15]. based [II17, Ili19a, IAK15, JM15, JJM18, JGX15, JZ19b, KKT15, KA15a, KS18, KCSD^{+}18, KGS^{+}15, KAG15, KSM^{+}19, Khl18, Kia15, KSL19, KML19, KE18, KEE17, KSS19, Lah18, LLM^{+}18, LY15a, LYC15, LPJ18, LZL15a, LS17a, LW18a, LZZL15, LL17a, Li17c, LZjy18, LY18c, LSZ^{+}18, LQZ^{+}18, LLz18, LWSX18a, LL19b, LSX19, LGCS19, LHLL19, LPS19, LZM^{+}18, LXC^{+}19, Lia15b, LF15, LLG^{+}15, LZY16, Liu17a, LCL17b, LXTF18, LD18b, LGL19c, LWW15c, LSM15, LJC16, LT18b, LCLZ19, Luo19, MZP15, MWLL17, MLSX18, MGL19, MGPL15, MBHS18, MRC^{+}18, MAR15, MS17b, MRE18, MRS18, MFF15, MRF18, MND^{+}15, Mil16, MS19c, MTB18, MKS19, MSL15, MPLS19, Naa15, NMNTBXVD15, NPT19, ND15, O’R18b, Oli17, bOqGyK^{+}15, PAE^{+}17, PJ18, PY19a, PY19b, PC16, Pla15, PS18b, PSP17, PP18b, QXW19, RF18, Rad17, RXZW17, RCTZ18, RR18b, RVE19, Rez15, RLGGAV15, RRS^{+}18]. based [SS18a, SM15a, SAA19, SSKA16, SDAB17, SST15, SD17, SSN15, SBG^{+}17, SHLL18, SCC17, SLLZ19, SIA15, SRAC16, SZRC15, SDWG19, TSJ16, TPL16, TWF^{+}18, TSZ19, TZ19a, TIT15, TTD18, TLZ^{+}19, Tom12, Tom15, Ton18, VK15, VK17a, VDI17, VS17, VK17b, VB15, WTD^{+}19, WYL15, WZL15, WDZ17, WHTW17, WLH^{+}18, WLWA18, WNF19, WL19b, WN16, WXM^{+}18, WWC^{+}18, WL17, WZLP18, WZK15, WF15c, XL15, XZB^{+}18, XY15b, XTXF18, XGQW19, XZLL17, YZD15, YZ15a, YZ19, YDCS15, YLP19, YZH^{+}19, YF17, YFLF17, YZ17b, YYP18, YLH^{+}19, Yüz16a, ZSA^{+}15, ZLW19, ZWF15, ZZL16a, ZYC18, ZZ18c, ZLZ^{+}19, ZCW^{+}19b, Zha15i, ZY15c, ZJYD17, ZHS18, ZHW^{+}18, ZVL19, ZS19b, ZPM16, lZA^{+}19, Gol15, LSAR15, NFS^{+}15, PS15c, ZW15d]. based-control [MTB18]. bases [Bag19, EGB15, JCW^{+}15, KM19, SK17a]. Bashforth [ZJJ18]. basic [CM16, YG15a, ZZ14, ZZ15c]. Basics [Pöt19]. Basin [FFGRSV19, SB15]. Basins [CN15c, NCS16, CN15d, SA16d]. basis [AD17, BMC19, BH17, Boy15, DA17, GN15, HAC18, Han15a, HH18a, KKT15, LLO16, LL16a, LLC19, MZP15, MRQZ19, NAM18, RPB15, SSK^{+}19, SR15d, SDGU16, Seg15, SS15a, SN18c, Su19, YDCS15, ZMBJ19, ZHS18]. Baskakov [AGN15, GO15a, GAS17, MG17, MS16b, MMS17b]. Basque [CST^{+}15]. bat [JLX15b]. batch [CLC^{+}19, CSA18, GYY18, Upa16, ZXDP18]. battery [BdS19]. Baxter [CCY19b, KC18b]. Baxter-like [CCY19b, KC18b]. Bayes [KKKD18]. Bayesian [DS16a, Fau17, HYJ19, LLC^{+}15]. bays [GHN^{+}16]. BBM [Li17b, YWS18]. BDF [ZRLX19]. be [ZM18a]. beach [DQW^{+}19]. Beale [WL16c]. beam [ALP18, ANGC17, BM18a, Che15f, HW18, KG17, LJL15a, WR15, WYL15, Wan16b]. beam-plasma [Wan16b]. beams [AJC^{+}18, AM15b, NND^{+}18, RM16]. beating [Pon18]. bed [MNYL18]. bee [fGlHyL^{+}15, MHZC15]. before [AELH15]. Behavior [HZCY17, ABL15, Amm15, ABGL17, Can15, DGZ15, GQC15, Gra15, Kan15, KMR18, LZL17, LAAZH19, LC15c, LJS18, MMS18, OL15, RG15, SA16b, SuSB15, SSZL17, TS15, TCS15, TD19, Tre15, WLJ15, WJNW15, YPZ18, ZDK15, ZMA^{+}18, ZWY19, ZMT15, ZLZW15, ZYZ^{+}16a, ZYZ16b]. Behavior-based [HZCY17]. Behavioral [HZYW15, Val18]. behaviors [CGW19, HL15b, WJNW15, YYX^{+}19, ZYC18, ZYZW16]. behaviour [Ant18, FLL17, GOH18, MDZ15]. behaviours [Moh19, MPHG15]. behind [GKN16]. Behnken [KVS15]. belief [DLD16, HCY^{+}18, HYJ19]. bell [HM17, MC15b, NR15, QZ15a, SGM^{+}16, TZTG17]. Bell-polynomial [SGM^{+}16]. bell-shaped [HM17]. Bellman [ZAP19, Fen15, Ied15, LW19c]. below [QBNF19]. Beltrami [Ers16a, MSI17]. BEM [ADSS16, JLL16, TPS15, Tch15, ZGZ16]. Bénard [MJ19, Zha16c, ZFL18]. benchmark [KKKD18]. benchmarking [CLL^{+}15]. Bendersky [You15]. Bending [SLW16b, JDL15, NND^{+}18]. benefit [WRZ19]. benefits [RXLM19]. Benjamin [Ghe16, HAC18, Kan16, ZP16]. benzenoid [LY15b, LY16a, LBRS18, ZZ16a]. benzenoids [ZZ16a]. Bergman [jJ15b, jJ15c]. Bermudan [JO15]. Bernoulli [AJC^{+}18, BN15, HASA15, HK18a, KOR19, Kim15c, KKD15, KG17, LJL15a, Luo15, LTZ^{+}17, Upa16, ZL18]. Bernstein [MAK15a, MAK16, ABQ15, AIK15, AFK15, AE17a, BM15b, BTIIE16, BA17, CZ16, CW18, DP15b, Gla19, HA19, IH15, KIAG16, KSS19, MRE18, MP19b, MS19c, MAK15b, MAK15c, PHM19, Sha15d, SIA15, TG17, TG19, WXW18, Yüz16a, ZX15, ZH15b].

Berrut [Zha17b]. Bertrand [AEP15].

Besov [Zha15h]. Bessel

[DSDL19, DDSL15, FMX15, KVZ16, KNT17, NR15, Yüz15, Yüz16b]. Best [AHG15, BN18, CN15d, IZ16, QM15a]. Beta [MG17, Sim15, GST19, SQL16]. Beta-type [Sim15]. Bethe [YwFW^{+}19]. better [DGL^{+}15]. between [AG15c, BKL^{+}17, CXL16b, Dan15b, DY17, DGNA15, DD16a, DY15, ES15, FHLL19, HKS19, JY19, Kim15a, KST15, KFHH15, LMZZ19, MMZ^{+}17, MCPO15, RXLM19, RSGFV16, SSN15, TLLD19, TSPF18, WF15a, WW16c, XGG18, YY15, ZQ18, ZWM16]. betweenness [ATV19]. Beurling [MR19, TQC15]. beyond [ZS17]. Bézier [AIK15, CNK^{+}18, DP15b, GLW17, KSS19, SA19, WGL15, Zha15c, ZH15b]. Bézier-summation-integral [AIK15]. BFGS [FFGRSV19]. BGS [HM15d]. BGS-Uzawa [HM15d]. Bhaskar [RZD^{+}19]. Bhatia [CN15d]. Bhatnagar [Kri18]. bi [BCST19, Chi15, CT18, DÇO15, KLZT15, MM18a, WG15b]. bi-convex [DÇO15]. bi-Hamiltonian [MM18a, WG15b]. bi-level [Chi15, KLZT15]. bi-parametric [BCST19]. bi-periodic [CT18]. bi-starlike [DÇO15]. Bicentric [LY19]. BiCR [AF18]. Bicubic [HH19, Arà16]. bicycle [TRRP15]. Bicyclic [HFK19, LWZ15, MR16b, QY15, FWZ19, FKA17, FT18, JFT17, JLW18, MRS19, Tep19, YW18a]. BID [ARB16]. bidiagonal [GGZR16]. bidirectional [RMB15]. BIE [MS15d]. bifunction [FPUK15]. bifurcated [PP18a]. bifurcating [GLN15, LLdM16]. Bifurcation [GYC19, Li15a, LX19b, PZ15, WZY^{+}19, ZZZ15a, ZZC^{+}19b, AYAA^{+}15, AVCCGAVDL15, CGC19, CL16c, EYE18, ERFE16, GSV18b, HCX^{+}17a, HLC19, JS15, LWL^{+}18, MHX17, PW19a, SSGG17, TXG15, WPL15, WL18a, WXLL19, WLZ15, WF15b, WLL16, XLL^{+}19, YL19a, YLL15, ZT16a, ZXW16, ZHW16, ZZH15]. Bifurcations [Cao19, HCX^{+}17b, LJK15, ZH15a, DU18, Fer15, HC18, Mon19, NSM16, TS15, XJZ^{+}15, XWZ15, ZXZB19]. big [HKW^{+}17]. Bihari [ZAP19]. Biharmonic [HH19, DH15, DLCT18, YBZ18]. Bijections [Yan18a]. bilateral [BM18b]. bilaterally [SPW15]. bilevel [CRCVMS18, MND^{+}15]. Bilinear [HGH19b, LZ17f, TZTG17]. Bilinearisation [DLjZ18]. Bilinearisation-reduction [DLjZ18]. bilinearization [VK19]. binarization [GHZ19, JM15]. Binary [SuSR15, TZTG17, DJLX17, HLT15, LWW19, RS15b, SuSR16, SPGP18, TST16, YCD17, YCLS19]. Binet [BR16a]. Bingham [FF18]. binomial [JDS18]. binormal [JKK15]. bio [CDFP17, ZDZ15]. bio-mathematical [ZDZ15]. bio-thermal [CDFP17]. biodiversity [Par18b]. bioeconomic [LYZL18, LXZL19]. Biogeography [Gol15]. Biogeography-Based [Gol15]. bioheat [GSP15]. biological [GGM15, Kun16, OB18, Świ15, ZZH15]. biology [EFPT18]. biomechanical [Pon17, SSG15b]. biophysical [UPMV18]. bioprocesses [WLZW15]. biorthogonal [BMR17]. bipancyclicity [LX17]. biparametric [GKM16]. Bipartite [Ili18, AVC^{+}19, BGMO18, CRC19, CL15a, CWD16, DMG16, DIS17, JYF19, Jin17, LLS15, LZ18e, LGL19a, PTO19, ZLJ15]. bipartiteness [LP16a]. bipolar [KL16, LL19c]. bipolarity [LLT19]. biquadratic [TW15a]. Birkhoff [ACD18, DDH18]. Birth [PRA^{+}19, DGN16, JLL15a, QLF15, Ter15, WCH19]. birth-death [DGN16]. Bistability [WRW16]. bisymmetric [HGM15b]. Bivariate [BTIIE16, BP17b, Erb16, ÖK19, BA17, CGPR15, GKN15b, MMPP18, MRR17, OR18a, TW15a, ZX15]. BJ [HM15d]. BJ-Uzawa [HM15d]. BKP [VK19, DT16]. Black [HH18a, RM18a]. blade [Ili19b]. blade-vortex [Ili19b]. Blasius [MS17a]. blasting [Hac17]. Blaze [GWPA18]. Blaze-DEM [GWPA18]. blind [CPP15]. Bloch [bBjJ16, jJ15b, Li16a, SzShG16, Zhu16]. Bloch-type [Zhu16]. Block [CL19b, CZL18, KZR17, MD18b, Sab18, YZ19, BM15a, BPR16, BFJ^{+}17, CFdCdS18, CKDC16, Den15, FZL19, Guo16, LW15c, LWK15, LR16b, LR17, LR19b, LJ15b, MS17a, MH15a, MAJR19, NE16, RBB19, RS17a, RP18, RR18a, SGKR19, SKK19, TZC19, Vis15, WLLK17, WLZ19a, YLL^{+}16, ZWF15, ZZL15b, MM04, SR15a]. Block-centered [KZR17, LR16b, LR17, LR19b, YLL^{+}16, ZWF15]. block-pulse [MH15a, SR15a]. blockage [Uzu17]. blocked [WM18b]. blocks [RSC16]. Blood [ARACCA^{+}18, AZL15, AAM19, BKKN16, GCM^{+}18, JV18, STVCC17]. Blow [BB15c, ES17a, FG19, LG16, MJ19, Mi17, ZJZ15, FAHZ16, HK15, LTL16, PS18b, YX19b, Zho15a, ZYZW16, ZLL18]. Blow-up [BB15c, ES17a, FG19, LG16, MJ19, Mi17, ZJZ15, FAHZ16, LTL16, PS18b, YX19b, Zho15a, ZYZW16]. blowflies [Tan15d]. Blowup [ZY15d]. bluff [KMB16]. blur [SHLW16]. BMAP [KKD17]. BMAP/PH/ [KKD17]. Board [Ano18o, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18p, Ano18q, Ano18r, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v]. boarding [YYX^{+}19]. bodies [MCO19]. body [KMB16, Kri19b, PPS25, Pon17, PP18a, RRR19, RRR20, SNS^{+}16, SSG15b, SPGP18]. Bogdanov [LWZ^{+}19b]. boiler [MNYL18]. boiling [JOÖ16]. Boiti [FMC15]. Boltzmann [Bor19, AHLR19, CCL^{+}15, DSSC19, GYZC15, KRF17, Kri18, Kri19b, LCS15, LSW16, MM15a, MHR18, NB15, RSC16, SK15, Sun18, Wan16b, WSLC17, WYD^{+}18]. Bona [Ghe16, HAC18, Kan16, ZP16]. Bond [ARB16, DDdF16, Dim17, GFWH17, HS15c, Pal17, Pal19, TZ18d]. bondage [HLL18]. bonded [ABN^{+}17]. bonds [LABK15, SPP17]. bones [MPR^{+}17]. Book [ZCML19]. boolean [LXD19, CWL16, DLY^{+}17a, GhDwZS17, LLL^{+}18b, TLL^{+}18, YMeFG19, ZLK^{+}19]. Boosting [SID19, VK17b]. boosts [Tan17a]. bordered [She18]. boreholes [WE19]. Borel [WXZ18]. Born [AZL15]. borne [Vil16, ZLJ15]. borrow [MZC19]. Borwein [WSdSY15]. Bose [ID19, QFS17, WJWX18]. both [HL18, LXT17, MRC^{+}18, MBM16, TZD^{+}19, WDLZ19, ZSZ17]. bottom [PMC16]. bouncing [How15a]. Bound [WW16b, CYY19, GW15c, GN19a, HSKHA15, JLW18, KL16, LY18d, WZ16a, WL18a, ZT16a, Zha19a, ZZ15d]. bound-and-reduce [GW15c]. boundaries [BE19, CK18, LLL18d]. Boundary [ABI19, BRD15, BÁR15, HAMA14, PK18a, YÖ18, AKS15, ABHR15, AN15, ANT^{+}16, AL18b, ASASM16, AAKT15, Ali15, AV18, AD17, AJP18, ANEP19, BE19, BMM19, BTH18, BST18, CM18, CDP19, CXL18, Cho19a, CLP^{+}18, Cui15, CZ15c, DMR15, DX17, DO18, EKS^{+}15, EM16, EMO18, EAY17, EM15, FFT15, GW16, Gao17c, GQC15, GR19a, GYZC15, HLIK19, HASC19, HL17, HXM18, Hua15, JSM16, JX19, KMBA18, KAI19, KD18, KPT16, KMZ19, KS19b, LLY15, LWT^{+}15, LJL15a, LZ17b, LCL17a, LLX18a, LBB18, LWS18, LLL18d, LDL15, MG15, MHX17, MC17, MKRS16, MGN18, MTB18, MS15d, MTJ15, NKS18, NE15, Pan15, PZ15, PAK15, PCS17, Qiu18, RRRR16, RRV15, RG19, RGA19, RMK19, SH15a, Sah19b, SM19, SR17, Sha12, SN18a, SN18b, Slo15, SDWG19, SH18b, Tha15, Ugu19, VK15, VK17a, WZ15b, WZM15]. boundary [WZZ18, WA18, WZF^{+}19, Won15, WSYL18, XP17, jXlZX18, XLW16, YBZ18, YU15, YWH15, ZAW15, Zha18e, ZjA19, ZG15b, ZY15d, ZZL16b, ZT16c]. boundary-layer [Zha18e]. boundary-value [Zha18e]. bounded [CYW18, Cho19b, DLXZ19, Els17, Gao17b, HR16, HMW16, Ili18, JZ19b, LB15a, LLLZ16, PL15c, PAK15, Ued19, WTZL18]. boundedly [CLC^{+}19]. Boundedness [BB16, NS15a, Cho19b, DLXZ19, LLLZ16, PYZS18, QDD15, XDH19, ZSS15, ZLZ16]. Bounding [LGS^{+}18, TNPL15]. Bounds [BGMO18, CDS15, LQG18, LG18, Pal17, Pal19, PPV15, TF18, ABQ15, AK16b, And18, ACP19, BCK15, CHL18, FK17, GLW18, Jah17, KwC15, LH17a, LY15c, LMZW16, Liu15g, LLX18b, LB19, Ma18, MZM17, MMG16, OP19, Sku15, SX15, WDB18, XL18a, YCZ15, YCZ19, Zha15c]. Boussinesq [KI15, eMBHA15, CCY19a, DF16, HMW16, LSW16, LFY18, SEKA16, UKE15, VK19, YMHW15, ZZC15, ZJJ18, dFGAN19]. Boussinesq-type [DF16]. box [CLHC15, CRZ19, FA15, SGFCRM^{+}18, HLLT15, KVS15, OL19]. box-constrained [CRZ19]. boxes [HP15b]. Bracketing [LJZ18, Raz15]. Brain [Tor16, AD15, FFL16]. braking [BKT16]. branch [AF15, CZ17a, HSKHA15]. branch-and-bound [HSKHA15]. branch-and-cut [AF15]. branched [SG19a]. Bratu [eMAGG19, ZI18]. Brazil [DTJ18]. breakage [LMF17]. breakdown [RLGGAV15]. breakdowns [KKD17]. breaking [HCH16, MMIA19]. breakpoint [GI16]. Bregman [BPR18, CCC^{+}17, LPZ19, SPW15]. Brenke [KM18a]. Brennan [MBHS18]. bridge [Kan15, MX17]. Bridged [Sta17]. bridges [ANGC17, ABB^{+}18, Hac17]. brief [dF15]. Bright [WTSQ15]. Brinkman [NM19, TA16]. broadcasting [GSV18a]. Brownian [RWY19, AwEE15, DWS19, LKZ18, LW19d, TB17a, WY19, XL18b, XL19, ZY15a]. BSCCB [JT16]. BSDEs [DWS19]. Bt [LTC16]. Bt-corn [LTC16]. Bt-crop [LTC16]. bubble [LTY16, LZX19, ZW19a]. bubble-sort [LTY16, ZW19a]. bubbles [ERFE16, GR17b]. Buckling [JA17, LZL15a, RM19b, AB18]. budget [DKN16, WF15c]. buffer [CSA18]. buffered [JMK17]. building [MNYL18, PK15, SKCS15]. bump [MPW15]. bundle [LPW15]. bundles [AB18]. Bundling [JO15]. Buongiorno [GJR^{+}15, SP15, SRP17]. buoyancy [Kim15a]. buoyancy-driven [Kim15a]. Burger [HEA15b]. Burgers’ [DWTT17, FLL^{+}19, GN19b, GfSmL16, LCS15, MA15c, NAM18, TSJ16, YGZ19, ZLLA19, ZCU15, ÇT16, EM16, HAC18, MDH15, PKS15, STL17, ZP16, ZZ15a]. buried [QBNF19]. Burning [LZH19]. Burnside [Che16b, LSM15]. Burst [CSA18]. Butterfly [NRY16]. BVP [Che19, IH15, KS16b]. BVPs [LWW15b, LJ19]. by-pass [TBSR16]. bypass [SCZ^{+}17]. byproduct [VVD15]. Byrnev [Pro16].

C [DSDL19, ARK15, WM18b]. Cable [LL19d, STK15, WYL15]. cable-stayed [WYL15]. cacti [Tu17, Wan17b, ZL16b]. cactoid [ZDJ19]. CaF [KSKW^{+}19]. Cahn [CJK17, LL19a, LG17, VD15]. Cairo [LY18b]. calcium [DZ17b, DZ17a]. calculate [FMSG15]. Calculating [SNS^{+}16, BM19d]. Calculation [Yor15, CL16c, FG18, Kan19, MWH^{+}18]. Calculations [PP18b, CNK^{+}18]. calculus [BCD^{+}19, BBD15, FÖB19, GT19, KOR19, MD18c, Tar15, TT18, VV18]. calibrated [MMB15]. Calibration [DTJ18]. call [LD17, LLZD18a, MZC19, ZLC^{+}18]. Calls [EG17]. cam [YLH^{+}19]. Camassa [HGL19, YX19b]. Cambodia [PG15a]. campaign [KIZ18]. Campanato [GR17a]. Can [BKAAM15, CCW19, WWC^{+}15, ZSW^{+}17]. cancer [KG15, KKS^{+}16]. candidates [Cim18]. Canonical [Li16b, AADMM16, GL15b, JG17]. cantilever [AM15b]. cap [ZWY19]. cap-and-trade [ZWY19]. capabilities [APMM18, GGAVGG^{+}18]. capability [WLX18, WLZ15, YZD15]. capacitated [CRCVMS18]. capacities [ZXDP18]. capacitive [KS16b, LMZZ19]. capacity [Vil16]. capital [KLZ19, LKZ18, RP19, SSM^{+}18, WWZ19]. Capturing [VG18, Gla19, IKHS16, ZM19]. Caputo [AHO18, AN15, AJP18, CGCK19, DB18, DSSC19, LXZ19a, WR19, YJW16, ZBBW17, ZZ19c]. Caputo-type [LXZ19a, WR19]. car [LZY16, RLGGAV15]. car-following [LZY16]. car-truck [LZY16]. Carathéodory [ALL18a]. Carbon [AB15b, Kia15, ZLL15b]. carbonation [Zha16b]. carbuncle [Kem18]. carcinoma [HWBL19]. cardiac [CK17, Cha19]. cardinal [Bej17]. care [DKN16]. Carlo [Ace19a, BG17, Har15, LGO17, LWZ16, UHLM15]. Carlson [DAS17]. Carlson-type [DAS17]. carotid [JV18]. Carreau [PK18a, HAMA14]. carrier [RSO^{+}18]. Cartesian [AK15b, Bac15, BDH16, FTF16, FM18, GLW18, GIK19, GSV15, LZ18e, LWM18]. CAS [GGAVGG^{+}18, Sae17]. cascade [CTT^{+}19, FFH^{+}18, Neu16, SJL16]. Cascading [LWSX18b]. case [Boy17, CD15d, CST^{+}15, Erb16, HK18b, HZK^{+}19, KDS16, KM18b, LP15a, NFS^{+}15, OLZ19, SG19b, WC15b, YG15a]. cases [ZMH18]. cash [SCB15]. Cassinian [BM19c]. Casson [EAY17, PP18a, WF15a]. Casteljau [Her19]. casting [BKLW15, FVMO15]. catacondensed [CXL16b]. catalytic [SB16]. catastrophe [SPP17]. categorical [CHL17]. caterpillars [LMS19]. Cattaneo [CP15d, LJL19, LCL17b]. Cauchy [AZL15, AHS18, BMM19, BCNN19, CLQ16, DHW^{+}19, DKKM15b, HPR16, Li19, MH19a, Par15, PTBN19, Rem18, SHA15b, SSA18, XCHW16, ZSS15, ZZ15b, Zha18a]. causal [Dok18]. cause [BF16a, SSGG17]. cavity [DeC19, GSDH15, GG15c, MS18b, SK15, SV17a, SP15, SRP17, SH18a, Zha15a]. Cayley [HKM18, LL16b, ML17, TZS17]. CD [HJ18]. CD4 [MK15b]. CDO [WL16d]. cell [AZL15, AM15a, AY16b, CS15a, EFPT18, Kim15a, Kim18, LY17, LXT17, NMNTBXVD15, PRMC19, WGLZ16, WJHA17, YL19a]. cell-average [AY16b]. cell-based [NMNTBXVD15]. cell-growth [AM15a]. cell-mediated [WGLZ16]. cell-to-cell [LXT17, WGLZ16]. cells [EA15a, HJ18, LZ19a, MK15b, NKKS16, PBMP15, SRJ^{+}15, SG18]. Cellular [RLH15, CM17b, HZCY17, KBR15, Ort15a, Sch16, SdR19, SZO16, XGZ17]. Cement [VB15]. Cement-based [VB15]. censored [AAHT18]. Center [Val15, WFZ17, HLLT15, LSZ15, MF17, MWY15, QCH19, YY17, ZS14]. center-Hölder [LSZ15]. centered [CR18, KZR17, LR16b, LR17, LR19b, YLL^{+}16, ZWF15]. centers [Gar16a, LLdM16, ZH19a]. Central [HCS19, TZH15, CCLPC19, EG17, LJD17, LY19, LJC16, WYLG19]. centrality [ATV16, ATV19, TLLD19, WLX18, XGG18]. Centralizer [ZCPM17]. centrifugal [BKLW15]. centrifuge [BKT16]. centroids [LLM^{+}18]. ceramic [NND^{+}18]. ceramic-metal [NND^{+}18]. ceramics [MWLL17]. Certain [ACK15, AED15, Bou18, GFWH17, GR15c, KR15, MWWM15, NZZ16, NS15a, Pac19, Sha15d, dF15]. certificateless [HKW^{+}17]. CFD [IR16, Ili19b, Ili19a]. CFD/CSD [Ili19b]. chain [ASZD18, CLT^{+}15, GJX^{+}19, MMS17a, PMSB15, WC15b]. chains [ARB16, CP15c, FXXC19, HHZ15, HLL16, LQZ16, LW18e, Sku15, WC17, Xu15b, YLWF15, YLWF16]. Challenges [FLS^{+}15, FPKF15]. change [ANU16, BKAAM15, FS16, GAT18b, HHP18, KCL19, KPP18, YWL19]. changes [FLL17, PG15a]. Changhee [JRSK16]. changing [CCC19, ZMC^{+}19b]. channel [AB15b, Alk19, CSZ15, CS18, CL19c, CM16, FTE19, HMHA15, LL18c, MSI17, PMSB15, SBG^{+}17, SS18c, SB16, SSM17, SH18b, Uzu17]. channels [BMT18, BKZA18, LLZD18b, LLZ^{+}19, SBV15, WLQ^{+}19]. Chaos [GNR^{+}19, TTZ^{+}19, CL18, DG15a, FPKF15, Li15a, LF15, NR18a, Pet15a, Vol15, YZJ18, ZZC^{+}19b, SGFCRM^{+}18]. chaos-based [FPKF15]. Chaotic [NR18b, Pai15a, CPCQ17, GCAA18, HGG16, HXW^{+}19a, JAN16, KBR15, LLCR15, LLC18, LZY^{+}17, LZKX19, MLSX18, Naa15, PSJJ19, SLZ^{+}16, SWZ^{+}19, SSPCPG^{+}16, SGY^{+}15, TIT15, WL18a, WSH^{+}18, WDG19a, WD19, YWS^{+}19, YJYH15, ZT16a, ZZLZ17, ZHWW17, ZPT^{+}15]. Chaplygin [SR15f, SSB15b, SAS19]. Chaplygin-type [SAS19]. character [TYD15]. characterisation [LLL18a]. Characteristic [LR17, ZZZ16, ASZD18, CC19, HN14, LR19c, LZ15d, NN19, TZ18d, Tur18, YCLS19, ZL16a]. characteristics [CGW16, LW19d, ZSK^{+}19]. characteristics-mixed [CGW16]. Characterization [AVC^{+}19, BCFR18, FT18, FLCSV18]. Characterizations [DYW15, JKK15, MS19a, Sha15f]. characterize [WLX18]. Characterizing [GTBC18, BCO15, Hsi15]. charge [BS19a, FW18b]. charged [Dan15b, LW19a]. charged-particle [LW19a]. Charlier [KA15a]. Chebyshev [AM15c, AKKB17, BAB18, BCM17, BdS19, Boy17, CCTV16, GDF^{+}17, HEA16, KBM16, LB15a, Nag17, NBK16, PPV15, SuRI15, Sak19, Sha15e, SPG17, Suk19, Tan15b, WK16a, WK16b, XYG^{+}18, ZHX^{+}17, ZX16]. Chebyshev-type [SPG17]. checkpoint [YJ19]. Chelyshkov [OS15]. Chemical [MNJM19, ADA18, AG17, Ban18, Ber17, CT17b, FSW16, FWLC15, FCW^{+}17, GJX^{+}19, GA18, LG19, Sah19b, SMG^{+}18, VS17, WZ19]. chemoimmunotherapy [RMCG19]. chemoprophylaxis [Bhu15]. chemostat [SSZL17, ZY18b]. chemotaxis [ZZZ16, ZZ17a]. chemotherapy [RLM19]. Chen [LK15, LZY^{+}17]. Chevalley [BCSG18]. CHF [Mah15]. chicken [Tan17a]. Chiellini [MR15a]. Childress [Ers16a]. China [PZLZ15, WXP16, XZG19, XSS^{+}17]. Chipot [HK15]. Chlodowsky [BA17]. Chlodowsky-type [BA17]. chloride [NKKK18]. choice [Bot15, BFJ^{+}17, CYY19, LXPD15, PABA15, Red19]. cholera [CMW17, TXL19]. Cholesky [LY15c, SFLM19]. choosability [CR19b]. Choquet [LCOS19]. chords [Xu19d]. chromatic [ACP19, WWW18, XLG18, YSY^{+}17, YW18b]. chromatography [BMRS18]. chronic [RMCG19]. Chua [FDPP18, LMZZ19, XZLL17]. Ciarlet [YBZ18]. cilia [Pon18, SSM17]. ciliary [FTE19]. cinlar [KÇ18a]. CIR [TZ18d]. circle [BCD15a, BMR17, CMR17, CMW15]. circuit [BB18b, JCW18, TVK^{+}16, TD18]. circuits [LMZZ19, Sta17]. circulant [CEG^{+}15, CZ19, JXW15, JH15, JZ15a, ZS15e]. circulants [LW18b]. Circular [CMS19, AYAA^{+}15, DDY15, MYH15, MWY15, Rad15, RM19a, RMK19, SV15a, SD19c]. circular-hole [MWY15]. circulation [BKT16, MNYL18, ZHWW15]. circumferentially [FFH^{+}15]. Ćirić [RZD^{+}19]. CISTs [QCH19]. city [GGAVGGRC15]. clamping [KK15b]. Clar [SZ19c, ZAY18]. Clarke [LZ15e, LLB16]. Class [LLL19c, AGZ15, ARB19, AAKT15, AM15h, AF15, AAMC19, BT15a, BE19, BAK19, Bej17, Bir15, BMR17, BMB15, CP19, CR15, CY16b, CCNQ19, Cas15, CTT^{+}19, CL16b, CXZZ18, CY15b, CJ17b, Dai16, DAQI15, DH15, DEK15, DZL15, DOZ18, tFyZZ15, FMX15, FMDZ19, FNS15, Flo18, FLYG18, GD15, GWWL17, GKN15a, Ghe16, GR19a, GYC19, HRS15, HS15a, How15b, How16, HM18b, HHY19, IB15, JCW18, JJ15a, JS16, KCK19, KMBA18, KBM16, KLS15, LLZ15, LWT^{+}15, LZ18c, Li19, LGK15, LHS18, LP15c, LWZZ18, Lv16, Lv18, MHN15, MFJZ15, MWRT17, MG19, Mil16, MMK17, MKS19, MK15b, MK15c, NB19a, NRT16, OS15, PAK15, QZ15b, Red19, RMA15, Rom19, RG19, RGA19, Sah19b, SY18a, SXS15, SGG15a, SGG15b, SJL16, SZ16, Tom12, Tom15, TBG16, VD16, VDI17, WqTjT15, Wan15e, WZAF15, WK16a, WZZ18, WDLZ19]. class [Won15, bWzZjH16, Wu17, WWHZ18, XZLZ17, XL15, XY15a, XLJC18, XzXqT^{+}15, XH19a, YWO18, YD18, YLCW18, YGZ19, YLNL15, Yu16, ZA16, ZSL15, Zha15d, ZZL16a, Zha18a, ZZ18a, ZjA19, ZWZ19a, ZZMS19, ZZL15a, Zha15i, ZM18b, ZLL18, dF15]. classes [BM16, BD15, FLT18, GLS16, JM18, LR16a, MH19a, MS19c, SHR18]. classic [Liu18c]. classical [AIIK15, BCFR18, LL17b, MMPP18, SD19a, VG18, ZY18b]. classification [CYZ17, CLR16b, CS16b, LYC15, LAVME18, Liu15a, LM15, WWZ15]. classified [LWSX18a]. classifier [VK17b]. Clausenian [SS15b]. cleaning [Kem16]. Clément [PVV^{+}15]. Clenshaw [LX19c, Ma15a]. Cleverly [CCW19]. clients [YSLS17]. Clifford [DRS16, HS17]. climate [BKAAM15, KCL19, MKMC15, YWL19]. clipping [dCM16, LCW19a]. close [ACC16, LH15, SKA16]. close-to-convex [LH15, SKA16]. Closed [LD19b, LXZ19b, XC15, CGJZ18, Nad15, SdR19]. Closed-form [LD19b]. Closed-loop [LXZ19b]. cloud [HKW^{+}17, JA17, KWH17]. Cluster [FYZ16]. Clusterability [NKL15]. Clustered [SRH15, CnTVsL18].

Clustered-gravitational [SRH15]. clustering [CHL17, GWC^{+}16, Suk19]. clusterings [ZCL18]. clusters [TDA17, ZL19c]. CMOS [TKDF15, TD18]. CNC [Han16]. CNLS [WXY15]. cnoidal [AM18a]. CNT [LZL15a, LZL17]. CNT-reinforced [LZL17]. Co [BMT19, LWZ18, TMM18, XZB^{+}18, MMK18, RC16, WW15, WLL16]. co-dimension [WLL16]. Co-dynamics [BMT19, TMM18]. Co-evolution [LWZ18]. co-infection [MMK18]. Co-Kriging [XZB^{+}18]. co-mutualism [WW15]. co-trees [RC16]. coalition [MCZ15]. coalitional [MCZ15]. coarsening [FGPR18]. coating [ABN^{+}17, WQ15]. coating/fiber [WQ15]. coaxial [FK18a]. cocirculating [EA15a]. code [DKM17, EEG^{+}16, GWPA18]. codebooks [GN19a, WNF19]. codes [CZ15b, WNF19, Wil18, ZK17]. codimension [Mon19, NSM16]. codimension-one [Mon19]. coding [YCD17]. Coefficient [Uça16, BdlRT18, ES17a, GY15, HGH19b, KI15, LP15a, PST18, SNL19, SM18, WTJ17, ZBKS16, ZEW19, Zho17]. coefficients [ACC16, BHRW16, CLR16b, CLS15, CLJ^{+}15, CNQRR18, Dai16, Dan18, DKBS15, DHS19, lDqCWzH18, GM19, HCY15, HLH17, KDS16, Kim15b, LL15d, LLY15, LQZ16, LWZ16, LW18e, LMMZ18, MH17c, NP18, OW18, Pol19, Prz15, RW17, RZMB17, RMA15, SR15b, SMI18, Ste19, TQ18, WR19, WL19a, XYG^{+}18, YSL19, YMMH15, YQ18, ZWF15, ZX16, ZLW18]. coercive [ZY15b]. Coevolution [BK18a, LGL^{+}19b, Tan17a, LSLL19, SLS16]. Coevolving [LJ17]. coexistence [VEF19, ZJ15a]. cognitive [APMM18, FXC17, HJP16, SM15a]. cographs [AVC^{+}19, CJ17a]. Cohen [EO15, LCAA17, PRC^{+}19, ZDPZ15]. coherence [Uzu17]. coherent [GGMPC15]. cohomology [BCSG18]. Coincidence [O’R18b, ARGF15, PN15b]. coindices [Gut17]. coined [WD19]. coinsurance [dCCGV15]. Coleman [AMCMFJ15]. collaboration [WWC^{+}18]. collapse [STVCC17]. collection [Hsi15]. collective [HCS19, KMR18, Lu16]. collective-risk [HCS19]. collinear [SPGP18]. Collision [GWK15, KSD^{+}15]. collisions [ID19]. Collocation [ER15, MH17a, SS18a, AMO18, AK18b, AD17, BAB18, BB15c, BLJ^{+}19, CCY19a, CXL18, CL16d, Ghe16, GIS15, GR19a, HEA15b, HEA16, KC19, KZ16, KPT16, KXZ19, LZ19d, LMMZ18, LFC^{+}19, LHWG16, MHN15, MZM16, Mic15b, MH16b, MH17c, Nag17, NAM18, OS15, Pan18, PJ15, Reu15, RME18, RG19, SR15b, SM19, SR16b, SK17a, SPS18, uIHA18, SRS17, Tan15b, YTH19, YAIN17, Yüz15, Yüz16a, Yüz18, ZZY^{+}17, ZX16]. collocations [VS19]. colony [DCA15, fGlHyL^{+}15, GWC^{+}16, MHZC15, SZRC15]. color [yWyLyY^{+}15]. colorable [ZLSX18]. colored [Che18b, CMM17, GST15b, WLL16, XYZ18]. coloring [CYZ17, CYY19, CJ17a, HWL^{+}19b, SY15b, WZ18a, WWW19, Yue16]. colorings [LZ18d, LT19, Liu15b]. colors [LLL19c]. Colpitts [RXLM19]. comaximal [SLF18]. Combination [KMA16, LLC18, Kim15b, RG18b, WYZ^{+}17]. combinations [PÖB15]. combinatorial [Bag15, BMRW19]. Combined [Tch15, AP19, KG15, MK15b, PKHD18, RS15b, RS16a, Rui18, YCH18, ZXZB19, ZCW^{+}19b, ZZ17b]. combining [MHZC15, VS19]. commensalism [GG16]. Comment [LZ19c, Pan15, PK18a, SR15d, dF15]. Comments [Dai16, Ort15b, Rho19, Rho14, dFY15]. commerce [BM17]. commercial [LLZ^{+}16]. commodity [GKDHR19]. common [ACS15, AK15d, Boi15, GS11, KT16, LVAB^{+}16, PG15b, PN15b, WJGM19, YZ17a, ZS15d]. communicable [XWDW17]. communicating [dSVR17]. communication [LSCK15, LLC18, LS19c, LGHW19, SWSW19, WZK15, ZZZY17, ZDF17]. communications [GSD16, ZZL^{+}18a]. Community [CLHC16, WWC^{+}18, ZCML19]. commutative [WL19c]. Commutator [CCD18]. Commuting [SML^{+}15, KC18b]. Compact [MDH15, WMSP19, ZCXX18, ACKS16, CC18b, CDL19, Cui15, DL18b, GfSmL16, HEA15a, HEA15b, cJzS15, LZCW18, MDHD18, Pan18, RLZ19, RW17, Rom19, RGA19, WZ15b, WMLG17, WZZ18, WJWX18, WW18, WR19, WLWD18, YMD15, YGZ19, ZZ18c, ZZX19, ZRLX19]. Compactly [SZ18a]. Compactness [ZZ16b, O’R18b, WL19a]. company [TB15]. Comparative [SBG^{+}17, BKZA18, DV15, Ely18, DPL18a, RS15a]. compare [CvRvdM18]. Comparing [CSZ19, CLSS19, CN15d, Vad19, BBFJ17]. Comparison [ASKK19, AH15, BVJ15, BL15, CN16b, DD16a, SdA18, SH15c, TLLD19, WR15, AG15c, GIS18, GL18b, KKKD18, MFF15, NCEGA15, ÖÖAD15, PL15b]. comparisons [DR19, KS15]. compartment [LWY^{+}19]. compassion [LZP18]. compatible [RSAF16]. Compensated [Her19, GJP18, Yue19]. Compensation [MGPL15, WLLL15]. competency [DCA15]. competing [PSM17, WYW16, ZJ15a]. Competition [Haf17, GG16, GCLG15, HCH16, KZ15, LLY19, SSZL17, WXLL19]. competitions [Par18b]. competitive [CWFL15, CXMP16, KN19, LLMS16, LYM18, MWY17, NLC19, PP15a, QD18, TLGX15, XYDL16]. complement [LZZT18, LpLxG18, NC19, RMSD16, Xu17b].

Complementarity

[BFIJ15, dCSS17, BFJ^{+}17, CW17, DLBQ19, LW15d, LVC15, NJTP19, TZF15, WZLP18, XL15, ZGL15, ZVL19]. complements [Gut17, JYSR18, LWG16, XLS18, YFY17]. Complete [AGN15, FT18, Tal15, Wan17a, BCD18, HH17, LZ18e, LGL19a, fLLPY16, Wil18, YCZ19]. completely [CE15, CP15b, Che15b, LS15a, LZjy18]. Completing [Tha18a]. completion [MTY16, WY15b, WL17, WLZ19b]. Complex [Mis15, PL15c, AF18, AQ15, AQ18, AAE19, AY16a, AJP18, BK18a, BAE15, CR15, CGC19, CZLS18, CZSH17, DWRZ17, DS19, DSM15, FZL19, Fan15b, FYZ16, GCAA18, GZLL17, HCX^{+}17b, HYZ^{+}18, JZ15b, LWQ18, LL15d, LWSX18b, LWL^{+}18, LWSX18a, LJYH19, LDS^{+}18, LML18b, LZKX19, Ma15a, MFC17, NS15b, PG15b, Pet15a, QZJY15, RVGS17, RG18a, SMI18, SJ16, SSA18, SCC17, SRCA17, SSPCPG^{+}16, SS15c, SQL16, yWyLyY^{+}15, WWDW17, WA18, WLX18, Wan19a, WRX^{+}19, WWSM19, WZS19, WHWW19, WZWD15, WZK15, XZS16, XLJC18, XKG15, XXZ^{+}18, YMD15, YWZS18, YYS17, YSF16, YQT15, ZSA^{+}15, ZLHH15, ZZL^{+}18a, ZLZ^{+}18, ZWHC15, ZYXF17, ZD17, ZSZ17, ZYC18, ZWZ19a, ZCML19, ZM15c, ZM18b, ZDF17]. complex-valued [CZLS18, CZSH17, GZLL17, HCX^{+}17b, LWL^{+}18, SCC17, WZS19, XZS16, ZLHH15, ZYXF17, ZSZ17, ZYC18]. complex-variable [XXZ^{+}18]. Complexity [BS19c, LYW17, AK18a, dCM16, Gon18, Han19, KJR^{+}18, LLS15, LYW18, NL18b, PBMP15]. complicated [Gia15a]. component [ACE19b, FS16, GMPP18, GYW15, GLW15, HXM18, SSS15, WZZT19, YL19a, YX19b]. components [CXZZ18, CnTVsL18, HTZY17, LZH^{+}18, RC16, SM15b]. composite [DNC^{+}17, Gia15a, Gia15b, GLW17, KCZK15, LZL15a, LZL17, Li16b, MTJ15, SSK15, YJW16, ZLL15b, ZM19]. Composites [VB15, MNNP19, WQ15]. composition [bBjJ16, JKK15, Mah18a, SNS^{+}16, SSB16, Zhu16]. Compositional [YSX15]. compound [AM15d, BKZA18, BE17, ZH17]. comprehensive [ÖÖAD15, WSXR19]. compressed [WNF19, ZWY15b]. Compressible [SBH17, ACG19, DMA^{+}19, DC16, GG18, GZW15, KFHH15, KSZ15, LR17, LTL16, LCW19b, NS17a, Pel17, QX17, RDG18, XP17, XLR17, ZZX19, ZM19]. compressible-fluid [GG18]. Compression [JJM18, Len15, AK15b, TMS15]. compression-expansion [AK15b]. comprise [Ber17]. Comput [MAK15a, MAK16, Sab18, Tom15, VK17a, WL18a]. Computation [AZ19, Bag19, BE17, DN16, DSDL19, EL19a, FMB^{+}19, KRS18, LLZD18a, PPS25, Pan15, PK18a, PABA15, PP15b, Pon17, RRR20, Rho19, SS17, She18, SN18a, XS19a, YLWF16, ZZ15c, ZLL15b, ZL19b, AQ18, AZL15, BB15a, Bas19, CSZ19, CH18c, DO17, FFN18, Fuk15c, Fuk15d, Fuk15b, GST19, HP17, KDS16, Miy19, NSM16, PSZ18, PH17, QLZ15, RRM16, SID19, SS15a, Sku15, SSPCPG^{+}16, TES15, Van15, WLZ15, XMX15, YWYS18, ZQ18, ZL15b, Ano19w]. Computational [ASA19, AB18, AELH15, FLS^{+}15, LZ17d, NBAGS19, PLL16, SST17, SAA19, VB15, ADG16, Cat19, FM17b, HAY19, HS15c, HHT19, Ili19b, IR16, IS16, JMK17, Kim18, MFF15, Par18a, PP19, Sah19b, WWR^{+}15, jXlZX18, dNdOX^{+}15]. computationally [SBV15]. computationally-efficient [SBV15]. Computations [BKS18, Bot15, Cal18, DP17, DG18a, Far18a, PS15a]. compute [ATV18, FG19, MGL19]. Computer [LCW^{+}18, wG15a, NMS^{+}18, PJ18, PH17, RCRL18, SKHA18, VD15, YSZ19].

Computing

[AT15b, AMCMFJ15, BDBKV19, BCd15b, CP15c, GST15a, LW17b, MPB^{+}17, SA18a, SGGD^{+}18, SKP16, WTZL18, YEMRV17, Zhu15, Ace19a, AJC^{+}18, ASZD18, dCM16, GGAVGG^{+}18, GJNC18, JXW15, Li17c, LCW19a, NJM^{+}18, SJ19a, SIA^{+}17, Tha18a, TB17b, USCA15, USCAAA15, WMS17, WYL^{+}19b, XZS16, XC15, Xu15b, YKZ15, Zad16]. CON [XY15b]. CON/SLK [XY15b]. concatenating [HRLGÁ18]. concavity [Dan15a, ZC19c]. concentrated [SLW16b]. Concentration [TES15, LWLM19]. concentrations [NFS^{+}15]. concept [AK16a, MS16a, Zha15f]. concepts [RL15]. Concerning [Rho19, CC17, MBM16, Rho14]. concrete [KKM15, KK15c, NKKK18, WR15, Zha16b]. condensate [WJWX18]. condensates [ID19, QFS17]. condensation [DMSS18, HH16]. condensing [RAH19a]. condition [ABHR15, BTH18, BF15, BST18, CGL15, DLQ17, DLH15, GXLL19, GS15b, HMGC18, KKT15, LSZ15, LDH18, MHX17, MVS19, PAM19, PLL16, Sap18, SR17, Sil16, SGMH16, Slo15, TPCL15, WZM15, WK16b, WwQqTW18, YBZ18, Yue19, ZS14]. Conditional [CD15d, GL18b, Cim18, HV19a, LHG18, QY17, TZ18d, TZS17, WX18, YFLF17, ZZX18]. conditions [AN15, ANT^{+}16, AL18b, AYD19, AM15e, AK15c, AR16b, AS15, ANEP19, BMM19, BB15d, BDSS15, CS17, CGW19, CM18, CDP19, CKG^{+}16, CZM18, CZ17b, Cui15, CZ15c, CS15b, Dan15a, DHQ19, DMR15, DO18, EKS^{+}15, EM16, FT17, Fer15, FG19, GZK16, GIS18, GXCY19, HLIK19, HL17, HST17, HS15c, JM18, JYF19, Kun16, LLWX19, LJZ18, LLX18a, LY15e, LZY^{+}17, LWS18, MM16, MMV16, MS15d, NE15, O’R18b, Qiu18, RGA19, RG18a, SR15b, SAM^{+}16, SMI18, SH18b, Tha15, Tua15, WZ15b, WZ16a, WTUZ15, XZ18, YCH19, YYM19, ZW15a, ZAW15, ZY15b, ZjA19, ZCX19, ZZW17, ZY18b, ZM18a, ZWL18, ZT16c, AAC16]. conducting [AKR15]. Conduction [YM18, KBK^{+}16, LP15a, LCL17a, Mah15, MPPP19, MR16a, Su19, TPS15, Xu19c, YDWZ19, YWYS18, ZT15, ZL19a, ZL19b]. conductive [CKG^{+}16]. conductivity [HL15d, KBK^{+}16, LP15a, WQ15, YX15]. conduit [RM19a, RMK19]. cone [Alz15, AK15b, BFIJ15, LW15d, Mah19, MRR17, SBH16, TZF15]. coneigenvalues [LCJ15]. cones [MYH15]. configurations [CCLPC19, FFH^{+}18, LY19, MSdC^{+}18, WZC17]. Conflict [Ino16, GOH18]. confluent [ALQ15, EJF15, EJF16, II18]. conformal [AK19a, ZBR15, ZQ18]. Conformance [SCS^{+}15]. conforming [cCpSkZ15, FA18, SLW16a, SW17b, ZSY17, ZCZM15]. conformity [SLLZ19]. congestion [Cao19, DKN16, NR18b, NR18a, PW19b]. conic [SKA16]. conical [MRF18]. conjecture [CSZ19, Dim17, YG15a, YCYL18]. conjectures [II17]. conjugacy [DLH15]. Conjugate [CFR18, FLMR18, HM18a, AKA19, CL16a, DCW15, Dai16, DLH15, LP15c, MS18b, PBP19, PG15b, RMA15, WSdSY15, WAS19, XM15, YLNL15]. conjugate-mixed [PBP19]. connected [Dob19, GW16, LYW17, LYW18, NS15b, RZH^{+}19, SJL17, TX18, WYL19a, ZM18a, ZWL18]. connectedness [HTM18, LLS15, LML18b]. Connecting [BDE^{+}15]. Connection [PR15b, GDL19, GXCY19, LLW17, ML17, MNJW19]. connections [CFSGM17]. Connectivity [CMT^{+}18, QGH19, BGMO18, BM19b, DD16a, DDdF16, Dim17, GFWH17, GLW18, GA18, LL15a, LY16a, LTY16, LMYW17, LW18b, LGL19a, LZLG19, LLTW19, Pal17, Pal19, TZS17, WM18a, WCD17, WC18, YFLZ18, ZW19a, ZH19b]. consecutive [DMP19]. Consensus [LZjy18, LGHW19, Sha16a, WS19a, XH16, CLZW19, FW18a, GFE19, HW17b, MM15b, MGCAPP16, SMG19, SY18a, YJMH18, ZLM^{+}18, ZJ15b, ZYLY17, ZJYD17, ZSL^{+}18]. conservation [AIIK15, BCI15, BDH16, Gla19, GL15b, HSSH18, KRB15, Liu17a, LFY18, Maj18, NKS18, QZL^{+}17, XLW15, ZY15c, dlRGB16]. Conservative [Ach19, GS17a, QFS17, WX19a, AO15c, DC16, GWK16, HP15a, DPL18a, PZ17, WJWX18, jWgX18, WDG19b, XZ19, ZZY^{+}17]. Conserved [SZ17, CWG15]. conserving [BBCI18, HGL19, LZ18b, ZLW18]. consideration [WWDW17]. considerations [Kem18, MMB15, VPS15]. considered [RCW^{+}18]. considering [KN19, LHHS16, RHJ18, TD19]. Consimilarity [LCJ15]. consistency [GXH18, SRK^{+}19]. consistent [ABGL17, AY16b, PSSX17, ZSX15]. constacyclic [ZK17]. Constant [CFG15, ACC16, CDK17b, CD15d, DKBS15, DHS19, GLM19, GM19, HSSH18, Kar19, KB17, LLY15, LLW19, LLLZ16, LY18d, LSY15, MWH^{+}18, Oni18, SSB18, SSB15a, UGY16, WL16b, ZS15b, Zha17b, ZYZW16]. constants [CC17, MPP17, Oka18a, You15]. constituent [MC15c]. constitutive [CH18c, DQ18]. constrained [ARS15, CL16a, CRZ19, CXW^{+}19, FA15, FLMR18, GdS18, GHL19, Gar16b, HH19, JPTL15, KCK19, LY18a, PZ16a, QFTT15, SNAR15, SPW15, WCZ15, WZ16a, WTUZ15, XHW18, dSH17]. Constraint [ARS15, De 18, SCR^{+}15, FMDZ19, HSKHA15, HW17c, SM17, ZZ15c]. constraints [CYSY16, Cov19, DKN16, DPSS16, LHJ15, LMZ^{+}19, PMS16, PS18b, SMG19, SSB15b, SBLZ17, TTD17, XCDZ19, ZZ14, ZLM^{+}18, ZXS^{+}18, ZX19]. construct [OW18, ZDZ15]. constructed [YFL16]. Constructing [CYCW15, DGN16, CT17d, Mao17, QCH19, GHM18, Lin15, ZCU17].

Construction [AU15, BCMT15b, FKV16, Kra18a, LZ16, LXZ19c, SST18, SuSR16, TMPG17, TLL16, ABGS15, ANGC17, ABGL17, Bra15, HCH16, HQJ^{+}15, KAH15, MFC17, Sko17, ZLD18]. Constructions [HTZY17, CZ15b, GN19a, WNF19]. Constructive [LZY^{+}17, JJ16, RRRR16]. consumer [Wan19b]. consumer-resource [Wan19b]. consumption [GNR^{+}19, HW17c, MPHG15, SBLZ17]. contact [ABD16, CS16a, HMZ17, MG19, Tch15]. contagion [PJ18, WL16d]. contained [CCY19b]. containing [GR17b, RMD18, Xu19d, YYL19, ZY18a]. Containment [LS19c, HWY15a, KZ19, SSCX18]. contaminant [ddGL15]. contaminants [LXZL19]. contaminated [WXP16]. content [NFS^{+}15, WZM15]. Context [PCA16, AJ17, TLL^{+}18]. Context-oriented [PCA16]. context-sensitive [TLL^{+}18]. continuation [Bot15, ERFE16, HR18, Kue15]. continued [CY15a, LSY15]. continuity [Han18]. Continuous [Liu17b, MLS19, ZBME16, AJMN18, ABI19, BB15b, BM15b, BJ19, CPST16, CCQS16, CGAL19, CY15c, CST^{+}15, CXW^{+}19, DY17, FVMO15, GI16, Gao17a, Gao18, Gat18a, Gav15, HMGC18, JLL15a, KWT15, LSZ15, Liu16b, LZG16, LZHC18, LGL19c, NRS19, PMSB15, PLL16, QZW19, Sch16, SK17a, SZO16, Sku15, TLL16, Tan18, TZ18c, TSZ19, TZ19a, Tan17a, WZYS15, WMZ16, WL19b, YSL19, ZLDZ17, Zha19b, ZCAN19, ZS19a, ZS14, ZCC15]. continuous-stage [ABI19, Tan18, TZ18c, TSZ19, TZ19a]. Continuous-time [MLS19, CCQS16, CXW^{+}19, Gao18, LZG16, LZHC18, LGL19c, WZYS15, WMZ16, ZLDZ17, ZCAN19]. continuum [RHJ18]. Contour [SH18a, CD16a]. contract [WWC^{+}18]. contractible [QGH19]. contraction [PN15b]. Contractions [Rho19, LZCW19, RDLG16, Rho14]. contractive [IZ16]. contracts [EG17]. control [AHP19, ASS16, AM19, BP17a, BAGN^{+}19, BMT19, Bhu15, CMW17, Cao19, CXP16, CKG^{+}16, CCQS16, CYSY16, CWL16, CGZ17, CLCW17, CXZZ18, CH18b, CLZW19, CCL^{+}16, Chi15, CL19c, CXW^{+}19, DHX^{+}19, DG15a, DQH18, DS16a, DHY^{+}18, DKL15, DH17, DWS19, FW18a, FYZ16, FYSC18, FXXC19, FNS15, FM16, GW15a, GSZ15, GhDwZS17, GZC19, GXLL19, GMOP18, GWWL17, GSPH19, GGM15, GMPP18, GRS^{+}15, GSD16, GWXZ18, Han16, HYL19a, HA19, HWY15a, HWY15b, HW17b, HHR17, HHGC16, HL16, HZLY17, HZK^{+}19, HLC19, JGX15, JGKL16, KRS18, KG15, KLKU18, KML19, KRA^{+}18b, KH17, KPP17, LZL19a, LPKS16, LY15d, Li15a, LL15b, LWPL16, LSZY17, LLL^{+}18b, LSLZ18, LQC^{+}18, LZ19d, LYH19, LGC19b, LSZ^{+}19a, LPZL16, LXC^{+}19, LGK15, LMW18, LZHC18, LS19c, LXZL19, LMZ^{+}19, LWSO19, LMZZ19, LZ17f, LTZ^{+}17, LCLZ19, Luo19, MFJZ15, MJL16, Ma17b, MLSX18, MA16, MAR15, MPS15, MGN18, MTB18, NPT19, OA18]. control [Pai15b, PY19a, PY19b, PZLZ15, PRL^{+}15, PW19b, PFA15, QTTL19, RSC^{+}17, RVGS17, RXZW17, RZHY16, RWW18, RLM19, RRS^{+}18, RPJ^{+}18, SAJ18, SSKA16, SKKL17, SJWK18, SJ19a, SZ15, SSCX18, SYZP16, SY17a, SLLC19, SMG19, SLZ^{+}16, SY18a, SS19b, SWZ^{+}19, SJ18, SML^{+}15, SRCA17, SMZ^{+}17, SWSW19, SS15c, SML18, SZT17, SLS^{+}18, SWZL19, SC17, TVK^{+}16, TTZ^{+}19, TFLZ19, Tan15b, TPL16, TTD18, TMM18, TLL^{+}18, WZAF15, Wan16a, WHTW17, WLH^{+}18, WSH^{+}18, WRX^{+}19, WSXR19, WXXY19, WCH15, WWHZ18, WSYL18, XZLZ17, XD19, XTXW15, XZL15, XH16, XPZ18, XLJC18, XCDZ19, XZCC15, XCW19, XZHX17, XCL18, XSSW19, XGQW19, XZLL17, YW15, YWZS18, YLP19, YYS17, YFW^{+}17, Yil16, YLW19, ZZ15a, ZZL^{+}18a, ZLLZ16, ZLDZ17, ZMMZ15, ZGH16, ZSS16, ZYXF17, ZZZY17, ZLM^{+}18, ZXS^{+}18, ZW19b, ZCCZ19, ZZMS19, ZY15c, ZJYD17, ZSXW18, ZWO^{+}19, ZMC^{+}19b, ZPT^{+}15, ZLK^{+}19, ZPM16, ZYWW16, ZXX^{+}19, ZZ16c]. Controllability [BGRT15, CS17, DG15b, LY15e, AS15, BT15a, CR19a, DX17, GZK16, LZ15e, LSL19b, LL15f, LLB16, Mah18b, PZ19, YL17, YW16, ZLK^{+}19]. controllable [KRS18]. controlled [CGC19, FLYG18, RVE19]. Controller [QKGW18, Zha19b, AADLF15, KSM^{+}19, LLX^{+}19, SA16a, SRSZ15, WCZY17]. controllers [LLL^{+}18b, Pai15a, PZ19, WCB19]. Controlling [HCX^{+}17a, ARV18]. controls [CR19a, LZT^{+}18]. convecting [Kud19a]. convection [AKAA15, BS17, CKG19, Cho19b, CVA18, DN15a, DBV18, DX17, GS17a, GJR^{+}15, Har16, HM19a, HRS17, LZ15d, Liu16a, Liu17b, LZ17e, MGK^{+}16, MS18a, Mah19, MN17, MW18, MS18b, Mis16, MRR17, MKN15b, PBP19, PCS17, RNM15, RW17, SHA15a, SP15, SRP17, SMR15, SM18, WSZ16, WSLC17, WW18, WX16, YM18, YWYS18, ZZY18, Zha18e, ZC19b, ZCZM15, ZX16]. convection-diffusion [AKAA15, BS17, CKG19, CVA18, DN15a, LZ15d, Liu16a, LZ17e, Mis16, WSZ16, WSLC17, WW18, WX16, ZC19b, ZCZM15]. convection-diffusion-reaction [SHA15a]. convection-dominated [MW18]. convection-reaction-diffusion [RW17]. convective [BST18, Cho18, PMV16, Zha15a]. convective-radiative [PMV16]. convectively [HAMA14, PK18a]. Convergence [CWG15, CNK^{+}18, Cui15, DN15b, DN16, EJF16, Ewa16, Fan17, GZZL17, GS17b, GIS18, Guo16, GS11, How15b, HMGC18, LP15a, LC15a, LZCW18, Lia15a, LSZ15, MHN15, Maj18, MSS15, PSSX17, SRS17, TRP15, WK16a, WY16, Yun15, ZS15d, ZL16a, ZYSL19, ZS14, AG15c, ACMT15b, ACMT15a, AM15g, AM15f, AG15d, AK15c, AG16, AM17, ABMA19, AO15c, BF15, Bor19, BLS15, BBGR15, BCT15, Cat19, Che16c, CZ17b, CDL15, CJ16, CEHVT15, CLMT15, CCGV19, CS15b, Dai16, DGZ16, EHVV15, tFyZZ15, FZLT15, GLM19, GN15, GR19a, Haz16, HVR16, HVR17, IAK15, JSG15, KKB15, KG19, LJ15a, LR19c, LYZ19, LYXL15, LB19, LFC^{+}19, LMMR15, LpLxG18, MA15b, MS17b, MSHG16, Nak15, PZ17, PZ16a, PV16, Pro16, PIP19, PV19, RA15, RG19, RM19a, Sha15e, Sha16b, SZ18b, SGJS16, SGMH16, SPKC16, Sun18, TM19, TY18]. convergence [WK16b, WH18, WLWD18, XY15a, XY17, XY18, YLW15, YSL19, Yu16, Yue19, ZRLX19, Zha15h, ZHX^{+}17, ZZX18, ZZQ18, ZYX19, ZLZW15]. convergences [You15]. Convergent [CJ17b, AS19a, AU15, Boi15, CPP15, DN15a, Fan18, LHJ15]. conversion [GZ15a, SM15a]. converters [SB16]. Converting [AF18]. Convex [ZSA^{+}15, ABGS15, AGZ15, BFIJ15, CL16a, Che15c, Cho19a, CJ17b, DÇO15, DLY17b, GGM18, GWK15, HZ16, HP15c, Isc16, Kim15b, LH15, LB19, MG15, MCZ15, MNNA15, NSR^{+}19, PCW15, PHSB18, RHP18, SISÖ15, SHLW16, SKA16, WJZF19, Wil18]. Convexity [SN15c, ZM17, CNNA16, WwQqTW18]. convolution [DMH19, HK18a, KAH15, LR16a, Li17e, Li19, Mac15, ÖK19, Boy15]. convolution-type [Li17e]. Convolutions [DD15]. convolved [SR16a, YZ17a]. cooled [GSDH15]. Cooperation [KZT19, AAA16, BK18a, CCW17, CYGL18, CCW19, CZ18b, CZM^{+}19, DZD^{+}16, DHY^{+}18, Du19, FK18b, GSD^{+}17, HZS15, HHL^{+}19, LZP18, LY18c, LJG^{+}19, LSX19, LMW^{+}19, LWWZ17, LJ17, LWZ18, Pla15, RCW^{+}18, SLS16, SLLZ19, Tan15e, WWW^{+}17, WJ19, XDG^{+}19, YC18, ZAZ19, ZY16, ZSL17, ZL19c]. Cooperative [HM15a, Ma17b, CBGGV17, GLR^{+}18, HWY15b, HHR17, HCY^{+}18, LYH19, Liu15d, MWY17, WJNW15, YWL19]. coordinate [BPR16]. coordinates [MB15a]. copper [VVA^{+}15]. coprime [HHZ15]. core [Kur18, Ma18, MFF15, RD15, ZMH18, ZC18, MS19a]. core-EP [MS19a]. cores [GIK19]. corn [LTC16]. corneal [Len15]. Corner [JLL16, TTZX15]. corona [VBG^{+}17]. corrected [BF16a, EMO18, GNW15, HMF15, YG15b]. Correction [Pal19, Bag19, CY15a, LS19b, RAV15, SJW16, Yüz16b]. corrector [How15b, How16, JDG18, KDG19, MS17a]. correctors [Tha18a]. correlation [Zhu15]. Corrigendum [DN16, DSDL19, LLZD18a, MAK15a, NCS16, PPS25, PP15b, Pon17, RRR20, SN18a, Tom15, YLWF16, ZZ15c]. corrupted [LJL15a]. COS [LOOGB18]. cosh [ZXZB19]. cosine [AQ15, SIA^{+}17]. Cosmological [Dim16]. cosmology [SR15f, SSB15b]. Cospectrality [AH18b]. Cost [SH15b, SJ19a, AE17b, BMT19, FK18b, GYY18, JNGB18, KM19, Kur19, LXPD15, MGPL15, PMSB15, TMM18, WYW16, WSXR19, YJMH18, ZY16]. cost-effectiveness [BMT19, WSXR19]. COSTs [SCR^{+}15, LW15a, NM15]. cotangent [MR19]. Coulson [QZL18a]. Coulson-type [QZL18a]. countable [O’R18b]. counter [BAGN^{+}19]. counter-terrorism [BAGN^{+}19]. Counterexamples [II17]. Countering [San19]. counterpart [SSS15]. counterparty [ASV17]. Counting [SZ19c, Flo18]. Country [CST^{+}15]. Couple [HM19a, SMG^{+}18, SCM19]. couple-stress [SCM19]. Coupled [KKM15, ANU16, ABS15, AN15, AZ14, ARSV15, AM15h, AZ19, BK18b, Bas19, CCY19a, CZ15c, DBV18, EFPT18, EYE18, Esl16, FFG^{+}16, FJ19, GW15a, GWXZ18, HRS15, HSS15, HM18a, HXW^{+}19a, HM18b, HHG19, ID19, IZ16, Ili19b, KRB15, LdCM19, LJW^{+}15, LCS15, LK19a, Lin19, LGL16, LXL17, LCW19b, LJ19, LCLZ19, MX17, MDH15, MNJM19, NB19a, PKS15, RWW15, RS15a, SV17b, Tho17, TPCL15, WWZD16, WJWX18, WXWX18, WYD^{+}18, WD16, XM15, XCL18, YWYS18, ZPZ^{+}19, ZZL17, ZWHC15, ZZC15, ZLW15, ZMA^{+}18, ZCQ15, dSSB^{+}16]. Coupling [IR16, SSPCPG^{+}16, ZLZ^{+}18, AY16a, BK18c, FYZ16, Hua15, LGL16, LMZZ19, LGHW19, MMZ^{+}17, NS17a, NRS19, PAJ^{+}19, RVGS17, RXLM19, SRCA17, TNMR18, Wu15, XJW17, YFW^{+}17, ZQG16]. Coupon [CJ17a]. Courant [SAES19]. Cournot [ANP16, AAA15b, AAA16, AE17b, CMG18, DLJW15, Els17, MS15b, Ued19, ZZC^{+}19b]. cover [LYW17, LZ18e, LZ18d, RZH^{+}19, Tu17, ZTW19]. coverage [ARV18, KIZ18, LZ15c]. covering [BGMS18, IMTJ19, PBMP15]. Cowell [LZ17b]. CPU [FFGRSV19]. CPU-based [GG18]. CQ [QLZ15]. CQ-like [QLZ15]. crack [DL15a, MWY15]. cracks [ABN^{+}17]. Craig [KSL19, Mis16]. Cramer [SWY18]. Crandall [BB15a]. Crane [Boy17]. Crank [BTWY15, LCL17b, QFS17, SLC18, ZJJ18]. crashes [LWSX18b]. creating [AKL19]. creation [IÖ19, LLT19]. credit [CLT^{+}15, JYVS15, MBM16, SCB15]. crises [WH15]. Criss [MZP15]. Criss-cross [MZP15]. criteria [ABLZ15, AH15, CDJ19, CCD18, DWZ15, DZ15, FAHZ16, GHW19, LLP17, LHS18, MJ19, MGMM17, RDG18, RCRL18, SSZ^{+}18, SLZ^{+}16, SCS^{+}15, SH15c, WL16c, WYCZ17, Zha17c]. criterion [CZS15, CXW18, Cho19b, GR17a, LZ18a, LWW15a, LTL16, MR19, QKGW18, Qia16a, Qia16b, WZL15, WDZ17, XCCL18, YZ17b]. Critical [AL19, CSZ19, DH15, Kar19, WZZT19, WC15b, YL18, Zha16a]. crop [LTC16]. Cross [LLY19, CT15, CT17d, LLWW19, MZP15, NB19a, PZ16b, PL18, YSF16, YW18c, WHX15]. cross-coupled [NB19a]. Cross-diffusion [LLY19, LLWW19, PZ16b, PL18, YW18c]. Cross-Hill [WHX15]. cross-sample [YSF16]. cross-section [CT15, CT17d]. crossed [KC18c, WM18a]. crossing [PTO19]. crossings [Zha19a]. crossover [FTF16, MHZC15, bOqGyK^{+}15]. crowd [HZYW15]. crowding [DSB15]. cryptography [UK18]. crystal [AKR15, Qia16a, Qia16b]. crystallization [SZO16]. crystals [DL16, QX17]. CS [NMNTBXVD15]. CSD [Ili19b]. CT [SKCS15]. CTL [WJHA17]. Cu [PBP19]. Cu-water [PBP19]. cubature [BGMZ17, ZWY^{+}15a]. cube [KC18c]. cubed [KNK15]. cubed-sphere [KNK15]. Cubes [CYCW15, CT17b, GSLC18, Lai15, WM18a]. cubic [AKR15, Bej17, BSZ16, cCpSkZ15, dCM16, DH19a, FHK^{+}19, GK16b, IAW18, KCK19, Kor18, LQZ^{+}18, LLz18, LCW19a, LLdM15, Lu15, MD17, MKN15b, SR15e, SR16b, TSJ16, TW15a, Val15, VLCT18, WTSQ15, XX19a, YY17, YAIN17, ZZG18, ZM17]. cubic-quintic [WTSQ15]. cuboid [LW19d]. Cuckoo [LF15, WW16c]. culling [KSN15]. culture [SZRC15]. culture-ant [SZRC15]. cumulative [BBD15, GST19, MRS18]. cups [How15a]. curl [WSC19]. current [ATV19, CLR15, CM17b, EAY17, Hsi15, KCZK15, KCWK15, KM17a, LW15b]. currents [BTP17]. currents-water [BTP17]. Curtis [LX19c, Ma15a]. Curtis-type [LX19c]. curvature [HYL^{+}19b, KSS19, Len15, Lu15]. curve [Cal18, DJLX17, eFJY17, GI16, Han15a, Khl18, KKG18, LLJ19, RRTR15, Sah19a, SR15d, SDGU16, Suk19, TÖGE18, YLH^{+}19]. curved [AB15b, Cat17, HMHA15, KHN19, KG16, MLS19]. curves [AM16, CM17b, DP15b, Erb16, GT15b, GLW17, HQJ^{+}15, JJ16, KSS19, MPR^{+}17, NEK18, RS15c, Sah19a, SHH15, UGY16, WGL15, XzXqT^{+}15, Zha15c, ZBME16]. customer [CRCVMS18, RRS^{+}18, SJ19b, SRAC16]. customers [CLC^{+}19, JBS15, PL15a, PD15]. cut [AF15, YL18]. cutting [Chi15, TTZX15, ZHWW15]. CVaR [NM15]. CWENO [NKS18]. CWT [Mah15]. cyanobacteria [NFS^{+}15]. cyanotoxin [NFS^{+}15]. cyber [LWY^{+}19]. Cycle [Li17f, CH19, CRC15, CLT^{+}15, Fer15, LWM18, MHZC15, SD19a, XLS18]. cycles [Ber17, CFSGM17, Che18b, DY17, GLN15, GYC19, Lai15, LZS15, Liu15b, LLL19c, LT19, LLdM15, LLdM16, MY19, San16, Sun15a, Xu19d, YY17, YD18, YYL19]. Cyclic [Rho19, ZK17, BCD18, BPR16, HGG16, HCH16, IZ16, JH17, LZCW19, Par18b, RDLG16, Rho14, SdR19, WTD^{+}19]. Cyclicity [MFOP17, MF17]. cylinders [Dan15b, FK18a, ZWM16]. cylindrical [KZA19, KA15c, RMM19, RSC16, RS15d, RM19a, RMK19, TPS15]. cytokine [HJ18].

D [AIP15, DHX^{+}19, Kaj18, OA15, ÖK19, WL18a, YZ17b, ZWaY18, ADSS16, ABGL17, AV15, AO15c, BAV18, BKW18, BZ17, BFR18, BP17b, BS19b, CWG15, CSV19, CLR16b, Che15e, CW17, CL18, CDL18, CJ16, CVA18, DA19, DL16, Ers16b, FK17, JOÖ16, JDL15, KCWK15, Kan16, KNK15, LYDH18, LZL15b, LTL16, LMMZ18, LDL15, LT18b, MGK^{+}16, MN17, MM18a, MS15d, MKN15b, NB19b, ODP17, Qia16a, QFTT15, RSC16, SA19, SZL^{+}18, SKCS15, SH19, SDWG19, TSA15, VG18, WL16c, WW18, WDG19b, WD19, XJW17, YLL18, YB15, YL19c, ZT16a, ZWF15, ZLDZ16, Zha16c, NKP16]. D-Kaup [MM18a]. D-MORPH [CL18]. D. [AIP15, Kaj18]. d.c. [JG17]. Dafermos [MCO19]. damage [HK18b, KKM15, KK15c]. damages [WR15]. damped [ABL15, ACMT15b, BGRT15, BGST16, Che15d, DNC^{+}17, GST15b, ID19, JX19, LSZF18, RS15d, TTP19, WFZ17, ACMT15a, VLCT18]. Damping [TTV15, BDGS15, LLX18a, MR15a, YWL18, YZZ18, ZGL15, ZWaY18, Zho15a]. dams [SRBA15, YJYH15]. Darbo [SSA19]. Darboux [LXZ19c, RM15, TÖGE18, Xu15c, XLW15]. Darcy [KZR17, NM19, WZWJ18, XLR17]. dark [HGH19b, MDH16, WTSQ15]. data [AGC19, ATV18, ACD18, BÁR15, CHL17, CLHC15, CCL^{+}16, DDH18, DJLX17, Elm15, GWWL17, GSPH19, GCM^{+}18, GWY15, HKW^{+}17, JCW^{+}15, KŚK^{+}18, KRA^{+}18b, LOOGB18, LWPL16, LXC^{+}19, LS19c, LLX^{+}19, MH19b, OLZ19, PL15b, PTBN19, PP19, QCH19, RSLL^{+}19, SKKL17, SD17, SY15a, SKCS15, SS15c, TB15, TDA17, WZAF15, WJNW15, Wan18, WSH^{+}18, WS19b, WCB19, WHWW19, WW19, XZB^{+}18, XS19a, Yos19, ZZLZ17, ZZL^{+}18a, ZJJ18, ZCL18]. data-driven [LOOGB18, XS19a]. databases [GTBC18]. datapath [BS15a]. Davey [GML18]. Dawson [AQ18]. dc [NJTP19, GMPP18, KSKW^{+}19]. DCA [TTD18]. DCell [QCH19]. dd [KMP15]. dd-adaptivity [KMP15]. DDEs [LZ16]. DDM [ZLW18]. DEA [HDG15, KWH17, YZ15a]. Dead [ZMH18, XCDZ19, ZWO^{+}19]. Dead-core [ZMH18]. dead-zone [ZWO^{+}19]. death [DGN16, JLL15a, PRA^{+}19, WCH19]. Deblurring [LJL16, DHW^{+}19, DMR15, SPW15]. debris [Vid18]. debunking [TD19]. Debye [YZJ18]. Decay [CZM18, BR15, CVJ19, FM17a, HW18, MAG15, SAJ18, ZC19a, ZZ15d, ZL17b]. Decay-rate-dependent [CZM18]. deceleration [TRRP15]. decentralization [QZL18b]. Decentralized [CHL15, GZC19, WWHZ18, XZLZ17, KLKU18]. deception [LXTF18]. deciding [LLS15]. decision [APMM18, GLL15a, GOH18, HJP16, HLFZ19, KEE17, SCB15]. decision-making [GLL15a, GOH18]. decisions [CLT^{+}15, ZW15c]. decomposition [Alz15, AL18d, BNY15, BK18c, CL19a, CL19b, FA15, GGZR16, HWZ17, HTZY17, KBR15, KZZ15, KK17, KMP15, Kyr17, LQZ^{+}18, LLz18, LLX18b, RCM15, SN15a, SN15b, SHHC18, SD19c, ZZ18c]. decompositions [LZ18d]. decompression [JJM18]. deconvolution [CPP15, LLC^{+}15, LSZ^{+}19b]. decoupled [LL17a, ZJJ18]. Decoupling [AP18a]. decrease [PMSB15]. decycling [CRC19]. dedicated [PZ19, WWR^{+}15]. deep [KSA19, Tsy19, GS19]. default [WL16d]. defeating [LS15b]. Defect [GNW15, SJW16, RAV15]. defection [FK18b]. defense [FHV15, QTTL19]. defibrillation [CK17]. deficient [MZP15]. defined [AIP15, AK15b, CIPT15, Gar16a, HLT16, Kaj18, TES15]. defining [DBV18]. definite [ACKS16, ANEP19, tFyZ15, FM17a, HL16, LWS15, SY18c, WZLP18, YZ15c]. definitions [SI15]. deflection [KS16b]. deflections [AJC^{+}18]. deformable [BVJ15]. deformation [WE19]. deformed [Xu15c]. deforming [RL15, TSPF18]. Degasperis [GYW15, WD16, ZZG18]. degeneracy [PP12, PP15b]. Degenerate [DKKM15b, KKMS16, DKKM15a, DX17, Erb16, FQMR19, Kim15c, KKD15, KH17, MN17, PS18a, ZJZ15, ZY15d]. degradation [PWFY18]. Degree [BA17, CD15a, GLW17, AHG15, ARB16, APPE17, ATX19, CYZ17, CDS15, CGJZ18, CYW18, DGNA15, FT18, FKV16, Gou18, GXCY19, Gut16, GFD16, HLL16, Ili18, JYF19, LLWX19, LYS17, LG18, LCM19, LZZT18, LLdM15, LLdM16, MB19, Rad17, RCTZ18, SZ16, TF18, TDS15, WZL15, WWW19, WYL^{+}19b, XZW^{+}15, YSY^{+}17, YD18]. degree-and-distance-based [GFD16]. degree-balanced [RCTZ18]. Degree-based [CD15a, CDS15, DGNA15]. degree-Kirchhoff [HLL16, MB19]. degrees [Zha19a]. Delay [HH18b, MAR15, MS15b, PAM19, SZ15, XKWG15, YKJY17, ZWHC15, ZHJ^{+}17, AO15b, ARV18, AK19b, BÇS15, BT15a, BDGS15, BB15d, BDSS15, BB16, BS15b, BGST16, CLL16, Cao19, CCNQ19, CDK17b, CJN19, CSWJ15, CJ18, CD16a, CGZ17, CLCW17, CXZZ18, CWLA18, CSL19, DN15a, DAQI15, DWZ15, DKBS15, DZ17a, FP18a, Fan15b, FLL15, FM16, GWW15, GS15a, GSZ15, GHW19, GLM19, GGHL15, GCMR18, GFG15, GJZ19, GRS^{+}15, GZD19, GM19, GSV18b, GHM15, HD17, HFCS16, HW18, HT16, JHWZ15, JS19, JGX15, KBR15, KN19, KSN15, KKL18, LP15a, LC15b, LZ16, LSZY17, LZ17b, LZCW18, LW18d, LSR18, LSLZ18, LZ18c, LQC^{+}18, LWZ^{+}19b, LTQF19, LLW19, LSL19a, LLLZ16, LLZ17, LXT17, LZ15b, LC15d, Liu15c, LY16b, LYM18, LZ18f, LXW^{+}18, LMW18, LZ19f, LMD19, LWZZ18, LZH^{+}18, LJHW19, MFJZ15, MC15a, MJL16, MLSX18, MLG18, Mil16, MSS15, Miy17, MDZ17, MO15b, Nas19]. delay [NT16, NHWZ15, Obr19, Pai15a, Pai15b, PSM17, QKGW18, QBC19, QZW19, QD18, RSC^{+}17, RN19, RZHY16, Reu15, RLM19, SuRI15, SAA19, SM15b, SR17, SLZ^{+}16, SDP16, SZFS17, SM18, SAT15, SWZL19, TTZ^{+}19, TG16, TCS15, TNPL15, Wan17a, Wan15d, WLZ16, WGLZ16, WDZ17, WLZY17, WCZY17, WZZT19, WY16, WSYL18, WLWD18, WNW^{+}19, XZLZ17, XJZ^{+}15, XWZ15, XCL18, XL18b, XDH19, YW15, YL17, YLCW18, YLL15, ZAP19, ZLHH15, ZZLZ17, ZLW19, ZS15c, ZZC16, ZHWW17, Zha17a, ZCXX18, ZYC18, ZYSL19, ZCAN19, ZS19a, ZC19a, ZRLX19, Zha15g, ZZX18, ZZL18b, ZPM16, ZSL^{+}18, ZXX^{+}19]. Delay-dependent [SZ15, XKWG15, YKJY17, ZWHC15, ZHJ^{+}17, CXZZ18, KKL18, LZ16, SLZ^{+}16, XZLZ17, XJZ^{+}15]. delay-differential [FP18a]. Delay-range-dependent [MAR15]. delay-type [SuRI15]. delayed [ADY16, AVCCGAVDL15, BW15b, BDSS18, CX19, DHX^{+}19, DKP16, DHS19, EA15a, FYZ16, FZG19, GWW15, GXLL19, GXH18, GIS15, HCX^{+}17b, HCX^{+}17a, HLC19, JAU15, LL15c, LX16b, LZjy18, LW18d, LGC19a, LYH19, LXZL19, MHX17, MPS15, PAM19, PW19b, Pos17, PRC^{+}19, QG16, QZCJ19, QLF15, RP19, SAJ18, SLZ17, TLX19, TS15, WYCZ17, WM18b, WY18, WSH^{+}18, WXLL19, Wu15, XZZ^{+}18, XZ15, XKG15, YWZS18, YFW^{+}17, YLW16, ZLJ15, ZYXF17, ZCX19, ZZH15, ZWZT18, ZDF17]. delayed-HIV [EA15a]. Delays [BF16a, AY16a, BPC15, BP17a, BJ19, CGC19, CHL15, CZS15, CZSH17, CLPY18, CG19, CCZ15, CCL^{+}16, DK15a, DCR15, DLXZ19, DHMT15, Du15, EO15, FYSC18, GSPH19, GZLL17, GWXZ18, GSVK17, HS15a, HXW^{+}19a, JYS16, JGKL16, JSW^{+}18, KBR15, Kim15b, LLP17, LY15d, LJ15a, LSAR15, LC16a, LCAA17, LSZ^{+}18, LZ18a, LSMT18, LWL^{+}18, LSLZ18, LSL19a, Lia15a, LS16, LLZ^{+}16, LZG16, LGL16, LYZL18, LGHW19, MRC^{+}18, MA16, MAR15, MH17a, MHZ18, NR16, NLC19, PY19a, PY19b, PW19a, PRMC19, RZSS15, RMB15, SSZ^{+}18, Sha16a, SZ15, SCC17, SD19a, SML18, SLS^{+}18, TXG15, TDL^{+}17, TZD^{+}19, WPL15, WHX18, WDLZ19, WCHW19, WCH15, WZY^{+}19, XYDL16, XZ18, XCCL18, XLL^{+}19, XXZY19, YWL15, YZH^{+}19, Yu16, ZZZ15a, ZXW16, ZDPZ15, ZLW15, ZYXF17, ZZZY17, ZSZ17, ZZ18a, ZWBC18, ZW19b, ZXX^{+}19]. delineation [SCS^{+}15]. delivery [HSKHA15]. delta [BW15a, HLC^{+}17, QFS17]. delta-Eddington-type [HLC^{+}17]. demand [AAA15b, Chi15, WQK15]. demands [PL15a]. DEMATEL [WC15b]. demicontractive [AT15a]. demographic [GLM16]. Demography [IYM19]. Dempster [DJW19]. denaturation [DLS19]. dendrimers [LL18b, SIA16]. Deng [BHD18]. dengue [ATN^{+}19, SG19b, XZ15, YG15a, ZXC19]. Denoising [GS19, CCC^{+}17, CCX15, Laz17]. denominator [Han15c, TW15a]. dense [YM18]. densities [CDP19]. density [BMC19, CNQRR18, GG19b, GIT15, KS19c, LDSD19, SK15, Tan15d, ZS16c, CLHC16]. density-dependent [Tan15d]. Density-Ordered [CLHC16]. Dependence [ZW15a, CY15c, Ely15, Liu17b, MCPO15]. dependent [Abd15, ABS15, AO15b, AAM19, BHM15, CL19a, CCGP16, CHL15, CZM18, CXZZ18, CLT^{+}15, CVA18, CP15d, DH19b, DSB15, DKZ16, DL17c, Dol15, GS17a, GSZ15, GW15b, GR19a, GR15b, HLIK19, HL15a, HS15b, HG18, HH18b, HYZ^{+}18, HLH17, HL15d, JX19, KCZK15, KLZ19, KB17, KI15, KKL18, LdCM19, LZ16, LW18g, LKZ18, LTQF19, Liu17b, LS19b, LFC^{+}19, LYY15, MDZ15, MWRT17, MAR15, Mil16, PGLG17, PYZS18, PW19a, PL18, RRRR16, RP19, RG19, RZW18, SZ15, SPP17, SMG19, SL15b, SLZ^{+}16, SY19, SJL16, SAT15, SCM19, TRP15, TMPG17, Tan15d, WPL15, Won15, Wu17, XZLZ17, XD15, XJZ^{+}15, XKWG15, XX19b, YKJY17, ZZLZ17, ZLLZ16, ZWHC15, ZHJ^{+}17, ZL17a, ZjA19, ZCCZ19, ZXD15, ZT16c, dFGAN19]. depending [ZSL17]. deployment [GGM15, SZRC15]. depot [ZLLP18]. depth [DeC19]. derivation [FT17, GYW15]. Derivative [HMT15b, ABHR15, ASHSC17, AP19, AG15c, AKKB17, Ata16, CP19, CC18b, DCW15, DQ18, Gao18, GR19a, GT15c, GST15b, Li18, LJL19, LCL17b, MWWM15, Maa15, NHWZ15, RR15, RR18a, RG19, SA16d, SGMH16, SKHA18, WZ15a, WZ16a, Won15, YJW16, Zha15c, ZL19a, ZL19b]. derivative-dependent [Won15]. derivative-free [ASHSC17, AKKB17, DCW15, SA16d, WZ16a]. Derivatives [BLK19, Qi15, ARACCA^{+}18, AHS18, Akt15, ARL15, AM18b, AG15d, AJP18, Ben18, BZ17, DB18, GDS16, Kat15, LQ15, LJ19, Mis16, RA15, Sow18, SQL16, Tua15, XYG^{+}18, XC15, ZT15, ZLD18, ZW18, ZZ19c]. descent [BPR16, DLH15, Fio16, FNS15, GS17b, NRAEC19, SDK15]. describe [DTJ18]. describing [AAS^{+}17, GR17b, SWA^{+}19, WJWX18, Zha15a]. description [DLS19, DH17, KSV15, yWyLyY^{+}15]. descriptor [CTT^{+}19, Du17a, GZC19, KGS^{+}15, LZ18c, WMZ16, BGSCGL19]. descriptors [PL15b]. Design [ACTV17, SFC^{+}17, XCY18, ZWY^{+}15a, AM19, BCM17, DZC15, DCJ19, DL15b, eFJY17, Han15a, KVS15, LW17a, LZ18c, LSZ^{+}19b, LWWW15, LXTF18, LLX^{+}19, NT16, QG16, QKGW18, RXZW17, SA16a, SRSZ15, SR15d, SBG^{+}17, TPL16, Wan19a, XZZ^{+}18, YLCW18, YYX^{+}19, YMeFG19, ZZL16a, ZwX16]. Designing [PHSB18]. designs [DL18a, LLL17a, LLL19b, NPT19, TZ18b, TZC19, YFL16]. destabilized [BT15b]. destabilizing [CGW19]. destructive [CLR15]. detect [DLCV17]. Detecting [NMS^{+}18, Pet15a]. Detection [ABB^{+}18, CLHC16, DMS^{+}18, Du17a, GWK15, KGS^{+}15, KPP18, LW17a, LY17, LSZ^{+}19a, LPS19, LZHC18, RCRL18, XCY18, ZSA^{+}15, ZLLZ16]. detector [YLL^{+}16]. deteriorated [LZ17c]. deteriorating [BS15b, MTY16, SCB15]. deterioration [RZR16]. Determinant [dFY15, APS16, DÇO15, FSW16, Net15]. Determinantal [CN15a, KR15, Kyr15, SR16a, CN19a, GLS16, KM15c, Kyr17]. Determinants [HHZ15, ZL15b, ZS15e]. Determination [BB18b, HWBL19, Dan15a, YX15]. determine [CD15b]. determined [DL17a, TS18, XLS18]. Determining [CLL^{+}15, Dan18, FLT18]. Deterministic [JA15, WNF19, CM16, KM18b, Vid18, WR15]. detrended [FMRRT15]. Developable [WJGM19]. developed [ZWM16]. Developing [CN19b, BCMT15a, GKN15b, PH17]. Development [FFG^{+}16, FFTB16, HS15c, KPG16, MC15c, SM15c, Cho19b, Haf17, KSP16, VDV15]. deviated [Mus18]. deviating [Can15, IT15, JLX15a, PKB15]. deviation [DJLX17]. device [ANU16, YLL^{+}16]. devices [DLCV17, dNdOX^{+}15]. DG [LZ19c, Bac15, BB19a, MO16]. diabetes [EOP17]. diagnosabilities [WX18]. diagnosability [GL18b, LHG18, WY17, WM18a]. diagnosis [EOP17, GL18b, LGCS19, RCRL18]. diagnostic [SR15f]. Diagonal [KK16a, HWL^{+}19b]. diagonalizable [CCY19b]. diagonally [LZZT18]. diameter [DLZ18, HWL^{+}19b, MBW19, RGH^{+}15, SISV19, WWH16, WLQZ18]. Diamond [LT15, BMGMC17]. Diamond- [LT15]. diaphragm [KS16b]. diarrhea [BMT19]. Dickman [DM18]. Dickson [KAS16]. dielectric [FW18b, QBNF19]. dielectric-magnetic [QBNF19]. diffeomorphisms [ZQ18]. Difference [Sap18, Ach19, AK16a, ANT^{+}16, ABGL17, AO15c, BS16, BD15, Bas19, BBFJ17, BKL^{+}17, BJ19, CXL15, CÖ15, CLS15, CLJ^{+}15, CKSV19, Cui15, DGZ15, DJL17, DFP15, DCR15, DBJC17, Ely15, EdLC18, Far18a, GGHL15, GQC15, GN15, GLL15b, HQ15, HJMB18, HEA15a, HEA15b, HK15, HFCS16, HV15, HP15a, HP16, HJKP18, JDG18, KZR17, KXZ19, KAS16, LL15e, LCQ15, LR16b, LR17, Li17b, LZCW18, LR19b, Liu15e, Liu15a, LZ17d, LXZ19c, LJC16, LT18b, MVS19, Mig15, Mig16, MNR19, MMZ19, Nam19, Oni18, Pos17, PCS17, QXW19, Reh18, RzS15, RW17, RGA19, SAA19, SR15b, SMI18, SR17, SzSC19, SDIS15, Su19, Sug17, SzShG16, SX16, WXY15, WC16, WMLG17, WSLC17, WJWX18, jWgX18, WDCJ19, WDG19b, WR19, WLQ^{+}19, Wei17, WBL17, WLWD18, XZ19, XW19, YMD15, YGZ19, YU15, Yos19, ZEF19, ZB17, ZZ18c]. difference/local [Wei17]. differences [ALN16, KKT15, LK15, NT19, SR15e, SS15b, YLL^{+}16, ZI18, ZWF15]. differencing [BKW18]. different [BMDS17, BS15b, HH16, Isc16, JYVS15, Kan19, Kun16, LSL19a, Luk15, MGL19, ND15, PP18b, SAM^{+}16, VS19, ZXDP18]. differentiable [DLY17b, HVR16, HVR17]. Differential [CGPT15, CW18, DLJT16, DGRR16, KSA19, AB15a, AED15, ALL18a, AÖ15a, ABLZ15, AÖ16, AZL16, AOH17, AHO18, AN15, AM15a, AL18b, ARB19, AEB18, AZ18, AZ14, AAKT15, AIG17, AP19, APL15, ALP15, AK19b, AAH19, AZ19, BD17, BÇS15, BEM15, BT15a, BS16, BMO15, BBF^{+}18, Bas19, BD18, BAE15, BDGS15, BB15d, BDSS15, BB16, BDSS18, BB19b, BGST16, BS15c, BR17, BCD^{+}19, CT17a, CS16a, Can15, CLL16, CPST16, CCNQ19, CR19a, CCJ15, CLX17, CDK17b, CJN19, CZP15, CDJ19, CKW15, CC15, CW17, CH18b, CG19, CCZ15, CMW15, CNQRR18, CKSV19, Cov19, DK15a, DK15b, DTM19, DOR18, DZW15, DK16, DZ15, DPSS16, DO15, Du15, DB17, DJ18, EAAM^{+}15, EA15a, Ely18, FP18a, Fan15a, Fan17, FG19, FY18, FTF16, FY19, FZG19, GS15a, GW15b]. differential [Gao16, Gao17a, Gao19, GGHL15, GCMR18, GD15, GAB17, GK15, GQC15, GFG15, GBJ19, GR15a, GLN15, Gra15, GJZ19, GE16, GZ15b, HLZ15, HPN15, HV19a, HJKP18, HZLY17, Hsi15, HL15b, HH18b, tHBSfY19, IT15, JDG18, JS19, JLX15a, JS15, Jia15, JYS16, JISI18, KJ15b, Koc15, KK17, KDG19, KAS16, LJZ18, LSfW15, LQ15, LJW^{+}15, LSAR15, LC15b, LL16a, Li17a, LH17c, LW18d, LZ18a, LSR18, LZCL18, LK19a, LXZ19a, LLW19, LY15e, LBB18, LWW15a, LWW15b, Liu15g, LWWW15, LFL^{+}15, Liu16b, LDH18, LZ18f, LMMZ18, LZ19f, LWSO19, Liu19, LMD19, LLdM15, LVC15, LRH15, LXZ19c, LZ19g, LO15, MWWM15, Ma17a, MZM15, MRF18, MYHS18, MMV16, Mil16, MH17c, MSS15, MD17, MGMM17, MDZ17, Mus18, Nas19, NRT16, NL18a, NE15, Obr19, OS15, Oka18b, ÖAD15, Özb15, PAE^{+}17, PKB15, Par15, PWS17, PW19a, PZ15, PTBN19, Pi15, PS18b, PS19, PK18b]. differential [QXW19, QZ15b, QTF^{+}16, QYF^{+}18, QZW19, RN18, Reh17, RT18a, Reu15, RME18, RRV15, RXL19, SWUE19, SuR15, Sae17, SAA19, SR15b, SR15c, Sak19, SM16, SR17, SHR18, SS19a, SXS15, SPS16, SPST17, SPS18, SN18a, SN18b, SGKR19, SAS19, SM18, SAT15, SE18, TB17a, TSJ16, TZ19b, Tho17, Tom12, Tom15, UBT18, Val15, VRC15, VDV15, VDI17, WNMC15, WqTjT15, WIF15, WZ15a, Wan15d, WLZW15, WA15b, WWZD16, WZ16b, Wan17c, WwQqTW18, Wan18, WXWX18, WY19, WZS19, WY16, WWMX16, WL19c, WZ18b, XPG15, XHW15, Xu17a, XDH19, Xu19b, XL19, YL17, YR15, YW16, YMS16, YWO18, YSL19, YYM19, YG15b, YMMH15, YJSW16, YLH^{+}19, Yüz15, Yüz16a, ZAP19, ZBBW17, ZGGY19, ZSL15, ZS15b, Zha15f, ZAW15, ZY15b, ZLWW15, ZMQ16, ZGH17, Zha17a, ZMZ18, ZYSL19, ZS19a, ZC19a, ZCQ15, ZSY15, ZCX17, ZZX18, ZYX19, ZYJ15, Zho15b, Zho15d, Zho15c]. differential [ZYZW16, ZZ19c, ZK18, dSH15, uRIS17]. differential-algebraic [AAH19]. differential-difference [CKSV19, GQC15, LXZ19c, SAA19]. Differential-recurrence [CW18]. differentiated [ANP16, AEP15]. differentiating [LD19b]. Differentiation [KA15b, CZLM18, GBJ19, LWQ18, ZZC^{+}19b]. differentiator [Ort15b]. differentiators [LD19b]. Difficulties [Pet15a]. Difficulty [NEK18]. Diffusion [AYAA^{+}15, SSM^{+}15, AKAA15, AAA^{+}15a, ARL15, AG15a, Ali15, AHY19, ADG16, Ata16, BMM19, BAB18, Bea15, BS17, BKW18, Bho15, Boy15, BHM15, dCCGV15, CH18a, CL19a, CG15, CM18, CKG19, CLJ^{+}15, CGW16, CZL18, CDZ^{+}18, CC18b, CWL19, CY18, CDL19, CD15d, CL16d, CJ16, CVA18, Cui15, DN15a, DBK15, DSSC19, EFPT18, FZLT15, FQMR19, GS17a, GN15, GPHAPRW16, GHZ19, HW19, HFCS16, HGH19a, HAY19, HAH19, HV19b, HSS15, HYW^{+}15, HHY19, IJ15, IWL17, JM15, JL15, cJzS15, KDH19, KD16, KK16b, Kri18, LRF^{+}18, LLY15, LL15e, LC16a, LCAA17, LZCW18, LW18g, LW18f, LLTL18, LLY19, LZ19d, LW19b, LL19c, LRL18, LZ15d, LDD15, LLG^{+}15, Liu15c, Liu16a, LZ17e, LLWW19, LFC^{+}19, LM15, MHX17, MN17, MHM15, MW18, Mis16, Mou15, NBAGS19, PZ16b, PSY^{+}17, PL18, PST18, PJ15, Pol19, PeO18, PCS17]. diffusion [QG19, RLZ19, RzS15, RW17, RZW18, SHA15a, SDS18, SA16b, SZ19a, SZL^{+}18, TTZ^{+}19, TTD17, TXG15, THM19, TDL^{+}17, VD16, WSZ16, WSLC17, WZZ18, WA18, WM18b, WW18, Wan19b, WR19, WGSX19, WX16, Wei17, WSYL18, WNW^{+}19, XW19, XX19b, YMS16, YW19a, YW19b, YAIN17, YW18c, ZMBJ19, ZNV19, ZPZ^{+}19, ZWF15, ZLZ^{+}18, ZJZ15, ZS16c, ZSY17, ZZY18, ZCXX18, ZC19b, ZRLX19, ZZH15, ZCZM15, ZZL18b, ZHW^{+}18, ZWZT18, ZMH18, ZEW19, ZBL15, ZTW16, ZX16, ZS17, ZT16c]. diffusion-elastodynamics [HSS15]. diffusion-wave [AHY19, HAY19, HAH19, LW18f, SZL^{+}18, Wei17]. Diffusional [KA15c]. diffusive [CX19, DWTT17, GXH18, HD17, HM19a, HL15d, KSN15, fLLPY16, MGK^{+}16, STJS16, SL15b, TS15, XZ15]. diffusively [ZWHC15]. diffusivity [BHM15, CLJ^{+}15, CD15d, Ste19, ZFL18]. digamma [QM15b]. digital [CMW15, Han19]. digraph [ZDJ19]. Digraphs [KST16a, CMT^{+}18, DLZ18, FKA17, HFK19, LS15a, LWM17, LH19c, ML17, MR16b, YW18a]. dihedrants [LXCK17]. Dijkstra [GGAVGGRC15]. dilated [MPPP19]. dilemma [BM17, CZM^{+}19, ESS16, ESES16, GlH^{+}19, GSD^{+}17, HCS19, HI15a, HCY^{+}18, LGL^{+}19b, NT19, RCW^{+}18, SLS16, Tan15e, Tan17a, Tan17b, WJ19, WL19b, XDG^{+}19, YFLF17, ZAZ19, ZHY16, ZSL17, Sch16]. dilemmas [BK18a, LZP18, LSX19, LSLL19, Pla16, Pla17, SLLZ19]. Dimension [BLM^{+}16, BGMS18, CP15d, FMSG15, GPRCTRY19, KKTY17, RHP18, TTP19, WLL16, YEMRV17, ZT16b, ZEW19, ZTW16, dSSB^{+}16]. Dimensional [YJSW16, eMAGG19, ARB19, AHY19, AND17, AE17a, AD19, AAMC19, AADMM16, BGMZ17, BN15, BB19a, BKKN16, BRD15, BBGR15, BTWY15, CL15b, CGW16, CDL19, CS18, DA17, DPL^{+}18b, DCS15, DL18b, FFN18, FMC15, GZG^{+}15, GfSmL16, GS17c, GYC19, HAC18, HGH19a, HS18, HSS15, HZ17b, HM19b, HSSH18, ID19, cJzS15, KE16, KM17a, Kok15, KLS15, KL16, KNT17, Kra18a, KSV15, KKG18, LJL18, LL18a, LW19b, LL19c, LW18h, LHZM19, LJC16, MWWM15, MZM16, MH19a, MRE18, MMIA19, MFF15, MMV16, MH15a, MH15b, MS19c, Mis16, MD17, MB17, MKS19, MPB^{+}17, NE16, NSM16, OW18, PeO18, QDH^{+}18, RN18, RC17a, RDG18, RSLL^{+}19, SK15, SM19, Sha15d, SBV15, SS19a, SJ18, SGJS16, SV17b, SGY^{+}15, TPS15, TMS15, VK19, VBG^{+}17, WDLL15, WTXZ16, WN16, WYD^{+}18, XYG^{+}18, XW19, YCH19, YMS16, YNMW16, YWH15, Yüz18]. dimensional [uDuI18, ZB17, ZH15a, ZW15b, ZF18, ZBH^{+}15, ZHX^{+}17, ZG15b, ZGZ16, ZMH18, ZPT^{+}15, ZLW18, ZT16c]. dimensionality [GIS18]. dimensions [ACE19b, CKG19, DKKR16, JL15, KL15, LL16c, MLT17, NS17a, PRD19, SdA18, VV18, YÖ18, ZZY^{+}17, Zhu18]. Dimensionwise [Rah19b]. dimer [LY18b]. dipolar [KMR18, MCO19]. Dirac [AOS19, Fuk15c, Fuk15d, Fuk15b, GWY15, tHBSfY19, Tha15]. Dirac-type [tHBSfY19]. Direct [EEG^{+}16, GR15c, HS15b, KBR15, KSN15, AHY19, ARR16, CYLW15, GAT18b, GE16, Hua15, HYW^{+}15, JISI18, LWQ18, SDWG19, ZVL19, BG17]. directed [ADMV19b, GDL19, LY15a, ZZL^{+}18a]. Direction [Sah19a, DLH15, GZ18, HW19, KM15b, LYXL15, MGK^{+}16, MN17]. directional [Tua15]. directions [CLZW19, FW18a, Ma17b, NRAEC19]. directly [DBV18]. Dirichlet [ASA19, BB15a, Cho19a, MHX17, ML15, ZT16c]. disc [WZY^{+}19]. discharge [GWPA18, ZBKS16]. discipline [KGGAV^{+}18]. Disconjugacy [AÖ15a]. disconnected [GRV15, MK15a]. discontinuities [AV18, Kac15, LS17a]. discontinuity [SS18b]. Discontinuous [AYD19, DKKR16, AKK15, AHY19, ACG19, BCG19, BVJ15, CZ18a, CG15, CDL18, DG15a, Dol15, DS16b, DF16, EKS^{+}15, EO18, FY19, GWK16, GM17a, Gia15a, Gia15b, GGHO18, GLN15, GZW15, GWY15, Hua15, HYW^{+}15, JS16, KCK16, KG16, KFHH15, LJD17, LHLL19, Pis18, PRC^{+}19, QLLY18, QZL^{+}17, Sar17, SSG^{+}15a, SM18, TSA15, UKK17, WSZ16, Wei17, YCH18, YX19a, ZY15c, ZZC19a]. discount [SCB15]. discounted [BCd15b, CLW18]. discouragement [SJ19a]. Discrete [DSB15, GG15a, HASC19, HV19b, LLC19, Pai15b, ZV17, AADLF15, AK15a, AO19, AD17, AD19, BH15, CÁGGLP16, CCNQ19, CHRD16, CCC19, CZS15, CL19c, DN15b, DN16, DS16a, DRS16, DLjZ18, DKBS15, DHS19, Du17a, Fan15b, Fau17, FLL15, Fen15, FXXC19, FLL17, GW16, GHQ17, GWK15, GZW15, HYW^{+}15, HYZ^{+}18, IB15, IR16, JJM18, KBR15, Kim15b, KWT15, Li15a, LSCK15, LW16, LCY16, LR16a, LX16b, LW17a, LWL^{+}18, LLZD18b, LX19b, LSZ^{+}19b, LLZ^{+}19, LZG16, LLL18c, Luo19, MZ15, MFJZ15, MC15a, MJL16, MZ17, Mah17, MMMMB17, MRQZ19, MS18b, Mon19, NR16, Nak18, NMNTBXVD15, NT16, NPT19, NSM16, PR15b, PTBN19, SSZ^{+}18, SSCX18, SLLC19, SFSR18, SDP16, SMZ^{+}17, Tan17a, TZD^{+}19, Upa16, WXZ18, WJZF19, WZZT19, Wu15, WV18, XPZ18, XZCC15, Xu15c, XZHX17, YWL15, YLL15, Yil16, YZ16, ZLDZ16, Zha15e, ZZL16a, ZSZ17, ZCW19a, ZG15a, ZZQ18]. Discrete [Boy15, SH15b]. discrete-approximation [Mah17]. Discrete-time [DSB15, Pai15b, AADLF15, CÁGGLP16, CL19c, Du17a, Fan15b, FXXC19, GHQ17, IB15, Kim15b, LSCK15, LCY16, LX16b, LW17a, LLZD18b, LX19b, LLZ^{+}19, LZG16, Luo19, MFJZ15, MJL16, Mon19, NR16, Nak18, NT16, SSZ^{+}18, SSCX18, SLLC19, SFSR18, SDP16, SMZ^{+}17, Upa16, WJZF19, XPZ18, XZHX17, YWL15, YZ16, ZZL16a, ZCW19a]. discretely [AM15b]. discretisation [BR17]. Discretization [GXH18, Ort15b, ABS15, ASS16, AK18b, BB18a, BFR18, DO18, GSV18b, KKS^{+}16, LT18a, MKS19, Por18, Pöt19, SR19, TM19, ZY15c]. discretized [CJN19, PS18a]. Discrimination [DSM15]. Disease [RS19, VEF19, ADY16, CDHK15, GRS^{+}15, LWPL16, LXJ^{+}17, MMG17, QTXY16, Sah15, WCH19, XWDW17, XZ15, ZHW16]. diseases [ARV18, BMT19, FJ19, TMM18, Tor16, Vil16, ZLJ15, ZSW^{+}17, ZG18]. disjoint [BLNSF17, GSV15, GSV18a, LYW18, MY19, Mao17, QY17, Yor15]. disk [Che15f, DS15, WJL19]. disk-beam [Che15f]. disks [BM19c, NSR^{+}19]. dislocated [SGY^{+}15]. disordered [WCZY17]. disparity [TES15]. dispersal [GZD19, Luo19, RWY19, WYZ^{+}17]. dispersals [XYDL16]. dispersed [KA15c].

Dispersion [BRV18, HGH19a, KKG18, QA19, SRP17, Zha18b, ZBH^{+}15]. dispersity [GWPA18]. dispersive [BMRS18, GLL15b, KR16, QBNF19]. Displacement [WE19, AM15b, GZW15, LR17, XS19b, YCH18, ZZ19a]. displacement-potential [AM15b]. displacements [JH17]. dissipation [GS17c, JL15, Kud16, LFY18, Pan15, Sha12, ZFL18, ZSXW18]. Dissipative [SRSZ15, WLH^{+}18, ANU16, CGZ17, CCD18, JIS15, KML19, LKZ18, XZ18, XZZ^{+}18]. Dissipativity [LXC^{+}19, SSKA16, DHX^{+}19, JSW^{+}18, NR16, SSZ^{+}18, WHWW19]. Dissipativity-based [LXC^{+}19, DHX^{+}19]. Distance [ACS15, CS16b, CXL16a, GPRCTRY19, LL16b, AH18b, CMM17, DGNA15, DV15, DESS15, Gat18a, GDZ18, Gut16, GFD16, HP17, HZL18, ID15, II17, JHL19, Jur17, KS19a, KMA16, LR19a, LWM17, LWW19, LD18b, MS15c, ROGS16, TDS15, WL17, XX19a, XLS18, YwFW^{+}19, ZL16b, ZDJ19]. Distance-based [CS16b, CXL16a, GPRCTRY19, DGNA15, GDZ18, LD18b]. distances [CCLW18, GTBC18]. distinct [HTW19]. distinguish [FK18b]. distinguishing [BRM^{+}19, XLG18, YSY^{+}17]. distorted [ZSY18]. Distributed [CLZW19, GSD16, HWY15a, HW17b, KMZ19, LLZ^{+}19, LZM^{+}18, SM17, SSCX18, ZLM^{+}18, ZYLY17, Ali15, AM18b, BL15, BL18, CGCK19, Can15, Cao19, CZS15, CWLA18, DAQI15, DQ18, HWY15b, KBR15, LJ15a, LL15b, LZjy18, LWL^{+}18, LTQF19, NND^{+}18, PS19, SHR18, SY18a, SWSW19, TZD^{+}19, Wan17a, WLZ16, WGLZ16, XH16, XCL18, YWZS18]. Distributed-observer-based [LZM^{+}18]. distributed-order [AM18b, DQ18]. Distribution [MR19, MS15c, BEM15, DDM16, DD16b, DWT17, DL15b, GdS18, GSDH15, Gat18a, GST19, IAK15, JS19, JPY16, JDS18, LXJ^{+}17, LJHA18, Liu19, LO19, LD19c, LTZ^{+}17, MRS18, MSdC^{+}18, NI19, ÖAD15, ÖÖAD15, PVV^{+}15, SJ16, ZY18b]. Distributional [LL17b]. distributions [CLL^{+}15, Zha19c]. district [BES19]. Disturbance [GWSC16, KSM^{+}19, Wan19a, XGQW19, YLP19, LDSD19, LNWS18, LMZ^{+}19, zMnZyZ15, QKGW18, YmZHC15]. Disturbance-observer-based [XGQW19, YLP19]. disturbances [CWL16, CPCQ17, DLXZ19, FW18a, GhDwZS17, HW17b, LSMT18, LLX^{+}19, NPT19, PP18b, TCS15, WXXY19, ZZL17, ZJYD17, ZWO^{+}19]. disturbed [FLYG18]. Div [EL19a, EL19b, EL19c, dFGAN19]. divergence [CDK15, JSM16, Kem16, PMS16, Zha15h]. divergence-free [Zha15h]. Diverse [HYZ^{+}18]. Diversity [Moh19, ATV18]. dividend [ZH17, ZXD15]. dividing [TH19]. division [Kim18]. divisor [HHZ15]. Dixit [TES15]. Dixon [QYF^{+}18]. DMLPG [ID19]. DMP [DY15]. DNA [DLS19]. do [Ber17, BF16a]. Dochev [Pro16]. document [GHZ19]. Does [SLS16, SSGG17, Tan17b]. domain [AYAA^{+}15, AG15a, AL18d, ARK15, BNY15, BK18c, BRD15, BZ17, BL15, CL19a, CDL18, FLS^{+}15, Gao17b, HMW16, KL16, KK17, KMP15, MSHG16, PWS17, SLLC19, SHHC18, TPS15, XCHW16, YWZ18, ZH15b, ZXC19]. domains [AG16, Bea15, BB19a, BDE^{+}15, BÁR15, BVJ15, DA19, EMM17, GS17a, Gia15a, GOA19, HAC18, LB15a, OR18a, WHSS15, WWH17]. domestic [Haf17, MPHG15]. Dominance [HLFZ19]. Dominance-based [HLFZ19]. Dominant [DM18, BM19d, LZZT18, YLZT17]. dominated [CGW16, MW18]. dominating [GHM18, LYW18, SRGW15]. Domination [TZC19, BF16b, BS19c, CMSZ18, DH19a, Du17b, MCY19, SRT18, WWL18, YZYgG18, YWLL18]. dominations [HLL18]. donation [CCW17, CCW19]. Doob [ALL18a]. door [WWC^{+}15]. doped [ZLH16]. dose [SNS^{+}16]. dot [AÇ18]. Double [JBS15, LK19b, PW19a, Pet15b, ZWZ19a, BS19a, GKN15a, GKN15b, HM19a, HW17b, JS15, JLX15b, Kia15, Lam15, LZM^{+}18, LLZ^{+}16, LYZL18, LXZL19, LX19d, MLSX18, MGK^{+}16, PHRB18, SJ19b, Sun15b, YDWZ19, YG18, YWLL18, ZXZB19]. double-diffusive [MGK^{+}16]. double-implicit [YG18]. double-integral [LZM^{+}18]. double-integrator [HW17b]. double-Newton [GKN15a, GKN15b]. Double-step [ZWZ19a]. double-subpopulation [JLX15b]. double-walled [Kia15]. doubly [GMOP18]. Douglas [BCH15]. down [Alk19, LD17, LLZD18a]. down-and-out [LD17, LLZD18a]. downhole [DWT17]. downside [SBLZ17]. downward [PHRB18]. DPSS [HMX15]. DQ [Tom15, Tom12]. DQ-based [Tom15, Tom12]. drawings [OFwCW19]. Drazin [CLT18, HLT15, HLT16, Kyr15, MD15c, Mos16, MD18b, RMSD16, SS15a, TB17b, Vis15, WMS17]. Drazin-inverse [TB17b]. drift [KPP18, MZS18, SS18b, TY18, WXM^{+}18, YG18, ZZL18b]. drift-implicit [YG18]. drilling [MTB18, ZWM16, ZHWW15]. Drinfeld [SKBR18]. drive [XXZ^{+}18]. drive-response [XXZ^{+}18]. driven [AYAA^{+}15, BS15a, BLJ^{+}19, CS17, DWS19, GG15c, JMK17, Kim15a, Laz17, LOOGB18, LXTF18, LSLL19, Liu19, MDZ15, Maj18, MZM15, MP18, PRMC19, RWY19, SK15, SS18c, SP15, SZPH16, TM19, TB17a, TLZ17, WJ19, WY19, WLL16, XS19a, XPG15, XDG^{+}19, XL19, YZ18b]. drop [BBGR15]. droplet [FS16, KRF17]. dropout [WSH^{+}18, XCW19]. dropouts [HZK^{+}19, HXW^{+}19b, LCY16, SS19b, TPL16]. dropping [SDP16]. drought [lZA^{+}19]. Drude [QBNF19]. drug [EA15a, GNR^{+}19, MSM15]. drug-eluting [MSM15]. Dry [ZCW^{+}19b, LZ19c, MO16]. drying [ZKV^{+}15]. DS [LXJ^{+}17]. DS-I-A [LXJ^{+}17]. DSDSA [GS19]. DSG3 [NMNTBXVD15]. DSMC [FFG^{+}16]. DTM [SSB18]. dual [CK17, CW18, DJL17, GG19a, JLL16, JPTL15, KHN19, PMSB15, QCH19, RD15, RCTZ18, YP18]. dual-channel [PMSB15]. dual-CISTs [QCH19]. dual-phase-lagging [DJL17]. duality [JG17]. duals [MK15c]. duct [HG18]. due [FTE19, Far18b, HRS17, KKG18, Pon18, XY15b]. Duffing [LR15]. Dufour [PMV16]. duopolies [MS15b]. duopoly [ANP16, AEP15, AAA15b, AAA16, CMG18, Els17, PL15c, Ued19]. duplication [MB15b]. during [CZ18b, Kim18]. Durrmeyer [AIP15, DD16b, GG15b, GAS17, KA15a, KIAG16, UDA15]. dust [Mis15]. dust-acoustic [Mis15]. dusty [CS19, HRS17, MSI17, SBH17]. Dwell [LZLW17, HXW^{+}19b, LW17a, XXW^{+}19, ZLLZ16]. dwell-time [HXW^{+}19b, XXW^{+}19, ZLLZ16]. Dynamic [AAA15b, BLA15, CCQS16, GLYT16, LYZL18, LMMR15, MN16, NS17b, TPS15, TPL16, ZLJ15, ABL15, ALP18, AH15, BK18a, BAE15, BCd15b, CXP16, CY15c, CST^{+}15, DWZ15, FHV15, GWC^{+}16, GXH18, Gon18, Hac17, HO15, HRS15, HI15b, Hua18, JAN16, KWH17, Kar19, KSM^{+}19, LMZ^{+}19, LGL19c, MG19, Pac19, Pai15b, PSSX17, RZR16, Reh18, RXQ19, RS15d, SZ19b, SRBA15, SH15c, TT18, UMF19, WR15, Wan17a, WA15a, WCH19, WRX^{+}19, XCY18, Yua15, ZWK16, ZCW^{+}19b, ZWL16]. Dynamical [ADMV19a, AADF18, HP15c, LJYH19, LXZL19, PRMC19, SSZL17, VLCT18, ZMA^{+}18, ZYC18, AY16a, BGRT15, BM18a, CCTV15, CL16c, ES17b, Fan15b, GG16, GDL16, HST17, LHS18, LO19, LHX15, MWRT17, Moh19, Mon19, MNJM19, RVGS17, SSKA16, Slo15, SS15c, SLWX15, WDLL15, WY18, Wan19a, WHWW19, WZZT19, Wu15, bWzZjH16, XLJC18, XCCL18, YWZS18, YYS17, ZZL^{+}18a, ZWHC15, Zha15a, ZW15b, ZZ16b]. Dynamically [NB15]. Dynamics [AMN19, CCTV16, DLJW15, DHMT15, Els17, FG15, HJ18, JAU15, LC15d, LB16a, LC16b, LYM18, PW19b, Pu15, SP19, SK19, SZJ19, SZT17, XD15, YW19a, ZYLY19, ANP16, ATN^{+}19, AA16, ARR16, AM15f, BCST19, BCMT15b, BCMT15a, BMT19, BRM^{+}19, dCCGV15, CMG18, CSWJ15, CWD^{+}19, CM16, DSB15, DWTT17, Du19, DZ17b, DZ17a, EA15a, ET15, GKN15a, GKM16, GKN16, HD17, Hsi15, HXS^{+}16, HZCY17, IT15, JGGK15, KZL19, Kha17, KCL19, KSA19, LKM17, LZZL15, LZL^{+}19b, LW19a, DDF18, LXT17, LZ17d, fLLPY16, MM15b, MTMB17, MM16, MS15b, zMnZyZ15, MPB^{+}17, OLZ19, PRL^{+}15, PL15c, PSP17, Pöt19, QLF15, RRS^{+}18, SSB15a, SA16c, SSGG17, SGM^{+}16, SY18a, SRAC16, SG17, SSM^{+}18, SZFS17, SG19b, SWA^{+}19, TMM18, TBSR16, TRRP15, Vid18, Wan15a, WRW16, WLZ16, WGLZ16, WHC18, WD19, WC17, WV18, XLJC18, XWDW17, XJW17, XGZ17, XXZ^{+}18, YSLS17, YLZT17]. dynamics [ZLZ^{+}18, ZS15a, ZWY19, ZJ15b, ZLC^{+}15]. dynamos [WZY^{+}19]. dysentery [BMT19].

e-commerce [BM17]. e-learning [DCA15]. early [SRK18]. earthquake [FMRRT15, GKDHR19]. Ebola [ADY16, DCK^{+}18]. eccentric [WC18, XX19a]. eccentricity [Du17b, HZL18, LZ19b]. echo [PAM19]. Eckhaus [WTSQ15]. eco [CDHK15, FHPV18, LPZL16, MLG18, Sah15]. eco-epidemiological [CDHK15, FHPV18, MLG18, Sah15]. eco-epidemiology [LPZL16]. ecological [CB19, Moh19, dSR15]. economic [HS15c, LTC16, MMB15, TSA15, WH15, ZZH15]. economics [LZL19a]. ecosystem [HJ15, QTTL19]. Eddington [HLC^{+}17].

Eddy [TT16, CLR15, KCZK15, KCWK15, KÇ18a, KM17a]. Eden [ADMV19b]. Edge [AV15, LX17, PJ18, QY17, WWW18, AC19, BM19b, CYY19, CXL16b, Che18b, CMM17, Far18b, GFWH17, GHZ19, HLX17, HHT19, KST15, LLM^{+}18, Laz17, LY15a, LYS17, LWSM16, LT19, Mao17, PP12, PP15b, SY18a, WLQZ18, WZ18a, WWW19, XYZ18, YFLZ18, YL18, YW18b, Zha19a, ZM18a]. Edge-based [PJ18, LLM^{+}18]. edge-bipancyclicity [LX17]. edge-colored [CMM17, XYZ18]. edge-coloring [WWW19]. edge-colorings [LT19]. edge-connected [ZM18a]. edge-connectivity [BM19b, YFLZ18]. edge-directed [LY15a]. edge-disjoint [Mao17]. Edge-fault-tolerant [LX17]. edge-neighbor-scattering [LWSM16]. Edge-partition [WWW18]. edge-preserving [GHZ19]. edge-rupture [LYS17]. edge-Szeged [WLQZ18]. edge-version [GFWH17]. edge-Wiener [CXL16b, KST15]. edges [QGH19, Rad17, SLW16b]. edit [DV15]. Editorial [Ano18o, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18p, Ano18q, Ano18r, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v]. editors [RSGFV16]. EEG [BMDS17]. Effect [HZS15, JAU15, KK15c, LY18c, MB15b, Pan15, Sha12, Ued19, ATK^{+}16, AO15b, ATN^{+}19, DTM19, DSB15, ES15, FG18, HD17, Har16, HM19a, HYJ19, LLY19, LJG^{+}19, LXD19, LFY18, MSI17, MMS17a, MS18b, MRR17, Mou15, MKN15b, PHRB18, PKHD18, PZ16b, Pon17, QZL18b, RHJ18, SRP17, SSG15b, SWSS18, VD16, WQ15, XP17, XY15b, YYX^{+}19, ZBKS16, ZHW16, ZCLS18, ZCFW19, ZL19c]. Effective [PGLG17, KMBA18, QW19, TMM18, jXlZX18]. effectiveness [BMT19, WSXR19]. effector [DDD19, SRJ^{+}15]. Effects [AAM19, dCCGV15, GST15b, KCL19, LZP18, LJR19, QTXY16, SH18b, GWW15, JZ15b, KG15, KG16, LSR18, LX19b, LWLM19, MGK^{+}16, MS18a, PMV16, RZSS15, RNM15, RMD18, SMR15, SBH16, WF15a, WXP16, ZLH16, ZML^{+}17, ZL19a, ZL19b, ZWZT18, dSVR17]. efficacy [EA15a]. Efficiency [GJNRG15, FXC17, LA15, LMMR15, VS19]. Efficient [BKW18, Bog18, CCY19a, CSV19, CJ18, DZW15, DG18a, EMO18, JO15, Kac15, Kan19, KM15a, KDTP18, Kue15, LWQ18, LW19a, LX19a, LYDH18, MA15c, PS15c, QTF^{+}16, Sku15, Tan15b, Tan15a, Tu17, WNN16, WGL15, XMX15, ABC^{+}16, AQ18, AS19a, AJC^{+}18, BAM16, CJ15, CWL19, DA19, Deh16, DCS^{+}19, EA15b, FWLC15, FCW^{+}17, FP18b, GTBC18, GN19b, How15b, How16, HLL18, KG19, KD18, KBM16, LPZ19, LH19a, LLL18d, LS17b, MGL19, MD18a, MS17b, MP19a, OW18, PSZ18, RSKB16, RP18, RB15, Rub19, SBV15, SHC^{+}17, SKBR18, SGKR19, TKDF15, TTD18, USCA15, WMLG17, WJWX18, jWgX18, Wu17, WLFH17, YZZ18]. effort [MS15c]. Egalitarian [Pla16]. Eggenberger [DDM16, DD16b]. egress [GRV15]. Ehrlich [Pro16]. EIC [YZZ16]. EIC-graphs [YZZ16]. eigenfunctions [Deh16, KKT19]. eigenpairs [GG15a]. eigenparameter [Tha15, ZjA19]. eigenparameter-dependent [ZjA19]. Eigenvalue [BM19d, BFIJ15, LH17a, LQ15, LJ15b, AK19a, BFJ^{+}17, CSZ19, CEP15, Gia15a, Gia15b, GGHO18, HKL15, HLZ15, HYB17, HM19b, Ili18, JYSR18, KG18, Liu15g, LLL17b, LPW15, Miy17, NJTP19, OA15, Pan18, SWL16, SD19b, TTP19, WS19b, WD15, XL18a, XBC19, YKZ15, YFY17, YZYgG18, ZY19, ZwX16, ZCEW18, dCSS17]. eigenvalue/eigenvector [GGHO18]. Eigenvalues [Won15, dF15, CD15b, Deh16, HTW19, JFT17, LGS^{+}18, LLL19a, Lin15, LLTW19, PP18b, RG18a, Tha15, WDB18, YY15, ZW15a, Zha18a, dFKMM15]. eigenvector [GGHO18]. eight [FKV16]. eighth [BAM16, CN15b, CN16b, IH15, KKB15, NCS16, Sha15e, SA16c, SA16d]. eighth-order [BAM16, NCS16]. Eikonal [DO18]. Einstein [ID19, JJ15a, QFS17, WJWX18]. Eisenhart [ZH19a]. either [Mah15]. EIV [ZZ15c, ZZ14, ZZHT17]. EJIIM [BW15a]. elastic [AZL15, AJC^{+}18, BKKN16, CD16b, lDqCWzH18, FFL16, GT15b, KFHH15, Lia15b, MC15c, NB19b, Rad15, RM16, SM15c, Ton18, TSPF18, WL16a, WE19, ZLL15b, ZZC15]. elastic-fusion-coupled [ZZC15]. elastic-viscoelastic [FFL16]. elastically [Kia15]. elasticity [AL18a, AAGZ19, BLK19, BCG19, cCpSkZ15, Che19, MC17, SCY19, Tch15, WHSS15, WZWD15, ZZLY15, ZF18, ZGZ16]. elastodynamic [DWRZ17]. elastodynamics [AJ17, HSS15, JLL16]. elastomers [Sin17]. elastoplastic [CSV19, NND^{+}18]. elastoplasticity [MHCF15]. electric [Hsi15, LW15b, VVA^{+}15]. electrical [ASKK19, LW18a, MNJM19]. electrically [BLJ^{+}19]. electricity [MPHG15]. electrified [Alk19]. electro [FW18b, PPS25, RRR19, RRR20]. electro-osmosis [FW18b]. electrochemical [KSKW^{+}19]. electrohydrodynamic [RM19a, RMK19, SKS17]. electromagnetic [BRV18, Fau17, KSD^{+}15, MMZ^{+}17, WMZ19]. electromechanical [PP18b]. electron [Wan16b]. electron-acoustic [Wan16b]. electronic [CNK^{+}18]. electronics [LPS19]. electrorheological [WF15a]. elegance [MD18c]. Element [QDH^{+}18, AGZ15, Ade19, AJ17, ALZ16, BB18a, BY19, BdS19, BRV18, BTWY15, CKG19, CH18b, CLR15, CD16b, DFJ16, DRS18, GS17a, GM15a, GWK16, Gia15b, GWK15, HYB17, HL16, HZLY17, HLJ18, Hua15, IR16, JMW15, JAN16, KCZK15, KM15a, Kok15, KMZ19, KK17, Laa18, LX16a, LZL15a, LCW15, LJD17, LCL17a, LH17c, LZ18b, LML18a, LD19a, LL19b, LZ17e, LS19b, LT18a, LZ17f, MC15c, MKN15b, NRS19, Pan15, QZW19, QFTT15, RRW^{+}15, SA16b, Sha12, SWL16, SLW16a, SW17b, SW17a, SZL^{+}18, SY18b, SLL19, SJW16, SFC^{+}17, SW17c, SDWG19, SH18b, TM19, Ton18, UKE15, WZM15, WQ15, WZWJ18, WA18, WZF^{+}19, WSC19, XJW17, XLR17, XS19b, YZ18b, YCH18, YJK19, YZJ18, hY18, YCLS19, ZL15a, ZLL15b, ZZZ16, ZT16b, ZL17a, ZZ17a, ZZ18b, ZF18, ZC19b, ZZ19a, ZBH^{+}15, ZY15c, ZCZM15, ZZC19a, ZS19b, ZCPM17, ZSF17]. element-characteristic [YCLS19]. Element-free [QDH^{+}18, LZL15a, LCW15, LCL17a, LD19a, SW17c, WZF^{+}19, XJW17, hY18, ZLL15b]. elemental [SNS^{+}16]. elementary [CS19, JGBY16, Rui18, RLH15]. elements [ABGS15, AGZ15, AAGZ19, AJ17, AM15c, AV15, AP18a, cCpSkZ15, DEdlT17, HK18b, HMT15a, KM15a, KG16, KAMD18, KK16b, LY18d, MBHS18, MM17, MD15b, NKP16, SST18, SCY19, ZZCY16, ZS19b, dFGAN18, dFGAN19]. eliminating [ACRZ18]. elimination [APS16, SHC^{+}17, She18]. ellipsoid [RS17b]. Elliptic [Luo15, Cov19, DH15, Gia15a, Gia15b, HL16, HZLY17, HM19b, Khl18, Kue15, LL15d, LWZ16, LML18a, LMMZ18, Mah15, ODP17, SZ16, TTK15, WZ15b, YCZ19, ZZZ15b, ZT16b, ZQG16]. elliptical [Bea15]. Elman [HWBL19, KŚK^{+}18]. eluting [MSM15]. Embedded [KGS^{+}15, ARK15, FKR15, Jai18, MSM15, Pan15, Sha12]. embedding [ZCML19]. embeddings [RC16]. EMC [NKKS16]. EMD [LLz18]. Emden [IAW18, KF15, MFMK17, SK17a, Tsi19]. Emergence [CZ18b, CCW17]. emergencies [TD19]. emergency [CFS18, CZ18b, WWC^{+}15, ZLLP18]. EMI [LLz18]. emission [LWZ19a, LXZL19, MNYL18]. emissions [ZWY19]. emitting [RMD18]. emotional [LWGS17]. Empirical [YYP18, IR18, LQZ^{+}18, QZL18b]. Enabling [KK16b, SHHC18]. enclave [SI15]. Enclosing [DCT19]. enclosure [MKN15b, PBP19]. enclosures [WW16a]. encompassing [RWH15]. encryption [FMB^{+}18, FMB^{+}19, GMOGCCC15, LZKX19, SGFCRM^{+}18, WDG19a]. end [DW17]. ending [NSN^{+}18]. endometrial [HWBL19]. endpoint [Boy17]. enemies [GGM15]. energies [BM16, BHD18, CL15a, CL16b, LQW^{+}15, MMG16, Wan16c]. Energy [BBCI18, BCR17, BCI15, FM17a, WYL15, AG15b, BES19, CWG15, CLY17, CCY19a, CvRvdM18, DMG15, DM16, DMG16, DGF17, FXC17, FKA17, Far18b, GT15b, GWX17, GR15b, HFK19, HP16, IB19, Jah17, JJ16, KS19b, LZ18b, LJL18, LH19a, LW19a, LPHH15, LWS18, MDZ15, MDHD18, MLT17, MSdC^{+}18, MR16b, MRS19, QZL18a, RGH^{+}15, SG19a, SLW15, XLYZ19, YWS18, YW18a, Yos19, ZLX19, ZCCL18]. Energy-based [WYL15]. energy-conserved [CWG15]. Energy-conserving [BBCI18]. energy-dependent [GR15b]. energy-preserving [CLY17, CCY19a, LJL18, LH19a, LW19a, LWS18, MDHD18, YWS18, ZLX19, ZCCL18]. engine [MZ15]. engineering [GH17, NCEGA15, WWR^{+}15]. Enhanced [VK17b, APMM18, BJM16, KSL19, RM16, SNAR15, YLW19]. enhancement [FXC17, LJR19]. enhances [BR17, SLLZ19, WWW^{+}17, ZSL17].

Enjoying [CBGGV17]. Enlarging

[AG15d, MSHG16]. ENO

[Liu17a, RR18b, SdA18]. Enrichment [NB19b, AGZ15]. ensemble [FTF16, VK17b, ZCL18, ZLK^{+}19]. ensemble-based [VK17b]. ensuring [CS15b]. entire [BLNSF17]. entrance [ZWL16]. entries [LH17a, SS15a, SS17]. Entropies [CDK17a, Ras15, CD15a, CDS15, GDZ18, ID15, LQW^{+}15]. Entropy [GIT15, CH18c, GWK16, GDRP^{+}19, GST15b, HYJ19, RC17a, WTD^{+}19, YSF16]. entry [GCH^{+}19]. Enumeration [FZZ18, Yan17]. envelope [LLz18, PK15]. envelopment [SY15a, TB15]. Environment [GSD^{+}17, DKDD15, JAU15, LZT^{+}18, LZ19a, QZJY15, SPP17, YYZ19, ZSL17, ZYZ15, ZZL16b]. Environmental [CJG^{+}18, MKMC15, SK19]. environments [GGM15, HHL^{+}19, LC15d, WLS15, WXP16, ZW15c]. EOQ [JYVS15]. EP [MS19a, MD15b]. Epidemic [WYZ^{+}17, ADY16, BW15b, CKW17, CJG^{+}18, CCYT19, CJ18, DTJ18, DAQI15, FLMC15, ELP^{+}18, GLM16, KM15d, LLL16a, Li15b, LZT^{+}18, LYZ19, Liu15c, LZ15c, LC16b, LJS18, NY18, QLF15, QZJY15, SJ16, SZFS17, Wan17a, WLZ16, WWK19, ZLZ^{+}18, Zha15e, ZG18, ZWZT18, ZWL^{+}19]. epidemics [PRL^{+}15, WGSX19]. epidemiological [BP17a, CDHK15, FHPV18, MLG18, Sah15, SKHA18, SHEBS19, XWZ15]. epidemiology [LPZL16]. epilepsy [PSJJ19]. Epistatic [FTF16]. EPQ [MBM16]. equal [Rad17, ZS19b, DGLE17]. equal-order [ZS19b]. equalities [LLP17]. equality [GZ15a, WCZ15, ZZLZ17, ZZ14, ZZ15c]. equals [AK18a]. equation [AKAA15, eMA15, eMBHA15, Ach19, AM18a, ABS15, AM15a, AAA^{+}15a, AM15c, AG15a, ACE19b, AAS^{+}17, Ali15, AR15, AHY19, AL18d, ACG19, Ata16, AO15c, AHK^{+}19, BCO15, BAV18, BMO15, BBCI18, BT15b, BAB18, Bas19, Bea15, BS17, BDGS15, BC15, BRD15, Boy17, BCI15, BC16, CVJ19, CCLS16, CLY17, CFR18, CCNQ19, CLX17, CJN19, CZP15, CGL15, CLJ^{+}15, CCD18, CC18b, CXL18, CCY19b, CG19, CCZ15, CJK17, ÇT16, Cov19, Cui15, DS19, DT16, DBK15, DM18, DFJ16, DWTT17, DSSC19, DQ18, DC16, DO18, DN15c, EM16, EO18, EJF15, EJF16, ES17a, ET15, Fan15a, Fan18, FCM17, FZLT15, FLL17, FM18, FLL^{+}19, GGL15, GM15a, Gao17a, Gao17b, Gao17c, Gao19, GMS19, GGHL15, GAB17, GSP15, GS15b, GN19b, GVA16, GT15c, GLL15b, GfSmL16, GR15b, GR17b, GB19, GM19, HAC18, HJMB18]. equation [HA16, HEA15a, HEA16, HLZ15, HK15, HW18, HPN15, HLIK19, HP15a, HAY19, HAH19, HGL19, HL15b, HLLT15, HGM15b, HXM18, HCY15, HYW^{+}15, Hua16, HC18, HGH19b, II18, IS17, IT15, cJzS15, JZ19a, JLY18, JS19, JSLW17, JRM15, KKT15, KRB15, KCK16, KZ16, KAU15, KS16a, KM15b, KK16a, KXZ19, KD16, Kor18, KNT17, Kri18, KSV15, KR16, Kud16, Kud19a, KK17, KC18b, KMP15, LJZ18, LL19a, LP15a, LRF^{+}18, LSfW15, LJL15a, LL15e, LG16, LL16a, Li17b, LG17, LZ18b, LJL18, LW18g, LM18, LZCL18, LLTL18, LR19b, LZ19d, LH19b, LJL19, LSZF18, LWW15a, Lin19, LLG^{+}15, Liu15a, LSW16, Liu16a, LCL17b, LDH18, LLL18c, LW18h, LWS18, Liu18a, LL19d, LG19, LZZ^{+}19, LR15, DPL18a, LFY18, LXZ19c, LZ19g, LM15, LHWG16, LT18b, MDH16, MZM16, MZ17, MM04, MS17b, MMIA19, MVS19, MVRV19, MMV16, MS16a, MR16a, Mis15]. equation [MD17, MMK17, MO15a, MS15d, MTJ15, MA15c, Mus18, Nag17, Nam19, NAM18, OPP18, OW18, ÖK19, OB18, Pac19, PKB15, Par15, PW19a, PG15b, PTBN19, Pi15, PJ15, PVL19, PFA15, PeO18, Pu15, QXW19, QDD15, QA19, RAH19a, RG15, Rem18, RZW18, S.15, SWUE19, Sab18, Sae17, SR15b, SJD^{+}18, SDS18, SLC18, Sar17, SR17, SSA19, SR15e, SR16b, SzSC19, SS19a, SLW16a, SY17b, SW17b, SW17a, SZL^{+}18, SPG17, SJW16, SPS16, SPST17, Sin18, SKBR18, SAT15, Sun15b, SW17c, Sun18, SE18, TES15, TSJ16, TZTG17, TLHC17, TTLX17, THM19, TTXZ16, TTK15, UKE15, UBT18, VD16, VD15, VK19, VDV15, VS17, WW16b, WNMC15, WTSQ15, Wan15d, WTXZ16, WMLG17, WTJ17, WWH17, WA18, WXW18, WW18, WDG19b, WJL19, WX16, Wei17, WWMX16, WMSP19, WD16, XCHW16, XX19b, Xu15c, YMD15, YL17, YMS16, YWS18, YLL18]. equation [Yan18b, YW19b, YWRP16, YWZ18, YB15, YZ15b, YZ15c, Yüz16b, ZV17, ZNV19, ZEF19, ZB17, ZP16, ZA16, ZWF15, ZJZ15, ZH15a, ZZC15, ZY15a, ZM15a, ZSX15, ZLWW15, ZWLZ15, ZMQ16, ZZY^{+}17, ZYY17, ZZ17a, ZZY18, ZCXX18, ZZG18, ZZ18c, ZJY18, ZXZB19, Zha19c, ZCU15, ZZL18b, ZHW^{+}18, ZMH18, ZEW19, ZY15d, Zho15b, Zho15a, Zho15e, Zho17, ZW15e, dlRGB16]. Equations [DGRR16, GSV18b, AB15a, ABC^{+}16, AED15, AF18, ALL18a, AK16a, AR16a, AÖ15a, ABLZ15, ABL15, AÖ16, AZL16, AOH17, AHO18, AN15, ANT^{+}16, ASHSC17, AL18b, AD16, ARB19, AEB18, AZ18, AH15, ARL15, ASA19, AZ14, AAKT15, AIG17, AST18, AM18b, AP19, APL15, ALP15, ARSV15, ABGL17, AIIK15, ACSS15, AR16b, AG16, AKKB17, AM18c, ABMA19, AP15, AK19b, AM15h, AE17a, AD17, AD19, AAMC19, AAH19, AKA19, AZ19, Bab16, BD17, Bag19, BÇS15, BW15a, BEM15, BM15a, BS16, BBF^{+}18, BE19, BB18a, BD18, BAM16, BSZ16, BB15c, BB15d, BDSS15, BB16, BB19b, BdS19, Bog16, BKLW15, BGST16, BFR18, BH15, BHM15, BJ19, BS19b, BdlRT18, BC16, BS15c, BR17, BK18d, BCD^{+}19, CT17a, CWG15, CS16a, Can15, CP15a, CP19, CPST16, CDK17b]. equations [CÖ15, CLS15, CDJ19, CJ15, CKW15, CL15b, Che16c, CZL18, CH18b, CXL18, CWL19, CY15b, CDL19, CXW18, CB17, CDL18, CN15c, CN19b, CY15c, CLMT15, CNQRR18, DCW15, DK15b, Dar16, DN15b, DN16, DTM19, DA17, DA19, DL17b, DGZ16, DOR18, DZW15, DGZ15, DWZ15, DCS15, DLjZ18, DL18b, DK16, DLQ17, DFP15, DCR15, DZ15, DJ15, DPSS16, DZL15, DNC^{+}17, lDqCWzH18, DO15, DLCT18, Du15, DC16, DF16, DB17, DJ18, ER15, EHS17, EAAM^{+}15, EGB15, Ers16b, Esl16, EMM17, EA15b, EE18, FP18a, FK17, Fan17, FZLZ19, FMX15, FWLC15, FB19, FY18, FY19, GRV15, GY15, GS15a, GW15b, GDL16, Gao16, GML18, GHL19, GLM19, GWK16, GD15, GK15, GZK16, GQC15, GS17b, GFG15, GBJ19, GN15, GR15a, Gra15, GJZ19, Gua19, GE16, GZ15b, GS17c, GG19a, GS11, HQ15, HEA15b, HO15]. equations [HFCS16, HRS15, HV19a, HARS15, HSAR19, HMF15, HJKP18, HL17, HH15, HS18, HS15b, HV19b, HMT15a, HZLY17, Hsi15, HM18a, HH18b, HM18b, HHY19, HMSS19, HHG19, IAW18, JJ15a, JDG18, JS15, Jia15, JYS16, JT17, JX19, JSG15, JISI18, JAN16, Kan15, Kan16, KJ15a, Kar19, KE16, KM17b, KNBA15, KDH19, KAH15, KL15, KJ15b, Koc15, KK16b, Kri18, KDG19, KAS16, KF15, LdCM19, LQ15, LJW^{+}15, LLY15, LC15b, LCS15, LWS15, LG15, Li16b, LWZ16, LR16a, Li17e, Li17a, LH17c, LZCW18, LW18d, LZ18a, LL18a, LSR18, LLX18a, LK19a, LXZ19a, LLW19, LR19c, Li19, Lia15a, LY15e, LBB18, LWW15b, LMF17, LVV18, LFL^{+}15, Liu16b, LZH16, Liu18c, LZ18f, LMMZ18, LZ19f, LHZM19, LFC^{+}19, Liu19, LMD19, LWW15c, LRH15, LJC16, MHN15, MDZ15, MG15, Ma17a, MX17, MDHD18, MD18a, Mah18b, MRE18, MPW15, MR15a, MHM15, MZM15, MYHS18, MAG15, Mi17]. equations [MYH15, Mic15a, Mic15b, Mig15, Mig16, MNR19, MMZ19, Mil16, MH17a, MHZ18, MZS18, MH15a, MH15b, MH16b, MH16a, MH17c, MH17b, MS19c, Mis16, MSS15, MGMM17, MDZ17, MDH15, MK16, MKS19, MKN^{+}15a, MO15b, Mou15, NH17, NE16, NBK16, NB19a, Nas19, NRT16, NL18a, NI19, NCS16, NA19, NWNA15, NWN15, Obr19, OR18a, OS15, Oka18b, Oni18, OL19, Özb15, PS18a, PAE^{+}17, PZ17, PHM19, PMC15, PR15a, PWS17, PYZS18, PZ15, PST18, PRD19, Pol19, Por18, PBBF18, Pos17, Pöt19, PS19, PKS15, PKB18, PK18b, QLLY18, QZ15b, QTF^{+}16, QYF^{+}18, QZW19, Qiu18, RDLG16, RZD^{+}19, RWW15, RLZ19, RS15a, RC17a, Reh17, Reh18, RWH15, RzS15, RW17, RT18a, Reu15, RME18, RXL19, Rub19, RM18b, RC17b, SAC18, SuR15, SuRI15, SAA19, SR15a, SR15c, Sak19, Sap18, SMI18, SEKA16]. equations [SM16, SHR18, SN15a, SN15b, SH16, SSK15, SB15, SSB16, Sha16b, SA16c, SSA18, Sha15f, SM15d, SGM^{+}16, SC18b, SK17a, SJ18, SXS15, SS16, SGJS16, Sil16, SK17b, SPS18, SN18a, SN18b, SAS19, SRS17, SR19, SWY18, SY18c, SG15, Ste19, SDIS15, Su16, SM18, Sug17, SI19, SH15c, SzShG16, SX16, TMS15, TRP15, TMPG17, TZ18a, Tan15b, Tan15c, TZ19b, Tha18b, Tho17, Tom12, Tom15, Ton18, TBG16, TY16, TA16, USCA15, USCAAA15, Vab15, VRC15, VDV15, VDI17, WW16a, Wan15c, WZ15b, Wan15e, WZ15a, WHSS15, WM15, WL15, WZ16b, WFZ17, WSLC17, Wan17c, WZZ18, WJWX18, jWgX18, WXZ18, WXWX18, WDCJ19, WX19a, WY19, WR19, WZS19, WNN16, WN16, WAS19, WG15b, WX16, WY16, WL19c, WN15, WLFH17, WBL17, WSZ18, Xy16, XY15a, XY17, XM15, XHW18, XYG^{+}18, XW19, XPG15, XHW15, Xu17a, XL18b, Xu19c]. equations [Xu19b, XL19, XLW15, YÖ18, YDWZ19, YLLH19, YR15, YWO18, YGZ19, YTH19, YSL19, YYM19, YAIN17, YZ17b, Yil16, YG15b, YMMH15, YJSW16, YL19b, YMHW15, YNMW16, hY18, YL19c, Yüz15, Yüz16a, Yüz18, ZAP19, ZMBJ19, uDuI18, ZA16, Zei16, ZZ15a, ZBBW17, ZGGY19, ZS15d, ZSL15, ZS15b, Zha15f, ZAW15, ZY15b, Zha15b, ZWK16, Zha16d, ZMQ16, ZZCY16, ZGH17, Zha17a, ZSY17, ZL17a, Zha17c, ZJJ18, ZW18, Zha18b, ZKKZ18, ZMZ18, ZC19b, ZYSL19, ZS19a, ZC19a, ZRLX19, ZBH^{+}15, ZCQ15, ZSY15, ZZ15d, ZCX17, ZL17b, ZZX18, ZLX19, ZYX19, ZS19b, ZYJ15, Zho15d, ZWT15, ZCC15, ZW15d, Zho15c, ZTW16, ZH16, ZYZW16, ZX16, ZLW18, ZCCL18, ZSY18, ZZ19c, ZSF17, ZLL18, ZT16c, ZK18, dFGAN19, dSH17, dSH15, fP15, uRIS17]. Equi [Cho18]. Equi-Gradient [Cho18]. equidistant [Pal15, Zha17b]. Equilibria [LLT19, SSM^{+}18, BCd15b]. Equilibrium [Bor19, YLW16, AT15a, BMRS18, CP15c, FPUK15, HH16, HHP18, JK15, Jeo16, LHJ15, MN15, QLLY18, SMR15, SML^{+}15, SPGP18, SCM19]. equipped [LWS18]. equivalence [BdlRT18, LLL^{+}18b, LCOS19, Zho15d]. Equivalent [CDP19, GWSC16, SNS^{+}16]. equivalent-input-disturbance [GWSC16]. erbium [ZLH16]. erbium-doped [ZLH16]. Erdélyi [Dar16]. Erdos [WYL19a]. Ergodic [ZJO^{+}18]. Ergodicity [HC18, YYZ19]. Ermakov [MR15a, PST18]. erosion [SFC^{+}17]. Erratum [EL19a, GR19a, MAK16, NN19, Sab18, VK17a]. Error [ALZ16, AHY19, DBJC17, GZW15, IZ16, JZ19a, LCW15, LB19, Oka18a, PSY^{+}17, SAES19, dFGAN18, AQ15, AQ18, ASA19, And18, Bac15, Bag19, BÇS15, BBG19, BCNN19, BCG19, BF15, Bor15b, Boy15, CG15, CST18, CKW15, CH18b, CDL15, DA17, DKL15, EGB15, FK17, FM17b, GI15, GGHO18, HL16, HZLY17, KwC15, KAS16, LX16a, LZ17f, MH19a, MW18, MS15c, MNK19, NY15, RzS15, SSS17, SSA18, SY17b, SW17c, SC17, WXY15, WZ16a, WSZ16, WW18, WSC19, XH19b, YZ18b, YCH18, Yil16, YG15b, Yos19, Yüz15, ZL17a, ZCZM15, ZCCL18]. errors [PWFY18, Pöt19]. escape [WDLL15, WDLS18]. escaping [Nav17]. Essential [Zhu16]. essentially [HJMB18, WYLG19]. estimate [CCD18, GI15, GG19b, Yos19, YJW16, ZJZ15, ZCCL18]. estimated [GDS16]. Estimates [Mor15, Ade19, ASA19, ALZ16, Bac15, BCNN19, CG15, CGW16, CH18b, CCGV19, CKDC16, GG15a, GR15c, HL16, HZLY17, IZ16, JZ19a, LZ17f, NY15, NNA15a, Oka18a, PSY^{+}17, SY17b, WSZ16, WW18, YCH18, ZL17a, ZCZM15]. Estimating [ÖAD15, ÖÖAD15, RZMB17, SJ16, DJLX17]. Estimation [GdS18, Han19, MS15a, RRTR15, SLLC19, UPMV18, Wan15e, ddGL15, BÇS15, CÁGGLP16, CST^{+}15, CSL19, DLQ17, DCJ19, FDPP18, FvdMS15, Gav15, GZ18, GMPP18, HXW^{+}19b, KS16b, KSP16, KAS16, LX16a, LPJ18, LW15b, LW18a, LSCK15, LLZD18b, LLZ^{+}19, LGC19a, LGC19b, LDSD19, LPS19, LXW^{+}18, MRC^{+}18, MC15b, MC15c, MS15c, PWFY18, Pop17, QZL18b, SKKL17, SDP16, WL18a, WWMX16, XS19a, Yüz15, ZT16a]. Estimations [Ele15, GG15b, MHM15, MZM15, MNK19]. estimator [BBG19, BCG19, GGHO18, LW18a, XZZ^{+}18, XH19b, ZWY^{+}15a]. estimators [Zam19]. Estrada [LQG18, LGQL19, Sha15c, WX19b]. estuary [ddGL15]. ETA [CST^{+}15]. ethnic [BMDS17]. Euclidean [CGG^{+}18, CKDC16]. Euler [AJC^{+}18, AM18b, ASS16, BS16, BN15, CJ16, DKKM15a, GLM19, Kim15c, KL16, KG17, LJL15a, Li16b, LYZ19, LL19c, LT18a, LSY15, Luo15, PBBF18, QLLY18, SHN16, TM19, Xu19a, YG15b, YNMW16, ZGGY19, ZH19a, ZSY15]. Eulerian [GSDH15, JOÖ16]. European [ASV17, HS15c, LGO17, MZC19, RPB15, ZLC^{+}18]. evacuation [CZ18b, GKDHR19, GW19, HZYW15, HZ17c, WWC^{+}15]. Evaluating [BKZA18, GM17b, Gra18]. Evaluation [KWH17, SLC18, CXK19, Cho19b, DP15b, DKZ16, JGBY16, LWQ18, RLGGAV15, TG17, YZD15, YF17, ZHY16]. Evaluations [Xu19a, VMLO19]. evaporation [KRF17]. Even [ABLZ15, ABC^{+}16, JZ15a]. Even-order [ABLZ15, JZ15a]. Event [CXW^{+}19, PY19a, PY19b, SMZ^{+}17, XPZ18, XZHX17, ZZZY17, ZDF17, CL19c, GSD16, LLC18, MLSX18, SRCA17, SWZL19, TFLZ19, WS19a, WRX^{+}19, WXXY19, XD19, YLP19]. Event-based [PY19a, PY19b]. Event-triggered [CXW^{+}19, SMZ^{+}17, XPZ18, XZHX17, ZZZY17, ZDF17, CL19c, GSD16, LLC18, MLSX18, SWZL19, TFLZ19, WS19a, WRX^{+}19, WXXY19, XD19, YLP19]. event-triggered/self-triggered [WS19a]. Eventually [CEP15, LLL17b]. every [JLL16, ZY18a]. Evolution [LMW^{+}19, ACSS15, CS17, CGPT15, CST^{+}15, DK15a, DOZ18, FFN18, FTF16, GZK16, GT15c, GE16, GSD^{+}17, HARS15, HZYW15, HZS15, KSN15, Kur19, Li16b, LZP18, LY18c, LY15e, LZ15e, LL15f, LLB16, LWZ18, Mah18b, ÖAD15, Rem18, SS18b, SS16, SG18, TBG16, Wan15c, WM15, WG15b, XZL15, XL18b, XLW15, ZCLS18]. Evolutionary [BLA15, RSGFV16, Sch16, XTXF18, XFTX19, YF17, BK18a, BH15, CDZ^{+}18, DLY^{+}17a, DZY^{+}18, GMOP18, KCSD^{+}18, LJR19, LJ17, MWL^{+}18, SWA^{+}19, TES15, Tan15e, YLZT17, ZWY19]. evolves [KZT19]. Evolving [XS15, HZ17c, Sha15c, ZXC19]. Exact [BdlRT18, CJN19, Esl16, GCMR18, Kud19a, Kun16, ZS15e, eMA15, AIG17, CMW15, GGHL15, HARS15, KZ15, LL18a, LFY18, MZ15, Nas19, RMA15, Rui18, S.15, STVCC17, Sun15b, VDV15, YMHW15, ZA16, ZSL15, ZXZB19]. exactly [JFT17]. example [Rem18, ZMQ16]. examples [BCSG18]. exchange [MKMC15, MTCV^{+}16, RXLM19]. exchanged [ZH19b]. exchanger [PHRB18, RSC16]. exchangers [GJR^{+}15]. excitable [GO15b, PRA^{+}19]. exclave [SI15]. exdeg [GA17]. execution [dNdOX^{+}15]. exist [RP18]. Existence [AN15, AL18b, AK15b, DZ16, DO15, FFT15, GW16, GT15c, GZLL17, IJ15, IKMW15, KE16, KHN19, KAH15, LSfW15, LZ15e, LL15f, Lv18, MZM15, NRT16, SZ19b, SAT15, TES15, TNMR18, Vol15, WwQqTW18, XPG15, ZLL15c, ZZL15a, dSH15, BH15, CS16a, CZ15c, Dan15a, DL16, Ers16a, JS16, JX19, KM19, LY17, MDZ15, MHM15, MS16a, Neu16, NE15, RM15, SSA19, SNL19, Yos19, ZLW15, ZY18b, Zho15a]. Exit [CFS18, WDLL15, WDLS18]. exits [CFS18]. Exogenous [SSGG17]. expanded [LFL^{+}15]. Expanding [AM15e, CSZ15, ZW15b]. Expansion [BMC19, AQ15, AK15b, BMR17, CL18, DQW^{+}19, EAAM^{+}15, FLS18, MRF18, ML15, Mil16, MV15, Par15, Zha15h]. expansions [BN15, BE17, CE15, CP15b, Che15a, Che16b, Che16a, CP17, CC17, CKSV19, Kan19, MH19a, WL16b, ZG15b]. Expectation [ARACCA^{+}18, WJ19]. expectations [ZAZ19]. Expected [WGL18, LjL15b, Nad15, VEF19]. experimental [NFS^{+}15, ÖÖAD15, PL15b, Rog19]. Experiments [BB15b, FM17b]. Explicit [CXL16b, CCY19b, FR16, HARS15, JH15, lJdW16, LY18d, QZ15a, Tsi19, Vab15, XZW^{+}15, ZAW15, ZXZB19, BdS19, BLJ^{+}19, Boy17, BMRS18, EKO19, FMC15, FCM17, FKR15, LL15e, Mil16, Oka18a, PST18, Pis18, RSKB15, TZ18a, TTXZ16, WSZ16, ZMZ18, ZLX19, ZYX19]. explicitly [Zha18b]. exploited [MMS18]. Exploring [BS15a, CDZ^{+}18, LW18d]. exponent [Our16, WLL16]. Exponential [BDGS15, BDSS18, CZS15, GWWL17, HT16, LJ15a, LCAA17, LSMT18, MDZ17, QBC19, RVGS17, SSZ^{+}18, SCC17, Sun18, TY16, XKG15, ZYXF17, ZHWW17, ZZ18a, Zho15b, Ace19a, BPC15, BDSS15, BKW18, CDP19, CZM18, CnTVsL18, DKBS15, EÖ17, FvdMS15, GWW15, GZLL17, JHWZ15, JZ19a, JYS16, KE18, LY15a, LZZ15, LSZ^{+}18, LZ18a, LZ15b, LZ17e, LDH18, Liu18a, LMD19, LS17b, MRC^{+}18, MDG18, Nis15, PRA^{+}19, RN19, Rez15, RM16, SID19, SN18a, SN18b, TSJ16, TF18, yWyLyY^{+}15, WHX18, WDLZ19, WCH15, XC15, YSL19, YMMH15, Yu16, ZZL^{+}18a, ZLW15, ZZC16, ZZX18, ZZ17b]. Exponentially [ZNV19, Arm19, EM15, FR16, Li19]. exponentials [Zha18d]. exponentiation [BLP17]. exponents [DH15]. exposure [Bhu15, SG17]. expression [HWBL19, KSS19, PBB18]. expressions [MDH16, Mor17, Nad15, NMM16, QZ15a]. Extended [BCM17, KSV15, LSZ^{+}18, XZB^{+}18, XZ18, XZZ^{+}18, ZA16, AAGZ19, AP18b, CJ18, DHX^{+}19, DGF17, DLY17b, LZ17b, LQG18, LL18a, LGQL19, LDH18, LCLZ19, ND15, SSZ^{+}18, SHLL18, YMS16]. extendible [DVJ18]. Extending [AM17, DEdlT17]. Extension [Cat17, GGAVGGRC15, SQL16, AGM16, ACTV17, BA17, BB19a, CZLM18, DZ15, Eft15, GKM16, HW17a, LB15a, MCO19, RF18]. Extensions [Che15c, Luo15, MH19b, NMS^{+}18]. exterior [LDL15, WWH17, YÖ18, ZQG16]. external [AAE19, DF18, FW18a, GST15b, HW17b, KA15c, LY17, TTV15, VVA^{+}15, WXXY19, YmZHC15]. Extinction [CYLW15, CXMP16, KL15, ZMT15, AELH15, CDHK15, LXJ^{+}17, LJHA18, SRJ^{+}15, SRK18, Zha15e]. extra [WY17, WM18a, YFLZ18]. Extracting [TDA17]. extraction [AGC19, CNK^{+}18]. extraneous [GKN16]. extrapolated [BS19b, LT18b, RW17]. extrapolation [Dok18, Li17c, LJC16, MH19a, SN18a, SN18b, WZ15b, YKZ15]. extrema [Van15]. Extremal [CL16b, DM16, FKA17, GT15b, GA17, LQZ16, LBRS18, LW18e, MRS19, Tep19, WX19b, ZL16b, BF16b, CWD16, CWZ17, GL18a, LLWX19, LWW19, PWS17, Wan17b, WWL18, WC18, YKLS19, ZY18a, ZAW15]. extreme [MNYL18]. extremism [San19]. eye [TSPF18]. eye-fluid-structure [TSPF18].

Face [GS19, BLNSF17, BST18, LVAB^{+}16, LTC16, SHLL18, TWF^{+}18]. Facebook [DW17]. faces [LVAB^{+}16]. facial [KSS19]. facilitates [RCW^{+}18]. facilities [ZLLP18]. facility [CRCVMS18, Mat15, MK15c, Pal15]. Factor [NL18b, BGS19, RSLL^{+}19, ZGL15]. factoring [SD17]. Factorization [Gol19, TPCL15, DR19, HGM15a, LY15c, SFLM19, TA16]. factors [Bor15a, WC15b, XSS^{+}17]. fading [LLZD18b, SP19]. Faedo [Mus18]. failure [HDG15, PK15, ZLDZ16]. failures [MCPO15, ZPZ^{+}19, ZW19b]. Faint [CnTVsL18]. fairing [WZL15]. Fairness [WZL15, ZCLS18]. Falkner [Li17d, RSKB16, RR18a]. Falkner-type [Li17d, RSKB16]. falsi} [SM15d]. Families [MH19b, AT15a, AKKB17, BCMT15b, BCMT15a, CN16b, SCY19]. Family [HM17, AAGZ19, Akt15, ACDT18, AM15f, AK15d, BAM16, BC16, BCT15, CCTV16, CFSGM17, CN15b, CN15d, CN16a, CLMT15, CGMT16, CSTH17, Dob19, eFJY17, GT15a, GHL19, GKN15b, GKN16, GPHAPRW16, GS11, HH15, KKB15, KM15a, KDG19, LKM17, Li17d, LMMR15, NBK16, PV16, PIP19, QZ15a, RS15c, SDGU16, Sha15d, SSK15, SA16c, SK17b, SAS19, WAS19, Xu15c, ZS15d, ZX15]. Fan [BHD18]. Farmer [PV19]. Fast [ASZD18, AV15, CYCW15, CMW15, DG18b, DPSS16, Khl18, Laz17, MSL15, MWH^{+}18, Naa15, Pal15, SDWG19, TWF^{+}18, ZLC^{+}18, ZM15c, AKK15, CLJ^{+}15, De 17, Fuk15a, Fuk15c, Fuk15d, Fuk15b, HM19b, JGBY16, KMZ19, LPZ19, LZ18b, OA18, PSP17, SNS^{+}16, UKS^{+}16, XS15, XW19, ZZL18b, ZMH18, ZWY15b]. faster [NSR^{+}19]. fatal [ZWL^{+}19]. fate [Tan15e]. Fault [Du17a, GSLC18, HYL19a, MLSX18, YYS17, CTT^{+}19, DH17, DCJ19, LX17, LW17a, LZX19, LGCS19, LDSD19, LPS19, LZHC18, RHP18, RCRL18, SJWK18, WLH^{+}18, XCY18, YLW19, ZLLZ16]. Fault-tolerant [YYS17, DH17, RHP18, SJWK18, WLH^{+}18, YLW19]. faults [KLKU18, KRA^{+}18b, LMW18, QY17, SKKL17, SJWK18, TFLZ19]. Favard [KTA15]. FD [GM17b, ODP17, PeO18, Su19]. FDM [LB19, WD16]. fear [Pla17, ZCFW19]. feasibility [QLZ15]. feasible [JPTL15]. Feature [BGSCGL19, VEF19, YSD^{+}17]. features [TH19]. feed [SS19b]. feed-forward [SS19b]. Feedback [MA16, PFA15, XZL15, YSWZ17, CR19a, CXP16, CLXW17, CCQS16, CLCW17, DHX^{+}19, DAQI15, DH17, GZC19, HYL19a, KSM^{+}19, LLCR15, LLZ15, LJL15a, LZT^{+}18, LMZ^{+}19, Luo19, Pai15b, RXZW17, SAJ18, SZ15, SLZ^{+}16, SS19b, SWSW19, SML18, SLL^{+}16, TTZ^{+}19, TPL16, TLL^{+}18, Upa16, WLH^{+}18, XCDZ19, XZCC15, XCL18, XZLL17, ZLDZ17, ZMMZ15, ZHWW17, ZW19b, ZXX^{+}19]. feedforward [EOP17]. Fejér [SISÖ15]. Feline [CYLW15]. FEM [NMNTBXVD15, AV15, SY19]. FeO [KSKW^{+}19]. Fermi [Fuk15c, Fuk15d, Fuk15b, NT19]. ferromagnetic [Ant18]. fever [PTA16, SG19b, TMM18, ZXC19]. few [Sza15, Zha19a]. Feynman [BHD18]. FFT [CDL15]. FG [LZL15a]. FGM [ABN^{+}17]. FGP [Lac15]. fiber [HGH19b, LHHS16, LZWZ19, WQ15, ZLH16].

Fibonacci [AAF15, BBD15, CT18, MH16b, Sta17, WW17, YZ18a, YZ17a]. fibration [AMN19]. fibre [KKG18]. fibre-reinforced [KKG18]. Fick [HSS15]. Fickian [LR16b]. fidelity [XZB^{+}18]. Field [RXLM19, AM15b, BTP17, BGP15, DS15, DWS19, HLJ18, ILPE17, JJ15a, JDL15, Kyr15, Liu16a, LRH15, MMZ^{+}17, MGK^{+}16, MPW15, PPS25, PL15b, RRR19, RRR20, SV17a, VVA^{+}15, WE19, WMZ19, ZHS18]. fields [Ant18, CZ15b, LL19b, QLLY18, YCD17]. fifth [AM18a, BdlRT18, CEHVT15, GfSmL16, HGH19b, Kud16, LZH16, SGMH16, Tal15, WTJ17]. fifth-order [BdlRT18, CEHVT15, GfSmL16, HGH19b, LZH16, Tal15, WTJ17]. filariasis [Bhu15]. Filippov [QTXY16, QTTL19]. filled [GYY15, GSDH15, SP15, SRP17]. film [BTH18, DS15, ZZ18b]. films [Alk19, BBGR15]. Filon [FMX15]. filter [DN15c, Gao18, GGZR16, LW18a, LW17a, LY18a, LSZ^{+}19b, LXTF18, Nak18, PZ16a, WCZ15, XCY18, ZWY^{+}15a]. filtering [AM19, AAHT18, CHL15, Kim19, LCY16, MC15a, PY19b, SZPH16, XXW^{+}19, YKJY17, YCX^{+}19, ZLDZ16, ZML^{+}17]. filters [AAF15, ARACCA^{+}18, BL15, GMPP18, MGPL15]. finance [TVK^{+}16, WZ18b]. financial [HDG15, YSX15, ZLS16]. find [CN15c, HSAR19, KD18, MN15, NCS16]. finders [GKN15a, GKN16, LKM17, vLtTBI18]. Finding [CRC15, ZWXW15, AU15, AG15d, GL15a, How15b, KC18b, PH17, PV16, SB15, SM15d, SRBA15, ZA16]. Finite [Bas19, BR16b, BDH16, CYSY16, CGZ17, FMDZ19, HWY15b, HJKP18, HS15a, LX16a, LW17a, LXD19, LLLZ16, LLZ17, MFJZ15, RSC^{+}17, RZHY16, SJWK18, SA16b, SYZP16, SGJS16, SJL16, SH18b, WZM15, WZZL16, WSLC17, WXX18, WDG19a, WRX^{+}19, WSYL18, WNW^{+}19, YSL18, YZ18b, ZXS^{+}18, ZCCZ19, ZBH^{+}15, ZJ15b, Zha15i, ZFZK15, ZWO^{+}19, AT15a, ABGS15, AGZ15, Ach19, Ade19, AJ17, ALZ16, AP18a, BB18a, BY19, BdS19, BGMS18, BTWY15, CXL15, CKG19, CHL15, cCpSkZ15, CZ15b, CGW16, CR18, CH18b, CCL^{+}16, CZ17b, CSA18, CD16b, DJL17, DKZ16, DFJ16, DRS18, DC16, DEdlT17, EdLC18, FDPP18, FZLT15, GW16, GY15, GS17a, GM15a, GGHL15, Gia15b, GN15, GfSmL16, GS11, HJMB18, HEA15a, HEA15b, HYB17, HK15, HFCS16, HWY15a, HP16, HW17b, HK18b, HV19b, HL16, HZLY17, HLJ18, Ied15, JBS15, JDG18, JMW15, JAN16, KKT15, KCZK15, KZR17, KNK15]. finite [KM15a, Kok15, KXZ19, KAMD18, KK16b, KK17, KRA^{+}18b, KS19c, Laa18, LSCK15, LCQ15, LR16b, LR17, LL17b, LJD17, LH17c, LZCW18, LW18d, LZ18b, LSR18, LR19b, LZ19c, LL19b, LZ17e, LY18d, LMW18, LS19b, LT18a, LZ17f, LT18b, MC15a, MJL16, MBHS18, Maj18, MVS19, MO16, MRQZ19, MC15c, MKN15b, Nak18, NKP16, NRS19, PSY^{+}17, Pla15, PP19, QZCJ19, QZW19, QFTT15, RC17a, RGA19, SR15e, SY17a, SzSC19, SWL16, SLW16a, SW17b, SW17a, SZL^{+}18, SY18b, SLL19, SJW16, STA^{+}18, SFC^{+}17, Su19, SCY19, TM19, Ton18, TZH15, UKE15, VS17, WZ15b, WMLG17, WLZY17, WZWJ18, WJWX18, WYLG19, WLQ^{+}19, WSC19, Wei17, WLWD18, XLR17, XCL18, XS19b, YMD15, YCH18, YJK19, YW19b, YZJ18, YDLK15, Yos19, YLZT17, YLL^{+}16, YWH15, ZEF19, ZB17, ZI18, ZWF15, ZL15a, ZS15d, ZZZ16, ZSS16, ZT16b, ZZCY16, ZML^{+}17, ZSY17, ZL17a, ZZ18c, ZZ18b, ZC19b, ZZ19a, ZY15c]. finite [ZCZM15, ZL19a, ZL19b, ZZC19a, ZS19b, ZSY18, ZSF17, dFGAN18, dFGAN19]. finite-buffer [CSA18]. Finite-difference [WSLC17, JDG18]. finite-horizon [ZML^{+}17]. Finite-time [CYSY16, CGZ17, FMDZ19, HWY15b, HS15a, LW17a, LXD19, LLLZ16, LLZ17, MFJZ15, RSC^{+}17, RZHY16, SJWK18, SYZP16, SJL16, WZZL16, WXX18, WDG19a, WRX^{+}19, WSYL18, WNW^{+}19, YSL18, ZXS^{+}18, ZCCZ19, ZJ15b, ZFZK15, ZWO^{+}19, CHL15, CCL^{+}16, DKZ16, FDPP18, HWY15a, HW17b, KRA^{+}18b, LSCK15, LL17b, LMW18, MC15a, MJL16, QZCJ19, WLZY17, XCL18, YDLK15, ZSS16]. Finite-volume [BR16b, STA^{+}18]. finitely [HHZ15]. Fink [Nis15]. FIR [LW18a]. fire [CFS18]. Firefly [ZBKS16, FPKF15, KSK^{+}15, YZMM15]. First [KI15, KH17, Tua15, AG15c, BEM15, CL15a, CC18b, CNQRR18, IKMW15, Kud19b, LZS15, LLdM15, MS17b, MRE18, MMDG16, NEK18, Oni18, RSKB15, UHN^{+}18, Xu15b, YY15, Yüz16b, Zho15d]. first-order [Oni18, RSKB15]. Fiscal [SD19a, QZL18b]. Fish [MM16, MMS18, PSM17]. Fisher [Ata16, Gat18a, HEA15b, QDD15, WJL19]. Fisher-KPP [QDD15]. fisheries [GLR^{+}18]. fission [WTJ17]. fitness [DHY^{+}18, ZHY16]. Fitted [TG16, FR16, LD18a, Li18, ZNV19]. fitting [GI16, KF19, LWPL16, MH19b, RS17b, SV15a]. FitzHugh [BTWY15, FLL17, GO15b, IWL17]. five [CH19, LAAZH19, SuSR16]. five-vertex [CH19]. Fixed [BT15b, ACS15, AKS15, Alf17, AM15h, Boi15, BST18, dCCGV15, CZ17a, DL17b, GK15, GKN16, Han19, HRS15, Jeo16, LR19a, LCAA17, LLTW19, MSL15, Nak18, PN15b, RDLG16, SSA19, SDK15, STD15, TTP16, WF15a, WC15a, WX19b, YLW19, YZYgG18, ZS16a, ZH19a, ZCW19a, ZYLY17]. fixed-lag [Nak18]. fixed-rate [dCCGV15]. fixed-time [LCAA17, YLW19, ZCW19a, ZYLY17]. Flag [DL18a, TZ18b]. Flag-transitive [DL18a, TZ18b]. Flat [MB15a, SSN15]. flexible [Abd15, AGM16, Ber15, CSV19, KRA^{+}18b, Pai15b, XJW17, ZCAN19]. flight [XCW19]. flights [PRD19]. FLLOP [MHCF15]. floating [WL16a]. Flocking [SWZL19]. Floquet [AK16a]. Flow [AwEE15, ZS16b, ATV19, AJMN18, AB15b, ATK^{+}16, ÁL15a, AAM19, BKKN16, BAR18, BGP15, Bot15, BLJ^{+}19, BVJ15, CSZ15, CKKF16, CT15, CT17d, DS15, DU18, DL15a, DL15b, EAY17, EM15, ERFE16, Ers16a, FFG^{+}16, FFH^{+}18, FFH^{+}15, FF18, GGAVGGRC15, GAT18b, GCM^{+}18, GYZC15, HM15a, HH16, HHP18, HAMA14, HP17, HRS17, HG18, Jai18, JMW15, JV18, KHN19, KKD17, KCKZ15, KFHH15, KSP16, KSZ15, LCQ15, LR16b, LWLM19, LZY16, Liu18b, LCW19b, LLL18d, DPL18a, MSI17, Mah19, MS17a, MKG15, MP18, MKRS16, MSL15, MNJM19, Neu16, NKNA16, PPS25, PMC16, Pan15, PK18a, Pel17, Pon18, PP18a, QX17, RRR19, RRR20, RP16, RSC16, RM19a, RMK19, RSO^{+}18, SST17, SS18a, SK15, SSB18, SSN15, SS18c, Sha12, SV17a, SSM17, SMG^{+}18, STVCC17, SH15d, SH18b, WF15a, WXM^{+}18, XP17]. flow [XYZK15, YJK19, Zei16, ZYLY19, ZWM16]. Flower [IAW18]. flows [BBG19, BTH18, BKL^{+}18, CS15a, Cho18, CK18, DMA^{+}19, DMSS18, EEG^{+}16, Ers16b, EdLC18, GG18, Ili19a, KPG16, KMB16, Laa18, LTL16, LXL17, LS19b, MZS18, NB15, Qia16a, Qia16b, RDG18, Tha18a, VG18, WL16c, WLQ^{+}19, WYD^{+}18, Xu19e, Zei16, Zha18e, ZZ18b, ZZX19, ZM19]. flowshop [PGLG17]. Fluctuating [WW16c, Mou15]. fluctuation [FMRRT15, HRS17, KSK^{+}15, ZCW^{+}19b]. fluctuations [LYZL18, UPMV18]. Fluid [Ili19a, TBSR16, TT16, ZS16b, AHLR19, Awa15, BG18, BBG19, BY19, BGP15, Bot15, CSZ15, CCL^{+}15, CKKF16, CT15, CT17d, DBV18, DWTT17, EAY17, FTE19, FF18, GG18, HAMA14, HMHA15, HRS17, HG18, IR16, Jai18, KHN19, Kud19a, KS19b, Laa18, LX16a, LWLM19, LXL17, LCW19b, LLL18d, LFY18, MSI17, Mah19, MKN15b, NM19, PPS25, PK18a, Pon18, PP18a, RRR19, RRR20, RP16, RVE19, RL15, RSC16, RM19a, SST17, SSN15, SGM^{+}16, SBH16, STK15, SCM19, TSPF18, WF15a, WL16a, Xu19e, Zha15a, ZLLA19, ZWM16]. fluid-saturated [RL15]. Fluid-structure [Ili19a, BY19, LX16a, STK15]. fluidized [MNYL18]. Fluids [ZS16b, DQ18, Kim15a, NS17a, RDG18, RG18b, WTJ17]. fluoranthene [LY15b, LY16a]. fluoranthene-type [LY15b, LY16a]. flux [Maj18, MZS18, PRA^{+}19, Pov15, SS18b, WXM^{+}18, XLR17]. flux-splitting [Maj18]. fluxes [LWLM19, Liu17a, QZL^{+}17]. FM [SJ19b, SJ19b]. FM/FM/1 [SJ19b]. foam [NDZ15]. Fock [Fau17]. Fokker [GDL16, GAB17, MW18, SR15e, XYG^{+}18, ZJY18]. fold [APK15]. folded [NMNTBXVD15, QY17]. folds [SH18b]. folk [BMDS17]. Follow [DDF18]. follower [WV18]. followers [Ask18]. following [CL19c, FW18a, LZY16, MGCAPP16]. followings [BGS19]. follows [LZZ15]. FOM [LH19b]. food [Liu15d, LJHA18, MMS17a, Sah15, SSB15a]. food-limited [Liu15d]. foot [CDFP17, LWPL16]. foraging [AADLF15, BS15a]. force [BKT16, Dan15b, GST15b, KZL19, KM15a, KPP18, Kri19b, KG17, LW16, LH19a, Ma17a, SLW16b, TTV15]. Forced [HPN15, YWL19, AÖ15a, AÖ16, LL19b, SH15c]. forces [YD18]. Forchheimer [KZR17, LCW19b, XLR17]. Forcing [ZZ16a, AL19, Bot15, DH19a, GM19, HQ15]. forecast [FCW^{+}17, WF15c]. Forecasting [YS16, CLHC15, Dok18, Lah18, LX19d, ZCW^{+}19b]. foreign [ARR16, Haf17]. forest [CYW18]. forests [GL18a, LLST19, WTD^{+}19]. Foreword [BKS18]. Form [SRBA15, BB19b, DPD18, EAAAS15, GS17a, HP17, Ili18, JSM16, Liu18a, LD19b, MNNP19, Nad15, PSPÁV15, SH16, TG19, VK19, XZW^{+}15, Khl18]. formalism [BHD18]. formation [DQW^{+}19, FVMO15, IWL17, PL18, TLZ17, WXXY19]. forms [AADMM16, CN15a, CN19a, GT15b, HGH19b, JS15]. formula [BR16a, Ben18, BLM^{+}16, Che18a, DeC19, DJLX17, EAAAS15, GK17, KDS16, LSM15, MP19b, PVV^{+}15, SdR19, SRK^{+}19, SKP16, Xu17b]. formulae [CRC19, Dan15b, PPV15, Reh17]. Formulas [Aca15, DHS19, ACK15, AHS18, BGMZ17, Che15a, Che16b, CP17, KA15b, PR15b, QM15a, QZL18a, SHA15b, SHN16, WCS16, XC15, YC15, ZLD18]. formulated [TPS15]. Formulation [AAC16, ABI19, BFR18, BM18b, KCWK15, KSL19, RSKB15, WYL15, WC17, ZF18]. formulations [Oli17]. Fornberg [RG15]. Forward [WY19, GR15a, SS19b, WZ18b, ZL15c]. Forward-backward [WY19, GR15a, WZ18b]. forward-start [ZL15c]. forwards [TNPL15]. foundation [AJC^{+}18, Rad15, RM16]. foundations [LZL15a, ZLL15b]. Four [CCL^{+}16, Sah19b, TTZX15, AKT19, CH19, DSAB16, HASC19, LA15, MDH15, MKS19, Mon19, SDAB17, SBL16, SuSR16, SGY^{+}15, TH19, WWH16, WWW19]. four-dimensional [SGY^{+}15]. four-parametric [LA15]. four-point [SBL16]. four-strategy [AKT19]. Four-Tank [CCL^{+}16]. four-vertex [CH19]. Fourier [ALQ15, BN15, BDE^{+}15, Boy15, CZLM18, DP15b, EO18, HZ17a, HGL19, JZ19a, MRF18, TQ18, WX19a, ZS17]. Fourier-type [HZ17a]. Fourier/discontinuous [EO18]. fourth [Ach19, AM15f, BLK19, BCMT15b, BCMT15a, CFG15, CLY17, Che16c, CN15d, DL18b, GKM16, HASC19, JSLW17, KKB15, KAI19, KNBA15, LKM17, LSfW15, LW18f, LJC16, MDHD18, Mah17, MAJR19, MK16, MMK17, MKS19, NBK16, PK18b, RLZ19, RW17, TBG16, XW19, XLW16, ZY15b, ZLX19]. fourth-order [Ach19, AM15f, BCMT15b, BCMT15a, CLY17, Che16c, DL18b, GKM16, HASC19, JSLW17, KAI19, LKM17, LW18f, LJC16, MDHD18, Mah17, MAJR19, RLZ19, RW17, TBG16, XW19, ZLX19]. Fowler [KF15]. FPGA [GCAA18]. FR [PR15a]. Fractal [KVZ16, SuSB15, ZTW16, BRD15, BÁR15, FMSG15, GT19, KCK19, RRT15, TWF^{+}18, VV18]. fractality [MTCV^{+}16]. fraction [CY15a, LSY15, Tan15e]. Fractional [AMO18, BC15, CDK17b, CCX15, DOR18, GBJ19, KS16a, LZ19d, LWW15a, MFC17, TT18, YW15, ZBBW17, ZZ15b, ZY15a, AB15a, AKAA15, ALL18a, ABHR15, AHO18, ARACCA^{+}18, AN15, AAA^{+}15a, ANT^{+}16, AL18b, AEB18, AYD19, ARL15, ASA19, AZ14, AG15a, AAKT15, AT15b, AM18b, APL15, ALP15, AHY19, ALN16, AE17a, AJP18, Ata16, AS19b, AZ19, BGRT15, BT15a, BMO15, BD15, BAB18, BD18, Bea19, BS17, Ben18, BTWY15, BCD^{+}19, BBD15, CVJ19, CGCK19, CH18a, CLL16, CLX17, CZP15, CXP16, CZLS18, Che15c, CLJ^{+}15, CKW15, CWHY15, CC15, CW17, CGWW17, CZL18, CC18b, CWL19, CG19, CDL19, CJ16, CS16c, CVMGPPGC15, DK15b, DG15a, Dar16, DB18, DZW15, DCS15, DFP15, DPSS16, DOZ18, DWS19, DQW^{+}19, EAAM^{+}15, ES16, ET15, Esl16, FZLT15, Fen15, FBF18, FÖB19, FY18]. fractional [GCAA18, GDM16, Gao18, GD15, GK15, GZK16, GFG15, GS15b, Gra15, GT15c, GZ15b, GR17b, HLZ15, HMZ17, HW19, HFCS16, HH18a, HGH19a, HPN15, HA19, HL17, HS18, HHGC16, HAY19, HAH19, HLLT15, HYW^{+}15, HCX^{+}17b, HCX^{+}17a, HLC19, HHG19, IKMW15, Isc16, JDG18, cJzS15, JLX15a, Jia15, JKK15, KSM^{+}19, Kat15, KAI19, KDH19, KI15, Koc15, KPT16, KDG19, KS19b, Lac15, LL19a, LdCM19, LQ15, LJW^{+}15, LLY15, LL15e, Li17b, LLC18, LZCW18, LW18d, LZ18b, LW18g, LKZ18, LW18f, LLTL18, LK19a, LGC19a, LW19b, LXZ19a, LGC19b, LJL19, LY15e, LSZF18, LWW15b, LZ15e, LLG^{+}15, Liu16b, LCL17b, LLL18c, LNWS18, LXZ19b, LL19d, LFC^{+}19, LJ19, LO19, LL15f, LFY18, LZ19g, LM15, LHX15, LHWG16, MWWM15, MDHD18, MD18a, Mah18a, Mah18b, MGPPGCCV15, MYHS18, MS19c, MS16a, MO15a, MTJ15, Nag17, NH17, NE16, Nas19, NL18a, NS17b, NE15]. fractional [Ort15b, OL19, PAM19, PAE^{+}17, PHM19, PWS17, PZ15, PTBN19, PJ15, PS19, Pov15, PKS15, PKB18, PRC^{+}19, Pu15, QXW19, QW19, RWW15, RS15a, RG15, RzS15, RW17, RZW18, Rui18, S.15, SuR15, SuRI15, Sae17, SDS18, STB16, SHR18, STS15, SISÖ15, SzSC19, SZL^{+}18, SY18b, SXS15, SS16, SSM^{+}15, SMM18, Sin18, SKBR18, SKHA18, Sow18, SML18, SAT15, SzShG16, TB17a, TZ19b, Tar15, THM19, UBT18, VV18, VRC15, VBG^{+}17, WW16b, Wan15c, Wan15e, WXY15, WIF15, WZ15a, WLZW15, WM15, WL15, WZ16b, WFZ17, jWgX18, WXW18, WXWX18, WY18, WDCJ19, WX19a, WR19, WXLL19, Wei17, WLFX15, bWzZjH16, Wu17, WBL17, XZL15, XHW18, XYG^{+}18, XZ19, XW19, XX19b, XXZ^{+}18, XL18b, XLL^{+}19, XL19, YL17, YWL18, YW16, YMS16, YLHS17, YWO18, YZH^{+}19, YAIN17, YmZHC15, YJW16, YWZ18, ZMBJ19, ZV17, ZNV19, ZEF19, ZT15, ZDK15, Zha15f, ZAW15]. fractional [ZLWW15, ZSZ17, ZZY18, ZYC18, ZLC^{+}18, ZLD18, ZW18, ZWBC18, Zha18b, ZMZ18, ZJY18, ZLLA19, ZRLX19, ZBH^{+}15, ZCQ15, ZHX^{+}17, ZCX17, ZZL18b, ZLX19, ZL19a, ZL19b, ZZC19a, ZEW19, ZYJ15, Zho15c, ZX16, ZZ19c, uRIS17, GR15b]. Fractional-order [AMO18, CCX15, DOR18, KS16a, AKAA15, CZLS18, CWHY15, CC15, GCAA18, Gao18, HW19, KSM^{+}19, LJW^{+}15, LLC18, LK19a, LGC19a, LGC19b, LNWS18, LXZ19b, NS17b, SSM^{+}15, WY18, WXLL19, bWzZjH16, XLL^{+}19, YLHS17, YZH^{+}19, YmZHC15, ZT15, Zha15f, ZSZ17, ZYC18, ZWBC18]. fractionally [RS15d]. fracture [CR18, ES17b, GLA^{+}17, MPR^{+}17]. fractured [LCW19b, RRT15]. fragile [DHX^{+}19, KRA^{+}18b, LSCK15, LGC19a, LTZ^{+}17, MRC^{+}18, RSC^{+}17, TTZ^{+}19, ZSS16, ZPM16, ZXX^{+}19]. fragmentation [YWS^{+}19]. frame [FMB^{+}18, FMB^{+}19, ILP^{+}15, Kri19a]. frame-like [Kri19a]. framelets [SZ18a]. frames [Far16, Kri19a, LZ15a, MFC17, Sko17]. Framework [HWL^{+}19a, CH18c, GZG^{+}15, JJ16, LAVME18, LdCC15, Liu15g, NB15, YPZ18, ZX19]. Fréchet [AK15b, SGMH16]. Fredholm [Bag19, Sab18, AM15c, ASA19, Bab16, BS16, BSZ16, CJ15, DA17, DGZ16, ER15, EGB15, EMM17, FK17, GG15a, KJ15a, LHZM19, MM04, MS17b, MRE18, Mic15a, Mic15b, MH15a, MH15b, MH16b, MH16a, OR18a, OA15, PMC15, Par15, RME18, SR15a, SR15b, Tan15c, YTH19, Yüz15, Yüz16a, uDuI18]. Free [SRP17, ASHSC17, AKKB17, BE19, BKL^{+}18, CP19, Cim18, CK18, DCW15, FA18, Fuk15a, wG15a, HRS17, HMT15b, LZL15a, LCW15, LHJ15, LCL17a, LY18a, LD19a, LL19b, LGL19c, Ma17a, Maa15, MGN18, MB17, NMNTBXVD15, Pan15, QDH^{+}18, RA15, Sha12, SA16d, SLWX15, SW17c, WZ16a, WZF^{+}19, WL19b, XJW17, hY18, ZZLZ17, ZLW19, ZLL15b, ZHJ^{+}17, Zha18e, Zha15h, ZZL16b]. free-matrix-based [ZLW19]. free-surface [BKL^{+}18]. free-weighting-matrix [ZHJ^{+}17]. freely [CCL^{+}15]. frequencies [SI19]. frequency [BM18a, BDE^{+}15, BL15, BL18, CM17b, DH19b, GB19, KDS16, KCKZ15, LW18a, NC16, Oli17, SY17a, TPS15, Zad16]. frequency-current [CM17b]. frequency-independent [Zad16]. Fresnelet [FMB^{+}19, FMB^{+}18]. fretwidth [ACC15]. Freud [AHM16]. Freud-like [AHM16]. friction [WZY^{+}19]. frictional [MG19]. frictionless [HMZ17]. frictions [JX19]. Friedrichs [BR16b, DB16]. Fries [ZAY18]. Frobenius [HM18b]. frog [LF15]. front [Zil15]. fronts [ZZM18]. FRP [MC15c]. frustration [CRC19]. frustum [MRR17]. FS3D [EEG^{+}16]. FSPL [MR16a]. FSSP [UHN^{+}18]. Full [vLtTBI18, CNQRR18, LL19c, Lia15b, LMZ^{+}19, PL15b, Pla15, Por18, SB16, XCDZ19, ZXS^{+}18]. full-field [PL15b]. full-size [SB16]. fullerene [GDRP^{+}19]. fullerenes [AOS17, BHMS15, SZ19c]. Fully [LLL18c, ZWM16, BFR18, HYW^{+}15, Ili19b, LRL18]. fully-coupled [Ili19b]. Function [JZ15b, YSWZ17, Abd15, AQ15, AQ18, AU15, ABMA19, AM18d, ACTV15, AAA15b, Bag15, BB15a, BT15a, BM19d, BH17, Bor15b, CEMR16, CST18, CP15b, Che15a, Che15b, Che16b, Che16a, CNQRR18, DK15b, Dan18, DD19, DDH18, DMH19, Ele15, Fuk15d, GYY15, GKN15b, GDS16, GST15a, GST19, GN15, Hag15, HSAR19, KCWK15, KCK19, Kha17, KwC15, KZZ15, KKG18, LR19a, LKM17, LY18a, LXJ^{+}17, LSM15, MZ15, Mah18a, MRS18, Miy19, Mor15, ND15, NEK18, PSPÁV15, Qi15, QM15b, RLH15, SC18a, SIA^{+}17, SS15a, SS17, SH16, SJ18, SS15b, SuSR15, SGG15a, SN18c, Su19, SGY^{+}15, SQL16, TCY15, TZ18d, VDI17, Wan15e, WL16b, WHTW17, WJZF19, WLLL15, XS19a, YC15, YCZ15, YDCS15, ZLDZ17, ZCU17]. function-to [LKM17]. Functional [Pol19, AH15, BD17, CC15, CCZ15, Dis15, Du15, Fan15a, FB19, FZG19, GHW19, GD15, HA16, HRS15, HSAR19, JS15, JRM15, KB17, KAH15, KCKZ15, KKL18, LZ18a, LSR18, LDH18, LO19, MZM15, MH17a, MHZ18, MO15a, NHWZ15, OS15, PPV15, QZ15b, Reu15, RXL19, SWUE19, SSA19, SL15b, SH15c, WPL15, Wan17c, WZS19, WN16, XHW15, XL19, ZPRZ18, Zho15b, ZK18]. functional-differential [CCZ15]. functional-integro-differential [QZ15b]. functionally [DWRZ17, HSS15, LZL17, MB17, NND^{+}18, Rad15, WZWD15, ZLL15b]. functionals [AR16a, AM18b, CEMR16, CWLA18, LJHW19]. functions [AKAA15, AGN15, ACK15, AIK15, AL18c, ALQ15, ASA19, ACKS16, ABA18, ALN16, AD17, APK15, Bab16, BM15a, BB15b, BM15b, BAB18, BMC19, BÁR15, Boy15, Boy17, BC16, CCLS16, CEMR16, CS16a, CM17a, CPST16, CE15, CCH15, Che15c, CP15b, Che15b, CGAL19, CNNA16, DCT19, Dan15a, DA17, DDSL15, DSDL19, DOR18, DT16, DÇO15, DLY17b, DMH19, Dzi15, EFL18, GA15, Gra18, GG15a, GJNRG15, HAC18, HR16, HLC^{+}17, HH18a, HHZ15, IKHS16, Isc16, JNGB18, JGBY16, KKT15, KIAG16, KA15b, KVZ16, KNT17, LS17a, LH15, LL16a, LW18d, LLC19, LDSD19, Li19, Liu18a, MM04, MS17b, MS19b, MRQZ19, MNNA15, MH15a, MH15b, MH16a, MH17b, NR15, NL18a, NMM16, NMS^{+}18, NLC19, NAM18, NNA15b, NB19b, ÖK19, RZD^{+}19, Ras15, Rom19, SS18a, Sab18, SR15a, SA18b, SAM^{+}16, SDGU16]. functions [SR15e, SISÖ15, SSN15, Sim15, SGG15a, SGG15b, TLX19, Uça16, Vil16, WA15a, WCS16, WWZ15, WBL17, XC15, Yüz16b, Zad16, ZMBJ19, ZS16c, ZL18, Zha15i, ZHS18, dlRGB16]. functions-based [ZHS18]. fundamental [AV18, DLCT18, TLHC17]. fundus [ARACCA^{+}18]. funnel [Hus19]. Further [CLQ16, Du17b, EYE18, GWW15, JHWZ15, JSW^{+}18, Kim15b, LS15a, WCH19, Qia16a, SGG15a, VDV15]. fused [LK19b]. fusion [KŚK^{+}18, WTJ17, ZZC15]. future [ZAZ19]. Fuzzy [CXP16, MJSA19, APMM18, CHL17, CLZW19, DH17, DZC15, GSPH19, HLFZ19, KBR15, KB15, KSM^{+}19, KRA^{+}18b, LY15d, LZX15, LSLZ18, LZ18c, LGC19b, LXC^{+}19, LO15, MA16, PY19a, PY19b, RXZW17, RMB15, Rez15, WWHZ18, XD19, XCDZ19, ZLDZ17, Zha19b, ZZMS19, ZYZ15, ZW15c, ZWO^{+}19]. FX [ZBL15].

G [KGGAV^{+}18]. GA [Gar16b]. Gabor [ARACCA^{+}18, LZ15a, PHSB18]. Gaddum [MCS^{+}19, YCYL18]. gain [LX16b, YQT15, ZZC16, ZML^{+}17]. gains [CLW18]. galactic [Nav17]. Galerkin [Pan15, AJMN18, ADSS16, AKK15, AM15c, AHY19, ACG19, AD19, BCG19, BTWY15, BVJ15, CG15, CKG19, CH18b, CDL18, DWRZ17, DK15c, DBK15, DKKR16, Dol15, DS16b, DF16, EKS^{+}15, EO18, ET15, FY19, GM15a, GML18, GWK16, GM17a, Gia15a, Gia15b, GGHO18, GZW15, HSS15, HZLY17, Hua15, HYW^{+}15, IH15, KCK16, KBK^{+}16, KG16, KFHH15, KK17, LCW15, LWZ16, LJD17, LCL17a, LJL18, LD19a, LHLL19, Lia15a, LZ17e, LW18h, LL19d, LYY15, MMV16, MB17, Mus18, NRS19, Par15, PK18b, QLLY18, QZL^{+}17, QDH^{+}18, QZW19, SH15a, Sar17, SSN15, Sha12, SW17c, UKE15, UKK17, WSZ16, WZWJ18, WZWD15, Wei17, XJW17, YCH18, YX19a, YWZ18, hY18, ZZCY16, ZC19b, ZY15c, ZZC19a]. Galerkin-finite [LJD17]. Galerkin-least [DBK15]. Galerkin-Petrov [AJMN18]. Galfenol [KCK18]. gallery [SHLL18]. gallery-based [SHLL18]. Game [GLR^{+}18, ANP16, AEP15, AKT19, Ask18, BM17, hCWW^{+}15, CCW17, CDZ^{+}18, CCW19, CZM^{+}19, DY17, DLJW15, ESS16, ESES16, Els17, FXC17, GW19, GSD^{+}17, HHR17, IÖ19, KEE17, LGL^{+}19b, LJR19, LWWZ17, MS15b, PL15c, SLS16, TES15, Tan15e, Ued19, WWW^{+}17, WHC18, WJ19, WWC^{+}18, WH15, XTXF18, XFTX19, XDG^{+}19, YF17, YFLF17, YLZT17, ZAZ19, ZHY16, ZSL17, ZWY19, ZZC^{+}19b]. Game-theoretical [GLR^{+}18]. games [AAA15b, AAA16, AE17b, BCd15b, CBGGV17, CYGL18, CP15c, DLD16, DJW19, DLY^{+}17a, Du19, ES15, EEST16, GH17, HI15a, KCSD^{+}18, LWZ19a, LZ17d, LGL19c, LJ17, MWL^{+}18, MCZ15, MN15, NT19, NR18b, NR18a, PSSX17, RCW^{+}18, RSGFV16, Tan17a, Tan17b, WWN^{+}15, WJNW15, WLWA18, WZ18b, WV18, ZWL16]. GameTheory [CBGGV17]. Gamma [BB15a, KAG15, AM15d, Bag15, CE15, CP15b, Che15a, Che15b, Che16b, Che16a, LSM15, YC15]. Gander [PIP19]. GAOR [MLW18]. Gap [KwC15, FVMO15, NMNTBXVD15, SWSS18, YY15]. GARCH [lZA^{+}19]. Gardens [ADMV19b]. Gardens-of-Eden [ADMV19b]. Gardner [FCM17, dlRGB16]. Gargantini [PV19]. gas [BKT16, CS19, DKL15, ERFE16, ET15, FFG^{+}16, FS16, GR17b, KSV15, PFH^{+}15, RZMB17, SBH17, SR15f, SSB15b, TB15, WLJ15, WXM^{+}18, TD18]. gas-kinetic [PFH^{+}15]. gas-liquid [KSV15, WXM^{+}18]. gasdynamic [AIIK15]. Gate [BS19a]. Gate-MOSFET [BS19a]. Gauss [ASKK19, ALQ15, Dan15a, EHS17, HZ17a, LZH16, LZZ^{+}19, She18, TTLX17]. Gauss-type [HZ17a]. Gaussian [DZ17a, HC18, Jar15, LPKS16, LO19, NLC19, NKL15, Tha15, WDLL15, WCS16, WLL16, YZ18b]. Gaussian-wavelet-type [NLC19]. GBS [BM15b, SIA15]. GCA [HDG15]. GCI [YLCF17]. GCI-groups [YLCF17]. GDN [ZZX18]. GDN-stability [ZZX18]. gene [HWBL19, PBB18]. General [DAS17, HW18, KAG15, PSPÁV15, SAJ18, ZL19a, ZL19b, ABGS15, AZ14, AVC^{+}19, AK15c, AZ19, BE19, BMC19, BM18a, Boy17, BFIJ15, CM18, CZ17c, CT17d, DAQI15, DHS19, DC16, EAAAS15, FÖB19, JAN16, KZR17, KAS16, LY16a, LY18a, LRH15, MR15b, MJSA19, MAG15, PP15a, PR15b, PZ19, Pro16, QZL18a, Rah19b, RGA19, SJ19a, SG15, STD15, TYD15, TLZ^{+}19, Won15, YKJY17, Zha15f, ZS16c, ZSY17, ZC19a, ZG15a, KJR^{+}18]. general-constrained [LY18a]. generalised [PBB18]. generalist [JAU15].

Generalization [SV15b, Xu17b, ACTV15, Dis15, DLY17b, EHS17, Isc16, KTA15, MM18a, OTE15, SL15a, SGG15b, ŚKK^{+}18, YZ17a, ZWT15]. Generalizations [Dzi15, GDF^{+}17, MMM18, Shi15a]. Generalized [Aca15, bBjJ16, BLM^{+}16, CIPT15, Ely15, ES16, FPUK15, GT15a, GLL15a, HRS15, II18, Jeo16, jJ15b, JCX19, KM18a, Kim19, LLP17, Li17e, LZX19, Mac15, MGCAPP16, MS19b, Muk16, OB18, RT18a, RT18b, TW15b, WW17, YLLH19, ZY19, AED15, ACK15, ALQ15, AK15a, AO15c, BS16, BR16a, Ben18, BPR16, BMB15, BS15c, BCD^{+}19, CL19b, CC16, CM15a, CMSZ18, CC19, CM15b, Cim18, CCGV19, DMP19, DLF15, DK15b, DS19, DT16, DYW15, DGZ15, Den15, DG18a, DÖA15, DM18, DD15, EAAAS15, ES17a, tFyZ15, FZLZ19, Fen15, FLT18, Flo18, Fuk15c, GLW18, Hag15, HAC18, HA16, HEA15a, HEA15b, HEA16, HLZ15, HP15a, HM18a, HXLC15, HWXC17, HM18b, ILP^{+}15, JK15, JPY16, JSW^{+}18, Kat15, KM15b, KwC15, KR16, Kud16, KF15, Lam15, LW15d, LQW^{+}15, LY15c, LTY16, LZ17b, LMYW17, LGL19a, LZLG19]. generalized [LZ19e, LWW15b, LL17c, Liu15a, LSW16, LLX18b, LJ15b, LFY18, LHX15, MWWM15, MGL19, Mal15, MZM17, MYH15, MP19b, MDG18, MD15c, Mos16, MD18b, NL18b, PSZ18, PÖB15, QDD15, RM18a, Rez15, RMSD16, RC17b, SAC18, SFLM19, STB16, SS17, STS15, SHLW16, SSZ^{+}18, Sil16, SGG15a, SGY^{+}15, TXM15, TTP16, Tur18, UDA15, WCS16, WTXZ16, WK16b, WYCZ17, WTJ17, WXLL19, WHWW19, WG15b, XL18a, XM15, XzXqT^{+}15, YLCW18, YW18b, YZ18a, YX19b, YwFW^{+}19, YZ17a, YYP18, YWZ18, YZ15b, ZLW19, ZPZ^{+}19, ZH15a, ZLL15a, Zha16d, ZHJ^{+}17, ZCXX18, ZXZB19, ZL17b, ZH19b, ZS15e, ZC18, ZZL15b, Zhu16]. generalizing [BC16]. generally [XKG15, XGQW19, YSL18, ZKKZ18]. generated [AAKT15, TZS17]. generates [DLH15]. Generating [CS16a, ZCU17, CM17a, Liu15b, Sim15, Vil16, ZZ17b]. generation [BB15b, CMW15, MPHG15, OFwCW19, RLGGAV17, Rom19, SGFCRM^{+}18, Tan15a, XS15, YG15a, ZCL18, ZBME16]. generations [Ter15]. generator [BB18b, Kok15]. Genetic [BKZA18, CWD^{+}19, DKM17, FFN18, FLCSV18, HXW^{+}19b, IKHS16, JPZ^{+}19, KLZT15, LABK15, LS15b, LZG16, MMG15, Sch16, SBG^{+}17, ZZZ15a, ZXDP18]. Genocchi [HASA15]. genome [DKM17]. genomics [AGC19]. genuine [GM15b]. genus [LG18]. Geo [KGGAV^{+}18]. geodesic [BM19c]. geodetic [GIK19]. geographic [GSV17, Ort15a]. geographical [KEE17]. Geometric [ZZW17, Ewa16, GFWH17, HARS15, LW19d, MMM18, RS16b, SG19a]. geometric-arithmetic [MMM18, RS16b, SG19a]. geometrical [CK18]. geometrically [LZL17]. geometries [BBGR15, GCM^{+}18, MCZ15, TBSR16]. Geometry [Den15, KG16, CZ15b, GLS19, HW17a, KZA19, KSA19, LB15a, SR15f, WLQ^{+}19]. geostrophic [CCD18]. Gersgorin [KG18]. Gersgorin-type [KG18]. Gezhouba [YJYH15]. Ghorbani [Das16, DIS17]. Gibbs [ACRZ18, BCD15a]. Ginzburg [KL15, YMD15]. girth [LLTW19]. given [BF16b, CWD16, CWZ17, CGJZ18, CZ17c, DLZ18, DMH19, Gou18, HLX17, HHT19, LCM19, LP16a, MBW19, Pos17, SISV19, WWL18, WLQZ18, ZL16b, ZWZ19b]. giving [GlH^{+}19, uRAD19]. Glaisher [WL16b, You15]. Global [CZLS18, CWFL15, CGAL19, DGZ15, EA15a, EO15, FW18a, FLMC15, FLYG18, GS17c, HQ15, HXS^{+}16, JGGK15, JL15, JX19, KM15d, LJW^{+}15, LLZ15, LZT^{+}18, LL16c, MLT17, MLG18, MDG18, Raz15, SSB15a, SSS17, SL15b, Tan15d, WZAF15, WLZ16, WHX18, WTUZ15, XWZ15, XGZ17, YSWZ17, YSLS17, YL19a, Yu16, ZS15a, Zha16c, ZG18, ZFL18, Zho15a, ZLL18, CFSGM17, Dai16, FHV15, FIG^{+}19, GYY15, GIS18, IT15, JG17, KwC15, KM18b, LSZ^{+}18, LH19b, fLLPY16, Naa15, bOqGyK^{+}15, Prz15, SCZ^{+}17, Wan17a, WHX15, WF15c, bWzZjH16, Yos19, ZLW15, ZLZ16, ZWaY18, ZBME16]. Globally [WDLZ19, WCH15, JLL15b, KG19, LS15b, TTK15, YSL19]. gluing [Kue15]. GMDH [SBG^{+}17]. GMHSS [DS19]. GMRES [LH19b, Xu19e]. go [CD15d]. goal [PABA15]. Godunov [CT15, CT17d, Zei16]. Godunov-type [CT15, CT17d, Zei16]. Golden [AAF15, CIPT15]. Goldstein [PRD19]. Gompertz [Zha15g]. good [LHG18, LT19, SC17, WY17, WX18]. goods [AEP15, BM19a, hCWW^{+}15, CCW17, CYGL18, CCW19, Du19, GlH^{+}19, HI15b, LJR19, LWWZ17, WWW^{+}17, WHC18]. Gordon [Tho17, jWgX18, WX19a, XZ19, Can16, JSLW17, LLL18c, MVRV19, Nag17, PFA15, RS15a, SLC18, SLW15, SC18b, Sun15b, ZXZB19]. Gordon-Maxwell [SC18b]. Gordon-Schrödinger [SLW15]. Gorges [YJYH15]. Gosper [Che16b]. Goto [UHN^{+}18]. Gould [KR15]. Gould-Hopper [KR15]. governance [HCS19]. governed [CH18b, HZLY17, LZ19d, Tan15b]. governing [CW17, LW15a, RM19a, RMK19, ZLC^{+}18]. government [Haf17]. Gower [QG19, STJS16]. GPU [EdLC18, GG18, GWK15, GWPA18, PS15c]. GPU-based [PS15c]. GPU/CPU [GG18]. GPU/CPU-based [GG18]. GPUs [FFGRSV19]. Grad [EL19a, dFGAN19, EL19b, EL19c]. Grad-Div [EL19a, dFGAN19]. grade [Awa15, DBV18, SST17]. graded [BZ16, DWRZ17, HSS15, LZL17, MB17, NND^{+}18, Rad15, SzSC19, WZWD15, ZLL15b]. Gradient [Bej17, Cho18, FLMR18, AB18, AKA19, BPR16, CL16a, CGWW17, CRZ19, DCW15, Dai16, DLH15, Dor16, HM18a, LP15c, MKRS16, Nak15, Rad15, RMA15, WSdSY15, WAS19, WY15b, XP17, XS19a, YDWZ19, YLNL15, YUK19, ZJZ15, dSRTZ18, CFR18]. gradients [LHHS16, Tha18b]. grafts [TBSR16]. granular [YM18]. granularity [YwFW^{+}19]. Graovac [Das16, DIS17, KST19]. Graph [ADG19, DESS15, GDZ18, GWXZ18, WTD^{+}19, BM19a, BCD18, CRC15, CHL18, CC19, CLWM19, DMG15, DV15, DSM15, DFSC15, Far18b, FHLL19, FTF16, GFWH17, GKS16, GDL19, GFD16, GDF^{+}17, tHBSfY19, HH17, HL18, ID15, II17, JYF19, KB18, KM15a, KST15, KS19a, LQW^{+}15, LMYW17, LGS^{+}18, LGL19a, LZLG19, LAAZH19, MCS^{+}19, MB15b, MMG16, PTO19, PJ18, QGH19, RC16, Rét19, SLF18, SY15b, SX15, TS18, Tra18, WY17, WGL18, WZ18a, WCD17, WZC17, XZC16a, XLYZ19, YL18, ZW19a, Ino16]. graph-based [BM19a]. Graph-theoretic [GWXZ18]. graph-theoretical [KM15a]. graphene [ZZLY15]. Graphic [YYL19, ZY18a]. Graphical [CGPR15]. graphics [PH17]. Graphs [CH19, KMS18, STD15, SISV19, AÇ18, ADMV19b, Alf17, AG15b, AK18a, APPE17, AVC^{+}19, ASZD18, ACP19, AC19, AH18b, ATX19, BLNSF17, BM19b, BRM^{+}19, BKKBS19, BS19c, CYZ17, CYY19, CRC19, CL15a, CL16b, CWD16, CWZ17, CZ17a, CLL17, CMSZ18, Che18b, CYW18, CR19b, CC18c, CLSS19, CT17b, CT17c, DM16, DMG16, Das16, DL17a, DGF17, DH19a, DVJ18, DSAB16, DSZY19, DIS17, Dob19, ES17b, FWZ19, FT18, FZZ18, FHK^{+}19, GW16, GLW18, GA17, GA18, GS16, GS18a, GDRP^{+}19, GPRCTRY19, GIK19, GHM18, GZhL^{+}19, GXCY19, GWX17, GL18b, Gut17, Gut18, HLX17, HHZ18, HV19b, HTW19, HLL18, HD16a, HLW15, HWL^{+}19b, HHT19, HKM18, Ili18, IB19, Jah17, JFT17, JQL15, JLW18, JYSR18, Jin17, JHL19, Jur17, KKTY17, KST16b, KMA16, LLM^{+}18, LLMS16, LLS15, LTY16, LWG16, LYW17, Li17f, LYS17, LLW17, LH17b, LWM17, LQG18]. graphs [LG18, LLL18a, LZ18e, LGQL19, LS19a, LCM19, LZ19b, LAAZH19, LL16b, LP16a, LW17b, LD18b, LZH19, LLTW19, LWZ15, LZ17g, MB19, MNJW19, MBW19, MK15a, MSdC^{+}18, MRS19, MD18c, NZZ16, OGC19, QZD16, QY15, SDAB17, Sha15c, SWG^{+}17, Sta17, SX15, SRT18, SJL17, SGDX18, SD19b, TLLD19, TZC19, Tep19, TX18, TDS15, TZS17, Tur18, VZ18, WTZL18, WTD^{+}19, Wan16c, WLQZ18, WWW19, WX18, WYL19a, WC18, WL18b, WS19c, XZC16b, XX19a, XLG18, XGG18, Xu19d, Yan17, YFL16, YSY^{+}17, YW18b, YFH19, YEMRV17, YQ15, YQT15, YFY17, YQD17, YLQ17, YHY^{+}18, YKLS19, YSZ19, Yue16, YWLL18, Zha19a, ZWL18, ZLSX18, YZZ16]. gravitational [LJK15, QLLY18, SRH15]. gravity [Dim16, MS18a]. gray [CLHC15]. Greed [Pla17]. greedy [ZTW19]. Greeks [JO15]. green [WC15b, CCLS16, KKG18, LJD17, PSPÁV15]. grey [LX19d]. Grid [BBGR15, JO15, RAV15, Boy15, How15a, KVS15, LR16b, PBMP15, Qiu18, RDG18, RG18b, SH16, SY17b, VVD15, YX19a, ZWF15]. gridding [GYZC15, LB15a]. grids [AP18a, Bac15, BNY15, BDH16, CJK17, CLR15, FM18, KCZK15, KJR^{+}18, LR19b, LR19c, PSY^{+}17, RC17a, RV16, SCY19, TZH15, Wil18, YW19b]. GRLW [HEA15a, HEA16, KZ16, LM18]. Gröbner [SS15a]. Gronwall [Fen15, LW19c, ZAP19]. Gross [Kri18, MDH16, WMLG17, WJWX18]. Grossberg [EO15, LCAA17, PRC^{+}19, ZDPZ15]. grossone [CPS18, Mar16, De 18]. ground [Hac17]. groundwater [YCLS19]. Group [HI15a, LM15, MFMK17, Ste19, eMA15, ACP19, AC19, CSZ15, DHW^{+}19, FHV15, Fio15, Fio16, GLL15a, GS16, GZD19, HAC18, HI15b, HH17, lJdW16, KM15d, Liu15a, Luo19, LJL16, QTTL19, RWY19, SJD^{+}18, TRRP15, WY17, WWK19]. Group-separations [HI15a]. groups [ACKS16, BMDS17, Bra15, DL18a, GS18a, HZYW15, JK15, LLL19b, ML17, RS19, YLCF17]. growing [Arm19]. Growth [YSWZ17, AM15a, BBD15, JAU15, LMF17, LDH18, PS16, RRTR15, SAES19, SD19a, TSA15, TY16, TÖGE18, XSLW17, XWZ15, YYM19, ZK18]. Grüss [APTS15, LT15, STS15]. Guaranteed [YJMH18, XD19]. Guo [RZD^{+}19]. Gupta [Mal15]. gymnast [ZLWS15]. gyrostat [ILPE17]. gyrostatic [RMD18].

H1N1 [LLL16a]. H7N9 [XSS^{+}17]. Haar [Bag19, AND17, EGB15, Hsi15, PJ15, SuR15]. Hadamard [SISÖ15, AT15b, BGMZ17, CGCK19, Che15c, Gav15, MWWM15, NNA15a, PWS17]. haemodynamics [Tor16]. Hager [WAS19]. Halanay [YWL15]. Halanay-type [YWL15]. Haldane [DDD19, Kha17]. Hale [BB19b]. half [ABLZ15, ABL15, CDJ19, CMT^{+}18, CC18c, DK16, Fuk15b, Gao17c, HV19a, HXM18, MSM15, MKS19, NR15, Reh17, SST17, SZJ19, WLZ19a, YJY16, ZDK15]. half-arc-transitive [CC18c]. half-boundary [Gao17c]. half-embedded [MSM15]. half-integral [NR15]. half-line [HXM18]. half-linear [ABLZ15, ABL15, CDJ19, DK16, HV19a, Reh17]. half-planes [YJY16]. half-quadratic [WLZ19a]. half-saturated [SZJ19]. half-step [MKS19]. Hall [EAY17, FAHZ16, WL16c, YZ17b, YL19c]. Hall-MHD [YZ17b]. Halley [AKKB17, CCTV16, KBM16, LSZ15, NBK16, Sha15e, WK16a, WK16b]. Halley-like [NBK16]. Hamilton [Hua16, Ied15, WYL19a]. Hamilton-connected [WYL19a]. Hamiltonian [ABI19, Ely18, FLYG18, HMD19, HGL19, KC18c, LWS18, LJS17, LHX15, Lv16, MD15a, MD18a, MM18a, NC16, SMZ15, SZ17, SZ16, VZ18, WG15b, YY17, ZZZ15b, ZY19]. Hamiltonian-conserving [HGL19]. Hamiltonian/skew [ZY19]. Hamiltonicity [JYF19]. Hammerstein [DN16, AD16, AST18, CB17, DN15b, DA19, MRQZ19, Mic15b, MM18b, MO15b, SR15a]. Hamming [ZGH17]. hand [ACMT15b, ACMT15a, Kac15, LWPL16, YZ19]. Handling [PMS16, AY16a, AP19, CCW19, ZLZ^{+}19]. Hankel [CLQ16, CSZ19, DÇO15, IS16, Mah18a, NCEGA15, XMX15, ZCEW18]. Hansen [KKB15]. hantavirus [XZG19]. haptotaxis [MT15]. Harary [HLW15]. hard [MZC19, Sun18, ZYH15]. hard-to-borrow [MZC19]. hardening [VB15]. Hardness [LLW17]. hardware [KTD15, KDTP18]. Harmonic [APK15, Sin17, Che18a, Ele15, GST15b, KM17a, Lam15, LR16a, MNNA15, MV15, PAK15, SA18b, ZQ18, uRAD19]. harmonically [Che15c, Isc16]. harmony [MJSA19, ACC15]. Hartwig [DKM15]. harvest [SSB15a]. harvesting [AO15b, AA16, DL17d, EFL18, HL15a, LLZ^{+}16, LB16b, LYM18, PSM17, QD18]. Hassell [XZCC15, ZS15c]. Hausdorff [FMSG15]. having [CNNA16, Kac15]. Hayat [PK18a]. HB [PW19b]. HB-AFT [PW19b]. HBV [XSLW17]. HBVMs [ABI19]. HCNNs [Yu16]. HD [SPGP18]. HDMR [BB15b]. health [DKN16]. Heat [BST18, Mah15, ZT15, ZS16b, AwEE15, ABS15, AB15b, ATK^{+}16, BKLW15, CKG^{+}16, DJL17, DBV18, DL15a, EAY17, EM15, GJR^{+}15, HJMB18, Har16, HLIK19, HMHA15, HCY15, HLH17, KHN19, Kun16, LP15a, LCL17a, MVS19, MPPP19, MKMC15, MR16a, NM19, PHRB18, PMV16, Pan15, Pov15, RZMB17, RMM19, RSC16, Sha12, SV17a, Ste19, Su19, TPS15, TY16, XCHW16, Xu19c, YDWZ19, YM18, YWYS18, Yan18b, YX15, ZjW19, ZL19a, ZL19b, ZS17]. heat/mass [HMHA15]. heated [GSDH15, HAMA14, Mah15, PK18a]. heater [MS18b]. heating [AKR15, BES19, KK15b]. Hedging [WL16d]. height [ATK^{+}16]. Heiles [ILP^{+}15]. Heine [PR15b]. Heisenberg [Pu15]. Hele [ÁL15a, CS15a, CK18, Kim15a]. helical [Ers16b, NKNA16, TBSR16]. helicopter [Ili19b, LZZL15]. helioseismology [CDP19]. Helmholtz [AAMC19, BRD15, CXL18, KMZ19, SJD^{+}18, TLHC17, WWH17]. help [IKHS16]. hematocrit [AAM19]. hematopoiesis [DZ16, GS15a]. hemivariational [LL15f, MG19]. Hénon [ILP^{+}15, dSSB^{+}16]. Hepatitis [KIZ18, WM18b]. hepatocyte [XSLW17]. herd [TS15]. Hermicity [HWZ17, RWH15]. Hermite [ACD18, AH18a, AK18b, BGMZ17, BCD15a, Che15c, DDH18, Gav15, GR15b, Han15b, Lu15, LYY15, MWLL17, NNA15a, PC16, SISÖ15, ZjW19]. Hermite-Legendre [ZjW19]. Hermite-type [MWLL17]. Hermitian [CY16b, DS19, tFyZ15, Liu15f, Liu18c, XL18a, YQ15, YLQ17, ZLL15a, ZY19, ZWT15, ZZL15b]. heroin [FLMC15, WWK19]. Hessenberg [MR15b]. heteroclinic [CFSGM17]. Heterogeneity [Lu19, NR18a, LWSX18b]. Heterogeneous [HCY^{+}18, DLJW15, HS15a, HZYW15, HZS15, KLW15, KKG18, LL15b, LZM^{+}18, LYDH18, LZY16, MM15b, Mah19, MS15b, MKMC15, WY18, YJMH18, ZSK^{+}19, ZZL16b, dNdOX^{+}15]. Heun [EJF15, EJF16, II18, Ras15]. heuristic [DKN16, KM19, WHX15, YJYH15]. heuristic-adjusted [YJYH15]. Heuristical [Kem18]. heuristics [MFF15]. hexagonal [CXL16b, LQZ16, MB19, MSI17, Rad17, RV16, Van15, YLWF15, YLWF16, ZAY18]. HFMD [LWPL16, WXP16]. hierarchical [AKK15, AP18a]. Hierarchy [WDZ17, CJ17b, Kud19b, MM18a, STL17, WG15b, ZW15b]. High [Cho18, DJL17, HMD19, JSM16, KNK15, KK15b, KS19c, LJD17, LW18h, MK16, Por18, QLLY18, TSZ19, Ton18, WLQ^{+}19, YCZ19, ZB17, Zha17a, ZZM18, ZLD18, AAH19, BCMT17, BB18b, BS15a, BM18a, Cat19, Che16c, CN19b, CLMT15, Deh16, DG18b, FMDZ19, FR16, FFTB16, GCH^{+}19, GL15a, HAC18, HM19b, cJzS15, KCKZ15, LZ16, LMMR15, MH17c, NSM16, PZ17, PP19, PV19, RSLL^{+}19, RME18, RV16, SAES19, SFC^{+}17, SzShG16, VVA^{+}15, WWH17, WW18, WR19, WXXY19, Xu19c, YGZ19, YU15, YLH^{+}19, ZZ18c, ZZX19, ZY15c]. high-accuracy [PZ17]. high-dimensional [NSM16, RSLL^{+}19]. high-frequency [KCKZ15]. high-index [Deh16]. High-order [DJL17, HMD19, KNK15, LW18h, Por18, ZB17, ZLD18, BCMT17, Che16c, DG18b, FMDZ19, cJzS15, LZ16, MH17c, PP19, PV19, RME18, WW18, WR19, WXXY19, YGZ19, YLH^{+}19, ZZ18c, ZZX19]. high-performance [FFTB16]. High-speed [KK15b, GCH^{+}19, YLH^{+}19]. Higher [ASHSC17, AKKB17, KBM16, SN15a, USCAAA15, ZCU15, AZL15, ABA18, BÁR15, CDL15, DK16, DS16b, GKN15b, GDS16, HQ15, Han18, HK18a, KD18, Liu16b, LMMZ18, LD19b, PRD19, SEKA16, SGM^{+}16, Tan15a, XY15a, XY17, XC15, YWH15, ZAP19, ZLH16, ZMH18, ZZ16c, ZJO^{+}18]. higher-dimensional [YWH15]. Higher-order [AKKB17, KBM16, ZCU15, AZL15, ABA18, CDL15, GKN15b, HK18a, LD19b, SEKA16, SGM^{+}16, Tan15a, XC15, ZLH16, ZZ16c]. Highlighting [FP18b]. highlights [ATV16]. Highly [FG18, BAM16, CZ15a, FMX15, FLS18, GNW15, HZ17a, LX19a, LX19c, LP16b, SM19, XMX15, YMMH15, ZYY17, ZYX19]. highway [Hac17]. Hilbert [BAK19, GFG15, HMS15, HXM18, LW18c, MCI16, SD19c, Wan15c, XL19]. Hilfer [AEB18, GT15c, WZ15a]. Hill [CCLS16, WHX15, Zho15e]. Hille [Kar19]. Hilliard [CJK17]. Hindmarsh [BK18b, GJX^{+}19, LWZ^{+}19b, PRA^{+}19]. Hirota [RC17b, TZTG17]. historical [WJNW15]. history [Kan15]. Hitting [BKKBS19, WGL18]. HIV [EA15a, FQMR19, LXT17, MK15b, PC17, SG17, WLZW15, WGLZ16, WJHA17, XGZ17, YSLS17]. HIV-1 [FQMR19, LXT17, WGLZ16]. HJB [BH15]. HLLC [DPL18a, Pel17, SLdLD16]. HLLC-type [DPL18a, Pel17]. HMGB1 [WM18b]. HMM [LSZ^{+}19a]. HODLR [SDWG19]. Hohenberg [BAV18, KR16, ZEF19]. hold [OLZ19]. Hölder [HMGC18, LSZ15, SGMH16, ZS14]. hole [MWY15]. Holling [CY18, NS17b, WPL15]. Holm [HGL19, YX19b]. Homoclinic [ZY15b, CFSGM17, LSfW15, Lv16, SNL19]. homogeneous [BMM19, BAK19, BB15c, CGPR15, HV15, LJYH19, LZL15b, LLdM16, Mah19, PBP19, PSPÁV15, Sun18, TM18, WL16d]. Homogenization [RL19, HK18b, RL15]. homotopy [HR18, Liu16a, SRJ^{+}15]. Hood [HMT15a]. Hopf [WL18a, AYAA^{+}15, AMN19, AVCCGAVDL15, CGC19, DKP16, JS15, LWL^{+}18, MHX17, PW19a, TS15, TXG15, WPL15, WF15b, XJZ^{+}15, XWZ15, YL19a, YLL15, ZT16a, ZXW16, ZZH15]. Hopfield [LZ15b, RN19, ZwX16]. hopper [GWPA18, KR15]. Hopping [FFGRSV19, MZ17]. horizon [Ied15, ZML^{+}17]. horizons [MN16]. horizontal [BKLW15, GLYT16, Kim15a, Tha18b]. hormetic [LX19b]. hormonal [LY17]. Hosoya [DSM15, GDRP^{+}19]. host [HJ18, MMK18, WLS15, WW16c]. hot [KRF17]. house [dCCGV15]. household [MPHG15]. housing [HW15]. hp [DS16b, JA17]. hp-cloud [JA17]. hp-discontinuous [DS16b]. HSLA [DDY15, JDL15]. HSS [LZ19e, RWW15, YLW15, ZCC15]. HSS-like [RWW15]. hull [NSR^{+}19]. Human [CLR16b, dNdOX^{+}15, CDFP17, HJP16, LWGS17, SNS^{+}16, SG18, STVCC17, SCZ^{+}17, SH18b, VEF19]. Humbert [WW17]. humoral [BF16a, LXT17]. humour [TSPF18]. Hungarian [RKQ19]. Hunter [GLW15]. hunting [LLL16b]. Hurwicz [YJK19]. Hurwitz [BB15a, SGG15a, TCY15]. Huxley [ÇT16, HEA15b]. Hybrid [BS16, EH15, JISI18, LTQF19, LS16, LXTF18, MR15b, Pan18, PCS17, WYLG19, AK16a, AJ17, AY16a, BL18, CXL15, DN15a, EH14, FR16, Gar16b, GG18, GYZC15, HDG15, HM15b, IS16, JIS15, JDS18, KSN15, KXZ19, KLZT15, Li17a, LZ17e, LYZL18, LHS18, LXZL19, MHZC15, MM04, MS17b, MAJR19, NFS^{+}15, NRS19, PGLG17, Rad15, RS17a, RXL19, Sab18, SH15b, SGKR19, SDK15, SGY^{+}15, TH19, WS19a, YFW^{+}17, YMMH15, YJYH15, ZZ19a, ZCW^{+}19b, SR15a]. Hybrid-delay-dependent [LTQF19]. Hybrid-driven-based [LXTF18]. hybridizable [KCK16]. hybridized [CDL18]. Hybridizing [PTO19]. hydraulic [CS18, RRT15]. hydrodynamic [FK18a, PPS25, RRR19, RRR20, SS18a]. hydrodynamical [Pu15]. Hydrodynamics [GSP15, FLL^{+}19, JMK17, YL19c]. hydroelastic [WL16a]. hydrogen [FG18]. hydrogeophysical [SH19]. Hydrological [KSP16, WF15c]. hydromagnetic [RMD18]. hydrostatic [BKL^{+}18, EdLC18]. Hyers [AO19, BCO15, Nam19, Oni18, WL15, ZSS15]. hyper [BGMZ17, LZKX19, SZ16, Tra18, WL18a, WD19, ZT16a]. hyper-chaotic [WL18a, WD19, ZT16a]. hyper-elliptic [SZ16]. hyper-rectangle [BGMZ17]. hyper-Wiener [Tra18]. Hyperbolic [BNS16, CN19a, DGRR16, Ade19, AHP19, ACSS15, Bac15, BB19a, BDH16, CGM16, CH18b, CN15a, GY15, GLL15a, HSSH18, Kem16, LJC16, MZM16, Mar16, MMV16, MTB18, MFMK17, NKS18, OFwCW19, QZL^{+}17, QDH^{+}18, SR16b, SW17a, SuSR15, SuSR16, SPS18, Vab15, YU15, Yüz16b, Yüz18, ZZM18, ZY15c, ZHS18]. Hyperchaotic [GMOGCCC15, MPB^{+}17, WZK15]. Hypercomplex [Fau17]. hypercube [EMM17]. hypercubes [CMM17, LX17, LHG18, QY17, XYZ18, YFLZ18, ZH19b]. hyperdissipation [YL19c]. hyperelastic [FLCSV18, Laa18]. hyperfactorial [WL16b]. hypergeometric [ACK15, ALQ15, BMB15, Cas15, II18, JDS18, KA15b, MNNA15, MMS17b, SS15b, WCS16]. hypergraphs [GL18a, GLQ^{+}18, KLQY18, LZ18d, LMZW16, ZM18a]. hyperholomorphic [BRD15]. Hypermonogenic [AOS19]. hypernetwork [CLHC16]. hyperplanes [BGMZ17]. hypersingular [DO17, LWQ18, LX19a, LX19c]. hypersurface [Kac15]. hypothese [AG15c]. Hypothesis [MR19]. hysteresis [BMC19, NEK18]. hysteretic [YZZ18].

I-type [RM18b]. i.e [LHX15]. IC [Liu15b]. IC-colorings [Liu15b]. ice [GZG^{+}15]. IDAEs [Pis18]. Idea [Rui18]. ideal [DP15a, GG18, Haz16, Liu15b, RDG18, SLF18]. ideals [AL19]. identical [CLZW19, LLCR15, XLJC18, dSVR17]. Identification [HL15d, LWSX18a, MRQZ19, DLJT16, GPRCTRY19, NEK18]. identify [FHPV18, WR15, WC15b]. identifying [CCYT19, Lai15, SJ18, XX19b]. Identities [ZL18, DÖA15, HASA15, KK15a, YZ18a]. identity [Bag15, Net15]. IDRA [LZZL15]. II [CY15a, Kim18, LB15a, NS17b, Sug17]. III [DDdF16, WPL15]. Ikebe [MR15b]. ill [CPST16, GS17b, JSG15, SGJS16]. ill-posed [CPST16, GS17b, JSG15, SGJS16]. illicit [GNR^{+}19]. illustrative [TB15]. ILU [Xu19e]. Image [FLMR18, GK16a, HGM15a, LY15a, LZKX19, ACK15, ASKK19, AHLR19, AY16b, ABB^{+}18, CC18a, CCX15, DG18a, DG18b, DMR15, EGH^{+}19, FMB^{+}18, FMB^{+}19, Guo16, HW19, Han19, KGS^{+}15, LL16a, LL17a, SPW15, SGFCRM^{+}18, SDK15, Świ15, yWyLyY^{+}15, WDG19a, WLZ19a, ZPRZ18, ZHW^{+}18]. images [ARACCA^{+}18, ABB^{+}18, DHW^{+}19, GHZ19, LJL16, TM18, TH19]. imaginary [AL18d]. imaginary-time [AL18d]. imbalance [KCKZ15, ZLZ^{+}19]. imitation [BN18, ZSW^{+}17]. IMLS [ZLL15b]. immanantal [YQ18]. immeasurable [Zha19b]. immersed [ÁL15a, GYZC15, LLL18d]. immersed-boundary [GYZC15]. Immune [dNdOX^{+}15, BF16a, CWFL15, DDD19, DHMT15, KN19, KPLC15, KLZT15, QZJY15, RLM19, SRJ^{+}15, WGLZ16, WJHA17, YJ19]. immunity [LS15b, LXT17]. immuno [RLM19]. immuno-chemotherapy [RLM19]. Impact [BK18b, BKT16, hCWW^{+}15, HL15c, MKRS16, MPHG15, WWN^{+}15, ZCFW19, AE17b, DMSS18, FFL16, LW15b, MSI17, SHHC18, SG19b, Tan15e, WGSX19, WWC^{+}15]. impacts [BBGR15]. impatient [Amm15, CLC^{+}19]. Impedance [NKKS16, AL18a, ASKK19]. impeller [CKKF16]. imperfect [LPJ18]. imperfections [RM19b]. Implementation [GSP15, VS17, ABMBR15, ARK15, CSV19, EdLC18, FFGRSV19, wG15a, JSG15, PLL16, UHN^{+}18, XW19, Xu19e, YWRP16, YCD17]. implementations [PS15c]. implemented [KDTP18, TKDF15, TD18]. Implementing [Ma15a]. Implicit [Laa18, Obr19, SC17, TZ18a, YDCS15, ASS16, BdS19, BMRS18, CLJ^{+}15, CWL19, CDL18, CJ16, DC16, FM18, GLM19, Gua19, HJMB18, HW19, HP15a, Ied15, JIS15, KLZ19, KK16b, LDH18, LW18h, LT18a, MK16, MMK17, MKS19, OL19, TM19, VBG^{+}17, WW16a, WSZ16, XZ19, Xu19e, YCH19, YG18, ZB17, ZZL18b]. implicit-explicit [BdS19, BMRS18, WSZ16]. implied [CLL^{+}15]. Imposing [ANEP19]. imprecise [Sku15]. improper [CR19b, GGAVGG^{+}18]. improve [AADLF15, XCCL18]. Improved [AP15, CXZZ18, GHW19, GQC15, LLP17, MNK19, NJTP19, RZSS15, RR18b, Sha15e, SGM^{+}16, WLX18, WLZ19a, WJZF19, Zha15c, ABLZ15, Dan15a, KZL19, LCW15, Li17d, LQZ^{+}18, LLz18, LCW19a, LMZ^{+}19, LJHW19, MFF15, RAV15, SSK^{+}19, SSB16, SM15d, WK16a, WZC^{+}18, WZF^{+}19, WZK15, ZHY16, ARS15, UKE15]. Improvement [wG15a, Zam19, CXW18, MY19, Oka18b, ZBKS16]. improvements [DBJC17]. Improving [AD15, AM15c, DGZ16, DMS^{+}18, GGAVGG^{+}18, GE16, KZ19, NRY16]. impulse [GWW15, LPZL16, LZKX19, Nak18]. impulses [AB15a, AOH17, AHO18, FZG19, GD15, JS16, JYS16, LK19a, LZ19f, LWSO19, RRRR16, SZ19b, YLHS17]. Impulsive [ANT^{+}16, GZ15b, LYH19, ZMC19a, ZS19a, AÖ16, AZ18, CLL16, CR19a, CWL16, DL17d, DG15b, GWW15, GhDwZS17, GZK16, KSN15, LSAR15, LSR18, LZCL18, LXD19, LLW19, LWW15b, LC15d, Liu16b, LZ18f, LXZL19, LD19c, MPS15, RZSS15, RVGS17, RZHY16, SM15b, SLLC19, SS16, SML18, SAT15, SLL^{+}16, TLGX15, WIF15, WA15a, WLG17, WZS19, WCH15, WNW^{+}19, XTXW15, XD15, XZL15, XWDW17, XXZY19, YL17, YWO18, YWZS18, YFW^{+}17, YYZ19, ZLHH15, ZSL15, ZS15b, Zha15f, ZYXF17, Zha17a, ZSY15, ZCX17]. impurities [SZO16]. incidence [BW15b, CKW17, LC16b, LJS18, SZJ19, TZC19, WLZ16, uRAD19]. incident [ARB16]. Inclination [MGK^{+}16]. inclined [Alk19]. Including [CFSGM17, Cas15, JAN16, PN15a, STVCC17]. inclusion [ABHR15, BCH15, CL16a, SDK15]. inclusions [AL18b, BT15a, CLL16, DK15a, GMOP18, LZ15e, LWSO19, LLB16, Mah17, MM18b, NE15, TPS15, TB17a, WIF15, YW16]. incommensurate [HCX^{+}17a, MGPPGCCV15]. incompatible [AMCMFJ15]. incomplete [AQ15, BB15a, JGKL16, KPP17, LSZY17, SY17a, SQL16]. incompletely [CCW17]. incompressible [AJMN18, BBG19, BFR18, Bot15, BS19b, DL16, Ers16b, HP17, JMW15, Laa18, LXL17, NS17a, NKNA16, Qiu18, SJW16, WYD^{+}18, WLFH17, Xu19e, YCH18, YL19c, ZFL18, ZSF17]. incorporating [CYLW15, LZ15c, YSLS17, ZCFW19, ZWL^{+}19]. incorporation [BAE15]. increase [Far18b]. increased [AKL19, NT19]. increasing [Bor15a, DCR15, Li19]. Incremental [RWW18, LS19b]. Indefinite [NHWZ15, CR15, FZL19, LJ15b, PS15a, PS15b, SL15a, TXM15, WA15b]. independence [ZC19c]. Independent [CDK17a, CYCW15, Bou18, BS19c, CYW18, Kar19, LMW^{+}19, PAM19, SRT18, WTZL18, Wan16a, Zad16]. Index [QYF^{+}18, ATV18, ADA18, AK18a, AOS17, AG17, AAH19, BBF^{+}18, BKL^{+}17, CXL16b, CLL17, CT17b, CZ17c, DD16a, Das16, Deh16, DV15, DDdF16, DIS17, Dim17, Ely18, FT18, GA17, GKS16, Gou18, HLX17, HKS19, HD16a, HLL16, HHT19, JLW18, KST15, KST16a, KKST16, KST16b, KKST17, KMS18, KST19, KZ19, LL15a, LY15b, LQG18, LBRS18, LGQL19, LCW^{+}18, LP16a, LL18b, LA15, LMMR15, MB19, MBW19, MMDG16, MMM18, QTF^{+}16, Rét19, RS16b, Sha15c, SWG^{+}17, Shi15a, STD15, SGDX18, SISV19, Tep19, Tra18, Wan17b, WWW18, WLQZ18, WC18, YCYL18, YW18b, YwFW^{+}19, YKLS19, ZWZ19b, ZCW^{+}19b]. index-aware [BBF^{+}18]. indexes [II17]. India [SG19b]. indicator [Hua16, ZSXW18]. indicators [HDG15, WLX18]. indicatrix [Sah19a]. indices [ARB16, AMCMFJ15, BHMS15, BCD18, BF16b, CXL16a, CWD16, DGNA15, DESS15, DSAB16, GFWH17, JQL15, LLMS16, LQW^{+}15, LY16a, LZ19b, LP15b, Pal17, Pal19, QZD16, Rad17, Rét19, SDAB17, SIA16, SX15, WWL18, WX19b, WCD17]. indirect [CYLW15]. Individual [LZ17d, hCWW^{+}15, GCLG15, LWSX18b]. individualized [SKCS15]. individuals [LSX19]. Induced [FW18b, DMH19, FTE19, FLL17, GJX^{+}19, GRS^{+}15, LWSX18b, LLY19, MMZ^{+}17, MM17, PY19a, PZ16b, QTXY16, QGH19, SSM17, STK15, SH18b, VVA^{+}15]. Induced-charge [FW18b]. induction [KK15b, PKHD18, Sar17]. industrial [ZKV^{+}15]. industry [KWH17, Sah19b]. Inequalities [CP15b, Che15a, Che15b, AED15, AÖ15a, AÖ16, APTS15, BGMZ17, BAK19, Bou18, CDK15, Che16a, CNNA16, DAS17, DD19, DK16, ES16, Fen15, GJNRG15, GDF^{+}17, GSVK17, HT16, Ied15, Isc16, KB15, KwC15, Lam15, LQW^{+}15, LW19c, LT15, LL15f, MNNA15, MPP17, Nis15, NNA15a, NNA15b, QM15a, QM15b, STB16, STS15, SISÖ15, Wan15e]. inequality [Che15c, DAS17, DLY17b, Gav15, GZ15a, HKS19, Jan15, JPTL15, KKAK15, KST15, LZH^{+}18, LJHW19, LjL15b, MG19, SZ19a, SLS^{+}18, Uça16, WZ16a, WwQqTW18, XZZ^{+}18, XCCL18, XS19b, YWL15, ZAP19, ZLW19, Zha18a, Zha15i]. Inertia [YQT15, JT17]. Inertial [BCH15]. inertias [YJMH18]. inexact [AM15g, Bot15, CC16, DLF15, HMX15, HM15b, RMA15, SL15a, WCZ15, WZ16a, ZM15c, ZM18b, BF15]. inf [dFGAN18, dFGAN19]. inf-sup [dFGAN19]. infected [LLL16b, SG18, WLS15]. infection [AVCCGAVDL15, FQMR19, GXH18, MMK18, MK15b, PRMC19, SSGG17, TXL19, WLZW15, WJHA17, WM18b, WSXR19, XSLW17, XGZ17, YW19a]. infections [IYM19, WXP16, WLG17, XZG19]. Infectious [FJ19, ARV18, GRS^{+}15, QTXY16, ZG18]. inference [YCD17, YYP18]. Inferring [WWW^{+}17]. Infinigons [Mar16]. Infinite [Dob19, BT15a, BB19a, Boy15, CLL16, CJ17b, DK15a, EO15, Gao17c, GSV18b, JYS16, Kan19, KA15c, LJ15a, LSR18, MAG15, Oka18a, PWS17, WN16, XL18b, YL17, YWZ18]. infinite-dimensional [WN16]. Infinitely [Lv16, CTT^{+}19, ZZL15a]. infinitesimals [Cal18, GGM18]. infinities [Cal18]. Infinity [OP19, BGP15, GK17, GLN15, LGK15, Mi17, PSPÁV15]. inflatable [SRBA15]. inflation [EG17]. inflow [Neu16, SH18b]. Influence [AJ15, GB19, HWL^{+}19a, SMI18, XLL^{+}19, BAR18, HS15c, LY17, LMW^{+}19, OTH16, PG15a, SK19, WLJ15]. Influences [Uzu17]. influential [BGS19, LWSX18a]. influenza [CX19, LLL16a, MMK18, SZJ19, XSS^{+}17, ZSW^{+}17]. information [CCW19, DGL^{+}15, Fan15b, GST15b, JGKL16, KEE17, LMW^{+}19, MS15a, MWL^{+}18, Ued19, ZLZ^{+}18, ZLLP18]. informed [AJ15]. infrastructure [SSM^{+}18]. Inheritance [Ban18]. inhibitor [YJ19]. inhomogeneous [Cho19a, Fio16, HGH19b, HL15d, LZWZ19]. Initial [LLY15, MTJ15, eMA15, AOH17, AHO18, AMO18, DHQ19, Ghe16, HXM18, LJZ18, LL15b, LD18a, Li18, MVS19, RSKB15, RS17a, RR18a, SM19, Sow18, TES15, TG16, Tan15e, VK15, VK17a, WZ15a, ZJJ18, ZZC19a, MAJR19]. Initial-boundary [LLY15, HXM18, SM19]. initial-value [RSKB15, RR18a]. initial-value-problem [eMA15]. initialization [TKDF15]. initializing [KDTP18]. initials [MWRT17]. initials-dependent [MWRT17]. initiated [RPJ^{+}18]. injecting [LZKX19]. injection [MKN15b, RP19]. injection/suction [MKN15b]. Injective [SY15b]. Inner [PHRB18, HP15b, PS15a, RVGS17, SRCA17, TLZ^{+}19, XCL18]. inner-outer [TLZ^{+}19]. innovation [YZD15]. innovative [lJdW16, JH17, SFC^{+}17]. inpainting [LL17a]. Input [GZD19, LHS18, AL15b, CLXW17, CYSY16, DZC15, DCJ19, FLYG18, GW15a, GWSC16, GXLL19, LSLZ18, LC15d, LLX^{+}19, MAR15, NT16, NHWZ15, Pai15a, Pai15b, QBC19, SZ15, SLLC19, SML^{+}15, WDLZ19, XD19, XZLL17, ZZLZ17, ZZL17, ZFZK15]. input-delay-dependent-free-matrix [ZZLZ17]. Input-to-state [GZD19, AL15b, GW15a, QBC19, ZFZK15]. Input-to-state-KL-stability [LHS18]. inputs [CSL19, LLZD18b, LL18c, ZCW19a]. insecticide [SWA^{+}19]. insight [LXW^{+}18, SRK^{+}19]. insights [FP18b, ZGH16]. inspired [HJP16]. Instability [GO15b, AYAA^{+}15, ELP^{+}18, HM19a, HGH19b, Kim15a, LLY19, PZ16b, SKS17]. instance [ZLZ^{+}19]. instance-based [ZLZ^{+}19]. instantaneous [AB15a, AOH17, AHO18, GD15, PP18b, SZ19b, YWO18]. insulated [PK15]. insurance [dCCGV15, KM19, PP15a]. Integer [MCY19, FNS19, Fuk15b, JXW15, NC16, SD17]. integer-factoring [SD17]. integers [Acu15, AFK15, BM15b, KS18, KAG15, Liu15e, SIA15]. Integrability [VK19, Fer15, FG19, RM18b, Tal15, WTXZ16]. Integrable [HA16, MR15a, Bra15, LXZ19c, MM18a, UBT18, Xu15c, XLW15]. Integral [EJF15, KI15, WWH16, ZC18, AED15, AQ18, AR16a, ABHR15, AIK15, AN15, AD16, AM15c, AST18, AE17a, AD17, AD19, AAMC19, Bab16, Bag19, BM15a, BSZ16, BB15c, BC15, BST18, Bou18, BHD18, CM17a, CGL15, CJ15, CD16a, CXL18, CZ15c, DAS17, Dar16, DN15b, DN16, DA17, DA19, DL17b, DGZ16, DÖA15, DJ15, ER15, EGB15, EMM17, FK17, FMX15, FLS18, Fuk15c, Fuk15d, Fuk15b, GT15a, GLM19, GVA16, GM15b, GG19a, GSVK17, HA16, HSAR19, HS18, HT16, Hus19, Jan15, KKT15, KJ15a, KE16, KAH15, KKD15, KSN15, LZS15, LW16, Li16a, Li17c, Li17e, LW18c, Li19, LW19c, LZM^{+}18, Liu18a, LHZM19, LJHW19, LP16b, LCOS19, MHN15, Ma15a, Ma17a, MM04, MS17b, MRE18, MS19b, MMV16, Mic15a, Mic15b, MNNA15, MH17a, MHZ18, MM18b, MH15a, MH15b, MH16b]. integral [MH16a, MH17b, MS19c, MS16a, MO15a, MO15b, NCEGA15, NE16, NR15, NNA15b, NE15, OR18a, OA15, ÖK19, PHM19, PMC15, Pan18, PPV15, QZL18a, RAH19a, RDLG16, RZD^{+}19, RMM19, Rui18, Sab18, SAC18, SR15a, Sak19, STB16, SSA19, SZ19b, Sha15d, SSA18, SS15b, SRS17, SR19, Tan15b, Tan15c, TTK15, Wan15e, WDZ17, WXXY19, WX16, Xy16, XZZ^{+}18, YÖ18, YCZ19, YTH19, YYM19, ZAP19, uDuI18, ZLW19, ZAW15, ZZMS19, Zho15d, ZW15e, fP15]. integral-algebraic [BM15a, SR19]. integral-type [Li16a]. integrality [DVJ18]. integrals [AHS18, BGMZ17, Che15c, CZ15a, DO17, ES16, GGAVGG^{+}18, HZ17a, Isc16, KB15, KMP19, Kan19, Kat15, Kud19b, LWQ18, LX19a, LX19c, LLdM15, LCOS19, MH19a, MP19b, STB16, STS15, SISÖ15, SHA15b, XMX15]. integrated [CLT^{+}15, HDG15, KEE17, SZT17, XTXW15, ZGH16]. integrating [SGKR19]. integration [BVJ15, GBJ19, GPHAPRW16, GBS15, JSLW17, JISI18, KS16a, KOR19, KPLC15, PS18b, Reh18, RXQ19, SSS17, SH18a, Tan15a, Tan15c, Tsi19, YWH15, ZX19, ZGZ16]. integration/differentiation [GBJ19]. integrator [HW17b, JZ19a, MAJR19, PAM19]. integrators [KE18, TSZ19, TZ19a]. integrity [PG15a]. integro [AED15, ANT^{+}16, AL18b, AEB18, AIG17, AP19, ALP15, BT15a, BS16, Bog16, CR19a, CH18b, DZ15, DO15, GAB17, HJKP18, HZLY17, tHBSfY19, KJ15b, KAS16, LX15, LX18, LY15e, LBB18, LFL^{+}15, LMD19, Ma17a, MMV16, MH17c, NL18a, OS15, Oka18b, Par15, PWS17, QXW19, QZ15b, RME18, SR15b, SR15c, Sak19, SM16, SPS16, SPST17, SAT15, Tom12, Tom15, WZ16b, WC16, WXWX18, WY16, Xu17a, Xu19b, YL17, YYM19, Yüz15, Yüz16a, ZM15b, ZM17, dSH15]. integro-difference [WC16]. integro-differential [AED15, AL18b, AEB18, AIG17, AP19, ALP15, BT15a, BS16, CR19a, CH18b, DO15, GAB17, HJKP18, HZLY17, tHBSfY19, KJ15b, LY15e, LBB18, LFL^{+}15, LMD19, Ma17a, MMV16, MH17c, NL18a, OS15, Oka18b, Par15, PWS17, QXW19, RME18, SR15c, Sak19, SM16, SPS16, SPST17, SAT15, Tom12, Tom15, WZ16b, WXWX18, WY16, Xu17a, Xu19b, YL17, YYM19, Yüz15, Yüz16a, dSH15]. integro-differential-difference [KAS16, SR15b]. integro-parabolic [Bog16]. integro-splines [LX18]. integrodifference [Pöt19]. integrodifferential [ASA19, CGL15]. integrodiffetential [Pac19]. intelligence [CYGL18, FLS^{+}15, SA16a, WL19b, ZCW^{+}19b]. intensity [YYWH19, YLZT17]. interaction [AKR15, AADF18, BY19, Dan15b, GG19a, HL15c, HHL^{+}19, Ili19b, Ili19a, KFHH15, KZT19, LX16a, LJK15, MM15a, MMIA19, Moh19, STK15, TLZ17, TSPF18, Wan19a, ZAZ19, ZZC15]. interactions [BR17, CCL^{+}15, CS19, DDD19, HJ18, LLT19, LZWZ19, MZS18, Mou15, WW16c, YJ19]. interarrival [MCPO15]. interception [MS15c]. interconnected [CHL15, FXXC19, GZC19, KLKU18, LW17a, RZHY16, XZLZ17]. interdependent [LWZ18, WWW^{+}17, ZL19c]. interest [CLR16a]. interface [Ade19, ÁL15a, CM18, DL15a, Hua15, Laa18, LZ17a, MKMC15, NY15, PSY^{+}17, uIHA18, WHSS15, WQ15, ZW15a]. interfaces [Kim15a, RZMB17, ZM19]. interference [HYJ19]. Interior [Sar17, ALP18, Che19, GWY15, JPTL15, PCS17, WS19b]. interior-point [JPTL15]. interleukin [DDD19]. intermediate [AKT19, JA17]. intermittent [FYZ16, FYSC18, GWXZ18, JZ19b, LZjy18, LS19c, LCLZ19, WHTW17, ZHWW17, ZZC^{+}19b]. internal [BR15, Har16, KRB15, LY17, NM19, RS15d]. internally [GSDH15]. International [ZS16b]. Internet [Cao19, PW19b]. interpolant [Han15c]. interpolants [Lu15, SST15]. Interpolating [Rom19, SW17c, LCL17a, SA19, SN15c, TW15a, WZC^{+}18, WZF^{+}19]. Interpolation [LLG^{+}15, ACD18, Arà16, Bej17, BCD15a, BK18d, Cat17, Dis15, DQW^{+}19, Erb16, GW15b, Han15b, Han18, HP17, KCK19, LLO16, LX15, LY15a, LC15a, LQZ^{+}18, LLz18, LY18d, LHZM19, MZM16, RPB15, RV16, Seg18, SSN15, SJ18, SRK^{+}19, SV15a, WD16, XZB^{+}18, YLH^{+}19, ZYY17, Zha17b, ZZG18, ZjW19, Zhu18]. interpolations [AH18a, YDCS15]. interpolator [LW15b]. Interpolatory [ZZS^{+}19, ACRZ18, DJLX17, LC18, SDGU16, SuSB15, TTZX15]. interpolatory-corner [TTZX15]. interpretations [HARS15]. interruptions [JBS15, PL15a]. intersecting [KH17, ZZW17]. intersection [Far18a, HW17a, Wil18]. intersection-complete [Wil18]. interspecific [WXLL19]. Interval [NJM^{+}18, BH17, CHL15, CZS15, CCQS16, CGZ17, DPL^{+}18b, Eft15, GSVK17, Hua18, JSW^{+}18, LLP17, LH17a, Li17c, NMS^{+}18, PZ15, Pop17, RG18a, Sha15f, SLS^{+}18, YWS^{+}19, ZMQ16, dSH17]. interval-valued [DPL^{+}18b]. intervals [LLL19a, Oka18a]. intervention [XTXW15]. interventions [WSXR19]. intracellular [DZ17b, DZ17a, LXT17]. intraguild [HD17]. Introducing [Ano19w, AKT19]. Introduction [Gol15, HSAR19]. intrusion [Xu15a]. intuitionistic [HLFZ19]. intuitive [APMM18]. invalidate [ZSW^{+}17]. invariance [HST17]. Invariant [DZL15, WXZ18, WL19c, AAF15, ASM16, BGSCGL19, GMPP18, KKD15, LAVME18, LPHH15, RRW^{+}15, Rui18, RC17b, SJD^{+}18, SS18b, TQC15, yWyLyY^{+}15]. invariants [DMG15, GFD16, LD18b, LLJ19, Van15]. Invasibility [WWW15]. invasion [KKS^{+}16, Li15b, QFTT15]. inventory [BS15b, CLT^{+}15, PL15a, WQK15]. Inverse [CZ18a, GH17, GWY15, HLIK19, ZDJ19, BR16a, BCNN19, CFdCdS18, CKDC16, DYW15, DDM16, DBV18, DD16b, Fuk15d, GZZL17, HP17, HLT15, tHBSfY19, JH15, lJdW16, JDS18, KSN15, Kur18, Kyr15, Kyr17, LLMS16, LC15a, LW18g, LHHS16, Liu18b, Ma18, MS19a, MC17, MWH^{+}18, MD15c, MD18b, OP19, PSZ18, PS15b, PLL16, RMSD16, SdR19, She18, SJ18, Slo15, TB17b, Vis15, WMS17, Wan18, WS19b, WD15, XBC19, YX15, ZJY18, ZjA19, ZY19, ZwX16, ZCPM17, ddGL15]. inverses [CEG^{+}15, CLT18, Den15, FLT18, HLT16, HS17, MGL19, MZM17, Mos16, PS15a, SS15a, SS17, SKP16, XZS16, ZL15b, ZS15e, ZC18]. Inversion [Awa15, SKK19, ASKK19, AM18d, CH18a, De 17, GST19, Lia15b, MSH15, NCEGA15, RZW18, SH19, Yan18a, dFY15]. inversions [JH17]. invert [LG19]. inverting [MR15b]. investigate [Ant18]. investigating [Rui18].

Investigation [ATK^{+}16, ANGC17, BBD15, ŚKK^{+}18, ZWL^{+}19, BTP17, DDD19, FLL17, GR15b, IR18, IS17, NKNA16, RG15, SKS17, WLJ15, WX19a]. investigations [ASASM16, GCH^{+}19]. investment [DLJW15, HW17c, LZZ15, WRZ19]. investment-consumption [HW17c]. investments [ARR16]. investor [XTXF18, XFTX19]. inviscid [PMC16, RDG18]. Invited [TT16]. involution [RD15]. involutions [BR16b]. involved [RR18b]. involving [AK16a, ANU16, All18b, AM18b, AJP18, Bag15, Ben18, Bor15a, DH15, GG16, HWZ17, HLT15, HSSH18, KAI19, KM18a, KK15a, LJ19, MNNA15, Mor17, MD15c, NMM16, SGG15b, WHSS15, bWzZjH16, XC15, Xu17b, YWO18, ZS16a, ZL18, dlRGB16]. Ion [CM16, BdS19, Wan16b]. ion- [Wan16b]. ions [KSKW^{+}19, Mis15]. iOS [DLCV17]. iota [FKA17, YW18a]. irrational [LXZ19b]. irregular [DA19, Dob19, Gut18, Prz15, WLQ^{+}19]. irregularities [MCS^{+}19]. irregularity [ADG19, ACP19, AC19, BLNSF17, CHL18, Rét19, TLLD19, XGG18]. isentropic [CT15, QLLY18, SS18b]. Ising [dFKMM15]. isochronous [TPCL15]. Isogemetric [ADSS16]. Isogeometric [HMT15a, RSAF16, CNK^{+}18, SDWG19]. isolated [XJZ^{+}15]. Isolation [dSVR17, Chy16]. isometries [MD15b]. isometry [YLX18]. isothermal [DL16, KF15, PBBF18, SBH16]. isotropic [NB19b, ZLL16, ZjW19]. issue [BKS18, DGRR16, LLL^{+}18b, ZS16b]. issues [BCI15]. Itô [ALL18a, RN18, SAS19, XKWG15, YSL18, YYM19, ZKKZ18]. Itô-type [YSL18]. item [BS15b]. items [RZR16, SCB15]. Iterated [ESS16, CEMR16, ES15, GVA16, Pac19, Red19, XZC16b, ZAZ19, ZTW19]. iteration [AKS15, ABA18, CR15, CY16b, CG19, DLBQ19, EHS17, FFT15, Gon18, GJNC18, Gua19, HSAR19, HM15c, KM17a, KM17b, LCZ19, LZZ^{+}19, MR15b, NH17, PN15b, RWW15, SGMH16, SPKC16, TTLX17, TLZ^{+}19, WN16, WL17, WZLP18, WC17, XL15, ZLL15a, Zha15b, ZWZ19a, ZVL19, ZWT15, ZCC15]. iteration-complexity [Gon18]. iterations [Bot15, CCC^{+}17, GK17, Jar15, LPZ19, PWS17, VS19, ZAW15, ZCU17]. Iterative [AHO18, ARGF15, HWL^{+}19a, KC18b, LL15b, LWSO19, MM18b, AOH17, ASHSC17, AJC^{+}18, ACDT18, AG15c, ABMA19, ACTV15, ACTV17, BCST19, BCMT15b, BCMT17, BCT15, CCTV15, CP15a, CP19, CB17, CEHVT15, CGMT16, CS15b, EAAM^{+}15, GL15a, Guo16, HP15c, HM15d, HM18b, HM18c, HMT15b, JK15, JLX15a, JSG15, KJ15a, KZZ15, LL15c, Lia15b, LA15, LT18b, Maa15, MC17, MSHG16, Mic15a, NWNA15, NWN15, PG15b, PKB18, PIP19, PV19, RAV15, RS17b, RB15, SN15a, SS19b, SGJS16, SMM18, SLTH15, SG15, TTP16, USCA15, USCAAA15, UKS^{+}16, WSZ18, XY18, Xu15b, YW15, YJK19, YZZ18, Yun15, ZYH15, ZD17, Zha18a, ZM16, ZS19b, ZZL15b, ZSF17]. iteratively [ASKK19]. Ito [WXW18]. IV [Dim17]. IVP [RSKB16]. IVPs [FR16, Kac15].

Jacobi [Ied15, BS17, Hua16, Lia15b, LZZ^{+}19, SK17b, Sin18, SRS17, TTLX17, WD15, WW19, XBC19, dFKMM15]. Jacobi-collocation [SRS17]. Jacobi-type [Lia15b]. Jacobian [MSL15, YG15a]. Jacobsthal [TW15b]. Jain [DD16b, OTE15]. Jakimovski [SV15b]. Janowski [SA18b]. Jaulent [GR15b]. Jeffreys [dDSI19]. Jensen [CDK15, KKAK15]. jerk [RHJ18]. Jimbo [LL18a]. job [Abd15, ZXDP18]. jobs [MTY16, RKQ19, WYW16]. join [CC19]. joining [SJ19b]. joins [OGC19]. Joint [CLW18, RZR16, SH19, LYXL15, SHLL18, YLX18]. Jordan [Nis15, She18]. Josephson [ZMA^{+}18]. judgment [Lu19]. Jumarie [DQW^{+}19]. jump [AV18, dCCGV15, CHL15, CYSY16, CL16d, DHX^{+}19, GHQ17, HXW^{+}19a, HZK^{+}19, JGX15, JGKL16, JSW^{+}18, Kim19, KML19, KPP17, LSCK15, LWZ16, LSZY17, LQC^{+}18, LSZ^{+}19a, LDS^{+}18, LXC^{+}19, LZHC18, MC15a, MJL16, NY15, SYZP16, SY17a, SMZ^{+}17, SWSW19, VMLO19, Wan16a, WMZ16, WLZY17, WHWW19, XKG15, XSSW19, XGQW19, YSL18, YKJY17, YLP19, YYS17, ZLDZ16, ZS16c, ZSS16, ZKKZ18, ZCAN19, ZCCZ19, ZMC^{+}19b, ZBL15, ZYWW16, ZDF17, ZXX^{+}19]. jump-diffusion [dCCGV15, ZBL15]. jumping [CZM18, GLQ^{+}18, KBR15, WLH^{+}18, XKWG15]. jumps [CLW18, Gao19, KLZ19, LKZ18, LC15c, LB15b, LB16a, LB16b, LXZL19, MHM15, MZM15, MDZ17, PYZS18, QD18, TRP15, YYWH19, Yue19, ZYZ16b]. junction [WHSS15, ZMA^{+}18]. junctions [BMT18]. Jungck [PN15b]. just [HTW19]. juvenile [LPL17].

K-SP [SNAR15]. Kac [PRD19]. Kadomstev [VK19]. Kadomtsev [EO18, RC17b]. Kalantari [GK17]. Kalman [AAF15, AAHT18, Gao18, ZWY^{+}15a]. Kamenev [BMO15]. Kamenev-type [BMO15]. Kansal [CN15d]. Kantorovich [Acu15, AFK15, BA17, CZ16, DDM16, DPD18, IAK15, Kaj18, MNK19, SIA15, Sil16]. Kanwar [CN15d]. Karamata [Reh18]. Karman [AJC^{+}18]. Kato [WL16c]. Katugampola [VV18, ZBBW17]. Kaup [KI15, MM18a, Xu15c]. Kawahara [HP15a]. Kawarada [Bea19, PS18a]. KdV [eMA15, Bas19, BdlRT18, HHG19, KXZ19, Kud19b, OW18, RC17b, WTXZ16, ZM15a, ZZ18c]. KdV-like [ZM15a, WTXZ16]. Keeler [LR19a]. keep [KZT19]. keeping [Kri19a, WWN^{+}15]. Keller [OL19, ZZZ16]. Kernel [LS17a, ASASM16, AP19, BMDS17, BAV18, DWTT17, DWT17, FB19, GQC15, GFG15, HS15b, KMBA18, LWT^{+}15, LLW19, Li19, ÖK19, Pan18, Par15, QXW19, SAA19, TG16, WZ16b, XMX15, YTH19, ZW15e]. Kernel-based [LS17a, BMDS17]. kernels [And18, ANEP19, BAK19, FMX15, FDPP18, FÖB19, KKT19, LX19a, LX19c, NL18a, PHM19, RF18, Xu17a]. kernels-based [FDPP18]. keyhole [PKHD18]. keys [ZXDP18]. Khattri [CN16a]. killed [Arm19]. kilns [ZKV^{+}15]. kind [AM15c, DA17, DGZ16, FK17, HK18a, LL15b, LR15, MS17b, Mic15a, Mic15b, PMC15, Sak19, TZS17, YCZ19, YTH19, YX19a, Yüz16b, ZL15b, ZX16]. kinds [KLQY18, LLL^{+}18b, LPZL16, MRE18, ZC19c]. kinetic [FWLC15, PFH^{+}15, ŚKK^{+}18, dFKMM15]. kinetics [MP18, NH17]. King [Eft15, GKM16, LMMR15, MNK19, RA15]. Kinkelin [WL16b, You15]. Kirchhoff [CHRD16, FT18, HLX17, HLL16, HHT19, LP16a, MB19, QZD16, YCYL18, Zha16a, ZZL15a, Zho15a]. Kite [DL17a, TS18]. KL [LHS18]. Klein [Can16, Nag17, SLC18, SLW15, SC18b, Tho17, jWgX18, WX19a, XZ19]. knapsack [MFF15]. Kneser [BM19b, BHRW16, DB17]. knots [KF19, LX18, ZS16a]. Knowledge [WWDW17, BAE15, KPP17, RLGGAV15, WWSM19, WWC^{+}18]. knowledge-based [RLGGAV15]. known [LSZ^{+}19a, QG16]. knuckleball [ALDHA^{+}17]. Kobayashi [QBNF19]. Kober [Dar16]. Koiter [FA18]. Kolmogoroff [Fio15]. Konhauser [ÖK19]. Kopel [CMG18]. Korea [LLL16a]. Korovkin [BLS15]. Korteweg [AO15c, AM18a, DKKR16, Gao17b, HXM18, Kud16, Mis15, SGM^{+}16, SzSC19, WTJ17]. Kotera [WXW18]. KP [eMA15, ZZC15, ZSX15]. KPP [QDD15].

Krasnoselskii [AK15b, BZ15, GK15]. Krasovskii [CWLA18, GHW19, KKL18, LJHW19, NHWZ15]. Krein [GLS19, LSZ^{+}19b, ZZL16a]. Kriging [XZB^{+}18]. Kronecker [AZ19, AZ14, AY16a, BLK19]. Krook [Kri18]. Kruger [SPGP18]. Krugman [TES15]. Krylov [BKW18, LHLL19, VS17]. Krylov-based [BKW18, VS17]. Kudryashov [GR17b, TTXZ16]. Kummer [GST15a]. Kundu [WTSQ15]. Kuramoto [BT15b]. Kurchatov [AR16b, CP19, CSTH17, HVR17]. Kurchatov-type [CSTH17, HVR17]. Kutta [ABI19, BdS19, FT17, TZ18c, TSZ19, TZ19a, AAH19, CWL19, FT16, FKR15, HMD19, LZ16, Li18, MD15a, QZL^{+}17, RN18, TLL16, Tan18, Tsi19, WY16, YNMW16, ZS15b, Zha17a, ZS19a, ZZX18]. Kutta-Nyström [FT16]. Kuznetsov [PKB18].

L [BG18, FFGRSV19, KBK15]. L-BFGS [FFGRSV19]. L-shaped [KBK15]. labeling [GZhL^{+}19, MK15a, YHY^{+}18]. labelings [LMS19]. labelling [DSZY19]. labyrinth [WLJ15, ZBKS16]. Lacunary [Haz16, NP18]. LAD [LK19b]. ladder [KSD^{+}15]. lag [Nak18, SAES19]. lagging [DJL17]. Lagrange [AM18b, Erb16, IKMW15, KM15a, LZZL15, LZY^{+}17, LY18d, LHZM19, NS15a, SBL16, WLZ19b]. Lagrangian [BFR18, GSDH15, JOÖ16, JM18, KBK15]. Lagrangians [SK17b]. Laguerre [ÖK19, DOR18, YWZ18]. Lainiotis [AAF15]. Lake [KSK^{+}15]. Lakshmikantam [RZD^{+}19]. Lambert [Miy19]. laminar [ATK^{+}16, SH15d]. laminar-turbulence [SH15d]. laminated [AM15b, KCWK15, LZL17, MRF18]. Landau [KL15, RZMB17, YMD15]. Lane [MFMK17, SK17a, Tsi19]. Langevin [WL15]. language [ARK15]. Laplace [Awa15, CH18a, CCJ15, GIT15, JPY16, LXZ19b, Mi17, MSH15, NCEGA15]. Laplacian [FFT15, HLZ15, All18b, CHRD16, CCC19, CZ17a, CT17c, DMG15, DM16, DMG16, DVJ18, DFSC15, GW16, GS15b, HTW19, HLL16, Jia15, LS15a, LH17b, LWM17, LGS^{+}18, LMZW16, LPHH15, Lv18, MB15b, SD19b, Wu17, XZC16a, XZC16b, XHW18, XLS18, YQ15, YFY17, YLQ17]. Laplacian-energy-like [LPHH15]. Laplacians [GFE19, HL18, MB19, ZDK15]. Large [FFTB16, CYZ17, CK17, CSZ19, CLSS19, DCW15, FA15, GT15a, GZ18, JYF19, KÇ18a, KST16a, LG15, LWGS17, LCW^{+}18, Mi17, Neu16, NY15, PR15a, QTF^{+}16, TTD18, WHC18, WWHZ18, XZLZ17, ZW15d, ZCEW18]. Large-scale [FFTB16, CLSS19, DCW15, FA15, LG15, LWGS17, PR15a, WWHZ18, XZLZ17, ZW15d]. largest [CL15a, Ili18, SD19b, WS19c, YKZ15]. laser [PKHD18]. Lasso [LK19b]. last [SK17b]. Lattice [DSSC19, NB15, SK15, Tar15, AHLR19, CCL^{+}15, DPL^{+}18b, DZD^{+}16, FZZ18, GYZC15, HMSS19, KRF17, Kri18, Kri19b, LLMW18, LCS15, LY18b, LSW16, MHR18, Wan16b, WSLC17, WL19a, WYD^{+}18, WZZT19, ZZ16b, Bor19, RSC16]. lattice-Boltzmann [CCL^{+}15]. lattices [AHG15, BGMS18, BP17b, BCFR18, LPHH15, LP15b, WWW^{+}17]. Lavrentiev [JSG15]. law [CP15d, Liu17a, OTH16, ZS17]. laws [AIIK15, AZAM15, BDH16, Gla19, GL15b, HSSH18, KRB15, LFY18, Maj18, Mos16, NKS18, QZL^{+}17, XLW15, ZY15c, dlRGB16]. Lax [BP17b, Kud19b, TQC15, XLW15, ZY15c]. Lax-type [BP17b]. Layer [BKZA18, AL18a, ABN^{+}17, BKL^{+}18, DS15, EAY17, EM15, FXC17, GQC15, HAMA14, HHL^{+}19, KKG18, MS18a, MKRS16, Pan15, PK18a, SKS17, Sha12, SMR15, SV17b, WL16a, XP17, XCHW16, Zha18e, ZG15b, ZXD15]. layered [ZZLY15, ZLL16]. layers [LZ15d, LZ17e, PCS17, SN18a, SN18b, YDWZ19, YWH15, ZL16a]. layout [Mat15, Pal15]. LBV [And18]. leader [Ask18, FW18a, DDF18, MGCAPP16, WV18]. leader-follower [WV18]. leader-following [FW18a, MGCAPP16]. leaderless [HW17b]. leaderships [BM19a]. leading [HGG16, YLLH19]. leads [NT19]. leaf [TH19]. leafless [YL18]. leakage [LC16a, MRC^{+}18, NR16, RZSS15, SM15b, WCHW19, ZSZ17]. leaky [LWWL19, PAM19]. Leapfrog [AOS17]. leaping [LF15]. Learning [BRM^{+}19, APMM18, DCA15, HHL^{+}19, KSS19, KSA19, LL15b, DDF18, LWSO19, MNYL18, bOqGyK^{+}15, SS19b, TD18, Tsy19, WWDW17, XY15b, YW15, ZLZ^{+}19, ZZMS19]. Least [AM16, MO15b, Nak18, RS17b, YZ15b, CPP15, DBK15, FB19, FLL^{+}19, HGM15b, HM18a, HXLC15, HM18b, JT17, JYSR18, JCW^{+}15, KJ15a, KF19, LY15a, LCW15, LLC19, MLW18, VS19, WZC^{+}18, WZF^{+}19, YSD^{+}17, YFY17, YZYgG18, YWH15, ZS16a, ZJJ15, ZWLZ15, ZZHT17, fP15]. least-square [LCW15, WZC^{+}18, WZF^{+}19]. Least-squares [Nak18, YZ15b, CPP15, HXLC15, JT17, KF19, LLC19]. leaves [TH19]. Lebesgue [Zha17b]. Lecar [LTUZ15]. Leffler [CZLS18, KKMS16, LJW^{+}15, LW18d, LK19a, ÖK19, SC18a, YLHS17, Zha19c]. left [LQ15, LJ19, RBB19]. left-looking [RBB19]. leg [QZ15b, TZ18a, Wan17c, ZZ19c]. Legendre [DN16, SR15a, AKAA15, AMO18, AM18b, DN15b, DOR18, HS18, IH15, KS16a, LW18h, MYHS18, PS19, RME18, SS18a, SR15b, SR15c, Xy16, ZMQ16, ZjW19]. Legendre-collocation [AMO18]. Lehmer [TCY15]. Leidenfrost [KRF17]. Lemma [KJR^{+}18]. length [DDdF16, Dim17, HRLGÁ18, MWL^{+}18, SH18b]. length- [MWL^{+}18]. lengths [KPG16]. Leon [FMC15]. Lerch [SGG15a]. Leslie [QG19, STJS16]. leukemia [RMCG19]. Level [EGH^{+}19, ALZ16, BS15a, CCGP16, Chi15, DH19b, DL15b, FM18, GSV15, GSV18a, HDG15, KSK^{+}15, KLZT15, KMP15, Lac15, LWZ16, LL16a, LXL17, LJR19, MBM16, MVS19, MC15c, MMK17, MKS19, SSS17, WEH18, WV18]. level-disjoint [GSV15, GSV18a]. level-set [DH19b, LXL17]. levels [Mou15]. Levenberg [Che16c, HMF15, WZ16a]. Levi [TZTG17]. Leviatan [SV15b]. Levy [MDZ15, CWD^{+}19, CDL15, GDL16, HC18, KPP18, LZZ15, LC15d, LB15b, LB16a, LB16b, LXZL19, Liu19, Maj18, MHM15, QD18, WDLS18, XPG15, XHW15, ZYZ16b]. Lévy-type [KPP18]. Lexicographic [CPS18, BS19c, HHZ18, LH19c, Mao17, SDAB17]. liability [HW17c]. library [WWR^{+}15]. Lichiardopol [MY19]. lid [GG15c, SK15, SP15]. lid-driven [GG15c, SK15, SP15]. Lie [eMA15, ACE19b, ASM16, BCSG18, Bra15, CSZ15, DOZ18, EHdLS16, HAC18, Nas19, SJD^{+}18, TTXZ16]. Lie-group [HAC18]. Liénard [JS16, RM18b, SK17b, Sun15a, TPCL15, YD18, YL19b]. Liénard-type [SK17b, TPCL15]. life [Elm15, MHZC15, MMB15]. life-cycle [MHZC15]. Lifetime [GCLG15, Nad15]. lift [ZH19a]. light [MHR18]. like [AHM16, ACMT15b, AM15g, AG16, AM18c, CZP15, CXL16a, CCY19b, DSZY19, EHVV15, GKN16, GOA19, GJNRG15, HVR16, HYJ19, KM17b, Kri19a, KC18b, LH19b, LPHH15, LZ17g, NBK16, QLZ15, RWW15, SLWX15, WTXZ16, WH18, WGSX19, WN16, WF15b, WBL17, LM19, Zha15b, ZM15a]. likelihood [WWMX16, YYP18]. Limit [Ber17, HV15, LLdM15, LLdM16, YD18, DY17, Fer15, GLN15, GYC19, Jar15, KL16, KCKZ15, LZS15, QX17, Sun15a, YY17]. limited [Dan18, DGL^{+}15, GR19b, Liu15d, LWZ18, QTXY16]. limited-information [DGL^{+}15]. limiting [ZSK^{+}19]. Limits [CCGP16]. Line [WCZ15, BFJ^{+}17, BK18d, CMR17, GKS16, HMF15, HXM18, LWQ18, LG15, LZ15a, LW16, LGL19a, NZZ16, NKKS16, NRAEC19, PZ16a, RMA15, SX15, TX18]. line-search [NRAEC19]. Linear [AHP19, ACP19, HGL19, Kri18, LZ18f, SSG^{+}15a, SDP16, WZ18b, ZS16a, AwEE15, AF18, Ade19, ABLZ15, ABL15, Agr15, ANGC17, AZ14, AIG17, ASS16, ARSV15, ACC16, AJP18, AZ19, BS16, BMO15, BDSS15, BDSS18, BB19b, BCG19, CT17a, CR15, CCNQ19, CFSGM17, CJN19, CXP16, CLXW17, CDJ19, CWHY15, CYSY16, dCM16, CW17, CCL^{+}16, CL19c, CvRvdM18, CPS18, CNQRR18, CXW^{+}19, DLBQ19, DQH18, DHQ19, DB16, DBV18, DK16, DKBS15, DHS19, DZ15, DL17c, DLXZ19, Du17a, DEdlT17, ER15, EAAM^{+}15, Ely18, EMM17, FP18a, FZL19, Fio16, FFTB16, GA15, GZC19, GGHL15, GMOGCCC15, GCMR18, GSP15, GLN15, GSD16, GZ15a, GSVK17, HN14, HV15, HV19a, HHZ18, HP15b, HST17, HLL16, HM18b, JT16, JT17, JLL15b, JKK15, KJ15a, KRS18, KOR19, KBK15, KPT16, KAMD18, KLZT15, Lac15, LLP17, LWT^{+}15]. linear [LSZ^{+}19b, LL19b, LLW19, LSL19a, Lia15a, LZ17c, LLZ17, LZL15b, LB19, MM15b, MTY16, MGL19, MCY19, MM04, MC17, MW18, MLW18, MH15b, MH17c, MH17b, MK16, MKS19, MH19b, Nak18, NN19, Oni18, PMV16, PHM19, Pop17, Pos17, PK18b, QZW19, RNM15, RP18, RLZ19, Ras15, Reh17, Rem18, Sab18, SCR^{+}15, SM16, SHR18, SR15e, SZ19a, SMZ15, Sha15f, SY17a, SLLC19, SG15, SGG15a, SML18, Sug17, Sun18, SLS^{+}18, SCY19, TM19, Tch15, TN19, TB17b, VD16, VMLO19, VK15, VK17a, WY15a, WZ16a, WS16, WHTW17, WZZ18, WH18, WDG19b, WYLG19, WLZ19a, WZLP18, WN15, XPZ18, jXlZX18, XSSW19, Xu19d, YPZ18, YZ19, YZ16, YYP18, Yüz15, Yüz18, ZLLZ16, ZZ14, ZS15b, ZZ15c, ZGL15, ZMQ16, ZD17, ZZHT17, ZF18, ZWZ19a, ZJYD17, ZZL18b, ZM15c, ZM18b, ZYJ15, ZZL15b, dSH17, vLtTBI18]. Linear-quadratic [WZ18b, ASS16]. linearisation [MRR17]. linearity [Kur19]. Linearization [MGMM17]. Linearized [Li17b, BPR18, Gua19, ZZQ18]. Linearly [BMRS18, BS19b, HP15a, PW19a]. lines [SDS18, YRA15]. linguistic [YZD15]. link [DJW19, GM15b, LSX19, ZLWS15]. linked [FHM16, SCB15]. links [Chy16, LSCK15, SWSW19]. Liouville [AYD19, ARL15, CZ18a, CD15c, Deh16, FFT15, GZZL17, HL17, KMT16, MNR19, PSPÁV15, PS15b, WBL17, YW16, ZW15a, ZLD18, ZjA19]. Liouville-like [WBL17]. Lipschitz [ALL18a, BÁR15, GJNRG15, KG19, NT16, NPT19, TY18, TTK15, XPG15, YCH19, YLCW18, YSL19, ZZL16a]. liquid [Alk19, BBGR15, DL16, GR17b, Jai18, KSV15, QX17, Qia16a, Qia16b, RMD18, WXM^{+}18]. Lissajous [Erb16]. List [HWL^{+}19b, CRCVMS18, CCW17, Xu19d]. Lithium [BdS19]. Lithium-ion [BdS19]. Liu [GCAA18]. LMI [CLXW17, LSAR15, MRC^{+}18, WZYS15]. LMI-based [LSAR15, MRC^{+}18]. load [DF18, MSL15, VVA^{+}15, WR15]. loaded [WE19]. loading [FFL16]. loads [NND^{+}18]. Local [ABMA19, BGSCGL19, BF15, CLY17, DL16, HL15a, KBK^{+}16, LM18, NAM18, PZ16a, PeO18, RXQ19, SMR15, YX19b, YHY^{+}18, AJMN18, AM17, ACG19, AD19, AAMC19, BES19, CEHVT15, DWRZ17, DS16b, ES16, FFGRSV19, GI16, GZW15, HVR17, HSS15, Hua15, KGS^{+}15, KK16b, LJD17, LL17c, LF15, MA15b, MSHG16, MS16a, MB17, MRR17, PS16, QZL18b, RRV15, RLH15, RRW^{+}15, SMI18, STB16, SST15, SGMH16, SC17, WZ16a, WSZ16, WZWD15, Wei17, WYZ^{+}17, YMS16, YWS18, YCH18, YDCS15, ZWF15, ZSL17, ZM15b, ZM17, ZZL15b, DKKR16, SST18]. Local-in-space [YX19b]. local-support [GI16]. Localization [LVV18, AK15b, CEP15, LLL17b]. localizations [KG18, KMR18]. Localized [ZZBS16, ZHS18]. locally [ACKS16, BLM^{+}16, GSLC18, JM15, TTK15]. locating [GHM18]. location [CRCVMS18, GWC^{+}16, GKDHR19, MK15c, ZLLP18]. location-allocation [GKDHR19]. location-routing [ZLLP18]. Lockage [YJYH15]. locking [FA18]. locking-free [FA18]. locomotive [Hsi15]. locus [CGG^{+}18, SWA^{+}19]. Lodge [MV15]. Log [ZC19c]. Log-concavity [ZC19c]. Logarithmic [ROGS16, AD17, AAMC19, CZ15a, CXW18, ZBL15]. logical [ZZ16c]. logistic [CCH15, EYE18, HM15a, HC18, KSN15, LDD15, LB15b, Wan15d, XD15, XSLW17, XWZ15]. Logit [LWWZ17]. Loizou [PV19]. lollipop [Tur18]. Long [Che18b, Kan15, AG15a, BAB18, BS19b, GM15a, GHN^{+}16, HEA15a, HEA16, SSG^{+}15a, SW17c, TT18, WXZ18, WDG19b, YWRP16]. Long-time [Kan15, BS19b]. long-wave [WDG19b]. long-wave-short-wave [WXZ18]. longest [PP12, PP15b]. longest-edge [PP12, PP15b]. longitudinal [ZZBS16]. looking [RBB19]. loop [LXZ19b]. Lorentz [LW16, LH19a, QBNF19]. Lorenz [LK15, SNL19, WF15b, ZS15a]. Lorenz-like [WF15b]. lose [DZD^{+}16]. loss [SZ18b, SBLZ17, WRZ19]. losses [CLW18, XPZ18]. lossy [QBNF19]. lot [MMB15]. Lotka [GIS15, HJ15, KZ15, Liu15d, LC15d, MWY17, QD18, SRK18, Vad19, Wan15f, YWL19, ZJ15a, ZJO^{+}18]. Love [GM19]. Low [Pel17, STA^{+}18, HZ16, KDS16, LL17a, QX17, SRH15, WHM^{+}19, WL17, XS19b, Zhu15, LWK15]. Low-Mach [STA^{+}18]. low-rank [HZ16, WHM^{+}19, WL17, Zhu15]. lower [ABQ15, AK16b, GLW18, Jah17, JLW18]. lowest [ZS19b]. loxodromic [Nam19]. LPS [Rub19]. LQG [AM19]. LRBF [ZHS18]. LRBF-PSM [ZHS18]. Lu [LK15, ZLZ16]. lubrication [ACG19]. Lucas [CT18, SR16a, YZ18a, YZ17a, ZS15e]. Lump [MMIA19]. Lump-type [MMIA19]. Lupas [Sha15d, DP17, IAK15, KS18]. Lur’e [LZH^{+}18, RVGS17, SLZ^{+}16, ZZLZ17, ZCX19, YWS^{+}19]. Lurie [ZG15a]. Lyapunov [AR16a, AÖ15a, AÖ16, CWLA18, DK16, DG15b, GDM16, GHW19, ILP^{+}15, KKL18, LZY^{+}17, LJHW19, NHWZ15, WHTW17, WJZF19, WLL16, WBL17, ZLDZ17]. Lyapunov-Krasovskii [LJHW19]. Lyapunov-type [DK16]. lymphatic [Bhu15]. lymphocytic [RMCG19]. Lyness [DGZ15]. Lyra [SR15f].

M [Amm15, PK18a, SJ19a, YLW16, ZWD15, Amm15, SJ19a, YLW16, ZWD15]. M/M/1 [Amm15, YLW16, ZWD15]. M/M/1/K [SJ19a]. MAC [LR19c]. Mach [Pel17, QX17, STA^{+}18]. machine [CS16b, GYY18, HSKHA15, KLW15, KSK^{+}15, KSS19, MNYL18, PN15a, XY15b]. machinery [HYL^{+}19b]. machines [ZXDP18]. machining [Han16, MJSA19]. Maclaurin [SHN16]. macro [HS15c]. macro-economic [HS15c]. Macroeconomic [TT18]. macroeconomics [Vol15]. made [VB15]. Madja [WL16c]. magic [MK15a, MWH^{+}18]. magnetic [Ant18, BMC19, BGP15, DS15, JL15, MJ19, MGK^{+}16, MS18a, MNJM19, NEK18, PPS25, PRA^{+}19, QBNF19, RRR19, RRR20, RPJ^{+}18, Sar17, SV17a]. Magneto [JMK17, GS17c, LL16c, OTH16, SS18a, SMG^{+}18, YL19c]. magneto-hydrodynamic [SS18a]. magneto-hydrodynamics [YL19c]. Magneto-hydrodynamics-driven [JMK17]. magneto-micropolar [GS17c, LL16c]. magneto-porous [SMG^{+}18]. magneto-thermoelastic [OTH16]. magnetoelastic [GB19]. magnetoelasticity [GB19]. Magnetohydrodynamic [SS18c, LTL16, WLFH17, ZSF17]. Magnetohydrodynamics [EL19a, BLJ^{+}19, EL19b, EL19c, Kem16, SJW16, YJK19]. magnetoplasma [Mis15]. magnetostrictive [KCK18]. magnitude [FMRRT15]. Maheshwari [CN15b]. Maheshwari-based [CN15b]. Mahony [HAC18, ZP16, Ghe16, Kan16]. main [JFT17]. Mainardi [BT15a]. mainland [WXP16, XZG19]. majorant [BF15, Sil16]. majorization [CGJZ18, GA17]. majorizing [AK15c]. makespan [GYY18]. making [GLL15a, GOH18, HJP16]. malaria [CWFL15, FG15, KCL19]. male [YSLS17]. management [CLT^{+}15, LZL19a, LWW15a, MN16, PCW15, SZT17, WC15b]. manifold [ACE19b, Val15, WFZ17]. Manifolds [CB19, ACD18, BF15, NKP16, SA19]. manipulators [MND^{+}15]. Mann [SPKC16]. manufacturer [Haf17, ZWY19]. manufacturing [CLHC15]. many [HHZ15, Kan19, LLL19c, Lv16, MN15, RP18, ZZL15a]. map [BB15b, ZQ18]. mapping [DBV18, DLY17b, PL15b]. mappings [AT15a, AK19a, ARGF15, APK15, BCO15, Nak15, PAK15, PN15b, SPKC16, SKA16, TTP16, ZBR15]. Maps [BMGMC17, APMM18, CB17, EYE18, HP15c, IZ16, Len15, MF17, MFOP17, MPB^{+}17, Naa15, NJM^{+}18, dSSB^{+}16]. marching [WSZ16]. margin [CS16b]. marked [GCLG15]. market [FHM16, TDA17, ZLS16]. marketing [GLR^{+}18, RRS^{+}18, SRAC16]. markets [PP15a, YSX15]. Markov [Bir15, CCR18, CYSY16, CP15c, DLCV17, DHX^{+}19, GHQ17, HXW^{+}19a, JSW^{+}18, LSCK15, LSZY17, LQC^{+}18, LSZ^{+}19a, LDS^{+}18, SPP17, SYZP16, SY17a, SWZ^{+}19, Sku15, SMZ^{+}17, SC17, VMLO19, WMZ16, WLH^{+}18, WRX^{+}19, WHWW19, WC17, Xu15b, XSSW19, YYS17, YCD17, ZCCZ19, ZXD15]. Markov-dependent [SPP17, ZXD15]. Markovian [CHL15, CGZ17, CZM18, Fan17, HL15b, HZK^{+}19, JGX15, JGKL16, KBR15, KKD17, Kim19, KML19, KPP17, LXC^{+}19, LXJ^{+}17, LYM18, LZHC18, MC15a, MJL16, Obr19, OL15, PD15, QG16, QZCJ19, Sha16a, Shi15b, SWSW19, Wan15b, Wan16a, WLZY17, WZS19, WL16d, XKWG15, XKG15, XGQW19, YSL18, YKJY17, YLP19, ZPZ^{+}19, ZLDZ16, ZSS16, ZKKZ18, ZCAN19, ZC19a, ZMC^{+}19b, ZH16, ZYWW16, ZDF17, ZXX^{+}19]. Marquardt [Che16c, HMF15, WZ16a]. marriage [PZJ17]. Maruyama [LC15b, LYZ19, LT18a, RN19, RT18a, YG15b, ZGGY19, ZSY15]. mask [HM17]. masks [CZ17b]. mass [HMHA15, PMV16, Ste19, ZLW18]. mass-conserving [ZLW18]. Massively [MHCF15]. Master [LPKS16, LG19, SLZ^{+}16, VS17]. master-slave [SLZ^{+}16]. matched [LZ17a]. Matching [YL18, BNY15, CLC^{+}19, CLR16b, CL15a, CL16b, CWD16, CWZ17, FHLL19, LWG16, QZD16, ZL16b]. Matchings [CDK17a, FZZ18, Jin17, SGDX18, WTZL18]. Material [DL15b, ANU16, HK18b, HL15d, OTH16, SA18a, WWR^{+}15, XRB17]. materials [DWRZ17, DL15a, DNC^{+}17, FLCSV18, HSS15, KMR18, LWWL19, NB19b, NEK18, Van15, WZWD15, YWYS18, Zha16b]. Math [MAK15a, MAK16, Sab18, Tom15, VK17a, WL18a]. Mathematica [ABMBR15]. Mathematical [Ant18, BG17, CLP^{+}18, KN19, MK15b, PPS25, Par18a, PBB18, Pon18, QZL18b, RRR19, RRR20, SG17, SI15, TXL19, ZGH16, Zha16b, uRAD19, ARV18, AAE19, BKAAM15, CDHK15, CWFL15, CC17, CD16b, DW17, GNR^{+}19, GRS^{+}15, HTZY17, KCL19, LHJ15, NSN^{+}18, Neu16, PZLZ15, RMCG19, San19, SP15, SRP17, SRAC16, SCZ^{+}17, Tor16, Vid18, YJ19, ZZZ15a, ZDZ15, ZHWW15]. Mathematics [Ano19w, AZ19, Bag19, DN16, DSDL19, EL19a, FMB^{+}19, LLZD18a, PPS25, Pan15, PK18a, PP15b, Pon17, RRR20, Rho19, SN18a, YLWF16, ZZ15c, ZL19b]. Mathieu [GGL15, MS19b, SI19]. Mathieu-type [MS19b]. MATLAB [AV15, CSV19, DRS18, Kok15]. matrices [AF18, AAKT15, APS16, AV15, AH18b, BR16a, BCR17, CÁGGLP16, CEG^{+}15, CSZ19, CN15a, CLT18, CEP15, CKDC16, DK15b, DP17, Den15, EÖ17, FLT18, Fio15, GLS16, GGZR16, HWZ17, HHZ15, JXW15, JH15, JZ15a, JH17, KWGRP16, Kur18, LH17a, Lin15, LCJ15, LLL17b, LZZT18, MR15b, NC19, OP19, PÖB15, Pos17, RM15, RG18a, SZ18a, SS17, She18, SML^{+}15, SPST17, SPS18, SKK19, SCN15, Tan15a, TXM15, Tan15c, Vis15, WY15a, WS16, WZC^{+}18, WNF19, WW19, WZLP18, XZS16, XBC19, YG15a, Yor15, Zha18a, ZY19, ZL15b, ZS15e, ZVL19, ZwX16, ZCEW18, dF15, dFKMM15, dFY15]. Matrix [BPC15, DLD16, TG17, TG19, YQD17, AKAA15, Ace19a, AZ14, AU15, AM15h, AE17a, AZ19, BS16, BD15, CCGP16, CCY19b, CD16b, CT18, CFdCdS18, DLBQ19, DGF17, DS19, DR19, DG15b, EEST16, GGMPC15, GBJ19, Gua19, GZ15a, Hag15, HZ16, HS18, HS17, HGM15a, HM18a, HM18b, IB19, IÖ19, lJdW16, JT17, KS16a, KM15b, KOR19, KA15c, KZZ15, KK15a, KK16a, KC18b, KKL18, LWS15, LW18d, LK19b, LH19b, Liu18c, LZZ^{+}19, LZH^{+}18, MNNP19, Miy19, MSL15, NE16, NB19a, PSZ18, PG15b, PSPÁV15, PS19, RWH15, RM16, SC18a, SIA^{+}17, SID19, SMZ15, SPS16, SWY18, SY18c, SDWG19, TTLX17, TF18, WD15, WY15b, WL17, WZLP18, WLZ19b, WSZ18, XL15, XM15, XLYZ19, YQ15, YLQ17, YZ15b, Yua15, YZ15c, ZZLZ17, ZLW19, ZGL15, ZWLZ15, Zha16d, ZHJ^{+}17, ZWL16, ZDJ19, Zhu15, ZT16c, ZZ16c, dCSS17]. Matrix-Tree [YQD17]. matrix-vector [GZ15a]. max [CCGV19]. max-product [CCGV19]. Maximal [FLT18, HP15b, HFK19, KST19, MR16b, RGH^{+}15, SG19a, Wil18, ZLSX18]. maximally [ZM18a]. maximise [SISV19]. Maximization [ARACCA^{+}18, HWL^{+}19a, AJ15, Els17]. Maximizing [Gou18]. Maximum [ARL15, HV19b, RzS15, WXY15, WWMX16, CYZ17, CYW18, FT18, GYY18, HM15a, HSKHA15, HLW15, KST16a, LPW15, SWG^{+}17, SX16, TDS15, WYW16, WWW19, XX19a, YSY^{+}17]. maximum-flow [HM15a]. Maximum-norm [WXY15]. Maxwell [MM15a, CWG15, CDL18, NI19, PMS16, SC18b, YZJ18, ZCCL18]. MCC [ZZ19a]. McKean [WWMX16]. MDCK [Kim18]. Mdm2 [PBB18]. Mean [DWS19, FHM16, LKZ18, LRH15, RN19, WWZD16, ZSY15, ZYX19, Bor15b, BCD^{+}19, Fio15, Gao19, KHN19, LC15b, LMD19, MS15a, NM15, PYZS18, RN18, SS19a, WDLL15, WDLS18, Xu15b, ZBL15, uRAD19]. Mean-field [DWS19, LRH15]. mean-reversion [ZBL15]. mean-semivariance-CVaR [NM15]. Mean-square [LKZ18, ZSY15, LC15b, PYZS18, RN18]. means [ACTV15, ABB^{+}18, BE17, CIPT15, DGN16, CVMGPPGC15, Jar15, RRTR15]. measles [BMT19, PRL^{+}15]. measure [ATV16, AKT19, BPC15, BMR17, CS17, DR19, HSAR19, KAH15, RAH19a, ZLWW15]. measured [ZZL17]. measurement [CSL19, HYJ19, LPJ18, PWFY18, SDP16, XXW^{+}19]. measurement-delay [CSL19, SDP16]. measurements [DPL^{+}18b, JZ19b]. measures [ADG19, CDK15, DSM15, DESS15, DCS^{+}19, II17, KE16, LZX19, SH19, WXZ18, WL19c]. measuring [DV15, MD18c]. Mechanical [AL18a, AZL15, FFL16, KK15c, LHHS16, TTV15, VPS15, YLW19]. mechanics [IR16, ZJJ15]. Mechanism [Kim18, Chy16, DZD^{+}16, FHV15, HXW^{+}19b, KTD15, LY18c, LMW^{+}19, TKDF15, WWDW17, WWSM19, ZHY16]. Media [KIZ18, WE19, ARV18, CDP19, DK15c, EB16, GO15b, GZW15, Har16, HZ17b, KS19b, LR17, LWGS17, LYDH18, LRL18, LZ15c, Liu18b, LCW19b, MHR18, PRA^{+}19, RL15, Zha18e]. Median [SM15a]. mediated [BF16a, WGLZ16, ZY16]. medical [APMM18, QTXY16]. medium [GB19, Jai18, KSD^{+}15, MSI17, NM19, Pan15, RNM15, SSB18, Sha12, SMG^{+}18, SCM19, YJW16, YCLS19, ZLL16]. Mehler [PR15b]. Meir [LR19a]. Mellin [DL17b, Kat15]. melodic [BDBKV19]. melt [SZO16]. Melting [AAE19, PHRB18]. membership [CGAL19]. membrane [AZL15, CM16]. membranes [Xu19b]. memductance [SAM^{+}16]. memories [MAG15]. Memory [CFdCdS18, SLLZ19, YFLF17, ASHSC17, AKKB17, CCTV15, Che15f, GSPH19, Kur19, LG16, LLX18a, Liu19, LA15, LTZ^{+}17, PC17, PSP17, RXZW17, RMB15, SP19, SLTH15, TT18, UKS^{+}16, WC17, Xu19e, ZL19c, ZL19a, ZL19b]. Memory-based [SLLZ19, YFLF17, RXZW17]. memory-saving [PSP17]. Memory-usage [CFdCdS18]. memristive [CZLS18, FYSC18, LC16a, LGC19a, LGC19b, MPS15, SAM^{+}16, TDL^{+}17, WCHW19, XZ18, ZZZY17, ZZ18a]. memristor [BPC15, CLPY18, GZLL17, LSZ^{+}18, LCLZ19, SCC17, WDG19a, XZLL17, YZH^{+}19, ZMA^{+}18, ZYC18]. memristor-based [BPC15, CLPY18, GZLL17, LSZ^{+}18, LCLZ19, SCC17, XZLL17, YZH^{+}19, ZYC18]. MEMS [KS16b, Zho17]. Menger [ACS15]. Menten [MMS17a]. menus [RLGGAV17]. merging [WGL15]. Meromorphic [LL18a, LH15, SS17]. meromorphically [SGG15a]. mesh [BCNN19, BDH16, DN15a, GAT18b, JSM16, Kok15, KKS^{+}16, MK16, PP19, Raz15, TMS15, UKK17, YNMW16, ZL16a, ZZ18b]. meshes [BLM^{+}16, BZ16, DEdlT17, Gia15b, KK16b, LZ17e, PP12, PP15b, SzSC19, SZL^{+}18, STA^{+}18, ZSY17, ZSY18]. Meshfree [LHZM19, EMM17]. Meshless [HHG19, LZL17, uIHA18, uDuI18, AD17, AD19, BLJ^{+}19, DWRZ17, HS15b, HSS15, KC19, LRL18, LLG^{+}15, MZM16, MB17, SJ18, WZWD15, ZYY17, KBK^{+}16]. messages [LdCC15]. meta [GI16, MFF15]. meta-heuristics [MFF15]. meta-model [GI16]. metachronal [FTE19, Pon18]. metacirculant [CZ19]. Metadomotic [MMG15]. metagenomic [KZ19]. Metaheuristic [KM18b, MMB15]. metaheuristics [PGLG17]. metal [NND^{+}18]. metaphysiological [HXS^{+}16]. metapopulation [FJ19, MTMB17]. Method [KI15, LjL15b, Rog19, ABMBR15, eMA15, eMBHA15, AF18, AKS15, Ace19a, AJ17, AJC^{+}18, ASKK19, AD16, ARB19, AD15, AJMN18, ANGC17, AMO18, ASASM16, AHLR19, ACE19b, AAS^{+}17, AP19, ÁL15a, AV18, AM16, ALZ16, AHY19, AL18d, AND17, ACMT15b, ACMT15a, AM15e, AG15d, AK15c, AR16b, AM17, ACG19, AD17, AKA19, AADMM16, Bac15, BAV18, BS16, BBF^{+}18, BBG19, BAB18, Bas19, BD18, BY19, BMT18, BNY15, BdS19, BKW18, BAR18, BF15, Bog16, BZ17, BPR16, BES19, BLJ^{+}19, BTWY15, BBD15, BG17, CGPR15, CGJV16, CL19a, CG15, CCL^{+}15, CR15, CY16a, CCNQ19, CCJ15, CLX17, CC16, CL16a, CKG19, CLJ^{+}15, CKW15, CD15b, Che16c, CD16a, CGW16, CW17, CR18, CZLM18, CXL18, CG19, Chi15, CDL18, CN15d, ÇT16, Cim18, CD16b, CJ16, CEHVT15, CXW^{+}19, DCW15, DLF15, DWRZ17]. method [DLBQ19, DA19, DL17b, DGZ16, Deh16, DS19, DH19b, DBV18, Dol15, DS16b, DLH15, DFJ16, Dor16, DLCT18, DWTT17, DWT17, DQW^{+}19, DL15b, DF16, ER15, EHS17, Eft15, EKS^{+}15, EAAM^{+}15, EKO19, ET15, EMM17, EA15b, EE18, FA15, FSW16, FZLT15, FG19, FBF18, FY19, FGPR18, FP18b, FLL^{+}19, GY15, GS17a, GYY15, Gao16, GZZL17, GLM19, GWK16, GGM18, GM17a, GQC15, GS17b, GN15, GM17b, Gol19, GWK15, GN19b, GVA16, GAT18b, Gua19, GE16, GZW15, GYZC15, GG19b, Hag15, HAC18, HHP18, HEA15a, HEA15b, HEA16, HZ16, HYB17, HP17, HVR17, HS18, HHGC16, HAY19, HSKHA15, HSS15, Hsi15, HLLT15, Hua15, HM15b, HYW^{+}15, HM19b, HMGC18, HHG19, Ied15, ID19, JOÖ16, JA17, JSM16, JK15, JDG18, JLY18, JMW15, JSLW17, JSG15, JIS15, KKT15, KC19, KCZK15]. method [KLZ19, KBK^{+}16, KZ16, KRF17, KJ15a, KM15a, KM15b, KM17b, KBK15, KZA19, KNBA15, Khl18, KK16a, Koc15, Kok15, KXZ19, KFHH15, KVS15, Kri18, Kri19b, Kue15, LX15, LLMS16, LPZ19, LY15a, LWQ18, LOOGB18, LW15b, LCW15, LW15d, LZS15, LWK15, Li16b, LW16, LWZ16, LL16a, LR16b, LR17, LL17a, LZ17a, LCL17a, LZCW18, LZ18b, LJL18, LQZ^{+}18, LLz18, LSLZ18, LML18a, LD19a, LM19, LR19b, LZ19d, LCW19a, LL19b, LYH19, LLW19, LHLL19, LJL19, Lia15b, LYXL15, LRL18, LVV18, LSZ15, LLG^{+}15, LFL^{+}15, Liu15a, LY16b, Liu16a, LXL17, LW18h, LMMZ18, Liu18b, LL19d, LS19b, LFC^{+}19, LMD19, LLL18d, LWW15c, LHX15, LYY15, LHWG16, LT18b, LPW15, MZ15, Ma15a, MGL19, Maa15, MDHD18, MD18a, MA15b, Mah18b, MS17a, MS17b, MHR18, MMV16, MYH15, Mia17, MP19b, Mic15a, Mic15b, MH16b, MD17, MNNP19, MDZ17, MB17, MDH15, MSL15, MS15d, MO15b, MRR17]. method [Nag17, Nak15, NMNTBXVD15, ND15, NFS^{+}15, NHWZ15, NRAEC19, NAM18, NEK18, ODP17, OS15, Oka18b, OPP18, Pan15, Par15, PC16, PZJ17, PSY^{+}17, PJ15, PVL19, PSP17, PKS15, PK18b, QBNF19, QDH^{+}18, QFTT15, RKQ19, RNM15, RSKB15, RSKB16, RS17a, RWW15, Raz15, RW17, RXQ19, RVE19, Reu15, RMM19, ROGS16, RME18, RGA19, RM19a, RMK19, Rui18, SS18a, SAC18, SuR15, Sae17, SAA19, SH15a, SR15b, SR15c, SDS18, SM19, SLC18, Sar17, SA16b, SR15e, SSN15, SL15a, SH16, SR16b, Sha12, SB15, SM15d, SPW15, SLW16a, SY17b, SZL^{+}18, SY18b, SLL19, SK17a, SGJS16, SJW16, Sil16, SPS18, SGKR19, SDK15, SRS17, SHHC18, STK15, SNL19, SG15, Su19, SM18, SLL^{+}16, SW17c, SDWG19, TM19, TSJ16, TRP15, TMPG17, TZF15, TG16, Tch15, Tha18a, Tha15, Tho17, TLHC17, TLZ^{+}19, Tom12]. method [Tom15, Ton18, TTK15, UKE15, USCA15, USCAAA15, UKS^{+}16, VLCT18, VRC15, VK15, VK17a, VBG^{+}17, VDV15, VS19, WNMC15, WSdSY15, WZ15b, WZM15, WHSS15, WZ16a, WS16, WZ16b, WMLG17, WHTW17, WLLK17, WZWJ18, WA18, WXWX18, WZC^{+}18, WW18, WLWA18, WZF^{+}19, WX19a, WR19, WSC19, WJL19, WNN16, WN16, WZWD15, Wei17, WW19, WY15b, WLLL15, WZLP18, WL19c, WHX15, WC15b, WD16, WLWD18, Xy16, XZB^{+}18, XM15, jXlZX18, XHW18, XCHW16, XX19b, XLR17, YDWZ19, YR15, YZD15, YLW15, YBZ18, YCH18, Yan18b, YJK19, YTH19, YW19b, YX19a, YZJ18, YKZ15, YG15b, YS16, YJSW16, YWZ18, YP18, YZZ18, hY18, YCLS19, YWH15, Yüz15, Yüz16a, ZMBJ19, ZA16, ZWF15, ZGGY19, ZL15a, ZLL15b, ZGL15, ZLL15a, Zha15b, ZZLY15, ZT16b, ZZCY16, ZZY^{+}17, ZGH17, ZD17, ZL17a, ZZ17a, ZCXX18, ZZG18, Zha18b, ZZ18b, ZJY18, ZF18, ZC19b, ZYSL19, ZWZ19a, ZZ19a, ZCW^{+}19b]. method [ZjW19, ZZ19b, ZCU17, ZBH^{+}15, ZY15c, ZSY15, ZHS18, ZZQ18, ZM15c, ZQG16, ZVL19, ZLK^{+}19, ZS14, ZWT15, ZW15d, ZYWW16, ZX16, ZZL15b, ZW15e, ZT16c, ZZ16c, uRIS17, Bor19, DGLE17, JO15]. method-based [DCW15]. methodology [CPS18, Laa18, VDV15]. Methods [ABI19, DGRR16, ABC^{+}16, ABD16, ASHSC17, ALP18, AKK15, AHP19, Ali15, ACDT18, ADG16, AG15c, AM15g, AM15f, AG16, AKKB17, AM18c, ABMA19, AP15, ASV17, ACTV15, ACTV17, AP18b, AAH19, Bab16, BMDS17, BBCI18, BCMT15b, BCMT15a, BAM16, BCMT17, BSZ16, BB19a, BK18c, BM19d, BST18, BRV18, BDH16, BCT15, BK18d, CFR18, CCTV15, CP15a, CP19, CY16b, CJ18, CGWW17, CH18b, CWL19, CN15b, CN15c, CN16a, CN16b, CN19b, CvRvdM18, CGMT16, CSTH17, CRZ19, DN15b, DN16, DJL17, DL18b, DKKR16, DMR15, EB16, EHVV15, FT16, FT17, tFyZZ15, Fan15a, FMX15, FB19, FKR15, FR16, GM15a, GML18, GHL19, GAB17, GKN15b, GKM16, Gia15a, Gia15b, GGHO18, GPHAPRW16, GR19a, GL15a, GIS18, HKL15, HMD19, HH15, HVR16, HP15c, HL16, HZLY17, How15b, How16, HH18b, HM15d, HM15c, HXLC15, HM18c]. methods [HHT19, HMT15b, Jeo16, JZ19a, JISI18, KZR17, KCK16, KWGRP16, KM17a, KBM16, KSL19, KDTP18, KKS^{+}16, KG16, KD16, KMZ19, KC18b, KDG19, LWZ19a, LLL16a, LL15d, LC15b, LB15a, LG15, LZ16, LZ17b, Li17c, Li17d, LJD17, LSjL17, LKZ18, LD18a, Li18, LW19a, LX19a, Lia15a, LZH16, LZ17e, Liu17a, LLL18c, LZ18f, LZZ^{+}19, LS17b, LA15, LZ17f, LJS17, LK18, MHN15, MD15a, MSHG16, MRF18, MLW18, MH17a, MK16, MMK17, MKS19, NCEGA15, NE16, NBK16, NCS16, NSM16, NRS19, Obr19, OR18a, OB18, Pan18, PR15a, PL15b, Pel17, PH17, PV16, Pro16, PIP19, PV19, QFS17, QLLY18, QZ15b, QDD15, QZW19, Qiu18, RP18, RR18a, RN18, RN19, RS15a, RA15, RT18a, RB15, RG19, SST17, SuRI15, SM15a, Sap18, SHR18, SN15b, Sha15e, SSK15, SB15, Sha16b, SA16c, SLTH15, SE18, TZ18a, Tan15b, TLL16, Tan18, TZ18c]. methods [TSZ19, TZ19a, TTLX17, TG17, TG19, TB17b, Tsi19, UKK17, VS19, WCZ15, WDLL15, WY15a, WK16a, WSZ16, WK16b, Wan17c, WZZ18, WH18, WAS19, WY16, WTUZ15, XL15, XY15a, XY18, XS19b, YZ18b, YLNL15, Yos19, Yue19, Yun15, uDuI18, ZV17, ZNV19, Zam19, ZSA^{+}15, Zei16, ZJJ15, ZS15b, Zha17a, Zha18a, Zha18c, ZX19, ZLZ^{+}19, ZS19a, ZRLX19, ZCQ15, ZCZM15, ZZX18, ZYX19, ZM16, ZCC15, ZZ19c, ZSF17, dFGAN18, dSRTZ18, vLtTBI18]. Metric [GHM18, LZ17g, ACS15, GGM18, JJ15a, KKTY17, LR19a, Laz17, LZCW19, RDLG16, RHP18, YEMRV17]. Metric-locating-dominating [GHM18]. metrics [DCK^{+}18]. metro [YYX^{+}19]. Mexican [FMRRT15]. Mexico [MTCV^{+}16]. MFS [MC17]. MGPBiCG [XM15]. MHD [BLJ^{+}19, BAR18, BGP15, CXW18, EM15, FAHZ16, JL15, PMV16, PP18a, SSB18, WL16c, YZ17b, YB15, ZWaY18, ZL17b]. mHealth [CDZ^{+}18]. Michaelis [MMS17a]. micro [BM18a]. micro-polar [BM18a]. microchannel [ATK^{+}16, JOÖ16]. microcirculation [PSP17]. microorganisms [RMD18]. microphone [KS16b]. micropolar [BGP15, CSZ15, FTE19, GS17c, LL16c, Mah19, OTH16, RMD18]. microscale [BR17, DJL17]. Microscopic [DLS19]. microstructure [MPW15, RC17b]. microstructures [LHHS16, XRB17]. microtemperatures [OTH16]. microtubule [AB18]. microtubules [CD16b, ZZBS16]. microwave [Oli17]. mid [LX18]. mid-knots [LX18]. middle [Yan17]. migration [LY18c, NKKK18, RCW^{+}18].

Mikhailov [Tal15]. Mikhailov-Novikov-Wang [Tal15]. Mild [GD15, APL15, GT15c, SS16, Wan15c, YB15]. millions [MHCF15]. Milne [Gao16]. Milstein [ARB19, WL19c, YG18, ZYSL19]. Mimetic [DO18]. Minc [Mor15]. Mindlin [RCM15]. MinHash [KZ19]. minimal [BM16, DDdF16, Dim17, JJ16, LG17, LCW^{+}18, STD15, WZ19, XZW^{+}15]. minimax [Fuk15d, Fuk15b]. minimization [CL16a, HSKHA15, MGPL15, PTO19, XHW18, ZPRZ18, ZWY15b]. minimize [GYY18, HHT19, MTY16, WYW16]. Minimizing [LP16a, Far16, Lu15]. Minimum [GXCY19, ZYH15, AL19, ADA18, Bor15b, CWL16, CSL19, FNS15, HLX17, HRLGÁ18, JYF19, KM19, KKST16, KKST17, KS19a, PZ19, RZH^{+}19, YZ19, ZTW19]. minimum-cost [KM19]. minimum-time [CWL16]. minimum-variance [CSL19]. mining [KPLC15]. Minkowski [GTBC18, MSdC^{+}18]. minor [YQD17]. minority [YC18]. Minute [Lah18]. Minute-ahead [Lah18]. Miodek [GR15b]. Mirakjan [MDG18, MNK19]. Mirakyan [GG15b, Aca15, DPD18, KM18a]. misaligned [YW19b]. miscible [GZW15, Kim15a, LR17, YCH18, ZZ19a]. Mises [Gat18a]. miss [MS15c]. mission [LXPD15]. mitosis [AVCCGAVDL15]. Mitsch [DKM15]. Mittag [CZLS18, KKMS16, LJW^{+}15, LW18d, LK19a, ÖK19, SC18a, YLHS17, Zha19c]. Mittag-Leffler [CZLS18, KKMS16, LJW^{+}15, LW18d, LK19a, ÖK19, SC18a, YLHS17, Zha19c]. Miwa [LL18a]. Mixed [CLCW17, GM15a, HXW^{+}19b, KKTY17, SP15, SS15c, AÖ16, AH15, AP18a, AZAM15, BBG19, Bas19, BY19, BB16, BM18b, CCJ15, cCpSkZ15, CGW16, CLPY18, CH18b, DN15b, DN16, DA19, DKP16, DFJ16, FPUK15, FYSC18, Gao17a, HA16, HRS17, HL16, HZLY17, JK15, Jeo16, JL15, KR15, KwC15, KWT15, Li16a, LTC16, LFL^{+}15, LZ17f, MZ15, Mis16, MSS15, MS15d, MKN15b, NRS19, OS15, Özb15, PBP19, RNM15, SR15b, SLC18, SZ19b, Sha15d, SWL16, SML18, SCY19, Tan17a, TPCL15, Wan18, WS19b, WSC19, WL19b, WCH15, WW19, XLR17, YBZ18, YCH18, YFW^{+}17, YQ15, YQD17, YLQ17, YCLS19, ZDPZ15, ZZZ16, ZH17, ZZ17a, ZZ18a, ZZ18b, ZF18, ZZ19a, ZjW19, ZCL18, dFGAN18]. mixed-mode [DKP16]. mixed-type [AZAM15, HA16]. mixing [JMK17, SSS15]. mixture [FS16, KSV15, NKL15, Zei16]. mixtures [Par18a, Zha18d]. mKP [ZW15b]. MLDG [AD19]. MLPG [HSS15, MB17]. MLRS [LWWL19]. MM* [LHG18, WY17]. MMS [PW19a]. mobile [hCWW^{+}15, LWY^{+}19]. mobility [VEF19]. Möbius [DTM19, MB19, Nam19]. Mocanu [Dzi15]. modal [KSL19]. Mode [Wan16a, AADLF15, CTT^{+}19, CHL15, CPCQ17, CL19c, DeC19, DKP16, GZC19, HYL19a, HZK^{+}19, JGX15, JGKL16, KML19, LQZ^{+}18, LZ18c, LGCS19, LPS19, LGK15, LMW18, LZWZ19, Pai15a, RXZW17, WXXY19, XZLL17, ZLLZ16, ZMC19a, ZCCZ19, ZYLY17, ZJYD17, ZMC^{+}19b, ZYWW16]. mode-dependent [CHL15, ZLLZ16, ZCCZ19]. Mode-independent [Wan16a]. Model [FXXC19, IB15, Ino16, KAU15, KSP16, AYAA^{+}15, AKL19, ALDHA^{+}17, AM15a, ALP18, ADY16, AO15b, ARV18, AAE19, AKT19, Alk19, ALZ16, Ant18, ARR16, AELH15, Ask18, AVCCGAVDL15, BM17, BM19a, BW15b, BN18, BG18, BBF^{+}18, BP17a, BKAAM15, BKKN16, BS15b, BF16a, BKL^{+}18, BTWY15, BL15, BR15, BL18, BMRS18, CYLW15, CMW17, CKW17, CJG^{+}18, CCC^{+}17, CMG18, CGC19, CLR16b, CLHC15, CSWJ15, CC18a, CKG^{+}16, CWFL15, CDZ^{+}18, CR18, CX19, CLW18, CL19c, CD16b, CDFP17, CM16, CM17b, CT15, CT17d, DLCV17, DTJ18, DY17, DHQ19, DKP16, DAQI15, DMA^{+}19, DZ16, DME^{+}15, DCK^{+}18, DF18, DKZ16, DL16, DSSC19, DQW^{+}19, DKDD15, Ery15, FHV15, FLMC15, ELP^{+}18, FQMR19, FG15, FHPV18, FFL16, GI16, GLR^{+}18, GO15b, GJR^{+}15, GM17a, GXH18, GKDHR19, GNR^{+}19, GR19b, GLA^{+}17, GRS^{+}15, GLM16, GL18b, GHM15]. model [HGG16, HD17, HW19, HHR17, HW17c, HLJ18, HZCY17, HZ17c, HLC19, HW15, IR16, IWL17, JYVS15, KMP19, KZL19, KÇ18a, KNK15, KB17, KIZ18, KSL19, KTD15, KKS^{+}16, KPP18, KSZ15, KRA^{+}18b, KS19c, KM15d, LGO17, LZX15, LCS15, Li15b, LCQ15, LWPL16, LR16b, LL17b, LJD17, LZT^{+}18, LWL^{+}18, LW18f, LX19b, LLY19, LJYH19, LYZ19, LPZL16, LXT17, LRL18, LDD15, Liu15d, Liu15c, LB15b, LZ15c, LTUZ15, LB16a, LC16b, LSW16, LZY16, LB16b, Liu17b, LXJ^{+}17, LYM18, LHG18, LJHA18, LJS18, LCW19b, LX19d, LGL19c, DPL18a, LPL17, fLLPY16, MNYL18, MZC19, MBM16, MMS17a, MMS18, MTMB17, MMG17, zMnZyZ15, MLG18, MC15c, MCPO15, MT15, NSN^{+}18, Neu16, NEK18, OA18, PBP19, PWFY18, PSJJ19, PZLZ15, PN15a, PTA16, Pel17, PL15c, PL18, Pet15b, PC17, PBB18, PKB18, PRMC19, PCA16, QBNF19, QLF15, QZJY15, QTXY16, QD18, QG19, RP19, RM18a, RZMB17, RLM19]. model [RMCG19, RRTR15, RHJ18, STJS16, SLdLD16, San19, SS18b, SZ19a, SHLW16, SZ15, SB16, SK19, SP15, SRP17, SL15b, SZJ19, SLL19, SRAC16, SKHA18, SHEBS19, SZFS17, SKCS15, SCZ^{+}17, Su16, SWA^{+}19, SLL^{+}16, SZT17, SSZL17, SCM19, SH18b, Świ15, ŚKK^{+}18, TES15, TPS15, Tan15d, TS15, Ter15, Tha18a, TLZ17, TXL19, TD19, TSA15, TRRP15, UPMV18, VVD15, VVA^{+}15, Wan17a, WLZW15, Wan15a, WPL15, WLS15, WLZ16, WTXZ16, Wan16b, WC16, WGLZ16, WSLC17, WY17, WWDW17, WLG17, WJHA17, WM18b, WWK19, WCH19, WWZ19, WSXR19, WXLL19, WHM^{+}19, WXM^{+}18, WYD^{+}18, WWC^{+}18, WZZT19, WH15, WYZ^{+}17, WMZ19, XD15, XJZ^{+}15, XWDW17, XSLW17, XZ15, XWZ15, XGZ17, YSLS17, YW19a, YWL19, YL19a, YLL15, YZJ18, YRA15, YX15, YJW16, YJ19, YYZ19, Yua15, YW18c, ZZZ15a, ZZBS16, ZI18, ZXW16, ZZ14, ZZ15c, ZJ15a, ZLJ15, ZS15a, Zha15a, Zha15e]. model [ZMMZ15, ZS16c, ZHW16, ZH17, ZZ17a, ZZHT17, ZLC^{+}18, ZLLA19, ZCFW19, ZZ19b, ZYZ15, Zha15g, ZZQ18, ZWZT18, ZBL15, ZXD15, ZZL16b, ZYZ16b, ZHWW15, ZXC19, ZJO^{+}18, dFKMM15, dSR15, dDSI19, dNdOX^{+}15, lZA^{+}19, uRAD19]. model-free [LGL19c]. modeled [Sch16, WMSP19]. Modeling [BAE15, FS16, HK18b, KSKW^{+}19, Kha17, KCKZ15, LLZ^{+}16, LZY16, LWY^{+}19, LLL16b, MMK18, DFG^{+}19, RL15, SZO16, XZG19, APMM18, Elm15, How15a, HXS^{+}16, JOÖ16, JG17, KN19, KCK18, KCL19, Laa18, LZL17, LZZL15, LCW19b, MJSA19, NDZ15, Ort15a, PPS25, PJ18, PKHD18, Par18a, PZJ17, QFS17, RRR19, RRR20, Rub19, SBV15, SSM^{+}15, SWSS18, SG18, UHLM15, VPS15, XRB17, YUK19, ZT15, Zha16b, ZZM18, ZWL16, Zho17, ZK18]. Modelling [CVJ19, GSV18a, RCM15, RSO^{+}18, WXP16, CH18c, JV18, Mou15, RSLL^{+}19, SFC^{+}17, Tor16, VB15, YG15a]. models [ADMV19a, ABA18, ASV17, BTH18, BCMT17, BG17, dCCGV15, CJ18, Che15e, CL16d, Cim18, CvRvdM18, CGMT16, CST^{+}15, DQH18, DJL17, DMS^{+}18, EA15a, EO15, ES17b, GCMR18, GZ18, GZD19, HH16, HH18a, HGH19a, HL15a, HYL^{+}19b, JA15, JV18, KWH17, KM17a, Kri19b, LLL16a, LYH19, LWW15a, Luo19, MCY19, MW18, MKMC15, MK15b, ND15, NKL15, OLZ19, QW19, RWY19, SSM^{+}15, uIHA18, SC17, TTD17, TT18, Vad19, Vid18, WRW16, WL16d, XLR17, YYP18, YZZ18, ZZZ16, ZG18, ZHS18, ZLC^{+}15]. moderately [JA17]. modes [CHL17, LXPD15, WYL15, WZZL16]. Modification [CCC^{+}17, FNS15, AHM16, MH15b, MH16a, MH17b, NL18a, RR18a]. modifications [FT16, KKB15, WK16b]. Modified [EH14, EH15, Gua19, HV19a, JCW^{+}15, Lac15, LSjL17, RKQ19, SuRI15, SHR18, SR15f, SSB15b, VDV15, YSF16, eMAGG19, ACMT15b, ACMT15a, AKA19, BPR18, CFR18, CST18, Che16c, CC18b, CXL18, Cho19b, DCW15, DS19, DPD18, DCK^{+}18, Dim16, DWTT17, FZL19, GO15a, GO15b, GM15a, GKN15a, GKN16, GVA16, GG19b, HXM18, HWXC17, Kud16, LG17, LBRS18, LLz18, LWWL19, LT18a, MNYL18, Mat15, MH15a, Mis16, NT19, NR15, OTH16, PMC15, PeO18, QFS17, RG15, TSJ16, TXM15, TLHC17, TTK15, VRC15, XCHW16, YZ15a, YP18, ZBKS16, Zam19, ZYLY17, ZCC15, dDSI19, NRY16]. modify [SDK15]. modular [BLP17]. modulated [MNJM19, ZZBS16]. modulational [HGH19b]. Modulus [XL15, DLBQ19, Van15, WZLP18, ZVL19, dDSI19]. Modulus-based [XL15, DLBQ19, WZLP18, ZVL19]. moisture [KK15c, WZM15, YX15]. molecular [KMA16]. molten [KSKW^{+}19]. Moment [GG15b, TB17a, FWZ19, Vil16, WZC^{+}18, WLL16, XHW15, XDH19]. moment-generating [Vil16]. moments [ABQ15, CLW18, DG18a, DG18b, Gav15, HLH17, TZ18d]. momentum [CXK19, LM18]. momentum-preserving [LM18]. Monge [MG15]. monitoring [HCS19]. mono [HTZY17]. monoclinic [GB19]. Monod [DDD19, Kha17, SSZL17]. Monod-Haldane [DDD19]. monodomain [BTWY15]. monogenic [AÇ18]. monolith [SB16]. monolithic [BY19, SBV15]. monomial [Dim16, FvdMS15]. monomial-exponential [FvdMS15]. Monotone [AOH17, JLX15a, LL15c, Arà16, AKA19, Bog16, BCH15, CÖ15, CB17, DCW15, FPUK15, GHL19, LG15, PR15a, SGJS16, SPKC16, TZF15, WAS19, Yue19, ZSY17, ZW15d]. monotonic [CE15, CP15b, Che15b]. Monotonicity [Cas15, Che16a, AGN15, QDD15, WCD17, ZM17]. Monte [BG17, Ace19a, Har15, LGO17, LWZ16, UHLM15]. Moore [Kyr17, SS17, She18]. Morgan [DD15]. Morlet [PSP17]. MORPH [CL18]. Morphing [SKCS15]. morphisms [O’R18b]. Morrey [GR17a]. Morris [LTUZ15]. Morrison [Xu17b, SKP16]. Morse [NL18b, ZCX17]. mortality [KPP18, Tan15d]. mortar [CLR15]. mortgages [dCCGV15]. MOSFET [BS19a]. mosquito [MTMB17, ZGH16]. mosquitoes [SWA^{+}19]. Most [CWD^{+}19, MZP15, RP18, SG19a, Xu19d]. most-obtuse-angle [MZP15]. Mostar [Tep19]. motif [MB15b]. Motion [TSPF18, AwEE15, AAS^{+}17, CLR16b, DWS19, DQW^{+}19, FKV16, HMHA15, LKZ18, RWY19, TB17a, Wan15a, WDLS18, WY19, XL18b, XL19, ZY15a, Zha15a, ZHWW15]. motions [GDL16, SK15]. motivated [TES15]. Mould [FVMO15]. Mould-taper [FVMO15]. mouth [LWPL16]. movement [CFS18, PHRB18]. Moving [FLL^{+}19, GAT18b, UKK17, fP15, AwEE15, ABN^{+}17, JCW^{+}15, Laa18, LY15a, LCW15, LLL18d, WR15, WZC^{+}18, WZF^{+}19, ZZ18b, AM16]. MPI [SHHC18]. MRF [AD15]. MRI [AD15]. much [FK18b]. Multi [BNY15, BK18c, BKZA18, HZ16, LWZ16, NM15, RF18, WD16, YwFW^{+}19, AM15a, ARL15, Ali15, AM18c, BCO15, BLP17, BJM16, CL19a, CGPT15, CLZW19, CPS18, DK15b, Ery15, EE18, FW18a, FFGRSV19, FS16, GMOGCCC15, GKDHR19, GJNC18, GSD16, GZD19, HWY15a, HW17b, HL17, HAH19, HCY^{+}18, HLFZ19, ID19, KD18, KM15d, Lac15, LLY15, LL15b, LZjy18, LD18a, LY16b, LZWZ19, LS19c, LGL^{+}19b, LHZM19, LL18c, LSL19b, LGHW19, Luo19, MM15b, Ma17b, MGCAPP16, MKG15, Mat15, MFF15, NY18, OW18, PN15a, PABA15, RN18, RWY19, ROGS16, RCRL18, SM17, SSS17, SSCX18, SMG19, SS19a, SZL^{+}18, SY18a, Shi15b, uIHA18, SSPCPG^{+}16, SWZL19, TFLZ19, USCA15, USCAAA15, WHSS15, WM15, WWK19, WS19a, WXXY19, WV18, XH16, XZB^{+}18, jXlZX18, XLW16, Xu17a, YL17, YJMH18, ZLL16, ZLM^{+}18, ZG18, ZLLP18, ZLLA19, ZTW19, ZJ15b, ZYLY17]. multi [ZJYD17, ZZC19a, ZXD15, ZSL^{+}18, ZZC^{+}19b, dFKMM15]. multi-additive-quadratic [BCO15]. multi-agent [CLZW19, FW18a, GSD16, HWY15a, HW17b, HCY^{+}18, LL15b, LZjy18, LS19c, LSL19b, LGHW19, MM15b, Ma17b, MGCAPP16, SM17, SSCX18, SMG19, SY18a, SWZL19, TFLZ19, WS19a, WXXY19, XH16, YJMH18, ZLM^{+}18, ZJ15b, ZYLY17, ZJYD17, ZSL^{+}18]. multi-channel [LL18c]. multi-choice [PABA15]. multi-commodity [GKDHR19]. multi-component [FS16]. multi-CPU [FFGRSV19]. multi-criteria [RCRL18]. multi-depot [ZLLP18]. multi-dimensional [ID19, LHZM19, MFF15, RN18, SS19a]. multi-dispersal [GZD19]. Multi-distance [YwFW^{+}19]. multi-domains [WHSS15]. multi-exponentiation [BLP17]. multi-fidelity [XZB^{+}18]. multi-game [LGL^{+}19b]. multi-group [GZD19, KM15d, Luo19, RWY19, WWK19]. Multi-Layer [BKZA18, ZXD15]. multi-layered [ZLL16]. Multi-level [LWZ16, Lac15, SSS17, WV18]. multi-machine [PN15a]. multi-mode [LZWZ19]. multi-objective [CGPT15, CPS18, GKDHR19, MFF15, ROGS16]. multi-order [DK15b, ZZC19a]. multi-pantograph [AM15a, EE18]. multi-patch [BJM16]. multi-period [GKDHR19]. multi-point [HL17, KD18, jXlZX18, XLW16]. Multi-quadric [WD16]. multi-relationships [NY18]. multi-resolution [uIHA18]. Multi-scale [RF18, HLFZ19]. multi-scroll [GMOGCCC15, SSPCPG^{+}16]. multi-server [Shi15b]. multi-soliton [OW18]. multi-stability [ZZC^{+}19b]. Multi-stage [HZ16, NM15, ZG18]. multi-start [MKG15, ZTW19]. multi-state [Ery15]. multi-step [AM18c, GJNC18, LD18a, LY16b, USCA15, USCAAA15]. multi-synchronization [MGCAPP16]. multi-temperature [dFKMM15]. multi-term [ARL15, Ali15, HAH19, LLY15, SZL^{+}18, Xu17a, ZLLA19]. Multi-time-step [BNY15, BK18c]. multi-timestep [CL19a]. multi-valued [WM15, YL17]. multi-vehicle [GKDHR19]. multiagent [DHY^{+}18]. Multichromatic [HMSS19]. multicommodity [MKG15]. multicore [Xu19e]. Multicriteria [HJP16]. multicut [BB15b]. multicut-HDMR [BB15b]. multidelay [DZ15]. multidelay-integro-differential [DZ15]. Multidimensional [ACTV15, KSS19, ACTV17, CCTV15, EMM17, PHM19]. multiexponential [LLC^{+}15]. Multifractal [FMRRT15]. multigame [LJG^{+}19]. Multigrid [FGPR18, Bho15, CJK17, DFSC15, HYB17, Liu18b, LL15d]. multilayer [CM18]. multilayered [GB19]. Multilevel [AL18d, KB18]. multiobjective [JCX19]. multipath [TPL16]. multiperson [Pla16, Pla17]. Multiphase [TM18, EEG^{+}16, KRF17, Liu18b]. Multiphysics [GM17a, ZKV^{+}15]. Multiple [CY15a, DY17, GRS^{+}15, LPZL16, SC18b, Sun15a, ZZZ15b, ZCQ15, AAC16, Amm15, ABN^{+}17, BE19, BP17a, BCMT15a, CPCQ17, CN15c, DKKR16, GT15a, fGlHyL^{+}15, GKN15a, GKN16, GSVK17, Haz16, HWY15b, Kim15c, Kue15, LKM17, LSZ15, LMMZ18, LWWL19, Luo19, MPR^{+}17, MS17a, MN16, PRMC19, RKQ19, RDM19, SB15, SHHC18, TXG15, Uzu17, VLCT18, WLS15, WDZ17, WLG17, WJZF19, WZY^{+}19, WGL15, XZ18, XLL^{+}19, XXZY19, YZ19, YZH^{+}19, YWH15, ZLL15c, ZWBC18, ZG15a, ZPT^{+}15, ZH15b, dNdOX^{+}15]. Multiple-correction [CY15a]. multiple-root [GKN16]. multiple-scale [LMMZ18]. multiple-zero [GKN15a, LKM17]. multiplex [WGSX19, ZWXW15, ZWL^{+}19]. multiplication [Khl18]. multiplicative [Dor16, HZL18, LLZD18b, LLX18b, MD15a, MLT17, SHLW16, TM19, TNMR18, WWL18]. Multiplicity [CHRD16, Our16, CP15a, CD15b, GW16, Li17f, YL19b]. multiplied [RR15]. multiplier [WLZ19b]. multipliers [NS15a, SK17b]. Multipoint [BK18d, LWW15b, CCTV16, RB15, XLR17, YU15]. multiprocessor [YSZ19]. multiquadric [GW15b, GM17b]. multiresolution [AY16b]. Multiscale [lDqCWzH18, LC15a, RG18b, YWYS18, HCY15, JMW15, JV18, Liu16a, LCW19b, VVA^{+}15, Yil16, YSF16, ZKV^{+}15]. multiscale-homotopy [Liu16a]. multischeme [EKO19]. Multisensor [KŚK^{+}18]. multiset [GPRCTRY19]. multispecies [HJ15]. multisplitting [HM15c, WY15a, WS16]. Multistability [CZSH17, NLC19, TLX19, WRW16]. multistationarity [FSW16]. multistep [AHP19, LZ18f, Xy16, vLtTBI18]. multitude [AY16a]. multivalent [Dzi15]. multivariate [AL18c, AK15a, BB15b, Oli17, Rah19b, SST15, SS15a, Sko17, TG17, YS16, ZZ14, ZZ15c]. Multivector [HS17]. music [BMDS17, BDBKV19]. Mustafa [PK18a]. mutualism [HJ15, LB16b, WW15, WW16c]. mutualisms [WWW15]. mutually [dSVR17]. Mycielskian [LZLG19].

Naghdi [LJD17]. Nagumo [BTWY15, Fio15, FLL17, GO15b, HMSS19, IWL17, LH17c, ZSY18]. Nagumo-type [LH17c, ZSY18]. Nagy [GDF^{+}17]. nanobioelectronics [YRA15]. Nanofluid [SV17a, ATK^{+}16, GJR^{+}15, MS18a, MRR17, PBP19, SP15, SRP17, SMG^{+}18]. nanofluids [AwEE15, PMV16]. Nanomechanical [Kia15]. nanoparticles [KHN19]. nanophotonics [LHLL19]. nanoplates [RM19b]. nanostar [SIA16]. nanostructures [NZZ16]. nanotube [ZLL15b]. nanotube-reinforced [ZLL15b]. nanotubes [AB15b, Kia15]. narrow [BG18]. Nash [CP15c, MN15]. Natural [GGM15, ARR16, GJR^{+}15, Hua15, MGK^{+}16, Mah19, MS18b, MRR17]. nature [KHN19]. Navier [ABS15, ABGL17, BFR18, BS19b, DKKR16, DC16, Ers16b, FZLZ19, FFG^{+}16, HMT15a, KMB16, MPP17, Qiu18, RC17a, Rub19, SH16, YLL18, Yil16, ZL17a, Zha17c, ZW18, ZZ15d, ZS19b]. near [KRF17, LA15, VVA^{+}15]. nearest [Zhu15]. Necessary [CZ17b, LZY^{+}17, FM17a, FG19, GXLL19]. nectar [WWW15]. negative [BD17, PL15a, SS15b]. Nehari [Kar19]. Neighbor [RCW^{+}18, XLG18, YSY^{+}17, FJ19, LWSM16, LHG18, WY17, WX18]. Neighbor-considered [RCW^{+}18]. neighborhood [Abd15, ES17b, KTD15, MKG15, PTO19, RRTR15]. neighborhoods [HL15c]. neighbors [LWSX18a, NT19]. Nekrasov [LZZT18, OP19, SCN15]. nematic [DL16, Qia16a, Qia16b, Sin17]. Nemes [Che15a, Che16b, LSM15]. Nernst [HP16]. nested [CLR15]. nets [Val18]. Network [CDK17a, LPJ18, TKDF15, ADMV19a, AM19, BK18a, BK18b, CCH15, CDS15, CWD^{+}19, CXW^{+}19, DCS^{+}19, DZY^{+}18, EO15, FJ19, GJX^{+}19, HWBL19, HCX^{+}17b, HZ17c, IR18, JY19, KWH17, KPLC15, LJW^{+}15, LWL^{+}18, LJG^{+}19, LGL^{+}19b, NT19, NY18, PAE^{+}17, PY19a, PSJJ19, PRA^{+}19, PZJ17, RPJ^{+}18, SBG^{+}17, Świ15, Tan17b, TXG15, Wan17a, WMS17, WCH19, XZLZ17, XZS16, YCD17, YMeFG19, ZZZ15a, ZXW16, ZDPZ15, ZCW^{+}19b, ZCML19]. Network-based [LPJ18]. network-induced [PY19a]. networked [CXZZ18, CL19c, DHY^{+}18, DZY^{+}18, LLC18, LGHW19, MA16, MWL^{+}18, PY19a, PY19b, SM17, SZ15, SMZ^{+}17, TPL16, XCW19, XZHX17]. Networks [HWL^{+}19a, ATV16, ATV18, ATV19, AAE19, AY16a, APPE17, AJ15, Ban18, BPC15, BAE15, BMT18, BES19, CGG^{+}18, CCYT19, CD15a, CGC19, CZLS18, CZS15, CWL16, CZSH17, CLPY18, CWLA18, DKL15, EOP17, ES17b, Fan15b, FXC17, FSW16, FYZ16, FYSC18, GCAA18, GW15a, GhDwZS17, GHW19, GWWL17, GZLL17, GWXZ18, HM15a, HS15a, HXW^{+}19a, HXW^{+}19b, HHL^{+}19, HYJ19, JNGB18, JHWZ15, JSW^{+}18, JPZ^{+}19, JZ15b, KBR15, KSM^{+}19, KŚK^{+}18, KTD15, KDTP18, KSP16, LPJ18, LLMW18, LJ15a, LSCK15, LC16a, LCAA17, LWSX18b, LLL^{+}18b, LSZ^{+}18, LWSX18a, LZL^{+}19b, LJYH19, LLZ^{+}19, LXD19, LZX19, LTQF19, LGC19a, LGC19b, LSZ^{+}19a, LZM^{+}18, LDS^{+}18, DDF18, LZ15b, LZG16, LXTF18, LML18b, LSLL19, LMW^{+}19, LWY^{+}19, LL18c, LWZ18, LCLZ19, MRC^{+}18, Mao17, MPS15, NR16, NKS18, NLC19, PAM19, PVV^{+}15, PRC^{+}19, QBC19, QZJY15, QCH19, RZSS15, RSC^{+}17, RVGS17, RN19, RMB15, SAJ18, SAM^{+}16, SM15b]. networks [SJ16, SSZ^{+}18, Sha16a, SLZ17, SCC17, SWZ^{+}19, SRCA17, SSPCPG^{+}16, SS15c, SML18, SZRC15, SLWX15, TTZ^{+}19, TLX19, TTD18, TLL^{+}18, Tsy19, TDL^{+}17, TZD^{+}19, WY17, WWDW17, WYCZ17, WHX18, WLX18, WY18, WSH^{+}18, WRX^{+}19, WGSX19, WCHW19, WWSM19, WL19b, WZS19, WHWW19, WCH15, Wu15, XZ18, XLJC18, XZZ^{+}18, XKG15, XTXF18, XXZ^{+}18, XFTX19, XLL^{+}19, XXZY19, YWL15, YZH^{+}19, YYS17, YFLF17, YFW^{+}17, YMeFG19, ZZL^{+}18a, ZPZ^{+}19, ZLZ^{+}18, ZLJ15, ZWHC15, ZLW15, ZHJ^{+}17, ZYXF17, ZHWW17, ZZZY17, ZSZ17, ZYC18, ZZ18a, ZWBC18, ZCCZ19, ZW19a, ZWXW15, ZL19c, ZLK^{+}19, ZDF17, ZwX16, ZWL^{+}19, ZZ16c]. Neumann [BMM19, Che19, KNT17, LWS18, RGA19, TTD17, XH19b, Zha16a]. Neural [TKDF15, XZS16, YCD17, AADLF15, BPC15, BAE15, CZLS18, CZS15, CZSH17, CLPY18, CWLA18, EOP17, EO15, FYSC18, GHW19, GWWL17, GJX^{+}19, GMPP18, GZLL17, HS15a, HWBL19, HXW^{+}19a, HCX^{+}17b, JNGB18, JHWZ15, JSW^{+}18, KBR15, KSM^{+}19, KŚK^{+}18, KTD15, KSP16, LPJ18, LJ15a, LSCK15, LC16a, LCAA17, LSZ^{+}18, LWL^{+}18, LTQF19, LGC19a, LGC19b, LSZ^{+}19a, LZ15b, LXTF18, LMZ^{+}19, LCLZ19, MRC^{+}18, MPW15, MPS15, NR16, NLC19, PAM19, PAE^{+}17, PRC^{+}19, QBC19, RZSS15, RSC^{+}17, RN19, RMB15, SAJ18, SAM^{+}16, SM15b, SBG^{+}17, SSZ^{+}18, SLZ17, SCC17, SWZ^{+}19, SML18, Świ15, TTZ^{+}19, TLX19, TXG15, Tsy19, TDL^{+}17, TZD^{+}19, WYCZ17, WMS17, WHX18, WSH^{+}18, WCHW19, WCH15, Wil18, XZLZ17, XZ18, XZZ^{+}18, XLL^{+}19, XXZY19, YWL15, YZH^{+}19, YFW^{+}17, ZXW16, ZPZ^{+}19, ZDPZ15, ZHJ^{+}17, ZHWW17, ZZZY17, ZSZ17, ZYC18, ZZ18a, ZWBC18, ZW19b, ZCCZ19, ZwX16, TD18]. neurodynamic [SZ19a]. neurological [Tor16]. neuron [BR15, KDTP18, KS19c, LWZ^{+}19b, LTUZ15, ŚKK^{+}18, UPMV18, Uzu17, WMZ19, ZWBC18]. neuronal [RPJ^{+}18]. neurons [BK18b, CM17b, MMZ^{+}17, MNJM19, PRA^{+}19, PAJ^{+}19, TKDF15]. neutral [AZL16, BEM15, BT15a, BB19b, Can15, CLL^{+}15, CZM18, CCZ15, DZ15, Du15, FP18a, Fan15a, GD15, JYS16, LC15b, LTQF19, LSL19a, LGK15, LWZZ18, MZM15, Mig15, Mig16, Mil16, MDZ17, Nas19, Obr19, QZW19, RZSS15, SKKL17, SJWK18, SXS15, SH15c, TZ18a, Wan15d, WY16, XKWG15, XHW15, Yu16, ZXW16, ZLW15, ZS15c, ZYJ15, Zho15b, ZXX^{+}19]. neutral-type [LSL19a, LGK15, XKWG15, ZXW16]. neutrons [SNS^{+}16]. Neville [APS16]. Newell [MM18a, Xu15c]. Newton [ASKK19, AL18c, ACMT15b, AM15g, AK15c, AG16, AM17, AM18c, BF15, Bot15, CGPR15, CGJV16, GKN15a, GKN15b, GKN16, HVR17, HP15c, HGM15a, HM15b, KNBA15, LW15d, LZH16, MR15b, MYH15, NAM18, PC16, SSK15, SSB16, Sil16, TZF15, Tch15, WN16, YNMW16, ZN15, Zha15i]. Newton-based [HGM15a]. Newton-like [ACMT15b, AM15g, AG16, AM18c, GKN16, WN16]. Newton-Traub [SSK15]. Newton-type [KNBA15, Zha15i]. Newtonian [PPS25, RRR20, BG18, CCL^{+}15, GCM^{+}18, HG18, ILPE17, Jai18, JV18, KPG16, LX16a, Pon18, RRR19, SSN15]. next [CGPT15, YG15a]. niche [HCH16]. Nicholson [Tan15d]. Nicolson [ZJJ18, BTWY15, LCL17b, QFS17, SLC18]. Nicolson/Adams [ZJJ18]. NID [QBNF19]. nilpotent [SZ16, YY17]. Nine [YY17]. Niño [MP19a]. Nizhnik [ZH15a]. no [CD15d, MNYL18]. no-go [CD15d]. no-slip [BTH18]. nodal [HLJ18, tHBSfY19, KG16, PS15b, WLX18]. node [CR18, Erb16, JLL16, KC19, LZjy18, LL19b, Ton18]. node-adaptive [KC19]. node-based [LZjy18, LL19b, Ton18]. node-centered [CR18]. nodes [JY19, WY18, Zha17b]. Noise [LAVME18, AK19b, AAH19, BPZB18, CWD^{+}19, DHW^{+}19, ES15, FLL17, GSZ15, GST15b, HMD19, HC18, KMP19, LW15b, LY17, LC15d, LO19, LT18a, MD15a, Maj18, MLT17, SHLW16, SRH15, TM19, WL19c, WLL16, XPG15, ZHY16, ZWY^{+}15a, ZJO^{+}18]. noises [DZ17a, LLZD18b, MX17, SMG19, WDLL15, YZ18b]. noisy [LD19b, LJL16]. Non [AY16b, AJP18, DHX^{+}19, Ers16b, EdLC18, Gar16a, GLQ^{+}18, HHR17, KAI19, LSCK15, LC18, LGC19a, LW19b, LTZ^{+}17, MC17, MRR17, TQ18, XYZK15, ZPM16, ZXX^{+}19, AwEE15, ALL18a, AOH17, AHO18, ANGC17, ACRZ18, BG18, BMM19, BAK19, BNY15, BB15d, BKL^{+}18, BCFR18, CY16b, CTT^{+}19, CÖ15, CYSY16, CLZW19, CGAL19, CD15d, Cho19a, CLR15, CXW^{+}19, DS15, DS16a, DBV18, DL16, DZ17a, tFyZ15, eFJY17, FM17a, FNS15, GRV15, Gao17a, GSP15, GG18, GWPA18, GCM^{+}18, GG19a, HGG16, HJMB18, HH16, HHP18, HJ15, HVR16, HVR17, HSKHA15, HS15a, HSS15, HC18, Jai18, JS15, JV18, KG19, Kor18, KRA^{+}18b, Kur19, LZS15, LB15a, LG15, LR16b, LW18f, LR19c, LC15c, Liu15d, LJK15, LLL18c, LB19, LR15, MRC^{+}18, MKRS16, MH17a, MK16, MH19b, NC16, NS15b, NKP16, PPS25, PMV16, PBP19, PCW15, PW19b, Pon18]. non [QTTL19, RRR19, RRR20, RSC^{+}17, RNM15, RSKB15, RAV15, RDG18, RRV15, SMI18, SR15e, SSN15, SMR15, SuSR15, SuSR16, SCM19, TTZ^{+}19, TST16, TZF15, TLLD19, Ter15, VZ18, VD16, VG18, WDLL15, WYLG19, WL19a, WL16d, Wu15, WLL16, WYZ^{+}17, WLL^{+}19, XLJC18, XPG15, XGG18, YWO18, YZ18b, YLL18, YSL19, ZSS15, ZSK^{+}19, ZWHC15, ZSS16, ZSY17, ZW19b, ZjW19, Zha15g, ZZ16b, ZZW17, ZZL18b, ZZ17b, ZTW16, dFGAN18, dSH17, Khl18]. non- [BKL^{+}18]. Non-Adjacent [Khl18]. non-analytical [Kor18]. non-autonomous [DS16a, GRV15, HGG16, HJ15, HS15a, LC15c, Liu15d, LR15, WL19a, YLL18, ZSS15, Zha15g, ZZ16b, ZTW16]. non-availability [HSKHA15]. non-classical [VG18]. non-conforming [ZSY17]. Non-consistent [AY16b]. non-constant [CD15d]. non-convex [Cho19a, PCW15]. non-delay [BB15d]. non-delayed [Wu15]. non-destructive [CLR15]. non-differentiable [HVR16, HVR17]. non-diffusively [ZWHC15]. non-equilibrium [HH16, HHP18, SMR15, SCM19]. non-Fick [HSS15]. non-Fickian [LR16b]. Non-fragile [DHX^{+}19, LSCK15, LGC19a, LTZ^{+}17, ZPM16, ZXX^{+}19, KRA^{+}18b, MRC^{+}18, RSC^{+}17, TTZ^{+}19, ZSS16]. non-Gaussian [DZ17a, HC18, WDLL15, WLL16, YZ18b]. non-globally [KG19, YSL19]. non-gravitational [LJK15]. non-Hamiltonian [VZ18]. non-Hermitian [CY16b, tFyZ15]. non-homogeneous [BMM19, BAK19, PBP19, WL16d]. Non-hydrostatic [EdLC18]. non-ideal [GG18, RDG18]. non-identical [CLZW19, XLJC18]. non-infinitely [CTT^{+}19]. non-instantaneous [AOH17, AHO18, YWO18]. non-integer [NC16]. non-interpolatory [ACRZ18]. non-isothermal [DL16]. non-isotropic [ZjW19]. Non-iterative [MC17, RAV15]. Non-jumping [GLQ^{+}18]. Non-linear [AJP18, AwEE15, ANGC17, CYSY16, CXW^{+}19, DBV18, GSP15, MK16, MH19b, PMV16, RNM15, SR15e, VD16, ZZL18b, dSH17]. non-linearity [Kur19]. non-Lipschitz [ALL18a, XPG15]. non-local [RRV15, SMI18, WYZ^{+}17]. non-matching [BNY15]. non-minimum [FNS15]. non-monotone [CÖ15, LG15, TZF15]. non-necessary [FM17a]. non-nested [CLR15]. non-Newtonian [PPS25, RRR20, BG18, GCM^{+}18, Jai18, JV18, Pon18, RRR19, SSN15]. non-oscillation [Gao17a]. non-oscillatory [HJMB18, WYLG19]. non-overlapping [Ter15]. Non-permutation [XYZK15]. non-polyhedral [LB19]. Non-polynomial [KAI19, LW19b, LW18f]. non-quadratic [CGAL19]. non-rational [LZS15]. non-renewable [WLL^{+}19]. non-repairable [NS15b]. non-resonance [JS15]. non-self-centrality [TLLD19, XGG18]. non-self-intersecting [ZZW17]. Non-similarity [MRR17]. Non-smooth [Gar16a, LLL18c, PW19b, QTTL19]. non-standard [RSKB15]. Non-stationary [Ers16b, SuSR15, SuSR16, TST16, ZSK^{+}19, ZZ17b]. non-strict [ZW19b]. non-symmetrical [BG18]. non-tensor [LB15a, MH19b]. non-tensor-product [NKP16]. Non-uniform [LC18, BCFR18, DS15, eFJY17, GG19a, LR19c]. non-uniformity [GWPA18]. Non-vanishing [TQ18, MH17a]. non-zero [MKRS16]. Non-zero-sum [HHR17]. nonautonomous [BR17, CXMP16, EO15, Kan16, TLGX15, WHX18, Zha15e]. noncanonical [DJ18, GJZ19]. noncentral [GST19]. noncompactness [HSAR19, KE16, KAH15, RAH19a]. Nonconforming [SZL^{+}18, ABGS15, AGZ15, AAGZ19, SWL16, SW17a, SLL19, SY19, SSJ15, SCY19, XS19b, ZCZM15]. nonconvex [GW15c, JLL15b, LPW15]. Nondegeneracy [CKG^{+}16]. nondestructive [WZM15]. Nonexistence [DFP15]. nonexpansive [JK15, Nak15, SPKC16, TTP16]. nonfragile [XCY18]. nonhomogeneous [DL15a, Kim19, LVV18, Oni18, SC18b, ZLL15c]. noninstantaneous [LWSO19, SZ19b]. nonintegrable [MZ17]. Noninvasive [JV18]. nonisothermal [SHA15a]. Nonlinear [ANP16, AM18d, AAHT18, CFR18, CD15d, De 18, GHZ19, KF19, LRF^{+}18, LZLW17, MWY17, NM19, NND^{+}18, PS18b, RS15d, SV17b, WL16a, WV18, YSWZ17, YRA15, ABC^{+}16, AED15, AS19a, Ach19, AR16a, AÖ16, AJ17, ASHSC17, AJC^{+}18, ALP18, ACE19b, AIG17, AP19, ABA18, ACRZ18, ACDT18, AL18d, Arà16, AKKB17, ACTV15, ACTV17, AM15h, AAA15b, AD19, Ata16, AKA19, Bag19, BCST19, BW15b, BLK19, BMM19, BBCI18, BAM16, BCMT17, BK18c, BB15c, BdS19, Bog16, BGST16, Bou18, BS15c, BK18d, BHRW16, CLY17, CKW17, CCY19a, Can15, CP15a, CP19, CGW19, CZP15, CXP16, CKG19, CJ15, CLJ^{+}15, CKW15, CD15b, CL15b, CXMP16, Che16c, CL16c, CXZZ18, CG19, CY15b, CDL19, CN15c, CN19b, CLP^{+}18, CLMT15, CGMT16, CVMGPPGC15, CZ15c, DCW15, Dai16, DA17, DWZ15, DLjZ18, DL18b, DFJ16, DOZ18, DO15]. nonlinear [DC16, ER15, Ely15, EGB15, ES17a, Esl16, EA15b, FW18a, FMDZ19, FNS15, GWW15, GS15a, GSZ15, GWSC16, GHL19, GLM19, GMOP18, GK15, GS17b, GFG15, GS15b, GR19a, Gra15, GVA16, GR17b, GHM15, HQ15, HAC18, HJMB18, HM19a, HPN15, HARS15, HA19, HSAR19, HMF15, HWY15a, HJKP18, HH15, HAY19, How16, HZ17b, HZK^{+}19, Hua15, HM15b, HGH19b, HHY19, HMT15b, IJ15, IB15, ID19, IS17, JSM16, JZ19a, JLX15a, JPTL15, JGX15, JZ19b, JAN16, KAU15, KE16, KOR19, KNBA15, KLKU18, KJ15b, KFHH15, KG18, KSV15, KMP15, LPKS16, LZL17, LW15a, LY15d, LLZ15, LSAR15, LC15b, LWS15, LG15, LG16, Li16b, LR16b, LH17c, LLC18, LZCW18, LZjy18, LZ18a, LZ18b, LJL18, LY18a, LR19b, LW19c, LSZF18, LRL18, Lin19, Liu15a, LS16, LZY16, Liu16a, LZH16, LLL18c, LWS18, Liu18b, LJHA18, LL19d, LX19d, LLX^{+}19, LGHW19, LM15, MZ15, MZ17, Maa15]. nonlinear [MDHD18, MAR15, MM16, MYHS18, MAG15, MS19c, Mis15, Mus18, Nag17, NE16, NBK16, NB19a, NRT16, NL18a, NCS16, NA19, ND15, NHWZ15, NWNA15, NWN15, OLZ19, PHM19, PMC16, PKB15, PR15a, PZ16a, PFH^{+}15, PTBN19, Pol19, PKB18, PK18b, QZ15b, QW19, Qiu18, RAH19a, RWW15, RR18b, RWW18, RXQ19, RMA15, RL15, RHJ18, RB15, RG19, RGA19, RM19a, RMK19, S.15, SS18a, SWUE19, SAJ18, SuR15, Sae17, SH15a, SR15a, SR15c, SJD^{+}18, SKKL17, SJWK18, SM19, SJ19b, SN15a, SN15b, SSK15, SB15, SSB16, Sha16b, SA16c, SM15d, SzSC19, SLW15, SLW16a, SW17b, SW17a, SY19, SPG17, SK17a, SPST17, SRCA17, SRS17, SWSW19, SJL16, Su16, SH15c, Sun15a, SWZL19, TVK^{+}16, TSJ16, TZ18a, TLX19, Tan15d, TCS15, TPCL15, TBG16, TY16, TTK15, TSPF18, USCA15, USCAAA15, VDV15, VDI17, WYL15, WZAF15, WCZ15]. nonlinear [WLJ15, WZZL16, WLZ16, WZ16a, WSZ16, WMZ16, WSLC17, Wan17c, WLH^{+}18, WXX18, WXWX18, WDLZ19, WCB19, WNN16, WN16, WAS19, WG15b, WZK15, XZLZ17, XD19, XL15, XY15a, XY17, XY18, XCDZ19, XXW^{+}19, XLW16, XZHX17, Xu19a, YW15, YLHS17, YWO18, YLCW18, YWZS18, YX19a, YG18, YLP19, YmZHC15, YDLK15, YLW19, YMMH15, YX15, hY18, Yüz16a, ZAP19, ZNV19, ZA16, ZZL17, ZMC19a, ZGGY19, ZAW15, ZMMZ15, Zha15b, ZZC16, ZWK16, ZZL16a, ZSS16, ZZY^{+}17, ZYY17, ZLM^{+}18, ZXS^{+}18, ZMZ18, ZYLY19, ZJ15b, ZFZK15, ZCQ15, ZLX19, ZWO^{+}19, ZYX19, ZVL19, Zho15d, ZW15d, ZH16, ZSL^{+}18, ZZ19c, dCSS17]. nonlinearities [AH15, LMW18, MC15a, MR15a, MPS15, NC16, Özb15, SMZ^{+}17, WDLZ19, WWHZ18, ZG15a, ZWO^{+}19]. nonlinearity [BB16, KG19, LLLZ16, MPP17, NRT16, OPP18, Pai15a, Ste19, WW16b, WTSQ15, Won15, XGQW19, XZLL17]. nonlinearly [FYZ16, How15a]. Nonlocal [FCM17, LVC15, TTD17, TLZ17, WIF15, WZ15a, AN15, AL18b, AB18, AS15, CS17, CR19a, CY18, CD16b, Cui15, DLjZ18, GZK16, HLIK19, HZ17b, LWT^{+}15, Li17b, LY15e, Mah18b, MVS19, PS18b, WZM15, WM18b, XYDL16, XZ15, YU15, ZZLY15, ZHW16, ZHS18, ZWZT18, ZMT15, ZY15d, ZLL18]. nonlocality [Dim16]. nonmonotone [BK18d, HMF15, LP15c]. nonmonotonic [TLX19]. Nonnegative [SY18c, DR19, HGM15a, Lin15, SCR^{+}15, YKZ15, YZ15c]. nonnegatively [FLMR18]. nonnegativity [FWLC15, FCW^{+}17, LH17c]. nonnesting [Yan18a]. nonobtuse [BCK15, Cih15]. Nonoscillation [Sug17, SI19]. nonperiodic [ZY15b]. nonpersistent [PD15]. nonsingular [CY16b, MR15b, WZC^{+}18, ZLL15a]. nonsmooth [BPZB18, GGM18, LL19d, LWW15c, MG19, Tua15, ZJJ18, ZX19]. Nonstandard [ZI18, GGHL15, GCMR18, GK17, ZEF19]. nonstationary [BPR18, KMP15]. nonsymmetric [FZL19, HWXC17, YZ19, ZZG17]. Nontrivial [ZCX17, MLS19]. nonuniform [CJK17, LR19b, NND^{+}18, SWZ^{+}19, TZH15]. nonuniqueness [DGLE17]. nonvanishing [BCSG18]. nonzero [DHQ19]. Nordhaus [MCS^{+}19, YCYL18]. Nörlund [Bor15a]. norm [CS15b, CKDC16, GK17, HM18b, Li16a, OP19, RzS15, TY16, WXY15, YSD^{+}17, ZL16a, ZZHT17, Zhu16]. Normal [JS15, Zha18e, BBGR15, MRS18, WYL15]. normality [HKM18]. Normalized [Han15a, SR15d, BZ17, HLL16, HL18, LH17b, LGS^{+}18, MB19, MB15b, SD19b, XZC16a, XZC16b]. normed [Haz16]. norms [CC18a, IÖ19, JZ15a]. Northern [NFS^{+}15]. nosocomial [WLG17]. nosocomial-pathogens-infections [WLG17]. Note [AC19, AA16, Ano19w, CWZ17, EL19b, EL19c, GR17a, GDL19, LLS15, ND15, Par15, Shi15a, YZZ16, ADA18, ABI19, AP18b, CWL19, CT18, DGLE17, DKM15, Dor16, FA15, HMGC18, Jan15, JZ15a, JYS16, KKD15, LLWX19, Li17a, LLL17b, Lu15, MMDG16, Net15, Qia16a, RDLG16, Sha16b, Tan18, Van15, VV18, WL18a, Yua19, ZS15d, Zha15d, EL19a]. notes [SC18a]. notice [AZ19, EH15, FMB^{+}19, Rho19]. Novel [CS16c, KSL19, KKL18, LQW^{+}15, PKHD18, SLZ^{+}16, XCCL18, ZCX19, AKS15, APMM18, BCST19, BBFJ17, CLHC15, CD15b, CDL19, DQW^{+}19, EAAM^{+}15, HST17, HLC19, IS16, Kim15b, KJ15b, KTD15, LLMS16, LCS15, LL19b, LRL18, LWWW15, LJHW19, MLG18, OLZ19, PS19, SZ19a, SSK15, SA16d, SLL^{+}16, TVK^{+}16, WQK15, WYD^{+}18, XZZ^{+}18, ZCW^{+}19b]. Novelty [FIG^{+}19]. Novikov [Tal15, ZH15a]. nowhere [ACP19]. nowhere-zero [ACP19]. nozzle [CT15, CT17d, HH16, SFC^{+}17, VG18]. NPDEs [Rui18]. NSDDEs [TY18]. nuclear [ACKS16]. nucleation [BG17]. null [AK15d]. Number [ATX19, XYZ18, AL19, ACP19, AM18c, BGS19, BF16b, CYY19, CRC19, CWD16, CWZ17, CZ17a, CMSZ18, Che18a, CZ17c, Du17b, DKDD15, FHLL19, GZ18, GDL19, GXCY19, KCSD^{+}18, LLST19, LG18, LWSM16, LZH19, MV15, Pel17, PBMP15, PZ19, QX17, Rad17, STA^{+}18, STD15, TLLD19, TZC19, WF15a, WYW16, WTD^{+}19, WC15a, WWL18, XLG18, YG15a, YSY^{+}17, Yor15, YZYgG18, ZL16b]. numbering [Ano19w]. numbers [BMRW19, DLQ17, Ele15, GA18, GLQ^{+}18, HLL18, IMTJ19, Jin17, KKD15, Lam15, LWG16, LLL19c, ML17, MNJW19, Qi15, QZD16, Rez15, SAES19, Sta17, WTZL18, WWH17, ZL18]. numerator [Han15c]. Numeric [AZAM15]. Numeric-analytic [AZAM15]. Numerical [AZL15, ASASM16, AG15a, AK19a, Ali15, AST18, AK19b, Awa15, Bab16, BM15a, Bea15, Bea19, Bog16, BKS18, BPR18, CH18a, CKW15, CZLM18, CC18b, Che19, CVA18, CDFP17, DK15b, DO17, DKKR16, DGRR16, DMSS18, ERFE16, FLL17, FFH^{+}18, FFL16, FFH^{+}15, Gao17a, GCH^{+}19, GJR^{+}15, GGM18, GFG15, GJP18, GSV18b, HH16, HHP18, HGH19a, IH15, IS17, JLY18, KLZ19, KRF17, KDH19, KCK18, KFHH15, LWZ19a, LZL19a, LHHS16, LBB18, LCW19b, LZ19g, MM04, MRE18, MPW15, MYHS18, MH15b, MH16a, MS19c, MR16a, MD17, MT15, Nag17, NE16, NL18a, NDZ15, NS17a, OPP18, OA15, PS18a, PG15a, PS18b, Pöt19, PKS15, RM18a, RME18, Sab18, Sak19, SDS18, SSS17, SR15e, SKS17, SR16b, SSA18, SPST17, SMG^{+}18, SWSS18, SD19c, SH15d, Tan15c, Tho17, WLJ15, WDLL15, Wan15d]. Numerical [WDLS18, WXW18, WXWX18, WLZ15, WXM^{+}18, XW19, Xu17a, Xu19b, YTH19, YWRP16, YJW16, Yüz15, Yüz16b, ZS16b, ZW18, ZZG18, ZYZ15, ZSXW18, ZHS18, ZXL15, ZH16, ZW15e, ABMBR15, ADSS16, ABGL17, ADG16, AP15, ASV17, AD17, AAMC19, BAV18, BMM19, Bas19, BTP17, BCM17, BMT18, BAR18, Bho15, BC15, BS19a, BKL^{+}18, BCI15, BS15c, Cal18, CLX17, Cha19, CJ18, CD15b, CS18, Cho19a, CLP^{+}18, DN15a, DA17, DA19, DCS15, DME^{+}15, DMS^{+}18, DZ15, Dol15, DNC^{+}17, lDqCWzH18, DWT17, DQW^{+}19, EFPT18, EEG^{+}16, Fan17, FY18, FKR15, FP18b, FM18, GS15a, GHN^{+}16, GCMR18, GN15, GR15a, GN19b, GLA^{+}17, GIS18, GAT18b, GLS19, GG15c, HW17a, HZ17a, HYW^{+}15, ID19, JOÖ16, cJzS15, JSLW17, Kem18, KXZ19, Kor18, KR16, Kue15, KAS16, Laa18, LP15a, LRF^{+}18, LW18f, LYDH18, LZ15b, Liu17a, LDH18]. numerical [LZ19f, LD19b, LXZ19b, LJ19, MWLL17, MZ17, MN17, MC17, MVRV19, MLG18, MWY15, Mic15a, MDH15, MK16, MA15c, NCEGA15, NKNA16, Obr19, Oli17, OL19, PBP19, PMC15, Par15, PYZS18, Pis18, RR18a, RG15, RAV15, RMK19, SWUE19, SR15a, SR15c, SN15b, SKBR18, SG18, SM18, TMPG17, TTD17, TTP16, Tha18a, UKE15, USCA15, USCAAA15, VG18, WHSS15, WX19a, WWZ19, WN16, YMS16, Yan18b, YYM19, YU15, YZMM15, YLL^{+}16, YCLS19, ZPRZ18, ZP16, ZBBW17, ZSL15, Zha16b, ZLC^{+}18, ZLD18, Zha18b, ZLLA19, ZZ19b, ZCU15, Zho15b, uRIS17, TT16]. numerically [AD19, MD18a, MAJR19, ZHX^{+}17]. numerics [KSA19]. Numerov [GZZL17]. NURBS [MPR^{+}17, RSAF16, ZZW17]. nutrient [JA15]. nutrient-phytoplankton-zooplankton [JA15]. NVIDIA [GWK15]. Nyman [MR19]. Nyström [ABI19, DL17b, FK17, FT16, FT17, OR18a, Oka18b, RF18, TZ18c, TSZ19, TZ19a]. Nyström-solver [FK17].

O [ATK^{+}16]. object [DRS18, KGS^{+}15]. objective [CGPT15, CPS18, GKDHR19, Lac15, Mat15, MFF15, ROGS16]. objectives [Haf17]. objects [CGK^{+}17, DCA15]. oblique [GCH^{+}19, WHM^{+}19]. observability [CGCK19]. observable [CTT^{+}19]. observation [LSZ^{+}19a]. observational [SSB15b]. observations [AAHT18, Luo19, SRK18]. observed [GOH18, WZ18b]. Observer [CL19c, JZ19b, KML19, WHTW17, CVMGPPGC15, DZC15, Du17a, DCJ19, LZ18c, LGCS19, LPS19, LZM^{+}18, LGK15, MAR15, MTB18, NT16, NPT19, QG16, XGQW19, YLCW18, YLP19, ZCW19a, ZPM16]. Observer-based [CL19c, JZ19b, KML19, WHTW17, MAR15, NPT19, ZPM16]. observers [CTT^{+}19]. obstacle [BM18b, LW15a, WEH18, XS19b]. obtained [CM17b, NC16, YG15a]. obtaining [VDV15]. obtuse [MZP15]. occlusions [LVAB^{+}16]. occupied [HW15]. occurrence [HZK^{+}19]. occurring [CCL^{+}16, HXW^{+}19a, LMW18, MC15a, TA16, ZML^{+}17]. octagonal [LW18e]. Odd [RC16, WZ18a, LLL19c, SH15c]. ODE [HPR18, SC17]. ODEs [ASM16, AADMM16, BHRW16, EHdLS16, RRM16, RS17a, USCA15, USCAAA15, WW16a]. oFEM [DRS18]. off [SH15b]. offers [SCB15]. offset [BKAAM15]. Ohm [OTH16]. oil [GLYT16].

Oldroyd [LS19b, ZZQ18]. olecranon [GLA^{+}17]. OLEDs [FFG^{+}16]. oligopoly [AE17b, BN18]. Olver [S.15]. OMP [YLX18]. On-Surface [AAC16]. One [Ask18, GR15a, ZZ19c, eMAGG19, AAKT15, AADF18, BB19a, BCSG18, BKKN16, BST18, Cat17, CL16b, CS18, CP15d, EKO19, FvdMS15, GfSmL16, HJKP18, HM15b, KIAG16, KL16, KNT17, Kra18a, LGO17, LH17a, LB16a, Mon19, NT16, NPT19, PVL19, QZ15b, RBB19, RSKB15, RP18, RS15d, SWA^{+}19, TMS15, TZ18a, TY18, Wan17c, YLNL15, YNMW16, uDuI18, ZJ15a, ZT16b, Zha15g, ZY18b, ZEW19, Zhu18]. one-dimensional [eMAGG19, BB19a, BKKN16, GfSmL16, KL16, KNT17, Kra18a, YNMW16, uDuI18]. One-leg [ZZ19c, QZ15b, TZ18a, Wan17c]. one-locus [SWA^{+}19]. one-phase [BST18]. one-prey [LB16a]. one-sided [NT16, NPT19, TY18]. one-step [EKO19, HM15b, RSKB15]. one-way [PVL19]. onion [SLWX15]. onion-like [SLWX15]. Online [MTY16, GZ15a, MNYL18]. only [AG15c, JYSR18, XWDW17]. onset [Kim15a, MS18a, SH18b]. onto [BBGR15]. Open [GKS16, BTP17, CS18, Far16, MK15a]. OpenMP [Xu19e]. operating [DKDD15, KSKW^{+}19, PN15a]. operation [UMF19]. Operational [SPS16, AKAA15, AE17a, DK15b, HS18, KS16a, KOR19, Nad15, NE16, PS19, SAC18, SR15c, SPST17]. operations [DSAB16, GFWH17, GDZ18, LAAZH19, SDAB17, SY15b]. operator [AIP15, AOS19, AG16, AS19b, GG15a, GWY15, GS11, Hus19, jJ15c, Kaj18, KKT19, Kra18a, Li16a, LR16a, LZ18b, LVV18, MDH15, Pan18, PSPÁV15, RAH19a, Rem18, SMM18, SGG15a, SGG15b, ZS15d, Zha17b]. operator-valued [GG15a]. operatorial [AK15b]. Operators [Aca15, MAK16, Acu15, Agr15, AIK15, AFK15, AT15b, AK15a, AY16b, bBjJ16, BM15b, BTIIE16, BA17, Bir15, CZ16, CLL16, Cat17, CCH15, CCGV19, CKSV19, CS15b, DPL^{+}18b, DFSC15, DDM16, DD16b, DPD18, DJ18, Fau17, GA15, GO15a, GLL15a, GM15b, GR15c, GG15b, GAS17, HO15, HVR16, HVR17, tHBSfY19, IAK15, jJ15b, JKK15, KA15a, KIAG16, KS18, KTA15, KAG15, KM18a, KT16, KMT16, Liu15g, LO15, Mal15, MG17, MS16b, MMS17b, MDG18, MCI16, MAK15a, MAK15b, MAK15c, MNK19, OTE15, PS15b, Ras15, RV16, Sha15d, SGJS16, SIA15, SV15b, TB17a, TA16, UDA15, WIF15, YY15, ZZ15b, ZX15, Zhu16]. opinion [WRW16, ZLC^{+}15]. opponent [FK18b]. opportunities [FLS^{+}15]. optical [Fio16, HGH19b]. Optimal [AHS18, ARR16, Arm19, BP17a, BAGN^{+}19, DQH18, DL17d, EFL18, EG17, GMOP18, GZ18, HMS15, HW17c, KMP19, LXPD15, LW18a, LB16b, Ma18, MHZ18, PP15a, Prz15, QD18, RLSGFMMF16, RLM19, RRS^{+}18, SA16a, SCR^{+}15, SS18b, SHA15b, SHN16, SBLZ17, THM19, WSXR19, WRZ19, Yua15, Zha17b, ZSZW16, ZCCL18, AHP19, ASS16, AA16, AM15f, BCMT15b, BCMT15a, BAM16, BMT19, CH18b, CN15b, CN16b, CLT^{+}15, CXW^{+}19, DS16a, DL15b, GN19a, GTBC18, GR19b, HA19, HHGC16, HL16, HZLY17, HGM15b, HW15, IKMW15, KKB15, KG15, KH17, LVAB^{+}16, LZL19a, LKM17, LZZ15, LWPL16, LZ19d, LPZL16, LTC16, LS15b, LXZL19, LZ17f, MLS19, NCS16, OA18, PRL^{+}15, RZR16, SA16c, SA16d, SY17b, SDP16, Tan15b, TMM18, TTXZ16, UMF19, WY15a, WNF19, Yil16, ZZ15a, ZZS^{+}19, ZZX19]. optimality [CKG^{+}16, Tua15, WTUZ15]. optimisation [SA18a, SFC^{+}17, ZSA^{+}15]. optimisation-based [ZSA^{+}15]. Optimising [RMM19]. Optimization [BLA15, EH15, Luk15, Mah17, MSdC^{+}18, MMB15, SZRC15, AADLF15, ARS15, BS15a, CY16a, CK17, Chi15, CS16c, DP15a, DH19b, DCA15, DPSS16, DKL15, DGL^{+}15, EH14, FA15, FSW16, FIG^{+}19, GI16, GYY15, GWC^{+}16, GdS18, Gar16b, GGM18, GLYT16, GIS18, HLJ18, JPTL15, JCX19, JG17, KKKD18, KSN15, KLZT15, KM18b, LYC15, LY18a, LABK15, LB19, LWWL19, LPW15, MJSA19, MND^{+}15, MMG15, DFG^{+}19, Naa15, NM15, NS15a, NRAEC19, ÖÖAD15, bOqGyK^{+}15, PAE^{+}17, PZ16a, PHSB18, Pet15b, QFTT15, SH15b, SM17, SJ19a, SBG^{+}17, SRH15, SSS15, SZT17, TIT15, TTV15, Tua15, VVD15, WYW16, WSdSY15, WCZ15, Wan16a, WLWA18, WHX15, WTUZ15, YZMM15, YJYH15, ZBKS16, ZN15, ZZMS19, ZWL16, ZBME16, dSRTZ18]. optimize [ZKV^{+}15]. optimized [HWBL19, NFS^{+}15, SGKR19]. optimizer [MHZC15]. Optimizing [KC18c, PP19, Bog18]. optimum [GWX17, KLW15]. Option [CL16d, CLL^{+}15, CW17, CDL15, LW15a, ZLS16, ZLC^{+}18]. options [AM15d, ASV17, CXL15, GM17b, GBS15, HS15c, JO15, LD17, LLZD18a, LGO17, MZC19, MBHS18, RPB15, ZBL15]. oral [BDBKV19]. orbit [Dan15a, JBS15, SJ19b]. orbits [BR15, NJM^{+}18, SNL19]. order [AKAA15, ABC^{+}16, AKS15, Ach19, AM18a, AZL15, ABLZ15, ABL15, AÖ16, AZL16, ASHSC17, AZ18, AH15, AMO18, ASASM16, AADLF15, AAKT15, Ali15, AK18a, AT15b, AM18b, APL15, ABA18, Alz15, ACDT18, AM15f, AKKB17, AS15, AE17a, AJP18, AAH19, ASM16, BCST19, BEM15, BLK19, BMO15, BD15, BBF^{+}18, BMM19, BCMT15b, BCMT15a, BAM16, BCMT17, BDGS15, BÁR15, BGST16, BS19b, BdlRT18, BMGMC17, BHRW16, CFG15, CLY17, CS16a, Can15, CP19, CZLS18, CDJ19, CCX15, CWHY15, CC15, Che16c, CGWW17, CZLM18, CC18b, CC18c, CDL15, CN15b, CN15c, CN15d, CN16a, CN16b, CN19b, CLT^{+}15, CJ16, CLT18, CEHVT15, CLMT15, CNQRR18, CS16c, CVMGPPGC15, CZ15c, DK15b, DG15a, DQH18, DD16a, DGZ16, DOR18, DWZ15, DJL17, DG18b, DL18b, DÇO15, DMA^{+}19, DK16, DS16b, DNC^{+}17, DQ18, DB17, DJ18]. order [EAAM^{+}15, EKO19, FT17, FMDZ19, FFT15, FBF18, FR16, FM18, Fuk15d, Fuk15b, FK18b, GCAA18, GO15a, GDM16, GFE19, GY15, Gao18, GZK16, GKN15a, GKN15b, GKM16, GKN16, GFG15, GR15a, GPHAPRW16, GR19b, GL15a, GJZ19, GfSmL16, HQ15, HJMB18, HEA15a, HEA15b, Han18, HW19, HMD19, HGH19a, HA19, HWY15a, HK18a, HASC19, HPR16, HS18, HAY19, HAH19, HGH19b, IAW18, IH15, cJzS15, JZ15a, JYS16, JSLW17, JISI18, KAU15, KSM^{+}19, KS16a, KNK15, KAI19, KNBA15, KD18, KBM16, KSL19, Koc15, KCKZ15, Kud16, KS19c, KS19b, KPLC15, LKM17, LW15b, LL15a, LSfW15, LJW^{+}15, LZS15, LL15c, LL15e, LY15b, LWK15, LZ16, LW16, LZ17b, LJD17, LLC18, LZjy18, LZ18a, LW18f, LD18a, LK19a, LGC19a, LGC19b, LHLL19, Liu15e, Liu16b, LZH16, LCL17b, LW18h, LMMZ18, LNWS18, LS19c, LD19b, LXZ19b, LSL19b, LMMR15, LJC16, LT18b, Lv16, Lv18, Maa15]. order [MDHD18, Mah17, MGPPGCCV15, ML15, MMZ19, MH17c, MS19c, MS16a, MGMM17, MAJR19, MK16, MMK17, MKS19, MSL15, MO15a, Mos16, Mus18, NE16, NBK16, NR15, NKS18, NC16, NCS16, NS17b, NEK18, Oni18, Ort15b, OA18, Özb15, PAE^{+}17, PHM19, PY19b, Pi15, PS15b, PVL19, PP19, Por18, PS19, PRC^{+}19, PK18b, PV19, QLLY18, RD15, RRM16, RSKB15, RSKB16, RS17a, RR18a, RLZ19, RN18, Rem18, RA15, RW17, RME18, RMK19, RZW18, SWUE19, SST18, SEKA16, SM16, SHR18, SN15a, SCB15, SSCX18, Sha15e, SB15, Sha16b, SA16c, SA16d, SGM^{+}16, SWL16, SSM^{+}15, SMM18, SGMH16, STD15, SM18, Sug17, SH15c, SzShG16, Tal15, Tan15a, TZF15, TSZ19, TZ19b, TBG16, Tua15, UHLM15, Ugu19, USCAAA15, WA15b, WK16a, WTJ17, WLLK17, WXWX18, WY18, WW18, WR19, WLQ^{+}19, WXLL19, WXXY19]. order [WLFX15, bWzZjH16, XY15a, XYZK15, XY17, XW19, XC15, XLW16, Xu19c, XLL^{+}19, XS19b, YW15, YLHS17, YWO18, YLCW18, YFH19, YGZ19, YZH^{+}19, YU15, YUK19, YmZHC15, YJMH18, YLH^{+}19, Yüz18, ZAP19, ZPRZ18, ZB17, ZT15, Zha15f, ZY15b, ZWK16, ZLH16, ZZCY16, Zha17a, ZSZ17, ZYC18, ZLD18, ZZ18c, ZWBC18, ZZS^{+}19, ZZX19, ZCU15, ZJ15b, ZY15c, ZYLY17, ZZL18b, ZZQ18, ZLX19, ZZC19a, ZS19b, ZZ16c, ZJO^{+}18]. order-linked [SCB15]. order-size [CLT^{+}15]. ordered [GLL15a, CLHC16]. Ordering [AG17, GA18, Wan16c, Kur18, LWG16, LGC19b, LD18b, SCB15, YW18a]. orders [Cat19, DY15, DKM15, HLT16, HCX^{+}17a, IKMW15, KKB15, MM17, MD15c, MCI16, RV16]. ordinal [BRM^{+}19]. ordinary [AK19b, CS16a, CZ17a, CNQRR18, DZW15, FY19, JISI18, LJZ18, LXZ19a, LMZW16, MGMM17, Nas19, RRV15, ZMQ16]. organization [LLT19]. organized [HYZ^{+}18]. organizing [KTD15, KDTP18]. orientation [Han16]. oriented [DRS18, GWX17, KST16b, LWZ15, MRS19, PCA16, QY15, Wan16c]. Orlicz [bBjJ16, CNNA16, jJ15b, jJ15c]. Orlicz-convexity [CNNA16]. Orness [DPL^{+}18b]. Orthoexponential [BÇS15]. orthogonal [AHG15, AM15d, BGP15, BP17b, BCFR18, CMR17, CCR18, GGMPC15, MMMMB17, MMPP18, MRQZ19, OA15, Rah19b, RS17b, SST18, Sza15, WZF^{+}19, ZZ18c]. orthonormal [MS19c]. orthotropic [ABN^{+}17, NB19b]. Osada [HMGC18, ZS14]. oscillating [CLS15, lDqCWzH18, HCY15]. Oscillation [AR15, BD17, BGST16, CLS15, DWZ15, DJ18, GS15a, Ma17a, SMZ15, WWZ19, AZL16, BMO15, Ban18, BJ19, CDJ19, Ely15, Gao17a, GJZ19, HPN15, HV19a, Wan15d, ZWK16]. Oscillation-free [Ma17a]. Oscillations [CÖ15, BF16a, DKP16, PFH^{+}15]. oscillator [BC15, BHD18, GST15b, Sun15a]. oscillators [GCAA18, NC16]. Oscillatory [ABL15, Can15, ABLZ15, CZ15a, CXK19, FMX15, FLS18, FKR15, FR16, GNW15, Gra15, HJMB18, HZ17a, JIS15, Kan19, Li17d, LD18a, Li18, LX19a, LX19c, LP16b, WYLG19, XMX15, uDuI18]. osculatory [LLO16]. Oseen [LR19c, TA16]. Osher [CGM16]. osmosis [FW18b, DFG^{+}19]. Ostrovsky [KRB15]. Ostrowski [CDK15, LT15]. other [ACD18, KKKD18]. outbreak [DTJ18, LLL16a]. outer [DYW15, GPRCTRY19, Pop17, TLZ^{+}19, XZS16]. outerplanar [YKLS19]. Outline [IKHS16]. Output [DH17, MM15b, XLJC18, YSWZ17, CXP16, CLXW17, CCQS16, CLCW17, FMDZ19, FNS15, FLYG18, GZC19, HYL19a, KSM^{+}19, LLZ15, LZM^{+}18, MAR15, Pai15b, PY19a, SYZP16, SWSW19, TTZ^{+}19, TPL16, TLL^{+}18, ZZL17, ZHWW17]. Output-Feedback [YSWZ17, CCQS16, LLZ15, TTZ^{+}19]. output-feedback-based [TPL16]. outputs [YCX^{+}19]. oval [BM19c]. over-constrained [dSH17]. over-relaxation [Gon18]. overfly [Tsy19]. Overlapping [ZGL15, CLR15, DHW^{+}19, LpLxG18, LJL16, Ter15]. overview [EEG^{+}16]. OWA [DPL^{+}18b]. owner [HW15]. owner-occupied [HW15].

p53 [PBB18]. p53-Mdm2 [PBB18]. Pachpatte [BAK19]. Pachpatte-type [BAK19]. Pacific [FMRRT15]. package [CBGGV17, DRS18, RCRL18]. packages [MHCF15]. packet [HZK^{+}19, HXW^{+}19b, LCY16, SS19b, SDP16, TPL16, WSH^{+}18, XPZ18, XCW19]. Padé [AL18c, MVRV19, NH17, SSB18]. page [LYC15]. PageRank [ATV16, GJNC18, HM15c, SHC^{+}17, TLZ^{+}19]. Pages [DN16]. Painlevé [RC17b]. pair [Kud19b, LABK15, Mis15, PZJ17]. pairings [UK18]. pairs [BP17b, FKR15, WJGM19, XLW15, Yor15]. Pairwise [KS15, NT19, KWGRP16, Pos17]. Paley [Ort15b]. pan [LCOS19]. pan-integrals [LCOS19]. panconnectedness [KC18c]. panel [MSdC^{+}18, TDA17]. Panleukopenia [CYLW15]. pantograph [AM15a, BÇS15, EE18, MHM15, RWY19, Reu15, YMMH15, ZGGY19, ZH16]. paper [DY17, Pan15, PK18a, TT16, ZS15d]. papers [ZS16b]. papillomavirus [SG18]. parabola [AG15d]. parabolic [AP15, AP18a, BNY15, BK18c, Bog16, CCJ15, DN15a, FGPR18, HZLY17, KK17, LZL19a, LZ17a, Lia15a, LFL^{+}15, LWSO19, MN17, MK16, MMK17, MKS19, RR15, Sap18, SMI18, SR17, SY17b, SW17b, SJ18, SN18a, SN18b, uIHA18, Slo15, SSJ15, TMS15, Vab15, VK15, VK17a, WLWD18, YX19a, ZZCY16, ZZ17a, ZKKZ18, ZMT15, ZY15d, Zho17, ZLW18, ZLL18]. Parallel [AGC19, BLP17, Cha19, CYCW15, HM15c, PVV^{+}15, WY15a, ZS19b, ADMV19a, BGP15, Dan15b, Fuk15a, JCW18, KK16b, Liu15b, MHCF15, MKN15b, SSS15, TD18, WS16, Xu19e, Zad16, ZPRZ18, ZXDP18, FFGRSV19]. parallelised [CSZ19]. parallelism [SD17]. parallelogram [ZZ16a]. paramagnet [Pu15]. Parameter [DK17, FvdMS15, PN15a, Akt15, BH17, BZ16, CÁGGLP16, CGCK19, CPST16, CCGP16, Dan18, DLJT16, Ely15, Fuk15c, HH18a, KLS15, LSfW15, LSAR15, LABK15, LS15b, LXJ^{+}17, NBK16, RR15, Red19, SR17, SRH15, SS15b, SuSB15, SWSW19, VVD15, YLNL15, ZG15b, dSR15, dSSB^{+}16]. Parameterized [Pop17, BBF^{+}18, CC16, DLF15, HR18, SL15a, Tan15c, ZLL15a, ZM15c, ZM18b]. parameters [ATK^{+}16, CZM18, Dor16, eFJY17, FDPP18, GZ18, HASC19, HM18c, KBR15, KA15b, LLCR15, LZKX19, LX19d, Luk15, MC15c, ÖAD15, ÖÖAD15, PN15a, RR18b, SN18c, UPMV18, VS17, WLH^{+}18, XJZ^{+}15, XKWG15, YWO18, YJW16, ZH15b]. Parametric [Lin19, AM16, ACDT18, BCST19, BBF^{+}18, CXP16, CGMT16, HP15b, JLL15b, LA15, Pop17, SJ19b, SDGU16, UKS^{+}16, WWZ15, XZW^{+}15, ZCCZ19, ZCX19, ZBME16]. Parametrized [Jar15]. parareal [Wu17, ARK15]. parasitism [WW16c]. Pareto [CGPT15, WYW16, Zil15]. Parisian [LLZD18a, LD17]. Parity [CYCW15, WC15a]. Parseval [DÖA15, MFC17, SZ18a, Sko17]. Part [Boy15, LB15a]. Partial [DGRR16, Mah18b, AB15a, ASA19, AGM16, BD18, BR17, BMGMC17, CW17, Chy16, Cov19, CT17b, DOR18, DSM15, DY15, DKM15, EAAM^{+}15, FY18, GAB17, GE16, GZ15b, GS17c, HPN15, HJKP18, JL15, JYS16, Kur18, LL16a, Liu15a, LY16b, LMMZ18, MJ19, MM17, MD15b, RD15, SuR15, SPS16, SPST17, SPS18, SN18a, SN18b, SQL16, VDV15, VDI17, WZZL16, WWZD16, YL17, YFW^{+}17, YJSW16, YYP18, YL19c, Zha16c, ZK18]. Partial-approximate [Mah18b]. Partially [KMB16, CY16a, CLZW19, DSB15, LSZ^{+}19a, LYZ19, LZHC18, WC16, WZ18b, ZGGY19, ZZMS19, ZDF17]. Partially-averaged [KMB16]. participants [Lu16]. participating [MHR18]. Particle [GI16, GSP15, TIT15, AM19, CS16c, DGL^{+}15, FLL^{+}19, GSDH15, GWPA18, KLZT15, LW19a, LWLM19, ÖÖAD15, SBG^{+}17, WL19b, YJYH15]. Particle-based [GI16]. particles [Ant18, Can16, CCL^{+}15, Nav17, SK15, ZHWW15]. particular [Cat17, HP17, Rét19]. particulate [SFC^{+}17]. partiteness [HLX17, HHT19]. Partition [EB16, CLHC16, LX18, WWW18, SCN15]. partitioned [CY16a, HMD19, MD15a, TLL16, XL18a]. Partitioning [LWM18, KB18, LSLZ18, PBMP15]. partitions [Bog18, CYW18, GSV15, GSV18a, Yan18a]. partly [FK18a, QG16, Zha19b]. parts [Pos17, TH19]. Pascal [BNS16]. pass [TBSR16]. passage [Xu15b]. passenger [YYX^{+}19]. passification [WCHW19]. Passive [FM16, CLCW17, SS15c, WZK15]. Passivity [HXW^{+}19a, JGX15, WCHW19, NR16, RSC^{+}17, SM15b]. Passivity-based [HXW^{+}19a]. past [EAY17, HRS17, Kan15, Mah19, MP18, Pan15, SSB18, Sha12]. Pasternak [LZL15a]. Patch [Li15b, BJM16, LL17a, LZ15c, Wan19b]. patch-based [LL17a]. patches [dSVR17]. patchy [GGM15, LZT^{+}18]. Path [GBS15, IB19, BHD18, CH19, GFE19, GLYT16, HP17, HRLGÁ18, JY19, LZ18e, DPL18a, MKG15, RZH^{+}19, RXQ19, SA19, SNAR15, XLS18]. path-conservative [DPL18a]. pathogen [SK19]. pathogens [WLG17]. Paths [Lai15, Che18b, DDdF16, Dim17, GL18a, HRLGÁ18, Mao17, MSdC^{+}18, QY17, San16, XYZ18]. Pathwise [KG19]. patient [JV18]. patient-specific [JV18]. Patrick [KKB15]. patronization [CRCVMS18]. Pattern [IWL17, LLWW19, PL18, SZFS17, HD17, HYZ^{+}18, MBM16, PZ16b, TLZ17, VK17b]. patterned [lJdW16, JH17]. patterns [BRM^{+}19, FFN18, HYZ^{+}18, RLSGFMMF16, TS15, ZZH15]. pay [FK18b]. payment [BS15b, PSSX17, WRZ19]. payoff [EEST16, Kur19, LJR19, NT19]. payoffs [DLD16]. PCHIP [ADS16]. PCM [PHRB18]. PDE [ASV17, HH19, MTB18, QFTT15, RR15, RC17b].

PDE-constrained [QFTT15]. PDEs [AND17, Dol15, EHdLS16, FBF18, GPHAPRW16, Kue15, RRM16, RAV15, TNMR18, USCAAA15, UKK17, VS19, ZHX^{+}17]. Peaceman [LZ17a]. peak [DLXZ19, Lu16]. peak-bounded [DLXZ19]. Peano [GI15, NB15]. Pearcey [LP16b]. PEC [QBNF19]. pedestrian [CFS18, GW19, HZCY17, HZ17c, KZL19, WWC^{+}15]. peloton [TRRP15, Tre15]. Pempinelli [FMC15]. penalized [LK19b, SBL16]. Penalty [EL19a, LW15a, ALP18, CW17, EL19b, EL19c, GGHO18, KCWK15, LY18a, MZ15, Sar17]. Penalty-Projection [EL19a]. penalty-type [GGHO18]. pencil [WJGM19, Yua15]. pencils [Wan18, dCSS17]. pendent [CZ17c, DDdF16, Dim17, JYSR18, SWG^{+}17, STD15]. penetrative [Har16]. Pennes [GSP15]. Penrose [Kyr17, SS17, She18]. pension [WRZ19]. pentagonal [LY18b]. pentavalent [YFH19]. people [BMDS17]. percentage [FHM16, SM15a]. perception [ZY16]. Perceptron [BKZA18]. perfect [CZ15b, FZZ18, SGDX18, YZZ16]. perforated [RL19]. Performance [EOP17, NCEGA15, Upa16, BS15a, CCQS16, FFTB16, Liu17a, XD19, YLW19]. peridynamic [TTD17]. Period [TSA15, GKDHR19, MBM16, SCB15, WQK15]. Periodic [BR15, Che15d, GGL15, HJ15, Liu15c, LD19c, SPGP18, XSLW17, ZS15c, ZMMZ15, ASASM16, AA16, ALN16, AM18c, ACC16, CL16c, CT18, DT16, DZ16, DFP15, DNC^{+}17, DO15, GO15b, GST15b, HV15, HPR16, HL15b, HHY19, IT15, JA15, JS16, JIS15, KLS15, LL15c, LZ15a, LG17, LXJ^{+}17, LL18c, LLdM16, Lv18, Ma17a, MPW15, MMZ19, NJM^{+}18, PPS25, PW19b, PP18a, RRR19, RRR20, Seg15, SXS15, SLL^{+}16, UMF19, WA15a, WTXZ16, WC16, XZCC15, YL19b, ZLW15, dSR15, dSH15]. periodically [FYZ16, GWXZ18, TN19, WHTW17, ZXC19]. Periodicity [Liu15e, AK16a, ACSS15]. Perishable [PL15a]. Peristaltic [FF18, AB15b, HMHA15]. Permanence [OL15, Zha15e]. permanent [ILPE17]. permanental [LQZ16, LW18e, LW17b, SR16a, WL18b, WS19c]. permeability [FK18a, Liu16a]. permeable [BAR18, EAY17, GG19a, Mah19]. permissible [BS15b]. permission [HHR17]. permit [LWZ19a]. PermonCube [MHCF15]. permutable [Pos17]. permutation [NL18b, WC15a, XYZK15, Yor15]. Perry [DCW15]. Persistence [LYH19, BB16, CDHK15, Kur19, VEF19]. Persistent [LSR18, HXW^{+}19b, XXW^{+}19]. person [DLY^{+}17a, LWZ19a, MN15]. personalized [RLGGAV17]. perspective [CDHK15, DZY^{+}18, Fau17, FNS19, IÖ19]. Perspectives [FPKF15]. persymmetric [AAKT15]. pertaining [SKHA18]. Perturbation [DQH18, EAY17, MZM17, eMAGG19, DHQ19, FYZ16, GG15a, JT16, LY15c, LLX18b, LJHA18, Luo19, Ma18, TIT15, XL18a, YZ19, YWH15, ZDPZ15, ZJO^{+}18]. perturbation-based [TIT15]. Perturbations [SZ16, CXZZ18, DGZ15, DJ15, JGX15, LD19c, MS19a, TLGX15, Wan17a, bWzZjH16, XH19a, ZYXF17]. perturbative [ACC16]. perturbed [AK18b, BZ16, CMR17, CJ16, CVA18, DN15a, Fan18, GQC15, GN19b, HV19a, KS16b, KK17, KH17, LSAR15, LB15a, LDS^{+}18, Liu18a, MN17, PCS17, SAA19, SR17, SN18a, SPGP18, SN18b, SRCA17, SM18, TG16, TNPL15, VD16, WLH^{+}18, ZY15a, dFKMM15]. perturbed-quadrifolium [LB15a]. perturbing [YY17]. pest [LPZL16, LTC16, LWW15a, SZT17, XTXW15]. pesticides [LTC16]. Petersen [CMSZ18, CZ19, YW18b]. Petri [Val18]. Petrov [KBK^{+}16, AJMN18, AM15c, DWRZ17, HSS15, MB17, WZWD15, YWZ18, ZY15c]. Petrovsky [LG16]. Petviashvili [VK19, EO18, RC17b]. Pfaffians [FZZ18]. PGM [LB19]. PH [KKD17, NC19]. PH-matrices [NC19]. Phase [MMZ^{+}17, BM19c, BB18b, BST18, DJL17, FNS15, FS16, GOA19, GAT18b, HHP18, HLC^{+}17, HLJ18, JLL15a, KZA19, Kri19b, KSZ15, LXL17, DPL18a, MS15a, MZS18, Par18a, Pel17, RSO^{+}18, SAES19, Tha18a, WXM^{+}18, Zei16, ZM19, ZHS18]. phase-field [ZHS18]. phenomena [RL15, SBV15]. phenomenological [ALDHA^{+}17]. phenomenon [ACRZ18, BCD15a, Kem18, KL15, LS15b, MMIA19, SZO16, VBG^{+}17]. Phillips [Sha15d]. phonation [SH18b]. phononic [RCM15]. phosphate [JMK17]. phosphate-buffered [JMK17]. photons [MHR18]. PHSS [Zha15b]. PHSS-like [Zha15b]. physical [FXC17, RZR16, SSM^{+}18]. physiological [HGG16, Li15a, Pon18]. physiology [SH15a]. phytoplankton [JA15, YYZ19]. PI [CVMGPPGC15, MBW19]. Picard [CG19, NH17, SuR15, Sae17, ZI18]. PID [SA16a, SY18a]. Piecewise [CG19, Han15b, LLW19, And18, Arà16, CFSGM17, CGAL19, DQW^{+}19, Fuk15b, Gao17a, GMOGCCC15, GLN15, Han15c, Han18, LLz18, Liu16b, NJM^{+}18, Pis18, RM16, TLX19, WA15a, WHTW17, WJZF19, XH19a, YSL19, ZS15b, ZYZW16, BCM17]. PIES [BZ17]. piezoelectric [KAMD18, LLL19a, MWLL17, ZLL16]. piezoelectric-type [LLL19a]. piezothermoelastic [KS19b]. piezoviscous [RP16]. PIM [LLG^{+}15]. Pinney [MR15a]. Pinning [ZZL^{+}18a, FYZ16, PRC^{+}19, RVGS17, YFW^{+}17]. pipe [MPPP19, PHRB18]. pipeline [AAS^{+}17]. Pisanski [KST19]. Pisarenko [GG19b]. Pitaevskii [MDH16, WMLG17, WJWX18]. pitches [ALMCDH^{+}18]. pivoting [APS16]. pivots [GTBC18, RBB19]. placement [SCZ^{+}17]. placenta [ZDZ15]. Planar [YSWZ17, CB19, DHS19, FG19, JJ16, LZS15, LL19c, OGC19, Van15, Wan15f, WFZ17, XLG18, Xu19d, YSY^{+}17, YKLS19, ZLSX18]. Planck [GDL16, GAB17, HP16, MW18, SR15e, XYG^{+}18, ZJY18]. plane [AOS19, BLNSF17, CGG^{+}18, Che19, Chi15, EM16, HWL^{+}19b, KM15a, Mar16, Neu16, SSA18, Vol15, dSR15]. planes [YJY16]. plankton [LYZL18]. planning [Cov19, GKDHR19, GLYT16, MND^{+}15, WLLL15]. plans [WRZ19]. plant [LPZL16, TH19, WW15]. plant-pest-predator [LPZL16]. plaques [MGN18]. plasma [Wan16b]. Plasmodium [KCL19]. plate [AAMC19, EAY17, JDL15, SLW16b, WL16a, Zho15a]. plates [DDY15, GG19a, JA17, LZL15a, LZL17, MC15c, MB17, Muk16, NMNTBXVD15, Rad15, RCM15, RL19, SSN15, ZLL15b]. platform [KGS^{+}15, YYX^{+}19]. player [ESS16, ESES16]. players [AMCMFJ15, DLJW15, ES15, KZT19, Pla15]. plot [CLHC15]. ply [MRF18]. PMC [LHG18, WY17]. PMHSS [CR15, LL15d]. PMHSS-multigrid [LL15d]. pneumococcus [MMK18]. Pneumonia [TMM18]. Pochhammer [Bag15]. pockets [CT17c]. POD [LJC16, LT18b]. Poincaré [MPB^{+}17]. point [ACS15, AKS15, ALMCDH^{+}18, ACE19b, Alf17, ACRZ18, AM15g, ARGF15, AM15h, AADMM16, BAR18, Boi15, BGP15, BLJ^{+}19, CFG15, CY16b, CL19b, CFSGM17, CC16, CM15a, CSZ19, DLF15, EB16, tFyZ15, tFyZZ15, FZLZ19, FB19, GK15, GKN15a, GKN16, GOA19, HM17, Han19, Har15, HL17, HM15d, HMX15, HM15b, HWXC17, HM18c, JA17, JSM16, Jeo16, JPTL15, KM17a, KM15c, KMBA18, KD18, Khl18, KKG18, LR19a, LW15c, LWK15, LSjL17, LZ17c, LZ19e, LL17c, LJ15b, MP18, Mia17, NH17, ODP17, PZ15, PN15b, RPB15, RDLG16, RS15b, RS16a, ROGS16, SSA19, SL15a, SJ18, SBL16, SuSR15, SN15c, SuSB15, SuSR16, SDK15, TTZX15, TST16, TXM15, WLLK17, jXlZX18, XLW16, YCH19, YLW15, YY17, YW19b, Yun15, ZV17, ZLL15a, ZZG17, ZM16, ZWY15b, LLG^{+}15]. points [BT15b, CL16c, CVA18, CNQRR18, DU18, Erb16, Ers16a, FHPV18, GRV15, GKN16, HRS15, IZ16, PN15b, SML^{+}15, SPGP18, SZ16, TW15b, TTP16, WHSS15]. pointwise [Gon18]. Poiseuille [KPG16]. PoisFFT [Fuk15a]. Poisson [BPZB18, CSA18, Fuk15a, HP16, KLZ19, KL16, LL19c, LDL15, LJL16, MLT17, PYZS18, SZPH16, TRP15, ZB17, ZLL15c, ZH17, ZZL15a]. Pol [Sun15a]. Polak [AKA19]. polar [BM18a, Jai18, KZZ15, LLX18b, yWyLyY^{+}15]. polarity [ADA18, AG17, HD16a, LdCC15, LL18b]. polarization [IR18, Kim18]. pole [LY16b]. poles [BM19d]. policies [BS15b, DL17d, PMSB15, RZR16, SCB15, ZPZ^{+}19, ZSW^{+}17]. policy [RRS^{+}18, SRAC16, SD19a, UMF19, ZWD15]. PoliMIce [GZG^{+}15]. pollination [WWW15]. pollination-mutualisms [WWW15]. pollinator [WW15]. pollutants [DMSS18]. pollute [BKAAM15]. polluted [LC15d, YYZ19, ZYZ15]. pollution [FCW^{+}17, LWZ19a, RSLL^{+}19, SK19, YCLS19, ZBR15]. poly [DKKM15b, GWPA18]. poly-Cauchy [DKKM15b]. poly-dispersity [GWPA18]. Polya [AIK15, IAK15, DDM16, DD16b, GAS17, MRS18]. Pólya-based [MRS18]. Polyak [AKA19]. Polychromatic [LZ18d]. polygamma [CE15]. polygon [Cho19a, DJLX17]. polygonal [Gia15b]. polyharmonic [RV16]. polyhedra [GWK15]. Polyhedral [GS18a, GS16, GWPA18, LB19, VZ18]. polymatrix [DJW19]. polymeric [KA15c, RG18b]. polymonogenic [BÁR15].

Polynomial [GSV17, LLX18a, SV15a, AHG15, AK16b, ASZD18, BÇS15, Bor15b, Boy17, CGJV16, CC19, CL18, CJ17b, DSM15, Fer15, Fuk15c, GK17, HPR18, JCW^{+}15, KAI19, KK15a, LZS15, LQZ16, LW18e, LW18f, LCW19a, LW19b, LWSM16, LDH18, LMMZ18, LLdM15, LLdM16, MRF18, Mig16, NRT16, PC16, Pis18, PS16, Pro16, PIP19, PV19, SGM^{+}16, SK17a, TNMR18, Tur18, WYL^{+}19b, WTUZ15, XH19a, XZW^{+}15, YZJ18, YQ18, ZS16a]. polynomially [YYM19]. polynomials [ABQ15, AHG15, AM15c, Akt15, AM15d, AHM16, AE17a, BN15, BS17, BMR17, BMB15, BP17b, BCFR18, Cas15, CMR17, CCR18, dCM16, CM15b, CW18, DDSL15, DSDL19, DD15, DKKM15a, DKKM15b, Far18a, GT15a, GGMPC15, Gla19, Han15a, HASA15, HK18a, HS18, IH15, JRSK16, KA15a, KM15c, KR15, KM18a, Kim15c, KKD15, KKMS16, KSS19, KAS16, LY15a, LB15a, LAAZH19, LW17b, Luo15, MMMMB17, MMPP18, MYHS18, MC15b, MS19c, Muk16, NR15, ÖK19, PR15b, QZ15a, QZL18a, Rah19b, RLSGFMMF16, RT18b, SR16a, Sak19, SR15d, SIA16, Sim15, Sin18, Sza15, TZTG17, TW15b, TG17, VLCT18, WW17, WXW18, YZ17a, Yüz16a, ZZ16a, ZHX^{+}17, ZZ17b, ZH15b, ZC19c]. Polynomiography [GK17]. polyomino [ARB16, HLL16]. polyoptimization [PN15a]. polyphenyl [CXL16a, YLWF15, YLWF16]. polytope [AGZ15]. polytopes [ABGS15, Cih15, RHP18]. polytopic [CWHY15, YCX^{+}19]. Pompeiu [ES16]. poor [GlH^{+}19]. Popularity [XDG^{+}19, CZM^{+}19]. Popularity-driven [XDG^{+}19]. population [AA16, CJ18, CST^{+}15, DP15a, ESES16, GLM16, GHM15, KSN15, KS19c, KM15d, LMF17, LGL^{+}19b, MDZ15, MS15a, MTMB17, PYZS18, QLF15, Sch16, SRK18, TRP15, TMPG17, Ter15, XD15, XJZ^{+}15, YLZT17]. populations [DL17d, HZS15, MMS18, MM16, Pla16, RS19, WHC18]. poroelastic [DK15c]. poroelasticity [GM17a, NRS19]. porous [CSZ15, EM15, EB16, FM17a, FM17b, FK18a, GZW15, GG15c, Har16, Jai18, LR17, Liu18b, MSI17, MKN15b, Pan15, PP18a, RNM15, RL15, RSC16, SST17, SSB18, SKS17, SS18c, Sha12, SRP17, SMR15, SMG^{+}18, SCM19, YWYS18, Zha16b, Zha18e]. port [FLYG18]. port-controlled [FLYG18]. portable [RLGGAV15]. portfolio [HD16b, HW15, JPY16, KM19, LZX15, MN16, NM15, PCW15, SBLZ17]. posed [CPST16, GS17b, JSG15, SGJS16]. posedness [GS17c, JL15, LL16c, Zha16c, ZFL18]. position [AVC^{+}19, CXK19, Dan15a, LJL15a, LS19a, LS19c, WWC^{+}15]. Positive [CD15c, HLZ15, LZCL18, MG15, WLZW15, YW18c, Agr15, ACKS16, ANEP19, BD17, CCQS16, DZ16, DM18, DL17c, DO15, tFyZ15, FM17a, FFT15, HL16, JXW15, KM19, LLY15, LWS15, LX16b, LQC^{+}18, Liu19, PWS17, PZ15, QZCJ19, Ras15, SC18b, SXS15, WA15b, WZLP18, XHW18, YQ15, YZ15c, ZLL15c, ZZC16]. positive-definite [tFyZ15]. Positivity [HJMB18, KE18, ZSY18, GOA19, LZ19c, MO16, TMPG17, ZZ19b, Zhu18]. Positivity-preserving [HJMB18, Zhu18]. possessing [BCSG18]. possession [HKW^{+}17]. post [LR19c]. post-processing [LR19c]. posteriori [Bac15, BBG19, BCNN19, BCG19, CG15, CH18b, GGHO18, HL16, MW18, XH19b, ZL17a, ZCZM15]. potential [eMBHA15, AM15b, AHM16, BW15a, Bhu15, BHD18, Can16, GR15b, KSP16, LL19a, QBNF19, SDWG19, XJZ19, YY15]. potentials [Cha19, Nav17, QFS17, Sun18, WE19]. Poussin [AÖ15a]. Powell [GK16b, MHZC15]. power [AMCMFJ15, BLK19, BD18, BB18b, BS15a, BL15, BL18, CW17, CLWM19, DSM15, DZ17a, GMPP18, GG19b, LPS19, LW19c, MGPL15, DFG^{+}19, OPP18, PN15a, PP18b, RCRL18, Ste19, SRGW15, TTD18, WC17, XTXF18, ZLS16]. power-based [XTXF18]. power-performance [BS15a]. power-type [OPP18]. powers [JXW15, LGS^{+}18, LL16b, Zha19a]. practical [WZAF15]. practice [BLA15, WWC^{+}18]. pre [Bhu15, CLT18, HLT16, MD15c, SG17]. pre-exposure [Bhu15, SG17]. pre-order [CLT18]. pre-orders [HLT16, MD15c]. precedence [GYY18]. precession [UGY16]. precipitation [WF15c]. Precise [Fuk15c, Fuk15d, Fuk15b, CL16c]. Precisely [CCYT19]. precision [DEdlT17, Her19, Ton18]. Preconditioned [WS16, ZM16, Bho15, CC16, DLBQ19, HXLC15, HM18c, LWK15, MLW18, SL15a, ZVL19, CFR18]. preconditioner [CM15a, CWL19, tFyZ15, FZL19, HMX15, LZ17c, LZ19e, LJ15b, LpLxG18, RBB19, TXM15, ZZG17]. preconditioners [CL19b, FZLZ19, HWXC17, LW15c, LL17c, WLZ19a]. Preconditioning [ZZ15a, AL18d, CZL18, Pel17]. predation [GG16, HD17, LPL17]. predator [AYAA^{+}15, AO15b, AADF18, CSWJ15, CY18, DSB15, FHV15, GNR^{+}19, GIS15, HL15a, HXS^{+}16, HCX^{+}17a, HYZ^{+}18, HLC19, JAU15, JGGK15, Kha17, KB17, LPZL16, LC15c, LB16a, LLZ^{+}16, LJHA18, LPL17, LD19c, fLLPY16, MHX17, MMS18, MMG17, zMnZyZ15, Mou15, NS17b, OL15, PZ16b, PL18, SP19, SSB15a, SL15b, SZT17, TS15, TLZ17, Vad19, Wan15a, WPL15, WXLL19, WLL^{+}19, XD15, XZCC15, XWDW17, YW18c, ZS15c, ZMMZ15, ZHW16, ZCFW19, ZZL16b, ZJO^{+}18]. predators [LLL16b, SSB15a, ZJ15a]. predecessors [ADMV19b]. Predetermining [AM18c]. Predicting [Lu16, MPLS19, SRK18, KSK^{+}15, NFS^{+}15, YS16]. prediction [CST^{+}15, HDG15, LHHS16, Rog19, Upa16, ZBKS16, lZA^{+}19]. Predictive [XCW19, DME^{+}15]. predictor [How15b, How16, JDG18, KDG19, MS17a]. predictor-corrector [How15b, How16, JDG18, KDG19]. Preface [Ano15, Ano16, DGRR16, ZS16b]. preference [BMDS17, LY18c]. preferences [CRCVMS18, GH17]. Preferential [ZSW^{+}17, ZL19c]. preinvex [NNA15b]. Preisach [BMC19, NEK18]. Preliminary [BTP17]. premise [Zha19b]. premium [PP15a]. preprocessed [GJNC18]. prescribed [Lin15]. presence [AwEE15, DS15, FS16, JYS16, JZ19b, KHN19, LW15b, MSI17, PSM17, YWL18]. Presenting [MN15]. preservation [LH17c, SN15c, SSM^{+}18]. preserve [Bir15, LZ16]. preserved [FWLC15, FCW^{+}17, YDWZ19]. preserving [BB18a, BMGMC17, CLY17, CCY19a, GHZ19, HJMB18, Han15c, Han18, HP16, HKW^{+}17, KMS18, KE18, LJL18, LM18, LZ19c, LH19a, LW19a, LWS18, MDHD18, MO16, QDD15, SHH15, TMPG17, TZ18c, WMSP19, YWS18, YUK19, Yos19, ZZ19b, ZLX19, ZCCL18, ZSY18, Zhu18]. Pressure [BTP17, RRT15, CXW18, EdLC18, HG18, LS19b, MKRS16, DFG^{+}19, QTXY16, SS18c, SCM19, WE19, XP17]. pressure-dependent [HG18]. pressure-driven [SS18c]. pressure-loaded [WE19]. preventive [DKN16]. Prey [MMG17, AYAA^{+}15, AO15b, AADF18, CSWJ15, CY18, DSB15, FHV15, GG16, GNR^{+}19, GIS15, HL15a, HXS^{+}16, HCX^{+}17a, HYZ^{+}18, HLC19, JAU15, JGGK15, Kha17, KB17, LC15c, LB16a, LLZ^{+}16, LJHA18, LLL16b, LPL17, LD19c, fLLPY16, MHX17, MMS17a, MMS18, zMnZyZ15, Mou15, NS17b, OL15, PZ16b, PL18, Sah15, SP19, SSB15a, SL15b, SZT17, TS15, TLZ17, Vad19, Wan15a, WPL15, WXLL19, WLL^{+}19, XD15, XZCC15, XWDW17, YW18c, ZJ15a, ZS15c, ZMMZ15, ZHW16, ZCFW19, ZZL16b, ZJO^{+}18]. prey-based [GNR^{+}19]. prey-one [AADF18]. prey-predation [GG16]. prey-predator [LLZ^{+}16, Mou15, OL15, WXLL19, ZMMZ15, ZCFW19]. price [dCCGV15, Cim18, Lah18, LZZ15]. prices [CLL^{+}15]. Pricing [AM15d, CLR16a, CDL15, LD17, LLZD18a, MZC19, MBHS18, PMSB15, RPB15, SPP17, ZW15c, ZL15c, dCCGV15, CXL15, CW17, CL16d, CLT^{+}15, HS15c, JO15, PP15a, RZR16, TZ18d, ZLC^{+}18, ZBL15]. Primal [CK17, RCTZ18, JPTL15, YP18]. Primal-dual [CK17, JPTL15, YP18]. primaries [SPGP18]. prime [JHL19, LLL18a]. primitive [JHL19]. Principal [YQD17, LKM17, MH19a]. principal-value [MH19a]. principle [ARL15, CCC19, HV19b, JV18]. principles [CH18c]. priori [BAE15, CH18b, FM17b, HZLY17, YCH18, Yil16]. Priority [DKDD15, JBS15]. priors [EGH^{+}19]. prismatic [BKZA18, KAMD18]. prisoner [CZM^{+}19, ESS16, ESES16, GSD^{+}17, HI15a, NT19, RCW^{+}18, SLS16, Tan15e, Tan17a, Tan17b, WJ19, WL19b, XDG^{+}19, YFLF17, ZAZ19, ZHY16, ZSL17, BM17, Sch16]. Privacy [HKW^{+}17]. Privacy-preserving [HKW^{+}17]. Probabilistic [PK15, ACS15, BCD^{+}19, CST^{+}15, CNQRR18, Haz16, NR16, Sch16, SHEBS19, TLL^{+}18, WR15, GGAVGGRC15]. probabilities [DKZ16, HZK^{+}19, KPP17, LZZ15, LSZY17, LSZ^{+}19a, LZHC18, SY17a, YKJY17]. probability [CLSS19, CNQRR18, Fau17, JDS18, LDSD19, LRH15, MSH15, Nad15, Rah19b, SJ16, WDLL15, WXZ18, WDLS18, YCD17, ZS16c, ZXL15]. probable [CWD^{+}19]. problem [eMA15, eMAGG19, AAGZ19, AOH17, AHO18, ABS15, ACE19b, AK19a, BE19, BMM19, BM19c, BY19, Bea19, BCNN19, BM18a, Boi15, BST18, Bor15b, BM18b, BFJ^{+}17, BZ16, CÁGGLP16, CG15, CFSGM17, CRCVMS18, CS19, CGPT15, CW17, Che19, Cho19a, CLP^{+}18, DLBQ19, DKN16, Dis15, DWS19, DQ18, FM17b, GWC^{+}16, GS18b, GZW15, GWY15, GG15c, HM15a, HZ16, HMZ17, HSKHA15, HXM18, HM15d, Hua15, HMX15, HM15c, HM18c, HM19b, HSSH18, KCZK15, KCWK15, KZA19, KLZT15, KH17, LLMS16, LZL19a, LW15a, LW15d, LL15c, LWK15, LYW17, LW18g, LY18b, LZ19d, Lia15b, LJK15, LY16b, LYW18, Liu18b, LML18b, LVC15, LZ17f, LDL15, MCY19, MGCAPP16, MPPP19, MKG15, Mat15, MFF15, ML15, MGN18, MMB15, MK15c, MTJ15, NY15, NJTP19, Our16, Pan18, PABA15, PZ15, PTBN19, PS15b, Pov15, PCS17, QDH^{+}18, QLZ15]. problem [QFTT15, RKQ19, RCTZ18, Rem18, RM19a, RMK19, SSK^{+}19, SNAR15, SH18a, SPW15, SWL16, SY18b, SY19, SJ18, SPGP18, Slo15, SSS15, SSJ15, STVCC17, Su19, TES15, TZF15, Tha18b, TCY15, TLZ^{+}19, TTP19, Tsi19, Tu17, Val15, WZM15, WQK15, WLLK17, WZWJ18, Wan18, WEH18, WS19b, WSC19, WD15, Won15, XL15, XL18a, XCHW16, XH19b, XBC19, XS19b, YÖ18, YWYS18, YCH18, YX15, Yua15, Yun15, ZY18a, ZL15a, ZJJ15, ZGL15, Zha16a, ZFL18, Zha18e, ZLLP18, ZJY18, ZF18, ZH19a, ZZ19a, ZTW19, ZM19, Zha15i, ZZL16b, ZWY15b, ZwX16, dFGAN18, ddGL15, BFIJ15, SH15b]. problems [ABD16, AT15a, AKS15, Ade19, ABHR15, AJMN18, AYD19, AMO18, ASASM16, AHP19, ADMV19b, AAKT15, Ali15, All18b, ASS16, ÁL15a, AV18, AAC16, And18, AA16, ADG16, AK18b, AP18a, ACTV15, ACTV17, ARS15, AF15, AJP18, Awa15, Bac15, BM16, BB19a, BNY15, BK18c, BRD15, BÁR15, BCH15, BVJ15, CFG15, CZ18a, CL19a, CY16b, CY16a, CL19b, CM18, CCJ15, CSV19, CHRD16, CC16, CL16a, CKG^{+}16, CKG19, CM15a, Che15d, CGW16, CH18b, CD15c, CJ16, CVA18, DLF15, DN15a, Deh16, DPSS16, tFyZ15, tFyZZ15, FZLZ19, FPUK15, Far16, FFT15, Flo18, FKR15, FGPR18, GW16, GS17a, GW15c, GZZL17, GdS18, Gar16b, GNW15, Ghe16, Gia15a, Gia15b, Gol15, HN14, HKL15, HLZ15, HYB17, HA19, HLIK19, HASC19, HPR16, HHGC16, HSS15, HL16, HZLY17, tHBSfY19, HM15b, HXLC15, HWXC17, IKMW15, JSM16]. problems [JK15, Jeo16, JIS15, JG17, KZR17, KKKD18, KM17a, KOR19, KMBA18, KAI19, KD18, KPT16, KEE17, KG18, KMZ19, KMT16, Lac15, LLO16, LX16a, LL15d, LQ15, LLY15, LWT^{+}15, LW15c, LL17b, LZ17b, LZ17a, LSjL17, LD18a, Li18, LML18a, LD19a, LW19b, LWW19, LZ19e, LBB18, LRL18, LL17c, LZ15d, LZ17e, Liu18a, LJ15b, LjL15b, MZ15, MG15, MN17, MK15a, MC17, MHCF15, MW18, MF17, Mia17, MLW18, MG19, Miy17, MS15c, MS15d, NCEGA15, NB19b, NN19, ODP17, OA15, bOqGyK^{+}15, PSY^{+}17, PS18b, RRRR16, RSKB15, RS17a, RR18a, RMM19, RRV15, RG19, RGA19, SH15a, Sah19b, SCR^{+}15, SM19, SA16b, SZ19a, SL15a, SHC^{+}17, SDK15, SBLZ17, Sow18, SDWG19, Tan15b, TG16, Tch15, Ugu19, VK15, VK17a, WSdSY15, WqTjT15, WZ15a, WSZ16, WZWD15, WZLP18, WTUZ15, Wu17, WLWD18, jXlZX18, XY15b, XLW16, YLW15, YBZ18, YX19a]. problems [YU15, YP18, YWH15, ZBR15, ZSS15, ZS16a, ZW15a, ZZ15b, ZLL15a, ZL16a, ZT16b, ZZG17, ZZM18, Zha18c, ZjA19, ZY19, ZCZM15, ZQG16, ZM16, ZZC19a, ZVL19, dCSS17, MAJR19, GR19a]. procedure [ACTV15, DLJT16, SGM^{+}16, ZZ19a]. Procesi [GYW15, WD16, ZZG18]. Process [EH15, Agr15, Arm19, BBD15, BG17, CR19a, EH14, FFG^{+}16, GCLG15, GG19b, JLL15a, KM19, LPKS16, LZZ15, NBAGS19, SZO16, SSM^{+}15, SWSS18, SHHC18, TZ18d, YZ19, ZWL^{+}19]. processed [XY15b]. processes [CDL15, DGN16, GR17b, HJP16, KCKZ15, LMF17, Liu19, MDZ15, SZPH16, XHW15, ZL19a, ZL19b]. processing [AY16b, ABB^{+}18, HGM15a, LR19c, Świ15, XRB17, ZXDP18]. processor [ZSK^{+}19]. produced [Seg18, TH19]. Producing [Eft15, DY17, JA15, YYZ19]. product [AÇ18, AY16a, AK15b, BCD18, BS19c, CCGV19, DPSS16, FTF16, GLW18, GIK19, HHZ18, jJ15b, jJ15c, Jur17, KAI19, LZ18e, LWM18, LH19c, Mao17, MH19b, NKP16, PS15a, SDAB17, TXM15, ZZC^{+}19b]. product-type [jJ15b, jJ15c]. production [Cov19, FFG^{+}16, GS15a, WLLL15, YL19a]. Products [DDSL15, DSDL19, ANP16, CP17, GSV15, LMYW17]. profile [ZXC19]. profiles [Neu16, YW19a]. profiling [PVV^{+}15]. Profit [DGLE17, Els17, KLW15, LTC16]. Programing [De 18, SJ19b]. programmable [KTD15]. programming [Alz15, CPS18, De 18, GW15c, HRS15, JLL15b, KLZT15, Lac15, LGL19c, MZ15, MCY19, NJTP19, PCW15, PABA15, ROGS16, SCR^{+}15, UMF19, WqTjT15, WTUZ15, Gol15]. programs [CJ17b, CRZ19, GRS^{+}15, KBK15, LHJ15]. projectiles [GCH^{+}19]. Projection [EL19a, SST18, AJMN18, AKA19, Chi15, CKSV19, CRZ19, DN15b, DN16, EL19b, EL19c, GHL19, GVA16, Nak15, VS17, WY15b, Xu19e, Zha18c, ZZQ18, ZLZW15, ZSZW16, ZYZ^{+}16a, dFGAN18]. projection/GMRES [Xu19e]. projections [BPR16, HZLY17]. projective [CZ15b, JZ15b, SGY^{+}15, bWzZjH16, XXZ^{+}18]. prolate [LDL15]. proliferation [WJHA17]. Promote [DHY^{+}18, CCW19, SLS16]. promotes [CCW17, CZM^{+}19, DZD^{+}16, Du19, GSD^{+}17, LSX19, LWWZ17, Moh19, SRJ^{+}15, WJ19, WWSM19, XDG^{+}19]. Promoting [YC18]. Promotion [CYGL18, CDZ^{+}18]. prone [LVAB^{+}16, MPLS19]. Prony [LW18a].

Proof [YG15a, BH15, LZY^{+}17, PP12, PP15b]. propagating [RPJ^{+}18]. propagation [BRV18, GJX^{+}19, GYW15, GB19, KSD^{+}15, Kun16, LYDH18, MHR18, MR16a, PBBF18, SSG^{+}15a, SV17b, ZWZT18]. Proper [CMM17, CLL17, GDL19, GXCY19, XYZ18, Zad16, ZZ18c]. Properties [DB17, LO19, Pac19, WLFX15, AZL15, ANU16, ABLZ15, AIK15, AR15, ALN16, BN15, BHRW16, Che16a, CW18, CK18, Dai16, DDdF16, Dim17, EA15a, FLMC15, HMHA15, JRSK16, Jar15, KVZ16, Kri19a, KZ15, LHHS16, LZ17g, MZ17, Mal15, MNR19, MS16b, MMS17b, MNJM19, NZZ16, Rét19, RS16b, SC18a, SSS17, SLF18, SBH16, SGG15a, Vis15, WLJ15, WZ19, YFL16, YDLK15, ZWD15, ZMZ18, Zha19c, ZM17, Zha15h, ZWL^{+}19]. property [GYW15, PP12, PP15b, WHWW19, YLX18, ZJO^{+}18]. prophylaxis [SG17]. proportional [CG19, CCZ15, GLL15a, HS15a, MHZ18, Reu15, Yu16]. Proposition [Cho19b]. propulsion [FTE19]. protecting [AKT19]. protection [PCA16]. protein [MPLS19, PBB18]. protocols [GFE19, LZjy18]. protocooperative [Tre15]. prototype [RLGGAV17]. prototypical [DKP16]. provable [HKW^{+}17]. provably [KG16]. provided [SSB15a]. provision [QZL18b]. provisions [HI15b]. proximal [GOA19, LPW15, PL15d, ROGS16, ZWY15b]. proximity [CCX15, IZ16]. PRP [ZW15d]. PRP-based [ZW15d]. Prüfer [HV19a]. pruning [SNAR15]. PSB [CY16a]. Pseudo [HPR16, APL15, ALP15, Bot15, CPST16, CZP15, CH18b, HGL19, Jan15, MFC17, MRR17, PVL19, SPG17, Xy16, XBC19, ZHS18, ZSZW16, ZLL18, DG18b]. pseudo-differential [CPST16]. pseudo-hyperbolic [CH18b]. pseudo-Jacobi [XBC19]. pseudo-parabolic [ZLL18]. pseudo-spectral [HGL19, MRR17, PVL19, SPG17, Xy16, ZHS18]. pseudo-splines [MFC17]. pseudo-transient [Bot15]. pseudoinverses [XZS16]. pseudoinversion [SKK19]. pseudoplastic [HMHA15]. Pseudospectral [WJL19, ZjW19, GBJ19, GSV18b, JZ19a, Tan15a, Tan15c]. psi [CE15, Ele15, YCZ15]. PSM [ZHS18]. PSO [BS15b, Gar16b, NFS^{+}15]. PSS [CY16b, LZ17c, ZZG17]. PSS2A [PN15a]. Psychological [WWC^{+}18]. PU [ZM16]. Public [HHL^{+}19, BM19a, hCWW^{+}15, CCW17, CYGL18, CCW19, Du19, GlH^{+}19, HI15b, LWGS17, LJR19, LWWZ17, QZL18b, WWW^{+}17, WHC18]. Publisher [Ano19w]. Publishing [CCW17]. Pullback [DK15a]. Pulsatile [PP18a]. Pulse [Sab18, BM15a, MH15a, NE16, QLF15, MM04, SR15a]. pulsed [BKT16]. pulses [LWW15a, Ter15]. punishing [YC18]. Pure [LTC16, LMZ^{+}19, MZM15, XCDZ19]. pure-feedback [LMZ^{+}19, XCDZ19]. purely [HP15c, MMV16]. pursuit [SSK^{+}19]. put [LW15a]. pyroelectric [YJY16].

QG [GR17a]. QP [LHJ15, LY18a]. QP-free [LHJ15, LY18a]. quad [WSC19, PAM19]. quadrangulation [SST15]. Quadratic [CÁGGLP16, De 18, ET15, LHWG16, AJ17, ASS16, Alz15, AADF18, BCO15, BBF^{+}18, CGAL19, CL16d, CRZ19, Dar16, DGZ15, DEdlT17, GW15c, Gar16a, GKS16, GYC19, Han15c, JLL15b, JSG15, KK16a, KAMD18, LX18, LY16b, LY18d, LFC^{+}19, MS16a, MPP17, NJTP19, PÖB15, RAH19a, SHH15, SST15, SZ18b, SGJS16, TPCL15, TTP19, UKE15, Val15, WqTjT15, WL18a, WLZ19a, WD19, WZ18b, XH19a, ZT16a, Zha15c, ZZZ15b, BFIJ15]. Quadrature [DLJT16, AHS18, APTS15, Bas19, CZ15a, GSV17, HZ17a, How15b, How16, LX19c, Ma15a, MRF18, MP19b, MD17, SHA15b, SHN16, TSJ16, Tho17, uRIS17]. quadrature-based [How15b, How16].

Quadratures [CM17a, GL15a, HMS15]. quadric [WD16]. quadrifolium [LB15a]. quadrilateral [HL18, LH17b, LY19]. Qualifications [De 18]. Qualitative [AR16a, MWWM15, BMM19]. quality [RZR16]. quantification [DK15c]. quantitative [CCGV19]. quantities [SZ17]. quantity [KPLC15, RZR16]. quantization [XXW^{+}19, XCY18, ZML^{+}17]. quantizations [CL19c]. Quantized [KLKU18, RXZW17, SWSW19, XCL18, CLXW17, FYSC18, KSM^{+}19, KML19, LCY16, MA16, YCX^{+}19]. quantum [AR15, CLSS19, HYJ19, NNA15a, NNA15b]. quantum-like [HYJ19]. Quartic [KD16, AHM16, CGJV16, Han15c, Nav17, ZM15b]. Quasi [CKSV19, DT16, FHK^{+}19, Har15, LGC19b, RV16, WY18, YLW19, AAGZ19, ALN16, ACC16, CCD18, DNC^{+}17, FWZ19, GW15b, HHZ15, Ied15, Isc16, JK15, JPTL15, JLL15a, KwC15, KLS15, LX16a, LHHS16, MKS19, PN15b, ROGS16, SST15, SZL^{+}18, TZ18b, VG18, WTXZ16, WD16, YLQ17, ZBR15, ZN15, ZYY17, ZZG18, ZH15b]. quasi- [JK15, YLQ17, VG18]. quasi-analytical [LHHS16]. Quasi-Bernstein [ZH15b]. quasi-bicyclic [FWZ19]. quasi-birth [JLL15a]. quasi-conformal [ZBR15]. quasi-contraction [PN15b]. quasi-convex [Isc16]. quasi-coprime [HHZ15]. quasi-distance [ROGS16]. quasi-geostrophic [CCD18]. quasi-interpolants [SST15]. Quasi-interpolation [RV16, GW15b, WD16, ZYY17, ZZG18]. quasi-linear [MKS19]. Quasi-Monte [Har15]. quasi-Newton [ZN15]. quasi-Newtonian [LX16a]. Quasi-periodic [DT16, ALN16, ACC16, DNC^{+}17, KLS15, WTXZ16]. Quasi-projection [CKSV19]. quasi-residual [TZ18b]. Quasi-semiregular [FHK^{+}19].

Quasi-state [LGC19b].

Quasi-synchronization [WY18, LGC19b]. quasi-variational [Ied15]. quasi-Wilson [AAGZ19, SZL^{+}18]. quasiconformal [SA18b]. quasilinear [TMS15, ZMT15]. quasilinearization [VRC15, VK15, VK17a]. Quaternion [ACKS16, yWyLyY^{+}15, BDE^{+}15, HWZ17, Kyr15, Kyr17, LGC19b, LCJ15, PAM19, QBC19, RWH15, SWY18, TLX19, TZD^{+}19, XzXqT^{+}15, ZWLZ15, Zha16d]. Quaternion-valued [ACKS16, LGC19b, PAM19, QBC19, TLX19, TZD^{+}19]. quaternionic [EÖ17, ZJJ15]. quaternions [LCJ15]. Quenching [Zho17, Bea19]. queue [Amm15, CSA18, KKD17, SJ19a, SJ19b, Upa16, YLW16, ZWD15]. queueing [KGGAV^{+}18, DKDD15]. queues [JBS15, PD15, Shi15b]. quicker [You15]. Quickest [KPP18]. QuickhullDisk [NSR^{+}19]. Quintic [AK18b, LX15, Bej17, KAI19, KD16, Tho17, WTSQ15, YWL18]. quipus [GKS16]. quotient [DBJC17, Mor15]. quotient-difference [DBJC17].

R [Pan15, CBGGV17]. R&D

[ZZ19b, ZZC^{+}19b]. Rabinowitz [GS15b]. Rachford [BCH15, LZ17a]. Radial [LL16a, XHW18, ZMBJ19, ZGZ16, AD17, BH17, Boy15, DA17, GN15, HAC18, HH18a, KKT15, LLC19, RPB15, SJ18, SN18c, Su19, YDCS15, ZHS18]. radially [HLLT15]. radiating [SPGP18]. Radiation [AAC16, MS18b, SBH16, XP17, DME^{+}15, DMS^{+}18, RMD18, RPJ^{+}18, SMG^{+}18, UHLM15, YWYS18]. Radiative [HLC^{+}17, CKG^{+}16, EM15, HCY15, PMV16, SE18]. radiative-conductive [CKG^{+}16]. radical [YF17]. radii [KLQY18, LMZW16]. Radio [Oli17, FXC17]. Radius [AS19b, CZ17a, CLWM19, DGF17, HLW15, HMGC18, LWM17, LCM19, LSZ15, XLYZ19, ZL16b, ZS14]. Rainbow [ML17, BS19c, Che18b, LLS15, LLW17, LT19, MNJW19, SRT18, SJL17]. Ramanujan [Che15a, Che18a, LSM15, MV15]. ramp [AKR15]. ramp-type [AKR15]. Ramsey [Jin17, LLL19c]. Randić [AG15b, CZ17c, DD16a, GA18, LL15a, LY15b, LBRS18, Shi15a, STD15]. Random [BCD^{+}19, ES17b, MX17, AK19b, CÁGGLP16, CCNQ19, CCJ15, CCL^{+}16, CNQRR18, DLY^{+}17a, DKDD15, Gat18a, Gav15, GDL19, LHHS16, MA16, MPS15, MT15, NRAEC19, PTBN19, PRD19, SS19b, SHEBS19, UPMV18, Wan17a, WQK15, WGL18, WL19a, WYZ^{+}17, YCD17, ZZ16b, ZWL16, ZTW16, ZXDP18]. random-keys [ZXDP18]. randomized [VMLO19]. randomly [HXW^{+}19a, LMW18, MC15a, WZYS15, ZML^{+}17, ZMC^{+}19b]. range [GLS16, GLS19, LLT19, MM15a, MAR15, Zha18a]. Rank [JT17, AL19, AP18b, BCSG18, FHLL19, HZ16, LL17a, LWZ15, QY15, WHM^{+}19, WL17, ZYH15, Zhu15]. rank-two [AP18b]. Rank/inertia [JT17]. Ranking [Rez15]. rankings [AGM16]. rapidly [DEK15, lDqCWzH18, HCY15]. rarefied [FFG^{+}16]. Rate [IAK15, YSWZ17, BR15, CKW17, dCCGV15, CP15a, CLR16a, CZM18, CDL15, Fan18, GS15a, GS17b, KSP16, LC15a, SZ18b, XGQW19, ZYSL19, ZZ15d, ZL17b, ZMH18, uRAD19]. rates [AK15a, HL15a, JGKL16, Mi17, QG16, XKG15, YSL18, ZKKZ18, ZDF17]. ratio [CSA18, HL15a, HYZ^{+}18, KB17, LKM17, PL18, QM15a, SL15b, WPL15, Zam19]. ratio-dependent [HL15a, HYZ^{+}18, PL18, SL15b, WPL15]. Rational [dCM16, Far16, ZEF19, ZM15a, AQ18, CLC^{+}19, FKV16, Fuk15d, Fuk15b, Gra18, Han15c, Han18, IKHS16, IS17, KCK19, LZS15, LLz18, LCW19a, LLdM15, MZM16, MRQZ19, OW18, RSKB16, SS18a, SS15a, TW15a, TF18, YF17, Zad16, Zha15c, Zha17b, ZX15, Zhu18]. rationality [Els17, HZYW15, PL15c, Ued19]. Rationalized [Bag19, EGB15]. Ratios [CIPT15, AF15, Che15b, GS18b, KS19b, LWW19, SK15]. Raviart [YBZ18]. Rayleigh [GM19]. Rayleigh-Love [GM19]. Razumikhin [CC15]. Razumikhin-type [CC15]. RBES [RLGGAV17]. RBF [GM17b, KC19, ODP17, PeO18, Su19]. RBF-collocation [KC19]. RBF-FD [GM17b, ODP17]. RBFs [ZYY17]. re [SSGG17, SY18c, YZ15c]. re-infection [SSGG17]. Re-nonnegative [SY18c, YZ15c]. Re-positive [YZ15c]. Reachability [ZZ16c]. reachable [LXW^{+}18, TNPL15]. reactances [BB18b]. Reaction [ZHW^{+}18, AAA^{+}15a, Ata16, Ban18, Bea15, BKW18, Bho15, BHM15, CG15, CDL19, CD15d, CJ16, EFPT18, FSW16, GN15, GPHAPRW16, KK16b, LRF^{+}18, LC16a, LCAA17, LZCW18, LRL18, Mou15, PST18, Pol19, PCS17, RW17, SHA15a, SZ19a, SMG^{+}18, TTZ^{+}19, TDL^{+}17, VD16, WZZ18, WA18, WM18b, WSYL18, WNW^{+}19, ZNV19, ZPZ^{+}19, ZRLX19, ZTW16]. reaction-convection-diffusion [PCS17]. reaction-diffusion [Ata16, Bea15, Bho15, BHM15, CG15, CDL19, CD15d, CJ16, EFPT18, KK16b, LC16a, LCAA17, Mou15, PST18, Pol19, SZ19a, TTZ^{+}19, TDL^{+}17, WZZ18, WSYL18, ZPZ^{+}19, ZRLX19, ZTW16]. reactions [Mah19]. reactive [CVJ19, KB18]. reactor [Sah19b]. reactors [Ber17]. reagent [JMK17]. Real [PG15b, WWZ15, AL18d, BCSG18, CCYT19, CMR17, CM15b, DO17, GBS15, HS17, LS15a, LZ15a, LWSX18a, LCW19a, LABK15, Lin15, QZL18a, RWH15, TES15, VB15, WLJ15, ZWZ19a]. real- [AL18d]. real-parameter [LABK15]. real-rootedness [CM15b]. real-valued [ZWZ19a]. real-world [LWSX18a]. realistic [GGAVGGRC15]. realization [GCAA18, SGJS16, SMM18, YYL19, ZY18a]. reappearance [XSS^{+}17]. rearrangements [EFL18]. Reasoning [LWGS17]. reciprocators [FK18b]. reciprocity [HL15c, KWT15, NT19, Tan17b, WWN^{+}15]. reckoning [TTP16]. Recognition [TH19, GS19, KSS19, SHLL18, TWF^{+}18, VK17b]. Recognizing [LVAB^{+}16]. recommendation [ZLZ^{+}19]. recommender [DR19, HRLGÁ18]. reconstructed [SCS^{+}15]. Reconstruction [DDH18, HLH17, KWGRP16, Liu16a, PS16, YLH^{+}19, ALMCDH^{+}18, ASKK19, CTT^{+}19, DG18b, DS16b, FHPV18, GK16a, Guo16, KNK15, LDSD19, MPR^{+}17, SDK15, UHN^{+}18]. reconstructions [WYLG19]. recovering [WW19]. recovery [DG18a, FNS19, LYXL15, YLX18, YP18]. recruitment [MM16]. rectangle [BGMZ17]. rectangles [GSV17]. rectangular [BM18a, cCpSkZ15, CLT18, HG18, KM15a, LR19b, LZ17f, MWY15, MB17, PSY^{+}17, SLW16b]. rectilinear [CGK^{+}17]. recurrence [CW18, WN15]. Recurrent [WMS17, WHX18, WZC17]. recursive [CFdCdS18, KTD15, LW18b, NMS^{+}18, PBMP15]. recursive-partitioning [PBMP15]. red [AZL15, PW19b]. Redistribution [Du19]. reduce [GW15c]. Reduced [Far18a, Gao18, SLWX15, UHLM15, YLCW18, AP18b, CVMGPPGC15, DQH18, DMA^{+}19, KSL19, LHLL19, LJC16, LT18b, MM18a, NRAEC19, OA18, PY19b, SB16, Xu15c, YUK19, YJSW16, ZZ18c]. Reduced-order [YLCW18, DQH18, LHLL19, LJC16, LT18b, PY19b]. reduced-space [NRAEC19]. reducibility [KLS15]. reducible [Yue16]. reduction [AHG15, BBF^{+}18, BZ15, CJ16, DHQ19, DLjZ18, FMC15, GLW17, GR19b, GLA^{+}17, GIS18, IB15, KAU15, Kud19b, QTF^{+}16, QYF^{+}18, SY15a, TTP19]. Reductions [RRM16, TB15]. redundancy [SSS15]. redundant [LLZ^{+}19, MND^{+}15]. reference [DZD^{+}16, FXXC19, ILP^{+}15]. Refinable [CEMR16, HR16, Rom19]. Refined [JDL15, BLM^{+}16, BJM16, SST15]. refinement [BDH16, KKAK15, KKS^{+}16, PP12, PP15b, Raz15, ZWF15]. reflection [CT17a, YJY16]. reflective [DMR15]. reformulated [JQL15, MMDG16]. refuge [JAU15, KB17, ZCFW19]. regard [Rét19]. regarding [Du17b]. regime [AHK^{+}19, HW17c, KRF17, LJS18, ZJO^{+}18]. regimes [GYW15, LL18c]. region [FLS18, LP16b, PZ16a, FMRRT15]. Regional [CGCK19]. regions [CJN19, LZ19c, MO16, MPLS19, PBMP15, SKA16]. registration [AHLR19, HW19]. regression [CLR16b, CL18, FFTB16, GZ18, Lah18, LK19b, SY15a, SZ18b, YSD^{+}17, ZBKS16, Zam19]. Regula [SM15d]. Regular [Ugu19, VZ18, APS16, Deh16, GWX17, Jin17, LYW17, PSPÁV15, SZL^{+}18, TLZ^{+}19, WX18]. Regularity [WL19a, YLL18, CCD18, CXW18, HMW16, Qia16a, Qia16b, RG18a, YZ17b, Zha17c, ZWaY18]. Regularization [ASS16, BPZB18, CPST16, CZLM18, DGZ16, FLMR18, JSG15, KRS18, RR15, Red19, SGJS16, TZF15, THM19, XCHW16, XX19b]. Regularized [WZC^{+}18, ASKK19, CPP15, GM15a, Gon18, HEA15a, HEA16, MC17, SHLW16, SW17c, WDG19b, YWRP16, ZZ15a, ZN15]. regulariztion [LL17a]. regularly [BHRW16]. regulated [ACC15]. regulation [FLYG18, ZWY19]. regulatory [CWD^{+}19, HXW^{+}19b, JPZ^{+}19, LZG16, ZZZ15a]. reinfection [ATN^{+}19]. reinforced [KKG18, LZL15a, LZL17, MT15, NKKK18, WR15, ZLL15b]. reinforcement [ZZMS19]. reinsurance [MLS19]. Reissner [RCM15]. rejection [GWSC16, KSM^{+}19, LNWS18]. relapse [ELP^{+}18]. related [APK15, CP15b, Che16b, CP17, DAS17, DB18, DD19, DWS19, FFN18, GHM18, KR15, LL17b, MR19, PAK15, RRTR15, SDAB17, SGG15a, WL16b, Xu15c, ZZ16a]. relatedness [ES15]. relation [CXL16b, DY15]. Relations [DGNA15, XGG18, DÖA15, EJF15, GYY18, HLT15, JCX19]. Relationship [FHLL19, NC16]. Relationships [DY15, NY18]. relative [BF15, Bor15b, Els17]. relatively [Nak15]. relativistic [AR15, CY15b, MM15a]. Relaxation [KL16, AHP19, AL18d, DB16, tFyZZ15, Fan15a, Gon18, HZ16, JLL15b, Kim19, Kri19b, LJS17, MTJ15, RNM15, SRBA15, WZLP18, YJW16, Yun15, Bor19]. relaxations [CJ17b]. relaxed [tFyZ15, FZLZ19, HMX15, LZ17c, LZH^{+}18, Zha15b, ZZG17]. relaxing [WwQqTW18]. relay [FXC17]. release [CGPT15, KA15c, MBM16]. Reliability [NS15b, DMP19, Elm15, LXPD15, LH19c, PWFY18, ZXL15]. Reliable [LMW18, MC15a, SAM^{+}16, SKKL17, KWGRP16, LPKS16, RS15a, SRSZ15, SWZ^{+}19, YSZ19, ZPZ^{+}19]. ReLU [LWWL19]. Remanufacturing [WLLL15]. Remark [MMM18, LB15b, LTL16, TZ19b, YLX18, YZ17b, Zha15g, ZL19b]. Remarks [DIS17, OGC19, Qia16b, Yos19, ZWaY18, AZL16, KST16b, LZZ^{+}19, Sza15, Tom12, Tom15]. remission [KG15]. removal [KGGAV^{+}18, BPZB18, KMS18, SHLW16, WHM^{+}19]. renewable [WLL^{+}19]. Renewal [Zha18d, KGGAV^{+}18, CLW18, MM16]. reorthogonalization [LG19]. repair [KLW15, MJSA19]. repairable [NS15b, ZXL15]. repairmen [KLW15]. repairs [YLW16]. repeated [CYSY16, EEST16, HZK^{+}19, HI15a, SMZ^{+}17]. repetitive [SSKA16]. replacement [HL15c]. replenishment [PMSB15, RZR16]. Replicator [WHC18, TES15]. reporting [HCS19]. Representation [KNT17, Pos17, AZ18, ALQ15, CN19a, GI15, KKT19, Kra18a, MMS17b, SDGU16, SHLL18, TG17]. Representations [PS15a, ZL19a, CGPR15, CN15a, DYW15, GT15a, Kyr15, Kyr17, MS19b, MD18b, SR16a, Vis15, ZC18, ZL19b]. Representing [YCH19]. Reproducing [BAV18, DWT17, ASASM16, AP19, And18, DWTT17, FB19, GQC15, GFG15, KMBA18, LWT^{+}15, LLW19, SAA19, TG16]. reproduction [GCLG15, YG15a]. repulsive [CVA18]. Reputation [CZM^{+}19, LSX19, WWW^{+}17]. Reputation-based [CZM^{+}19, LSX19]. required [PZ19]. requirement [PLL16]. research [FLS^{+}15, FPKF15]. Researches [YSZ19]. resembled [DL15b]. reservation [DKDD15]. reservoir [NFS^{+}15, RRT15]. Residual [NI19, BÇS15, BD18, BF15, IH15, LG15, LjL15b, TZ18b, Yüz16b, ZW15d]. residuals [VS19]. resilience [AAE19]. resistance [HZL18, LTC16, SWA^{+}19, TDS15]. resistance-distance [HZL18]. Resistances [JY19]. resistant [Vid18]. resistivity [Sar17]. Resolution [BR17, Cho18, MT15, uIHA18, Tan17b, Ino16]. resolvent [AR16a, Far18b, TB17a]. resolves [LSLL19, LGL^{+}19b]. Resolving [GlH^{+}19, SRGW15, TT16]. Resolving-power [SRGW15]. resonance [JS15, Jia15, RS15d, Uzu17, WXZ18, XJZ19, YWL18]. resonances [FG18]. Resonant [Har16, ZLH16]. resonator [PFH^{+}15]. Resource [YZ15a, DCK^{+}18, QTXY16, WHC18, Wan19b]. resources [ARR16, LJ17, LWZ18, WLL^{+}19]. respect [AH18b, BMR17, GA17, Jur17, KA15b, LXPD15, LQZ16, LWG16, LW18e, Tep19, Wan17b, YW18a, ZL16b]. respecting [PKHD18]. Response [KS19b, BN18, Ber15, BF16a, hCWW^{+}15, DDD19, FFL16, Hac17, Kha17, KB17, KCK18, LXZ19b, Nak18, RCM15, RS15d, SL15b, SH15d, WPL15, WJHA17, XXZ^{+}18]. responses [CWFL15, LWGS17, MRF18, RRT15, WGLZ16]. Restarted [LH19b]. resting [LZL15a, Rad15, ZLL15b]. restoration [CC18a, GK16a, WLZ19a, ZHW^{+}18]. restoring [YD18]. restrained [SLW16b]. Restricted [BMRW19, BGMO18, BM19b, LpLxG18, SPGP18, SG15, YLX18, ZGL15]. result [BMO15, HW18, LLW17, LXZ19a, MAG15, Mis16, ND15, WwQqTW18, YCYL18, YL19b]. resultant [QYF^{+}18]. resulting [Che15b]. results [AB15a, AN15, BS19c, CLL16, CHRD16, CLS15, CLQ16, Du17b, ES17a, GW16, GWW15, GS17b, JHWZ15, JSW^{+}18, KS18, Kim15b, LRF^{+}18, LS15a, LWM17, LL15f, MCS^{+}19, MRC^{+}18, MCO19, MHZ18, MPP17, MAK15a, MAK15c, Our16, RDLG16, RZD^{+}19, SLZ17, SAT15, WM15, WJZF19]. retail [ZW15c]. Retailer [SCB15, MBM16]. retarded [AED15, Bou18, CÖ15, JS15, LW19c, DFG^{+}19, Pi15]. retinal [ARACCA^{+}18]. RETRACTED [Rho14]. Retraction [AZ19, EH15, FMB^{+}19, Rho19]. retrial [KGGAV^{+}18, DKDD15, JBS15, PL15a, PD15, SJ19a, SJ19b, Shi15b, Upa16]. returns [FHM16]. reversal [NEK18]. Reverse [Mos16, ZHWW15]. Reversibility [HPR18]. reversible [CFSGM17, CN19a, SdR19, TZ19a, XH19a]. reversion [ZBL15]. Review [HYL^{+}19b, WWSM19, FPKF15, Tor16]. revised [JLW18]. revisited [AAMCSV18, CD15a, CCR18, DESS15]. revolution [WF15a]. Reynolds [ACG19, WF15a]. RHC [GXLL19]. rhythms [HGG16]. rib [ATK^{+}16]. rib-microchannel [ATK^{+}16]. Ribiére [AKA19]. riblet [DMA^{+}19]. Riccati [DLQ17, Gua19, KS16a, PST18, Sug17, WSZ18]. Riccati-type [Sug17]. Rich [CSWJ15, Wan15a, Moh19]. Richards [KMP15, RRTR15]. Richards-type [RRTR15]. Richardson [SN18a, SN18b]. Rickart [MM17]. Riemann [ARL15, AAMCSV18, BKLW15, CS19, DT16, HL17, HXM18, HSSH18, KSZ15, MR19, STVCC17, Tha18b, WBL17, YW16, ZLD18, ZM19]. Riemannian [BF15, Fio16, QA19, SA19]. Riesz [CDL19, MDHD18, ZLD18, Zha18b, ZLX19]. Rift [PTA16]. right [ACMT15b, ACMT15a, Kac15, LQ15, LJ19, SLF18, YZ19]. right-hand [ACMT15a, Kac15, YZ19]. rigid [BCSG18, HWY15b, XJW17]. rigid-flexible [XJW17]. Rigorous [GYW15, LY15c]. ring [RSO^{+}18, SLF18]. rings [MM17, RD15]. Ripa [SLdLD16]. Risk [PCW15, ASV17, CLL^{+}15, CLW18, DKZ16, GLL15a, HCS19, HHR17, HD16b, KCKZ15, LZX15, LL17b, LY18c, PK15, RP19, SPP17, YLZT17, ZH17, ZXD15, HW17c]. risk-dominant [YLZT17]. risks [WL16d]. Ritz [FY18, ZLL15b, ZZLY15]. RKN [LD18a]. RLW [HP15a, WMSP19]. Ro [BHD18]. Ro-vibrational [BHD18]. robbers [WWW15]. Robin [All18b, CL19a, RGA19, Sap18, SR17, YÖ18]. robot [ZLWS15]. robotic [dSH17]. Robust [CTT^{+}19, CWHY15, CCL^{+}16, DCJ19, GSPH19, KRA^{+}18b, LCY16, LSLZ18, LZ18c, LGCS19, LGK15, LNWS18, MJL16, NPT19, QZCJ19, RSLL^{+}19, SM15b, SY18a, SSS15, SLS^{+}18, TN19, VMLO19, WLZY17, XZLL17, XXZY19, YCX^{+}19, YSD^{+}17, YmZHC15, ZSS16, ZCZM15, ZMC^{+}19b, ZYWW16, BAR18, CLX17, CXP16, GLYT16, GN19b, JAN16, Kim19, KSZ15, LAVME18, LSZ^{+}18, OA18, PAM19, PC16, SH15b, XKWG15, XS19a, XCY18, Zam19, ZWY^{+}15a]. robustness [JCX19]. rock [DY17]. rock-scissors-paper [DY17]. rocks [ES17b]. rods [Xu19b]. Roe [Kem16, Pel17]. Roe-type [Kem16, Pel17]. Roesser [AAS^{+}17, Che15e, IKMW15]. ROF [CCC^{+}17]. Rogue [ZLH16, ZSX15]. ROI [BPZB18]. Role [ARV18, KB17, Sah15, ZY16, De 18, KN19, PC17, SG17]. roles [WL19b]. Roman [YWLL18]. room [CFS18, GW19]. Root [GL15a, GKN16, How15b, LKM17, MWH^{+}18, PH17, PV16, Raz15, SA16d, WWZ15, ZWY^{+}15a, vLtTBI18]. root-finders [vLtTBI18]. root-finding [How15b, PH17, PV16]. Rooted [GSV15]. rootedness [CM15b]. roots [BCMT15a, CP15a, dCM16, CN15c, LCW19a, LSZ15, NCS16, QZL18a, SB15, SM15d, VLCT18]. Rose [BK18b, GJX^{+}19, LWZ^{+}19b, PRA^{+}19]. Rosenau [AO15c, HP15a, WMSP19]. Rossby [LFY18, ZYLY19]. Rossiter [DeC19]. rotary [MTB18]. rotating [Che15f, DDY15, DS15, HYL^{+}19b, ILP^{+}15, SMR15, YJY16, Zha15a, ZWM16]. Rotation [OTH16, BGSCGL19, Far16, HYL^{+}19b, KK15b, LAVME18, PAK15]. rotation-minimizing [Far16]. rotation/curvature [HYL^{+}19b]. rotational [HQJ^{+}15]. rotationally [SLW16b]. Rotations [DRS16, ILPE17]. rotor [WLJ15]. rough [HLFZ19, San16]. routines [FFTB16]. routing [GWC^{+}16, HRLGÁ18, ZLLP18]. row [Pal15]. RPIM [MWLL17]. Ruin [LZZ15, DKZ16, LL17b, MSH15, RP19]. rule [AZL15, KPLC15, Li17c, LW18c, LX19c, MZP15, NT19, QXW19, SdR19, SWY18, XDG^{+}19]. Ruled [LLJ19]. rules [AGC19, APTS15, BN18, DAQI15, GGAVGG^{+}18, HZ17a, Red19, RLH15]. Rumor [TD19, BGS19, LJYH19, LWY^{+}19, Lu19]. run [GHN^{+}16]. run-up [GHN^{+}16]. Runge [ABI19, AAH19, BdS19, CWL19, FT16, FT17, FKR15, HMD19, LZ16, Li18, LS15b, MD15a, QZL^{+}17, RN18, TLL16, Tan18, TZ18c, TSZ19, TZ19a, Tsi19, WY16, YNMW16, ZS15b, Zha17a, ZS19a, ZZX18]. runtime [AD15]. rupture [LYS17, LG18, WYL^{+}19b]. Ruscheweyh [LO15]. Rushton [CKKF16]. RVT [SHEBS19]. s [ABGL17, DH17, GSPH19, KSM^{+}19, KRA^{+}18b, LSLZ18, LZWZ19, Zha19b, ZLW18]. s-consistent [ABGL17]. S-DDM [ZLW18]. Sabin [GK16b]. SABR [LGO17]. saddle [CY16b, CL19b, CC16, CM15a, DLF15, tFyZ15, tFyZZ15, FZLZ19, HM15d, HMX15, HM15b, HWXC17, HM18c, KM17a, LW15c, LWK15, LSjL17, LZ17c, LZ19e, LL17c, LJ15b, Mia17, SL15a, TXM15, WLLK17, YLW15, Yun15, ZLL15a, ZZG17, ZM16]. saddle-point [HM15b, TXM15, YLW15]. Saddlepoint [Gat18a]. Sadia [PK18a]. Saeed [WL18a]. Sage [ABMBR15]. Sakiadis [MS17a]. Salagean [LO15]. same [ATN^{+}19, CFSGM17, GKS16, ND15]. sample [DLQ17, OLZ19, YSF16]. Sampled [LLX^{+}19, WHWW19, ZZLZ17, CCL^{+}16, Dan18, GWWL17, GSPH19, KRA^{+}18b, LXC^{+}19, LS19c, MLSX18, OLZ19, SKKL17, SS15c, WZAF15, WSH^{+}18, WCB19, YCX^{+}19, ZZL^{+}18a]. Sampled-data [LLX^{+}19, WHWW19, ZZLZ17, CCL^{+}16, GWWL17, GSPH19, KRA^{+}18b, LXC^{+}19, OLZ19, SKKL17, SS15c, WZAF15, WSH^{+}18, WCB19, ZZL^{+}18a]. samples [PS16, Rog19]. Sampling [AQ15, CCGV19, MS15a, RLSGFMMF16, SWZ^{+}19, YWS^{+}19, ZZZY17]. sand [DQW^{+}19]. Sard [HMS15]. Sathre [Mor15]. satisfied [Kor18]. satisfying [SGMH16, YCH19]. saturable [WW16b]. saturated [Jai18, LC16b, MKN15b, NM19, RL15, SZJ19, SML^{+}15]. saturating [LLZ17, SCM19]. saturation [FM16, LLY19, LLX^{+}19, MFJZ15, MJL16, QG16, QKGW18, SJWK18, SLLC19, XGQW19]. saving [PSP17]. savings [FHM16]. Sawada [WXW18]. Saxton [GLW15]. scalability [SB16]. scalar [AAH19, Bac15, BB15d, CS16a, CYSY16, Gla19, HZK^{+}19, SMZ^{+}17]. scalarization [ROGS16]. Scale [AKK15, BGSCGL19, AO19, CK17, CLSS19, DCW15, DKP16, DNC^{+}17, FA15, FFTB16, HLFZ19, KÇ18a, Kar19, LG15, LWGS17, LMMZ18, LSL19b, PR15a, QTF^{+}16, RF18, SLWX15, TTD18, Wan17a, WL19b, WWHZ18, XZLZ17, ZWZ19a, ZW15d]. scale-free [SLWX15, WL19b]. scaled [LCL17a, MSH15, NAM18, ZJYD17]. scales [AR16a, AY16a, DWZ15, Gol19, HO15, LT15, NDZ15, Pac19, Reh18, SZ19b, SZ17, WA15a, WHX18, XH16, ZWK16]. Scaling [SCN15, DLH15, OP19, SSN15, WSdSY15]. scalogram [BBFJ17]. scattered [ACD18, JCW^{+}15, MH19b]. scattering [ANU16, AAC16, KDS16, LG18, LWSM16, QBNF19, RDM19]. scenario [JYVS15]. Schauder [BZ15]. Scheduling [GYY18, Abd15, Chy16, HSKHA15, MTY16, MMB15, PGLG17, WYW16, XYZK15, XY15b, YJYH15, ZXDP18]. scheme [Ach19, AU15, ACRZ18, ABGL17, ADS16, AD19, AAMC19, BCST19, Bho15, CXL15, CDL19, Cho18, CT15, CT17d, DN15a, DC16, EO18, EdLC18, GGHL15, GR15a, GfSmL16, HJMB18, HEA15b, HW19, HK15, HFCS16, HP15a, HP16, HKW^{+}17, HJKP18, Hua16, JAN16, KC19, KB18, KNK15, KG19, KE18, LP15a, LLCR15, LCQ15, LW18f, LZ19c, LR19c, LYZ19, LCL17b, LWWL19, LT18a, LJC16, LT18b, Maj18, MN17, MO16, Mis16, NRY16, OL19, PZ17, PFH^{+}15, PP12, PP15b, PCS17, PN15b, QZL^{+}17, QXW19, RR18b, RC17a, Red19, RS15b, RS16a, SLdLD16, SMI18, SLC18, SR17, SzSC19, SWL16, SPG17, SN15c, SuSB15, SSJ15, TST16, TY18, TTP16, WXY15, WJWX18, jWgX18, WDG19b, WRX^{+}19, WZK15, XW19, XCL18, Xu19c, YMD15, YNMW16, ZEF19, ZSY17, ZZY18, ZJJ18, ZZ18c, ZZMS19, ZZL18b, ZLX19, ZZ17b, ZLW18, ZSY18]. schemes [AP18b, AK19b, AO15c, Bej17, BR16b, BS19b, BVJ15, BMRS18, CWG15, CCY19a, CGM16, CCGP16, CZ17b, Cho19b, CL19c, Cui15, DB16, Ewa16, eFJY17, FM18, GCMR18, GIS18, GLL15b, HM17, HGL19, Kem16, LL15e, Li17b, Lin19, Liu15a, DPL18a, MVS19, MVRV19, MH19b, NI19, NS17a, PSSX17, PP19, RF18, RLZ19, RS15c, RzS15, SdA18, SSS17, SN15a, SBL16, SuSR15, SuSR16, STA^{+}18, SzShG16, TZH15, Vab15, WYLG19, WLQ^{+}19, YGZ19, YG18, YU15, Yos19, ZB17, ZZS^{+}19, ZZX19, ZCCL18]. schlieren [GG18]. Scholes [HH18a, RM18a]. schooling [MMS18]. Schrödinger [ZZL15a, ACE19b, AL18d, AHK^{+}19, BW15a, BBCI18, CLY17, CCY19a, CY15b, DLjZ18, ES17a, Esl16, Fau17, GS15b, GR15b, HGH19b, ID19, IS17, JZ19a, KAU15, KKT19, Kor18, KNT17, Kra18a, LdCM19, LZ18b, LJL18, LSZF18, Lin19, LW18h, MZ17, MLT17, RWW15, SLW15, SLW16a, SPG17, WW16b, jWgX18, WX19a, YY15, ZA16, ZLL15c, ZZY^{+}17]. Schrödinger-type [ZA16]. Schur [LZZT18, LpLxG18, NC19, RMSD16, SX15, Xu17b]. Schur-bounds [SX15]. Schurer [Acu15, AFK15, BM15b, KIAG16, MS16b, SIA15, CZ16, KS18]. Schwartz [MBHS18]. Schwarz [AL18d, MNNP19, ZGL15, Zha18a]. science [CST^{+}15, DZY^{+}18, WWR^{+}15]. scissors [DY17]. scroll [GMOGCCC15, SSPCPG^{+}16]. SCW [WZ16b, ZW15e]. SDD [CEP15, LLL17b]. SDE [GLM16]. SDEs [Prz15, YG18, Yue19]. SDFEM [LZ15d, Luk15, ZL16a]. SDP [CJ17b]. SE [HDG15]. SE-DEA [HDG15]. seal [WLJ15]. seal-rotor [WLJ15]. search [Abd15, BFJ^{+}17, BK18d, DLH15, FIG^{+}19, fGlHyL^{+}15, HMF15, KKST16, KVS15, LCW^{+}18, LF15, LWWW15, MZ15, MHZC15, MKG15, MJSA19, NRAEC19, PTO19, PZ16a, RRTR15, SRH15, VVD15, WCZ15, ACC15]. searches [FFGRSV19, GTBC18, LG15, RMA15]. Searching [LS15b]. seasonal [WXP16, WF15c, ZSW^{+}17]. seasonal-influenza [ZSW^{+}17]. Secant [AM15e, MA15b, BK18d, CFR18, EHVV15, HVR16, WCZ15, ACMT15a]. secant-like [EHVV15, HVR16]. Second [DÇO15, DNC^{+}17, GFE19, LL15e, LY15b, LSL19b, ML15, ABL15, AÖ16, AZL16, AZ18, AM15c, ASASM16, AADLF15, Alz15, AS15, ASM16, BMM19, BDGS15, BS19b, BHRW16, Can15, CDJ19, CC18b, CLWM19, CNQRR18, CZ15c, DA17, DGZ16, DWZ15, DJL17, DBV18, DQ18, FK17, FM17a, FT18, FJ19, FR16, FM18, GY15, HWY15a, HK18a, HPR16, JYS16, LL15c, LZ17b, LZjy18, LZ18a, LD18a, LZ19b, Liu15e, LCL17b, LS19c, Lv16, Lv18, MRE18, Mic15a, Mic15b, MMZ19, MGMM17, MSL15, Mus18, Özb15, PMC15, Pi15, PS15b, RRM16, RSKB16, RR18a, RN18, Rem18, Rom19, SWUE19, SM16, SSCX18, SWL16, Sug17, SD19b, TZF15, Tua15, WA15b, WS19c, YCZ19, YTH19, YJMH18, Yüz18, ZPRZ18, ZZCY16, ZJ15b, ZYLY17, ZZL18b, ZZQ18]. second-maximum [FT18]. second-neighbor [FJ19]. Second-order [DNC^{+}17, GFE19, LL15e, LSL19b, ABL15, AZ18, ASASM16, Alz15, ASM16, BMM19, BS19b, CDJ19, CNQRR18, FR16, HWY15a, HPR16, JYS16, LZ17b, LZjy18, LZ18a, LD18a, Liu15e, LS19c, Lv16, Lv18, MGMM17, RSKB16, RN18, Rem18, SSCX18, Sug17, TZF15, Tua15, YJMH18, Yüz18, ZPRZ18, ZJ15b, ZYLY17, ZZL18b]. section [AAF15, CT15, CT17d, FFH^{+}18, PP12, PP15b]. sections [SV15a]. Sector [DKP16, ARR16, JZ19b, LLLZ16]. sector-bounded [JZ19b]. Sector-delayed-Hopf-type [DKP16]. sectorial [CLL16, WIF15, ZZ15b]. sectors [Gar16a]. secure [WZK15]. security [FXC17]. sedentary [WC16]. seed [LTC16]. seeking [Sha16a]. Segel [ZZZ16]. segment [ZWZ19b]. segmentation [ARACCA^{+}18, AD15, EGH^{+}19, LL16a, TM18, TDA17, YSX15, ZPRZ18]. segments [CGK^{+}17, WGL15]. segregation [SZO16]. SEI [DTJ18]. Seidel [EHS17, LZZ^{+}19, TTLX17]. Seidel-type [LZZ^{+}19, TTLX17]. SEIR [AELH15, DTJ18, DAQI15, DCK^{+}18, XWZ15]. selected [ZS16b]. Selection [HH18a, SCR^{+}15, YLZT17, CFS18, CRZ19, DZD^{+}16, FXC17, HDG15, HD16b, HW15, KPLC15, LZX15, LWWZ17, QTXY16, SBLZ17, TIT15, XS19a, YF17, YFLF17, YSD^{+}17, dSRTZ18]. Selective [PSM17, zMnZyZ15]. Self [ARS15, DZD^{+}16, LLT19, ACC15, AADF18, Far18a, HYZ^{+}18, KB18, KTD15, KDTP18, Liu15g, STL17, SLTH15, TLLD19, TBG16, WWDW17, XGG18, ZSX15, ZZW17]. self-accelerators [SLTH15]. Self-Adaptive [ARS15, DZD^{+}16]. self-adjoint [Liu15g]. self-adjointness [TBG16]. self-consistent [ZSX15]. self-intersection [Far18a]. self-learning [WWDW17]. Self-organization [LLT19]. self-organized [HYZ^{+}18]. self-organizing [KTD15, KDTP18]. self-regulated [ACC15]. self-similar [STL17]. self-triggered [WS19a]. self-tuning [KB18]. Semi [Ber15, CM18, FM18, Yil16, Bej17, BB19a, BFR18, BCFR18, CLJ^{+}15, CJ17b, DC16, tFyZZ15, GLM19, GZW15, HN14, HXW^{+}19a, JK15, JIS15, KD18, KJ15b, KML19, LQC^{+}18, LDS^{+}18, LT18a, NN19, PS18a, Pis18, QZCJ19, RLZ19, SZ18b, Tch15, VK15, VK17a, WMZ16, WZZ18, WH18, WRX^{+}19, WHWW19, Xu15c, XSSW19, XGQW19, YLW15, YLP19, YYS17, YWZ18, ZCAN19, ZCCZ19, ZYWW16]. Semi-analytical [Ber15, CM18, KD18, KJ15b]. semi-cardinal [Bej17]. semi-classical [BCFR18]. semi-convergence [tFyZZ15, WH18, YLW15]. Semi-discrete [Yil16, GZW15, Xu15c]. semi-discretized [PS18a]. semi-explicit [Pis18]. semi-groups [JK15]. Semi-implicit [FM18, CLJ^{+}15, DC16, GLM19, JIS15, LT18a]. semi-infinite [BB19a, CJ17b, YWZ18]. semi-Lagrangian [BFR18]. semi-linear [RLZ19, VK15, VK17a, WZZ18]. semi-Markov [HXW^{+}19a, LQC^{+}18, LDS^{+}18, WMZ16, WRX^{+}19, WHWW19, XSSW19, YYS17, ZCCZ19]. semi-Markovian [KML19, QZCJ19, XGQW19, YLP19, ZCAN19, ZYWW16]. semi-smooth [Tch15]. semi-supervised [SZ18b]. semiaxis [DO17]. semiclassical [AHK^{+}19, GGMPC15]. semiconductor [YLL^{+}16]. Semigroups [HO15, AÇ18, Liu19]. semilinear [AS15, BCI15, CZP15, DX17, GZK16, HFCS16, JX19, SY17b, TM19, XH19b, ZRLX19, ZZX18]. Semilocal [SGMH16, WK16b, AM17, EHVV15, MA15b, Pro16]. semipositone [WLZW15]. semiregular [FHK^{+}19]. semivariance [NM15]. Semrl [DKM15]. senior [YSLS17]. sense [DSSC19, GG19b, HMS15]. sensing [WNF19, ZWY15b]. sensitive [TLL^{+}18]. Sensitivity [SFLM19, PTA16, RL15, VVD15]. sensitization [HYL^{+}19b]. sensitized [DeC19]. Sensor [LPS19, LLZ^{+}19, LMW18, SZRC15, ZLDZ16]. sensors [Kia15]. separable [Abd15, CY16a, Pol19]. separated [ANT^{+}16]. separation [AKK15, BM19c, HZ16, HL15c, Par18a]. separations [HI15a]. Separatrix [FHPV18]. septal [PBMP15]. septic [KZ16]. sequence [Abd15, AAF15, Agr15, BD15, CS15b, LCM19, LS15b, NL18b, PGLG17, XZCC15, ZWZ19b]. sequence-dependent [Abd15, PGLG17]. sequences [AK15c, Bor15a, Cat19, CGJZ18, CT18, DEK15, Dok18, Flo18, Gou18, Haz16, KK15a, LSY15, Yan18a, YZ18a, YYL19, ZY18a]. sequencing [DCA15, ARS15]. sequential [AN15, AL18b, MNYL18, ZCL18]. Serial [ZPRZ18, GYY18]. serial-batch [GYY18]. Series [Mor17, ABMBR15, AP19, BB15a, BD18, BBFJ17, BRM^{+}19, Boy17, Dan18, DQW^{+}19, FMRRT15, GDS16, HA19, KNT17, MS19b, NS15b, PHM19, Par15, RSLL^{+}19, Rog19, SK17a, SSS15, YSX15, YS16, YSF16, SM16]. series-parallel [SSS15]. serotype [ATN^{+}19, VEF19]. server [Shi15b]. servers [CLC^{+}19, DKDD15]. service [JBS15, KWH17, PL15a, ZW15c]. set [ABD16, BFJ^{+}17, CHL17, CK17, CYW18, DH19b, DH17, DL15b, FM18, GRV15, JLL15a, JCX19, KS19a, LL16a, LXL17, LXW^{+}18, Sha15f, TLL^{+}18, Tua15, bWzZjH16, Yan18a]. set-theoretic [DH17]. set-valued [CHL17, JCX19, Tua15, bWzZjH16]. Sets [CDK17a, AGM16, AVC^{+}19, CGG^{+}18, CLHC15, DZL15, EGH^{+}19, GHM18, HGG16, HR16, Har15, HP15b, HLFZ19, LYW18, MM17, NS15a, O’R18b, Pop17, Rog19, RG18a, SRGW15, TNPL15, WTZL18, YCH19]. sets-based [EGH^{+}19]. setting [AGM16, LZ15c, Pöt19, Prz15, TES15]. settling [CCL^{+}15]. setup [Abd15, PGLG17, SLS16]. setups [KWT15]. seven [SZ16]. seventh [ACDT18, YD18]. seventh-order [ACDT18]. Several [QW19, SLTH15, CLS15, CC17, CN15c, CN16b, NS17a]. sextic [LL19a]. sexually [IYM19, WCH19]. SGPA [SM15a]. shadows [DMS^{+}18]. Shafer [DJW19, Nis15]. shallow [BMT18, BKLW15, EdLC18, GWK16, GYW15, KNK15, Kem16, KK16b, LCQ15, NB15, SSG^{+}15a, Tha18b, WLQ^{+}19, WMSP19, hY18]. shallow-water [BMT18]. Shape [Han15c, Han18, PL15b, SHH15, BH17, DL15b, EGH^{+}19, ERFE16, GWPA18, HH18a, HP15b, HLJ18, LX19d, SA19, SN18c, WC15a, YLH^{+}19, ZH15b]. Shape-preserving [Han15c, Han18]. shaped [GHN^{+}16, HM17, JA17, KBK15]. shaping [Pai15b]. Shapley [AMCMFJ15, YZ15a]. shared [Xu19e]. shared-memory [Xu19e]. sharing [LMW^{+}19, XFTX19]. Sharma [Pan15, S.15]. Sharp [ABQ15, AK16b, LMZW16, YCZ15, ZY18b, GI15, MRS18, RD15, ZZM18]. Sharpening [Nis15]. Shaw [ÁL15a, CS15a, CK18, Kim15a]. Shear [GB19, AAM19, BKZA18, CS15a, KMR18, NMNTBXVD15, dDSI19]. shear-dependent [AAM19]. sheared [WHM^{+}19]. shearlets [BPZB18]. sheet [BAR18, EM15, HAMA14, MP18, MWY15, PMV16, PK18a, SS18a, ZZLY15]. Sheffer [KK15a]. shelf [MMB15]. shelf-life [MMB15]. shell [FA18, KAMD18, MRF18, RSAF16, RS15d]. shells [Muk16]. Sherman [SKP16, Xu17b]. shift [CM15a, CN15a, CN19a, DZD^{+}16, HWXC17, LCZ19, LL17c, LG19, Rem18, TQC15]. shift-splitting [CM15a, HWXC17, LCZ19, LL17c]. shifted [BHD18, HS18]. shipwrecks [KZL19]. Shishkin [DN15a, LZ17e, ZL16a]. Shock [Gla19, Ery15, MCPO15, PBBF18]. shocks [VG18]. shooting [MS17a]. shop [Abd15, XYZK15]. short [BB18b, Hac17, MM15a, WXZ18]. short-circuit [BB18b]. short-span [Hac17]. shortage [JYVS15]. Shortcut [CGG^{+}18]. shortest [CRC15, SNAR15, WL17]. should [FK18b]. shrinkage [KKM15]. shrinking [BAR18, MP18, PMV16]. Shubik [AMCMFJ15]. SI [LZT^{+}18]. side [ACMT15b, ACMT15a, Cat17, ZBKS16]. sided [CLJ^{+}15, CLW18, DCS15, FZLT15, KCKZ15, NT16, NPT19, TY18]. sides [Kac15, YZ19]. Sierpinski [Cal18, DSZY19, LZ17g]. Sierpiński-like [DSZY19, LZ17g]. sieving [SD17]. sign [AU15, APS16, CFG15, CCC19, Hag15, KZZ15]. sign-changing [CCC19]. Signal [DZ17a, BM19d, HTZY17, LW18a, LY17, LLz18, LLTL18, RXLM19, Suk19, SD19c]. signals [BMDS17, LLC^{+}15, LD19b, WYCZ17, YMeFG19]. signed [HFK19, HTW19, LGHW19, ZLWW15]. significantly [WWC^{+}15]. signless [CZ17a, CT17c, LWM17, LMZW16, XLS18, YFY17]. Signorini [LD19a, ZL15a, Zha18c]. silence [WWN^{+}15]. SIMD [SD17]. SIMD-based [SD17]. similar [STL17]. similarities [LK15, Val18]. similarity [DV15, DCS^{+}19, GTBC18, JJ15a, MRR17, SH19]. Simple [BDSS15, AIG17, BL18, CD16b, GH17, JA17, LLCR15, NCS16, SM15d]. simplest [cCpSkZ15, RP18, VDV15]. simplex [wG15a, PS15c, SCY19]. simplex-type [wG15a]. simplices [BCK15]. simplicial [TG17]. Simplified [Dan15a, GK16a, HHP18, KS15, LYC15, LZ19e, STK15, ZS15a, Zhu15]. simply [YBZ18]. Simpson [DLY17b, Li17c, LW18c, WNMC15]. simulate [DMA^{+}19, GGAVGGRC15]. simulated [Mat15, Pal15, PTO19, RRTR15]. Simulating [CCL^{+}15, PFH^{+}15, LXL17, SPP17, ZZX19]. Simulation [BG17, How15a, NKKK18, ZZ18b, BS19a, BL18, Cha19, CC18b, CDFP17, DKL15, DWT17, DQW^{+}19, FFG^{+}16, FFH^{+}18, FFH^{+}15, GJR^{+}15, GZG^{+}15, GAT18b, GE16, HSAR19, ID19, KÇ18a, KRF17, KDH19, KKS^{+}16, KFHH15, Kri18, LGO17, LYDH18, LZY16, MHR18, DFG^{+}19, MD17, NBAGS19, NB15, Oli17, Par18a, PL15b, PBMP15, RAH19a, RSC16, RG18b, SK15, SH15b, SMG^{+}18, SCZ^{+}17, SH15d, TVK^{+}16, TPS15, VBG^{+}17, VG18, WXWX18, WXM^{+}18, Xu19e, YLL^{+}16, ZLC^{+}18, ZWY19, lZA^{+}19]. simulation-optimization [SH15b]. simulations [BMM19, BBGR15, CS18, CLP^{+}18, EEG^{+}16, GCM^{+}18, GYZC15, KMB16, LJ19, LLL18d, MLG18, NMS^{+}18, PG15a, PMS16, PP19, RVE19, TT16, UHLM15, ZKV^{+}15, ZHS18]. simulator [GG18]. Simultaneous [HWZ17, LDSD19, LZHC18, RZW18, UDA15, ZLLZ16, GSV18a, Pro16, PIP19, PV19, RZR16]. simultaneously [TM18]. Sinc [Nag17, Oka18a, PJ15, SR19, AQ15, Dan18, KXZ19, Tha15, Oka18b]. Sinc-Chebyshev [Nag17]. sinc-Gaussian [Tha15]. Sine [LLL18c, Sun15b, AK16b, JSLW17, MVRV19, PFA15, RS15a]. Sine-Gordon [Sun15b, JSLW17, MVRV19, PFA15, RS15a]. Sinelshchikov [GR17b, TTXZ16]. Single [XY15b, CXW^{+}19, DKBS15, DL17d, DMH19, FLYG18, HSKHA15, Kri19b, DPL18a, MKG15, Pal15, SHLL18, SB16, SML^{+}15, Uzu17, WQK15, YY15, ZZLY15, ZYZ15, SM16]. single-channel [SB16]. single-layered [ZZLY15]. Single-machine [XY15b, HSKHA15]. single-network [CXW^{+}19]. single-period [WQK15]. single-phase [Kri19b]. single-relaxation-time [Kri19b]. single-row [Pal15]. single-species [DL17d, ZYZ15]. single-velocity [DPL18a]. single-well [YY15]. Singular [ACE19b, ASM16, Boy15, DHQ19, GGZR16, Li19, AD16, AHS18, AST18, AK15a, AD17, Bea15, CY16b, CC16, CM15a, Che15e, CLCW17, DB18, DTM19, DH15, tFyZZ15, FLL15, FM16, GHQ17, GR19a, JGX15, KMBA18, KML19, KPP17, Kyr17, Lah18, LR16a, LX16b, Li17e, LZCL18, LX19a, LZ17c, LBB18, LXW^{+}18, LMW18, LWZZ18, MHN15, MFJZ15, MJL16, ML15, NL18a, NS17b, PHM19, Pan18, Par15, QXW19, RG19, RGA19, RMK19, SWUE19, SH15a, SHA15b, SSA18, SPST17, SRS17, SZ16, TB17b, Wan15b, Wan16a, WZ16b, WLH^{+}18, XXW^{+}19, XZHX17, YLW15, YKJY17, YY17, YTH19, YmZHC15, YWH15, ZMC19a, ZLL15a, ZLWW15, ZMQ16, ZCW19a, ZMC^{+}19b, ZY15d, ZW15e]. singular-logarithmic [AD17]. singular-ordinary [ZMQ16]. singularities [Boy17, CZ15a, DL17b, NB19b, ODP17]. Singularity [YLQ17, LWZ^{+}19b, YL19b, ZS16a]. Singularly [KS16b, AK18b, BZ16, CJ16, CVA18, DN15a, Fan18, GQC15, GN19b, KK17, KH17, LSAR15, LDS^{+}18, Liu18a, MN17, PCS17, SAA19, SR17, SN18a, SN18b, SRCA17, SM18, TG16]. sinh [ZXZB19]. sinh-cosh-Gordon [ZXZB19]. sink [Har16]. SIQR [LJS18]. SIR [BW15b, LYZ19, LC16b, SHEBS19, WGSX19, WYZ^{+}17]. SIR-like [WGSX19]. SIR-type [SHEBS19]. SIRS [CKW17, Wan17a]. SIRVS [Zha15e]. SIS [GLM16, KM15d, LZ15c, ZYZ16b]. Sivashinsky [BT15b]. six [DPL18a, SN15c]. six-equation [DPL18a]. Sixth [PVL19, BCST19, EKO19, GKN15a, GKN16, HJMB18, LMMR15, NBK16, RMK19, Sha15e, SB15, WK16a]. Sixth-order [PVL19, GKN15a, GKN16, HJMB18, LMMR15, NBK16, RMK19, WK16a]. Size [Pet15b, CLT^{+}15, GLM16, GXCY19, HI15b, KM15d, QLZ15, RSKB15, RSKB16, RRW^{+}15, SB16, SGKR19, YZMM15, ZCLS18, ZSZW16, ZYZ^{+}16a]. sizes [AO19, WTZL18, ZXDP18]. skeleton [SKCS15]. Skew [JH17, TZ18d, DS19, tFyZ15, GWX17, JH15, Kyr15, LZL15a, LWZ15, QY15, Wan16c, YZ15b, ZLL15a, ZS15e, ZWT15, ZZL15b]. skew-Hamiltonian [ZY19]. skew-Hermitian [DS19, tFyZ15, ZLL15a, ZWT15, ZZL15b]. skew-rank [LWZ15, QY15]. skew-symmetric [YZ15b]. slab [QBNF19, SHA15a]. slag [KSKW^{+}19]. slanted [RMD18]. slantlet [XS15]. slantlet-Walsh [XS15]. slave [LPKS16, SLZ^{+}16]. Sliding [CPCQ17, HZK^{+}19, JGKL16, AADLF15, CTT^{+}19, CL19c, GZC19, HYL19a, JGX15, KML19, LZ18c, LGCS19, LPS19, LGK15, LMW18, Pai15a, RXZW17, WXXY19, XZLL17, ZYLY17, ZJYD17, ZMC^{+}19b, ZYWW16]. sliding-mode [LMW18, XZLL17]. slip [BTH18, EAY17, KPG16, Pon17, Qiu18, SSB18, SSG15b]. slip/no [BTH18]. slip/no-slip [BTH18]. SLK [XY15b]. slope [KVS15]. slow [WLH^{+}18]. slowly [BB15b]. Smagorinsky [ALZ16, SLL19]. Small [Yos19, BM19d, CLHC15, DLQ17, EOP17, GST15a, Har15, HRS17, HZYW15, KLS15, LLT19, LSAR15, RR15, Rog19, ZS16c]. small-signal [BM19d]. small-time [ZS16c]. small-world [EOP17]. Smallest [CSZ19, WS19c, ZCEW18]. smart [GGAVGGRC15]. smoking [uRAD19]. smooth [And18, GI16, Gar16a, LL15d, LLL18c, LL19d, MJ19, PW19b, QTTL19, SDGU16, Seg15, Seg18, Tch15, WL16c]. Smoothed [GSP15, LL19b, NMNTBXVD15, Ton18]. Smoothed-Particle [GSP15]. smoother [Nak18, ZZL16a]. Smoothing [KPT16, LW15d, LLTL18, WLZ19b, Zha15i]. smoothness [BJM16, Hua16, LXZ19a, NRY16, ZSXW18]. Sneyd [Mis16]. snowdrift [YF17]. Sobolev [DFJ16, LCW15, LR19b, LT18b, MMMMB17, PR15b, SHA15b]. Social [HWL^{+}19a, SSM^{+}18, AJ15, BK18a, CST^{+}15, HCS19, HCY^{+}18, KZL19, LWGS17, LZP18, LSX19, LSLL19, LGL^{+}19b, LWY^{+}19, Lu19, Pla16, Pla17, SLLZ19]. socle [LLL19b, TZ18b]. soft [KDS16, NEK18]. software [Xu15a]. Sokolov [SKBR18]. solar [DME^{+}15, DMS^{+}18, MSdC^{+}18]. SOLD [LK18]. solid [GSDH15, KAMD18, LWLM19, RZMB17, SFC^{+}17, TSPF18]. solid-gas [RZMB17]. solid-shell [KAMD18]. solidification [SWSS18]. solids [BZ17, LL19b]. Solitary [KZ16, AM18a, GR17b, LFY18, MZ17, NRT16, TNMR18, VDI17, Wan16b, WMSP19]. Solitary-wave [KZ16]. Soliton [CY15b, WTJ17, Bas19, ES17a, GL15b, ID19, MDH16, MM18a, OW18, WTSQ15, YRA15, ZZC15, MMIA19].

Solitons

[KRB15, HGH19b, KSD^{+}15, LZWZ19]. Solomon [CGM16]. solubility [Liu17b]. solute [KA15c]. Solution [ADMV19b, CGL15, MK15a, NB19a, RWH15, YCYL18, eMAGG19, AKS15, AM15a, AG15a, AP15, AD17, Bab16, Bag19, BAV18, BM15a, BS17, BM18a, BC15, Boy17, BCI15, BVJ15, CVJ19, CM18, CL16c, CCZ15, CJK17, Cho19a, CN19b, CNQRR18, DK15b, DA17, DA19, DJW19, DHS19, DKKR16, DL16, Dol15, EFPT18, EGB15, FLL17, FKR15, FM18, GMS19, GFG15, GN15, GT15c, GR15b, GR17b, GS11, HN14, HH16, HHP18, HJ15, HH18a, HSAR19, HP15b, HG18, HLLT15, HGM15b, HM18b, HHG19, IÖ19, JMK17, Kac15, KCK16, KE16, KOR19, KZA19, KNBA15, KHN19, KK16b, KMT16, Kue15, KG17, KH17, LZL19a, LP15a, LG16, LBB18, Liu15f, Liu18c, LDH18, LZ19g, MDH16, MM04, MRE18, MC17, MHCF15, ML15, MH15b, MH16a, MS19c, MR16a, MDH15, MO15a, MO15b, MRR17, Nag17, NL18a]. solution [Neu16, NN19, OL19, PS18a, PMC15, Par15, PG15b, Pis18, Pop17, RR18a, RM18a, Rem18, Reu15, RME18, RSO^{+}18, Sab18, SDS18, SM19, SLC18, SSB18, SR15e, SZ19b, SS18c, SR16b, SSA18, Sha15f, SH18a, SS16, SPS16, SPST17, Sin18, SHEBS19, STVCC17, SLL^{+}16, Tan15c, Ton18, TB17b, USCA15, USCAAA15, VS19, XZCC15, XSLW17, YB15, ZEF19, ZS15d, Zha15f, ZLL15b, ZLW15, ZMMZ15, ZMQ16, ZZHT17, ZZG18, Zha18c, ZM19, ZCU15, ZSY15, ZYZ15, ZXL15, ZG15b, Zho15b]. Solutions [AIIK15, BM16, GS15b, GM19, WN15, eMA15, eMBHA15, ALL18a, AM18a, AOS19, AL18b, AZ18, AZ14, AIG17, Ali15, AST18, ASS16, APL15, ALP15, AV18, AK15b, ABN^{+}17, AZAM15, AZ19, BÇS15, BEM15, Bas19, Bea15, BB15c, BRD15, BZ17, BHM15, BdlRT18, BHRW16, CFG15, CRCVMS18, CDK17b, CZP15, Che15d, CL15b, CCY19b, CD15c, CY15b, CB17, CY15c, Cov19, DT16, DGZ15, DH15, DBV18, DZ16, DFP15, DCR15, DM18, Dim16, DZL15, DO15, DLCT18, DB17, EJF15, EJF16, EM15, ES17a, Esl16, Fan17, FMC15, FCM17, FFT15, GGL15, GO15b, GS15a, GD15, Ghe16, Gra15, GL15b, HA16, HLZ15, HK15, HGH19a, HPN15, HARS15, HPR16, HZ17b, HHY19, IKMW15, II18, IH15, IS17, JJ15a, JLX15a, JS16, JT17, JG17, KZ16, KD18, KAH15, Kor18, KNT17, Kra18a]. solutions [KZ15, KR16, Kud16, Kud19a, KC18b, KH17, LJZ18, LdCM19, LSfW15, LWS15, LL18a, LZCL18, LVV18, LZ15b, Liu15c, Liu16b, LLL18c, LL19d, LFY18, LXZ19c, LD19c, Lv16, Lv18, MDZ15, MG15, MWWM15, Ma17a, MJ19, MPW15, MMIA19, MHM15, MZM15, MLT17, MYHS18, Mi17, Mig15, Mig16, MNR19, MMZ19, Mil16, MM18b, MS16a, MSS15, Miy17, Nas19, NRT16, NE15, OW18, ÖK19, PHM19, PKB15, PWS17, PYZS18, PW19b, PZ15, PST18, Pla16, PS18b, Pol19, PFA15, Pos17, QZ15b, RS15a, RAV15, Reh17, Rui18, RC17b, S.15, STJS16, SR15a, SR15c, SJD^{+}18, Sak19, SHA15a, SS18b, SEKA16, SC18b, SXS15, SPGP18, STL17, SY18c, Sun15b, SX16, TES15, TNMR18, Tho17, TLHC17, TY16, TTXZ16, UKE15, UBT18, VD16, VD15, VDV15, VDI17, WW16a, Wan15c, WTSQ15, WLZW15, WA15b, WL16c, WTXZ16, XHW18, XPG15, YTH19, YU15]. solutions [YRA15, YL19b, YMHW15, YZ15b, YZ15c, YW18c, Yüz15, Yüz16b, ZBR15, ZA16, ZBBW17, ZSL15, ZH15a, ZZC15, ZAW15, ZLL15c, ZY15b, ZZZ15b, ZS15c, ZM15a, ZWLZ15, ZLL16, Zha18e, ZMZ18, ZLLA19, ZXZB19, ZZL15a, ZCQ15, ZZ15d, ZCX17, ZL17b, ZMT15, Zho15a, Zho15e, ZW15e, ZLL18, ZS17, dSH15, uRIS17]. Solvability [Jia15, LJ19, RAH19a, HMZ17, HSAR19]. solvable [BCSG18, RM18b, Zha18b]. solve [ABD16, CCNQ19, CKW15, CJ16, DKN16, EA15b, FB19, KEE17, KAS16, MH15a, SAC18, SNAR15, SSJ15, Yüz16a]. solver [AS19a, BKLW15, FK17, FCW^{+}17, Fuk15a, KSZ15, LWW15c, XZ19]. solvers [AKK15, DPSS16, SA16d, SAS19, SC17]. solves [HCS19]. Solving [AM15h, BR16a, CNQRR18, Dim17, FBF18, GW15b, Gia15b, KMBA18, KLZT15, KMP15, LWT^{+}15, Liu18a, Mat15, MH17b, PAE^{+}17, PJ15, RSKB15, WWH17, Yan18b, dSH17, AKAA15, ABC^{+}16, eMA15, eMBHA15, AAGZ19, AM15c, AMO18, ASA19, AAS^{+}17, ABA18, ACDT18, AND17, AR16b, AG16, AKKB17, AM18c, ABMA19, ACTV15, ACTV17, AE17a, AD19, AAMC19, AKA19, BCST19, BS16, BAM16, BCMT17, BSZ16, BAR18, Bog16, BS15c, BK18d, CP19, CCJ15, CKG19, CJ15, CG19, CLMT15, CGMT16, DCW15, DLF15, DS19, DZW15, DL18b, DWTT17, EHS17, EAAM^{+}15, EMM17, EE18, FT16, FT17, GWC^{+}16, Gao16, GAB17, Gol15, GS18b, GZ15a, GfSmL16, HAC18, HEA15a, HEA15b, HEA16, HA19, HR18, HH15, HS18, HHGC16, HAY19, HAH19, How16, Hsi15, HM18a, HM15d, HM15b, HM19b, IAW18, IS16, JLY18]. solving [JLL15b, JIS15, JAN16, KKKD18, KJ15a, KS16a, KM17a, KM17b, KAI19, KJ15b, Koc15, KDG19, LJD17, LZ17a, Li18, LZH16, LLL18c, LHZM19, LG19, LjL15b, Maa15, MKG15, MH16b, MH17c, MS19c, MAJR19, MKN^{+}15a, NE16, NJTP19, NAM18, NWNA15, NWN15, PVL19, PeO18, PKS15, QDH^{+}18, RSKB16, RS17a, RLZ19, RB15, SST17, SS18a, SWUE19, SAA19, SH15a, Sah19b, SHR18, SSN15, SZ19a, SN15b, SSK15, SSB16, Sha16b, SHC^{+}17, Sil16, SDK15, Sow18, SG15, Su16, Tan15b, Tom12, Tom15, VS17, WNMC15, WZM15, WHSS15, WZ16b, Wan17c, WA18, WNN16, WZLP18, XY15a, XY17, XY18, XM15, jXlZX18, XYG^{+}18, YMD15, YDWZ19, YR15, YMS16, YGZ19, YWRP16, YJSW16, Yüz18, ZMBJ19, ZV17, ZNV19, ZL15a, ZMQ16, ZGH17, ZD17, ZHX^{+}17, ZVL19, ZWT15, ZX16, ZT16c]. Some [Ade19, AZL16, ABA18, CLL16, CL15b, CHL18, CNNA16, CK18, Fen15, HASA15, HZL18, HXLC15, Jah17, JRSK16, JJ15a, KS18, KST16b, Kur18, LR19a, LR16a, LWM17, LWW19, LW19c, LZCW19, LT19, LZZ^{+}19, LSM15, Mal15, MNNA15, MKN^{+}15a, MCI16, MAK15a, MAK15c, NY15, NNA15a, NNA15b, QM15a, QM15b, Rho14, Rho19, SC18a, Seg18, SN15b, SA16d, SLF18, SPKC16, SGG15a, SKA16, SzShG16, Tom12, Tom15, WM15, WDB18, YFL16, YZ18a, YDLK15, You15, ZWL18, ZSF17, dFY15, AA16, ADG16, BGMZ17, BGMO18, BR16a, BS19c, CEG^{+}15, Cat17, CWD16, CC17, CY18, CKDC16, DY15, EAAAS15, EJF16, FNS15, GFD16, HARS15, HHZ18, HLT16, HP15c, jJ15c, KA15b, KE16, KK15a, KMA16, LX18, Li17f, LGQL19, Liu18c, MNJW19, Mah18a, MK15a, MF17, MFOP17, MMG16, MMM18, MS19c, MS16a, MPP17, Mor17, NMM16]. some [PÖB15, PL15d, PFA15, RDLG16, Rét19, STB16, Sha16b, SIA16, SY15b, Sun15a, YP18, Zha18a, ZMZ18, ZC19c, dFKMM15]. Soner [DO18]. sonic [VG18]. SOR [KM17b, LM19]. SOR-like [LM19, KM17b]. Soret [MRR17, PMV16]. sort [LTY16, LZX19, ZW19a]. sorting [TD18, TRRP15]. sound [FM17a]. source [GHZ19, Har16, HLIK19, KKG18, LW18g, LM15, NM19, Pan15, RZW18, Sha12, Slo15, SM18, THM19, TTK15, XX19b, ZY15d, ddGL15]. source/sink [Har16]. sources [EB16, ZSX15, ZMT15]. South [FMRRT15]. sowing [LTC16]. SP [GL18b, SNAR15]. Space [LJL19, AG15c, AK15c, BD18, Boy15, CPST16, Cha19, CRC15, CLJ^{+}15, CGW16, CZL18, CXK19, CDL19, CP15d, DCS15, DKKR16, DF16, DMH19, FZLT15, FGPR18, GR17a, GT15b, GFG15, GTBC18, GLS19, Han15a, HFCS16, HMS15, jJ15c, KDH19, KK16b, LL15e, Li16a, LW18g, LSZ^{+}19b, LZ19d, LSZF18, LLG^{+}15, LLJ19, MDHD18, MDH15, MKS19, MS15d, NS17a, NRAEC19, Ort15b, RR15, RWW15, Rem18, RZW18, SST17, SDS18, SA19, SR15d, SST15, Seg15, SHA15b, SLW15, SzShG16, TMS15, THM19, TÖGE18, WXY15, WX19a, XYG^{+}18, XW19, XX19b, YX19b, ZDK15, ZZL16a, ZT16b, ZLC^{+}18, ZJY18, ZRLX19, ZBH^{+}15, ZZ16b, ZZL18b, ZLX19, ZEW19, dSSB^{+}16]. space-based [ZZL16a]. space-dependent [RZW18, XX19b]. space-fractional [CLJ^{+}15, FZLT15, HFCS16, KDH19, LSZF18, MDHD18, THM19, XW19, ZBH^{+}15]. Space-time [LJL19, BD18, Cha19, FGPR18]. spacecraft [HWY15b, KRA^{+}18b]. spaces [ACS15, APTS15, ACKS16, ALP15, AK15b, AM15g, Bab16, bBjJ16, BD15, BCM17, BLM^{+}16, CLL16, CB17, CEHVT15, DL17b, FPUK15, Haz16, JK15, jJ15b, jJ15c, KT16, LR19a, LCW15, LL15c, Li16a, LZCW19, LRH15, Luk15, Mah18a, MSHG16, MDG18, MCI16, Nak15, PS15a, RDLG16, Rah19b, RV16, Wan15c, WIF15, XL19, Zha15h, Zhu16]. Spain [NFS^{+}15]. span [Hac17]. Spanning [CYCW15, LY18b, WZC17, ATX19, HLL16, MB19, RCTZ18, WTD^{+}19, Yan17]. Sparse [CL18, GS19, SN18c, CYW18, DVJ18, FNS19, HZ16, LYXL15, RG18b, SHLL18, WJNW15, YLX18, YP18, Yue16]. sparsity [DHW^{+}19, LJL16]. Spatial [KWT15, MZ17, AYAA^{+}15, ATV18, BDE^{+}15, CSWJ15, hCWW^{+}15, CYGL18, CZM^{+}19, Du19, GW19, GSD^{+}17, Hac17, HD17, HL15c, Li15b, LZP18, LJG^{+}19, LSX19, LJ17, NT19, QZL18b, Sch16, SLS16, SZFS17, TS15, Tan15e, Tan17a, Tan17b, WWN^{+}15, Wan15a, WJNW15, WJ19, WYZ^{+}17, WLL^{+}19, XDG^{+}19, YW19a, ZHW16, ZYLY19, ZZH15]. spatial-temporal [ZYLY19]. spatially [BHM15, CJ18, FXXC19, RW17, SN18c]. Spatiotemporal [HD17, Naa15]. SPDE [MP19a]. SPDEs [KG19, LT18a, TM19, YZ18b]. Special [LPW15, ZWLZ15, BKS18, CKDC16, DGRR16, EKO19, GLS16, JISI18, KR15, LO15, MNJW19, MK16, RSKB16, SST15, TW15b, WQ15, WJGM19, ZS16b, ZL15b]. species [CXMP16, DL17d, GG16, GCLG15, MWY17, Moh19, PSM17, Par18b, SSZL17, XYDL16, XWDW17, ZYZ15]. specific [AG15b, ARK15, JV18]. specification [RRW^{+}15]. spectra [CT17c, DL17a, LS15a, PÖB15, XLS18, dFY15]. Spectral [ABS15, ACE19b, AKA19, JYF19, KLQY18, LB15a, LG15, MHR18, RS16b, WYL19a, ZEW19, Bog18, BFJ^{+}17, CCLS16, CZ18a, CZ17a, CHL18, CLWM19, DN15b, DN16, DGF17, DLH15, Ely15, EE18, FWZ19, GWK16, Ghe16, GWY15, GG19b, GDF^{+}17, HEA16, HGL19, HLW15, HHG19, JZ15a, KMT16, LWM17, LML18a, LZ19d, LCM19, LJL19, LMZW16, LLL18c, LW18h, LL19d, LYY15, MHN15, MMM18, MRR17, PVL19, RNM15, RCM15, SR15b, SM19, SSS17, SPG17, SJ18, Wan18, WX19a, WS19b, WW19, Xy16, XLYZ19, YTH19, YWZ18, ZLWW15, ZL16b, ZZY^{+}17, ZJY18, ZHS18]. Spectrally [FK17]. spectrum [BCD18, CCY19b, Lah18, LW15b, LH17b, MB15b, SMG19, TS18, XZC16a, XZC16b]. Speculations [ZDZ15]. speed [GCH^{+}19, GYW15, KK15b, SD17, XY18, YLH^{+}19, ZL19b]. speeds [WC16, ZL19a]. speedup [dNdOX^{+}15]. SPH [FP18b, Maa15, SRK^{+}19]. sphere [ACD18, Jai18, KDS16, KNK15, LC15a, Pov15, XCHW16]. spheres [KF15]. spherical [CCL^{+}15, CPST16, GWPA18, JJ15a, RMM19, SM15c, SH18a, YJW16]. spheroid [LDL15]. SPI [lZA^{+}19]. SPI-based [lZA^{+}19]. spillover [QZL18b, ZZC^{+}19b]. spin [KSD^{+}15, WJWX18]. spin-1 [WJWX18]. spinning [ALMCDH^{+}18]. Spinor [KL15]. Spiral [dSR15, PRA^{+}19]. spiro [YLWF15, YLWF16]. Spline [And18, FP18a, BSZ16, BLM^{+}16, CL16d, DBK15, ET15, HH19, HMT15a, HQJ^{+}15, IAW18, KZ16, KCK19, KAI19, KPT16, KD16, KF19, LX15, LQZ^{+}18, LLz18, LW19b, Lin19, LFC^{+}19, LHWG16, Mic15b, MD17, RG19, SST15, SR15e, SR16b, TMS15, TSJ16, TW15a, Tho17, UKE15, WWZ15, XzXqT^{+}15, YAIN17, ZV17, ZEF19, ZZG18, ZM17, GR19a]. spline-nonstandard [ZEF19]. splines [AAMC19, Bej17, BLM^{+}16, BJM16, GK16b, LX18, MFC17, PS16, RV16, SDS18, SHH15, Seg18, SuSR16, ZS16a, ZM15b, Zhu18]. Split [LMD19, ARB19, BB18a, Boi15, CL16a, EO18, EÖ17, HH18b, MDZ17, QLZ15, SPW15, SDK15, TRP15, Yue19, ZJJ15, CCC^{+}17, LPZ19]. Split-step [LMD19, ARB19, HH18b, MDZ17, TRP15, Yue19]. Splitting [SE18, AP15, BCH15, CWG15, CM15a, ÇT16, DLBQ19, DS19, EMO18, tFyZ15, FZLZ19, GAB17, GJNC18, HL16, HM15c, HWXC17, KM17a, LPZ19, Li16b, LCZ19, LL17c, MH19a, Maj18, QFS17, Rub19, WMLG17, WZLP18, XL15, ZLL15a, ZZZ16, ZZ17a, ZWZ19a, ZM16, ZWT15, ZCCL18, ZZL15b]. splittings [AHK^{+}19, TLZ^{+}19]. sporadic [TZ18b]. sports [FLS^{+}15]. spread [DCK^{+}18, Ort15a]. spreaders [LWSX18a]. Spreading [WC16, BGS19, FJ19, LJYH19, LWY^{+}19, Lu19, NY18, QZJY15, RS19, SJ16, TD19, WLX18, ZLZ^{+}18, ZWL^{+}19]. spring [JLY18]. SQP [Zhu15]. square [ATX19, BCD^{+}19, DZD^{+}16, DBK15, FB19, LCW15, LC15b, LKZ18, LMD19, MWH^{+}18, NKNA16, PYZS18, RN18, RN19, RS17b, SP15, UBT18, WWZD16, WZC^{+}18, WZF^{+}19, ZWY^{+}15a, ZSY15, ZYX19, fP15]. square-root [ZWY^{+}15a]. squared [AGN15, Bor15b]. squares [CPP15, FLL^{+}19, HGM15b, HM18a, HXLC15, HM18b, JT17, JCW^{+}15, KJ15a, KF19, LY15a, LLC19, MLW18, MO15b, Nak18, YSD^{+}17, YZ15b, YWH15, ZS16a, ZJJ15, ZWLZ15, AM16]. Squeeze [RP16]. squircle [LB15a]. squircle-bounded [LB15a]. Srivastava [Mal15, SGG15b]. SSOR [WH18]. SSOR-like [WH18]. Stabbing [CGK^{+}17]. stabilities [AB15a]. Stability [AM18a, ARB19, AL15b, ACDT18, AAH19, AVCCGAVDL15, BB15d, BC16, CGW19, CGC19, CDK17b, Che15e, CD16a, CS18, CS15a, CGMT16, DKBS15, FHV15, FZLT15, Gao19, GMS19, GK15, GN15, GSVK17, ILPE17, JS19, Kor18, Kri19b, LC16a, LX16b, LQC^{+}18, LR19c, LL19c, LZL15b, LZG16, LZLW17, LFC^{+}19, Liu19, LWZZ18, LZH^{+}18, LJHW19, MM15a, MKMC15, NA19, PRC^{+}19, RWY19, RXL19, SS19a, SFSR18, SCM19, TS15, TZD^{+}19, WPL15, WW15, Wan15f, WSZ16, WMZ16, WW18, WCB19, WXLL19, WZS19, WF15b, WLWD18, XJZ^{+}15, YG18, YZ16, YNMW16, ZLHH15, ZLDZ17, ZHW16, ZSZ17, ZKKZ18, ZZQ18, ALL18a, AO19, BCO15, BDGS15, BDSS15, BDSS18, BB19b, BM19d, BS19b, CJN19, CZS15, CWHY15, CC15, CZM18, CWLA18, CSTH17, DAQI15, DZ15, DLY^{+}17a, EO15, ELP^{+}18, GWW15, GW15a, GHW19, GLL15b, GZLL17, GZD19, HN14, HM19a, HL15a, HS15a, HT16, HH18b]. stability [Hua18, ILP^{+}15, IS17, JHWZ15, JYS16, JSW^{+}18, JX19, JPZ^{+}19, KHN19, Kim15b, KVS15, KM15d, KKL18, LLP17, LJW^{+}15, LSAR15, LC15b, LZ16, Li17a, LZCW18, LW18d, LSZ^{+}18, LZ18a, LZT^{+}18, LSR18, LLX18a, LXD19, LK19a, LSL19a, Lia15a, LZY^{+}17, LZ15b, LTUZ15, LGL16, LDH18, LHS18, LZ19f, LLX^{+}19, LMD19, LHX15, MWY17, MRC^{+}18, Mis16, MDZ17, Nam19, NM19, NHWZ15, NN19, Oni18, PAM19, Pi15, QBC19, QA19, RZSS15, RN18, RN19, RM16, SMI18, SA16b, SEKA16, SZ19b, SSZ^{+}18, SS18c, SLZ17, SL15b, SCC17, SLL^{+}16, SD19c, TB17a, Tan15d, TN19, VMLO19, WL15, WDZ17, WYCZ17, WLZY17, WXX18, WDCJ19, WJZF19, WNW^{+}19, XKWG15, XCCL18, XPG15, XWZ15, XHW15, Xu17a, XXZY19, YSL18, YWL15, YLHS17, YL19a, YSL19, YmZHC15, YMMH15, YB15, ZSS15, ZAP19, ZLW19, ZSL15, ZS15b, ZLW15, ZHJ^{+}17, ZZZY17, ZZ18a, ZCAN19, ZC19a, ZG15a, ZFZK15, ZSY15, ZZX18, Zho15b]. stability [ZZC^{+}19b]. stabilizability [DL17c, GXLL19, TN19]. stabilizable [YMeFG19]. Stabilization [Che15f, EL19a, FZG19, LJL15a, LZLW17, Luo19, MWL^{+}18, WZYS15, WCZY17, YSWZ17, ZLWS15, AJMN18, CLXW17, CZLS18, CWHY15, Che15e, CXZZ18, CGAL19, DLY^{+}17a, DN15c, EL19b, EL19c, FMDZ19, GSPH19, JZ19b, LLZ15, LSMT18, LLZ17, LS16, LML18b, QZCJ19, SZ15, SFSR18, SJL16, TLL^{+}18, Wan15b, WZZL16, WMZ16, WLZY17, WXX18, WDLZ19, WCH15, YSL18, Yil16, ZMC19a, ZZZY17, ZCAN19, dFGAN19, SST18]. stabilization/continuous [AJMN18]. stabilized [BBG19, Qiu18, ZS19b]. stabilizers [PN15a]. Stable [BCMT17, Ghe16, KCSD^{+}18, QZ15b, CWD^{+}19, CLMT15, DG18b, JPZ^{+}19, KG16, MMK17, OLZ19, RSKB15, RS17a, RLZ19, RC17a, RVE19, SBG^{+}17, Tan15a, Ton18, WFZ17, Wan17c, WZWJ18, Xu19e, YZH^{+}19, YU15, YDLK15, Zha18b, dFGAN18, dFGAN19]. Stacked [GS19]. Stäckel [MB15a]. Stackelberg [Ask18, FXC17, LGL19c, PL15c]. stage [AO15b, ABI19, FHV15, GPHAPRW16, HZ16, HWL^{+}19a, Kha17, KB17, LWPL16, LPL17, NM15, TLL16, Tan18, TZ18c, TSZ19, TZ19a, WL17, ZG18, ZZC^{+}19b]. stage-structure [Kha17]. stage-structured [AO15b, FHV15, LPL17]. stages [WLS15]. Stagnation [MP18, BAR18, BGP15]. stagnation-point [BGP15]. stake [ZCLS18]. Stancu [BA17, CZ16, MAK15a, AIP15, Acu15, BM15b, DDM16, Kaj18, KS18, KTA15, MMS17b, MAK15c]. stand [FFH^{+}18]. stand-alone [FFH^{+}18]. standard [CKKF16, FGPR18, LJS18, RSKB15]. standby [LXPD15, MJSA19]. Star [RD15, WWW19, CYY19, LLST19, LTY16, WWW18, Yue16, ZW19a]. Stark [FG18]. starlike [APK15, DÇO15, SA18b, Uça16]. starlikeness [AS19b]. start [MKG15, ZTW19, ZL15c]. starting [Tsi19]. State [GhDwZS17, Ino16, LLZD18b, SML18, AO15b, AL15b, BB18b, BL18, DHX^{+}19, DL17c, DC16, Ery15, GW15a, GSZ15, GZD19, HXW^{+}19a, HI15b, HXW^{+}19b, JCW18, LPJ18, LL15b, LSCK15, LLZ^{+}19, LGC19a, LGC19b, LHS18, LMZ^{+}19, Luo19, MRC^{+}18, NHWZ15, PAM19, PW19a, QBC19, QLLY18, RRRR16, SKKL17, SHA15a, SJ16, SZ15, SJL16, SAT15, SLL^{+}16, VS17, WW16b, WA18, WLH^{+}18, XD15, XCDZ19, XJW17, XZLL17, ZLDZ17, ZMMZ15, ZLM^{+}18, ZXS^{+}18, ZFZK15, ZXL15]. state-dependent [DL17c, GSZ15, PW19a, RRRR16, SJL16, SAT15, XD15]. statefinder [SR15f]. states [Ber17, Fau17, HSSH18, QG19, VD15]. Static [NMNTBXVD15, TLL^{+}18, Ber15, CLCW17, GZC19, LL19b, MN15, SYZP16, WLWA18]. station [DFG^{+}19]. Stationary [LXJ^{+}17, LJHA18, BW15a, DU18, Ers16a, Ers16b, GG19b, LL18c, LD19c, Qiu18, Red19, SLL19, SJW16, SuSR15, SuSR16, TST16, WLFH17, YJK19, ZSK^{+}19, ZY18b, ZZ17b, ZS19b]. statistic [ZWY^{+}15a]. Statistical [KBS17, ZLZW15, ZYZ^{+}16a, ASKK19, BLS15, EGH^{+}19, KVS15, SY18c, Zil15]. Statistics [DZ17b]. stay [DZD^{+}16]. stayed [WYL15]. Steady [BL18, HI15b, LCL17a, QG19, VD15, Ber17, BB18b, CSL19, EM15, MGN18, Neu16, RDG18, SST17, SHA15a, VG18, WA18, ZXL15, ZSF17]. Steady-state [HI15b, SHA15a, WA18, ZXL15]. Steady-states [QG19]. steam [HH16, HHP18]. steel [DDY15, JDL15]. steepest [Fio16, FNS15, GS17b, SDK15]. Stefan [BST18, KZA19, RMM19].

Steffensen [CP19, Hag15, KB15, LMMR15]. Stein [LH19b]. Stein-like [LH19b]. Steiner [Gut16, Mao17, Tra18, ZWZ19b]. Steklov [AK19a, Our16]. STELLA [ARK15]. stencil [ARK15]. stencil-based [ARK15]. stenosed [BG18]. stenoses [JV18, SCZ^{+}17]. stenosis [PPS25, PP18a, RRR19, RRR20]. stenotic [Pon17, SSG15b]. stent [MSM15]. Step [Pet15b, ARB19, AO19, AM18c, BNY15, BK18c, DLBQ19, EO18, EKO19, EA15b, FR16, GJNC18, HH18b, HM15b, LGO17, LC15b, LD18a, LY16b, LMD19, MDZ17, MKS19, PBBF18, QLZ15, RSKB15, RSKB16, RS17a, RP18, RR18a, RN19, RT18a, SGKR19, TRP15, USCA15, USCAAA15, VBG^{+}17, WLFH17, YZMM15, Yue19, ZWZ19a, ZYX19, ZSZW16, ZYZ^{+}16a]. step-size [QLZ15, RSKB15, RSKB16, SGKR19, ZSZW16]. Stepanov [CZP15]. Stepanov-like [CZP15]. Steplength [CRZ19, PL15d, dSRTZ18]. steplengths [GLL15b]. stepping [BES19, BS19b, KK16b, Xu19e]. steps [AM18c]. stepsize [Gon18, Oni18]. Stepwise [Gut18, SY15a]. Stewartson [GML18]. stick [CLR16b]. stick-model [CLR16b]. Stieltjes [DF18]. stiff [FY19, Li16b, NH17, TZ18a, YG15b, ZZ19c]. stiffened [AM15b, NMNTBXVD15]. stiffly [Xu19e]. Stiglitz [TES15]. Still [UK18]. Stirling [BMRW19]. stirred [CKKF16]. Stochastic [AM15a, CWLA18, DLY^{+}17a, ELP^{+}18, GSZ15, Hac17, JPZ^{+}19, LLL16a, LML18b, MD15a, NR16, SWUE19, Wan15b, WSH^{+}18, XL18b, XJZ19, ZPZ^{+}19, ALL18a, ARB19, AEB18, AL15b, Alz15, AS15, AELH15, AAH19, BT15a, CÁGGLP16, CKW17, CJG^{+}18, CLR16a, CGZ17, CZM18, Cov19, DS16a, DK15c, DJ15, DWS19, Du15, Fan15a, Fan17, FMDZ19, FYZ16, FQMR19, FZG19, GW15a, GDL16, Gao17b, GXLL19, GLM19, GR15a, GJP18, GE16, GWXZ18, GZD19, HKL15, HJ15, HMD19, HL15b, HH18b, HC18, JAU15, JA15, JYS16, JZ19b, JZ15b, KMP19, KLZ19, KBK15, KSA19, KRA^{+}18b, LPKS16, LY15d, LJ15a, LC15b, LWZ16, Li17a, LKZ18, LLZD18b, LLZ^{+}19, LXD19, LZ19a, LYZ19, LSL19a, LC15c, LZ15b, LDD15, Liu15d, LC15d, LB15b, LZL15b, LZ15c, LB16a, LC16b, LB16b, LGL16, LDH18, LYZL18, LYM18, LJHA18, LJS18, LXZL19, Liu19, LGL19c, LMD19, LL15f, LRH15, LjL15b, LLB16, LD19c, LTZ^{+}17, Luo19, MDZ15]. stochastic [MHM15, MZM15, MLT17, MLG18, Mil16, MSS15, MGMM17, MDZ17, NBAGS19, NT19, NM15, Nak18, Obr19, OL15, PS18a, PCW15, PYZS18, QG16, QKGW18, QD18, RN18, RN19, RT18a, RWY19, RRTR15, RXL19, SSKA16, SM15b, SRJ^{+}15, SMG19, SAS19, SZPH16, SSZL17, Świ15, ŚKK^{+}18, TTZ^{+}19, TB17a, TRP15, TLGX15, TMPG17, TFLZ19, TIT15, Vad19, Vid18, Wan15c, WDLL15, Wan16a, WWZD16, WJHA17, WCZY17, WDLS18, WY19, WZS19, WWMX16, WL19c, WZ18b, WSZ18, XKWG15, XSLW17, XPG15, XHW15, XDH19, XL19, YL17, YSL18, YPZ18, YCX^{+}19, YSL19, YYWH19, YG15b, YDLK15, YMMH15, YLZT17, YYZ19, ZZL^{+}18a, ZGGY19, ZDPZ15, ZJ15a, ZY15a, ZYXF17, ZXS^{+}18, ZKKZ18, ZYSL19, ZCAN19, ZCCZ19, ZZ19b, Zha19c, ZC19a, ZJ15b, ZFZK15, ZSY15, ZYZ15, Zha15g, ZZ16b, ZY18b, ZYX19, ZBL15, Zho15b, ZTW16, ZH16, ZPM16, ZYZ16b, ZSL^{+}18, ZL15c, ZXX^{+}19, JO15]. stochasticity [GLM16]. stock [Cim18, FHM16, HI15b, Lah18, LZZ15, MZC19, MM16, MTCV^{+}16]. stock-recruitment [MM16]. Stokes [DKKR16, ZZ15d, ABS15, ABGL17, BFR18, BS19b, DL15b, DQ18, DC16, Ers16b, FZLZ19, FFG^{+}16, HMT15a, Jai18, KMB16, LML18a, MPP17, Qiu18, RC17a, Rub19, SH16, TA16, WZWJ18, YLL18, Yil16, ZL17a, Zha17c, ZW18, ZS19b, dFGAN18]. Stokes-Voigt [YLL18]. Stokes/DSMC [FFG^{+}16]. Stolarsky [Jan15]. stone [PG15a]. stopping [Arm19]. storage [HKW^{+}17, UMF19]. Störmer [LZ17b]. Störmer-Cowell [LZ17b]. STP [WLWA18, ZLK^{+}19]. Strain [VB15, AB18, CX19, FG15, HK18b, KM15a, KMR18, LL19b]. Strain-hardening [VB15]. strains [CWFL15]. Strang [ÇT16, EMO18]. Strategic [GOH18]. strategies [BPC15, CZL18, CDZ^{+}18, fGlHyL^{+}15, HR18, KCSD^{+}18, LB15a, LPZL16, LTC16, LLL16b, MN15, PP15a, SNAR15, XTXW15, XXZ^{+}18, YLW16, YJYH15, ZGH16]. Strategy [WLWA18, ZSL17, AKT19, CK17, hCWW^{+}15, EFL18, HW17c, HLC19, KWT15, MLS19, QZJY15, QTTL19, RXZW17, SJ19b, SHC^{+}17, SZT17, Tan17a, Tan17b, TXL19, WS19a, WRZ19, WL19b, XDG^{+}19, YLZT17, ZLWS15, ZH17, ZXD15]. Strategy-updating [ZSL17]. stratification [RNM15]. stratified [LFY18]. Stratonovich [MH17b]. stream [HRS17, SH16]. streamfunction [BFR18]. strength [ACP19, AC19, BLNSF17, Moh19, SSPCPG^{+}16]. Stress [AM15b, AAE19, BKZA18, KM15a, SMG^{+}18, SCM19, VVA^{+}15]. stressed [YJY16]. stresses [HM19a]. stretching [EM15, HAMA14, MP18, PMV16, Pan15, PK18a, SS18a, Sha12]. stretching/shrinking [MP18, PMV16]. strict [ZW19b]. strictly [APS16, Zad16]. strikes [NSN^{+}18]. string [DF18]. strip [QBNF19]. stripe [WHM^{+}19]. Strong [GLM19, GIK19, LYZ19, Nak15, TM19, TY18, Val18, YW18b, YSL19, Yue19, CP15c, DL16, GR17a, Jur17, LML18b, MMS17a, Mos16, PSSX17, WZ18a, ZMH18]. stronger [JGGK15]. Strongly [Wil18, Boi15, GS11, JPTL15, RM19a, RMK19, Sun15a, YW19b, ZS15d]. Structural [ES17b, LTUZ15, DDdF16, Dim17, DL15a, PG15a, WZ19, YwFW^{+}19, Yua15].

Structure [YUK19, ZW19a, BK18a, BK18b, BB18a, BY19, CLZW19, CNK^{+}18, DB16, EKO19, Ili19a, Kha17, KB17, KFHH15, LX16a, LWPL16, LWWW15, LGL16, LZWZ19, MCO19, MCZ15, RRT15, STK15, Ter15, TSA15, TSPF18, WMSP19, Yos19, ZYZ15, ZCML19, ZMC^{+}19b, dSSB^{+}16]. structure-preserving [BB18a, WMSP19, Yos19]. Structured [DLQ17, AO15b, AA16, BK18b, CMW17, FHV15, FG15, JT16, LYZ19, LPL17, Par18b, RS19, WWK19, XGZ17, YLLH19, YW19a, ZZHT17, ZG18, ZwX16]. structures [AZ14, ASM16, AZ19, DH19b, Den15, DLD16, DNC^{+}17, FMSG15, KCWK15, LS19a, MLS19, Pai15b, PP19, RM18b, SZ19c, VB15, WG15b, ZAY18, ZGZ16, dSR15]. student [NSN^{+}18]. studies [AK19a, Guo16, Ili19b, MSHG16, WWZ19]. Study [BG18, CKKF16, CGWW17, MS17a, ADSS16, ADG16, BKZA18, BGMS18, Boy17, CDHK15, CNK^{+}18, CST^{+}15, ERFE16, FFL16, GHN^{+}16, GWPA18, IR16, JMK17, KM15c, Kim18, KKS^{+}16, LL17b, LH17c, LW18d, Liu17a, LZWZ19, DPL18a, LMMR15, MC17, MFF15, MMM18, MK16, MKN15b, NFS^{+}15, PBP19, PMC16, Pon17, QBNF19, RS15a, SK17a, SS16, SSG15b, SG19b, TDA17, VK19, WC15b, YG15a, ZS16a, Zha16b, ZYJ15]. studying [PH17]. Sturm [AYD19, CZ18a, CD15c, Deh16, FFT15, GZZL17, MNR19, PSPÁV15, PS15b, ZW15a, ZjA19]. Sturmian [Özb15]. style [LD17, LLZD18a]. sub [CD15c, DSSC19, HZS15, cJzS15, JPTL15, LW19b, OA18, PeO18, SA16b, WR19, YAIN17]. sub-diffusion [DSSC19, cJzS15, LW19b, PeO18, SA16b, WR19, YAIN17]. sub-feasible [JPTL15]. sub-optimal [OA18]. sub-populations [HZS15]. sub-superlinear [CD15c]. subadditive [Jur17, MK15c]. subcells [LZ19c, MO16]. Subclass [SA18b, DEK15, LH15]. subclasses [SKA16]. subcubic [CR19b]. subdifferential [LZ15e, LLB16]. subdiffusion [HLLT15, LHWG16, OL19, YWZ18]. subdivision [ACRZ18, ADS16, CCGP16, CZ17b, DJLX17, Ewa16, eFJY17, HM17, LC18, MH19b, NZZ16, NRY16, RS15b, RS15c, RS16a, RSAF16, SBL16, SuSR15, SN15c, SuSB15, SuSR16, SX15, TTZX15, TST16, ZZS^{+}19, ZZ17b]. subdivisions [XZC16a]. subdomains [SHHC18]. subdominant [DM18]. subglottal [SH18b]. subgradient [LWW15c]. subgraphs [BKKBS19, LVAB^{+}16]. subgrid [BR17, KÇ18a]. Subharmonic [Zho15e]. SubIval [Sow18]. subject [CCL^{+}16, HXW^{+}19b, Kun16, LSZ^{+}19a, Liu15f, LXTF18, MFJZ15, NT16, NPT19, SMG19, TLGX15, WYCZ17, WHWW19]. subjected [AKR15, NND^{+}18, Pov15]. submerged [SSG^{+}15a]. submultiplicative [Jur17]. subordinations [LO15]. subpopulation [JLX15b]. subquadratic [Lv16]. subsets [GHM18, LZ15a]. subsidy [ZSW^{+}17]. subsonic [Ber15]. subspace [CRC15, LHLL19, Rui18]. subspaces [TQC15]. substepping [LS17b]. Substitution [SGFCRM^{+}18]. substructure [ZW19a]. substructures [LLMW18]. subsystems [CGW19, JPZ^{+}19, NS15b, YZ16, ZMC19a]. subtransient [BB18b]. subtree [LWW19, YwFW^{+}19]. Subtrees [YLWF15, YLWF16]. success [Nad15]. Successive [Bor19, Du15, PWS17]. such [AZ14, AZ19]. suction [MKN15b]. Sufficient [RG18a, ZM18a, CS15b, GXLL19, LZY^{+}17, ZWL18]. Sugeno [CGAL19, DAS17, LZ18c]. suitable [DGN16]. sum [ACP19, AF15, DD16a, DJW19, Du17b, GA17, GA18, GS18b, HHR17, IÖ19, KKST17, LL15a, LQZ16, LGS^{+}18, LW18e, MR19, NC16, NP18, NE15, WX19b, WL18b, XY15b, XLG18, YSY^{+}17, Zha19a]. sum-Balaban [KKST17]. sum-connectivity [DD16a, GA18, LL15a]. sum-of-processed [XY15b]. sum-of-ratios [AF15, GS18b]. summability [Bor15a, KBS17]. summation [AIK15, GM15b, Sha15d]. summation-integral [GM15b]. sums [AGN15, AH18a, ALN16, BCK15, FvdMS15, TYD15, WS19c, XX19a, Xu19a, Zha18d]. Sumudu [GT19, PKB18]. sun [SA16a]. sup [dFGAN18, dFGAN19]. Super [HTM18, WZZT19, CZLM18, IWL17, MK15a, MMG17, UKS^{+}16]. Super-critical [WZZT19]. super-diffusion [IWL17]. super-fast [UKS^{+}16]. super-order [CZLM18]. super-predator [MMG17]. Superconvergence [Bac15, HL16, SLW16a, SY18b, SY19, XS19b, Bej17, DFJ16, LX18, MHZ18, SW17b, SW17a]. Superconvergent [SST15]. superelements [SM15c]. Supereulerian [DLZ18]. superhydrophobic [FK18a]. superlinear [CD15c, LZCL18]. superlinearly [LHJ15]. superorganism [Tre15]. superposition [ZG15a]. superquadratic [Che15d, ZZZ15b]. supersingular [Li17c]. supervised [SZ18b]. SUPG [GS17a]. Supplementary [ZL19b]. supplementation [HHY19]. supplier [Haf17, KPLC15, MBM16, SCB15]. supply [CLT^{+}15, DKL15, PMSB15, WC15b]. supplying [BL15, BL18]. support [APMM18, CS16b, GI16, GlH^{+}19, KSK^{+}15, Lah18, MJSA19, Rom19, YSD^{+}17, ZBKS16]. supported [Kia15, SZ18a, YBZ18]. supports [JA17]. Suppressing [DG15a]. supraglottal [SH18b]. sure [LZ15b, LDH18, RN19, XDH19]. surely [MZM15]. surface [AwEE15, BKL^{+}18, DMA^{+}19, EFPT18, HRS17, KRF17, KHN19, Kud19a, KS19b, Kun16, LWQ18, LG17, LC18, LMZ^{+}19, MKRS16, Pan15, PBMP15, SDGU16, Sha12, SBH17, SCS^{+}15, VVA^{+}15, WJGM19, AAC16]. Surfaces [AK15d, BM19c, Bra15, DP15b, FK18a, GT15b, HH19, HQJ^{+}15, LLJ19, RSAF16, WF15a, XZW^{+}15, YCH19, ZQ18, ZZW17]. surrogate [XRB17]. surrounded [CD16b]. survey [Cat19, Che18b, KSK^{+}15]. Survival [LZ19a, YYZ19]. susceptible [LLL16b]. suspension [Kan15, LWLM19, MX17, YM18]. Sustainable [ZLLP18, Haf17]. Suzuki [TTP16]. SV [CHL17]. SV- [CHL17]. SVEIR [WLZ16]. SVM [MPLS19, NFS^{+}15]. SVM-based [NFS^{+}15]. swallowtail [FLS18]. swaps [CLR16a, YYWH19, ZL15c]. swarm [CYGL18, CS16c, DGL^{+}15, KLZT15, LYC15, ÖÖAD15, SA16a, SBG^{+}17, TIT15, WL19b, YJYH15]. Sweep [AAS^{+}17]. Swift [BAV18, KR16, ZEF19]. swimmer [BTP17]. swimming [BTP17]. switch [XCW19]. Switched [LZLW17, CGW19, DL17c, DCJ19, GWW15, HXW^{+}19b, JPZ^{+}19, LLZ15, LX16b, LSZ^{+}18, LSMT18, LXD19, LGCS19, LPS19, LLLZ16, LLZ17, Liu15e, MFJZ15, RZHY16, RWW18, SSKA16, SKKL17, SFSR18, SJL16, TN19, TCS15, WZYS15, WZZL16, WXX18, WSH^{+}18, WJZF19, WCB19, XPZ18, XZZ^{+}18, XXW^{+}19, XXZY19, YZ16, YMeFG19, ZLLZ16, ZLDZ17, ZMC19a, ZZC16, ZFZK15]. switches [XJZ^{+}15]. Switching [DK17, GGHL15, CGW19, CGZ17, DL17c, Fan17, GHQ17, HL15b, HW17c, HXW^{+}19b, JZ15b, Kur19, LZ19a, LXJ^{+}17, LYM18, LJS18, Obr19, OL15, PAJ^{+}19, QG16, QZCJ19, Sha16a, SSCX18, SWZ^{+}19, SJL16, WXX18, WRX^{+}19, WZS19, YMeFG19, ZPZ^{+}19, ZC19a, ZH16, ZJO^{+}18]. switchings [Wan15b]. Sylvester [DS19, HGM15b, HM18a, HM18b, KM15b, XM15, YZ15b, Zha16d, ZWT15, ZCC15]. Sylvester-conjugate [HM18a, XM15]. Sylvester-transpose [HM18b]. symbol [Bag15]. Symbolic [FT17, PH17, CH18c, NSM16, RRM16]. symbols [JKK15]. Symmetric [Cov19, DH15, GY15, TZ19a, AF18, AM15b, ADSS16, ÁL15a, BBGR15, BFJ^{+}17, CGJV16, CR15, CL19b, DLQ17, ESS16, GDL16, HLLT15, JJ15a, JGBY16, LH17a, LW15d, LXCK17, LM19, LXL17, Tch15, YX19b, YWRP16, YZ15b, ZD17, ZF18, ZWZ19a, ZM15c, ZM18b]. symmetric-triangular [CL19b]. symmetrical [BG18, Can16]. Symmetries [LMF17, STL17, eMBHA15, ACE19b, CS16a, CD15d, EHdLS16, FCM17, RM18b, TTXZ16, dlRGB16]. Symmetry [BHM15, FMC15, HCH16, ZMZ18, AADMM16, Cim18, DOZ18, KMA16, Kri19a, Nas19, SJD^{+}18, SZ18a, XHW18]. Symplectic [LJS17, CWG15, Ely15, Fio15, Fio16, HMD19, MD15a, TLL16, TSZ19, ZYY17]. Symplecticity [TZ18c]. Symplecticity-preserving [TZ18c]. Symposium [ZS16b]. synapse [MNJM19]. synapse-modulated [MNJM19]. synaptic [PC17]. synched [FFGRSV19]. Synchronizability [PAJ^{+}19, SLWX15]. Synchronization [AY16a, CVMGPPGC15, DK17, Fan15b, FYSC18, GCAA18, LSZ^{+}19a, LMZZ19, MGPPGCCV15, MPS15, RMB15, SRCA17, Wu15, YWZS18, YFW^{+}17, ZZL17, ZWBC18, BK18b, BPC15, CPCQ17, CLPY18, DZ17b, FYZ16, GWWL17, GJX^{+}19, HXW^{+}19a, JZ15b, KBR15, LPKS16, LCAA17, LLC18, LTQF19, LGC19b, LZM^{+}18, LDS^{+}18, LS16, LCLZ19, MMZ^{+}17, MLSX18, MGCAPP16, PRC^{+}19, RVGS17, SAJ18, SAM^{+}16, SLZ^{+}16, SSPCPG^{+}16, SS15c, SML18, SGY^{+}15, TTZ^{+}19, TDL^{+}17, WWZD16, WHX18, WY18, WSH^{+}18, WDG19a, WRX^{+}19, WHWW19, WZK15, XLJC18, XKG15, XXZ^{+}18, XCL18, YWS^{+}19, YYS17, ZZLZ17, ZZL^{+}18a, ZPZ^{+}19, ZDPZ15, ZWHC15, ZYXF17, ZHWW17, ZCCZ19, ZCX19, ZDF17]. synchronizing [GWXZ18, LLCR15]. synchronous [ADMV19b, BB18b]. synthesis [BS15a, LQC^{+}18, ZZC16, Zha19b]. System [SM16, dNdOX^{+}15, AN15, ARSV15, APMM18, ABMA19, AS15, KGGAV^{+}18, AADF18, ASM16, BDSS15, Bho15, BL15, BL18, CCY19a, Cao19, CLC^{+}19, CDHK15, CHL15, CKW15, CXMP16, CCL^{+}16, CY18, Che15f, CS18, CLT^{+}15, CZ15c, DMP19, DCT19, DS16a, DHY^{+}18, DSB15, DCR15, DHMT15, DX17, Ery15, EA15b, FAHZ16, FM17a, FDPP18, FMC15, FK18a, FM16, GG16, Gao18, GIS15, GL15b, GYW15, GLW15, HGG16, HW18, HMF15, HP16, HMW16, HJKP18, HRLGÁ18, HST17, Hsi15, HXS^{+}16, HCX^{+}17a, HYZ^{+}18, HCY^{+}18, HSSH18, IJ15, ID19, ILP^{+}15, JGGK15, JL15, JT16, KLZ19, KRS18, KLW15, Kha17, KN19, KNBA15, KL16, KZ15, LXPD15, Li15a, LJW^{+}15, LZS15, LW16, LY17, LKZ18, LH19a, LK19a, LL19c, LZY^{+}17, LC15c, Liu15e, LL16c, LLZ^{+}16, LYZL18, LMW18, LXZL19, LZKX19, LCW19b, LLWW19, LWW15c, LD19c, MWWM15, MM15a, MHX17, MJ19, MGPPGCCV15, MJSA19, MLT17]. system [MAG15, MH17b, MTB18, MFMK17, Moh19, Mon19, NS15b, NWN15, NS17b, OL15, OB18, PL15a, Par18b, PN15a, PW19b, PG15b, PZ16b, Pet15a, PZ19, PP18b, QKGW18, RWH15, RLGGAV15, SA16a, Sah15, SP19, SR15a, SR15c, SKKL17, SS18b, Seg15, SSB15a, SSA18, SLW15, SN18a, SN18b, SRK18, SFC^{+}17, SWY18, SY18c, SNL19, SDIS15, SCZ^{+}17, Su16, SX16, SLS^{+}18, TVK^{+}16, Tal15, TLGX15, TIT15, WLJ15, WLZW15, WA15b, Wan15f, WA15a, Wan16b, WL18a, Wan19b, WD19, WNN16, WAS19, WZK15, WF15b, WLL16, WLL^{+}19, XYDL16, XH16, XZCC15, XZ19, XCW19, Xu15a, XJZ19, YWL18, YY17, YX19b, YmZHC15, YSZ19, Yüz15, ZT16a, ZSK^{+}19, ZLW15, ZLL15c, ZS15a, ZZZ15b, ZS15c, ZLZ16, Zha16c, ZLH16, Zha16d, ZWaY18, ZW19b, ZWY19, ZCX19, ZZL15a, ZZH15, ZCQ15, ZHX^{+}17, ZMT15, ZXL15, ZPT^{+}15, dSVR17, KI15]. Systems [HL17, LZLW17, YSWZ17, ABC^{+}16, AK16a, ASHSC17, ASASM16, AHP19, ADMV19b, AIG17, AL15b, AM19, ACDT18, AK15b, Ant18, AND17, AM18c, ACC16, AKA19, AZAM15, BCST19, BLK19, BGRT15, BAE15, BR16a, BDSS18, BR16b, BKW18, BK18d, CT17a, CÁGGLP16, CGCK19, CEMR16, CR15, CS17, CGW19, CFSGM17, CGM16, CTT^{+}19, CXP16, CLXW17, CWHY15, CC15, CXL16a, CCQS16, CXL16b, CYSY16, Che16c, CL16c, CD16a, CGZ17, CLCW17, CPCQ17, CZM18, CXZZ18, CLZW19, CCL^{+}16, CGAL19, CD15d, CB19, CL19c, CN19b, CVMGPPGC15, CXW^{+}19, CSL19, DHX^{+}19, DG15a, DQH18, DHQ19, DB18, DTM19, DR19, DH15, DB16, DBV18, DKBS15, DHS19, DKKR16, DL17c, DLXZ19, DH17, DOZ18, DZC15, Du17a, DCJ19, DG15b, EFPT18, Ely15, Ely18, EKO19, EHdLS16, EE18, FP18a, Fan15a, FW18a, FZL19, FMDZ19, FLL15, FXXC19, Fer15, FG19, Fio16]. systems [FNS15, FY19, FLYG18, FZG19, GRV15, GDM16, GWW15, GW15a, GSZ15, GWSC16, GDL16, GHQ17, GZC19, GMOGCCC15, GSPH19, GLN15, GWPA18, GSD16, GWXZ18, GYC19, GSVK17, HPR18, HMD19, HRS15, HV15, HR18, HWY15a, HW17b, HT16, How16, HZK^{+}19, HMT15b, IB15, IKMW15, JLX15a, JGX15, JS16, JGKL16, JZ19b, JAN16, KLKU18, Kim15b, Kim19, KLS15, KML19, KEE17, KPP17, KKL18, LPKS16, LLP17, LK15, LLCR15, LY15d, LLZ15, LL15b, LSAR15, LY15b, LG15, LY16a, LCY16, LX16b, LSZY17, Li17d, LW17a, LLC18, LZjy18, LBRS18, LSMT18, LSLZ18, LZ18c, LQC^{+}18, LLZD18b, LM19, LSZ^{+}19b, LGCS19, LPS19, LSL19a, LZ17c, LXC^{+}19, LLLZ16, LLZ17, LC15d, LGK15, LZL15b, LS16, LGL16, LHS18, LXW^{+}18, LNWS18, LZHC18, LS19c, LXZ19b, LMZ^{+}19, LHZM19, LLX^{+}19, LGL19c, LJ19, LLdM15, LVC15, LWZZ18, LZH^{+}18, LJHW19, LSL19b, LJS17, LGHW19, LTZ^{+}17, Lv16, Lv18, MFJZ15, MM15b, MC15a, MD15a, MJL16, Ma17b, MWRT17]. systems [MWY17, MLSX18, Maa15, MA16, MAR15, MB15a, MRQZ19, MH17c, MGMM17, Nak18, NBK16, NT16, NPT19, NHWZ15, NSM16, OLZ19, OB18, Pai15a, PY19a, PY19b, PR15a, PST18, Pla15, Pop17, QG16, QZCJ19, QTF^{+}16, Rad17, RN18, RZHY16, RWW18, RXLM19, RXQ19, RRV15, RB15, SRSZ15, SSKA16, SJWK18, SM17, SZ19b, SMZ15, SSCX18, SSK15, SSB16, Sha15f, SYZP16, SY17a, SLLC19, SMG19, SLZ^{+}16, SY18a, SFSR18, SML^{+}15, SPGP18, SGKR19, SR19, SSS15, SZPH16, SDP16, SZ17, SMZ^{+}17, SWSW19, SJL16, SGY^{+}15, SZ16, SWZL19, TFLZ19, TPL16, TZ19a, Tan17a, TNMR18, Tha15, TN19, TCS15, TPCL15, TB17b, TNPL15, TTV15, TTXZ16, USCA15, USCAAA15, Val15, VMLO19, WZAF15, WDLL15, WY15a, WZYS15, Wan15b, WZZL16, WZ16a, Wan16a, WWZD16, WS16, WMZ16, WDZ17, WHTW17, WLZY17, WCZY17, WH18, WLH^{+}18, WXX18, WDLS18, WXWX18, WS19a, Wan19a, WDG19a, WDLZ19]. systems [WLZ19a, WJZF19, WL19a, WL19b, WCB19, WXXY19, bWzZjH16, WWHZ18, WSYL18, WNW^{+}19, XZLZ17, XD19, XZL15, XY15a, XY17, XY18, XPZ18, XCDZ19, XKWG15, XXW^{+}19, XCY18, XCCL18, XH19a, XZHX17, XCL18, XSSW19, XDH19, XGQW19, XZLL17, YW15, YSL18, YPZ18, YKJY17, YLHS17, YD18, YLCW18, YWZS18, YZ19, YL19a, YWS^{+}19, YCX^{+}19, YLP19, YUK19, YDLK15, YLW19, YZ16, YJMH18, ZSS15, ZLHH15, ZZLZ17, ZLW19, ZLDZ16, ZLLZ16, ZLDZ17, ZZL17, ZAY18, ZMC19a, Zha15a, ZW15b, ZZC16, ZZL16a, ZSS16, ZD17, ZLM^{+}18, ZXS^{+}18, Zha19b, ZWZ19a, ZCAN19, ZX19, ZCW19a, ZZMS19, ZG15a, ZJ15b, ZFZK15, ZZ16b, ZYLY17, ZJYD17, ZWO^{+}19, ZMC^{+}19b, ZM15c, ZM18b, ZPM16, ZYWW16, ZSL^{+}18, ZZL15b, ZXX^{+}19, dSH17, fP15]. Szász [MMS17b, SV15b, Aca15, DPD18, GR15c, GG15b, KA15a, KM18a, MS16b, MNK19, KTA15, MDG18]. Szeged [BKL^{+}17, CT17b, HKS19, JLW18, Wan17b, WLQZ18]. Szego [CMR17]. Szokefalvi [GDF^{+}17].

T [PK18a, CFSGM17, DH17, GSPH19, KSM^{+}19, KRA^{+}18b, LSLZ18, MK15b, PRMC19, WJHA17, Zha19b]. T-cell [PRMC19, WJHA17]. T-cells [MK15b]. T-point [CFSGM17]. T-S [DH17, GSPH19, KSM^{+}19, Zha19b]. T.E.A.M. [KKKD18]. tableaux [WC15a]. tables [HLFZ19]. tactics [BAGN^{+}19]. tag [ZY16]. tag-mediated [ZY16]. Tail [JPY16]. tailored [YW19b]. Takagi [CGAL19, LZ18c]. Takens [LWZ^{+}19b]. Takeover [BLA15]. Takes [TKDF15]. Taking [RPJ^{+}18, SWSS18]. tamed [TY18]. tangent [Qi15, Sah19a]. tanh [ND15]. tanh-function [ND15]. Tanimoto [DV15]. Tank [CCL^{+}16, UMF19]. Tanner [CY18]. taper [FVMO15]. tapered [PPS25, RRR19, RRR20, SM15c]. tardiness [XYZK15]. tardy [WYW16]. target [EA15a, SCS^{+}15, WRZ19]. targeted [ZSW^{+}17]. task [Świ15]. Tasso [S.15]. Tau [AD16, EE18, SLC18]. Tavakoli [WL18a]. Taylor [Sab18, ABMBR15, AP19, BS16, Ben18, DQW^{+}19, EAAAS15, GDS16, GIS15, HMT15a, KK17, MM04, Mil16, Muk16, Par15]. Taylor-series [Par15]. TB [SSGG17]. Tchebychef [Bor15b]. TCP [PW19b]. TCP-RED [PW19b]. TD [JLL16]. Teaching [bOqGyK^{+}15, ZL19c]. Teaching-learning [bOqGyK^{+}15]. technique [AOH17, AL18d, CST^{+}15, DQW^{+}19, GZK16, GLA^{+}17, IH15, JM15, KMBA18, KD18, LL15c, LHLL19, LXZ19b, MC15c, NWNA15, NWN15, PAE^{+}17, PZ16a, PeO18, PKB18, RG18b, SM16, SN15a, SN15b, SH16, SZO16, SN18a, SN18b, SHEBS19, SCN15, TPCL15, YMS16, ZZ18c]. Techniques [SCR^{+}15, AHO18, ABMA19, BCNN19, HS15b, KKKD18, Kim19, LK18, MKN^{+}15a, MA15c, Oli17, OA18, Pel17, RC17b, uIHA18]. technology [AM19, KK15b, TKDF15, TD18, Zho17]. telegraph [MZM16, PRD19, SR16b, Yüz16b]. telegraphic [MD17]. Teller [RZMB17]. temperature [DWT17, HRS17, JDL15, Liu17b, PG15a, SFC^{+}17, SCM19, Tha18b, dFKMM15]. temperature-dependent [Liu17b]. tempered [ZLD18]. templates [BDBKV19]. temples [PG15a]. temporal [AKL19, CVJ19, ZHW16, ZYLY19]. temptations [BKAAM15]. ten [Xu19a]. tens [MHCF15]. tensile [MWY15]. tension [GLA^{+}17, SuSB15]. tensor [DPSS16, HW19, LB15a, MH19b, NKP16, RVE19, Van15, WC17, YLLH19]. tensor-based [RVE19]. tensorial [TG19]. tensors [CLQ16, LLL19a, WDB18, YLLH19, YKZ15]. tenth [RS17a]. Term [SM16, ARL15, Ali15, APL15, BD17, BDGS15, BB15d, CLY17, CZP15, DLH15, DX17, GM19, HQ15, HAH19, KCWK15, LLY15, LC16a, Li17b, LM15, NR16, RZW18, SZL^{+}18, SM18, Tan15d, TY16, TDL^{+}17, WCHW19, Xu17a, ZZ15a, ZLLA19, ZZL18b, ZLL18]. terminal [ZCAN19, ZYLY17]. termination [Zha15i]. terms [AY16a, Bot15, CXW18, DMG15, KAI19, KKT19, KNT17, LRF^{+}18, LCAA17, Ma17a, QM15b, RMSD16, ZSS15, Zha16c]. ternary [ACRZ18, Par18a, RS15c, RS16a, SuSB15]. terrorism [BAGN^{+}19]. test [BCM17, Che19, FFH^{+}18]. testing [CLR15, SFC^{+}17, WZM15]. Tetravalent [CC18c, FHK^{+}19, LXCK17]. text [JM15]. textiles [YX15]. textural [TM18]. texture [LAVME18]. TFETI [SHHC18]. th [XHW15, ZWK16]. th-order [ZWK16]. their [AK18a, AADMM16, BD15, BCMT15b, BCMT15a, BR16a, BMRW19, BM19d, DL17a, DOR18, DHS19, DU18, GLL15a, GKN15a, GKM16, GKN16, GKS16, GS18a, GR19a, Gut17, HARS15, HZL18, HL18, JO15, JRSK16, Jar15, LW19c, MH19b, OB18, PS16, RSGFV16, RG19, SA16d, Sim15, SX15, TTD17, WGL18, XLS18, YZ18b, YWL19, ZS16a]. them [AM15c, RP18, TW15b]. theorem [AK15b, AM15h, Boy15, BCFR18, BZ15, CCR18, CZ15c, GRV15, GK15, GS11, LC15a, MY19, Pro16, Reh18, Sil16, TA16, WYL19a, YQD17, Yua19, ZS15d]. theorems [ACS15, BLS15, CC15, Ely15, KBS17, Kar19, KCKZ15, Kur18, LR19a, LZCW19, MDG18, SMZ15, SPKC16, Sug17]. theoretic [DH17, GWXZ18, WH15]. Theoretical [HHT19, Kim15a, Oka18b, SK17a, YYM19, ZYZW16, GLR^{+}18, GG15c, KM15a]. Theory [BKS18, CPS18, JDS18, WE19, AK16a, Alf17, AR15, AB18, BLA15, CCLS16, CLSS19, DJW19, GDM16, GDF^{+}17, Han19, HP17, HHY19, JDL15, JG17, KEE17, LF15, LpLxG18, Özb15, SSA19, TWF^{+}18, WWC^{+}18, ZZLY15, ZCX17, ZL19a, ZL19b]. therapeutic [Kun16]. therapy [MK15b, XGZ17]. thermal [ABB^{+}18, BES19, CDFP17, DLS19, KBK^{+}16, KK15b, Liu17b, MR16a, SRP17, SMR15, SMG^{+}18, Ste19, SCM19, WQ15, WYD^{+}18, YX15, YJW16, ZFL18]. thermally [AKR15, PK15]. Thermistor [SY18b, SY19]. Thermo [Rad15, SM15c, Ton18]. Thermo-elastic [Rad15, SM15c, Ton18]. Thermoelastic [AKR15, DDY15, DL15a, FM17a, FM17b, HW18, OTH16, SH18a]. thermoelasticity [FM17a, Pov15]. thermographic [ABB^{+}18]. thermoviscoelastic [ARSV15]. theta [DT16, LZH19, LMD19, RLZ19, TY18, Yue19]. thick [JA17, LZL15a, PBP19, ZLL15b]. thickness [AwEE15, DDY15]. thin [AL18a, AAMC19, BTH18, RDM19, WL16a, ZZ18b]. Third [LZH16, Maa15, RR18a, AKS15, AH15, Awa15, BGST16, CS16a, CP19, CN15c, DB17, DJ18, FFT15, GPHAPRW16, GJZ19, IAW18, JISI18, NKS18, SST17, Sak19, SM18, Ugu19, WZ15b, ZX16]. Third- [LZH16]. third-order [BGST16, DB17, DJ18, FFT15, GJZ19, Ugu19]. Three [AKT19, CZ15b, KKG18, MB17, ALMCDH^{+}18, ADG16, AG15c, ATX19, AADMM16, BBGR15, CC18a, DKP16, DCS15, DDdF16, Dim17, DLH15, ESS16, ESES16, EA15b, FFN18, GG16, GKN16, GPHAPRW16, GZG^{+}15, GYC19, HSSH18, JFT17, KL15, LWK15, Li18, LSL19a, LL19c, LA15, MWY17, MRE18, MH15a, MD17, MKS19, Moh19, Par18b, PZ15, RS15d, SK15, SdA18, SBV15, SJ18, SPGP18, TPS15, TST16, VBG^{+}17, WLLK17, YL19a, ZWBC18, ZXL15, ZGZ16, ZPT^{+}15, YJYH15]. three- [LA15]. three-body [SPGP18]. three-component [YL19a]. three-derivative [Li18]. Three-dimensional [KKG18, MB17, AADMM16, BBGR15, FFN18, GZG^{+}15, GYC19, MRE18, MH15a, MD17, SK15, SBV15, SJ18, TPS15, VBG^{+}17, ZGZ16, ZPT^{+}15]. three-dimensions [SdA18]. three-level [MKS19]. three-neuron [ZWBC18]. three-order [LWK15, WLLK17]. three-player [ESS16, ESES16]. three-point [GKN16, PZ15, TST16]. three-point-based [ALMCDH^{+}18]. three-species [Moh19]. three-stage [GPHAPRW16]. three-step [EA15b]. Three-strategy [AKT19]. three-term [DLH15]. three-time-scale [DKP16]. three-to-one [RS15d]. Threshold [LXT17, LJS18, QTTL19, WGLZ16, ZWD15, ZYZ16b, ASZD18, DM16]. thresholding [ZYH15]. thresholds [CCYT19]. thrown [How15a]. tidy [BKKBS19]. TIFIM [HWL^{+}19a]. Tight [BCK15, GL18a]. Tikhonov [KRS18, RR15, Red19, XX19b]. Time [AHK^{+}19, BLA15, GR19b, GMPP18, LdCM19, LLTL18, LZLW17, LFY18, LYY15, Pov15, SH15b, UMF19, WXXY19, ZL17b, AKAA15, AS19a, AAF15, AR16a, AJ17, ABS15, AAA^{+}15a, ADY16, ARL15, ASA19, AADLF15, AG15a, AY16a, AO19, AL18d, AK18b, AP15, BPC15, BAB18, BP17a, BD18, BS17, BNY15, BK18c, BKW18, BBFJ17, BRM^{+}19, BES19, BJ19, BS19b, BVJ15, BL15, BL18, CÁGGLP16, CWG15, CGCK19, CL19a, CM18, Cha19, CJ18, CHL15, CCQS16, CYSY16, CWL16, CD16a, CGZ17, CZSH17, CLCW17, CZL18, CCL^{+}16, CDL18, CL19c, CLT^{+}15, CVA18, CP15d, CXW^{+}19, Dan18, DS16a, DKP16, DAQI15, DWZ15, DL18b, DSB15, DKZ16, DLXZ19, Dol15, DHMT15, DOZ18, Du17a, DSSC19, DZ17a, DF16, ET15, Fan15b, FMDZ19, FDPP18, FLL15, FYZ16, FXXC19, FGPR18, FM16, GS17a, GWW15, GSZ15, GW15b, GHQ17, Gao18, GHW19, Gat18a]. time [GSPH19, Gol19, GPHAPRW16, GST15b, GLL15b, GZ15a, GZLL17, GWXZ18, GZD19, GSVK17, HO15, HFCS16, HW18, HH18a, HGH19a, HLIK19, HWY15a, HWY15b, HW17b, HK18b, HAH19, HSKHA15, HS15a, HT16, HS15b, HXW^{+}19a, HYW^{+}15, HXW^{+}19b, HLH17, HHG19, HL15d, IB15, Ied15, Ino16, JHWZ15, JGX15, JGKL16, JSW^{+}18, JX19, JAN16, KCZK15, Kan15, Kar19, KM17a, KN19, KI15, Kim15b, KSN15, KKS^{+}16, KK16b, Kri19b, KRA^{+}18b, KKL18, LLP17, LGO17, LY15d, LJ15a, LLY15, LSCK15, LCY16, LC16a, LX16b, LSZY17, LL17b, LCAA17, LW17a, LZCW18, LW18d, LSZ^{+}18, LZ18a, LSMT18, LW18g, LSLZ18, LZ18c, LQC^{+}18, LLZD18b, LX19b, LWZ^{+}19b, LSZ^{+}19b, LLZ^{+}19, LXD19, LSL19a, LJL19, LLLZ16, LLZ17, LT15, LLZ^{+}16, LY16b, LZG16, LGL16, LYZL18, LLL18c, LXW^{+}18, LMW18, LZHC18, LXZ19b, LL19d, LS19b, LFC^{+}19, LGL19c, LWZZ18, LJHW19, LSL19b, LM15, LHWG16, Luo19, MFJZ15, MC15a, MTY16, MJL16, MLSX18, MRC^{+}18, MAR15]. time [MLG18, MLS19, Mil16, MDH15, MKS19, MS15d, Mon19, NR16, Nag17, Nak18, NT16, NLC19, NHWZ15, Pac19, Pai15a, Pai15b, PSM17, PY19b, PSSX17, PJ15, Prz15, QG16, QKGW18, QBC19, QZCJ19, QFS17, QLF15, QW19, RSC^{+}17, RP19, RMB15, RG15, Reh18, RSLL^{+}19, RZHY16, RW17, RLM19, Rog19, RZW18, Rub19, Rui18, SJWK18, SM15b, SSS17, SZ19b, SSZ^{+}18, Sha16a, SSCX18, SYZP16, SLLC19, SzSC19, SLZ^{+}16, SCC17, SZL^{+}18, SY18b, SFSR18, SY19, Sku15, SDP16, SZ17, SMZ^{+}17, SZFS17, SJL16, SzShG16, SLS^{+}18, SWZL19, TTZ^{+}19, TXG15, TNPL15, TDL^{+}17, Upa16, WDLL15, WZYS15, WPL15, WA15a, WLS15, WZZL16, WSZ16, WMZ16, WMLG17, WDZ17, WYCZ17, WLZY17, WXX18, WHX18, WM18b, WDLS18, WXW18, WDG19a, WRX^{+}19, WDLZ19, WR19, WJZF19, WCHW19, WCH15, WZY^{+}19, WWC^{+}15, Wu17, WSYL18, WV18, WNW^{+}19, XH16, XZ18, XPZ18, XYG^{+}18, XXW^{+}19, XCCL18, XX19b, XWZ15]. time [XZHX17, XCL18, XLL^{+}19, Xu19e, XXZY19, YSL18, YWL15, YLCW18, YSX15, YDLK15, YS16, YSF16, YLW19, Yu16, YZ16, YWZ18, ZNV19, ZEF19, ZZZ15a, ZLHH15, ZLW19, ZLLZ16, ZLDZ17, ZDPZ15, ZJZ15, ZWHC15, ZZC16, ZWK16, ZS16c, ZZL16a, ZSS16, ZHJ^{+}17, ZYXF17, ZHWW17, ZL17a, ZZY18, ZYC18, ZZ18a, ZLC^{+}18, ZXS^{+}18, ZW18, ZWBC18, ZJY18, ZW19b, ZLLA19, ZCAN19, ZX19, ZCW19a, ZCCZ19, ZCX19, ZJ15b, ZFZK15, ZY15c, ZYLY17, ZZL18b, ZZQ18, ZWO^{+}19, ZPM16, ZX16, ZSL^{+}18, ZXX^{+}19, ZT16c, dFGAN19]. time-consistent [PSSX17]. time-delay [CD16a, CGZ17, DAQI15, FM16, GWW15, GSZ15, HT16, LSL19a, LLZ17, LXW^{+}18, LMW18, MFJZ15, MC15a, NT16, NHWZ15, Pai15a, Pai15b, QKGW18, RLM19, TTZ^{+}19, TNPL15, WDZ17, ZPM16]. time-delayed [ADY16, QG16, WM18b, ZYXF17]. time-delays [MAR15]. Time-dependent [LYY15, ABS15, CP15d, GS17a, GW15b, HLIK19, HS15b, HLH17, HL15d, JX19, KCZK15, LW18g, LS19b, LFC^{+}19, Mil16, SY19, Wu17, ZL17a, ZT16c, dFGAN19]. time-domain [CDL18]. time-efficient [AS19a]. Time-fractional [LLTL18, LFY18, Pov15, AKAA15, ARL15, ASA19, BS17, HH18a, HHG19, LLY15, LM15, PJ15, RG15, RW17, RZW18, Rui18, SzSC19, SY18b, WXW18, WR19, XX19b, YWZ18, ZEF19, ZLLA19, ZX16]. time-frequency [BL18]. time-harmonic [KM17a]. Time-invariant [GMPP18]. time-irregular [Prz15]. Time-limited [GR19b]. time-marching [WSZ16]. time-proportional [Yu16]. time-space [CZL18, ZLC^{+}18, ZJY18, ZZL18b]. time-splitting [AP15, CWG15, QFS17, Rub19, WMLG17]. time-step [LGO17]. time-stepping [BS19b, Xu19e]. time-variable [LZ18a]. time-variant [LSZ^{+}19b]. Time-varying [WXXY19, BPC15, CM18, CZSH17, DLXZ19, FLL15, FYZ16, GHW19, GSPH19, GMPP18, GZ15a, GWXZ18, GZD19, GSVK17, HW18, JHWZ15, JGX15, JGKL16, JSW^{+}18, Kim15b, KKL18, LLP17, LY15d, LJ15a, LC16a, LSZY17, LCAA17, LSZ^{+}18, LSMT18, LSLZ18, LZ18c, LQC^{+}18, LGL16, LWZZ18, LJHW19, MJL16, MRC^{+}18, NR16, NLC19, PY19b, RSC^{+}17, RZHY16, SM15b, SSZ^{+}18, Sha16a, SCC17, SLS^{+}18, SWZL19, TDL^{+}17, WLS15, WYCZ17, WLZY17, WCHW19, XCCL18, YLCW18, ZLW19, ZDPZ15, ZWHC15, ZZC16, ZHJ^{+}17, ZYXF17, ZHWW17, ZZ18a, ZW19b, ZCAN19, ZCX19, ZXX^{+}19]. time-varying-delay [SLZ^{+}16]. Time/Cost [SH15b]. timed [Val18]. timed-arc [Val18]. times [Abd15, Her19, PGLG17, SJ19a, WGL18, XY15b, Xu15b]. timestep [CL19a]. timestepping [LS17b]. tipping [FHPV18]. tissue [Kun16, PBMP15]. tobacco [PZLZ15]. Toda [BP17b]. Toeplitz [WLZ19b]. tolerance [BF15, GSLC18, LZX19, Xu15a]. tolerant [DH17, HYL19a, LX17, MLSX18, RHP18, SJWK18, WLH^{+}18, YYS17, YLW19]. tomography [ASKK19, BPZB18]. tool [BBFJ17, Han16]. tools [PH17]. topography [PMC16, SSG^{+}15a, ZYLY19]. Topological [Alf17, BCD18, SX15, BHMS15, CXL16a, CWD16, DGNA15, DV15, DESS15, FPUK15, GK16a, II17, LLMS16, LQW^{+}15, LP15b, NZZ16, PP19, Rad17, WCD17]. topologies [Sha16a, SSCX18, SWZ^{+}19, SLWX15, WRX^{+}19]. Topology [DH19b, CLZW19, HLJ18, JZ15b, LDS^{+}18, LWWL19, SA18a, Wan19a, YYS17]. torque [Dan15b]. Torrey [SzShG16]. Torvik [KCK16, ZV17]. Total [DH19a, DHW^{+}19, MNJW19, ZZHT17, ASV17, KGGAV^{+}18, CCX15, CLW18, DSZY19, Gat18a, HLL18, JK15, KM15d, Li17f, LLW17, LGL19a, LJL16, MTY16, MCY19, SHLW16, SJL17, TLLD19, WX19b, XLG18, XGG18, YSY^{+}17]. Touchard [CM15b]. tournament [CGPT15]. tournaments [BGMO18, MY19]. toxic [fLLPY16]. toxicant [LC15d]. toxicity [PSM17]. toxin [JA15, YYZ19]. toxin-producing [YYZ19]. Traceability [TX18]. Traces [tHBSfY19]. Tracking [ZZMS19, CL19c, CXW^{+}19, FXXC19, FNS15, HWY15b, LY15d, PY19a, SA16a, TFLZ19, WZAF15, WWHZ18, ZLWS15, ZLM^{+}18, ZXS^{+}18, ZYLY17, ZWO^{+}19, ZYZ^{+}16a]. tractions [JLL16]. Tractrix [Han16]. Tractrix-based [Han16]. Trade [SH15b, CLT^{+}15, JYVS15, MBM16, ZWY19].

Trade-off [SH15b]. tradeoff [BS15a]. trading [LWZ19a]. traditions [BDBKV19]. traffic [Chi15, GGAVGGRC15, LZY16, RHJ18, YS16]. train [HRLGÁ18]. Trained [Świ15, ŚKK^{+}18]. training [LP15c, PAE^{+}17]. Trajectories [ALMCDH^{+}18]. trajectory [MND^{+}15, ZLWS15]. tranches [WL16d]. transaction [LW15a, NM15]. transboundary [LWZ19a]. Transcendental [HA19]. Transfer [ZS16b, AwEE15, AB15b, ATK^{+}16, BM19d, CKG^{+}16, DJL17, DBV18, EAY17, EM15, GSP15, HLC^{+}17, HMHA15, HCY15, LRL18, PMV16, Pan15, RZMB17, Sha12, SV17a, Ste19, SE18, WLZ15, YM18, YWYS18, YX15, Zad16, ZjW19]. transform [ALQ15, Awa15, BDE^{+}15, CH18a, CCJ15, DTM19, De 17, FBF18, GT19, IS16, JJM18, JLL15a, LXZ19b, Mah18a, MSH15, Pos17, PLL16, PSP17, PKB18, SD19c, yWyLyY^{+}15, XS15, YMS16, YJSW16]. Transformation [BC15, RRW^{+}15, CCNQ19, CMR17, KPT16, KMT16, LXZ19c, MGL19, SB15, Sug17, WTSQ15, Xu15c, XLW15]. transformations [BD15, BdlRT18, DGN16, Ely18, HZL18, PS18b, RS17b, RM15]. transformed [GFE19, Ma15a]. transformer [RCRL18]. transforming [AADMM16]. transforms [CM17a, CXK19, DÖA15, GIT15, Kat15, KSN15, NCEGA15]. Transient [Amm15, KBK^{+}16, LLC^{+}15, ZZLY15, Bot15, CGW16, DGN16, JCW18, KSL19, LL18c, SC17, YPZ18, dFGAN18]. Transition [WW16c, BRM^{+}19, Ino16, JGKL16, KPP17, LSZY17, LZHC18, QG16, SY17a, SH15d, VPS15, XKG15, XGQW19, YSL18, YKJY17, ZKKZ18, ZDF17]. transitional [SSM^{+}18]. transitions [HYZ^{+}18, LLWW19]. transitive [CMT^{+}18, CC18c, DL18a, FHK^{+}19, HLL18, LLL18a, TZ18b, XX19a, YFH19]. translates [DMH19]. translation [BGSCGL19]. Transmission [KJR^{+}18, PRL^{+}15, AYD19, ADY16, ATN^{+}19, AVCCGAVDL15, Dob19, FG15, HYB17, IYM19, KIZ18, KCL19, LZL^{+}19b, NKKS16, PC17, RL19, SSGG17, SG17, WGLZ16, WWDW17, WS19b, WWSM19, XZ15, YG15a, YRA15, YJY16]. transmissions [CYLW15, LXT17]. transmitted [IYM19, WCH19]. transmutation [KKT19, KMT16, Kra18a]. transonic [DMSS18, HHP18]. Transparent [EKS^{+}15]. transport [BS19a, BES19, KSKW^{+}19, KK15c, MSM15, SBV15, UHLM15]. transpose [HM18b]. transversal [MKRS16]. transverse [DS15]. transversely [ZLL16]. Trapezoid [CCLPC19]. trapezoidal [DD19, QXW19, Rez15]. trapped [CLY17, Mis15]. Trasona [NFS^{+}15]. Traub [SSK15, SSB16, VLCT18]. travel [Chi15]. Traveling [BW15b, CY18, XYDL16, AM18a, Alk19, Esl16, GO15b, Kud19b, MWY17, MKRS16, Sun15b, UKK17, WL16a, WC16, WZZT19, YMHW15, ZH15a, ZXZB19]. travelling [HMSS19, STJS16, STL17]. travelling-wave [STL17]. treat [FLMC15]. treat-age [FLMC15]. treatment [BAK19, CC17, JLL16, NKS18, STA^{+}18, SG19b, WLG17, ZP16]. treatments [DDD19, HJ18]. Tree [YQD17, CXL16a, KK16b, LWM18, QGH19, RCTZ18, RGH^{+}15, ZY18a, CLHC16]. tree-like [CXL16a]. Trees [CYCW15, KST19, ADA18, AG17, ATX19, BM16, BF16b, CCLW18, CGJZ18, CZ17c, DV15, DDdF16, Dim17, Gou18, Gut16, GFD16, HLL16, LLWX19, LL15a, LY18b, LWW19, LCM19, LCW^{+}18, LWSM16, LYW18, MB19, RC16, SG19a, SGDX18, WWH16, WWL18, WX19b, WYL^{+}19b, WZC17, WZ19, Yan17, YwFW^{+}19, ZWZ19b, ZC19c]. treewidth [WTZL18]. Trefftz [EKS^{+}15]. Trend [MTCV^{+}16]. Trends [BKS18]. tri [DP15a, UKS^{+}16]. tri-parametric [UKS^{+}16]. tri-population [DP15a]. trial [Liu18a, OB18]. triangle [BRV18, Cat17, HL18, LZWZ19, OLZ19, WGL18]. triangle-based [BRV18]. triangles [BNS16]. triangular [AP18a, CL19b, DK15b, FZLZ19, KJR^{+}18, LW15c, LML18a, LZ17e, LY18d, LJ15b, Ort15a, PP12, PP15b, RC17a, ZLL15a, ZL16a, ZH15b]. triangulations [CR18, Cih15, GK16b, XZC16b]. triaxial [SPGP18]. Tribonacci [Flo18, JH15, RT18b]. Tricobsthal [RT18b]. tricyclic [JQL15]. tridiagonal [dF15, dFY15]. trigamma [QM15b]. triggered [CL19c, CXW^{+}19, GSD16, LLC18, MLSX18, SRCA17, SMZ^{+}17, SWZL19, TFLZ19, WS19a, WRX^{+}19, WXXY19, XD19, XPZ18, XZHX17, YLP19, ZZZY17, ZDF17]. triggered/self [WS19a]. Trigonometric [AH18a, GW15b, Han15a, Han15b, JRM15, LQZ^{+}18, NMM16, SR15d, SHH15, TYD15, YAIN17, Zha15d, ZL18]. Trigonometrically [LD18a, Li18]. trinomial [BD17, NP18]. triparametric [LKM17]. tripartite [PJ18]. triple [LZWZ19, WHSS15, XZZ^{+}18, XJZ19]. triple-S [LZWZ19]. triple-triangle [LZWZ19]. triple-well [XJZ19]. Triply [LG17]. Tripoly [Ask18]. trivariate [SM19]. truck [LZY16]. truncated [Dan18, LYZ19, ZI18, ZGGY19, ZYSL19]. trust [HHR17, Lu19, PZ16a]. trust-region [PZ16a]. tube [Pon18, SSM17]. tubes [HP17]. tumor [DDD19, DHMT15, HJ18, KN19, LX19b, LZ19a, QFTT15, SRJ^{+}15, YJ19, ZK18]. tumor-CD [HJ18]. tumor-effector-interleukin- [DDD19]. tumor-immune [KN19, YJ19]. tumour [RLM19]. tumour-immune [RLM19]. tuning [KB18]. tunnel [FFH^{+}18]. Turán [LLST19, Yua19]. turbidostat [YLL15]. turbine [FFH^{+}15]. turbulence [AKL19, HYL^{+}19b, SH15d]. Turbulent [VPS15, Ili19a, MKRS16, MKMC15, MS18b, XP17]. Turing [LLY19, PZ16b]. turning [CVA18]. Tutte [LAAZH19]. TV [CCC^{+}17, CC18a, Laz17]. TV-based [CC18a]. TVD [SdA18, YNMW16]. twelve [GYC19]. twisted [GSLC18, JRSK16, Lai15, LHG18]. Twitter [LdCC15]. Two [ABHR15, AKL19, APPE17, BSZ16, CR15, CPST16, ESES16, FKR15, GLW18, GN19a, HWL^{+}19a, HS18, HSS15, HSSH18, JHL19, LW15c, MLW18, PBBF18, QDD15, Qiu18, SIA^{+}17, Suk19, WSZ18, YMS16, ZW15b, Zha18c, ABD16, AT15a, Akt15, ACE19b, Alk19, APS16, APL15, ALZ16, AHY19, AO19, ARR16, AND17, AM15g, ARGF15, AP18b, AE17a, Ask18, AD19, AAMC19, AADF18, BM16, BN15, BB18b, BS15b, BRD15, BKL^{+}18, Bou18, BZ16, BTWY15, CFG15, CCL^{+}15, CFS18, CGC19, CTT^{+}19, CKG19, CLJ^{+}15, CL15b, CXMP16, CGW16, CW17, CXZZ18, CX19, CDL19, CLW18, CL19c, DLBQ19, Dan15a, DS15, Dan15b, DA17, Den15, DCS15, DL18b, DNC^{+}17, FZLT15, FvdMS15, FG15, FR16, Gar16a, GKN15a, GCLG15, GOA19, GAT18b, GYW15, GLW15, GS17c, HRLGÁ18, HTW19, HZ17b, HXM18, HDG15, HWL^{+}19b, HHL^{+}19, JSM16, JJ15a, cJzS15, JH17, JYSR18]. two [JY19, KIAG16, KLQY18, KE16, KMBA18, KZA19, KT16, Kok15, KCKZ15, KMB16, KSV15, KSZ15, KMP15, Laa18, LWZ19a, LC15b, LG15, LW16, LR16b, LBRS18, LLL^{+}18b, LJL18, LL18a, LW19b, LZ15c, LB16a, LXL17, LYW18, LW18h, LMZZ19, LSL19b, DPL18a, LJC16, MWWM15, MZM16, MMZ^{+}17, MH19a, MBM16, MS15a, MVS19, MLT17, MMV16, Mic15b, MZS18, MH15b, MS19c, Mis16, MGMM17, MDH15, MMK17, MKS19, MN15, MPB^{+}17, NE16, NBK16, NDZ15, PSM17, PAJ^{+}19, Pel17, PÖB15, PD15, PeO18, QDH^{+}18, RBB19, RS17a, RN19, RS15a, RC17a, RDG18, RT18a, RSO^{+}18, SM19, SdA18, SHR18, SSN15, SZ15, SH16, SG19a, Sha15d, Shi15a, SY17b, SJ18, SI19, SSZL17, Tha18a, TH19, WYW16, Wan17a, WDLL15, WPL15, WL16a, WEH18, Wan19b, WXM^{+}18, WYD^{+}18, WZY^{+}19, WZZT19, WL17, WLL16, WLFH17, XYG^{+}18, XW19, XCHW16, XCCL18, Xu19d, YÖ18]. two [YX19b, YX19a, YY15, YLZT17, Yüz18, ZV17, ZB17, Zei16, ZSK^{+}19, ZXW16, ZJ15a, ZLWS15, ZZY^{+}17, ZF18, ZH19a, ZM19, ZBH^{+}15, ZHX^{+}17, ZYX19, ZG15b, ZGZ16, ZWM16, ZC18, ZLW18, ZZC^{+}19b, ZZL15b, Zhu18, ZT16c, dSVR17]. two- [MKS19, SdA18, ZGZ16]. two-asset [CW17]. two-by-two [Den15, ZZL15b]. two-channel [CL19c]. two-component [ACE19b, GYW15, GLW15, HXM18, WZZT19, YX19b]. two-determinant [APS16]. Two-dimensional [HS18, YMS16, AHY19, AND17, AE17a, AD19, AAMC19, BN15, BTWY15, CGW16, CDL19, DA17, DL18b, GS17c, HZ17b, cJzS15, KE16, Kok15, LW19b, LW18h, LJC16, MWWM15, MZM16, MH19a, MMV16, MH15b, MS19c, Mis16, MKS19, NE16, PeO18, QDH^{+}18, RDG18, SM19, Sha15d, WDLL15, WYD^{+}18, XYG^{+}18, XW19, Yüz18, ZB17, ZF18, ZBH^{+}15, ZHX^{+}17, ZG15b, ZLW18, ZT16c]. Two-distance-primitive [JHL19]. two-fluid [Laa18]. Two-grid [Qiu18, LR16b, SH16, SY17b, YX19a]. two-layer [BKL^{+}18, DS15, HHL^{+}19, WL16a, XCHW16]. two-level [ALZ16, HDG15, KMP15, MBM16, MVS19, MMK17]. two-link [ZLWS15]. Two-parameter [CPST16, BZ16, NBK16, ZG15b]. two-patch [Wan19b]. two-person [MN15]. two-phase [BB18b, GAT18b, KSZ15, LXL17, DPL18a, MS15a, MZS18, Pel17, RSO^{+}18, Tha18a, WXM^{+}18, Zei16, ZM19]. two-phase-like [GOA19]. two-point [AM15g, CFG15, GKN15a, JSM16, KMBA18]. two-predator [LB16a]. two-processor [ZSK^{+}19]. two-scale [DNC^{+}17, Wan17a]. two-sector [ARR16]. two-sided [CLJ^{+}15, CLW18, DCS15, FZLT15, KCKZ15]. two-species [GCLG15, SSZL17]. Two-stage [HWL^{+}19a, WL17, ZZC^{+}19b]. Two-step [PBBF18, DLBQ19, FR16, LC15b, RS17a, RT18a, WLFH17, ZYX19]. two-strain [CX19, FG15]. two-strategy [YLZT17]. two-term [APL15]. two-time-scale [LSL19b]. Two-walks [APPE17]. two-warehouse [BS15b, MBM16]. Type [Rho19, AKR15, ACS15, ABD16, ALL18a, AÖ15a, AÖ16, AIK15, AN15, AHS18, ASA19, AAS^{+}17, APTS15, AT15b, AP19, AKKB17, AP18b, AZAM15, BGMZ17, bBjJ16, BMO15, BM15b, BTIIE16, BAK19, BA17, Bir15, BP17b, BLS15, BS15c, CZ16, CZ18a, CDK15, CHRD16, CC15, Che15f, CSTH17, CZ19, CT15, CT17d, DAS17, DN15b, DN16, DL17b, DKP16, DÖA15, DK16, DKKM15a, DF16, ES16, FT17, Fen15, GRV15, GO15a, wG15a, Gao17a, Gav15, GS17b, GKN15a, GGHO18, GM15b, GR15c, GAS17, HA16, HJ15, HLC^{+}17, HZ17a, HVR17, tHBSfY19, Hus19, IKMW15, IAW18, Isc16, Jan15, JH15, jJ15b, jJ15c, JRM15, KA15a, KIAG16, KB15, KS18, KTA15, KAG15, Kem16, Kha17, KNBA15, KM18a, KAH15, Kim15c, KBM16, KKD17, KG18, KPP18, LY15b, LG16, LY16a, Li16a, LR16a, Li17d, Li17e, LH17c, LW18d, LLL19a, LXZ19a]. type [Li19, LW19c, LSL19a, Lia15b, LT15, LZ15e, LGK15, LX19c, LWSO19, LZZ^{+}19, LWZZ18, DPL18a, LMMR15, LLB16, MWLL17, MCS^{+}19, MRC^{+}18, MMIA19, MMS17a, MS19b, Mia17, Mig15, Mig16, MNR19, MDG18, MNK19, NLC19, Nis15, OPP18, PR15a, PWS17, Pel17, PS15c, PV16, PIP19, PV19, QZL18a, RAH19a, RZD^{+}19, RZSS15, RSKB16, RA15, Rho14, RRTR15, RM18b, SuRI15, SEKA16, SR17, STS15, SISÖ15, SHA15b, SHN16, Sha15d, SSA18, SPG17, SK17a, SIA15, SMM18, Sim15, SK17b, SHEBS19, SAS19, SV15b, Sug17, TTD17, Tch15, TTLX17, TPCL15, UKE15, VK19, WF15a, WPL15, WL16c, WK16a, WR19, XKWG15, XYG^{+}18, YSL18, YWL15, YCYL18, ZAP19, ZA16, Zei16, ZXW16, ZBBW17, ZZZ15b, ZS15c, Zha16a, Zha15i, ZS15e, Zho15a, ZSY18, Zhu16, ZSF17, uRAD19, NS17b]. type-II [NS17b]. types [MGL19, MK16, Mon19, PD15]. Typhoid [TMM18]. Typical [CM17b, HHP18].

U [GHN^{+}16]. U-shaped [GHN^{+}16]. Ulam [AB15a, AO19, BCO15, Nam19, Oni18, WL15, ZSS15, ZAP19]. ultimate [DLXZ19, WL18a, ZT16a]. Ultrametrics [DKM17]. UMV [HYL19a]. unbalanced [RKQ19]. Unbiased [CSL19]. Unbounded [DCR15, KBR15, LS16, NLC19, Obr19, OR18a]. uncapacitated [MK15c]. Uncertain [Gao17c, GKDHR19, HD16b, CLZW19, Chi15, Gao16, GhDwZS17, Gao19, GMS19, HZK^{+}19, Hua18, JLY18, JS19, JGX15, LJZ18, LLCR15, LLZ15, LCY16, LC16a, LLC18, LSZ^{+}18, LSZ^{+}19b, LXW^{+}18, LNWS18, LZ19g, LTZ^{+}17, Pai15a, Pai15b, RZSS15, RVGS17, SRSZ15, Sha16a, SS19a, SWZ^{+}19, SRCA17, SGY^{+}15, SLS^{+}18, WNMC15, WZAF15, WQK15, WLZY17, WWHZ18, XD19, XKWG15, XZZ^{+}18, XCY18, XKG15, XGQW19, YSL18, YR15, YZD15, Yan18b, YCX^{+}19, YmZHC15, ZSS16, ZLS16, ZGH17, ZLLP18, ZKKZ18, ZYWW16]. uncertainties [CXP16, CWHY15, CPCQ17, DZC15, NPT19, SML18, ZCCZ19]. Uncertainty [HYJ19, PTA16, APMM18, AE17b, DK15c, DJW19, HM15a, PN15a, SSS15, ZCX19, ZW15c]. uncomplete [YKJY17]. Unconditional [DFJ16, GLL15b, SY17b, SW17b, SW17a, Yos19]. Unconditionally [RLZ19, MMK17]. unconstrained [CY16a, DP15a, NRAEC19, Pet15b, WSdSY15, ZN15, dSRTZ18]. Uncovering [WJNW15]. underactuated [ZLWS15]. underdamped [XJZ19]. undetermined [SNL19]. Unicyclic [JFT17, WS19c, CL15a, LCM19, QZD16, SGDX18, TDS15, Wan16c, WLQZ18, YFY17]. Unification [HW17a]. Unified [AG16, CC17, ZMQ16, ARB16, BAK19, FBF18, HST17, LP15b, Liu15a, SGG15b, Su16, YPZ18]. Uniform [CGW16, CJ16, ZRLX19, BDSS15, Boy15, BCFR18, DS15, DL15a, DLXZ19, ERFE16, eFJY17, GLQ^{+}18, GG19a, ILPE17, KLQY18, LX18, LC18, LR19c, LMZW16, PBMP15, SR17, SM18, YLL18]. uniformity [GWPA18]. Uniformly [DN15a, WA15a, YZH^{+}19]. unilateral [ABD16, ZX19]. unimodal [GS15a]. union [BLNSF17]. unique [DL16]. Uniqueness [AB15a, BH15, CZ15c, Dan15a, GR17a, GZLL17, IJ15, MZM15, SZ19b]. unit [BMR17, CMR17, CM16, Mor17, PMSB15, SS15b, YQT15]. unitary [LL16b]. unites [ZXL15]. unity [EB16, ZYZ^{+}16a]. univalent [SGG15a]. univariate [KM18b]. university [NSN^{+}18]. Unknown [DZC15, NT16, YSWZ17, CLZW19, CSL19, Dan18, DCJ19, FW18a, Kac15, KZT19, LY15d, LLZD18b, LZHC18, Ma17b, THM19, Wan15e, XCDZ19, ZCW19a, ZZMS19, ZDF17]. unknowns [MHCF15]. unobservable [YLW16]. unpredictability [BKAAM15]. unrelated [ZXDP18]. unreliable [KKD17, LSCK15, NMS^{+}18, SWSW19]. unremitting [MBM16]. unsolved [BM16]. Unstable [dSSB^{+}16, CGW19, WZZL16, YZ16, ZMC19a]. unstable-mode [ZMC19a]. Unsteady [DS15, HRS17, Awa15, EAY17, MSM15, Pan15, Sha12, WW18, WX16, Xu19c, Yil16, ZZ15a, ZCU15, ZG15b]. Unstructured [Ort15a, PP19, SdA18, STA^{+}18]. update [AP18b, KK16a, WJ19]. updating [CCYT19, LMW^{+}19, MC15c, TIT15, XDG^{+}19, Yua15, YZZ18, ZSL17]. upon [KMS18, LWW15c]. Upper [MMG16, Zha19a, ZZ15d, ABQ15, AK16b, CYY19, WDB18]. urban [ATV16, ATV19, TES15]. Urmia [KSK^{+}15]. urokinase [KKS^{+}16]. Urysohn [Dar16, Hus19]. Usage [EL19a, CFdCdS18, EL19b, EL19c]. USAOR [WS16]. Use [QBNF19, SD17, Kim15b, MWH^{+}18, UK18, dNdOX^{+}15]. used [KTD15, Kri19b, Świ15]. user [HHR17]. Using [AGM16, AE17a, CvRvdM18, Fan15a, SSN15, ZM19, AZL15, AR16a, ARACCA^{+}18, AM15b, ASASM16, AZ14, AM19, AK15b, AK15c, ABMA19, AM18d, AAMC19, AZ19, BGMZ17, BMDS17, BM15a, BT15a, BM15b, BAE15, BMC19, BZ17, BKZA18, BL18, BCD^{+}19, CCL^{+}15, CLR16b, CTT^{+}19, CD16a, CDZ^{+}18, CPS18, DK15b, DT16, DLY17b, EGH^{+}19, EdLC18, FSW16, FMSG15, FBF18, FY18, Gao18, GJR^{+}15, GK16a, GDZ18, GDS16, Gia15b, GWPA18, GG19a, HH16, HHP18, HR18, HK18b, HSS15, HGM15a, HDG15, Ili19a, IKHS16, JA17, KBK^{+}16, KZ16, KRF17, KJ15a, KRS18, KM15a, KMBA18, KAH15, KSK^{+}15, KŚK^{+}18, KVS15, KSP16, KAS16, LVAB^{+}16, LLC^{+}15, LR19a, LLMS16, LY15a, LPJ18, LWT^{+}15, LL16a, LCL17a, LG17, LZ19c, LDSD19, LHHS16, LZKX19, LO15, MBHS18, MM04, MS15a, MHCF15, MPS15, MHR18, MO16]. using [MMG15, MRQZ19, MP19b, MMS17b, MD17, MSH15, MMB15, MRR17, Nag17, NE16, NC16, NJM^{+}18, NMNTBXVD15, NKKS16, NSM16, NEK18, OP19, ÖÖAD15, OA18, PAE^{+}17, Pan15, Par15, PHSB18, PVV^{+}15, PP19, PN15b, QFTT15, RAH19a, RSKB15, RS15a, RS17b, RZMB17, Rog19, RL15, RCM15, RME18, RPJ^{+}18, RG18b, RSO^{+}18, Sab18, SNS^{+}16, Sak19, SHH15, SSB18, SM16, SR15e, SN15a, SN15b, SHLL18, Sha12, SRP17, SuSR15, SSM^{+}15, SGFCRM^{+}18, Sin18, SHEBS19, SFC^{+}17, SLTH15, STK15, SKP16, SH19, SRBA15, SLS^{+}18, SCM19, TPS15, Tan15c, Ton18, Tua15, UKE15, UMF19, Vil16, WC15b, XTXW15, YTH19, Yüz16b, ZBR15, Zam19, ZT15, Zei16, ZLL15b, ZLWS15, ZZLY15, ZYY17, ZjW19, ZZMS19, ZYLY17, ZLK^{+}19, ZYWW16, ZW15e, lZA^{+}19]. utility [GLL15a, KCSD^{+}18]. utilizing [AD17, GSVK17, NT19]. Uvarov [AHM16]. Uzawa [CY16b, CC16, DLF15, HM15d, HM15b, LWK15, LSjL17, Mia17, SL15a, WLLK17, YLW15, ZM15c, ZM18b, ZSF17]. Uzawa-HSS [YLW15]. Uzawa-PSS [CY16b]. Uzawa-type [Mia17, ZSF17].

V [GHN^{+}16]. V-shaped [GHN^{+}16]. vacation [MJSA19]. vacations [Amm15, Shi15b]. vaccination [AKT19, Bhu15, CMW17, DAQI15, TXL19, Zha15e]. vaccinations [CX19]. vacuum [LTL16]. valency [JHL19, LLL18a]. valent [AS19b]. validated [JGBY16, WW16a]. validation [FFG^{+}16, GJP18, IR16, SCS^{+}15]. Validity [SB16]. Vallée [AÖ15a]. Valley [PTA16]. Valuation [ZLS16, LW15a, YYWH19]. Value [ABI19, MAJR19, eMA15, AKS15, AOH17, AHO18, AMO18, ASASM16, AAKT15, Ali15, AV18, ASV17, AJP18, BRD15, BÁR15, Cho19a, CLP^{+}18, CZ15c, EHS17, FFT15, GW16, Ghe16, GR19a, GGZR16, HN14, HASC19, HXM18, JSM16, KM17b, KMBA18, KAI19, KD18, KKST16, KKST17, KST19, KPT16, Kyr17, LL15a, LLY15, LWT^{+}15, LZ17b, LD18a, Li18, LS19a, LBB18, MG15, MH19a, MC17, MCZ15, MYH15, MKN^{+}15a, MTJ15, NN19, PZ15, RRRR16, RSKB15, RS17a, RR18a, RRV15, RG19, RGA19, RMK19, SH15a, Sah19b, SM19, Sow18, TES15, TG16, Ugu19, VK15, VK17a, WZ15a, WZZ18, Won15, jXlZX18, XX19a, XLW16, YZ15a, YU15, Zha15b, Zha18e, HW17c]. Value-at-Risk [HW17c]. value/boundary [MTJ15]. valued [ACKS16, CHL17, CZLS18, CZSH17, DPL^{+}18b, GG15a, GZLL17, HCX^{+}17b, JCX19, LZ15a, LWL^{+}18, LGC19b, PAM19, QBC19, SCC17, TLX19, TZD^{+}19, Tua15, WM15, WZS19, bWzZjH16, XZS16, YL17, ZLHH15, ZYXF17, ZSZ17, ZYC18, ZWZ19a, Zha15h]. values [Bab16, CL16b, CWD16, GST15a, Har15, MRS19, SQL16, YDCS15]. Vandermonde [BR16a]. vanishing [MH17a, MHZ18, TQ18]. vapour [DMSS18]. variability [CJG^{+}18, dSSB^{+}16]. variable [AwEE15, Ali15, AT15b, BJ19, BdlRT18, CCNQ19, CLJ^{+}15, CT15, CT17d, DDY15, DWRZ17, Du15, ES17a, eFJY17, GY15, GZC19, GGM18, GA17, GLL15b, HGH19a, HA19, HAY19, HAH19, HKS19, HM18c, HGH19b, JSM16, KBK^{+}16, KF19, LWQ18, LP15a, LZ18a, LMMZ18, LZ19f, MS18a, MKG15, MH17c, MK16, OW18, Our16, PTO19, PST18, Pol19, RSKB15, RSKB16, RW17, RRTR15, SR15b, SY15a, SV17a, SBH16, SMM18, Ste19, SN18c, TMS15, TB15, TF18, VS17, WTJ17, WA18, WR19, WZWD15, XYG^{+}18, XXZ^{+}18, YZMM15, ZWF15, ZEW19, ZX16, Zho17, ZLW18]. variable-coefficient [BdlRT18, HGH19b, WTJ17]. variable-distributed [Ali15]. variable-order [HAY19, HAH19]. variable-type [SMM18]. variables [Akt15, Bou18, FvdMS15, Gav15, KIAG16, LZH16, Mic15b, XRB17, YZD15, Zha19b]. variably [NAM18]. variance [CLR16a, CSL19, JPY16, Rez15, ZL15c]. variant [AIP15, ATV19, ACC15, CCTV16, FZLZ19, JLX15b, Kaj18, LSZ^{+}19b, LZ19e, LMMR15, RR15, YMHW15]. Variants [WLLK17, CR15]. variation [CCX15, DHW^{+}19, GLW15, Hac17, Lu15, LJL16, SHLW16]. Variational [MG19, CL16a, GJNRG15, Ied15, JMW15, KwC15, LjL15b, SZ19a, SDK15, WW16a, WwQqTW18, XS19b, Yil16, Zha15i, ZCQ15]. Variational-hemivariational [MG19]. variational-like [GJNRG15]. variations [KOR19, LVAB^{+}16, ZML^{+}17]. various [ANEP19, LWLM19, LP15b, Liu17a, RMB15, SK15, SB15]. Varley [XZCC15, ZS15c]. varying [BPC15, BB15b, BHRW16, CM18, CZSH17, DLXZ19, DEK15, DKDD15, FLL15, FYZ16, GHW19, GSPH19, GMPP18, GLM16, GZ15a, GWXZ18, GZD19, GSVK17, HW18, JHWZ15, JGX15, JGKL16, JSW^{+}18, Kim15b, KM15d, KKL18, LLP17, LY15d, LJ15a, LC16a, LSZY17, LCAA17, LSZ^{+}18, LSMT18, LSLZ18, LZ18c, LQC^{+}18, LGL16, LWZZ18, LJHW19, MJL16, MRC^{+}18, MMMMB17, NR16, NLC19, PY19b, RSC^{+}17, RZHY16, SM15b, SSZ^{+}18, Sha16a, SLZ^{+}16, SCC17, SLS^{+}18, SWZL19, TDL^{+}17, WLS15, WYCZ17, WLZY17, WCHW19, WXXY19, XCCL18, YLCW18, ZLW19, ZDPZ15, ZWHC15, ZZC16, ZHJ^{+}17, ZYXF17, ZHWW17, ZZ18a, ZW19b, ZYLY19, ZCAN19, ZCX19, ZXX^{+}19]. Vector [AZ19, LZ15a, LGC19b, Ace19a, AZ14, AK15b, ACMT15b, ACMT15a, AM18d, CS16b, FPUK15, GTBC18, GZ15a, GG19b, KSK^{+}15, Lah18, NS15a, Tua15, Vil16, YSD^{+}17, ZBKS16, ZLJ15, Zha15h]. vector-borne [Vil16, ZLJ15]. Vector-valued [LZ15a, Zha15h]. vectorial [Vil16]. vectors [CGJZ18, Dan15a]. vegetation [LLWW19]. vehicle [GKDHR19]. veins [STVCC17]. velocity [AwEE15, BGP15, DPL18a, Pon17, SSG15b]. venous [Tor16]. Verification [VVA^{+}15, WN16]. Verified [Miy17, Miy19, Liu15g]. version [AK15b, DLH15, GFWH17, GG19b, LpLxG18, MFF15, RBB19, YQD17]. versions [LBRS18]. versus [Ask18, CJN19, DH19a, GMPP18, LB15a]. Vertex [LLM^{+}18, LLTW19, Rad17, CCLW18, CH19, CRC19, CMT^{+}18, CYW18, GPRCTRY19, Gou18, HLX17, HTM18, HLL18, Ili18, KMS18, KS19a, LLS15, LYW17, LZ18e, LP16a, MBW19, MK15a, RZH^{+}19, Tu17, ZTW19]. vertex- [LLS15]. Vertex-based [LLM^{+}18]. Vertex-connectivity [LLTW19].

Vertex-degree-based [Rad17]. vertex-transitive [HLL18]. vertex-weighted [CCLW18]. vertical [EAY17, HRS17, IYM19, MRR17, PMV16, RDM19, SSN15, SKS17, SBH16, SBH17]. vertically [RRT15]. vertices [ATX19, CZ17a, CZ17c, GHM18, JYSR18, Lai15, SWG^{+}17, SG19a, STD15, ZY18a]. very [FY19]. Veselov [ZH15a]. vessel [ARACCA^{+}18, BKKN16]. vessels [CKKF16]. VF [DA19]. via [eMA15, eMBHA15, AÖ15a, ABHR15, ANGC17, AM15c, AK19a, APTS15, AM15d, AM15h, BR16a, BRM^{+}19, BdlRT18, BBD15, CCYT19, CMR17, Che15c, CLCW17, CWLA18, CLHC16, CN15a, Cim18, DH19b, DLY17b, EE18, FYZ16, FYSC18, GZC19, GHW19, GK17, GZK16, GA17, GS18a, GS19, GZ15a, GWXZ18, GSV18b, HT16, Isc16, KE16, Kim15b, KSN15, KZ19, KK16b, Laz17, LPKS16, LZS15, LJL15a, LZjy18, LGK15, LS19c, LHZM19, LMZZ19, LZH^{+}18, LJHW19, LCLZ19, Luo19, Mah18b, MS17a, MH15b, MH16a, MH17b, NL18a, NRAEC19, NNA15b, PS18a, PWFY18, PW19a, PW19b, PJ15, Pos17, PKB18, QKGW18, RZD^{+}19, RXZW17, Rem18, RM18b, SAJ18, SA18a, SKKL17, SISÖ15, SY18a, SPST17, SPS18, SDK15, SRCA17, SNL19, Sug17, SWZL19, TTZ^{+}19, TB17a, WZAF15, WZYS15, WW17, WHTW17, WCZY17, WXWX18, WCH15, WLZ19b, XZZ^{+}18, XHW18, YWZS18, Yan18b]. via [YFW^{+}17, ZAP19, ZZLZ17, ZLDZ17, ZYH15, ZYXF17, ZHWW17, ZCQ15, ZCX17, ZHS18, ZZ16c, uRAD19]. Viability [XL19]. vibrating [BM18a]. vibration [ANGC17, AB18, Che15d, CvRvdM18, Hac17, JLY18, LL19b, NMNTBXVD15, Pai15b, RM19b, Sin18, ZLL15b, ZGZ16]. Vibrational [MRF18, BHD18, YWL18]. vibrations [MB17, STK15, SH18b]. vibro [RL19]. vibro-acoustic [RL19]. ViennaMaterials [WWR^{+}15]. view [FB19, KM15c]. viewed [Cal18, MGCAPP16]. viewpoint [Han19, Wan16a]. views [MN16]. violent [San19]. viral [AVCCGAVDL15, GXH18, YW19a]. Virtual [ZF18, WWC^{+}18, ZDZ15]. virtualization [Chy16]. VirtuaSchlieren [GG18]. virus [CYLW15, DTJ18, LZL^{+}19b, LXT17, MMK18, PJ18, WLS15, WGLZ16, WM18b, WSXR19, XGZ17]. virus-to-cell [LXT17]. viruses [SKHA18]. viscoelastic [Alk19, DH19b, DQ18, FFL16, HMZ17, KKG18, TY16, Xu19b]. viscoelasticity [CLP^{+}18]. viscoleastic [TSPF18]. viscosity [AAM19, HG18, Jeo16, MJ19, SDK15, SCM19, Zha16c]. viscous [BKKN16, EM15, GYZC15, LCS15, Li17b, MDH15, NKNA16, Pan15, Sha12, WZY^{+}19]. Visualization [Zil15]. vitreous [TSPF18]. vivax [KCL19]. vivo [PSP17]. Vlasov [MM15a, PMS16, WWMX16]. vocal [SH18b]. voids [VVA^{+}15]. Voigt [AQ15, YLL18, ZZ15d]. Voigt/complex [AQ15]. Volatility [YYWH19, CLR16a, ZBL15, ZL15c]. volcanic [Zei16]. voltage [KSKW^{+}19]. Volterra [Tom15, AR16a, AD16, AP19, AD19, Bab16, BM15a, BB15c, CJ15, Dar16, DA17, DJ15, ER15, FMX15, FDPP18, GLM19, GIS15, HJ15, KE16, KJ15b, KZ15, Liu15d, LC15d, MHN15, MWY17, Ma17a, MRE18, MS19b, Mic15a, MH17a, MHZ18, MH15a, MH16b, MH16a, MH17c, MH17b, MS16a, MO15b, NA19, Oka18b, OPP18, PHM19, QD18, RME18, SR15c, SM16, SRK18, SRS17, SR19, Tan15b, Tan15c, Tom12, Vad19, Wan15f, Wan17c, Xy16, Xu19b, YWL19, YYM19, Yüz15, Yüz16a, ZJ15a, ZY15a, ZW15e, ZJO^{+}18]. Volume [DN16, BR16b, BDH16, CR18, DC16, EFPT18, EdLC18, FZLT15, GY15, GfSmL16, HR16, KNK15, KA15c, KS19c, LZ19c, LXL17, Maj18, MO16, Mor17, PSY^{+}17, PP19, RC17a, SA16b, STA^{+}18, SCS^{+}15, TZH15, WZ15b, WYLG19, YCLS19, ZL15a, ZSY17, ZY15c, ZSY18]. volume-of-fluid [LXL17]. volume-surface [EFPT18]. volumes [HV19b]. Volumetric [Alz15]. voluntary [GSD^{+}17, HI15b, LWWZ17]. Voronovskaya [BLS15]. vortex [Ili19b, RSO^{+}18]. vortex-ring [RSO^{+}18]. vorticity [BFR18, YZ17b]. vorticity-streamfunction [BFR18]. Voyce [DD15]. Vreman [DN15c]. Vries [AM18a, AO15c, Gao17b, HXM18, Kud16, Mis15, SGM^{+}16, SzSC19, WTJ17]. vs [KM18b, ZAY18]. vulnerabilities [DLCV17]. vulnerable [DLCV17].

W [GPHAPRW16]. Ważewski [GRV15]. WAFOM [Har15]. wages [TES15]. waiting [YYX^{+}19]. wake [BTP17]. walk [CGJZ18, Gat18a, MT15]. walks [APPE17, AAMCSV18, WGL18]. wall [HMHA15, KPG16, PBP19]. walled [Kia15]. Wallis [CP17, QM15a]. Walsh [SM16, XS15]. Wang [Tal15]. warehouse [BS15b, MBM16]. Water [GG19a, RDM19, SEKA16, ATK^{+}16, BTP17, BMT18, BKLW15, BG17, DKL15, DMSS18, EB16, GCH^{+}19, GWK16, GYW15, KNK15, Kem16, KSK^{+}15, LCQ15, Mou15, NB15, PBP19, PVV^{+}15, RG15, SSG^{+}15a, Tha18b, UMF19, WLQ^{+}19, WMSP19, WF15c, hY18, ZBR15]. water-Al [ATK^{+}16]. waterborne [SK19]. Wave [GJX^{+}19, YJY16, ZWZT18, Ach19, AM18a, ANU16, AHY19, BB18a, BRV18, BCI15, CS19, CL15b, DCS15, DL18b, DZL15, DNC^{+}17, lDqCWzH18, Esl16, Fan18, FTE19, GO15b, GM15a, Gao17c, GMS19, GHN^{+}16, GR17b, GB19, GG19a, HEA15a, HEA16, HAY19, HAH19, JZ19a, JX19, KKT15, KZ16, Kud19b, LZ18b, LW18f, LLX18a, LL19c, LYDH18, Liu15a, LWS18, MDHD18, MD18a, MAG15, MZS18, MS15d, PVL19, Por18, QA19, RPJ^{+}18, RDM19, STJS16, SEKA16, SSG^{+}15a, SZL^{+}18, STL17, SV17b, Sun15b, SW17c, VDI17, WTXZ16, WWH17, WXZ18, WDG19b, Wei17, WMSP19, YWRP16, YJW16, YMHW15, ZH15a, ZYY17, ZZM18, ZXZB19, ZLX19, ZL19a, ZL19b]. waveform [AL18d, Fan15a, Lia15b, LJS17]. waveforms [BB18b, PP18b]. wavefronts [MWY17]. Wavelet [JLL15a, PK18b, SH15a, VK15, VK17a, Bag19, BBFJ17, CCY19a, CXL18, EGB15, HAH19, Kri19a, Liu16a, Mah18a, NLC19, PLL16, PSP17, QFS17, SuR15, SuRI15, SK17a, SPS18, Sko17, Zha15h, BGSCGL19]. wavelet-based [PSP17]. Wavelets [HHGC16, AND17, CKW15, Hsi15, KOR19, Kri19a, OA15, SR15c, WXWX18, XYG^{+}18, ZX16, GR15b]. waves [AOS19, BW15b, CY18, DWTT17, HMSS19, KRB15, KSV15, Kud19a, LFY18, MZ17, MKRS16, Mis15, PRA^{+}19, PBBF18, RG15, Sin17, UKK17, Wan15f, Wan16b, WL16a, WC16, WZZT19, WYZ^{+}17, XYDL16, YWL19, YX19b, ZZBS16, ZSX15, ZLH16, ZYLY19]. wavy [PBP19, SRP17, SBH16, SBH17]. way [PVL19]. WCNS [ZSXW18]. Weak [AK15c, HMZ17, ZZCY16, AR16b, Boy17, CKG19, CZ19, GR17a, JS15, KG17, LWZ16, Liu18a, Neu16, NB19b, RN18, SS18b, WZWJ18, WYL^{+}19b, ZT16b, ZC19b, ZL17b]. weak-form [Liu18a]. weak-strong [GR17a]. weaker [AM15e, JGGK15, LR19a]. weakly [AD16, AST18, DHS19, DC16, LX19a, MHN15, NL18a, PHM19, Pan18, QXW19, SPST17, SRS17, WZ16b, YTH19, ZF18, ZW15e]. wear [MCPO15]. web [PCA16, LYC15]. wedge [HRS17, SSB18]. Weibull [Elm15, NBAGS19, ÖAD15, ÖÖAD15]. Weierstrass [AAMCSV18, CP17, PV16]. Weierstrass-type [PV16]. weight [AHM16, ABMA19, ACTV15, CCC19, Gou18, LKM17, LSX19, LLdM16, MDG18, RZH^{+}19, WA15b, WX19b, Xu19a, ZTW19]. weight-degree [LLdM16]. weight-function [ACTV15]. weight-homogeneous [LLdM16]. Weighted [APTS15, APL15, ALP15, BLS15, CZP15, CLT18, DD19, HLT15, Kyr17, MD15b, MD15c, Alf17, bBjJ16, Bog16, CCLW18, CN15a, CN19a, DL17b, DCS15, GW16, GLL15a, GT15b, GfSmL16, HJMB18, HP17, HLT16, HT16, HZ17c, JJM18, jJ15b, jJ15c, JT17, KBS17, KRS18, Kyr15, Laz17, LT15, Liu17a, LjL15b, MTY16, MCY19, MBW19, MLW18, NS15b, PPV15, PS16, RR18b, Red19, She18, SY18a, TYD15, WMS17, WX19b, WYLG19, XYZK15, XLYZ19, Zha17c, ZZ16b, ZDJ19, Zhu16]. weighted-EP [MD15b]. weighting [GKN15b, NEK18, WY15a, ZHJ^{+}17]. weights [KDTP18, LP15a, PS15b, RR18b]. weir [ZBKS16]. Weissler [HK15]. Welch [GN19a]. welding [PKHD18]. well [Can16, DB16, GWK16, GS17c, JL15, LCQ15, LJD17, LZ19c, LL16c, LFY18, MO16, QLLY18, RRT15, Tha18a, WLQ^{+}19, XJZ19, YY15, hY18, Zha16c, ZFL18]. well-balanced [DB16, LCQ15, LJD17, LZ19c, MO16, QLLY18, Tha18a, WLQ^{+}19, hY18]. well-posedness [GS17c, JL15, LL16c, Zha16c, ZFL18]. wellbore [WXM^{+}18]. Wellposedness [SLW15]. wells [BW15a, GLYT16]. Wendroff [ZY15c]. WENO [BDH16, Hua16, KS19c, LCQ15, WLQ^{+}19]. Werner [RA15]. Werner-type [RA15]. Western [DCK^{+}18]. wheel [TZS17]. which [Bir15, ID19, LZ16, PAK15, RP18]. white [MX17, ZJO^{+}18]. Whitham [RG15]. Whitney [BRV18]. whose [AK18a, JYSR18]. wide [GG19b]. widening [BCT15]. Wiener [Ort15b, ADA18, AK18a, AOS17, AG17, BKL^{+}17, CXL16b, CT17b, GKS16, Gou18, HKS19, HD16a, KJR^{+}18, KST15, KST16a, KST16b, KMS18, LL18b, PWFY18, SZPH16, SISV19, Tra18, ZWZ19b]. Wilf [CP17]. willingness [WJ19]. Wilson [AAGZ19, SZL^{+}18, SKBR18]. win [DZD^{+}16]. win-stay-lose-shift [DZD^{+}16]. wind [FFH^{+}18, STK15]. wind-induced [STK15]. window [Khl18, XY15b]. windowed [BBFJ17]. Windows [Nav17, PHSB18]. wings [Ber15]. Winner [TKDF15]. wireless [SZRC15, TTD18]. wiring [GLA^{+}17]. Wirtinger [SLS^{+}18]. with/without [AM19]. WITHDRAWN [EL19c]. within [AJ17, hCWW^{+}15, FLS^{+}15, KBK15, Mah15, MMK18, SSB18, WLS15, YX15]. within-host [MMK18, WLS15]. without [ASHSC17, AM19, AKKB17, BDGS15, CH19, GYY18, GS15b, JPZ^{+}19, LY18a, LML18b, PLL16, SWG^{+}17, WNW^{+}19, ZY15b]. wood [WZM15, ZKV^{+}15]. Woodbury [Xu17b]. work [KGGAV^{+}18]. workers [TES15]. working [Her19]. world [EOP17, LWSX18a]. Wreath [BCD18]. Wright [BC16]. wrong [UK18].

X [KGGAV^{+}18].

Yang [CCY19b, KC18b, LZY^{+}17]. yielding [JSG15, SGJS16]. Young [Van15]. yourself [ZWXW15]. Yukawa [Tch15].

Zagreb [Pal19, BF16b, DSAB16, JQL15, LZ19b, MMDG16, Pal17, SDAB17, SIA16, SGDX18, WWL18]. Zakharov [PZ17, PKB18, Tho17, XZ19]. Zarei [WL18a]. Zassenhaus [AHK^{+}19]. Zernike [DG18a, DG18b, LB15a, RLSGFMMF16]. Zero [DJW19, JIS15, AL19, ACP19, AG15d, GKN15a, HHR17, IÖ19, KT16, LKM17, LLP17, MKRS16, MS15c, OLZ19, ZZLZ17, ZFL18]. Zero-dissipative [JIS15]. zeroeffort [MS15c]. Zero-sum [DJW19]. zeros [AGN15, BEM15, BMB15, Cas15, CCR18, CS15b, DCT19, Dan15a, DSM15, JRSK16, MC15b, Pro16, PIP19, PV19, Zad16]. zeroth [DD16a, LL15a, STD15]. zerothorder [DD16a, LL15a, STD15]. zeta [BB15a, SGG15a, TCY15]. zeta-function [TCY15]. Zhang [WAS19]. Zika [DTJ18, LZL^{+}19b, WSXR19]. ZNN [GZ15a]. zonally [ZYLY19]. zone [XCDZ19, ZWO^{+}19]. zooming [LY15a]. zooplankton [JA15, YYZ19]. Zufiria [SEKA16]. Zygmund [bBjJ16].

References

Anita:2016:NSP

[AA16] Laura-Iulia Anita and Sebastian Anita. Note on some periodic optimal harvesting problems for age-structured population dynamics. Applied Mathematics and Computation, 276(??):21–30, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315016136.

Ahmad:2015:TFR

[AAA^{+}15a] B. Ahmad, M. S. Alhothuali, H. H. Alsulami, M. Kirane, and S. Timoshin. On a time fractional reaction diffusion equation. Applied Mathematics and Computation, 257(??):199–204, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300314009436.

Askar:2015:DCD

[AAA15b] S. S. Askar, Ahmad M. Alshamrani, and K. Alnowibet. Dynamic Cournot duopoly games with nonlinear demand function. Applied Mathematics and Computation, 259(??): 427–437, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315002726.

Askar:2016:ACC

[AAA16] S. S. Askar, Ahmad M. Alshamrani, and K. Alnowibet. The arising of cooperation in Cournot duopoly games. Applied Mathematics and Computation, 273(??):535–542, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013697.

Alzubaidi:2016:FAS

[AAC16] Hasan Alzubaidi, Xavier Antoine, and Chokri Chniti. Formulation and accuracy of On-Surface Radiation Conditions for acoustic multiple scattering problems. Applied Mathematics and Computation, 277(??):82–100, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300114.

Aybar:2018:DAT

[AADF18] I. Kusbeyzi Aybar, O. O.

Aybar, M. Dukarić, and B. Fercec. Dynamical analysis of a two prey-one predator system with quadratic self interaction. Applied Mathematics and Computation, 333(??):118–132, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303047.

Alanis:2015:BFO

[AADLF15] Alma Y. Alanis, Nancy Arana-Daniel, and Carlos Lopez-Franco. Bacterial foraging optimization algorithm to improve a discrete-time neural second order sliding mode controller. Applied Mathematics and Computation, 271(??):43–51, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315011686.

Azad:2016:PSB

[AADMM16] H. Azad, Ahmad Y. Al-Dweik, F. M. Mahomed, and M. T. Mustafa. A point symmetry based method for transforming ODEs with threedimensional symmetry algebras to their canonical forms. Applied Mathematics and Computation, 289(??):444–463, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303447.

Alalwan:2019:MCN

[AAE19] Najlaa Alalwan, Alex Arenas, and Ernesto Estrada. “Melting” of complex networks. A mathematical model of complex networks resilience to external stress. Applied Mathematics and Computation, 362 (??):Article 124579, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305715.

Adam:2015:GSF

[AAF15] Maria Adam, Nicholas Assimakis, and Alfonso Farina. Golden section, Fibonacci sequence and the time invariant Kalman and Lainiotis filters. Applied Mathematics and Computation, 250(??):817–831, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015446.

Achchab:2019:EFN

[AAGZ19] Boujemâa Achchab, Abdellatif Agouzal, Allal Guessab, and Yassine Zaim. An extended family of nonconforming quasi-Wilson elements for solving elasticity problem. Applied Mathematics and Computation, 344–345(??):1–19, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308464.

Avaji:2019:SAH

[AAH19] M. Avaji, A. Jodayree Akbarfam, and A. Haghighi. Stability analysis of high order Runge–Kutta methods for index 1 stochastic differential-algebraic equations with scalar noise. Applied Mathematics and Computation, 362(??): Article 124544, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305272.

Arthur:2018:NKF

[AAHT18] Joseph Arthur, Adam Attarian, Franz Hamilton, and Hien Tran. Nonlinear Kalman filtering for censored observations. Applied Mathematics and Computation, 316(??):155–166, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305532.

Aleroev:2015:OCP

[AAKT15] M. Aleroev, T. Aleroev, M. Kirane, and Y-F. Tang. On one class of persymmetric matrices generated by boundary value problems for differential equations of fractional order. Applied Mathematics and Computation, 268(??):151–163, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008577.

Ameenuddin:2019:ESD

[AAM19] Mohammed Ameenuddin, Mohan Anand, and Mehrdad Massoudi. Effects of shear-dependent viscosity and hematocrit on blood flow. Applied Mathematics and Computation, 356(??):299–311, September 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319302292.

Assari:2019:NSS

[AAMC19] Pouria Assari, Fatemeh Asadi-Mehregan, and Salvatore Cuomo. A numerical scheme for solving a class of logarithmic integral equations arisen from twodimensional Helmholtz equations using local thin plate splines. Applied Mathematics and Computation, 356(??):157–172, September 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319302498.

Almaguer:2018:RWW

[AAMCSV18] F-Javier Almaguer, Omar González Amezcua, Javier Morales-Castillo, and Roberto Soto-Villalobos. Riemann and Weierstrass walks revisited. Applied Mathematics and Computation, 319(??):518–526, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303600.

Aliev:2017:SMS

[AAS^{+}17] Fikret A. Aliev, N. A. Aliev, N. A. Safarova, R. M. Tagiev, and M. F. Rajabov. Sweep method for solving the Roesser type equation describing the motion in the pipeline. Applied Mathematics and Computation, 295(??):16–23, February 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305926.

Abbas:2015:UUS

[AB15a] Saïd Abbas and Mouffak Benchohra. Uniqueness and Ulam stabilities results for partial fractional differential equations with not instantaneous impulses. Applied Mathematics and Computation, 257(??):190–198, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314009175.

Akbar:2015:CNA

[AB15b] Noreen Sher Akbar and Adil Wahid Butt. Carbon nanotubes analysis for the peristaltic flow in curved channel with heat transfer. Applied Mathematics and Computation, 259(??):231–241, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002404.

Aria:2018:CVB

[AB18] A. Imani Aria and H. Biglari.

Computational vibration and buckling analysis of microtubule bundles based on nonlocal strain gradient theory. Applied Mathematics and Computation, 321(??):313–332, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307592.

Alzahrani:2018:SHO

[ABA18] Abdullah Khamis Hassan Alzahrani, Ramandeep Behl, and Ali Saleh Alshomrani. Some higherorder iteration functions for solving nonlinear models. Applied Mathematics and Computation, 334(??):80–93, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303011.

Asdrubali:2018:DTB

[ABB^{+}18] Francesco Asdrubali, Giorgio Baldinelli, Francesco Bianchi, Danilo Costarelli, Antonella Rotili, Marco Seracini, and Gianluca Vinti. Detection of thermal bridges from thermographic images by means of image processing approximation algorithms. Applied Mathematics and Computation, 317(??):160–171, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306161.

Abbasbandy:2016:NEM

[ABC^{+}16] Saeid Abbasbandy, Parisa Bakhtiari, Alicia Cordero, Juan R. Torregrosa, and Taher Lotfi. New efficient methods for solving nonlinear systems of equations with arbitrary even order. Applied Mathematics and Computation, 287– 288(??):94–103, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302880.

Abdelmaguid:2015:NSF

[Abd15] Tamer F. Abdelmaguid.

A neighborhood search function for flexible job shop scheduling with separable sequence-dependent setup times. Applied Mathematics and Computation, 260(??):188–203, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003732.

Abide:2016:ATA

[ABD16] Stéphane Abide, Mikaël Barboteu, and David Danan. Analysis of two active set type methods to solve unilateral contact problems. Applied Mathematics and Computation, 284(??):286–307, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302041.

Amodio:2017:ACN

[ABGL17] Pierluigi Amodio, Yuri Blinkov, Vladimir Gerdt, and Roberto La Scala. Algebraic construction and numerical behavior of a new s-consistent difference scheme for the 2D Navier–Stokes equations. Applied Mathematics and Computation, 314(??):408–421, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304502.

Achchab:2015:GAC

[ABGS15] Boujemâa Achchab, Khalid Bouihat, Allal Guessab, and Gerhard Schmeisser. A general approach to the construction of nonconforming finite elements on convex polytopes. Applied Mathematics and Computation, 268(??):916–923, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009121.

Agarwal:2015:TFD

[ABHR15] Ravi P. Agarwal, Dumitru Baleanu, Vahid Hedayati, and Shahram Rezapour. Two fractional derivative inclusion problems via integral boundary condition. Applied Mathematics and Computation, 257(??):205–212, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401457X.

Amodio:2019:NCS

[ABI19] Pierluigi Amodio, Luigi Brugnano, and Felice Iavernaro. A note on the continuous-stage Runge– Kutta(–Nyström) formulation of Hamiltonian Boundary Value Methods (HBVMs). Applied Mathematics and Computation, 363(??): Article 124634, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306265.

Agarwal:2015:OBS

[ABL15] Ravi P. Agarwal, Martin Bohner, and Tongxing Li. Oscillatory behavior of second-order half-linear damped dynamic equations. Applied Mathematics and Computation, 254(??):408–418, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017524.

Agarwal:2015:EOH

[ABLZ15] Ravi P. Agarwal, Martin Bohner, Tongxing Li, and Chenghui Zhang. Evenorder half-linear advanced differential equations: improved criteria in oscillatory and asymptotic properties. Applied Mathematics and Computation, 266(??):481–490, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006244.

Argyros:2019:LCI

[ABMA19] Ioannis K. Argyros, Ramandeep Behl, J. A. Tenreiro Machado, and Ali Saleh Alshomrani. Local convergence of iterative methods for solving equations and system of equations using weight function techniques. Applied Mathematics and Computation, 347(??):891–902, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308476.

Abad:2015:AIN

[ABMBR15] A. Abad, R. Barrio, M. Marco-Buzunariz, and M. Rodríguez. Automatic implementation of the numerical Taylor series method: a Mathematica and Sage approach. Applied Mathematics and Computation, 268(??):227–245, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008231.

Ayatollahi:2017:ASM

[ABN^{+}17] Mojtaba Ayatollahi, Rasul Bagheri, Mahsa Nourazar, Mojtaba Mahmoudi Monfared, and S. Mahmoud Mousavi. Analytic solutions of multiple moving cracks in an orthotropic layer bonded to an orthotropic FGM coating. Applied Mathematics and Computation, 293(??):394–403, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305100.

Adell:2015:SUL

[ABQ15] José A. Adell, Jorge Bustamante, and José M. Quesada. Sharp upper and lower bounds for the moments of Bernstein polynomials. Applied Mathematics and Computation, 265(??):723–732, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007444.

Agroum:2015:SDT

[ABS15] Rahma Agroum, Christine Bernardi, and Jamil Satouri. Spectral discretization of the timedependent Navier–Stokes problem coupled with the heat equation. Applied Mathematics and Computation, 268(??):59–82, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008280.

Akgunes:2018:DPG

[AÇ18] Nihat Akgünes and Büsra Çagan. On the dot product of graphs over monogenic semigroups. Applied Mathematics and Computation, 322(??):1–5, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317307853.

Anholcer:2019:NGE

[AC19] Marcin Anholcer and Sylwia Cichacz. Note on the group edge irregularity strength of graphs. Applied Mathematics and Computation, 350(??):237–241, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300153.

Acar:2015:AFG

[Aca15] Tuncer Acar. Asymptotic formulas for generalized Szász–Mirakyan operators. Applied Mathematics and Computation, 263(??):233–239, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005226.

Amaya:2015:HSA

[ACC15] Ivan Amaya, Jorge Cruz, and Rodrigo Correa. Harmony Search algorithm: a variant with self-regulated fretwidth. Applied Mathematics and Computation, 266(??):1127–1152, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315008218.

Arnal:2016:PAQ

[ACC16] Ana Arnal, Fernando Casas, and Cristina Chiralt. A perturbative algorithm for quasi-periodic linear systems close to constant coefficients. Applied Mathematics and Computation, 273(??):398–409, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013442.

Allasia:2018:HBI

[ACD18] Giampietro Allasia, Roberto Cavoretto, and Alessandra De Rossi. Hermite– Birkhoff interpolation on scattered data on the sphere and other manifolds. Applied Mathematics and Computation, 318 (??):35–50, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317303065.

Amiri:2018:SAP

[ACDT18] Abdolreza Amiri, Alicia Cordero, M. Taghi Darvishi, and Juan R. Torregrosa. Stability analysis of a parametric family of seventh-order iterative methods for solving nonlinear systems. Applied Mathematics and Computation, 323(??):43–57, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308226.

Acebron:2019:MCM

[Ace19a] Juan A. Acebrón. A Monte Carlo method for computing the action of a matrix exponential on a vector. Applied Mathematics and Computation, 362 (??):Article 124545, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305284.

Albares:2019:SPT

[ACE19b] P. Albares, J. M. Conde, and P. G. Estévez. Spectral problem for a twocomponent nonlinear Schrödinger equation in 2 + 1 dimensions: Singular manifold method and Lie point symmetries. Applied Mathematics and Computation, 355(??):585–594, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302061.

Arregui:2019:LDG

[ACG19] Iñigo Arregui, J. Jesús Cendán, and María González. A local discontinuous Galerkin method for the compressible Reynolds lubrication equation. Applied Mathematics and Computation, 349(??): 337–347, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318311214.

Achouri:2019:CFD

[Ach19]  Talha Achouri.  Con-

servative finite difference scheme for the nonlinear fourth-order wave equation. Applied Mathematics and Computation, 359(??):121–131, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319303212.

Agarwal:2015:CIF

[ACK15] Praveen Agarwal, Mehar Chand, and Erkinjon Tulkinovich Karimov. Certain image formulas of generalized hypergeometric functions. Applied Mathematics and Computation, 266(??):763–772, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007328.

Alpay:2016:QVP

[ACKS16] Daniel Alpay, Fabrizio Colombo, David P. Kimsey, and Irene Sabadini. Quaternion-valued positive definite functions on locally compact Abelian groups and nuclear spaces. Applied Mathematics and Computation, 286(??):115–125, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302351.

Argyros:2015:CDS

[ACMT15a] Ioannis K. Argyros, Alicia Cordero, Á. Alberto Magreñán, and J. R. Torregrosa. On the convergence of a Damped Secant method with modified right-hand side vector. Applied Mathematics and Computation, 252(??):315–323, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031401683X.

Argyros:2015:CDN

[ACMT15b] Ioannis K. Argyros, Alicia Cordero, Alberto Magreñán, and Juan R. Torregrosa. On the convergence of a damped Newton-like method with modified right hand side vector. Applied Mathematics and Computation, 266(??):927–936, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007900.

Anholcer:2019:LBN

[ACP19] Marcin Anholcer, Sylwia Cichacz, and Jakub Przybylo. Linear bounds on nowhere-zero group irregularity strength and nowhere-zero group sum chromatic number of graphs. Applied Mathematics and Computation, 343(??): 149–155, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031830835X.

Amat:2018:NPT

[ACRZ18] S. Amat, A. Choutri, J. Ruiz, and S. Zouaoui. On a nonlinear 4-point ternary and non-interpolatory subdivision scheme eliminating the Gibbs phenomenon. Applied Mathematics and Computation, 320(??):16–26, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317306136.

Abdou:2015:DTC

[ACS15] Afrah A. N. Abdou, Yeol Je Cho, and Reza Saadati. Distance type and common fixed point theorems in Menger probabilistic metric type spaces. Applied Mathematics and Computation, 265(??): 1145–1154, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315006761.

Andrade:2015:APH

[ACSS15] Filipe Andrade, Claudio Cuevas, Clessius Silva, and Herme Soto. Asymptotic periodicity for hyperbolic evolution equations and applications. Applied Mathematics and Computation, 269(??):169–195, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009650.

Artidiello:2015:MGI

[ACTV15] Santiago Artidiello, Alicia Cordero, Juan R. Torregrosa, and Maria P. Vassileva. Multidimensional generalization of iterative methods for solving nonlinear problems by means of weight-function procedure. Applied Mathematics and Computation, 268(??):1064–1071, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009364.

Artidiello:2017:DME

[ACTV17] S. Artidiello, Alicia Cordero, Juan R. Torregrosa, and M. P. Vassileva. Design and multidimensional extension of iterative methods for solving nonlinear problems. Applied Mathematics and Computation, 293(??):194–203, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S009630031630529X.

Acu:2015:SSK

[Acu15] Ana Maria Acu. Stancu– Schurer–Kantorovich operators based on q-integers. Applied Mathematics and Computation, 259(??): 896–907, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315003379.

Ahmadvand:2015:IRM

[AD15] Ali Ahmadvand and Mohammad Reza Daliri. Improving the runtime of MRF based method for MRI brain segmentation. Applied Mathematics and Computation, 256(??):808–818, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000831.

Ahmadabadi:2016:TAM

[AD16] M. Nili Ahmadabadi and H. Laeli Dastjerdi. Tau approximation method for the weakly singular Volterra–Hammerstein integral equations. Applied Mathematics and Computation, 285(??):241–247, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302399.

Assari:2017:MDC

[AD17] Pouria Assari and Mehdi Dehghan. A meshless discrete collocation method for the numerical solution of singular-logarithmic boundary integral equations utilizing radial basis functions. Applied Mathematics and Computation, 315(??):424–444, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305349.

Assari:2019:MLD

[AD19] Pouria Assari and Mehdi Dehghan. A meshless local discrete Galerkin (MLDG) scheme for numerically solving two-dimensional nonlinear Volterra integral equations. Applied Mathematics and Computation, 350(??):249–265, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300219.

Ali:2018:NCT

[ADA18] Akbar Ali, Zhibin Du, and Muhammad Ali. A note on chemical trees with minimum Wiener polarity index. Applied Mathematics and Computation, 335(??):231–236, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303692.

Adewole:2019:ALH

[Ade19] Matthew O. Adewole.

Approximation of linear hyperbolic interface problems on finite element: Some new estimates. Applied Mathematics and Computation, 349(??):245–257, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311081.

Appadu:2016:CST

[ADG16] A. R. Appadu, J. K.

Djoko, and H. H. Gidey. A computational study of three numerical methods for some advectiondiffusion problems. Applied Mathematics and Computation, 272 (part 3) (??):629–647, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315004154.

Abdo:2019:GIM

[ADG19] Hosam Abdo, Darko Dimitrov, and Ivan Gutman. Graph irregularity and its measures. Applied Mathematics and Computation, 357(??):317–324, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302954.

Aledo:2019:DAP

[ADMV19a] Juan A. Aledo, Luis G. Diaz, Silvia Martinez, and Jose C. Valverde. Dynamical attraction in parallel network models. Applied Mathematics and Computation, 361(??):874–888, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304503.

Aledo:2019:SPG

[ADMV19b] Juan A. Aledo, Luis G. Diaz, Silvia Martinez, and Jose C. Valverde.

Solution to the predecessors and Gardens-of-Eden problems for synchronous systems over directed graphs. Applied Mathematics and Computation, 347(??):22–28, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309512.

Arandiga:2016:PSS

[ADS16] F. Aràndiga, R. Donat, and M. Santágueda. The PCHIP subdivision scheme. Applied Mathematics and Computation, 272 (part 1)(??):28–40, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500990X.

Aimi:2016:IAS

[ADSS16] A. Aimi, M. Diligenti, M. L. Sampoli, and A. Sestini. Isogemetric analysis and symmetric Galerkin BEM: a 2D numerical study. Applied Mathematics and Computation, 272 (part 1)(??):173–186, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315011534.

Al-Darabsah:2016:TDE

[ADY16] Isam Al-Darabsah and Yuan Yuan. A timedelayed epidemic model for Ebola disease transmission. Applied Mathematics and Computation, 290(??):307–325, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303526.

Asgari:2017:UOM

[AE17a] M. Asgari and R. Ezzati. Using operational matrix of two-dimensional Bernstein polynomials for solving two-dimensional integral equations of fractional order. Applied Mathematics and Computation, 307(??):290–298, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317301832.

Askar:2017:ICU

[AE17b] S. S. Askar and M. F. Elettreby. The impact of cost uncertainty on Cournot oligopoly games. Applied Mathematics and Computation, 312(??):169–176, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303740.

Ahmed:2018:HFS

[AEB18] Hamdy M. Ahmed and Mahmoud M. El-Borai. Hilfer fractional stochastic integro-differential equations. Applied Mathematics and Computation, 331(??):182–189, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301747.

Abdeldaim:2015:GCR

[AED15] A. Abdeldaim and A. A.

El-Deeb. On generalized of certain retarded nonlinear integral inequalities and its applications in retarded integro-differential equations. Applied Mathematics and Computation, 256(??):375–380, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000612.

Artalejo:2015:SSM

[AELH15] J. R. Artalejo, A. Economou, and M. J. Lopez-Herrero. The stochastic SEIR model before extinction: Computational approaches. Applied Mathematics and Computation, 265(??): 1026–1043, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315007729.

Ahmed:2015:BDG

[AEP15] E. Ahmed, A. A. Elsadany, and Tonu Puu. On Bertrand duopoly game with differentiated goods. Applied Mathematics and Computation, 251(??):169–179, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015835.

Ashtiani:2015:BCA

[AF15] Alireza M. Ashtiani and Paulo A. V. Ferreira. A branch-and-cut algorithm for a class of sum-of-ratios problems. Applied Mathematics and Computation, 268(??):596–608, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500870X.

Abe:2018:CBM

[AF18] Kuniyoshi Abe and Seiji Fujino. Converting BiCR method for linear equations with complex symmetric matrices. Applied Mathematics and Computation, 321(??):564–576, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307452.

Agrawal:2015:BSK

[AFK15] P. N. Agrawal, Zoltán Finta, and A. Sathish Kumar. Bernstein–Schurer– Kantorovich operators based on q-integers. Applied Mathematics and Computation, 256(??): 222–231, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314017676.

Alavizadeh:2015:NSF

[AG15a] S. R. Alavizadeh and F. M. Maalek Ghaini. Numerical solution of fractional diffusion equation over a long time domain. Applied Mathematics and Computation, 263(??):240–250, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500497X.

Alikhani:2015:RES

[AG15b] Saeid Alikhani and Nima Ghanbari. Randić energy of specific graphs. Applied Mathematics and Computation, 269(??):722–730, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010346.

Argyros:2015:BCC

[AG15c] Ioannis K. Argyros and Santhosh George. Ball convergence comparison between three iterative methods in Banach space under hypothese only on the first derivative. Applied Mathematics and Computation, 266(??): 1031–1037, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315008127.

Argyros:2015:ECB

[AG15d] Ioannis K. Argyros and Santhosh George. Enlarging the convergence ball of the method of parabola for finding zero of derivatives. Applied Mathematics and Computation, 256(??):68– 74, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000442.

Argyros:2016:UCD

[AG16] Ioannis K. Argyros and Santhosh George. Unified convergence domains of Newton-like methods for solving operator equations. Applied Mathematics and Computation, 286(??):106–114, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302521.

Ashrafi:2017:OCT

[AG17] Ali Reza Ashrafi and Ali Ghalavand. Ordering chemical trees by Wiener polarity index. Applied Mathematics and Computation, 313(??):301–312, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300317304071.

Agapito:2019:PEA

[AGC19] Giuseppe Agapito, Pietro Hiram Guzzi, and Mario Cannataro. Parallel extraction of association rules from genomics data. Applied Mathematics and Computation, 350(??):434–446, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306471.

Aledo:2016:UES

[AGM16] Juan A. Aledo, Jose A. Gámez, and David Molina. Using extension sets to aggregate partial rankings in a flexible setting. Applied Mathematics and Computation, 290(??):208–223, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303769.

Abel:2015:CMZ

[AGN15] Ulrich Abel, Wolfgang Gawronski, and Thorsten Neuschel. Complete monotonicity and zeros of sums of squared Baskakov functions. Applied Mathematics and Computation, 258(??):130–137, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000946.

Agratini:2015:SPL

[Agr15]  Octavian Agratini.  A

sequence of positive linear operators associated with an approximation process. Applied Mathematics and Computation, 269(??):23–28, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009625.

Achchab:2015:NCN

[AGZ15] Boujemâa Achchab, Allal Guessab, and Yassine Zaim. A new class of nonconforming finite elements for the enrichment of Q_{1} element on convex polytope. Applied Mathematics and Computation, 271(??):657–668, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012771.

Akin:2015:CCT

[AH15] Elvan Akin and Taher S. Hassan. Comparison criteria for third order functional dynamic equations with mixed nonlinearities. Applied Mathematics and Computation, 268(??):169–185, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008279.

Annaby:2018:TSH

[AH18a] M. A. Annaby and H. A. Hassan. Trigonometric sums by Hermite interpolations. Applied Mathematics and Computation, 330(??):213–224, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301553.

Aouchiche:2018:CGR

[AH18b] Mustapha Aouchiche and Pierre Hansen. Cospectrality of graphs with respect to distance matrices. Applied Mathematics and Computation, 325(??):309–321, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308846.

Ait-Haddou:2015:BPD

[AHG15] Rachid Ait-Haddou and Ron Goldman. Best polynomial degree reduction on q-lattices with applications to q-orthogonal polynomials. Applied Mathematics and Computation, 266(??):267–276, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500692X.

Auzinger:2019:TAZ

[AHK^{+}19] Winfried Auzinger, Harald Hofstätter, Othmar Koch, Karolina Kropielnicka, and Pranav Singh. Time adaptive Zassenhaus splittings for the Schrödinger equation in the semiclassical regime. Applied Mathematics and Computation, 362(??): Article 124550, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305338.

Alahyane:2019:LBM

[AHLR19] M. Alahyane, A. Hakim, A. Laghrib, and S. Raghay. A lattice Boltzmann method applied to the fluid image registration. Applied Mathematics and Computation, 349(??):421–438, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311123.

Arceo:2016:PAU

[AHM16] Alejandro Arceo, Edmundo J. Huertas, and Francisco Marcellán. On polynomials associated with an Uvarov modification of a quartic potential Freud-like weight. Applied Mathematics and Computation, 281(??):102–120, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300480.

Agarwal:2018:ITI

[AHO18] Ravi Agarwal, S. Hristova, and D. O’Regan. Iterative techniques for the initial value problem for Caputo fractional differential equations with noninstantaneous impulses. Applied Mathematics and Computation, 334(??): 407–421, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318303084.

Albi:2019:LMM

[AHP19] G. Albi, M. Herty, and L. Pareschi. Linear multistep methods for optimal control problems and applications to hyperbolic relaxation systems. Applied Mathematics and Computation, 354(??):460–477, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301195.

Akhmedov:2018:OQF

[AHS18] D. M. Akhmedov, A. R.

Hayotov, and Kh. M. Shadimetov. Optimal quadrature formulas with derivatives for Cauchy type singular integrals. Applied Mathematics and Computation, 317(??): 150–159, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031730629X.

An:2019:EAD

[AHY19] Na An, Chaobao Huang, and Xijun Yu. Error analysis of direct discontinuous Galerkin method for two-dimensional fractional diffusion-wave equation. Applied Mathematics and Computation, 349(??):148–157, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311093.

Ali:2017:SAE

[AIG17] Liaqat Ali, Saeed Islam, and Taza Gul. A simple algorithm for exact solutions of systems of linear and nonlinear integro-differential equations. Applied Mathematics and Computation, 307(??):311–320, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301844.

Anco:2015:SGE

[AIIK15] Stephen Anco, Nail H. Ibragimov, Ksenia V. Imamutdinova, and Elena N. Karimova. Solutions of gasdynamic equations associated with classical and new conservation laws. Applied Mathematics and Computation, 268(??):52–58, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004725.

Agrawal:2015:APB

[AIK15] P. N. Agrawal, Nurhayat Ispir, and Arun Kajla. Approximation properties of Bézier-summation-integral type operators based on Polya–Bernstein functions. Applied Mathematics and Computation, 259(??):533–539, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003082.

Abel:2015:DVO

[AIP15] Ulrich Abel, Mircea Ivan, and Radu Paltanea. The Durrmeyer variant of an operator defined by D. D. Stancu. Applied Mathematics and Computation, 259(??):116–123, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001952.

AskariSichani:2015:IMI

[AJ15] Omid AskariSichani and Mahdi Jalili. Influence maximization of informed agents in social networks. Applied Mathematics and Computation, 254(??):229–239, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314018001.

Agrawal:2017:QTF

[AJ17] Manish Agrawal and C. S. Jog. A quadratic time finite element method for nonlinear elastodynamics within the context of hybrid finite elements. Applied Mathematics and Computation, 305(??):203–220, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300851.

Ahmad:2018:EIM

[AJC^{+}18] Fayyaz Ahmad, T. S. Jang, Juan A. Carrasco, Shafiq Ur Rehman, Zulfiqar Ali, and Nukhaze Ali. An efficient iterative method for computing deflections of Bernoulli– Euler–von Karman beams on a nonlinear elastic foundation. Applied Mathematics and Computation, 334(??):269–287, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830211X.

Ahmed:2018:LPS

[AJMN18] Naveed Ahmed, Volker John, Gunar Matthies, and Julia Novo. A local projection stabilization/continuous Galerkin-Petrov method for incompressible flow problems. Applied Mathematics and Computation, 333(??):304–324, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302698.

Atanackovic:2018:NLB

[AJP18] Teodor M. Atanacković, Marko Janev, and Stevan Pilipović. Non-linear boundary value problems involving Caputo derivatives of complex fractional order. Applied Mathematics and Computation, 334(??):326–342, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318303448.

Anastassiou:2015:ARM

[AK15a] George A. Anastassiou and Merve Kester. L_{p} approximation with rates by multivariate generalized discrete singular operators. Applied Mathematics and Computation, 265(??):652–666, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006979.

Andras:2015:ELS

[AK15b] Sz. András and J. J. Kolumbán. Existence and localization of solutions for operatorial systems defined on Cartesian product of Fréchet spaces using a new vector version of Krasnoselskii’s cone compressionexpansion theorem. Applied Mathematics and Computation, 265(??):40– 50, August 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315005950.

Argyros:2015:WCC

[AK15c] I. K. Argyros and S. K. Khattri. Weak convergence conditions for the Newton’s method in Banach space using general majorizing sequences. Applied Mathematics and Computation, 263(??):59– 72, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004749.

Atalay:2015:SFC

[AK15d] Gülnur Saffak Atalay and Emin Kasap. Surfaces family with common null asymptotic. Applied Mathematics and Computation, 260(??):135–139, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003811.

Adivar:2016:FTB

[AK16a] Murat Adivar and Halis Can Koyuncuoglu. Floquet theory based on new periodicity concept for hybrid systems involving q-difference equations. Applied Mathematics and Computation, 273(??):1208–1233, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315012059.

Alzer:2016:SUL

[AK16b] Horst Alzer and Man Kam Kwong. Sharp upper and lower bounds for a sine polynomial. Applied Mathematics and Computation, 275(??):81–85, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015222.

Alizadeh:2018:GWW

[AK18a] Yaser Alizadeh and Sandi Klavzar. On graphs whose Wiener complexity equals their order and on Wiener index of asymmetric graphs. Applied Mathematics and Computation, 328(??):113–118, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300602.

Arora:2018:AQH

[AK18b] Shelly Arora and Inderpreet Kaur. Applications of quintic Hermite collocation with time discretization to singularly perturbed problems. Applied Mathematics and Computation, 316(??):409–421, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730591X.

Alhejaili:2019:NSS

[AK19a] Weaam Alhejaili and Chiu-Yen Kao. Numerical studies of the Steklov eigenvalue problem via conformal mappings. Applied Mathematics and Computation, 347(??):785–802, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310221.

Asai:2019:NSO

[AK19b] Y. Asai and P. E. Kloeden. Numerical schemes for ordinary delay differential equations with random noise. Applied Mathematics and Computation, 347(??):306–318, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310075.

Awwal:2019:SMP

[AKA19] Aliyu Muhammed Awwal, Poom Kumam, and Auwal Bala Abubakar. Spectral modified Polak–Ribiére–Polyak projection conjugate gradient method for solving monotone systems of nonlinear equations. Applied Mathematics and Computation, 362(??): Article 124514, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930493X.

Abbasbandy:2015:AOM

[AKAA15] Saeid Abbasbandy, Saeed Kazem, Mohammed S. Alhuthali, and Hamed H. Alsulami. Application of the operational matrix of fractional-order Legendre functions for solving the time-fractional convection-diffusion equation. Applied Mathematics and Computation, 266(??):31–40, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006013.

Aizinger:2015:SSF

[AKK15] Vadym Aizinger, Dmitri Kuzmin, and Lukas Korous. Scale separation in fast hierarchical solvers for discontinuous Galerkin methods. Applied Mathematics and Computation, 266(??):838–849, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006645.

Argyros:2017:HOD

[AKKB17] Ioannis K. Argyros, Munish Kansal, Vinay Kanwar, and Sugandha Bajaj. Higher-order derivativefree families of Chebyshev– Halley type methods with or without memory for solving nonlinear equations. Applied Mathematics and Computation, 315(??):224–245, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730512X.

Aggul:2019:TAC

[AKL19] Mustafa Aggul, Songul Kaya, and Alexander E. Labovsky. Two approaches to creating a turbulence model with increased temporal accuracy. Applied Mathematics and Computation, 358(??):25–36, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302589.

Abbas:2015:TIT

[AKR15] Ibrahim A. Abbas, Rajneesh Kumar, and Leena Rani. Thermoelastic interaction in a thermally conducting cubic crystal subjected to ramptype heating. Applied Mathematics and Computation, 254(??):360–369, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031401772X.

Abushammala:2015:NFP

[AKS15] Mariam Abushammala, S. A. Khuri, and A. Sayfy. A novel fixed point iteration method for the solution of third order boundary value problems. Applied Mathematics and Computation, 271(??):131–141, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315012102.

Aktas:2015:PDF

[Akt15] Rabia Aktas. On parameter derivatives of a family of polynomials in two variables. Applied Mathematics and Computation, 256(??):769–777, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001010.

Alam:2019:TSF

[AKT19] Muntasir Alam, Kazuki Kuga, and Jun Tanimoto. Three-strategy and fourstrategy model of vaccination game introducing an intermediate protecting measure. Applied Mathematics and Computation, 346(??):408–422, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308816.

Alvarez:2015:IIM

[ÁL15a] Juan Ruiz Álvarez and Zhilin Li. The immersed interface method for axissymmetric problems and application to the Hele– Shaw flow. Applied Mathematics and Computation, 264(??):179–197, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315004506.

Alwan:2015:SIS

[AL15b] Mohamad S. Alwan and Xinzhi Liu. Stability and input-to-state stability for stochastic systems and applications. Applied Mathematics and Computation, 268(??):450–461, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008516.

Abdallaoui:2018:MIT

[AL18a] Athmane Abdallaoui and Keddour Lemrabet. Mechanical impedance of a thin layer in asymmetric elasticity. Applied Mathematics and Computation, 316(??):467–479, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305969.

Ahmad:2018:ESS

[AL18b] Bashir Ahmad and Rodica Luca. Existence of solutions for sequential fractional integro-differential equations and inclusions with nonlocal boundary conditions. Applied Mathematics and Computation, 339(??):507–515, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305836.

Akal:2018:NPA

[AL18c] C. Akal and A. Lukashov.

Newton–Padé approximations for multivariate functions. Applied Mathematics and Computation, 334(??):367–374, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303515.

Antoine:2018:MPT

[AL18d] X. Antoine and E. Lorin.

Multilevel preconditioning technique for Schwarz waveform relaxation domain decomposition method for real- and imaginarytime nonlinear Schrödinger equation. Applied Mathematics and Computation, 336(??):403–417, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304016.

Alfaro:2019:CIM

[AL19] Carlos A. Alfaro and Jephian C.-H. Lin. Critical ideals, minimum rank and zero forcing number. Applied Mathematics and Computation, 358(??):305–313, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319303315.

Aguirre-Lopez:2017:PMA

[ALDHA^{+}17] Mario A. Aguirre-López, O. Díaz-Hernández, F-Javier Almaguer, Javier Morales-Castillo, and Gerardo J. Escalera Santos. A phenomenological model for the aerodynamics of the knuckleball. Applied Mathematics and Computation, 311(??):58–65, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302898.

Alfuraidan:2017:TAW

[Alf17] M. R. Alfuraidan. Topological aspects of weighted graphs with application to fixed point theory. Applied Mathematics and Computation, 314(??):287–292, December 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317304459.

Alikhanov:2015:NMS

[Ali15] Anatoly A. Alikhanov.

Numerical methods of solutions of boundary value problems for the multiterm variable-distributed order diffusion equation. Applied Mathematics and Computation, 268(??):12– 22, October 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315008267.

Alkharashi:2019:MTV

[Alk19] Sameh A. Alkharashi. A model of two viscoelastic liquid films traveling down in an inclined electrified channel. Applied Mathematics and Computation, 355(??):553–575, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301985.

Abouagwa:2018:CAS

[ALL18a] Mahmoud Abouagwa, Jicheng Liu, and Ji Li. Carathéodory approximations and stability of solutions to non-Lipschitz stochastic fractional differential equations of Itô–Doob type. Applied Mathematics and Computation, 329(??): 143–153, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318300997.

Allaoui:2018:RPI

[All18b] Mostafa Allaoui. Robin problems involving the p(x)-Laplacian. Applied Mathematics and Computation, 332(??):457–468, September 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031830225X.

Aguirre-Lopez:2018:TRS

[ALMCDH^{+}18] Mario A. Aguirre-López, Javier Morales-Castillo, O. Díaz-Hernández, Gerardo J. Escalera Santos, and F-Javier Almaguer. Trajectories reconstruction of spinning baseball pitches by three-point-based algorithm. Applied Mathematics and Computation, 319(??):2–12, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300243.

Area:2016:QPP

[ALN16] Iván Area, Jorge Losada, and Juan J. Nieto. On quasi-periodic properties of fractional sums and fractional differences of periodic functions. Applied Mathematics and Computation, 273(??):190–200, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501320X.

Alvarez:2015:WPA

[ALP15] Edgardo Alvarez, Carlos Lizama, and Rodrigo Ponce. Weighted pseudo antiperiodic solutions for fractional integrodifferential equations in Banach spaces. Applied Mathematics and Computation, 259(??):164–172, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002350.

Ahn:2018:IPM

[ALP18] Jeongho Ahn, Seulip Lee, and Eun-Jae Park. C^{0} interior penalty methods for a dynamic nonlinear beam model. Applied Mathematics and Computation, 339(??):675–684, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306088.

Al-Lail:2015:FTR

[ALQ15] Mohammed H. Al-Lail and Asghar Qadir. Fourier transform representation of the generalized hypergeometric functions with applications to the confluent and Gauss hypergeometric functions. Applied Mathematics and Computation, 263(??):392–397, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005457.

Alzalg:2015:VBD

[Alz15] Baha Alzalg. Volumetric barrier decomposition algorithms for stochastic quadratic second-order cone programming. Applied Mathematics and Computation, 265(??): 494–508, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031500630X.

An:2016:EET

[ALZ16] Rong An, Yuan Li, and Yuqing Zhang. Error estimates of two-level finite element method for Smagorinsky model. Applied Mathematics and Computation, 274(??): 786–800, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031501543X.

Ahmad:2015:SAS

[AM15a] Iftikhar Ahmad and Areej Mukhtar. Stochastic approach for the solution of multi-pantograph differential equation arising in cell-growth model. Applied Mathematics and Computation, 261(??):360–372, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004270.

Ahmed:2015:SAS

[AM15b] S. Reaz Ahmed and Partha Modak. Stress analysis of symmetric and anti-symmetric discretely stiffened laminated cantilever beams using displacement-potential field. Applied Mathematics and Computation, 258(??):465–482, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001174.

Akhavan:2015:IPG

[AM15c] S. Akhavan and K. Maleknejad. Improving Petrov– Galerkin elements via Chebyshev polynomials and solving Fredholm integral equation of the second kind by them. Applied Mathematics and Computation, 271(??):352–364, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315012096.

Aprahamian:2015:PAO

[AM15d] Hrayer Aprahamian and Bacel Maddah. Pricing Asian options via compound gamma and orthogonal polynomials. Applied Mathematics and Computation, 264(??):21–43, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004993.

Argyros:2015:EAS

[AM15e] Ioannis K. Argyros and Á. Alberto Magreñán. Expanding the applicability of the secant method under weaker conditions. Applied Mathematics and Computation, 266(??):1000–1012, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315008188.

Argyros:2015:COF

[AM15f] Ioannis K. Argyros and Á. Alberto Magreñán. On the convergence of an optimal fourth-order family of methods and its dynamics. Applied Mathematics and Computation, 252(??):336–346, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016063.

Argyros:2015:CIT

[AM15g] Ioannis Konstantinos Argyros and Ángel Alberto Magreñán. On the convergence of inexact two-point Newton-like methods on Banach spaces. Applied Mathematics and Computation, 265(??):893–902, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007584.

Asgari:2015:SCN

[AM15h] Mohammad Sadegh Asgari and Baharak Mousavi. Solving a class of nonlinear matrix equations via the coupled fixed point theorem. Applied Mathematics and Computation, 259(??):364–373, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002374.

Amirfakhrian:2016:APC

[AM16] M. Amirfakhrian and H. Mafikandi. Approximation of parametric curves by Moving Least Squares method. Applied Mathematics and Computation, 283(??):290–298, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301503.

Argyros:2017:EAL

[AM17] Ioannis K. Argyros and Á. Alberto Magreñán. Extending the applicability of the local and semilocal convergence of Newton’s method. Applied Mathematics and Computation, 292(??):349–355, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304428.

Adams:2018:SSC

[AM18a] Ronald Adams and Stefan C. Mancas. Stability of solitary and cnoidal traveling wave solutions for a fifth order Korteweg–de Vries equation. Applied Mathematics and Computation, 321(??):745–751, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307786.

Almeida:2018:ELL

[AM18b] Ricardo Almeida and M. Luísa Morgado. The Euler–Lagrange and Legendre equations for functionals involving distributedorder fractional derivatives. Applied Mathematics and Computation, 331(??):394–403, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301875.

Argyros:2018:PNP

[AM18c] I. K. Argyros and St. Maruster. Predetermining the number of periodic steps in multi-step Newton-like methods for solving equations and systems of equations. Applied Mathematics and Computation, 329(??):420–431, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301267.

Arnas:2018:NFI

[AM18d] David Arnas and Daniele Mortari. Nonlinear function inversion using kvector. Applied Mathematics and Computation, 320(??):754–768, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306987.

Alyazidi:2019:LCD

[AM19] Nezar M. Alyazidi and Magdi S. Mahmoud. On LQG control design for network systems with/ without acknowledgments using a particle filtering technology. Applied Mathematics and Computation, 359(??):52–70, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319302887.

Alonso-Meijide:2015:CBC

[AMCMFJ15] J. M. Alonso-Meijide, B. Casas-Méndez, and M. G. Fiestras-Janeiro. Computing Banzhaf–Coleman and Shapley–Shubik power indices with incompatible players. Applied Mathematics and Computation, 252(??):377–387, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016658.

Ammar:2015:TAQ

[Amm15] Sherif I. Ammar. Transient analysis of an M/M/1 queue with impatient behavior and multiple vacations. Applied Mathematics and Computation, 260 (??):97–105, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031500380X.

Ates:2019:DHF

[AMN19] Osman Ates, Marian Ioan Munteanu, and Ana Irina Nistor. Dynamics on S^{3} and the Hopf fibration. Applied Mathematics and Computation, 347(??):429–441, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310014.

Al-Mdallal:2018:FOL

[AMO18] Qasem M. Al-Mdallal and Ahmed S. Abu Omer. Fractional-order Legendrecollocation method for solving fractional initial value problems. Applied Mathematics and Computation, 321(??):74–84, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307014.

Ahmad:2015:ERC

[AN15] Bashir Ahmad and Sotiris K. Ntouyas. Existence results for a coupled system of Caputo type sequential fractional differential equations with nonlocal integral boundary conditions. Applied Mathematics and Computation, 266(??):615–622, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500747X.

Arbabi:2017:TDH

[AND17] Somayeh Arbabi, Akbar Nazari, and Mohammad Taghi Darvishi. A two-dimensional Haar wavelets method for solving systems of PDEs. Applied Mathematics and Computation, 292(??):33– 46, January 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316304702.

Andrzejczak:2018:SRK

[And18]  Grzegorz  Andrzejczak.

Spline reproducing kernels on R and error bounds for piecewise smooth LBV problems. Applied Mathematics and Computation, 320(??):27–44, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317306422.

Azarnavid:2019:IVB

[ANEP19] Babak Azarnavid, Mohammad Nabati, Mahdi Emamjome, and Kourosh Parand. Imposing various boundary conditions on positive definite kernels. Applied Mathematics and Computation, 361(??):453–465, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304540.

Akbari:2017:INL

[ANGC17] M. R. Akbari, M. Nimafar, D. D. Ganji, and H. Karimi Chalmiani. Investigation on non-linear vibration in arched beam for bridges construction via AGM method. Applied Mathematics and Computation, 298(??):95–110, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630666X.

Anonymous:2015:P

[Ano15]  Anonymous.  Preface.

Applied Mathematics and Computation, 267(??):1–2, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315009455.

Anonymous:2016:P

[Ano16]  Anonymous.  Preface.

Applied Mathematics and Computation, 272 (part 1) (??):1, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315014423.

Anonymous:2018:EBa

[Ano18a]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 322 (??):ii, April 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317308639.

Anonymous:2018:EBb

[Ano18b]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 323(??):ii, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308986.

Anonymous:2018:EBc

[Ano18c]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 324 (??):ii, May 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031830016X.

Anonymous:2018:EBd

[Ano18d]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 325(??):ii, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300407.

Anonymous:2018:EBe

[Ano18e]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 326 (??):ii, June 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318300857.

Anonymous:2018:EBf

[Ano18f]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 327(??):ii, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301085.

Anonymous:2018:EBg

[Ano18g]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 328 (??):ii, July 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318301425.

Anonymous:2018:EBh

[Ano18h]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 329(??):ii, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301978.

Anonymous:2018:EBi

[Ano18i]  Anonymous.  Edito-

rial Board. Applied Mathematics and Computation, 330(??):ii, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302479.

Anonymous:2018:EBj

[Ano18j]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 331(??):ii, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303138.

Anonymous:2018:EBk

[Ano18k]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 332(??):ii, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303345.

Anonymous:2018:EBl

[Ano18l]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 333(??):ii, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303783.

Anonymous:2018:EBm

[Ano18m]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 334(??):ii, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304284.

Anonymous:2018:EBn

[Ano18n]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 335(??):ii, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304594.

Anonymous:2018:EBo

[Ano18o]  Anonymous.  Editorial

board. Applied Mathematics and Computation, 336(??):ii, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304831.

Anonymous:2018:EBp

[Ano18p]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 337(??):1–634, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305617.

Anonymous:2018:EBq

[Ano18q]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 338(??):1–902, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318306374.

Anonymous:2018:EBr

[Ano18r]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 339(??):1–916, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318307689.

Anonymous:2019:EBa

[Ano19a]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 340(??):ii, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308075.

Anonymous:2019:EBb

[Ano19b]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 342(??):ii, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308762.

Anonymous:2019:EBc

[Ano19c]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 343(??):ii, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309366.

Anonymous:2019:EBd

[Ano19d]  Anonymous.  Edito-

rial Board. Applied Mathematics and Computation, 344–345(??):ii, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309640.

Anonymous:2019:EBe

[Ano19e]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 346 (??):ii, April 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318310580.

Anonymous:2019:EBf

[Ano19f]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 347(??):ii, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310944.

Anonymous:2019:EBg

[Ano19g]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 348 (??):ii, May 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319300050.

Anonymous:2019:EBh

[Ano19h]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 349(??):ii, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300319300566.

Anonymous:2019:EBi

[Ano19i]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 350 (??):ii, June 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319300967.

Anonymous:2019:EBj

[Ano19j]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 351(??):ii, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301249.

Anonymous:2019:EBk

[Ano19k]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 352 (??):ii, July 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319301729.

Anonymous:2019:EBl

[Ano19l]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 353(??):ii, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302139.

Anonymous:2019:EBm

[Ano19m]  Anonymous.  Edito-

rial Board. Applied Mathematics and Computation, 354(??):ii, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302395.

Anonymous:2019:EBn

[Ano19n]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 355(??):ii, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302929.

Anonymous:2019:EBo

[Ano19o]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 356(??):ii, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303182.

Anonymous:2019:EBp

[Ano19p]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 357(??):ii, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303686.

Anonymous:2019:EBq

[Ano19q]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 358(??):ii, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930400X.

Anonymous:2019:EBr

[Ano19r]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 359(??):ii, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304291.

Anonymous:2019:EBs

[Ano19s]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 360(??):ii, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304631.

Anonymous:2019:EBt

[Ano19t]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 361(??):ii, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305624.

Anonymous:2019:EBu

[Ano19u]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 362 (??):Article 124650, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319306423.

Anonymous:2019:EBv

[Ano19v]  Anonymous.  Editorial

Board. Applied Mathematics and Computation, 363 (??):Article 124705, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306976.

Anonymous:2019:PNI

[Ano19w]  Anonymous.  Publisher

note: Introducing article numbering to Applied Mathematics and Computation. Applied Mathematics and Computation, 362 (??):Article 124651, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306435.

Agliari:2016:NDC

[ANP16] A. Agliari, A. K. Naimzada, and N. Pecora. Nonlinear dynamics of a Cournot duopoly game with differentiated products. Applied Mathematics and Computation, 281(??):1–15, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300455.

Ahmad:2016:IFI

[ANT^{+}16] Bashir Ahmad, Sotiris K.

Ntouyas, Jessada Tariboon, Ahmed Alsaedi, and Hamed H. Alsulami. Impulsive fractional q-integro-difference equations with separated boundary conditions. Applied Mathematics and Computation, 281(??): 199–213, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316300601.

Anthonys:2018:MMI

[Ant18] Gehan Anthonys. Mathematical model to investigate the behaviour of the systems of ferromagnetic particles under the magnetic fields. Applied Mathematics and Computation, 320(??):654–676, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306860.

Afzal:2016:ADD

[ANU16] Muhammad Afzal, Rab Nawaz, and Aasim Ullah. Attenuation of dissipative device involving coupled wave scattering and change in material properties. Applied Mathematics and Computation, 290(??):154–163, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303721.

Agarwal:2015:DLV

[AÖ15a] Ravi P. Agarwal and Abdullah Özbekler. Disconjugacy via Lyapunov and Vallée–Poussin type inequalities for forced differential equations. Applied Mathematics and Computation, 265(??):456–468, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006554.

Al-Omari:2015:ESD

[AO15b] Jafar Fawzi M. Al-Omari.

The effect of state dependent delay and harvesting on a stage-structured predator–prey model. Applied Mathematics and Computation, 271(??): 142–153, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S009630031501200X.

Atouani:2015:CCD

[AO15c] Noureddine Atouani and Khaled Omrani. On the convergence of conservative difference schemes for the 2D generalized Rosenau–Korteweg de Vries equation. Applied Mathematics and Computation, 250(??):832–847, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014817.

Agarwal:2016:LTI

[AÖ16] Ravi P. Agarwal and Abdullah Özbekler. Lyapunov type inequalities for second order forced mixed nonlinear impulsive differential equations. Applied Mathematics and Computation, 282(??):216–225, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301187.

Anderson:2019:HUS

[AO19] Douglas R. Anderson and Masakazu Onitsuka. Hyers–Ulam stability for a discrete time scale with two step sizes. Applied Mathematics and Computation, 344–345(??):128– 140, March 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318308804.

Agarwal:2017:MIT

[AOH17] Ravi Agarwal, D. O’Regan, and S. Hristova. Monotone iterative technique for the initial value problem for differential equations with non-instantaneous impulses. Applied Mathematics and Computation, 298(??):45–56, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300316306075.

Andova:2017:LFW

[AOS17] Vesna Andova, Damir Orlić, and Riste Skrekovski. Leapfrog fullerenes and Wiener index. Applied Mathematics and Computation, 309(??):281–288, September 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317302278.

Adan:2019:HSP

[AOS19] Alí Guzmán Adán, Heikki Orelma, and Franciscus Sommen. Hypermonogenic solutions and plane waves of the Dirac operator in R^{p} × R^{q}. Applied Mathematics and Computation, 346(??):1– 14, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308452.

Arraras:2015:IAT

[AP15] A. Arrarás and L. Portero. Improved accuracy for time-splitting methods for the numerical solution of parabolic equations. Applied Mathematics and Computation, 267(??):294–303, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003872.

Arraras:2018:DMF

[AP18a] A. Arrarás and L. Portero. Decoupling mixed finite elements on hierarchical triangular grids for parabolic problems. Applied Mathematics and Computation, 319(??):662–680, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305039.

Asadbeigi:2018:NER

[AP18b] Leila Asadbeigi and Mahmoud Paripour. A note on extended reduced rank-two Abaffian update schemes in the ABStype methods. Applied Mathematics and Computation, 326(??):105–107, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300195.

Alvandi:2019:CRK

[AP19] Azizallah Alvandi and Mahmoud Paripour. The combined reproducing kernel method and Taylor series for handling nonlinear Volterra integrodifferential equations with derivative type kernel. Applied Mathematics and Computation, 355(??): 151–160, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319301274.

Aydogan:2015:HMR

[APK15] Melike Aydoğan, Yasar Polatoğlu, and Yasemin Kahramaner. Harmonic mappings related to the m-fold starlike functions. Applied Mathematics and Computation, 267(??): 805–809, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314013824.

Alvarez-Pardo:2015:WPA

[APL15] Edgardo Alvarez-Pardo and Carlos Lizama. Weighted pseudo almost automorphic mild solutions for two-term fractional order differential equations. Applied Mathematics and Computation, 271(??): 154–167, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315012138.

Amirkhani:2018:NMD

[APMM18] Abdollah Amirkhani, Elpiniki I. Papageorgiou, Mohammad R. Mosavi, and Karim Mohammadi. A novel medical decision support system based on fuzzy cognitive maps enhanced by intuitive and learning capabilities for modeling uncertainty. Applied Mathematics and Computation, 337(??):545–561, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304417.

Allen-Perkins:2017:TWD

[APPE17]  Alfonso  Allen-Perkins,

Juan Manuel Pastor, and Ernesto Estrada. Twowalks degree assortativity in graphs and networks. Applied Mathematics and Computation, 311(??):262–271, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303235.

Alonso:2016:ASS

[APS16] P. Alonso, J. M. Peña, and M. L. Serrano. Almost strictly sign regular matrices and Neville elimination with two-determinant pivoting. Applied Mathematics and Computation, 289(??):426–434, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303514.

Aljinovic:2015:WQR

[APTS15] A. Aglić Aljinović, J. Pecarić, and S. Tipurić-Spuzević. Weighted quadrature rules via Grüss type inequalities for weighted L^{p} spaces. Applied Mathematics and Computation, 264(??):1–12, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004312.

Abrarov:2015:SIC

[AQ15] S. M. Abrarov and B. M. Quine. Sampling by incomplete cosine expansion of the sinc function: Application to the Voigt/complex error function. Applied Mathematics and Computation, 258(??):425–435, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001046.

Abrarov:2018:RAD

[AQ18] Sanjar M. Abrarov and Brendan M. Quine. A rational approximation of the Dawson’s integral for efficient computation of the complex error function. Applied Mathematics and Computation, 321(??):526–543, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317307312.

Aliyev:2015:OPE

[AR15] Ziyatkhan S. Aliyev and Humay Sh. Rzayeva. Oscillation properties for the equation of the relativistic quantum theory. Applied Mathematics and Computation, 271(??):308–316, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012163.

Adivar:2016:QAN

[AR16a] Murat Adivar and Youssef N.

Raffoul. Qualitative analysis of nonlinear Volterra integral equations on time scales using resolvent and Lyapunov functionals. Applied Mathematics and Computation, 273(??):258–266, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013259.

Argyros:2016:KMS

[AR16b] Ioannis K. Argyros and Hongmin Ren. On the Kurchatov method for solving equations under weak conditions. Applied Mathematics and Computation, 273(??):98–113, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501303X.

Arandiga:2016:NAM

[Arà16] F. Aràndiga. A nonlinear algorithm for monotone piecewise bicubic interpolation. Applied Mathematics and Computation, 272 (part 1)(??):100–113, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010826.

Aguirre-Ramos:2018:BVS

[ARACCA^{+}18] Hugo Aguirre-Ramos, Juan Gabriel Avina-Cervantes, Ivan Cruz-Aceves, José Ruiz-Pinales, and Sergio Ledesma. Blood vessel segmentation in retinal fundus images using Gabor filters, fractional derivatives, and expectation maximization. Applied Mathematics and Computation, 339(??):556–567, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318306234.

Ali:2016:BID

[ARB16] Akbar Ali, Zahid Raza, and Akhlaq Ahmad Bhatti. Bond incident degree (BID) indices of polyomino chains: a unified approach. Applied Mathematics and Computation, 287–288(??):28–37, September 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316302624.

Ahmadian:2019:SAS

[ARB19] D. Ahmadian, O. Farkhondeh Rouz, and L. V. Ballestra. Stability analysis of split-step θ-Milstein method for a class of ndimensional stochastic differential equations. Applied Mathematics and Computation, 348(??): 413–424, May 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318309068.

Ariza-Ruiz:2015:IAC

[ARGF15] D. Ariza-Ruiz and J. Garcia-Falset. Iterative approximation to a coincidence point of two mappings. Applied Mathematics and Computation, 259(??):762–776, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500301X.

Arteaga:2015:SBI

[ARK15] Andrea Arteaga, Daniel Ruprecht, and Rolf Krause. A stencil-based implementation of Parareal in the C++ domain specific embedded language STELLA. Applied Mathematics and Computation, 267(??):727–741, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017093.

Al-Refai:2015:MPM

[ARL15] Mohammed Al-Refai and Yuri Luchko. Maximum principle for the multiterm time-fractional diffusion equations with the Riemann–Liouville fractional derivatives. Applied Mathematics and Computation, 257(??):40–51, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300314017883.

Armerin:2019:OSK

[Arm19]  F. Armerin.  Optimal

stopping of a killed exponentially growing process. Applied Mathematics and Computation, 353(??):208–214, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301043.

Antoci:2016:ODT

[ARR16] Angelo Antoci, Stefania Ragni, and Paolo Russu. Optimal dynamics in a two-sector model with natural resources and foreign direct investments. Applied Mathematics and Computation, 273(??):290–307, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012333.

Asafuddoula:2015:ISA

[ARS15] Md Asafuddoula, Tapabrata Ray, and Ruhul Sarker. An Improved Self-Adaptive Constraint Sequencing approach for constrained optimization problems. Applied Mathematics and Computation, 253(??):23– 39, February 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314016865.

Alves:2015:AAC

[ARSV15] Margareth S. Alves, Jaime E. Muñoz Rivera, Mauricio Sepúlveda, and Octavio Vera. About analyticity for the coupled system of linear thermoviscoelastic equations. Applied Mathematics and Computation, 270(??):943–952, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011480.

AlBasir:2018:RMC

[ARV18] Fahad Al Basir, Santanu Ray, and Ezio Venturino. Role of media coverage and delay in controlling infectious diseases: a mathematical model. Applied Mathematics and Computation, 337(??):360–371, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S009630031830451X.

Arora:2015:ACS

[AS15] Urvashi Arora and N. Sukavanam. Approximate controllability of second order semilinear stochastic system with nonlocal conditions. Applied Mathematics and Computation, 258(??):111–119, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001502.

Abro:2019:NTE

[AS19a] Hameer Akhtar Abro and Muhammad Mujtaba Shaikh. A new timeefficient and convergent nonlinear solver. Applied Mathematics and Computation, 355(??):516–536, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930205X.

Aydogan:2019:RSV

[AS19b] S. M. Aydogan and F. M. Sakar. Radius of starlikeness of pvalent λ-fractional operator. Applied Mathematics and Computation, 357(??):374–378, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310415.

Al-Smadi:2019:CAS

[ASA19] Mohammed Al-Smadi and Omar Abu Arqub. Computational algorithm for solving Fredholm timefractional partial integrodifferential equations of Dirichlet functions type with error estimates. Applied Mathematics and Computation, 342(??):280–294, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307926.

Al-Smadi:2016:NIS

[ASASM16]  Mohammed  Al-Smadi,

Omar Abu Arqub, Nabil Shawagfeh, and Shaher Momani. Numerical investigations for systems of second-order periodic boundary value problems using reproducing kernel method. Applied Mathematics and Computation, 291(??):137–148, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303733.

Ahmad:2017:HOD

[ASHSC17] F. Ahmad, F. Soleymani, F. Khaksar Haghani, and S. Serra-Capizzano. Higher order derivativefree iterative methods with and without memory for systems of nonlinear equations. Applied Mathematics and Computation, 314(??):199–211, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304630.

Askar:2018:TSG

[Ask18]  S. S. Askar.  Tripoly

Stackelberg game model: One leader versus two followers. Applied Mathematics and Computation, 328(??):301–311, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300626.

Ahmad:2019:CSI

[ASKK19] Sanwar Ahmad, Thilo Strauss, Shyla Kupis, and Taufiquar Khan. Comparison of statistical inversion with iteratively regularized Gauss Newton method for image reconstruction in electrical impedance tomography. Applied Mathematics and Computation, 358(??):436–448, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302711.

Ayub:2016:SIS

[ASM16] Muhammad Ayub, Sadia Sadique, and F. M. Mahomed. Singular invariant structures for Lie algebras admitted by a system of second-order ODEs. Applied Mathematics and Computation, 281(??):137–147, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300042.

Alt:2016:RIE

[ASS16] Walter Alt, Christopher Schneider, and Martin Seydenschwanz. Regularization and implicit Euler discretization of linearquadratic optimal control problems with bangbang solutions. Applied Mathematics and Computation, 287–288 (??):104–124, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302788.

Allouch:2018:NSW

[AST18] C. Allouch, D. Sbibih, and M. Tahrichi. Numerical solutions of weakly singular Hammerstein integral equations. Applied Mathematics and Computation, 329(??):118–128, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300699.

Arregui:2017:PMN

[ASV17] Iñigo Arregui, Beatriz Salvador, and Carlos Vázquez. PDE models and numerical methods for total value adjustment in European and American options with counterparty risk. Applied Mathematics and Computation, 308(??):31–53, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301790.

Andelic:2018:FAC

[ASZD18] M. Andelić, S. K. Simić, D. Zivković, and E. Ć. Dolićanin. Fast algorithms for computing the characteristic polynomial of threshold and chain graphs. Applied Mathematics and Computation, 332(??):329–337, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301899.

Abkar:2015:NAT

[AT15a] A. Abkar and M. Tavakkoli. A new algorithm for two finite families of demicontractive mappings and equilibrium problems. Applied Mathematics and Computation, 266(??):491–500, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500702X.

Almeida:2015:CHT

[AT15b] Ricardo Almeida and Delfim F. M. Torres. Computing Hadamard type operators of variable fractional order. Applied Mathematics and Computation, 257(??):74–88, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017251.

Atangana:2016:NFD

[Ata16] Abdon Atangana. On the new fractional derivative and application to nonlinear Fisher’s reactiondiffusion equation. Applied Mathematics and Computation, 273(??):948–956, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013533.

Akbari:2016:IRH

[ATK^{+}16] Omid Ali Akbari, Davood Toghraie, Arash Karimipour, Mohammad Reza Safaei, Marjan Goodarzi, Habibollah Alipour, and Mahidzal Dahari. Investigation of rib’s height effect on heat transfer and flow parameters of laminar water-Al_{2}O_{3} nanofluid in a rib-microchannel. Applied Mathematics and Computation, 290(??):135–153, November 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S009630031630371X.

Anggriani:2019:ERS

[ATN^{+}19] N. Anggriani, H. Tasman, M. Z. Ndii, A. K. Supriatna, E. Soewono, and E. Siregar. The effect of reinfection with the same serotype on dengue transmission dynamics. Applied Mathematics and Computation, 349(??):62–80, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310695.

Agryzkov:2016:NHN

[ATV16] Taras Agryzkov, Leandro Tortosa, and Jose F. Vicent. New highlights and a new centrality measure based on the Adapted PageRank Algorithm for urban networks. Applied Mathematics and Computation, 291(??):14–29, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304076.

Agryzkov:2018:ACD

[ATV18] Taras Agryzkov, Leandro Tortosa, and Jose F. Vicent. An algorithm to compute data diversity index in spatial networks. Applied Mathematics and Computation, 337(??):54–62, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303941.

Agryzkov:2019:VCF

[ATV19] Taras Agryzkov, Leandro Tortosa, and Jose F. Vicent. A variant of the current flow betweenness centrality and its application in urban networks. Applied Mathematics and Computation, 347(??):600–615, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310063.

Aye:2019:NVD

[ATX19] Win Min Aye, Tao Tian, and Liming Xiong. Number of vertices of degree three in spanning 3-trees in square graphs. Applied Mathematics and Computation, 357(??):258–262, September 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S009630031930270X.

Alaidarous:2015:CCS

[AU15] Eman S. Alaidarous and Malik Zaka Ullah. Construction of a convergent scheme for finding matrix sign function. Applied Mathematics and Computation, 260(??):242–248, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500363X.

Anjam:2015:FMA

[AV15] I. Anjam and J. Valdman. Fast MATLAB assembly of FEM matrices in 2D and 3D: Edge elements. Applied Mathematics and Computation, 267(??):252–263, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004191.

Alves:2018:AMF

[AV18] Carlos J. S. Alves and Svilen S. Valtchev. On the application of the method of fundamental solutions to boundary value problems with jump discontinuities. Applied Mathematics and Computation, 320(??):61–74, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306392.

Anand:2019:CGP

[AVC^{+}19] Bijo S. Anand, Ullas Chandran S. V., Manoj Changat, Sandi Klavzar, and Elias John Thomas. Characterization of general position sets and its applications to cographs and bipartite graphs. Applied Mathematics and Computation, 359(??):84– 89, October 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319303522.

Avila-Vales:2015:SHB

[AVCCGAVDL15] Eric Avila-Vales, Noé Chan-Chí, Gerardo E. García-Almeida, and Cruz Vargas-De-León. Stability and Hopf bifurcation in a delayed viral infection model with mitosis transmission. Applied Mathematics and Computation, 259(??):293–312, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002416.

Awais:2015:ANI

[Awa15] M. Awais. Applications of the numerical inversion of the Laplace transform to unsteady problems of the third grade fluid. Applied Mathematics and Computation, 250(??):228–234, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014842.

Abdel-wahed:2015:FHT

[AwEE15] M. S. Abdel-wahed, E. M. A. Elbashbeshy, and T. G. Emam. Flow and heat transfer over a moving surface with non-linear velocity and variable thickness in a nanofluids in the presence of Brownian motion. Applied Mathematics and Computation, 254(??):49–62, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017482.

Ali:2016:SCD

[AY16a] M. Syed Ali and J. Yogambigai. Synchronization of complex dynamical networks with hybrid coupling delays on time scales by handling multitude Kronecker product terms. Applied Mathematics and Computation, 291(??):244–258, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630426X.

Arandiga:2016:NCC

[AY16b] Francesc Aràndiga and Dionisio F. Yáñez. Nonconsistent cell-average multiresolution operators with application to image processing. Applied Mathematics and Computation, 272 (part 1)(??):208–222, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011303.

Abid:2015:DDI

[AYAA^{+}15] Walid Abid, Radouane Yafia, M. A. Aziz-Alaoui, Habib Bouhafa, and Azgal Abichou. Diffusion driven instability and Hopf bifurcation in spatial predator– prey model on a circular domain. Applied Mathematics and Computation, 260(??):292–313, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003847.

Akdogan:2019:DFS

[AYD19] Zülfigar Akdogan, Ali Yakar, and Mustafa Demirci. Discontinuous fractional Sturm–Liouville problems with transmission conditions. Applied Mathematics and Computation, 350 (??):1–10, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031831110X.

Al-Zhour:2014:GVS

[AZ14] Zeyad Abdel Aziz Al-Zhour. The general (vector) solutions of such linear (coupled) matrix fractional differential equations by using Kronecker structures. Applied Mathematics and Computation, 232(??):498–510, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314001167. See

retraction notice [AZ19].

Akgol:2018:ARS

[AZ18] S. Dogru Akgöl and A. Zafer. Asymptotic representation of solutions for second-order impulsive differential equations. Applied Mathematics and Computation, 333(??):53–60, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302674.

AzizAl-Zhour:2019:RNG

[AZ19] Zeyad Abdel AzizAl-Zhour. Retraction notice to “The general (Vector) solutions of such linear (coupled) matrix fractional differential equations by using Kronecker structures” [Applied Mathematics and Computation 232 (2014) 498–510]. Applied Mathematics and Computation, 361(??):889, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304357.

See [AZ14].

Az-Zobi:2015:NAS

[AZAM15] Emad A. Az-Zo’bi, Kamal Al Dawoud, and Mohammad Marashdeh. Numeric-analytic solutions of mixed-type systems of balance laws. Applied Mathematics and Computation, 265(??):133–143, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005901.

Ademiloye:2015:NCE

[AZL15] A. S. Ademiloye, L. W. Zhang, and K. M. Liew. Numerical computation of the elastic and mechanical properties of red blood cell membrane using the higher-order Cauchy– Born rule. Applied Mathematics and Computation, 268(??):334–353, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008528.

Agarwal:2016:SRO

[AZL16] Ravi P. Agarwal, Chenghui Zhang, and Tongxing Li. Some remarks on oscillation of second order neutral differential equations. Applied Mathematics and Computation, 274(??):178–181, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S009630031501440X.

Baxhaku:2017:DAB

[BA17] Behar Baxhaku and Purshottam Narain Agrawal. Degree of approximation for bivariate extension of Chlodowskytype q-Bernstein–Stancu– Kantorovich operators. Applied Mathematics and Computation, 306(??):56– 72, August 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317301030.

Babenko:2016:NMS

[Bab16] Vira Babenko. Numerical methods for solution of Volterra and Fredholm integral equations for functions with values in Lspaces. Applied Mathematics and Computation, 291(??):354–372, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304623.

Baseri:2018:CMF

[BAB18] A. Baseri, S. Abbasbandy, and E. Babolian. A collocation method for fractional diffusion equation in a long time with Chebyshev functions. Applied Mathematics and Computation, 322(??):55–65, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308305.

Baccouch:2015:SPE

[Bac15] Mahboub Baccouch. Superconvergence and a posteriori error estimates of the DG method for scalar hyperbolic problems on Cartesian grids. Applied Mathematics and Computation, 265(??):144–162, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005974.

Bellamine:2015:MCD

[BAE15] Fethi Bellamine, A. Almansoori, and A. Elkamel. Modeling of complex dynamic systems using differential neural networks with the incorporation of a priori knowledge. Applied Mathematics and Computation, 266(??):515–526, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007535.

Bagdasaryan:2015:CII

[Bag15] Armen Bagdasaryan. A combinatorial identity involving gamma function and Pochhammer symbol. Applied Mathematics and Computation, 256(??):125–130, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000223.

Baghani:2019:CRH

[Bag19] Omid Baghani. A correction to “Rationalized Haar wavelet bases to approximate solution of nonlinear Fredholm integral equations with error analysis [Applied Mathematics and Computation 265 (2015) 304–312]”. Applied Mathematics and Computation, 352(??):249–257, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300426. See

[EGB15].

Bayon:2019:OCC

[BAGN^{+}19] L. Bayón, P. Fortuny Ayuso, P. J. García-Nieto, J. M. Grau, and M. M. Ruiz. Optimal control of counterterrorism tactics. Applied Mathematics and Computation, 347(??):477–491, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309962.

Batbold:2019:UTH

[BAK19] Tserendorj Batbold, Laith E. Azar, and Mario Krnić. A unified treatment of Hilbert–Pachpatte-type inequalities for a class of non-homogeneous kernels. Applied Mathematics and Computation, 343(??):167–182, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308269.

Behl:2016:NHE

[BAM16] Ramandeep Behl, Ioannis K. Argyros, and S. S. Motsa. A new highly efficient and optimal family of eighth-order methods for solving nonlinear equations. Applied Mathematics and Computation, 282(??):175–186, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301138.

Banaji:2018:IOC

[Ban18] Murad Banaji. Inheritance of oscillation in chemical reaction networks. Applied Mathematics and Computation, 325(??):191–209, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308718.

Blaya:2015:BVPb

[BÁR15] Ricardo Abreu Blaya, Rafael Ávila Ávila, and Juan Bory Reyes. Boundary value problems with higher order Lipschitz boundary data for polymonogenic functions in fractal domains. Applied Mathematics and Computation, 269(??):802–808, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501067X.

Bhatti:2018:RNM

[BAR18] M. M. Bhatti, M. Ali Abbas, and M. M. Rashidi. A robust numerical method for solving stagnation point flow over a permeable shrinking sheet under the influence of MHD. Applied Mathematics and Computation, 316(??):381–389, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305830.

Bashan:2019:MAN

[Bas19] Ali Bashan. A mixed algorithm for numerical computation of soliton solutions of the coupled KdV equation: Finite difference method and differential quadrature method. Applied Mathematics and Computation, 360(??):42–57, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303753.

Bakhtiari:2018:RKM

[BAV18] P. Bakhtiari, S. Abbasbandy, and R. A. Van Gorder. Reproducing kernel method for the numerical solution of the 1D Swift–Hohenberg equation. Applied Mathematics and Computation, 339(??):113–131, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305563.

Bailey:2015:CCI

[BB15a] D. H. Bailey and J. M.

Borwein. Crandall’s computation of the incomplete gamma function and the Hurwitz zeta function, with applications to Dirichlet L-series. Applied Mathematics and Computation, 268(??):462–477, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008292.

Baran:2015:EAM

[BB15b] M. Baran and L. K. Bieniasz. Experiments with an adaptive multicut-HDMR map generation for slowly varying continuous multivariate functions. Applied Mathematics and Computation, 258(??):206–219, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001599.

Benitez:2015:BCS

[BB15c] R. Benítez and V. J. Bolós.

Blow-up collocation solutions of nonlinear homogeneous Volterra integral equations. Applied Mathematics and Computation, 256(??):754–768, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000879.

Berezansky:2015:SCS

[BB15d] Leonid Berezansky and Elena Braverman. Stability conditions for scalar delay differential equations with a non-delay term. Applied Mathematics and Computation, 250(??):157–164, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014635.

Berezansky:2016:BPD

[BB16] Leonid Berezansky and Elena Braverman. Boundedness and persistence of delay differential equations with mixed nonlinearity. Applied Mathematics and Computation, 279(??):154–169, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300157.

Bauer:2018:SPS

[BB18a] Werner Bauer and Jörn Behrens. A structurepreserving split finite element discretization of the split wave equations. Applied Mathematics and Computation, 325(??):375–400, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308949.

Berhausen:2018:DHP

[BB18b] Sebastian Berhausen and Andrzej Boboń. Determination of high power synchronous generator subtransient reactances based on the waveforms for a steady state two-phase short-circuit. Applied Mathematics and Computation, 319(??):538–550, February 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317304058.

Benacchio:2019:EDM

[BB19a] Tommaso Benacchio and Luca Bonaventura. An extension of DG methods for hyperbolic problems to one-dimensional semiinfinite domains. Applied Mathematics and Computation, 350(??):266–282, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311184.

Berezansky:2019:SLN

[BB19b] Leonid Berezansky and Elena Braverman. On stability of linear neutral differential equations in the Hale form. Applied Mathematics and Computation, 340(??):63–71, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306490.

Barletti:2018:ECM

[BBCI18] L. Barletti, L. Brugnano, G. Frasca Caccia, and F. Iavernaro. Energyconserving methods for the nonlinear Schrödinger equation. Applied Mathematics and Computation, 318(??):3–18, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317302709.

Buyukkilic:2015:ICG

[BBD15] F. Buyukkilic, Z. Ok Bayrakdar, and D. Demirhan. Investigation of cumulative growth process via Fibonacci method and fractional calculus. Applied Mathematics and Computation, 265(??):237–244, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006438.

Banagaaya:2018:IAP

[BBF^{+}18] Nicodemus Banagaaya, Peter Benner, Lihong Feng, Peter Meuris, and Wim Schoenmaker. An index-aware parametric model order reduction method for parameterized quadratic differential– algebraic equations. Applied Mathematics and Computation, 319(??): 409–424, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S009630031730276X.

Bolos:2017:WSD

[BBFJ17] V. J. Bolós, R. Benítez, R. Ferrer, and R. Jammazi. The windowed scalogram difference: a novel wavelet tool for comparing time series. Applied Mathematics and Computation, 312(??):49–65, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303478.

Barrios:2019:NPE

[BBG19] Tomás P. Barrios, Edwin M. Behrens, and María González. New a posteriori error estimator for an stabilized mixed method applied to incompressible fluid flows. Applied Mathematics and Computation, 351(??):37– 47, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300293.

Brambilla:2015:GCA

[BBGR15] P. Brambilla, P. Brivio, A. Guardone, and G. Romanelli. Grid convergence assessment for adaptive grid simulations of normal drop impacts onto liquid films in axi-symmetric and three-dimensional geometries. Applied Mathematics and Computation, 267(??):487–497, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001290.

Bai:2016:GWC

[bBjJ16] Hong bin Bai and Zhi jie Jiang. Generalized weighted composition operators from Zygmund spaces to Bloch–Orlicz type spaces. Applied Mathematics and Computation, 273(??):89–97, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501293X.

Blaszczyk:2015:FOE

[BC15] Tomasz Blaszczyk and Mariusz Ciesielski. Fractional oscillator equation — transformation into integral equation and numerical solution. Applied Mathematics and Computation, 257(??):428–435, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017834.

Brzdek:2016:SFE

[BC16] Janusz Brzdek and Liviu Cadariu. Stability for a family of equations generalizing the equation of p-Wright affine functions. Applied Mathematics and Computation, 276(??):158–171, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015921.

Berriochoa:2015:GPH

[BCD15a] E. Berriochoa, A. Cachafeiro, and J. Díaz. Gibbs phenomenon in the Hermite interpolation on the circle. Applied Mathematics and Computation, 253(??):274–286, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017172.

Burkov:2015:CED

[BCd15b] Andriy Burkov and Brahim Chaib-draa. Computing equilibria in discounted dynamic games. Applied Mathematics and Computation, 269(??):863–884, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500987X.

Belardo:2018:WPC

[BCD18] Francesco Belardo, Matteo Cavaleri, and Alfredo Donno. Wreath product of a complete graph with a cyclic graph: Topological indices and spectrum. Applied Mathematics and Computation, 336(??):288–300, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304181.

Burgos:2019:RFG

[BCD^{+}19] C. Burgos, J.-C. Cortés, A. Debbouche, L. Villafuerte, and R.-J. Villanueva. Random fractional generalized Airy differential equations: a probabilistic analysis using mean square calculus. Applied Mathematics and Computation, 352(??):15– 29, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300487.

Branquinho:2018:CTS

[BCFR18] A. Branquinho, Y. Chen, G. Filipuk, and M. N. Rebocho. A characterization theorem for semiclassical orthogonal polynomials on non-uniform lattices. Applied Mathematics and Computation, 334(??):356–366, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303400.

Bird:2019:PDG

[BCG19] Robert E. Bird, William M. Coombs, and Stefano Giani. A posteriori discontinuous Galerkin error estimator for linear elasticity. Applied Mathematics and Computation, 344–345(??):78–96, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307537.

Bot:2015:IDR

[BCH15] Radu Ioan Bot, Ernö Robert Csetnek, and Christopher Hendrich. Inertial Douglas–Rachford splitting for monotone inclusion problems. Applied Mathematics and Computation, 256(??):472–487, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315000314.

Brugnano:2015:ECI

[BCI15] L. Brugnano, G. Frasca Caccia, and F. Iavernaro. Energy conservation issues in the numerical solution of the semilinear wave equation. Applied Mathematics and Computation, 270(??):842–870, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011340.

Brandts:2015:TBA

[BCK15] Jan Brandts, Apo Cihangir, and Michal Krízek. Tight bounds on angle sums of nonobtuse simplices. Applied Mathematics and Computation, 267(??):397–408, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002155.

Beccari:2017:PEC

[BCM17] Carolina Vittoria Beccari, Giulio Casciola, and Marie-Laurence Mazure. Piecewise Extended Chebyshev spaces: a numerical test for design. Applied Mathematics and Computation, 296(??):239–256, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305975.

Behl:2015:DFO

[BCMT15a] Ramandeep Behl, Alicia Cordero, S. S. Motsa, and Juan R. Torregrosa. On developing fourth-order optimal families of methods for multiple roots and their dynamics. Applied Mathematics and Computation, 265(??):520–532, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006025.

Behl:2015:CFO

[BCMT15b] Ramandeep Behl, Alicia Cordero, Sandile S. Motsa, and Juan R. Torregrosa. Construction of fourth-order optimal families of iterative methods and their dynamics. Applied Mathematics and Computation, 271(??):89–101, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315011765.

Behl:2017:SHO

[BCMT17] Ramandeep Behl, Alicia Cordero, Sandile S. Motsa, and Juan R. Torregrosa. Stable high-order iterative methods for solving nonlinear models. Applied Mathematics and Computation, 303(??):70–88, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300371.

Bergam:2019:AMT

[BCNN19] A. Bergam, A. Chakib, A. Nachaoui, and M. Nachaoui. Adaptive mesh techniques based on a posteriori error estimates for an inverse Cauchy problem. Applied Mathematics and Computation, 346(??):865–878, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308567.

Bahyrycz:2015:ECM

[BCO15] Anna Bahyrycz, Krzysztof Ciepliński, and Jolanta Olko. On an equation characterizing multiadditive-quadratic mappings and its Hyers–Ulam stability. Applied Mathematics and Computation, 265(??):448–455, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006542.

Bravo:2017:EM

[BCR17] Diego Bravo, Florencia Cubría, and Juan Rada. Energy of matrices. Applied Mathematics and Computation, 312(??):149–157, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303636.

Bahsi:2015:OPS

[BÇS15] M. Mustafa Bahsi, Mehmet Çevik, and Mehmet Sezer. Orthoexponential polynomial solutions of delay pantograph differential equations with residual error estimation. Applied Mathematics and Computation, 271(??):11–21, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011571.

Bermudez:2018:NER

[BCSG18] J. M. Ancochea Bermúdez, R. Campoamor-Stursberg, and F. Oviaño García. New examples of rank one solvable real rigid Lie algebras possessing a nonvanishing Chevalley cohomology. Applied Mathematics and Computation, 339(??):422–430, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306015.

Bahl:2019:NBP

[BCST19] Ashu Bahl, Alicia Cordero, Rajni Sharma, and Juan R. Torregrosa. A novel biparametric sixth order iterative scheme for solving nonlinear systems and its dynamics. Applied Mathematics and Computation, 357(??):147–166, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302796.

Budzko:2015:NFI

[BCT15] Dzmitry Budzko, Alicia Cordero, and Juan R. Torregrosa. A new family of iterative methods widening areas of convergence. Applied Mathematics and Computation, 252(??):405–417, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016828.

Baliarsingh:2015:CFO

[BD15] P. Baliarsingh and S. Dutta. On the classes of fractional order difference sequence spaces and their matrix transformations. Applied Mathematics and Computation, 250(??):665–674, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014969.

Baculikova:2017:OFT

[BD17] B. Baculíková and J. Dzurina. Oscillation of functional trinomial differential equations with positive and negative term. Applied Mathematics and Computation, 295(??):47–52, February 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306014.

Bayrak:2018:NAS

[BD18] Mine Aylin Bayrak and Ali Demir. A new approach for space-time fractional partial differential equations by residual power series method. Applied Mathematics and Computation, 336(??):215–230, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303503.

Bereg:2019:CMT

[BDBKV19] Sergey Bereg, José-Miguel Díaz-Báñez, Nadine Kroher, and Inmaculada Ventura. Computing melodic templates in oral music traditions. Applied Mathematics and Computation, 344–345(??):219–229, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308580.

Bie:2015:CSF

[BDE^{+}15] H. De Bie, N. De Schepper, T. A. Ell, K. Rubrecht, and S. J. Sangwine. Connecting spatial and frequency domains for the quaternion Fourier transform. Applied Mathematics and Computation, 271(??):581–593, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012813.

Berezansky:2015:ESS

[BDGS15] Leonid Berezansky, Alexander Domoshnitsky, Mikhail Gitman, and Valery Stolbov. Exponential stability of a second order delay differential equation without damping term. Applied Mathematics and Computation, 258(??):483–488, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001460.

Buchmuller:2016:FVW

[BDH16] Pawel Buchmüller, Jürgen Dreher, and Christiane Helzel. Finite volume WENO methods for hyperbolic conservation laws on Cartesian grids with adaptive mesh refinement. Applied Mathematics and Computation, 272 (part 2)(??):460–478, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315003926.

Bruzon:2018:ESE

[BdlRT18] M. S. Bruzón, R. de la Rosa, and R. Tracinà. Exact solutions via equivalence transformations of variable-coefficient fifthorder KdV equations. Applied Mathematics and Computation, 325(??): 239–245, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317309049.

Bermejo:2019:IER

[BdS19] R. Bermejo and P. Galán del Sastre. An implicitexplicit Runge–Kutta– Chebyshev finite element method for the nonlinear lithium-ion battery equations. Applied Mathematics and Computation, 361(??):398–420, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304047.

Berezansky:2015:SUE

[BDSS15] Leonid Berezansky, Josef Diblík, Zdenek Svoboda, and Zdenek Smarda. Simple uniform exponential stability conditions for a system of linear delay differential equations. Applied Mathematics and Computation, 250(??):605–614, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014921.

Berezansky:2018:ESL

[BDSS18] Leonid Berezansky, Josef Diblík, Zdenek Svoboda, and Zdenek Smarda. Exponential stability of linear delayed differential systems. Applied Mathematics and Computation, 320(??):474–484, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307026.

Buric:2017:CAA

[BE17] Tomislav Burić and Neven Elezović. Computation and analysis of the asymptotic expansions of the compound means. Applied Mathematics and Computation, 303(??):48–54, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317300334.

Barbu:2019:FBP

[BE19] L. Barbu and C. Enache.

A free boundary problem with multiple boundaries for a general class of anisotropic equations. Applied Mathematics and Computation, 362(??): Article 124551, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930534X.

Beauregard:2015:NSS

[Bea15] Matthew A. Beauregard.

Numerical solutions to singular reaction-diffusion equation over elliptical domains. Applied Mathematics and Computation, 254 (??):75–91, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314017470.

Beauregard:2019:NAF

[Bea19] Matthew A. Beauregard.

Numerical approximations to a fractional Kawarada quenching problem. Applied Mathematics and Computation, 349(??):14– 22, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310816.

Bejancu:2017:GSC

[Bej17] Aurelian Bejancu. Gradient superconvergence for a class of semi-cardinal interpolation schemes with cubic and quintic Bsplines. Applied Mathematics and Computation, 308(??):142–148, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302072.

Baker:2015:DZA

[BEM15] Faroq A. Baker and Hassan A. El-Morshedy. The distribution of zeros of all solutions of first order neutral differential equations. Applied Mathematics and Computation, 259(??):777–789, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002982.

Benjemaa:2018:TFI

[Ben18] Mondher Benjemaa. Taylor’s formula involving generalized fractional derivatives. Applied Mathematics and Computation, 335(??):182–195, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303588.

Berci:2015:SAS

[Ber15] M. Berci. Semi-analytical static aeroelastic analysis and response of flexible subsonic wings. Applied Mathematics and Computation, 267(??):148–169, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315005573.

Berezowski:2017:LCD

[Ber17] Marek Berezowski. Limit cycles that do not comprise steady states of chemical reactors. Applied Mathematics and Computation, 312(??):129–133, November 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317302904.

Borsche:2019:LTS

[BES19] Raul Borsche, Matthias Eimer, and Norbert Siedow. A local time stepping method for thermal energy transport in district heating networks. Applied Mathematics and Computation, 353(??):215–229, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930089X.

Bittencourt:2015:LCA

[BF15] Tiberio Bittencourt and Orizon Pereira Ferreira. Local convergence analysis of Inexact Newton method with relative residual error tolerance under majorant condition in Riemannian manifolds. Applied Mathematics and Computation, 261(??):28–38, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500394X.

Bodnar:2016:DDC

[BF16a] Marek Bodnar and Urszula Foryś. Delays do not cause oscillations in a corrected model of humoral mediated immune response. Applied Mathematics and Computation, 289(??):7–21, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303046.

Borovicanin:2016:EZI

[BF16b] Bojana Borovićanin and Boris Furtula. On extremal Zagreb indices of trees with given domination number. Applied Mathematics and Computation, 279(??):208–218, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300170.

Bras:2015:QEC

[BFIJ15] Carmo P. Brás, Masao Fukushima, Alfredo N. Iusem, and Joaquim J. Júdice. On the Quadratic Eigenvalue Complementarity Problem over a general convex cone. Applied Mathematics and Computation, 271(??):594–608, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315012321.

Bras:2017:BAS

[BFJ^{+}17] Carmo P. Brás, Andreas Fischer, Joaquim J.

Júdice, Klaus Schönefeld, and Sarah Seifert. A block active set algorithm with spectral choice line search for the symmetric eigenvalue complementarity problem. Applied Mathematics and Computation, 294(??):36–48, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305653.

Bonaventura:2018:FSL

[BFR18] Luca Bonaventura, Roberto Ferretti, and Lorenzo Rocchi. A fully semi-Lagrangian discretization for the 2D incompressible Navier–Stokes equations in the vorticitystreamfunction formulation. Applied Mathematics and Computation, 323(??):132–144, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308123.

Bykov:2017:MMW

[BG17] N. Y. Bykov and Yu. E. Gorbachev. Mathematical models of water nucleation process for the Direct Simulation Monte Carlo method. Applied Mathematics and Computation, 296(??):215–232, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306026.

Bali:2018:SNN

[BG18] Rekha Bali and Nivedita Gupta. Study of non-Newtonian fluid by K– L model through a nonsymmetrical stenosed narrow artery. Applied Mathematics and Computation, 320(??):358–370, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306483.

Balbuena:2018:BKR

[BGMO18] C. Balbuena, D. González-Moreno, and M. Olsen. Bounds on the k-restricted arc connectivity of some bipartite tournaments. Applied Mathematics and Computation, 331(??):54– 60, August 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318301486.

Boyadzhiev:2018:SCD

[BGMS18] D. Boyadzhiev, D. N. Georgiou, A. C. Megaritis, and F. Sereti. A study of a covering dimension of finite lattices. Applied Mathematics and Computation, 333(??):276–285, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302145.

Bachar:2017:NCF

[BGMZ17] Mostafa Bachar, Allal Guessab, Osama Mohammed, and Yassine Zaim. New cubature formulas and Hermite– Hadamard type inequalities using integrals over some hyperplanes in the ddimensional hyper-rectangle. Applied Mathematics and Computation, 315(??): 347–362, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317305441.

Borrelli:2015:MOS

[BGP15] Alessandra Borrelli, Giulia Giantesio, and Maria Cristina Patria. MHD orthogonal stagnation-point flow of a micropolar fluid with the magnetic field parallel to the velocity at infinity. Applied Mathematics and Computation, 264(??):44–60, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005160.

Balachandran:2015:CFD

[BGRT15] K. Balachandran, V. Govindaraj, M. Rivero, and J. J. Trujillo. Controllability of fractional damped dynamical systems. Applied Mathematics and Computation, 257(??):66–73, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017135.

Bodaghi:2019:NFI

[BGS19] Amirhosein Bodaghi, Sama Goliaei, and Mostafa Salehi. The number of followings as an influential factor in rumor spreading. Applied Mathematics and Computation, 357(??):167–184, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300319302814.

Barajas-Garcia:2019:STR

[BGSCGL19] Carolina Barajas-García, Selene Solorza-Calderón, and Everardo Gutiérrez-López. Scale, translation and rotation invariant Wavelet Local Feature Descriptor. Applied Mathematics and Computation, 363(??): Article 124594, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305867.

Bohner:2016:OTO

[BGST16] Martin Bohner, Said R. Grace, Ilgin Sager, and Ercan Tunç. Oscillation of third-order nonlinear damped delay differential equations. Applied Mathematics and Computation, 278(??):21–32, March 31, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315300242.

Boulaaras:2015:NPE

[BH15] Salah Boulaaras and Mohamed Haiour. A new proof for the existence and uniqueness of the discrete evolutionary HJB equations. Applied Mathematics and Computation, 262 (??):42–55, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315004099.

Biazar:2017:ISP

[BH17] Jafar Biazar and Mohammad Hosami. An interval for the shape parameter in radial basis function approximation. Applied Mathematics and Computation, 315(??):131–149, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305088.

Boukabcha:2018:RVE

[BHD18] H. Boukabcha, M. Hachama, and A. Diaf. Rovibrational energies of the shifted Deng–Fan oscillator potential with Feynman path integral formalism. Applied Mathematics and Computation, 321(??):121–129, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300317307439.

Bradshaw-Hajek:2015:SSR

[BHM15] B. H. Bradshaw-Hajek and R. J. Moitsheki. Symmetry solutions for reaction-diffusion equations with spatially dependent diffusivity. Applied Mathematics and Computation, 254(??):30–38, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017998.

Baca:2015:TIF

[BHMS15] Martin Baca, Jarmila Horváthová, Martina Mokrisová, and Alzbeta Suhányiová. On topological indices of fullerenes. Applied Mathematics and Computation, 251(??):154–161, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016014.

Bhowmik:2015:MPN

[Bho15] Samir Kumar Bhowmik. A multigrid preconditioned numerical scheme for a reaction-diffusion system. Applied Mathematics and Computation, 254(??):266–276, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017160.

Burkotova:2016:APK

[BHRW16] Jana Burkotová, Michael Hubner, Irena Rachunková, and Ewa B. Weinmüller. Asymptotic properties of Kneser solutions to nonlinear second order ODEs with regularly varying coefficients. Applied Mathematics and Computation, 274(??):65–82, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014253.

Bhunu:2015:APP

[Bhu15] C. P. Bhunu. Assessing the potential of preexposure vaccination and chemoprophylaxis in the control of lymphatic filariasis. Applied Mathematics and Computation, 250(??):571–579, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015409.

Birou:2015:CMT

[Bir15] Marius Mihai Birou. A class of Markov type operators which preserve e_{j}, j ≥ 1. Applied Mathematics and Computation, 250(??):1–11, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014532.

Braverman:2019:ODE

[BJ19] Elena Braverman and William T. Johnson. On oscillation of difference equations with continuous time and variable delays. Applied Mathematics and Computation, 355(??):449–457, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301924.

Buchegger:2016:ARM

[BJM16] Florian Buchegger, Bert Jüttler, and Angelos Mantzaflaris. Adaptively refined multi-patch Bsplines with enhanced smoothness. Applied Mathematics and Computation, 272 (part 1)(??): 159–172, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031500836X.

Bandyopadhyay:2018:CCN

[BK18a] Abhirup Bandyopadhyay and Samarjit Kar. Coevolution of cooperation and network structure in social dilemmas in evolutionary dynamic complex network. Applied Mathematics and Computation, 320(??):710–730, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306768.

Bandyopadhyay:2018:INS

[BK18b] Abhirup Bandyopadhyay and Samarjit Kar. Impact of network structure on synchronization of Hindmarsh–Rose neurons coupled in structured network. Applied Mathematics and Computation, 333(??):194–212, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302650.

Benes:2018:MTS

[BK18c] Michal Benes and Jaroslav Kruis. Multi-time-step domain decomposition and coupling methods for nonlinear parabolic problems. Applied Mathematics and Computation, 319(??):444–460, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302783.

Burdakov:2018:MSI

[BK18d] Oleg Burdakov and Ahmad Kamandi. Multipoint secant and interpolation methods with nonmonotone line search for solving systems of nonlinear equations. Applied Mathematics and Computation, 338(??):412–420, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304508.

Basic:2015:MMC

[BKAAM15] Aleksandra Murks Basic, Salahuddin M. Kamal, Mansour Almazroui, and Fahad M. Al-Marzouki. A mathematical model for the climate change: Can unpredictability offset the temptations to pollute? Applied Mathematics and Computation, 265(??):187–195, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006128.

Bresar:2019:HST

[BKKBS19] Bostjan Bresar, Tim Kos, Rastislav Krivos-Bellus, and Gabriel Semanisin. Hitting subgraphs in P_{4}- tidy graphs. Applied Mathematics and Computation, 352(??):211–219, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300918.

Berntsson:2016:ODM

[BKKN16] Fredrik Berntsson, Matts Karlsson, Vladimir Kozlov, and Sergey A. Nazarov. A one-dimensional model of viscous blood flow in an elastic vessel. Applied Mathematics and Computation, 274(??):125–132, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014289.

Bonamy:2017:DBS

[BKL^{+}17] Marthe Bonamy, Martin Knor, Borut Luzar, Alexandre Pinlou, and Riste Skrekovski. On the difference between the Szeged and the Wiener index. Applied Mathematics and Computation, 312(??):202–213, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031730348X.

Bohacek:2018:NHF

[BKL^{+}18] Jan Bohacek, Abdellah Kharicha, Andreas Ludwig, Menghuai Wu, Ebrahim Karimi-Sibaki, Armin Paar, Michael Brandner, Leonel Elizondo, and Thomas Trickl. A (non-)hydrostatic freesurface numerical model for two-layer flows. Applied Mathematics and Computation, 319(??):301–317, February 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317302126.

Bohacek:2015:ARS

[BKLW15] Jan Bohácek, Abdellah Kharicha, Andreas Ludwig, and Menghuai Wu.

An approximate Riemann solver for shallow water equations and heat advection in horizontal centrifugal casting. Applied Mathematics and Computation, 267(??):179–194, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004865.

Brugnano:2018:FSI

[BKS18] Luigi Brugnano, Dmitri E.

Kvasov, and Yaroslav D. Sergeyev. Foreword to the special issue “Recent Trends in Numerical Computations: Theory and Algorithms”. Applied Mathematics and Computation, 318(??):1–2, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730632X.

Bogovalov:2016:IPB

[BKT16] S. V. Bogovalov, V. A.

Kislov, and I. V. Tronin. Impact of the pulsed braking force on the axial circulation in a gas centrifuge. Applied Mathematics and Computation, 272 (part 3)(??):670–675, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501070X.

Bhatt:2018:EKB

[BKW18] H. P. Bhatt, A. Q. M. Khaliq, and B. A. Wade. Efficient Krylov-based exponential time differencing method in application to 3D advection-diffusion-reaction systems. Applied Mathematics and Computation, 338(??):249–259, December 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318305137.

Bonakdari:2018:EAS

[BKZA18] Hossein Bonakdari, Zohreh Sheikh Khozani, Amir Hossein Zaji, and Navid Asadpour. Evaluating the apparent shear stress in prismatic compound channels using the genetic algorithm based on multi-layer perceptron: a comparative study. Applied Mathematics and Computation, 338(??):387–399, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305046.

Bula:2015:CFD

[BL15] D. Bula and M. Lewandowski.

Comparison of frequency domain and time domain model of a distributed power supplying system with active power filters (APFs). Applied Mathematics and Computation, 267(??):771–779, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016142.

Bula:2018:SSS

[BL18] D. Bula and M. Lewandowski.

Steady state simulation of a distributed power supplying system using a simple hybrid time-frequency model. Applied Mathematics and Computation, 319(??):195–202, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301327.

Bravo:2015:TTE

[BLA15] Yesnier Bravo, Gabriel Luque, and Enrique Alba. Takeover time in evolutionary dynamic optimization: From theory to practice. Applied Mathematics and Computation, 250(??):94–104, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014829.

Bourantas:2019:EMP

[BLJ^{+}19] G. C. Bourantas, V. C. Loukopoulos, G. R. Joldes, A. Wittek, and K. Miller. An explicit meshless point collocation method for electrically driven magnetohydrodynamics (MHD) flow. Applied Mathematics and Computation, 348(??):215–233, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310282.

Bakhtiari:2019:DFO

[BLK19] Mehrdad Bakhtiari, Aouni A. Lakis, and Youcef Kerboua. Derivatives of fourth order Kronecker power systems with applications in nonlinear elasticity. Applied Mathematics and Computation, 362 (??):Article 124501, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.

com/science/article/
pii/S0096300319304801.

Bracco:2016:GSS

[BLM^{+}16] Cesare Bracco, Tom Lyche, Carla Manni, Fabio Roman, and Hendrik Speleers. Generalized spline spaces over T - meshes: Dimension formula and locally refined generalized B-splines. Applied Mathematics and Computation, 272 (part 1) (??):187–198, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315010747.

Baca:2017:EFI

[BLNSF17] Martin Baca, Marcela Lascsáková, Maria Naseem, and Andrea Semanicová-Fenovcíková. On entire face irregularity strength of disjoint union of plane graphs. Applied Mathematics and Computation, 307(??):232–238, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301625.

Borges:2017:PAM

[BLP17] Fábio Borges, Pedro Lara, and Renato Portugal. Parallel algorithms for modular multi-exponentiation. Applied Mathematics and Computation, 292(??): 406–416, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031630474X.

Braha:2015:WSC

[BLS15] Naim L. Braha, Valdete Loku, and H. M. Srivastava. Λ^{2} -weighted statistical convergence and Korovkin and Voronovskaya type theorems. Applied Mathematics and Computation, 266(??):675–686, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315007390.

Balakumar:2015:NSV

[BM15a] V. Balakumar and K. Murugesan. Numerical solution of Volterra integralalgebraic equations using block pulse functions. Applied Mathematics and Computation, 263(??):165–170, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315004932.

Barbosu:2015:ACF

[BM15b] Dan Barbosu and Carmen Violeta Muraru. Approximating B-continuous functions using GBS operators of Bernstein– Schurer–Stancu type based on q-integers. Applied Mathematics and Computation, 259(??):80–87, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002106.

Bai:2016:SUP

[BM16] Yongqiang Bai and Hongping Ma. Solutions to unsolved problems on the minimal energies of two classes of trees. Applied Mathematics and Computation, 286(??):49–56, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630248X.

Bahbouhi:2017:PDG

[BM17] Jalal Eddine Bahbouhi and Najem Moussa. Prisoner’s dilemma game model for e-commerce. Applied Mathematics and Computation, 292(??):

128–144, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316304489.

Bhattacharyya:2018:GSD

[BM18a] A. Bhattacharyya and B. Mukhopadhyay. The general solution for dynamical problem of rectangular micro-polar beam vibrating at high frequency. Applied Mathematics and Computation, 332(??):376–389, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302418.

Bouchlaghem:2018:AMF

[BM18b] Mohammed Bouchlaghem and El Bekkaye Mermri. Analysis of a mixed formulation of a bilateral obstacle problem. Applied Mathematics and Computation, 320(??):45–55, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306379.

Bahbouhi:2019:GBM

[BM19a] Jalal Eddine Bahbouhi and Najem Moussa. A graph-based model for public goods with leaderships. Applied Mathematics and Computation, 349(??):53–61, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310907.

Balbuena:2019:REC

[BM19b] C. Balbuena and X. Marcote. The p-restricted edge-connectivity of Kneser graphs. Applied Mathematics and Computation, 343(??):258–267, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308592.

Barg:2019:PSP

[BM19c] Michael C. Barg and Amanda J. Mangum. A phase separation problem and geodesic disks on Cassinian oval surfaces. Applied Mathematics and Computation, 354(??):192–205, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300319301419.

Bezerra:2019:EMC

[BM19d] Licio Hernanes Bezerra and Nelson Martins. Eigenvalue methods for calculating dominant poles of a transfer function and their applications in smallsignal stability. Applied Mathematics and Computation, 347(??):113–121, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830955X.

Bracciali:2015:ZCG

[BMB15] Cleonice F. Bracciali and Juan José Moreno-Balcázar. On the zeros of a class of generalized hypergeometric polynomials. Applied Mathematics and Computation, 253(??):151–158, February 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314017378.

Bhattacharjee:2019:EPD

[BMC19] Arindam Bhattacharjee, Atanu K. Mohanty, and Anindya Chatterjee. Expansion of Preisach density in magnetic hysteresis using general basis functions. Applied Mathematics and Computation, 341(??):418–427, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300318307811.

Bajoulvand:2017:AFM

[BMDS17] Atena Bajoulvand, Ramtin Zargari Marandi, Mohammad Reza Daliri, and Seyed Hojjat Sabzpoushan. Analysis of folk music preference of people from different ethnic groups using kernel-based methods on EEG signals. Applied Mathematics and Computation, 307(??):62–70, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301534.

Burgos:2017:MPD

[BMGMC17] M. Burgos, A. C. Márquez-García, and A. Morales-Campoy. Maps preserving the diamond partial order. Applied Mathematics and Computation, 296(??):137–147, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300316306166.

Barbu:2019:QAN

[BMM19] Tudor Barbu, Alain Miranville, and Costica Morosanu. A qualitative analysis and numerical simulations of a nonlinear second-order anisotropic diffusion problem with non-homogeneous Cauchy– Neumann boundary conditions. Applied Mathematics and Computation, 350(??):170–180, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300128.

Baleanu:2015:KTO

[BMO15] Dumitru Baleanu, Octavian G. Mustafa, and Donal O’Regan. A Kamenev-type oscillation result for a linear 1 + αorder fractional differential equation. Applied Mathematics and Computation, 259(??):374–378, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002337.

Borrego-Morell:2017:AEC

[BMR17] J. Borrego-Morell and F. R. Rafaeli. An asymptotic expansion for a class of biorthogonal polynomials with respect to a measure on the unit circle. Applied Mathematics and Computation, 313(??):52–73, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730382X.

Burger:2018:LIE

[BMRS18] Raimund Bürger, Pep Mulet, Lihki Rubio, and Mauricio Sepúlveda. Linearly implicit-explicit schemes for the equilibrium dispersive model of chromatography. Applied Mathematics and Computation, 317(??):172–186, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306264.

Benyi:2019:RSN

[BMRW19] Beáta Bényi, Miguel Méndez, José L. Ramírez, and Tanay Wakhare. Restricted r-Stirling numbers and their combinatorial applications. Applied Mathematics and Computation, 348(??):186–205, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031831021X.

Bellamoli:2018:NMJ

[BMT18] Francesca Bellamoli, Lucas O. Müller, and Eleuterio F. Toro. A numerical method for junctions in networks of shallow-water channels. Applied Mathematics and Computation, 337(??):179–189, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304430.

Berhe:2019:CDM

[BMT19] Hailay Weldegiorgis Berhe, Oluwole Daniel Makinde, and David Mwangi Theuri. Co-dynamics of measles and dysentery diarrhea diseases with optimal control and cost-effectiveness analysis. Applied Mathematics and Computation, 347(??):903–921, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310233.

Bayad:2015:APF

[BN15] Abdelmejid Bayad and Luis Navas. Algebraic properties and Fourier expansions of two-dimensional Apostol–Bernoulli and Apostol–Euler polynomials. Applied Mathematics and Computation, 265(??):883–892, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007596.

Baiardi:2018:OMB

[BN18] Lorenzo Cerboni Baiardi and Ahmad K. Naimzada. An oligopoly model with best response and imitation rules. Applied Mathematics and Computation, 336(??):193–205, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303874.

Belbachir:2016:HPT

[BNS16] Hacène Belbachir, László Németh, and László Szalay. Hyperbolic Pascal triangles. Applied Mathematics and Computation, 273(??):453–464, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013338.

Benes:2015:MTS

[BNY15] Michal Benes, Ales Nekvinda, and Manoj Kumar Yadav. Multi-time-step domain decomposition method with non-matching grids for parabolic problems. Applied Mathematics and Computation, 267(??): 571–582, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315000855.

Boglaev:2016:NSN

[Bog16] Igor Boglaev. Numerical solving nonlinear integroparabolic equations by the monotone weighted average method. Applied Mathematics and Computation, 274(??):152–162, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315014757.

Bogosel:2018:EAO

[Bog18] Beniamin Bogosel. Efficient algorithm for optimizing spectral partitions. Applied Mathematics and Computation, 333(??):61–75, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302686.

Boikanyo:2015:SCA

[Boi15] Oganeditse A. Boikanyo.

A strongly convergent algorithm for the split common fixed point problem. Applied Mathematics and Computation, 265(??):844–853, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007614.

Ouyang:2015:TLB

[bOqGyK^{+}15] Hai bin Ouyang, Li qun Gao, Xiang yong Kong, De xuan Zou, and Steven Li. Teaching-learning based optimization with global crossover for global optimization problems. Applied Mathematics and Computation, 265(??): 533–556, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315006281.

Bor:2015:ANS

[Bor15a] Hüseyin Bor. Absolute Nörlund summability factors involving almost increasing sequences. Applied Mathematics and Computation, 259(??): 828–830, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315003100.

Borges:2015:PFA

[Bor15b] Carlos F. Borges. On polynomial function approximation with minimum mean squared relative error and a problem of Tchebychef. Applied Mathematics and Computation, 258(??):22–28, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001605.

Boraey:2019:AAS

[Bor19] Mohammed A. Boraey. An Asymptotically Adaptive Successive Equilibrium Relaxation approach for the accelerated convergence of the Lattice Boltzmann Method. Applied Mathematics and Computation, 353(??):29–41, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300785.

Botti:2015:CFT

[Bot15] L. Botti. A choice of forcing terms in inexact Newton iterations with application to pseudo-transient continuation for incompressible fluid flow computations. Applied Mathematics and Computation, 266(??):713–737, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007675.

Boudeliou:2018:CNN

[Bou18] Ammar Boudeliou. On certain new nonlinear retarded integral inequalities in two independent variables and applications. Applied Mathematics and Computation, 335(??):103–111, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830359X.

Boyd:2015:FEA

[Boy15] John P. Boyd. A Fourier error analysis for radial basis functions and the Discrete Singular Convolution on an infinite uniform grid, Part 1: Error theorem and diffusion in Fourier space. Applied Mathematics and Computation, 264(??):132–140, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005378.

Boyd:2017:CEI

[Boy17] John P. Boyd. The Crane equation uu_{xx} = −2: the general explicit solution and a case study of Chebyshev polynomial series for functions with weak endpoint singularities. Applied Mathematics and Computation, 301(??):214–223, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307469.

Bashier:2017:OCE

[BP17a] Eihab B. M. Bashier and Kailash C. Patidar. Optimal control of an epidemiological model with multiple time delays. Applied Mathematics and Computation, 292(??):47–56, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304398.

Bracciali:2017:BOP

[BP17b] Cleonice F. Bracciali and Teresa E. Pérez. Bivariate orthogonal polynomials, 2D Toda lattices and Laxtype pairs. Applied Mathematics and Computation, 309(??):142–155, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302400.

Bao:2015:MMS

[BPC15] Haibo Bao, Ju H. Park, and Jinde Cao. Matrix measure strategies for exponential synchronization and anti-synchronization of memristor-based neural networks with timevarying delays. Applied Mathematics and Computation, 270(??):543–556, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315011200.

Bonettini:2016:CBC

[BPR16] Silvia Bonettini, Marco Prato, and Simone Rebegoldi. A cyclic block coordinate descent method with generalized gradient projections. Applied Mathematics and Computation, 286(??):288–300, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302818.

Buccini:2018:NAN

[BPR18] Alessandro Buccini, Yonggi Park, and Lothar Reichel. Numerical aspects of the nonstationary modified linearized Bregman algorithm. Applied Mathematics and Computation, 337(??):372–385, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304533.

Bubba:2018:NRA

[BPZB18] Tatiana A. Bubba, Federica Porta, Gaetano Zanghirati, and Silvia Bonettini. A nonsmooth regularization approach based on shearlets for Poisson noise removal in ROI tomography. Applied Mathematics and Computation, 318(??):131–152, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306215.

Bula:2015:PON

[BR15] I. Bula and M. A.

Radin. Periodic orbits of a neuron model with periodic internal decay rate. Applied Mathematics and Computation, 266(??):293–303, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007213.

BenTaher:2016:SSG

[BR16a] R. Ben Taher and M. Rachidi.

Solving some generalized Vandermonde systems and inverse of their associate matrices via new approaches for the Binet formula. Applied Mathematics and Computation, 290(??):267–280, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303770.

Betancourt:2016:FVS

[BR16b] Fernando Betancourt and Christian Rohde. Finitevolume schemes for Friedrichs systems with involutions. Applied Mathematics and Computation, 272 (part 2) (??):420–439, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315003641.

Bunder:2017:RSM

[BR17] J. E. Bunder and A. J. Roberts. Resolution of subgrid microscale interactions enhances the discretisation of nonautonomous partial differential equations. Applied Mathematics and Computation, 304(??):164–179, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300747.

Bracken:2015:CIS

[Bra15]  Paul Bracken.  On

the construction of integrable surfaces on Lie groups. Applied Mathematics and Computation, 261(??):167–175, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004142.

Blaya:2015:BVPa

[BRD15] Ricardo Abreu Blaya, Juan Bory Reyes, and Ramón M. Rodríguez Dagnino. Boundary value problems for hyperholomorphic solutions of two dimensional Helmholtz equation in a fractal domain. Applied Mathematics and Computation, 261(??):183–191, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004178.

Borges:2019:LDT

[BRM^{+}19] João B. Borges, Heitor S.

Ramos, Raquel A. F. Mini, Osvaldo A. Rosso, Alejandro C. Frery, and Antonio A. F. Loureiro. Learning and distinguishing time series dynamics via ordinal patterns transition graphs. Applied Mathematics and Computation, 362(??): Article 124554, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305375.

Bonazzoli:2018:DAT

[BRV18] Marcella Bonazzoli, Francesca Rapetti, and Chiara Venturini. Dispersion analysis of triangle-based Whitney element methods for electromagnetic wave propagation. Applied Mathematics and Computation, 319(??):274–286, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302084.

Bhadauria:2015:ABF

[BS15a] Saumya Bhadauria and Anirban Sengupta. Adaptive bacterial foraging driven datapath optimization: Exploring powerperformance tradeoff in high level synthesis. Applied Mathematics and Computation, 269(??): 265–278, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315009613.

Bhunia:2015:APT

[BS15b] A. K. Bhunia and Ali Akbar Shaikh. An application of PSO in a two-warehouse inventory model for deteriorating item under permissible delay in payment with different inventory policies. Applied Mathematics and Computation, 256(??):831–850, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017986.

Bulbul:2015:NAS

[BS15c] Berna Bülbül and Mehmet Sezer. A numerical approach for solving generalized Abel-type nonlinear differential equations. Applied Mathematics and Computation, 262(??):169–177, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005159.

Balci:2016:HET

[BS16] Mehmet Ali Balci and Mehmet Sezer. Hybrid Euler–Taylor matrix method for solving of generalized linear Fredholm integro-differential difference equations. Applied Mathematics and Computation, 273(??):33–41, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013235.

Behroozifar:2017:ASB

[BS17] M. Behroozifar and A. Sazmand. An approximate solution based on Jacobi polynomials for time-fractional convectiondiffusion equation. Applied Mathematics and Computation, 296(??):1–17, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305987.

Blokhin:2019:NAN

[BS19a] Alexander Blokhin and Boris Semisalov. A new approach to numerical simulation of charge transport in double Gate-MOSFET. Applied Mathematics and Computation, 342(??):206–223, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308026.

Breckling:2019:LTS

[BS19b] Sean Breckling and Sidney Shields. The longtime L_{2} and H_{1} stability of linearly extrapolated second-order timestepping schemes for the 2D incompressible Navier– Stokes equations. Applied Mathematics and Computation, 342(??):263–279, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830794X.

Brezovnik:2019:CRI

[BS19c] Simon Brezovnik and Tadeja Kraner Sumenjak. Complexity of krainbow independent domination and some results on the lexicographic product of graphs. Applied Mathematics and Computation, 349(??):214–220, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310506.

Bollati:2018:HBI

[BST18] J. Bollati, J. Semitiel, and D. A. Tarzia. Heat balance integral methods applied to the one-phase Stefan problem with a convective boundary condition at the fixed face. Applied Mathematics and Computation, 331(??):1–19, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318301644.

Bellour:2016:TCS

[BSZ16] A. Bellour, D. Sbibih, and A. Zidna. Two cubic spline methods for solving Fredholm integral equations. Applied Mathematics and Computation, 276(??):1–11, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501557X.

Balasubramaniam:2015:ACC

[BT15a] P. Balasubramaniam and P. Tamilalagan. Approximate controllability of a class of fractional neutral stochastic integro-differential inclusions with infinite delay by using Mainardi’s function. Applied Mathematics and Computation, 256(??):232–246, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000491.

Bartha:2015:FPD

[BT15b] Ferenc A. Bartha and Warwick Tucker. Fixed points of a destabilized Kuramoto–Sivashinsky equation. Applied Mathematics and Computation, 266(??):339–349, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007067.

Bayada:2018:ANM

[BTH18] G. Bayada, M. EL Alaoui Talibi, and M. Hilal. About new models of slip/no-slip boundary condition in thin film flows. Applied Mathematics and Computation, 338(??): 828–841, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318305332.

Bascanbaz-Tunca:2016:BBT

[BTIIE16] Gülen Bascanbaz-Tunca, Hatice Gül Ince-Ilarslan, and Aysegül Erençin. Bivariate Bernstein type operators. Applied Mathematics and Computation, 273(??):543–552, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501379X.

Beaumont:2017:PNI

[BTP17] Fabien Beaumont, Redha Taïar, and Guillaume Polidori. Preliminary numerical investigation in open currents-water swimming: Pressure field in the swimmer wake. Applied Mathematics and Computation, 302(??):48–57, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307652.

Bu:2015:CNA

[BTWY15] Weiping Bu, Yifa Tang, Yingchuan Wu, and Jiye Yang. Crank–Nicolson ADI Galerkin finite element method for twodimensional fractional FitzHugh– Nagumo monodomain model. Applied Mathematics and Computation, 257(??): 355–364, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314012594.

Bublik:2015:CDG

[BVJ15] Ondrej Bublík, Jan Vimmr, and Alena Jonásová. Comparison of discontinuous Galerkin time integration schemes for the solution of flow problems with deformable domains. Applied Mathematics and Computation, 267(??):329–340, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500377X.

Bai:2015:ESS

[BW15a] Jiejing Bai and Li Wang.

EJIIM for the stationary Schrödinger equations with delta potential wells. Applied Mathematics and Computation, 254(??):113–124, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017561.

Bai:2015:TWD

[BW15b] Zhenguo Bai and Shi-Liang Wu. Traveling waves in a delayed SIR epidemic model with nonlinear incidence. Applied Mathematics and Computation, 263(??):221–232, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005068.

Wu:2016:CGF

[bWzZjH16] Zeng bao Wu, Yun zhi Zou, and Nan jing Huang. A class of global fractionalorder projective dynamical systems involving setvalued perturbations. Applied Mathematics and Computation, 277(??):23– 33, March 20, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315300217.

Bean:2019:MMF

[BY19] Maranda Bean and Son-Young Yi. A monolithic mixed finite element method for a fluidstructure interaction problem. Applied Mathematics and Computation, 363 (??):Article 124615, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319306071.

Burton:2015:RKT

[BZ15] T. A. Burton and Bo Zhang. On the reduction of Krasnoselskii’s theorem to Schauder’s theorem. Applied Mathematics and Computation, 250(??): 339–351, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314014684.

Brdar:2016:GMT

[BZ16] Mirjana Brdar and Helena Zarin. On graded meshes for a two-parameter singularly perturbed problem. Applied Mathematics and Computation, 282(??):97– 107, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300698.

Boltuc:2017:ADS

[BZ17] Agnieszka Boltuć and Eugeniusz Zieniuk. Approximation of the derivatives of solutions in a normalized domain for 2D solids using the PIES method. Applied Mathematics and Computation, 293(??):138–155, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305136.

Caballero-Aguila:2016:QEP

[CÁGGLP16] R. Caballero-Águila, I. García-Garrido, and J. Linares-Pérez. Quadratic estimation problem in discretetime stochastic systems with random parameter matrices. Applied Mathematics and Computation, 273(??):308–320, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013375.

Caldarola:2018:SCV

[Cal18] Fabio Caldarola. The Sierpinski curve viewed by numerical computations with infinities and infinitesimals. Applied Mathematics and Computation, 318(??):321–328, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730437X.

Candan:2015:OBS

[Can15]  T. Candan.  Oscilla-

tory behavior of second order nonlinear neutral differential equations with distributed deviating arguments. Applied Mathematics and Computation, 262(??):199–203, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315004531.

Candemir:2016:KGP

[Can16] Nuray Candemir. Klein– Gordon particles in symmetrical well potential. Applied Mathematics and Computation, 274(??): 531–538, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315015210.

Cao:2019:BIC

[Cao19] Yang Cao. Bifurcations in an Internet congestion control system with distributed delay. Applied Mathematics and Computation, 347(??):54–63, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309731.

Castillo:2015:MZC

[Cas15] Kenier Castillo. Monotonicity of zeros for a class of polynomials including hypergeometric polynomials. Applied Mathematics and Computation, 266(??):183–193, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006827.

Catinas:2017:ESP

[Cat17]  Teodora Catinas.  Ex-

tension of some particular interpolation operators to a triangle with one curved side. Applied Mathematics and Computation, 315(??):286–297, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305210.

Catinas:2019:SHC

[Cat19] Emil Catinas. A survey on the high convergence orders and computational convergence orders of sequences. Applied Mathematics and Computation, 343(??):1–20, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306453.

Chidume:2017:IAA

[CB17] C. E. Chidume and A. U. Bello. An iterative algorithm for approximating solutions of Hammerstein equations with monotone maps in Banach spaces. Applied Mathematics and Computation, 313(??):408–417, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304253.

Ching:2019:MBP

[CB19] Atheeta Ching and Stephen Baigent. Manifolds of balance in planar ecological systems. Applied Mathematics and Computation, 358(??):204–215, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303352.

Cano-Berlanga:2017:ECG

[CBGGV17] Sebastián Cano-Berlanga, José-Manuel Giménez-Gómez, and Cori Vilella. Enjoying cooperative games: the R package GameTheory. Applied Mathematics and Computation, 305(??): 381–393, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317301145.

Chen:2015:RTS

[CC15] Boshan Chen and Jiejie Chen. Razumikhin-type stability theorems for functional fractional-order differential systems and applications. Applied Mathematics and Computation, 254(??):63–69, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016646.

Chao:2016:GPP

[CC16] Zhen Chao and Guo-Liang Chen. A generalized preconditioned parameterized inexact Uzawa method for singular saddle point problems. Applied Mathematics and Computation, 284(??):219–225, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301710.

Chen:2017:UTS

[CC17] Chao-Ping Chen and Junesang Choi. Unified treatment of several asymptotic expansions concerning some mathematical constants. Applied Mathematics and Computation, 305(??):348–363, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300978.

Chang:2018:AAT

[CC18a] Qianshun Chang and Zengyan Che. An adaptive algorithm for TV-based model of three norms L_{q}(q = 1/2, 1, 2) in image restoration. Applied Mathematics and Computation, 329(??):251–265, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300614.

Chen:2018:NSS

[CC18b] Y. Chen and Chang-Ming Chen. Numerical simulation with the second order compact approximation of first order derivative for the modified fractional diffusion equation. Applied Mathematics and Computation, 320(??):319–330, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305313.

Cheng:2018:THA

[CC18c] Huiwen Cheng and Li Cui.

Tetravalent half-arc-transitive graphs of order p^{5}. Applied Mathematics and Computation, 332(??):506–518, September 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318302571.

Chen:2019:CPG

[CC19] Yu Chen and Haiyan Chen. The characteristic polynomial of a generalized join graph. Applied Mathematics and Computation, 348(??):456–464, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310543.

Campagna:2017:MTR

[CCC^{+}17] Rosanna Campagna, Serena Crisci, Salvatore Cuomo, Livia Marcellino, and Gerardo Toraldo. Modification of TV–ROF denoising model based on Split Bregman iterations. Applied Mathematics and Computation, 315(??):453–467, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305520.

Chehabi:2019:APD

[CCC19] Hamza Chehabi, Omar Chakrone, and Mohammed Chehabi. On the antimaximum principle for the discrete p-Laplacian with sign-changing weight. Applied Mathematics and Computation, 342(??): 112–117, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318307847.

Chen:2018:CEA

[CCD18] Jianwen Chen, Zhi-Min Chen, and Bo-Qing Dong. Commutator estimate and its application to regularity criteria of the dissipative quasi-geostrophic equation. Applied Mathematics and Computation, 329(??):84–91, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300912.

Charina:2016:LLP

[CCGP16] Maria Charina, Costanza Conti, Nicola Guglielmi, and Vladimir Protasov. Limits of level and parameter dependent subdivision schemes: a matrix approach. Applied Mathematics and Computation, 272 (part 1)(??):20–27, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012011.

Coroianu:2019:MPG

[CCGV19] Lucian Coroianu, Danilo Costarelli, Sorin G. Gal, and Gianluca Vinti. The max-product generalized sampling operators: convergence and quantitative estimates. Applied Mathematics and Computation, 355(??):173–183, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301869.

Chen:2015:ANO

[CCH15] Zhixiang Chen, Feilong Cao, and Jinjie Hu. Approximation by network operators with logistic activation functions. Applied Mathematics and Computation, 256(??):565–571, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500079X.

Casaban:2015:RLT

[CCJ15] M.-C. Casabán, J.-C.

Cortés, and L. Jódar. A random Laplace transform method for solving random mixed parabolic differential problems. Applied Mathematics and Computation, 259(??):654–667, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002921.

Cao:2015:SIT

[CCL^{+}15] Chuansheng Cao, Sheng Chen, Jing Li, Zhaohui Liu, Lu Zha, Sheng Bao, and Chuguang Zheng. Simulating the interactions of two freely settling spherical particles in Newtonian fluid using lattice-Boltzmann method. Applied Mathematics and Computation, 250(??):533–551, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015471.

Cheng:2016:RFT

[CCL^{+}16] Jun Cheng, Shiqiang Chen, Zhijun Liu, Hailing Wang, and Jin Li. Robust finite-time sampleddata control of linear systems subject to random occurring delays and its application to fourtank system. Applied Mathematics and Computation, 281(??):55–76, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300467.

Corbera:2019:TCC

[CCLPC19] Montserrat Corbera, Josep M. Cors, Jaume Llibre, and Ernesto Pérez-Chavela. Trapezoid central configurations. Applied Mathematics and Computation, 346(??):127–142, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309329.

Cabada:2016:GFS

[CCLS16] Alberto Cabada, José A. Cid, and Lucía López-Somoza. Green’s functions and spectral theory for the Hill’s equation. Applied Mathematics and Computation, 286(??):88–105, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300316302405.

Cai:2018:DVW

[CCLW18] Qingqiong Cai, Fuyuan Cao, Tao Li, and Hua Wang. On distances in vertex-weighted trees. Applied Mathematics and Computation, 333(??): 435–442, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318302984.

Caraballo:2019:ARV

[CCNQ19] Tomás Caraballo, J.-C.

Cortés, and A. Navarro-Quiles. Applying the random variable transformation method to solve a class of random linear differential equation with discrete delay. Applied Mathematics and Computation, 356(??):198–218, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302553.

Chen:2015:SCA

[cCpSkZ15] Shao chun Chen, Yan ping Sun, and Ji kun Zhao. The simplest conforming anisotropic rectangular and cubic mixed finite elements for elasticity. Applied Mathematics and Computation, 265(??):292–303, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005883.

Chen:2016:DOF

[CCQS16] Xiaoming Chen, Mou Chen, Wenhai Qi, and Jun Shen. Dynamic output-feedback control for continuous-time interval positive systems under L_{1} performance. Applied Mathematics and Computation, 289(??):48–59, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303009.

Castillo:2018:MTZ

[CCR18] K. Castillo, M. S. Costa, and F. R. Rafaeli. On Markov’s theorem on zeros of orthogonal polynomials revisited. Applied Mathematics and Computation, 339(??):374–389, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318306143.

Campos:2015:MDA

[CCTV15] Beatriz Campos, Alicia Cordero, Juan R. Torregrosa, and Pura Vindel. A multidimensional dynamical approach to iterative methods with memory. Applied Mathematics and Computation, 271(??):701–715, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012941.

Campos:2016:DMV

[CCTV16] Beatriz Campos, Alicia Cordero, Juan R. Torregrosa, and Pura Vindel. Dynamics of a multipoint variant of Chebyshev– Halley’s family. Applied Mathematics and Computation, 284(??):195–208, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302016.

Chen:2017:PDL

[CCW17] Qiao Chen, Tong Chen, and Yongjie Wang. Publishing the donation list incompletely promotes the emergence of cooperation in public goods game. Applied Mathematics and Computation, 310(??):48–56, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302692.

Chen:2019:CHD

[CCW19] Qiao Chen, Tong Chen, and Yongjie Wang. Cleverly handling the donation information can promote cooperation in public goods game. Applied Mathematics and Computation, 346(??):363–373, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309421.

Chen:2015:FOT

[CCX15] Dali Chen, YangQuan Chen, and Dingyu Xue. Fractional-order total variation image denoising based on proximity algorithm. Applied Mathematics and Computation, 257(??):537–545, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315000260.

Cai:2019:EEP

[CCY19a] Jiaxiang Cai, Juan Chen, and Bin Yang. Efficient energy-preserving wavelet collocation schemes for the coupled nonlinear Schrödinger–Boussinesq system. Applied Mathematics and Computation, 357(??):1–11, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302668.

Chen:2019:ESY

[CCY19b] Dongmei Chen, Zhibing Chen, and Xuerong Yong. Explicit solutions of the Yang–Baxterlike matrix equation for a diagonalizable matrix with spectrum contained in {1, α, 0}. Applied Mathematics and Computation, 348(??):523–530, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310865.

Cai:2019:PIE

[CCYT19] Shi-Min Cai, Xuan-Hao Chen, Xi-Jun Ye, and Ming Tang. Precisely identifying the epidemic thresholds in real networks via asynchronous updating. Applied Mathematics and Computation, 361(??):377–388, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304412.

Cheng:2015:ASN

[CCZ15] Xue Cheng, Zhong Chen, and Qingpu Zhang. An approximate solution for a neutral functional-differential equation with proportional delays. Applied Mathematics and Computation, 260(??):27–34, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003458.

Cao:2015:DBE

[CD15a] Shujuan Cao and Matthias Dehmer. Degree-based entropies of networks revisited. Applied Mathematics and Computation, 261(??):141–147, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315003604.

Chen:2015:NNM

[CD15b] Xiao-Ping Chen and Hua Dai. A novel numerical method to determine the algebraic multiplicity of nonlinear eigenvalues. Applied Mathematics and Computation, 271(??):344–351, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012424.

Cheng:2015:PSS

[CD15c] Xiyou Cheng and Guowei Dai. Positive solutions of sub-superlinear Sturm– Liouville problems. Applied Mathematics and Computation, 261(??): 351–359, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315004440.

Cherniha:2015:NRD

[CD15d] Roman Cherniha and Vasyl’ Davydovych. Nonlinear reaction-diffusion systems with a nonconstant diffusivity: Conditional symmetries in nogo case. Applied Mathematics and Computation, 268(??):23–34, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007985.

Chen:2016:SAT

[CD16a] Xiao-Ping Chen and Hua Dai. Stability analysis of time-delay systems using a contour integral method. Applied Mathematics and Computation, 273(??):390–397, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013326.

Civalek:2016:SMM

[CD16b] Ömer Civalek and Cigdem Demir. A simple mathematical model of microtubules surrounded by an elastic matrix by nonlocal finite element method. Applied Mathematics and Computation, 289(??):335–352, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303435.

Copetti:2017:NAS

[CDFP17] M. I. M. Copetti, J. Durany, J. R. Fernández, and L. Poceiro. Numerical analysis and simulation of a bio-thermal model for the human foot. Applied Mathematics and Computation, 305(??):103–116, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300930.

Chakraborty:2015:MSE

[CDHK15] Kunal Chakraborty, Kunal Das, Samadyuti Haldar, and T. K. Kar. A mathematical study of an eco-epidemiological system on disease persistence and extinction perspective. Applied Mathematics and Computation, 254(??):99–112, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017706.

Chatzarakis:2019:NOC

[CDJ19] G. E. Chatzarakis, J. Dzurina, and I. Jadlovská. New oscillation criteria for second-order halflinear advanced differential equations. Applied Mathematics and Computation, 347(??):404–416, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309718.

Cerone:2015:JOT

[CDK15] Pietro Cerone, Sever S. Dragomir, and Eder Kikianty. Jensen–Ostrowski type inequalities and applications for f-divergence measures. Applied Mathematics and Computation, 266(??):304–315, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006955.

Cao:2017:NEB

[CDK17a] Shujuan Cao, Matthias Dehmer, and Zhe Kang. Network entropies based on independent sets and matchings. Applied Mathematics and Computation, 307(??):265–270, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730125X.

Cermak:2017:FDE

[CDK17b] Jan Cermák, Zuzana Doslá, and Tomás Kisela. Fractional differential equations with a constant delay: Stability and asymptotics of solutions. Applied Mathematics and Computation, 298(??):336–350, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306841.

Chiu:2015:PAO

[CDL15] Chun-Yuan Chiu, Tian-Shyr Dai, and Yuh-Dauh Lyuu. Pricing Asian option by the FFT with higher-order error convergence rate under Lévy processes. Applied Mathematics and Computation, 252(??):418–437, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016476.

Christophe:2018:IHD

[CDL18]  Alexandra  Christophe,

Stéphane Descombes, and Stéphane Lanteri. An implicit hybridized discontinuous Galerkin method for the 3D time-domain Maxwell equations. Applied Mathematics and Computation, 319(??): 395–408, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317302758.

Cheng:2019:NCA

[CDL19] Xiujun Cheng, Jinqiao Duan, and Dongfang Li. A novel compact ADI scheme for two-dimensional Riesz space fractional nonlinear reaction-diffusion equations. Applied Mathematics and Computation, 346(??):452–464, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309317.

Chabassier:2019:EBC

[CDP19] J. Chabassier, M. Duruflé, and V. Péron. Equivalent boundary conditions for acoustic media with exponential densities. Application to the atmosphere in helioseismology. Applied Mathematics and Computation, 361(??):177–197, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.

com/science/article/
pii/S0096300319303534.

Chen:2015:BDB

[CDS15] Zengqiang Chen, Matthias Dehmer, and Yongtang Shi. Bounds for degreebased network entropies. Applied Mathematics and Computation, 265(??): 983–993, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315007791.

Chen:2018:EDS

[CDZ^{+}18] Yi Chen, Shuai Ding, Handong Zheng, Youtao Zhang, and Shanlin Yang. Exploring diffusion strategies for mHealth promotion using evolutionary game model. Applied Mathematics and Computation, 336(??):148–161, November 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318303886.

Chen:2015:AEC

[CE15] Chao-Ping Chen and Neven Elezović. Asymptotic expansions and completely monotonic functions associated with the gamma, psi and polygamma functions. Applied Mathematics and Computation, 269(??): 232–241, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315009637.

Carmona:2015:ISC

[CEG^{+}15] A. Carmona, A. M. Encinas, S. Gago, M. J. Jiménez, and M. Mitjana. The inverses of some circulant matrices. Applied Mathematics and Computation, 270(??):785–793, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011406.

Cordero:2015:LCF

[CEHVT15] A. Cordero, J. A. Ezquerro, M. A. Hernández-Verón, and J. R. Torregrosa. On the local convergence of a fifth-order iterative method in Banach spaces. Applied Mathematics and Computation, 251(??):396–403, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300314016166.

Calabro:2016:RFF

[CEMR16] F. Calabrò, A. Corbo Esposito, G. Mantica, and T. Radice. Refinable functions, functionals, and iterated function systems. Applied Mathematics and Computation, 272 (part 1)(??):199–207, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007821.

Cvetkovic:2015:ESM

[CEP15] Lj. Cvetković, M. Erić, and J. M. Peña. Eventually SDD matrices and eigenvalue localization. Applied Mathematics and Computation, 252(??):535–540, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401666X.

Cosme:2018:MUA

[CFdCdS18] Iria C. S. Cosme, Isaac F.

Fernandes, João L. de Carvalho, and Samuel Xavier de Souza. Memory-usage advantageous block recursive matrix inverse. Applied Mathematics and Computation, 328(??):

125–136, July 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318300742.

Cabada:2015:CSS

[CFG15] Alberto Cabada and Carlos Fernández-Gómez. Constant sign solutions of twopoint fourth order problems. Applied Mathematics and Computation, 263(??):122–133, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004269.

Caliciotti:2018:PNC

[CFR18] Andrea Caliciotti, Giovanni Fasano, and Massimo Roma. Preconditioned Nonlinear Conjugate Gradient methods based on a modified secant equation. Applied Mathematics and Computation, 318(??):196–214, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305805.

Cao:2018:ESP

[CFS18] Shuchao Cao, Libi Fu, and Weiguo Song. Exit selection and pedestrian movement in a room with two exits under fire emergency. Applied Mathematics and Computation, 332(??):136–147, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302212.

Carmona:2017:IHC

[CFSGM17] V. Carmona, F. Fernández-Sánchez, and E. García-Medina. Including homoclinic connections and T-point heteroclinic cycles in the same global problem for a reversible family of piecewise linear systems. Applied Mathematics and Computation, 296(??):33–41, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306063.

Cao:2015:NPE

[CG15] Jiwei Cao and Zhihao Ge. A new posteriori error estimates for a discontinuous Galerkin method for a reactiondiffusion problem. Applied Mathematics and Computation, 256(??):735–741, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000466.

Chen:2019:PPI

[CG19] Zhong Chen and QianQian Gou. Piecewise Picard iteration method for solving nonlinear fractional differential equation with proportional delays. Applied Mathematics and Computation, 348(??):465–478, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830924X.

Cherifi:2019:GNQ

[CGAL19] Abdelmadjid Cherifi, Kevin Guelton, Laurent Arcese, and Valter J. S. Leite. Global non-quadratic Dstabilization of Takagi– Sugeno systems with piecewise continuous membership functions. Applied Mathematics and Computation, 351(??):23–36, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300402.

Cao:2019:SHB

[CGC19] Jinde Cao, Luca Guerrini, and Zunshui Cheng. Stability and Hopf bifurcation of controlled complex networks model with two delays. Applied Mathematics and Computation, 343(??):21–29, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307999.

Cai:2019:ROH

[CGCK19] Ruiyang Cai, Fudong Ge, YangQuan Chen, and Chunhai Kou. Regional observability for Hadamard–Caputo time fractional distributed parameter systems. Applied Mathematics and Computation, 360(??):190–202, November 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319303832.

Caceres:2018:SSL

[CGG^{+}18] José Cáceres, Delia Garijo, Antonio González, Alberto Márquez, María Luz Puertas, and Paula Ribeiro. Shortcut sets for the locus of plane Euclidean networks. Applied Mathematics and Computation, 334(??):192–205, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303229.

Campos:2016:NMS

[CGJV16] Beatriz Campos, Antonio Garijo, Xavier Jarque, and Pura Vindel. Newton method for symmetric quartic polynomial. Applied Mathematics and Computation, 290(??):326–335, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303927.

Chen:2018:MCW

[CGJZ18] Ya-Hong Chen, Daniel Gray, Ya-Lei Jin, and Xiao-Dong Zhang. On majorization of closed walk vectors of trees with given degree sequences. Applied Mathematics and Computation, 336(??):326–337, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304107.

Claverol:2017:SSR

[CGK^{+}17] Mercè Claverol, Delia Garijo, Matias Korman, Carlos Seara, and Rodrigo I. Silveira. Stabbing segments with rectilinear objects. Applied Mathematics and Computation, 309(??):359–373, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302369.

Chaoui:2015:SIE

[CGL15] A. Chaoui and A. Guezane-Lakoud. Solution to an integrodifferential equation with integral condition. Applied Mathematics and Computation, 266(??):903–908, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007808.

Castro:2016:AOS

[CGM16] Manuel J. Castro, José M.

Gallardo, and Antonio Marquina. Approximate Osher–Solomon schemes for hyperbolic systems. Applied Mathematics and Computation, 272 (part 2) (??):347–368, January 1, 2016. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315008851.

Cordero:2016:SAP

[CGMT16] Alicia Cordero, José M. Gutiérrez, Á. Alberto Magreñán, and Juan R. Torregrosa. Stability analysis of a parametric family of iterative methods for solving nonlinear models. Applied Mathematics and Computation, 285(??):26– 40, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302144.

Calcines:2015:GRH

[CGPR15] José M. García Calcines, José M. Gutiérrez, Luis J. Hernández Paricio, and M. Teresa Rivas Rodríguez. Graphical representations for the homogeneous bivariate Newton’s method. Applied Mathematics and Computation, 269(??):988–1006, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010243.

Chaves-Gonzalez:2015:DEP

[CGPT15] José M. Chaves-González and Miguel A. Pérez-Toledano. Differential evolution with Pareto tournament for the multiobjective next release problem. Applied Mathematics and Computation, 252(??):1–13, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016257.

Chen:2016:UEC

[CGW16] Huanzhen Chen, Lei Gao, and Hong Wang. Uniform estimates for characteristicsmixed finite method for transient advection-dominated diffusion problems in twodimensional space. Applied Mathematics and Computation, 280(??):86–102, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300315.

Cao:2019:SCS

[CGW19] Zhengbao Cao, Lijun Gao, and Gang Wang. Stability conditions of switched nonlinear systems with unstable subsystems and destabilizing switching behaviors. Applied Mathematics and Computation, 363(??): Article 124558, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305417.

Chen:2017:SFO

[CGWW17] Yuquan Chen, Qing Gao, Yiheng Wei, and Yong Wang. Study on fractional order gradient methods. Applied Mathematics and Computation, 314(??): 310–321, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317304745.

Chen:2017:FTD

[CGZ17] Guici Chen, Yu Gao, and Shasha Zhu. Finitetime dissipative control for stochastic interval systems with time-delay and Markovian switching. Applied Mathematics and Computation, 310(??):169–181, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317302850.

Campos:2018:NIL

[CH18a] Rafael G. Campos and Adolfo Huet. Numerical inversion of the Laplace transform and its application to fractional diffusion. Applied Mathematics and Computation, 327(??):70– 78, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830047X.

Chen:2018:PPE

[CH18b] Hongbo Chen and Tianliang Hou. A priori and a posteriori error estimates of H_{1}-Galerkin mixed finite element methods for optimal control problems governed by pseudo-hyperbolic integrodifferential equations. Applied Mathematics and Computation, 328(??): 100–112, July 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318300638.

Cheviakov:2018:SCF

[CH18c] A. F. Cheviakov and J. Heß. A symbolic computation framework for constitutive modelling based on entropy principles. Applied Mathematics and Computation, 324(??):105–118, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308573.

Cao:2019:GFV

[CH19] Shujuan Cao and Shenwei Huang. Graphs without five-vertex path and four-vertex cycle. Applied Mathematics and Computation, 363(??): Article 124602, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305946.

Chamakuri:2019:PST

[Cha19] Nagaiah Chamakuri. Parallel and space-time adaptivity for the numerical simulation of cardiac action potentials. Applied Mathematics and Computation, 353(??):406–417, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300803.

Chen:2015:IAEb

[Che15a] Chao-Ping Chen. Inequalities and asymptotic expansions associated with the Ramanujan and Nemes formulas for the gamma function. Applied Mathematics and Computation, 261(??):337–350, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004610.

Chen:2015:ICM

[Che15b] Chao-Ping Chen. Inequalities and completely monotonic functions associated with the ratios of functions resulting from the gamma function. Applied Mathematics and Computation, 259(??):790–799, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003148.

Chen:2015:EHH

[Che15c] Feixiang Chen. Extensions of the Hermite– Hadamard inequality for harmonically convex functions via fractional integrals. Applied Mathematics and Computation, 268(??):121–128, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008322.

Chen:2015:PSS

[Che15d]  Guanwei Chen.  Pe-

riodic solutions of superquadratic damped vibration problems. Applied Mathematics and Computation, 270(??):794–801, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315011510.

Chen:2015:SAS

[Che15e] Shyh-Feng Chen. Stability analysis and stabilization of 2-D singular Roesser models. Applied Mathematics and Computation, 250(??):779–791, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015070.

Chentouf:2015:SMT

[Che15f] Boumediène Chentouf. Stabilization of memory type for a rotating diskbeam system. Applied Mathematics and Computation, 258(??):227–236, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000788.

Chen:2016:MPI

[Che16a] Chao-Ping Chen. Monotonicity properties, inequalities and asymptotic expansions associated with the gamma function. Applied Mathematics and Computation, 283(??):385–396, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301515.

Chen:2016:AEG

[Che16b]  Chao-Ping Chen.  On

the asymptotic expansions of the gamma function related to the Nemes, Gosper and Burnside formulas. Applied Mathematics and Computation, 276(??):417–431, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315016057.

Chen:2016:HOM

[Che16c] Liang Chen. A highorder modified Levenberg– Marquardt method for systems of nonlinear equations with fourth-order convergence. Applied Mathematics and Computation, 285(??):79–93, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316302314.

Chen:2018:RFH

[Che18a] Chao-Ping Chen. Ramanujan’s formula for the harmonic number. Applied Mathematics and Computation, 317(??):121–128, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306112.

Chen:2018:LRP

[Che18b] He Chen. Long rainbow paths and rainbow cycles in edge colored graphs — a survey. Applied Mathematics and Computation, 317(??):187–192, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0096300317306252.

Chen:2019:NTN

[Che19] Y. Z. Chen. Numerical test for the Neumann problem of interior BVP of plane elasticity. Applied Mathematics and Computation, 361(??):61–73, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930414X.

Chiou:2015:CPP

[Chi15] Suh-Wen Chiou. A cutting plane projection method for bi-level area traffic control optimization with uncertain travel demand. Applied Mathematics and Computation, 266(??):390–403, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006256.

Chen:2015:DFT

[CHL15] Zhaohui Chen, Qi Huang, and Zhang Liu. Decentralized finite-time H_{∞} filtering for interconnected Markovian jump system with interval modedependent delays. Applied Mathematics and Computation, 258(??):138–154, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000971.

Cao:2017:FSM

[CHL17] Fuyuan Cao, Joshua Zhexue Huang, and Jiye Liang. A fuzzy SV-k-modes algorithm for clustering categorical data with setvalued attributes. Applied Mathematics and Computation, 295(??):1– 15, February 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316305938.

Chen:2018:SNS

[CHL18] Xiaodan Chen, Yaoping Hou, and Fenggen Lin. Some new spectral bounds for graph irregularity. Applied Mathematics and Computation, 320(??):331–340, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306665.

Chourushi:2018:HRE

[Cho18] T. Chourushi. A high resolution equi-gradient scheme for convective flows. Applied Mathematics and Computation, 338(??):115–122, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304788.

Choi:2019:NSI

[Cho19a] Hyung Jun Choi. A numerical solution for the inhomogeneous Dirichlet boundary value problem on a non-convex polygon. Applied Mathematics and Computation, 341:31–45, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307252.

Chourushi:2019:PMC

[Cho19b] T. Chourushi. Proposition of modified convection boundedness criterion and its evaluation for the development of bounded schemes. Applied Mathematics and Computation, 346(??):710–739, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308506.

Chakrone:2016:MRL

[CHRD16] O. Chakrone, El Miloud Hssini, M. Rahmani, and O. Darhouche. Multiplicity results for a p-Laplacian discrete problems of Kirchhoff type. Applied Mathematics and Computation, 276(??): 310–315, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315015891.

Chydzinski:2016:ASM

[Chy16] Andrzej Chydzinski. Analysis of the scheduling mechanism for virtualization of links with partial isolation. Applied Mathematics and Computation, 281(??):39–54, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300479.

Cihangir:2015:PNT

[Cih15] Apo Cihangir. On 0 /1polytopes with nonobtuse triangulations. Applied Mathematics and Computation, 267(??):17–27, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315009285.

Cimpoiasu:2018:NCA

[Cim18] Rodica Cimpoiasu. New candidates for arbitragefree stock price models via generalized conditional symmetry method. Applied Mathematics and Computation, 333(??): 460–466, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318302960.

Craciun:2015:GGR

[CIPT15] Ioana Craciun, Daniela Inoan, Dorian Popa, and Lucian Tudose. Generalized golden ratios defined by means. Applied Mathematics and Computation, 250(??):221–227, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014702.

Chen:2015:EAS

[CJ15] Zhong Chen and Wei Jiang. An efficient algorithm for solving nonlinear Volterra–Fredholm integral equations. Applied Mathematics and Computation, 259(??):614–619, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002799.

Clavero:2016:UCO

[CJ16] C. Clavero and J. C. Jorge. Uniform convergence and order reduction of the fractional implicit Euler method to solve singularly perturbed 2D reaction-diffusion problems. Applied Mathematics and Computation, 287– 288(??):12–27, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302673.

Chen:2017:CCC

[CJ17a] He Chen and Zemin Jin. Coupon coloring of cographs. Applied Mathematics and Computation, 308(??):90–95, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317302059.

Chuong:2017:CHS

[CJ17b] T. D. Chuong and V. Jeyakumar. Convergent hierarchy of SDP relaxations for a class of semiinfinite convex polynomial programs and applications. Applied Mathematics and Computation, 315(??):381–399, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305374.

Chang:2018:ENM

[CJ18] Lili Chang and Zhen Jin. Efficient numerical methods for spatially extended population and epidemic models with time delay. Applied Mathematics and Computation, 316(??):138–154, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305799.

Cai:2018:EVS

[CJG^{+}18] Yongli Cai, Jianjun Jiao, Zhanji Gui, Yuting Liu, and Weiming Wang. Environmental variability in a stochastic epidemic model. Applied Mathematics and Computation, 329(??):210–226, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301036.

Choi:2017:MSC

[CJK17] Yongho Choi, Darae Jeong, and Junseok Kim. A multigrid solution for the Cahn–Hilliard equation on nonuniform grids. Applied Mathematics and Computation, 293(??):320–333, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305215.

Cermak:2019:EVD

[CJN19] Jan Cermák, Jirí Jánský, and Ludek Nechvátal. Exact versus discretized stability regions for a linear delay differential equation. Applied Mathematics and Computation, 347(??):712–722, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310002.

Ji:2015:HOC

[cJzS15] Cui cui Ji and Zhi zhong Sun. The high-order compact numerical algorithms for the two-dimensional fractional sub-diffusion equation. Applied Mathematics and Computation, 269(??):775–791, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010103.

Chamakuri:2017:PDA

[CK17] Nagaiah Chamakuri and Karl Kunisch. Primal-dual active set strategy for large scale optimization of cardiac defibrillation. Applied Mathematics and Computation, 292(??):178–193, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304738.

Curt:2018:SGP

[CK18] Paula Curt and Mirela Kohr. Some geometrical properties of free boundaries in the Hele– Shaw flows. Applied Mathematics and Computation, 323(??):86–94, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308330.

Cvetkovic:2016:ENE

[CKDC16] Ljiljana Cvetković, Vladimir Kostić, Ksenija Doroslovacki, and Dragana Lj. Cvetković. Euclidean norm estimates of the inverse of some special block matrices. Applied Mathematics and Computation, 284(??):12–23, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301679.

Chebotarev:2016:NOC

[CKG^{+}16] Alexander Yu. Chebotarev, Andrey E. Kovtanyuk, Gleb V. Grenkin, Nikolai D. Botkin, and Karl-Heinz Hoffmann. Nondegeneracy of optimality conditions in control problems for a radiativeconductive heat transfer model. Applied Mathematics and Computation, 289(??):371–380, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303459.

Cheichan:2019:WGF

[CKG19] Mohammed S. Cheichan, Hashim A. Kashkool, and Fuzheng Gao. A weak Galerkin finite element method for solving nonlinear convection-diffusion problems in two dimensions. Applied Mathematics and Computation, 354(??):149–163, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930147X.

Chara:2016:SFF

[CKKF16] Zdenek Chara, Bohus Kysela, Jiri Konfrst, and Ivan Fort. Study of fluid flow in baffled vessels stirred by a Rushton standard impeller. Applied Mathematics and Computation, 272 (part 3)(??):614–628, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008255.

Costarelli:2019:QPO

[CKSV19] D. Costarelli, A. Krivoshein, M. Skopina, and G. Vinti. Quasi-projection operators with applications to differential-difference expansions. Applied Mathematics and Computation, 363(??):Article 124623, December 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319306150.

Chen:2015:NAS

[CKW15] Yiming Chen, Xiaohong Ke, and Yanqiao Wei. Numerical algorithm to solve system of nonlinear fractional differential equations based on wavelets method and the error analysis. Applied Mathematics and Computation, 251(??):475–488, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016117.

Cai:2017:SSE

[CKW17] Yongli Cai, Yun Kang, and Weiming Wang. A stochastic SIRS epidemic model with nonlinear incidence rate. Applied Mathematics and Computation, 305(??):221–240, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300991.

Chen:2015:BUG

[CL15a] Lin Chen and Jinfeng Liu. The bipartite unicyclic graphs with the first ⌊(n − 3)/4⌋ largest matching energies. Applied Mathematics and Computation, 268(??):644–656, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008966.

Chen:2015:SNN

[CL15b] Yiren Chen and Rui Liu. Some new nonlinear wave solutions for two (3 + 1)-dimensional equations. Applied Mathematics and Computation, 260(??):397–411, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004129.

Che:2016:CGM

[CL16a] Haitao Che and Meixia Li. The conjugate gradient method for split variational inclusion and constrained convex minimization problems. Applied Mathematics and Computation, 290(??):426–438, November 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316303782.

Chen:2016:EVM

[CL16b] Lin Chen and Jinfeng Liu. Extremal values of matching energies of one class of graphs. Applied Mathematics and Computation, 273(??):976–992, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013673.

Chen:2016:PCB

[CL16c] Y. M. Chen and J. K. Liu. A precise calculation of bifurcation points for periodic solution in nonlinear dynamical systems. Applied Mathematics and Computation, 273(??):1190–1195, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315012114.

Christara:2016:OPJ

[CL16d] Christina C. Christara and Nat Chun-Ho Leung. Option pricing in jump diffusion models with quadratic spline collocation. Applied Mathematics and Computation, 279(??):28–42, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300333.

Cheng:2018:SPC

[CL18] Kai Cheng and Zhenzhou Lu. Sparse polynomial chaos expansion based on D-MORPH regression. Applied Mathematics and Computation, 323(??):17–30, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308263.

Canuto:2019:MTR

[CL19a] Claudio Canuto and Andrea Lo Giudice. A multitimestep Robin–Robin domain decomposition method for time dependent advection–diffusion problems. Applied Mathematics and Computation, 363(??):Article 124596, December 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319305880.

Cao:2019:BTP

[CL19b] Yang Cao and Sen Li. Block triangular preconditioners based on symmetric-triangular decomposition for generalized saddle point problems. Applied Mathematics and Computation, 358(??):262–277, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319303273.

Chu:2019:OBM

[CL19c] Xiaoan Chu and Muguo Li. Observer-based model following sliding mode tracking control of discretetime linear networked systems with two-channel event-triggered schemes and quantizations. Applied Mathematics and Computation, 355(??):428–448, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302036.

Chai:2019:BMS

[CLC^{+}19] Xudong Chai, Liwei Liu, Baoxian Chang, Tao Jiang, and Zhen Wang. On a batch matching system with impatient servers and boundedly rational customers. Applied Mathematics and Computation, 354(??):308–328, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930102X.

Chen:2017:MPC

[CLCW17] Jian Chen, Chong Lin, Bing Chen, and Qing-Guo Wang. Mixed H_{∞} and passive control for singular systems with time delay via static output feedback. Applied Mathematics and Computation, 293(??):244–253, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305240.

Chang:2015:NGF

[CLHC15] Che-Jung Chang, Der-Chiang Li, Yi-Hsiang Huang, and Chien-Chih Chen. A novel gray forecasting model based on the box plot for small manufacturing data sets. Applied Mathematics and Computation, 265(??):400–408, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500613X.

Cheng:2016:CDH

[CLHC16] Qing Cheng, Zhong Liu, Jincai Huang, and Guangquan Cheng. Community detection in hypernetwork via Density-Ordered Tree partition. Applied Mathematics and Computation, 276(??):384–393, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315300278.

Chen:2015:FSI

[CLJ^{+}15] S. Chen, F. Liu, X. Jiang, I. Turner, and V. Anh. A fast semi-implicit difference method for a nonlinear two-sided spacefractional diffusion equation with variable diffusivity coefficients. Applied Mathematics and Computation, 257(??):591–601, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314011254.

Celis:2015:DBR

[CLL^{+}15] Oliver Salazar Celis, Lingzhi Liang, Damiaan Lemmens, Jacques Tempère, and Annie Cuyt. Determining and benchmarking risk neutral distributions implied from option prices. Applied Mathematics and Computation, 258(??):372–387, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001642.

Cao:2016:SRI

[CLL16] Jianxin Cao, Yiping Luo, and Guanghui Liu. Some results for impulsive fractional differential inclusions with infinite delay and sectorial operators in Banach spaces. Applied Mathematics and Computation, 273(??):237–257, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013107.

Chen:2017:PIG

[CLL17] Lin Chen, Xueliang Li, and Jinfeng Liu. The kproper index of graphs. Applied Mathematics and Computation, 296(??):57– 63, March 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300316306233.

Cordero:2015:SFH

[CLMT15] Alicia Cordero, Taher Lotfi, Katayoun Mahdiani, and Juan R. Torregrosa. A stable family with high order of convergence for solving nonlinear equations. Applied Mathematics and Computation, 254(??):240–251, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314018025.

Cipolatti:2018:BVP

[CLP^{+}18] R. Cipolatti, I.-S. Liu, L. A. Palermo, M. A. Rincon, and R. M. S. Rosa. A boundary value problem arising from nonlinear viscoelasticity: Mathematical analysis and numerical simulations. Applied Mathematics and Computation, 335(??):237–247, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303527.

Chen:2018:ASM

[CLPY18] Chuan Chen, Lixiang Li, Haipeng Peng, and Yixian Yang. Adaptive synchronization of memristorbased BAM neural networks with mixed delays. Applied Mathematics and Computation, 322(??):100–110, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308196.

Chen:2016:FRC

[CLQ16] Haibin Chen, Guoyin Li, and Liqun Qi. Further results on Cauchy tensors and Hankel tensors. Applied Mathematics and Computation, 275(??):50–62, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015532.

Christophe:2015:MEA

[CLR15]  Alexandra  Christophe,

Yann Le Bihan, and Francesca Rapetti. A mortar element approach on overlapping non-nested grids: Application to eddy current non-destructive testing. Applied Mathematics and Computation, 267(??):71–82, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007572.

Cao:2016:PVS

[CLR16a] Jiling Cao, Guanghua Lian, and Teh Raihana Nazirah Roslan. Pricing variance swaps under stochastic volatility and stochastic interest rate. Applied Mathematics and Computation, 277(??):72–81, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315300151.

Chan:2016:HMC

[CLR16b] C. K. Chan, W. P. Loh, and I. Abd Rahim. Human motion classification using 2D stick-model matching regression coefficients. Applied Mathematics and Computation, 283(??):70– 89, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301436.

Chatzarakis:2015:ORD

[CLS15] G. E. Chatzarakis, M. Lafci, and I. P. Stavroulakis. Oscillation results for difference equations with several oscillating coefficients. Applied Mathematics and Computation, 251(??):81–91, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401577X.

Choi:2019:CLS

[CLSS19] Hayoung Choi, Hosoo Lee, Yifei Shen, and Yuanming Shi. Comparing large-scale graphs based on quantum probability theory. Applied Mathematics and Computation, 358(??):1–15, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302693.

Chung:2015:AOC

[CLT^{+}15] Kun-Jen Chung, Jui-Jung Liao, Pin-Shou Ting, Shy-Der Lin, and H. M. Srivastava. The algorithm for the optimal cycle time and pricing decisions for an integrated inventory system with order-size dependent trade credit in supply chain management. Applied Mathematics and Computation, 268(??):322–333, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008206.

Coll:2018:WDI

[CLT18] C. Coll, M. Lattanzi, and N. Thome. Weighted G-Drazin inverses and a new pre-order on rectangular matrices. Applied Mathematics and Computation, 317(??):12–24, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305982.

Cheung:2018:JMT

[CLW18] Eric C. K. Cheung, Haibo Liu, and Gordon E. Willmot. Joint moments of the total discounted gains and losses in the renewal risk model with two-sided jumps. Applied Mathematics and Computation, 331(??):358–377, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302108.

Chen:2019:SRS

[CLWM19] Yuanyuan Chen, Dan Li, Zhiwen Wang, and Jixiang Meng. A_{α}-spectral radius of the second power of a graph. Applied Mathematics and Computation, 359(??):418–425, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303790.

Cen:2017:RNM

[CLX17] Zhongdi Cen, Anbo Le, and Aimin Xu. A robust numerical method for a fractional differential equation. Applied Mathematics and Computation, 315(??):445–452, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305623.

Chang:2017:LAI

[CLXW17] Xiao-Heng Chang, Zhi-Min Li, Jun Xiong, and Yi-Ming Wang. LMI approaches to input and output quantized feedback stabilization of linear systems. Applied Mathematics and Computation, 315(??):162–175, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S009630031730499X.

Cai:2017:LEP

[CLY17] Jiaxiang Cai, Hua Liang, and Bin Yang. Local energy-preserving algorithms for nonlinear fourth-order Schrödinger equation with trapped term. Applied Mathematics and Computation, 296 (??):23–32, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316306099.

Chen:2019:DFC

[CLZW19] Jiaxi Chen, Junmin Li, Rui Zhang, and Chengzhou Wei. Distributed fuzzy consensus of uncertain topology structure multiagent systems with nonidentical partially unknown control directions. Applied Mathematics and Computation, 362(??): Article 124581, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305739.

Chen:2015:GSS

[CM15a] Cai-Rong Chen and Chang-Feng Ma. A generalized shift-splitting preconditioner for singular saddle point problems. Applied Mathematics and Computation, 269(??):947–955, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010759.

Chow:2015:RRG

[CM15b] Chak-On Chow and Toufik Mansour. On the realrootedness of generalized Touchard polynomials. Applied Mathematics and Computation, 254(??):204–209, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401769X.

Correale:2016:DAM

[CM16] T. G. Correale and L. H. A.

Monteiro. On the dynamics of axonal membrane: Ion channel as the basic unit of a deterministic model. Applied Mathematics and Computation, 291(??):292–302, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300316304362.

Campos:2017:QIT

[CM17a] Rafael G. Campos and Francisco Marcellán. Quadratures and integral transforms arising from generating functions. Applied Mathematics and Computation, 297(??):8–18, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306592.

Correale:2017:TFC

[CM17b] T. G. Correale and L. H. A. Monteiro. Typical frequencycurrent curves of neurons obtained from a model based on cellular automaton. Applied Mathematics and Computation, 304(??):136–141, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300607.

Carr:2018:SAS

[CM18] Elliot J. Carr and Nathan G. March. Semi-analytical solution of multilayer diffusion problems with timevarying boundary conditions and general interface conditions. Applied Mathematics and Computation, 333(??):286–303, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302765.

Canovas:2018:DKC

[CMG18] J. S. Cánovas and M. Muñoz-Guillermo. On the dynamics of Kopel’s Cournot duopoly model. Applied Mathematics and Computation, 330(??):292–306, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830153X.

Cheng:2017:PDE

[CMM17] Eddie Cheng, Colton Magnant, and Dhruv Medarametla. Proper distance in edge-colored hypercubes. Applied Mathematics and Computation, 313(??):384–391, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303776.

Castillo:2017:POP

[CMR17] Kenier Castillo, Francisco Marcellán, and Jorge Rivero. On perturbed orthogonal polynomials on the real line and the unit circle via Szegő’s transformation. Applied Mathematics and Computation, 302(??):97–110, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300267.

Cakic:2019:CB

[CMS19] Nenad Cakić, Toufik Mansour, and Armend Sh. Shabani. Circular bargraphs. Applied Mathematics and Computation, 347(??):803–807, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310208.

Chen:2018:DNG

[CMSZ18] Lily Chen, Yingbin Ma, Yongtang Shi, and Yan Zhao. On the [1, 2]- domination number of generalized Petersen graphs. Applied Mathematics and Computation, 327(??):1–7, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318300262.

Chen:2018:CHV

[CMT^{+}18] Laihuan Chen, Jixiang Meng, Yingzhi Tian, Xiaodong Liang, and Fengxia Liu. Connectivity of half vertex transitive digraphs. Applied Mathematics and Computation, 316(??):25–29, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730557X.

Cieslinski:2015:FED

[CMW15] Jan L. Cieśliński, Leonid V.

Moroz, and Cezary J. Walczyk. Fast exact digital differential analyzer for circle generation. Applied Mathematics and Computation, 271(??):68–79, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011601.

Cai:2017:ASM

[CMW17] Li-Ming Cai, Chairat Modnak, and Jin Wang. An age-structured model for cholera control with vaccination. Applied Mathematics and Computation, 299(??):127–140, April 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306816.

Chien:2015:DRH

[CN15a] M. T. Chien and H. Nakazato. Determinantal representations of hyperbolic forms via weighted shift matrices. Applied Mathematics and Computation, 258(??):172–181, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001721.

Chun:2015:AFM

[CN15b] Changbum Chun and Beny Neta. An analysis of a family of Maheshwaribased optimal eighth order methods. Applied Mathematics and Computation, 253(??):294–307, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017184.

Chun:2015:BAS

[CN15c] Changbum Chun and Beny Neta. Basins of attraction for several third order methods to find multiple roots of nonlinear equations. Applied Mathematics and Computation, 268(??):129–137, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008498. See

corrigendum [NCS16].

Chun:2015:CBA

[CN15d] Changbum Chun and Beny Neta. Comparing the basins of attraction for Kanwar– Bhatia–Kansal family to the best fourth order method. Applied Mathematics and Computation, 266(??):277–292, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006931.

Chun:2016:AKO

[CN16a] Changbum Chun and Beny Neta. An analysis of a Khattri’s 4th order family of methods. Applied Mathematics and Computation, 279(??):198–207, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S009630031630025X.

Chun:2016:CSF

[CN16b] Changbum Chun and Beny Neta. Comparison of several families of optimal eighth order methods. Applied Mathematics and Computation, 274(??):762–773, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015386.

Chien:2019:HFA

[CN19a] M. T. Chien and H. Nakazato. Hyperbolic forms admit reversible weighted shift determinantal representation. Applied Mathematics and Computation, 362 (??):Article 124595, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305879.

Chun:2019:DHO

[CN19b] Changbum Chun and Beny Neta. Developing high order methods for the solution of systems of nonlinear equations. Applied Mathematics and Computation, 342(??):178–190, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308105.

Cimrman:2018:CSI

[CNK^{+}18] Robert Cimrman, Matyás Novák, Radek Kolman, Miroslav Tuma, Jiří Plešek, and Jiří Vackár. Convergence study of isogeometric analysis based on Bézier extraction in electronic structure calculations. Applied Mathematics and Computation, 319(??):138–152, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301273.

Cristescu:2016:SIF

[CNNA16] Gabriela Cristescu, Muhammad Aslam Noor, Khalida Inayat Noor, and Muhammad Uzair Awan. Some inequalities for functions having Orliczconvexity. Applied Mathematics and Computation, 273(??):226–236, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013065.

Cortes:2018:SSO

[CNQRR18] J.-C. Cortés, A. Navarro-Quiles, J.-V. Romero, and M.-D. Roselló. Solving second-order linear differential equations with random analytic coefficients about ordinary points: A full probabilistic solution by the first probability density function. Applied Mathematics and Computation, 331(??):33–45, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301619.

Cuyt:2018:FCC

[CnTVsL18] Annie Cuyt, Min nan Tsai, Marleen Verhoye, and Wen shin Lee. Faint and clustered components in exponential analysis. Applied Mathematics and Computation, 327(??):93–103, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307804.

Chatzarakis:2015:ODE

[CÖ15] G. E. Chatzarakis and Özkan Öcalan. Oscillations of difference equations with non-monotone retarded arguments. Applied Mathematics and Computation, 258(??):60– 66, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001423.

Covei:2019:SSE

[Cov19] Dragos-Patru Covei. Symmetric solutions for an elliptic partial differential equation that arises in stochastic production planning with production constraints. Applied Mathematics and Computation, 350(??):190–197, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300232.

Candela:2015:RMR

[CP15a] V. Candela and R. Peris.

The rate of multiplicity of the roots of nonlinear equations and its application to iterative methods. Applied Mathematics and Computation, 264(??):417–430, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005548.

Chen:2015:IAEa

[CP15b] Chao-Ping Chen and Richard B. Paris. Inequalities, asymptotic expansions and completely monotonic functions related to the gamma function. Applied Mathematics and Computation, 250(??):514–529, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031401515X.

Clempner:2015:CSN

[CP15c] Julio B. Clempner and Alexander S. Poznyak. Computing the strong Nash equilibrium for Markov chains games. Applied Mathematics and Computation, 265(??):911–927, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500781X.

Conti:2015:TDC

[CP15d] Monica Conti and Vittorino Pata. On the timedependent Cattaneo law in space dimension one. Applied Mathematics and Computation, 259(??):32– 44, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002192.

Chen:2017:AEP

[CP17] Chao-Ping Chen and Richard B. Paris. On the asymptotic expansions of products related to the Wallis, Weierstrass, and Wilf formulas. Applied Mathematics and Computation, 293(??):30–39, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304982.

Candela:2019:CTO

[CP19] V. Candela and R. Peris.

A class of third order iterative Kurchatov–Steffensen (derivative free) methods for solving nonlinear equations. Applied Mathematics and Computation, 350(??):93–104, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311032.

Chen:2017:SMS

[CPCQ17] Xiangyong Chen, Ju H.

Park, Jinde Cao, and Jianlong Qiu. Sliding mode synchronization of multiple chaotic systems with uncertainties and disturbances. Applied Mathematics and Computation, 308(??):161–173, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302163.

Cornelio:2015:CLS

[CPP15] Anastasia Cornelio, Federica Porta, and Marco Prato. A convergent least-squares regularized blind deconvolution approach. Applied Mathematics and Computation, 259(??):173–186, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002362.

Cococcioni:2018:LMO

[CPS18] Marco Cococcioni, Massimo Pappalardo, and Yaroslav D. Sergeyev. Lexicographic multi-objective linear programming using grossone methodology: Theory and algorithm. Applied Mathematics and Computation, 318(??):298–311, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303703.

Cao:2016:TPR

[CPST16] Hui Cao, Sergei V. Pereverzyev, Ian H. Sloan, and Pavlo Tkachenko. Two-parameter regularization of ill-posed spherical pseudo-differential equations in the space of continuous functions. Applied Mathematics and Computation, 273(??): 993–1005, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315014046.

Cao:2015:TVP

[CR15] Yang Cao and Zhi-Ru Ren. Two variants of the PMHSS iteration method for a class of complex symmetric indefinite linear systems. Applied Mathematics and Computation, 264(??):61–71, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500507X.

Chen:2018:NCF

[CR18] Shuangshuang Chen and Hongxing Rui. A nodecentered finite volume method for a fracture model on triangulations. Applied Mathematics and Computation, 327(??):55– 69, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300456.

Cardinali:2019:CII

[CR19a] T. Cardinali and P. Rubbioni. The controllability of an impulsive integro-differential process with nonlocal feedback controls. Applied Mathematics and Computation, 347(??):29–39, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309573.

Chen:2019:AIC

[CR19b] Min Chen and André Raspaud. Acyclic improper choosability of subcubic graphs. Applied Mathematics and Computation, 356(??):92–98, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302280.

Chao:2015:FSC

[CRC15] Fugang Chao, Han Ren, and Ni Cao. Finding a shortest cycle in a subspace of the cycle space of a graph. Applied Mathematics and Computation, 268(??):393–398, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008346.

Cao:2019:NFB

[CRC19] Fayun Cao, Han Ren, and Hanlin Chen. New formulae for the bipartite vertex frustration and decycling number of graphs. Applied Mathematics and Computation, 347(??):101–112, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309561.

Casas-Ramirez:2018:ASB

[CRCVMS18] Martha-Selene Casas-Ramírez, José-Fernando Camacho-Vallejo, and Iris-Abril Martínez-Salazar. Approximating solutions to a bilevel capacitated facility location problem with customer’s patronization toward a list of preferences. Applied Mathematics and Computation, 319(??):369–386, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302357.

Crisci:2019:SSG

[CRZ19] Serena Crisci, Valeria Ruggiero, and Luca Zanni. Steplength selection in gradient projection methods for box-constrained quadratic programs. Applied Mathematics and Computation, 356(??): 312–327, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300319302462.

Chesnokov:2015:SAS

[CS15a] Alexander A. Chesnokov and Irina V. Stepanova. Stability analysis of shear flows in a Hele–Shaw cell. Applied Mathematics and Computation, 265(??):320–328, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006360.

Cui:2015:SCE

[CS15b] Huanhuan Cui and Menglong Su. On sufficient conditions ensuring the norm convergence of an iterative sequence to zeros of accretive operators. Applied Mathematics and Computation, 258(??):67– 71, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500140X.

Campoamor-Stursberg:2016:GFE

[CS16a] R. Campoamor-Stursberg.

Generating functions and existence of contact symmetries of third order scalar ordinary differential equations. Applied Mathematics and Computation, 273(??):1179–1189, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012126.

Chen:2016:DBM

[CS16b] Yan-Cheng Chen and Chao-Ton Su. Distancebased margin support vector machine for classification. Applied Mathematics and Computation, 283(??):141–152, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301357.

Couceiro:2016:NFO

[CS16c] Micael Couceiro and Seenith Sivasundaram. Novel fractional order particle swarm optimization. Applied Mathematics and Computation, 283(??):36–54, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300911.

Cao:2017:CMD

[CS17] Yueju Cao and Jitao Sun. Controllability of measure driven evolution systems with nonlocal conditions. Applied Mathematics and Computation, 299(??):119–126, April 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307135.

Chentouf:2018:SAN

[CS18] Boumediène Chentouf and Nejib Smaoui. Stability analysis and numerical simulations of a one dimensional open channel hydraulic system. Applied Mathematics and Computation, 321(??):498–511, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307671.

Chaudhary:2019:RPE

[CS19] J. P. Chaudhary and L. P. Singh. Riemann problem and elementary wave interactions in dusty gas. Applied Mathematics and Computation, 342(??):147–165, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307951.

Chydzinski:2018:BRF

[CSA18] Andrzej Chydzinski, Dominik Samociuk, and Blazej Adamczyk. Burst ratio in the finite-buffer queue with batch Poisson arrivals. Applied Mathematics and Computation, 330(??):225–238, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301231.

Cui:2019:USM

[CSL19] Beibei Cui, Xinmin Song, and Xiyu Liu. Unbiased steady minimum-variance estimation for systems with measurement-delay and unknown inputs. Applied Mathematics and Computation, 356(??): 379–391, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319302437.

Cortes:2015:PEP

[CST^{+}15] Juan-Carlos Cortés, Francisco-J. Santonja, Ana-C. Tarazona, Rafael-J. Villanueva, and Javier Villanueva-Oller. A probabilistic estimation and prediction technique for dynamic continuous social science models: the evolution of the attitude of the Basque Country population towards ETA as a case study. Applied Mathematics and Computation, 264(??):13–20, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004476.

Ceretani:2018:AME

[CST18] Andrea N. Ceretani, Natalia N. Salva, and Domingo A. Tarzia. Approximation of the modified error function. Applied Mathematics and Computation, 337(??):591–606, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304715.

Cordero:2017:FKT

[CSTH17] Alicia Cordero, Fazlollah Soleymani, Juan R. Torregrosa, and F. Khaksar Haghani. A family of Kurchatov-type methods and its stability. Applied Mathematics and Computation, 294(??):264–279, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305811.

Cermak:2019:EFM

[CSV19] M. Cermák, S. Sysala, and J. Valdman. Efficient and flexible MAT-LAB implementation of 2D and 3D elastoplastic problems. Applied Mathematics and Computation, 355(??):595–614, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301584.

Chang:2015:RDS

[CSWJ15] Lili Chang, Gui-Quan Sun, Zhen Wang, and Zhen Jin. Rich dynamics in a spatial predator–prey model with delay. Applied Mathematics and Computation, 256(??):540–550, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500082X.

Cao:2015:FMF

[CSZ15] Limei Cao, Xinhui Si, and Liancun Zheng. The flow of a micropolar fluid through a porous expanding channel: a Lie group analysis. Applied Mathematics and Computation, 270(??):242–250, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010280.

Chen:2019:SEL

[CSZ19] Yang Chen, Jakub Sikorowski, and Mengkun Zhu. Smallest eigenvalue of large Hankel matrices at critical point: Comparing conjecture with parallelised computation. Applied Mathematics and Computation, 363(??): Article 124628, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306204.

Cuong:2015:GTS

[CT15] Dao Huy Cuong and Mai Duc Thanh. A Godunov-type scheme for the isentropic model of a fluid flow in a nozzle with variable crosssection. Applied Mathematics and Computation, 256(??):602–629, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000387.

Cicek:2016:SSM

[ÇT16] Y. Çiçek and G. Tanoğlu.

Strang splitting method for Burgers–Huxley equation. Applied Mathematics and Computation, 276(??):454–467, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300175.

Cabada:2017:LDE

[CT17a] Alberto Cabada and F. Adrián F. Tojo. On linear differential equations and systems with reflection. Applied Mathematics and Computation, 305(??):84–102, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031730053X.

Crepnjak:2017:SIW

[CT17b] Matevz Crepnjak and Niko Tratnik. The Szeged index and the Wiener index of partial cubes with applications to chemical graphs. Applied Mathematics and Computation, 309(??):324–333, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302552.

Cui:2017:SSL

[CT17c] Shu-Yu Cui and Gui-Xian Tian. The spectra and the signless Laplacian spectra of graphs with pockets. Applied Mathematics and Computation, 315(??):363–371, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305179.

Cuong:2017:CGT

[CT17d] Dao Huy Cuong and Mai Duc Thanh. Constructing a Godunov-type scheme for the model of a general fluid flow in a nozzle with variable cross-section. Applied Mathematics and Computation, 305(??):136–160, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300917.

Coskun:2018:NBP

[CT18] Arzu Coskun and Necati Taskara. A note on the bi-periodic Fibonacci and Lucas matrix sequences. Applied Mathematics and Computation, 320(??):400–406, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306598.

Chan:2019:RFR

[CTT^{+}19] Joseph Chang Lun Chan, Chee Pin Tan, Hieu Trinh, Md Abdus Samad Kamal, and Yeong Shiong Chiew. Robust fault reconstruction for a class of non-infinitely observable descriptor systems using two sliding mode observers in cascade. Applied Mathematics and Computation, 350(??):78–92, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311329.

Cui:2015:CAC

[Cui15] Ming Rong Cui. Convergence analysis of compact difference schemes for diffusion equation with nonlocal boundary conditions. Applied Mathematics and Computation, 260(??):227–241, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003446.

Clavero:2018:NAT

[CVA18] C. Clavero and J. Vigo-Aguiar. Numerical approximation of 2D time dependent singularly perturbed convection-diffusion problems with attractive or repulsive turning points. Applied Mathematics and Computation, 317(??): 223–233, January 15, 2018.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317306173.

C:2019:MTD

[CVJ19] Ewin Sánchez C. and Pedro Vega-Jorquera. Modelling temporal decay of aftershocks by a solution of the fractional reactive equation. Applied Mathematics and Computation, 340(??):43–49, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307367.

Cruz-Victoria:2015:SNF

[CVMGPPGC15] Juan C. Cruz-Victoria, Rafael Martínez-Guerra, Claudia A. Pérez-Pinacho, and Gian Carlo Gómez-Cortés. Synchronization of nonlinear fractional order systems by means of PI^{rα} reduced order observer. Applied Mathematics and Computation, 262(??):224–231, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004397.

Civin:2018:UEM

[CvRvdM18] D. Civin, N. F. J. van Rensburg, and A. J. van der Merwe. Using energy methods to compare linear vibration models. Applied Mathematics and Computation, 321(??):602–613, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307816.

Chen:2017:PPM

[CW17] Wen Chen and Song Wang. A power penalty method for a 2D fractional partial differential linear complementarity problem governing two-asset American option pricing. Applied Mathematics and Computation, 305(??):174–187, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300954.

Chudy:2018:DRP

[CW18] Filip Chudy and Pawel Woźny. Differentialrecurrence properties of dual Bernstein polynomials. Applied Mathematics and Computation, 338(??):520–536, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305289.

Chen:2016:EVS

[CWD16] Hanlin Chen, Renfang Wu, and Hanyuan Deng. The extremal values of some topological indices in bipartite graphs with a given matching number. Applied Mathematics and Computation, 280(??):103–109, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630042X.

Chen:2019:MPD

[CWD^{+}19] Xiaoli Chen, Fengyan Wu, Jinqiao Duan, Jürgen Kurths, and Xiaofan Li. Most probable dynamics of a genetic regulatory network under stable Lévy noise. Applied Mathematics and Computation, 348(??):425–436, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310464.

Chen:2015:GAM

[CWFL15] Hongyan Chen, Wendi Wang, Rui Fu, and Jianfeng Luo. Global analysis of a mathematical model on malaria with competitive strains and immune responses. Applied Mathematics and Computation, 259(??):132–152, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002738.

Cai:2015:CTS

[CWG15] Jiaxiang Cai, Yushun Wang, and Yuezheng Gong. Convergence of time-splitting energyconserved symplectic schemes for 3D Maxwell’s equations. Applied Mathematics and Computation, 265(??):51–67, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005895.

Chen:2015:RSS

[CWHY15] Liping Chen, Ranchao Wu, Yigang He, and Lisheng Yin. Robust stability and stabilization of fractional-order linear systems with polytopic uncertainties. Applied Mathematics and Computation, 257(??):274–284, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017640.

Chen:2016:MTC

[CWL16] Hongwei Chen, Bo Wu, and Jianquan Lu. A minimum-time control for Boolean control networks with impulsive disturbances. Applied Mathematics and Computation, 273(??):477–483, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013132.

Chen:2019:NEP

[CWL19] Hao Chen, Xiaoli Wang, and Xiaolin Li. A note on efficient preconditioner of implicit Runge–Kutta methods with application to fractional diffusion equations. Applied Mathematics and Computation, 351(??):116–123, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300319300505.

Chen:2018:SSD

[CWLA18] Yonggang Chen, Zidong Wang, Yurong Liu, and Fuad E. Alsaadi. Stochastic stability for distributed delay neural networks via augmented Lyapunov– Krasovskii functionals. Applied Mathematics and Computation, 338(??): 842–868, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318304764.

Chen:2017:NEG

[CWZ17] Ya-Hong Chen, Hua Wang, and Xiao-Dong Zhang. Note on extremal graphs with given matching number. Applied Mathematics and Computation, 308(??):149–156, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031730190X.

Chen:2019:DDI

[CX19] Zhenwu Chen and Zhiting Xu. A delayed diffusive influenza model with two-strain and two vaccinations. Applied Mathematics and Computation, 349(??):439–453, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311263.

Chen:2019:EOT

[CXK19] Ruyun Chen, Shuhuang Xiang, and Xuesong Kuang. On evaluation of oscillatory transforms from position to momentum space. Applied Mathematics and Computation, 344–345(??):183–190, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830883X.

Cen:2015:HFD

[CXL15] Zhongdi Cen, Aimin Xu, and Anbo Le. A hybrid finite difference scheme for pricing Asian options. Applied Mathematics and Computation, 252(??):229–239, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016610.

Chen:2016:DBT

[CXL16a] Ailian Chen, Xianzhu Xiong, and Fenggen Lin. Distance-based topological indices of the treelike polyphenyl systems. Applied Mathematics and Computation, 281(??): 233–242, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316300662.

Chen:2016:ERB

[CXL16b] Ailian Chen, Xianzhu Xiong, and Fenggen Lin. Explicit relation between the Wiener index and the edge-Wiener index of the catacondensed hexagonal systems. Applied Mathematics and Computation, 273(??):1100–1106, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315014149.

Chen:2018:WCM

[CXL18] Xiangling Chen, Ziqing Xie, and Jianshu Luo. A wavelet collocation method for boundary integral equations of the modified Helmholtz equation. Applied Mathematics and Computation, 321(??):300–312, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307361.

Chen:2016:ETS

[CXMP16] Fengde Chen, Xiangdong Xie, Zhanshuai Miao, and Liqiong Pu. Extinction in two species nonautonomous nonlinear competitive system. Applied Mathematics and Computation, 274(??):119–124, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014198.

Chang:2016:FRD

[CXP16] Xiao-Heng Chang, Jun Xiong, and Ju H. Park. Fuzzy robust dynamic output feedback control of nonlinear systems with linear fractional parametric uncertainties. Applied Mathematics and Computation, 291(??):213–225, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304313.

Chi:2018:LIR

[CXW18] Meiling Chi, Fuyi Xu, and Yonghong Wu. A logarithmic improvement of regularity criterion for the MHD equations in terms of the pressure. Applied Mathematics and Computation, 327(??):46–54, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300286.

Cui:2019:ETS

[CXW^{+}19] Lili Cui, Xiangpeng Xie, Xiaowei Wang, Yanhong Luo, and Jingbo Liu. Event-triggered singlenetwork ADP method for constrained optimal tracking control of continuoustime non-linear systems. Applied Mathematics and Computation, 352(??): 220–234, July 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319300839.

Chen:2018:IDD

[CXZZ18] Guoliang Chen, Jianwei Xia, Guangming Zhuang, and Junsheng Zhao. Improved delay-dependent stabilization for a class of networked control systems with nonlinear perturbations and two delay components. Applied Mathematics and Computation, 316(??):1–17, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305337.

Cao:2015:MCC

[CY15a] Xiaodong Cao and Xu You.

Multiple-correction and continued fraction approximation (II). Applied Mathematics and Computation, 261(??):192–205, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004208.

Cheng:2015:SSC

[CY15b] Yongkuan Cheng and Yangxin Yao. Soliton solutions to a class of relativistic nonlinear Schrödinger equations. Applied Mathematics and Computation, 260(??):342–350, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003690.

Cichon:2015:CDS

[CY15c] Mieczysław Cichoń and Ahmet Yantir. On continuous dependence of solutions of dynamic equations. Applied Mathematics and Computation, 252(??):473–483, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017019.

Cao:2016:PPM

[CY16a] Huiping Cao and Lan Yao. A partitioned PSB method for partially separable unconstrained optimization problems. Applied Mathematics and Computation, 290(??):164–177, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303800.

Cao:2016:CUP

[CY16b] Yang Cao and Shi-Chao Yi. A class of Uzawa-PSS iteration methods for nonsingular and singular non-Hermitian saddle point problems. Applied Mathematics and Computation, 275(??):41–49, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015519.

Cheng:2018:TWS

[CY18] Hongmei Cheng and Rong Yuan. Traveling waves of some Holling–Tanner predator–prey system with nonlocal diffusion. Applied Mathematics and Computation, 338(??):1–11, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303679.

Chang:2015:FPA

[CYCW15] Yu-Huei Chang, Jinn-Shyong Yang, Jou-Ming Chang, and Yue-Li Wang. A fast parallel algorithm for constructing independent spanning trees on parity cubes. Applied Mathematics and Computation, 268(??):489–495, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008620.

Chen:2018:PCB

[CYGL18] Ya-Shan Chen, Han-Xin Yang, Wen-Zhong Guo, and Geng-Geng Liu. Promotion of cooperation based on swarm intelligence in spatial public goods games. Applied Mathematics and Computation, 320(??):614–620, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730721X.

Cai:2015:EFP

[CYLW15] Yongli Cai, Yuan Yuan, Xinze Lian, and Weiming Wang. Extinction in a Feline Panleukopenia virus model incorporating direct and indirect transmissions. Applied Mathematics and Computation, 258(??):358–366, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000351.

Chen:2016:FTA

[CYSY16] Mengshen Chen, Xiaofei Yang, Hao Shen, and Fengqi Yao. Finite-time asynchronous H_{∞} control for Markov jump repeated scalar non-linear systems with input constraints. Applied Mathematics and Computation, 275(??):172–180, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015726.

Chen:2018:VPS

[CYW18] Min Chen, Weiqiang Yu, and Weifan Wang. On the vertex partitions of sparse graphs into an independent vertex set and a forest with bounded maximum degree. Applied Mathematics and Computation, 326(??):117–123, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300092.

Cai:2019:UBC

[CYY19] Jiansheng Cai, Chunhua Yang, and Jiguo Yu. An upper bound for the choice number of star edge coloring of graphs. Applied Mathematics and Computation, 348(??):588–593, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310634.

Cai:2017:CCG

[CYZ17] Jiansheng Cai, Guiying Yan, and Xia Zhang. The classification of f-coloring of graphs with large maximum degree. Applied Mathematics and Computation, 313(??):119–121, November 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317303715.

Chen:2015:QHO

[CZ15a] Ruyun Chen and Xiaoliang Zhou. On quadrature of highly oscillatory integrals with logarithmic singularities. Applied Mathematics and Computation, 265(??):973–982, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500795X.

Chen:2015:TCP

[CZ15b] Shangdi Chen and Xiaolian Zhang. Three constructions of perfect authentication codes from projective geometry over finite fields. Applied Mathematics and Computation, 253(??):308–317, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300314017494.

Cui:2015:EUT

[CZ15c] Yujun Cui and Yumei Zou. An existence and uniqueness theorem for a second order nonlinear system with coupled integral boundary value conditions. Applied Mathematics and Computation, 256(??):438–444, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001009.

Cai:2016:AKT

[CZ16] Qing-Bo Cai and Guorong Zhou. On (p, q)-analogue of Kantorovich type Bernstein– Stancu–Schurer operators. Applied Mathematics and Computation, 276(??):12– 20, March 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315016070.

Chen:2017:OSL

[CZ17a] Mingzhu Chen and Bo Zhou.

On ordinary and signless Laplacian spectral radius of graphs with fixed number of branch vertices. Applied Mathematics and Computation, 305(??):166–173, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300929.

Cheng:2017:NCC

[CZ17b] Li Cheng and Xinlong Zhou. Necessary conditions for the convergence of subdivision schemes with finite masks. Applied Mathematics and Computation, 303(??):34–41, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300206.

Cui:2017:GRI

[CZ17c] Qing Cui and Lingping Zhong. The general Randić index of trees with given number of pendent vertices. Applied Mathematics and Computation, 302(??):111–121, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300292.

Cai:2018:ISP

[CZ18a] Jinming Cai and Zhaowen Zheng. Inverse spectral problems for discontinuous Sturm–Liouville problems of Atkinson type. Applied Mathematics and Computation, 327(??):22– 34, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300237.

Cheng:2018:ECD

[CZ18b] Yuan Cheng and Xiaoping Zheng. Emergence of cooperation during an emergency evacuation. Applied Mathematics and Computation, 320(??):485–494, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307002.

Cui:2019:PTC

[CZ19] Li Cui and Jin-Xin Zhou.

Petersen type n-circulant and weak metacirculant. Applied Mathematics and Computation, 358(??): 251–261, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319303418.

Chen:2018:BPS

[CZL18] Hao Chen, Tongtong Zhang, and Wen Lv. Block preconditioning strategies for time-space fractional diffusion equations. Applied Mathematics and Computation, 337(??):34–40, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304041.

Chen:2018:NDF

[CZLM18] Baoqin Chen, Zhenyu Zhao, Zhi Li, and Zehong Meng. Numerical differentiation by a Fourier extension method with super-order regularization. Applied Mathematics and Computation, 334(??):1–10, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303096.

Chang:2018:GML

[CZLS18] Wenting Chang, Song Zhu, Jinyu Li, and Kaili Sun. Global Mittag-Leffler stabilization of fractional-order complexvalued memristive neural networks. Applied Mathematics and Computation, 338(??):337–345, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305307.

Chen:2018:DRD

[CZM18] Weimin Chen, Baoyong Zhang, and Qian Ma. Decay-rate-dependent conditions for exponential stability of stochastic neutral systems with Markovian jumping parameters. Applied Mathematics and Computation, 321(??):93–105, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307336.

Chu:2019:RBP

[CZM^{+}19] Chen Chu, Yao Zhai, Chunjiang Mu, Die Hu, Tong Li, and Lei Shi. Reputation-based popularity promotes cooperation in the spatial prisoner’s dilemma game. Applied Mathematics and Computation, 362(??): Article 124493, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304722.

Chang:2015:WPA

[CZP15] Yong-Kui Chang, Mei-Juan Zhang, and Rodrigo Ponce. Weighted pseudo almost automorphic solutions to a semilinear fractional differential equation with Stepanovlike weighted pseudo almost automorphic nonlinear term. Applied Mathematics and Computation, 257(??):158–168, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314011345.

Chen:2015:ESC

[CZS15] Hao Chen, Shouming Zhong, and Jinliang Shao. Exponential stability criterion for interval neural networks with discrete and distributed delays. Applied Mathematics and Computation, 250(??):121–130, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014647.

Chen:2017:MCV

[CZSH17] Xiaofeng Chen, Zhenjiang Zhao, Qiankun Song, and Jin Hu. Multistability of complex-valued neural networks with timevarying delays. Applied Mathematics and Computation, 294(??):18–35, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305574.

Dastjerdi:2017:NSN

[DA17] H. Laeli Dastjerdi and M. Nili Ahmadabadi. The numerical solution of nonlinear two-dimensional Volterra–Fredholm integral equations of the second kind based on the radial basis functions approximation with error analysis. Applied Mathematics and Computation, 293(??):545–554, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305586.

Dastjerdi:2019:EMN

[DA19] H. Laeli Dastjerdi and M. Nili Ahmadabadi. An efficient method for the numerical solution of Hammerstein mixed VF integral equations on 2D irregular domains. Applied Mathematics and Computation, 344–345(??):46–56, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308634.

Dai:2016:CNC

[Dai16] Zhifeng Dai. Comments on a new class of nonlinear conjugate gradient coefficients with global convergence properties. Applied Mathematics and Computation, 276(??):297–300, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015878.

Danchick:2015:SEU

[Dan15a] Roy Danchick. Simplified existence and uniqueness conditions for the zeros and the concavity of the F and G functions of improved Gauss orbit determination from two position vectors. Applied Mathematics and Computation, 269(??):279–287, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009194.

Danov:2015:AFI

[Dan15b] Krassimir D. Danov. Asymptotic formulae for the interaction force and torque between two charged parallel cylinders. Applied Mathematics and Computation, 256(??):642–655, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001113.

Danchick:2018:DUB

[Dan18] Roy Danchick. On determining the unknown bandparameter and truncated sinc series coefficients from a time sampled bandlimited function. Applied Mathematics and Computation, 334(??):60–79, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302911.

DelaSen:2015:SSE

[DAQI15] M. De la Sen, S. Alonso-Quesada, and A. Ibeas. On the stability of an SEIR epidemic model with distributed timedelay and a general class of feedback vaccination rules. Applied Mathematics and Computation, 270(??):953–976, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315011558.

Darwish:2016:EKF

[Dar16] Mohamed Abdalla Darwish. On Erdélyi– Kober fractional Urysohn– Volterra quadratic integral equations. Applied Mathematics and Computation, 273(??):562–569, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013910.

Das:2016:GGI

[Das16] Kinkar Ch. Das. On the Graovac–Ghorbani index of graphs. Applied Mathematics and Computation, 275(??):353–360, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501588X.

Daraby:2017:GRI

[DAS17] Bayaz Daraby, Hassan Ghazanfary Asll, and Ildar Sadeqi. General related inequalities to Carlsontype inequality for the Sugeno integral. Applied Mathematics and Computation, 305(??):323–329, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301017.

Despres:2016:SWB

[DB16] Bruno Després and Christophe Buet. The structure of well-balanced schemes for Friedrichs systems with linear relaxation. Applied Mathematics and Computation, 272 (part 2)(??): 440–459, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315005470.

Dzurina:2017:PKS

[DB17] J. Dzurina and B. Baculíková.

Properties of Kneser solutions for third-order differential equations. Applied Mathematics and Computation, 307(??):165–169, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301662.

Dassios:2018:CRF

[DB18] Ioannis K. Dassios and Dumitru I. Baleanu. Caputo and related fractional derivatives in singular systems. Applied Mathematics and Computation, 337(??):583–590, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304089.

Du:2017:AQD

[DBJC17] Peibing Du, Roberto Barrio, Hao Jiang, and Lizhi Cheng. Accurate quotient-difference algorithm: Error analysis, improvements and applications. Applied Mathematics and Computation, 309(??):245–271, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302394.

Dhawan:2015:GLS

[DBK15] S. Dhawan, Samir Kumar Bhowmik, and Sheo Kumar. Galerkin-least square B-spline approach toward advection-diffusion equation. Applied Mathematics and Computation, 261(??):128–140, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315004063.

Dewasurendra:2018:MDD

[DBV18] Mangalagama Dewasurendra, Mathew Baxter, and Kuppalapalle Vajravelu. A method of directly defining the inverse mapping for solutions of non-linear coupled systems arising in convection heat transfer in a second grade fluid. Applied Mathematics and Computation, 339(??):750–757, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305733.

Dumbser:2016:CWN

[DC16] Michael Dumbser and Vincenzo Casulli. A conservative, weakly nonlinear semi-implicit finite volume scheme for the compressible Navier–Stokes equations with general equation of state. Applied Mathematics and Computation, 272 (part 2)(??): 479–497, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315010978.

Dharshini:2015:ACO

[DCA15] A. Priya Dharshini, S. Chandrakumarmangalam, and G. Arthi. Ant colony optimization for competency based learning objects sequencing in elearning. Applied Mathematics and Computation, 263(??):332–341, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005299.

Calvo-Garrido:2015:EJD

[dCCGV15] María del Carmen Calvo-Garrido and Carlos Vázquez. Effects of jump-diffusion models for the house price dynamics in the pricing of fixed-rate mortgages, insurance and coinsurance. Applied Mathematics and Computation, 271(??):730–742, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315012898.

Du:2019:RFE

[DCJ19] Dongsheng Du, Vincent Cocquempot, and Bin Jiang. Robust fault estimation observer design for switched systems with unknown input. Applied Mathematics and Computation, 348(??):70–83, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310087.

Diaz:2018:MSM

[DCK^{+}18] Paul Diaz, Paul Constantine, Kelsey Kalmbach, Eric Jones, and Stephen Pankavich. A modified SEIR model for the spread of Ebola in Western Africa and metrics for resource allocation. Applied Mathematics and Computation, 324(??):141–155, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308214.

Chen:2016:RCC

[dCM16] Xiao diao Chen and Weiyin Ma. Rational cubic clipping with linear complexity for computing roots of polynomials. Applied Mathematics and Computation, 273(??):1051–1058, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315014058.

Deniz:2015:SHD

[DÇO15] Erhan Deniz, Murat Çaglar, and Halit Orhan. Second Hankel determinant for bi-starlike and biconvex functions of order β. Applied Mathematics and Computation, 271(??):301–307, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501228X.

Diblik:2015:UIS

[DCR15] Josef Diblík, Radoslav Chupác, and Miroslava Ruzicková. Unbounded increasing solutions of a system of difference equations with delays. Applied Mathematics and Computation, 251(??):489–498, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016075.

Deng:2015:WNA

[DCS15] Kaiying Deng, Minghua Chen, and Tieli Sun. A weighted numerical algorithm for two and three dimensional twosided space fractional wave equations. Applied Mathematics and Computation, 257(??):264–273, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314011333.

Dehmer:2019:ENS

[DCS^{+}19] Matthias Dehmer, Zengqiang Chen, Yongtang Shi, Yusen Zhang, Shailesh Tripathi, Modjtaba Ghorbani, Abbe Mowshowitz, and Frank Emmert-Streib. On efficient network similarity measures. Applied Mathematics and Computation, 362(??): Article 124521, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305004.

daCosta:2017:CEP

[dCSS17] A. Pinto da Costa, A. Seeger, and F. M. F. Simões. Complementarity eigenvalue problems for nonlinear matrix pencils. Applied Mathematics and Computation, 312(??):134–148, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300317303260.

Dahne:2019:EAZ

[DCT19] J. Dahne, M. F. Ciappina, and W. Tucker. Enclosing all zeros of a system of analytic functions. Applied Mathematics and Computation, 348(??):513–522, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318307550.

Dai:2015:MPC

[DCW15] Zhifeng Dai, Xiaohong Chen, and Fenghua Wen. A modified Perry’s conjugate gradient methodbased derivative-free method for solving large-scale nonlinear monotone equations. Applied Mathematics and Computation, 270(??):378–386, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010693.

Djordjevic:2015:CGM

[DD15] Gospava B. Djordjević and Snezana S. Djordjević. Convolutions of the generalized Morgan–Voyce polynomials. Applied Mathematics and Computation, 259(??):106–115, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001940.

Das:2016:CBZ

[DD16a] Kinkar Ch. Das and Matthias Dehmer. Comparison between the zerothorder Randić index and the sum-connectivity index. Applied Mathematics and Computation, 274(??):585–589, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015180.

Dhamija:2016:JDO

[DD16b] Minakshi Dhamija and Naokant Deo. Jain– Durrmeyer operators associated with the inverse Pólya–Eggenberger distribution. Applied Mathematics and Computation, 286(??):15–22, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302089.

Delavar:2019:WTI

[DD19] M. Rostamian Delavar and S. S. Dragomir. Weighted trapezoidal inequalities related to the area balance of a function with applications. Applied Mathematics and Computation, 340(??):5–14, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318307380.

Das:2019:IMH

[DDD19] Parthasakha Das, Pritha Das, and Samhita Das. An investigation on Monod-Haldane immune response based tumor-effector-interleukin-2 interactions with treatments. Applied Mathematics and Computation, 361(??):536–551, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304254.

Dimitrov:2016:SPT

[DDdF16] Darko Dimitrov, Zhibin Du, and Carlos M. da Fonseca. On structural properties of trees with minimal atom-bond connectivity index III: Trees with pendent paths of length three. Applied Mathematics and Computation, 282(??):276–290, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301229.

Lillo:2018:F

[DDF18] Silvana De Lillo, Marina Dolfin, and Gioia Fioriti. “Follow the leader” learning dynamics on networks. Applied Mathematics and Computation, 332(??):316–328, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302248.

deSouzaParolin:2015:ECS

[ddGL15] Radael de Souza Parolin, Antônio José da Silva Neto, Pedro Paulo Gomes Watts Rodrigues, and Orestes Llanes Santiago. Estimation of a contaminant source in an estuary with an inverse problem approach. Applied Mathematics and Computation, 260(??):331–341, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315003689.

DellAccio:2018:RFH

[DDH18] Francesco Dell’Accio, Filomena Di Tommaso, and Kai Hormann. Reconstruction of a function from Hermite–Birkhoff data. Applied Mathematics and Computation, 318(??):51– 69, February 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317303727.

Deo:2016:SKO

[DDM16] Naokant Deo, Minakshi Dhamija, and Dan Miclaus. Stancu–Kantorovich operators based on inverse Pólya–Eggenberger distribution. Applied Mathematics and Computation, 273(??):281–289, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013405.

deDeus:2019:ASM

[dDSI19] Hilbeth P. Azikri de Deus, Thales A. Barbosa Pinto Silva, and Mikhail Itskov. An analysis on the shear modulus of the modified Jeffreys model. Applied Mathematics and Computation, 346(??):649–665, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308956.

Dattoli:2015:PBF

[DDSL15] Giuseppe Dattoli, Emanuele Di Palma, Elio Sabia, and Silvia Licciardi. Products of Bessel functions and associated polynomials. Applied Mathematics and Computation, 266(??):507–514, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007092. See

corrigendum [DSDL19].

Dai:2015:TAR

[DDY15] Hong-Liang Dai, Ting Dai, and Xiang Yan. Thermoelastic analysis for rotating circular HSLA steel plates with variable thickness. Applied Mathematics and Computation, 268(??):1095–1109, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009297.

DeMicheli:2017:FAI

[De 17] Enrico De Micheli. A fast algorithm for the inversion of Abel’s transform. Applied Mathematics and Computation, 301(??):12– 24, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307354.

DeLeone:2018:NPG

[De 18] Renato De Leone. Nonlinear programming and Grossone: Quadratic programing and the role of constraint qualifications. Applied Mathematics and Computation, 318(??):290–297, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302114.

DeChant:2019:CDS

[DeC19] Lawrence DeChant. A cavity depth sensitized Rossiter mode formula. Applied Mathematics and Computation, 347(??): 143–148, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/
pii/S0096300318309147.

Duque:2017:ELF

[DEdlT17] Daniel Duque, Pep Español, and Jaime Arturo de la Torre. Extending linear finite elements to quadratic precision on arbitrary meshes. Applied Mathematics and Computation, 301(??):201–213, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307366.

Dehghan:2016:EMA

[Deh16] M. Dehghan. An efficient method to approximate eigenfunctions and highindex eigenvalues of regular Sturm–Liouville problems. Applied Mathematics and Computation, 279(??):249–257, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300261.

Djurcic:2015:SCR

[DEK15] Dragan Djurcić, Nebojsa Elez, and Ljubisa D. R. Kocinac. On a subclass of the class of rapidly varying sequences. Applied Mathematics and Computation, 251(??):626–632, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016324.

Deng:2015:GSG

[Den15]  Chunyuan Deng.  Ge-

ometry structures of the generalized inverses of block two-by-two matrices. Applied Mathematics and Computation, 250(??):479–491, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015136.

Dehmer:2015:GDM

[DESS15] Matthias Dehmer, Frank Emmert-Streib, and Yongtang Shi. Graph distance measures based on topological indices revisited. Applied Mathematics and Computation, 266(??):623–633, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006967.

daFonseca:2015:ECC

[dF15] Carlos M. da Fonseca.

Eigenvalues of a certain class of tridiagonal matrices: a brief comment. Applied Mathematics and Computation, 262(??):111–112, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004695.

Dumbser:2016:STD

[DF16] M. Dumbser and M. Facchini. A space–time discontinuous Galerkin method for Boussinesqtype equations. Applied Mathematics and Computation, 272 (part 2)(??): 336–346, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315008334.

Dilna:2018:SSM

[DF18] N. Dilna and M. Feckan.

The Stieltjes string model with external load. Applied Mathematics and Computation, 337(??):335–349, November 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300318304351.

Michele:2019:MSO

[DFG^{+}19] F. Di Michele, E. Felaco, I. Gasser, A. Serbinovskiy, and H. Struchtrup. Modeling, simulation and optimization of a pressure retarded osmosis power station. Applied Mathematics and Computation, 353(??):189–207, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031930061X.

deFrutos:2018:EAP

[dFGAN18] Javier de Frutos, Bosco García-Archilla, and Julia Novo. Error analysis of projection methods for non inf–sup stable mixed finite elements. The transient Stokes problem. Applied Mathematics and Computation, 322(??):154–173, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308287.

deFrutos:2019:GDS

[dFGAN19] Javier de Frutos, Bosco García-Archilla, and Julia Novo. Grad-div stabilization for the timedependent Boussinesq equations with inf-sup stable finite elements. Applied Mathematics and Computation, 349(??):281–291, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311238.

Dongyang:2016:USA

[DFJ16] Shi Dongyang, Yan Fengna, and Wang Junjun. Unconditional superconvergence analysis of a new mixed finite element method for nonlinear Sobolev equation. Applied Mathematics and Computation, 274(??):182–194, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012229.

daFonseca:2015:MTK

[dFKMM15] Carlos M. da Fonseca, Saïd Kouachi, Dan A. Mazilu, and Irina Mazilu. A multitemperature kinetic Ising model and the eigenvalues of some perturbed Jacobi matrices. Applied Mathematics and Computation, 259(??):205–211, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002519.

Diblik:2015:NPS

[DFP15] J. Diblík, M. Feckan, and M. Pospísil. Nonexistence of periodic solutions and S-asymptotically periodic solutions in fractional difference equations. Applied Mathematics and Computation, 257(??):230–240, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016415.

DellAcqua:2015:AMG

[DFSC15] Pietro Dell’Acqua, Antonio Frangioni, and Stefano Serra-Capizzano. Accelerated multigrid for graph Laplacian operators. Applied Mathematics and Computation, 270(??):193–215, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010887.

daFonseca:2015:SCT

[dFY15] Carlos M. da Fonseca and Fatih Yilmaz. Some comments on k-tridiagonal matrices: Determinant, spectra, and inversion. Applied Mathematics and Computation, 270(??): 644–647, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315011443.

Danca:2015:SCD

[DG15a] Marius-F. Danca and Roberto Garrappa. Suppressing chaos in discontinuous systems of fractional order by active control. Applied Mathematics and Computation, 257(??):89–102, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016312.

Dubey:2015:CIM

[DG15b] Bhaskar Dubey and Raju K.

George. Controllability of impulsive matrix Lyapunov systems. Applied Mathematics and Computation, 254(??):327–339, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017950.

Deng:2018:ECG

[DG18a] An-Wen Deng and Chih-Ying Gwo. Efficient computations for generalized Zernike moments and image recovery. Applied Mathematics and Computation, 339(??):294–301, December 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318305873.

Deng:2018:FSA

[DG18b] An-Wen Deng and Chih-Ying Gwo. Fast and stable algorithms for highorder Pseudo Zernike moments and image reconstruction. Applied Mathematics and Computation, 334(??):239–253, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303059.

Das:2017:SRE

[DGF17] Kinkar Ch. Das, Ivan Gutman, and Boris Furtula. On spectral radius and energy of extended adjacency matrix of graphs. Applied Mathematics and Computation, 296(??):116–123, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306361.

Du:2015:ABP

[DGL^{+}15] Wen-Bo Du, Yang Gao, Chen Liu, Zheng Zheng, and Zhen Wang. Adequate is better: particle swarm optimization with limitedinformation. Applied Mathematics and Computation, 268(??):832–838, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008437.

Dahlberg:2017:NNE

[DGLE17] Joen Dahlberg, Maud Göthe-Lundgren, and Stefan Engevall. A note on the nonuniqueness of the Equal Profit Method. Applied Mathematics and Computation, 308(??):84– 89, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317301923.

Crescenzo:2016:CTB

[DGN16] Antonio Di Crescenzo, Virginia Giorno, and Amelia G. Nobile. Constructing transient birthdeath processes by means of suitable transformations. Applied Mathematics and Computation, 281(??):152–171, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300674.

Das:2015:RBD

[DGNA15] Kinkar Ch. Das, Ivan Gutman, and Mohammad J. Nadjafi-Arani. Relations between distance-based and degree-based topological indices. Applied Mathematics and Computation, 270(??):142–147, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011169.

Dumbser:2016:PSI

[DGRR16] Michael Dumbser, Gregor Gassner, Christian Rohde, and Sabine Roller. Preface to the special issue “Recent Advances in Numerical Methods for Hyperbolic Partial Differential Equations”. Applied Mathematics and Computation, 272 (part 2)(??):235–236, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501485X.

Deng:2015:GBS

[DGZ15] Guifeng Deng, Fengjie Geng, and Yun Zhang. Global behavior of solutions of the generalized Lyness difference equations under quadratic perturbations. Applied Mathematics and Computation, 259(??):579–586, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003008.

Debbar:2016:ICO

[DGZ16] R. Debbar, H. Guebbai, and Z. Zereg. Improving the convergence order of the regularization method for Fredholm integral equations of the second kind. Applied Mathematics and Computation, 289(??):204–213, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303253.

Deng:2015:SSC

[DH15] Zhiying Deng and Yisheng Huang. Symmetric solutions for a class of singular biharmonic elliptic systems involving critical exponents. Applied Mathematics and Computation, 264(??):323–334, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005615.

Dong:2017:OFF

[DH17] Jiuxiang Dong and Junteng Hou. Output feedback fault-tolerant control by a set-theoretic description of T-S fuzzy systems. Applied Mathematics and Computation, 301(??):117–134, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307433.

Davila:2019:TFV

[DH19a] Randy Davila and Michael A. Henning. Total forcing versus total domination in cubic graphs. Applied Mathematics and Computation, 354(??):385–395, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S009630031930164X.

Delgado:2019:TOF

[DH19b] G. Delgado and M. Hamdaoui. Topology optimization of frequency dependent viscoelastic structures via a level-set method. Applied Mathematics and Computation, 347(??):522–541, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309871.

Dong:2015:DTI

[DHMT15] Yueping Dong, Gang Huang, Rinko Miyazaki, and Yasuhiro Takeuchi. Dynamics in a tumor immune system with time delays. Applied Mathematics and Computation, 252(??):99–113, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016282.

Daraghmeh:2019:BMR

[DHQ19] Adnan Daraghmeh, Carsten Hartmann, and Naji Qatanani. Balanced model reduction of linear systems with nonzero initial conditions: Singular perturbation approximation. Applied Mathematics and Computation, 353(??):295–307, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300992.

Diblik:2019:FGS

[DHS19] J. Diblík, H. Halfarová, and J. Safarík. Formulas for the general solution of weakly delayed planar linear discrete systems with constant coefficients and their analysis. Applied Mathematics and Computation, 358(??):363–381, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302760.

Ding:2019:TVO

[DHW^{+}19] Meng Ding, Ting-Zhu Huang, Si Wang, Jin-Jin Mei, and Xi-Le Zhao. Total variation with overlapping group sparsity for deblurring images under Cauchy noise. Applied Mathematics and Computation, 341:128–147, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318307288.

Dai:2019:NFE

[DHX^{+}19] Mingcheng Dai, Zhengguo Huang, Jianwei Xia, Bo Meng, Jian Wang, and Hao Shen. Non-fragile extended dissipativity-based state feedback control for 2-D Markov jump delayed systems. Applied Mathematics and Computation, 362(??): Article 124571, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305545.

Deng:2018:PCN

[DHY^{+}18] Wenfeng Deng, Keke Huang, Chunhua Yang, Hongqiu Zhu, and Zhaofei Yu. Promote of cooperation in networked multiagent system based on fitness control. Applied Mathematics and Computation, 339(??):779–804, December 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318306416.

Dimitrijevic:2016:CSM

[Dim16] Ivan Dimitrijevic. Cosmological solutions in modified gravity with monomial nonlocality. Applied Mathematics and Computation, 285(??):195–203, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302302.

Dimitrov:2017:SPT

[Dim17]  Darko Dimitrov.  On

structural properties of trees with minimal atombond connectivity index IV: Solving a conjecture about the pendent paths of length three. Applied Mathematics and Computation, 313(??):418–430, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304265.

Disibuyuk:2015:FGI

[Dis15]  Çetin Disibüyük. A func-

tional generalization of the interpolation problem. Applied Mathematics and Computation, 256(??):247–251, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315000168.

Dimitrov:2017:RGG

[DIS17] Darko Dimitrov, Barbara Ikica, and Riste Skrekovski. Remarks on the Graovac–Ghorbani index of bipartite graphs. Applied Mathematics and Computation, 293(??): 370–376, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316305422.

Djordjevic:2015:BSV

[DJ15] Jasmina Djordjević and Svetlana Janković. Backward stochastic Volterra integral equations with additive perturbations. Applied Mathematics and Computation, 265(??): 903–910, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315007018.

Dzurina:2018:OTO

[DJ18] Jozef Dzurina and Irena Jadlovská. Oscillation of third-order differential equations with noncanonical operators. Applied Mathematics and Computation, 336(??):394–402, November 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318303618.

Deng:2017:HOF

[DJL17] Dingwen Deng, Yaolin Jiang, and Dong Liang. High-order finite difference methods for a second order dual-phase-lagging models of microscale heat transfer. Applied Mathematics and Computation, 309(??):31–48, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302199.

Deng:2017:FED

[DJLX17] Chongyang Deng, Wenbiao Jin, Yajuan Li, and Huixia Xu. A formula for estimating the deviation of a binary interpolatory subdivision curve from its data polygon. Applied Mathematics and Computation, 304(??):10– 19, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300498.

Deng:2019:ZSP

[DJW19] Xinyang Deng, Wen Jiang, and Zhen Wang. Zerosum polymatrix games with link uncertainty: a Dempster–Shafer theory solution. Applied Mathematics and Computation, 340(??):101–112, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830746X.

Dac:2015:PAD

[DK15a] Nguyen Van Dac and Tran Dinh Ke. Pullback attractor for differential evolution inclusions with infinite delays. Applied Mathematics and Computation, 265(??):667–680, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500716X.

Damarla:2015:NSM

[DK15b] Seshu Kumar Damarla and Madhusree Kundu. Numerical solution of multi-order fractional differential equations using generalized triangular function operational matrices. Applied Mathematics and Computation, 263(??):189–203, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005093.

Delgado:2015:SGA

[DK15c] Paul Delgado and Vinod Kumar. A stochastic Galerkin approach to uncertainty quantification in poroelastic media. Applied Mathematics and Computation, 266(??):328–338, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006001.

Dhar:2016:LTI

[DK16] Sougata Dhar and Qingkai Kong. Lyapunov-type inequalities for higher order half-linear differential equations. Applied Mathematics and Computation, 273(??):114–124, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013284.

Danca:2017:PSS

[DK17] Marius-F. Danca and Nikolay Kuznetsov. Parameter switching synchronization. Applied Mathematics and Computation, 313(??):94–102, November 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317303879.

Diblik:2015:SES

[DKBS15] J. Diblík, D. Ya. Khusainov, J. Bastinec, and A. S. Sirenko. Stability and exponential stability of linear discrete systems with constant coefficients and single delay. Applied Mathematics and Computation, 269(??):9–16, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009492.

Dudin:2015:PRQ

[DKDD15] Alexander Dudin, Chesoong Kim, Sergey Dudin, and Olga Dudina. Priority retrial queueing model operating in random environment with varying number and reservation of servers. Applied Mathematics and Computation, 269(??):674–690, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010607.

Dolgy:2015:BTD

[DKKM15a] Dmitry V. Dolgy, Dae San Kim, Taekyun Kim, and Toufik Mansour. Barnestype degenerate Euler polynomials. Applied Mathematics and Computation, 261(??):388–396, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004592.

Dolgy:2015:DPC

[DKKM15b] Dmitry V. Dolgy, Dae San Kim, Taekyun Kim, and Toufik Mansour. Degenerate poly-Cauchy polynomials. Applied Mathematics and Computation, 269(??):637–646, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010395.

Diehl:2016:NSN

[DKKR16] Dennis Diehl, Jenny Kremser, Dietmar Kröner, and Christian Rohde. Numerical solution of Navier– Stokes–Korteweg systems by Local Discontinuous Galerkin methods in multiple space dimensions. Applied Mathematics and Computation, 272 (part 2) (??):309–335, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315013181.

Domschke:2015:ABE

[DKL15] Pia Domschke, Oliver Kolb, and Jens Lang. Adjoint-based error control for the simulation and optimization of gas and water supply networks. Applied Mathematics and Computation, 259(??):1003–1018, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003343.

Dolinar:2015:NPO

[DKM15] Gregor Dolinar, Bojan Kuzma, and Janko Marovt. A note on partial orders of Hartwig, Mitsch, and Semrl. Applied Mathematics and Computation, 270(??):711–713, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315011224.

Dragovich:2017:UGC

[DKM17] Branko Dragovich, Andrei Yu. Khrennikov, and Natasa Z. Misić. Ultrametrics in the genetic code and the genome. Applied Mathematics and Computation, 309(??):350–358, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302564.

Davari:2016:HAS

[DKN16] Soheil Davari, Kemal Kilic, and Siamak Naderi. A heuristic approach to solve the preventive health care problem with budget and congestion constraints. Applied Mathematics and Computation, 276(??):442–453, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015751.

DeMaesschalck:2016:SDH

[DKP16] P. De Maesschalck, E. Kutafina, and N. Popović. Sector-delayed-Hopf-type mixed-mode oscillations in a prototypical threetime-scale model. Applied Mathematics and Computation, 273(??):337–352, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013211.

Dimitrova:2016:EFT

[DKZ16] Dimitrina S. Dimitrova, Vladimir K. Kaishev, and Shouqi Zhao. On the evaluation of finite-time ruin probabilities in a dependent risk model. Applied Mathematics and Computation, 275(??):268–286, February 15, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315015842.

Ding:2015:TAN

[DL15a] Sheng-Hu Ding and Xing Li. Thermoelastic analysis of nonhomogeneous structural materials with an interface crack under uniform heat flow. Applied Mathematics and Computation, 271(??):22–33, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315011595.

Duan:2015:MDR

[DL15b] Xianbao Duan and Feifei Li. Material distribution resembled level set method for optimal shape design of Stokes flow. Applied Mathematics and Computation, 266(??):21–30, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005913.

Ding:2016:LEU

[DL16] Shijin Ding and Quanrong Li. Local existence of unique strong solution to non-isothermal model for incompressible nematic liquid crystals in 3D. Applied Mathematics and Computation, 290(??):487–506, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630409X.

Das:2017:KGD

[DL17a] Kinkar Ch. Das and Muhuo Liu. Kite graphs determined by their spectra. Applied Mathematics and Computation, 297(??):74–78, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306397.

DeBonis:2017:NMI

[DL17b] M. C. De Bonis and C. Laurita. A Nyström method for integral equations with fixed singularities of Mellin type in weighted L ^{p} spaces. Applied Mathematics and Computation, 303(??):55– 69, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300358.

Ding:2017:SSP

[DL17c] Xiuyong Ding and Xiu Liu. On stabilizability of switched positive linear systems under state-dependent switching. Applied Mathematics and Computation, 307(??):92–101, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301789.

Dou:2017:OIH

[DL17d] Jiawei Dou and Shundong Li. Optimal impulsive harvesting policies for single-species populations. Applied Mathematics and Computation, 292(??):145–155, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304659.

Dai:2018:FTV

[DL18a] Shaojun Dai and Shangzhao Li. Flag-transitive 4 − (v, k, 3) designs and PSL(2, q) groups. Applied Mathematics and Computation, 332(??):167–171, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301772.

Deng:2018:TFO

[DL18b] Dingwen Deng and Dong Liang. The time fourthorder compact ADI methods for solving twodimensional nonlinear wave equations. Applied Mathematics and Computation, 329(??):188–209, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301048.

Dai:2019:PTS

[DLBQ19] Ping-Fan Dai, Jicheng Li, Jianchao Bai, and Jinming Qiu. A preconditioned two-step modulusbased matrix splitting iteration method for linear complementarity problem. Applied Mathematics and Computation, 348(??):542–551, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310531.

Dou:2018:AMF

[DLCT18] Fangfang Dou, Zi-Cai Li, C. S. Chen, and Zhaolu Tian. Analysis on the method of fundamental solutions for biharmonic equations. Applied Mathematics and Computation, 339(??):323–345, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305745.

DOrazio:2017:MAM

[DLCV17] Christian J. D’Orazio, Rongxing Lu, Kim-Kwang Raymond Choo, and Athanasios V. Vasilakos. A Markov adversary model to detect vulnerable iOS devices and vulnerabilities in iOS apps. Applied Mathematics and Computation, 293(??):523–544, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305549.

Deng:2016:MGP

[DLD16] Xinyang Deng, Qi Liu, and Yong Deng. Matrix games with payoffs of belief structures. Applied Mathematics and Computation, 273(??):868–879, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014071.

Dai:2015:NGP

[DLF15] Lifang Dai, Maolin Liang, and Hongtao Fan. A new generalized parameterized inexact Uzawa method for solving saddle point problems. Applied Mathematics and Computation, 265(??):414–430, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006384.

Dong:2015:NVT

[DLH15] Xiao Liang Dong, Hong Wei Liu, and Yu Bo He. New version of the three-term conjugate gradient method based on spectral scaling conjugacy condition that generates descent search direction. Applied Mathematics and Computation, 269(??):606–617, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009868.

DeRosa:2016:DQB

[DLJT16] Maria A. De Rosa, Maria Lippiello, Ram Jiwari, and Stefania Tomasiello. A differential quadrature based procedure for parameter identification. Applied Mathematics and Computation, 290(??):460–466, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303903.

Ding:2015:DCI

[DLJW15] Zhanwen Ding, Qiang Li, Shumin Jiang, and Xuedi Wang. Dynamics in a Cournot investment game with heterogeneous players. Applied Mathematics and Computation, 256(??):939–950, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000922.

Deng:2018:BRA

[DLjZ18] Xiao Deng, Senyue Lou, and Da jun Zhang. Bilinearisationreduction approach to the nonlocal discrete nonlinear Schrödinger equations. Applied Mathematics and Computation, 332(??):477–483, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302340.

Diao:2017:SCN

[DLQ17] Huai-An Diao, Dongmei Liu, and Sanzheng Qiao. Structured condition numbers and small sample condition estimation of symmetric algebraic Riccati equations. Applied Mathematics and Computation, 314(??):80–97, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317304411.

delaRosa:2016:SCL

[dlRGB16] R. de la Rosa, M. L.

Gandarias, and M. S. Bruzón. On symmetries and conservation laws of a Gardner equation involving arbitrary functions. Applied Mathematics and Computation, 290(??):125–134, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630368X.

Debowski:2019:MDD

[DLS19] Mateusz Debowski, Mirosław Lachowicz, and Zuzanna Szymańska. Microscopic description of DNA thermal denaturation. Applied Mathematics and Computation, 361(??):47–60, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319303881.

Ding:2019:UUB

[DLXZ19] Yucai Ding, Hui Liu, Hui Xu, and Shouming Zhong. On uniform ultimate boundedness of linear systems with time-varying delays and peak-bounded disturbances. Applied Mathematics and Computation, 349(??):381–392, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311299.

Ding:2017:SSS

[DLY^{+}17a] Xueying Ding, Haitao Li, Qiqi Yang, Yingrui Zhou, Ahmed Alsaedi, and Fuad E. Alsaadi. Stochastic stability and stabilization of n-person random evolutionary Boolean games. Applied Mathematics and Computation, 306(??):1–12, August 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301248.

Du:2017:GSI

[DLY17b] Tingsong Du, Yujiao Li, and Zhiqiao Yang. A generalization of Simpson’s inequality via differentiable mapping using extended (s, m)-convex functions. Applied Mathematics and Computation, 293(??):358–369, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305409.

Dong:2018:SDG

[DLZ18] Changchang Dong, Juan Liu, and Xindong Zhang. Supereulerian digraphs with given diameter. Applied Mathematics and Computation, 329(??):5– 13, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300754.

Das:2016:ELEa

[DM16] Kinkar Ch. Das and Seyed Ahmad Mojallal. Extremal Laplacian energy of threshold graphs. Applied Mathematics and Computation, 273(??):267–280, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501334X.

Diblik:2018:DSP

[DM18] Josef Diblík and Rigoberto Medina. Dominant and subdominant positive solutions to generalized Dickman equation. Applied Mathematics and Computation, 333(??):169–186, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302716.

Deolmi:2019:ROM

[DMA^{+}19] G. Deolmi, S. Müller, M. Albers, P. S. Meysonnat, and W. Schröder. A reduced order model to simulate compressible flows over an actuated riblet surface. Applied Mathematics and Computation, 348(??):234–256, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310336.

Diaz:2015:NPS

[DME^{+}15] F. Díaz, G. Montero, J. M. Escobar, E. Rodríguez, and R. Montenegro. A new predictive solar radiation numerical model. Applied Mathematics and Computation, 267(??):596–603, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315000508.

Das:2015:LET

[DMG15] Kinkar Ch. Das, Seyed Ahmad Mojallal, and Ivan Gutman. On Laplacian energy in terms of graph invariants. Applied Mathematics and Computation, 268(??):83–92, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008450.

Das:2016:ELEb

[DMG16] Kinkar Ch. Das, Seyed Ahmad Mojallal, and Ivan Gutman. On energy and Laplacian energy of bipartite graphs. Applied Mathematics and Computation, 273(??):759–766, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013983.

Dung:2019:ATS

[DMH19] Dinh Dũng, Charles A.

Micchelli, and Vu Nhat Huy. Approximation by translates of a single function of functions in space induced by the convolution with a given function. Applied Mathematics and Computation, 361(??):777–787, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304990.

Dafnis:2019:RGC

[DMP19] Spiros D. Dafnis, Frosso S. Makri, and Andreas N. Philippou. The reliability of a generalized consecutive system. Applied Mathematics and Computation, 359(??):186–193, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319303443.

Donatelli:2015:AMI

[DMR15] Marco Donatelli, David Martin, and Lothar Reichel. Arnoldi methods for image deblurring with anti-reflective boundary conditions. Applied Mathematics and Computation, 253(??):135–150, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017123.

Diaz:2018:ISD

[DMS^{+}18] F. Díaz, H. Montero, D. Santana, G. Montero, E. Rodriguez, L. Mazorra Aguiar, and A. Oliver. Improving shadows detection for solar radiation numerical models. Applied Mathematics and Computation, 319(??):71–85, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300644.

Dykas:2018:NAI

[DMSS18] Slawomir Dykas, Miroslaw Majkut, Krystian Smolka, and Michal Strozik. Numerical analysis of the impact of pollutants on water vapour condensation in atmospheric air transonic flows. Applied Mathematics and Computation, 338(??):432–450, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305368.

Das:2015:UCH

[DN15a] Abhishek Das and Srinivasan Natesan. Uniformly convergent hybrid numerical scheme for singularly perturbed delay parabolic convectiondiffusion problems on Shishkin mesh. Applied Mathematics and Computation, 271(??):168–186, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315012187.

Das:2015:CAD

[DN15b] Payel Das and Gnaneshwar Nelakanti. Convergence analysis of discrete Legendre spectral projection methods for Hammerstein integral equations of mixed type. Applied Mathematics and Computation, 265(??):574–601, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007316. See

corrigendum [DN16].

Dunca:2015:VFB

[DN15c] Argus A. Dunca and Monika Neda. On the vreman filter based stabilization for the advection equation. Applied Mathematics and Computation, 269(??):379–388, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501005X.

Das:2016:CCA

[DN16] Payel Das and Gnaneshwar Nelakanti. Corrigendum to: “Convergence analysis of discrete Legendre spectral projection methods for Hammerstein integral equations of mixed type”, Applied Mathematics and Computation Volume 265, 15 August 2015, Pages 574–601. Applied Mathematics and Computation, 281(??):394–395, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031530028X. See

[DN15b].

Dong:2017:SOT

[DNC^{+}17] Hao Dong, Yufeng Nie, Junzhi Cui, Yatao Wu, and Zihao Yang. Secondorder two-scale analysis and numerical algorithm for the damped wave equations of composite materials with quasi-periodic structures. Applied Mathematics and Computation, 298(??):201–220, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306919.

doNascimento:2015:UMH

[dNdOX^{+}15] T. M. do Nascimento, J. M. de Oliveira, M. P. Xavier, A. B. Pigozzo, R. W. dos Santos, and M. Lobosco. On the use of multiple heterogeneous devices to speedup the execution of a computational model of the human immune system. Applied Mathematics and Computation, 267(??):304–313, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003860.

Dorociakova:2015:EPP

[DO15] Bozena Dorociaková and Rudolf Olach. Existence of positive periodic solutions to nonlinear integro-differential equations. Applied Mathematics and Computation, 253(??):287–293, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314013113.

DeBonis:2017:NCH

[DO17] Maria Carmela De Bonis and Donatella Occorsio. Numerical computation of hypersingular integrals on the real semiaxis. Applied Mathematics and Computation, 313(??):367–383, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304216.

Dumett:2018:MDE

[DO18] Miguel A. Dumett and Jorge E. Ospino. Mimetic discretization of the eikonal equation with Soner boundary conditions. Applied Mathematics and Computation, 335(??):25–37, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303394.

Dernek:2015:NIP

[DÖA15] Nese Dernek, Eyüp Ömer Ölçüçü, and Fatih Aylikçi. New identities and Parseval type relations for the generalized integral transforms L_{4n}, P_{4n}, F_{s,2n}, and F_{c,2n}. Applied Mathematics and Computation, 269(??):536–547, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315010176.

Dobrynin:2019:IFC

[Dob19] Andrey A. Dobrynin. Infinite family of 2connected transmission irregular graphs. Applied Mathematics and Computation, 340(??):1–4, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318307562.

Dokuchaev:2018:CES

[Dok18] Nikolai Dokuchaev. On causal extrapolation of sequences with applications to forecasting. Applied Mathematics and Computation, 328(??):276–286, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300596.

Dolejsi:2015:AAD

[Dol15] Vít Dolejsí. Anisotropic hp-adaptive discontinuous Galerkin method for the numerical solution of time dependent PDEs. Applied Mathematics and Computation, 267(??):682–697, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017603.

Dordevic:2016:NMP

[Dor16] Snezana S. Dordević. A note on a multiplicative parameters gradient method. Applied Mathematics and Computation, 283(??):90–107, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301424.

Dehestani:2018:FOL

[DOR18] H. Dehestani, Y. Ordokhani, and M. Razzaghi. Fractional-order Legendre–Laguerre functions and their applications in fractional partial differential equations. Applied Mathematics and Computation, 336(??):433–453, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830420X.

Dorjgotov:2018:LSA

[DOZ18] Khongorzul Dorjgotov, Hiroyuki Ochiai, and Uuganbayar Zunderiya. Lie symmetry analysis of a class of time fractional nonlinear evolution systems. Applied Mathematics and Computation, 329(??):105–117, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300791.

Das:2015:ITP

[DP15a] Kedar Nath Das and Raghav Prasad Parouha. An ideal tri-population approach for unconstrained optimization and applications. Applied Mathematics and Computation, 256(??):666–701, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001083.

Delgado:2015:AEB

[DP15b] Jorge Delgado and J. M. Peña. Accurate evaluation of Bézier curves and surfaces and the Bernstein–Fourier algorithm. Applied Mathematics and Computation, 271(??):113–122, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S009630031501142X.

Delgado:2017:ACL

[DP17] Jorge Delgado and J. M.

Peña. Accurate computations with Lupas matrices. Applied Mathematics and Computation, 303(??):171–177, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300395.

Dhamija:2018:AKF

[DPD18] Minakshi Dhamija, Ram Pratap, and Naokant Deo. Approximation by Kantorovich form of modified Szász–Mirakyan operators. Applied Mathematics and Computation, 317(??):109–120, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306240.

Lorenzo:2018:HTP

[DPL18a] M. De Lorenzo, M. Pelanti, and Ph. Lafon. HLLC-type and path-conservative schemes for a single-velocity sixequation two-phase flow model: A comparative study. Applied Mathematics and Computation, 333(??):95–117, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830273X.

DeMiguel:2018:OML

[DPL^{+}18b] L. De Miguel, D. Paternain, I. Lizasoain, G. Ochoa, and H. Bustince. Orness measurements for lattice m-dimensional intervalvalued OWA operators. Applied Mathematics and Computation, 339(??):59– 62, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031830571X.

Dolgov:2016:FTP

[DPSS16] Sergey Dolgov, John W. Pearson, Dmitry V. Savostyanov, and Martin Stoll. Fast tensor product solvers for optimization problems with fractional differential equations as constraints. Applied Mathematics and Computation, 273(??):604–623, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012783.

Duan:2018:SSP

[DQ18] Jun-Sheng Duan and Xiang Qiu. Stokes’ second problem of viscoelastic fluids with constitutive equation of distributedorder derivative. Applied Mathematics and Computation, 331(??):130–139, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301309.

Daraghmeh:2018:OCL

[DQH18] Adnan Daraghmeh, Naji Qatanani, and Carsten Hartmann. Optimal control of linear systems with balanced reduced-order models: Perturbation approximations. Applied Mathematics and Computation, 337(??):102–118, November 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318303916.

Du:2019:NMN

[DQW^{+}19] Mingjing Du, Xiaohua Qiao, Biao Wang, Yulan Wang, and Bo Gao. A novel method for numerical simulation of sand motion model in beach formation based on fractional Taylor–Jumarie series expansion and piecewise interpolation technique. Applied Mathematics and Computation, 347(??):15–21, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309597.

DelCorso:2019:ANM

[DR19] Gianna M. Del Corso and Francesco Romani. Adaptive nonnegative matrix factorization and measure comparisons for recommender systems. Applied Mathematics and Computation, 354(??):164–179, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300621.

DeRidder:2016:RDC

[DRS16] H. De Ridder, T. Raeymaekers, and F. Sommen. Rotations in discrete Clifford analysis. Applied Mathematics and Computation, 285(??):114–140, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300316302272.

Dudzinski:2018:OOO

[DRS18] Michael Dudzinski, Marco Rozgić, and Marcus Stiemer. oFEM: An object oriented finite element package for Matlab. Applied Mathematics and Computation, 334(??):117–140, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730824X.

Dandapat:2015:UTL

[DS15] B. S. Dandapat and S. K.

Singh. Unsteady two-layer film flow on a non-uniform rotating disk in presence of uniform transverse magnetic field. Applied Mathematics and Computation, 258(??):545–555, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002386.

Dassios:2016:BOC

[DS16a] Ioannis K. Dassios and Krzysztof J. Szajowski. Bayesian optimal control for a non-autonomous stochastic discrete time system. Applied Mathematics and Computation, 274(??):556–564, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014642.

Dolejsi:2016:HDG

[DS16b] Vít Dolejsí and Pavel Solin. hp-discontinuous Galerkin method based on local higher order reconstruction. Applied Mathematics and Computation, 279(??):219–235, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300248.

Dehghan:2019:GMH

[DS19] Mehdi Dehghan and Akbar Shirilord. A generalized modified Hermitian and skew-Hermitian splitting (GMHSS) method for solving complex Sylvester matrix equation. Applied Mathematics and Computation, 348(??):632–651, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310385.

Deng:2016:ZIF

[DSAB16] Hanyuan Deng, D. Sarala, S. K. Ayyaswamy, and S. Balachandran. The Zagreb indices of four operations on graphs. Applied Mathematics and Computation, 275(??):422–431, February 15, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031501560X.

Dhar:2015:DTD

[DSB15] Joydip Dhar, Harkaran Singh, and Harbax Singh Bhatti. Discrete-time dynamics of a system with crowding effect and predator partially dependent on prey. Applied Mathematics and Computation, 252(??):324–335, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016750.

Dattoli:2019:CPB

[DSDL19] G. Dattoli, E. Sabia, E. Di Palma, and S. Licciardi. Corrigendum to “Products of Bessel functions and associated polynomials” [Applied Mathematics and Computation, 266(C) (2015) 507– 514]. Applied Mathematics and Computation, 362 (??):Article 124516, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304953.

See [DDSL15].

dosSantos:2015:EAP

[dSH15] José Paulo Carvalho dos Santos and Hernán R. Henríquez. Existence of S-asymptotically ωperiodic solutions to abstract integro-differential equations. Applied Mathematics and Computation, 256(??):109–118, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000132.

ditSandretto:2017:SCS

[dSH17] Julien Alexandre dit Sandretto and Milan Hladík. Solving over-constrained systems of non-linear interval equations — and its robotic application. Applied Mathematics and Computation, 313(??): 180–195, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.

com/science/article/
pii/S0096300317303892.

Dehmer:2015:DPG

[DSM15] Matthias Dehmer, Yongtang Shi, and Abbe Mowshowitz. Discrimination power of graph measures based on complex zeros of the partial Hosoya polynomial. Applied Mathematics and Computation, 250(??):352–355, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014143.

daSilva:2015:SPS

[dSR15] Rodrigo A. da Silva and Paulo C. Rech. Spiral periodic structures in a parameter plane of an ecological model. Applied Mathematics and Computation, 254(??):9–13, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017767.

diSerafino:2018:SSG

[dSRTZ18] Daniela di Serafino, Valeria Ruggiero, Gerardo Toraldo, and Luca Zanni. On the steplength selection in gradient methods for unconstrained optimization. Applied Mathematics and Computation, 318(??):176–195, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304988.

dosSantos:2016:UDV

[dSSB^{+}16] Vagner dos Santos, José D. Szezech, Jr., Murilo S. Baptista, Antonio M. Batista, and Iberê L. Caldas. Unstable dimension variability structure in the parameter space of coupled Hénon maps. Applied Mathematics and Computation, 286(??):23–28, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302491.

Du:2019:LBM

[DSSC19] Rui Du, Dongke Sun, Baochang Shi, and Zhenhua Chai. Lattice Boltzmann model for time subdiffusion equation in Caputo sense. Applied Mathematics and Computation, 358(??):80–90, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300319302966.

daSilva:2017:IES

[dSVR17] D. J. Pamplona da Silva, R. P. Villar, and L. C. Ramos. Isolation effects in a system of two mutually communicating identical patches. Applied Mathematics and Computation, 315(??):494–499, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305544.

Deng:2019:DTL

[DSZY19] Xingchao Deng, Zhiwei Shao, Huan Zhang, and Weihua Yang. The (d, 1)-total labelling of Sierpiński-like graphs. Applied Mathematics and Computation, 361(??): 484–492, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL https://www.sciencedirect.
com/science/article/
pii/S0096300319304527.

Demiray:2016:QPS

[DT16] Seçil Demiray and Filiz Tascan. Quasi-periodic solutions of (3 + 1) generalized BKP equation by using Riemann theta functions. Applied Mathematics and Computation, 273(??):131–141, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013363.

Dantas:2018:CSS

[DTJ18] Eber Dantas, Michel Tosin, and Americo Cunha Jr. Calibration of a SEIR– SEI epidemic model to describe the Zika virus outbreak in Brazil. Applied Mathematics and Computation, 338(??):231–248, December 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318305125.

Dassios:2019:MTE

[DTM19] Ioannis Dassios, Georgios Tzounas, and Federico Milano. The Möbius transform effect in singular systems of differential equations. Applied Mathematics and Computation, 361(??):338–353, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304497.

Du:2015:SAN

[Du15] Bo Du. Successive approximation of neutral functional stochastic differential equations with variable delays. Applied Mathematics and Computation, 268(??):609–615, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008875.

Du:2017:FDD

[Du17a] Dongsheng Du. Fault detection for discrete-time linear systems based on descriptor observer approach. Applied Mathematics and Computation, 293(??):575–585, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305550.

Du:2017:FRR

[Du17b] Zhibin Du. Further results regarding the sum of domination number and average eccentricity. Applied Mathematics and Computation, 294(??):299–309, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305744.

Didov:2018:ASP

[DU18] A. A. Didov and M. Yu. Uleysky. Analysis of stationary points and their bifurcations in the ABCflow. Applied Mathematics and Computation, 330(??):56–64, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301346.

Du:2019:RPC

[Du19] Jinming Du. Redistribution promotes cooperation in spatial public goods games under aspiration dynamics. Applied Mathematics and Computation, 363(??): Article 124629, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319306216.

Dehmer:2015:CAT

[DV15] Matthias Dehmer and Kurt Varmuza. A comparative analysis of the Tanimoto index and graph edit distance for measuring the topological similarity of trees. Applied Mathematics and Computation, 259(??):242–250, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002222.

Del-Vecchio:2018:LIP

[DVJ18] Renata R. Del-Vecchio and Átila Arueira Jones. Laplacian integrality in P_{4}-sparse and P_{4}-extendible graphs. Applied Mathematics and Computation, 330(??):307–315, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301565.

DeLegge:2017:EFM

[DW17] Anthony DeLegge and Hannah Wangler. Is this the end for Facebook? a mathematical analysis. Applied Mathematics and Computation, 305(??):364–380, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301182.

Dai:2017:CVM

[DWRZ17] Baodong Dai, Dandan Wei, Hongping Ren, and Zhu Zhang. The complex variable meshless local Petrov–Galerkin method for elastodynamic analysis of functionally graded materials. Applied Mathematics and Computation, 309(??):17–26, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302266.

Douissi:2019:MFA

[DWS19] Soukaina Douissi, Jiaqiang Wen, and Yufeng Shi. Mean-field anticipated BS-DEs driven by fractional Brownian motion and related stochastic control problem. Applied Mathematics and Computation, 355(??):282–298, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301821.

Du:2017:RKM

[DWT17] Ming-Jing Du, Yu-Lan Wang, and Chao-Lu Temuer. Reproducing kernel method for numerical simulation of downhole temperature distribution. Applied Mathematics and Computation, 297(??):19–30, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306439.

Du:2017:MRK

[DWTT17] Ming-Jing Du, Yu-Lan Wang, Chao-Lu Temuer, and Dan Tian. A modified reproducing kernel method for solving Burgers’ equation arising from diffusive waves in fluid dynamics. Applied Mathematics and Computation, 315(??):500–506, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730526X.

Deng:2015:OCS

[DWZ15] Xun-Huan Deng, Qi-Ru Wang, and Zhan Zhou. Oscillation criteria for second order nonlinear delay dynamic equations on time scales. Applied Mathematics and Computation, 269(??):834–840, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315010656.

Du:2017:BCS

[DX17] Runmei Du and Fengdan Xu. On the boundary controllability of a semilinear degenerate system with the convection term. Applied Mathematics and Computation, 303(??):113–127, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300164.

Deng:2015:RBD

[DY15] Chunyuan Deng and Anqi Yu. Relationships between DMP relation and some partial orders. Applied Mathematics and Computation, 266(??):41–53, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315006402.

Daoxiang:2017:MLC

[DY17] Zhang Daoxiang and Ping Yan. Multiple limit cycles for the continuous model of the rock-scissors-paper game between bacteriocin producing bacteria. Applied Mathematics and Computation, 295(??):136–140, February 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304829.

Deng:2015:CRP

[DYW15] Chunyuan Deng, Anqi Yu, and Yimin Wei. Characterizations and representations of the (P, Q)- outer generalized inverse. Applied Mathematics and Computation, 269(??): 432–442, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315009765.

Ding:2015:ENS

[DZ15] Jianwan Ding and Chengjian Zhang. An extension of numerical stability criteria for linear neutral multidelay-integro-differential equations. Applied Mathematics and Computation, 265(??): 347–351, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315006311.

Diagana:2016:EPA

[DZ16] Toka Diagana and Hui Zhou. Existence of positive almost periodic solutions to the hematopoiesis model. Applied Mathematics and Computation, 274(??):644–648, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013715.

Duan:2017:SPA

[DZ17a] Wei-Long Duan and Chunhua Zeng. Signal power amplification of intracellular calcium dynamics with non-Gaussian noises and time delay. Applied Mathematics and Computation, 292(??):400–405, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304854.

Duan:2017:SAS

[DZ17b] Wei-Long Duan and Chunhua Zeng. Statistics for anti-synchronization of intracellular calcium dynamics. Applied Mathematics and Computation, 293(??):611–616, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300316304799.

Du:2015:UIO

[DZC15] Xiao-Kun Du, Hui Zhao, and Xiao-Heng Chang. Unknown input observer design for fuzzy systems with uncertainties. Applied Mathematics and Computation, 266(??):108–118, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006633.

Deng:2016:SAW

[DZD^{+}16] Xinyang Deng, Zhipeng Zhang, Yong Deng, Qi Liu, and Shuhua Chang. Selfadaptive win-stay-lose-shift reference selection mechanism promotes cooperation on a square lattice. Applied Mathematics and Computation, 284(??):322–331, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302028.

Dziok:2015:GMM

[Dzi15] Jacek Dziok. Generalizations of multivalent Mocanu functions. Applied Mathematics and Computation, 269(??):965–971, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010905.

Dong:2015:ISS

[DZL15] Yaying Dong, Shunli Zhang, and Shanbing Li. Invariant sets and solutions to a class of wave equations. Applied Mathematics and Computation, 253(??):369–376, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017548.

Deng:2015:EAS

[DZW15] Jingwei Deng, Lijing Zhao, and Yujiang Wu. Efficient algorithms for solving the fractional ordinary differential equations. Applied Mathematics and Computation, 269(??):196–216, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009674.

Du:2018:NEA

[DZY^{+}18] Wenbo Du, Mingyuan Zhang, Wen Ying, Matjaž Perc, Ke Tang, Xianbin Cao, and Dapeng Wu. The networked evolutionary algorithm: a network science perspective. Applied Mathematics and Computation, 338(??):25–32, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304909.

Elaiw:2015:GPD

[EA15a] A. M. Elaiw and N. A.

Almuallem. Global properties of delayed-HIV dynamics models with differential drug efficacy in cocirculating target cells. Applied Mathematics and Computation, 265(??):1067–1089, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007870.

Esmaeili:2015:ETS

[EA15b] H. Esmaeili and M. Ahmadi. An efficient threestep method to solve system of nonlinear equations. Applied Mathematics and Computation, 266 (??):1093–1101, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007006.

El-Ajou:2015:GFG

[EAAAS15] Ahmad El-Ajou, Omar Abu Arqub, and Mohammed Al-Smadi. A general form of the generalized Taylor’s formula with some applications. Applied Mathematics and Computation, 256(??):851–859, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500048X.

El-Ajou:2015:NEI

[EAAM^{+}15] Ahmad El-Ajou, Omar Abu Arqub, Shaher Momani, Dumitru Baleanu, and Ahmed Alsaedi. A novel expansion iterative method for solving linear partial differential equations of fractional order. Applied Mathematics and Computation, 257(??):119–133, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017822.

El-Aziz:2017:PAU

[EAY17] Mohamed Abd El-Aziz and Aishah S. Yahya. Perturbation analysis of unsteady boundary layer slip flow and heat transfer of Casson fluid past a vertical permeable plate with Hall current. Applied Mathematics and Computation, 307(??):146–164, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301455.

Exner:2016:PUM

[EB16] Pavel Exner and Jan Brezina. Partition of unity methods for approximation of point water sources in porous media. Applied Mathematics and Computation, 273(??):21–32, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012862.

Escalante:2018:NHP

[EdLC18] C. Escalante, T. Morales de Luna, and M. J. Castro. Non-hydrostatic pressure shallow flows: GPU implementation using finite volume and finite difference scheme. Applied Mathematics and Computation, 338(??):607–623, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305241.

Ezz-Eldien:2018:SSM

[EE18] S. S. Ezz-Eldien. On solving systems of multipantograph equations via spectral tau method. Applied Mathematics and Computation, 321(??):63– 73, March 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317307038.

Eisenschmidt:2016:DNS

[EEG^{+}16]  Kathrin  Eisenschmidt,

Moritz Ertl, Hassan Gomaa, Corine Kieffer-Roth, Christian Meister, Philipp Rauschenberger, Martin Reitzle, Karin Schlottke, and Bernhard Weigand. Direct numerical simulations for multiphase flows: an overview of the multiphase code FS3D. Applied Mathematics and Computation, 272 (part 2)(??): 508–517, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315007195.

Elshobaky:2016:PMR

[EEST16] Entisarat M. Elshobaky, Essam El-Seidy, and Mohammad A. Taha. The payoff matrix of repeated asymmetric 2 × 2 games. Applied Mathematics and Computation, 274(??):393–402, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014654.

Fang:2017:FNU

[eFJY17] Mei e Fang, Byeongseon Jeong, and Jungho Yoon. A family of non-uniform subdivision schemes with variable parameters for curve design. Applied Mathematics and Computation, 313(??):1–11, November 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317303752.

Emamizadeh:2018:OHS

[EFL18]  Behrouz  Emamizadeh,

Amin Farjudian, and Yichen Liu. Optimal harvesting strategy based on rearrangements of functions. Applied Mathematics and Computation, 320(??):677–690, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306951.

Egger:2018:ANS

[EFPT18] H. Egger, K. Fellner, J.-F. Pietschmann, and B. Q. Tang. Analysis and numerical solution of coupled volume-surface reaction-diffusion systems with application to cell biology. Applied Mathematics and Computation, 336(??):351–367, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303497.

Eftekhari:2015:PIE

[Eft15] Tahereh Eftekhari. Producing an interval extension of the King method. Applied Mathematics and Computation, 260(??): 288–291, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315003938.

Ewald:2017:OCC

[EG17] Christian-Oliver Ewald and Johannes Geißler. Optimal contracts for central bankers: Calls on inflation. Applied Mathematics and Computation, 292(??):57–62, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304416.

Erfanian:2015:RHW

[EGB15] M. Erfanian, M. Gachpazan, and H. Beiglo. Rationalized Haar wavelet bases to approximate solution of nonlinear Fredholm integral equations with error analysis. Applied Mathematics and Computation, 265(??):304–312, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006268. See

correction [Bag19].

Eltanboly:2019:LSB

[EGH^{+}19] Ahmed Eltanboly, Mohammed Ghazal, Hassan Hajjdiab, Ahmed Shalaby, Andy Switala, Ali Mahmoud, Prasanna Sahoo, Magdi El-Azab, and Ayman El-Baz. Level setsbased image segmentation approach using statistical shape priors. Applied Mathematics and Computation, 340(??):164–179, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318307240.

Elhosseini:2014:MHA

[EH14] Mostafa A. Elhosseini and Amira Y. Haikal. Modified hybrid algorithm for process optimization. Applied Mathematics and Computation, 232(??):1209–1224, April 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314001623. See

retraction notice [EH15].

Elhosseini:2015:RNM

[EH15] Mostafa A. Elhosseini and Amira Y. Haikal. Retraction notice to “Modified Hybrid Algorithm for Process Optimization” [AMC 232 (2014) 1209– 1224]. Applied Mathematics and Computation, 266(??):1177, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009467. See

[EH14].

Estevez:2016:LSL

[EHdLS16] P. G. Estévez, F. J. Herranz, J. de Lucas, and C. Sardón. Lie symmetries for Lie systems: Applications to systems of ODEs and PDEs. Applied Mathematics and Computation, 273(??):435–452, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013168.

Edalatpour:2017:GGS

[EHS17] Vahid Edalatpour, Davod Hezari, and Davod Khojasteh Salkuyeh. A generalization of the Gauss– Seidel iteration method for solving absolute value equations. Applied Mathematics and Computation, 293(??):156–167, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630515X.

Ezquerro:2015:ASC

[EHVV15] J. A. Ezquerro, M. A.

Hernández-Verón, and A. I. Velasco. An analysis of the semilocal convergence for secant-like methods. Applied Mathematics and Computation, 266(??):883–892, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007912.

El-Jaick:2015:IRS

[EJF15] Léa Jaccoud El-Jaick and Bartolomeu D. B. Figueiredo. Integral relations for solutions of the confluent Heun equation. Applied Mathematics and Computation, 256(??):885–904, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001265.

El-Jaick:2016:CAS

[EJF16] Léa Jaccoud El-Jaick and Bartolomeu D. B. Figueiredo. Convergence and applications of some solutions of the confluent Heun equation. Applied Mathematics and Computation, 284(??):234–259, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301953.

Eremin:2019:EOS

[EKO19] Alexey S. Eremin, Nikolai A. Kovrizhnykh, and Igor V. Olemskoy. An explicit one-step multischeme sixth order method for systems of special structure. Applied Mathematics and Computation, 347(??):853–864, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310270.

Egger:2015:TBC

[EKS^{+}15] Herbert Egger, Fritz Kretzschmar, Sascha M. Schnepp, Igor Tsukerman, and Thomas Weiland. Transparent boundary conditions for a discontinuous Galerkin Trefftz method. Applied Mathematics and Computation, 267(??):42–55, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008073.

Erkmen:2019:ENU

[EL19a] Dilek Erkmen and Alexander E. Labovsky. Erratum to: ‘Note On the Usage of Grad-Div Stabilization for the Penalty-Projection Algorithm in Magnetohydrodynamics, Applied Mathematics and Computation (2019) 97–101. Applied Mathematics and Computation, 349(??):R1, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300281. This

article was withdrawn because of accidental duplication of [EL19b].

Erkmen:2019:NUG

[EL19b] Dilek Erkmen and Alexander E. Labovsky. Note on the usage of grad–div stabilization for the penalty– projection algorithm in magnetohydrodynamics. Applied Mathematics and Computation, 349(??):48– 52, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310889.

Erkmen:2019:WNU

[EL19c] Dilek Erkmen and Alexander E. Labovsky. WITH-DRAWN: Note on the usage of grad–div stabilization for the penalty– projection algorithm in magnetohydrodynamics. Applied Mathematics and Computation, 349(??):97– 101, May 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318310804.

See erratum [EL19a]. This article was withdrawn because of accidental duplication of [EL19b].

Elezovic:2015:EPF

[Ele15]  Neven Elezović.  Es-

timations of psi function and harmonic numbers. Applied Mathematics and Computation, 258(??):192–205, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001617.

Elmahdy:2015:NAW

[Elm15] Emad E. Elmahdy. A new approach for Weibull modeling for reliability life data analysis. Applied Mathematics and Computation, 250(??):708–720, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014027.

Fatini:2018:SSI

[ELP^{+}18] Mohamed El Fatini, Aadil Lahrouz, Roger Pettersson, Adel Settati, and Regragui Taki. Stochastic stability and instability of an epidemic model with relapse. Applied Mathematics and Computation, 316(??):326–341, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730588X.

Elsadany:2017:DCD

[Els17] A. A. Elsadany. Dynamics of a Cournot duopoly game with bounded rationality based on relative profit maximization. Applied Mathematics and Computation, 294(??):253–263, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305781.

Elyseeva:2015:GOT

[Ely15]  Julia Elyseeva.  Gen-

eralized oscillation theorems for symplectic difference systems with nonlinear dependence on spectral parameter. Applied Mathematics and Computation, 251(??):92–107, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015744.

Elyseeva:2018:CIT

[Ely18] Julia Elyseeva. The comparative index and transformations of linear Hamiltonian differential systems. Applied Mathematics and Computation, 330(??):185–200, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301280.

Ene:2015:ASS

[EM15] Remus-Daniel Ene and Vasile Marinca. Approximate solutions for steady boundary layer MHD viscous flow and radiative heat transfer over an exponentially porous stretching sheet. Applied Mathematics and Computation, 269(??):389–401, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009509.

Egidi:2016:ABC

[EM16] Nadaniela Egidi and Pierluigi Maponi. Artificial boundary conditions for the Burgers equation on the plane. Applied Mathematics and Computation, 286(??):1–14, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302508.

Abd-el-Malek:2015:NES

[eMA15] Mina B. Abd el Malek and Amr M. Amin. New exact solutions for solving the initial-value-problem of the KdV– KP equation via the Lie group method. Applied Mathematics and Computation, 261(??):408–418, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004361.

Abd-el-Malek:2019:MPS

[eMAGG19] Mina B. Abd el Malek, Amr Abdelrazek, Mohammed Ghazy, and Gehad Gamal. A modified perturbation solution to the one-dimensional Bratu problem. Applied Mathematics and Computation, 354(??):296–304, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301304.

Abd-el-Malek:2015:NSS

[eMBHA15] Mina B. Abd el Malek, Nagwa A. Badran, Hossam S. Hassan, and Heba H. Abbas. New solutions for solving Boussinesq equation via potential symmetries method. Applied Mathematics and Computation, 251(??): 225–232, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314015872.

Esmaeilbeigi:2017:MMS

[EMM17] Mohsen Esmaeilbeigi, Farshid Mirzaee, and Davoud Moazami. A meshfree method for solving multidimensional linear Fredholm integral equations on the hypercube domains. Applied Mathematics and Computation, 298(??):236–246, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306889.

Einkemmer:2018:EBC

[EMO18] Lukas Einkemmer, Martina Moccaldi, and Alexander Ostermann. Efficient boundary corrected Strang splitting. Applied Mathematics and Computation, 332(??):76–89, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301711.

Esteves:2015:GAS

[EO15] Salete Esteves and José J. Oliveira. Global asymptotic stability of nonautonomous Cohen–Grossberg neural network models with infinite delays. Applied Mathematics and Computation, 265(??):333–346, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005652.

Erdogdu:2017:ESQ

[EÖ17] Melek Erdogdu and Mustafa Özdemir. On exponential of split quaternionic matrices. Applied Mathematics and Computation, 315(??):468–476, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317305581.

Einkemmer:2018:SSF

[EO18] Lukas Einkemmer and Alexander Ostermann. A split step Fourier/ discontinuous Galerkin scheme for the Kadomtsev– Petviashvili equation. Applied Mathematics and Computation, 334(??): 311–325, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318303254.

Erkaymaz:2017:PSW

[EOP17] Okan Erkaymaz, Mahmut Ozer, and Matjaz Perc. Performance of smallworld feedforward neural networks for the diagnosis of diabetes. Applied Mathematics and Computation, 311(??):22–28, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302989.

Ebrahimi:2015:CML

[ER15] Nehzat Ebrahimi and Jalil Rashidinia. Collocation method for linear and nonlinear Fredholm and Volterra integral equations. Applied Mathematics and Computation, 270(??):156–164, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010875.

Erb:2016:BLI

[Erb16] Wolfgang Erb. Bivariate Lagrange interpolation at the node points of Lissajous curves — the degenerate case. Applied Mathematics and Computation, 289(??):409–425, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303265.

Er-Riani:2016:ASG

[ERFE16] Mustapha Er-Riani, Youness Filali, and Mustapha El Jarroudi. Axisymmetric shape of gas bubbles in a uniform flow: Numerical study of bifurcation by analytic continuation. Applied Mathematics and Computation, 274(??):306–317, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014368.

Ershkov:2016:AES

[Ers16a] Sergey V. Ershkov. About existence of stationary points for the Arnold– Beltrami–Childress (ABC) flow. Applied Mathematics and Computation, 276(??):379–383, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300266.

Ershkov:2016:NSH

[Ers16b] Sergey V. Ershkov. Nonstationary helical flows for incompressible 3D Navier– Stokes equations. Applied Mathematics and Computation, 274(??):611–614, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015295.

Eryilmaz:2015:AMS

[Ery15] Serkan Eryilmaz. Assessment of a multi-state system under a shock model. Applied Mathematics and Computation, 269(??):1–8, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315009133.

El-Seidy:2015:ENA

[ES15] Essam El-Seidy. The effect of noise and average relatedness between players in iterated games. Applied Mathematics and Computation, 269(??):343–350, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009728.

Erden:2016:GPT

[ES16] Samet Erden and Mehmet Zeki Sarikaya. Generalized Pompeiu type inequalities for local fractional integrals and its applications. Applied Mathematics and Computation, 274(??):282–291, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014745.

Escorcia:2017:BRS

[ES17a] J. Escorcia and E. Suazo. Blow-up results and soliton solutions for a generalized variable coefficient nonlinear Schrödinger equation. Applied Mathematics and Computation, 301(??):155–176, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307445.

Estrada:2017:RNG

[ES17b] Ernesto Estrada and Matthew Sheerin. Random neighborhood graphs as models of fracture networks on rocks: Structural and dynamical analysis. Applied Mathematics and Computation, 314(??):360–379, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304319.

El-Seidy:2016:TPT

[ESES16] Essam El-Seidy, Entisarat M. Elshobaky, and Karim M. Soliman. Two population three-player prisoner’s dilemma game. Applied Mathematics and Computation, 277(??):44– 53, March 20, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315300357.

Eslami:2016:ETW

[Esl16]  Mostafa Eslami.  Ex-

act traveling wave solutions to the fractional coupled nonlinear Schrödinger equations. Applied Mathematics and Computation, 285(??):141–148, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302326.

El-Seidy:2016:IST

[ESS16] Essam El-Seidy and Karim. M. Soliman. Iterated symmetric three-player prisoner’s dilemma game. Applied Mathematics and Computation, 282(??):117–127, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301011.

Esen:2015:ATF

[ET15] A. Esen and O. Tasbozan. An approach to time fractional gas dynamics equation: Quadratic B-spline Galerkin method. Applied Mathematics and Computation, 261(??):330–336, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004452.

Ewald:2016:CGS

[Ewa16] Tobias Ewald. Convergence of geometric subdivision schemes. Applied Mathematics and Computation, 272 (part 1)(??): 41–52, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315009881.

Elsadany:2018:FAB

[EYE18] A. A. Elsadany, A. M.

Yousef, and Amr Elsonbaty. Further analytical bifurcation analysis and applications of coupled logistic maps. Applied Mathematics and Computation, 338(??):305–313, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830496X.

Fallahati:2015:NDM

[FA15] Farzane Fallahati and Masoud Allame. A note on the decomposition method for largescale box constrained optimization. Applied Mathematics and Computation, 265(??):568–569, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006815.

Ferchichi:2018:CLF

[FA18] Hanen Ferchichi and Saloua Mani Aouadi. A conforming locking-free approximation for a Koiter shell. Applied Mathematics and Computation, 339(??):367–373, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318306052.

Fan:2016:BCN

[FAHZ16] Jishan Fan, Bashir Ahmad, Tasawar Hayat, and Yong Zhou. On blow-up criteria for a new Hall– MHD system. Applied Mathematics and Computation, 274(??):20–24, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315014241.

Fan:2015:UWR

[Fan15a]  Zhencheng Fan.  Us-

ing waveform relaxation methods to approximate neutral stochastic functional differential equation systems. Applied Mathematics and Computation, 263(??):151–164, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004968.

Fang:2015:SCD

[Fan15b]  Mei Fang.  Synchro-

nization for complex dynamical networks with time delay and discretetime information. Applied Mathematics and Computation, 258(??):1– 11, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001381.

Fan:2017:CNS

[Fan17] Zhencheng Fan. Convergence of numerical solutions to stochastic differential equations with Markovian switching. Applied Mathematics and Computation, 315(??):176–187, December 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317305222.

Fan:2018:CRA

[Fan18] Xiaoming Fan. On convergent rate of the attractor for a singularly perturbed wave equation. Applied Mathematics and Computation, 316(??):370–380, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305143.

Farouki:2016:RRM

[Far16] Rida T. Farouki. Rational rotation-minimizing frames — recent advances and open problems. Applied Mathematics and Computation, 272 (part 1)(??):80–91, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005937.

Farouki:2018:RDP

[Far18a] Rida T. Farouki. Reduced difference polynomials and self-intersection computations. Applied Mathematics and Computation, 324(??):174–190, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308755.

Farrugia:2018:IRE

[Far18b] Alexander Farrugia. The increase in the resolvent energy of a graph due to the addition of a new edge. Applied Mathematics and Computation, 321(??):25–36, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307191.

Faustino:2017:HFS

[Fau17]  N. Faustino.  Hyper-

complex Fock states for discrete electromagnetic Schrödinger operators: a Bayesian probability perspective. Applied Mathematics and Computation, 315(??):531–548, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305416.

Ferreira:2019:LSP

[FB19] José Claudinei Ferreira and Maria Caruline Baquião. A least square point of view to reproducing kernel methods to solve functional equations. Applied Mathematics and Computation, 357(??):206–221, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031930284X.

Fernandez:2018:SPF

[FBF18] Arran Fernandez, Dumitru Baleanu, and Athanassios S. Fokas. Solving PDEs of fractional order using the unified transform method. Applied Mathematics and Computation, 339(??):727–737, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318306271.

Fei:2017:NSE

[FCM17] Jinxi Fei, Weiping Cao, and Zhengyi Ma. Nonlocal symmetries and explicit solutions for the Gardner equation. Applied Mathematics and Computation, 314(??):293–298, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304538.

Feng:2017:NPE

[FCW^{+}17] Fan Feng, Xuebin Chi, Zifa Wang, Jie Li, Jinrong Jiang, and Wenyi Yang. A nonnegativity preserved efficient chemical solver applied to the air pollution forecast. Applied Mathematics and Computation, 314(??):44–57, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304204.

Fedele:2018:VKB

[FDPP18] Giuseppe Fedele, Luigi D’Alfonso, Gilberto Pin, and Thomas Parisini. Volterra’s kernels-based finite-time parameters estimation of the Chua system. Applied Mathematics and Computation, 318(??):121–130, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305908.

Feng:2015:SNG

[Fen15]  Qinghua Feng.  Some

new generalized Gronwall– Bellman type discrete fractional inequalities. Applied Mathematics and Computation, 259(??):403–411, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003069.

Fercec:2015:ICL

[Fer15] Brigita Fercec. On integrability conditions and limit cycle bifurcations for polynomial systems. Applied Mathematics and Computation, 263(??):94–106, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004701.

Fusi:2018:PAF

[FF18] L. Fusi and A. Farina. Peristaltic axisymmetric flow of a Bingham fluid. Applied Mathematics and Computation, 320(??):1–15, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306380.

Farber:2016:DVC

[FFG^{+}16] K. Farber, P. Farber, J. Gräbel, S. Krick, J. Reitz, and P. Ueberholz. Development and validation of a coupled Navier– Stokes/DSMC simulation for rarefied gas flow in the production process for OLEDs. Applied Mathematics and Computation, 272 (part 3)(??):648–656, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006578.

Ferreiro-Ferreiro:2019:BHS

[FFGRSV19] Ana M. Ferreiro-Ferreiro, José A. García-Rodríguez, Luis Souto, and Carlos Vázquez. Basin hopping with synched multi L-BFGS local searches. Parallel implementation in multi-CPU and GPUs. Applied Mathematics and Computation, 356(??): 282–298, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300319301444.

Furst:2015:NSC

[FFH^{+}15] Jirí Fürst, Jaroslav Fort, Jan Halama, Jirí Holman, Jan Karel, Vladimír Prokop, and David Trdlicka. Numerical simulation of circumferentially averaged flow in a turbine. Applied Mathematics and Computation, 267(??):498–505, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001186.

Fort:2018:NSF

[FFH^{+}18] J. Fořt, J. Fürst, J. Halama, V. Hric, P. Louda, and M. Luxa D. Šimurda. Numerical simulation of flow through cascade in wind tunnel test section and stand-alone configurations. Applied Mathematics and Computation, 319(??):633–646, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305015.

Fu:2016:NSM

[FFL16] Yong Fu, Qihong Fang, and Jia Li. Numerical study on the mechanical response of brain under the impact loading based on elasticviscoelastic model. Applied Mathematics and Computation, 289(??):258–271, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303034.

Falcon:2018:ACG

[FFN18] O. J. Falcón, R. M. Falcón, and J. Núñez. Algebraic computation of genetic patterns related to threedimensional evolution algebras. Applied Mathematics and Computation, 319(??):510–517, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303466.

Feng:2015:EIP

[FFT15] Xingfang Feng, Hanying Feng, and Huixuan Tan. Existence and iteration of positive solutions for third-order Sturm–Liouville boundary value problems with plaplacian. Applied Mathematics and Computation, 266(??):634–641, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007493.

Frank:2016:LSL

[FFTB16] Alvaro Frank, Diego Fabregat-Traver, and Paolo Bientinesi. Large-scale linear regression: Development of high-performance routines. Applied Mathematics and Computation, 275(??):411–421, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015805.

Forouzannia:2015:DAS

[FG15] Farinaz Forouzannia and A. Gumel. Dynamics of an age-structured twostrain model for malaria transmission. Applied Mathematics and Computation, 250(??):860–886, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314013526.

Fernandez:2018:HAC

[FG18] Francisco M. Fernández and Javier Garcia. Highly accurate calculation of the resonances in the Stark effect in hydrogen. Applied Mathematics and Computation, 317(??):101–108, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031730601X.

Fercec:2019:BMC

[FG19] Brigita Fercec and Jaume Giné. Blow-up method to compute necessary conditions of integrability for planar differential systems. Applied Mathematics and Computation, 358(??):16–24, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302838.

Gao:2015:ABC

[fGlHyL^{+}15] Wei feng Gao, Ling ling Huang, San yang Liu, Felix T. S. Chan, Cai Dai, and Xian Shan. Artificial bee colony algorithm with multiple search strategies. Applied Mathematics and Computation, 271(??):269–287, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012370.

Franco:2018:MMB

[FGPR18] Sebastião Romero Franco, Francisco José Gaspar, Marcio Augusto Villela Pinto, and Carmen Rodrigo. Multigrid method based on a space-time approach with standard coarsening for parabolic problems. Applied Mathematics and Computation, 317(??):25–34, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305945.

Feng:2019:QSA

[FHK^{+}19] Yan-Quan Feng, Ademir Hujdurović, István Kovács, Klavdija Kutnar, and Dragan Marusic. Quasisemiregular automorphisms of cubic and tetravalent arc-transitive graphs. Applied Mathematics and Computation, 353(??): 329–337, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319300633.

Feng:2019:RBR

[FHLL19] Zhimin Feng, Jing Huang, Shuchao Li, and Xiaobing Luo. Relationship between the rank and the matching number of a graph. Applied Mathematics and Computation, 354(??):411–421, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301596.

Feng:2016:MPR

[FHM16] Ling Feng, Zhigang Huang, and Xuerong Mao. Mean percentage of returns for stock market linked savings accounts. Applied Mathematics and Computation, 273(??):1130–1147, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012874.

Francomano:2018:SRI

[FHPV18] Elisa Francomano, Frank M.

Hilker, Marta Paliaga, and Ezio Venturino. Separatrix reconstruction to identify tipping points in an eco-epidemiological model. Applied Mathematics and Computation, 318(??):80– 91, February 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317304733.

Falconi:2015:SGD

[FHV15] Manuel Falconi, Marcelo Huenchucona, and Claudio Vidal. Stability and global dynamic of a stagestructured predator–prey model with group defense mechanism of the prey. Applied Mathematics and Computation, 270(??):47–61, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010310.

Fister:2019:NSG

[FIG^{+}19] Iztok Fister, Andres Iglesias, Akemi Galvez, Javier Del Ser, Eneko Osaba, Iztok Fister, Matjaž Perc, and Mitja Slavinec. Novelty search for global optimization. Applied Mathematics and Computation, 347(??):865–881, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310269.

Fiori:2015:KNM

[Fio15] Simone Fiori. Kolmogoroff– Nagumo mean over the affine symplectic group of matrices. Applied Mathematics and Computation, 266(??):820–837, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006876.

Fiori:2016:RSD

[Fio16] Simone Fiori. A Riemannian steepest descent approach over the inhomogeneous symplectic group: Application to the averaging of linear optical systems. Applied Mathematics and Computation, 283(??):251–264, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301217.

Feng:2019:IDS

[FJ19] Shanshan Feng and Zhen Jin. Infectious diseases spreading on a metapopulation network coupled with its second-neighbor network. Applied Mathematics and Computation, 361(??):87–97, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303923.

Fairbairn:2017:SAN

[FK17] Abigail I. Fairbairn and Mark A. Kelmanson. Spectrally accurate Nyströmsolver error bounds for 1-D Fredholm integral equations of the second kind. Applied Mathematics and Computation, 315(??):211–223, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300317304952.

Filippov:2018:HPS

[FK18a] Anatoly Filippov and Yulia Koroleva. On a hydrodynamic permeability of a system of coaxial partly porous cylinders with superhydrophobic surfaces. Applied Mathematics and Computation, 338(??):346–362, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830523X.

Fukutomi:2018:HMC

[FK18b] Masao Fukutomi and Shun Kurokawa. How much cost should reciprocators pay in order to distinguish the opponent’s cooperation from the opponent’s defection? Applied Mathematics and Computation, 336(??):301–314, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304132.

Farooq:2017:EIE

[FKA17] Rashid Farooq, Mehtab Khan, and Faiz Ahmad. Extremal iota energy of bicyclic digraphs. Applied Mathematics and Computation, 303(??):24–33, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730036X.

Franco:2015:TNE

[FKR15] J. M. Franco, Y. Khiar, and L. Rández. Two new embedded pairs of explicit Runge–Kutta methods adapted to the numerical solution of oscillatory problems. Applied Mathematics and Computation, 252(??):45–57, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016294.

Ferjancic:2016:CRM

[FKV16] Karla Ferjancic, Marjeta Krajnc, and Vito Vitrih. Construction of G^{3} rational motion of degree eight. Applied Mathematics and Computation, 272 (part 1)(??):127–138, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011297.

Fernandez:2018:GAC

[FLCSV18] J. R. Fernández, J. A. López-Campos, A. Segade, and J. A. Vilán. A genetic algorithm for the characterization of hyperelastic materials. Applied Mathematics and Computation, 329(??):239–250, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301024.

Feng:2015:NAA

[FLL15] Zhiguang Feng, Wenxing Li, and James Lam. New admissibility analysis for discrete singular systems with timevarying delay. Applied Mathematics and Computation, 265(??):1058–1066, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007997.

Ford:2017:NIN

[FLL17] Neville J. Ford, Pedro M. Lima, and Patricia M. Lumb. Numerical investigation of noise induced changes to the solution behaviour of the discrete FitzHugh–Nagumo equation. Applied Mathematics and Computation, 293(??):448–460, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305306.

Fu:2019:MLS

[FLL^{+}19] Fangyan Fu, Jiao Li, Jun Lin, Yanjin Guan, Fuzheng Gao, Cunsheng Zhang, and Liang Chen. Moving least squares particle hydrodynamics method for Burgers’ equation. Applied Mathematics and Computation, 356(??):362–378, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302474.

Lv:2016:CAG

[fLLPY16] Yun fei Lv, Tongtong Li, Yongzhen Pei, and Rong Yuan. A complete analysis of the global dynamics of a diffusive predator and toxic prey model. Applied Mathematics and Computation, 291(??):182–196, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300316304118.

Fang:2015:GAP

[FLMC15] Bin Fang, Xue-Zhi Li, Maia Martcheva, and Li-Ming Cai. Global asymptotic properties of a heroin epidemic model with treat-age. Applied Mathematics and Computation, 263(??):315–331, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005135.

Favati:2018:IRN

[FLMR18] P. Favati, G. Lotti, O. Menchi, and F. Romani. Image regularization by nonnegatively constrained conjugate gradient. Applied Mathematics and Computation, 327(??):35–45, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300249.

Florek:2018:CGT

[Flo18]  Wojciech Florek.  A

class of generalized Tribonacci sequences applied to counting problems. Applied Mathematics and Computation, 338(??):789–808, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305022.

Fister:2015:CIS

[FLS^{+}15] Iztok Fister, Jr., Karin Ljubic, Ponnuthurai Nagaratnam Suganthan, Matjaz Perc, and Iztok Fister. Computational intelligence in sports: Challenges and opportunities within a new research domain. Applied Mathematics and Computation, 262(??):178–186, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004300.

Ferreira:2018:AES

[FLS18] Chelo Ferreira, José L.

López, and Ester Pérez Sinusía. The asymptotic expansion of the swallowtail integral in the highly oscillatory region. Applied Mathematics and Computation, 339(??):820–836, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318305666.

Ferreyra:2018:MCM

[FLT18] D. E. Ferreyra, F. E. Levis, and N. Thome. Maximal classes of matrices determining generalized inverses. Applied Mathematics and Computation, 333(??):42–52, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302832.

Fu:2018:GOR

[FLYG18] Baozeng Fu, Shihua Li, Jun Yang, and Lei Guo. Global output regulation for a class of single input port-controlled Hamiltonian disturbed systems. Applied Mathematics and Computation, 325(??):322–331, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300055.

Fu:2016:PCS

[FM16] Lei Fu and Yuechao Ma. Passive control for singular time-delay system with actuator saturation. Applied Mathematics and Computation, 289(??):181–193, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302995.

Fareh:2017:EDP

[FM17a] Abdelfeteh Fareh and Salim A. Messaoudi. Energy decay for a porous thermoelastic system with thermoelasticity of second sound and with a non-necessary positive definite energy. Applied Mathematics and Computation, 293(??):493–507, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305355.

Fernandez:2017:PTP

[FM17b] J. R. Fernández and M. Masid. A porous thermoelastic problem: an a priori error analysis and computational experiments. Applied Mathematics and Computation, 305(??):117–135, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300966.

Frolkovic:2018:SIS

[FM18] Peter Frolkovic and Karol Mikula. Semi-implicit second order schemes for numerical solution of level set advection equation on Cartesian grids. Applied Mathematics and Computation, 329(??):129–142, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300948.

Farwa:2018:FAI

[FMB^{+}18] Shabieh Farwa, Nazeer Muhammad, Nargis Bibi, Sajjad A. Haider, Syed R. Naqvi, and Sheraz Anjum. Fresnelet approach for image encryption in the algebraic frame. Applied Mathematics and Computation, 334(??):343–355, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302868. See

retraction notice [FMB^{+}19].

Farwa:2019:RNF

[FMB^{+}19] Shabieh Farwa, Nazeer Muhammad, Nargis Bibi, Sajjad A. Haider, Syed R. Naqvi, and Sheraz Anjum. Retraction notice to “Fresnelet approach for image encryption in the algebraic frame” [Applied Mathematics and Computation 334 (2018) 343–355]. Applied Mathematics and Computation, 341(??):451, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300318308038.

See [FMB^{+}18].

Fei:2015:SRE

[FMC15] Jinxi Fei, Zhengyi Ma, and Yuanming Chen. Symmetry reduction and explicit solutions of the (2 + 1)- dimensional Boiti–Leon– Pempinelli system. Applied Mathematics and Computation, 268(??): 432–438, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031500867X.

Fang:2019:FTS

[FMDZ19] Liandi Fang, Li Ma, Shihong Ding, and Dean Zhao. Finite-time stabilization for a class of highorder stochastic nonlinear systems with an output constraint. Applied Mathematics and Computation, 358(??):63–79, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302759.

Flores-Marquez:2015:MDF

[FMRRT15] E. L. Flores-Márquez, A. Ramírez-Rojas, and L. Telesca. Multifractal detrended fluctuation analysis of earthquake magnitude series of Mexican South Pacific Region. Applied Mathematics and Computation, 265(??):1106–1114, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008103.

Fernandez-Martinez:2015:HCH

[FMSG15] M. Fernández-Martínez and M. A. Sánchez-Granero. How to calculate the Hausdorff dimension using fractal structures. Applied Mathematics and Computation, 264(??):116–131, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005172.

Fang:2015:FMC

[FMX15] Chunhua Fang, Junjie Ma, and Meiying Xiang. On Filon methods for a class of Volterra integral equations with highly oscillatory Bessel kernels. Applied Mathematics and Computation, 268(??):783–792, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008929.

Firman:2015:MSD

[FNS15] Firman, Janson Naiborhu, and Roberd Saragih. Modification of a steepest descent control for output tracking of some class non-minimum phase nonlinear systems. Applied Mathematics and Computation, 269(??):497–506, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501019X.

Fukshansky:2019:API

[FNS19] Lenny Fukshansky, Deanna Needell, and Benny Sudakov. An algebraic perspective on integer sparse recovery. Applied Mathematics and Computation, 340(??):31–42, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318306465.

Fernandez:2019:FCG

[FÖB19] Arran Fernandez, Mehmet Ali Özarslan, and Dumitru Baleanu. On fractional calculus with general analytic kernels. Applied Mathematics and Computation, 354(??):248–265, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301493.

far:2015:MLS

[fP15] Mashallah Matin far and Masoumeh Pourabd. Moving least square for systems of integral equations. Applied Mathematics and Computation, 270(??):879–889, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315011546.

Fabiano:2018:SAS

[FP18a] R. H. Fabiano and Catherine Payne. Spline approximation for systems of linear neutral delaydifferential equations. Applied Mathematics and Computation, 338(??): 774–788, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318305526.

Francomano:2018:HNI

[FP18b] E. Francomano and M. Paliaga.

Highlighting numerical insights of an efficient SPH method. Applied Mathematics and Computation, 339(??):888–898, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830626X.

Fister:2015:RCB

[FPKF15] Iztok Fister, Jr., Matjaz Perc, Salahuddin M. Kamal, and Iztok Fister. A review of chaos-based firefly algorithms: Perspectives and research challenges. Applied Mathematics and Computation, 252(??):155–165, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314016518.

Farajzadeh:2015:GME

[FPUK15] A. P. Farajzadeh, S. Plubtieng, K. Ungchittrakool, and D. Kumtaeng. Generalized mixed equilibrium problems with generalized α–η-monotone bifunction in topological vector spaces. Applied Mathematics and Computation, 265(??):313–319, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006293.

Feng:2019:ASH

[FQMR19] Tao Feng, Zhipeng Qiu, Xinzhu Meng, and Libin Rong. Analysis of a stochastic HIV-1 infection model with degenerate diffusion. Applied Mathematics and Computation, 348(??):437–455, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310488.

Franco:2016:EEF

[FR16] J. M. Franco and L. Rández. Explicit exponentially fitted two-step hybrid methods of high order for second-order oscillatory IVPs. Applied Mathematics and Computation, 273(??):493–505, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013739.

Furfaro:2016:MDP

[FS16] Damien Furfaro and Richard Saurel. Modeling droplet phase change in the presence of a multi-component gas mixture. Applied Mathematics and Computation, 272 (part 2)(??): 518–541, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315002830.

Felix:2016:AMC

[FSW16] Bryan Félix, Anne Shiu, and Zev Woodstock. Analyzing multistationarity in chemical reaction networks using the determinant optimization method. Applied Mathematics and Computation, 287–288(??):60–73, September 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300316302806.

Famelis:2016:MRK

[FT16] I. Th. Famelis and Ch. Tsitouras. On modifications of Runge–Kutta-Nyström methods for solving y^{(4)} = f(x, y). Applied Mathematics and Computation, 273(??):726–734, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013661.

Famelis:2017:SDR

[FT17] Ioannis Th. Famelis and Ch. Tsitouras. Symbolic derivation of Runge– Kutta–Nyström type order conditions and methods for solving y′′′ = f(x, y). Applied Mathematics and Computation, 297(??):50–60, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630635X.

Fei:2018:CCB

[FT18] Junqi Fei and Jianhua Tu. Complete characterization of bicyclic graphs with the maximum and second-maximum degree Kirchhoff index. Applied Mathematics and Computation, 330(??):118–124, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301279.

Farooq:2019:PMF

[FTE19] A. A. Farooq, D. Tripathi, and Thanaa Elnaqeeb. On the propulsion of micropolar fluid inside a channel due to ciliary induced metachronal wave. Applied Mathematics and Computation, 347(??):225–235, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308877.

Fister:2016:EAC

[FTF16] Iztok Fister, Aleksandra Tepeh, and Iztok Fister, Jr. Epistatic arithmetic crossover based on Cartesian graph product in ensemble differential evolution. Applied Mathematics and Computation, 283(??):181–194, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630145X.

Fuka:2015:PFP

[Fuk15a] V. Fuka. PoisFFT — a free parallel fast Poisson solver. Applied Mathematics and Computation, 267(??):356–364, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003057.

Fukushima:2015:PFCb

[Fuk15b] Toshio Fukushima. Precise and fast computation of Fermi–Dirac integral of integer and half integer order by piecewise minimax rational approximation. Applied Mathematics and Computation, 259(??):708–729, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003033.

Fukushima:2015:PFC

[Fuk15c] Toshio Fukushima. Precise and fast computation of generalized Fermi– Dirac integral by parameter polynomial approximation. Applied Mathematics and Computation, 270(??):802–807, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011509.

Fukushima:2015:PFCa

[Fuk15d] Toshio Fukushima. Precise and fast computation of inverse Fermi–Dirac integral of order 1/2 by minimax rational function approximation. Applied Mathematics and Computation, 259(??):698–707, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003094.

Fermo:2015:PEM

[FvdMS15] L. Fermo, C. van der Mee, and S. Seatzu. Parameter estimation of monomialexponential sums in one and two variables. Applied Mathematics and Computation, 258(??):576–586, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002131.

Florio:2015:MTA

[FVMO15] B. J. Florio, M. Vynnycky, S. L. Mitchell, and S. B. G. O’Brien. Mouldtaper asymptotics and air gap formation in continuous casting. Applied Mathematics and Computation, 268(??):1122–1139, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009236.

Fan:2018:GLF

[FW18a] Ming-Can Fan and Yue Wu. Global leaderfollowing consensus of nonlinear multi-agent systems with unknown control directions and unknown external disturbances. Applied Mathematics and Computation, 331(??):274–286, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301723.

Feng:2018:ICE

[FW18b] Huicheng Feng and Teck Neng Wong. Induced-charge electro-osmosis in dielectric annuli. Applied Mathematics and Computation, 333(??):133–144, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S009630031830287X.

Feng:2015:NPE

[FWLC15] Fan Feng, Zifa Wang, Jie Li, and Gregory R. Carmichael. A nonnegativity preserved efficient algorithm for atmospheric chemical kinetic equations. Applied Mathematics and Computation, 271(??):519–531, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012692.

Fang:2019:SMQ

[FWZ19] Longfei Fang, Bing Wang, and Mingqing Zhai. On the spectral moment of quasi-bicyclic graphs. Applied Mathematics and Computation, 363(??): Article 124601, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305934.

Fang:2017:SGB

[FXC17] He Fang, Li Xu, and Kim-Kwang Raymond Choo. Stackelberg game based relay selection for physical layer security and energy efficiency enhancement in cognitive radio networks. Applied Mathematics and Computation, 296(??):153–167, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306208.

Feng:2019:MRT

[FXXC19] Hongyan Feng, Huiling Xu, Shengyuan Xu, and Weimin Chen. Model reference tracking control for spatially interconnected discrete-time systems with interconnected chains. Applied Mathematics and Computation, 340(??):50–62, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307422.

Firoozjaee:2018:NAF

[FY18] M. A. Firoozjaee and S. A. Yousefi. A numerical approach for fractional partial differential equations by using Ritz approximation. Applied Mathematics and Computation, 338(??):698–710, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304521.

Fortin:2019:ADG

[FY19] A. Fortin and D. Yakoubi.

An adaptive discontinuous Galerkin method for very stiff systems of ordinary differential equations. Applied Mathematics and Computation, 358(??):330–347, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302875.

Feng:2018:SMN

[FYSC18] Yuming Feng, Xinsong Yang, Qiang Song, and Jinde Cao. Synchronization of memristive neural networks with mixed delays via quantized intermittent control. Applied Mathematics and Computation, 339(??):866–873, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306489.

Feng:2016:CSN

[FYZ16] Jianwen Feng, Pan Yang, and Yi Zhao. Cluster synchronization for nonlinearly time-varying delayed coupling complex networks with stochastic perturbation via periodically intermittent pinning control. Applied Mathematics and Computation, 291(??):52–68, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304015.

Fu:2019:SSF

[FZG19] Xiaozheng Fu, Quanxin Zhu, and Yingxin Guo. Stabilization of stochastic functional differential systems with delayed impulses. Applied Mathematics and Computation, 346(??):776–789, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309299.

Fan:2019:MBP

[FZL19] Hong-Tao Fan, Yan-Jun Zhang, and Ya-Jing Li. A modified block preconditioner for complex nonsymmetric indefinite linear systems. Applied Mathematics and Computation, 358(??):455–467, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303406.

Feng:2015:SCN

[FZLT15] L. B. Feng, P. Zhuang, F. Liu, and I. Turner. Stability and convergence of a new finite volume method for a twosided space-fractional diffusion equation. Applied Mathematics and Computation, 257(??):52–65, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017147.

Fan:2019:VRT

[FZLZ19] Hongtao Fan, Xinyun Zhu, Yajing Li, and Bing Zheng. A variant of relaxed triangular splitting preconditioners for generalized saddle point problems from Navier–Stokes equations. Applied Mathematics and Computation, 362 (??):Article 124495, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304746.

Feng:2018:EPM

[FZZ18] Xing Feng, Lianzhu Zhang, and Mingzu Zhang. Enumeration of perfect matchings of lattice graphs by Pfaffians. Applied Mathematics and Computation, 338(??):400–411, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305150.

Gadjiev:2015:AAF

[GA15] Akif D. Gadjiev and Rashid A. Aliev. Approximation of analytic functions in annulus by linear operators. Applied Mathematics and Computation, 252(??):438–445, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016798.

Ghalavand:2017:EGR

[GA17] A. Ghalavand and A. R. Ashrafi. Extremal graphs with respect to variable sum exdeg index via majorization. Applied Mathematics and Computation, 303(??):19–23, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300152.

Ghalavand:2018:OCG

[GA18] Ali Ghalavand and Ali Reza Ashrafi. Ordering chemical graphs by Randić and sum-connectivity numbers. Applied Mathematics and Computation, 331(??):160–168, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301590.

Gaviraghi:2017:ASM

[GAB17] B. Gaviraghi, M. Annunziato, and A. Borzì. Analysis of splitting methods for solving a partial integrodifferential Fokker–Planck equation. Applied Mathematics and Computation, 294(??):1–17, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305537.

Gao:2016:MMS

[Gao16] Rong Gao. Milne method for solving uncertain differential equations. Applied Mathematics and Computation, 274(??):

774–785, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315015416.

Gao:2017:NON

[Gao17a] Jianfang Gao. Numerical oscillation and nonoscillation for differential equation with piecewise continuous arguments of mixed type. Applied Mathematics and Computation, 299(??):16–27, April 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630707X.

Gao:2017:SKV

[Gao17b] Peng Gao. The stochastic Korteweg–de Vries equation on a bounded domain. Applied Mathematics and Computation, 310(??):97–111, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302837.

Gao:2017:UWE

[Gao17c]  Rong Gao.  Uncertain

wave equation with infinite half-boundary. Applied Mathematics and Computation, 304(??):28–40, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307299.

Gao:2018:ROK

[Gao18] Zhe Gao. Reduced order Kalman filter for a continuous-time fractionalorder system using fractionalorder average derivative. Applied Mathematics and Computation, 338(??):55–71, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304946.

Gao:2019:SMU

[Gao19]  Rong Gao.  Stabil-

ity in mean for uncertain differential equation with jumps. Applied Mathematics and Computation, 346(??):15–22, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308555.

Garcia:2016:NSQ

[Gar16a] Isaac A. García. Nonsmooth quadratic centers defined in two arbitrary sectors. Applied Mathematics and Computation, 289(??):37–47, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303241.

Garg:2016:HPG

[Gar16b] Harish Garg. A hybrid PSO–GA algorithm for constrained optimization problems. Applied Mathematics and Computation, 274(??):292–305, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014630.

Gupta:2017:ABT

[GAS17] Vijay Gupta, Ana Maria Acu, and Daniel Florin Sofonea. Approximation of Baskakov type Pólya–Durrmeyer operators. Applied Mathematics and Computation, 294(??):318–331, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305720.

Gatto:2018:SAD

[Gat18a] R. Gatto. Saddlepoint approximation to the distribution of the total distance of the von Mises– Fisher continuous time random walk. Applied Mathematics and Computation, 324(??):285–294, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308895.

Gros:2018:MMM

[GAT18b] E. Gros, G. Anjos, and J. Thome. Moving mesh method for direct numerical simulation of twophase flow with phase change. Applied Mathematics and Computation, 339(??):622–635, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306180.

Gavrea:2015:HHT

[Gav15]  Bogdan Gavrea.  A

Hermite–Hadamard type inequality with applications to the estimation of moments of continuous random variables. Applied Mathematics and Computation, 254(??):92–98, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017901.

Gupta:2019:IMF

[GB19] S. Gupta and N. Bhengra. Influence of magnetoelasticity on frequency equation of shear wave propagation in a multilayered magnetoelastic anisotropic monoclinic medium. Applied Mathematics and Computation, 355(??):366–384, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301997.

Gholami:2019:FPI

[GBJ19] Saeid Gholami, Esmail Babolian, and Mohammad Javidi. Fractional pseudospectral integration/differentiation matrix and fractional differential equations. Applied Mathematics and Computation, 343(??):314–327, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307586.

Grillo:2015:PIR

[GBS15] Sebastian Grillo, Gerardo Blanco, and Christian E. Schaerer. Path integration for real options. Applied Mathematics and Computation, 265(??):120–132, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005731.

G:2018:SFR

[GCAA18] Soriano-Sánchez A. G., Posadas-Castillo C., Platas-Garza M. A., and Arellano-Delgado A. Synchronization and FPGA realization of complex networks with fractionalorder Liu chaotic oscillators. Applied Mathematics and Computation, 332(??):250–262, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302303.

Gao:2019:NIO

[GCH^{+}19] Jian-Guo Gao, Zhi-Hua Chen, Zhen-Gui Huang, Wei-Tao Wu, and Yu-Jie Xiao. Numerical investigations on the oblique water entry of high-speed projectiles. Applied Mathematics and Computation, 362 (??):Article 124547, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305302.

Gomez-Corral:2015:LRM

[GCLG15] A. Gómez-Corral and M. López-García. Lifetime and reproduction of a marked individual in a two-species competition process. Applied Mathematics and Computation, 264(??):223–245, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005238.

Guerra:2018:DAA

[GCM^{+}18] Telma Guerra, Catarina Catarino, Tânia Mestre, Sara Santos, Jorge Tiago, and Adélia Sequeira. A data assimilation approach for non-Newtonian blood flow simulations in 3D geometries. Applied Mathematics and Computation, 321(??):176–194, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317307282.

Garcia:2018:ENN

[GCMR18] M. A. García, M. A.

Castro, J. A. Martín, and F. Rodríguez. Exact and nonstandard numerical schemes for linear delay differential models. Applied Mathematics and Computation, 338(??):314–336, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305186.

Gautam:2015:MSC

[GD15] Ganga Ram Gautam and Jaydev Dabas. Mild solutions for class of neutral fractional functional differential equations with not instantaneous impulses. Applied Mathematics and Computation, 259(??):480–489, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002696.

Gutman:2017:GS

[GDF^{+}17] Ivan Gutman, Kinkar Ch.

Das, Boris Furtula, Emina Milovanović, and Igor Milovanović. Generalizations of Szőkefalvi Nagy and Chebyshev inequalities with applications in spectral graph theory. Applied Mathematics and Computation, 313(??):235–244, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303764.

Gao:2016:FPE

[GDL16] Ting Gao, Jinqiao Duan, and Xiaofan Li. Fokker– Planck equations for stochastic dynamical systems with symmetric Lévy motions. Applied Mathematics and Computation, 278(??):1–20, March 31, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300108.

Gu:2019:NDP

[GDL19] Ran Gu, Bo Deng, and Rui Li. Note on directed proper connection number of a random graph. Applied Mathematics and Computation, 361(??):169–174, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304217.

Gallegos:2016:LTF

[GDM16] Javier A. Gallegos and Manuel A. Duarte-Mermoud. On the Lyapunov theory for fractional order systems. Applied Mathematics and Computation, 287– 288(??):161–170, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302892.

Ghorbani:2019:HEF

[GDRP^{+}19] Modjtaba Ghorbani, Matthias Dehmer, Mina Rajabi-Parsa, Frank Emmert-Streib, and Abbe Mowshowitz. Hosoya entropy of fullerene graphs. Applied Mathematics and Computation, 352(??):88–98, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300335.

Ghosh:2016:TSA

[GDS16] Suchismita Ghosh, Anish Deb, and Gautam Sarkar. Taylor series approach for function approximation using ‘estimated’ higher derivatives.

Applied Mathematics and Computation, 284(??):89– 101, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301801.

Gao:2018:EDA

[GdS18] Shujun Gao and Clarence W. de Silva. Estimation distribution algorithms on constrained optimization problems. Applied Mathematics and Computation, 339(??):308–322, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306027.

Ghorbani:2018:GOB

[GDZ18] Modjtaba Ghorbani, Matthias Dehmer, and Samaneh Zangi. Graph operations based on using distance-based graph entropies. Applied Mathematics and Computation, 333(??):547–555, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303072.

Guias:2016:ISD

[GE16] Flavius Guias and Pavel Eremeev. Improving the stochastic direct simulation method with applications to evolution partial differential equations. Applied Mathematics and Computation, 289(??):353–370, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303423.

Gutman:2016:SDDb

[GFD16] Ivan Gutman, Boris Furtula, and Kinkar Ch. Das. On some degree-and-distance-based graph invariants of trees. Applied Mathematics and Computation, 289(??):1–6, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302909.

Gambuzza:2019:SOC

[GFE19] Lucia Valentina Gambuzza, Mattia Frasca, and Ernesto Estrada. Secondorder consensus protocols based on transformed d-path Laplacians. Applied Mathematics and Computation, 343(??):183–194, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308166.

Ghasemi:2015:NSN

[GFG15] M. Ghasemi, M. Fardi, and R. Khoshsiar Ghaziani. Numerical solution of nonlinear delay differential equations of fractional order in reproducing kernel Hilbert space. Applied Mathematics and Computation, 268(??):815–831, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007936.

Guo:2016:FOF

[GfSmL16] Yan Guo, Yu feng Shi, and Yi min Li. A fifth-order finite volume weighted compact scheme for solving one-dimensional Burgers’ equation. Applied Mathematics and Computation, 281(??):172–185, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300704.

Gao:2017:EVA

[GFWH17] Wei Gao, Mohammad Reza Farahani, Shaohui Wang, and Mohamad Nazri Husin. On the edge-version atombond connectivity and geometric arithmetic indices of certain graph operations. Applied Mathematics and Computation, 308(??):11–17, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301571.

Grubisic:2015:DPE

[GG15a] Luka Grubisić and Antonia Grbić. Discrete perturbation estimates for eigenpairs of Fredholm operator-valued functions. Applied Mathematics and Computation, 267(??): 632–647, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315000247.

Gupta:2015:MEN

[GG15b] Vijay Gupta and G. C.

Greubel. Moment estimations of new Szász– Mirakyan–Durrmeyer operators. Applied Mathematics and Computation, 271(??):540–547, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315012734.

Gutt:2015:LDP

[GG15c] Robert Gutt and Teodor Grosan. On the lid-driven problem in a porous cavity. A theoretical and numerical approach. Applied Mathematics and Computation, 266(??):1070–1082, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031500819X.

Gakkhar:2016:TSD

[GG16] Sunita Gakkhar and Komal Gupta. A three species dynamical system involving prey-predation, competition and commensalism. Applied Mathematics and Computation, 273(??):54–67, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315012722.

Gori:2018:VHG

[GG18] Giulio Gori and Alberto Guardone. VirtuaSchlieren: A hybrid GPU/CPU-based schlieren simulator for ideal and non-ideal compressiblefluid flows. Applied Mathematics and Computation, 319(??):647–661, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305027.

Gupta:2019:WWI

[GG19a] Sourav Gupta and R. Gayen.

Water wave interaction with dual asymmetric non-uniform permeable plates using integral equations. Applied Mathematics and Computation, 346(??):436–451, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309287.

Gutierrez-Gutierrez:2019:MVP

[GG19b] Jesús Gutiérrez-Gutiérrez.

A modified version of the Pisarenko method to estimate the power spectral density of any asymptotically wide sense stationary vector process. Applied Mathematics and Computation, 362(??): Article 124526, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305053.

Galan-Garcia:2018:ICC

[GGAVGG^{+}18] José L. Galán-García, Gabriel Aguilera-Venegas, María Á. Galán-García, Pedro Rodríguez-Cielos, and Iván Atencia-Mc. Killop. Improving CAS capabilities: New rules for computing improper integrals. Applied Mathematics and Computation, 316(??):525–540, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307585.

Galan-Garcia:2015:NPE

[GGAVGGRC15] José L. Galán-García, Gabriel Aguilera-Venegas, María Á. Galán-García, and Pedro Rodríguez-Cielos. A new Probabilistic Extension of Dijkstra’s Algorithm to simulate more realistic traffic flow in a smart city. Applied Mathematics and Computation, 267(??):780–789, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016087.

Garba:2015:SES

[GGHL15] S. M. Garba, A. B. Gumel, A. S. Hassan, and J. M.- S. Lubuma. Switching from exact scheme to nonstandard finite difference scheme for linear delay differential equation. Applied Mathematics and Computation, 258(??):388–403, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001204.

Giani:2018:PEE

[GGHO18] Stefano Giani, Luka Grubisić, Harri Hakula, and Jeffrey S. Ovall. An a posteriori estimator of eigenvalue/eigenvector error for penalty-type discontinuous Galerkin methods. Applied Mathematics and Computation, 319(??):562–574, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304587.

Gadella:2015:PAA

[GGL15] M. Gadella, H. Giacomini, and L. P. Lara. Periodic analytic approximate solutions for the Mathieu equation. Applied Mathematics and Computation, 271(??):436–445, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012369.

Ghosh:2015:NED

[GGM15] Bapan Ghosh, Frédéric Grognard, and Ludovic Mailleret. Natural enemies deployment in patchy environments for augmentative biological control. Applied Mathematics and Computation, 266(??):982–999, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008024.

Gaudioso:2018:NIV

[GGM18] Manlio Gaudioso, Giovanni Giallombardo, and Marat Mukhametzhanov. Numerical infinitesimals in a variable metric method for convex nonsmooth optimization. Applied Mathematics and Computation, 318(??):312–320, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305180.

Garza:2015:MAS

[GGMPC15] Lino G. Garza, Luis E.

Garza, Francisco Marcellán, and Natalia C. Pinzón-Cortés. A matrix approach for the semiclassical and coherent orthogonal polynomials. Applied Mathematics and Computation, 256(??):459–471, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001034.

Gutierrez-Gutierrez:2016:SVD

[GGZR16] Jesús Gutiérrez-Gutiérrez and Marta Zárraga-Rodríguez. Singular value decomposition for bidiagonal filter matrices. Applied Mathematics and Computation, 289(??):214–219, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303344.

Garcia:2017:IEP

[GH17] A. Garcia and K. W.

Hipel. Inverse engineering preferences in simple games. Applied Mathematics and Computation, 311(??):184–194, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303041.

Gao:2017:SAB

[GhDwZS17] Bo Gao, Zheng hong Deng, Da wei Zhao, and Qun Song. State analysis of Boolean control networks with impulsive and uncertain disturbances. Applied Mathematics and Computation, 301(??):187–192, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307160.

Gheorghiu:2016:SSC

[Ghe16] C. I. Gheorghiu. Stable spectral collocation solutions to a class of Benjamin Bona Mahony initial value problems. Applied Mathematics and Computation, 273(??):1090–1099, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014290.

Gao:2019:AFP

[GHL19] Peiting Gao, Chuanjiang He, and Yang Liu. An adaptive family of projection methods for constrained monotone nonlinear equations with applications. Applied Mathematics and Computation, 359(??):1–16, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319302723.

Gyo:2015:NDP

[GHM15] István Győri, Ferenc Hartung, and Nahed A. Mohamady. On a nonlinear delay population model. Applied Mathematics and Computation, 270(??):909–925, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011467.

Gonzalez:2018:MLD

[GHM18] Antonio González, Carmen Hernando, and Mercè Mora. Metric-locating-dominating sets of graphs for constructing related subsets of vertices. Applied Mathematics and Computation, 332(??):449–456, September 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318302261.

Garayshin:2016:ANS

[GHN^{+}16] V. V. Garayshin, M. W. Harris, D. J. Nicolsky, E. N. Pelinovsky, and A. V. Rybkin. An analytical and numerical study of long wave runup in U-shaped and Vshaped bays. Applied Mathematics and Computation, 279(??):187–197, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300054.

Gao:2017:AAD

[GHQ17] Xianwen Gao, Hangfeng He, and Wenhai Qi. Admissibility analysis for discrete-time singular Markov jump systems with asynchronous switching. Applied Mathematics and Computation, 313(??): 431–441, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317304034.

Gao:2019:ISC

[GHW19] Zhen-Man Gao, Yong He, and Min Wu. Improved stability criteria for the neural networks with timevarying delay via new augmented Lyapunov– Krasovskii functional. Applied Mathematics and Computation, 349(??): 258–269, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318310786.

Guo:2019:NEP

[GHZ19] Jiebin Guo, Chuanjiang He, and Xiaoting Zhang. Nonlinear edge-preserving diffusion with adaptive source for document images binarization. Applied Mathematics and Computation, 351(??):8– 22, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930030X.

Gavrea:2015:SEP

[GI15] Ioan Gavrea and Mircea Ivan. A sharp estimate for the Peano error representation. Applied Mathematics and Computation, 252(??):14–19, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016713.

Galvez:2016:PBM

[GI16] Akemi Gálvez and Andrés Iglesias. Particle-based meta-model for continuous breakpoint optimization in smooth local-support curve fitting. Applied Mathematics and Computation, 275(??):195–212, February 15, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315015520.

Giani:2015:ACD

[Gia15a] Stefano Giani. hp-adaptive composite discontinuous Galerkin methods for elliptic eigenvalue problems on complicated domains. Applied Mathematics and Computation, 267(??):604–617, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000454.

Giani:2015:SEE

[Gia15b] Stefano Giani. Solving elliptic eigenvalue problems on polygonal meshes using discontinuous Galerkin composite finite element methods. Applied Mathematics and Computation, 267(??):618–631, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000259.

Gledel:2019:SGC

[GIK19] Valentin Gledel, Vesna Irsic, and Sandi Klavzar. Strong geodetic cores and Cartesian product graphs. Applied Mathematics and Computation, 363(??): Article 124609, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306010.

Gokmen:2015:TCA

[GIS15] Elcin Gokmen, Osman Rasit Isik, and Mehmet Sezer. Taylor collocation approach for delayed Lotka–Volterra predator– prey system. Applied Mathematics and Computation, 268(??):671–684, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315008917.

Grishagin:2018:CCN

[GIS18] Vladimir Grishagin, Ruslan Israfilov, and Yaroslav Sergeyev. Convergence conditions and numerical comparison of global optimization methods based on dimensionality reduction schemes. Applied Mathematics and Computation, 318(??):270–280, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304496.

Gzyl:2015:EDA

[GIT15] Henryk Gzyl, Pierluigi Novi Inverardi, and Aldo Tagliani. Entropy and density approximation from Laplace transforms. Applied Mathematics and Computation, 265(??):225–236, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006372.

Gu:2018:PMS

[GJNC18] Chuanqing Gu, Xianglong Jiang, Ying Nie, and Zhibing Chen. A preprocessed multi-step splitting iteration for computing PageRank. Applied Mathematics and Computation, 338(??):72–86, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304429.

Gutierrez:2015:ETV

[GJNRG15] C. Gutiérrez, B. Jiménez, V. Novo, and G. Ruiz-Garzón. Efficiency through variational-like inequalities with Lipschitz functions. Applied Mathematics and Computation, 259(??):438–449, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500274X.

Graillat:2018:NVC

[GJP18] Stef Graillat, Fabienne Jézéquel, and Romain Picot. Numerical validation of compensated algorithms with stochastic arithmetic. Applied Mathematics and Computation, 329(??):339–363, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300985.

Garoosi:2015:NSN

[GJR^{+}15] Faroogh Garoosi, Leila Jahanshaloo, Mohammad Mehdi Rashidi, Arash Badakhsh, and Mohammed E. Ali. Numerical simulation of natural convection of the nanofluid in heat exchangers using a Buongiorno model. Applied Mathematics and Computation, 254(??):183–203, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017779.

Ge:2019:WPS

[GJX^{+}19] Mengyan Ge, Ya Jia, Ying Xu, Lulu Lu, Huiwen Wang, and Yunjie Zhao. Wave propagation and synchronization induced by chemical autapse in chain Hindmarsh–Rose neural network. Applied Mathematics and Computation, 352(??):136–145, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300761.

Grace:2019:OTO

[GJZ19] Said R. Grace, Irena Jadlovská, and Agacik Zafer. On oscillation of third-order noncanonical delay differential equations. Applied Mathematics and Computation, 362 (??):Article 124530, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305090.

Ge:2015:SAK

[GK15] Fudong Ge and Chunhai Kou. Stability analysis by Krasnoselskii’s fixed point theorem for nonlinear fractional differential equations. Applied Mathematics and Computation, 257(??):308–316, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016427.

Gazzola:2016:IRR

[GK16a] Silvia Gazzola and Anna Karapiperi. Image reconstruction and restoration using the simplified topological ϵ-algorithm. Applied Mathematics and Computation, 274(??): 539–555, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S009630031501509X.

Groselj:2016:CSP

[GK16b] Jan Groselj and Marjeta Krajnc. C^{1} cubic splines on Powell–Sabin triangulations. Applied Mathematics and Computation, 272 (part 1)(??):114–126, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500925X.

Gdawiec:2017:PPI

[GK17] Krzysztof Gdawiec and Wiesław Kotarski. Polynomiography for the polynomial infinity norm via Kalantari’s formula and nonstandard iterations. Applied Mathematics and Computation, 307(??):17– 30, August 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317301492.

Ghasemi:2019:UMO

[GKDHR19] Peiman Ghasemi, Kaveh Khalili-Damghani, Ashkan Hafezalkotob, and Sadigh Raissi. Uncertain multiobjective multi-commodity multi-period multi-vehicle location-allocation model for earthquake evacuation planning. Applied Mathematics and Computation, 350(??):105–132, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311226.

Geum:2016:BEK

[GKM16] Young Hee Geum, Young Ik Kim, and Á. Alberto Magreñán. A biparametric extension of King’s fourth-order methods and their dynamics. Applied Mathematics and Computation, 282(??):254–275, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301230.

Geum:2015:CTP

[GKN15a] Young Hee Geum, Young Ik Kim, and Beny Neta. A class of two-point sixthorder multiple-zero finders of modified double-Newton type and their dynamics. Applied Mathematics and Computation, 270(??):387–400, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315010942.

Geum:2015:DHO

[GKN15b] Young Hee Geum, Young Ik Kim, and Beny Neta. On developing a higherorder family of double-Newton methods with a bivariate weighting function. Applied Mathematics and Computation, 254(??):277–290, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017913.

Geum:2016:SOF

[GKN16] Young Hee Geum, Young Ik Kim, and Beny Neta. A sixth-order family of threepoint modified Newtonlike multiple-root finders and the dynamics behind their extraneous fixed points. Applied Mathematics and Computation, 283(??):120–140, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301400.

Ghebleh:2016:OQS

[GKS16] M. Ghebleh, A. Kanso, and D. Stevanović. Open quipus with the same Wiener index as their quadratic line graph. Applied Mathematics and Computation, 281(??): 130–136, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316300406.

Graca:2015:RFH

[GL15a] Mário M. Graça and Pedro M. Lima. Root finding by high order iterative methods based on quadratures. Applied Mathematics and Computation, 264(??):466–482, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005597.

Guo:2015:CSC

[GL15b] Rui Guo and Yue-Feng Liu. The canonical AB system: conservation laws and soliton solutions. Applied Mathematics and Computation, 259(??):153–163, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002088.

Gu:2018:EHF

[GL18a] Ran Gu and Rui Li. On extremal hypergraphs for forests of tight paths. Applied Mathematics and Computation, 325(??): 291–296, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317309037.

Guo:2018:CDS

[GL18b] Jia Guo and Mei Lu. Conditional diagnosability of the SP_{n} graphs under the comparison diagnosis model. Applied Mathematics and Computation, 336(??):249–256, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304120.

Greenfield:2017:NMT

[GLA^{+}17] J. R. F. Greenfield, P. Lestriez, C. Arand, D. Gruszka, T. Nowak, P. M. Rommens, and R. Taiar. A numerical model of the tension band wiring technique for olecranon fracture reduction. Applied Mathematics and Computation, 297(??):31–38, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306142.

Glaubitz:2019:SCB

[Gla19] Jan Glaubitz. Shock capturing by Bernstein polynomials for scalar conservation laws. Applied Mathematics and Computation, 363(??): Article 124593, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305855.

Gao:2019:RPG

[GlH^{+}19] Bo Gao, Xuan liu, Shuxia Hou, Danyang Jia, and Mingjing Du. Resolving public goods dilemma by giving the poor more support. Applied Mathematics and Computation, 362 (??):Article 124529, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305089.

Gao:2015:GOW

[GLL15a] Jianwei Gao, Ming Li, and Huihui Liu. Generalized ordered weighted utility proportional averaginghyperbolic absolute risk aversion operators and their applications to group decision-making. Applied Mathematics and Computation, 252(??):114–132, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016634.

Guo:2015:USA

[GLL15b] Geyang Guo, Shujuan Lü, and Bo Liu. Unconditional stability of alternating difference schemes with variable time steplengths for dispersive equation. Applied Mathematics and Computation, 262(??): 249–259, July 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315004956.

Greenhalgh:2016:SSE

[GLM16] D. Greenhalgh, Y. Liang, and X. Mao. SDE SIS epidemic model with demographic stochasticity and varying population size. Applied Mathematics and Computation, 276(??):218–238, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315016161.

Gao:2019:SCS

[GLM19] Jianfang Gao, Hui Liang, and Shufang Ma. Strong convergence of the semiimplicit Euler method for nonlinear stochastic Volterra integral equations with constant delay. Applied Mathematics and Computation, 348(??):385–398, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308919.

Gouveia:2015:LCB

[GLN15] Márcio R. A. Gouveia, Jaume Llibre, and Douglas D. Novaes. On limit cycles bifurcating from the infinity in discontinuous piecewise linear differential systems. Applied Mathematics and Computation, 271(??):365–374, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012400.

Gu:2018:NJN

[GLQ^{+}18] Ran Gu, Xueliang Li, Zhongmei Qin, Yongtang Shi, and Kang Yang. Non-jumping numbers for 5-uniform hypergraphs. Applied Mathematics and Computation, 317(??):234–251, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305659.

Gamez:2018:GTM

[GLR^{+}18] Manuel Gámez, Inmaculada López, Carmelo Rodríguez, Zoltán Varga, and József Garay. Gametheoretical model for marketing cooperative in fisheries. Applied Mathematics and Computation, 329(??):325–338, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830119X.

Guterman:2016:DRS

[GLS16] Alexander Guterman, Rute Lemos, and Graça Soares. On the C-determinantal range for special classes of matrices. Applied Mathematics and Computation, 275(??):86–94, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015404.

Guterman:2019:MGK

[GLS19] Alexander Guterman, Rute Lemos, and Graça Soares. More on geometry of Krein space C-numerical range. Applied Mathematics and Computation, 352(??):258–269, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300384.

Guo:2015:VTC

[GLW15] Zhengguang Guo, Shengrong Liu, and Weiming Wang. On a variation of the two-component Hunter–Saxton system. Applied Mathematics and Computation, 259(??):45– 52, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001770.

Gospodarczyk:2017:DRC

[GLW17] Przemysław Gospodarczyk, Stanisław Lewanowicz, and Paweł Woźny. Degree reduction of composite Bézier curves. Applied Mathematics and Computation, 293(??):40–48, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304994.

Gao:2018:TLB

[GLW18] Hui Gao, Benjian Lv, and Kaishun Wang. Two lower bounds for generalized 3connectivity of Cartesian product graphs. Applied Mathematics and Computation, 338(??):290–304, December 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318303199.

Gong:2016:DOR

[GLYT16] Zhaohua Gong, Ryan Loxton, Changjun Yu, and Kok Lay Teo. Dynamic optimization for robust path planning of horizontal oil wells. Applied Mathematics and Computation, 274(??):711–725, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015283.

Gao:2015:MGF

[GM15a] Yali Gao and Liquan Mei. Mixed Galerkin finite element methods for modified regularized long wave equation. Applied Mathematics and Computation, 258(??):267–281, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001654.

Gupta:2015:AGS

[GM15b] Vijay Gupta and Neha Malik. Approximation for genuine summationintegral type link operators. Applied Mathematics and Computation, 260(??):321–330, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003914.

Ge:2017:MDG

[GM17a] Zhihao Ge and Mengxia Ma. Multiphysics discontinuous Galerkin method for a poroelasticity model. Applied Mathematics and Computation, 301(??):78– 94, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630738X.

Golbabai:2017:NME

[GM17b] Ahmad Golbabai and Ehsan Mohebianfar. A new method for evaluating options based on multiquadric RBF-FD method. Applied Mathematics and Computation, 308(??): 130–141, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317301935.

Gupta:2019:SRL

[GM19] Nishi Gupta and Md. Maqbul. Solutions to Rayleigh-love equation with constant coefficients and delay forcing term. Applied Mathematics and Computation, 355(??): 123–134, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319301638.

Gao:2018:GMD

[GML18] Yali Gao, Liquan Mei, and Rui Li. Galerkin methods for the Davey–Stewartson equations. Applied Mathematics and Computation, 328(??):144–161, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300675.

Garcia-Martinez:2015:HEB

[GMOGCCC15] M. García-Martínez, L. J.

Ontañón-García, E. Campos-Cantón, and S. Celikovský. Hyperchaotic encryption based on multiscroll piecewise linear systems. Applied Mathematics and Computation, 270(??):413–424, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010929.

Gasinski:2018:OCD

[GMOP18] Leszek Gasiński, Stanisław Migórski, Anna Ochal, and Zijia Peng. Optimal control for doubly nonlinear evolutionary inclusions. Applied Mathematics and Computation, 321(??):244–254, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307403.

Grabowski:2018:TIT

[GMPP18] Dariusz Grabowski, Marcin Maciazek, Marian Pasko, and Anna Piwowar. Timeinvariant and time-varying filters versus neural approach applied to DC component estimation in control algorithms of active power filters. Applied Mathematics and Computation, 319(??):203–217, February 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317301339.

Gao:2019:SSU

[GMS19] Rong Gao, Nana Ma, and Gaoji Sun. Stability of solution for uncertain wave equation. Applied Mathematics and Computation, 356(??):469–478, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301882.

Golbabai:2015:SCR

[GN15] Ahmad Golbabai and Ahmad Nikpour. Stability and convergence of radial basis function finite difference method for the numerical solution of the reaction–diffusion equations. Applied Mathematics and Computation, 271(??):567–580, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012709.

Gao:2019:TCA

[GN19a] You Gao and Min-Yao Niu. Two constructions of asymptotically optimal codebooks according to the Welch bound. Applied Mathematics and Computation, 348(??):167–174, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310361.

Gowrisankar:2019:ERN

[GN19b] S. Gowrisankar and Srinivasan Natesan. An efficient robust numerical method for singularly perturbed Burgers’ equation. Applied Mathematics and Computation, 346(??):385–394, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309159.

Ginoux:2019:CPP

[GNR^{+}19] Jean-Marc Ginoux, Roomila Naeck, Yusra Bibi Ruhomally, Muhammad Zaid Dauhoo, and Matjaz Perc. Chaos in a predator– prey-based mathematical model for illicit drug consumption. Applied Mathematics and Computation, 347(??):502–513, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830969X.

Gerisch:2015:DCA

[GNW15] A. Gerisch, A. Naumann, and J. Wensch. Defect corrected averaging for highly oscillatory problems. Applied Mathematics and Computation, 261(??):90–103, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003951.

Gal:2015:AAO

[GO15a] Sorin G. Gal and Bogdan D. Opris. Approximation with an arbitrary order by modified Baskakov type operators. Applied Mathematics and Computation, 265(??):329–332, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006517.

Gani:2015:IPT

[GO15b] M. Osman Gani and Toshiyuki Ogawa. Instability of periodic traveling wave solutions in a modified FitzHugh– Nagumo model for excitable media. Applied Mathematics and Computation, 256(??):968–984, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001411.

Gregorio:2019:TPL

[GOA19] R. M. Gregório, P. R.

Oliveira, and C. D. S. Alves. A two-phase-like proximal point algorithm in domains of positivity. Applied Mathematics and Computation, 343(??):67–89, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308336.

Garcia:2018:SAD

[GOH18] Amanda Garcia, Amer Obeidi, and Keith W. Hipel. Strategic advice for decision-making under conflict based on observed behaviour. Applied Mathematics and Computation, 332(??):96–104, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302042.

Golafshani:2015:IBB

[Gol15] Emadaldin Mohammadi Golafshani. Introduction of Biogeography-Based Programming as a new algorithm for solving problems. Applied Mathematics and Computation, 270(??): 1–12, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315010814.

Golinski:2019:FMT

[Gol19] Tomasz Goliński. Factorization method on time scales. Applied Mathematics and Computation, 347(??):354–359, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309809.

Goncalves:2018:PIC

[Gon18] M. L. N. Gonçalves. On the pointwise iterationcomplexity of a dynamic regularized ADMM with over-relaxation stepsize. Applied Mathematics and Computation, 336(??):315–325, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303928.

Goubko:2018:MWI

[Gou18] Mikhail Goubko. Maximizing Wiener index for trees with given vertex weight and degree sequences. Applied Mathematics and Computation, 316(??):102–114, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305386.

Gonzalez-Pinto:2016:FTS

[GPHAPRW16] S. González-Pinto, D. Hernández-Abreu, S. Pérez-Rodríguez, and R. Weiner. A family of three-stage third order AMF-w-methods for the time integration of advection diffusion reaction PDEs. Applied Mathematics and Computation, 274(??):565–584, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013454.

Gil-Pons:2019:DBV

[GPRCTRY19] Reynaldo Gil-Pons, Yunior Ramírez-Cruz, Rolando Trujillo-Rasua, and Ismael G. Yero. Distancebased vertex identification in graphs: the outer multiset dimension. Applied Mathematics and Computation, 363(??): Article 124612, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306046.

Geng:2015:IRK

[GQC15] F. Z. Geng, S. P. Qian, and M. G. Cui. Improved reproducing kernel method for singularly perturbed differential-difference equations with boundary layer behavior. Applied Mathematics and Computation, 252(??):58–63, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314016397.

Gong:2015:OON

[GR15a] Benxue Gong and Hongxing Rui. One order numerical scheme for forwardbackward stochastic differential equations. Applied Mathematics and Computation, 271(??):220–231, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315012084.

Gupta:2015:IHW

[GR15b] A. K. Gupta and S. Saha Ray. An investigation with Hermite Wavelets for accurate solution of Fractional Jaulent–Miodek equation associated with energydependent Schrödinger potential. Applied Mathematics and Computation, 270(??):458–471, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011133.

Gupta:2015:DEC

[GR15c] Vijay Gupta and Themistocles M. Rassias. Direct estimates for certain Szász type operators. Applied Mathematics and Computation, 251(??):469–474, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016105.

Gala:2017:NWS

[GR17a] Saddek Gala and Maria Alessandra Ragusa. Note on the weak-strong uniqueness criterion for the β-QG in Morrey–Campanato space. Applied Mathematics and Computation, 293(??):65–71, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305069.

Gupta:2017:SWS

[GR17b] A. K. Gupta and S. Saha Ray. On the solitary wave solution of fractional Kudryashov– Sinelshchikov equation describing nonlinear wave processes in a liquid containing gas bubbles. Applied Mathematics and Computation, 298(??):1– 12, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300316306610.

Goura:2019:EBS

[GR19a] V. M. K. Prasad Goura and Pradip Roul. Erratum to: B-spline collocation methods and their convergence for a class of nonlinear derivative dependent singular boundary value problems. Applied Mathematics and Computation, 361(??):198–201, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304151.

Goyal:2019:TLO

[GR19b] Pawan Goyal and Martin Redmann. Timelimited H_{2}-optimal model order reduction. Applied Mathematics and Computation, 355(??):184–197, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301754.

Grace:2015:OBS

[Gra15] Said R. Grace. On the oscillatory behavior of solutions of nonlinear fractional differential equations. Applied Mathematics and Computation, 266(??):259–266, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006864.

Graillat:2018:AAE

[Gra18] Stef Graillat. An accurate algorithm for evaluating rational functions. Applied Mathematics and Computation, 337(??):487–493, November 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031830448X.

Greenhalgh:2015:APC

[GRS^{+}15] David Greenhalgh, Sourav Rana, Sudip Samanta, Tridip Sardar, Sabyasachi Bhattacharya, and Joydev Chattopadhyay. Awareness programs control infectious disease — multiple delay induced mathematical model. Applied Mathematics and Computation, 251(??):539–563, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016233.

Gabor:2015:WTT

[GRV15] Grzegorz Gabor, Sebastian Ruszkowski, and Jirí Vítovec. Ważewski type theorem for nonautonomous systems of equations with a disconnected set of egress points. Applied Mathematics and Computation, 265(??):358–369, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500644X.

Gurudwan:2011:CTC

[GS11] Niyati Gurudwan and B. K. Sharma. Convergence theorem for the common solution for a finite family of ϕ-strongly accretive operator equations. Applied Mathematics and Computation, 217(15):6748–6754, April 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300311001329. See

note [ZS15d].

Gao:2015:OAN

[GS15a] Jianfang Gao and Fuyi Song. Oscillation analysis of numerical solutions for nonlinear delay differential equations of hematopoiesis with unimodal production rate. Applied Mathematics and Computation, 264(??):72–84, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500524X.

Gou:2015:SNS

[GS15b] Tian-Xiang Gou and Hong-Rui Sun. Solutions of nonlinear Schrödinger equation with fractional Laplacian without the Ambrosetti–Rabinowitz condition. Applied Mathematics and Computation, 257(??):409–416, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314012600.

Ghorbani:2016:AGP

[GS16] Modjtaba Ghorbani and Mahin Songhori. On the automorphism group of polyhedral graphs. Applied Mathematics and Computation, 282(??):237–243, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300316300753.

Ganesan:2017:ASF

[GS17a] Sashikumaar Ganesan and Shweta Srivastava. ALE-SUPG finite element method for convection–diffusion problems in time-dependent domains: Conservative form. Applied Mathematics and Computation, 303(??):128–145, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300401.

George:2017:CRR

[GS17b] Santhosh George and M. Sabari. Convergence rate results for steepest descent type method for nonlinear ill-posed equations. Applied Mathematics and Computation, 294(??):169–179, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305690.

Guo:2017:GWP

[GS17c] Yana Guo and Haifeng Shang. Global wellposedness of two-dimensional magneto-micropolar equations with partial dissipation. Applied Mathematics and Computation, 313(??):392–407, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304290.

Ghorbani:2018:PGT

[GS18a] Modjtaba Ghorbani and Mahin Songhori. Polyhedral graphs via their automorphism groups. Applied Mathematics and Computation, 321(??):1–10, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307270.

Gruzdeva:2018:SSR

[GS18b] Tatiana V. Gruzdeva and Alexander S. Strekalovsky. On solving the sum-of-ratios problem. Applied Mathematics and Computation, 318(??):260–269, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305350.

Gorgel:2019:FRD

[GS19] Pelin Görgel and Ahmet Simsek. Face recognition via Deep Stacked Denoising Sparse Autoencoders (DSDSA). Applied Mathematics and Computation, 355(??):325–342, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930181X.

Guinaldo:2016:DAC

[GSD16] M. Guinaldo, J. Sánchez, and S. Dormido. Distributed adaptive control of linear multi-agent systems with event-triggered communications. Applied Mathematics and Computation, 274(??):195–207, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014411.

Guo:2017:EPE

[GSD^{+}17] Hao Guo, Chen Shen, Dameng Dai, Mi Zhang, Chen Chu, and Lei Shi. Environment promotes the evolution of cooperation in spatial voluntary prisoner’s dilemma game. Applied Mathematics and Computation, 315(??):47–53, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305052.

Garoosi:2015:ELA

[GSDH15] F. Garoosi, M. R. Safaei, M. Dahari, and K. Hooman. Eulerian–Lagrangian analysis of solid particle distribution in an internally heated and cooled airfilled cavity. Applied Mathematics and Computation, 250(??):28–46, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031401443X.

Guo:2018:FTL

[GSLC18] Litao Guo, Guifu Su, Wenshui Lin, and Jinsong Chen. Fault tolerance of locally twisted cubes. Applied Mathematics and Computation, 334(??):401–406, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302881.

Ghazanfarian:2015:ISP

[GSP15] J. Ghazanfarian, R. Saghatchi, and D. V. Patil. Implementation of smoothedparticle hydrodynamics for non-linear Pennes’ bioheat transfer equation. Applied Mathematics and Computation, 259(??):21–31, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002167.

Ge:2019:RHS

[GSPH19] Chao Ge, Yanpen Shi, Ju H. Park, and Changchun Hua. Robust H_{∞} stabilization for T-S fuzzy systems with time-varying delays and memory sampleddata control. Applied Mathematics and Computation, 346(??):500–512, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309500.

Gil:2015:CKF

[GST15a] Amparo Gil, Javier Segura, and Nico M. Temme. Computing the Kummer function U(a, b, z) for small values of the arguments. Applied Mathematics and Computation, 271(??):532–539, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315012837.

Guo:2015:ECN

[GST15b] Yong-Feng Guo, Ya-Jun Shen, and Jian-Guo Tan. Effects of colored noise and external periodic force on the time derivative of information entropy for a damped harmonic oscillator. Applied Mathematics and Computation, 252(??):20–26, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016385.

Gil:2019:CIC

[GST19] Amparo Gil, Javier Segura, and Nico M. Temme. On the computation and inversion of the cumulative noncentral beta distribution function. Applied Mathematics and Computation, 361(??):74–86, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304072.

Gregor:2015:RLD

[GSV15] Petr Gregor, Riste Skrekovski, and Vida Vukasinović. Rooted level-disjoint partitions of Cartesian products. Applied Mathematics and Computation, 266(??):244–258, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006839.

Gentile:2017:PAQ

[GSV17] M. Gentile, A. Sommariva, and M. Vianello. Polynomial approximation and quadrature on geographic rectangles. Applied Mathematics and Computation, 297(??):159–179, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305094.

Gregor:2018:MSB

[GSV18a] Petr Gregor, Riste Skrekovski, and Vida Vukasinović. Modelling simultaneous broadcasting by leveldisjoint partitions. Applied Mathematics and Computation, 325(??):15–23, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308779.

Gyllenberg:2018:EID

[GSV18b] Mats Gyllenberg, Francesca Scarabel, and Rossana Vermiglio. Equations with infinite delay: Numerical bifurcation analysis via pseudospectral discretization. Applied Mathematics and Computation, 333(??):490–505, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302856.

Gyurkovics:2017:SAL

[GSVK17] É. Gyurkovics, G. Szabó-Varga, and K. Kiss. Stability analysis of linear systems with interval timevarying delays utilizing multiple integral inequalities. Applied Mathematics and Computation, 311(??):164–177, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302928.

Gao:2015:SCN

[GSZ15] Ming Gao, Li Sheng, and Weihai Zhang. Stochastic H_{2}/H_{∞} control of nonlinear systems with timedelay and state-dependent noise. Applied Mathematics and Computation, 266(??):429–440, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007146.

Gaboury:2015:GMI

[GT15a] Sebastien Gaboury and Richard Tremblay. Generalized multiple integral representations for a large family of polynomials with applications. Applied Mathematics and Computation, 261(??):39–47, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004026.

Garay:2015:ECW

[GT15b] Óscar J. Garay and Yoshihiko Tazawa. Extremal curves for weighted elastic energy in surfaces of 3-space forms. Applied Mathematics and Computation, 251(??):349–362, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015914.

Gu:2015:EMS

[GT15c] Haibo Gu and Juan J. Trujillo. Existence of mild solution for evolution equation with Hilfer fractional derivative. Applied Mathematics and Computation, 257(??):344–354, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014581.

Golmankhaneh:2019:STF

[GT19] Alireza K. Golmankhaneh and Cemil Tunç. Sumudu transform in fractal calculus. Applied Mathematics and Computation, 350(??):386–401, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300347.

Gomez-Toston:2018:COP

[GTBC18] Carolina Gómez-Tostón, Manuel Barrena, and Álvaro Cortés. Characterizing the optimal pivots for efficient similarity searches in vector space databases with Minkowski distances. Applied Mathematics and Computation, 328(??):203–223, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300493.

Guan:2019:MAL

[Gua19] Jinrui Guan. Modified alternately linearized implicit iteration method for M-matrix algebraic Riccati equations. Applied Mathematics and Computation, 347(??):442–448, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309998.

Guo:2016:CSB

[Guo16] Xue-Ping Guo. Convergence studies on block iterative algorithms for image reconstruction. Applied Mathematics and Computation, 273(??):525–534, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013703.

Gutman:2016:SDDa

[Gut16]  Ivan Gutman.  On

Steiner degree distance of trees. Applied Mathematics and Computation, 283(??):163–167, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301497.

Gutman:2017:CGT

[Gut17]  Ivan Gutman.  On

coindices of graphs and their complements. Applied Mathematics and Computation, 305(??): 161–165, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317301054.

Gutman:2018:SIG

[Gut18] Ivan Gutman. Stepwise irregular graphs. Applied Mathematics and Computation, 325(??):234–238, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300043.

Grammont:2016:MIP

[GVA16] Laurence Grammont, Paulo B. Vasconcelos, and Mario Ahues. A modified iterated projection method adapted to a nonlinear integral equation. Applied Mathematics and Computation, 276(??):432–441, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300072.

Gao:2015:ISS

[GW15a] Shang Gao and Boying Wu. On input-to-state stability for stochastic coupled control systems on networks. Applied Mathematics and Computation, 262(??):90–101, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004580.

Gao:2015:STD

[GW15b] Wenwu Gao and Zongmin Wu. Solving timedependent differential equations by multiquadric trigonometric quasi-interpolation. Applied Mathematics and Computation, 253(??): 377–386, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314016622.

Gao:2015:NBR

[GW15c] Yuelin Gao and Fei Wei. A new bound-and-reduce approach of nonconvex quadratic programming problems. Applied Mathematics and Computation, 250(??):298–308, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014520.

Galewski:2016:EMR

[GW16] Marek Galewski and Renata Wieteska. Existence and multiplicity results for boundary value problems connected with the discrete p(·)- Laplacian on weighted finite graphs. Applied Mathematics and Computation, 290(??):376–391, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303873.

Guan:2019:TPR

[GW19] Junbiao Guan and Kaihua Wang. Towards pedestrian room evacuation with a spatial game. Applied Mathematics and Computation, 347(??):492–501, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318309767.

Gao:2016:ACO

[GWC^{+}16] Shangce Gao, Yirui Wang, Jiujun Cheng, Yasuhiro Inazumi, and Zheng Tang. Ant colony optimization with clustering for solving the dynamic location routing problem. Applied Mathematics and Computation, 285(??):149–173, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302363.

Govender:2015:CDC

[GWK15] Nicolin Govender, Daniel N.

Wilke, and Schalk Kok. Collision detection of convex polyhedra on the NVIDIA GPU architecture for the discrete element method. Applied Mathematics and Computation, 267(??):810–829, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314013794.

Gassner:2016:WBE

[GWK16] Gregor J. Gassner, Andrew R. Winters, and David A. Kopriva. A well balanced and entropy conservative discontinuous Galerkin spectral element method for the shallow water equations. Applied Mathematics and Computation, 272 (part 2)(??):291–308, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009261.

Govender:2018:SSN

[GWPA18] Nicolin Govender, Daniel N. Wilke, Patrick Pizette, and Nor-Edine Abriak. A study of shape nonuniformity and poly-dispersity in hopper discharge of spherical and polyhedral particle systems using the Blaze-DEM GPU code. Applied Mathematics and Computation, 319(??):318–336, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302217.

Gao:2016:DRN

[GWSC16] Fang Gao, Min Wu, Jinhua She, and Weihua Cao. Disturbance rejection in nonlinear systems based on equivalent-input-disturbance approach. Applied Mathematics and Computation, 282(??): 244–253, May 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316301175.

Gao:2015:FRE

[GWW15] Lijun Gao, Dandan Wang, and Gang Wang. Further results on exponential stability for impulsive switched nonlinear timedelay systems with delayed impulse effects. Applied Mathematics and Computation, 268(??):186–200, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008048.

Ge:2017:ESC

[GWWL17] Chao Ge, Bingfang Wang, Xian Wei, and Yajuan Liu. Exponential synchronization of a class of neural networks with sampleddata control. Applied Mathematics and Computation, 315(??):150–161, December 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317305064.

Guo:2017:ROG

[GWX17] Lifeng Guo, Ligong Wang, and Peng Xiao. 5-regular oriented graphs with optimum skew energy. Applied Mathematics and Computation, 301(??):43–59, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307408.

Guo:2018:GTA

[GWXZ18] Beibei Guo, Yinhu Wu, Yu Xiao, and Chiping Zhang. Graph-theoretic approach to synchronizing stochastic coupled systems with time-varying delays on networks via periodically intermittent control. Applied Mathematics and Computation, 331(??):341–357, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301851.

Guo:2015:IPI

[GWY15] Yongxia Guo, Guangsheng Wei, and Ruoxia Yao. Inverse problem for interior spectral data of discontinuous Dirac operator. Applied Mathematics and Computation, 268(??):775–782, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008723.

Guan:2019:MDS

[GXCY19] Xiaxia Guan, Lina Xue, Eddie Cheng, and Weihua Yang. Minimum degree and size conditions for the proper connection number of graphs. Applied Mathematics and Computation, 352(??):205–210, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300797.

Geng:2018:DDC

[GXH18] Yan Geng, Jinhu Xu, and Jiangyong Hou. Discretization and dynamic consistency of a delayed and diffusive viral infection model. Applied Mathematics and Computation, 316(??):282–295, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305921.

Gao:2019:NSR

[GXLL19] Rong Gao, Juanjuan Xu, Wuquan Li, and Xiaohua Liu. A necessary and sufficient RHC stabilizability condition for stochastic control with delayed input. Applied Mathematics and Computation, 360(??):122–130, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319303741.

Gan:2015:SFV

[GY15] Xiao-Ting Gan and Jun-Feng Yin. Symmetric finite volume method for second order variable coefficient hyperbolic equations. Applied Mathematics and Computation, 268(??):1015–1028, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009224.

Guo:2019:BAC

[GYC19] Laigang Guo, Pei Yu, and Yufu Chen. Bifurcation analysis on a class of threedimensional quadratic systems with twelve limit cycles. Applied Mathematics and Computation, 363 (??):Article 124577, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305697.

Guo:2015:RDP

[GYW15] Fei Guo, Li Yan, and Run Wang. Rigorous derivation and propagation speed property for a twocomponent Degasperis– Procesi system in shallow water regimes. Applied Mathematics and Computation, 259(??):980–986, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002970.

Gao:2015:NFF

[GYY15] Yuelin Gao, Yongjian Yang, and Mi You. A new filled function method for global optimization. Applied Mathematics and Computation, 268(??):685–695, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008711.

Geng:2018:SPR

[GYY18] Zhichao Geng, Jinjiang Yuan, and Junling Yuan. Scheduling with or without precedence relations on a serial-batch machine to minimize makespan and maximum cost. Applied Mathematics and Computation, 332(??):1–18, September 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318301668.

Guo:2015:HAG

[GYZC15] Xixiong Guo, Jieke Yao, Chengwen Zhong, and Jun Cao. A hybrid adaptivegridding immersed-boundary lattice Boltzmann method for viscous flow simulations. Applied Mathematics and Computation, 267(??):529–553, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001149.

Guo:2015:ZSO

[GZ15a] Dongsheng Guo and Yunong Zhang. ZNN for solving online time-varying linear matrix-vector inequality via equality conversion. Applied Mathematics and Computation, 259(??):327–338, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500260X.

Guo:2015:IFP

[GZ15b] Tian Liang Guo and Kan-Jian Zhang. Impulsive fractional partial differential equations. Applied Mathematics and Computation, 257(??):581–590, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314008017.

Gillard:2018:OED

[GZ18] Jonathan Gillard and Anatoly Zhigljavsky. Optimal estimation of direction in regression models with large number of parameters. Applied Mathematics and Computation, 318(??):281–289, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303624.

Gao:2019:DSO

[GZC19] Rui Gao, Ding Zhai, and Jun Cheng. Decentralized static output feedback sliding mode control for interconnected descriptor systems via linear sliding variable. Applied Mathematics and Computation, 357(??):185–198, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302851.

Guo:2019:ISS

[GZD19] Ying Guo, Wei Zhao, and Xiaohua Ding. Input-to-state stability for stochastic multi-group models with multi-dispersal and time-varying delay. Applied Mathematics and Computation, 343(??):114–127, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306246.

Gori:2015:PSF

[GZG^{+}15] G. Gori, M. Zocca, M. Garabelli, A. Guardone, and G. Quaranta. PoliM-Ice: a simulation framework for three-dimensional ice accretion. Applied Mathematics and Computation, 267(??):96–107, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315007055.

Guan:2019:ALG

[GZhL^{+}19] Xiaxia Guan, Shurong Zhang, Rong hua Li, Lin Chen, and Weihua Yang. Anti-k-labeling of graphs. Applied Mathematics and Computation, 363(??): Article 124549, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305326.

Ge:2016:ACS

[GZK16] Fu-Dong Ge, Hua-Cheng Zhou, and Chun-Hai Kou. Approximate controllability of semilinear evolution equations of fractional order with nonlocal and impulsive conditions via an approximating technique. Applied Mathematics and Computation, 275(??):107–120, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015581.

Guo:2017:EUE

[GZLL17] Runan Guo, Ziye Zhang, Xiaoping Liu, and Chong Lin. Existence, uniqueness, and exponential stability analysis for complexvalued memristor-based BAM neural networks with time delays. Applied Mathematics and Computation, 311(??):100–117, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317303090.

Guo:2015:EAS

[GZW15] Hui Guo, Qinghua Zhang, and Jichao Wang. Error analysis of the semidiscrete local discontinuous Galerkin method for compressible miscible displacement problem in porous media. Applied Mathematics and Computation, 259(??):88–105, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001228.

Gao:2017:CNM

[GZZL17] Qin Gao, Quanting Zhao, Xuan Zheng, and Yonghui Ling. Convergence of Numerov’s method for inverse Sturm–Liouville problems. Applied Mathematics and Computation, 293(??):1–17, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305021.

Hamid:2016:ISG

[HA16] Haydar Abdel Hamid and Waad Al Sayed. Integrable solutions of a generalized mixed-type functional integral equation. Applied Mathematics and Computation, 276(??):356–366, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300126.

Hassani:2019:TBS

[HA19] Hossein Hassani and Zakieh Avazzadeh. Transcendental Bernstein series for solving nonlinear variable order fractional optimal control problems. Applied Mathematics and Computation, 362 (??):Article 124563, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300319305466.

Haciefendioglu:2017:SDR

[Hac17]  Kemal  Haciefendioglu.

Stochastic dynamic response of short-span highway bridges to spatial variation of blasting ground vibration. Applied Mathematics and Computation, 292(??):194–209, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304775.

Hajiketabi:2018:LGM

[HAC18] M. Hajiketabi, S. Abbasbandy, and F. Casas. The Lie-group method based on radial basis functions for solving nonlinear high dimensional generalized Benjamin–Bona–Mahony– Burgers equation in arbitrary domains. Applied Mathematics and Computation, 321(??):223–243, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307609.

Hafezalkotob:2017:CDM

[Haf17] Ashkan Hafezalkotob. Competition of domestic manufacturer and foreign supplier under sustainable development objectives of government. Applied Mathematics and Computation, 292(??):294–308, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304374.

Haghani:2015:GSM

[Hag15] F. Khaksar Haghani. A generalized Steffensen’s method for matrix sign function. Applied Mathematics and Computation, 260(??):249–256, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003653.

Heydari:2019:WAS

[HAH19] Mohammad Hossein Heydari, Zakieh Avazzadeh, and Malih Farzi Haromi. A wavelet approach for solving multi-term variableorder time fractional diffusionwave equation. Applied Mathematics and Computation, 341(??):215–228, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300318307483.

Hayat:2014:BLF

[HAMA14] T. Hayat, Sadia Asad, M. Mustafa, and A. Alsaedi. Boundary layer flow of Carreau fluid over a convectively heated stretching sheet. Applied Mathematics and Computation, 246(??):12–22, November 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314010406. See

comment [PK18a].

Han:2015:NBS

[Han15a] Xuli Han. Normalized B-basis of the space of trigonometric polynomials and curve design. Applied Mathematics and Computation, 251(??):336–348, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016026. See

comment [SR15d].

Han:2015:PTH

[Han15b]  Xuli Han.  Piecewise

trigonometric Hermite interpolation. Applied Mathematics and Computation, 268(??):616–627, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009091.

Han:2015:SPP

[Han15c] Xuli Han. Shape-preserving piecewise rational interpolant with quartic numerator and quadratic denominator. Applied Mathematics and Computation, 251(??):258–274, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015999.

Han:2016:TBT

[Han16] Chang Yong Han. Tractrixbased tool orientation control for 5-axis CNC machining. Applied Mathematics and Computation, 272 (part 1)(??):92–99, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500939X.

Han:2018:SPP

[Han18] Xuli Han. Shape-preserving piecewise rational interpolation with higher order continuity. Applied Mathematics and Computation, 337(??):ii, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304223.

Han:2019:ECD

[Han19] Sang-Eon Han. Estimation of the complexity of a digital image from the viewpoint of fixed point theory. Applied Mathematics and Computation, 347(??):236–248, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830941X.

Harase:2015:QMC

[Har15]  Shin Harase.  Quasi-

Monte Carlo point sets with small t-values and WAFOM. Applied Mathematics and Computation, 254(??):318–326, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000028.

Harfash:2016:RPC

[Har16] Akil J. Harfash. Resonant penetrative convection in porous media with an internal heat source/sink effect. Applied Mathematics and Computation, 281(??):323–342, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300066.

Hassan:2015:EES

[HARS15] M. M. Hassan, M. A.

Abdel-Razek, and A. A.- H. Shoreh. Explicit exact solutions of some nonlinear evolution equations with their geometric interpretations. Applied Mathematics and Computation, 251(??):243–252, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015781.

He:2015:SNI

[HASA15] Yuan He, Serkan Araci, H. M. Srivastava, and Mehmet Acikgoz. Some new identities for the Apostol–Bernoulli polynomials and the Apostol– Genocchi polynomials. Applied Mathematics and Computation, 262(??):31– 41, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004518.

Heidarkhani:2019:DFO

[HASC19] Shapour Heidarkhani, Ghasem A. Afrouzi, Amjad Salari, and Giuseppe Caristi. Discrete fourth-order boundary value problems with four parameters. Applied Mathematics and Computation, 346(??):167–182, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309056.

Heydari:2019:CMS

[HAY19] Mohammad Hossein Heydari, Zakieh Avazzadeh, and Yin Yang. A computational method for solving variable-order fractional nonlinear diffusionwave equation. Applied Mathematics and Computation, 352(??):235–248, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031930092X.

Hazarika:2016:LIC

[Haz16] Bipan Hazarika. Lacunary ideal convergence of multiple sequences in probabilistic normed spaces. Applied Mathematics and Computation, 279(??):139–153, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300369.

Huang:2018:EBS

[HC18] Zaitang Huang and Junfei Cao. Ergodicity and bifurcations for stochastic logistic equation with non-Gaussian Lévy noise. Applied Mathematics and Computation, 330(??):1– 10, August 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318300778.

Han:2016:SBC

[HCH16] Xiaozhuo Han, Baoying Chen, and Cang Hui. Symmetry breaking in cyclic competition by niche construction. Applied Mathematics and Computation, 284(??):66– 78, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301758.

He:2019:CGB

[HCS19] Nanrong He, Xiaojie Chen, and Attila Szolnoki. Central governance based on monitoring and reporting solves the collective-risk social dilemma. Applied Mathematics and Computation, 347(??):334–341, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310038.

Chen:2015:IIR

[hCWW^{+}15] Mei huan Chen, Li Wang, Juan Wang, Shi wen Sun, and Cheng yi Xia. Impact of individual response strategy on the spatial public goods game within mobile agents. Applied Mathematics and Computation, 251(??):192–202, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015847.

Huang:2017:CBD

[HCX^{+}17a] Chengdai Huang, Jinde Cao, Min Xiao, Ahmed Alsaedi, and Fuad E. Alsaadi. Controlling bifurcation in a delayed fractional predator–prey system with incommensurate orders. Applied Mathematics and Computation, 293(??):293–310, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305288.

Huang:2017:BDF

[HCX^{+}17b] Chengdai Huang, Jinde Cao, Min Xiao, Ahmed Alsaedi, and Tasawar Hayat. Bifurcations in a delayed fractional complex-valued neural network. Applied Mathematics and Computation, 292(??):210–227, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304672.

Huang:2015:MAR

[HCY15] Jizu Huang, Liqun Cao, and Chao Yang. A multiscale algorithm for radiative heat transfer equation with rapidly oscillating coefficients. Applied Mathematics and Computation, 266(??):149–168, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315006657.

Huang:2018:HCB

[HCY^{+}18] Keke Huang, Xiaofang Chen, Zhaofei Yu, Chunhua Yang, and Weihua Gui. Heterogeneous cooperative belief for social dilemma in multiagent system. Applied Mathematics and Computation, 320(??):572–579, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307178.

Hua:2016:WPI

[HD16a] Hongbo Hua and Kinkar Ch. Das. On the Wiener polarity index of graphs. Applied Mathematics and Computation, 280(??): 162–167, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316300431.

Huang:2016:UPS

[HD16b] Xiaoxia Huang and Hao Di. Uncertain portfolio selection with background risk. Applied Mathematics and Computation, 276(??):284–296, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300060.

Han:2017:SDS

[HD17] Renji Han and Binxiang Dai. Spatiotemporal dynamics and spatial pattern in a diffusive intraguild predation model with delay effect. Applied Mathematics and Computation, 312(??):177–201, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730365X.

Huang:2015:HAU

[HDG15] Chao Huang, Chong Dai, and Miao Guo. A hybrid approach using two-level DEA for financial failure prediction and integrated SE-DEA and GCA for indicators selection. Applied Mathematics and Computation, 251(??):431–441, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016099.

Hammad:2015:OCFa

[HEA15a] D. A. Hammad and M. S.

El-Azab. A 2N order compact finite difference method for solving the generalized regularized long wave (GRLW) equation. Applied Mathematics and Computation, 253(??):248–261, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401724X.

Hammad:2015:OCFb

[HEA15b] D. A. Hammad and M. S. El-Azab. 2N order compact finite difference scheme with collocation method for solving the generalized Burger’s– Huxley and Burger’s– Fisher equations. Applied Mathematics and Computation, 258(??):296–311, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001629.

Hammad:2016:CCS

[HEA16] D. A. Hammad and M. S. El-Azab. Chebyshev– Chebyshev spectral collocation method for solving the generalized regularized long wave (GRLW) equation. Applied Mathematics and Computation, 285(??):228–240, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630234X.

Hermes:2019:CCA

[Her19] Danny Hermes. Compensated de Casteljau algorithm in K times the working precision. Applied Mathematics and Computation, 357(??):57–74, September 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319302541.

Hao:2016:FDS

[HFCS16] Zhaopeng Hao, Kai Fan, Wanrong Cao, and Zhizhong Sun. A finite difference scheme for semilinear space-fractional diffusion equations with time delay. Applied Mathematics and Computation, 275(??):238–254, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015738.

Hafeez:2019:BSD

[HFK19] Sumaira Hafeez, Rashid Farooq, and Mehtab Khan. Bicyclic signed digraphs with maximal energy. Applied Mathematics and Computation, 347(??):702–711, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310142.

Housiadas:2018:ASF

[HG18] Kostas D. Housiadas and Georgios C. Georgiou. Analytical solution of the flow of a Newtonian fluid with pressure-dependent viscosity in a rectangular duct. Applied Mathematics and Computation, 322(??):123–128, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308111.

Hadaeghi:2016:NAS

[HGG16] Fatemeh Hadaeghi, Seyed Mohammad Reza Hashemi Golpayegani, and Laura Gardini. A non-autonomous system leading to cyclic chaotic sets to model physiological rhythms. Applied Mathematics and Computation, 281(??):343–355, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300613.

Haq:2019:NSV

[HGH19a] Sirajul Haq, Abdul Ghafoor, and Manzoor Hussain. Numerical solutions of variable order time fractional (1 + 1)- and (1 + 2)-dimensional advection dispersion and diffusion models. Applied Mathematics and Computation, 360(??):107–121, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930387X.

Huang:2019:BFM

[HGH19b] Qian-Min Huang, Yi-Tian Gao, and Lei Hu. Bilinear forms, modulational instability and dark solitons for a fifth-order variable-coefficient nonlinear Schrödinger equation in an inhomogeneous optical fiber. Applied Mathematics and Computation, 352(??):270–278, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300360.

Hong:2019:LHC

[HGL19] Qi Hong, Yuezheng Gong, and Zhongquan Lv. Linear and Hamiltonianconserving Fourier pseudospectral schemes for the Camassa–Holm equation. Applied Mathematics and Computation, 346(??):86– 95, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309093.

Hu:2015:IPU

[HGM15a] Li-Ying Hu, Gong-De Guo, and Chang-Feng Ma. Image processing using Newton-based algorithm of nonnegative matrix factorization. Applied Mathematics and Computation, 269(??):956–964, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010899.

Hu:2015:LSA

[HGM15b] Li-Ying Hu, Gong-De Guo, and Chang-Feng Ma. The least squares antibisymmetric solution and the optimal approximation solution for Sylvester equation. Applied Mathematics and Computation, 259(??):212–219, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002441.

Herceg:2015:FMS

[HH15] Djordje Herceg and Dragoslav Herceg. A family of methods for solving nonlinear equations. Applied Mathematics and Computation, 259(??):882–895, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003331.

Halama:2016:NSS

[HH16] Jan Halama and Vladimír Hric. Numerical solution of steam flow in a nozzle using different nonequilibrium condensation models. Applied Mathematics and Computation, 272 (part 3)(??):657–669, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006918.

Huang:2017:AGC

[HH17] Xueyi Huang and Qiongxiang Huang. Automorphism group of the complete alternating group graph. Applied Mathematics and Computation, 314(??):58–64, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304605.

Haq:2018:SSP

[HH18a] Sirajul Haq and Manzoor Hussain. Selection of shape parameter in radial basis functions for solution of time-fractional Black– Scholes models. Applied Mathematics and Computation, 335(??):248–263, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303631.

Hu:2018:DDS

[HH18b] Peng Hu and Chengming Huang. Delay dependent stability of stochastic split-step θ methods for stochastic delay differential equations. Applied Mathematics and Computation, 339(??):651–662, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306301.

Han:2019:BBS

[HH19] Xuli Han and Jing Han.

Bicubic B-spline surfaces constrained by the biharmonic PDE. Applied Mathematics and Computation, 361(??):766–776, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319304904.

Hussain:2019:MSM

[HHG19] Manzoor Hussain, Sirajul Haq, and Abdul Ghafoor. Meshless spectral method for solution of time-fractional coupled KdV equations. Applied Mathematics and Computation, 341(??):321–334, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300318307732.

Heydari:2016:WMS

[HHGC16] M. H. Heydari, M. R.

Hooshmandasl, F. M. Maalek Ghaini, and C. Cattani. Wavelets method for solving fractional optimal control problems. Applied Mathematics and Computation, 286(??):139–154, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630251X.

Huang:2019:PCT

[HHL^{+}19] Changwei Huang, Wenchen Han, Haihong Li, Hongyan Cheng, Qionglin Dai, and Junzhong Yang. Public cooperation in two-layer networks with asymmetric interaction and learning environments. Applied Mathematics and Computation, 340(??):305–313, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307355.

Halama:2018:NST

[HHP18] Jan Halama, Vladimír Hric, and Marek Pátý. Numerical solution of transonic flow of steam with non-equilibrium phase change using typical and simplified method. Applied Mathematics and Computation, 319(??):499–509, February 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300317303454.

Helil:2017:NZS

[HHR17] Nurmamat Helil, Azhar Halik, and Kaysar Rahman. Non-zero-sum cooperative access control game model with user trust and permission risk. Applied Mathematics and Computation, 307(??):299–310, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301777.

Huang:2019:TCM

[HHT19] Guixian Huang, Weihua He, and Yuanyao Tan. Theoretical and computational methods to minimize Kirchhoff index of graphs with a given edge k-partiteness. Applied Mathematics and Computation, 341(??):348–357, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300318307744.

Huang:2019:STP

[HHY19] Haochuan Huang, Rui Huang, and Jingxue Yin. The supplementation of the theory of periodic solutions for a class of nonlinear diffusion equations. Applied Mathematics and Computation, 346(??):753–766, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309445.

Hu:2015:DMA

[HHZ15] Shuangnian Hu, Shaofang Hong, and Jianrong Zhao. Determinants of matrices associated with arithmetic functions on finitely many quasi-coprime divisor chains. Applied Mathematics and Computation, 258(??):502–508, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001058.

He:2018:LAS

[HHZ18] Shengjie He, Rong-Xia Hao, and Liancui Zuo. The linear (n − 1)-arboricity of some lexicographic product graphs. Applied Mathematics and Computation, 338(??):141–151, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318304958.

Huang:2015:GSB

[HI15a] Yuankan Huang and Takehiro Inohara. Groupseparations based on the repeated prisoners’ dilemma games. Applied Mathematics and Computation, 256(??):267–275, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000545.

Huang:2015:SSS

[HI15b] Yuankan Huang and Takehiro Inohara. Steady-state stock and group size: an approach of dynamic voluntary provisions of public goods. Applied Mathematics and Computation, 270(??):505–510, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501108X.

Han:2015:PSS

[HJ15] Qixing Han and Daqing Jiang. Periodic solution for stochastic nonautonomous multispecies Lotka–Volterra mutualism type ecosystem. Applied Mathematics and Computation, 262(??):204–217, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005007.

Hu:2018:DTC

[HJ18] Xiaochuan Hu and Sophia R.-J. Jang. Dynamics of tumor-CD_{4}^{+}-cytokine-host cells interactions with treatments. Applied Mathematics and Computation, 321(??):700–720, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307828.

Hecht:2018:FDS

[HJKP18] Frederic Hecht, Temur Jangveladze, Zurab Kiguradze, and Olivier Pironneau. Finite difference scheme for one system of nonlinear partial integro-differential equations. Applied Mathematics and Computation, 328(??):287–300, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300730.

Hajipour:2018:PPS

[HJMB18] Mojtaba Hajipour, Amin Jajarmi, Alaeddin Malek, and Dumitru Baleanu. Positivity-preserving sixthorder implicit finite difference weighted essentially non-oscillatory scheme for the nonlinear heat equation. Applied Mathematics and Computation, 325(??):146–158, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308858.

Homenda:2016:MDM

[HJP16] Wladysław Homenda, Agnieszka Jastrzebska, and Witold Pedrycz. Multicriteria decision making inspired by human cognitive processes. Applied Mathematics and Computation, 290(??):392–411, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303502.

Hani:2015:FDS

[HK15] Houda Hani and Moez Khenissi. On a finite difference scheme for blow up solutions for the Chipot– Weissler equation. Applied Mathematics and Computation, 268(??):1199–1216, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009418.

He:2018:HOC

[HK18a] Yuan He and Taekyun Kim. A higher-order convolution for Bernoulli polynomials of the second kind. Applied Mathematics and Computation, 324(??):51–58, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308731.

Heczko:2018:MMD

[HK18b] Jan Heczko and Radek Kottner. Modeling of material damage using finite elements and time homogenization in case of finite strain. Applied Mathematics and Computation, 319(??):264–273, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301881.

Hakula:2015:AMS

[HKL15] H. Hakula, V. Kaarnioja, and M. Laaksonen. Approximate methods for stochastic eigenvalue problems. Applied Mathematics and Computation, 267(??):664–681, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017731.

Hujdurovic:2018:NCG

[HKM18] Ademir Hujdurović, Klavdija Kutnar, and Dragan Marusic. On normality of n-Cayley graphs. Applied Mathematics and Computation, 332(??):469–476, September 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318302273.

Hrinakova:2019:IBV

[HKS19] Katarína Hrináková, Martin Knor, and Riste Skrekovski. An inequality between variable Wiener index and variable Szeged index. Applied Mathematics and Computation, 362 (??):Article 124557, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305405.

He:2017:PPC

[HKW^{+}17] Debiao He, Neeraj Kumar, Huaqun Wang, Lina Wang, and Kim-Kwang Raymond Choo. Privacypreserving certificateless provable data possession scheme for big data storage on cloud. Applied Mathematics and Computation, 314(??):31–43, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304599.

Heggerud:2015:LSA

[HL15a] Christopher M. Heggerud and Kunquan Lan. Local stability analysis of ratio-dependent predator– prey models with predator harvesting rates. Applied Mathematics and Computation, 270(??):349–357, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315011170.

Hu:2015:ABS

[HL15b] Guixin Hu and Yanfang Li. Asymptotic behaviors of stochastic periodic differential equation with Markovian switching. Applied Mathematics and Computation, 264(??):403–416, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004919.

Huang:2015:ISI

[HL15c] Shasha Huang and Dang Luo. Impact of separation of interaction and replacement neighborhoods on spatial reciprocity. Applied Mathematics and Computation, 253(??):318–323, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017597.

Hussein:2015:ITD

[HL15d] M. S. Hussein and D. Lesnic.

Identification of the timedependent conductivity of an inhomogeneous diffusive material. Applied Mathematics and Computation, 269(??):35–58, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315009583.

Hou:2016:SPE

[HL16] Tianliang Hou and Li Li. Superconvergence and a posteriori error estimates of splitting positive definite mixed finite element methods for elliptic optimal control problems. Applied Mathematics and Computation, 273(??):1196–1207, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315012072.

Henderson:2017:SRL

[HL17] Johnny Henderson and Rodica Luca. Systems of Riemann–Liouville fractional equations with multi-point boundary conditions. Applied Mathematics and Computation, 309(??):303–323, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031730228X.

Huang:2018:NLB

[HL18] Jing Huang and Shuchao Li. The normalized Laplacians on both ktriangle graph and kquadrilateral graph with their applications. Applied Mathematics and Computation, 320(??):213–225, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730663X.

Han:2017:RTD

[HLC^{+}17] Weimin Han, Feixiao Long, Wenxiang Cong, Xavier Intes, and Ge Wang. Radiative transfer with delta-Eddington-type phase functions. Applied Mathematics and Computation, 300(??):70–78, May 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300316307172.

Huang:2019:NSB

[HLC19] Chengdai Huang, Huan Li, and Jinde Cao. A novel strategy of bifurcation control for a delayed fractional predator– prey model. Applied Mathematics and Computation, 347(??):808–838, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310051.

Huang:2019:DBR

[HLFZ19] Bing Huang, Huaxiong Li, Guofu Feng, and Xianzhong Zhou. Dominancebased rough sets in multiscale intuitionistic fuzzy decision tables. Applied Mathematics and Computation, 348(??):487–512, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310658.

Huntul:2017:RTD

[HLH17] M. J. Huntul, D. Lesnic, and M. S. Hussein. Reconstruction of timedependent coefficients from heat moments. Applied Mathematics and Computation, 301(??):233–253, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307627.

Hazanee:2019:ITD

[HLIK19] A. Hazanee, D. Lesnic, M. I. Ismailov, and N. B. Kerimov. Inverse timedependent source problems for the heat equation with nonlocal boundary conditions. Applied Mathematics and Computation, 346(??):800–815, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309251.

Hu:2018:NFE

[HLJ18] Xianliang Hu, Yixin Li, and Hangjie Ji. A nodal finite element approximation of a phase field model for shape and topology optimization. Applied Mathematics and Computation, 339(??):663–674, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306155.

Huang:2016:NLD

[HLL16] Jing Huang, Shuchao Li, and Xuechao Li. The normalized Laplacian, degree-Kirchhoff index and spanning trees of the linear polyomino chains. Applied Mathematics and Computation, 289(??):324–334, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303332.

Hu:2018:TBN

[HLL18] Fu-Tao Hu, Lu Li, and Jia-Bao Liu. The total bondage numbers and efficient total dominations of vertex-transitive graphs. Applied Mathematics and Computation, 332(??):35–41, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302194.

Hu:2015:CBM

[HLLT15] Xiuling Hu, Hong-Lin Liao, F. Liu, and I. Turner. A center Box method for radially symmetric solution of fractional subdiffusion equation. Applied Mathematics and Computation, 257(??):467–486, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000296.

Hernandez:2015:WBR

[HLT15] A. Hernández, M. Lattanzi, and N. Thome. Weighted binary relations involving the Drazin inverse. Applied Mathematics and Computation, 253(??):215–223, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017639.

Hernandez:2016:SNP

[HLT16] A. Hernández, M. Lattanzi, and N. Thome. On some new pre-orders defined by weighted Drazin inverses. Applied Mathematics and Computation, 282(??):108–116, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300649.

Huang:2015:GMH

[HLW15] Fei Huang, Xueliang Li, and Shujing Wang. On graphs with maximum Harary spectral radius. Applied Mathematics and Computation, 266(??):937–945, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007882.

He:2017:MKI

[HLX17] Weihua He, Hao Li, and Shuofa Xiao. On the minimum Kirchhoff index of graphs with a given vertex k-partiteness and edge kpartiteness. Applied Mathematics and Computation, 315(??):313–318, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305283.

Han:2015:PSE

[HLZ15] Zhenlai Han, Hongling Lu, and Chao Zhang. Positive solutions for eigenvalue problems of fractional differential equation with generalized plaplacian. Applied Mathematics and Computation, 257(??):526–536, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000272.

Hafezalkotob:2015:CMF

[HM15a] Ashkan Hafezalkotob and Ahmad Makui. Cooperative maximum-flow problem under uncertainty in logistic networks. Applied Mathematics and Computation, 250(??):593–604, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014556.

Huang:2015:NIU

[HM15b] Na Huang and Chang-Feng Ma. The nonlinear inexact Uzawa hybrid algorithms based on onestep Newton method for solving nonlinear saddlepoint problems. Applied Mathematics and Computation, 270(??):291–311, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315010772.

Huang:2015:PMI

[HM15c] Na Huang and Chang-Feng Ma. Parallel multisplitting iteration methods based on M-splitting for the PageRank problem. Applied Mathematics and Computation, 271(??):337–343, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012345.

Huang:2015:BUB

[HM15d] Na Huang and Changfeng Ma. The BGS-Uzawa and BJ-Uzawa iterative methods for solving the saddle point problem. Applied Mathematics and Computation, 256(??):94–108, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500020X.

Hameed:2017:FPA

[HM17] Rabia Hameed and Ghulam Mustafa. Family of a-point b-ary subdivision schemes with bell-shaped mask. Applied Mathematics and Computation, 309(??):289–302, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302576.

Hu:2018:CGL

[HM18a] Jingjing Hu and Changfeng Ma. Conjugate gradient least squares algorithm for solving the generalized coupled Sylvesterconjugate matrix equations. Applied Mathematics and Computation, 334(??):174–191, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830300X.

Huang:2018:IAL

[HM18b] Baohua Huang and Changfeng Ma. An iterative algorithm for the least Frobenius norm least squares solution of a class of generalized coupled Sylvester-transpose linear matrix equations. Applied Mathematics and Computation, 328(??):58–74, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830033X.

Huang:2018:PIM

[HM18c] Na Huang and Chang-Feng Ma. The preconditioned iterative methods with variable parameters for saddle point problem. Applied Mathematics and Computation, 334(??):152–167, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302996.

Harfash:2019:CSE

[HM19a] Akil J. Harfash and Ghazi Abed Meften. Couple stresses effect on instability and nonlinear stability in a double diffusive convection. Applied Mathematics and Computation, 341(??):301–320, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318307598.

Huang:2019:NFM

[HM19b] Ruihao Huang and Lin Mu. A new fast method of solving the high dimensional elliptic eigenvalue problem. Applied Mathematics and Computation, 356(??):338–346, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302425.

Han:2019:HOS

[HMD19] Minggang Han, Qiang Ma, and Xiaohua Ding. Highorder stochastic symplectic partitioned Runge– Kutta methods for stochastic Hamiltonian systems with additive noise. Applied Mathematics and Computation, 346(??): 575–593, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031830907X.

He:2015:CLM

[HMF15] Yedan He, Changfeng Ma, and Bin Fan. A corrected Levenberg–Marquardt algorithm with a nonmonotone line search for the system of nonlinear equations. Applied Mathematics and Computation, 260(??):159–169, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003902.

Hueso:2018:NO

[HMGC18] José L. Hueso, Eulalia Martínez, D. K. Gupta, and Fabricio Cevallos. A note on “Convergence radius of Osada’s method under Hölder continuous condition”. Applied Mathematics and Computation, 321(??):689–699, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307762. See

[ZS14].

Hina:2015:PMP

[HMHA15] S. Hina, M. Mustafa, T. Hayat, and Naif D. Alotaibi. On peristaltic motion of pseudoplastic fluid in a curved channel with heat/mass transfer and wall properties. Applied Mathematics and Computation, 263(??):378–391, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005305.

Hayotov:2015:OQS

[HMS15] Abdullo R. Hayotov, Gradimir V. Milovanović, and Kholmat M. Shadimetov. Optimal quadratures in the sense of Sard in a Hilbert space. Applied Mathematics and Computation, 259(??):637–653, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002945.

Hupkes:2019:MTW

[HMSS19] Hermen Jan Hupkes, Leonardo Morelli, Petr Stehlík, and Vladimír Svígler. Multichromatic travelling waves for lattice Nagumo equations. Applied Mathematics and Computation, 361(??): 430–452, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304382.

Hosseini:2015:IAN

[HMT15a] Babak S. Hosseini, Matthias Möller, and Stefan Turek. Isogeometric analysis of the Navier–Stokes equations with Taylor–Hood Bspline elements. Applied Mathematics and Computation, 267(??):264–281, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031500418X.

Hueso:2015:DFI

[HMT15b] José L. Hueso, Eulalia Martínez, and Carles Teruel. Derivative free iterative methods for nonlinear systems. Applied Mathematics and Computation, 259(??):955–966, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003318.

He:2016:RBS

[HMW16] Fangyi He, Caochuan Ma, and Yanan Wang. On regularity for the Boussinesq system in a bounded domain. Applied Mathematics and Computation, 281(??):148–151, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300637.

Huang:2015:IRD

[HMX15] Na Huang, Chang-Feng Ma, and Ya-Jun Xie. An inexact relaxed DPSS preconditioner for saddle point problem. Applied Mathematics and Computation, 265(??):431–447, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006426.

Han:2017:WSF

[HMZ17] Jiangfeng Han, Stanisław Migórski, and Huidan Zeng. Weak solvability of a fractional viscoelastic frictionless contact problem. Applied Mathematics and Computation, 303(??):1– 18, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300176.

Hajmohammadi:2014:SCV

[HN14] M. R. Hajmohammadi and S. S. Nourazar. On the solution of characteristic value problems arising in linear stability analysis; semi analytical approach. Applied Mathematics and Computation, 239(??):126–132, July 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314006006. See

erratum [NN19].

Hamza:2015:SOA

[HO15] Alaa E. Hamza and Karima M. Oraby. Semigroups of operators and abstract dynamic equations on time scales. Applied Mathematics and Computation, 270(??): 334–348, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315010322.

Howe:2015:SMT

[How15a] J. Andrew Howe. Simulation modeling of a thrown ball bouncing nonlinearly across a grid of cups. Applied Mathematics and Computation, 253(??):61–71, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314017159.

Howk:2015:CCE

[How15b] Cory L. Howk. Convergence of a class of efficient quadrature-based predictor-corrector methods for root-finding. Applied Mathematics and Computation, 252(??): 189–200, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314016853.

Howk:2016:CEQ

[How16] Cory L. Howk. A class of efficient quadraturebased predictor-corrector methods for solving nonlinear systems. Applied Mathematics and Computation, 276(??):394–406, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315300205.

He:2015:LIC

[HP15a] Dongdong He and Kejia Pan. A linearly implicit conservative difference scheme for the generalized Rosenau–Kawahara–RLW equation. Applied Mathematics and Computation, 271(??):323–336, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012394.

Hladik:2015:MIB

[HP15b] Milan Hladík and Evgenija D. Popova. Maximal inner boxes in parametric AE-solution sets with linear shape. Applied Mathematics and Computation, 270(??):606–619, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315010589.

Honorato:2015:DAS

[HP15c] Gerardo Honorato and Sergio Plaza. Dynamical aspects of some convex acceleration methods as purely iterative algorithm for Newton’s maps. Applied Mathematics and Computation, 251(??):507–520, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016154.

He:2016:EPF

[HP16] Dongdong He and Kejia Pan. An energy preserving finite difference scheme for the Poisson–Nernst– Planck system. Applied Mathematics and Computation, 287–288 (??):214–223, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303058.

Henderson:2017:IDW

[HP17] Nélio Henderson and Luciana Pena. The inverse distance weighted interpolation applied to a particular form of the path tubes method: theory and computation for advection in incompressible flow. Applied Mathematics and Computation, 304(??):114–135, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300711.

Harikrishnan:2015:FOS

[HPN15] S. Harikrishnan, P. Prakash, and J. J. Nieto. Forced oscillation of solutions of a nonlinear fractional partial differential equation.

Applied Mathematics and Computation, 254(??):14– 19, March 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314017287.

Henriquez:2016:PAP

[HPR16] Hernán R. Henríquez, Michelle Pierri, and Vanessa Rolnik. Pseudo Sasymptotically periodic solutions of second-order abstract Cauchy problems. Applied Mathematics and Computation, 274(??):590–603, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015246.

Han:2018:RPS

[HPR18] Maoan Han, Tatjana Petek, and Valery G. Romanovski. Reversibility in polynomial systems of ODE’s. Applied Mathematics and Computation, 338(??):44–54, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304685.

Hai:2015:GAN

[HQ15] D. D. Hai and C. Qian. Global attractivity in nonlinear difference equations of higher order with a forcing term. Applied Mathematics and Computation, 264(??):198–207, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005482.

Hu:2015:CBS

[HQJ^{+}15] Gang Hu, Xinqiang Qin, Xiaomin Ji, Guo Wei, and Suxia Zhang. The construction of λμ-B-spline curves and its application to rotational surfaces. Applied Mathematics and Computation, 266(??):194–211, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006803.

Hakenberg:2016:VSB

[HR16] Jan Hakenberg and Ulrich Reif. On the volume of sets bounded by refinable functions. Applied Mathematics and Computation, 272 (part 1)(??):2–19, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008310.

Hauenstein:2018:ASS

[HR18] Jonathan D. Hauenstein and Margaret H. Regan. Adaptive strategies for solving parameterized systems using homotopy continuation. Applied Mathematics and Computation, 332(??):19–34, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301930.

Hernando:2018:RST

[HRLGÁ18] Antonio Hernando, Eugenio Roanes-Lozano, and Alberto Garcia-Álvarez. A recommender system for train routing: When concatenating two minimum length paths is not the minimum length path. Applied Mathematics and Computation, 319(??):486–498, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303442.

Harjani:2015:GCF

[HRS15] J. Harjani, J. Rocha, and K. Sadarangani. Generalized coupled fixed points and its application to a class of systems of functional equations arising in dynamic programming. Applied Mathematics and Computation, 268(??):839–843, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500908X.

Hossain:2017:UMC

[HRS17] Md. Anwar Hossain, Nepal C. Roy, and Sadia Siddiqa. Unsteady mixed convection dusty fluid flow past a vertical wedge due to small fluctuation in free stream and surface temperature. Applied Mathematics and Computation, 293(??):480–492, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305434.

Hien:2015:FTS

[HS15a] Le Van Hien and Doan Thai Son. Finite-time stability of a class of nonautonomous neural networks with heterogeneous proportional delays. Applied Mathematics and Computation, 251(??):14–23, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015768.

Hon:2015:DMK

[HS15b] Y. C. Hon and R. Schaback. Direct meshless kernel techniques for time-dependent equations. Applied Mathematics and Computation, 258(??):220–226, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314018013.

Huang:2015:DCA

[HS15c] Jia-Ping Huang and Ushio Sumita. Development of computational algorithms for pricing European bond options under the influence of macro-economic conditions. Applied Mathematics and Computation, 251(??):453–468, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314016038.

Hitzer:2017:MMM

[HS17] Eckhard Hitzer and Stephen Sangwine. Multivector and multivector matrix inverses in real Clifford algebras. Applied Mathematics and Computation, 311(??):375–389, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317303259.

Hesameddini:2018:TDS

[HS18] Esmail Hesameddini and Mehdi Shahbazi. Twodimensional shifted Legendre polynomials operational matrix method for solving the twodimensional integral equations of fractional order. Applied Mathematics and Computation, 322(??):40– 54, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308068.

Hazarika:2019:ASF

[HSAR19] Bipan Hazarika, H. M. Srivastava, Reza Arab, and Mohsen Rabbani. Application of simulation function and measure of noncompactness for solvability of nonlinear functional integral equations and introduction to an iteration algorithm to find solution. Applied Mathematics and Computation, 360(??):131–146, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303467.

Hsiao:2015:HWM

[Hsi15] Chun-Hui Hsiao. A Haar wavelets method of solving differential equations characterizing the dynamics of a current collection system for an electric locomotive. Applied Mathematics and Computation, 265(??):928–935, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007833.

Hfaiedh:2015:BBM

[HSKHA15] Walid Hfaiedh, Chérif Sadfi, Imed Kacem, and Atidel Hadj-Alouane. A branch-and-bound method for the single-machine scheduling problem under a non-availability constraint for maximum delivery time minimization. Applied Mathematics and Computation, 252(??):496–502, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016361.

Hosseini:2015:TDA

[HSS15] Seyed Mahmoud Hosseini, Jan Sladek, and Vladimir Sladek. Two dimensional analysis of coupled non-Fick diffusionelastodynamics problems in functionally graded materials using meshless local Petrov–Galerkin (MLPG) method. Applied Mathematics and Computation, 268(??):937–946, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009212.

Hwang:2018:TDR

[HSSH18] Jinah Hwang, Myoungin Shin, Suyeon Shin, and Woonjae Hwang. Two dimensional Riemann problem for a 2 × 2 system of hyperbolic conservation laws involving three constant states. Applied Mathematics and Computation, 321(??):49–62, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307440.

Horvath:2017:NUA

[HST17] Zoltán Horváth, Yunfei Song, and Tamás Terlaky. A novel unified approach to invariance conditions for a linear dynamical system. Applied Mathematics and Computation, 298(??):351–367, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306051.

Hien:2016:EST

[HT16] Le Van Hien and Hieu Trinh. Exponential stability of time-delay systems via new weighted integral inequalities. Applied Mathematics and Computation, 275(??):335–344, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015787.

Hu:2018:SVC

[HTM18] Xiaomin Hu, Yingzhi Tian, and Jixiang Meng. Super R^{k}-vertex-connectedness.

Applied Mathematics and Computation, 339(??): 805–811, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318305708.

Hou:2019:SGJ

[HTW19] Yaoping Hou, Zikai Tang, and Dijian Wang. On signed graphs with just two distinct Laplacian eigenvalues. Applied Mathematics and Computation, 351(??):1–7, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300116.

Huang:2017:CMC

[HTZY17] Chao Huang, Lihui Tan, Qian Zhang, and Lihua Yang. Constructions of ϵ-mono-components and mathematical analysis on signal decomposition algorithm. Applied Mathematics and Computation, 293(??):555–564, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305318.

Huang:2015:DCL

[Hua15] Hongying Huang. The direct coupling of local discontinuous Galerkin and natural boundary element method for nonlinear interface problem in R^{3}. Applied Mathematics and Computation, 262(??):232–248, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004944.

Huang:2016:WSN

[Hua16]  Cong Huang.  WENO

scheme with new smoothness indicator for Hamilton– Jacobi equation. Applied Mathematics and Computation, 290(??):21–32, November 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316303319.

Huang:2018:IAU

[Hua18] Youqin Huang. An interval algorithm for uncertain dynamic stability analysis. Applied Mathematics and Computation, 338(??):552–566, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305253.

Huseyin:2019:AIF

[Hus19] Anar Huseyin. Approximation of the integral funnel of the Urysohn type integral operator. Applied Mathematics and Computation, 341(??):277–287, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318307604.

Hasil:2015:LPH

[HV15] Petr Hasil and Michal Veselý. Limit periodic homogeneous linear difference systems. Applied Mathematics and Computation, 265(??):958–972, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007845.

Hasil:2019:MPA

[HV19a] Petr Hasil and Michal Veselý. Modified Prüfer angle and conditional oscillation of perturbed linear and half-linear differential equations. Applied Mathematics and Computation, 361(??):788–809, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319304928.

Hosek:2019:DAD

[HV19b] Radim Hosek and Jonás Volek. Discrete advection– diffusion equations on graphs: Maximum principle and finite volumes. Applied Mathematics and Computation, 361(??):630–644, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304795.

Hernandez-Veron:2016:BCS

[HVR16] M. A. Hernández-Verón and M. J. Rubio. On the ball of convergence of secant-like methods for non-differentiable operators. Applied Mathematics and Computation, 273(??):506–512, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013399.

Hernandez-Veron:2017:LCN

[HVR17] M. A. Hernández-Verón and M. J. Rubio. On the local convergence of a Newton–Kurchatovtype method for nondifferentiable operators. Applied Mathematics and Computation, 304(??):1–9, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300188.

Hui:2015:NOP

[HW15] Eddie C. M. Hui and Guangchen Wang. A new optimal portfolio selection model with owneroccupied housing. Applied Mathematics and Computation, 270(??):714–723, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315011315.

Hauenstein:2017:UEI

[HW17a] Jonathan D. Hauenstein and Charles W. Wampler. Unification and extension of intersection algorithms in numerical algebraic geometry. Applied Mathematics and Computation, 293(??):226–243, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305185.

He:2017:DFT

[HW17b] Xiaoyan He and Qingyun Wang. Distributed finitetime leaderless consensus control for doubleintegrator multi-agent systems with external disturbances. Applied Mathematics and Computation, 295(??):65–76, February 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630604X.

Hu:2017:OIC

[HW17c] Fengxia Hu and Rongming Wang. Optimal investment-consumption strategy with liability and regime switching model under Value-at-Risk constraint. Applied Mathematics and Computation, 313(??):103–118, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302862.

Hao:2018:GDR

[HW18] Jianghao Hao and Peipei Wang. General decay result for thermoelastic beam equation system with time-varying delay. Applied Mathematics and Computation, 334(??):168–173, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303291.

Han:2019:ADI

[HW19] Huan Han and Zhanqing Wang. An alternating direction implicit scheme of a fractional-order diffusion tensor image registration model. Applied Mathematics and Computation, 356(??):105–118, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302255.

Hu:2019:DEC

[HWBL19] Hongping Hu, Haiyan Wang, Yanping Bai, and Maoxing Liu. Determination of endometrial carcinoma with gene expression based on optimized Elman neural network. Applied Mathematics and Computation, 341(??):204–214, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031830777X.

He:2019:TTS

[HWL^{+}19a] Qiang He, Xingwei Wang, Zhencheng Lei, Min Huang, Yuliang Cai, and Lianbo Ma. TIFIM: a two-stage iterative framework for influence maximization in social networks. Applied Mathematics and Computation, 354(??):338–352, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301602.

Huang:2019:LCD

[HWL^{+}19b] Danjun Huang, Yiqiao Wang, Jing Lv, Yanping Yang, and Weifan Wang. List coloring and diagonal coloring for plane graphs of diameter two. Applied Mathematics and Computation, 363 (??):Article 124562, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305454.

Huang:2017:GMS

[HWXC17] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. The generalized modified shiftsplitting preconditioners for nonsymmetric saddle point problems. Applied Mathematics and Computation, 299(??):95–118, April 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307147.

He:2015:DFT

[HWY15a] Xiaoyan He, Qingyun Wang, and Wenwu Yu. Distributed finite-time containment control for second-order nonlinear multi-agent systems. Applied Mathematics and Computation, 268(??): 509–521, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315008826.

He:2015:FTD

[HWY15b] Xiaoyan He, Qingyun Wang, and Wenwu Yu. Finite-time distributed cooperative attitude tracking control for multiple rigid spacecraft. Applied Mathematics and Computation, 256(??):724–734, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000934.

He:2017:SDQ

[HWZ17] Zhuo-Heng He, Qing-Wen Wang, and Yang Zhang. Simultaneous decomposition of quaternion matrices involving η-Hermicity with applications. Applied Mathematics and Computation, 298(??):13–35, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306609.

Huang:2015:SNP

[HXLC15] Zheng-Ge Huang, Zhong Xu, Quan Lu, and Jing-Jing Cui. Some new preconditioned generalized AOR methods for generalized least-squares problems. Applied Mathematics and Computation, 269(??):87–104, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009819.

Hu:2018:RHA

[HXM18] Bei-Bei Hu, Tie-Cheng Xia, and Wen-Xiu Ma. Riemann–Hilbert approach for an initial-boundary value problem of the two-component modified Korteweg–de Vries equation on the half-line. Applied Mathematics and Computation, 332(??): 148–159, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318302224.

Hu:2016:GDP

[HXS^{+}16] Jiang-Hong Hu, Ya-Kui Xue, Gui-Quan Sun, Zhen Jin, and Juan Zhang. Global dynamics of a predator–prey system modeling by metaphysiological approach. Applied Mathematics and Computation, 283(??):369–384, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301527.

Hu:2019:PBS

[HXW^{+}19a] Xiaohui Hu, Jianwei Xia, Yunliang Wei, Bo Meng, and Hao Shen. Passivitybased state synchronization for semi-Markov jump coupled chaotic neural networks with randomly occurring time delays. Applied Mathematics and Computation, 361(??):32–41, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304096.

Huang:2019:MHL

[HXW^{+}19b] Zhengguo Huang, Jianwei Xia, Jing Wang, Yunliang Wei, Zhen Wang, and Jian Wang. Mixed H_{∞}/l_{2}– l_{∞} state estimation for switched genetic regulatory networks subject to packet dropouts: a persistent dwell-time switching mechanism. Applied Mathematics and Computation, 355(??):198–212, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301912.

Yuan:2018:WBE

[hY18]  Xu hua Yuan.  A

well-balanced element-free Galerkin method for the nonlinear shallow water equations. Applied Mathematics and Computation, 331(??):46–53, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318300900.

Han:2017:NMF

[HYB17] Jiayu Han, Yidu Yang, and Hai Bi. A new multigrid finite element method for the transmission eigenvalue problems. Applied Mathematics and Computation, 292(??):96–106, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630460X.

Huang:2019:UMB

[HYJ19] Zhiming Huang, Lin Yang, and Wen Jiang. Uncertainty measurement with belief entropy on the interference effect in the quantum-like Bayesian networks. Applied Mathematics and Computation, 347(??):417–428, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310105.

Hao:2019:FTC

[HYL19a] Li-Ying Hao, Ying Yu, and Hui Li. Fault tolerant control of UMV based on sliding mode output feedback. Applied Mathematics and Computation, 359(??):433–455, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303571.

Huang:2019:RST

[HYL^{+}19b] Xianbei Huang, Wei Yang, Yaojun Li, Baoyun Qiu, Qiang Guo, and Liu Zhuqing. Review on the sensitization of turbulence models to rotation/ curvature and the application to rotating machinery. Applied Mathematics and Computation, 341:46–69, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307410.

Huang:2015:NMB

[HYW^{+}15] Chaobao Huang, Xijun Yu, Cheng Wang, Zhenzhen Li, and Na An. A numerical method based on fully discrete direct discontinuous Galerkin method for the time fractional diffusion equation. Applied Mathematics and Computation, 264(??):483–492, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S009630031500555X.

Huang:2018:DSO

[HYZ^{+}18] Tousheng Huang, Hongju Yang, Huayong Zhang, Xuebing Cong, and Ge Pan. Diverse self-organized patterns and complex pattern transitions in a discrete ratio-dependent predator– prey system. Applied Mathematics and Computation, 326(??):141–158, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300250.

Han:2016:MSC

[HZ16] Le Han and Qin Zhang. Multi-stage convex relaxation method for low-rank and sparse matrix separation problem. Applied Mathematics and Computation, 284(??):175–184, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301849.

He:2017:NAF

[HZ17a] Guo He and Chuanlin Zhang. On the numerical approximation for Fourier-type highly oscillatory integrals with Gauss-type quadrature rules. Applied Mathematics and Computation, 308(??):96–104, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301959.

Hu:2017:NAS

[HZ17b] Yahong Hu and Quanyong Zhu. New analytic solutions of two-dimensional nonlocal nonlinear media. Applied Mathematics and Computation, 305(??):53– 61, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300310.

Huang:2017:WEN

[HZ17c] Keke Huang and Xiaoping Zheng. A weighted evolving network model for pedestrian evacuation. Applied Mathematics and Computation, 298(??):57– 64, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306567.

Huang:2017:BBC

[HZCY17] Keke Huang, Xiaoping Zheng, Yuan Cheng, and Yeqing Yang. Behaviorbased cellular automaton model for pedestrian dynamics. Applied Mathematics and Computation, 292(??):417–424, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304325.

Hu:2019:SMC

[HZK^{+}19] Jun Hu, Panpan Zhang, Yonggui Kao, Hongjian Liu, and Dongyan Chen. Sliding mode control for Markovian jump repeated scalar nonlinear systems with packet dropouts: the uncertain occurrence probabilities case. Applied Mathematics and Computation, 362(??): Article 124574, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305570.

Hong:2018:STM

[HZL18] Yunchao Hong, Zhongxun Zhu, and Amu Luo. Some transformations on multiplicative eccentricity resistance-distance and their applications. Applied Mathematics and Computation, 323(??):75–85, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308445.

Hou:2017:NEP

[HZLY17] Tianliang Hou, Jiaqi Zhang, Yanzhong Li, and Yueting Yang. New elliptic projections and a priori error estimates of H^{1}-Galerkin mixed finite element methods for optimal control problems governed by parabolic integro-differential equations. Applied Mathematics and Computation, 311(??):29–46, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302886.

Huang:2015:EHS

[HZS15] Keke Huang, Xiaoping Zheng, and Yunpeng Su. Effect of heterogeneous sub-populations on the evolution of cooperation. Applied Mathematics and Computation, 270(??):681–687, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315011121.

Huang:2015:BEE

[HZYW15] Keke Huang, Xiaoping Zheng, Yeqing Yang, and Tao Wang. Behavioral evolution in evacuation crowd based on heterogeneous rationality of small groups. Applied Mathematics and Computation, 266(??):501–506, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500689X.

Ispir:2015:RCL

[IAK15] Nurhayat Ispir, Purshottam Narain Agrawal, and Arun Kajla. Rate of convergence of Lupas Kantorovich operators based on Polya distribution. Applied Mathematics and Computation, 261(??):323–329, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004130.

Iqbal:2018:NCB

[IAW18] Muhammad Kashif Iqbal, Muhammad Abbas, and Imtiaz Wasim. New cubic B-spline approximation for solving third order Emden–Flower type equations. Applied Mathematics and Computation, 331(??):319–333, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301905.

Ibrir:2015:MRC

[IB15] S. Ibrir and M. Bettayeb.

Model reduction of a class of discrete-time nonlinear systems. Applied Mathematics and Computation, 250(??):78–93, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014386.

Ilic:2019:PMP

[IB19] Aleksandar Ilić and Milan Basić. Path matrix and path energy of graphs. Applied Mathematics and Computation, 355(??):537–541, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930195X.

Ilic:2015:DBG

[ID15] Aleksandar Ilić and Matthias Dehmer. On the distance based graph entropies. Applied Mathematics and Computation, 269(??):647–650, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501022X.

Ilati:2019:DMN

[ID19] Mohammad Ilati and Mehdi Dehghan. DMLPG method for numerical simulation of soliton collisions in multi-dimensional coupled damped nonlinear Schrödinger system which arises from Bose–Einstein condensates. Applied Mathematics and Computation, 346(??):244–253, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308828.

Ieda:2015:IMF

[Ied15] Masashi Ieda. An implicit method for the finite time horizon Hamilton–Jacobi– Bellman quasi-variational inequalities. Applied Mathematics and Computation, 265(??):163–175, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004890.

Islam:2015:NSE

[IH15] Md. Shafiqul Islam and Md. Bellal Hossain. Numerical solutions of eighth order BVP by the Galerkin residual technique with Bernstein and Legendre polynomials. Applied Mathematics and Computation, 261(??):48–59, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004051.

Ilic:2017:CCG

[II17] Aleksandar Ilić and Milovan Ilić. Counterexamples to conjectures on graph distance measures based on topological indexes. Applied Mathematics and Computation, 296(??):148–152, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306221.

Ishkhanyan:2018:GCH

[II18] T. A. Ishkhanyan and A. M. Ishkhanyan. Generalized confluent hypergeometric solutions of the Heun confluent equation. Applied Mathematics and Computation, 338(??):507–514, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305423.

Ibrahim:2015:EUA

[IJ15] Rabha W. Ibrahim and Jay M. Jahangiri. Existence and uniqueness of an attractive nonlinear diffusion system. Applied Mathematics and Computation, 257(??):169–177, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314009370.

Irshad:2016:OCU

[IKHS16] Misbah Irshad, Sobia Khalid, Malik Zawwar Hussain, and Muhammad Sarfraz. Outline capturing using rational functions with the help of genetic algorithm. Applied Mathematics and Computation, 274(??):661–678, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013466.

Idczak:2015:EOS

[IKMW15] Dariusz Idczak, Rafal Kamocki, Marek Majewski, and Stanisław Walczak. Existence of optimal solutions to Lagrange problems for Roesser type systems of the first and fractional orders. Applied Mathematics and Computation, 266(??):809–819, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007730.

Ilic:2018:BGB

[Ili18] Aleksandar Ilić. Bipartite graphs with bounded vertex degree and largest eigenvalue of the form r. Applied Mathematics and Computation, 330(??):288–291, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301620.

Ilie:2019:FSI

[Ili19a]  Marcel Ilie.  Fluid-

structure interaction in turbulent flows; a CFD based aeroelastic algorithm using LES. Applied Mathematics and Computation, 342(??):309–321, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307683.

Ilie:2019:FCC

[Ili19b]  Marcel Ilie.  A fully-

coupled CFD/CSD computational approach for aeroelastic studies of helicopter blade-vortex interaction. Applied Mathematics and Computation, 347(??):122–142, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309433.

Inarrea:2015:LSG

[ILP^{+}15] M. Iñarrea, V. Lanchares, J. F. Palacián, A. I. Pascual, J. P. Salas, and P. Yanguas. Lyapunov stability for a generalized Hénon–Heiles system in a rotating reference frame. Applied Mathematics and Computation, 253(??):159–171, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017263.

Inarrea:2017:SPR

[ILPE17] M. Iñarrea, V. Lanchares, A. I. Pascual, and A. Elipe. Stability of the permanent rotations of an asymmetric gyrostat in a uniform Newtonian field. Applied Mathematics and Computation, 293(??):404–415, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305367.

Izquierdo-Marquez:2019:NCA

[IMTJ19] Idelfonso Izquierdo-Marquez and Jose Torres-Jimenez. New covering array numbers. Applied Mathematics and Computation, 353(??):134–146, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300864.

Inohara:2016:STT

[Ino16] Takehiro Inohara. State transition time analysis in the Graph Model for Conflict Resolution. Applied Mathematics and Computation, 274(??):372–382, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014824.

Izgi:2019:NPS

[IÖ19] Burhaneddin Izgi and Murat Özkaya. A new perspective to the solution and creation of zero sum matrix game with matrix norms. Applied Mathematics and Computation, 341:148–159, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307276.

Iqbal:2016:CDE

[IR16] Naveed Iqbal and Cornelia Rauh. Coupling of discrete element model (DEM) with computational fluid mechanics (CFD): a validation study. Applied Mathematics and Computation, 277(??):154–163, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315300254.

Interian:2018:EIN

[IR18] Ruben Interian and Celso C.

Ribeiro. An empirical investigation of network polarization. Applied Mathematics and Computation, 339(??):636–650, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306325.

Irfan:2016:NCH

[IS16] Nagma Irfan and A. H.

Siddiqi. A novel computational hybrid approach in solving Hankel transform. Applied Mathematics and Computation, 281(??):121–129, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300285.

Islas:2017:NIS

[IS17] A. Islas and C. M.

Schober. Numerical investigation of the stability of the rational solutions of the nonlinear Schrödinger equation. Applied Mathematics and Computation, 305(??):17– 26, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300863.

Iscan:2016:GDT

[Isc16] Imdat Iscan. On generalization of different type inequalities for harmonically quasi-convex functions via fractional integrals. Applied Mathematics and Computation, 275(??):287–298, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015763.

Ivanov:2015:GDP

[IT15] Anatoli F. Ivanov and Sergei I. Trofimchuk. On global dynamics in a periodic differential equation with deviating argument. Applied Mathematics and Computation, 252(??):446–456, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016695.

Iqbal:2017:PFS

[IWL17] Naveed Iqbal, Ranchao Wu, and Biao Liu. Pattern formation by superdiffusion in FitzHugh– Nagumo model. Applied Mathematics and Computation, 313(??):245–258, November 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317303843.

Ito:2019:DST

[IYM19] Hiromu Ito, Taro Yamamoto, and Satoru Morita. Demography of sexually transmitted infections with vertical transmission. Applied Mathematics and Computation, 348(??):363–370, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310439.

Ilchev:2016:EEA

[IZ16] A. Ilchev and B. Zlatanov. Error estimates for approximation of coupled best proximity points for cyclic contractive maps. Applied Mathematics and Computation, 290(??): 412–425, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300316303939.

Jang:2015:DSN

[JA15] Sophia R.-J. Jang and Edward J. Allen. Deterministic and stochastic nutrient-phytoplankton-zooplankton models with periodic toxin producing phytoplankton. Applied Mathematics and Computation, 271(??):52–67, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315011212.

Jafari:2017:BMT

[JA17] Nasrin Jafari and Mojtaba Azhari. Buckling of moderately thick arbitrarily shaped plates with intermediate point supports using a simple hp-cloud method. Applied Mathematics and Computation, 313(??):196–208, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317303910.

Jahanbani:2017:SNL

[Jah17] Akbar Jahanbani. Some new lower bounds for energy of graphs. Applied Mathematics and Computation, 296(??):233–238, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306178.

Jaiswal:2018:NNL

[Jai18] Bharat Raj Jaiswal. A non-Newtonian liquid sphere embedded in a polar fluid saturated porous medium: Stokes flow. Applied Mathematics and Computation, 316(??):488–503, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730560X.

Jang:2015:NPS

[Jan15] Lee-Chae Jang. A note on the pseudo Stolarsky type inequality for the ḡintegral. Applied Mathematics and Computation, 269(??):809–815, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010097.

Jog:2016:TFE

[JAN16] C. S. Jog, Manish Agrawal, and Arup Nandy. The time finite element as a robust general scheme for solving nonlinear dynamic equations including chaotic systems. Applied Mathematics and Computation, 279(??):43–61, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315016082.

Jarczyk:2015:PML

[Jar15] Justyna Jarczyk. Parametrized means and limit properties of their Gaussian iterations. Applied Mathematics and Computation, 261(??):81–89, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003999.

Jana:2015:DGP

[JAU15] Debaldev Jana, Rashmi Agrawal, and Ranjit Kumar Upadhyay. Dynamics of generalist predator in a stochastic environment: Effect of delayed growth and prey refuge. Applied Mathematics and Computation, 268(??):1072–1094, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008796.

Jain:2015:DOF

[JBS15] Madhu Jain, Amita Bhagat, and Chandra Shekhar. Double orbit finite retrial queues with priority customers and service interruptions. Applied Mathematics and Computation, 253(??):324–344, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017202.

Joldes:2015:MML

[JCW^{+}15] Grand Roman Joldes, Habibullah Amin Chowdhury, Adam Wittek, Barry Doyle, and Karol Miller. Modified moving least squares with polynomial bases for scattered data approximation. Applied Mathematics and Computation, 266(??):893–902, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315007924.

Jakubowska-Ciszek:2018:ATS

[JCW18] A. Jakubowska-Ciszek and J. Walczak. Analysis of the transient state in a parallel circuit of the class RL_{β}C_{α}. Applied Mathematics and Computation, 319(??):287–300, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302102.

Jiang:2019:GMR

[JCX19] Ling Jiang, Jinde Cao, and Lianglin Xiong. Generalized multiobjective robustness and relations to set-valued optimization. Applied Mathematics and Computation, 361(??):599–608, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304710.

Jhinga:2018:NFD

[JDG18] Aman Jhinga and Varsha Daftardar-Gejji. A new finite-difference predictorcorrector method for fractional differential equations. Applied Mathematics and Computation, 336(??):418–432, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318304065.

Jiang:2015:RPT

[JDL15] Hao-Jie Jiang, Hong-Liang Dai, and Shu-Zhi Li. Refined plate theory for bending analysis of a HSLA steel plate under 3D temperature field. Applied Mathematics and Computation, 250(??):497–513, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014970.

Jodlbauer:2018:HBI

[JDS18] Herbert Jodlbauer, Matthias Dehmer, and Sonja Strasser. A hybrid binomial inverse hypergeometric probability distribution: Theory and applications. Applied Mathematics and Computation, 338(??):33–43, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304880.

Jeong:2016:GVA

[Jeo16] Jae Ug Jeong. Generalized viscosity approximation methods for mixed equilibrium problems and fixed point problems. Applied Mathematics and Computation, 283(??):168–180, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300321.

Javarsineh:2017:UBG

[JFT17] Mehrnoosh Javarsineh and Gholam Hossein Fath-Tabar. Unicyclic and bicyclic graphs with exactly three Q-main eigenvalues. Applied Mathematics and Computation, 315(??):603–614, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304460.

Jin:2017:MGS

[JG17] Zhong Jin and David Y. Gao. On modeling and global solutions for d.c. optimization problems by canonical duality theory. Applied Mathematics and Computation, 296(??):168–181, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306087.

Jiang:2016:AVF

[JGBY16] Hao Jiang, Stef Graillat, Roberto Barrio, and Canqun Yang. Accurate, validated and fast evaluation of elementary symmetric functions and its application. Applied Mathematics and Computation, 273(??):1160–1178, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012151.

Jana:2015:GDP

[JGGK15] Soovoojeet Jana, Abhijit Ghorai, Srabani Guria, and T. K. Kar. Global dynamics of a predator, weaker prey and stronger prey system. Applied Mathematics and Computation, 250(??):235–248, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014726.

Jiang:2016:SMC

[JGKL16] Baoping Jiang, Cunchen Gao, Yonggui Kao, and Zhen Liu. Sliding mode control of Markovian jump systems with incomplete information on timevarying delays and transition rates. Applied Mathematics and Computation, 290(??):66–79, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303472.

Jiang:2015:PBS

[JGX15] Baoping Jiang, Cunchen Gao, and Jing Xie. Passivity based sliding mode control of uncertain singular Markovian jump systems with time-varying delay and nonlinear perturbations. Applied Mathematics and Computation, 271(??):187–200, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011996.

Jiang:2015:EIM

[JH15] Xiaoyu Jiang and Kicheon Hong. Explicit inverse matrices of Tribonacci skew circulant type matrices. Applied Mathematics and Computation, 268(??):93–102, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315007341.

Jiang:2017:SCD

[JH17] Xiaoyu Jiang and Kicheon Hong. Skew cyclic displacements and inversions of two innovative patterned matrices. Applied Mathematics and Computation, 308(??):174–184, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302060.

Jin:2019:TDP

[JHL19] Wei Jin, Yang Huang, and Wei Jun Liu. Twodistance-primitive graphs with prime valency. Applied Mathematics and Computation, 357(??): 310–316, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300319302607.

Ji:2015:FRE

[JHWZ15] Meng-Di Ji, Yong He, Min Wu, and Chuan-Ke Zhang. Further results on exponential stability of neural networks with timevarying delay. Applied Mathematics and Computation, 256(??):175–182, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000120.

Jiang:2015:SFD

[Jia15] Weihua Jiang. Solvability of fractional differential equations with p-Laplacian at resonance. Applied Mathematics and Computation, 260(??):48– 56, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003410.

Jin:2017:ARN

[Jin17] Zemin Jin. Anti-Ramsey numbers for matchings in 3-regular bipartite graphs. Applied Mathematics and Computation, 292(??):114–119, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304751.

Jikantoro:2015:ZDS

[JIS15] Y. D. Jikantoro, F. Ismail, and N. Senu. Zerodissipative semi-implicit hybrid method for solving oscillatory or periodic problems. Applied Mathematics and Computation, 252(??):388–396, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016749.

Jikantoro:2018:HMD

[JISI18] Y. D. Jikantoro, F. Ismail, N. Senu, and Z. B. Ibrahim. Hybrid methods for direct integration of special third order ordinary differential equations. Applied Mathematics and Computation, 320(??):452–463, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306926.

Jasim:2015:SNS

[JJ15a] M. K. Jasim and Zainab John. Some new similarity solutions of Einstein field equations for spherical symmetric metric of class two. Applied Mathematics and Computation, 253(??):242–247, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314017299.

Jiang:2015:GPT

[jJ15b] Zhi jie Jiang. Generalized product-type operators from weighted Bergman– Orlicz spaces to Bloch– Orlicz spaces. Applied Mathematics and Computation, 268(??):966–977, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008814.

Jiang:2015:PTO

[jJ15c]  Zhi jie Jiang.  On

a product-type operator from weighted Bergman– Orlicz space to some weighted type spaces. Applied Mathematics and Computation, 256(??):37– 51, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000399.

Johnson:2016:CFM

[JJ16] Michael J. Johnson and Hakim S. Johnson. A constructive framework for minimal energy planar curves. Applied Mathematics and Computation, 276(??):172–181, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015490.

Jahedi:2018:CDB

[JJM18] S. Jahedi, F. Javadi, and M. J. Mehdipour. Compression and decompression based on discrete weighted transform. Applied Mathematics and Computation, 335(??):133–145, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303667.

Jeong:2015:IMT

[JK15] Jae Ug Jeong and Soo Hwan Kim. An iterative method for total quasiϕ-asymptotically nonexpansive semi-groups and generalized mixed equilibrium problems in Banach spaces. Applied Mathematics and Computation, 254(??):39–48, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017962.

Jung:2015:CBC

[JKK15] Sungeun Jung, Yoenha Kim, and Eungil Ko. Characterizations of binormal composition operators with linear fractional symbols on H^{2}. Applied Mathematics and Computation, 261(??):252–263, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004105.

Ji:2015:GWP

[JL15] Ruihong Ji and Shan Li. Global well-posedness for MHD system with mixed partial dissipation and magnetic diffusion in 2 + 1/2 dimensions. Applied Mathematics and Computation, 258(??):325–341, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001101.

Jiang:2015:WTQ

[JLL15a] Shuxia Jiang, Guy Latouche, and Yuanyuan Liu. Wavelet transform for quasi-birth–death process with a continuous phase set. Applied Mathematics and Computation, 252(??):354–376, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016774.

Jiao:2015:PLR

[JLL15b] Hongwei Jiao, Sanyang Liu, and Nan Lu. A parametric linear relaxation algorithm for globally solving nonconvex quadratic programming. Applied Mathematics and Computation, 250(??):973–985, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015550.

Ji:2016:CTA

[JLL16] Duofa Ji, Weidong Lei, and Hongjun Li. Corner treatment by assigning dual tractions to every node for elastodynamics in TD–BEM. Applied Mathematics and Computation, 284(??):125–135, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301783.

Ji:2018:LBR

[JLW18] Shengjin Ji, Mengmeng Liu, and Jianliang Wu. A lower bound of revised Szeged index of bicyclic graphs. Applied Mathematics and Computation, 316(??):480–487, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306021.

Jian:2015:MIS

[JLX15a] Hui Jian, Bin Liu, and Songfa Xie. Monotone iterative solutions for nonlinear fractional differential systems with deviating arguments. Applied Mathematics and Computation, 262(??):1– 14, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004464.

Jun:2015:DSV

[JLX15b] Luo Jun, Liu Liheng, and Wu Xianyi. A doublesubpopulation variant of the bat algorithm. Applied Mathematics and Computation, 263(??):361–377, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315004920.

Jia:2018:NMS

[JLY18] Lifen Jia, Waichon Lio, and Xiangfeng Yang. Numerical method for solving uncertain spring vibration equation. Applied Mathematics and Computation, 337(??):419–427, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304545.

Jacobs:2015:LAD

[JM15] B. A. Jacobs and E. Momoniat. A locally adaptive, diffusion based text binarization technique. Applied Mathematics and Computation, 269(??):464–472, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010139.

Jamal:2018:ACA

[JM18] Sameerah Jamal and Nkosingiphile Mnguni. Approximate conditions admitted by classes of the Lagrangian L = (1/2)(−u^{′2}+u^{2})+ϵ^{i}G_{i}(u, u′, u′′). Applied Mathematics and Computation, 335(??):65– 74, October 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318303382.

Jeon:2017:MHD

[JMK17] Hyoseung Jeon, Mehrdad Massoudi, and Jeongho Kim. Magneto-hydrodynamics-driven mixing of a reagent and a phosphate-buffered solution: a computational study. Applied Mathematics and Computation, 298(??):261–271, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306944.

Jiang:2015:FEV

[JMW15] Yu Jiang, Liquan Mei, and Huiming Wei. A finite element variational multiscale method for incompressible flow. Applied Mathematics and Computation, 266(??):374–384, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315006797.

Jafarian:2018:ANN

[JNGB18] Ahmad Jafarian, Safa Measoomy Nia, Alireza Khalili Golmankhaneh, and Dumitru Baleanu. On artificial neural networks approach with new cost functions. Applied Mathematics and Computation, 339(??):535–545, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306192.

Jain:2015:SGB

[JO15] Shashi Jain and Cornelis W. Oosterlee. The Stochastic Grid Bundling Method: Efficient pricing of Bermudan options and their Greeks. Applied Mathematics and Computation, 269(??):412–431, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010073.

Jafari:2016:NMB

[JOÖ16] Rahim Jafari and Tuba Okutucu-Özyurt. 3D numerical modeling of boiling in a microchannel by arbitrary Lagrangian– Eulerian (ALE) method. Applied Mathematics and Computation, 272 (part 3) (??):593–603, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315003471.

Jian:2015:SSF

[JPTL15] Jin-Bao Jian, Hua-Qin Pan, Chun-Ming Tang, and Jian-Ling Li. A strongly sub-feasible primaldual quasi interior-point algorithm for nonlinear inequality constrained optimization. Applied Mathematics and Computation, 266(??):560–578, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007158.

Jiang:2016:TVP

[JPY16] Chun-Fu Jiang, Hong-Yi Peng, and Yu-Kuan Yang. Tail variance of portfolio under generalized Laplace distribution. Applied Mathematics and Computation, 282(??):187–203, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300893.

Jiao:2019:SSA

[JPZ^{+}19] Ticao Jiao, Ju H. Park, Guangdeng Zong, Jian Liu, and Yu Chen. Stochastic stability analysis of switched genetic regulatory networks without stable subsystems. Applied Mathematics and Computation, 359(??):261–277, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303479.

Ji:2015:RZI

[JQL15] Shengjin Ji, Yongke Qu, and Xia Li. The reformulated Zagreb indices of tricyclic graphs. Applied Mathematics and Computation, 268(??):590–595, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008395.

Jung:2015:FET

[JRM15] Soon-Mo Jung, Michael Th.

Rassias, and Cristinel Mortici. On a functional equation of trigonometric type. Applied Mathematics and Computation, 252(??):294–303, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314016737.

Jang:2016:SPT

[JRSK16] Lee-Chae Jang, C. S. Ryoo, J. J. Seo, and Hyuck In Kwon. Some properties of the twisted Changhee polynomials and their zeros. Applied Mathematics and Computation, 274(??):169–177, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315014034.

Jiang:2015:NFN

[JS15] Heping Jiang and Yongli Song. Normal forms of non-resonance and weak resonance double Hopf bifurcation in the retarded functional differential equations and applications. Applied Mathematics and Computation, 266 (??):1102–1126, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007961.

Jiang:2016:EDP

[JS16] Fangfang Jiang and Jitao Sun. On the existence of discontinuous periodic solutions for a class of Liénard systems with impulses. Applied Mathematics and Computation, 291(??):259–265, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304349.

Jia:2019:SDU

[JS19] Lifen Jia and Yuhong Sheng. Stability in distribution for uncertain delay differential equation. Applied Mathematics and Computation, 343(??):49–56, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308154.

Jidesh:2015:QCY

[JSG15] P. Jidesh, Vorkady S.

Shubha, and Santhosh George. A quadratic convergence yielding iterative method for the implementation of Lavrentiev regularization method for illposed equations. Applied Mathematics and Computation, 254(??):148–156, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314017512.

Jiang:2017:FOA

[JSLW17] Chaolong Jiang, Jianqiang Sun, Haochen Li, and Yifan Wang. A fourthorder AVF method for the numerical integration of sine-Gordon equation. Applied Mathematics and Computation, 313(??):144–158, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317303673.

Jain:2016:HAV

[JSM16] M. K. Jain, Sachin Sharma, and R. K. Mohanty. High accuracy variable mesh method for nonlinear two-point boundary value problems in divergence form. Applied Mathematics and Computation, 273(??):885–896, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013727.

Jiao:2018:FRD

[JSW^{+}18] Shiyu Jiao, Hao Shen, Yunliang Wei, Xia Huang, and Zhen Wang. Further results on dissipativity and stability analysis of Markov jump generalized neural networks with time-varying interval delays. Applied Mathematics and Computation, 336(??):338–350, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304168.

Jiang:2016:ASP

[JT16] Zhao-Lin Jiang and Xia Tang. Analysis of the structured perturbation for the BSCCB linear system. Applied Mathematics and Computation, 277(??):1–9, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300187.

Jiang:2017:RIA

[JT17] Bo Jiang and Yongge Tian.

Rank/inertia approaches to weighted least-squares solutions of linear matrix equations. Applied Mathematics and Computation, 315(??):400–413, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300317305404.

Jurkiewicz:2017:ADS

[Jur17] Marcin Jurkiewicz. Average distance is submultiplicative and subadditive with respect to the strong product of graphs. Applied Mathematics and Computation, 315(??):278–285, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304381.

Jonasova:2018:NAC

[JV18] Alena Jonásová and Jan Vimmr. Noninvasive assessment of carotid artery stenoses by the principle of multiscale modelling of non-Newtonian blood flow in patient-specific models. Applied Mathematics and Computation, 319(??):598–616, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304939.

Wang:2018:ECDa

[jWgX18] Jun jie Wang and Ai guo Xiao. An efficient conservative difference scheme for fractional Klein–Gordon– Schrödinger equations. Applied Mathematics and Computation, 320(??): 691–709, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317305866.

Jiao:2019:GES

[JX19] Zhe Jiao and Yong Xu.

Global existence and stability for semilinear wave equations damped by time-dependent boundary frictions. Applied Mathematics and Computation, 354(??):282–295, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301365.

Xie:2018:ECM

[jXlZX18] Lie jun Xie, Cai lian Zhou, and Song Xu. An effective computational method for solving linear multi-point boundary value problems. Applied Mathematics and Computation, 321(??):255–266, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307051.

Jiang:2015:CPI

[JXW15] Zhaolin Jiang, Hongxia Xin, and Hongwei Wang. On computing of positive integer powers for rcirculant matrices. Applied Mathematics and Computation, 265(??):409–413, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006396.

Jiang:2019:RBT

[JY19] Zhuozhuo Jiang and Weigen Yan. Resistances between two nodes of a path network. Applied Mathematics and Computation, 361(??):42–46, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319303935.

Jiang:2019:SCH

[JYF19] Gui-Sheng Jiang, Gui-Dong Yu, and Yi Fang. Spectral conditions and Hamiltonicity of a balanced bipartite graph with large minimum degree. Applied Mathematics and Computation, 356(??):137–143, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930253X.

Jiang:2016:NES

[JYS16] Feng Jiang, Hua Yang, and Yi Shen. A note on exponential stability for second-order neutral stochastic partial differential equations with infinite delays in the presence of impulses. Applied Mathematics and Computation, 287–288 (??):125–133, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302715.

Jiang:2018:LEG

[JYSR18] Guisheng Jiang, Guidong Yu, Wei Sun, and Zheng Ruan. The least eigenvalue of graphs whose complements have only two pendent vertices. Applied Mathematics and Computation, 331(??):112–119, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301589.

Jaggi:2015:EMA

[JYVS15] Chandra K. Jaggi, V. S. S. Yadavalli, Mona Verma, and Anuj Sharma. An EOQ model with allowable shortage under trade credit in different scenario. Applied Mathematics and Computation, 252(??):541–551, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016944.

Jiang:2015:NSN

[JZ15a] Zhaolin Jiang and Jianwei Zhou. A note on spectral norms of even-order rcirculant matrices. Applied Mathematics and Computation, 250(??):368–371, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015422.

Jin:2015:FPS

[JZ15b] Yunguo Jin and Shouming Zhong. Function projective synchronization in complex networks with switching topology and stochastic effects. Applied Mathematics and Computation, 259(??):730–740, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002805.

Ji:2019:EEE

[JZ19a] Bingquan Ji and Luming Zhang. Error estimates of exponential wave integrator Fourier pseudospectral methods for the nonlinear Schrödinger equation. Applied Mathematics and Computation, 343(??):100–113, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308191.

Jiang:2019:OBS

[JZ19b] Yan Jiang and Junyong Zhai. Observerbased stabilization of sector-bounded nonlinear stochastic systems in the presence of intermittent measurements. Applied Mathematics and Computation, 346(??):740–752, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308993.

Kajla:2015:SDT

[KA15a] Arun Kajla and P. N. Agrawal. Szász–Durrmeyer type operators based on Charlier polynomials. Applied Mathematics and Computation, 268(??):1001–1014, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009108.

Kang:2015:DFS

[KA15b] Hongchao Kang and Congpei An. Differentiation formulas of some hypergeometric functions with respect to all parameters. Applied Mathematics and Computation, 258(??):454–464, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001708.

Khamene:2015:DRD

[KA15c] Zeynab Mousavi Khamene and Mohammad J. Abdekhodaie. Diffusional release of a dispersed solute from a cylindrical polymeric matrix into an infinite external volume. Applied Mathematics and Computation, 259(??):676–685, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002635.

Kacewicz:2015:ESI

[Kac15] Bolesław Kacewicz. Efficient solution of IVPs with right-hand sides having discontinuities on an unknown hypersurface. Applied Mathematics and Computation, 265(??):469–485, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006566.

Karsli:2015:GGT

[KAG15] Harun Karsli, P. N.

Agrawal, and Meenu Goyal. General gamma type operators based on qintegers. Applied Mathematics and Computation, 251(??):564–575, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016178.

Khosravi:2015:ESF

[KAH15] Hassan Khosravi, Reza Allahyari, and Ali Shole Haghighi. Existence of solutions of functional integral equations of convolution type using a new construction of a measure of noncompactness on L^{p}(R_{+}). Applied Mathematics and Computation, 260(??):140–147, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003409.

Khalid:2019:NPQ

[KAI19] Nauman Khalid, Muhammad Abbas, and Muhammad Kashif Iqbal. Nonpolynomial quintic spline for solving fourth-order fractional boundary value problems involving product terms. Applied Mathematics and Computation, 349(??):393–407, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311275.

Kajla:2018:KVO

[Kaj18] Arun Kajla. The Kantorovich variant of an operator defined by D. D. Stancu. Applied Mathematics and Computation, 316(??):400–408, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305726.

Kpeky:2018:NLQ

[KAMD18] Fessal Kpeky, Farid Abed-Meraim, and El Mostafa Daya. New linear and quadratic prismatic piezoelectric solid-shell finite elements. Applied Mathematics and Computation, 319(??):355–368, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302345.

Kang:2015:LTB

[Kan15] Jum-Ran Kang. Longtime behavior of a suspension bridge equations with past history. Applied Mathematics and Computation, 265(??):509–519, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005871.

Kang:2016:AAN

[Kan16]  Jum-Ran Kang.  At-

tractors for autonomous and nonautonomous 3D Benjamin–Bona–Mahony equations. Applied Mathematics and Computation, 274(??):343–352, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501437X.

Kang:2019:ECA

[Kan19] Hongchao Kang. Efficient calculation and asymptotic expansions of many different oscillatory infinite integrals. Applied Mathematics and Computation, 346(??):305–318, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308531.

Karpuz:2019:HNT

[Kar19]  Basak Karpuz.  Hille–

Nehari theorems for dynamic equations with a time scale independent critical constant. Applied Mathematics and Computation, 346(??):336–351, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308348.

Kurkcu:2016:NAE

[KAS16] Ömür Kivanç Kürkçü, Ersin Aslan, and Mehmet Sezer. A numerical approach with error estimation to solve general integro-differential-difference equations using Dickson polynomials. Applied Mathematics and Computation, 276(??):324–339, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300138.

Katugampola:2015:MTG

[Kat15] Udita N. Katugampola.

Mellin transforms of generalized fractional integrals and derivatives. Applied Mathematics and Computation, 257(??):566–580, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017214.

Karasozen:2015:MOR

[KAU15] Bülent Karasözen, Canan Akkoyunlu, and Murat Uzunca. Model order reduction for nonlinear Schrödinger equation. Applied Mathematics and Computation, 258(??):509–519, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001538.

Kaluszka:2015:STI

[KB15] Marek Kaluszka and Michal Boczek. Steffensen type inequalities for fuzzy integrals. Applied Mathematics and Computation, 261(??):176–182, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004166.

Khajanchi:2017:RCP

[KB17] Subhas Khajanchi and Sandip Banerjee. Role of constant prey refuge on stage structure predator– prey model with ratio dependent functional response. Applied Mathematics and Computation, 314(??):193–198, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031730468X.

Kalayci:2018:RST

[KB18] Tahir Emre Kalayci and Roberto Battiti. A reactive self-tuning scheme for multilevel graph partitioning. Applied Mathematics and Computation, 318(??):227–244, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305829.

Ketabchi:2015:ALM

[KBK15] Saeed Ketabchi and Malihe Behboodi-Kahoo. Augmented Lagrangian method within L-shaped method for stochastic linear programs. Applied Mathematics and Computation, 266(??):12–20, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006232.

Karagiannakis:2016:TTC

[KBK^{+}16] N. P. Karagiannakis, G. C. Bourantas, A. N. Kalarakis, E. D. Skouras, and V. N. Burganos. Transient thermal conduction with variable conductivity using the Meshless Local Petrov–Galerkin method. Applied Mathematics and Computation, 272 (part 3)(??):676–686, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002842.

Kim:2016:HOE

[KBM16] Young Ik Kim, Ramandeep Behl, and S. S. Motsa. Higher-order efficient class of Chebyshev– Halley type methods. Applied Mathematics and Computation, 273(??): 1148–1159, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031501231X.

Kalpana:2015:DDD

[KBR15] M. Kalpana, P. Balasubramaniam, and K. Ratnavelu. Direct delay decomposition approach to synchronization of chaotic fuzzy cellular neural networks with discrete, unbounded distributed delays and Markovian jumping parameters. Applied Mathematics and Computation, 254(??):291–304, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017949.

Kadak:2017:SWS

[KBS17] Uğur Kadak, Naim L.

Braha, and H. M. Srivastava. Statistical weighted B-summability and its applications to approximation theorems. Applied Mathematics and Computation, 302(??):80–96, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730019X.

Kara:2018:CSS

[KÇ18a] R. Kara and M. Çaglar.

Çinlar subgrid scale model for large eddy simulation. Applied Mathematics and Computation, 322(??):89– 99, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308159.

Kumar:2018:IMF

[KC18b] Ashim Kumar and João R.

Cardoso. Iterative methods for finding commuting solutions of the Yang– Baxter-like matrix equation. Applied Mathematics and Computation, 333(??):246–253, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302595.

Kung:2018:OHP

[KC18c] Tzu-Liang Kung and Hon-Chan Chen. Optimizing Hamiltonian panconnectedness for the crossed cube architecture. Applied Mathematics and Computation, 331(??):287–296, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830167X.

Kaennakham:2019:ANA

[KC19] S. Kaennakham and N. Chuathong. An automatic node-adaptive scheme applied with a RBF-collocation meshless method. Applied Mathematics and Computation, 348(??):102–125, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310403.

Karaaslan:2016:ASB

[KCK16] Mehmet Fatih Karaaslan, Fatih Celiker, and Muhammet Kurulay. Approximate solution of the Bagley–Torvik equation by hybridizable discontinuous Galerkin methods. Applied Mathematics and Computation, 285(??):51–58, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302247.

Kholmetska:2018:NMG

[KCK18]  Ielizaveta  Kholmetska,

Jan Chleboun, and Pavel Krejcí. Numerical modeling of Galfenol magnetostrictive response. Applied Mathematics and Computation, 319(??): 527–537, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317303648.

Katiyar:2019:NCR

[KCK19] S. K. Katiyar, A. K. B Chand, and G. Saravana Kumar. A new class of rational cubic spline fractal interpolation function and its constrained aspects. Applied Mathematics and Computation, 346(??):319–335, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309020.

Korolev:2015:MHF

[KCKZ15] V. Yu. Korolev, A. V. Chertok, A. Yu. Korchagin, and A. I. Zeifman. Modeling high-frequency order flow imbalance by functional limit theorems for two-sided risk processes. Applied Mathematics and Computation, 253(??):224–241, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017391.

Kim:2019:ECC

[KCL19] Jung Eun Kim, Yongin Choi, and Chang Hyeong Lee. Effects of climate change on Plasmodium vivax malaria transmission dynamics: a mathematical modeling approach. Applied Mathematics and Computation, 347(??):616–630, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309743.

Kang:2018:SSA

[KCSD^{+}18] Bingyi Kang, Gyan Chhipi-Shrestha, Yong Deng, Kasun Hewage, and Rehan Sadiq. Stable strategies analysis based on the utility of Z-number in the evolutionary games. Applied Mathematics and Computation, 324(??):202–217, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308597.

Kang:2015:FPF

[KCWK15] Tong Kang, Tao Chen, Yanfang Wang, and Kwang Ik Kim. A T -ψ formulation with the penalty function term for the 3D eddy current problem in laminated structures. Applied Mathematics and Computation, 271(??):618–641, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501214X.

Kang:2015:FEM

[KCZK15] Tong Kang, Tao Chen, Huai Zhang, and Kwang Ik Kim. A − ϕ finite element method with composite grids for timedependent eddy current problem. Applied Mathematics and Computation, 267(??):365–381, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002775.

Korkmaz:2016:QQB

[KD16] Alper Korkmaz and Idris Dag. Quartic and quintic B-spline methods for advection-diffusion equation. Applied Mathematics and Computation, 274(??):208–219, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315014666.

Kheybari:2018:ETF

[KD18] S. Kheybari and M. T. Darvishi. An efficient technique to find semianalytical solutions for higher order multi-point boundary value problems. Applied Mathematics and Computation, 336(??):76–93, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304004.

Kumar:2019:NFP

[KDG19] Manoj Kumar and Varsha Daftardar-Gejji. A new family of predictorcorrector methods for solving fractional differential equations. Applied Mathematics and Computation, 363(??): Article 124633, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306253.

Kheybari:2019:NSS

[KDH19] Samad Kheybari, Mohammad Taghi Darvishi, and Mir Sajjad Hashemi. Numerical simulation for the space-fractional diffusion equations. Applied Mathematics and Computation, 348(??):57–69, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310154.

Kariotou:2016:AFA

[KDS16] Foteini Kariotou, Michael Doschoris, and Dimitris E. Sinikis. An algebraic formula for the accelerated computation of the low frequency scattering coefficients: the case of the acoustically soft sphere. Applied Mathematics and Computation, 275(??):13–23, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015428.

Kolasa:2018:EMI

[KDTP18] Marta Kolasa, Rafal Dlugosz, Tomasz Talaśka, and Witold Pedrycz. Efficient methods of initializing neuron weights in self-organizing networks implemented in hardware. Applied Mathematics and Computation, 319(??):31–47, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300619.

Kazemi:2016:ESS

[KE16] M. Kazemi and R. Ezzati. Existence of solution for some nonlinear two-dimensional Volterra integral equations via measures of noncompactness. Applied Mathematics and Computation, 275(??):165–171, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015684.

Korkut:2018:PPS

[KE18] Sila Ö. Korkut and Utku Erdogan. Positivity preserving scheme based on exponential integrators. Applied Mathematics and Computation, 320(??):731–739, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307269.

Kose:2017:IAB

[KEE17] Erkan Kose, Mehmet Erbas, and Erkan Ersen. An integrated approach based on game theory and geographical information systems to solve decision problems. Applied Mathematics and Computation, 308(??):105–114, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301947.

Kemm:2016:RTS

[Kem16] Friedemann Kemm. Roetype schemes for shallow water magnetohydrodynamics with hyperbolic divergence cleaning. Applied Mathematics and Computation, 272 (part 2)(??):385–402, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007031.

Kemm:2018:HNC

[Kem18] Friedemann Kemm. Heuristical and numerical considerations for the carbuncle phenomenon. Applied Mathematics and Computation, 320(??):596–613, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306355.

Kycia:2015:GEF

[KF15] Radosław Antoni Kycia and Galina Filipuk. On the generalized Emden–Fowler and isothermal spheres equations. Applied Mathematics and Computation, 265(??):1003–1010, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007717.

Kovacs:2019:NLS

[KF19] Péter Kovács and Andrea M. Fekete. Nonlinear least-squares spline fitting with variable knots.

Applied Mathematics and Computation, 354(??): 490–501, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319301559.

Kosik:2015:NSI

[KFHH15] Adam Kosík, Miloslav Feistauer, Martin Hadrava, and Jaromír Horácek. Numerical simulation of the interaction between a nonlinear elastic structure and compressible flow by the discontinuous Galerkin method. Applied Mathematics and Computation, 267(??):382–396, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002453.

Khajanchi:2015:CEO

[KG15] Subhas Khajanchi and Dibakar Ghosh. The combined effects of optimal control in cancer remission. Applied Mathematics and Computation, 271(??):375–388, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315012308.

Kopriva:2016:GEN

[KG16] David A. Kopriva and Gregor J. Gassner. Geometry effects in nodal discontinuous Galerkin methods on curved elements that are provably stable. Applied Mathematics and Computation, 272 (part 2)(??):274–290, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315011029.

Kundu:2017:AWS

[KG17] Bidisha Kundu and Ranjan Ganguli. Analysis of weak solution of Euler–Bernoulli beam with axial force. Applied Mathematics and Computation, 298(??):247–260, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306877.

Kostic:2018:GTL

[KG18] V. Kostić and D. Gardasević. On the Gersgorintype localizations for nonlinear eigenvalue problems. Applied Mathematics and Computation, 337(??):158–178, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304211.

Khan:2019:PCE

[KG19] Feroz Khan and Siqing Gan. Pathwise convergence of an efficient scheme for SPDEs with non-globally Lipschitz nonlinearity. Applied Mathematics and Computation, 353(??):114–133, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301146.

Atencia-McKillop:2018:GXG

[KGGAV^{+}18] Ivan Atencia-Mc. Killop, José L. Galán-García, Gabriel Aguilera-Venegas, Pedro Rodríguez-Cielos, and Mángeles Galán-García. A Geo[X] / G[X] / 1 retrial queueing system with removal work and total renewal discipline. Applied Mathematics and Computation, 319(??):245–253, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300317301364.

Kapela:2015:EPL

[KGS^{+}15] Rafal Kapela, Karol Gugala, Pawel Sniatala, Aleksandra Swietlicka, and Krzysztof Kolanowski. Embedded platform for local image descriptor based object detection. Applied Mathematics and Computation, 267(??):419–426, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031500209X.

Kurina:2017:FAA

[KH17] Galina Kurina and Nguyen Thi Hoai. First asymptotic approximations to a solution of singularly perturbed optimal control problem with intersecting solutions of degenerate problem. Applied Mathematics and Computation, 292(??):356–374, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304763.

Khajanchi:2017:MDS

[Kha17] Subhas Khajanchi. Modeling the dynamics of stage-structure predator– prey system with Monod– Haldane type response function. Applied Mathematics and Computation, 302(??):122–143, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300279.

Khleborodov:2018:FEC

[Khl18] Denis Khleborodov. Fast elliptic curve point multiplication based on window Non-Adjacent Form method. Applied Mathematics and Computation, 334(??):41–59, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302935.

Khan:2019:ESH

[KHN19] A. U. Khan, S. T. Hussain, and S. Nadeem. Existence and stability of heat and fluid flow in the presence of nanoparticles along a curved surface by mean of dual nature solution. Applied Mathematics and Computation, 353(??):66– 81, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300591.

Kilic:2015:FIM

[KI15] Bulent Kilic and Mustafa Inc. The first integral method for the time fractional Kaup–Boussinesq System with time dependent coefficient. Applied Mathematics and Computation, 254(??):70–74, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031401755X.

Kiani:2015:NSB

[Kia15] Keivan Kiani. Nanomechanical sensors based on elastically supported double-walled carbon nanotubes. Applied Mathematics and Computation, 270(??):216–241, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501036X.

Kajla:2016:IBS

[KIAG16] Arun Kajla, Nurhayat Ispir, P. N. Agrawal, and Meenu Goyal. q Bernstein– Schurer–Durrmeyer type operators for functions of one and two variables. Applied Mathematics and Computation, 275(??):372–385, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015507.

Kim:2015:TAO

[Kim15a] Min Chan Kim. Theoretical analysis on the onset of buoyancy-driven instability of horizontal interfaces between miscible fluids in a Hele–Shaw cell. Applied Mathematics and Computation, 265(??):207–224, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005986.

Kim:2015:FRS

[Kim15b] Sung Hyun Kim. Further results on stability analysis of discretetime systems with timevarying delays via the use of novel convex combination coefficients. Applied Mathematics and Computation, 261(??):104–113, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315004002.

Kim:2015:BTM

[Kim15c] Tae Kyun Kim. Barnes’ type multiple degenerate Bernoulli and Euler polynomials. Applied Mathematics and Computation, 258(??):556–564, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002209.

Kim:2018:MMI

[Kim18] Myles Kim. Mechanism of MDCK II cell polarization during the cell division: a computational study. Applied Mathematics and Computation, 317(??):1–11, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306124.

Kim:2019:GRT

[Kim19] Sung Hyun Kim. Generalized relaxation techniques for robust H_{∞} filtering of nonhomogeneous Markovian jump systems. Applied Mathematics and Computation, 347(??):542–556, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309494.

Khan:2018:MCC

[KIZ18] Muhammad Altaf Khan, Saeed Islam, and Gul Zaman. Media coverage campaign in Hepatitis B transmission model. Applied Mathematics and Computation, 331(??):378–393, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302029.

Karimi:2015:NIM

[KJ15a] Saeed Karimi and Meisam Jozi. A new iterative method for solving linear Fredholm integral equations using the least squares method. Applied Mathematics and Computation, 250(??):744–758, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015161.

Kim:2015:NSA

[KJ15b] Kyunghoon Kim and Bongsoo Jang. A novel semi-analytical approach for solving nonlinear Volterra integro-differential equations. Applied Mathematics and Computation, 263(??):25–35, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004622.

Klavzar:2018:GTL

[KJR^{+}18] Sandi Klavzar, D. Azubha Jemilet, Indra Rajasingh, Paul Manuel, and N. Parthiban. General Transmission Lemma and Wiener complexity of triangular grids. Applied Mathematics and Computation, 338(??):101–114, December 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318304739.

Kim:2015:MAS

[KK15a] Dae San Kim and Taekyun Kim. A matrix approach to some identities involving Sheffer polynomial sequences. Applied Mathematics and Computation, 253(??):83–101, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300314017020.

Koudela:2015:HSR

[KK15b] Lukás Koudela and Václav Kotlan. High-speed rotation induction heating in thermal clamping technology. Applied Mathematics and Computation, 267(??):445–455, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001393.

Kruis:2015:EMD

[KK15c] Jaroslav Kruis and Tomás Koudelka. Effect of mechanical damage on moisture transport in concrete. Applied Mathematics and Computation, 267(??):139–147, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005858.

Kim:2016:DUM

[KK16a] Young-Jin Kim and Hyun-Min Kim. Diagonal update method for a quadratic matrix equation. Applied Mathematics and Computation, 283(??):208–215, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301199.

Krause:2016:ELT

[KK16b] D. Krause and R. Krause. Enabling local time stepping in the parallel implicit solution of reactiondiffusion equations via space–time finite elements on shallow tree meshes. Applied Mathematics and Computation, 277(??):164–179, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300059.

Kumar:2017:DDT

[KK17] Sunil Kumar and B. V. Rathish Kumar. A domain decomposition Taylor Galerkin finite element approximation of a parabolic singularly perturbed differential equation. Applied Mathematics and Computation, 293(??):508–522, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305264.

Khan:2015:RJI

[KKAK15] M. Adil Khan, G. Ali Khan, T. Ali, and A. Kilicman. On the refinement of Jensen’s inequality. Applied Mathematics and Computation, 262(??):128–135, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004634.

Kansal:2015:NMH

[KKB15] Munish Kansal, V. Kanwar, and Saurabh Bhatia. New modifications of Hansen–Patrick’s family with optimal fourth and eighth orders of convergence. Applied Mathematics and Computation, 269(??):507–519, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010231.

Kim:2015:NDB

[KKD15] Dae San Kim, Taekyun Kim, and Dmitry V. Dolgy. A note on degenerate Bernoulli numbers and polynomials associated with p-adic invariant integral on Z_{p}. Applied Mathematics and Computation, 259(??):198–204, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002684.

Kim:2017:AUB

[KKD17] Chesoong Kim, V. I. Klimenok, and A. N. Dudin. Analysis of unreliable BMAP/PH/N type queue with Markovian flow of breakdowns. Applied Mathematics and Computation, 314(??):154–172, December 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317304484.

Kundu:2018:TDG

[KKG18] Santimoy Kundu, Alka Kumari, and Shishir Gupta. Three-dimensional Green’s function approach for analysis of dispersion and attenuation curve in fibre-reinforced heterogeneous viscoelastic layer due to a point source. Applied Mathematics and Computation, 338(??):376–386, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318303230.

Karban:2018:BAS

[KKKD18] Pavel Karban, Petr Kropík, Václav Kotlan, and Ivo Dolezel. Bayes approach to solving T.E.A.M. benchmark problems 22 and 25 and its comparison with other optimization techniques. Applied Mathematics and Computation, 319(??):681–692, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305040.

Kwon:2018:NLK

[KKL18] W. Kwon, Baeyoung Koo, and S. M. Lee. Novel Lyapunov–Krasovskii functional with delay-dependent matrix for stability of time-varying delay systems. Applied Mathematics and Computation, 320(??):149–157, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306641.

Koudelka:2015:CSD

[KKM15] Tomás Koudelka, Jaroslav Kruis, and Jirí Madera. Coupled shrinkage and damage analysis of autoclaved aerated concrete. Applied Mathematics and Computation, 267(??):427–435, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001691.

Kim:2016:DML

[KKMS16] Dae San Kim, Taekyun Kim, Toufik Mansour, and Jong-Jin Seo. Degenerate Mittag-Leffler polynomials. Applied Mathematics and Computation, 274(??):258–266, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014769.

Kolbe:2016:STD

[KKS^{+}16] Niklas Kolbe, Jana Kat’uchová, Nikolaos Sfakianakis, Nadja Hellmann, and Mária Lukácová-Medvid’ová. A study on time discretization and adaptive mesh refinement methods for the simulation of cancer invasion: the urokinase model. Applied Mathematics and Computation, 273(??):353–376, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010784.

Knor:2016:SMV

[KKST16] M. Knor, J. Kranjc, Riste Skrekovski, and Aleksandra Tepeh. A search for the minimum value of Balaban index. Applied Mathematics and Computation, 286(??):301–310, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302739.

Knor:2017:MVS

[KKST17] Martin Knor, Jaka Kranjc, Riste Skrekovski, and Aleksandra Tepeh. On the minimum value of sum-Balaban index. Applied Mathematics and Computation, 303(??):203–210, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300553.

Kadalbajoo:2015:RBF

[KKT15] Mohan K. Kadalbajoo, Alpesh Kumar, and Lok Pati Tripathi. A radial basis functions based finite differences method for wave equation with an integral condition. Applied Mathematics and Computation, 253(??):8–16, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017500.

Khmelnytskaya:2019:RTK

[KKT19] Kira V. Khmelnytskaya, Vladislav V. Kravchenko, and Sergii M. Torba. A representation of the transmutation kernels for the Schrödinger operator in terms of eigenfunctions and applications. Applied Mathematics and Computation, 353(??):274–281, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301286.

Kelenc:2017:MMD

[KKTY17] Aleksander Kelenc, Dorota Kuziak, Andrej Taranenko, and Ismael G. Yero. Mixed metric dimension of graphs. Applied Mathematics and Computation, 314(??):429–438, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317304782.

Kim:2015:EPS

[KL15] Kwang Ik Kim and Zuhan Liu. Extinction phenomenon for Spinor Ginzburg– Landau equations in three dimensions. Applied Mathematics and Computation, 256(??):786–807, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500096X.

Kong:2016:RLO

[KL16] Haiyue Kong and Yeping Li. Relaxation limit of the one-dimensional bipolar Euler–Poisson system in the bound domain. Applied Mathematics and Computation, 274(??):1–13, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014381.

Khan:2018:QAD

[KLKU18] Wakeel Khan, Yan Lin, Sarmad Ullah Khan, and Nasim Ullah. Quantized adaptive decentralized control for interconnected nonlinear systems with actuator faults. Applied Mathematics and Computation, 320(??):175–189, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306318.

Kang:2018:SRT

[KLQY18] Liying Kang, Lele Liu, Liqun Qi, and Xiying Yuan. Spectral radii of two kinds of uniform hypergraphs. Applied Mathematics and Computation, 338(??):631–659, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305034.

Kong:2015:RCD

[KLS15] Yuedong Kong, Xuezhu Lu, and Yanling Shi. On reducibility for a class of n-dimensional quasiperiodic systems with a small parameter. Applied Mathematics and Computation, 264(??):272–278, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005706.

Ke:2015:OAP

[KLW15] Jau-Chuan Ke, Tzu-Hsin Liu, and Chia-Huang Wu. An optimum approach of profit analysis on the machine repair system with heterogeneous repairmen. Applied Mathematics and Computation, 253(??):40–51, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016804.

Kang:2019:NAB

[KLZ19] Ting Kang, Qiang Li, and Qimin Zhang. Numerical analysis of the balanced implicit method for stochastic age-dependent capital system with Poisson jumps. Applied Mathematics and Computation, 353(??):166–177, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309202.

Kuo:2015:SBL

[KLZT15] R. J. Kuo, Y. H. Lee, Ferani E. Zulvia, and F. C. Tien. Solving bi-level linear programming problem through hybrid of immune genetic algorithm and particle swarm optimization algorithm. Applied Mathematics and Computation, 266 (??):1013–1026, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008061.

Kaveh:2015:EFE

[KM15a] A. Kaveh and M. S. Massoudi. Efficient finite element analysis using graph-theoretical force method; rectangular plane stress and plane strain Lagrange family elements. Applied Mathematics and Computation, 266(??):72–94, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006888.

Ke:2015:ADM

[KM15b] Yi-Fen Ke and Chang-Feng Ma. Alternating direction method for generalized Sylvester matrix equation AXB + CY D = E. Applied Mathematics and Computation, 260(??):106–125, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003665.

Keleshteri:2015:SAP

[KM15c] Marzieh Eini Keleshteri and Nazim I. Mahmudov. A study on q-Appell polynomials from determinantal point of view. Applied Mathematics and Computation, 260(??):351–369, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003112.

Kuniya:2015:GSM

[KM15d] Toshikazu Kuniya and Yoshiaki Muroya. Global stability of a multi-group SIS epidemic model with varying total population size. Applied Mathematics and Computation, 265(??):785–798, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007559.

Ke:2017:DSI

[KM17a] Yi-Fen Ke and Chang-Feng Ma. The dimensional splitting iteration methods for solving saddle point problems arising from time-harmonic eddy current models. Applied Mathematics and Computation, 303(??):146–164, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300516.

Ke:2017:SLI

[KM17b] Yi-Fen Ke and Chang-Feng Ma. SOR-like iteration method for solving absolute value equations. Applied Mathematics and Computation, 311(??):195–202, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303363.

Khatri:2018:GSM

[KM18a] Kejal Khatri and Vishnu Narayan Mishra. Generalized Szász–Mirakyan operators involving Brenke type polynomials. Applied Mathematics and Computation, 324(??):228–238, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308317.

Kvasov:2018:MVD

[KM18b] Dmitri E. Kvasov and Marat S. Mukhametzhanov. Metaheuristic vs. deterministic global optimization algorithms: the univariate case. Applied Mathematics and Computation, 318(??):245–259, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303028.

Katsikis:2019:HPE

[KM19] Vasilios N. Katsikis and Spyridon D. Mourtas. A heuristic process on the existence of positive bases with applications to minimum-cost portfolio insurance in C[a, b]. Applied Mathematics and Computation, 349(??):221–244, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311056.

Koorepazan-Moftakhar:2016:CDS

[KMA16] F. Koorepazan-Moftakhar and A. R. Ashrafi. Combination of distance and symmetry in some molecular graphs. Applied Mathematics and Computation, 281(??):223–232, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300741.

Krajnovic:2016:PAN

[KMB16] Sinisa Krajnović, Guglielmo Minelli, and Branislav Basara. Partially-averaged Navier–Stokes simulations of two bluff body flows. Applied Mathematics and Computation, 272 (part 3) (??):692–706, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315004555.

Khaleghi:2018:SCS

[KMBA18] M. Khaleghi, M. Talebi Moghaddam, E. Babolian, and S. Abbasbandy. Solving a class of singular two-point boundary value problems using new effective reproducing kernel technique. Applied Mathematics and Computation, 331(??):264–273, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301887.

Kong:2019:OBQ

[KML19] Chuifeng Kong, Yuechao Ma, and Deyou Liu. Observer-based quantized sliding mode dissipative control for singular semi-Markovian jump systems. Applied Mathematics and Computation, 362(??): Article 124539, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305181.

Kuraz:2015:SNN

[KMP15] Michal Kuraz, Petr Mayer, and Pavel Pech. Solving the nonlinear and nonstationary Richards equation with two-level adaptive domain decomposition (ddadaptivity). Applied Mathematics and Computation, 267(??):207–222, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500449X.

Kaluza:2019:OAS

[KMP19] Andrzej Kaluza, Pawel M.

Morkisz, and Pawel Przybylowicz. Optimal approximation of stochastic integrals in analytic noise model. Applied Mathematics and Computation, 356(??):74–91, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302231.

Kudryashov:2018:CBS

[KMR18] N. A. Kudryashov, R. V. Muratov, and P. N. Ryabov. The collective behavior of shear strain localizations in dipolar materials. Applied Mathematics and Computation, 338(??):152–163, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304934.

Knor:2018:GPW

[KMS18] Martin Knor, Snjezana Majstorović, and Riste Skrekovski. Graphs preserving Wiener index upon vertex removal. Applied Mathematics and Computation, 338(??):12–24, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304648.

Kravchenko:2016:LTA

[KMT16] Vladislav V. Kravchenko, Samy Morelos, and Sergii M. Torba. Liouville transformation, analytic approximation of transmutation operators and solution of spectral problems. Applied Mathematics and Computation, 273(??):321–336, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013430.

Kravcenko:2019:DFB

[KMZ19] Michal Kravcenko, Michal Merta, and Jan Zapletal. Distributed fast boundary element methods for Helmholtz problems. Applied Mathematics and Computation, 362 (??):Article 124503, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304825.

Khajanchi:2019:MMT

[KN19] Subhas Khajanchi and Juan J. Nieto. Mathematical modeling of tumorimmune competitive system, considering the role of time delay. Applied Mathematics and Computation, 340(??):180–205, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830732X.

Khan:2015:NFO

[KNBA15] Waseem Asghar Khan, Khalida Inayat Noor, Kaleemulah Bhatti, and Faryal Aijaz Ansari. A new fourth order Newtontype method for solution of system of nonlinear equations. Applied Mathematics and Computation, 270(??):724–730, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011376.

Katta:2015:HOF

[KNK15] Kiran K. Katta, Ramachandran D. Nair, and Vinod Kumar. High-order finite volume shallow water model on the cubedsphere: 1D reconstruction scheme. Applied Mathematics and Computation, 266(??):316–327, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315005111.

Kravchenko:2017:RSO

[KNT17] Vladislav V. Kravchenko, Luis J. Navarro, and Sergii M. Torba. Representation of solutions to the one-dimensional Schrödinger equation in terms of Neumann series of Bessel functions. Applied Mathematics and Computation, 314(??):173–192, December 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317304575.

Koca:2015:MSD

[Koc15] Ilknur Koca. A method for solving differential equations of q-fractional order. Applied Mathematics and Computation, 266(??):1–5, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006669.

Koko:2015:MMG

[Kok15] Jonas Koko. A Matlab mesh generator for the two-dimensional finite element method. Applied Mathematics and Computation, 250(??):650–664, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015148.

Korkmaz:2018:SSN

[Kor18] Alper Korkmaz. Stability satisfied numerical approximates to the nonanalytical solutions of the cubic Schrödinger equation. Applied Mathematics and Computation, 331(??):210–231, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830208X.

Keshavarz:2019:BWO

[KOR19] E. Keshavarz, Y. Ordokhani, and M. Razzaghi. The Bernoulli wavelets operational matrix of integration and its applications for the solution of linear and nonlinear problems in calculus of variations. Applied Mathematics and Computation, 351(??):83–98, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310841.

Kountouriotis:2016:DLN

[KPG16] Zacharias Kountouriotis, Maria Philippou, and Georgios C. Georgiou. Development lengths in Newtonian Poiseuille flows with wall slip. Applied Mathematics and Computation, 291(??):98–114, December 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316304210.

Kuo:2015:IAR

[KPLC15] R. J. Kuo, C. M. Pai, R. H. Lin, and H. C. Chu. The integration of association rule mining and artificial immune network for supplier selection and order quantity allocation. Applied Mathematics and Computation, 250(??):958–972, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015215.

Kwon:2017:CSM

[KPP17] Nam Kyu Kwon, In Seok Park, and PooGyeon Park. H_{∞} control for singular Markovian jump systems with incomplete knowledge of transition probabilities. Applied Mathematics and Computation, 295(??):126–135, February 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305641.

Krawiec:2018:QDC

[KPP18] Michal Krawiec, Zbigniew Palmowski, and Lukasz Plociniczak. Quickest drift change detection in Lévytype force of mortality model. Applied Mathematics and Computation, 338(??):421–431, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305277.

Kolk:2016:STS

[KPT16] Marek Kolk, Arvet Pedas, and Enn Tamme. Smoothing transformation and spline collocation for linear fractional boundary value problems. Applied Mathematics and Computation, 283(??):234–250, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301552.

Khan:2015:DAC

[KR15] Subuhi Khan and Mumtaz Riyasat. Determinantal approach to certain mixed special polynomials related to Gould-Hopper polynomials. Applied Mathematics and Computation, 251(??):599–614, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016130.

Kudryashov:2016:ANS

[KR16] Nikolay A. Kudryashov and Pavel N. Ryabov. Analytical and numerical solutions of the generalized dispersive Swift– Hohenberg equation. Applied Mathematics and Computation, 286(??): 171–177, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316302740.

Kravchenko:2018:CTO

[Kra18a] Vladislav V. Kravchenko.

Construction of a transmutation for the onedimensional Schrödinger operator and a representation for solutions. Applied Mathematics and Computation, 328(??):75–81, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300584.

Kumar:2018:RFT

[KRA^{+}18b] S. Vimal Kumar, R. Raja, S. Marshal Anthoni, Jinde Cao, and Zhengwen Tu. Robust finite-time nonfragile sampled-data control for T–S fuzzy flexible spacecraft model with stochastic actuator faults. Applied Mathematics and Computation, 321(??):483–497, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307749.

Kara:2015:SCL

[KRB15] A. H. Kara, Polina Razborova, and Anjan Biswas. Solitons and conservation laws of coupled Ostrovsky equation for internal waves. Applied Mathematics and Computation, 258(??):95– 99, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001253.

Karami:2017:NSD

[KRF17] Naser Karami, Mohammad Hassan Rahimian, and Mohsen Farhadzadeh. Numerical simulation of droplet evaporation on a hot surface near Leidenfrost regime using multiphase lattice Boltzmann method. Applied Mathematics and Computation, 312(??):91–108, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303399.

Krivovichev:2018:LBG

[Kri18] Gerasim V. Krivovichev.

Linear Bhatnagar–Gross– Krook equations for simulation of linear diffusion equation by lattice Boltzmann method. Applied Mathematics and Computation, 325(??):102–119, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730886X.

Krivoshein:2019:FLW

[Kri19a] A. V. Krivoshein. From frame-like wavelets to wavelet frames keeping approximation properties and symmetry. Applied Mathematics and Computation, 344–345(??):204– 218, March 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318308646.

Krivovichev:2019:SAB

[Kri19b] Gerasim V. Krivovichev. Stability analysis of body force action models used in the single-relaxation-time single-phase lattice Boltzmann method. Applied Mathematics and Computation, 348(??):25–41, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310300.

Katta:2018:CCL

[KRS18] Ravinder Katta, G. D. Reddy, and N. Sukavanam. Computation of control for linear approximately controllable system using weighted Tikhonov regularization. Applied Mathematics and Computation, 317(??):252–263, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305635.

Koczkodaj:2015:PCS

[KS15] W. W. Koczkodaj and J. Szybowski. Pairwise comparisons simplified. Applied Mathematics and Computation, 253(??):387–394, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017238.

Kashkari:2016:FOL

[KS16a] Bothayna S. H. Kashkari and Muhammed I. Syam. Fractional-order Legendre operational matrix of fractional integration for solving the Riccati equation with fractional order. Applied Mathematics and Computation, 290(??):281–291, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303745.

Khader:2016:SPB

[KS16b] M. M. Khader and N. H.

Sweilam. Singularly perturbed BVP to estimation of diaphragm deflection in MEMS capacitive microphone: an application of ADM. Applied Mathematics and Computation, 281(??):214–222, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300273.

Kanat:2018:SAR

[KS18] K. Kanat and M. Sofyalioglu. Some approximation results for Stancu type Lupas–Schurer operators based on (p, q)- integers. Applied Mathematics and Computation, 317(??):129–142, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305970.

Knor:2019:MDV

[KS19a] Martin Knor and Riste Skrekovski. On the minimum distance in a k-vertex set in a graph. Applied Mathematics and Computation, 356(??):99–104, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302577.

Kumar:2019:RFO

[KS19b] Rajneesh Kumar and Poonam Sharma. Response of fractional order on energy ratios at the boundary surface of fluid–piezothermoelastic media. Applied Mathematics and Computation, 358(??):194–203, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303029.

Kumar:2019:HOW

[KS19c] Santosh Kumar and Paramjeet Singh. High order WENO finite volume approximation for population density neuron model. Applied Mathematics and Computation, 356(??):173–189, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302218.

Kuhnel:2019:DGS

[KSA19] Line Kühnel, Stefan Sommer, and Alexis Arnaudon. Differential geometry and stochastic dynamics with deep learning numerics. Applied Mathematics and Computation, 356(??):411–437, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302516.

Kavitha:2015:CPE

[KSD^{+}15] L. Kavitha, B. Srividya, S. Dhamayanthi, V. Senthil Kumar, and D. Gopi. Collision and propagation of electromagnetic solitons in an antiferromagnetic spin ladder medium. Applied Mathematics and Computation, 251(??):643–668, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031401621X.

Kisi:2015:SWL

[KSK^{+}15] Ozgur Kisi, Jalal Shiri, Sepideh Karimi, Shahaboddin Shamshirband, Shervin Motamedi, Dalibor Petković, and Roslan Hashim. A survey of water level fluctuation predicting in urmia lake using support vector machine with firefly algorithm. Applied Mathematics and Computation, 270(??):731–743, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0096300315011418.

Kolanowski:2018:MDF

[KŚK^{+}18]  Krzysztof  Kolanowski,

Aleksandra Świetlicka, Rafal Kapela, Janusz Pochmara, and Andrzej Rybarczyk. Multisensor data fusion using Elman neural networks. Applied Mathematics and Computation, 319(??):236–244, February 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317301352.

Karimi-Sibaki:2019:MET

[KSKW^{+}19] E. Karimi-Sibaki, A. Kharicha, M. Wu, A. Ludwig, and J. Bohacek. Modeling electrochemical transport of ions in the molten CaF_{2}– FeO slag operating under a DC voltage. Applied Mathematics and Computation, 357(??):357–373, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300213.

Kim:2019:NMM

[KSL19] Jin-Gyun Kim, Jaho Seo, and Jae Hyuk Lim. Novel modal methods for transient analysis with a reduced order model based on enhanced Craig–Bampton formulation. Applied Mathematics and Computation, 344–345(??):30–45, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308579.

Karthick:2019:DRF

[KSM^{+}19] S. A. Karthick, R. Sakthivel, Y. K. Ma, S. Mohanapriya, and A. Leelamani. Disturbance rejection of fractional-order T-S fuzzy neural networks based on quantized dynamic output feedback controller. Applied Mathematics and Computation, 361(??):846–857, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304941.

Knupp:2015:DIA

[KSN15] Diego C. Knupp, Wagner F. Sacco, and Antônio J. Silva Neto. Direct and inverse analysis of diffusive logistic population evolution with time delay and impulsive culling via integral transforms and hybrid optimization. Applied Mathematics and Computation, 250(??):105–120, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014350.

Kostic:2016:HFR

[KSP16] Srdan Kostić, Milan Stojković, and Stevan Prohaska. Hydrological flow rate estimation using artificial neural networks: Model development and potential applications. Applied Mathematics and Computation, 291(??):373–385, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304441.

Kucukoglu:2019:MBP

[KSS19] Irem Kucukoğlu, Buket Simsek, and Yilmaz Simsek. Multidimensional Bernstein polynomials and Bézier curves: Analysis of machine learning algorithm for facial expression recognition based on curvature. Applied Mathematics and Computation, 344–345(??):150–162, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308725.

Knor:2015:IBE

[KST15] Martin Knor, Riste Skrekovski, and Aleksandra Tepeh. An inequality between the edge-Wiener index and the Wiener index of a graph. Applied Mathematics and Computation, 269(??):714–721, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009698.

Knor:2016:DLM

[KST16a] Martin Knor, Riste Skrekovski, and Aleksandra Tepeh. Digraphs with large maximum Wiener index. Applied Mathematics and Computation, 284(??): 260–267, July 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316301990.

Knor:2016:SRW

[KST16b] Martin Knor, Riste Skrekovski, and Aleksandra Tepeh. Some remarks on Wiener index of oriented graphs. Applied Mathematics and Computation, 273(??):

631–636, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315013752.

Knor:2019:TMV

[KST19] Martin Knor, Riste Skrekovski, and Aleksandra Tepeh. Trees with the maximal value of Graovac–Pisanski index. Applied Mathematics and Computation, 358(??):287–292, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303224.

Kudryashov:2015:ETD

[KSV15] Nikolay A. Kudryashov, Dmitry I. Sinelshchikov, and Alexandr K. Volkov. Extended two dimensional equation for the description of nonlinear waves in gas-liquid mixture. Applied Mathematics and Computation, 268(??):581–589, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008760.

Kuila:2015:RAR

[KSZ15] Sahadeb Kuila, T. Raja Sekhar, and D. Zeidan. A robust and accurate Riemann solver for a compressible two-phase flow model. Applied Mathematics and Computation, 265(??):681–695, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007109.

Kim:2016:ACZ

[KT16] Jong Kyu Kim and Truong Minh Tuyen. Approximation common zero of two accretive operators in Banach spaces. Applied Mathematics and Computation, 283(??):265–281, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301412.

Karaisa:2015:STG

[KTA15] Ali Karaisa, Durhasan Turgut Tollu, and Yasin Asar. Stancu type generalization of q-Favard–Szàsz operators. Applied Mathematics and Computation, 264(??):249–257, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005639.

Kolasa:2015:NRA

[KTD15] Marta Kolasa, Tomasz Talaska, and Rafal Dlugosz. A novel recursive algorithm used to model hardware programmable neighborhood mechanism of self-organizing neural networks. Applied Mathematics and Computation, 267(??):314–328, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003823.

Kudryashov:2016:SGM

[Kud16] Nikolay A. Kudryashov.

On solutions of generalized modified Korteweg– de Vries equation of the fifth order with dissipation. Applied Mathematics and Computation, 280(??):39–45, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300327.

Kudryashov:2019:ESE

[Kud19a] Nikolay A. Kudryashov.

Exact solutions of the equation for surface waves in a convecting fluid. Applied Mathematics and Computation, 344–345 (??):97–106, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318308658.

Kudryashov:2019:LPF

[Kud19b] Nikolay A. Kudryashov. Lax pair and first integrals of the traveling wave reduction for the KdV hierarchy. Applied Mathematics and Computation, 350(??):323–330, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300438.

Kuehn:2015:EGN

[Kue15] Christian Kuehn. Efficient gluing of numerical continuation and a multiple solution method for elliptic PDEs. Applied Mathematics and Computation, 266(??):656–674, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007511.

Kundu:2016:EAP

[Kun16] Balaram Kundu. Exact analysis for propagation of heat in a biological tissue subject to different surface conditions for therapeutic applications. Applied Mathematics and Computation, 285(??):204–216, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302387.

Kurata:2018:STC

[Kur18] Hiroshi Kurata. Some theorems on the core inverse of matrices and the core partial ordering. Applied Mathematics and Computation, 316(??):43–51, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730543X.

Kurokawa:2019:HMC

[Kur19] Shun Kurokawa. How memory cost, switching cost, and payoff nonlinearity affect the evolution of persistence. Applied Mathematics and Computation, 341(??):174–192, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318307641.

Kostic:2015:NAG

[KVS15] Srdan Kostić, Nebojsa Vasović, and Dusko Sunarić. A new approach to grid search method in slope stability analysis using Box–Behnken statistical design. Applied Mathematics and Computation, 256(??):425–437, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000363.

Korkut:2016:FPB

[KVZ16] L. Korkut, D. Vlah, and V. Zupanović. Fractal properties of Bessel functions. Applied Mathematics and Computation, 283(??):55–69, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301369.

Khan:2015:GFG

[KwC15] Suhel Ahmad Khan and Jia wei Chen. Gap function and global error bounds for generalized mixed quasi variational inequalities. Applied Mathematics and Computation, 260(??):71–81, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003707.

Karanik:2016:RMA

[KWGRP16] Marcelo Karanik, Leonardo Wanderer, Jose Antonio Gomez-Ruiz, and Jose Ignacio Pelaez. Reconstruction methods for AHP pairwise matrices: How reliable are they? Applied Mathematics and Computation, 279(??):103–124, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630008X.

Kao:2017:ECS

[KWH17] Han-Ying Kao, Dong-Jyun Wu, and Chia-Hui Huang. Evaluation of cloud service industry with dynamic and network DEA models. Applied Mathematics and Computation, 315(??):188–202, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300317305209.

Kokubo:2015:SRD

[KWT15] Satoshi Kokubo, Zhen Wang, and Jun Tanimoto. Spatial reciprocity for discrete, continuous and mixed strategy setups. Applied Mathematics and Computation, 259(??):552–568, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003124.

Kong:2019:HNM

[KXZ19] Desong Kong, Yufeng Xu, and Zhoushun Zheng. A hybrid numerical method for the KdV equation by finite difference and sinc collocation method. Applied Mathematics and Computation, 355(??):61–72, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301353.

Kyrchei:2015:DRW

[Kyr15] Ivan Kyrchei. Determinantal representations of the W -weighted Drazin inverse over the quaternion skew field. Applied Mathematics and Computation, 264(??):453–465, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005962.

Kyrchei:2017:WSV

[Kyr17] Ivan Kyrchei. Weighted singular value decomposition and determinantal representations of the quaternion weighted Moore–Penrose inverse. Applied Mathematics and Computation, 309(??):1– 16, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317302321.

Kudryashov:2015:APE

[KZ15] Nikolay A. Kudryashov and Anastasia S. Zakharchenko. Analytical properties and exact solutions of the Lotka– Volterra competition system. Applied Mathematics and Computation, 254(??):219–228, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017743.

Karakoc:2016:SWS

[KZ16] S. Battal Gazi Karakoç and Halil Zeybek. Solitarywave solutions of the GRLW equation using septic B-spline collocation method. Applied Mathematics and Computation, 289(??):159–171, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303307.

Koslicki:2019:IMC

[KZ19] David Koslicki and Hooman Zabeti. Improving Min-Hash via the containment index with applications to metagenomic analysis. Applied Mathematics and Computation, 354(??):206–215, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031930116X.

Khalid:2019:AMS

[KZA19] Muhammad Zeeshan Khalid, Muhammad Zubair, and Majid Ali. An analytical method for the solution of two phase Stefan problem in cylindrical geometry. Applied Mathematics and Computation, 342(??):295–308, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306343.

Kang:2019:ISF

[KZL19] Zengxin Kang, Lei Zhang, and Kun Li. An improved social force model for pedestrian dynamics in shipwrecks. Applied Mathematics and Computation, 348(??):355–362, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310427.

Kang:2017:BCF

[KZR17] Zhijiang Kang, Danhui Zhao, and Hongxing Rui. Block-centered finite difference methods for general Darcy–Forchheimer problems. Applied Mathematics and Computation, 307(??):124–140, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301479.

Kurokawa:2019:CEM

[KZT19] Shun Kurokawa, Xiudeng Zheng, and Yi Tao. Cooperation evolves more when players keep the interaction with unknown players. Applied Mathematics and Computation, 350(??):209–216, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311044.

Kielbasinski:2015:IAP

[KZZ15] Andrzej Kielbasiński, Pawel Zieliński, and Krystyna Zietak. On iterative algorithms for the polar decomposition of a matrix and the matrix sign function. Applied Mathematics and Computation, 270(??):483–495, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010590.

Lotfi:2015:NTF

[LA15] Taher Lotfi and Paria Assari. New three- and four-parametric iterative with memory methods with efficiency index near 2. Applied Mathematics and Computation, 270 (??):1004–1010, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010723.

Laadhari:2018:IFE

[Laa18] Aymen Laadhari. Implicit finite element methodology for the numerical modeling of incompressible two-fluid flows with moving hyperelastic interface. Applied Mathematics and Computation, 333(??):376–400, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302558.

Liao:2019:BTP

[LAAZH19] Yunhua Liao, M. A. Aziz-Alaoui, Junchan Zhao, and Yaoping Hou. The behavior of Tutte polynomials of graphs under five graph operations and its applications. Applied Mathematics and Computation, 363 (??):Article 124641, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306332.

Lim:2015:APB

[LABK15] Ting Yee Lim, Mohammed Azmi Al-Betar, and Ahamad Tajudin Khader. Adaptive pair bonds in genetic algorithm: an application to real-parameter optimization. Applied Mathematics and Computation, 252(??):503–519, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016841.

Lachhwani:2015:MFA

[Lac15] Kailash Lachhwani. Modified FGP approach for multi-level multi objective linear fractional programming problems. Applied Mathematics and Computation, 266(??):1038–1049, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315008085.

Lahmiri:2018:MAS

[Lah18] Salim Lahmiri. Minuteahead stock price forecasting based on singular spectrum analysis and support vector regression. Applied Mathematics and Computation, 320(??):444–451, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306859.

Lai:2015:PCI

[Lai15] Pao-Lien Lai. Paths and cycles identifying vertices in twisted cubes. Applied Mathematics and Computation, 259(??):620–627, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500291X.

Lampret:2015:ADI

[Lam15]  Vito Lampret.  Accu-

rate double inequalities for generalized harmonic numbers. Applied Mathematics and Computation, 265(??):557–567, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006359.

Legaz-Aparicio:2018:NRR

[LAVME18] Álvar-Ginés Legaz-Aparicio, Rafael Verdú-Monedero, and Kjersti Engan. Noise robust and rotation invariant framework for texture analysis and classification. Applied Mathematics and Computation, 335(??):124–132, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303308.

Lazzaro:2017:FWT

[Laz17] Damiana Lazzaro. Fast weighted TV denoising via an edge driven metric. Applied Mathematics and Computation, 297(??):61–73, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306373.

Li:2015:SMN

[LB15a] Shan Li and John P. Boyd. Spectral methods in non-tensor geometry, Part II: Chebyshev versus Zernike polynomials, gridding strategies and spectral extension on squirclebounded and perturbedquadrifolium domains. Applied Mathematics and Computation, 269(??): 759–774, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315009856.

Liu:2015:RSL

[LB15b] Meng Liu and Chuanzhi Bai. A remark on a stochastic logistic model with Lévy jumps. Applied Mathematics and Computation, 251(??):521–526, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016269.

Liu:2016:DSO

[LB16a] Meng Liu and Chuanzhi Bai. Dynamics of a stochastic one-prey twopredator model with Lévy jumps. Applied Mathematics and Computation, 284(??):308–321, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301448.

Liu:2016:OHS

[LB16b] Meng Liu and Chuanzhi Bai. Optimal harvesting of a stochastic mutualism model with Lévy jumps. Applied Mathematics and Computation, 276(??):301–309, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501591X.

Liu:2019:EBN

[LB19] Yulan Liu and Shujun Bi. Error bounds for nonpolyhedral convex optimization and applications to linear convergence of FDM and PGM. Applied Mathematics and Computation, 358(??):418–435, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319303364.

Lima:2018:NSI

[LBB18] Pedro M. Lima, Azzeddine Bellour, and Mikhail V. Bulatov. Numerical solution of integro-differential equations arising from singular boundary value problems. Applied Mathematics and Computation, 336(??):1–15, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318303709.

Li:2018:EBS

[LBRS18] Fengwei Li, Hajo Broersma, Juan Rada, and Yuefang Sun. Extremal benzenoid systems for two modified versions of the Randić index. Applied Mathematics and Computation, 337(??):1–13, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304247.

Li:2015:MIS

[LC15a] Ming Li and Feilong Cao.

Multiscale interpolation on the sphere: Convergence rate and inverse theorem. Applied Mathematics and Computation, 263(??):134–150, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004907.

Li:2015:MSS

[LC15b] Xiuping Li and Wanrong Cao. On meansquare stability of twostep Maruyama methods for nonlinear neutral stochastic delay differential equations. Applied Mathematics and Computation, 261(??):373–381, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004294.

Liu:2015:ABS

[LC15c] Qun Liu and Qingmei Chen. Asymptotic behavior of a stochastic nonautonomous predator– prey system with jumps. Applied Mathematics and Computation, 271(??): 418–428, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315010954.

Liu:2015:DSD

[LC15d] Qun Liu and Qingmei Chen. Dynamics of stochastic delay Lotka– Volterra systems with impulsive toxicant input and Lévy noise in polluted environments. Applied Mathematics and Computation, 256(??):52–67, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000235.

Li:2016:SAR

[LC16a] Ruoxia Li and Jinde Cao. Stability analysis of reaction-diffusion uncertain memristive neural networks with timevarying delays and leakage term. Applied Mathematics and Computation, 278(??):54–69, March 31, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300169.

Liu:2016:DSS

[LC16b] Qun Liu and Qingmei Chen. Dynamics of a stochastic SIR epidemic model with saturated incidence. Applied Mathematics and Computation, 282(??):155–166, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301254.

Li:2018:NUI

[LC18] Xin Li and Yubo Chang.

Non-uniform interpolatory subdivision surface. Applied Mathematics and Computation, 324(??): 239–253, May 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317308172.

Li:2017:EFTb

[LCAA17] Ruoxia Li, Jinde Cao, Ahmad Alsaedi, and Fuad Alsaadi. Exponential and fixed-time synchronization of Cohen– Grossberg neural networks with time-varying delays and reaction-diffusion terms. Applied Mathematics and Computation, 313(??):37–51, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303855.

Ling:2015:CQC

[LCJ15] Si-Tao Ling, Xue-Han Cheng, and Tong-Song Jiang. Consimilarity of quaternions and coneigenvalues of quaternion matrices. Applied Mathematics and Computation, 270(??):984–992, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011492.

Li:2017:SHC

[LCL17a] Qinghua Li, Shenshen Chen, and Xinmei Luo. Steady heat conduction analyses using an interpolating element-free Galerkin scaled boundary method. Applied Mathematics and Computation, 300(??):103–115, May 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307330.

Liu:2017:SOC

[LCL17b] Zhengguang Liu, Aijie Cheng, and Xiaoli Li. A second order Crank– Nicolson scheme for fractional Cattaneo equation based on new fractional derivative. Applied Mathematics and Computation, 311(??):361–374, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303338.

Luo:2019:ESA

[LCLZ19] Mengzhuo Luo, Jun Cheng, Xinzhi Liu, and Shouming Zhong. An extended synchronization analysis for memristor-based coupled neural networks via aperiodically intermittent control. Applied Mathematics and Computation, 344–345(??):163–182, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308610.

Li:2019:SRT

[LCM19] Dan Li, Yuanyuan Chen, and Jixiang Meng. The A_{α}-spectral radius of trees and unicyclic graphs with given degree sequence. Applied Mathematics and Computation, 363(??): Article 124622, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319306149.

Lv:2019:ECI

[LCOS19] Huadong Lv, Ya Chen, Yao Ouyang, and Hongxia Sun. On the equivalence of the Choquet integral and the pan-integrals from above. Applied Mathematics and Computation, 361(??):15–21, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304035.

Li:2015:WBF

[LCQ15] Gang Li, Valerio Caleffi, and Zhengkun Qi. A wellbalanced finite difference WENO scheme for shallow water flow model. Applied Mathematics and Computation, 265(??):1–16, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005123.

Li:2015:NLB

[LCS15] Qianhuan Li, Zhenhua Chai, and Baochang Shi. A novel lattice Boltzmann model for the coupled viscous Burgers’ equations. Applied Mathematics and Computation, 250(??):948–957, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015598.

Li:2015:EAS

[LCW15] Xiaolin Li, Hao Chen, and Yan Wang. Error analysis in Sobolev spaces for the improved moving leastsquare approximation and the improved element-free Galerkin method. Applied Mathematics and Computation, 262(??):56–78, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004282.

Lin:2018:CSL

[LCW^{+}18] Wenshui Lin, Jianfeng Chen, Zhixi Wu, Darko Dimitrov, and Linshan Huang. Computer search for large trees with minimal ABC index. Applied Mathematics and Computation, 338(??):207–220, December 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318305009.

Li:2019:IRC

[LCW19a] Shi Li, Guang Chen, and Yigang Wang. An improved rational cubic clipping method for computing real roots of a polynomial. Applied Mathematics and Computation, 349(??):207–213, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311019.

Liu:2019:NAM

[LCW19b] Wei Liu, Jintao Cui, and Zhifeng Wang. Numerical analysis and modeling of multiscale Forchheimer– Forchheimer coupled model for compressible fluid flow in fractured media aquifer system. Applied Mathematics and Computation, 353(??):7–28, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300517.

Li:2016:RQF

[LCY16] Zhi-Min Li, Xiao-Heng Chang, and Lu Yu. Robust quantized H_{∞} filtering for discrete-time uncertain systems with packet dropouts. Applied Mathematics and Computation, 275(??):361–371, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015714.

Li:2019:ASSb

[LCZ19] Zhizhi Li, Risheng Chu, and Huai Zhang. Accelerating the shift-splitting iteration algorithm. Applied Mathematics and Computation, 361(??):421–429, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319304588.

Le:2017:PAS

[LD17] Nhat-Tan Le and Duy-Minh Dang. Pricing American-style Parisian down-and-out call options. Applied Mathematics and Computation, 305(??):330–347, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301194. See

corrigendum [LLZD18a].

Li:2018:TFM

[LD18a] Jiyong Li and Shuo Deng. Trigonometrically fitted multi-step RKN methods for second-order oscillatory initial value problems. Applied Mathematics and Computation, 320(??):740–753, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730735X.

Liu:2018:ODB

[LD18b] Muhuo Liu and Kinkar Ch. Das. On the ordering of distance-based invariants of graphs. Applied Mathematics and Computation, 324(??):191–201, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308743.

Li:2019:AEF

[LD19a] Xiaolin Li and Haiyun Dong. Analysis of the element-free Galerkin method for Signorini problems. Applied Mathematics and Computation, 346 (??):41–56, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318309196.

Liu:2019:CFH

[LD19b] Chein-Shan Liu and Leiting Dong. Closedform higher-order numerical differentiators for differentiating noisy signals. Applied Mathematics and Computation, 359(??):386–403, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303108.

Lu:2019:PSS

[LD19c] Chun Lu and Xiaohua Ding. Periodic solutions and stationary distribution for a stochastic predator–prey system with impulsive perturbations. Applied Mathematics and Computation, 350(??):313–322, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300323.

Lima:2015:PAF

[LdCC15] Ana Carolina E. S. Lima, Leandro Nunes de Castro, and Juan M. Corchado. A polarity analysis framework for Twitter messages. Applied Mathematics and Computation, 270(??):756–767, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315011145.

Lenzi:2019:TDS

[LdCM19] E. K. Lenzi, A. S. M. de Castro, and R. S. Mendes. Time dependent solutions for fractional coupled Schrödinger equations. Applied Mathematics and Computation, 346(??):622–632, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309482.

Liu:2015:ASL

[LDD15] Meng Liu, Meiling Deng, and Bo Du. Analysis of a stochastic logistic model with diffusion. Applied Mathematics and Computation, 266(??):169–182, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006670.

Liu:2018:ASE

[LDH18] Linna Liu, Feiqi Deng, and Ting Hou. Almost sure exponential stability of implicit numerical solution for stochastic functional differential equation with extended polynomial growth condition. Applied Mathematics and Computation, 330(??):201–212, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301334.

Luo:2015:DAA

[LDL15] Xuqiong Luo, Qikui Du, and Libin Liu. A D-N alternating algorithm for exterior 3-D Poisson problem with prolate spheroid boundary. Applied Mathematics and Computation, 269(??):252–264, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009820.

Dong:2018:MNA

[lDqCWzH18] Qiao li Dong, Li qun Cao, Xin Wang, and Ji zu Huang. Multiscale numerical algorithms for elastic wave equations with rapidly oscillating coefficients. Applied Mathematics and Computation, 336(??):16–35, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303990.

Liang:2018:LLS

[LDS^{+}18] Kun Liang, Mingcheng Dai, Hao Shen, Jing Wang, Zhen Wang, and Bo Chen. L_{2}–L_{∞} synchronization for singularly perturbed complex networks with semi-Markov jump topology. Applied Mathematics and Computation, 321(??):450–462, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307385.

Li:2019:SDE

[LDSD19] Tao Li, Zhuxiang Dai, Gongfei Song, and Haiping Du. Simultaneous disturbance estimation and fault reconstruction using probability density functions. Applied Mathematics and Computation, 362 (??):Article 124561, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305442.

Lenarduzzi:2015:CCM

[Len15] Licia Lenarduzzi. Compression of corneal maps of curvature. Applied Mathematics and Computation, 252(??):77–87, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401635X.

Liu:2015:CSA

[LF15] Xueying Liu and Meiling Fu. Cuckoo search algorithm based on frog leaping local search and chaos theory. Applied Mathematics and Computation, 266 (??):1083–1092, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S009630031500822X.

Liu:2019:SCA

[LFC^{+}19] Jun Liu, Hongfei Fu, Xiaochao Chai, Yanan Sun, and Hui Guo. Stability and convergence analysis of the quadratic spline collocation method for timedependent fractional diffusion equations. Applied Mathematics and Computation, 346(??):633–648, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309123.

Liu:2015:NEM

[LFL^{+}15] Yang Liu, Zhichao Fang, Hong Li, Siriguleng He, and Wei Gao. A new expanded mixed method for parabolic integro-differential equations. Applied Mathematics and Computation, 259(??):600–613, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002817.

Lu:2018:TFG

[LFY18] Changna Lu, Chen Fu, and Hongwei Yang. Timefractional generalized Boussinesq equation for Rossby solitary waves with dissipation effect in stratified fluid and conservation laws as well as exact solutions. Applied Mathematics and Computation, 327(??):104–116, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300316.

Li:2015:SRM

[LG15] Xiangli Li and Xiao Guo.

Spectral residual methods with two new nonmonotone line searches for large-scale nonlinear systems of equations. Applied Mathematics and Computation, 269(??):59–69, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010012.

Li:2016:BSN

[LG16] Fushan Li and Qingyong Gao. Blow-up of solution for a nonlinear Petrovsky type equation with memory. Applied Mathematics and Computation, 274(??):383–392, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315014800.

Li:2017:TPM

[LG17] Yibao Li and Shimin Guo. Triply periodic minimal surface using a modified Allen–Cahn equation. Applied Mathematics and Computation, 295(??):84–94, February 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306038.

Li:2018:BSN

[LG18] Yinkui Li and Ruijuan Gu. Bounds for scattering number and rupture degree of graphs with genus. Applied Mathematics and Computation, 337(??):315–328, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304326.

Liu:2019:SIR

[LG19] Yong Liu and Chuanqing Gu. A shift and invert reorthogonalization Arnoldi algorithm for solving the chemical master equation. Applied Mathematics and Computation, 349(??):1– 13, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310683.

Li:2019:NFS

[LGC19a] Ruoxia Li, Xingbao Gao, and Jinde Cao. Nonfragile state estimation for delayed fractionalorder memristive neural networks. Applied Mathematics and Computation, 340(??):221–233, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307458.

Li:2019:QSE

[LGC19b] Ruoxia Li, Xingbao Gao, and Jinde Cao. Quasistate estimation and quasisynchronization control of quaternion-valued fractionalorder fuzzy memristive neural networks: Vector ordering approach. Applied Mathematics and Computation, 362(??): Article 124572, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300319305557.

Li:2019:RFD

[LGCS19] Jian Li, Xinxin Guo, Cong Chen, and Qingyu Su. Robust fault diagnosis for switched systems based on sliding mode observer. Applied Mathematics and Computation, 341(??):193–203, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318307501.

Lu:2019:CSN

[LGHW19] Jianquan Lu, Xing Guo, Tingwen Huang, and Zhen Wang. Consensus of signed networked multiagent systems with nonlinear coupling and communication delays. Applied Mathematics and Computation, 350(??):153–162, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300141.

Liu:2015:RIC

[LGK15] Zhen Liu, Cunchen Gao, and Yonggui Kao. Robust H-infinity control for a class of neutral-type systems via sliding mode observer. Applied Mathematics and Computation, 271(??):669–681, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012254.

Liu:2016:SSC

[LGL16] Yan Liu, Ying Guo, and Wenxue Li. The stability of stochastic coupled systems with time delays and time-varying coupling structure. Applied Mathematics and Computation, 290(??):507–520, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304003.

Li:2019:GCL

[LGL19a] Yinkui Li, Ruijuan Gu, and Hui Lei. The generalized connectivity of the line graph and the total graph for the complete bipartite graph. Applied Mathematics and Computation, 347(??):645–652, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318310129.

Liu:2019:CMG

[LGL^{+}19b] Chen Liu, Hao Guo, Zhibin Li, Xiaoyuan Gao, and Shudong Li. Coevolution of multi-game resolves social dilemma in network population. Applied Mathematics and Computation, 341(??):402–407, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318307914.

Liu:2019:SGM

[LGL19c] Xikui Liu, Yingying Ge, and Yan Li. Stackelberg games for model-free continuous-time stochastic systems based on adaptive dynamic programming. Applied Mathematics and Computation, 363 (??):Article 124568, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031930551X.

Leitao:2017:OTS

[LGO17] Álvaro Leitao, Lech A. Grzelak, and Cornelis W. Oosterlee. On a one timestep Monte Carlo simulation approach of the SABR model: Application to European options. Applied Mathematics and Computation, 293(??):461–479, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305252.

Li:2019:EEI

[LGQL19] Jing Li, Nan Gao, Lu Qiao, and Sujuan Liu. On the extended Estrada index of some graphs. Applied Mathematics and Computation, 355(??):311–318, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301791.

Li:2018:BSP

[LGS^{+}18] Jianxi Li, Ji-Ming Guo, Wai Chee Shiu, S. Burcu Bozkurt Altindag, and Durmus Bozkurt. Bounding the sum of powers of normalized Laplacian eigenvalues of a graph. Applied Mathematics and Computation, 324(??):82–92, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300317308561.

Li:2015:NSM

[LH15] Jin-Dong Li and Guang-Xin Huang. On new subclass of meromorphic close-to-convex functions. Applied Mathematics and Computation, 266(??):424–428, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006992.

Leng:2017:EBS

[LH17a] Huinan Leng and Zhiqing He. Eigenvalue bounds for symmetric matrices with entries in one interval. Applied Mathematics and Computation, 299(??):58–65, April 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307111.

Li:2017:NLS

[LH17b] Deqiong Li and Yaoping Hou. The normalized Laplacian spectrum of quadrilateral graphs and its applications. Applied Mathematics and Computation, 297(??):180–188, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306579.

Li:2017:SNP

[LH17c] Xianping Li and Weizhang Huang. A study on nonnegativity preservation in finite element approximation of Nagumo-type nonlinear differential equations. Applied Mathematics and Computation, 309(??):49–67, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302229.

Li:2019:EEPa

[LH19a] Haochen Li and Qi Hong.

An efficient energy-preserving algorithm for the Lorentz force system. Applied Mathematics and Computation, 358(??):161–168, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303236.

Li:2019:RGF

[LH19b] Sheng-Kun Li and Ting-Zhu Huang. Restarted global FOM and GMRES algorithms for the Steinlike matrix equation X + M(X) = C. Applied Mathematics and Computation, 348(??):206–214, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310324.

Liu:2019:RLP

[LH19c] Qinghai Liu and Yanmei Hong. The reliability of lexicographic product digraphs. Applied Mathematics and Computation, 358(??):449–454, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303285.

Liu:2018:GNC

[LHG18] Huiqing Liu, Xiaolan Hu, and Shan Gao. The g-good neighbor conditional diagnosability of twisted hypercubes under the PMC and MM* model. Applied Mathematics and Computation, 332(??):484–492, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318302157.

Lim:2016:NPF

[LHHS16] Jae Hyuk Lim, Milan Henry, Do-Soon Hwang, and Dongwoo Sohn. Numerical prediction of fiber mechanical properties considering random microstructures using inverse analysis with quasi-analytical gradients. Applied Mathematics and Computation, 273(??):201–216, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013144.

Li:2015:SCQ

[LHJ15] Jianling Li, Renshuai Huang, and Jinbao Jian. A superlinearly convergent QP-free algorithm for mathematical programs with equilibrium constraints. Applied Mathematics and Computation, 269(??):885–903, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010036.

Li:2019:ROD

[LHLL19] Kun Li, Ting-Zhu Huang, Liang Li, and Stéphane Lanteri. A reduced-order discontinuous Galerkin method based on a Krylov subspace technique in nanophotonics. Applied Mathematics and Computation, 358(??):128–145, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303133.

Liu:2018:ISK

[LHS18] Bin Liu, David J. Hill, and Zhijie Sun. Input-to-state-KL-stability and criteria for a class of hybrid dynamical systems. Applied Mathematics and Computation, 326(??):124–140, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300080.

Luo:2016:QSC

[LHWG16] Wei-Hua Luo, Ting-Zhu Huang, Guo-Cheng Wu, and Xian-Ming Gu. Quadratic spline collocation method for the time fractional subdiffusion equation. Applied Mathematics and Computation, 276(??):252–265, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300084.

Luo:2015:NMD

[LHX15] Shao-Kai Luo, Jin-Man He, and Yan-Li Xu. A new method of dynamical stability, i.e. fractional generalized Hamiltonian method, and its applications. Applied Mathematics and Computation, 269(??):77–86, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009662.

Liu:2019:MAS

[LHZM19] Hongyan Liu, Jin Huang, Wei Zhang, and Yanying Ma. Meshfree approach for solving multi-dimensional systems of Fredholm integral equations via barycentric Lagrange interpolation. Applied Mathematics and Computation, 346(??):295–304, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308907.

Li:2015:BCD

[Li15a] Li Li. Bifurcation and chaos in a discrete physiological control system. Applied Mathematics and Computation, 252(??): 397–404, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314016403.

Li:2015:PIS

[Li15b]  Li Li.  Patch inva-

sion in a spatial epidemic model. Applied Mathematics and Computation, 258(??):342–349, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001587.

Li:2016:ITO

[Li16a] Hao Li. On an integraltype operator from the Bloch space to mixed norm spaces. Applied Mathematics and Computation, 273(??):624–630, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013648.

Li:2016:CES

[Li16b]  Shoufu Li.  Canonical

Euler splitting method for nonlinear composite stiff evolution equations. Applied Mathematics and Computation, 289(??):220–236, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303228.

Li:2017:NSH

[Li17a] Bing Li. A note on stability of hybrid stochastic differential equations. Applied Mathematics and Computation, 299(??):45– 57, April 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031630710X.

Li:2017:LDS

[Li17b] Can Li. Linearized difference schemes for a BBM equation with a fractional nonlocal viscous term. Applied Mathematics and Computation, 311(??):240–250, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730320X.

Li:2017:EMB

[Li17c] Jin Li. The extrapolation methods based on Simpson’s rule for computing supersingular integral on interval. Applied Mathematics and Computation, 310(??):204–214, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302382.

Li:2017:FIF

[Li17d]  Jiyong Li.  A fam-

ily of improved Falknertype methods for oscillatory systems. Applied Mathematics and Computation, 293(??):345–357, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305410.

Li:2017:GCT

[Li17e] Pingrun Li. Generalized convolution-type singular integral equations. Applied Mathematics and Computation, 311(??):314–323, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0096300317303375.

Li:2017:CMS

[Li17f] Yinkui Li. Cycle multiplicity of some total graphs. Applied Mathematics and Computation, 292(??):107–113, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304222.

Li:2018:TFT

[Li18]  Jiyong Li.  Trigono-

metrically fitted threederivative Runge–Kutta methods for solving oscillatory initial value problems. Applied Mathematics and Computation, 330(??):103–117, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300304.

Li:2019:SIE

[Li19] Pingrun Li. Singular integral equations of convolution type with Cauchy kernel in the class of exponentially increasing functions. Applied Mathematics and Computation, 344–345(??):116–127, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830852X.

Liang:2015:CAS

[Lia15a] Hui Liang. Convergence and asymptotic stability of Galerkin methods for linear parabolic equations with delays. Applied Mathematics and Computation, 264(??):160–178, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005664.

Liao:2015:ABJ

[Lia15b]  Wenyuan Liao.  An

adjoint-based Jacobi-type iterative method for elastic full waveform inversion problem. Applied Mathematics and Computation, 267(??):56–70, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007869.

Lin:2015:ACN

[Lin15] Matthew M. Lin. An algorithm for constructing nonnegative matrices with prescribed real eigenvalues. Applied Mathematics and Computation, 256(??):582–590, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000478.

Lin:2019:PSS

[Lin19]  Bin Lin.  Parametric

spline schemes for the coupled nonlinear Schrödinger equation. Applied Mathematics and Computation, 360(??):58–69, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303340.

Liu:2015:UGC

[Liu15a] Hanze Liu. The unified group classification method for the generalized nonlinear wave equation and its partial difference schemes. Applied Mathematics and Computation, 251(??):203–210, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015392.

Liu:2015:PAG

[Liu15b] Li-Min Liu. A parallel algorithm for generating ideal IC-colorings of cycles. Applied Mathematics and Computation, 250(??):615–627, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014945.

Liu:2015:PSE

[Liu15c] Pan-Ping Liu. Periodic solutions in an epidemic model with diffusion and delay. Applied Mathematics and Computation, 265(??):275–291, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006451.

Liu:2015:ASN

[Liu15d]  Qun Liu.  Analysis

of a stochastic nonautonomous food-limited Lotka–Volterra cooperative model. Applied Mathematics and Computation, 254(??):1–8, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.

com/science/article/
pii/S0096300314018037.

Liu:2015:PSO

[Liu15e] Wanping Liu. Periodicity of a second-order switched difference system over integers. Applied Mathematics and Computation, 250(??):733–743, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401546X.

Liu:2015:HSI

[Liu15f] Xifu Liu. The Hermitian solution of AXA^{∗} = B subject to CXC^{∗} ≥ D. Applied Mathematics and Computation, 270(??):890–898, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011583.

Liu:2015:FVE

[Liu15g] Xuefeng Liu. A framework of verified eigenvalue bounds for selfadjoint differential operators. Applied Mathematics and Computation, 267(??):341–355, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003628.

Liu:2016:RPF

[Liu16a] Tao Liu. Reconstruction of a permeability field with the wavelet multiscalehomotopy method for a nonlinear convectiondiffusion equation. Applied Mathematics and Computation, 275(??):432–437, February 15, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315016173.

Liu:2016:PCS

[Liu16b]  Yuji Liu.  On piece-

wise continuous solutions of higher order impulsive fractional differential equations and applications. Applied Mathematics and Computation, 287– 288(??):38–49, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316302429.

Liu:2017:NSP

[Liu17a] Hongxia Liu. A numerical study of the performance of alternative weighted ENO methods based on various numerical fluxes for conservation law. Applied Mathematics and Computation, 296(??):182–197, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630621X.

Liu:2017:CDT

[Liu17b]  Yan Liu.  Continuous

dependence for a thermal convection model with temperature-dependent solubility. Applied Mathematics and Computation, 308(??):18–30, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301753.

Liu:2018:SSPa

[Liu18a] Chein-Shan Liu. Solving singularly perturbed problems by a weak-form integral equation with exponential trial functions. Applied Mathematics and Computation, 329(??):154–174, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318300961.

Liu:2018:NMM

[Liu18b]  Tao Liu.  A nonlin-

ear multigrid method for inverse problem in the multiphase porous media flow. Applied Mathematics and Computation, 320(??):271–281, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306677.

Liu:2018:AHS

[Liu18c] Xin Liu. The η-anti-Hermitian solution to some classic matrix equations. Applied Mathematics and Computation, 320(??):264–270, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306616.

Liu:2019:SDS

[Liu19] Kai Liu. Stability in distribution for stochastic differential equations with memory driven by positive semigroups and Lévy processes. Applied Mathematics and Computation, 362 (??):Article 124580, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305727.

Li:2015:ECA

[LJ15a] Liangliang Li and Jigui Jian. Exponential pconvergence analysis for stochastic BAM neural networks with timevarying and infinite distributed delays. Applied Mathematics and Computation, 266(??):860–873, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315008036.

Lu:2015:EAG

[LJ15b] Junfeng Lu and Yiming Jin. Eigenvalue analysis of a generalized indefinite block triangular preconditioner for generalized saddle point problems. Applied Mathematics and Computation, 271(??):389–399, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012710.

Luo:2017:CAR

[LJ17] Chao Luo and Zhipeng Jiang. Coevolving allocation of resources and cooperation in spatial evolutionary games. Applied Mathematics and Computation, 311(??):47–57, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302977.

Liu:2019:SNS

[LJ19] Xiping Liu and Mei Jia. Solvability and numerical simulations for BVPs of fractional coupled systems involving left and right fractional derivatives. Applied Mathematics and Computation, 353(??):230–242, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301092.

Luo:2016:ROE

[LJC16] Zhendong Luo, Shiju Jin, and Jing Chen. A reducedorder extrapolation central difference scheme based on POD for two-dimensional fourth-order hyperbolic equations. Applied Mathematics and Computation, 289(??):396–408, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303411.

Li:2017:HOW

[LJD17] Maojun Li, Yuting Jiang, and Haiyun Dong. High order well-balanced central local discontinuous Galerkin-finite element methods for solving the Green–Naghdi model. Applied Mathematics and Computation, 315(??): 113–130, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317305118.

Jiang:2016:EGI

[lJdW16] Zhao lin Jiang and Dan dan Wang. Explicit group inverse of an innovative patterned matrix. Applied Mathematics and Computation, 274(??):220–228, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014836.

Li:2019:EMC

[LJG^{+}19] Zhibin Li, Danyang Jia, Hao Guo, Yini Geng, Chen Shen, Zhen Wang, and Xuelong Li. The effect of multigame on cooperation in spatial network. Applied Mathematics and Computation, 351(??):162–167, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311202.

Liu:2018:SDE

[LJHA18] Qun Liu, Daqing Jiang, Tasawar Hayat, and Bashir Ahmad. Stationary distribution and extinction of a stochastic predator– prey model with additional food and nonlinear perturbation. Applied Mathematics and Computation, 320(??):226–239, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306586.

Long:2019:SAS

[LJHW19] Fei Long, Lin Jiang, Yong He, and Min Wu. Stability analysis of systems with time-varying delay via novel augmented Lyapunov-Krasovskii functionals and an improved integral inequality. Applied Mathematics and Computation, 357(??):325–337, September 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319302802.

Liu:2015:BNG

[LJK15] Xia Liu and H. Jardón-Kojakhmetov. Bifurcations of a non-gravitational interaction problem. Applied Mathematics and Computation, 251(??): 253–257, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314015987.

Li:2015:SEB

[LJL15a] Lei Li, Xinchun Jia, and Jiankang Liu. Stabilization of an Euler–Bernoulli beam equation via a corrupted boundary position feedback. Applied Mathematics and Computation, 270(??):648–653, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011273.

Lu:2015:MWE

[LjL15b] Fang Lu and Sheng jie Li.

Method of weighted expected residual for solving stochastic variational inequality problems. Applied Mathematics and Computation, 269(??):651–663, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010371.

Lv:2016:DPN

[LJL16] Xiao-Guang Lv, Le Jiang, and Jun Liu. Deblurring Poisson noisy images by total variation with overlapping group sparsity. Applied Mathematics and Computation, 289(??):132–148, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302296.

Li:2018:GEP

[LJL18] Haochen Li, Chaolong Jiang, and Zhongquan Lv. A Galerkin energypreserving method for two dimensional nonlinear Schrödinger equation. Applied Mathematics and Computation, 324(??):16– 27, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317308457.

Li:2019:STS

[LJL19] Hui Li, Wei Jiang, and Wenya Li. Space-time spectral method for the Cattaneo equation with time fractional derivative. Applied Mathematics and Computation, 349(??):325–336, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311111.

Liu:2019:EEL

[LJR19] Run-Ran Liu, Chun-Xiao Jia, and Zhihai Rong. Effects of enhancement level on evolutionary public goods game with payoff aspirations. Applied Mathematics and Computation, 350(??):242–248, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300177.

Lu:2017:SWR

[LJS17] Yi Lu, Yao-Lin Jiang, and Bo Song. Symplectic waveform relaxation methods for Hamiltonian systems. Applied Mathematics and Computation, 292(??):228–239, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304830.

Liu:2018:TBS

[LJS18] Qun Liu, Daqing Jiang, and Ningzhong Shi. Threshold behavior in a stochastic SIQR epidemic model with standard incidence and regime switching. Applied Mathematics and Computation, 316(??): 310–325, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317305933.

Li:2015:GML

[LJW^{+}15] Hong-Li Li, Yao-Lin Jiang, Zuolei Wang, Long Zhang, and Zhidong Teng. Global Mittag-Leffler stability of coupled system of fractional-order differential equations on network. Applied Mathematics and Computation, 270(??):269–277, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501098X.

Li:2019:DAR

[LJYH19] Jiarong Li, Haijun Jiang, Zhiyong Yu, and Cheng Hu. Dynamical analysis of rumor spreading model in homogeneous complex networks. Applied Mathematics and Computation, 359(??):374–385, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303789.

LeMezo:2018:BSO

[LJZ18] Thomas Le Mézo, Luc Jaulin, and Benoît Zerr. Bracketing the solutions of an ordinary differential equation with uncertain initial conditions. Applied Mathematics and Computation, 318(??):70–79, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304976.

Leonov:2015:DSA

[LK15] G. A. Leonov and N. V.

Kuznetsov. On differences and similarities in the analysis of Lorenz, Chen, and Lu systems. Applied Mathematics and Computation, 256(??):334–343, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017937.

Lukas:2018:ATS

[LK18] Petr Lukás and Petr Knobloch. Adaptive techniques in SOLD methods. Applied Mathematics and Computation, 319(??):24–30, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300425.

Li:2019:MLS

[LK19a] Hui Li and YongGui Kao. Mittag-Leffler stability for a new coupled system of fractional-order differential equations with impulses. Applied Mathematics and Computation, 361(??):22–31, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304114.

Li:2019:DFL

[LK19b] Mei Li and Lingchen Kong. Double fused lasso penalized LAD for matrix regression. Applied Mathematics and Computation, 357(??):119–138, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302590.

Lee:2017:DTF

[LKM17] Min-Young Lee, Young Ik Kim, and Á. Alberto Magreñán. On the dynamics of a triparametric family of optimal fourthorder multiple-zero finders with a weight function of the principal m^{th} root of a function-to function ratio. Applied Mathematics and Computation, 315(??):564–590, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305568.

Li:2018:MSD

[LKZ18] Qiang Li, Ting Kang, and Qimin Zhang. Meansquare dissipative methods for stochastic agedependent capital system with fractional Brownian motion and jumps. Applied Mathematics and Computation, 339(??):63–80, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305769.

Li:2015:AVZ

[LL15a] Jing Li and Yiyang Li. The asymptotic value of the zeroth-order Randić index and sumconnectivity index for trees. Applied Mathematics and Computation, 266 (??):1027–1030, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008097.

Li:2015:ILC

[LL15b] Jinsha Li and Junmin Li. Iterative learning control approach for a kind of heterogeneous multiagent systems with distributed initial state learning. Applied Mathematics and Computation, 265(??):1044–1057, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008164.

Li:2015:MIT

[LL15c] Qiang Li and Yongxiang Li. Monotone iterative technique for second order delayed periodic problem in Banach spaces. Applied Mathematics and Computation, 270(??):654–664, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011261.

Li:2015:APM

[LL15d] Shishun Li and Wenqiang Li. The analysis of PMHSS-multigrid methods for elliptic problems with smooth complex coefficients. Applied Mathematics and Computation, 265(??):196–206, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006414.

Li:2015:SOE

[LL15e] Wei Li and Can Li.

Second-order explicit difference schemes for the space fractional advection diffusion equation. Applied Mathematics and Computation, 257(??):446–457, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314015525.

Lu:2015:ECR

[LL15f] Liang Lu and Zhenhai Liu. Existence and controllability results for stochastic fractional evolution hemivariational inequalities. Applied Mathematics and Computation, 268(??):1164–1176, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009352.

Li:2016:RBF

[LL16a] Shuling Li and Xiaolin Li. Radial basis functions and level set method for image segmentation using partial differential equation. Applied Mathematics and Computation, 286(??):29–40, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316302442.

Liu:2016:DPU

[LL16b] Xiaogang Liu and Binlong Li. Distance powers of unitary Cayley graphs. Applied Mathematics and Computation, 289(??):272–280, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303320.

Liu:2016:GWP

[LL16c] Yujun Liu and Shan Li. Global well-posedness for magneto-micropolar system in 2 1/2 dimensions. Applied Mathematics and Computation, 280(??):72–85, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300029.

Li:2017:DMI

[LL17a] Fang Li and Xiaoguang Lv. A decoupled method for image inpainting with patch-based low rank regulariztion. Applied Mathematics and Computation, 314(??):334–348, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730440X.

Li:2017:DSF

[LL17b] Shuanming Li and Yi Lu.

Distributional study of finite-time ruin related problems for the classical risk model. Applied Mathematics and Computation, 315(??):319–330, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305155.

Ling:2017:NLG

[LL17c] Si-Tao Ling and Qing-Bing Liu. New local generalized shift-splitting preconditioners for saddle point problems. Applied Mathematics and Computation, 302(??):58–67, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730022X.

Li:2018:MES

[LL18a] Hui Li and Ye-Zhou Li. Meromorphic exact solutions of two extended (3+1)-dimensional Jimbo– Miwa equations. Applied Mathematics and Computation, 333(??):369–375, September 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318302807.

Liu:2018:WPI

[LL18b] Guoliang Liu and Guodong Liu. Wiener polarity index of dendrimers. Applied Mathematics and Computation, 322(??):151–153, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308202.

Livinska:2018:TSR

[LL18c] Hanna Livinska and Eugene Lebedev. On transient and stationary regimes for multi-channel networks with periodic inputs. Applied Mathematics and Computation, 319(??):13–23, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300255.

Lee:2019:FAC

[LL19a] Seunggyu Lee and Dongsun Lee. The fractional Allen–Cahn equation with the sextic potential. Applied Mathematics and Computation, 351(??):176–192, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300463.

Li:2019:NNB

[LL19b] Y. Li and G. R. Liu. A novel node-based smoothed finite element method with linear strain fields for static, free and forced vibration analyses of solids. Applied Mathematics and Computation, 352(??):30– 58, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300529.

Li:2019:SPD

[LL19c] Yeping Li and Li Lu. Stability of planar diffusion wave for the three dimensional full bipolar Euler–Poisson system. Applied Mathematics and Computation, 356(??):392–410, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302206.

Liu:2019:GSM

[LL19d] Haiyu Liu and Shujuan Lü. Galerkin spectral method for nonlinear time fractional cable equation with smooth and nonsmooth solutions. Applied Mathematics and Computation, 350(??):32–47, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311330.

Lu:2016:ACS

[LLB16] Liang Lu, Zhenhai Liu, and Maojun Bin. Approximate controllability for stochastic evolution inclusions of Clarke’s subdifferential type. Applied Mathematics and Computation, 286(??):201–212, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302703.

Lai:2015:TMS

[LLC^{+}15] Wei Lai, Xianming Liu, Weimin Chen, Xiaohua Lei, Xiaosheng Tang, and Zhigang Zang. Transient multiexponential signals analysis using Bayesian deconvolution. Applied Mathematics and Computation, 265(??):486–493, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315006499.

Li:2018:CET

[LLC18] Qiaoping Li, Sanyang Liu, and Yonggang Chen. Combination event-triggered adaptive networked synchronization communication for nonlinear uncertain fractional-order chaotic systems. Applied Mathematics and Computation, 333(??):521–535, September 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318302753.

Li:2019:DLS

[LLC19] Siqing Li, Leevan Ling, and Ka Chun Cheung. Discrete least-squares radial basis functions approximations. Applied Mathematics and Computation, 355(??):542–552, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302000.

Leung:2015:SAF

[LLCR15] Andrew Y. T. Leung, Xian-Feng Li, Yan-Dong Chu, and Xiao-Bo Rao. A simple adaptive-feedback scheme for identical synchronizing chaotic systems with uncertain parameters. Applied Mathematics and Computation, 253(??):172–183, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017305.

Llibre:2015:LCC

[LLdM15] Jaume Llibre, Bruno D.

Lopes, and Jaime R. de Moraes. Limit cycles of cubic polynomial differential systems with rational first integrals of degree 2. Applied Mathematics and Computation, 250(??):887–907, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015513.

Llibre:2016:LCB

[LLdM16] J. Llibre, B. D. Lopes, and J. R. de Moraes. Limit cycles bifurcating from the periodic annulus of the weight-homogeneous polynomial centers of weight-degree 2. Applied Mathematics and Computation, 274(??):47–54, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315014307.

Liu:2015:MMB

[LLG^{+}15] Q. Liu, F. Liu, Y. T. Gu, P. Zhuang, J. Chen, and I. Turner. A meshless method based on Point Interpolation Method (PIM) for the space fractional diffusion equation. Applied Mathematics and Computation, 256(??):930–938, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001241.

Liu:2019:RIA

[LLJ19] Huili Liu, Yixuan Liu, and Seoung Dal Jung. Ruled invariants and associated ruled surfaces of a space curve. Applied Mathematics and Computation, 348(??):479–486, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031831052X.

Lee:2016:SME

[LLL16a] Hyojung Lee, Sunmi Lee, and Chang Hyeong Lee. Stochastic methods for epidemic models: an application to the 2009 H1N1 influenza outbreak in Korea. Applied Mathematics and Computation, 286(??):232–249, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302697.

Lu:2016:MHS

[LLL16b] Yang Lu, Dan Li, and Shengqiang Liu. Modeling of hunting strategies of the predators in susceptible and infected prey. Applied Mathematics and Computation, 284(??):268–285, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301977.

Li:2017:AID

[LLL17a] Shangzhao Li, Weijun Liu, and Xianhua Li. On the automorphisms of 2 − (v, k, 1) designs. Applied Mathematics and Computation, 292(??):120–127, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303988.

Liu:2017:NES

[LLL17b] Qiong Liu, Zhengbo Li, and Chaoqian Li. A note on eventually SDD matrices and eigenvalue localization. Applied Mathematics and Computation, 311(??):19–21, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302990.

Li:2018:CAT

[LLL18a] Jing Jian Li, Bo Ling, and Guodong Liu. A characterisation on arctransitive graphs of prime valency. Applied Mathematics and Computation, 325(??):227–233, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308834.

Li:2018:EIT

[LLL^{+}18b] Meilin Li, Jianquan Lu, Jungang Lou, Yang Liu, and Fuad E. Alsaadi. The equivalence issue of two kinds of controllers in Boolean control networks. Applied Mathematics and Computation, 321(??):633–640, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307841.

Liu:2018:FDS

[LLL18c] Zeting Liu, Shujuan Lü, and Fawang Liu. Fully discrete spectral methods for solving time fractional nonlinear sine– Gordon equation with smooth and non-smooth solutions. Applied Mathematics and Computation, 333(??):213–224, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830242X.

Lo:2018:EIB

[LLL18d] D. C. Lo, C-P Lee, and I-F Lin. An efficient immersed boundary method for fluid flow simulations with moving boundaries. Applied Mathematics and Computation, 328(??):312–337, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300353.

Li:2019:EIP

[LLL19a] Chaoqian Li, Yajun Liu, and Yaotang Li. Ceigenvalues intervals for piezoelectric-type tensors. Applied Mathematics and Computation, 358(??): 244–250, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319303248.

Li:2019:DAA

[LLL19b] Shangzhao Li, Weijun Liu, and Xianhua Li. 2 − (v, k, 1) designs admitting automorphism groups with socle Sz(q). Applied Mathematics and Computation, 351(??):153–161, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300451.

Liu:2019:CRN

[LLL19c] Meng Liu, Yusheng Li, and Qizhong Lin. Class Ramsey numbers of odd cycles in many colors. Applied Mathematics and Computation, 363(??): Article 124613, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306058.

Lin:2016:FTB

[LLLZ16] Xiangze Lin, Xueling Li, Shihua Li, and Yun Zou. Finite-time boundedness for switched systems with sector bounded nonlinearity and constant time delay. Applied Mathematics and Computation, 274(??):25–40, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014721.

Lan:2018:VBE

[LLM^{+}18] Yongxin Lan, Tao Li, Yuede Ma, Yongtang Shi, and Hua Wang. Vertexbased and edge-based centroids of graphs. Applied Mathematics and Computation, 331(??):445–456, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302182.

Lang:2016:NMA

[LLMS16] Rongling Lang, Tao Li, Desen Mo, and Yongtang Shi. A novel method for analyzing inverse problem of topological indices of graphs using competitive agglomeration. Applied Mathematics and Computation, 291(??):115–121, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304283.

Lei:2018:ALN

[LLMW18] Hui Lei, Tao Li, Yuede Ma, and Hua Wang. Analyzing lattice networks through substructures. Applied Mathematics and Computation, 329(??):297–314, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301140.

Lamnii:2016:NBO

[LLO16] A. Lamnii, M. Lamnii, and F. Oumellal. A new basis for osculatory interpolation problems and applications. Applied Mathematics and Computation, 283(??):355–368, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301795.

Lee:2017:ISC

[LLP17] Seok Young Lee, Won Il Lee, and PooGyeon Park.

Improved stability criteria for linear systems with interval time-varying delays: Generalized zero equalities approach. Applied Mathematics and Computation, 292(??):336–348, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304453.

Li:2015:NCD

[LLS15] Shasha Li, Xueliang Li, and Yongtang Shi. Note on the complexity of deciding the rainbow (vertex- ) connectedness for bipartite graphs. Applied Mathematics and Computation, 258(??):155–161, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500168X.

Lan:2019:TNS

[LLST19] Yongxin Lan, Tao Li, Yongtang Shi, and Jianhua Tu. The Turán number of star forests. Applied Mathematics and Computation, 348(??):270–274, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318310452.

Lachowicz:2019:SOS

[LLT19] Miroslaw Lachowicz, Henryk Leszczyński, and Krzysztof A. Topolski. Self-organization with small range interactions: Equilibria and creation of bipolarity. Applied Mathematics and Computation, 343(??):156–166, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308294.

Li:2018:TFD

[LLTL18] Yuanlu Li, Fawang Liu, Ian W. Turner, and Tao Li. Time-fractional diffusion equation for signal smoothing. Applied Mathematics and Computation, 326(??):108–116, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300201.

Liu:2019:VCE

[LLTW19] Ruifang Liu, Hong-Jian Lai, Yingzhi Tian, and Yang Wu. Vertexconnectivity and eigenvalues of graphs with fixed girth. Applied Mathematics and Computation, 344–345(??):141–149, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308324.

Li:2017:HRT

[LLW17] Wenjing Li, Xueliang Li, and Di Wu. Hardness result for the total rainbow k-connection of graphs. Applied Mathematics and Computation, 305(??):27– 31, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300942.

Li:2019:PRK

[LLW19] Xiuying Li, Haixia Li, and Boying Wu. Piecewise reproducing kernel method for linear impulsive delay differential equations with piecewise constant arguments. Applied Mathematics and Computation, 349(??):304–313, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311159.

Liu:2019:PTV

[LLWW19] Chen Liu, Li Li, Zhen Wang, and Ruiwu Wang. Pattern transitions in a vegetation system with cross-diffusion. Applied Mathematics and Computation, 342(??):255–262, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308178.

Lan:2019:NET

[LLWX19] Yongxin Lan, Tao Li, Hua Wang, and Chengyi Xia. A note on extremal trees with degree conditions. Applied Mathematics and Computation, 341:70–79, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318307409.

Li:2018:PSW

[LLX18a] Chan Li, Jin Liang, and Ti-Jun Xiao. Polynomial stability for wave equations with acoustic boundary conditions and boundary memory damping. Applied Mathematics and Computation, 321(??):593–601, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308019.

Liu:2018:NMP

[LLX18b] Na Liu, Wei Luo, and Qingxiang Xu. New multiplicative perturbation bounds for the generalized polar decomposition. Applied Mathematics and Computation, 339(??):242–258, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305812.

Liu:2019:SDC

[LLX^{+}19] Wenhui Liu, Junwei Lu, Shengyuan Xu, Yongmin Li, and Zhengqiang Zhang. Sampled-data controller design and stability analysis for nonlinear systems with input saturation and disturbances. Applied Mathematics and Computation, 360(??):14–27, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303066.

Li:2015:IBV

[LLY15] Zhiyuan Li, Yikan Liu, and Masahiro Yamamoto. Initial-boundary value problems for multi-term timefractional diffusion equations with positive constant coefficients. Applied Mathematics and Computation, 257(??):381–397, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016051.

Li:2019:CDI

[LLY19] Qiang Li, Zhijun Liu, and Sanling Yuan. Crossdiffusion induced Turing instability for a competition model with saturation effect. Applied Mathematics and Computation, 347(??):64–77, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309457.

Li:2015:GOF

[LLZ15] Erpeng Li, Lijun Long, and Jun Zhao. Global output-feedback stabilization for a class of switched uncertain nonlinear systems. Applied Mathematics and Computation, 256(??):551–564, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000533.

Liu:2016:MAP

[LLZ^{+}16] Chao Liu, Na Lu, Qingling Zhang, Jinna Li, and Peiyong Liu. Modeling and analysis in a prey-predator system with commercial harvesting and double time delays. Applied Mathematics and Computation, 281(??):77–101, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630039X.

Lin:2017:FTS

[LLZ17] Xiangze Lin, Shihua Li, and Yun Zou. Finite-time stabilization of switched linear time-delay systems with saturating actuators. Applied Mathematics and Computation, 299(??):66–79, April 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307159.

Li:2018:IEMb

[LLz18] Hongyi Li, Ling Li, and Di zhao. An improved EMD method with modified envelope algorithm based on C^{2} piecewise rational cubic spline interpolation for EMI signal decomposition. Applied Mathematics and Computation, 335(??):112–123, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303205.

Li:2019:DSE

[LLZ^{+}19] Qian Li, Xinzhi Liu, Qingxin Zhu, Shouming Zhong, and Dian Zhang. Distributed state estimation for stochastic discrete-time sensor networks with redundant channels. Applied Mathematics and Computation, 343(??):230–246, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308233.

Le:2018:CPA

[LLZD18a] Nhat-Tan Le, Xiaoping Lu, Song-Ping Zhu, and Duy-Minh Dang. Corrigendum to “Pricing American-style Parisian down-and-out call options” [Applied Mathematics and Computation 305 (2017) 330–347]. Applied Mathematics and Computation, 321(??):812, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307750. See

[LD17].

Li:2018:SES

[LLZD18b] Yueyang Li, Shuai Liu, Maiying Zhong, and Steven X. Ding. State estimation for stochastic discretetime systems with multiplicative noises and unknown inputs over fading channels. Applied Mathematics and Computation, 320(??):116–130, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306288.

Lukashchuk:2015:GCN

[LM15] S. Yu. Lukashchuk and A. V. Makunin. Group classification of nonlinear time-fractional diffusion equation with a source term. Applied Mathematics and Computation, 257(??):335–343, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016191.

Li:2018:LMP

[LM18] Qi Li and Liquan Mei. Local momentum-preserving algorithms for the GRLW equation. Applied Mathematics and Computation, 330(??):77–92, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301358.

Li:2019:ASSa

[LM19] Cheng-Liang Li and Chang-Feng Ma. An accelerated symmetric SORlike method for augmented systems. Applied Mathematics and Computation, 341(??):408–417, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318306428.

Liu:2019:SST

[LMD19] Linna Liu, Haoyi Mo, and Feiqi Deng. Splitstep theta method for stochastic delay integrodifferential equations with mean square exponential stability. Applied Mathematics and Computation, 353(??):320–328, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300906.

Lin:2017:SPB

[LMF17] Fubiao Lin, Sergey V.

Meleshko, and Adrian E. Flood. Symmetries of population balance equations for aggregation, breakage and growth processes. Applied Mathematics and Computation, 307(??):193–203, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301595.

Li:2018:TSE

[LML18a] Jingliang Li, Heping Ma, and Huiyuan Li. A triangular spectral element method for elliptic and Stokes problems. Applied Mathematics and Computation, 321(??):195–208, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307245.

Liu:2018:SSPb

[LML18b] Yan Liu, Jingling Mei, and Wenxue Li. Stochastic stabilization problem of complex networks without strong connectedness. Applied Mathematics and Computation, 332(??):304–315, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302637.

Lotfi:2015:VSK

[LMMR15] T. Lotfi, Á. A. Magreñán, K. Mahdiani, and J. Javier Rainer. A variant of Steffensen–King’s type family with accelerated sixth-order convergence and high efficiency index: Dynamic study and approach. Applied Mathematics and Computation, 252(??):347–353, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016877.

Liu:2018:MSH

[LMMZ18] Guojun Liu, Wentao Ma, Hailong Ma, and Lin Zhu.

A multiple-scale higher order polynomial collocation method for 2D and 3D elliptic partial differential equations with variable coefficients. Applied Mathematics and Computation, 331(??):430–444, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301863.

Lozano:2019:ALC

[LMS19] Antoni Lozano, Mercè Mora, and Carlos Seara. Antimagic labelings of caterpillars. Applied Mathematics and Computation, 347(??):734–740, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310178.

Liu:2018:RFT

[LMW18] Yangfan Liu, Yuechao Ma, and Yanning Wang. Reliable finite-time slidingmode control for singular time-delay system with sensor faults and randomly occurring nonlinearities. Applied Mathematics and Computation, 320(??):341–357, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306707.

Liu:2019:ECI

[LMW^{+}19] Jinzhuo Liu, Haoran Meng, Wei Wang, Zhongwen Xie, and Qian Yu. Evolution of cooperation on independent networks: the influence of asymmetric information sharing updating mechanism. Applied Mathematics and Computation, 340(??):234–241, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830554X.

Li:2017:GCG

[LMYW17] Hengzhe Li, Yingbin Ma, Weihua Yang, and Yifei Wang. The generalized 3-connectivity of graph products. Applied Mathematics and Computation, 295(??):77–83, February 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306002.

Liu:2019:IAN

[LMZ^{+}19] Wei Liu, Qian Ma, Guangming Zhuang, Junwei Lu, and Yuming Chu. An improved adaptive neural dynamic surface control for pure-feedback systems with full state constraints and disturbance. Applied Mathematics and Computation, 358(??):37–50, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302620.

Lin:2016:SBO

[LMZW16] Hongying Lin, Biao Mo, Bo Zhou, and Weiming Weng. Sharp bounds for ordinary and signless Laplacian spectral radii of uniform hypergraphs. Applied Mathematics and Computation, 285(??):217–227, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302090.

Liu:2019:SCB

[LMZZ19] Zhilong Liu, Jun Ma, Ge Zhang, and Yin Zhang. Synchronization control between two Chua’s circuits via capacitive coupling. Applied Mathematics and Computation, 360(??):94–106, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319303911.

Liu:2018:RDR

[LNWS18] Rui-Juan Liu, Zhuo-Yun Nie, Min Wu, and Jinhua She. Robust disturbance rejection for uncertain fractional-order systems. Applied Mathematics and Computation, 322(??):79– 88, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308032.

Lupas:2015:SFD

[LO15] Alina Alb Lupas and Gheorghe Oros. On special fuzzy differential subordinations using Salagean and Ruscheweyh operators. Applied Mathematics and Computation, 261(??):119–127, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004014.

Loch-Olszewska:2019:PDD

[LO19] Hanna Loch-Olszewska. Properties and distribution of the dynamical functional for the fractional Gaussian noise. Applied Mathematics and Computation, 356(??):252–271, September 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319302450.

Leitao:2018:DDC

[LOOGB18] Álvaro Leitao, Cornelis W.

Oosterlee, Luis Ortiz-Gracia, and Sander M. Bohte. On the data-driven COS method. Applied Mathematics and Computation, 317(??):68–84, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306227.

Lekomtsev:2015:CSW

[LP15a] Andrey Lekomtsev and Vladimir Pimenov. Convergence of the scheme with weights for the numerical solution of a heat conduction equation with delay for the case of variable coefficient of heat conductivity. Applied Mathematics and Computation, 256(??):83–93, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000107.

Liu:2015:UAA

[LP15b] Jia-Bao Liu and Xiang-Feng Pan. A unified approach to the asymptotic topological indices of various lattices. Applied Mathematics and Computation, 270(??):62–73, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010632.

Livieris:2015:NCN

[LP15c] Ioannis E. Livieris and Panagiotis Pintelas. A new class of nonmonotone conjugate gradient training algorithms. Applied Mathematics and Computation, 266(??):404–413, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006773.

Liu:2016:MKI

[LP16a] Jia-Bao Liu and Xiang-Feng Pan. Minimizing Kirchhoff index among graphs with a given vertex bipartiteness. Applied Mathematics and Computation, 291(??):84–88, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303885.

Lopez:2016:PIH

[LP16b] José L. López and Pedro J.

Pagola. The Pearcey integral in the highly oscillatory region. Applied Mathematics and Computation, 275(??):404–410, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015829.

Liu:2015:ALE

[LPHH15] Jia-Bao Liu, Xiang-Feng Pan, Fu-Tao Hu, and Feng-Feng Hu. Asymptotic Laplacian-energy-like invariant of lattices. Applied Mathematics and Computation, 253(??):205–214, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016890.

Lee:2018:NBI

[LPJ18] Tae H. Lee, Ju H.

Park, and Hoyoul Jung.

Network-based H_{∞} state estimation for neural networks using imperfect measurement. Applied Mathematics and Computation, 316(??):205–214, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305854.

Lee:2016:MSS

[LPKS16] S. H. Lee, M. J. Park, O. M. Kwon, and R. Sakthivel. Master–slave synchronization for nonlinear systems via reliable control with Gaussian stochastic process. Applied Mathematics and Computation, 290(??):439–459, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303897.

Lu:2017:SSP

[LPL17] Yang Lu, Kasia A. Pawelek, and Shengqiang Liu. A stage-structured predator– prey model with predation over juvenile prey. Applied Mathematics and Computation, 297(??):115–130, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306427.

Lu:2018:CTR

[LpLxG18] Xin Lu, Xing ping Liu, and Tong xiang Gu. The convergence theory for the restricted version of the overlapping Schur complement preconditioner. Applied Mathematics and Computation, 339(??):410–421, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306039.

Li:2019:SFD

[LPS19] Jian Li, Kunpeng Pan, and Qingyu Su. Sensor fault detection and estimation for switched power electronics systems based on sliding mode observer. Applied Mathematics and Computation, 353(??):282–294, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301080.

Lv:2015:SBP

[LPW15] Jian Lv, Li-Ping Pang, and Jin-He Wang. Special backtracking proximal bundle method for nonconvex maximum eigenvalue optimization. Applied Mathematics and Computation, 265(??):635–651, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500750X.

Lazzaro:2019:FSM

[LPZ19] D. Lazzaro, E. Loli Piccolomini, and F. Zama. A fast splitting method for efficient Split Bregman iterations. Applied Mathematics and Computation, 357(??):139–146, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302735.

Liang:2016:MKO

[LPZL16] Xiyin Liang, Yongzhen Pei, Meixia Zhu, and Yunfei Lv. Multiple kinds of optimal impulse control strategies on plantpest-predator model with eco-epidemiology. Applied Mathematics and Computation, 287–288(??):1–11, September 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300316302843.

Li:2015:EPF

[LQ15] Jing Li and Jiangang Qi. Eigenvalue problems for fractional differential equations with right and left fractional derivatives. Applied Mathematics and Computation, 256(??):1– 10, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000041.

Li:2018:SAC

[LQC^{+}18] Lei Li, Wenhai Qi, Xiaoming Chen, Yonggui Kao, Xianwen Gao, and Yunliang Wei. Stability analysis and control synthesis for positive semi-Markov jump systems with time-varying delay. Applied Mathematics and Computation, 332(??):363–375, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301656.

Li:2018:BEE

[LQG18] Jing Li, Lu Qiao, and Nan Gao. Bounds of the extended Estrada index of graphs. Applied Mathematics and Computation, 317(??):143–149, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306367.

Li:2015:NIG

[LQW^{+}15] Xueliang Li, Zhongmei Qin, Meiqin Wei, Ivan Gutman, and Matthias Dehmer. Novel inequalities for generalized graph entropies — graph energies and topological indices. Applied Mathematics and Computation, 259(??):470–479, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002520.

Li:2016:EHC

[LQZ16] Wei Li, Zhongmei Qin, and Heping Zhang. Extremal hexagonal chains with respect to the coefficients sum of the permanental polynomial. Applied Mathematics and Computation, 291(??):30–38, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300316303964.

Li:2018:IEMa

[LQZ^{+}18] Hongyi Li, Xuyao Qin, Di Zhao, Jiaxin Chen, and Pidong Wang. An improved empirical mode decomposition method based on the cubic trigonometric B-spline interpolation algorithm. Applied Mathematics and Computation, 332(??):406–419, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301498.

Llibre:2015:NAK

[LR15] Jaume Llibre and Ana Rodrigues. A nonautonomous kind of Duffing equation. Applied Mathematics and Computation, 251(??):669–674, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015124.

Li:2016:SCE

[LR16a] Pingrun Li and Guangbin Ren. Some classes of equations of discrete type with harmonic singular operator and convolution. Applied Mathematics and Computation, 284(??):185–194, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301965.

Li:2016:TGB

[LR16b] Xiaoli Li and Hongxing Rui. A two-grid block-centered finite difference method for nonlinear non-Fickian flow model. Applied Mathematics and Computation, 281(??):300–313, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300650.

Li:2017:CBC

[LR17] Xiaoli Li and Hongxing Rui. Characteristic block-centered finite difference method for compressible miscible displacement in porous media. Applied Mathematics and Computation, 314(??):391–407, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304629.

Lakzian:2019:SFP

[LR19a] Hossein Lakzian and B. E.

Rhoades. Some fixed point theorems using weaker Meir–Keeler function in metric spaces with w-distance. Applied Mathematics and Computation, 342(??):18–25, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830537X.

Li:2019:BCF

[LR19b] Xiaoli Li and Hongxing Rui. A blockcentered finite difference method for the nonlinear Sobolev equation on nonuniform rectangular grids. Applied Mathematics and Computation, 363 (??):Article 124607, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305995.

Li:2019:SCC

[LR19c] Xiaoli Li and Hongxing Rui. Stability and convergence of characteristic MAC scheme and postprocessing for the Oseen equations on nonuniform grids. Applied Mathematics and Computation, 342(??):94–111, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307628.

Lenzi:2018:NDE

[LRF^{+}18] E. K. Lenzi, M. A. Ribeiro, M. E. K. Fuziki, M. K. Lenzi, and H. V. Ribeiro. Nonlinear diffusion equation with reaction terms: Analytical and numerical results. Applied Mathematics and Computation, 330(??):254–265, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301504.

Lu:2015:MFB

[LRH15] Wen Lu, Yong Ren, and Lanying Hu. Mean-field backward stochastic differential equations in general probability spaces. Applied Mathematics and Computation, 263(??):1– 11, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004658.

Lin:2018:NMM

[LRL18] Ji Lin, S. Y. Reutskiy, and Jun Lu. A novel meshless method for fully nonlinear advectiondiffusion-reaction problems to model transfer in anisotropic media. Applied Mathematics and Computation, 339(??):451–458, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306106.

Li:2015:FRD

[LS15a] Hong-Hai Li and Li Su.

Further results on digraphs with completely real Laplacian spectra. Applied Mathematics and Computation, 265(??): 630–634, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031500733X.

Lin:2015:SGO

[LS15b] Hongwei Lin and Linjie Sun. Searching globally optimal parameter sequence for defeating Runge phenomenon by immunity genetic algorithm. Applied Mathematics and Computation, 264(??):85–98, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005317.

Liu:2016:HSS

[LS16] Xinzhi Liu and Peter Stechlinski. Hybrid stabilization and synchronization of nonlinear systems with unbounded delays. Applied Mathematics and Computation, 280(??):140–161, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300236.

Lenarduzzi:2017:KBA

[LS17a] Licia Lenarduzzi and Robert Schaback. Kernelbased adaptive approximation of functions with discontinuities. Applied Mathematics and Computation, 307(??):113–123, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317301546.

Lord:2017:NES

[LS17b] G. J. Lord and D. Stone.

New efficient substepping methods for exponential timestepping. Applied Mathematics and Computation, 307(??):342–365, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301637.

Li:2019:PVS

[LS19a] Daniel Li Li and Erfang Shan. The position value and the structures of graphs. Applied Mathematics and Computation, 356(??):190–197, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302486.

Liu:2019:IPC

[LS19b] Cui Liu and Zhiyong Si. An incremental pressure correction finite element method for the time-dependent Oldroyd flows. Applied Mathematics and Computation, 351(??):99–115, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300359.

Liu:2019:CCS

[LS19c] Yifan Liu and Housheng Su. Containment control of second-order multiagent systems via intermittent sampled position data communication. Applied Mathematics and Computation, 362 (??):Article 124522, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305016.

Li:2015:LBS

[LSAR15] Xiaodi Li, Jianhua Shen, Haydar Akca, and R. Rakkiyappan. LMI-based stability for singularly perturbed nonlinear impulsive differential systems with delays of small parameter. Applied Mathematics and Computation, 250(??):798–804, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031401488X.

Li:2015:NFF

[LSCK15] Feng Li, Hao Shen, Mengshen Chen, and Qingkai Kong. Non-fragile finitetime ℓ_{2}–ℓ_{∞} state estimation for discretetime Markov jump neural networks with unreliable communication links. Applied Mathematics and Computation, 271(??):467–481, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012655.

Li:2015:EHS

[LSfW15] Tiexiang Li, Juntao Sun, and Tsung fang Wu. Existence of homoclinic solutions for a fourth order differential equation with a parameter. Applied Mathematics and Computation, 251(??):499–506, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015884.

Li:2017:MUM

[LSjL17] Chen Li, Xin-Hui Shao, and Chang jun Li. Modified Uzawa methods for saddle point problems. Applied Mathematics and Computation, 315(??): 507–515, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300317304897.

Li:2019:SNT

[LSL19a] Zhao-Yan Li, Shengnan Shang, and James Lam. On stability of neutral-type linear stochastic time-delay systems with three different delays. Applied Mathematics and Computation, 360(??):147–166, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319303583.

Long:2019:SOC

[LSL19b] Mingkang Long, Housheng Su, and Bo Liu. Secondorder controllability of two-time-scale multi-agent systems. Applied Mathematics and Computation, 343(??):299–313, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308117.

Liu:2019:ADC

[LSLL19] Chen Liu, Juan Shi, Tong Li, and Jinzhuo Liu. Aspiration driven coevolution resolves social dilemmas in networks. Applied Mathematics and Computation, 342(??):247–254, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308129.

Li:2018:RHC

[LSLZ18] Min Li, Feng Shu, Duyu Liu, and Shouming Zhong. Robust H_{∞} control of T–S fuzzy systems with input time-varying delays: A delay partitioning method. Applied Mathematics and Computation, 321(??):209–222, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307622.

Lu:2015:SNA

[LSM15] Dawei Lu, Lixin Song, and Congxu Ma. Some new asymptotic approximations of the gamma function based on Nemes’ formula, Ramanujan’s formula and Burnside’s formula. Applied Mathematics and Computation, 253(??):1–7, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314017317.

Li:2018:ESS

[LSMT18] Yanan Li, Yuangong Sun, Fanwei Meng, and Yazhou Tian. Exponential stabilization of switched time-varying systems with delays and disturbances. Applied Mathematics and Computation, 324(??):131–140, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308706.

Li:2018:PIE

[LSR18] Xiaodi Li, Jianhua Shen, and R. Rakkiyappan. Persistent impulsive effects on stability of functional differential equations with finite or infinite delay. Applied Mathematics and Computation, 329(??):14– 22, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300572.

Liu:2016:LBM

[LSW16] Fang Liu, Weiping Shi, and Fangfang Wu. A lattice Boltzmann model for the generalized Boussinesq equation. Applied Mathematics and Computation, 274(??):331–342, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015076.

Li:2019:RBA

[LSX19] Xiaopeng Li, Shiwen Sun, and Chengyi Xia. Reputation-based adaptive adjustment of link weight among individuals promotes the cooperation in spatial social dilemmas. Applied Mathematics and Computation, 361(??):810–820, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930503X.

Lu:2015:NSC

[LSY15] Dawei Lu, Lixin Song, and Yang Yu. New sequences with continued fraction towards Euler’s constant. Applied Mathematics and Computation, 259(??):12– 20, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001745.

Liu:2015:CRH

[LSZ15] Suzhen Liu, Yongzhong Song, and Xiaojian Zhou. Convergence radius of Halley’s method for multiple roots under center-Hölder continuous condition. Applied Mathematics and Computation, 265(??):1011–1018, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007894.

Li:2018:ERG

[LSZ^{+}18] Xiaoqing Li, Kun She, Shouming Zhong, Kaibo Shi, Wei Kang, Jung Cheng, and Yongbin Yu. Extended robust global exponential stability for uncertain switched memristor-based neural networks with time-varying delays. Applied Mathematics and Computation, 325(??):271–290, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308913.

Li:2019:SCM

[LSZ^{+}19a] Feng Li, Shuai Song, Jianrong Zhao, Shengyuan Xu, and Zhengqiang Zhang.

Synchronization control for Markov jump neural networks subject to HMM observation and partially known detection probabilities. Applied Mathematics and Computation, 360(??): 1–13, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319303145.

Li:2019:DFD

[LSZ^{+}19b] Yueyang Li, Xinmin Song, Zhijie Zhang, Dong Zhao, and Zhonghua Wang. H_{∞} deconvolution filter design for uncertain linear discrete time-variant systems: a Krein space approach. Applied Mathematics and Computation, 361(??):131–143, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304084.

Liang:2018:ADN

[LSZF18] Jiarui Liang, Songhe Song, Weien Zhou, and Hao Fu. Analysis of the damped nonlinear space-fractional Schrödinger equation. Applied Mathematics and Computation, 320(??): 495–511, March 1, 2018.

CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317306999.

Li:2017:CMJ

[LSZY17] Lingchun Li, Mouquan Shen, Guangming Zhang, and Shen Yan. H_{∞} control of Markov jump systems with time-varying delay and incomplete transition probabilities. Applied Mathematics and Computation, 301(??):95–106, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307615.

Liu:2015:DWO

[LT15] Wenjun Liu and Adnan Tuna. Diamond-α weighted Ostrowski type and Grüss type inequalities on time scales. Applied Mathematics and Computation, 270(??):251–260, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315010541.

Lord:2018:MSI

[LT18a] Gabriel J. Lord and Antoine Tambue. A modified semi-implicit Euler– Maruyama scheme for finite element discretization of SPDEs with additive noise. Applied Mathematics and Computation, 332(??):105–122, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301796.

Luo:2018:ROE

[LT18b] Zhendong Luo and Fei Teng. A reduced-order extrapolated finite difference iterative scheme based on POD method for 2D Sobolev equation. Applied Mathematics and Computation, 329(??):374–383, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301243.

Liu:2019:SRC

[LT19] Zi-Long Liu and Fang Tian. Some rainbow cycles in r-good edge-colorings. Applied Mathematics and Computation, 362(??): Article 124566, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300319305491.

Liang:2016:PBC

[LTC16] Juhua Liang, Sanyi Tang, and Robert A. Cheke. Pure Bt-crop and mixed seed sowing strategies for optimal economic profit in the face of pest resistance to pesticides and Bt-corn. Applied Mathematics and Computation, 283(??):6– 21, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301266.

Liu:2016:RBC

[LTL16] Yujun Liu, Tong Tang, and Shan Li. A remark on blow-up criterion for 3D compressible magnetohydrodynamic flows with vacuum. Applied Mathematics and Computation, 290(??):189–200, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303484.

Li:2019:HDD

[LTQF19] Tao Li, Xiaoling Tang, Wei Qian, and Shumin Fei. Hybrid-delay-dependent approach to synchronization in distributed delay neutral neural networks. Applied Mathematics and Computation, 347(??):449–463, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309925.

Liu:2015:SSM

[LTUZ15] Zhenhai Liu, Salih Tatar, Süleyman Ulusoy, and Mustafa Zeki. Structural stability for the Morris– Lecar neuron model. Applied Mathematics and Computation, 270(??): 261–268, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315010863.

Li:2016:GCS

[LTY16] Shasha Li, Jianhua Tu, and Chenyan Yu. The generalized 3-connectivity of star graphs and bubblesort graphs. Applied Mathematics and Computation, 274(??):41–46, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014782.

Luo:2017:NFA

[LTZ^{+}17] Jinnan Luo, Wenhong Tian, Shouming Zhong, Kaibo Shi, Hao Chen, Xian-Ming Gu, and Wenqin Wang. Non-fragile asynchronous H_{∞} control for uncertain stochastic memory systems with Bernoulli distribution. Applied Mathematics and Computation, 312(??): 109–128, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317302916.

Lu:2015:NCV

[Lu15] Lizheng Lu. A note on curvature variation minimizing cubic Hermite interpolants. Applied Mathematics and Computation, 259(??):596–599, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002891.

Lu:2016:PPP

[Lu16] Peng Lu. Predicting peak of participants in collective action. Applied Mathematics and Computation, 274(??):318–330, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014770.

Lu:2019:HJS

[Lu19] Peng Lu. Heterogeneity, judgment, and social trust of agents in rumor spreading. Applied Mathematics and Computation, 350(??):447–461, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309536.

Lukas:2015:OPS

[Luk15] Petr Lukás. Optimization of parameters in SDFEM for different spaces of parameters. Applied Mathematics and Computation, 267(??):711–715, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401738X.

Luo:2015:EEA

[Luo15]  Qiu-Ming Luo.  El-

liptic extensions of the Apostol–Bernoulli and Apostol–Euler polynomials. Applied Mathematics and Computation, 261(??):156–166, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004038.

Luo:2019:SMG

[Luo19] Tianjiao Luo. Stabilization of multi-group models with multiple dispersal and stochastic perturbation via feedback control based on discrete-time state observations. Applied Mathematics and Computation, 354(??):396–410, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300670.

Lv:2016:IMH

[Lv16] Xiang Lv. Infinitely many homoclinic solutions for a class of subquadratic second-order Hamiltonian systems. Applied Mathematics and Computation, 290(??):298–306, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630385X.

Lv:2018:EPS

[Lv18] Xiang Lv. Existence of periodic solutions for a class of second-order p-Laplacian systems. Applied Mathematics and Computation, 338(??): 487–506, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318305216.

Lahasan:2016:RFP

[LVAB^{+}16] Badr Mohammed Lahasan, Ibrahim Venkat, Mohammed Azmi Al-Betar, Syaheerah Lebai Lutfi, and Philippe De Wilde. Recognizing faces prone to occlusions and common variations using optimal face subgraphs. Applied Mathematics and Computation, 283(??):316–332, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301588.

Loi:2015:NPD

[LVC15] Nguyen Van Loi, Mai Quoc Vu, and Pham Tuan Cuong. Nonlocal problem for differential complementarity systems. Applied Mathematics and Computation, 265(??):799–806, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007742.

Lisei:2018:LMS

[LVV18] Hannelore Lisei, Csaba Varga, and Orsolya Vas. Localization method for the solutions of nonhomogeneous operator equations. Applied Mathematics and Computation, 329(??):64–83, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300523.

Lesmana:2015:PAN

[LW15a] Donny Citra Lesmana and Song Wang. Penalty approach to a nonlinear obstacle problem governing American put option valuation under transaction costs. Applied Mathematics and Computation, 251(??):318–330, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015926.

Lewandowski:2015:AII

[LW15b] M. Lewandowski and J. Walczak. Analysis of the impact of interpolator’s order on the accuracy of electric current spectrum estimation method in the presence of noise. Applied Mathematics and Computation, 267(??):554–561, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000909.

Li:2015:TBT

[LW15c] Cui-Xia Li and Shi-Liang Wu. Two block triangular preconditioners for asymmetric saddle point problems. Applied Mathematics and Computation, 269(??):456–463, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010152.

Li:2015:GSN

[LW15d] Yuan-Min Li and Deyun Wei. A generalized smoothing Newton method for the symmetric cone complementarity problem. Applied Mathematics and Computation, 264(??):335–345, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005676.

Li:2016:DLI

[LW16] Haochen Li and Yushun Wang. A discrete line integral method of order two for the Lorentz force system. Applied Mathematics and Computation, 291(??):207–212, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304246.

Li:2017:FTF

[LW17a] Jian Li and Chun-Yu Wu. Finite-time fault detection filter design for discretetime interconnected systems with average dwell time. Applied Mathematics and Computation, 313(??):259–270, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317303831.

Liu:2017:CPP

[LW17b] Xiaogang Liu and Tingzeng Wu. Computing the permanental polynomials of graphs. Applied Mathematics and Computation, 304(??):103–113, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730070X.

Lewandowski:2018:OBF

[LW18a] Michal Lewandowski and Janusz Walczak. Optimal base frequency estimation of an electrical signal based on Prony’s estimator and a FIR filter. Applied Mathematics and Computation, 319(??):551–561, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730406X.

Li:2018:CCR

[LW18b] Hengzhe Li and Jiajia Wang. The λ_{3}- connectivity and κ_{3}-connectivity of recursive circulants. Applied Mathematics and Computation, 339(??): 738–749, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318306313.

Li:2018:SRA

[LW18c] Jin Li and Zhaoqing Wang. Simpson’s rule to approximate Hilbert integral and its application. Applied Mathematics and Computation, 339(??):390–397, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305691.

Li:2018:EDM

[LW18d] Mengmeng Li and Jin-Rong Wang. Exploring delayed Mittag-Leffler type matrix functions to study finite time stability of fractional delay differential equations. Applied Mathematics and Computation, 324(??):254–265, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308536.

Li:2018:EOC

[LW18e] Shuchao Li and Wei Wei. Extremal octagonal chains with respect to the coefficients sum of the permanental polynomial. Applied Mathematics and Computation, 328(??):45–57, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300547.

Li:2018:NPN

[LW18f] Xuhao Li and Patricia J. Y. Wong. A non-polynomial numerical scheme for fourth-order fractional diffusion-wave model. Applied Mathematics and Computation, 331(??):80–95, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301541.

Li:2018:ITD

[LW18g] Y. S. Li and T. Wei.

An inverse time-dependent source problem for a time– space fractional diffusion equation. Applied Mathematics and Computation, 336(??):257–271, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304193.

Liu:2018:HOI

[LW18h] Wenjie Liu and Boying Wu. High-order implicit Galerkin–Legendre spectral method for the twodimensional Schrödinger equation. Applied Mathematics and Computation, 324(??):59–68, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317308688.

Li:2019:EEPb

[LW19a] Ting Li and Bin Wang. Efficient energy-preserving methods for chargedparticle dynamics. Applied Mathematics and Computation, 361(??):703–714, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304874.

Li:2019:NPS

[LW19b] Xuhao Li and Patricia J. Y. Wong. Nonpolynomial spline approach in two-dimensional fractional sub-diffusion problems. Applied Mathematics and Computation, 357(??):222–242, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302528.

Li:2019:SNG

[LW19c] Zizun Li and Wu-Sheng Wang. Some nonlinear Gronwall–Bellman type retarded integral inequalities with power and their applications. Applied Mathematics and Computation, 347(??):839–852, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308853.

Liang:2019:BCG

[LW19d] Mingjie Liang and Bingyao Wu. Brownian cuboid and its geometric characteristics. Applied Mathematics and Computation, 358(??):278–286, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303297.

Li:2016:OCG

[LWG16] Hong-Hai Li, Qian-Qian Wu, and Ivan Gutman. On ordering of complements of graphs with respect to matching numbers. Applied Mathematics and Computation, 282(??):167–174, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300316300881.

Li:2017:RHE

[LWGS17] Xianghua Li, Zhen Wang, Chao Gao, and Lei Shi. Reasoning human emotional responses from large-scale social and public media. Applied Mathematics and Computation, 310(??):182–193, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302151.

Li:2015:ULM

[LWK15] Ji-Cheng Li, Na-Na Wang, and Xu Kong. Uzawa– Low method and preconditioned Uzawa–Low method for three-order block saddle point problem. Applied Mathematics and Computation, 269(??):626–636, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010292.

Li:2018:HBA

[LWL^{+}18] Li Li, Zhen Wang, Yuxia Li, Hao Shen, and Junwei Lu. Hopf bifurcation analysis of a complex-valued neural network model with discrete and distributed delays. Applied Mathematics and Computation, 330(??):152–169, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301310.

Li:2019:EPC

[LWLM19] Yubai Li, Wei-Tao Wu, Xianglei Liu, and Mehrdad Massoudi. The effects of particle concentration and various fluxes on the flow of a fluid–solid suspension. Applied Mathematics and Computation, 358(??):151–160, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302991.

Li:2017:SRD

[LWM17] Dan Li, Guoping Wang, and Jixiang Meng. Some results on the distance and distance signless Laplacian spectral radius of graphs and digraphs. Applied Mathematics and Computation, 293(??):218–225, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300316305203.

Liu:2018:PCP

[LWM18] Fengxia Liu, Baoyindureng Wu, and Jixiang Meng. Partitioning the Cartesian product of a tree and a cycle. Applied Mathematics and Computation, 332(??):90–95, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301802.

Li:2016:DFO

[LWPL16] Yong Li, Lianwen Wang, Liuyong Pang, and Sanhong Liu. The data fitting and optimal control of a hand, foot and mouth disease (HFMD) model with stage structure. Applied Mathematics and Computation, 276(??):61–74, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015933.

Lee:2018:EHL

[LWQ18] Cheuk-Yu Lee, Hui Wang, and Qing-Hua Qin. Efficient hypersingular line and surface integrals direct evaluation by complex variable differentiation method. Applied Mathematics and Computation, 316(??):256–281, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305787.

Li:2015:PDS

[LWS15] Lei Li, Qing-Wen Wang, and Shu-Qian Shen. On positive definite solutions of the nonlinear matrix equations X ± A^{∗}X^{q}A = Q. Applied Mathematics and Computation, 271 (??):556–566, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012205.

Liu:2018:NEP

[LWS18] Changying Liu, Xinyuan Wu, and Wei Shi. New energy-preserving algorithms for nonlinear Hamiltonian wave equation equipped with Neumann boundary conditions. Applied Mathematics and Computation, 339(??): 568–587, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300318306258.

Liu:2016:PAE

[LWSM16] Yong Liu, Zongtian Wei, Jiarong Shi, and Anchan Mai. A polynomial algorithm of edge-neighbor-scattering number of trees. Applied Mathematics and Computation, 283(??):1–5, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301242.

Liu:2019:ILC

[LWSO19] Shengda Liu, JinRong Wang, Dong Shen, and Donal O’Regan. Iterative learning control for differential inclusions of parabolic type with noninstantaneous impulses. Applied Mathematics and Computation, 350(??):48– 59, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318311196.

Li:2018:IIS

[LWSX18a] Chao Li, Li Wang, Shiwen Sun, and Chengyi Xia. Identification of influential spreaders based on classified neighbors in real-world complex networks. Applied Mathematics and Computation, 320(??):512–523, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306884.

Li:2018:CCI

[LWSX18b] Jie Li, Juan Wang, Shiwen Sun, and Chengyi Xia. Cascading crashes induced by the individual heterogeneity in complex networks. Applied Mathematics and Computation, 323(??):182–192, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308482.

Li:2015:SCL

[LWT^{+}15] Zhi-Yuan Li, Yu-Lan Wang, Fu-Gui Tan, Xiao-Hui Wan, Hao Yu, and Jun-Sheng Duan. Solving a class of linear nonlocal boundary value problems using the reproducing kernel. Applied Mathematics and Computation, 265(??):1098–1105, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007481.

Lin:2015:FDE

[LWW15a] Zeng Lin, JinRong Wang, and Wei Wei. Fractional differential equation models with pulses and criterion for pest management. Applied Mathematics and Computation, 257(??):398–408, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014623.

Lin:2015:MBG

[LWW15b] Zeng Lin, JinRong Wang, and Wei Wei. Multipoint BVPs for generalized impulsive fractional differential equations. Applied Mathematics and Computation, 258(??):608–616, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017536.

Long:2015:SNE

[LWW15c] Qiang Long, Changzhi Wu, and Xiangyu Wang. A system of nonsmooth equations solver based upon subgradient method. Applied Mathematics and Computation, 251(??): 284–299, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314015963.

Li:2019:SER

[LWW19] Shuchao Li, Hua Wang, and Shujing Wang. Some extremal ratios of the distance and subtree problems in binary trees. Applied Mathematics and Computation, 361(??): 232–245, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL https://www.sciencedirect.
com/science/article/
pii/S0096300319304163.

Liu:2019:MLR

[LWWL19] Yisi Liu, Xiaojun Wang, Lei Wang, and Dongliang Liu. A modified leaky ReLU scheme (MLRS) for topology optimization with multiple materials. Applied Mathematics and Computation, 352(??):188–204, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300475.

Liu:2015:NDS

[LWWW15] Jianjun Liu, Changzhi Wu, Guoning Wu, and Xiangyu Wang. A novel differential search algorithm and applications for structure design. Applied Mathematics and Computation, 268(??):246–269, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008176.

Lu:2017:LSP

[LWWZ17] Jinna Lu, Lu Wang, Yi-Ling Wang, and Xiaoguang Zhang. Logit selection promotes cooperation in voluntary public goods game. Applied Mathematics and Computation, 310(??):134–138, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302795.

Liu:2019:MCR

[LWY^{+}19] Wanping Liu, Xiao Wu, Wu Yang, Xiaofei Zhu, and Shouming Zhong. Modeling cyber rumor spreading over mobile social networks: a compartment approach. Applied Mathematics and Computation, 343(??):214–229, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308270.

Lu:2015:BOG

[LWZ15] Yong Lu, Ligong Wang, and Qiannan Zhou. Bicyclic oriented graphs with skew-rank 6. Applied Mathematics and Computation, 270(??):899–908, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315011613.

Li:2016:MLM

[LWZ16] Jingshi Li, Xiaoshen Wang, and Kai Zhang. Multilevel Monte Carlo weak Galerkin method for elliptic equations with stochastic jump coefficients. Applied Mathematics and Computation, 275(??): 181–194, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315015660.

Luo:2018:CEC

[LWZ18] Chao Luo, Xingyuan Wang, and Yuanjie Zheng.

Co-evolution of cooperation and limited resources on interdependent networks. Applied Mathematics and Computation, 316(??):174–185, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305696.

Lai:2019:NMT

[LWZ19a] Junyu Lai, Justin W. L. Wan, and Shuhua Zhang. Numerical methods for two person games arising from transboundary pollution with emission permit trading. Applied Mathematics and Computation, 350(??):11–31, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311147.

Li:2019:BTS

[LWZ^{+}19b] Yingying Li, Zhouchao Wei, Wei Zhang, Matjaz Perc, and Robert Repnik. Bogdanov–Takens singularity in the Hindmarsh– Rose neuron with time delay. Applied Mathematics and Computation, 354(??):180–188, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930150X.

Long:2018:SAC

[LWZZ18] Shaohua Long, Yunlong Wu, Shouming Zhong, and Dian Zhang. Stability analysis for a class of neutral type singular systems with time-varying delay. Applied Mathematics and Computation, 339(??):105–112, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305472.

Lang:2015:QBS

[LX15] Feng-Gong Lang and Xiao-Ping Xu. Quintic Bspline method for integro interpolation. Applied Mathematics and Computation, 263(??):353–360, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500538X.

Lee:2016:FEE

[LX16a] Hyesuk Lee and Shuhan Xu. Finite element error estimation for quasi-Newtonian fluid-structure interaction problems. Applied Mathematics and Computation, 274(??):93– 105, February 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315014228.

Li:2016:SGG

[LX16b] Shuo Li and Zhengrong Xiang. Stability, l_{1}- gain and l_{∞}-gain analysis for discrete-time positive switched singular delayed systems. Applied Mathematics and Computation, 275(??):95–106, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015556.

Li:2017:EFTa

[LX17] Pingshan Li and Min Xu. Edge-fault-tolerant edgebipancyclicity of balanced hypercubes. Applied Mathematics and Computation, 307(??):180–192, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301583.

Lang:2018:SSQ

[LX18] Feng-Gong Lang and Xiao-Ping Xu. On the superconvergence of some quadratic integro-splines at the midknots of a uniform partition. Applied Mathematics and Computation, 338(??):882–902, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305356.

Li:2019:EMH

[LX19a] Bin Li and Shuhuang Xiang. Efficient methods for highly oscillatory integrals with weakly singular and hypersingular kernels. Applied Mathematics and Computation, 362 (??):Article 124499, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304783.

Li:2019:BAH

[LX19b] Qian Li and Yanni Xiao.

Bifurcation analyses and hormetic effects of a discrete-time tumor model. Applied Mathematics and Computation, 363(??): Article 124618, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319306101.

Liu:2019:CCT

[LX19c] Guidong Liu and Shuhuang Xiang. Clenshaw–Curtistype quadrature rule for hypersingular integrals with highly oscillatory kernels. Applied Mathematics and Computation, 340(??):251–267, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830643X.

Liu:2019:NGF

[LX19d] Xiaomei Liu and Naiming Xie. A nonlinear grey forecasting model with double shape parameters and its application. Applied Mathematics and Computation, 360(??):203–212, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304059.

Liang:2019:DBS

[LXC^{+}19] Xingyue Liang, Jianwei Xia, Guoliang Chen, Huasheng Zhang, and Zhen Wang. Dissipativitybased sampled-data control for fuzzy Markovian jump systems. Applied Mathematics and Computation, 361(??):552–564, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319304400.

Li:2017:TSD

[LXCK17] Jingjian Li, Shangjin Xu, Mengyue Cao, and Zhe Kang. On tetravalent symmetric dihedrants. Applied Mathematics and Computation, 306(??):49–55, August 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301315.

Li:2019:FTS

[LXD19] Haitao Li, Xiaojing Xu, and Xueying Ding. Finitetime stability analysis of stochastic switched boolean networks with impulsive effect. Applied Mathematics and Computation, 347(??):557–565, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318309913.

Liu:2017:SDE

[LXJ^{+}17] Songnan Liu, Xiaojie Xu, Daqing Jiang, Tasawar Hayat, and Bashir Ahmad. Stationary distribution and extinction of the DS-I-A model disease with periodic parameter function and Markovian switching. Applied Mathematics and Computation, 311(??):66–84, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302813.

Liu:2017:CLS

[LXL17] Fushui Liu, Yang Xu, and Yikai Li. A coupled levelset and volume-of-fluid method for simulating axisymmetric incompressible two-phase flows. Applied Mathematics and Computation, 293(??):112–130, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630501X.

Levitin:2015:OCS

[LXPD15] Gregory Levitin, Liudong Xing, Sun Peng, and Yuanshun Dai. Optimal choice of standby modes in 1-out-of-N system with respect to mission reliability and cost. Applied Mathematics and Computation, 258(??):587–596, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002349.

Lin:2017:TDH

[LXT17] Jiazhe Lin, Rui Xu, and Xiaohong Tian. Threshold dynamics of an HIV-1 virus model with both virus-to-cell and cell-to-cell transmissions, intracellular delay, and humoral immunity. Applied Mathematics and Computation, 315(??):516–530, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305556.

Liu:2018:HDB

[LXTF18] Jinliang Liu, Jilei Xia, Engang Tian, and Shumin Fei. Hybrid-driven-based H_{∞} filter design for neural networks subject to deception attacks. Applied Mathematics and Computation, 320(??):158–174, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306276.

Liu:2018:NIR

[LXW^{+}18] Guobao Liu, Shengyuan Xu, Yunliang Wei, Zhidong Qi, and Zhengqiang Zhang. New insight into reachable set estimation for uncertain singular time-delay systems. Applied Mathematics and Computation, 320(??):769–780, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307348.

Li:2019:NSR

[LXZ19a] Binjie Li, Xiaoping Xie, and Shiquan Zhang. A new smoothness result for Caputo-type fractional ordinary differential equations. Applied Mathematics and Computation, 349(??):408–420, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318311135.

Liu:2019:CLT

[LXZ19b] Lu Liu, Dingyu Xue, and Shuo Zhang. Closedloop time response analysis of irrational fractionalorder systems with numerical Laplace transform technique. Applied Mathematics and Computation, 350(??):133–152, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309949.

Lu:2019:CSI

[LXZ19c] Rong-Wu Lu, Xi-Xiang Xu, and Ning Zhang. Construction of solutions for an integrable differentialdifference equation by Darboux–Bäcklund transformation. Applied Mathematics and Computation, 361(??):389–397, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304515.

Liu:2019:DAO

[LXZL19] Chao Liu, Xinying Xun, Qingling Zhang, and Yuanke Li. Dynamical analysis and optimal control in a hybrid stochastic double delayed bioeconomic system with impulsive contaminants emission and Lévy jumps.

Applied Mathematics and Computation, 352(??):99– 118, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300608.

Lee:2015:IZM

[LY15a] Yeon Ju Lee and Jungho Yoon. Image zooming method using edgedirected moving least squares interpolation based on exponential polynomials. Applied Mathematics and Computation, 269(??):569–583, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010085.

Li:2015:SOR

[LY15b] Fengwei Li and Qingfang Ye. Second order Randić index of fluoranthene-type benzenoid systems. Applied Mathematics and Computation, 268(??):534–546, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008371.

Li:2015:NRP

[LY15c] Hanyu Li and Yanfei Yang. New rigorous perturbation bounds for the generalized Cholesky factorization. Applied Mathematics and Computation, 259(??):668–675, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003021.

Li:2015:AFT

[LY15d] Junmin Li and Hongyun Yue. Adaptive fuzzy tracking control for stochastic nonlinear systems with unknown time-varying delays. Applied Mathematics and Computation, 256(??):514–528, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017652.

Liang:2015:CFI

[LY15e] Jin Liang and He Yang.

Controllability of fractional integro-differential evolution equations with nonlocal conditions. Applied Mathematics and Computation, 254(??):20– 29, March 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031500003X.

Li:2016:GCI

[LY16a] Fengwei Li and Qingfang Ye. The general connectivity indices of fluoranthene-type benzenoid systems. Applied Mathematics and Computation, 273(??):897–911, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315014010.

Liu:2016:MSM

[LY16b] Hao Liu and Yongxin Yuan. A multi-step method for partial quadratic pole assignment problem with time delay. Applied Mathematics and Computation, 283(??):29–35, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301151.

Li:2017:IIN

[LY17] Hongying Li and Chengli Yao. The influence of internal noise on the detection of hormonal signal with the existence of external noise in a cell system. Applied Mathematics and Computation, 314(??):1–6, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304344.

Li:2018:QFA

[LY18a] Jianling Li and Zhenping Yang. A QP-free algorithm without a penalty function or a filter for nonlinear general-constrained optimization. Applied Mathematics and Computation, 316(??):52–72, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305647.

Li:2018:STD

[LY18b] Shuli Li and Weigen Yan. Spanning trees and dimer problem on the Cairo pentagonal lattice. Applied Mathematics and Computation, 337(??):25–33, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304156.

Li:2018:EMM

[LY18c] Yan Li and Hang Ye. Effect of the migration mechanism based on risk preference on the evolution of cooperation. Applied Mathematics and Computation, 320(??):621–632, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307221.

Liu:2018:EBQ

[LY18d] Xuefeng Liu and Chun’guang You. Explicit bound for quadratic Lagrange interpolation constant on triangular finite elements. Applied Mathematics and Computation, 319(??): 693–701, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317305714.

Llibre:2019:BQC

[LY19] Jaume Llibre and Pengfei Yuan. Bicentric quadrilateral central configurations. Applied Mathematics and Computation, 362 (??):Article 124507, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304862.

Lee:2015:WPC

[LYC15] Ji-Hyun Lee, Wei-Chang Yeh, and Mei-Chi Chuang. Web page classification based on a simplified swarm optimization. Applied Mathematics and Computation, 270(??):13– 24, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315010425.

Liao:2018:EAN

[LYDH18] Wenyuan Liao, Peng Yong, Hatef Dastour, and Jianping Huang. Efficient and accurate numerical simulation of acoustic wave propagation in a 2D heterogeneous media. Applied Mathematics and Computation, 321(??):385–400, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307610.

Li:2019:PDC

[LYH19] Xiaodi Li, Xueyan Yang, and Tingwen Huang. Persistence of delayed cooperative models: Impulsive control method. Applied Mathematics and Computation, 342(??):130–146, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307756.

Liu:2018:DSD

[LYM18] Meng Liu, Jingyi Yu, and Partha Sarathi Mandal. Dynamics of a stochastic delay competitive model with harvesting and Markovian switching. Applied Mathematics and Computation, 337(??):329–334, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302170.

Li:2017:ERD

[LYS17] Fengwei Li, Qingfang Ye, and Yuefang Sun. On edge-rupture degree of graphs. Applied Mathematics and Computation, 292(??):282–293, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304787.

Li:2017:CAC

[LYW17] Yuchao Li, Zishen Yang, and Wei Wang. Complexity and algorithms for the connected vertex cover problem in 4regular graphs. Applied Mathematics and Computation, 301(??):107–114, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307305.

Liu:2018:ACT

[LYW18] Xianliang Liu, Zishen Yang, and Wei Wang. Algorithm and complexity of the two disjoint connected dominating sets problem on trees. Applied Mathematics and Computation, 337(??):408–418, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304466.

Liao:2015:ACA

[LYXL15] Anping Liao, Xiaobo Yang, Jiaxin Xie, and Yuan Lei. Analysis of convergence for the alternating direction method applied to joint sparse recovery. Applied Mathematics and Computation, 269(??):548–557, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010267.

Luo:2015:TDH

[LYY15] Xue Luo, Shing-Tung Yau, and Stephen S.-T. Yau. Time-dependent Hermite–Galerkin spectral method and its applications. Applied Mathematics and Computation, 264(??):378–391, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005500.

Li:2019:SCP

[LYZ19] Yan Li, Ming Ye, and Qimin Zhang. Strong convergence of the partially truncated Euler– Maruyama scheme for a stochastic age-structured SIR epidemic model. Applied Mathematics and Computation, 362(??): Article 124519, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304989.

Liu:2018:DAH

[LYZL18] Chao Liu, Longfei Yu, Qingling Zhang, and Yuanke Li. Dynamic analysis of a hybrid bioeconomic plankton system with double time delays and stochastic fluctuations. Applied Mathematics and Computation, 316(??):115–137, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305702.

Li:2015:VVG

[LZ15a] Yun-Zhang Li and Yan Zhang. Vector-valued Gabor frames associated with periodic subsets of the real line. Applied Mathematics and Computation, 253(??):102–115, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017007.

Liu:2015:ASE

[LZ15b] Linna Liu and Quanxin Zhu. Almost sure exponential stability of numerical solutions to stochastic delay Hopfield neural networks. Applied Mathematics and Computation, 266(??):698–712, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007651.

Liu:2015:SSE

[LZ15c] Wenbin Liu and Qiben Zheng. A stochastic SIS epidemic model incorporating media coverage in a two patch setting. Applied Mathematics and Computation, 262(??):160–168, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004762.

Liu:2015:ASC

[LZ15d] Xiaowei Liu and Jin Zhang. Analysis of the SDFEM for convectiondiffusion problems with characteristic layers. Applied Mathematics and Computation, 262(??): 326–334, July 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315005032.

Liu:2015:ECF

[LZ15e] Zhenhai Liu and Biao Zeng. Existence and controllability for fractional evolution inclusions of Clarke’s subdifferential type. Applied Mathematics and Computation, 257(??):178–189, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017111.

Li:2016:CHO

[LZ16] Dongfang Li and Chengjian Zhang. Construction of high-order Runge–Kutta methods which preserve delay-dependent stability of DDEs. Applied Mathematics and Computation, 280(??):168–179, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300229.

Li:2017:MPR

[LZ17a] Chuan Li and Shan Zhao.

A matched Peaceman– Rachford ADI method for solving parabolic interface problems. Applied Mathematics and Computation, 299(??):28–44, April 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300316307093.

Li:2017:EGS

[LZ17b] Cui Li and Chengjian Zhang. The extended generalized Störmer-Cowell methods for second-order delay boundary value problems. Applied Mathematics and Computation, 294(??):87–95, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316305665.

Liang:2017:ARD

[LZ17c] Zhao-Zheng Liang and Guo-Feng Zhang. Analysis of the relaxed deteriorated PSS preconditioner for singular saddle point linear systems. Applied Mathematics and Computation, 305(??):308–322, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317301157.

Liu:2017:CAG

[LZ17d] Chanjuan Liu and Enqiang Zhu. Computational ability in games: Individual difference and dynamics. Applied Mathematics and Computation, 313(??):313–320, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303880.

Liu:2017:GFE

[LZ17e] Xiaowei Liu and Jin Zhang. Galerkin finite element methods for convection-diffusion problems with exponential layers on Shishkin triangular meshes and hybrid meshes. Applied Mathematics and Computation, 307(??):244–256, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301741.

Lu:2017:EER

[LZ17f] Zuliang Lu and Shuhua Zhang. L^{∞}-error estimates of rectangular mixed finite element methods for bilinear optimal control problem. Applied Mathematics and Computation, 300(??):79–94, May 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307329.

Luo:2017:MPS

[LZ17g] Chunmei Luo and Liancui Zuo. Metric properties of Sierpiński-like graphs. Applied Mathematics and Computation, 296(??):124–136, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630618X.

Li:2018:ESC

[LZ18a] Cui Li and Chengjian Zhang. An exponential stability criterion for nonlinear second-order functional differential equations with time-variable delays. Applied Mathematics and Computation, 328(??):119–124, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300924.

Li:2018:FEC

[LZ18b] Meng Li and Yong-Liang Zhao. A fast energy conserving finite element method for the nonlinear fractional Schrödinger equation with wave operator. Applied Mathematics and Computation, 338(??):739–757, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304983.

Li:2018:RSM

[LZ18c] Rongchang Li and Qingling Zhang. Robust H_{∞} sliding mode observer design for a class of Takagi– Sugeno fuzzy descriptor systems with time-varying delay. Applied Mathematics and Computation, 337(??):153–157, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304119.

Li:2018:PCC

[LZ18d] Tingting Li and Xia Zhang. Polychromatic colorings and cover decompositions of hypergraphs. Applied Mathematics and Computation, 339(??):144–152, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305770.

Li:2018:KPV

[LZ18e] Zhao Li and Liancui Zuo.

The k-path vertex cover in Cartesian product graphs and complete bipartite graphs. Applied Mathematics and Computation, 331(??):69–79, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301735.

Liu:2018:LMM

[LZ18f] X. Liu and Y. M. Zeng. Linear multistep methods for impulsive delay differential equations. Applied Mathematics and Computation, 321(??):555–563, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307968.

Li:2019:SAT

[LZ19a] Dongxi Li and Yue Zhao. Survival analysis for tumor cells in stochastic switching environment. Applied Mathematics and Computation, 357(??):199–205, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302863.

Li:2019:SZE

[LZ19b] Jianping Li and Jianbin Zhang. On the second Zagreb eccentricity indices of graphs. Applied Mathematics and Computation, 352(??):180–187, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300876.

Li:2019:CPP

[LZ19c] Longxiang Li and Qinghe Zhang. Comment on “A positivity preserving and well-balanced DG scheme using finite volume subcells in almost dry regions”. Applied Mathematics and Computation, 359(??):476–477, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303728.

See [MO16].

Li:2019:FSC

[LZ19d] Shengyue Li and Zhaojie Zhou. Fractional spectral collocation method for optimal control problem governed by space fractional diffusion equation. Applied Mathematics and Computation, 350(??):331–347, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300268.

Liao:2019:GVS

[LZ19e] Li-Dan Liao and Guo-Feng Zhang. A generalized variant of simplified HSS preconditioner for generalized saddle point problems. Applied Mathematics and Computation, 346(??):790–799, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309470.

Liu:2019:ANS

[LZ19f] X. Liu and Y. M. Zeng. Analytic and numerical stability of delay differential equations with variable impulses. Applied Mathematics and Computation, 358(??):293–304, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031930339X.

Lu:2019:NAS

[LZ19g] Ziqiang Lu and Yuanguo Zhu. Numerical approach for solution to an uncertain fractional differential equation. Applied Mathematics and Computation, 343(??):137–148, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308221.

liu:2019:SBD

[lZA^{+}19] Qi liu, Guanlan Zhang, Shahzad Ali, Xiaopeng Wang, Guodong Wang, Zhenkuan Pan, and Jiahua Zhang. SPI-based drought simulation and prediction using ARMA– GARCH model. Applied Mathematics and Computation, 355(??):96–107, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301626.

Li:2018:PSS

[LZCL18] Qiuyue Li, Yaoming Zhou, Fuzhong Cong, and Hu Liu. Positive solutions to superlinear attractive singular impulsive differential equation. Applied Mathematics and Computation, 338(??):733–738, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305538.

Li:2018:CSC

[LZCW18] Lili Li, Boya Zhou, Xiaoli Chen, and Zhiyong Wang. Convergence and stability of compact finite difference method for nonlinear time fractional reaction– diffusion equations with delay. Applied Mathematics and Computation, 337(??):137–143, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303837.

Liang:2019:SNT

[LZCW19] Min Liang, Chuanxi Zhu, Chunfang Chen, and Zhaoqi Wu. Some new theorems for cyclic contractions in G_{b}-metric spaces and some applications. Applied Mathematics and Computation, 346(??):545–558, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308944.

Liu:2016:SAC

[LZG16] Tingting Liu, Xian Zhang, and Xiangyu Gao. Stability analysis for continuoustime and discrete-time genetic regulatory networks with delays. Applied Mathematics and Computation, 274(??):628–643, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015301.

Liu:2016:TFO

[LZH16] Zhongli Liu, Quan Zheng, and Chun-E Huang. Thirdand fifth-order Newton– Gauss methods for solving nonlinear equations with n variables. Applied Mathematics and Computation, 290(??):250–257, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303812.

Long:2018:SAL

[LZH^{+}18] Fei Long, Chuan-Ke Zhang, Yong He, Lin Jiang, Qing-Guo Wang, and Min Wu. Stability analysis of Lur’e systems with additive delay components via a relaxed matrix inequality. Applied Mathematics and Computation, 328(??):224–242, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300225.

Liu:2019:BNT

[LZH19] Huiqing Liu, Ruiting Zhang, and Xiaolan Hu. Burning number of theta graphs. Applied Mathematics and Computation, 361(??):246–257, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304242.

Liu:2018:SFD

[LZHC18] Xuan Liu, Ding Zhai, Da-Kuo He, and Xiao-Heng Chang. Simultaneous fault detection and control for continuoustime Markovian jump systems with partially unknown transition probabilities. Applied Mathematics and Computation, 337(??):461–468, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304661.

Li:2018:CSO

[LZjy18] Hongjie Li, Yinglian Zhu, Liu jing, and Wang ying. Consensus of second-order delayed nonlinear multiagent systems via nodebased distributed adaptive completely intermittent protocols. Applied Mathematics and Computation, 326(??):1–15, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300110.

Liu:2019:IEU

[LZKX19] Hongjun Liu, Yingqian Zhang, Abdurahman Kadir, and Yanqiu Xu. Image encryption using complex hyper chaotic system by injecting impulse into parameters. Applied Mathematics and Computation, 360(??):83–93, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303807.

Lei:2015:BFC

[LZL15a] Z. X. Lei, L. W. Zhang, and K. M. Liew. Buckling of FG–CNT reinforced composite thick skew plates resting on Pasternak foundations based on an element-free approach. Applied Mathematics and Computation, 266(??):773–791, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500778X.

Liu:2015:SLS

[LZL15b] Shutang Liu, Yongping Zhang, and Wei Li. Stability of linear stochastic 2-D homogeneous systems. Applied Mathematics and Computation, 261(??):419–430, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004567.

Lei:2017:MMG

[LZL17] Z. X. Lei, L. W. Zhang, and K. M. Liew. Meshless modeling of geometrically nonlinear behavior of CNT-reinforced functionally graded composite laminated plates. Applied Mathematics and Computation, 295(??):24–46, February 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031630577X.

Lapin:2019:NSP

[LZL19a] Alexander Lapin, Shuhua Zhang, and Sergey Lapin. Numerical solution of a parabolic optimal control problem arising in economics and management. Applied Mathematics and Computation, 361(??):715–729, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930476X.

Li:2019:ATD

[LZL^{+}19b] Li Li, Jie Zhang, Chen Liu, Hong-Tao Zhang, Yi Wang, and Zhen Wang. Analysis of transmission dynamics for Zika virus on networks. Applied Mathematics and Computation, 347(??):566–577, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310166.

Li:2019:GCM

[LZLG19] Shasha Li, Yan Zhao, Fengwei Li, and Ruijuan Gu. The generalized 3connectivity of the Mycielskian of a graph. Applied Mathematics and Computation, 347(??):882–890, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309792.

Liu:2017:SSN

[LZLW17] Lei Liu, Qi Zhou, Hongjing Liang, and Lijie Wang. Stability and stabilization of nonlinear switched systems under average dwell time. Applied Mathematics and Computation, 298(??):77–94, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306646.

Liang:2018:DOB

[LZM^{+}18] Hongjing Liang, Yu Zhou, Hui Ma, Qinghui Wu, and Zhandong Yu. Distributedobserver-based output synchronization for heterogeneous double-integral networks. Applied Mathematics and Computation, 337(??):514–534, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304776.

Li:2018:ECE

[LZP18] Yumeng Li, Jun Zhang, and Matjaz Perc. Effects of compassion on the evolution of cooperation in spatial social dilemmas. Applied Mathematics and Computation, 320(??):437–443, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306896.

Li:2015:LCP

[LZS15] Shimin Li, Yulin Zhao, and Zhaohong Sun. On the limit cycles of planar polynomial system with non-rational first integral via averaging method at any order. Applied Mathematics and Computation, 256(??):876–880, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001216.

Li:2018:GSS

[LZT^{+}18] Hong-Li Li, Long Zhang, Zhidong Teng, Yao-Lin Jiang, and Ahmadjan Muhammadhaji. Global stability of an SI epidemic model with feedback controls in a patchy environment. Applied Mathematics and Computation, 321(??):372–384, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730766X.

Liu:2019:AST

[LZWZ19] Wenjun Liu, Yujia Zhang, Abdul Majid Wazwaz, and Qin Zhou. Analytic study on triple-S, tripletriangle structure interactions for solitons in inhomogeneous multi-mode fiber. Applied Mathematics and Computation, 361(??):325–331, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304485.

Li:2015:FPS

[LZX15] Ting Li, Weiguo Zhang, and Weijun Xu. A fuzzy portfolio selection model with background risk. Applied Mathematics and Computation, 256(??):505–513, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315000211.

Li:2019:GMF

[LZX19] Xiang-Jun Li, Xue-Qian Zeng, and Jun-Ming Xu. Generalized measures of fault tolerance for bubble sort networks. Applied Mathematics and Computation, 363(??): Article 124630, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306228.

Liu:2016:MSC

[LZY16] Lan Liu, Liling Zhu, and Da Yang. Modeling and simulation of the car-truck heterogeneous traffic flow based on a nonlinear carfollowing model. Applied Mathematics and Computation, 273(??):706–717, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013740.

Liao:2017:CPL

[LZY^{+}17] Xiaoxin Liao, Guopeng Zhou, Qigui Yang, Yuli Fu, and Guanrong Chen. Constructive proof of Lagrange stability and sufficient — necessary conditions of Lyapunov stability for Yang–Chen chaotic system. Applied Mathematics and Computation, 309(??):205–221, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302175.

Li:2015:RPO

[LZZ15] Ping Li, Wu Zhao, and Wei Zhou. Ruin probabilities and optimal investment when the stock price follows an exponential Lévy process. Applied Mathematics and Computation, 259(??):1030–1045, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016968.

Liu:2019:SRJ

[LZZ^{+}19] Zhongyun Liu, Yang Zhou, Yuelan Zhang, Lu Lin, and Dongxiu Xie. Some remarks on Jacobi and Gauss–Seidel-type iteration methods for the matrix equation AXB = C. Applied Mathematics and Computation, 354(??):305–307, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301122.

Li:2015:IAM

[LZZL15] Yingjie Li, Dingxuan Zhao, Zhongjun Zhang, and Jingang Liu. An IDRA approach for modeling helicopter based on Lagrange dynamics. Applied Mathematics and Computation, 265(??):733–747, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007420.

Liu:2018:NDD

[LZZT18] Jianzhou Liu, Juan Zhang, Lixin Zhou, and Gen Tu. The Nekrasov diagonally dominant degree on the Schur complement of Nekrasov matrices and its applications. Applied Mathematics and Computation, 320(??):251–263, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306604.

Ma:2015:ICI

[Ma15a] Junjie Ma. Implementing the complex integral method with the transformed Clenshaw– Curtis quadrature. Applied Mathematics and Computation, 250(??):792–797, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314013320.

Magrenan:2015:NSL

[MA15b] Á. Alberto Magreñán and Ioannis K. Argyros. New semilocal and local convergence analysis for the secant method. Applied Mathematics and Computation, 262(??):298–307, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004774.

Mukundan:2015:ENT

[MA15c] Vijitha Mukundan and Ashish Awasthi. Efficient numerical techniques for Burgers’ equation. Applied Mathematics and Computation, 262(??):282–297, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004415.

Mahmoud:2016:FFC

[MA16] Magdi S. Mahmoud and Naif B. Almutairi. Feedback fuzzy control for quantized networked systems with random delays. Applied Mathematics and Computation, 290(??):80–97, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303496.

Ma:2017:OFS

[Ma17a] Junjie Ma. Oscillation-free solutions to Volterra integral and integro-differential equations with periodic force terms. Applied Mathematics and Computation, 294(??):294–298, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305689.

Ma:2017:CCM

[Ma17b] Qian Ma. Cooperative control of multi-agent systems with unknown control directions. Applied Mathematics and Computation, 292(??):240–252, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304465.

Ma:2018:OPB

[Ma18]  Haifeng Ma.  Optimal

perturbation bounds for the core inverse. Applied Mathematics and Computation, 336(??):176–181, November 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318303850.

Maatouk:2015:TOD

[Maa15] Khalil Maatouk. Third order derivative free SPH iterative method for solving nonlinear systems. Applied Mathematics and Computation, 270(??):557–566, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031501139X.

Machado:2015:GC

[Mac15] J. A. Tenreiro Machado. Generalized convolution. Applied Mathematics and Computation, 257(??):34– 39, April 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314013071.

Messaoudi:2015:GDR

[MAG15] Salim A. Messaoudi and Mohammad M. Al-Gharabli. A general decay result of a nonlinear system of wave equations with infinite memories. Applied Mathematics and Computation, 259(??):540–551, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002854.

Mahfouz:2015:HCW

[Mah15] F. M. Mahfouz. Heat conduction within an elliptic annulus heated at either CWT or CHF. Applied Mathematics and Computation, 266(??):357–368, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315006943.

Mahmudov:2017:OFO

[Mah17] Elimhan N. Mahmudov.

Optimization of fourthorder discrete-approximation inclusions. Applied Mathematics and Computation, 292(??):19–32, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304404.

Mahato:2018:CFH

[Mah18a] Kanailal Mahato. The composition of fractional Hankel wavelet transform on some function spaces. Applied Mathematics and Computation, 337(??):63–75, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304387.

Mahmudov:2018:PAC

[Mah18b] N. I. Mahmudov. Partialapproximate controllability of nonlocal fractional evolution equations via approximating method. Applied Mathematics and Computation, 334(??):227–238, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302972.

Mahdy:2019:AHH

[Mah19] A. Mahdy. Aspects of homogeneous–heterogeneous reactions on natural convection flow of micropolar fluid past a permeable cone. Applied Mathematics and Computation, 352(??):59–67, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300645.

Majee:2018:CFS

[Maj18] Ananta K. Majee. Convergence of a flux-splitting finite volume scheme for conservation laws driven by Lévy noise. Applied Mathematics and Computation, 338(??):669–675, December 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318305071.

Modebei:2019:BHI

[MAJR19] Mark I. Modebei, Rapheal B.

Adeniyi, Samuel N. Jator, and Higinio Ramos. A block hybrid integrator for numerically solving fourth-order Initial Value Problems. Applied Mathematics and Computation, 346(??):680–694, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318309548.

Mursaleen:2015:CSA

[MAK15a] M. Mursaleen, Khursheed J. Ansari, and Asif Khan. Corrigendum to: “Some approximation results by (p, q)-analogue of Bernstein–Stancu operators” [Appl. Math. Comput. 264 (2015) 392– 402]. Applied Mathematics and Computation, 269(??):744–746, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010401.

Mursaleen:2015:PAB

[MAK15b] M. Mursaleen, Khursheed J. Ansari, and Asif Khan. On (p, q)- analogue of Bernstein operators. Applied Mathematics and Computation, 266(??):874–882, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005524. See

erratum [MAK16].

Mursaleen:2015:SAR

[MAK15c] M. Mursaleen, Khursheed J. Ansari, and Asif Khan. Some approximation results by (p, q)- analogue of Bernstein– Stancu operators. Applied Mathematics and Computation, 264(??):392–402, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004543.

Mursaleen:2016:EPA

[MAK16] M. Mursaleen, Khursheed J. Ansari, and Asif Khan. Erratum to “On (p, q)-analogue of Bernstein Operators” [Appl. Math. Comput. 266 (2015) 874–882]. Applied Mathematics and Computation, 278(??):70–71, March 31, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300923. See

[MAK15b].

Malik:2015:SAP

[Mal15] Neha Malik. Some approximation properties for generalized Srivastava–Gupta operators. Applied Mathematics and Computation, 269(??):747–758, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010358.

Mao:2017:CED

[Mao17] Yaping Mao. Constructing edge-disjoint Steiner paths in lexicographic product networks. Applied Mathematics and Computation, 308(??):1–10, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317301893.

Majeed:2015:DRD

[MAR15] Raafia Majeed, Sohaira Ahmad, and Muhammad Rehan. Delay-range-dependent observer-based control of nonlinear systems under input and output time-delays. Applied Mathematics and Computation, 262(??):145–159, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500510X.

Margenstern:2016:IHP

[Mar16] Maurice Margenstern. Infinigons of the hyperbolic plane and grossone. Applied Mathematics and Computation, 278(??):45– 53, March 31, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316300194.

Matai:2015:SMO

[Mat15]  Rajesh Matai.  Solv-

ing multi objective facility layout problem by modified simulated annealing. Applied Mathematics and Computation, 261(??):302–311, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500421X.

Marciniak:2015:FCF

[MB15a] Krzysztof Marciniak and Maciej Blaszak. Flat coordinates of flat Stäckel systems. Applied Mathematics and Computation, 268(??):706–716, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008802.

Mehatari:2015:ENG

[MB15b] Ranjit Mehatari and Anirban Banerjee. Effect on normalized graph Laplacian spectrum by motif attachment and duplication. Applied Mathematics and Computation, 261(??):382–387, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004373.

Moghaddam:2017:TDF

[MB17] M. Rashidi Moghaddam and G. H. Baradaran. Three-dimensional free vibrations analysis of functionally graded rectangular plates by the meshless local Petrov–Galerkin (MLPG) method. Applied Mathematics and Computation, 304(??):153–163, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730067X.

Ma:2019:NLD

[MB19] Xiaoling Ma and Hong Bian. The normalized Laplacians, degree-Kirchhoff index and the spanning trees of hexagonal Möbius graphs. Applied Mathematics and Computation, 355(??):33– 46, August 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300319301560.

Madi:2018:PAO

[MBHS18] Sofiane Madi, Mohamed Cherif Bouras, Mohamed Haiour, and Andreas Stahel. Pricing of American options, using the Brennan– Schwartz algorithm based on finite elements. Applied Mathematics and Computation, 339(??):837–845, December 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318305174.

Majumder:2016:EMT

[MBM16] P. Majumder, U. K. Bera, and M. Maiti. An EPQ model for two-warehouse in unremitting release pattern with two-level trade credit period concerning both supplier and retailer. Applied Mathematics and Computation, 274(??):430–458, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014083.

Ma:2019:WVP

[MBW19] Gang Ma, Qiuju Bian, and Jianfeng Wang. The weighted vertex PI index of (n, m)-graphs with given diameter. Applied Mathematics and Computation, 354(??):329–337, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301481.

Ma:2015:RFT

[MC15a] Yuechao Ma and Hui Chen. Reliable finite-time H_{∞} filtering for discrete time-delay systems with Markovian jump and randomly occurring nonlinearities. Applied Mathematics and Computation, 268(??):897–915, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008486.

Mezo:2015:EZB

[MC15b] István Mező and Roberto B. Corcino. The estimation of the zeros of the Bell and r-Bell polynomials. Applied Mathematics and Computation, 250(??):727–732, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300314014337.

Mishra:2015:DFE

[MC15c] Asim Kumar Mishra and Sushanta Chakraborty. Development of a finite element model updating technique for estimation of constituent level elastic parameters of FRP plates. Applied Mathematics and Computation, 258(??):84– 94, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001551.

Marin:2017:NIR

[MC17] Liviu Marin and Corina Cipu. Non-iterative regularized MFS solution of inverse boundary value problems in linear elasticity: a numerical study. Applied Mathematics and Computation, 293(??):265–286, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305161.

Mosic:2016:SOO

[MCI16] Dijana Mosić and Dragana S. Cvetković-Ilić. Some orders for operators on Hilbert spaces. Applied Mathematics and Computation, 275(??):229–237, February 15, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315015611.

Marin:2019:EDR

[MCO19] M. Marin, A. Chirila, and M. I. A. Othman. An extension of Dafermos’s results for bodies with a dipolar structure. Applied Mathematics and Computation, 361(??):680–688, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304898.

Montoro-Cazorla:2015:SWM

[MCPO15] Delia Montoro-Cazorla and Rafael Pérez-Ocón. A shock and wear model with dependence between the interarrival failures. Applied Mathematics and Computation, 259(??): 339–352, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315001575.

Ma:2019:NGT

[MCS^{+}19] Yuede Ma, Shujuan Cao, Yongtang Shi, Matthias Dehmer, and Chengyi Xia. Nordhaus–Gaddum type results for graph irregularities. Applied Mathematics and Computation, 343(??):268–272, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308440.

Ma:2019:ILP

[MCY19] Yuede Ma, Qingqiong Cai, and Shunyu Yao. Integer linear programming models for the weighted total domination problem. Applied Mathematics and Computation, 358(??):146–150, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303261.

Meng:2015:CVG

[MCZ15] Fanyong Meng, Xiaohong Chen, and Qiang Zhang. A coalitional value for games on convex geometries with a coalition structure. Applied Mathematics and Computation, 266(??):605–614, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007419.

Ma:2015:SSP

[MD15a] Qiang Ma and Xiaohua Ding. Stochastic symplectic partitioned Runge–Kutta methods for stochastic Hamiltonian systems with multiplicative noise. Applied Mathematics and Computation, 252(??):520–534, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016993.

Mosic:2015:WPI

[MD15b] Dijana Mosić and Dragan S. Djordjević. Weighted partial isometries and weighted-EP elements in C^{∗}-algebras. Applied Mathematics and Computation, 265(??):17–30, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005640.

Mosic:2015:WPO

[MD15c] Dijana Mosić and Dragan S. Djordjević. Weighted pre-orders involving the generalized Drazin inverse. Applied Mathematics and Computation, 270(??):496–504, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011054.

Mittal:2017:NST

[MD17] R. C. Mittal and Sumita Dahiya. Numerical simulation of three-dimensional telegraphic equation using cubic B-spline differential quadrature method. Applied Mathematics and Computation, 313(??): 442–452, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317304277.

Macias-Diaz:2018:NEH

[MD18a] J. E. Macías-Díaz. A numerically efficient Hamiltonian method for fractional wave equations. Applied Mathematics and Computation, 338(??): 221–230, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300318304910.

Mosic:2018:BRG

[MD18b] Dijana Mosić and Dragan S. Djordjević. Block representations of the generalized Drazin inverse. Applied Mathematics and Computation, 331(??):200–209, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301929.

Mowshowitz:2018:CME

[MD18c] Abbe Mowshowitz and Matthias Dehmer. A calculus for measuring the elegance of abstract graphs. Applied Mathematics and Computation, 320(??):142–148, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306446.

Mishra:2018:GAT

[MDG18] Vishnu Narayan Mishra, Ankita R. Devdhara, and R. B. Gandhi. Global approximation theorems for the generalized Szàsz– Mirakjan type operators in exponential weight spaces. Applied Mathematics and Computation, 336(??):206–214, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303606.

Mohanty:2015:COM

[MDH15] R. K. Mohanty, Weizhong Dai, and Fei Han. Compact operator method of accuracy two in time and four in space for the numerical solution of coupled viscous Burgers’ equations. Applied Mathematics and Computation, 256(??):381–393, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000818.

Ma:2016:AED

[MDH16] Manjun Ma, Chi Dang, and Zhe Huang. Analytical expressions for dark soliton solution of a Gross–Pitaevskii equation. Applied Mathematics and Computation, 273(??):383–389, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315013508.

Macias-Diaz:2018:CFO

[MDHD18] J. E. Macías-Díaz, A. S. Hendy, and R. H. De Staelen. A compact fourthorder in space energypreserving method for Riesz space-fractional nonlinear wave equations. Applied Mathematics and Computation, 325(??):1– 14, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031730855X.

Ma:2015:EAB

[MDZ15] Weijun Ma, Baocang Ding, and Qimin Zhang. The existence and asymptotic behaviour of energy solutions to stochastic age-dependent population equations driven by Levy processes. Applied Mathematics and Computation, 256(??):656–665, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001095.

Mo:2017:ESS

[MDZ17] Haoyi Mo, Feiqi Deng, and Chaolong Zhang. Exponential stability of the split-step θ-method for neutral stochastic delay differential equations with jumps. Applied Mathematics and Computation, 315(??):85–95, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304472.

Mencinger:2017:CCP

[MF17] Matej Mencinger and Brigita Fercec. The center and cyclicity problems for some analytic maps. Applied Mathematics and Computation, 306(??):73–85, August 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301431.

Massopust:2017:FCP

[MFC17] Peter Massopust, Brigitte Forster, and Ole Christensen. Fractional and complex pseudo-splines and the construction of Parseval frames. Applied Mathematics and Computation, 314(??):12–24, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300317304368.

Mavrotas:2015:IVC

[MFF15] George Mavrotas, Kostas Florios, and José Rui Figueira. An improved version of a core based algorithm for the multiobjective multi-dimensional knapsack problem: a computational study and comparison with metaheuristics. Applied Mathematics and Computation, 270(??):25–43, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010735.

Ma:2015:FTC

[MFJZ15] Yuechao Ma, Lei Fu, Yanhui Jing, and Qingling Zhang. Finite-time H_{∞} control for a class of discrete-time switched singular time-delay systems subject to actuator saturation. Applied Mathematics and Computation, 261(??):264–283, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004257.

Mogorosi:2017:GAH

[MFMK17] Tshepo E. Mogorosi, Igor L.

Freire, Ben Muatjetjeja, and Chaudry Masood Khalique. Group analysis of a hyperbolic Lane– Emden system. Applied Mathematics and Computation, 292(??):156–164, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304714.

Mencinger:2017:CSA

[MFOP17] Matej Mencinger, Brigita Fercec, Regilene Oliveira, and Dusan Pagon. Cyclicity of some analytic maps. Applied Mathematics and Computation, 295(??):114–125, February 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305963.

Ma:2015:PCS

[MG15] Ruyun Ma and Hongliang Gao. Positive convex solutions of boundary value problems arising from Monge–Ampère equations. Applied Mathematics and Computation, 259(??):390–402, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002994.

Malik:2017:APB

[MG17] Neha Malik and Vijay Gupta. Approximation by (p, q)-Baskakov-beta operators. Applied Mathematics and Computation, 293(??):49–56, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305008.

Migorski:2019:VHI

[MG19] Stanislaw Migórski and Piotr Gamorski. Variationalhemivariational inequality for a class of dynamic nonsmooth frictional contact problems. Applied Mathematics and Computation, 346(??):465–479, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308713.

Martinez-Guerra:2016:GMS

[MGCAPP16] Rafael Martínez-Guerra, Christopher D. Cruz-Ancona, and Claudia A. Pérez-Pinacho. Generalized multi-synchronization viewed as a multi-agent leader-following consensus problem. Applied Mathematics and Computation, 282(??):226–236, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301205.

Maatki:2016:IEM

[MGK^{+}16] Chemseddine Maatki, Kaouther Ghachem, Lioua Kolsi, Ahmed Kadhim Hussein, Mohamed Naceur Borjini, and Habib Ben Aissia. Inclination effects of magnetic field direction in 3D double-diffusive natural convection. Applied Mathematics and Computation, 273(??):178–189, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012795.

Ma:2019:EMC

[MGL19] Jie Ma, Feng Gao, and Yongshu Li. An efficient method to compute different types of generalized inverses based on linear transformation. Applied Mathematics and Computation, 349(??):367–380, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318311251.

Mkhize:2017:LCS

[MGMM17] T. G. Mkhize, K. Govinder, S. Moyo, and S. V. Meleshko. Linearization criteria for systems of two second-order stochastic ordinary differential equations. Applied Mathematics and Computation, 301(??):25–35, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307457.

Miniak-Gorecka:2018:ACF

[MGN18]  Alicja  Miniak-Górecka

and Andrzej Nowakowski. Analysis of control for a free boundary problem of steady plaques in the artery. Applied Mathematics and Computation, 331(??):473–483, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301188.

Maciazek:2015:CBA

[MGPL15] Marcin Maciazek, Dariusz Grabowski, Marian Pasko, and Michal Lewandowski. Compensation based on active power filters — the cost minimization. Applied Mathematics and Computation, 267(??):648–654, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000065.

Martinez-Guerra:2015:SIF

[MGPPGCCV15] Rafael Martínez-Guerra, Claudia A. Pérez-Pinacho, Gian Carlo Gómez-Cortés, and Juan C. Cruz-Victoria. Synchronization of incommensurate fractional order system. Applied Mathematics and Computation, 262(??):260–266, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004403.

Mirzaee:2015:AMB

[MH15a] Farshid Mirzaee and Elham Hadadiyan. Applying the modified blockpulse functions to solve the three-dimensional Volterra– Fredholm integral equations. Applied Mathematics and Computation, 265(??):759–767, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007560.

Mirzaee:2015:NSL

[MH15b] Farshid Mirzaee and Elham Hadadiyan. Numerical solution of linear Fredholm integral equations via two-dimensional modification of hat functions. Applied Mathematics and Computation, 250(??):805–816, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015033. See

erratum [Sab18].

Mirzaee:2016:NSV

[MH16a] Farshid Mirzaee and Elham Hadadiyan. Numerical solution of Volterra– Fredholm integral equations via modification of hat functions. Applied Mathematics and Computation, 280(??):110–123, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300388.

Mirzaee:2016:AFC

[MH16b] Farshid Mirzaee and Seyede Fatemeh Hoseini. Application of Fibonacci collocation method for solving Volterra–Fredholm integral equations. Applied Mathematics and Computation, 273(??):637–644, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013776.

Ming:2017:CMV

[MH17a] Wanyuan Ming and Chengming Huang. Collocation methods for Volterra functional integral equations with non-vanishing delays. Applied Mathematics and Computation, 296(??):198–214, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306191.

Mirzaee:2017:SSL

[MH17b] Farshid Mirzaee and Elham Hadadiyan. Solving system of linear Stratonovich Volterra integral equations via modification of hat functions. Applied Mathematics and Computation, 293(??):254–264, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305112.

Mirzaee:2017:NCA

[MH17c] Farshid Mirzaee and Seyede Fatemeh Hoseini. A new collocation approach for solving systems of high-order linear Volterra integrodifferential equations with variable coefficients. Applied Mathematics and Computation, 311(??): 272–282, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317303326.

Ma:2019:AEE

[MH19a] Yanying Ma and Jin Huang. Asymptotic error expansions and splitting extrapolation algorithm for two classes of two-dimensional Cauchy principal-value integrals. Applied Mathematics and Computation, 357(??): 107–118, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300319302644.

Mustafa:2019:FNL

[MH19b] Ghulam Mustafa and Rabia Hameed. Families of non-linear subdivision schemes for scattered data fitting and their non-tensor product extensions. Applied Mathematics and Computation, 359(??):214–240, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303716.

Markopoulos:2015:MPS

[MHCF15] Alexandros Markopoulos, Vaclav Hapla, Martin Cermak, and Martin Fusek. Massively parallel solution of elastoplasticity problems with tens of millions of unknowns using PermonCube and FLLOP packages. Applied Mathematics and Computation, 267(??):698–710, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017585.

Mao:2015:EAE

[MHM15] Wei Mao, Liangjian Hu, and Xuerong Mao. The existence and asymptotic estimations of solutions to stochastic pantograph equations with diffusion and Lévy jumps. Applied Mathematics and Computation, 268(??):883–896, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008905.

Ma:2015:CAS

[MHN15] Xiaohua Ma, Chengming Huang, and Xin Niu. Convergence analysis of spectral collocation methods for a class of weakly singular Volterra integral equations. Applied Mathematics and Computation, 250(??):131–144, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014751.

McHardy:2018:SSL

[MHR18] Christopher McHardy, Tobias Horneber, and Cornelia Rauh. Spectral simulation of light propagation in participating media by using a lattice Boltzmann method for photons. Applied Mathematics and Computation, 319(??):59–70, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300632.

Ma:2017:HBD

[MHX17] Zhan-Ping Ma, Hai-Feng Huo, and Hong Xiang. Hopf bifurcation for a delayed predator– prey diffusion system with Dirichlet boundary condition. Applied Mathematics and Computation, 311(??):1–18, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303004.

Ming:2018:OSR

[MHZ18] Wanyuan Ming, Chengming Huang, and Longbin Zhao. Optimal superconvergence results for Volterra functional integral equations with proportional vanishing delays. Applied Mathematics and Computation, 320(??):292–301, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306732.

Ma:2015:HAB

[MHZC15] Lianbo Ma, Kunyuan Hu, Yunlong Zhu, and Hanning Chen. A hybrid artificial bee colony optimizer by combining with life-cycle, Powell’s search and crossover. Applied Mathematics and Computation, 252(??):133–154, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016373.

Mi:2017:BRL

[Mi17] Ling Mi. Blow-up rates of large solutions for infinity Laplace equations. Applied Mathematics and Computation, 298(??):36– 44, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306658.

Miao:2017:NUT

[Mia17] Shu-Xin Miao. A new Uzawa-type method for saddle point problems. Applied Mathematics and Computation, 300(??):95– 102, May 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0096300316307342.

Micula:2015:INM

[Mic15a] Sanda Micula. An iterative numerical method for Fredholm–Volterra integral equations of the second kind. Applied Mathematics and Computation, 270(??):935–942, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011662.

Micula:2015:SCM

[Mic15b] Sanda Micula. A spline collocation method for Fredholm–Hammerstein integral equations of the second kind in two variables. Applied Mathematics and Computation, 265(??):352–357, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006335.

Migda:2015:ASD

[Mig15] Janusz Migda. Approximative solutions to difference equations of neutral type. Applied Mathematics and Computation, 268(??):763–774, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008784.

Migda:2016:APS

[Mig16] Janusz Migda. Asymptotically polynomial solutions to difference equations of neutral type. Applied Mathematics and Computation, 279(??):16–27, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300017.

Milosevic:2016:EAA

[Mil16] Marija Milosević. An explicit analytic approximation of solutions for a class of neutral stochastic differential equations with timedependent delay based on Taylor expansion. Applied Mathematics and Computation, 274(??):745–761, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015088.

Misra:2015:CKV

[Mis15] A. P. Misra. Complex Korteweg–de Vries equation and nonlinear dustacoustic waves in a magnetoplasma with a pair of trapped ions. Applied Mathematics and Computation, 256(??):368–374, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500034X.

Mishra:2016:NSR

[Mis16] Chittaranjan Mishra. A new stability result for the modified Craig–Sneyd scheme applied to twodimensional convectiondiffusion equations with mixed derivatives. Applied Mathematics and Computation, 285(??):41–50, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302156.

Miyajima:2017:VSD

[Miy17] Shinya Miyajima. Verified solutions of delay eigenvalue problems. Applied Mathematics and Computation, 303(??):211–225, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300413.

Miyajima:2019:VCM

[Miy19] Shinya Miyajima. Verified computation for the matrix Lambert W function. Applied Mathematics and Computation, 362 (??):Article 124555, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305387.

Ma:2019:BCS

[MJ19] Liangliang Ma and Ruihong Ji. Blow-up criteria of smooth solutions for magnetic Bénard system with partial viscosity. Applied Mathematics and Computation, 346(??):816–831, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308683.

Ma:2016:RFT

[MJL16] Yuechao Ma, Xiaorui Jia, and Deyou Liu. Robust finite-time H_{∞} control for discrete-time singular Markovian jump systems with time-varying delay and actuator saturation. Applied Mathematics and Computation, 286(??):213–227, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302284.

Meena:2019:FMH

[MJSA19] Rakesh Kumar Meena, Madhu Jain, Sudeep Singh Sanga, and Assif Assad. Fuzzy modeling and harmony search optimization for machining system with general repair, standby support and vacation. Applied Mathematics and Computation, 361(??):858–873, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304552.

Marimuthu:2015:SSO

[MK15a] Gurusamy Thevar Marimuthu and Gurunathan Kumar. Solution to some open problems on E-super vertex magic labeling of disconnected graphs. Applied Mathematics and Computation, 268(??):657–663, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008772.

Mojaver:2015:MAC

[MK15b] Aida Mojaver and Hossein Kheiri. Mathematical analysis of a class of HIV infection models of CD4^{+} T-cells with combined antiretroviral therapy. Applied Mathematics and Computation, 259(??):258–270, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002647.

Monabbati:2015:CSD

[MK15c] Ehsan Monabbati and Hossein Taghizadeh Kakhki. On a class of subadditive duals for the uncapacitated facility location problem. Applied Mathematics and Computation, 251(??):118–131, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014477.

Mohanty:2016:HAI

[MK16] R. K. Mohanty and Deepti Kaur. High accuracy implicit variable mesh methods for numerical study of special types of fourth order non-linear parabolic equations. Applied Mathematics and Computation, 273(??):678–696, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013788.

Masri:2015:MSV

[MKG15] H. Masri, S. Krichen, and A. Guitouni. A multistart variable neighborhood search for solving the single path multicommodity flow problem. Applied Mathematics and Computation, 251(??):132–142, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014982.

Mihailovic:2015:SAT

[MKMC15] Dragutin Mihailović, Vladimir Kostić, Gordan Mimić, and Ljiljana Cvetković. Stability analysis of turbulent heat exchange over the heterogeneous environmental interface in climate models. Applied Mathematics and Computation, 265(??):79–90, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315005925.

Moosaei:2015:STS

[MKN^{+}15a] H. Moosaei, S. Ketabchi, M. A. Noor, J. Iqbal, and V. Hooshyarbakhsh. Some techniques for solving absolute value equations. Applied Mathematics and Computation, 268(??):696–705, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500853X.

Murthy:2015:PFE

[MKN15b] S. V. S. S. N. V. G. Krishna Murthy, B. V. Rathish Kumar, and Mohit Nigam. A parallel finite element study of 3D mixed convection in a fluid saturated cubic porous enclosure under injection/ suction effect. Applied Mathematics and Computation, 269(??):841–862, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009406.

Meysonnat:2016:ITT

[MKRS16] Pascal S. Meysonnat, Seong Ryong Koh, Benedikt Roidl, and Wolfgang Schröder. Impact of transversal traveling surface waves in a nonzero pressure gradient turbulent boundary layer flow. Applied Mathematics and Computation, 272 (part 2)(??):498–507, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007183.

Mohanty:2019:CTT

[MKS19] R. K. Mohanty, Deepti Kaur, and Swarn Singh. A class of two- and three-level implicit methods of order two in time and four in space based on half-step discretization for two-dimensional fourth order quasi-linear parabolic equations. Applied Mathematics and Computation, 352(??):68– 87, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300657.

Mi:2015:SOE

[ML15] Ling Mi and Bin Liu.

Second order expansion for the solution to a singular Dirichlet problem. Applied Mathematics and Computation, 270(??):401–412, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010917.

Ma:2017:RCN

[ML17] Yingbin Ma and Zaiping Lu. Rainbow connection numbers of Cayley digraphs on abelian groups. Applied Mathematics and Computation, 311(??):178–183, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317303223.

Meng:2018:GAN

[MLG18] Xinzhu Meng, Fei Li, and Shujing Gao. Global analysis and numerical simulations of a novel stochastic eco-epidemiological model with time delay. Applied Mathematics and Computation, 339(??):685–700, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318306040.

Meng:2019:CTO

[MLS19] Hui Meng, Pu Liao, and Tak Kuen Siu. Continuoustime optimal reinsurance strategy with nontrivial curved structures. Applied Mathematics and Computation, 363(??): Article 124585, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305776.

Ma:2018:FTS

[MLSX18] Dazhong Ma, Xiaoyu Li, Qiuye Sun, and Xiangpeng Xie. Fault tolerant synchronization of chaotic systems with time delay based on the double event-triggered sampled control. Applied Mathematics and Computation, 333(??):20–31, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302741.

Meng:2017:GES

[MLT17] Lixin Meng, Jingyu Li, and Jian Tao. Global energy solutions to a stochastic Schrödinger–Poisson system with multiplicative noise in two dimensions.

Applied Mathematics and Computation, 300(??):40– 59, May 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307184.

Miao:2018:TNP

[MLW18] Shu-Xin Miao, Yu-Hua Luo, and Guang-Bin Wang. Two new preconditioned GAOR methods for weighted linear least squares problems. Applied Mathematics and Computation, 324(??):93–104, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308664.

Maleknejad:2004:NSL

[MM04] K. Maleknejad and Y. Mahmoudi. Numerical solution of linear Fredholm integral equation by using hybrid Taylor and Block– Pulse functions. Applied Mathematics and Computation, 149(3):799–806, February 22, 2004. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See erratum [Sab18].

Ma:2015:SRV

[MM15a] Fanghua Ma and Xuan Ma. Stability of the relativistic Vlasov–Maxwell– Boltzmann system for short range interaction. Applied Mathematics and Computation, 265(??): 854–882, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315006608.

Ma:2015:OCH

[MM15b] Qian Ma and Guoying Miao. Output consensus for heterogeneous multiagent systems with linear dynamics. Applied Mathematics and Computation, 271(??):548–555, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011984.

Marinoschi:2016:FPD

[MM16] Gabriela Marinoschi and Angela Martiradonna. Fish populations dynamics with nonlinear stock-recruitment renewal conditions. Applied Mathematics and Computation, 277(??): 101–110, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300315300291.

Marovt:2017:SER

[MM17] Janko Marovt and Katja Mihelic. On sets of elements in Rickart rings induced by partial orders. Applied Mathematics and Computation, 315(??):555–563, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305660.

McAnally:2018:IGD

[MM18a] Morgan McAnally and Wen-Xiu Ma. An integrable generalization of the D-Kaup–Newell soliton hierarchy and its bi-Hamiltonian reduced hierarchy. Applied Mathematics and Computation, 323(??):220–227, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307774.

Minjibir:2018:IAS

[MM18b] M. S. Minjibir and I. Mohammed. Iterative algorithms for solutions of Hammerstein integral inclusions. Applied Mathematics and Computation, 320(??):389–399, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306690.

Mohammadi:2015:OEL

[MMB15] Maryam Mohammadi, Siti Nurmaya Musa, and Ardeshir Bahreininejad. Optimization of economic lot scheduling problem with backordering and shelflife considerations using calibrated metaheuristic algorithms. Applied Mathematics and Computation, 251(??):404–422, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015586.

Milovanovic:2016:NFR

[MMDG16] E. I. Milovanović, I. Z.

Milovanović, E. Ć. Dolićanin, and E. Glogić. A note on the first reformulated Zagreb index. Applied Mathematics and Computation, 273(??):16–20, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315013260.

Merino:2015:MOU

[MMG15] S. Merino, J. Martínez, and F. Guzmán. Metadomotic optimization using genetic algorithms. Applied Mathematics and Computation, 267(??): 170–178, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315004877.

Milovanovic:2016:UBS

[MMG16] Igor Milovanović, Emina Milovanović, and Ivan Gutman. Upper bounds for some graph energies. Applied Mathematics and Computation, 289(??):435–443, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630354X.

Mbava:2017:PPS

[MMG17] W. Mbava, J. Y. T. Mugisha, and J. W. Gonsalves. Prey, predator and super-predator model with disease in the super-predator. Applied Mathematics and Computation, 297(??):92–114, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306415.

Manafian:2019:LTS

[MMIA19] Jalil Manafian, Behnam Mohammadi-Ivatloo, and Mehdi Abapour. Lumptype solutions and interaction phenomenon to the (2 + 1)-dimensional Breaking Soliton equation. Applied Mathematics and Computation, 356(??):13–41, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302097.

Mohanty:2017:CTL

[MMK17] R. K. Mohanty, Sean Mc-Kee, and Deepti Kaur. A class of two-level implicit unconditionally stable methods for a fourth order parabolic equation. Applied Mathematics and Computation, 309(??):272–280, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302539.

Mbabazi:2018:MWH

[MMK18] Fulgensia Kamugisha Mbabazi, J. Y. T. Mugisha, and M. Kimathi. Modeling the within-host coinfection of influenza: A virus and pneumococcus. Applied Mathematics and Computation, 339(??):477–487, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305964.

Milovanovic:2018:RSS

[MMM18] E. I. Milovanović, I. Z. Milovanović, and M. M. Matejić. Remark on spectral study of the geometric-arithmetic index and some generalizations. Applied Mathematics and Computation, 334(??):206–213, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303187.

Manas-Manas:2017:AVD

[MMMMB17] Juan F. Mañas-Mañas, Francisco Marcellán, and Juan J. Moreno-Balcázar. Asymptotics for varying discrete Sobolev orthogonal polynomials. Applied Mathematics and Computation, 314(??):65–79, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304332.

Marcellan:2018:BCO

[MMPP18] Francisco Marcellán, Misael Marriaga, Teresa E. Pérez, and Miguel A. Piñar. On bivariate classical orthogonal polynomials. Applied Mathematics and Computation, 325(??):340–357, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317309074.

Manna:2017:MMT

[MMS17a] Debasis Manna, Alakes Maiti, and G. P. Samanta. A Michaelis–Menten type food chain model with strong Allee effect on the prey. Applied Mathematics and Computation, 311(??):390–409, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303417.

Mishra:2017:APB

[MMS17b] Vishnu Narayan Mishra, Ram N. Mohapatra, and Preeti Sharma. On approximation properties of Baskakov–Szász–Stancu operators using hypergeometric representation. Applied Mathematics and Computation, 294(??):77– 86, February 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316305446.

Manna:2018:APP

[MMS18] Debasis Manna, Alakes Maiti, and G. P. Samanta. Analysis of a predator– prey model for exploited fish populations with schooling behavior. Applied Mathematics and Computation, 317(??):35– 48, January 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317306100.

Merad:2016:GMT

[MMV16] A. Merad and J. Martín-Vaquero. A Galerkin method for two-dimensional hyperbolic integro-differential equation with purely integral conditions. Applied Mathematics and Computation, 291(??):386–394, December 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316304337.

Ma:2017:PSB

[MMZ^{+}17] Jun Ma, Lv Mi, Ping Zhou, Ying Xu, and Tasawar Hayat. Phase synchronization between two neurons induced by coupling of electromagnetic field. Applied Mathematics and Computation, 307(??):321–328, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730173X.

Migda:2019:APSb

[MMZ19] Janusz Migda, Małgorzata Migda, and Zenon Zbaszyniak. Asymptotically periodic solutions of second order difference equations. Applied Mathematics and Computation, 350(??):181–189, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300189.

Mohtadi:2015:PAF

[MN15] Mohammad Mahdi Mohtadi and Kazem Nogondarian. Presenting an algorithm to find Nash equilibrium in twoperson static games with many strategies. Applied Mathematics and Computation, 251(??):442–452, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015628.

Meucci:2016:DPM

[MN16] A. Meucci and M. Nicolosi. Dynamic portfolio management with views at multiple horizons. Applied Mathematics and Computation, 274(??):495–518, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501471X.

Majumdar:2017:ADN

[MN17] Anirban Majumdar and Srinivasan Natesan. Alternating direction numerical scheme for singularly perturbed 2D degenerate parabolic convection– diffusion problems. Applied Mathematics and Computation, 313(??):

453–473, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317304228.

Menasri:2015:TPR

[MND^{+}15] R. Menasri, A. Nakib, B. Daachi, H. Oulhadj, and P. Siarry. A trajectory planning of redundant manipulators based on bilevel optimization. Applied Mathematics and Computation, 250(??):934–947, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014763.

Mostaghimi:2019:CES

[MNJM19] Soudeh Mostaghimi, Fahimeh Nazarimehr, Sajad Jafari, and Jun Ma. Chemical and electrical synapsemodulated dynamical properties of coupled neurons under magnetic flow. Applied Mathematics and Computation, 348(??):42– 56, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031831004X.

Ma:2019:TRC

[MNJW19] Yingbin Ma, Kairui Nie, Fengxia Jin, and Cui Wang. Total rainbow connection numbers of some special graphs. Applied Mathematics and Computation, 360(??):213–220, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319303959.

Mursaleen:2019:IAE

[MNK19] M. Mursaleen, Ambreen Naaz, and Asif Khan. Improved approximation and error estimations by King type (p, q)-Szász– Mirakjan Kantorovich operators. Applied Mathematics and Computation, 348(??):175–185, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031831018X.

Mihai:2015:SII

[MNNA15] Marcela V. Mihai, Muhammad Aslam Noor, Khalida Inayat Noor, and Muhammad Uzair Awan. Some integral inequalities for harmonic hconvex functions involving hypergeometric functions. Applied Mathematics and Computation, 252(??):257–262, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016725.

Mityushev:2019:SAM

[MNNP19] V. Mityushev, W. Nawalaniec, D. Nosov, and E. Pesetskaya. Schwarz’s alternating method in a matrix form and its applications to composites. Applied Mathematics and Computation, 356(??):144–156, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302334.

Migda:2019:APSa

[MNR19] Janusz Migda and Magdalena Nockowska-Rosiak. Asymptotic properties of solutions to difference equations of Sturm– Liouville type. Applied Mathematics and Computation, 340(??):126–137, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318306404.

Ma:2018:MOS

[MNYL18] Yunpeng Ma, Peifeng Niu, Shanshan Yan, and Guoqiang Li. A modified online sequential extreme learning machine for building circulation fluidized bed boiler’s NO_{x} emission model. Applied Mathematics and Computation, 334(??):214–226, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301759.

Mollapourasl:2015:SFI

[MO15a] R. Mollapourasl and A. Ostadi. On solution of functional integral equation of fractional order. Applied Mathematics and Computation, 270(??):631–643, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315011248.

Mosleh:2015:LSA

[MO15b] Maryam Mosleh and Mahmood Otadi. Least squares approximation method for the solution of Hammerstein– Volterra delay integral equations. Applied Mathematics and Computation, 258(??):105–110, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001320.

Meister:2016:PPW

[MO16] Andreas Meister and Sigrun Ortleb. A positivity preserving and wellbalanced DG scheme using finite volume subcells in almost dry regions. Applied Mathematics and Computation, 272 (part 2)(??):259–273, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012023. See

comment [LZ19c].

Mohd:2019:DIS

[Moh19] Mohd Hafiz Mohd. Diversity in interaction strength promotes rich dynamical behaviours in a threespecies ecological system. Applied Mathematics and Computation, 353(??):243–253, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301055.

Monteiro:2019:DTD

[Mon19] L. H. A. Monteiro. A discrete-time dynamical system with four types of codimension-one bifurcations. Applied Mathematics and Computation, 354(??):189–191, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301389.

Mortici:2015:EFQ

[Mor15]  Cristinel Mortici.  Es-

timates of the function and quotient by Minc– Sathre. Applied Mathematics and Computation, 253(??):52–60, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017342.

Mortici:2017:SAS

[Mor17] Cristinel Mortici. Series associated to some expressions involving the volume of the unit ball and applications. Applied Mathematics and Computation, 294(??):121–138, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305227.

Mosic:2016:ROL

[Mos16] Dijana Mosić. Reverse order laws for the generalized strong Drazin inverses. Applied Mathematics and Computation, 284(??):37– 46, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301709.

Moussaoui:2015:RDE

[Mou15] Ali Moussaoui. A reactiondiffusion equations modelling the effect of fluctuating water levels on prey-predator interactions. Applied Mathematics and Computation, 268(??): 1110–1121, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315009303.

Merkin:2018:SPF

[MP18] J. H. Merkin and I. Pop.

Stagnation point flow past a stretching/shrinking sheet driven by Arrhenius kinetics. Applied Mathematics and Computation, 337(??):562–582, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304338.

Mena:2019:ESA

[MP19a] H. Mena and L. Pfurtscheller. An efficient SPDE approach for El Niño. Applied Mathematics and Computation, 352(??): 146–156, July 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319300888.

Miclaus:2019:NMA

[MP19b] Dan Miclaus and Laurian Ioan Piscoran. A new method for the approximation of integrals using the generalized Bernstein quadrature formula. Applied Mathematics and Computation, 340(??):146–155, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318306477.

Mukherjee:2017:CTD

[MPB^{+}17] Sayan Mukherjee, Sanjay Kumar Palit, Santo Banerjee, A. W. A. Wahab, MRK Ariffin, and D. K. Bhattacharya. Computing two dimensional Poincaré maps for hyperchaotic dynamics. Applied Mathematics and Computation, 301(??):140–154, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307603.

Motlagh:2015:AHE

[MPHG15] Omid Motlagh, Phillip Paevere, Tang Sai Hong, and George Grozev. Analysis of household electricity consumption behaviours: Impact of domestic electricity generation. Applied Mathematics and Computation, 270(??):165–178, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501084X.

Moreira:2019:PRP

[MPLS19] Carlos Alves Moreira, Eric Allison Philot, Angélica Nakagawa Lima, and Ana Ligia Scott. Predicting regions prone to protein aggregation based on SVM algorithm. Applied Mathematics and Computation, 359(??):502–511, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319302978.

Morosi:2017:NRC

[MPP17] Carlo Morosi, Mario Pernici, and Livio Pizzocchero. New results on the constants in some inequalities for the Navier– Stokes quadratic nonlinearity. Applied Mathematics and Computation, 308(??):54–72, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317301650.

Marusic-Paloka:2019:AAH

[MPPP19] Eduard Marusić-Paloka, Igor Pazanin, and Marija Prsa. Asymptotic analysis of the heat conduction problem in a dilated pipe. Applied Mathematics and Computation, 355(??):135–150, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301936.

Majeed:2017:NCA

[MPR^{+}17] Abdul Majeed, Abd Rahni Mt Piah, Muhammad Rafique, Johari Yap Abdullah, and Zainul Ahmad Rajion. NURBS curves with the application of multiple bones fracture reconstruction. Applied Mathematics and Computation, 315(??):70–84, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303739.

Mathiyalagan:2015:SDM

[MPS15] K. Mathiyalagan, Ju H.

Park, and R. Sakthivel. Synchronization for delayed memristive BAM neural networks using impulsive control with random nonlinearities. Applied Mathematics and Computation, 259(??): 967–979, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315003276.

Malyutina:2015:NAB

[MPW15] Elena Malyutina, Arcady Ponosov, and John Wyller. Numerical analysis of bump solutions for neural field equations with periodic microstructure. Applied Mathematics and Computation, 260(??):370–384, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003720.

Mancas:2015:IEE

[MR15a] Stefan C. Mancas and Haret C. Rosu. Integrable equations with Ermakov– Pinney nonlinearities and Chiellini damping. Applied Mathematics and Computation, 259(??):1– 11, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002179.

Marrero:2015:HIN

[MR15b] J. Abderramán Marrero and M. Rachidi. Hybrid Ikebe–Newton’s iteration for inverting general nonsingular Hessenberg matrices. Applied Mathematics and Computation, 268(??):413–421, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008656.

Mishra:2016:NSF

[MR16a] T. N. Mishra and K. N.

Rai. Numerical solution of FSPL heat conduction equation for analysis of thermal propagation. Applied Mathematics and Computation, 273(??):1006–1017, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014332.

Monsalve:2016:BDM

[MR16b] Juan Monsalve and Juan Rada. Bicyclic digraphs with maximal energy. Applied Mathematics and Computation, 280(??):124–131, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300376.

Maier:2019:DCS

[MR19] Helmut Maier and Michael Th.

Rassias. Distribution of a cotangent sum related to the Nyman–Beurling criterion for the Riemann Hypothesis. Applied Mathematics and Computation, 363(??): Article 124589, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305818.

Maharajan:2018:LBR

[MRC^{+}18] C. Maharajan, R. Raja, Jinde Cao, G. Rajchakit, Zhengwen Tue, and Ahmed Alsaedi. LMI-based results on exponential stability of BAM-type neural networks with leakage and both time-varying delays: A non-fragile state estimation approach. Applied Mathematics and Computation, 326(??):33–55, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300079.

Maleknejad:2018:NST

[MRE18] Khosrow Maleknejad, Jalil Rashidinia, and Tahereh Eftekhari. Numerical solution of three-dimensional Volterra–Fredholm integral equations of the first and second kinds based on Bernstein’s approximation. Applied Mathematics and Computation, 339(??):259–271, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305794.

Mehditabar:2018:VRA

[MRF18] Aref Mehditabar, Gholam H. Rahimi, and Keramat Malekzadeh Fard. Vibrational responses of antisymmetric angle-ply laminated conical shell by the methods of polynomial based differential quadrature and Fourier expansion based differential quadrature. Applied Mathematics and Computation, 320(??):580–595, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307166.

Mi:2019:IDH

[MRQZ19] Wen Mi, Hangmei Rao, Tao Qian, and Shouming Zhong. Identification of discrete Hammerstein systems by using adaptive finite rational orthogonal basis functions. Applied Mathematics and Computation, 361(??):354–364, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304539.

Motsa:2017:NSS

[MRR17] S. S. Motsa, Ch. Ram-Reddy, and Ch. Venkata Rao. Non-similarity solution for Soret effect on natural convection over the vertical frustum of a cone in a nanofluid using new bivariate pseudospectral local linearisation method. Applied Mathematics and Computation, 314(??):439–455, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304526.

Matic:2018:SPB

[MRS18] Ivan Matić, Rados Radoičić, and Dan Stefanica. A sharp Pólya-based approximation to the normal cumulative distribution function. Applied Mathematics and Computation, 322(??):111–122, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031730718X.

Monsalve:2019:EVE

[MRS19] Juan Monsalve, Juan Rada, and Yongtang Shi. Extremal values of energy over oriented bicyclic graphs. Applied Mathematics and Computation, 342(??):26–34, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307902.

Malik:2015:EPM

[MS15a] Sachin Malik and Rajesh Singh. Estimation of population mean using information on auxiliary attribute in twophase sampling. Applied Mathematics and Computation, 261(??):114–118, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003598.

Matsumoto:2015:DDC

[MS15b] Akio Matsumoto and Ferenc Szidarovszky. Delay dynamics of a Cournot game with heterogeneous duopolies. Applied Mathematics and Computation, 269(??):699–713, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009704.

Moldavskaya:2015:DZE

[MS15c] Elina Moldavskaya and Josef Shinar. Distribution of the zero-effort miss distance estimation error in interception problems. Applied Mathematics and Computation, 269(??):217–231, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009923.

Monegato:2015:STB

[MS15d] G. Monegato and L. Scuderi. A space–time BIE method for 2D wave equation problems with mixed boundary conditions. Applied Mathematics and Computation, 259(??):1046–1070, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002829.

Mishra:2016:CEL

[MS16a] Lakshmi Narayan Mishra and Mausumi Sen. On the concept of existence and local attractivity of solutions for some quadratic Volterra integral equation of fractional order. Applied Mathematics and Computation, 285(??):174–183, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301941.

Mishra:2016:APB

[MS16b] Vishnu Narayan Mishra and Preeti Sharma. On approximation properties of Baskakov–Schurer– Szász operators. Applied Mathematics and Computation, 281(??):381–393, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300339.

Majid:2017:SPC

[MS17a] Zanariah Abdul Majid and Phang Pei See. Study of predictor corrector block method via multiple shooting to Blasius and Sakiadis flow. Applied Mathematics and Computation, 314(??):469–483, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304514.

Maleknejad:2017:EMB

[MS17b] K. Maleknejad and E. Saeedipoor. An efficient method based on hybrid functions for Fredholm integral equation of the first kind with convergence analysis. Applied Mathematics and Computation, 304(??):93– 102, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300218.

Mahajan:2018:OCM

[MS18a] Amit Mahajan and Mahesh Kumar Sharma. The onset of convection in a magnetic nanofluid layer with variable gravity effects. Applied Mathematics and Computation, 339(??):607–621, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306283.

Miroshnichenko:2018:REC

[MS18b] Igor V. Miroshnichenko and Mikhail A. Sheremet. Radiation effect on conjugate turbulent natural convection in a cavity with a discrete heater. Applied Mathematics and Computation, 321(??):358–371, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730783X.

Ma:2019:CAP

[MS19a] Haifeng Ma and Predrag S.

Stanimirović. Characterizations, approximation and perturbations of the core-EP inverse. Applied Mathematics and Computation, 359(??):404–417, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930373X.

Mehrez:2019:GVF

[MS19b] Khaled Mehrez and Sergei M.

Sitnik. Generalized Volterra functions, its integral representations and applications to the Mathieutype series. Applied Mathematics and Computation, 347(??):578–589, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309779.

Mirzaee:2019:NSB

[MS19c] Farshid Mirzaee and Nasrin Samadyar. Numerical solution based on two-dimensional orthonormal Bernstein polynomials for solving some classes of two-dimensional nonlinear integral equations of fractional order. Applied Mathematics and Computation, 344–345(??):191–203, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308865.

Merino:2018:OED

[MSdC^{+}18] S. Merino, F. J. Sánchez, M. Sidrach de Cardona, F. Guzmán, R. Guzmán, J. Martınez, and P. J. Sotorrío. Optimization of energy distribution in solar panel array configurations by graphs and Minkowski’s paths. Applied Mathematics and Computation, 319(??):48–58, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300620.

Mnatsakanov:2015:ARP

[MSH15] Robert M. Mnatsakanov, Khachatur Sarkisian, and Artak Hakobyan. Approximation of the ruin probability using the scaled Laplace transform inversion. Applied Mathematics and Computation, 268(??):717–727, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008681.

Martinez:2016:ECD

[MSHG16] Eulalia Martínez, Sukhjit Singh, José L. Hueso, and Dharmendra K. Gupta. Enlarging the convergence domain in local convergence studies for iterative methods in Banach spaces. Applied Mathematics and Computation, 281(??):252–265, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300364.

Madhura:2017:IBE

[MSI17] K. R. Madhura, D. S.

Swetha, and S. S. Iyengar. The impact of Beltrami effect on dusty fluid flow through hexagonal channel in presence of porous medium. Applied Mathematics and Computation, 313(??):342–354, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300317304289.

Mokhlis:2015:FAS

[MSL15] H. Mokhlis, A. Shahriari, and J. A. Laghari. Fast and accurate second order load flow method based on fixed Jacobian matrix. Applied Mathematics and Computation, 269(??):584–593, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009947.

Mandal:2015:UAA

[MSM15] Akash Pradip Mandal, Sarifuddin, and Prashanta Kumar Mandal. An unsteady analysis of arterial drug transport from halfembedded drug-eluting stent. Applied Mathematics and Computation, 266(??):968–981, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008140.

Mishura:2015:CSM

[MSS15] Yuliya Mishura, Taras Shalaiko, and Georgiy Shevchenko. Convergence of solutions of mixed stochastic delay differential equations with applications. Applied Mathematics and Computation, 257(??):487–497, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000338.

Munoz:2015:NRR

[MT15] Ana I. Muñoz and J. Ignacio Tello. Numerical resolution of a reinforced random walk model arising in haptotaxis. Applied Mathematics and Computation, 256(??):415–424, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000570.

Mlayeh:2018:BBO

[MTB18] Rhouma Mlayeh, Samir Toumi, and Lotfi Beji. Backstepping boundary observer based-control for hyperbolic PDE in rotary drilling system. Applied Mathematics and Computation, 322(??):66–78, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308160.

Morales:2016:TFA

[MTCV^{+}16] Javier Morales, Víctor Tercero, José-Fernando Camacho-Vallejo, Alvaro E. Cordero, Luis E. López Nerio, and F-Javier Almaguer. Trend and fractality assessment of Mexico’s stock exchange. Applied Mathematics and Computation, 285(??):103–113, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316302077.

Mophou:2015:IVB

[MTJ15] G. Mophou, S. Tao, and C. Joseph. Initial value/ boundary value problem for composite fractional relaxation equation. Applied Mathematics and Computation, 257(??):134–144, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031401306X.

Manyombe:2017:MMP

[MTMB17] M. L. Mann Manyombe, B. Tsanou, J. Mbang, and S. Bowong. A metapopulation model for the population dynamics of anopheles mosquito. Applied Mathematics and Computation, 307(??):71–91, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301509.

Ma:2016:OSL

[MTY16] Ran Ma, Jiping Tao, and Jinjiang Yuan. Online scheduling with linear deteriorating jobs to minimize the total weighted completion time. Applied Mathematics and Computation, 273(??):570–583, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014095.

Mukhtar:2016:GTP

[Muk16] Faisal M. Mukhtar. Generalized Taylor polynomials for axisymmetric plates and shells. Applied Mathematics and Computation, 276(??):182–199, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315016045.

Muslim:2018:FGA

[Mus18]  M. Muslim.  Faedo–

Galerkin approximation of second order nonlinear differential equation with deviated argument. Applied Mathematics and Computation, 329(??):315–324, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300894.

Mortici:2015:RLH

[MV15] Cristinel Mortici and Mark B. Villarino. On the Ramanujan–Lodge harmonic number expansion. Applied Mathematics and Computation, 251(??):423–430, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016208.

Martin-Vergara:2019:PNS

[MVRV19] Francisca Martin-Vergara, Francisco Rus, and Francisco R. Villatoro. Padé numerical schemes for the sine-Gordon equation. Applied Mathematics and Computation, 358(??):232–243, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300319303303.

Martin-Vaquero:2019:TLF

[MVS19] Jesús Martín-Vaquero and Svajunas Sajavicius. The two-level finite difference schemes for the heat equation with nonlocal initial condition. Applied Mathematics and Computation, 342(??):166–177, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307975.

Matculevich:2018:PEA

[MW18] Svetlana Matculevich and Monika Wolfmayr. On the a posteriori error analysis for linear Fokker–Planck models in convectiondominated diffusion problems. Applied Mathematics and Computation, 339(??):768–778, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304673.

Moroz:2018:FCI

[MWH^{+}18] Leonid V. Moroz, Cezary J.

Walczyk, Andriy Hrynchyshyn, Vijay Holimath, and Jan L. Cieśliński. Fast calculation of inverse square root with the use of magic constant — analytical approach. Applied Mathematics and Computation, 316(??):245–255, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305763.

Mao:2018:SEN

[MWL^{+}18] Ying Mao, Liqing Wang, Yang Liu, Jianquan Lu, and Zhen Wang. Stabilization of evolutionary networked games with length-r information. Applied Mathematics and Computation, 337(??):428–441, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304363.

Ma:2017:NAP

[MWLL17] Jichao Ma, Gaofeng Wei, Dandan Liu, and Gongtian Liu. The numerical analysis of piezoelectric ceramics based on the Hermitetype RPIM. Applied Mathematics and Computation, 309(??):170–182, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302291.

Ma:2017:CID

[MWRT17] Jun Ma, Fuqiang Wu, Guodong Ren, and Jun Tang. A class of initialsdependent dynamical systems. Applied Mathematics and Computation, 298 (??):65–76, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316306622.

Ma:2015:QAS

[MWWM15] Qinghua Ma, Rongnian Wang, Junwei Wang, and Yicheng Ma. Qualitative analysis for solutions of a certain more generalized two-dimensional fractional differential system with Hadamard derivative. Applied Mathematics and Computation, 257(??):436–445, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014593.

Miao:2015:NAC

[MWY15] Changqing Miao, Yintao Wei, and Xiangqiao Yan. A numerical analysis of a center circularhole crack in a rectangular tensile sheet. Applied Mathematics and Computation, 250(??):356–367, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015185.

Ma:2017:NST

[MWY17] Zhaohai Ma, Xin Wu, and Rong Yuan. Nonlinear stability of traveling wavefronts for competitive– cooperative Lotka–Volterra systems of three species. Applied Mathematics and Computation, 315(??): 331–346, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317305295.

Ma:2017:RAC

[MX17] Qiaozhen Ma and Ling Xu. Random attractors for the coupled suspension bridge equations with white noises. Applied Mathematics and Computation, 306(??):38–48, August 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301236.

Ma:2019:ILT

[MY19] Fuhong Ma and Jin Yan. An improvement of Lichiardopol’s theorem on disjoint cycles in tournaments. Applied Mathematics and Computation, 347(??):162–168, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309603.

Miao:2015:GNM

[MYH15] Xin-He Miao, Jiantao Yang, and Shenglong Hu. A generalized Newton method for absolute value equations associated with circular cones. Applied Mathematics and Computation, 269(??):155–168, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009832.

Meng:2018:NSN

[MYHS18] Zhijun Meng, Mingxu Yi, Jun Huang, and Lei Song. Numerical solutions of nonlinear fractional differential equations by alternative Legendre polynomials. Applied Mathematics and Computation, 336(??):454–464, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303989.

Ma:2015:EPF

[MZ15] Cheng Ma and Liansheng Zhang. On an exact penalty function method for nonlinear mixed discrete programming problems and its applications in search engine advertising problems. Applied Mathematics and Computation, 271(??):642–656, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012382.

Ma:2017:SPN

[MZ17] Li-Yuan Ma and Zuo-Nong Zhu. Spatial properties and numerical solitary waves of a nonintegrable discrete nonlinear Schrödinger equation with nonlinear hopping. Applied Mathematics and Computation, 309(??):93–106, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S009630031730231X.

Ma:2019:PEC

[MZC19] Guiyuan Ma, Song-Ping Zhu, and Wenting Chen. Pricing European call options under a hard-to-borrow stock model. Applied Mathematics and Computation, 357(??): 243–257, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300319302784.

Mao:2015:EUA

[MZM15] Wei Mao, Quanxin Zhu, and Xuerong Mao. Existence, uniqueness and almost surely asymptotic estimations of the solutions to neutral stochastic functional differential equations driven by pure jumps. Applied Mathematics and Computation, 254(??):252–265, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017871.

Ma:2016:MCA

[MZM16] Wentao Ma, Baowen Zhang, and Hailong Ma. A meshless collocation approach with barycentric rational interpolation for two-dimensional hyperbolic telegraph equation. Applied Mathematics and Computation, 279(??):236–248, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300224.

Meng:2017:PBG

[MZM17] Lingsheng Meng, Bing Zheng, and Peilan Ma. Perturbation bounds of generalized inverses. Applied Mathematics and Computation, 296(??):88– 100, March 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316306154.

Ma:2015:CCA

[MZP15] Yanqin Ma, Lili Zhang, and Pingqi Pan. Crisscross algorithm based on the most-obtuse-angle rule and deficient basis. Applied Mathematics and Computation, 268(??):439–449, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315008619.

Minhajul:2018:WID

[MZS18] Minhajul, D. Zeidan, and T. Raja Sekhar. On the wave interactions in the drift–flux equations of two-phase flows. Applied Mathematics and Computation, 327(??):117–131, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300341.

Ngoc:2019:SNV

[NA19] Pham Huu Anh Ngoc and Tran The Anh. Stability of nonlinear Volterra equations and applications. Applied Mathematics and Computation, 341(??):1–14, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830585X.

Naanaa:2015:FCO

[Naa15] Anis Naanaa. Fast chaotic optimization algorithm based on spatiotemporal maps for global optimization. Applied Mathematics and Computation, 269(??):402–411, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010334.

Nadarajah:2015:CFE

[Nad15] Saralees Nadarajah. On closed form expressions for probability of operational success and expected lifetime. Applied Mathematics and Computation, 262(??):308–311, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005263.

Nagy:2017:NST

[Nag17] A. M. Nagy. Numerical solution of time fractional nonlinear Klein– Gordon equation using sinc-Chebyshev collocation method. Applied Mathematics and Computation, 310(??):139–148, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302734.

Nakajo:2015:SCG

[Nak15] Kazuhide Nakajo. Strong convergence for gradient projection method and relatively nonexpansive mappings in Banach spaces. Applied Mathematics and Computation, 271(??):251–258, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011522.

Nakamori:2018:LSF

[Nak18] Seiichi Nakamori. Leastsquares finite impulse response fixed-lag smoother and filter in linear discretetime stochastic systems. Applied Mathematics and Computation, 334(??):94– 106, October 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318303023.

Nojavan:2018:LVS

[NAM18] Hananeh Nojavan, Saeid Abbasbandy, and Maryam Mohammadi. Local variably scaled Newton basis functions collocation method for solving Burgers’ equation. Applied Mathematics and Computation, 330(??):23–41, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318301012.

Nam:2019:HUS

[Nam19] Young Woo Nam. Hyers– Ulam stability of loxodromic Möbius difference equation. Applied Mathematics and Computation, 356(??):119–136, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302346.

Nass:2019:LSA

[Nas19] Aminu M. Nass. Lie symmetry analysis and exact solutions of fractional ordinary differential equations with neutral delay. Applied Mathematics and Computation, 347(??):370–380, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309755.

Navarro:2017:WEP

[Nav17] Juan F. Navarro. Windows for escaping particles in quartic galactic potentials. Applied Mathematics and Computation, 303(??):190–202, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300541.

Neumann:2015:DAL

[NB15] Philipp Neumann and Hans-Joachim Bungartz. Dynamically adaptive lattice Boltzmann simulation of shallow water flows with the Peano framework. Applied Mathematics and Computation, 267(??):795–804, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014155.

Nashine:2019:SCC

[NB19a] Hemant Kumar Nashine and Snehasish Bose. Solution of a class of crosscoupled nonlinear matrix equations. Applied Mathematics and Computation, 362(??): Article 124534, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305132.

Noormohammadi:2019:EFW

[NB19b] N. Noormohammadi and B. Boroomand. Enrichment functions for weak singularities in 2D elastic problems with isotropic and orthotropic materials. Applied Mathematics and Computation, 350(??):402–415, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300244.

Nafidi:2019:SWD

[NBAGS19] A. Nafidi, M. Bahij, B. Achchab, and R. Gutiérrez-Sanchez. The stochastic Weibull diffusion process: Computational aspects and simulation. Applied Mathematics and Computation, 348(??):575–587, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310646.

Narang:2016:NTP

[NBK16] Mona Narang, Saurabh Bhatia, and V. Kanwar. New two-parameter Chebyshev–Halley-like family of fourth and sixthorder methods for systems of nonlinear equations. Applied Mathematics and Computation, 275(??):394–403, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015659.

Navarro:2016:AFR

[NC16] H. A. Navarro and L. Cveticanin. Amplitude-frequency relationship obtained using Hamiltonian approach for oscillators with sum of non-integer order nonlinearities. Applied Mathematics and Computation, 291(??):162–171, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304271.

Nedovic:2019:SCP

[NC19] Maja Nedović and Ljiljana Cvetković. The Schur complement of PH-matrices. Applied Mathematics and Computation, 362(??): Article 124541, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930520X.

Naeeni:2015:PCN

[NCEGA15] M. Raoofian Naeeni, R. Campagna, M. Eskandari-Ghadi, and Alireza A. Ardalan. Performance comparison of numerical inversion methods for Laplace and Hankel integral transforms in engineering problems. Applied Mathematics and Computation, 250(??):759–775, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014775.

Neta:2016:CBA

[NCS16] Beny Neta, Changbum Chun, and Melvin Scott. Corrigendum to “Basins of attraction for optimal eighth-order methods to find simple roots of nonlinear equations”. Applied Mathematics and Computation, 281(??):396–403, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300023. See

[CN15c].

Ni:2015:NSR

[ND15] Wei-Guo Ni and Chao-Qing Dai. Note on same result of different ansätz based on extended tanh-function method for nonlinear models. Applied Mathematics and Computation, 270(??):434–440, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011078.

Nemecek:2015:NMA

[NDZ15] J. Nemecek, F. Denk, and P. Zlámal. Numerical modeling of aluminium foam on two scales. Applied Mathematics and Computation, 267(??):506–516, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001162.

Ntouyas:2015:ESF

[NE15] Sotiris K. Ntouyas and Sina Etemad. On the existence of solutions for fractional differential inclusions with sum and integral boundary conditions. Applied Mathematics and Computation, 266(??):235–243, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006530.

Najafalizadeh:2016:NMS

[NE16] S. Najafalizadeh and R. Ezzati. Numerical methods for solving twodimensional nonlinear integral equations of fractional order by using twodimensional block pulse operational matrix. Applied Mathematics and Computation, 280(??):46– 56, April 20, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315300308.

Novak:2018:DIP

[NEK18] Miroslav Novak, Jakub Eichler, and Miloslav Kosek. Difficulty in identification of Preisach hysteresis model weighting function using first order reversal curves method in soft magnetic materials. Applied Mathematics and Computation, 319(??):469–485, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303053.

Neto:2015:NDI

[Net15] Antônio Francisco Neto. A note on a determinant identity. Applied Mathematics and Computation, 264(??):246–248, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500541X.

Neustupa:2016:SFT

[Neu16] Tomás Neustupa. A steady flow through a plane cascade of profiles with an arbitrarily large inflow: The mathematical model, existence of a weak solution. Applied Mathematics and Computation, 272 (part 3)(??):687–691, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006906.

Nieto:2015:HPO

[NFS^{+}15] P. J. García Nieto, J. R. Alonso Fernández, V. M. González Suárez, C. Díaz Muñiz, E. García-Gonzalo, and R. Mayo Bayón. A hybrid PSO optimized SVMbased method for predicting of the cyanotoxin content from experimental cyanobacteria concentrations in the trasona reservoir: a case study in Northern Spain. Applied Mathematics and Computation, 260(??):170–187, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003896.

Nahla:2017:PIP

[NH17] Abdallah A. Nahla and A. A. Hemeda. Picard iteration and Padé approximations for stiff fractional point kinetics equations. Applied Mathematics and Computation, 293(??):72–80, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305033.

Ning:2015:IDL

[NHWZ15] Chongyang Ning, Yong He, Min Wu, and Shaowu Zhou. Indefinite derivative Lyapunov–Krasovskii functional method for input to state stability of nonlinear systems with time-delay. Applied Mathematics and Computation, 270(??):534–542, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315011194.

Neoh:2019:RDS

[NI19] S. S. Neoh and F. Ismail. Residual distribution schemes for Maxwell’s equations. Applied Mathematics and Computation, 348(??):275–317, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310294.

Nishizawa:2015:SJT

[Nis15] Yusuke Nishizawa. Sharpening of Jordan’s type and Shafer–Fink’s type inequalities with exponential approximations. Applied Mathematics and Computation, 269(??):146–154, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009601.

Nepomuceno:2018:ICP

[NJM^{+}18] Erivelton G. Nepomuceno, Heitor M. Rodrigues Junior, Samir A. M. Martins, Matjaz Perc, and Mitja Slavinec. Interval computing periodic orbits of maps using a piecewise approach. Applied Mathematics and Computation, 336(??):67–75, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318303898.

Niu:2019:IDP

[NJTP19] Yi-Shuai Niu, Joaquim Júdice, Hoai An Le Thi, and Dinh Tao Pham. Improved dc programming approaches for solving the quadratic eigenvalue complementarity problem. Applied Mathematics and Computation, 353(??):95–113, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301158.

Nemecek:2018:SCM

[NKKK18] J. Nemecek, J. Kruis, T. Koudelka, and T. Krejcí. Simulation of chloride migration in reinforced concrete. Applied Mathematics and Computation, 319(??):575–585, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304903.

Nikolayev:2016:IAT

[NKKS16] Denys Nikolayev, Zdenek Kubík, Pavel Karban, and Jirí Skála. Impedance analysis of transmission line cells for EMC applications using Agros2D. Applied Mathematics and Computation, 289(??): 381–387, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316303368.

Nowakowska:2015:CAG

[NKL15] Ewa Nowakowska, Jacek Koronacki, and Stan Lipovetsky. Clusterability assessment for Gaussian mixture models. Applied Mathematics and Computation, 256(??):591–601, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016920.

Nobari:2016:NII

[NKNA16] M. R. H. Nobari, S. Kazemi, V. Nourian, and M. Adib. A numerical investigation of incompressible viscous flow in a helical square annulus. Applied Mathematics and Computation, 277(??):127–141, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300230.

Nguyen:2016:FEN

[NKP16] Thien Nguyen, Kestutis Karciauskas, and Jörg Peters. C^{1} finite elements on non-tensor-product 2 d and 3 d manifolds. Applied Mathematics and Computation, 272 (part 1)(??):148–158, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500884X.

Naumann:2018:TOC

[NKS18] Alexander Naumann, Oliver Kolb, and Matteo Semplice. On a third order CWENO boundary treatment with application to networks of hyperbolic conservation laws. Applied Mathematics and Computation, 325(??):252–270, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317309086.

Nemati:2018:NSN

[NL18a] S. Nemati and P. M. Lima. Numerical solution of nonlinear fractional integro-differential equations with weakly singular kernels via a modification of hat functions.

Applied Mathematics and Computation, 327(??):79– 92, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300511.

Niu:2018:FCP

[NL18b] Min Niu and Qiuxia Liu. Factor complexity and permutation complexity of the generalized Morse sequence. Applied Mathematics and Computation, 329(??):395–401, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301218.

Nie:2019:MAC

[NLC19] Xiaobing Nie, Jinling Liang, and Jinde Cao. Multistability analysis of competitive neural networks with Gaussianwavelet-type activation functions and unbounded time-varying delays. Applied Mathematics and Computation, 356(??): 449–468, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300319302279.

Najafi:2015:MSS

[NM15] Amir Abbas Najafi and Siamak Mushakhian. Multistage stochastic meansemivariance-CVaR portfolio optimization under transaction costs. Applied Mathematics and Computation, 256(??):445–458, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000806.

Nandal:2019:NSA

[NM19] Reena Nandal and Amit Mahajan. Nonlinear stability analysis of a fluid saturated anisotropic Darcy–Brinkman medium with internal heat source. Applied Mathematics and Computation, 358(??): 216–231, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319302243.

Nenezic:2016:NAS

[NMM16] Marija Nenezić, Branko Malesević, and Cristinel Mortici. New approximations of some expressions involving trigonometric functions. Applied Mathematics and Computation, 283(??):299–315, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301461.

Nguyen-Minh:2015:SFV

[NMNTBXVD15] N. Nguyen-Minh, T. Nguyen-Thoi, T. Bui-Xuan, and T. Vo-Duy. Static and free vibration analyses of stiffened folded plates using a cell-based smoothed discrete shear gap method (CS–FEM– DSG3). Applied Mathematics and Computation, 266(??):212–234, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006591.

Nepomuceno:2018:DUC

[NMS^{+}18] Erivelton G. Nepomuceno, Samir A. M. Martins, Bruno C. Silva, Gleison F. V. Amaral, and Matjaz Perc. Detecting unreliable computer simulations of recursive functions with interval extensions. Applied Mathematics and Computation, 329(??):408–419, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830122X.

Nowruzi:2019:ESC

[NN19] Hashem Nowruzi and S. Salman Nourazar. Erratum to “On the solution of characteristic value problems arising in linear stability analysis; semi analytical approach”. Applied Mathematics and Computation, 350(??):462, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031930044X. See

[HN14].

Noor:2015:SQE

[NNA15a] Muhammad Aslam Noor, Khalida Inayat Noor, and Muhammad Uzair Awan. Some quantum estimates for Hermite–Hadamard inequalities. Applied Mathematics and Computation, 251(??):675–679, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016221.

Noor:2015:SQI

[NNA15b] Muhammad Aslam Noor, Khalida Inayat Noor, and Muhammad Uzair Awan.

Some quantum integral inequalities via preinvex functions. Applied Mathematics and Computation, 269(??):242–251, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010000.

Nguyen:2018:NBE

[NND^{+}18] Dinh Kien Nguyen, Khoa Viet Nguyen, Van Manh Dinh, Buntara S. Gan, and Sergei Alexandrov. Nonlinear bending of elastoplastic functionally graded ceramic-metal beams subjected to nonuniform distributed loads. Applied Mathematics and Computation, 333(??):443–459, September 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318302819.

Ni:2018:LST

[NP18] He-Xia Ni and Hao Pan. On the lacunary sum of trinomial coefficients. Applied Mathematics and Computation, 339(??):272–285, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305861.

Nguyen:2019:ROO

[NPT19] Cuong M. Nguyen, Pubudu N. Pathirana, and Hieu Trinh. Robust observer and observer-based control designs for discrete one-sided Lipschitz systems subject to uncertainties and disturbances. Applied Mathematics and Computation, 353(??):42– 53, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300815.

Natalini:2015:BPM

[NR15] Pierpaolo Natalini and Paolo Emilio Ricci. Bell polynomials and modified Bessel functions of halfintegral order. Applied Mathematics and Computation, 268(??):270–274, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008504.

Nagamani:2016:SDP

[NR16] G. Nagamani and S. Ramasamy. Stochastic dissipativity and passivity analysis for discretetime neural networks with probabilistic time-varying delays in the leakage term. Applied Mathematics and Computation, 289(??):237–257, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303022.

Naimzada:2018:HCC

[NR18a] A. K. Naimzada and Roberto Raimondo. Heterogeneity and chaos in congestion games. Applied Mathematics and Computation, 335(??):278–291, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303217.

Naimzada:2018:CCG

[NR18b] Ahmad Kabir Naimzada and Roberto Raimondo. Chaotic congestion games. Applied Mathematics and Computation, 321(??): 333–348, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317307208.

Nino-Ruiz:2019:RSL

[NRAEC19] Elias D. Nino-Ruiz, Carlos Ardila, Jesus Estrada, and Jose Capacho. A reduced-space line-search method for unconstrained optimization via random descent directions. Applied Mathematics and Computation, 341:15–30, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318307343.

Niu:2019:CHM

[NRS19] Chunyan Niu, Hongxing Rui, and Ming Sun. A coupling of hybrid mixed and continuous Galerkin finite element methods for poroelasticity. Applied Mathematics and Computation, 347(??):767–784, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309950.

Navickas:2016:ESS

[NRT16] Z. Navickas, M. Ragulskis, and T. Telksnys. Existence of solitary solutions in a class of nonlinear differential equations with polynomial nonlinearity. Applied Mathematics and Computation, 283(??):333–338, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301680.

Novara:2016:ISA

[NRY16] Paola Novara, Lucia Romani, and Jungho Yoon. Improving smoothness and accuracy of Modified Butterfly subdivision scheme. Applied Mathematics and Computation, 272 (part 1)(??):64–79, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009844.

Nath:2015:BCS

[NS15a] Triloki Nath and S. R.

Singh. Boundedness of certain sets of Lagrange multipliers in vector optimization. Applied Mathematics and Computation, 271(??):429–435, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007432.

Negi:2015:RAN

[NS15b] Seema Negi and S. B.

Singh. Reliability analysis of non-repairable complex system with weighted subsystems connected in series. Applied Mathematics and Computation, 262(??):79–89, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004385.

Neusser:2017:NSC

[NS17a] J. Neusser and V. Schleper. Numerical schemes for the coupling of compressible and incompressible fluids in several space dimensions. Applied Mathematics and Computation, 304(??):65–82, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300528.

Nosrati:2017:DAF

[NS17b] Komeil Nosrati and Masoud Shafiee. Dynamic analysis of fractionalorder singular Holling type-II predator–prey system. Applied Mathematics and Computation, 313(??):159–179, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730379X.

Niu:2016:ACB

[NSM16] Wei Niu, Jian Shi, and Chenqi Mou. Analysis of codimension 2 bifurcations for high-dimensional discrete systems using symbolic computation methods. Applied Mathematics and Computation, 273(??):934–947, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013417.

Nannyonga:2018:AAM

[NSN^{+}18] Betty Nannyonga, Joseph Ssebuliba, Juliet Nakakawa, Betty Nabiyonga, and J. Y. T. Mugisha. To apprehend or not to apprehend: a mathematical model for ending student strikes in a university. Applied Mathematics and Computation, 339(??):588–606, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830599X.

Nguyen:2019:QFC

[NSR^{+}19] Linh Kieu Nguyen, Chanyoung Song, Joonghyun Ryu, Phan Thanh An, Nam-Dũng Hoang, and Deok-Soo Kim. QuickhullDisk: a faster convex hull algorithm for disks. Applied Mathematics and Computation, 363 (??):Article 124626, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319306186.

Nguyen:2016:UIO

[NT16] Minh Cuong Nguyen and Hieu Trinh. Unknown input observer design for one-sided Lipschitz discrete-time systems subject to time-delay. Applied Mathematics and Computation, 286(??):57–71, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316302454.

Nagashima:2019:SPF

[NT19] Keisuke Nagashima and Jun Tanimoto. A stochastic pairwise Fermi rule modified by utilizing the average in payoff differences of neighbors leads to increased network reciprocity in spatial prisoner’s dilemma games. Applied Mathematics and Computation, 361(??): 661–669, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL https://www.sciencedirect.
com/science/article/
pii/S0096300319304369.

Noor:2015:NITb

[NWN15] Muhammad Aslam Noor, Muhammad Waseem, and Khalida Inayat Noor. New iterative technique for solving a system of nonlinear equations. Applied Mathematics and Computation, 271(??):446–466, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315012060.

Noor:2015:NITa

[NWNA15] Muhammad Aslam Noor, Muhammad Waseem, Khalida Inayat Noor, and Muhammad Aamir Ali. New iterative technique for solving nonlinear equations. Applied Mathematics and Computation, 265(??):1115–1125, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007602.

Nie:2015:SEE

[NY15] Cunyun Nie and Haiyuan Yu. Some error estimates on the large jump asymptotic approach for the interface problem. Applied Mathematics and Computation, 263(??):268–279, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005020.

Nian:2018:ESM

[NY18] Fuzhong Nian and Shuanglong Yao. The epidemic spreading on the multi-relationships network. Applied Mathematics and Computation, 339(??):853–865, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305952.

Nadeem:2016:TPL

[NZZ16] Muhammad Faisal Nadeem, Sohail Zafar, and Zohaib Zahid. On topological properties of the line graphs of subdivision graphs of certain nanostructures. Applied Mathematics and Computation, 273(??):125–130, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013429.

Oliveira:2015:NAF

[OA15] Saulo P. Oliveira and Juarez S. Azevedo. Numerical approximation of 2D Fredholm integral eigenvalue problems by orthogonal wavelets. Applied Mathematics and Computation, 267(??):517–528, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315001150.

Oulghelou:2018:FRS

[OA18] M. Oulghelou and C. Allery.

A fast and robust suboptimal control approach using reduced order model adaptation techniques. Applied Mathematics and Computation, 333(??): 416–434, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318302728.

Orkcu:2015:EPPa

[ÖAD15] H. Hasan Örkcü, Ertuğrul Aksoy, and Mustafa Isa Doğan. Estimating the parameters of 3-p Weibull distribution through differential evolution. Applied Mathematics and Computation, 251(??):211–224, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015021.

Ozyapici:2018:GST

[OB18] Ali Ozyapici and Bülent Bilgehan. Generalized system of trial equation methods and their applications to biological systems. Applied Mathematics and Computation, 338(??):711–721, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305083.

Obradovic:2019:INM

[Obr19] Maja Obradović. Implicit numerical methods for neutral stochastic differential equations with unbounded delay and Markovian switching. Applied Mathematics and Computation, 347(??):664–687, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310117.

Oanh:2017:ARF

[ODP17] Dang Thi Oanh, Oleg Davydov, and Hoang Xuan Phu. Adaptive RBF-FD method for elliptic problems with point singularities in 2D. Applied Mathematics and Computation, 313(??):474–497, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304186.

Ouyang:2019:AGH

[OFwCW19] Peichang Ouyang, Robert W.

Fathauer, Kwok wai Chung, and Xinchang Wang. Automatic generation of hyperbolic drawings. Applied Mathematics and Computation, 347(??):653–663, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308312.

Ouyang:2019:RJP

[OGC19] Zhangdong Ouyang, Jun Ge, and Yichao Chen.

Remarks on the joins of 1-planar graphs. Applied Mathematics and Computation, 362(??): Article 124537, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305168.

Ozarslan:2019:BML

[ÖK19] Mehmet Ali Özarslan and Cemaliye Kürt. Bivariate Mittag-Leffler functions arising in the solutions of convolution integral equation with 2D–Laguerre– Konhauser polynomials in the kernel. Applied Mathematics and Computation, 347(??):631–644, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309834.

Okayama:2018:EEE

[Oka18a] Tomoaki Okayama. Error estimates with explicit constants for the sinc approximation over infinite intervals. Applied Mathematics and Computation, 319(??):125–137, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301261.

Okayama:2018:TAS

[Oka18b] Tomoaki Okayama. Theoretical analysis of a Sinc–Nyström method for Volterra integro-differential equations and its improvement. Applied Mathematics and Computation, 324(??):1–15, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308524.

Ouyang:2015:PAB

[OL15] Mengqian Ouyang and Xiaoyue Li. Permanence and asymptotical behavior of stochastic prey-predator system with Markovian switching. Applied Mathematics and Computation, 266(??):539–559, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007079.

Osman:2019:IKB

[OL19] S. A. Osman and T. A. M.

Langlands. An implicit Keller Box numerical scheme for the solution of fractional subdiffusion equations. Applied Mathematics and Computation, 348(??):609–626, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310622.

Oliveira:2017:RFM

[Oli17] Jorge F. Oliveira. Radio frequency and microwave numerical simulation techniques based on multivariate formulations. Applied Mathematics and Computation, 294(??):238–252, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305756.

Ou:2019:NAS

[OLZ19] Minghui Ou, Shan Liang, and Cheng Zeng. A novel approach to stable zero dynamics of sampleddata models for nonlinear systems in backward triangle sample and hold case. Applied Mathematics and Computation, 355(??):47–60, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S009630031930178X.

Onitsuka:2018:HUS

[Oni18] Masakazu Onitsuka. Hyers– Ulam stability of firstorder nonhomogeneous linear difference equations with a constant stepsize. Applied Mathematics and Computation, 330(??):143–151, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301383.

Orkcu:2015:EPPb

[ÖÖAD15] H. Hasan Örkcü, Volkan Soner Özsoy, Ertugrul Aksoy, and Mustafa Isa Dogan. Estimating the parameters of 3-p Weibull distribution using particle swarm optimization: a comprehensive experimental comparison. Applied Mathematics and Computation, 268(??):201–226, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008243.

Orera:2019:INB

[OP19] H. Orera and J. M. Peña.

Infinity norm bounds for the inverse of Nekrasov matrices using scaling matrices. Applied Mathematics and Computation, 358(??):119–127, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303091.

Okrasinska-Plociniczak:2018:NMV

[OPP18] Hanna Okrasińska-Plociniczak and Lukasz Plociniczak. Numerical method for Volterra equation with a power-type nonlinearity. Applied Mathematics and Computation, 337(??):442–451, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304454.

Occorsio:2018:NMB

[OR18a] Donatella Occorsio and Maria Grazia Russo. Nyström methods for bivariate Fredholm integral equations on unbounded domains. Applied Mathematics and Computation, 318(??):19–34, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317304964.

ORegan:2018:CMB

[O’R18b] Donal O’Regan. Coincidence for morphisms based on compactness conditions on countable sets. Applied Mathematics and Computation, 339(??):45–58, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305551.

Ortigoza:2015:UTC

[Ort15a] Gerardo M. Ortigoza. Unstructured triangular cellular automata for modeling geographic spread. Applied Mathematics and Computation, 258(??): 520–536, May 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315001484.

Ortigueira:2015:CDF

[Ort15b] Manuel D. Ortigueira.

Comments on “Discretization of fractional order differentiator over Paley– Wiener space”. Applied Mathematics and Computation, 270(??):44–46, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300315010991.

Oguz:2015:CCM

[OS15] Cem Oguz and Mehmet Sezer. Chelyshkov collocation method for a class of mixed functional integro-differential equations. Applied Mathematics and Computation, 259(??):943–954, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500329X.

Olgun:2015:GJO

[OTE15] A. Olgun, F. Tasdelen, and A. Erençin. A generalization of Jain’s operators. Applied Mathematics and Computation, 266(??):6–11, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006840.

Othman:2016:RMO

[OTH16] Mohamed I. A. Othman, Ramadan S. Tantawi, and Mohamed I. M. Hilal. Rotation and modified Ohm’s law influence on magnetothermoelastic micropolar material with microtemperatures. Applied Mathematics and Computation, 276(??):468–480, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300199.

Ourraoui:2016:MRS

[Our16] Anass Ourraoui. Multiplicity results for Steklov problem with variable exponent. Applied Mathematics and Computation, 277(??):34–43, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031530031X.

Osman:2018:EAC

[OW18] M. S. Osman and Abdul-Majid Wazwaz. An efficient algorithm to construct multi-soliton rational solutions of the (2 + 1)-dimensional KdV equation with variable coefficients. Applied Mathematics and Computation, 321(??):282–289, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307415.

Ozbekler:2015:STS

[Özb15]  A. Özbekler.  Stur-

mian theory for second order differential equations with mixed nonlinearities. Applied Mathematics and Computation, 259(??):379–389, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002957.

Patro:2015:CMC

[PABA15] Kanan K. Patro, M. M. Acharya, M. P. Biswal, and Srikumar Acharya. Computation of a multichoice goal programming problem. Applied Mathematics and Computation, 271(??):489–501, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315012667.

Pachpatte:2019:PCI

[Pac19] Deepak B. Pachpatte. Properties of certain iterated dynamic integrodiffetential equation on time scales. Applied Mathematics and Computation, 346(??):767–775, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309007.

Pakdaman:2017:SDE

[PAE^{+}17] M. Pakdaman, A. Ahmadian, S. Effati, S. Salahshour, and D. Baleanu. Solving differential equations of fractional order using an optimization technique based on training artificial neural network. Applied Mathematics and Computation, 293(??):81–95, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304593.

Pai:2015:CSM

[Pai15a] Ming-Chang Pai. Chaotic sliding mode controllers for uncertain time-delay chaotic systems with input nonlinearity. Applied Mathematics and Computation, 271(??):757–767, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315012965.

Pai:2015:DTD

[Pai15b] Ming-Chang Pai. Discretetime dynamic output feedback input shaping control of vibration in uncertain time-delay flexible structures. Applied Mathematics and Computation, 250(??):675–688, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015616.

Parastesh:2019:STN

[PAJ^{+}19] Fatemeh Parastesh, Hamed Azarnoush, Sajad Jafari, Boshra Hatef, Matjaž Perc, and Robert Repnik. Synchronizability of two neurons with switching in the coupling. Applied Mathematics and Computation, 350(??):217–223, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300190.

Polatoglu:2015:CHM

[PAK15] Yasar Polatoğlu, Melike Aydoğan, and Yasemin Kahramaner. On the class of harmonic mappings which is related to the class of bounded boundary rotation. Applied Mathematics and Computation, 267(??):790–794, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014301.

Palubeckis:2015:FSA

[Pal15] Gintaras Palubeckis. Fast simulated annealing for single-row equidistant facility layout. Applied Mathematics and Computation, 263(??):287–301, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005354.

Palacios:2017:BAZ

[Pal17] José Luis Palacios. Bounds for the augmented Zagreb and the atom-bond connectivity indices. Applied Mathematics and Computation, 307(??):141–145, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301807. See

correction [Pal19].

Palacios:2019:CBA

[Pal19] José Luis Palacios. Correction to “Bounds for the augmented Zagreb and the atom-bond connectivity indices”. Applied Mathematics and Computation, 362(??): Article 124559, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305429.

See [Pal17].

Pahnehkolaei:2019:DIR

[PAM19] Seyed Mehdi Abedi Pahnehkolaei, Alireza Alfi, and J. A. Tenreiro Machado. Delay independent robust stability analysis of delayed fractional quaternionvalued leaky integrator echo state neural networks with QUAD condition. Applied Mathematics and Computation, 359(??):278–293, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303856.

Pantokratoras:2015:CPE

[Pan15] Asterios Pantokratoras. Comment on the paper “Effect of viscous dissipation and heat source on unsteady boundary layer flow and heat transfer past a stretching surface embedded in a porous medium using element free Galerkin method, by R. Sharma, Applied Mathematics and Computation (2012) 219:976– 987”. Applied Mathematics and Computation, 266(??):70–71, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500661X. See

[Sha12].

Panigrahi:2018:HCM

[Pan18] Bijaya Laxmi Panigrahi.

Hybrid collocation methods for eigenvalue problem of a compact integral operator with weakly singular kernel. Applied Mathematics and Computation, 328(??):353–364, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830081X.

Paripour:2015:NNS

[Par15] Mahmoud Paripour. Note on numerical solution of the Fredholm singular integro-differential equation with Cauchy kernel by using Taylor-series expansion and Galerkin method. Applied Mathematics and Computation, 250(??):530–532, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014830.

Park:2018:MMC

[Par18a] Jang Min Park. Mathematical modeling and computational simulation of phase separation in ternary mixtures. Applied Mathematics and Computation, 330(??):11–22, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301000.

Park:2018:BAC

[Par18b] Junpyo Park. Balancedness among competitions for biodiversity in the cyclic structured three species system. Applied Mathematics and Computation, 320(??):425–436, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306756.

Piotrowska:2018:MAG

[PBB18] Monika J. Piotrowska, Agnieszka Bartlomiejczyk, and Marek Bodnar. Mathematical analysis of a generalised p53-Mdm2 protein gene expression model. Applied Mathematics and Computation, 328(??):26– 44, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300274.

Porubov:2018:TSS

[PBBF18] A. V. Porubov, R. S. Bondarenkov, D. Bouche, and A. L. Fradkov. Two-step shock waves propagation for isothermal Euler equations. Applied Mathematics and Computation, 332(??):160–166, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302285.

Petkovic:2015:ANI

[PBMP15] Marko D. Petković, Predrag R. Bakic, Andrew D. A. Maidment, and David Pokrajac. Asymptotic number of Z^{3}Δ cells covering C^{(∞)} surface on uniform grid and complexity of recursivepartitioning simulation of septal tissue regions. Applied Mathematics and Computation, 252(??): 263–272, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314016440.

Pal:2019:NSN

[PBP19] S. K. Pal, S. Bhattacharyya, and I. Pop. A numerical study on nonhomogeneous model for the conjugate-mixed convection of a Cu-water nanofluid in an enclosure with thick wavy wall. Applied Mathematics and Computation, 356(??):219–234, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302012.

Pavaloiu:2016:RAN

[PC16] Ion Pavaloiu and Emil Catinas. On a robust Aitken–Newton method based on the Hermite polynomial. Applied Mathematics and Computation, 287–288(??):224–231, September 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316302375.

Pinto:2017:RST

[PC17] Carla M. A. Pinto and Ana R. M. Carvalho. The role of synaptic transmission in a HIV model with memory. Applied Mathematics and Computation, 292(??):76–95, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304696.

Prokhorenko:2016:COW

[PCA16] Victor Prokhorenko, Kim-Kwang Raymond Choo, and Helen Ashman. Contextoriented web application protection model. Applied Mathematics and Computation, 285(??):59–78, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302260.

Prabha:2017:HDS

[PCS17] T. Prabha, M. Chandru, and V. Shanthi. Hybrid difference scheme for singularly perturbed reaction-convection-diffusion problem with boundary and interior layers. Applied Mathematics and Computation, 314(??):237–256, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304423.

Pang:2015:RMP

[PCW15] Li-Ping Pang, Shuang Chen, and Jin-He Wang. Risk management in portfolio applications of nonconvex stochastic programming. Applied Mathematics and Computation, 258(??):565–575, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002118.

Phung-Duc:2015:AAM

[PD15] Tuan Phung-Duc. Asymptotic analysis for Markovian queues with two types of nonpersistent retrial customers. Applied Mathematics and Computation, 265(??):768–784, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500764X.

Pelanti:2017:LMN

[Pel17] Marica Pelanti. Low Mach number preconditioning techniques for Roe-type and HLLC-type methods for a two-phase compressible flow model. Applied Mathematics and Computation, 310(??):112–133, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302588.

Pourbashash:2018:LRF

[PeO18] Hossein Pourbashash and Mahmood Khaksar e Oshagh. Local RBF–FD technique for solving the two-dimensional modified anomalous sub-diffusion equation. Applied Mathematics and Computation, 339(??):132–143, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305320.

Petkov:2015:DDC

[Pet15a] Boyan H. Petkov. Difficulties in detecting chaos in a complex system. Applied Mathematics and Computation, 260(??):35–47, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003392.

Petrovic:2015:ADS

[Pet15b] Milena J. Petrović. An accelerated double step size model in unconstrained optimization. Applied Mathematics and Computation, 250(??):309–319, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014799.

Porubov:2015:FCS

[PFA15] A. V. Porubov, A. L. Fradkov, and B. R. Andrievsky. Feedback control for some solutions of the sine-Gordon equation. Applied Mathematics and Computation, 269(??):17–22, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009595.

Peng:2015:SNA

[PFH^{+}15] Yehui Peng, Heying Feng, Jinjun Hou, Zhiguang Xiong, and Yingfu Guo. Simulating the nonlinear acoustic oscillations in a resonator by gaskinetic scheme. Applied Mathematics and Computation, 250(??):451–462, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015008.

Pasek:2015:NSI

[PG15a] Jan Pasek and Hansley Pravin Gaya. Numerical simulations of the influence of temperature changes on structural integrity of stone temples in Angkor, Cambodia. Applied Mathematics and Computation, 267(??):409–418, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002143.

Peng:2015:RIA

[PG15b] Xue Peng and Xiao-Xia Guo. Real iterative algorithms for a common solution to the complex conjugate matrix equation system. Applied Mathematics and Computation, 270(??):472–482, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010279.

Pan:2017:EMS

[PGLG17] Quan-Ke Pan, Liang Gao, Xin-Yu Li, and Kai-Zhou Gao. Effective metaheuristics for scheduling a hybrid flowshop with sequence-dependent setup times. Applied Mathematics and Computation, 303(??):89–112, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300127.

Petkovic:2017:SCC

[PH17] I. Petković and – D. Herceg. Symbolic computation and computer graphics as tools for developing and studying new root-finding methods. Applied Mathematics and Computation, 295(??):95–113, February 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305951.

Pan:2019:BSS

[PHM19] Yubin Pan, Jin Huang, and Yanying Ma. Bernstein series solutions of multidimensional linear and nonlinear Volterra integral equations with fractional order weakly singular kernels. Applied Mathematics and Computation, 347(??):149–161, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308889.

Pahamli:2018:IPD

[PHRB18] Y. Pahamli, M. J. Hosseini, A. A. Ranjbar, and R. Bahrampoury. Inner pipe downward movement effect on melting of PCM in a double pipe heat exchanger. Applied Mathematics and Computation, 316(??):30–42, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305271.

Perraudin:2018:DGW

[PHSB18] Nathanaël Perraudin, Nicki Holighaus, Peter L. Søndergaard, and Peter Balazs. Designing Gabor windows using convex optimization. Applied Mathematics and Computation, 330(??):266–287, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300560.

Pi:2015:SSO

[Pi15] Dingheng Pi. On the stability of a second order retarded differential equation. Applied Mathematics and Computation, 256(??):324–333, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000090.

Proinov:2019:CGT

[PIP19] Petko D. Proinov, Stoil I. Ivanov, and Miodrag S. Petković. On the convergence of Gander’s type family of iterative methods for simultaneous approximation of polynomial zeros. Applied Mathematics and Computation, 349(??):168–183, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318311020.

Pishbin:2018:NSS

[Pis18]  S. Pishbin.  The nu-

merical solution of the semi-explicit IDAEs by discontinuous piecewise polynomial approximation. Applied Mathematics and Computation, 339(??):81–92, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305678.

Pirkhedri:2015:STF

[PJ15] A. Pirkhedri and H. H. S.

Javadi. Solving the timefractional diffusion equation via sinc–Haar collocation method. Applied Mathematics and Computation, 257(??):317–326, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017718.

Pan:2018:EBM

[PJ18] Wei Pan and Zhen Jin.

Edge-based modeling of computer virus contagion on a tripartite graph. Applied Mathematics and Computation, 320(??): 282–291, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317306720.

Pasek:2015:PAF

[PK15] Jan Pasek and Petr Kesl.

Probabilistic assessment of failure risk of the building envelope thermally insulated from the inside. Applied Mathematics and Computation, 267(??):108–118, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007043.

Pantokratoras:2018:CPB

[PK18a] Asterios Pantokratoras and Muhammad Kamran. Comment on the paper “Boundary layer flow of Carreau fluid over a convectively heated stretching sheet, T. Hayat, Sadia Asad, M. Mustafa, A. Alsaedi, Applied Mathematics and Computation 246 (2014) 12–22”. Applied Mathematics and Computation, 329(??):227–229, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301206. See

[HAMA14].

Priyadarshi:2018:WGM

[PK18b] Gopal Priyadarshi and B. V. Rathish Kumar. Wavelet Galerkin method for fourth order linear and nonlinear differential equations. Applied Mathematics and Computation, 327(??):8–21, June 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300067.

Pandey:2015:ASN

[PKB15] D. N. Pandey, P. Kumar, and D. Bahuguna. Approximations of solutions for a nonlinear differential equation with a deviating argument. Applied Mathematics and Computation, 261(??):242–251, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003859.

Prakash:2018:NIT

[PKB18] Amit Prakash, Manoj Kumar, and Dumitru Baleanu. A new iterative technique for a fractional model of nonlinear Zakharov–kuznetsov equations via Sumudu transform. Applied Mathematics and Computation, 334(??):30–40, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302789.

Panek:2018:NAM

[PKHD18] David Pánek, Václav Kotlan, Roman Hamar, and Ivo Dolezel. Novel algorithm for modeling combined laser and induction welding respecting keyhole effect. Applied Mathematics and Computation, 319(??):254–263, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317301820.

Prakash:2015:NMS

[PKS15] Amit Prakash, Manoj Kumar, and Kapil K. Sharma. Numerical method for solving fractional coupled Burgers equations. Applied Mathematics and Computation, 260(??):314–320, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003422.

P:2015:PIS

[PL15a] Vijaya Laxmi P. and Soujanya M. L. Perishable inventory system with service interruptions, retrial demands and negative customers. Applied Mathematics and Computation, 262(??):102–110, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004646.

Pasialis:2015:SDM

[PL15b] V. Pasialis and G. Lampeas. Shape descriptors and mapping methods for full-field comparison of experimental to simulation data. Applied Mathematics and Computation, 256(??):203–221, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000119.

Peng:2015:CDA

[PL15c] Yu Peng and Qian Lu.

Complex dynamics analysis for a duopoly Stackelberg game model with bounded rationality. Applied Mathematics and Computation, 271(??): 259–268, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315012448.

Porta:2015:SSA

[PL15d] Federica Porta and Ignace Loris. On some steplength approaches for proximal algorithms. Applied Mathematics and Computation, 253(??):345–362, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017330.

Peng:2018:PFR

[PL18] Yahong Peng and Heyang Ling. Pattern formation in a ratio-dependent predator–prey model with cross-diffusion. Applied Mathematics and Computation, 331(??):307–318, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302066.

Platkowski:2015:ABF

[Pla15] Tadeusz Platkowski. Aspirationbased full cooperation in finite systems of players. Applied Mathematics and Computation, 251(??):46–54, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015860.

Platkowski:2016:ESM

[Pla16] Tadeusz Platkowski. Egalitarian solutions to multiperson social dilemmas in populations. Applied Mathematics and Computation, 284(??):226–233, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630203X.

Platkowski:2017:GFM

[Pla17] Tadeusz Platkowski. Greed and fear in multiperson social dilemmas. Applied Mathematics and Computation, 308(??):157–160, September 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317302096.

Postnikov:2016:CII

[PLL16] Eugene B. Postnikov, Elena A. Lebedeva, and Anastasia I. Lavrova. Computational implementation of the inverse continuous wavelet transform without a requirement of the admissibility condition. Applied Mathematics and Computation, 282(??):128–136, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301163.

Panda:2015:MAN

[PMC15] Srikumar Panda, S. C. Martha, and A. Chakrabarti. A modified approach to numerical solution of Fredholm integral equations of the second kind. Applied Mathematics and Computation, 271(??):102–112, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315011674.

Panda:2016:AAS

[PMC16] Srikumar Panda, S. C. Martha, and A. Chakrabarti. An alternative approach to study nonlinear inviscid flow over arbitrary bottom topography. Applied Mathematics and Computation, 273(??):165–177, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013247.

Pinto:2016:HDC

[PMS16] Martin Campos Pinto, Marie Mounier, and Eric Sonnendrücker. Handling the divergence constraints in Maxwell and Vlasov–Maxwell simulations. Applied Mathematics and Computation, 272 (part 2)(??):403–419, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010115.

Panda:2015:PRP

[PMSB15] S. Panda, N. M. Modak, S. S. Sana, and M. Basu. Pricing and replenishment policies in dualchannel supply chain under continuous unit cost decrease. Applied Mathematics and Computation, 256(??):913–929, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001137.

Pal:2016:SDE

[PMV16] Dulal Pal, Gopinath Mandal, and Kuppalapalle Vajravalu. Soret and Dufour effects on MHD convective-radiative heat and mass transfer of nanofluids over a vertical non-linear stretching/shrinking sheet. Applied Mathematics and Computation, 287–288 (??):184–200, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316302879.

Paszek:2015:PPP

[PN15a] Stefan Paszek and Adrian Nocoń. Parameter polyoptimization of PSS2A power system stabilizers operating in a multi-machine power system including the uncertainty of model parameters. Applied Mathematics and Computation, 267(??):750–757, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016671.

Proinov:2015:APC

[PN15b] Petko D. Proinov and Ivanka A. Nikolova. Approximation of point of coincidence and common fixed points of quasicontraction mappings using the Jungck iteration scheme. Applied Mathematics and Computation, 264(??):359–365, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005603.

Petik:2015:SSC

[PÖB15] Tuğba Petik, Halim Özdemir, and Julio Benítez. On the spectra of some combinations of two generalized quadratic matrices. Applied Mathematics and Computation, 268(??):978–990, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008747.

Polyanin:2019:FSS

[Pol19] Andrei D. Polyanin. Functional separable solutions of nonlinear reactiondiffusion equations with variable coefficients. Applied Mathematics and Computation, 347(??): 282–292, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031830972X.

Ponalagusamy:2017:CBA

[Pon17] R. Ponalagusamy. Corrigendum to “A biomechanical approach to study the effect of body acceleration and slip velocity through stenotic artery”

[applied mathematics and computation, 261 (2015) 148–155]. Applied Mathematics and Computation, 301(??):115–116, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307639. See

[SSG15b].

Ponalagusamy:2018:MAF

[Pon18] R. Ponalagusamy. Mathematical analysis of flow of non-Newtonian fluid due to metachronal beating of cilia in a tube and its physiological applications. Applied Mathematics and Computation, 337(??):535–544, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830465X.

Popova:2017:POE

[Pop17] Evgenija D. Popova. Parameterized outer estimation of AE-solution sets to parametric interval linear systems. Applied Mathematics and Computation, 311(??):353–360, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303430.

Portillo:2018:HOF

[Por18] A. M. Portillo. Highorder full discretization for anisotropic wave equations. Applied Mathematics and Computation, 323(??):1–16, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308275.

Pospisil:2017:RSD

[Pos17] Michal Pospísil. Representation of solutions of delayed difference equations with linear parts given by pairwise permutable matrices via Ztransform. Applied Mathematics and Computation, 294(??):180–194, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305793.

Potzsche:2019:NDI

[Pöt19] Christian Pötzsche. Numerical dynamics of integrodifference equations: Basics and discretization errors in a C^{0}- setting. Applied Mathematics and Computation, 354(??):422–443, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301377.

Povstenko:2015:TFT

[Pov15] Yuriy Povstenko. Timefractional thermoelasticity problem for a sphere subjected to the heat flux. Applied Mathematics and Computation, 257(??):327–334, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017275.

Perdomo:2012:NPD

[PP12] Francisco Perdomo and Ángel Plaza. A new proof of the degeneracy property of the longestedge n-section refinement scheme for triangular meshes. Applied Mathematics and Computation, 219(4):2342–2344, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300312008223. See

corrigendum [PP15b].

Pantelous:2015:OPP

[PP15a] Athanasios A. Pantelous and Eudokia Passalidou. Optimal premium pricing strategies for competitive general insurance markets. Applied Mathematics and Computation, 259(??):858–874, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500332X.

Perdomo:2015:CNP

[PP15b] Francisco Perdomo and Ángel Plaza. Corrigendum to “A new proof of the degeneracy property of the longest-edge n-section refinement scheme for triangular meshes” [Applied Mathematics and Computation 219 (4) (2012) 2342–2344]. Applied Mathematics and Computation, 260(??):412–413, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003975. See

[PP12].

Ponalagusamy:2018:PMF

[PP18a] R. Ponalagusamy and S. Priyadharshini. Pulsatile MHD flow of a Casson fluid through a porous bifurcated arterial stenosis under periodic body acceleration. Applied Mathematics and Computation, 333(??):325–343, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302844.

Pruski:2018:CPS

[PP18b] Piotr Pruski and Stefan Paszek. Calculations of power system electromechanical eigenvalues based on analysis of instantaneous power waveforms at different disturbances. Applied Mathematics and Computation, 319(??):104–114, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300759.

Pola:2019:OCH

[PP19] Fernanda Paula Barnosa Pola and Ives Renê Venturini Pola. Optimizing computational highorder schemes in finite volume simulations using unstructured mesh and topological data structures. Applied Mathematics and Computation, 342(??):1–17, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830612X.

Padma:2025:CMM

[PPS25] R. Padma, R. Ponalagusamy, and R. Tamil Selvi. Corrigendum to “Mathematical modeling of electro hydrodynamic non-Newtonian fluid flow through tapered arterial stenosis with periodic body acceleration and applied magnetic field” [Applied Mathematics and Computation, 362 (2019) 124453]. Applied Mathematics and Computation, 500(??):??, September 1, 2025. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300325001456. See

[RRR19, RRR20].

Pecaric:2015:BCF

[PPV15] J. Pecarić, M. Ribicić Penava, and A. Vukelić. Bounds for the Chebyshev functional and applications to the weighted integral formulae. Applied Mathematics and Computation, 268(??):957–965, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009157.

Papp:2015:FTM

[PR15a] Zoltan Papp and Sanja Rapajić. FR type methods for systems of largescale nonlinear monotone equations. Applied Mathematics and Computation, 269(??):816–823, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010577.

Pena:2015:CFG

[PR15b] A. Peña and M. L. Rezola. Connection formulas for general discrete Sobolev polynomials: Mehler–Heine asymptotics. Applied Mathematics and Computation, 261(??):216–230, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003987.

Parastesh:2019:BDS

[PRA^{+}19] Fatemeh Parastesh, Karthikeyan Rajagopal, Fawaz E. Alsaadi, Tasawar Hayat, V.- T. Pham, and Iqtadar Hussain. Birth and death of spiral waves in a network of Hindmarsh–Rose neurons with exponential magnetic flux and excitable media. Applied Mathematics and Computation, 354(??):377–384, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301456.

Pratap:2019:SPS

[PRC^{+}19] A. Pratap, R. Raja, J. Cao, C. P. Lim, and O. Bagdasar. Stability and pinning synchronization analysis of fractional order delayed Cohen– Grossberg neural networks with discontinuous activations. Applied Mathematics and Computation, 359(??):241–260, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303509.

Pogorui:2019:GKT

[PRD19] Anatoliy A. Pogorui and Ramón M. Rodríguez-Dagnino. Goldstein–Kac telegraph equations and random flights in higher dimensions. Applied Mathematics and Computation, 361(??):617–629, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304473.

Pang:2015:TDO

[PRL^{+}15] Liuyong Pang, Shigui Ruan, Sanhong Liu, Zhong Zhao, and Xinan Zhang. Transmission dynamics and optimal control of measles epidemics. Applied Mathematics and Computation, 256(??):131–147, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017573.

Prakash:2019:DAA

[PRMC19] M. Prakash, R. Rakkiyappan, A. Manivannan, and Jinde Cao. Dynamical analysis of antigen-driven T-cell infection model with multiple delays. Applied Mathematics and Computation, 354(??):266–281, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301547.

Proinov:2016:GSC

[Pro16] Petko D. Proinov. A general semilocal convergence theorem for simultaneous methods for polynomial zeros and its applications to Ehrlich’s and Dochev– Byrnev’s methods. Applied Mathematics and Computation, 284(??):102–114, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301746.

Przybylowicz:2015:OGA

[Prz15] Pawel Przybylowicz. Optimal global approximation of SDEs with timeirregular coefficients in asymptotic setting. Applied Mathematics and Computation, 270(??): 441–457, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315011108.

Petrovic:2015:RCI

[PS15a] Milos Z. Petrović and Predrag S. Stanimirović. Representations and computations of {2, 3˜} and {2, 4˜}- inverses in indefinite inner product spaces. Applied Mathematics and Computation, 254(??):157–171, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017615.

Pinasco:2015:NIP

[PS15b] Juan P. Pinasco and Cristian Scarola. A nodal inverse problem for second order Sturm–Liouville operators with indefinite weights. Applied Mathematics and Computation, 256(??):819–830, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001332.

Ploskas:2015:EGB

[PS15c] Nikolaos Ploskas and Nikolaos Samaras. Efficient GPU-based implementations of simplex type algorithms. Applied Mathematics and Computation, 250(??):552–570, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014714.

Ponnaian:2016:RSP

[PS16] Devaraj Ponnaian and Yugesh Shanmugam. Reconstruction of L-splines of polynomial growth from their local weighted average samples. Applied Mathematics and Computation, 273(??):1018–1024, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013946.

Padgett:2018:NSD

[PS18a] Joshua L. Padgett and Qin Sheng. Numerical solution of degenerate stochastic Kawarada equations via a semi-discretized approach. Applied Mathematics and Computation, 325(??):210–226, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308937.

Polyanin:2018:NPB

[PS18b] Andrei D. Polyanin and Inna K. Shingareva. Nonlinear problems with blowup solutions: Numerical integration based on differential and nonlocal transformations, and differential constraints. Applied Mathematics and Computation, 336(??):107–137, November 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318303977.

Pourbabaee:2019:NLO

[PS19] Marzieh Pourbabaee and Abbas Saadatmandi. A novel Legendre operational matrix for distributed order fractional differential equations. Applied Mathematics and Computation, 361(??):215–231, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304230.

Panahi:2019:NCN

[PSJJ19] Shirin Panahi, Touraj Shirzadian, Mahdi Jalili, and Sajad Jafari. A new chaotic network model for epilepsy. Applied Mathematics and Computation, 346(??):395–407, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309275.

Pal:2017:SHT

[PSM17] D. Pal, G. P. Samanta, and G. S. Mahapatra. Selective harvesting of two competing fish species in the presence of toxicity with time delay. Applied Mathematics and Computation, 313(??):74–93, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303818.

Postnikov:2017:FMS

[PSP17] Eugene B. Postnikov, Elena S. Stiukhina, and Dmitry E. Postnov. A fast memory-saving method for the Morlet wavelet-based transform and its application to in vivo assessment of microcirculation dynamics. Applied Mathematics and Computation, 305(??):251–261, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301042.

Pernas-Salomon:2015:GFG

[PSPÁV15]  René  Pernas-Salomón,

R. Pérez-Álvarez, and V. R. Velasco. General form of the Green’s function regular at infinity for the homogeneous Sturm–Liouville matrix operator. Applied Mathematics and Computation, 269(??):824–833, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010565.

Petrosyan:2017:CST

[PSSX17] Leon Petrosyan, Artem Sedakov, Hao Sun, and Genjiu Xu. Convergence of strong time-consistent payment schemes in dynamic games. Applied Mathematics and Computation, 315(??):96–112, December 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317304915.

Pereira:2018:RES

[PST18] Enrique Pereira, Erwin Suazo, and Jessica Trespalacios. Riccati–Ermakov systems and explicit solutions for variable coefficient reaction-diffusion equations. Applied Mathematics and Computation, 329(??):278–296, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300705.

Peng:2017:EEF

[PSY^{+}17] Jie Peng, Shi Shu, HaiYuan Yu, Chunsheng Feng, Mingxian Kan, and Ganghua Wang. Error estimates on a finite volume method for diffusion problems with interface on rectangular grids. Applied Mathematics and Computation, 311(??):335–352, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303272.

Pan:2018:ECG

[PSZ18] V. Y. Pan, F. Soleymani, and L. Zhao. An efficient computation of generalized inverse of a matrix. Applied Mathematics and Computation, 316(??):89–101, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305611.

Pedro:2016:USA

[PTA16] S. A. Pedro, H. E. Z. Tonnang, and S. Abelman. Uncertainty and sensitivity analysis of a Rift Valley fever model. Applied Mathematics and Computation, 279(??):170–186, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300030.

Phuong:2019:CPN

[PTBN19] Nguyen Duc Phuong, Nguyen Huy Tuan, Dumitru Baleanu, and Tran Bao Ngoc. On Cauchy problem for nonlinear fractional differential equation with random discrete data. Applied Mathematics and Computation, 362 (??):Article 124458, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304229.

Palubeckis:2019:HSA

[PTO19] Gintaras Palubeckis, Arunas Tomkevicius, and Armantas Ostreika. Hybridizing simulated annealing with variable neighborhood search for bipartite graph crossing minimization. Applied Mathematics and Computation, 348(??):84–101, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310257.

Pu:2015:DFH

[Pu15] Xueke Pu. Dynamics of a fractional hydrodynamical equation for the Heisenberg paramagnet. Applied Mathematics and Computation, 257(??):213–229, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314010662.

Proinov:2016:FWT

[PV16] Petko D. Proinov and Maria T. Vasileva. On a family of Weierstrasstype root-finding methods with accelerated convergence. Applied Mathematics and Computation, 273(??):957–968, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013995.

Proinov:2019:CHO

[PV19] Petko D. Proinov and Maria T. Vasileva. On the convergence of highorder Gargantini–Farmer– Loizou type iterative methods for simultaneous approximation of polynomial zeros. Applied Mathematics and Computation, 361(??):202–214, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304199.

Pleshkevich:2019:SOA

[PVL19] Alexander Pleshkevich, Dmitriy Vishnevskiy, and Vadim Lisitsa. Sixthorder accurate pseudospectral method for solving one-way wave equation. Applied Mathematics and Computation, 359(??):34–51, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930311X.

Pezzi:2015:PPW

[PVV^{+}15] Guilherme Peretti Pezzi, Evelyne Vaissié, Yann Viala, Denis Caromel, and Philippe Gourbesville. Parallel profiling of water distribution networks using the Clément formula. Applied Mathematics and Computation, 267(??):83–95, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315007080.

Pei:2019:DHB

[PW19a] Lijun Pei and Shuo Wang.

Double Hopf bifurcation of differential equation with linearly state-dependent delays via MMS. Applied Mathematics and Computation, 341(??):256–276, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300318307549.

Pei:2019:DPS

[PW19b] Lijun Pei and Shuo Wang.

Dynamics and the periodic solutions of the delayed non-smooth Internet TCP-RED congestion control system via HB-AFT. Applied Mathematics and Computation, 361(??):689–702, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304977.

Pan:2018:REA

[PWFY18] Donghui Pan, Yantao Wei, Houzhang Fang, and Wenzhi Yang. A reliability estimation approach via Wiener degradation model with measurement errors. Applied Mathematics and Computation, 320(??):131–141, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306410.

Pei:2017:SIP

[PWS17] Ke Pei, Guotao Wang, and Yanyan Sun. Successive iterations and positive extremal solutions for a Hadamard type fractional integro-differential equations on infinite domain. Applied Mathematics and Computation, 312(??):158–168, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303685.

Pan:2019:EBO

[PY19a] Yingnan Pan and Guang-Hong Yang. Event-based output tracking control for fuzzy networked control systems with networkinduced delays. Applied Mathematics and Computation, 346(??):513–530, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309044.

Pan:2019:EBR

[PY19b] Yingnan Pan and Guang-Hong Yang. Event-based reduced-order fuzzy filtering for networked control systems with timevarying delays. Applied Mathematics and Computation, 359(??):71–83, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303492.

Pei:2018:AMS

[PYZS18] Yongzhen Pei, Hongfu Yang, Qimin Zhang, and Fangfang Shen. Asymptotic mean-square boundedness of the numerical solutions of stochastic age-dependent population equations with Poisson jumps. Applied Mathematics and Computation, 320(??):524–534, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307294.

Peng:2015:BIP

[PZ15] Li Peng and Yong Zhou.

Bifurcation from interval and positive solutions of the three-point boundary value problem for fractional differential equations. Applied Mathematics and Computation, 257(??):458–466, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016245.

Pei:2016:LCT

[PZ16a] Yonggang Pei and Detong Zhu. Local convergence of a trust-region algorithm with line search filter technique for nonlinear constrained optimization. Applied Mathematics and Computation, 273(??):797–808, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014113.

Peng:2016:TIP

[PZ16b] Yahong Peng and Tonghua Zhang. Turing instability and pattern induced by cross-diffusion in a predator–prey system with Allee effect. Applied Mathematics and Computation, 275(??):1–12, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015696.

Pan:2017:CHA

[PZ17] Xintian Pan and Luming Zhang. On the convergence of a high-accuracy conservative scheme for the Zakharov equations. Applied Mathematics and Computation, 297(??):79– 91, March 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316306403.

Portal:2019:MNG

[PZ19] Alberto Portal and Pedro J. Zufiria. On the minimum number of general or dedicated controllers required for system controllability. Applied Mathematics and Computation, 355(??):417–427, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302085.

Pei:2017:APA

[PZJ17] Xin Pei, Xiu-Xiu Zhan, and Zhen Jin. Application of pair approximation method to modeling and analysis of a marriage network. Applied Mathematics and Computation, 294(??):280–293, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305707.

Pang:2015:MMA

[PZLZ15] Liuyong Pang, Zhong Zhao, Sanhong Liu, and Xinan Zhang. A mathematical model approach for tobacco control in China. Applied Mathematics and Computation, 259(??):497–509, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002787.

Quiceno:2019:ASD

[QA19] H. R. Quiceno and C. Arias. Analysis of the stability and dispersion for a Riemannian acoustic wave equation. Applied Mathematics and Computation, 341(??):288–300, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300318307616.

Qi:2019:EIS

[QBC19] Xingnan Qi, Haibo Bao, and Jinde Cao. Exponential input-to-state stability of quaternion-valued neural networks with time delay. Applied Mathematics and Computation, 358(??):382–393, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303339.

Qadeer:2019:UKP

[QBNF19] Neelam Qadeer, Nayab Bhatti, Qaisar Abbas Naqvi, and Muhammad Arshad Fiaz. Use of Kobayashi potential method and Lorentz–Drude model to study scattering from a PEC strip buried below a lossy dispersive NID dielectric-magnetic slab. Applied Mathematics and Computation, 362 (??):Article 124573, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305569.

Qin:2019:CDC

[QCH19] Xiao-Wen Qin, Jou-Ming Chang, and Rong-Xia Hao. Constructing dual-CISTs of DCell data center networks. Applied Mathematics and Computation, 362(??): Article 124546, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305296.

Qiu:2018:OHS

[QD18] Hong Qiu and Wenmin Deng. Optimal harvesting of a stochastic delay competitive Lotka– Volterra model with Lévy jumps. Applied Mathematics and Computation, 317(??):210–222, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305957.

Qin:2015:TBM

[QDD15] Wendi Qin, Deqiong Ding, and Xiaohua Ding. Two boundedness and monotonicity preserving methods for a generalized Fisher-KPP equation. Applied Mathematics and Computation, 252(??): 552–567, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S009630031401697X.

Qin:2018:EFG

[QDH^{+}18] Xinqiang Qin, Xianbao Duan, Gang Hu, Lijun Su, and Xing Wang. An element-free Galerkin method for solving the two-dimensional hyperbolic problem. Applied Mathematics and Computation, 321(??):106–120, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307397.

Qian:2017:CMC

[QFS17] Xu Qian, Hao Fu, and Songhe Song. Conservative modified Crank–Nicolson and time-splitting wavelet methods for modeling Bose–Einstein condensates in delta potentials. Applied Mathematics and Computation, 307(??):1–16, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301480.

Quiroga:2015:AMT

[QFTT15] A. A. I. Quiroga, D. Fernández, G. A. Torres, and C. V. Turner. Adjoint method for a tumor invasion PDEconstrained optimization problem in 2D using adaptive finite element method. Applied Mathematics and Computation, 270(??):358–368, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010930.

Qi:2016:ODS

[QG16] Wenhai Qi and Xianwen Gao. H_{∞} observer design for stochastic time-delayed systems with Markovian switching under partly known transition rates and actuator saturation. Applied Mathematics and Computation, 289(??):80–97, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303095.

Qiu:2019:SSL

[QG19] Huanhuan Qiu and Shangjiang Guo. Steady-states of a Leslie–Gower model with diffusion and advection. Applied Mathematics and Computation, 346(??):695–709, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308622.

Qin:2019:CGI

[QGH19] Chengfu Qin, Litao Guo, and Lexian Huang. Connectivity of the graph induced by contractible edges of a k-tree. Applied Mathematics and Computation, 353(??):1–6, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300669.

Qi:2015:DTF

[Qi15] Feng Qi. Derivatives of tangent function and tangent numbers. Applied Mathematics and Computation, 268(??):844–858, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009078.

Qian:2016:FNR

[Qia16a] Chenyin Qian. A further note on the regularity criterion for the 3D nematic liquid crystal flows. Applied Mathematics and Computation, 290(??):258–266, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303824.

Qian:2016:RRC

[Qia16b] Chenyin Qian. Remarks on the regularity criterion for the nematic liquid crystal flows in R^{3}. Applied Mathematics and Computation, 274(??):679–689, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315014691.

Qiu:2018:TGS

[Qiu18]  Hailong Qiu.  Two-

grid stabilized methods for the stationary incompressible Navier–Stokes equations with nonlinear slip boundary conditions. Applied Mathematics and Computation, 332(??):172–188, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830239X.

Qi:2018:CDT

[QKGW18] Wenhai Qi, Yonggui Kao, Xianwen Gao, and Yunliang Wei. Controller design for time-delay system with stochastic disturbance and actuator saturation via a new criterion. Applied Mathematics and Computation, 320(??):535–546, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306653.

Qiao:2015:DTD

[QLF15] Meihong Qiao, Anping Liu, and Urszula Foryś. The dynamics of a time delayed epidemic model on a population with birth pulse. Applied Mathematics and Computation, 252(??):166–174, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016762.

Qian:2018:HOW

[QLLY18] Shouguo Qian, Yu Liu, Gang Li, and Li Yuan. High order well-balanced discontinuous Galerkin methods for Euler equations at isentropic equilibrium state under gravitational fields. Applied Mathematics and Computation, 329(??):23–37, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300882.

Qu:2015:CSS

[QLZ15] Biao Qu, Binghua Liu, and Na Zheng. On the computation of the stepsize for the CQ-like algorithms for the split feasibility problem. Applied Mathematics and Computation, 262(??):218–223, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005147.

Qi:2015:SBA

[QM15a] Feng Qi and Cristinel Mortici. Some best approximation formulas and inequalities for the Wallis ratio. Applied Mathematics and Computation, 253(??):363–368, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016932.

Qi:2015:SIT

[QM15b] Feng Qi and Cristinel Mortici. Some inequalities for the trigamma function in terms of the digamma function. Applied Mathematics and Computation, 271(??):502–511, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012758.

Qin:2016:EIR

[QTF^{+}16] Xiaolin Qin, Juan Tang, Yong Feng, Bernhard Bachmann, and Peter Fritzson. Efficient index reduction algorithm for large scale systems of differential algebraic equations. Applied Mathematics and Computation, 277(??):10–22, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315016100.

Qin:2019:TCS

[QTTL19] Wenjie Qin, Xuewen Tan, Marco Tosato, and Xinzhi Liu. Threshold control strategy for a nonsmooth Filippov ecosystem with group defense. Applied Mathematics and Computation, 362(??): Article 124532, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305119.

Qin:2016:ELM

[QTXY16] Wenjie Qin, Sanyi Tang, Changcheng Xiang, and Yali Yang. Effects of limited medical resource on a Filippov infectious disease model induced by selection pressure. Applied Mathematics and Computation, 283(??):339–354, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301539.

Qin:2019:SEA

[QW19] Hongyu Qin and Fengyan Wu. Several effective algorithms for nonlinear time fractional models. Applied Mathematics and Computation, 363 (??):Article 124598, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305909.

Qi:2017:LMN

[QX17] Guohua Qi and Jiang Xu. The low Mach number limit for the compressible flow of liquid crystals. Applied Mathematics and Computation, 297(??):39–49, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306245.

Qiao:2019:ADS

[QXW19] Leijie Qiao, Da Xu, and Zhibo Wang. An ADI difference scheme based on fractional trapezoidal rule for fractional integrodifferential equation with a weakly singular kernel. Applied Mathematics and Computation, 354(??):103–114, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301201.

Qu:2015:BOG

[QY15] Hui Qu and Guihai Yu. Bicyclic oriented graphs with skew-rank 2 or 4. Applied Mathematics and Computation, 258(??):182–191, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315001678.

Qiao:2017:EDP

[QY17] Yalin Qiao and Weihua Yang. Edge disjoint paths in hypercubes and folded hypercubes with conditional faults. Applied Mathematics and Computation, 294(??):96–101, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305628.

Qin:2018:IRD

[QYF^{+}18] Xiaolin Qin, Lu Yang, Yong Feng, Bernhard Bachmann, and Peter Fritzson. Index reduction of differential algebraic equations by differential Dixon resultant. Applied Mathematics and Computation, 328(??):189–202, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308883.

Qi:2015:EEF

[QZ15a] Feng Qi and Miao-Miao Zheng. Explicit expressions for a family of the Bell polynomials and applications. Applied Mathematics and Computation, 258(??):597–607, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001964.

Qin:2015:SSO

[QZ15b] Tingting Qin and Chengjian Zhang. Stable solutions of one-leg methods for a class of nonlinear functionalintegro-differential equations. Applied Mathematics and Computation, 250(??):47–57, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015082.

Qi:2019:RFT

[QZCJ19] Wenhai Qi, Guangdeng Zong, Jun Cheng, and Ticao Jiao. Robust finite-time stabilization for positive delayed semi-Markovian switching systems. Applied Mathematics and Computation, 351(??):139–152, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311305.

Qi:2016:KIM

[QZD16] Xuli Qi, Bo Zhou, and Zhibin Du. The Kirchhoff indices and the matching numbers of unicyclic graphs. Applied Mathematics and Computation, 289(??):464–480, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303010.

Qin:2015:EAE

[QZJY15] Yang Qin, Xiaoxiong Zhong, Hao Jiang, and Yibin Ye. An environment aware epidemic spreading model and immune strategy in complex networks. Applied Mathematics and Computation, 261(??):206–215, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004245.

Qiao:2017:RKD

[QZL^{+}17] Dian-Liang Qiao, Peng Zhang, Zhi-Yang Lin, S. C. Wong, and Keechoo Choi. A Runge–Kutta discontinuous Galerkin scheme for hyperbolic conservation laws with discontinuous fluxes. Applied Mathematics and Computation, 292(??):309–319, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304684.

Qiao:2018:CTI

[QZL18a] Lu Qiao, Shenggui Zhang, and Jing Li. Coulsontype integral formulas for the general energy of polynomials with real roots. Applied Mathematics and Computation, 320(??):202–212, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306458.

Que:2018:SSE

[QZL18b] Wei Que, Yabin Zhang, and Shaobo Liu. The spatial spillover effect of fiscal decentralization on local public provision: Mathematical application and empirical estimation. Applied Mathematics and Computation, 331(??):416–429, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318302054.

Qin:2019:CGF

[QZW19] Hongyu Qin, Qifeng Zhang, and Shaohua Wan. The continuous Galerkin finite element methods for linear neutral delay differential equations. Applied Mathematics and Computation, 346(??):76–85, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309226.

Ren:2015:CKW

[RA15] Hongmin Ren and Ioannis K. Argyros. On the convergence of King– Werner-type methods of order 1 + \sqrt{2} free of derivatives. Applied Mathematics and Computation, 256(??):148–159, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000429.

Rad:2015:TEA

[Rad15] A. Behravan Rad. Thermoelastic analysis of functionally graded circular plates resting on a gradient hybrid foundation. Applied Mathematics and Computation, 256(??):276–298, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000405.

Rada:2017:VDB

[Rad17]  Juan Rada.  Vertex-

degree-based topological indices of hexagonal systems with equal number of edges. Applied Mathematics and Computation, 296(??):270–276, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306130.

Rabbani:2019:SNQ

[RAH19a] Mohsen Rabbani, Reza Arab, and Bipan Hazarika. Solvability of nonlinear quadratic integral equation by using simulation type condensing operator and measure of noncompactness. Applied Mathematics and Computation, 349(??):102–117, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310853.

Rahman:2019:DMO

[Rah19b] Sharif Rahman. Dimensionwise multivariate orthogonal polynomials in general probability spaces. Applied Mathematics and Computation, 362(??): Article 124538, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031930517X.

Rasa:2015:EHF

[Ras15]  Ioan Rasa.  Entropies

and Heun functions associated with positive linear operators. Applied Mathematics and Computation, 268(??):422–431, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008668.

Razi:2015:GAN

[RAV15] M. Razi, P. J. Attar, and P. Vedula. Grid adaptation and noniterative defect correction for improved accuracy of numerical solutions of PDEs. Applied Mathematics and Computation, 269(??):473–487, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010255.

Razbani:2015:GRB

[Raz15] M. A. Razbani. Global root bracketing method with adaptive mesh refinement. Applied Mathematics and Computation, 268(??):628–635, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009029.

Rostamy:2015:NEM

[RB15] Davoud Rostamy and Parisa Bakhtiari. New efficient multipoint iterative methods for solving nonlinear systems. Applied Mathematics and Computation, 266(??):350–356, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315007110.

Rafiei:2019:BVL

[RBB19] Amin Rafiei, Matthias Bollhöfer, and Fayssal Benkhaldoun. A block version of left-looking AINV preconditioner with one by one or two by two block pivots. Applied Mathematics and Computation, 350(??):366–385, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300207.

Ren:2016:OCC

[RC16] Han Ren and Fugang Chao. Odd components of co-trees and graph embeddings. Applied Mathematics and Computation, 286(??):228–231, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302776.

Ray:2017:ESF

[RC17a] Deep Ray and Praveen Chandrashekar. An entropy stable finite volume scheme for the two dimensional Navier–Stokes equations on triangular grids. Applied Mathematics and Computation, 314(??):257–286, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730471X.

Russo:2017:ASM

[RC17b] Matthew Russo and S. Roy Choudhury. Analytic solutions of a microstructure PDE and the KdV and Kadomtsev–Petviashvili equations by invariant Painlevé analysis and generalized Hirota techniques. Applied Mathematics and Computation, 311(??):228–239, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300735.

Rohan:2015:MRP

[RCM15] E. Rohan, R. Cimrman, and B. Miara. Modelling response of phononic Reissner–Mindlin plates using a spectral decomposition. Applied Mathematics and Computation, 258(??):617–630, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016919.

Roncero-Clemente:2018:MCC

[RCRL18] Carlos Roncero-Clemente and Eugenio Roanes-Lozano. A multi-criteria computer package for power transformer fault detection and diagnosis. Applied Mathematics and Computation, 319(??):153–164, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301285.

Ran:2018:PDB

[RCTZ18] Yingli Ran, Zhihao Chen, Shaojie Tang, and Zhao Zhang. Primal dual based algorithm for degreebalanced spanning tree problem. Applied Mathematics and Computation, 316(??):167–173, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305672.

Ren:2018:NCM

[RCW^{+}18] Yizhi Ren, Xiangyu Chen, Zhen Wang, Benyun Shi, Guanghai Cui, Ting Wu, and Kim-Kwang Raymond Choo. Neighborconsidered migration facilitates cooperation in prisoner’s dilemma games. Applied Mathematics and Computation, 323(??):95– 105, April 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317308251.

Rakic:2015:SSC

[RD15] Dragan S. Rakić and Dragan S. Djordjević. Star, sharp, core and dual core partial order in rings with involution. Applied Mathematics and Computation, 259(??):800–818, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002623.

Re:2018:AGA

[RDG18] B. Re, C. Dobrzynski, and A. Guardone. Assessment of grid adaptation criteria for steady, twodimensional, inviscid flows in non-ideal compressible fluids. Applied Mathematics and Computation, 319(??):337–354, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302333.

Radenovic:2016:NSR

[RDLG16] Stojan Radenović, Tatjana Dosenović, Tatjana Aleksić Lampert, and Zorana Goluboví’c. A note on some recent fixed point results for cyclic contractions in b-metric spaces and an application to integral equations. Applied Mathematics and Computation, 273(??):155–164, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013272.

Roy:2019:WWS

[RDM19] R. Roy, Soumen De, and B. N. Mandal. Water wave scattering by multiple thin vertical barriers. Applied Mathematics and Computation, 355(??):458–481, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301973.

Reddy:2019:CPC

[Red19] G. D. Reddy. A class of parameter choice rules for stationary iterated weighted Tikhonov regularization scheme. Applied Mathematics and Computation, 347(??):464–476, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309883.

Rehak:2017:AFS

[Reh17] Pavel Rehák. Asymptotic formulae for solutions of half-linear differential equations. Applied Mathematics and Computation, 292(??):165–177, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304581.

Rehak:2018:KIT

[Reh18] Pavel Rehák. The Karamata integration theorem on time scales and its applications in dynamic and difference equations. Applied Mathematics and Computation, 338(??):466–486, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305113.

Remizov:2018:ASC

[Rem18] Ivan D. Remizov. Approximations to the solution of Cauchy problem for a linear evolution equation via the space shift operator (second-order equation example). Applied Mathematics and Computation, 328(??):243–246, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300808.

Reti:2019:SPG

[Rét19] Tamás Réti. On some properties of graph irregularity indices with a particular regard to the σ-index. Applied Mathematics and Computation, 344–345(??):107–115, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308701.

Reutskiy:2015:NCM

[Reu15]  S. Yu. Reutskiy.  A

new collocation method for approximate solution of the pantograph functional differential equations with proportional delay. Applied Mathematics and Computation, 266(??):642–655, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007663.

Rezvani:2015:RGE

[Rez15]  S. Rezvani.  Ranking

generalized exponential trapezoidal fuzzy numbers based on variance. Applied Mathematics and Computation, 262(??):191–198, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004889.

Rabin:2018:MSK

[RF18] N. Rabin and D. Fishelov.

Multi-scale kernels for Nyström based extension schemes. Applied Mathematics and Computation, 319(??):165–177, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301297.

Ray:2015:NIT

[RG15] S. Saha Ray and A. K.

Gupta. A numerical investigation of time-fractional modified Fornberg–Whitham equation for analyzing the behavior of water waves. Applied Mathematics and Computation, 266(??):135–148, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006621.

Roy:2018:SRC

[RG18a] Falguni Roy and Dharmendra. K. Gupta. Sufficient regularity conditions for complex interval matrices and approximations of eigenvalues sets. Applied Mathematics and Computation, 317(??):193–209, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306148.

Ruttgers:2018:MSP

[RG18b] A. Rüttgers and M. Griebel. Multiscale simulation of polymeric fluids using the sparse grid combination technique. Applied Mathematics and Computation, 319(??):425–443, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302771.

Roul:2019:BSC

[RG19] Pradip Roul and V. M. K. Prasad Goura. Bspline collocation methods and their convergence for a class of nonlinear derivative dependent singular boundary value problems. Applied Mathematics and Computation, 341(??):428–450, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300318307835.

Roul:2019:CFD

[RGA19] Pradip Roul, V. M.

K. Prasad Goura, and Ravi Agarwal. A compact finite difference method for a general class of nonlinear singular boundary value problems with Neumann and Robin boundary conditions. Applied Mathematics and Computation, 350(??):283–304, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300013.

Renqian:2015:TDM

[RGH^{+}15] Suonan Renqian, Yunpeng Ge, Bofeng Huo, Shengjin Ji, and Qiangqiang Diao. On the tree with diameter 4 and maximal energy. Applied Mathematics and Computation, 268(??):364–374, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315008413.

Rongjun:2018:NAN

[RHJ18] Cheng Rongjun, Ge Hongxia, and Wang Jufeng. The nonlinear analysis for a new continuum model considering anticipation and traffic jerk effect. Applied Mathematics and Computation, 332(??):493–505, September 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318302583.

Rhoades:2014:RSC

[Rho14]  B. E. Rhoades.  RE-

TRACTED: Some comments concerning cyclic type contractions. Applied Mathematics and Computation, 233(??):505–506, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314002021.

See retraction notices [?, Rho19].

Rhoades:2019:RNS

[Rho19] B. E. Rhoades. Retraction notice to “Some Comments Concerning Cyclic Type Contractions” [Applied Mathematics and Computation 233 (2014) 505–506]. Applied Mathematics and Computation, 341(??):452, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031830818X.

See [Rho14].

Raza:2018:FTM

[RHP18] Hassan Raza, Sakander Hayat, and Xiang-Feng Pan. On the faulttolerant metric dimension of convex polytopes. Applied Mathematics and Computation, 339(??):158–171, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830568X.

Rabbani:2019:MHM

[RKQ19] Quazzafi Rabbani, Aamir Khan, and Abdul Quddoos. Modified Hungarian method for unbalanced assignment problem with multiple jobs. Applied Mathematics and Computation, 361(??):493–498, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300319304436.

Rohan:2015:MNP

[RL15] E. Rohan and V. Lukes. Modeling nonlinear phenomena in deforming fluidsaturated porous media using homogenization and sensitivity analysis concepts. Applied Mathematics and Computation, 267(??):583–595, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000843.

Rohan:2019:HVA

[RL19] E. Rohan and V. Lukes. Homogenization of the vibro-acoustic transmission on perforated plates. Applied Mathematics and Computation, 361(??): 821–845, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL https://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304709.

Roanes-Lozano:2015:PKB

[RLGGAV15] Eugenio Roanes-Lozano, José Luis Galán-García, and Gabriel Aguilera-Venegas. A portable knowledge-based system for car breakdown evaluation. Applied Mathematics and Computation, 267(??):758–770, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016464.

Roanes-Lozano:2017:PRP

[RLGGAV17] Eugenio Roanes-Lozano, José Luis Galán-García, and Gabriel Aguilera-Venegas. A prototype of a RBES for personalized menus generation. Applied Mathematics and Computation, 315(??):615–624, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307573.

Ruan:2015:AFL

[RLH15] Yuhong Ruan, Anwei Li, and Jing Huang. Analytic function of local rules of Elementary Cellular Automata. Applied Mathematics and Computation, 258(??):404–409, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001757.

Rihan:2019:OCT

[RLM19] F. A. Rihan, S. Lakshmanan, and H. Maurer. Optimal control of tumour-immune model with time-delay and immunochemotherapy. Applied Mathematics and Computation, 353(??):147–165, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301006.

Ramos-Lopez:2016:OSP

[RLSGFMMF16] D. Ramos-López, M. A. Sánchez-Granero, M. Fernández-Martínez, and A. Martínez-Finkelshtein. Optimal sampling patterns for Zernike polynomials. Applied Mathematics and Computation, 274(??): 247–257, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031501468X.

Ran:2019:USC

[RLZ19] Maohua Ran, Taibai Luo, and Li Zhang. Unconditionally stable compact theta schemes for solving the linear and semilinear fourth-order diffusion equations. Applied Mathematics and Computation, 342(??):118–129, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307987.

Rolania:2015:EDT

[RM15] D. Barrios Rolanía and D. Manrique. On the existence of Darboux transformations for banded matrices. Applied Mathematics and Computation, 253(??):116–125, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401707X.

Ruocco:2016:EEM

[RM16] Eugenio Ruocco and Vincenzo Mallardo. An enhanced exponential matrix approach aimed at the stability of piecewise beams on elastic foundation. Applied Mathematics and Computation, 285(??):8– 25, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302132.

Rao:2018:NSG

[RM18a] S. Chandra Sekhara Rao and Manisha. Numerical solution of generalized Black–Scholes model. Applied Mathematics and Computation, 321(??): 401–421, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317306938.

Ruiz:2018:ILT

[RM18b] A. Ruiz and C. Muriel. On the integrability of Liénard I-type equations via λsymmetries and solvable structures. Applied Mathematics and Computation, 339(??):874–887, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306222.

Roul:2019:NAM

[RM19a] Pradip Roul and Harshita Madduri. A new approximate method and its convergence for a strongly nonlinear problem governing electrohydrodynamic flow of a fluid in a circular cylindrical conduit. Applied Mathematics and Computation, 341(??):335–347, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300318307823.

Ruocco:2019:BVA

[RM19b] Eugenio Ruocco and Vincenzo Mallardo. Buckling and vibration analysis nanoplates with imperfections. Applied Mathematics and Computation, 357(??):282–296, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302310.

Rivaie:2015:NCN

[RMA15] Mohd Rivaie, Mustafa Mamat, and Abdelrhaman Abashar. A new class of nonlinear conjugate gradient coefficients with exact and inexact line searches. Applied Mathematics and Computation, 268(??):1152–1163, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009315.

Ratnavelu:2015:SFB

[RMB15] K. Ratnavelu, M. Manikandan, and P. Balasubramaniam. Synchronization of fuzzy bidirectional associative memory neural networks with various time delays. Applied Mathematics and Computation, 270(??):582–605, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009807.

Rodrigues:2019:MMC

[RMCG19] D. S. Rodrigues, P. F. A. Mancera, T. Carvalho, and L. F. Gonçalves. A mathematical model for chemoimmunotherapy of chronic lymphocytic leukemia. Applied Mathematics and Computation, 349(??):118–133, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031831049X.

Ramya:2018:AER

[RMD18] E. Ramya, M. Muthtamilselvan, and Deog Hee Doh. Absorbing/emitting radiation and slanted hydromagnetic effects on micropolar liquid containing gyrostatic microorganisms. Applied Mathematics and Computation, 324(??):69–81, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308548.

Rohaninasab:2018:NSH

[RME18] N. Rohaninasab, K. Maleknejad, and R. Ezzati. Numerical solution of highorder Volterra–Fredholm integro-differential equations by using Legendre collocation method. Applied Mathematics and Computation, 328(??): 171–188, July 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318300535.

Roul:2019:SON

[RMK19] Pradip Roul, Harshita Madduri, and Klaus Kassner. A sixth-order numerical method for a strongly nonlinear singular boundary value problem governing electrohydrodynamic flow in a circular cylindrical conduit. Applied Mathematics and Computation, 350(??):416–433, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311317.

Ribera:2019:OHB

[RMM19] H. Ribera, T. G. Myers, and M. M. MacDevette. Optimising the heat balance integral method in spherical and cylindrical Stefan problems. Applied Mathematics and Computation, 354(??):216–231, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301432.

Robles:2016:GDI

[RMSD16] J. Robles, M. F. Martínez-Serrano, and E. Dopazo. On the generalized Drazin inverse in Banach algebras in terms of the generalized Schur complement. Applied Mathematics and Computation, 284(??):162–168, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630176X.

Rathinasamy:2018:AMS

[RN18] Anandaraman Rathinasamy and Priya Nair. Asymptotic mean-square stability of weak secondorder balanced stochastic Runge–Kutta methods for multi-dimensional Itô stochastic differential systems. Applied Mathematics and Computation, 332(??):276–303, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302388.

Rathinasamy:2019:MSS

[RN19] A. Rathinasamy and J. Narayanasamy.

Mean square stability and almost sure exponential stability of two step Maruyama methods of stochastic delay Hopfield neural networks. Applied Mathematics and Computation, 348(??):126–152, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310373.

RamReddy:2015:SRM

[RNM15] Ch. RamReddy, P. A. Lakshmi Narayana, and S. S. Motsa. A spectral relaxation method for linear and non-linear stratification effects on mixed convection in a porous medium. Applied Mathematics and Computation, 268(??):991–1000, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009182.

Rogoza:2019:MPT

[Rog19] Walery Rogoza. Method for the prediction of time series using small sets of experimental samples. Applied Mathematics and Computation, 355(??):108–122, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301663.

Rocha:2016:LQD

[ROGS16] Rogério Azevedo Rocha, Paulo Roberto Oliveira, Ronaldo Malheiros Gregório, and Michael Souza. Logarithmic quasi-distance proximal point scalarization method for multiobjective programming. Applied Mathematics and Computation, 273(??): 856–867, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315014162.

Romani:2019:IRF

[Rom19] Lucia Romani. Interpolating m-refinable functions with compact support: the second generation class. Applied Mathematics and Computation, 361(??):735–746, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304837.

Rehor:2016:SFP

[RP16] Martin Rehor and Vít Prusa. Squeeze flow of a piezoviscous fluid. Applied Mathematics and Computation, 274(??):414–429, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014708.

Ramos:2018:HMS

[RP18] Higinio Ramos and Paul Popescu. How many kstep linear block methods exist and which of them is the most efficient and simplest one? Applied Mathematics and Computation, 316(??):296–309, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305878.

Ramsden:2019:TRD

[RP19] Lewis Ramsden and Apostolos D. Papaioannou. On the time to ruin for a dependent delayed capital injection risk model. Applied Mathematics and Computation, 352(??):119–135, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300372.

Rad:2015:PEA

[RPB15] Jamal Amani Rad, Kourosh Parand, and Luca Vincenzo Ballestra. Pricing European and American options by radial basis point interpolation. Applied Mathematics and Computation, 251(??):363–377, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015380.

Rostami:2018:TCI

[RPJ^{+}18] Zahra Rostami, Viet-Thanh Pham, Sajad Jafari, Fatemeh Hadaeghi, and Jun Ma. Taking control of initiated propagating wave in a neuronal network using magnetic radiation. Applied Mathematics and Computation, 338(??):123–140, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304922.

Rajan:2015:VTR

[RR15] M. P. Rajan and G. D.

Reddy. A variant of Tikhonov regularization for parabolic PDE with space derivative multiplied by a small parameter ϵ. Applied Mathematics and Computation, 259(??):412–426, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002714.

Ramos:2018:TDM

[RR18a] Higinio Ramos and M. A.

Rufai. Third derivative modification of k-step block Falkner methods for the numerical solution of second order initial-value problems. Applied Mathematics and Computation, 333(??):231–245, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318302790.

Rathan:2018:IWE

[RR18b] Samala Rathan and G. Naga Raju. Improved weighted ENO scheme based on parameters involved in nonlinear weights. Applied Mathematics and Computation, 331(??):120–129, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302078.

Ramirez:2016:RPS

[RRM16] J. Ramírez, J. L. Romero, and C. Muriel. Reductions of PDEs to second order ODEs and symbolic computation. Applied Mathematics and Computation, 291(??):122–136, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304234.

R:2019:MME

[RRR19] Padma R., Ponalagusamy R., and Tamil Selvi R. Mathematical modeling of electro hydrodynamic non-Newtonian fluid flow through tapered arterial stenosis with periodic body acceleration and applied magnetic field. Applied Mathematics and Computation, 362 (??):Article 124453, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304175.

See corrigenda [RRR20, PPS25].

R:2020:CMM

[RRR20] Padma R., Ponalagusamy R., and Tamil Selvi R. Corrigendum to “Mathematical modeling of electro hydrodynamic non-Newtonian fluid flow through tapered arterial stenosis with periodic body acceleration and applied magnetic field” [Applied Mathematics and Computation, 362 (2019) 124453]. Applied Mathematics and Computation, 373(??):Article 125031, May 15, 2020. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319310239. See

[RRR19].

Rachunkova:2016:CAB

[RRRR16] Irena Rachunková, Lukás Rachunek, András Rontó, and Miklós Rontó. A constructive approach to boundary value problems with state-dependent impulses. Applied Mathematics and Computation, 274(??):726–744, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015234.

Rosa:2018:OCC

[RRS^{+}18] S. Rosa, P. Rebelo, C. M. Silva, H. Alves, and P. G. Carvalho. Optimal control of the customer dynamics based on marketing policy. Applied Mathematics and Computation, 330(??):42–55, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301292.

Razminia:2015:PRV

[RRT15] Kambiz Razminia, Abolhassan Razminia, and Delfim F. M. Torres. Pressure responses of a vertically hydraulic fractured well in a reservoir with fractal structure. Applied Mathematics and Computation, 257(??):374–380, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017858.

Roman-Roman:2015:SMR

[RRTR15] P. Román-Román and F. Torres-Ruiz. A stochastic model related to the Richards-type growth curve. estimation by means of simulated annealing and variable neighborhood search. Applied Mathematics and Computation, 266(??):579–598, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007201.

Ronto:2015:NAN

[RRV15] András Rontó, Miklós Rontó, and Jana Varha. A new approach to nonlocal boundary value problems for ordinary differential systems. Applied Mathematics and Computation, 250(??):689–700, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015434.

Rudolf:2015:TIL

[RRW^{+}15] Florian Rudolf, Karl Rupp, Josef Weinbub, Andreas Morhammer, and Siegfried Selberherr. Transformation invariant local element size specification. Applied Mathematics and Computation, 267(??):195–206, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004853.

Ray:2015:CSA

[RS15a] S. Saha Ray and S. Sahoo. A comparative study on the analytic solutions of fractional coupled sine-Gordon equations by using two reliable methods. Applied Mathematics and Computation, 253(??):72–82, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017068.

Rehan:2015:CBP

[RS15b] Kashif Rehan and Shahid S. Siddiqi. A combined binary 6-point subdivision scheme. Applied Mathematics and Computation, 270(??):130–135, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010851.

Rehan:2015:FTS

[RS15c] Kashif Rehan and Shahid S.

Siddiqi. A family of ternary subdivision schemes for curves. Applied Mathematics and Computation, 270(??):114–123, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010796.

Rossikhin:2015:NDR

[RS15d] Yury A. Rossikhin and Marina V. Shitikova. Nonlinear dynamic response of a fractionally damped cylindrical shell with a three-to-one internal resonance. Applied Mathematics and Computation, 257(??):498–525, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000326.

Rehan:2016:CTP

[RS16a] Kashif Rehan and Muhammad Athar Sabri. A combined ternary 4-point subdivision scheme. Applied Mathematics and Computation, 276(??):278–283, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315300047.

Rodriguez:2016:SPG

[RS16b] José M. Rodríguez and José M. Sigarreta. Spectral properties of geometricarithmetic index. Applied Mathematics and Computation, 277(??):142–153, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315300345.

Ramos:2017:TOS

[RS17a] Higinio Ramos and Gurjinder Singh. A tenth order A-stable two-step hybrid block method for solving initial value problems of ODEs. Applied Mathematics and Computation, 310(??):75–88, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302722.

Reza:2017:LSE

[RS17b] Amit Reza and Anand S. Sengupta. Least square ellipsoid fitting using iterative orthogonal transformations. Applied Mathematics and Computation, 314(??):349–359, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304769.

Ramos:2019:DSP

[RS19] A. B. M. Ramos and P. H. T. Schimit. Disease spreading on populations structured by groups. Applied Mathematics and Computation, 353(??):265–273, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300700.

Riffnaller-Schiefer:2016:ISA

[RSAF16] A. Riffnaller-Schiefer, U. H.

Augsdörfer, and D. W. Fellner. Isogeometric shell analysis with NURBS compatible subdivision surfaces. Applied Mathematics and Computation, 272 (part 1)(??):139–147, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008942.

Rong:2016:LBS

[RSC16] Fumei Rong, Baochang Shi, and Xilin Cui. Lattice Boltzmann simulation of heat and fluid flow in 3D cylindrical heat exchanger with porous blocks. Applied Mathematics and Computation, 276(??):367–378, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315014277.

Rajavel:2017:FTN

[RSC^{+}17] S. Rajavel, R. Samidurai, Jinde Cao, Ahmed Alsaedi, and Bashir Ahmad. Finite-time nonfragile passivity control for neural networks with timevarying delay. Applied Mathematics and Computation, 297(??):145–158, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306452.

Rodriguez-Sanchez:2016:EGB

[RSGFV16] Rosa Rodriguez-Sánchez, J. A. García, and J. Fdez-Valdivia. Evolutionary games between authors and their editors. Applied Mathematics and Computation, 273(??):645–655, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013764.

Ramos:2015:SFO

[RSKB15] Higinio Ramos, Gurjinder Singh, V. Kanwar, and Saurabh Bhatia. Solving first-order initial-value problems by using an explicit non-standard Astable one-step method in variable step-size formulation. Applied Mathematics and Computation, 268(??):796–805, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009005.

Ramos:2016:EVS

[RSKB16] Higinio Ramos, Gurjinder Singh, V. Kanwar, and Saurabh Bhatia. An efficient variable step-size rational Falknertype method for solving the special secondorder IVP. Applied Mathematics and Computation, 291(??):39–51, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304040.

Reisen:2019:RFM

[RSLL^{+}19] Valdério Anselmo Reisen, Adriano Marcio Sgrancio, Céline Lévy-Leduc, Pascal Bondon, Edson Zambon Monte, Higor Henrique Aranda Cotta, and Flávio Augusto Ziegelmann. Robust factor modelling for highdimensional time series: an application to air pollution data. Applied Mathematics and Computation, 346(??):842–852, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830849X.

Rybdylova:2018:MTP

[RSO^{+}18] O. Rybdylova, S. S. Sazhin, A. N. Osiptsov, F. B. Kaplanski, S. Begga, and M. Heikal. Modelling of a two-phase vortexring flow using an analytical solution for the carrier phase. Applied Mathematics and Computation, 326(??):159–169, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0096300318300031.

Ren:2018:GTS

[RT18a] Quanwei Ren and Hongjiong Tian. Generalized twostep Maruyama methods for stochastic differential equations. Applied Mathematics and Computation, 332(??):48–57, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301681.

Rybolowicz:2018:GTG

[RT18b] Bernard Rybolowicz and Agnieszka Tereszkiewicz. Generalized Tricobsthal and generalized Tribonacci polynomials. Applied Mathematics and Computation, 325(??):297–308, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300018.

Rubino:2019:ETS

[Rub19] Samuele Rubino. An efficient time-splitting approximation of the Navier– Stokes equations with LPS modeling. Applied Mathematics and Computation, 348(??):318–337, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310397.

Rui:2018:IIS

[Rui18]  Weiguo Rui.  Idea of

invariant subspace combined with elementary integral method for investigating exact solutions of time-fractional NPDEs. Applied Mathematics and Computation, 339(??):153–157, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305988.

Rossini:2016:QIO

[RV16] Milvia Rossini and Elena Volontè. Quasi-interpolation operators on hexagonal grids with high approximation orders in spaces of polyharmonic splines. Applied Mathematics and Computation, 272 (part 1)(??):223–234, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010413.

Renhe:2019:STB

[RVE19] Marcelo Caniato Renhe, Marcelo Bernardes Vieira, and Claudio Esperança. A stable tensor-based method for controlled fluid simulations. Applied Mathematics and Computation, 343(??):195–213, February 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318308300.

Rakkiyappan:2017:ESL

[RVGS17] R. Rakkiyappan, G. Velmurugan, J. Nicholas George, and R. Selvamani. Exponential synchronization of Lur’e complex dynamical networks with uncertain inner coupling and pinning impulsive control. Applied Mathematics and Computation, 307(??):217–231, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301522.

Ren:2017:FOE

[RW17] Lei Ren and Yuan-Ming Wang. A fourth-order extrapolated compact difference method for timefractional convection-reaction-diffusion equations with spatially variable coefficients. Applied Mathematics and Computation, 312(??):1–22, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317303387.

Rehman:2015:SSR

[RWH15] Abdur Rehman, Qing-Wen Wang, and Zhuo-Heng He. Solution to a system of real quaternion matrix equations encompassing η-Hermicity. Applied Mathematics and Computation, 265(??):945–957, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007353.

Ran:2015:HLI

[RWW15] Yu-Hong Ran, Jun-Gang Wang, and Dong-Ling Wang. On HSS-like iteration method for the space fractional coupled nonlinear Schrödinger equations. Applied Mathematics and Computation, 271(??):482–488, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315012643.

Ren:2018:IHC

[RWW18] Yuanhong Ren, Weiqun Wang, and Yixiang Wang. Incremental H_{∞} control for switched nonlinear systems. Applied Mathematics and Computation, 331(??):251–263, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301814.

Ren:2019:SAS

[RWY19] Yong Ren, Kai Wang, and Huijin Yang. Stability analysis of stochastic pantograph multigroup models with dispersal driven by G-brownian motion. Applied Mathematics and Computation, 355(??):356–365, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301961.

Ruan:2019:SHS

[RXL19] Dehao Ruan, Liping Xu, and Jiaowan Luo. Stability of hybrid stochastic functional differential equations. Applied Mathematics and Computation, 346(??):832–841, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302376.

Ren:2019:FCB

[RXLM19] Guodong Ren, Yuxiong Xue, Yuwei Li, and Jun Ma. Field coupling benefits signal exchange between Colpitts systems. Applied Mathematics and Computation, 342(??):45– 54, February 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318307896.

Ren:2019:LAP

[RXQ19] Zhicong Ren, Wei Xu, and Yan Qiao. Local averaged path integration method approach for nonlinear dynamic systems. Applied Mathematics and Computation, 344–345(??):68–77, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308488.

Ran:2017:QFF

[RXZW17] Suzhen Ran, Yanmei Xue, Bo-Chao Zheng, and Zhenyou Wang. Quantized feedback fuzzy sliding mode control design via memory-based strategy. Applied Mathematics and Computation, 298(??):283–295, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306798.

Radenovic:2019:BGL

[RZD^{+}19] Stojan Radenović, Kastriot Zoto, Nebojsa Dedović, Vesna Sesum-Cavic, and Arslan Hojat Ansari. Bhaskar–Guo–Lakshmikantam– Ćirić type results via new functions with applications to integral equations. Applied Mathematics and Computation, 357(??):75–87, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302656.

Ran:2019:AAM

[RZH^{+}19] Yingli Ran, Zhao Zhang, Xiaohui Huang, Xiaosong Li, and Ding-Zhu Du. Approximation algorithms for minimum weight connected 3-path vertex cover. Applied Mathematics and Computation, 347(??):

723–733, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318310191.

Ren:2016:FTC

[RZHY16] Hangli Ren, Guangdeng Zong, Linlin Hou, and Yang Yi. Finite-time control of interconnected impulsive switched systems with time-varying delay. Applied Mathematics and Computation, 276(??):143–157, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315016185.

Ribeiro:2017:EHT

[RZMB17] Vinicius G. Ribeiro, Jorge G. Zabadal, Cíntia O. Monticelli, and Volnei Borges. Estimating heat transfer coefficients for solid-gas interfaces using the Landau– Teller model. Applied Mathematics and Computation, 301(??):135–139, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307421.

Rabbani:2016:JOD

[RZR16] Masoud Rabbani, Nadia Pourmohammad Zia, and Hamed Rafiei. Joint optimal dynamic pricing and replenishment policies for items with simultaneous quality and physical quantity deterioration. Applied Mathematics and Computation, 287– 288(??):149–160, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302661.

Ren:2015:MNE

[RzS15] Jincheng Ren and Zhi zhong Sun. Maximum norm error analysis of difference schemes for fractional diffusion equations. Applied Mathematics and Computation, 256(??):299–314, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000156.

Raja:2015:ISA

[RZSS15] R. Raja, Quanxin Zhu, S. Senthilraj, and R. Samidurai. Improved stability analysis of uncertain neutral type neural networks with leakage delays and impulsive effects. Applied Mathematics and Computation, 266 (??):1050–1069, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008115.

Ruan:2018:SIF

[RZW18] Zhousheng Ruan, Wen Zhang, and Zewen Wang. Simultaneous inversion of the fractional order and the space-dependent source term for the timefractional diffusion equation. Applied Mathematics and Computation, 328(??):365–379, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300468.

S:2015:NFS

[S.15] Cesar A. Gómez S. A nonlinear fractional Sharma– Tasso–Olver equation: New exact solutions. Applied Mathematics and Computation, 266(??):385–389, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300315006980.

Sabir:2016:OPC

[SA16a] Mirza Muhammad Sabir and Tariq Ali. Optimal PID controller design through swarm intelligence algorithms for sun tracking system. Applied Mathematics and Computation, 274(??):690–699, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501526X.

Sayevand:2016:FVE

[SA16b] K. Sayevand and F. Arjang. Finite volume element method and its stability analysis for analyzing the behavior of subdiffusion problems. Applied Mathematics and Computation, 290(??): 224–239, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300316303794.

Sharma:2016:NFO

[SA16c] Janak Raj Sharma and Himani Arora. A new family of optimal eighth order methods with dynamics for nonlinear equations. Applied Mathematics and Computation, 273(??):924–933, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014009.

Sharma:2016:SNO

[SA16d] Janak Raj Sharma and Himani Arora. Some novel optimal eighth order derivative-free root solvers and their basins of attraction. Applied Mathematics and Computation, 284(??):149–161, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301734.

Safonov:2018:CMT

[SA18a] Alexander Safonov and Andrew Adamatzky. Computing via material topology optimisation. Applied Mathematics and Computation, 318(??):109–120, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305817.

Sakar:2018:SQH

[SA18b] F. M. Sakar and M. Aydogan. Subclass of mquasiconformal harmonic functions in association with Janowski starlike functions. Applied Mathematics and Computation, 319(??):461–468, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303016.

Samir:2019:IBP

[SA19] Chafik Samir and Ines Adouani. C^{1} interpolating Bézier path on Riemannian manifolds, with applications to 3D shape space. Applied Mathematics and Computation, 348(??):371–384, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310348.

Sahihi:2019:CMB

[SAA19] Hussein Sahihi, Saeid Abbasbandy, and Tofigh Allahviranloo. Computational method based on reproducing kernel for solving singularly perturbed differential-difference equations with a delay. Applied Mathematics and Computation, 361(??):583–598, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319304758.

Sabzevari:2018:ENS

[Sab18] Mehdi Sabzevari. Erratum to “Numerical solution of linear Fredholm integral equation by using hybrid Taylor and Block–Pulse functions” [Appl. Math. Comput. 149 (2004) 799– 806]. Applied Mathematics and Computation, 339(??):899–915, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305824. See

[MM04].

Sadri:2018:NOM

[SAC18] K. Sadri, A. Amini, and C. Cheng. A new operational method to solve Abel’s and generalized Abel’s integral equations. Applied Mathematics and Computation, 317(??):49–67, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306185.

Saeed:2017:CPM

[Sae17] Umer Saeed. CAS Picard method for fractional nonlinear differential equation. Applied Mathematics and Computation, 307(??):102–112, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301558.

Sharma:2019:EGP

[SAES19] D. Sharma, S. Amiroudine, A. Erriguible, and T. K. Sengupta. Error growth and phase lag analysis for high Courant numbers. Applied Mathematics and Computation, 346(??):374–384, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309019.

Sahoo:2015:RAF

[Sah15] Banshidhar Sahoo. Role of additional food in ecoepidemiological system with disease in the prey. Applied Mathematics and Computation, 259(??):61– 79, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002180.

Sahiner:2019:DCT

[Sah19a] Burak Sahiner. Direction curves of tangent indicatrix of a curve. Applied Mathematics and Computation, 343(??):273–284, February 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318307938.

Sahlan:2019:FCA

[Sah19b] M. Nosrati Sahlan. Four computational approaches for solving a class of boundary value problems arising in chemical reactor industry. Applied Mathematics and Computation, 355(??):253–268, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300256.

Sader:2018:GDS

[SAJ18] Malika Sader, Abdujelil Abdurahman, and Haijun Jiang. General decay synchronization of delayed BAM neural networks via nonlinear feedback control. Applied Mathematics and Computation, 337(??):281–301, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304557.

Sakran:2019:NSI

[Sak19] M. R. A. Sakran. Numerical solutions of integral and integro-differential equations using Chebyshev polynomials of the third kind. Applied Mathematics and Computation, 351(??):66–82, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300396.

Sakthivel:2016:RAS

[SAM^{+}16] R. Sakthivel, R. Anbuvithya, K. Mathiyalagan, Yong-Ki Ma, and P. Prakash. Reliable antisynchronization conditions for BAM memristive neural networks with different memductance functions. Applied Mathematics and Computation, 275(??):213–228, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015623.

Sanahuja:2016:NRA

[San16] Silvia M. Sanahuja. New rough approximations for n-cycles and n-paths. Applied Mathematics and Computation, 276(??):96– 108, March 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315015544.

Santoprete:2019:CVE

[San19] Manuele Santoprete. Countering violent extremism: a mathematical model. Applied Mathematics and Computation, 358(??):314–329, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031930342X.

Sapa:2018:DMP

[Sap18]  Lucjan Sapa.  Differ-

ence methods for parabolic equations with Robin condition. Applied Mathematics and Computation, 321(??):794–811, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307701.

Sarkar:2017:IPD

[Sar17]  Tanmay Sarkar.  Inte-

rior penalty discontinuous Galerkin method for magnetic induction equation with resistivity. Applied Mathematics and Computation, 314(??):212–227, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304691.

Soheili:2019:FCT

[SAS19] Ali R. Soheili, Mohammad Amini, and Fazlollah Soleymani. A family of Chaplygin-type solvers for Itô stochastic differential equations. Applied Mathematics and Computation, 340(??):296–304, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307525.

Suganya:2015:ERI

[SAT15] S. Suganya, M. Mallika Arjunan, and J. J. Trujillo. Existence results for an impulsive fractional integro-differential equation with state-dependent delay. Applied Mathematics and Computation, 266(??):54–69, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006487.

Sharma:2015:SOT

[SB15] Rajni Sharma and Ashu Bahl. A sixth order transformation method for finding multiple roots of nonlinear equations and basin attractors for various methods. Applied Mathematics and Computation, 269(??):105–117, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009753.

Sharma:2016:VSA

[SB16] A. K. Sharma and E. Birgersson. Validity and scalability of an asymptotically reduced singlechannel model for full-size catalytic monolith converters. Applied Mathematics and Computation, 281(??):186–198, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300625.

Shaghaghi:2017:CAG

[SBG^{+}17] Saba Shaghaghi, Hossein Bonakdari, Azadeh Gholami, Isa Ebtehaj, and Maryam Zeinolabedini. Comparative analysis of GMDH neural network based on genetic algorithm and particle swarm optimization in stable channel design. Applied Mathematics and Computation, 313(??):271–286, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304241.

Siddiqa:2016:REI

[SBH16] Sadia Siddiqa, Naheed Begum, and M. Anwar Hossain. Radiation effects from an isothermal vertical wavy cone with variable fluid properties. Applied Mathematics and Computation, 289(??):149–158, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303071.

Siddiqa:2017:CDG

[SBH17] Sadia Siddiqa, Naheed Begum, and Md. Anwar Hossain. Compressible dusty gas along a vertical wavy surface. Applied Mathematics and Computation, 293(??):600–610, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630532X.

Si:2016:FPP

[SBL16] X. Si, J. Baccou, and J. Liandrat. On four-point penalized Lagrange subdivision schemes. Applied Mathematics and Computation, 281(??):278–299, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300303.

Song:2017:OCP

[SBLZ17] Jingjing Song, Xiuchun Bi, Rong Li, and Shuguang Zhang. Optimal consumption and portfolio selection problems under loss aversion with downside consumption constraints. Applied Mathematics and Computation, 299(??):80–94, April 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307032.

Sharma:2015:TCE

[SBV15] A. K. Sharma, E. Birgersson, and M. Vynnycky. Towards computationallyefficient modeling of transport phenomena in threedimensional monolithic channels. Applied Mathematics and Computation, 254(??):392–407, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000569.

Sune:2017:IOS

[SC17] Víctor Suñé and Juan Antonio Carrasco. Implicit ODE solvers with good local error control for the transient analysis of Markov models. Applied Mathematics and Computation, 293(??):96–111, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305045.

Sadeghi:2018:SNP

[SC18a] Amir Sadeghi and João R.

Cardoso. Some notes on properties of the matrix Mittag-Leffler function. Applied Mathematics and Computation, 338(??):624–630, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305265.

Shi:2018:MPS

[SC18b] Hongxia Shi and Haibo Chen. Multiple positive solutions for nonhomogeneous Klein–Gordon-Maxwell equations. Applied Mathematics and Computation, 337(??): 494–503, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318304697.

Shah:2015:RDO

[SCB15] Nita H. Shah and Leopoldo Eduardo Cárdenas-Barrón. Retailer’s decision for ordering and credit policies for deteriorating items when a supplier offers order-linked credit period or cash discount. Applied Mathematics and Computation, 259(??):569–578, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003045.

Shi:2017:ESC

[SCC17] Yanchao Shi, Jinde Cao, and Guanrong Chen. Exponential stability of complex-valued memristorbased neural networks with time-varying delays. Applied Mathematics and Computation, 313(??):222–234, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303909.

Schimit:2016:EAS

[Sch16] P. H. T. Schimit. Evolutionary aspects of spatial Prisoner’s Dilemma in a population modeled by continuous probabilistic cellular automata and genetic algorithm. Applied Mathematics and Computation, 290(??):178–188, November 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316303708.

Sunil:2019:SAC

[SCM19] Sunil, Shalu Choudhary, and Amit Mahajan. Stability analysis of a couplestress fluid saturating a porous medium with temperature and pressure dependent viscosity using a thermal non-equilibrium model. Applied Mathematics and Computation, 340(??):15–30, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307392.

Szulc:2015:STP

[SCN15] Tomasz Szulc, Ljiljana Cvetković, and Maja Nedović. Scaling technique for Partition–Nekrasov matrices. Applied Mathematics and Computation, 271(??):201–208, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012175.

Saito:2015:COS

[SCR^{+}15] G. Saito, H. W. Corley, Jay M. Rosenberger, Tai-Kuan Sung, and Alireza Noroziroshan. Constraint optimal selection techniques (COSTs) for nonnegative linear programming problems. Applied Mathematics and Computation, 251(??):586–598, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016129.

Starzynski:2015:CCV

[SCS^{+}15] Jacek Starzynski, Bartosz Chaber, Robert Szmurlo, Zuzanna Krawczyk, and Anna Zawadzka. Conformance criteria for validation of target volume surface reconstructed from delineation. Applied Mathematics and Computation, 267(??):456–464, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500137X.

Sun:2019:FNMb

[SCY19] Yan-Ping Sun, Shao-Chun Chen, and Yong-Qin Yang. The families of nonconforming mixed finite elements for linear elasticity on simplex grids. Applied Mathematics and Computation, 358(??):348–362, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302188.

Strocchi:2017:GMM

[SCZ^{+}17] M. Strocchi, C. Contarino, Q. Zhang, R. Bonmassari, and E. F. Toro. A global mathematical model for the simulation of stenoses and bypass placement in the human arterial system. Applied Mathematics and Computation, 300(??):21– 39, May 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306968.

Sengupta:2017:USB

[SD17] Binanda Sengupta and Abhijit Das. Use of SIMD-based data parallelism to speed up sieving in integer-factoring algorithms. Applied Mathematics and Computation, 293(??):204–217, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305148.

Sportelli:2019:FPD

[SD19a] Mario Sportelli and Luigi De Cesare. Fiscal policy delays and the classical growth cycle. Applied Mathematics and Computation, 354(??):9–31, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301341.

Sun:2019:SLN

[SD19b] Shaowei Sun and Kinkar Ch.

Das. On the second largest normalized Laplacian eigenvalue of graphs. Applied Mathematics and Computation, 348(??):531–541, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310701.

Sun:2019:NSC

[SD19c] Xiaoyun Sun and Pei Dang. Numerical stability of circular Hilbert transform and its application to signal decomposition. Applied Mathematics and Computation, 359(??):357–373, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303820.

SaviodeGoesMaciel:2018:CAU

[SdA18] Edisson Sávio de Góes Maciel and Cláudia Regina de Andrade. Comparison among unstructured TVD, ENO and UNO schemes in two- and three-dimensions. Applied Mathematics and Computation, 321(??):130–175, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307257.

Sarala:2017:ZIG

[SDAB17] D. Sarala, Hanyuan Deng, S. K. Ayyaswamy, and S. Balachandran. The Zagreb indices of graphs based on four new operations related to the lexicographic product. Applied Mathematics and Computation, 309(??):156–169, September 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317302370.

Schmitter:2016:FSI

[SDGU16] D. Schmitter, R. Delgado-Gonzalo, and M. Unser. A family of smooth and interpolatory basis functions for parametric curve and surface representation. Applied Mathematics and Computation, 272 (part 1)(??):53–63, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500911X.

Stevic:2015:SDE

[SDIS15] Stevo Stević, Josef Diblík, Bratislav Iricanin, and Zdenek Smarda. On the system of difference equations x_{n} = x_{n−1}y_{n−2}/(ay_{n−2}+by_{n−1}), y_{n} = y_{n−1}x_{n−2}/(cx_{n−2} + dx_{n−1}). Applied Mathematics and Computation, 270(??):688–704, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011285.

Sitthithakerngkiet:2015:HVA

[SDK15] Kanokwan Sitthithakerngkiet, Jitsupa Deepho, and Poom Kumam. A hybrid viscosity algorithm via modify the hybrid steepest descent method for solving the split variational inclusion in image reconstruction and fixed point problems. Applied Mathematics and Computation, 250(??):986–1001, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314015057.

Song:2016:LOE

[SDP16] Xinmin Song, Zhenhua Duan, and Ju H. Park. Linear optimal estimation for discrete-time systems with measurementdelay and packet dropping. Applied Mathematics and Computation, 284(??):115–124, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316301576.

Serrano:2019:CFI

[SdR19] D. Hernández Serrano and A. Martín del Rey. A closed formula for the inverse of a reversible cellular automaton with (2R + 1)-cyclic rule. Applied Mathematics and Computation, 357(??):23–34, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319302681.

Salehi:2018:NSS

[SDS18] Younes Salehi, Mohammad T. Darvishi, and William E. Schiesser. Numerical solution of space fractional diffusion equation by the method of lines and splines. Applied Mathematics and Computation, 336(??):465–480, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303710.

Sun:2019:FDI

[SDWG19] F. L. Sun, C. Y. Dong, Y. H. Wu, and Y. P. Gong. Fast direct isogeometric boundary element method for 3D potential problems based on HODLR matrix. Applied Mathematics and Computation, 359(??):17–33, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303121.

Sun:2018:SMD

[SE18] Jie Sun and Joseph A.

Eichholz. Splitting methods for differential approximations of the radiative transfer equation. Applied Mathematics and Computation, 322(??):140–150, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300317308081.

Segeth:2015:PBS

[Seg15]  Karel Segeth.  A pe-

riodic basis system of the smooth approximation space. Applied Mathematics and Computation, 267(??):436–444, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001526.

Segeth:2018:SSP

[Seg18] Karel Segeth. Some splines produced by smooth interpolation. Applied Mathematics and Computation, 319(??):387–394, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302746.

Seadawy:2016:WWS

[SEKA16] A. R. Seadawy, O. H. El-Kalaawy, and R. B. Aldenari. Water wave solutions of Zufiria’s higherorder Boussinesq type equations and its stability. Applied Mathematics and Computation, 280(??):57–71, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300145.

Smith:2017:DON

[SFC^{+}17] N. A. S. Smith, A. T.

Fry, L. E. Crocker, F. Cernuschi, and L. Lorenzoni. Design and optimisation of the nozzle of an innovative high temperature solid particulate erosion testing system using finite element modelling. Applied Mathematics and Computation, 301(??):60– 69, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630755X.

Samar:2019:SAG

[SFLM19] Mahvish Samar, Aamir Farooq, Hanyu Li, and Chunlai Mu. Sensitivity analysis for the generalized Cholesky factorization. Applied Mathematics and Computation, 362 (??):Article 124556, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305399.

Shi:2018:SSD

[SFSR18] Shuang Shi, Zhongyang Fei, Zhenpeng Shi, and Shunqing Ren. Stability and stabilization for discrete-time switched systems with asynchronism. Applied Mathematics and Computation, 338(??):515–519, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305381.

Srivastava:2015:IMS

[SG15] Shwetabh Srivastava and Dharmendra K. Gupta. An iterative method for solving general restricted linear equations. Applied Mathematics and Computation, 262(??):344–353, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005056.

Simpson:2017:MAR

[SG17] Lindsay Simpson and Abba B. Gumel. Mathematical assessment of the role of pre-exposure prophylaxis on HIV transmission dynamics. Applied Mathematics and Computation, 293(??):168–193, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304817.

Solis:2018:ANM

[SG18] Francisco J. Solis and Luz M. Gonzalez. Analytical and numerical modeling of the evolution of human papillomavirus infected cells. Applied Mathematics and Computation, 324(??):28–35, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308184.

Shao:2019:MGA

[SG19a] Yanling Shao and Yubin Gao. The maximal geometric-arithmetic energy of trees with at most two branched vertices. Applied Mathematics and Computation, 362 (??):Article 124528, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305077.

Srivastav:2019:AIT

[SG19b] Akhil Kumar Srivastav and Mini Ghosh. Assessing the impact of treatment on the dynamics of dengue fever: a case study of India. Applied Mathematics and Computation, 362 (??):Article 124533, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305120.

Sun:2018:AZI

[SGDX18] Xiaoling Sun, Yubin Gao, Jianwei Du, and Lan Xu. Augmented Zagreb index of trees and unicyclic graphs with perfect matchings. Applied Mathematics and Computation, 335(??):75–81, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303576.

Silva-Garcia:2018:SBG

[SGFCRM^{+}18] V. M. Silva-García, R. Flores-Carapia, C. Rentería-Márquez, B. Luna-Benoso, and M. Aldape-Pérez. Substitution box generation using Chaos: An image encryption application. Applied Mathematics and Computation, 332(??):123–135, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830184X.

Srivastava:2015:SFP

[SGG15a] H. M. Srivastava, S. Gaboury, and F. Ghanim. Some further properties of a linear operator associated with the λ-generalized Hurwitz–Lerch zeta function related to the class of meromorphically univalent functions. Applied Mathematics and Computation, 259(??):1019–1029, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003070.

Srivastava:2015:UCA

[SGG15b] H. M. Srivastava, S. Gaboury, and F. Ghanim. A unified class of analytic functions involving a generalization of the Srivastava– Attiya operator. Applied Mathematics and Computation, 251(??):35–45, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015811.

Solin:2018:RAC

[SGGD^{+}18] Pavel Solin, José Luis Galán-García, Vit Dolejsí, Philipp Neumann, Peter Svacek, Jaroslav Kruis, and Volker John. Recent advances in computing. Applied Mathematics and Computation, 319(??):1, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307476.

Shubha:2016:FDR

[SGJS16] Vorkady. S. Shubha, Santhosh George, P. Jidesh, and M. E. Shobha. Finite dimensional realization of a quadratic convergence yielding iterative regularization method for ill-posed equations with monotone operators. Applied Mathematics and Computation, 273(??):1041–1050, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315014022.

Singh:2019:EOA

[SGKR19] Gurjinder Singh, Arvind Garg, V. Kanwar, and Higinio Ramos. An efficient optimized adaptive step-size hybrid block method for integrating differential systems. Applied Mathematics and Computation, 362(??): Article 124567, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305508.

Shen:2016:IBP

[SGM^{+}16] Yu-Jia Shen, Yi-Tian Gao, Gao-Qing Meng, Yi Qin, and Xin Yu. Improved Bell-polynomial procedure for the higherorder Korteweg–de Vries equations in fluid dynamics. Applied Mathematics and Computation, 274(??):403–413, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014344.

Singh:2016:SLC

[SGMH16] Sukhjit Singh, Dharmendra Kumar Gupta, E. Martínez, and José L. Hueso. Semilocal and local convergence of a fifth order iteration with Fréchet derivative satisfying Hölder condition. Applied Mathematics and Computation, 276(??): 266–277, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315015647.

Sun:2015:AGH

[SGY^{+}15] Junwei Sun, Jinchao Guo, Cunxiang Yang, Anping Zheng, and Xuncai Zhang. Adaptive generalized hybrid function projective dislocated synchronization of new four-dimensional uncertain chaotic systems. Applied Mathematics and Computation, 252(??):304–314, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031401649X.

Sahlan:2015:WGM

[SH15a] M. Nosrati Sahlan and E. Hashemizadeh. Wavelet Galerkin method for solving nonlinear singular boundary value problems arising in physiology. Applied Mathematics and Computation, 250(??):260–269, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0096300314014933.

Said:2015:HSO

[SH15b] Samer S. Said and Mohamed Haouari. A hybrid simulation-optimization approach for the robust Discrete Time/Cost Tradeoff Problem. Applied Mathematics and Computation, 259(??):628–636, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002933.

Sun:2015:CCO

[SH15c] Yuangong Sun and Taher S.

Hassan. Comparison criteria for odd order forced nonlinear functional neutral dynamic equations. Applied Mathematics and Computation, 251(??):387–395, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016270.

Svacek:2015:NSA

[SH15d] Petr Svácek and Jaromír Horácek. Numerical simulation of aeroelastic response of an airfoil in flow with laminar-turbulence transition. Applied Mathematics and Computation, 267(??):28–41, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008887.

Shao:2016:TGM

[SH16] Xinping Shao and Danfu Han. A two-grid method with backtracking technique for the stream function form of the Navier– Stokes equations. Applied Mathematics and Computation, 274(??):649–660, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015209.

Sherief:2018:CIS

[SH18a] Hany H. Sherief and Eman M. Hussein. Contour integration solution for a thermoelastic problem of a spherical cavity. Applied Mathematics and Computation, 320(??):557–571, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307233.

Svacek:2018:FEA

[SH18b] Petr Svácek and Jaromír Horácek. Finite element approximation of flow induced vibrations of human vocal folds model: Effects of inflow boundary conditions and the length of subglottal and supraglottal channel on phonation onset. Applied Mathematics and Computation, 319(??):178–194, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301303.

Steklova:2019:JHI

[SH19] Klara Steklova and Eldad Haber. 3D joint hydrogeophysical inversion using similarity measures. Applied Mathematics and Computation, 357(??):338–356, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730646X.

Sharma:2012:EVD

[Sha12]  Rajesh Sharma.  Ef-

fect of viscous dissipation and heat source on unsteady boundary layer flow and heat transfer past a stretching surface embedded in a porous medium using element free Galerkin method. Applied Mathematics and Computation, 219(3):976–987, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300312007023. See

comment [Pan15].

Salari:2015:AAS

[SHA15a] Hashem Salari, Hassan Hassanzadeh, and Jalal Abedi. Approximate analytical solutions for steady-state nonisothermal convection-diffusion-reaction in a slab. Applied Mathematics and Computation, 264(??):141–159, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005421.

Shadimetov:2015:OQF

[SHA15b] Kh. M. Shadimetov, A. R. Hayotov, and D. M. Akhmedov. Optimal quadrature formulas for Cauchy type singular integrals in Sobolev space. Applied Mathematics and Computation, 263(??): 302–314, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315005287.

Shang:2015:EIE

[Sha15c]  Yilun Shang.  The

Estrada index of evolving graphs. Applied Mathematics and Computation, 250(??):415–423, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015045.

Sharma:2015:CFM

[Sha15d]  Honey Sharma.  On

certain family of mixed summation integral type two-dimensional q-Lupas– Phillips–Bernstein operators. Applied Mathematics and Computation, 259(??):741–752, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003136.

Sharma:2015:ICH

[Sha15e] Janak Raj Sharma. Improved Chebyshev–Halley methods with sixth and eighth order convergence. Applied Mathematics and Computation, 256(??):

119–124, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315000077.

Shary:2015:NCS

[Sha15f] Sergey P. Shary. New characterizations for the solution set to interval linear systems of equations. Applied Mathematics and Computation, 265(??):570–573, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006463.

Shang:2016:CSM

[Sha16a] Yilun Shang. Consensus seeking over Markovian switching networks with time-varying delays and uncertain topologies. Applied Mathematics and Computation, 273(??):1234–1245, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315011960.

Sharma:2016:NCO

[Sha16b] Janak Raj Sharma. A note on the convergence order of some recent methods for solving nonlinear equations. Applied Mathematics and Computation, 273(??):793–796, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501406X.

Shen:2017:EES

[SHC^{+}17] Zhao-Li Shen, Ting-Zhu Huang, Bruno Carpentieri, Xian-Ming Gu, and Chun Wen. An efficient elimination strategy for solving PageRank problems. Applied Mathematics and Computation, 298(??):111–122, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306385.

Sheng:2018:CWM

[She18] Xingping Sheng. Computation of weighted Moore– Penrose inverse through Gauss–Jordan elimination on bordered matrices. Applied Mathematics and Computation, 323(??):64–74, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308238.

Slama:2019:APS

[SHEBS19] Howida Slama, A. Hussein, Nabila A. El-Bedwhey, and Mustafa M. Selim. An approximate probabilistic solution of a random SIR-type epidemiological model using RVT technique. Applied Mathematics and Computation, 361(??):144–156, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304126.

Sarfraz:2015:SPC

[SHH15] Muhammad Sarfraz, Malik Zawwar Hussain, and Farsia Hussain. Shape preserving curves using quadratic trigonometric splines. Applied Mathematics and Computation, 265(??):1126–1144, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007626.

Sojka:2018:IEM

[SHHC18] Radim Sojka, David Horák, Václav Hapla, and Martin Cermák. The impact of enabling multiple subdomains per MPI process in the TFETI domain decomposition method. Applied Mathematics and Computation, 319(??):586–597, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304927.

Shi:2015:NTG

[Shi15a] Yongtang Shi. Note on two generalizations of the Randić index. Applied Mathematics and Computation, 265(??):1019–1025, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008000.

Shin:2015:AAM

[Shi15b]  Yang Woo Shin.  Al-

gorithmic approach to Markovian multi-server retrial queues with vacations. Applied Mathematics and Computation, 250(??):287–297, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014544.

Shang:2018:SGB

[SHLL18] Kun Shang, Zheng-Hai Huang, Wanquan Liu, and Zhi-Ming Li. A single gallery-based face recognition using extended joint sparse representation. Applied Mathematics and Computation, 320(??):99– 115, March 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317305192.

Shama:2016:CTG

[SHLW16] Mu-Ga Shama, Ting-Zhu Huang, Jun Liu, and Si Wang. A convex total generalized variation regularized model for multiplicative noise and blur removal. Applied Mathematics and Computation, 276(??):109–121, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315016069.

Shadimetov:2016:OQF

[SHN16] Kh. M. Shadimetov, A. R. Hayotov, and F. A. Nuraliev. Optimal quadrature formulas of Euler– Maclaurin type. Applied Mathematics and Computation, 276(??):340–355, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300102.

Semary:2018:MMS

[SHR18] Mourad S. Semary, Hany N.

Hassan, and Ahmed G. Radwan. Modified methods for solving two classes of distributed order linear fractional differential equations. Applied Mathematics and Computation, 323(??):106–119, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308299.

Suzuki:2015:MDE

[SI15] Shinji Suzuki and Takehiro Inohara. Mathematical definitions of enclave and exclave, and applications. Applied Mathematics and Computation, 268(??):728–742, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008954.

Sugie:2019:NME

[SI19] Jitsuro Sugie and Kazuki Ishibashi. Nonoscillation of Mathieu equations with two frequencies. Applied Mathematics and Computation, 346(??):491–499, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309469.

Sidharth:2015:GOB

[SIA15] Manjari Sidharth, Nurhayat Ispir, and P. N. Agrawal. GBS operators of Bernstein– Schurer–Kantorovich type based on q-integers. Applied Mathematics and Computation, 269(??): 558–568, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315009716.

Siddiqui:2016:ZIZ

[SIA16] Muhammad Kamran Siddiqui, Muhammad Imran, and Ali Ahmad. On Zagreb indices, Zagreb polynomials of some nanostar dendrimers. Applied Mathematics and Computation, 280(??):132–139, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300316300418.

Sastre:2017:TAC

[SIA^{+}17] Jorge Sastre, Javier Ibáñez, Pedro Alonso, Jesús Peinado, and Emilio Defez. Two algorithms for computing the matrix cosine function. Applied Mathematics and Computation, 312(??):66–77, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303077.

Sastre:2019:BCM

[SID19] J. Sastre, J. Ibáñez, and E. Defez. Boosting the computation of the matrix exponential. Applied Mathematics and Computation, 340(??):206–220, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307318.

Silva:2016:KTN

[Sil16] Gilson N. Silva. Kantorovich’s theorem on Newton’s method for solving generalized equations under the majorant condition. Applied Mathematics and Computation, 286(??):178–188, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630265X.

Simsek:2015:BTP

[Sim15] Yilmaz Simsek. Beta-type polynomials and their generating functions. Applied Mathematics and Computation, 254(??):172–182, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017792.

Singh:2017:HWA

[Sin17] S. S. Singh. Harmonic waves in anisotropic nematic elastomers. Applied Mathematics and Computation, 302(??):1–8, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300012.

Singh:2018:ASF

[Sin18] Harendra Singh. Approximate solution of fractional vibration equation using Jacobi polynomials. Applied Mathematics and Computation, 317(??):85–100, January 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730615X.

Set:2015:NIH

[SISÖ15] Erhan Set, Imdat Iscan, M. Zeki Sarikaya, and M. Emin Özdemir. On new inequalities of Hermite– Hadamard–Fejér type for convex functions via fractional integrals. Applied Mathematics and Computation, 259(??):875–881, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003355.

Sun:2019:GGD

[SISV19] Qiang Sun, Barbara Ikica, Riste Skrekovski, and Vida Vukasinović. Graphs with a given diameter that maximise the Wiener index. Applied Mathematics and Computation, 356(??):438–448, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302267.

Schaum:2016:ESP

[SJ16] Alexander Schaum and Roberto Bernal Jaquez. Estimating the state probability distribution for epidemic spreading in complex networks. Applied Mathematics and Computation, 291(??):197–206, December 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316304088.

Shivanian:2018:IPI

[SJ18] Elyas Shivanian and Ahmad Jafarabadi. An inverse problem of identifying the control function in two and three-dimensional parabolic equations through the spectral meshless radial point interpolation. Applied Mathematics and Computation, 325(??):82–101, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308871.

Sanga:2019:COA

[SJ19a] Sudeep Singh Sanga and Madhu Jain. Cost optimization and ANFIS computing for admission control of M/M/1/K queue with general retrial times and discouragement. Applied Mathematics and Computation, 363 (??):Article 124624, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306162.

Sanga:2019:FFD

[SJ19b] Sudeep Singh Sanga and Madhu Jain. FM/FM/1 double orbit retrial queue with customers’ joining strategy: a parametric nonlinear programing approach. Applied Mathematics and Computation, 362(??): Article 124542, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305259.

Sakkaravarthi:2018:LSA

[SJD^{+}18] K. Sakkaravarthi, A. G.

Johnpillai, A. Durga Devi, T. Kanna, and M. Lakshmanan. Lie symmetry analysis and group invariant solutions of the nonlinear Helmholtz equation. Applied Mathematics and Computation, 331(??):457–472, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301760.

Su:2016:FTS

[SJL16] Qingyu Su, Xiaolong Jia, and Honghai Liu. Finitetime stabilization of a class of cascade nonlinear switched systems under state-dependent switching. Applied Mathematics and Computation, 289(??):172–180, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303083.

Sun:2017:TRC

[SJL17] Yuefang Sun, Zemin Jin, and Fengwei Li. On total rainbow k-connected graphs. Applied Mathematics and Computation, 311(??):223–227, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317303089.

Si:2016:DCF

[SJW16] Zhiyong Si, Shujie Jing, and Yunxia Wang. Defect correction finite element method for the stationary incompressible magnetohydrodynamics equation. Applied Mathematics and Computation, 285(??):184–194, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302235.

Sakthivel:2018:FTF

[SJWK18] R. Sakthivel, Maya Joby, Chao Wang, and B. Kaviarasan. Finite-time fault-tolerant control of neutral systems against actuator saturation and nonlinear actuator faults. Applied Mathematics and Computation, 332(??):425–436, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302625.

Safdari:2015:LBS

[SK15] Arman Safdari and Kyung Chun Kim. Lattice Boltzmann simulation of the threedimensional motions of particles with various density ratios in lid-driven cavity flow. Applied Mathematics and Computation, 265(??):826–843, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007377.

Shiralashetti:2017:TSC

[SK17a] S. C. Shiralashetti and S. Kumbinarasaiah. Theoretical study on continuous polynomial wavelet bases through wavelet series collocation method for nonlinear Lane–Emden type equations. Applied Mathematics and Computation, 315(??):591–602, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305325.

Sinelshchikov:2017:JLM

[SK17b] Dmitry I. Sinelshchikov and Nikolay A. Kudryashov. On the Jacobi last multipliers and Lagrangians for a family of Liénardtype equations. Applied Mathematics and Computation, 307(??):257–264, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317301819.

Sharma:2019:DWP

[SK19] Sandeep Sharma and Nitu Kumari. Dynamics of a waterborne pathogen model under the influence of environmental pollution. Applied Mathematics and Computation, 346(??):219–243, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830910X.

Srivastava:2016:SSC

[SKA16] H. M. Srivastava, Muhammad Rafiullah Khan, and Muhammad Arif. Some subclasses of close-to-convex mappings associated with conic regions. Applied Mathematics and Computation, 285(??):94–102, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302259.

Singh:2018:ENA

[SKBR18] Jagdev Singh, Devendra Kumar, Dumitru Baleanu, and Sushila Rathore. An efficient numerical algorithm for the fractional Drinfeld–Sokolov–Wilson equation. Applied Mathematics and Computation, 335(??):12–24, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303436.

Starzynski:2015:MAB

[SKCS15] Jacek Starzynski, Zuzanna Krawczyk, Bartosz Chaber, and Robert Szmurlo. Morphing algorithm for building individualized 3D skeleton model from CT data. Applied Mathematics and Computation, 267(??):655–663, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017755.

Singh:2018:FEM

[SKHA18] Jagdev Singh, Devendra Kumar, Zakia Hammouch, and Abdon Atangana. A fractional epidemiological model for computer viruses pertaining to a new fractional derivative. Applied Mathematics and Computation, 316(??):504–515, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317305994.

Swietlicka:2018:IGA

[ŚKK^{+}18]  Aleksandra  Świetlicka,

Krzysztof Kolanowski, Rafal Kapela, Mirosław Galicki, and Andrzej Rybarczyk. Investigation of generalization ability of a trained stochastic kinetic model of neuron. Applied Mathematics and Computation, 319(??):115–124, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300760.

Stanimirovic:2019:IPB

[SKK19] Predrag S. Stanimirović, Vasilios N. Katsikis, and Dejan Kolundzija. Inversion and pseudoinversion of block arrowhead matrices. Applied Mathematics and Computation, 341(??):379–401, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300318307781.

Sakthivel:2017:RSE

[SKKL17] R. Sakthivel, S. A. Karthick, B. Kaviarasan, and Yongdo Lim. Reliable state estimation of switched neutral system with nonlinear actuator faults via sampleddata control. Applied Mathematics and Computation, 311(??):129–147, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302941.

Skopina:2017:CMP

[Sko17] M. Skopina. On construction of multivariate Parseval wavelet frames. Applied Mathematics and Computation, 301(??):1– 11, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307391.

Stanimirovic:2016:CIU

[SKP16] Predrag S. Stanimirović, Vasilios N. Katsikis, and Dimitrios Pappas. Computing {2, 4} and {2, 3}- inverses by using the Sherman–Morrison formula. Applied Mathematics and Computation, 273(??):584–603, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S009630031501365X.

Shankar:2017:NIE

[SKS17] B. M. Shankar, Jai Kumar, and I. S. Shivakumara. Numerical investigation of electrohydrodynamic instability in a vertical porous layer. Applied Mathematics and Computation, 310(??):15–39, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302849.

Skulj:2015:ECB

[Sku15]  Damjan Skulj.  Ef-

ficient computation of the bounds of continuous time imprecise Markov chains. Applied Mathematics and Computation, 250(??):165–180, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014672.

Shao:2015:GPP

[SL15a] Xin-Hui Shao and Chen Li. A generalization of preconditioned parameterized inexact Uzawa method for indefinite saddle point problems. Applied Mathematics and Computation, 269(??):691–698, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010309.

Shi:2015:GAS

[SL15b] Hong-Bo Shi and Yan Li. Global asymptotic stability of a diffusive predator– prey model with ratiodependent functional response. Applied Mathematics and Computation, 250(??):71–77, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401491X.

Saray:2018:EMC

[SLC18] Behzad Nemati Saray, Mehrdad Lakestani, and Carlo Cattani. Evaluation of mixed Crank–Nicolson scheme and Tau method for the solution of Klein– Gordon equation. Applied Mathematics and Computation, 331(??):169–181, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301577.

Sanchez-Linares:2016:HSR

[SLdLD16] C. Sánchez-Linares, T. Morales de Luna, and M. J. Castro Díaz. A HLLC scheme for Ripa model. Applied Mathematics and Computation, 272 (part 2)(??): 369–384, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315007687.

Shen:2018:SPC

[SLF18] Shouqiang Shen, Weijun Liu, and Lihua Feng. Some properties of comaximal right ideal graph of a ring. Applied Mathematics and Computation, 333(??):225–230, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830256X.

Sun:2016:NMA

[SLL^{+}16] Mingjing Sun, Yinli Liu, Sujuan Liu, Zuoliang Hu, and Lansun Chen. A novel method for analyzing the stability of periodic solution of impulsive state feedback model. Applied Mathematics and Computation, 273(??):425–434, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013314.

Shi:2019:NFE

[SLL19] Dongyang Shi, Minghao Li, and Zhenzhen Li. A nonconforming finite element method for the stationary Smagorinsky model. Applied Mathematics and Computation, 353(??):308–319, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301109.

Shen:2019:EDA

[SLLC19] Zixiang Shen, Chuandong Li, Hongfei Li, and Zhengran Cao. Estimation of the domain of attraction for discrete-time linear impulsive control systems with input saturation. Applied Mathematics and Computation, 362 (??):Article 124502, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304813.

Shu:2019:MBC

[SLLZ19] Feng Shu, Yaojun Liu, Xingwen Liu, and Xiaobing Zhou. Memorybased conformity enhances cooperation in social dilemmas. Applied Mathematics and Computation, 346(??):480–490, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309214.

Slodicka:2015:PIS

[Slo15] M. Slodicka. A parabolic inverse source problem with a dynamical boundary condition. Applied Mathematics and Computation, 256(??):529–539, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001356.

Shen:2016:DCS

[SLS16] Chen Shen, Jun Lu, and Lei Shi. Does coevolution setup promote cooperation in spatial prisoner’s dilemma game? Applied Mathematics and Computation, 290(??):201–207, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300316303538.

Sun:2018:RCU

[SLS^{+}18] Yonghui Sun, Ning Li, Mouquan Shen, Zhinong Wei, and Guoqiang Sun. Robust H_{∞} control of uncertain linear system with interval time-varying delays by using Wirtinger inequality. Applied Mathematics and Computation, 335(??):1–11, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830345X.

Soleymani:2015:SIM

[SLTH15] F. Soleymani, T. Lotfi, E. Tavakoli, and F. Khaksar Haghani. Several iterative methods with memory using selfaccelerators. Applied Mathematics and Computation, 254(??):452–458, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000594.

Shi:2015:WES

[SLW15] Qi-Hong Shi, Wan-Tong Li, and Shu Wang. Wellposedness in energy space for the nonlinear Klein– Gordon-Schrödinger system. Applied Mathematics and Computation, 251(??):55–64, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016002.

Shi:2016:SAC

[SLW16a] Dongyang Shi, Xin Liao, and Lele Wang. Superconvergence analysis of conforming finite element method for nonlinear Schrödinger equation. Applied Mathematics and Computation, 289(??):298–310, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303290.

Shi:2016:BRP

[SLW16b] Wencong Shi, X.-F. Li, and C. Y. Wang. Bending of a rectangular plate with rotationally restrained edges under a concentrated force. Applied Mathematics and Computation, 286(??):265–278, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S009630031630279X.

Sun:2015:RSD

[SLWX15] Shiwen Sun, Ruiqi Li, Li Wang, and Chengyi Xia. Reduced synchronizability of dynamical scalefree networks with onionlike topologies. Applied Mathematics and Computation, 252(??):249–256, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016981.

Shi:2016:NDD

[SLZ^{+}16] Kaibo Shi, Xinzhi Liu, Hong Zhu, Shouming Zhong, Yong Zeng, and Chun Yin. Novel delaydependent master-slave synchronization criteria of chaotic Lur’e systems with time-varying-delay feedback control. Applied Mathematics and Computation, 282(??):137–154, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300716.

Shao:2017:NSR

[SLZ17] Hanyong Shao, Huanhuan Li, and Chuanjie Zhu. New stability results for delayed neural networks. Applied Mathematics and Computation, 311(??):324–334, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303211.

Saha:2015:MBC

[SM15a] Apu Kumar Saha and Mrinmoy Majumder. Median based conversion of SGPA into percentage by cognitive methods. Applied Mathematics and Computation, 266 (??):1153–1162, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007699.

Samidurai:2015:RPA

[SM15b] Rajendran Samidurai and Raman Manivannan. Robust passivity analysis for stochastic impulsive neural networks with leakage and additive time-varying delay components. Applied Mathematics and Computation, 268(??):743–762, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315008978.

Shamloofard:2015:DTE

[SM15c] M. Shamloofard and M. R. Movahhedy. Development of thermo-elastic tapered and spherical superelements. Applied Mathematics and Computation, 265(??):380–399, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005688.

Shaw:2015:IRF

[SM15d] Soumen Shaw and Basudeb Mukhopadhyay. An improved Regula falsi method for finding simple roots of nonlinear equations. Applied Mathematics and Computation, 254(??):370–374, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016956.

Sekar:2016:SLS

[SM16] R. Chandra Guru Sekar and K. Murugesan. System of linear second order Volterra integrodifferential equations using Single Term Walsh Series technique. Applied Mathematics and Computation, 273(??):484–492, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013302.

Sakurama:2017:DCO

[SM17] Kazunori Sakurama and Masashi Miura. Distributed constraint optimization on networked multi-agent systems. Applied Mathematics and Computation, 292(??): 272–281, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316304386.

Subburayan:2018:UNM

[SM18] V. Subburayan and R. Mahendran. An ϵ-uniform numerical method for third order singularly perturbed delay differential equations with discontinuous convection coefficient and source term. Applied Mathematics and Computation, 331(??):404–415, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302091.

Samuel:2019:HAT

[SM19] F. M. Samuel and S. S. Motsa. A highly accurate trivariate spectral collocation method of solution for two-dimensional nonlinear initial-boundary value problems. Applied Mathematics and Computation, 360(??):221–235, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319303844.

Sithole:2018:NSC

[SMG^{+}18] Hloniphile Sithole, Hiranmoy Mondal, Sicelo Goqo, Precious Sibanda, and Sandile Motsa. Numerical simulation of couple stress nanofluid flow in magneto-porous medium with thermal radiation and a chemical reaction. Applied Mathematics and Computation, 339(??):812–819, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318306076.

Sheng:2019:CCS

[SMG19] Li Sheng, Pingping Ma, and Ming Gao. H_{∞} consensus control with spectrum constraints for stochastic multi-agent systems subject to (x, u, v)- dependent noises. Applied Mathematics and Computation, 362(??): Article 124560, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305430.

Sapagovas:2018:ICC

[SMI18] M. Sapagovas, T. Meskauskas, and F. Ivanauskas. Influence of complex coefficients on the stability of difference scheme for parabolic equations with non-local conditions. Applied Mathematics and Computation, 332(??):228–240, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302534.

Shuli:2015:CME

[SML^{+}15] Guo Shuli, Irene Moroz, Han Lina, Xin Wenfang, and Feng Xianjia. Commuting matrices, equilibrium points for control systems with single saturated input. Applied Mathematics and Computation, 259(??):987–1002, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002751.

Subramanian:2018:SFS

[SML18] K. Subramanian, P. Muthukumar, and S. Lakshmanan. State feedback synchronization control of impulsive neural networks with mixed delays and linear fractional uncertainties. Applied Mathematics and Computation, 321(??):267–281, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307373.

Sierociuk:2018:ARF

[SMM18] Dominik Sierociuk, Michal Macias, and Wiktor Malesza. Analog realization of fractional variable-type and -order iterative operator. Applied Mathematics and Computation, 336(??):138–147, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303655.

Shivakumara:2015:LTN

[SMR15] I. S. Shivakumara, A. L.

Mamatha, and M. Ravisha. Local thermal non-equilibrium effects on thermal convection in a rotating anisotropic porous layer. Applied Mathematics and Computation, 259(??):838–857, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003288.

Shao:2015:OTL

[SMZ15] Jing Shao, Fanwei Meng, and Zhaowen Zheng. Oscillation theorems for linear matrix Hamiltonian systems. Applied Mathematics and Computation, 253(??):402–409, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017627.

Song:2017:ETI

[SMZ^{+}17] Xiaona Song, Yunzhe Men, Jianping Zhou, Junjie Zhao, and Hao Shen. Event-triggered H_{∞} control for networked discretetime Markov jump systems with repeated scalar nonlinearities. Applied Mathematics and Computation, 298(??):123–132, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306580.

Shah:2015:HOI

[SN15a] Farooq Ahmed Shah and Muhammad Aslam Noor. Higher order iterative schemes for nonlinear equations using decomposition technique. Applied Mathematics and Computation, 266(??):414–423, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315006785.

Shah:2015:SNM

[SN15b] Farooq Ahmed Shah and Muhammad Aslam Noor. Some numerical methods for solving nonlinear equations by using decomposition technique. Applied Mathematics and Computation, 251(??):378–386, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015975.

Siddiqi:2015:CPS

[SN15c] Shahid S. Siddiqi and Tayyaba Noreen. Convexity preservation of six point C^{2} interpolating subdivision scheme. Applied Mathematics and Computation, 265(??): 936–944, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315004750.

Singh:2018:CRE

[SN18a] Maneesh Kumar Singh and Srinivasan Natesan. Corrigendum to “Richardson extrapolation technique for singularly perturbed system of parabolic partial differential equations with exponential boundary layers” [Applied Mathematics and Computation 333 (2018) 254– 275]. Applied Mathematics and Computation, 338(??):822–827, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304995. See

[SN18b].

Singh:2018:RET

[SN18b] Maneesh Kumar Singh and Srinivasan Natesan. Richardson extrapolation technique for singularly perturbed system of parabolic partial differential equations with exponential boundary layers. Applied Mathematics and Computation, 333(??):254–275, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302327. See

corrigendum [SN18a].

Stolbunov:2018:SRB

[SN18c] Valentin Stolbunov and Prasanth B. Nair. Sparse radial basis function approximation with spatially variable shape parameters. Applied Mathematics and Computation, 330(??):170–184, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830095X.

Sedeno-Noda:2015:EKS

[SNAR15] Antonio Sedeño-Noda and Sergio Alonso-Rodríguez. An enhanced K-SP algorithm with pruning strategies to solve the constrained shortest path problem. Applied Mathematics and Computation, 265(??):602–618, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007407.

Song:2019:EHO

[SNL19] Juan Song, Yanmin Niu, and Xiong Li. The existence of homoclinic orbits in the Lorenz system via the undetermined coefficient method. Applied Mathematics and Computation, 355(??):497–515, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302048.

Saeed:2016:CAD

[SNS^{+}16] A. Saeed, Sherif S. Nafee, Salem A. Shaheen, Gehan A. Raouf, Y. Al-Hadeethi, Salahuddin M. Kamal, and M. A. N. Razvi. Calculating the ambient dose equivalent of fast neutrons using elemental composition of human body. Applied Mathematics and Computation, 274(??):604–610, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315015271.

Sowa:2018:ASS

[Sow18] Marcin Sowa. Application of SubIval in solving initial value problems with fractional derivatives. Applied Mathematics and Computation, 319(??):86–103, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300656.

Sheremet:2015:MCL

[SP15] M. A. Sheremet and I. Pop. Mixed convection in a lid-driven square cavity filled by a nanofluid: Buongiorno’s mathematical model. Applied Mathematics and Computation, 266(??):792–808, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007766.

Sahoo:2019:DPP

[SP19] Banshidhar Sahoo and Swarup Poria. Dynamics of predator–prey system with fading memory. Applied Mathematics and Computation, 347(??):319–333, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830986X.

Shindin:2017:ACT

[SPG17] Sergey Shindin, Nabendra Parumasur, and Saieshan Govinder. Analysis of a Chebyshev-type pseudospectral scheme for the nonlinear Schrödinger equation. Applied Mathematics and Computation, 307(??):271–289, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301765.

Singh:2018:PSA

[SPGP18] Jagadish Singh, A. E.

Perdiou, Jessica Mrumun Gyegwe, and E. A. Perdios. Periodic solutions around the collinear equilibrium points in the perturbed restricted threebody problem with triaxial and radiating primaries for binary HD 191408, Kruger 60 and HD 155876 systems. Applied Mathematics and Computation, 325(??):358–374, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308342.

Song:2016:SCT

[SPKC16] Yisheng Song, Khanittha Promluang, Poom Kumam, and Yeol Je Cho. Some convergence theorems of the Mann iteration for monotone α-nonexpansive mappings. Applied Mathematics and Computation, 287– 288(??):74–82, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316302533.

Shao:2017:PSC

[SPP17] Jia Shao, Apostolos D. Papaioannou, and Athanasios A. Pantelous. Pricing and simulating catastrophe risk bonds in a Markov-dependent environment. Applied Mathematics and Computation, 309(??):68–84, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302254.

Singh:2016:OMA

[SPS16] Somveer Singh, Vijay Kumar Patel, and Vineet Kumar Singh. Operational matrix approach for the solution of partial integro-differential equation. Applied Mathematics and Computation, 283(??):195–207, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301473.

Singh:2018:AWC

[SPS18] Somveer Singh, Vijay Kumar Patel, and Vineet Kumar Singh. Application of wavelet collocation method for hyperbolic partial differential equations via matrices. Applied Mathematics and Computation, 320(??):407–424, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306719.

Singh:2017:NSN

[SPST17] Somveer Singh, Vijay Kumar Patel, Vineet Kumar Singh, and Emran Tohidi. Numerical solution of nonlinear weakly singular partial integro-differential equation via operational matrices. Applied Mathematics and Computation, 298(??):310–321, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306804.

Shi:2015:ASB

[SPW15] Baoli Shi, Zhi-Feng Pang, and Jun Wu. Alternating split Bregman method for the bilaterally constrained image deblurring problem. Applied Mathematics and Computation, 250(??):402–414, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015094.

Sun:2016:EPD

[SQL16] Zhongfeng Sun, Huizeng Qin, and Aijuan Li. Extension of the partial derivatives of the incomplete beta function for complex values. Applied Mathematics and Computation, 275(??):63–71, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015568.

Sahu:2015:HLB

[SR15a] P. K. Sahu and S. Saha Ray. Hybrid Legendre Block-Pulse functions for the numerical solutions of system of nonlinear Fredholm–Hammerstein integral equations. Applied Mathematics and Computation, 270(??):871–878, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315011637.

Sahu:2015:LSC

[SR15b] P. K. Sahu and S. Saha Ray. Legendre spectral collocation method for Fredholm integrodifferential-difference equation with variable coefficients and mixed conditions. Applied Mathematics and Computation, 268(??):575–580, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008991.

Sahu:2015:LWO

[SR15c] P. K. Sahu and S. Saha Ray. Legendre wavelets operational method for the numerical solutions of nonlinear Volterra integrodifferential equations system. Applied Mathematics and Computation, 256(??):715–723, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000958.

Sanchez-Reyes:2015:CNB

[SR15d] J. Sánchez-Reyes. Comment on “Normalized Bbasis of the space of trigonometric polynomials and curve design”. Applied Mathematics and Computation, 259(??):1071, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001198. See

[Han15a].

Sepehrian:2015:NSN

[SR15e] Behnam Sepehrian and Marzieh Karimi Radpoor. Numerical solution of non-linear Fokker– Planck equation using finite differences method and the cubic spline functions. Applied Mathematics and Computation, 262(??):187–190, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003768.

Singh:2015:MCGa

[SR15f] J. K. Singh and Sarita Rani. Modified Chaplygin gas cosmology with statefinder diagnostic in Lyra geometry. Applied Mathematics and Computation, 259(??):187–197, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500243X.

Sahin:2016:DPR

[SR16a] Adem Sahin and José L.

Ramírez. Determinantal and permanental representations of convolved Lucas polynomials. Applied Mathematics and Computation, 281(??):314–322, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630073X.

Sharifi:2016:NSH

[SR16b] Shokofeh Sharifi and Jalil Rashidinia. Numerical solution of hyperbolic telegraph equation by cubic B-spline collocation method. Applied Mathematics and Computation, 281(??):28–38, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300492.

Selvi:2017:PUD

[SR17] P. Avudai Selvi and N. Ramanujam. A parameter uniform difference scheme for singularly perturbed parabolic delay differential equation with Robin type boundary condition. Applied Mathematics and Computation, 296(??):101–115, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306257.

Sohrabi:2019:SDS

[SR19] S. Sohrabi and H. Ranjbar. On sinc discretization for systems of Volterra integral-algebraic equations. Applied Mathematics and Computation, 346(??):193–204, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308920.

Silva:2016:MMC

[SRAC16] César M. Silva, Silvério Rosa, Helena Alves, and Pedro G. Carvalho. A mathematical model for the customer dynamics based on marketing policy. Applied Mathematics and Computation, 273(??):42–53, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012886.

Streeter:2015:FFA

[SRBA15] M. Streeter, L. Rhode-Barbarigos, and S. Adriaenssens. Form finding and analysis of inflatable dams using dynamic relaxation. Applied Mathematics and Computation, 267(??):742–749, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017081.

Sivaranjani:2017:SNS

[SRCA17] K. Sivaranjani, R. Rakkiyappan, Jinde Cao, and Ahmed Alsaedi. Synchronization of nonlinear singularly perturbed complex networks with uncertain inner coupling via event triggered control. Applied Mathematics and Computation, 311(??):283–299, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302953.

Stephen:2015:RPD

[SRGW15] Sudeep Stephen, Bharati Rajan, Cyriac Grigorious, and Albert William. Resolving-power dominating sets. Applied Mathematics and Computation, 256(??):778–785, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500051X.

Shams:2015:CGS

[SRH15] Masumeh Shams, Esmat Rashedi, and Ahmad Hakimi. Clusteredgravitational search algorithm and its application in parameter optimization of a low noise amplifier. Applied Mathematics and Computation, 258(??):436–453, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315001733.

Sardanyes:2015:AEI

[SRJ^{+}15] Josep Sardanyés, Carla Rodrigues, Cristina Januário, Nuno Martins, Gabriel Gil-Gómez, and Jorge Duarte. Activation of effector immune cells promotes tumor stochastic extinction: a homotopy analysis approach. Applied Mathematics and Computation, 252(??):484–495, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016506.

Skvortsov:2018:PPE

[SRK18] Alex Skvortsov, Branko Ristic, and Alex Kamenev. Predicting population extinction from early observations of the Lotka– Volterra system. Applied Mathematics and Computation, 320(??):371–379, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306574.

Sigalotti:2019:NIC

[SRK^{+}19] Leonardo Di G. Sigalotti, Otto Rendón, Jaime Klapp, Carlos A. Vargas, and Fidel Cruz. A new insight into the consistency of the SPH interpolation formula. Applied Mathematics and Computation, 356(??):50–73, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930219X.

Sheremet:2017:FCP

[SRP17] Mikhail A. Sheremet, Cornelia Revnic, and Ioan Pop. Free convection in a porous wavy cavity filled with a nanofluid using Buongiorno’s mathematical model with thermal dispersion effect. Applied Mathematics and Computation, 299(??):1–15, April 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307081.

Sohrabi:2017:CAJ

[SRS17] S. Sohrabi, H. Ranjbar, and M. Saei. Convergence analysis of the Jacobicollocation method for nonlinear weakly singular Volterra integral equations. Applied Mathematics and Computation, 299(??):141–152, April 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306907.

Sakthivel:2015:DRC

[SRSZ15] R. Sakthivel, M. Rathika, Srimanta Santra, and Quanxin Zhu. Dissipative reliable controller design for uncertain systems and its application. Applied Mathematics and Computation, 263(??):107–121, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004609.

Sumenjak:2018:KRI

[SRT18] Tadeja Kraner Sumenjak, Douglas F. Rall, and Aleksandra Tepeh. On krainbow independent domination in graphs. Applied Mathematics and Computation, 333(??):353–361, September 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318302947.

Sendra:2015:GBC

[SS15a] Juana Sendra and J. Rafael Sendra. Gröbner basis computation of Drazin inverses with multivariate rational function entries. Applied Mathematics and Computation, 259(??):450–459, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002702.

Shpot:2015:CHF

[SS15b] M. A. Shpot and H. M. Srivastava. The Clausenian hypergeometric function _{3}F_{2} with unit argument and negative integral parameter differences. Applied Mathematics and Computation, 259(??):819–827, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003367.

Su:2015:MIP

[SS15c] Lei Su and Hao Shen. Mixed H_{∞}/passive synchronization for complex dynamical networks with sampled-data control. Applied Mathematics and Computation, 259(??):931–942, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003380.

Shu:2016:SMS

[SS16] Xiao-Bao Shu and Yajing Shi. A study on the mild solution of impulsive fractional evolution equations. Applied Mathematics and Computation, 273(??):465–476, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013521.

Sendra:2017:CMP

[SS17] J. Rafael Sendra and Juana Sendra. Computation of Moore–Penrose generalized inverses of matrices with meromorphic function entries. Applied Mathematics and Computation, 313(??):355–366, November 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317304198.

Saadatmandi:2018:CMB

[SS18a] Abbas Saadatmandi and Zeinab Sanatkar. Collocation method based on rational Legendre functions for solving the magnetohydrodynamic flow over a nonlinear stretching sheet. Applied Mathematics and Computation, 323(??):193–203, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308494.

Satapathy:2018:OSI

[SS18b] Purnima Satapathy and T. Raja Sekhar. Optimal system, invariant solutions and evolution of weak discontinuity for isentropic drift flux model. Applied Mathematics and Computation, 334(??):107–116, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302959.

Shankar:2018:MSP

[SS18c] B. M. Shankar and I. S. Shivakumara. Magnetohydrodynamic stability of pressure-driven flow in an anisotropic porous channel: Accurate solution. Applied Mathematics and Computation, 321(??):752–767, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317307798.

Sheng:2019:SMM

[SS19a] Yuhong Sheng and Gang Shi. Stability in mean of multi-dimensional uncertain differential equation. Applied Mathematics and Computation, 353(??):178–188, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301067.

Shi:2019:NAF

[SS19b] Xuanxuan Shi and Mouquan Shen. A new approach to feedback feedforward iterative learning control with random packet dropouts. Applied Mathematics and Computation, 348(??):399–412, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031831035X.

Sharma:2018:NSS

[SSA18] Vaishali Sharma, Amit Setia, and Ravi P. Agarwal. Numerical solution for system of Cauchy type singular integral equations with its error analysis in complex plane. Applied Mathematics and Computation, 328(??):338–352, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300298.

Sen:2019:DFP

[SSA19] Mausumi Sen, Dipankar Saha, and R. P. Agarwal. A Darbo fixed point theory approach towards the existence of a functional integral equation in a Banach algebra. Applied Mathematics and Computation, 358(??):111–118, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303030.

Sen:2015:GDA

[SSB15a] Moitri Sen, P. D. N. Srinivasu, and Malay Banerjee. Global dynamics of an additional food provided predator–prey system with constant harvest in predators. Applied Mathematics and Computation, 250(??):193–211, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S009630031401460X.

Singh:2015:MCGb

[SSB15b] J. K. Singh, N. K. Sharma, and A. Beesham. Modified Chaplygin gas cosmology with observational constraints. Applied Mathematics and Computation, 270(??):567–581, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501111X.

Sharma:2016:INT

[SSB16] Janak Raj Sharma, Rajni Sharma, and Ashu Bahl. An improved Newton–Traub composition for solving systems of nonlinear equations. Applied Mathematics and Computation, 290(??):98–110, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303691.

Sayyed:2018:ASM

[SSB18] S. R. Sayyed, B. B.

Singh, and Nasreen Bano. Analytical solution of MHD slip flow past a constant wedge within a porous medium using DTM–Padé. Applied Mathematics and Computation, 321(??):472–482, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307713.

Shao:2018:DCC

[SSCX18] Jinliang Shao, Lei Shi, Mengtao Cao, and Hong Xia. Distributed containment control for asynchronous discrete-time second-order multi-agent systems with switching topologies. Applied Mathematics and Computation, 336(??):47–59, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830393X.

Shankar:2015:LLW

[SSG^{+}15a] Ravi Shankar, Yan Sheng, Megan Golbek, Tucker Hartland, Peter Gerrodette, Sergei Fomin, and Vladimir Chugunov. Linear long wave propagation over discontinuous submerged shallow water topography. Applied Mathematics and Computation, 252(??):27–44, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015574.

Siddiqui:2015:BAS

[SSG15b] S. U. Siddiqui, S. R. Shah, and Geeta. A biomechanical approach to study the effect of body acceleration and slip velocity through stenotic artery. Applied Mathematics and Computation, 261(??):148–155, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003963. See

corrigendum [Pon17].

Sharomi:2017:ERI

[SSGG17] Oluwaseun Y. Sharomi, Mohammad A. Safi, Abba B. Gumel, and David J. Gerberry. Exogenous reinfection does not always cause backward bifurcation in TB transmission dynamics. Applied Mathematics and Computation, 298(??):322–335, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306774.

Song:2015:NSS

[SSJ15] Shicang Song, Ming Sun, and Liying Jiang. A nonconforming scheme to solve the parabolic problem. Applied Mathematics and Computation, 265(??):108–119, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005512.

Sharma:2015:NFC

[SSK15] Janak Raj Sharma, Rajni Sharma, and Nitin Kalra. A novel family of composite Newton-traub methods for solving systems of nonlinear equations. Applied Mathematics and Computation, 269(??):520–535, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010140.

Saha:2019:IAB

[SSK^{+}19] Tanay Saha, Shwetabh Srivastava, Swanand Khare, Predrag S. Stanimirović, and Marko D. Petković. An improved algorithm for basis pursuit problem and its applications. Applied Mathematics and Computation, 355(??):385–398, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301833.

Sakthivel:2016:DBR

[SSKA16] R. Sakthivel, T. Saravanakumar, B. Kaviarasan, and S. Marshal Anthoni. Dissipativity based repetitive control for switched stochastic dynamical systems. Applied Mathematics and Computation, 291(??):340–353, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304490.

Sierociuk:2015:DPM

[SSM^{+}15] Dominik Sierociuk, Tomas Skovranek, Michal Macias, Igor Podlubny, Ivo Petras, Andrzej Dzielinski, and Pawel Ziubinski. Diffusion process modeling by using fractionalorder models. Applied Mathematics and Computation, 257(??):2–11, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015501.

Siddiqui:2017:ACT

[SSM17] A. M. Siddiqui, Ayesha Sohail, and Khadija Maqbool. Analysis of a channel and tube flow induced by cilia. Applied Mathematics and Computation, 309(??):133–141, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302436.

Soares:2018:SIP

[SSM^{+}18] Helena Soares, Tiago Neves Sequeira, Pedro Macias Marques, Orlando Gomes, and Alexandra Ferreira-Lopes. Social infrastructure and the preservation of physical capital: Equilibria and transitional dynamics. Applied Mathematics and Computation, 321(??):614–632, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307658.

Seyedi:2015:UIS

[SSN15] S. H. Seyedi, B. N. Saray, and M. R. H. Nobari. Using interpolation scaling functions based on Galerkin method for solving non-Newtonian fluid flow between two vertical flat plates. Applied Mathematics and Computation, 269(??):488–496, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010218.

Soriano-Sanchez:2016:CSC

[SSPCPG^{+}16] A. G. Soriano-Sánchez, C. Posadas-Castillo, M. A. Platas-Garza, C. Cruz-Hernández, and R. M. López-Gutiérrez. Coupling strength computation for chaotic synchronization of complex networks with multi-scroll attractors. Applied Mathematics and Computation, 275(??):305–316, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015830.

Soltani:2015:RCO

[SSS15] Roya Soltani, Jalal Safari, and Seyed Jafar Sadjadi. Robust counterpart optimization for the redundancy allocation problem in seriesparallel systems with component mixing under uncertainty. Applied Mathematics and Computation, 271(??):80–88, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501125X.

Sengupta:2017:GSA

[SSS17] Tapan K. Sengupta, Aditi Sengupta, and Kumar Saurabh. Global spectral analysis of multi-level time integration schemes: Numerical properties for error analysis. Applied Mathematics and Computation, 304(??):41–57, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300346.

Sbibih:2015:SLQ

[SST15] D. Sbibih, A. Serghini, and A. Tijini. Superconvergent local quasiinterpolants based on special multivariate quadratic spline space over a refined quadrangulation. Applied Mathematics and Computation, 250(??):145–156, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014659.

Saadatmandi:2017:CMS

[SST17] Abbas Saadatmandi, Zeinab Sanatkar, and Seyed Pendar Toufighi. Computational methods for solving the steady flow of a third grade fluid in a porous half space. Applied Mathematics and Computation, 298(??):133–140, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306865.

Schieweck:2018:COE

[SST18] F. Schieweck, P. Skrzypacz, and L. Tobiska. Construction of L^{2}-orthogonal elements of arbitrary order for Local Projection Stabilization. Applied Mathematics and Computation, 337(??):76–86, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303965.

Shan:2018:ESE

[SSZ^{+}18] Yaonan Shan, Kun She, Shouming Zhong, Qishui Zhong, Kaibo Shi, and Can Zhao. Exponential stability and extended dissipativity criteria for generalized discrete-time neural networks with additive time-varying delays. Applied Mathematics and Computation, 333(??):145–168, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302820.

Sun:2017:DBS

[SSZL17] Shulin Sun, Yaru Sun, Guang Zhang, and Xinzhi Liu. Dynamical behavior of a stochastic two-species Monod competition chemostat model. Applied Mathematics and Computation, 298(??):153–170, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306634.

Stampfli:2017:BGC

[Sta17] Marx Stampfli. Bridged graphs, circuits and Fibonacci numbers. Applied Mathematics and Computation, 302(??):68–79, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307640.

Simmonds:2018:LMN

[STA^{+}18] Nicholas Simmonds, Panagiotis Tsoutsanis, Antonis F. Antoniadis, Karl W. Jenkins, and Adrian Gaylard. Low-Mach number treatment for finitevolume schemes on unstructured meshes. Applied Mathematics and Computation, 336(??): 368–393, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318304028.

Sarikaya:2016:GSI

[STB16] Mehmet Zeki Sarikaya, Tuba Tunc, and Hüseyin Budak. On generalized some integral inequalities for local fractional integrals. Applied Mathematics and Computation, 276(??):316–323, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315016203.

Su:2015:GFN

[STD15] Guifu Su, Jianhua Tu, and Kinkar Ch. Das. Graphs with fixed number of pendent vertices and minimal zeroth-order general Randić index. Applied Mathematics and Computation, 270(??):705–710, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315011157.

Stepanova:2019:GAV

[Ste19] Irina V. Stepanova. Group analysis of variable coefficients heat and mass transfer equations with power nonlinearity of thermal diffusivity. Applied Mathematics and Computation, 343(??):57–66, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308142.

Safuan:2016:TWS

[STJS16] Hamizah M. Safuan, Isaac N. Towers, Zlatko Jovanoski, and Harvinder S. Sidhu. On travelling wave solutions of the diffusive Leslie–Gower model. Applied Mathematics and Computation, 274(??): 362–371, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315014393.

Soltys:2015:AWI

[STK15] Robert Soltys, Michal Tomko, and Stanislav Kmet. Analysis of windinduced vibrations of an anchor cable using a simplified fluid-structure interaction method. Applied Mathematics and Computation, 267(??):223–236, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315004336.

Sinuvasan:2017:STW

[STL17] R. Sinuvasan, K. M.

Tamizhmani, and P. G. L. Leach. Symmetries, travelling-wave and selfsimilar solutions of the Burgers hierarchy. Applied Mathematics and Computation, 303(??):165–170, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300504.

Set:2015:GGT

[STS15] Erhan Set, Muharrem Tomar, and Mehmet Zeki Sarikaya. On generalized Grüss type inequalities for k-fractional integrals. Applied Mathematics and Computation, 269(??):29–34, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009388.

Spiller:2017:ESR

[STVCC17] C. Spiller, E. F. Toro, M. E. Vázquez-Cendón, and C. Contarino. On the exact solution of the Riemann problem for blood flow in human veins, including collapse. Applied Mathematics and Computation, 303(??):178–189, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300322.

Su:2016:UMS

[Su16] Qifang Su. A unified model for solving a system of nonlinear equations. Applied Mathematics and Computation, 290(??):46–55, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303630.

Su:2019:RBF

[Su19] LingDe Su. A radial basis function (RBF)-finite difference (FD) method for the backward heat conduction problem. Applied Mathematics and Computation, 354(??):232–247, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301390.

Sugie:2017:NTS

[Sug17] Jitsuro Sugie. Nonoscillation theorems for secondorder linear difference equations via the Riccatitype transformation, II. Applied Mathematics and Computation, 304(??): 142–152, July 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317300668.

Sukhorukova:2019:TCC

[Suk19] N. Sukhorukova. Two curve Chebyshev approximation and its application to signal clustering. Applied Mathematics and Computation, 356(??):42–49, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S009630031930222X.

Sun:2015:MLC

[Sun15a] Xianbo Sun. Multiple limit cycles of some strongly nonlinear Liénard–van der Pol oscillator. Applied Mathematics and Computation, 270(??):620–630, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315011042.

Sun:2015:NET

[Sun15b] Yunchuan Sun. New exact traveling wave solutions for double sine-Gordon equation. Applied Mathematics and Computation, 258(??):100–104, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500154X.

Sun:2018:ECL

[Sun18] Baoyan Sun. Exponential convergence for the linear homogeneous Boltzmann equation for hard potentials. Applied Mathematics and Computation, 339(??):701–726, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306167.

Saeed:2015:HWP

[SuR15] Umer Saeed and Mujeeb ur Rehman. Haar wavelet Picard method for fractional nonlinear partial differential equations. Applied Mathematics and Computation, 264(??):310–322, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005585.

Saeed:2015:MCW

[SuRI15] Umer Saeed, Mujeeb ur Rehman, and Muhammad Asad Iqbal. Modified Chebyshev wavelet methods for fractional delaytype equations. Applied Mathematics and Computation, 264(??):431–442, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005755.

Siddiqi:2015:FBT

[SuSB15] Shahid S. Siddiqi, Wardat us Salam, and Nadeem Ahmad Butt. Fractal behavior of ternary 4point interpolatory subdivision scheme with tension parameter. Applied Mathematics and Computation, 260(??):148–158, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003616.

Siddiqi:2015:BPP

[SuSR15] Shahid S. Siddiqi, Wardat us Salam, and Kashif Rehan. Binary 3-point and 4point non-stationary subdivision schemes using hyperbolic function. Applied Mathematics and Computation, 258(??):120–129, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500123X.

Siddiqi:2016:CBF

[SuSR16] Shahid S. Siddiqi, Wardat us Salam, and Kashif Rehan. Construction of binary four and five point non-stationary subdivision schemes from hyperbolic B-splines. Applied Mathematics and Computation, 280(??):30–38, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300200.

Sommariva:2015:PFI

[SV15a] Alvise Sommariva and Marco Vianello. Polynomial fitting and interpolation on circular sections. Applied Mathematics and Computation, 258(??):410–424, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001666.

Sucu:2015:GJI

[SV15b] Sezgin Sucu and Serhan Varma. Generalization of Jakimovski– Leviatan type Szasz operators. Applied Mathematics and Computation, 270(??):977–983, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011339.

Sheikholeslami:2017:NFH

[SV17a] M. Sheikholeslami and K. Vajravelu. Nanofluid flow and heat transfer in a cavity with variable magnetic field. Applied Mathematics and Computation, 298(??):272–282, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306932.

Smirnov:2017:NCW

[SV17b] Yury G. Smirnov and Dmitry V. Valovik. Nonlinear coupled wave propagation in a n-dimensional layer. Applied Mathematics and Computation, 294(??):146–156, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305719.

Shi:2017:USAb

[SW17a] Dongyang Shi and Junjun Wang. Unconditional superconvergence analysis for nonlinear hyperbolic equation with nonconforming finite element. Applied Mathematics and Computation, 305(??):1– 16, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300681.

Shi:2017:USAa

[SW17b] Dongyang Shi and Junjun Wang. Unconditional superconvergence analysis of conforming finite element for nonlinear parabolic equation. Applied Mathematics and Computation, 294(??):216–226, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305197.

Sun:2017:IEF

[SW17c] FengXin Sun and JuFeng Wang. Interpolating element-free Galerkin method for the regularized long wave equation and its error analysis. Applied Mathematics and Computation, 315(??):54–69, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305131.

Suandi:2019:OLM

[SWA^{+}19] Dani Suandi, Karunia Putra Wijaya, Mochamad Apri, Kuntjoro Adji Sidarto, Din Syafruddin, Thomas Götz, and Edy Soewono. A one-locus model describing the evolutionary dynamics of resistance against insecticide in anopheles mosquitoes. Applied Mathematics and Computation, 359(??):90–106, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319302322.

Shao:2017:MAI

[SWG^{+}17] Zehui Shao, Pu Wu, Yingying Gao, Ivan Gutman, and Xiujun Zhang. On the maximum ABC index of graphs without pendent vertices. Applied Mathematics and Computation, 315(??):298–312, December 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317305362.

Swietlicka:2015:TSM

[Świ15]  Aleksandra  Świetlicka.

Trained stochastic model of biological neural network used in image processing task. Applied Mathematics and Computation, 267(??):716–726, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300314017366.

Shi:2016:NNM

[SWL16] Dongyang Shi, Lele Wang, and Xin Liao. A new nonconforming mixed finite element scheme for second order eigenvalue problem. Applied Mathematics and Computation, 273(??):842–855, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014150.

Skrzypczak:2018:NMS

[SWSS18] Tomasz Skrzypczak, Ewa Wegrzyn-Skrzypczak, and Leszek Sowa. Numerical modeling of solidification process taking into account the effect of air gap. Applied Mathematics and Computation, 321(??):768–779, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308056.

Song:2019:QOF

[SWSW19] Xiaona Song, Mi Wang, Shuai Song, and Zhen Wang. Quantized output feedback control for nonlinear Markovian jump distributed parameter systems with unreliable communication links. Applied Mathematics and Computation, 353(??):371–395, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300840.

Sabir:2019:SNA

[SWUE19] Zulqurnain Sabir, Hafiz Abdul Wahab, Muhammad Umar, and Fevzi Erdogan. Stochastic numerical approach for solving second order nonlinear singular functional differential equation. Applied Mathematics and Computation, 363(??): Article 124605, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305971.

Song:2018:CRS

[SWY18] Guang-Jing Song, Qing-Wen Wang, and Shao-Wen Yu. Cramer’s rule for a system of quaternion matrix equations with applications. Applied Mathematics and Computation, 336(??):490–499, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303746.

Shi:2019:NRN

[SWZ^{+}19] Kaibo Shi, Jun Wang, Shouming Zhong, Xiaojun Zhang, Yajuan Liu, and Jun Cheng. New reliable nonuniform sampling control for uncertain chaotic neural networks under Markov switching topologies. Applied Mathematics and Computation, 347(??):169–193, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309846.

Sun:2019:FNMa

[SWZL19] Fenglan Sun, Rui Wang, Wei Zhu, and Yongfu Li. Flocking in nonlinear multi-agent systems with time-varying delay via event-triggered control. Applied Mathematics and Computation, 350(??):66– 77, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311172.

Su:2015:TIL

[SX15] Guifu Su and Lan Xu.

Topological indices of the line graph of subdivision graphs and their Schurbounds. Applied Mathematics and Computation, 253(??):395–401, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014283.

Sun:2016:SSD

[SX16] Taixiang Sun and Hongjian Xi. On the solutions of a system of difference equations with maximum. Applied Mathematics and Computation, 290(??):292–297, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303915.

Shu:2015:APP

[SXS15] Xiao-Bao Shu, Fei Xu, and Yajing Shi. Sasymptotically ω-positive periodic solutions for a class of neutral fractional differential equations. Applied Mathematics and Computation, 270(??):768–776, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011364.

Sharma:2015:SRD

[SY15a] Mithun J. Sharma and Song Jin Yu. Stepwise regression data envelopment analysis for variable reduction. Applied Mathematics and Computation, 253(??):126–134, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017044.

Song:2015:ICS

[SY15b] Jiamei Song and Jun Yue.

Injective coloring of some graph operations. Applied Mathematics and Computation, 264(??):279–283, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004439.

Shen:2017:FFA

[SY17a] Mouquan Shen and Dan Ye. A finite frequency approach to control of Markov jump linear systems with incomplete transition probabilities. Applied Mathematics and Computation, 295(??):53– 64, February 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031630594X.

Shi:2017:UOE

[SY17b] Dongyang Shi and Huaijun Yang. Unconditional optimal error estimates of a two-grid method for semilinear parabolic equation. Applied Mathematics and Computation, 310(??):40–47, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302540.

Shi:2018:RCC

[SY18a] Chong-Xiao Shi and Guang-Hong Yang. Robust consensus control for a class of multi-agent systems via distributed PID algorithm and weighted edge dynamics. Applied Mathematics and Computation, 316(??):73–88, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305301.

Shi:2018:SAF

[SY18b] Dongyang Shi and Huaijun Yang. Superconvergence analysis of finite element method for time-fractional thermistor problem. Applied Mathematics and Computation, 323(??):31–42, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308093.

Song:2018:NDR

[SY18c] Guangjing Song and Shaowen Yu. Nonnegative definite and Re-nonnegative definite solutions to a system of matrix equations with statistical applications. Applied Mathematics and Computation, 338(??):809–821, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305344.

Shi:2019:SAN

[SY19] Dongyang Shi and Huaijun Yang. Superconvergence analysis of nonconforming FEM for nonlinear time-dependent thermistor problem. Applied Mathematics and Computation, 347(??):210–224, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318308841.

Shen:2016:FTS

[SYZP16] Mouquan Shen, Shen Yan, Guangming Zhang, and Ju H. Park. Finitetime H_{∞} static output control of Markov jump systems with an auxiliary approach. Applied Mathematics and Computation, 273(??):553–561, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013892.

Shao:2015:DDS

[SZ15] Hanyong Shao and Zhengqiang Zhang. Delay-dependent state feedback stabilization for a networked control model with two additive input delays. Applied Mathematics and Computation, 265(??):748–758, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007456.

Sun:2016:PCH

[SZ16] Xianbo Sun and Liqin Zhao. Perturbations of a class of hyperelliptic Hamiltonian systems of degree seven with nilpotent singular points. Applied Mathematics and Computation, 289(??):194–203, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302685.

Song:2017:CQH

[SZ17] Chuan-Jing Song and Yi Zhang. Conserved quantities for Hamiltonian systems on time scales. Applied Mathematics and Computation, 313(??):24–36, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303867.

SanAntolin:2018:CSP

[SZ18a] A. San Antolín and R. A.

Zalik. Compactly supported Parseval framelets with symmetry associated to Ed(2)(Z) matrices. Applied Mathematics and Computation, 325(??):179–190, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308676.

Sheng:2018:CRS

[SZ18b] Baohuai Sheng and Hancan Zhu. The convergence rate of semisupervised regression with quadratic loss. Applied Mathematics and Computation, 321(??):11–24, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307324.

Sha:2019:NNR

[SZ19a] Chunlin Sha and Hongyong Zhao. A novel neurodynamic reactiondiffusion model for solving linear variational inequality problems and its application. Applied Mathematics and Computation, 346(??):57–75, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308890.

Shah:2019:EUS

[SZ19b] Syed Omar Shah and Akbar Zada. Existence, uniqueness and stability of solution to mixed integral dynamic systems with instantaneous and noninstantaneous impulses on time scales. Applied Mathematics and Computation, 359(??):202–213, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303327.

Shi:2019:CCS

[SZ19c] Lingjuan Shi and Heping Zhang. Counting Clar structures of (4, 6)-fullerenes. Applied Mathematics and Computation, 346(??):559–574, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308932.

Szablowski:2015:FRO

[Sza15] Pawel J. Szablowski. A few remarks on orthogonal polynomials. Applied Mathematics and Computation, 252(??):215–228, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016452.

Song:2017:PDS

[SZFS17] Li-Peng Song, Rong-Ping Zhang, Li-Ping Feng, and Qiong Shi. Pattern dynamics of a spatial epidemic model with time delay. Applied Mathematics and Computation, 292(??):390–399, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630443X.

Shi:2019:DAI

[SZJ19] Zhenfeng Shi, Xinhong Zhang, and Daqing Jiang. Dynamics of an avian influenza model with halfsaturated incidence. Applied Mathematics and Computation, 355(??): 399–416, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319301808.

Shi:2018:NQW

[SZL^{+}18] Z. G. Shi, Y. M. Zhao, F. Liu, F. L. Wang, and Y. F. Tang. Nonconforming quasi-Wilson finite element method for 2D multi-term time fractional diffusion-wave equation on regular and anisotropic meshes. Applied Mathematics and Computation, 338(??):274–289, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305149.

Shumylyak:2016:MIS

[SZO16] Liliya Shumylyak, Vladimir Zhikharevich, and Sergiy Ostapov. Modeling of impurities segregation phenomenon in the melt crystallization process by the continuous cellular automata technique. Applied Mathematics and Computation, 290(??):336–354, November 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316303836.

Song:2016:FSS

[SZPH16] Bo Song, Ya Zhang, Ju H. Park, and Huan Huang. L_{2}–L_{∞} filtering for stochastic systems driven by Poisson processes and Wiener processes. Applied Mathematics and Computation, 276(??):407–416, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031530014X.

Sun:2015:ODW

[SZRC15] Xuemei Sun, Yiming Zhang, Xu Ren, and Ke Chen. Optimization deployment of wireless sensor networks based on culture-ant colony algorithm. Applied Mathematics and Computation, 250(??):58–70, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014660.

Shen:2019:FDS

[SzSC19] Jinye Shen, Zhi zhong Sun, and Wanrong Cao. A finite difference scheme on graded meshes for time-fractional nonlinear Korteweg–de Vries equation. Applied Mathematics and Computation, 361(??):752–765, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304886.

Sun:2016:SHO

[SzShG16] Hong Sun, Zhi zhong Sun, and Guang hua Gao. Some high order difference schemes for the space and time fractional Bloch– Torrey equations. Applied Mathematics and Computation, 281(??):356–380, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300443.

Sun:2017:DAC

[SZT17] Kaibiao Sun, Tonghua Zhang, and Yuan Tian. Dynamics analysis and control optimization of a pest management predator– prey model with an integrated control strategy. Applied Mathematics and Computation, 292(??):253–271, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304842.

Tumuluri:2016:FTO

[TA16] Suman Kumar Tumuluri and T. Amaranath. A factorization theorem for operators occurring in the Stokes, Brinkman and Oseen equations. Applied Mathematics and Computation, 276(??):75–79, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315016148.

Talati:2015:CIF

[Tal15] Daryoush Talati. Complete integrability of the fifth-order Mikhailov-Novikov-Wang system. Applied Mathematics and Computation, 250(??):776–778, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401501X.

Tang:2015:ESG

[Tan15a] Xiaojun Tang. Efficient and stable generation of higher-order pseudospectral integration matrices. Applied Mathematics and Computation, 261(??):60– 67, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500404X.

Tang:2015:ECC

[Tan15b]  Xiaojun Tang.  Effi-

cient Chebyshev collocation methods for solving optimal control problems governed by Volterra integral equations. Applied Mathematics and Computation, 269(??):118–128, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009741.

Tang:2015:NSV

[Tan15c] Xiaojun Tang. Numerical solution of Volterra– Fredholm integral equations using parameterized pseudospectral integration matrices. Applied Mathematics and Computation, 270(??):744–755, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011327.

Tang:2015:GAS

[Tan15d] Yi Tang. Global asymptotic stability for Nicholson’s blowflies model with a nonlinear densitydependent mortality term. Applied Mathematics and Computation, 250(??): 854–859, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314014908.

Tanimoto:2015:IIC

[Tan15e] Jun Tanimoto. The impact of initial cooperation fraction on the evolutionary fate in a spatial prisoner’s dilemma game. Applied Mathematics and Computation, 263(??):171–188, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005019.

Tanimoto:2017:CDM

[Tan17a] Jun Tanimoto. Coevolution of discrete, mixed, and continuous strategy systems boosts in the spatial prisoner’s dilemma and chicken games. Applied Mathematics and Computation, 304(??):20– 27, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300231.

Tanimoto:2017:HDR

[Tan17b]  Jun Tanimoto.  How

does resolution of strategy affect network reciprocity in spatial prisoner’s dilemma games? Applied Mathematics and Computation, 301(??):36– 42, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316307123.

Tang:2018:NCS

[Tan18] Wensheng Tang. A note on continuous-stage Runge– Kutta methods. Applied Mathematics and Computation, 339(??):220–230, December 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031830609X.

Tarasov:2015:LFC

[Tar15] Vasily E. Tarasov. Lattice fractional calculus. Applied Mathematics and Computation, 257(??):12– 33, April 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314015562.

Toloo:2015:VRD

[TB15] Mehdi Toloo and Seddigheh Babaee. On variable reductions in data envelopment analysis with an illustrative application to a gas company. Applied Mathematics and Computation, 270(??):527–533, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009030.

Tamilalagan:2017:MSR

[TB17a] P. Tamilalagan and P. Balasubramaniam. Moment stability via resolvent operators of fractional stochastic differential inclusions driven by fractional Brownian motion. Applied Mathematics and Computation, 305(??):299–307, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301170.

Toutounian:2017:NMC

[TB17b] F. Toutounian and R. Buzhabadi. New methods for computing the Drazin-inverse solution of singular linear systems. Applied Mathematics and Computation, 294(??):343–352, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305732.

Tracina:2016:NSA

[TBG16] R. Tracinà, M. S. Bruzón, and M. L. Gandarias. On the nonlinear selfadjointness of a class of fourth-order evolution equations. Applied Mathematics and Computation, 275(??):299–304, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015817.

Totorean:2016:FDH

[TBSR16] Alin F. Totorean, Sandor I. Bernad, and Romeo F. Susan-Resiga. Fluid dynamics in helical geometries with applications for by-pass grafts. Applied Mathematics and Computation, 272 (part 3)(??): 604–613, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315006475.

Tchoualag:2015:CSB

[Tch15] L. Tchoualag. Combined symmetric BEM and semismooth Newton method for the contact problems in linear elasticity of Yukawa type. Applied Mathematics and Computation, 269(??):1007–1021, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010838.

Tian:2015:ABS

[TCS15] Yazhou Tian, Yuanli Cai, and Yuangong Sun. Asymptotic behavior of switched delay systems with nonlinear disturbances. Applied Mathematics and Computation, 268(??):522–533, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008541.

Tian:2015:PAL

[TCY15] Qing Tian, Shujuan Cao, and Chunxiao Yang. On a problem analogues to Lehmer’s and Hurwitz zeta-function. Applied Mathematics and Computation, 268(??):636–643, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500898X.

Talaska:2018:APS

[TD18] Tomasz Talaśka and Rafal Dlugosz. Analog, parallel, sorting circuit for the application in Neural Gas learning algorithm implemented in the CMOS technology. Applied Mathematics and Computation, 319(??):218–235, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317301340.

Tian:2019:RSM

[TD19] Yong Tian and Xuejun Ding. Rumor spreading model with considering debunking behavior in emergencies. Applied Mathematics and Computation, 363(??): Article 124599, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305910.

Trindade:2017:ECA

[TDA17] Graça Trindade, José G.

Dias, and Jorge Ambrósio. Extracting clusters from aggregate panel data: a market segmentation study. Applied Mathematics and Computation, 296(??):277–288, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306105.

Tu:2017:ASM

[TDL^{+}17] Zhengwen Tu, Nan Ding, Liangliang Li, Yuming Feng, Limin Zou, and Wei Zhang. Adaptive synchronization of memristive neural networks with time-varying delays and reaction-diffusion term. Applied Mathematics and Computation, 311(??):118–128, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730293X.

Tu:2015:UGM

[TDS15] Jianhua Tu, Junfeng Du, and Guifu Su. The unicyclic graphs with maximum degree resistance distance. Applied Mathematics and Computation, 268(??):859–864, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008449.

Tepeh:2019:EBG

[Tep19] Aleksandra Tepeh. Extremal bicyclic graphs with respect to Mostar index. Applied Mathematics and Computation, 355(??):319–324, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302073.

Terry:2015:PMB

[Ter15] Alan J. Terry. A population model with birth pulses, age structure, and non-overlapping generations. Applied Mathematics and Computation, 271(??):400–417, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012242.

Tabata:2015:ECS

[TES15] Minoru Tabata, Nobuoki Eshima, and Yuusuke Sakai. Existence and computation of solutions to the initial value problem for the replicator equation of evolutionary game defined by the Dixit–Stiglitz– Krugman model in an urban setting: Concentration of workers motivated by disparity in real wages. Applied Mathematics and Computation, 254(??):419–451, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000430.

Tsitouras:2018:BVD

[TF18] Ch. Tsitouras and I. Th.

Famelis. Bounds for variable degree rational L_{∞} approximations to the matrix exponential. Applied Mathematics and Computation, 338(??):363–375, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305290.

Tan:2019:AAE

[TFLZ19] Yushun Tan, Shumin Fei, Jinliang Liu, and Dan Zhang. Asynchronous adaptive eventtriggered tracking control for multi-agent systems with stochastic actuator faults. Applied Mathematics and Computation, 355(??):482–496, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302024.

Fan:2015:GRP

[tFyZ15] Hong tao Fan and Xin yun Zhu. A generalized relaxed positive-definite and skew-Hermitian splitting preconditioner for non-Hermitian saddle point problems. Applied Mathematics and Computation, 258(??):36–48, May 1, 2015. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315001514.

Fan:2015:SCC

[tFyZZ15] Hong tao Fan, Xin yun Zhu, and Bing Zheng. On semi-convergence of a class of relaxation methods for singular saddle point problems. Applied Mathematics and Computation, 261(??):68–80, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004075.

Tang:2016:FRK

[TG16] Z. Q. Tang and F. Z. Geng.

Fitted reproducing kernel method for singularly perturbed delay initial value problems. Applied Mathematics and Computation, 284(??):169–174, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301989.

Titi:2017:MMS

[TG17] Jihad Titi and Jürgen Garloff. Matrix methods for the simplicial Bernstein representation and for the evaluation of multivariate polynomials. Applied Mathematics and Computation, 315(??):246–258, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304770.

Titi:2019:MMT

[TG19] Jihad Titi and Jürgen Garloff. Matrix methods for the tensorial Bernstein form. Applied Mathematics and Computation, 346(??):254–271, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830763X.

Turkoglu:2019:RPL

[TH19] Muammer Turkoglu and Davut Hanbay. Recognition of plant leaves: an approach with hybrid features produced by dividing leaf images into two and four parts. Applied Mathematics and Computation, 352(??):1– 14, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300319300694.

Tharwat:2015:AED

[Tha15] M. M. Tharwat. Approximation of eigenvalues of Dirac systems with eigenparameter in all boundary conditions by sinc-Gaussian method. Applied Mathematics and Computation, 262(??):113–127, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004683.

Thanh:2018:CWB

[Tha18a] Mai Duc Thanh. Completing a well-balanced numerical method for a model of two-phase flows by computing correctors. Applied Mathematics and Computation, 322(??):6– 29, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308135.

Thanh:2018:RPS

[Tha18b] Mai Duc Thanh. The Riemann problem for the shallow water equations with horizontal temperature gradients. Applied Mathematics and Computation, 325(??):159–178, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308901.

Hu:2019:TIN

[tHBSfY19] Yi teng Hu, Natalia Pavlovna Bondarenko, Chung-Tsun Shieh, and Chuan fu Yang. Traces and inverse nodal problems for Dirac-type integro-differential operators on a graph. Applied Mathematics and Computation, 363(??): Article 124606, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305983.

Trong:2019:ORU

[THM19] Dang Duc Trong, Dinh Nguyen Duy Hai, and Nguyen Dang Minh. Optimal regularization for an unknown source of space-fractional diffusion equation. Applied Mathematics and Computation, 349(??):184–206, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310828.

Thoudam:2017:NSC

[Tho17] Roshan Thoudam. Numerical solutions of coupled Klein–Gordon–Zakharov equations by quintic Bspline differential quadrature method. Applied Mathematics and Computation, 307(??):50–61, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301601.

Tatsumi:2015:PSO

[TIT15] Keiji Tatsumi, Takeru Ibuki, and Tetsuzo Tanino. Particle swarm optimization with stochastic selection of perturbationbased chaotic updating system. Applied Mathematics and Computation, 269(??):904–929, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010206.

Talaska:2015:EIM

[TKDF15] Tomasz Talaśka, Marta Kolasa, Rafal Dlugosz, and Pierre-André Farine. An efficient initialization mechanism of neurons for winner takes all neural network implemented in the CMOS technology. Applied Mathematics and Computation, 267(??):119–138, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005949.

Tan:2015:NCS

[TLGX15] Ronghua Tan, Zhijun Liu, Shengliang Guo, and Huili Xiang. On a nonautonomous competitive system subject to stochastic and impulsive perturbations. Applied Mathematics and Computation, 256(??):702–714, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500106X.

Tian:2017:AMF

[TLHC17] Zhaolu Tian, Zi-Cai Li, Hung-Tsai Huang, and C. S. Chen. Analysis of the method of fundamental solutions for the modified Helmholtz equation. Applied Mathematics and Computation, 305(??):262–281, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300899.

Tang:2016:CSP

[TLL16] Wensheng Tang, Guangming Lang, and Xuqiong Luo. Construction of symplectic (partitioned) Runge–Kutta methods with continuous stage. Applied Mathematics and Computation, 286(??): 279–287, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316302764.

Tong:2018:SOF

[TLL^{+}18] Liyun Tong, Yang Liu, Jungang Lou, Jianquan Lu, and Fuad E. Alsaadi. Static output feedback set stabilization for context-sensitive probabilistic Boolean control networks. Applied Mathematics and Computation, 332(??):263–275, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302169.

Tang:2019:CBN

[TLLD19] Zikai Tang, Hechao Liu, Huimin Luo, and Hanyuan Deng. Comparison between the non-self-centrality number and the total irregularity of graphs. Applied Mathematics and Computation, 361(??): 332–337, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL https://www.sciencedirect.
com/science/article/
pii/S0096300319304564.

Tan:2019:MAD

[TLX19] Manchun Tan, Yunfeng Liu, and Desheng Xu. Multistability analysis of delayed quaternion-valued neural networks with nonmonotonic piecewise nonlinear activation functions. Applied Mathematics and Computation, 341(??):229–255, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300318307471.

Tian:2017:NID

[TLZ17] Canrong Tian, Zhi Ling, and Lai Zhang. Nonlocal interaction driven pattern formation in a prey– predator model. Applied Mathematics and Computation, 308(??):73–83, September 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317301911.

Tian:2019:GIO

[TLZ^{+}19] Zhaolu Tian, Yong Liu, Yan Zhang, Zhongyun Liu, and Maoyi Tian. The general inner-outer iteration method based on regular splittings for the PageRank problem. Applied Mathematics and Computation, 356(??):479–501, September 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319301766.

Thai:2018:MSS

[TM18] Duy Hoang Thai and Lucas Mentch. Multiphase segmentation for simultaneously homogeneous and textural images. Applied Mathematics and Computation, 335(??):146–181, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303412.

Tambue:2019:SCL

[TM19] Antoine Tambue and Jean Daniel Mukam. Strong convergence of the linear implicit Euler method for the finite element discretization of semilinear SPDEs driven by multiplicative or additive noise. Applied Mathematics and Computation, 346(??):23– 40, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308609.

Tilahun:2018:CDP

[TMM18] Getachew Teshome Tilahun, Oluwole Daniel Makinde, and David Malonza. Codynamics of pneumonia and typhoid fever diseases with cost effective optimal control analysis. Applied Mathematics and Computation, 316(??):438–459, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305246.

Tan:2017:CPP

[TMPG17] Jianguo Tan, Weiwei Men, Yongzhen Pei, and Yongfeng Guo. Construction of positivity preserving numerical method for stochastic age-dependent population equations. Applied Mathematics and Computation, 293(??):57– 64, January 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316305057.

Talwar:2015:NSC

[TMS15] Jyoti Talwar, R. K. Mohanty, and Swarn Singh. A new spline in compression approximation for one space dimensional quasilinear parabolic equations on a variable mesh. Applied Mathematics and Computation, 260(??):82– 96, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003719.

Thuan:2019:RSR

[TN19] Do Duc Thuan and Le Van Ngoc. Robust stability and robust stabilizability for periodically switched linear systems. Applied Mathematics and Computation, 361(??):112–130, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304187.

Telksnys:2018:ESS

[TNMR18] T. Telksnys, Z. Navickas, R. Marcinkevicius, and M. Ragulskis. Existence of solitary solutions in systems of PDEs with multiplicative polynomial coupling. Applied Mathematics and Computation, 320(??):380–388, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306872.

Trinh:2015:BFR

[TNPL15] H. Trinh, Phan T. Nam, Pubudu N. Pathirana, and H. P. Le. On backwards and forwards reachable sets bounding for perturbed time-delay systems. Applied Mathematics and Computation, 269(??):664–673, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010383.

Tug:2018:ADG

[TÖGE18] Gül Tug, Zehra Özdemir, Ismail Gök, and F. Nejat Ekmekci. Accretive Darboux growth along a space curve. Applied Mathematics and Computation, 316(??):516–524, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305891.

Tomasiello:2012:SRN

[Tom12] S. Tomasiello. Some remarks on a new DQbased method for solving a class of Volterra integro-differential equations. Applied Mathematics and Computation, 219(1):399–407, September 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300312006303. See

corrigendum [Tom15].

Tomasiello:2015:CSR

[Tom15] Stefania Tomasiello. Corrigendum to “Some remarks on a new DQbased method for solving a class of Volterra integro-differential equations” [Appl. Math. Comput. 219 (2012) 399– 407]. Applied Mathematics and Computation, 253(??):410, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315000089. See

[Tom12].

Tong:2018:HPS

[Ton18]  Yanchun Tong.  High

precision solution for thermo-elastic equations using stable node-based smoothed finite element method. Applied Mathematics and Computation, 336(??):272–287, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304090.

Toro:2016:BVH

[Tor16] Eleuterio F. Toro. Brain venous haemodynamics, neurological diseases and mathematical modelling. A review. Applied Mathematics and Computation, 272 (part 2)(??):542–579, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008474.

Tiwari:2015:FTI

[TPCL15] Ajey K. Tiwari, S. N. Pandey, V. K. Chandrasekar, and M. Lakshmanan. Factorization technique and isochronous condition for coupled quadratic and mixed Liénardtype nonlinear systems. Applied Mathematics and Computation, 252(??): 457–472, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314017032.

Tang:2016:DOF

[TPL16] Ze Tang, Ju H. Park, and Tae H. Lee. Dynamic output-feedback-based H_{∞} design for networked control systems with multipath packet dropouts. Applied Mathematics and Computation, 275(??):121–133, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015398.

Tadeu:2015:DST

[TPS15] António Tadeu, Joana Prata, and Nuno Simões. Dynamic simulation of three-dimensional heat conduction through cylindrical inclusions using a BEM model formulated in the frequency domain. Applied Mathematics and Computation, 261(??):397–407, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004427.

Tian:2018:NVF

[TQ18] Peng Tian and Hourong Qin. Non-vanishing Fourier coefficients of Δ_{k}. Applied Mathematics and Computation, 339(??):488–506, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305800.

Tan:2015:NAB

[TQC15] Lihui Tan, Tao Qian, and Qiuhui Chen. New aspects of Beurling–Lax shift invariant subspaces. Applied Mathematics and Computation, 256(??):257–266, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000053.

Tratnik:2018:SHW

[Tra18] Niko Tratnik. On the Steiner hyper-Wiener index of a graph. Applied Mathematics and Computation, 337(??):350–359, November 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318304442.

Trenchard:2015:PSP

[Tre15] Hugh Trenchard. The peloton superorganism and protocooperative behavior. Applied Mathematics and Computation, 270(??):179–192, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010619.

Tan:2015:CSS

[TRP15] Jianguo Tan, A. Rathinasamy, and Yongzhen Pei. Convergence of the split-step θ-method for stochastic age-dependent population equations with Poisson jumps. Applied Mathematics and Computation, 254(??):305–317, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401786X.

Trenchard:2015:DMB

[TRRP15] Hugh Trenchard, Erick Ratamero, Ashlin Richardson, and Matjaz Perc. A deceleration model for bicycle peloton dynamics and group sorting. Applied Mathematics and Computation, 251(??):24–34, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015537.

Tang:2015:SHB

[TS15] Xiaosong Tang and Yongli Song. Stability, Hopf bifurcations and spatial patterns in a delayed diffusive predator–prey model with herd behavior. Applied Mathematics and Computation, 254(??):375–391, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314018049.

Topcu:2018:KGD

[TS18] Hatice Topcu and Sezer Sorgun. The kite graph is determined by its adjacency spectrum. Applied Mathematics and Computation, 330(??):134–142, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301474.

Tramontana:2015:PAS

[TSA15] Fabio Tramontana, Iryna Sushko, and Viktor Avrutin. Period adding structure in a 2D discontinuous model of economic growth. Applied Mathematics and Computation, 253(??): 262–273, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300314017329.

Tsitouras:2019:ERK

[Tsi19]  Ch. Tsitouras.  Ex-

plicit Runge–Kutta methods for starting integration of Lane–Emden problem. Applied Mathematics and Computation, 354(??):353–364, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301511.

Tamsir:2016:ABE

[TSJ16] Mohammad Tamsir, Vineet K. Srivastava, and Ram Jiwari. An algorithm based on exponential modified cubic B-spline differential quadrature method for nonlinear Burgers’ equation. Applied Mathematics and Computation, 290(??):111–124, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303642.

Tuma:2018:MVH

[TSPF18] Karel Tuma, Judith Stein, Vít Prusa, and Elfriede Friedmann. Motion of the vitreous humour in a deforming eye-fluid-structure interaction between a nonlinear elastic solid and viscoleastic fluid. Applied Mathematics and Computation, 335(??):50–64, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303485.

Tan:2016:NSB

[TST16] Jieqing Tan, Jiaze Sun, and Guangyue Tong. A non-stationary binary three-point approximating subdivision scheme. Applied Mathematics and Computation, 276(??):37– 43, March 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315015945.

Tsygvintsev:2019:OAD

[Tsy19] Alexei Tsygvintsev. On the overfly algorithm in deep learning of neural networks. Applied Mathematics and Computation, 349(??):348–358, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311160.

Tang:2019:HOS

[TSZ19] Wensheng Tang, Yajuan Sun, and Jingjing Zhang. High order symplectic integrators based on continuous-stage Runge– Kutta–Nyström methods. Applied Mathematics and Computation, 361(??):670–679, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304965.

Tucker:2016:ERS

[TT16] Paul G. Tucker and James C. Tyacke. Eddy resolving simulations in aerospace — invited paper (Numerical Fluid 2014). Applied Mathematics and Computation, 272 (part 3)(??):582–592, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500171X.

Tarasov:2018:MML

[TT18] Vasily E. Tarasov and Valentina V. Tarasova. Macroeconomic models with long dynamic memory: Fractional calculus approach. Applied Mathematics and Computation, 338(??):451–465, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830506X.

Tao:2017:NDP

[TTD17] Yunzhe Tao, Xiaochuan Tian, and Qiang Du. Nonlocal diffusion and peridynamic models with Neumann type constraints and their numerical approximations. Applied Mathematics and Computation, 305(??):282–298, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300875.

Thi:2018:EDB

[TTD18] Hoai An Le Thi, Anh Son Ta, and Tao Pham Dinh. An efficient DCA based algorithm for power control in large scale wireless networks. Applied Mathematics and Computation, 318(??):215–226, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306197.

Tuan:2015:MIE

[TTK15] Nguyen Huy Tuan, Le Duc Thang, and Vo Anh Khoa. A modified integral equation method of the nonlinear elliptic equation with globally and locally Lipschitz source. Applied Mathematics and Computation, 265(??):245–265, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004348.

Tian:2017:JGS

[TTLX17] Zhaolu Tian, Maoyi Tian, Zhongyun Liu, and Tongyang Xu. The Jacobi and Gauss–Seidel-type iteration methods for the matrix equation AXB = C. Applied Mathematics and Computation, 292(??):63–75, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304647.

Thakur:2016:NIS

[TTP16] Balwant Singh Thakur, Dipti Thakur, and Mihai Postolache. A new iterative scheme for numerical reckoning fixed points of Suzuki’s generalized nonexpansive mappings. Applied Mathematics and Computation, 275(??):147–155, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015672.

Truhar:2019:ADQ

[TTP19] Ninoslav Truhar, Zoran Tomljanović, and Matea Puvaca. Approximation of damped quadratic eigenvalue problem by dimension reduction. Applied Mathematics and Computation, 347(??):40–53, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309135.

Truhar:2015:DOM

[TTV15] Ninoslav Truhar, Zoran Tomljanović, and Kresimir Veselić. Damping optimization in mechanical systems with external force. Applied Mathematics and Computation, 250(??):270–279, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014568.

Tu:2016:LSO

[TTXZ16] Jian-Min Tu, Shou-Fu Tian, Mei-Juan Xu, and Tian-Tian Zhang. On Lie symmetries, optimal systems and explicit solutions to the Kudryashov–Sinelshchikov equation. Applied Mathematics and Computation, 275(??):345–352, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501574X.

Tai:2019:CSS

[TTZ^{+}19] Weipeng Tai, Qingyong Teng, Youmei Zhou, Jianping Zhou, and Zhen Wang. Chaos synchronization of stochastic reactiondiffusion time-delay neural networks via nonfragile output-feedback control. Applied Mathematics and Computation, 354(??):115–127, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301328.

Tan:2015:FPI

[TTZX15] Jieqing Tan, Guangyue Tong, Li Zhang, and Jin Xie. Four point interpolatory-corner cutting subdivision. Applied Mathematics and Computation, 265(??):819–825, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007389.

Tu:2017:EAV

[Tu17]  Jianhua Tu.  Efficient

algorithm for the vertex cover P_{k} problem on cacti. Applied Mathematics and Computation, 311(??):217–222, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303351.

Tuan:2015:FSO

[Tua15] Nguyen Dinh Tuan. First and second-order optimality conditions for nonsmooth vector optimization using set-valued directional derivatives. Applied Mathematics and Computation, 251(??):300–317, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015938.

Tura:2018:CPG

[Tur18]  Fernando Tura.  The

characteristic polynomial of generalized lollipop graphs. Applied Mathematics and Computation, 337(??):119–136, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304053.

Tacha:2016:AAC

[TVK^{+}16] O. I. Tacha, Ch. K.

Volos, I. M. Kyprianidis, I. N. Stouboulos, S. Vaidyanathan, and V.-T. Pham. Analysis, adaptive control and circuit simulation of a novel nonlinear finance system. Applied Mathematics and Computation, 276(??):200–217, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315300035.

Tao:2015:BRC

[TW15a] Youtian Tao and Dongyin Wang. A bivariate rational cubic interpolating spline with biquadratic denominator. Applied Mathematics and Computation, 264(??):366–377, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005627.

Tereszkiewicz:2015:GJP

[TW15b] Agnieszka Tereszkiewicz and Izabela Wawreniuk. Generalized Jacobsthal polynomials and special points for them. Applied Mathematics and Computation, 268(??):806–814, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009054.

Tang:2018:FFR

[TWF^{+}18] Zhijie Tang, Xiaocheng Wu, Bin Fu, Weiwei Chen, and Hao Feng. Fast face recognition based on fractal theory. Applied Mathematics and Computation, 321(??):721–730, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307993.

Tian:2018:TCL

[TX18] Tao Tian and Liming Xiong. Traceability on 2-connected line graphs. Applied Mathematics and Computation, 321(??): 463–471, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317307427.

Tian:2015:HBA

[TXG15] Xiaohong Tian, Rui Xu, and Qintao Gan. Hopf bifurcation analysis of a BAM neural network with multiple time delays and diffusion. Applied Mathematics and Computation, 266(??):909–926, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007857.

Tian:2019:MAC

[TXL19] Xiaohong Tian, Rui Xu, and Jiazhe Lin. Mathematical analysis of a cholera infection model with vaccination strategy. Applied Mathematics and Computation, 361(??):517–535, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304576.

Tang:2015:MPP

[TXM15] Jia Tang, Ya-Jun Xie, and Chang-Feng Ma. A modified product preconditioner for indefinite and asymmetric generalized saddle-point matrices. Applied Mathematics and Computation, 268(??):303–310, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008139.

Truong:2016:EGN

[TY16] Le Xuan Truong and Nguyen Van Y. Exponential growth with L^{p}-norm of solutions for nonlinear heat equations with viscoelastic term. Applied Mathematics and Computation, 273(??):656–663, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013491.

Tan:2018:SCT

[TY18] Li Tan and Chenggui Yuan. Strong convergence of a tamed theta scheme for NSD-DEs with one-sided Lipschitz drift. Applied Mathematics and Computation, 338(??):588–606, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830540X.

Tian:2015:GTS

[TYD15] Qing Tian, Wei Yang, and Liping Ding. On the general trigonometric sums weighted by character sums. Applied Mathematics and Computation, 268(??):375–387, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007973.

Tan:2018:IEO

[TZ18a] Zengqiang Tan and Chengjian Zhang. Implicit–explicit one-leg methods for nonlinear stiff neutral equations. Applied Mathematics and Computation, 335(??):196–210, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303643.

Tang:2018:FTQ

[TZ18b] Lang Tang and Shenglin Zhou. Flag-transitive quasi-residual designs with sporadic socle. Applied Mathematics and Computation, 320(??):56–60, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306306.

Tang:2018:SPC

[TZ18c] Wensheng Tang and Jingjing Zhang. Symplecticitypreserving continuousstage Runge–Kutta–Nyström methods. Applied Mathematics and Computation, 323(??):204–219, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308433.

Tian:2018:SCP

[TZ18d] Yingxu Tian and Haoyan Zhang. Skew CIR process, conditional characteristic function, moments and bond pricing. Applied Mathematics and Computation, 329(??):230–238, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301152.

Tang:2019:SIB

[TZ19a] Wensheng Tang and Jingjing Zhang. Symmetric integrators based on continuousstage Runge–Kutta–Nyström methods for reversible systems. Applied Mathematics and Computation, 361(??):1–12, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304060.

Tang:2019:RFO

[TZ19b] Yongchao Tang and Tie Zhang. A remark on the q-fractional order differential equations. Applied Mathematics and Computation, 350(??):198–208, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300165.

Tang:2019:DNI

[TZC19] Lang Tang, Shenglin Zhou, and Jing Chen. Domination number of incidence graphs of block designs. Applied Mathematics and Computation, 363 (??):Article 124600, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305922.

Tu:2019:SAQ

[TZD^{+}19] Zhengwen Tu, Yongxiang Zhao, Nan Ding, Yuming Feng, and Wei Zhang. Stability analysis of quaternion-valued neural networks with both discrete and distributed delays. Applied Mathematics and Computation, 343(??):342–353, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318308282.

Tang:2015:NMR

[TZF15] Jingyong Tang, Jinchuan Zhou, and Liang Fang.

A non-monotone regularization Newton method for the second-order cone complementarity problem. Applied Mathematics and Computation, 271(??):743–756, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012357.

Touma:2015:CFV

[TZH15] R. Touma, D. Zeidan, and S. Habre. Central finite volume schemes on nonuniform grids and applications. Applied Mathematics and Computation, 262(??):15–30, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004488.

Tu:2017:KCC

[TZS17] Jianhua Tu, Yukang Zhou, and Guifu Su. A kind of conditional connectivity of Cayley graphs generated by wheel graphs. Applied Mathematics and Computation, 301(??):177–186, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S009630031630741X.

Tang:2017:BBP

[TZTG17] Yaning Tang, Weijian Zai, Siqiao Tao, and Qing Guan. Binary Bell polynomials, Hirota bilinear approach to Levi equation. Applied Mathematics and Computation, 293(??):565–574, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305173.

Ugurlu:2018:SIS

[UBT18] Ekin Ugurlu, Dumitru Baleanu, and Kenan Tas. On square integrable solutions of a fractional differential equation. Applied Mathematics and Computation, 337(??):144–152, November 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031830417X.

Ucar:2016:CIS

[Uça16] H. Esra Özkan Uçar. Coefficient inequality for qstarlike functions. Applied Mathematics and Computation, 276(??):122–126, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315016094.

Ulusoy:2015:SAG

[UDA15] Gülsüm Ulusoy, Emre Deniz, and Ali Aral. Simultaneous approximation with generalized Durrmeyer operators. Applied Mathematics and Computation, 260(??):126–134, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003744.

Zaheer-ud-Din:2018:MMO

[uDuI18] Zaheer ud Din and Siraj ul Islam. Meshless methods for one-dimensional oscillatory Fredholm integral equations. Applied Mathematics and Computation, 324(??):156–173, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308512.

Ueda:2019:EIA

[Ued19] Masahiko Ueda. Effect of information asymmetry in Cournot duopoly game with bounded rationality. Applied Mathematics and Computation, 362 (??):Article 124535, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305144.

Ugurlu:2019:RTO

[Ugu19]  Ekin Ugurlu.  Regu-

lar third-order boundary value problems. Applied Mathematics and Computation, 343(??):247–257, February 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318308257.

Uzunoglu:2016:NAC

[UGY16] Beyhan Uzunoglu, Ismail Gök, and Yusuf Yayli. A new approach on curves of constant precession. Applied Mathematics and Computation, 275(??):317–323, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015854.

Udagedara:2015:ROM

[UHLM15] Indika Udagedara, Brian Helenbrook, Aaron Luttman, and Stephen E. Mitchell. Reduced order modeling for accelerated Monte Carlo simulations in radiation transport. Applied Mathematics and Computation, 267(??):237–251, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315004324.

Umeo:2018:NRF

[UHN^{+}18] Hiroshi Umeo, Mitsuki Hirota, Youhei Nozaki, Keisuke Imai, and Takashi Sogabe. A new reconstruction and the first implementation of Goto’s FSSP algorithm. Applied Mathematics and Computation, 318(??):92–108, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730303X.

Siraj-ul-Islam:2018:MMR

[uIHA18] Siraj ul Islam, Nadeem Haider, and Imran Aziz. Meshless and multi-resolution collocation techniques for parabolic interface models. Applied Mathematics and Computation, 335(??):313–332, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830362X.

Uzunkol:2018:SWU

[UK18] Osmanbey Uzunkol and Mehmet Sabir Kiraz. Still wrong use of pairings in cryptography. Applied Mathematics and Computation, 333(??):467–479, September 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318302352.

Ucar:2015:NAN

[UKE15] Y. Ucar, B. Karaagac, and A. Esen. A new approach on numerical solutions of the Improved Boussinesq type equation using quadratic B-spline Galerkin finite element method. Applied Mathematics and Computation, 270(??):148–155, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010620.

Uzunca:2017:MMD

[UKK17] M. Uzunca, B. Karasözen, and T. Küçükseyhan.

Moving mesh discontinuous Galerkin methods for PDEs with traveling waves. Applied Mathematics and Computation, 292(??):9–18, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304726.

Ullah:2016:SFT

[UKS^{+}16] M. Zaka Ullah, S. Kosari, F. Soleymani, F. Khaksar Haghani, and A. S. Al-Fhaid. A superfast tri-parametric iterative method with memory. Applied Mathematics and Computation, 289(??):486–491, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303381.

Unami:2019:TPO

[UMF19] Koichi Unami, Osama Mohawesh, and Rasha M. Fadhil. Time periodic optimal policy for operation of a water storage tank using the dynamic programming approach. Applied Mathematics and Computation, 353(??):418–431, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301031.

Upadhyaya:2016:PPD

[Upa16] Shweta Upadhyaya. Performance prediction of a discrete-time batch arrival retrial queue with Bernoulli feedback. Applied Mathematics and Computation, 283(??): 108–119, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316301370.

Upadhyay:2018:EBP

[UPMV18] Ranjit Kumar Upadhyay, Chinmoy Paul, Argha Mondal, and Gajendra K. Vishwakarma. Estimation of biophysical parameters in a neuron model under random fluctuations. Applied Mathematics and Computation, 329(??):364–373, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301139.

urRahman:2019:MAG

[uRAD19] Ghaus ur Rahman, Ravi P. Agarwal, and Qamar Din.

Mathematical analysis of giving up smoking model via harmonic mean type incidence rate. Applied Mathematics and Computation, 354(??):128–148, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300682.

urRehman:2017:QMN

[uRIS17] Mujeeb ur Rehman, Amna Idrees, and Umer Saeed. A quadrature method for numerical solutions of fractional differential equations. Applied Mathematics and Computation, 307(??):38–49, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301649.

Ullah:2015:EMS

[USCA15] Malik Zaka Ullah, Stefano Serra-Capizzano, and Fayyaz Ahmad. An efficient multi-step iterative method for computing the numerical solution of systems of nonlinear equations associated with ODEs. Applied Mathematics and Computation, 250(??):249–259, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014787.

Ullah:2015:HOM

[USCAAA15] Malik Zaka Ullah, S. Serra-Capizzano, Fayyaz Ahmad, and Eman S. Al-Aidarous. Higher order multi-step iterative method for computing the numerical solution of systems of nonlinear equations: Application to nonlinear PDEs and ODEs. Applied Mathematics and Computation, 269(??):972–987, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010188.

Uzun:2017:IAC

[Uzu17] Rukiye Uzun. Influences of autapse and channel blockage on multiple coherence resonance in a single neuron. Applied Mathematics and Computation, 315(??):203–210, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305167.

Vabishchevich:2015:ESP

[Vab15] P. N. Vabishchevich. Explicit schemes for parabolic and hyperbolic equations. Applied Mathematics and Computation, 250(??):424–431, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014994.

Vadillo:2019:CSL

[Vad19] Fernando Vadillo. Comparing stochastic Lotka– Volterra predator–prey models. Applied Mathematics and Computation, 360(??):181–189, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303893.

Valls:2015:CPC

[Val15]  Claudia Valls.  Cen-

ter problem in the center manifold for quadratic and cubic differential systems in R^{3}. Applied Mathematics and Computation, 251(??):180–191, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314015896.

Valero:2018:SBS

[Val18] Valentín Valero. Strong behavioral similarities in timed-arc Petri nets. Applied Mathematics and Computation, 333(??): 401–415, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318302546.

Vannucci:2015:NCE

[Van15] Paolo Vannucci. A note on the computation of the extrema of Young’s modulus for hexagonal materials: an approach by planar tensor invariants. Applied Mathematics and Computation, 270(??):124–129, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010802.

Vorel:2015:CMR

[VB15] Jan Vorel and W. P. Boshoff. Computational modelling of real structures made of strainhardening cement-based composites. Applied Mathematics and Computation, 267(??):562–570, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000867.

Villa:2017:ITD

[VBG^{+}17] Andrea Villa, Luca Barbieri, Marco Gondola, Andres R. Leon-Garzon, and Roberto Malgesini. An implicit three-dimensional fractional step method for the simulation of the corona phenomenon. Applied Mathematics and Computation, 311(??):85– 99, October 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317303107.

Vaz:2015:SSS

[VD15] Cristina L. D. Vaz and Hugo A. C. Diniz. Steady states solutions of Allen– Cahn equation by computer algebra. Applied Mathematics and Computation, 252(??):240–248, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401580X.

Varatharajan:2016:EPA

[VD16] N. Varatharajan and Anirvan DasGupta. The effect of perturbed advection on a class of solutions of a non-linear reaction–diffusion equation. Applied Mathematics and Computation, 290(??):33–45, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303460.

Vitanov:2017:SWS

[VDI17] Nikolay K. Vitanov, Zlatinka I. Dimitrova, and Tsvetelina I. Ivanova. On solitary wave solutions of a class of nonlinear partial differential equations based on the function 1/coshn(αx + βt). Applied Mathematics and Computation, 315(??):372–380, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305258.

Vitanov:2015:MMS

[VDV15] Nikolay K. Vitanov, Zlatinka I. Dimitrova, and Kaloyan N. Vitanov. Modified method of simplest equation for obtaining exact analytical solutions of nonlinear partial differential equations: further development of the methodology with applications. Applied Mathematics and Computation, 269(??):363–378, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009790.

Vilches:2019:DPS

[VEF19] T. N. Vilches, L. Esteva, and C. P. Ferreira. Disease persistence and serotype coexistence: an expected feature of human mobility. Applied Mathematics and Computation, 355(??):161–172, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301651.

Vimercati:2018:NSN

[VG18] Davide Vimercati and Alberto Guardone. On the numerical simulation of non-classical quasi-1D steady nozzle flows: Capturing sonic shocks. Applied Mathematics and Computation, 319(??): 617–632, February 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317304940.

Vidurupola:2018:ADS

[Vid18] Sukhitha W. Vidurupola. Analysis of deterministic and stochastic mathematical models with resistant bacteria and bacteria debris for bacteriophage dynamics. Applied Mathematics and Computation, 316(??):215–228, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305738.

Villela:2016:AVC

[Vil16] Daniel A. M. Villela. Analysis of the vectorial capacity of vector-borne diseases using moment-generating functions. Applied Mathematics and Computation, 290(??):1–8, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303356.

Visnjic:2015:APD

[Vis15] Jelena Visnjić. On additive properties of the Drazin inverse of block matrices and representations. Applied Mathematics and Computation, 250(??):444–450, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015112.

Vijesh:2015:WBQ

[VK15] V. Antony Vijesh and K. Harish Kumar. Wavelet based quasilinearization method for semi-linear parabolic initial boundary value problems. Applied Mathematics and Computation, 266(??):1163–1176, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315007705.

See erratum [VK17a].

Vijesh:2017:EWB

[VK17a] V. Antony Vijesh and K. Harish Kumar. Erratum to “Wavelet based quasilinearization method for semi-linear parabolic initial boundary value problems” [appl. math. comput. 266 (2015) 1163– 1176]. Applied Mathematics and Computation, 314(??):484, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317303405. See

[VK15].

Volna:2017:EEB

[VK17b] Eva Volna and Martin Kotyrba. Enhanced ensemble-based classifier with boosting for pattern recognition. Applied Mathematics and Computation, 310(??):1–14, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302710.

Verma:2019:IBA

[VK19] Pallavi Verma and Lakhveer Kaur. Integrability, bilinearization and analytic study of new form of (3+1)-dimensional B-type Kadomstev–Petviashvili (BKP)-Boussinesq equation. Applied Mathematics and Computation, 346(??):879–886, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310245.

Vazquez-Lozano:2018:DAC

[VLCT18] J. Enrique Vázquez-Lozano, Alicia Cordero, and Juan R. Torregrosa. Dynamical analysis on cubic polynomials of Damped Traub’s method for approximating multiple roots. Applied Mathematics and Computation, 328(??):82–99, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830064X.

vanLith:2018:FLM

[vLtTBI18] Bart S. van Lith, Jan H. M. ten Thije Boonkkamp, and Wilbert L. IJzerman. Full linear multistep methods as rootfinders. Applied Mathematics and Computation, 320(??):190–201, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306239.

Vargas:2019:RSM

[VMLO19] Alessandro N. Vargas, Marcio A. F. Montezuma, Xinghua Liu, and Ricardo C. L. F. Oliveira. Robust stability of Markov jump linear systems through randomized evaluations. Applied Mathematics and Computation, 346(??): 287–294, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318308518.

Volna:2015:ECP

[Vol15]  Barbora Volná.  Exis-

tence of chaos in the plane R^{2} and its application in macroeconomics. Applied Mathematics and Computation, 258(??):237–266, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001277.

Vizinho:2015:TTM

[VPS15] Rui Vizinho, José Páscoa, and Miguel Silvestre. Turbulent transition modeling through mechanical considerations. Applied Mathematics and Computation, 269(??):308–325, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010024.

Vijesh:2015:MQM

[VRC15] V. Antony Vijesh, Rupsha Roy, and G. Chandhini. A modified quasilinearization method for fractional differential equations and its applications. Applied Mathematics and Computation, 266(??):687–697, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007638.

Vo:2017:IVP

[VS17] H. D. Vo and R. B.

Sidje. Implementation of variable parameters in the Krylov-based finite state projection for solving the chemical master equation. Applied Mathematics and Computation, 293(??):334–344, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305082.

Vorozhtsov:2019:ECD

[VS19] Evgenii V. Vorozhtsov and Vasily P. Shapeev. On the efficiency of combining different methods for acceleration of iterations at the solution of PDEs by the method of collocations and least residuals. Applied Mathematics and Computation, 363 (??):Article 124644, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319306368.

Verma:2018:NKF

[VV18] S. Verma and P. Viswanathan. A note on Katugampola fractional calculus and fractal dimensions. Applied Mathematics and Computation, 339(??):206–219, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318306003.

Vigonski:2015:VMS

[VVA^{+}15] Simon Vigonski, Mihkel Veske, Alvo Aabloo, Flyura Djurabekova, and Vahur Zadin. Verification of a multiscale surface stress model near voids in copper under the load induced by external high electric field. Applied Mathematics and Computation, 267(??):476–486, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001344.

Verwaeren:2015:SGP

[VVD15] Jan Verwaeren, Pieter Van der Weeën, and Bernard De Baets. A search grid for parameter optimization as a byproduct of model sensitivity analysis. Applied Mathematics and Computation, 261(??):8– 27, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003781.

VanCleemput:2018:RNH

[VZ18] Nico Van Cleemput and Carol T. Zamfirescu. Regular non-Hamiltonian polyhedral graphs. Applied Mathematics and Computation, 338(??):175–191, December 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031830479X.

Wang:2015:URP

[WA15a] Chao Wang and Ravi P.

Agarwal. Uniformly rdpiecewise almost periodic functions with applications to the analysis of impulsive Δ-dynamic system on time scales. Applied Mathematics and Computation, 259(??):271–292, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315002428.

Wang:2015:PSS

[WA15b] Fanglei Wang and Yukun An. On positive solutions for a second order differential system with indefinite weight. Applied Mathematics and Computation, 259(??):753–761, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002908.

Wang:2018:CVB

[WA18] Xue Wang and Whye-Teong Ang. A complex variable boundary element method for solving a steady-state advectiondiffusion-reaction equation. Applied Mathematics and Computation, 321(??):731–744, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307981.

Wang:2015:RDP

[Wan15a] Caiyun Wang. Rich dynamics of a predator–prey model with spatial motion. Applied Mathematics and Computation, 260(??):1–9, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003434.

Wang:2015:SSS

[Wan15b] Guoliang Wang. Stochastic stabilization of singular systems with Markovian switchings. Applied Mathematics and Computation, 250(??):390–401, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015410.

Wang:2015:AMS

[Wan15c]  JinRong Wang.  Ap-

proximate mild solutions of fractional stochastic evolution equations in Hilbert spaces. Applied Mathematics and Computation, 256(??):315–323, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000193.

Wang:2015:NON

[Wan15d] Qi Wang. Numerical oscillation of neutral logistic delay differential equation. Applied Mathematics and Computation, 258(??):49– 59, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001459.

Wang:2015:EUF

[Wan15e] Wu-Sheng Wang. Estimation of unknown function of a class of integral inequalities and applications in fractional integral equations. Applied Mathematics and Computation, 268(??):1029–1037, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009273.

Wang:2015:SPW

[Wan15f] Xiaohuan Wang. Stability of planar waves in a Lotka– Volterra system. Applied Mathematics and Computation, 259(??):313–326, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315002398.

Wang:2016:MIC

[Wan16a] Guoliang Wang. Modeindependent control of singular Markovian jump systems: a stochastic optimization viewpoint. Applied Mathematics and Computation, 286(??):155–170, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302466.

Wang:2016:LBM

[Wan16b] Huimin Wang. A lattice Boltzmann model for the ion- and electron-acoustic solitary waves in beamplasma system. Applied Mathematics and Computation, 279(??):62–75, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300078.

Wang:2016:OOU

[Wan16c] Wen-Huan Wang. Ordering of oriented unicyclic graphs by skew energies. Applied Mathematics and Computation, 284(??):136–148, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301722.

Wanduku:2017:CGA

[Wan17a] Divine Wanduku. Complete global analysis of a two-scale network SIRS epidemic dynamic model with distributed delay and random perturbations. Applied Mathematics and Computation, 294(??):49– 76, February 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316305616.

Wang:2017:ECR

[Wan17b] Shujing Wang. On extremal cacti with respect to the Szeged index. Applied Mathematics and Computation, 309(??):85–92, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302205.

Wang:2017:SOL

[Wan17c]  Wansheng Wang.  On

A-stable one-leg methods for solving nonlinear Volterra functional differential equations. Applied Mathematics and Computation, 314(??):380–390, December 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317304642.

Wang:2018:ISP

[Wan18]  Yu Ping Wang.  The

inverse spectral problem for differential pencils by mixed spectral data. Applied Mathematics and Computation, 338(??):537–543, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305411.

Wang:2019:DAC

[Wan19a] Jiqiang Wang. Disturbance attenuation of complex dynamical systems through interaction topology design. Applied Mathematics and Computation, 355(??):576–584, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301857.

Wang:2019:ADT

[Wan19b] Yuanshi Wang. Asymmetric diffusion in a twopatch consumer-resource system. Applied Mathematics and Computation, 361(??):258–273, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304370.

Waziri:2019:FHZ

[WAS19] Mohammed Yusuf Waziri, Kabiru Ahmed, and Jamilu Sabi’u. A family of Hager– Zhang conjugate gradient methods for system of monotone nonlinear equations. Applied Mathematics and Computation, 361(??):645–660, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304771.

Wu:2017:LFR

[WBL17] Guo-Cheng Wu, Dumitru Baleanu, and Wei-Hua Luo. Lyapunov functions for Riemann–Liouville-like fractional difference equations. Applied Mathematics and Computation, 314(??):228–236, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304320.

Wang:2015:PNP

[WC15a] Carol J. Wang and Zhongjin Chen. The parity of the number of permutation tableaux in a fixed shape. Applied Mathematics and Computation, 265(??):1090–1097, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500805X.

Wu:2015:CSU

[WC15b] Hsin-Hung Wu and Shih-Yu Chang. A case study of using DEMATEL method to identify critical factors in green supply chain management. Applied Mathematics and Computation, 256(??):394–403, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000557.

Wang:2016:SSP

[WC16] Jie Wang and Cui-Ping Cheng. Spreading speeds and periodic traveling waves of a partially sedentary integro-difference model. Applied Mathematics and Computation, 274(??): 459–479, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315014678.

Wu:2017:MCM

[WC17] Sheng-Jhih Wu and Moody T. Chu. Markov chains with memory, tensor formulation, and the dynamics of power iteration. Applied Mathematics and Computation, 303(??):226–239, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300383.

Wu:2018:EEC

[WC18] Yueyu Wu and Yaojun Chen. On the extremal eccentric connectivity index of graphs. Applied Mathematics and Computation, 331(??):61–68, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301528.

Wang:2019:SAN

[WCB19] Zhichuang Wang, Guoliang Chen, and Hezhen Ba. Stability analysis of nonlinear switched systems with sampled-data controllers. Applied Mathematics and Computation, 357(??):297–309, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302826.

Wu:2017:MTI

[WCD17] Renfang Wu, Hanlin Chen, and Hanyuan Deng. On the monotonicity of topological indices and the connectivity of a graph. Applied Mathematics and Computation, 298(??):188–200, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306853.

Wei:2015:GES

[WCH15] Linna Wei, Wu-Hua Chen, and Ganji Huang. Globally exponential stabilization of neural networks with mixed time delays via impulsive control. Applied Mathematics and Computation, 260(??):10–26, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003574.

Wang:2019:FDA

[WCH19] Yi Wang, Jinde Cao, and Gang Huang. Further dynamic analysis for a network sexually transmitted disease model with birth and death. Applied Mathematics and Computation, 363(??): Article 124635, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319306277.

Wang:2019:PPM

[WCHW19] Shengbo Wang, Yanyi Cao, Tingwen Huang, and Shiping Wen. Passivity and passification of memristive neural networks with leakage term and time-varying delays. Applied Mathematics and Computation, 361(??):294–310, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304424.

Wang:2016:AFG

[WCS16] Miao-Kun Wang, Yu-Ming Chu, and Ying-Qing Song. Asymptotical formulas for Gaussian and generalized hypergeometric functions. Applied Mathematics and Computation, 276(??):44–60, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015908.

Wang:2015:LSF

[WCZ15] Zhujun Wang, Li Cai, and Detong Zhu. Line search filter inexact secant methods for nonlinear equality constrained optimization. Applied Mathematics and Computation, 263(??):47– 58, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004671.

Wang:2017:SSD

[WCZY17] Guoliang Wang, Hongyang Cai, Qingling Zhang, and Chunyu Yang. Stabilization of stochastic delay systems via a disordered controller. Applied Mathematics and Computation, 314(??):98–109, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304356.

Wei:2015:IEP

[WD15] Ying Wei and Hua Dai. An inverse eigenvalue problem for Jacobi matrix. Applied Mathematics and Computation, 251(??):633–642, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314016348.

Wu:2016:MQQ

[WD16] Hui-Yuan Wu and Yong Duan. Multi-quadric quasi-interpolation method coupled with FDM for the Degasperis–Procesi equation. Applied Mathematics and Computation, 274(??): 83–92, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315013958.

Wang:2019:NDC

[WD19] Haijun Wang and Guili Dong. New dynamics coined in a 4-D quadratic autonomous hyper-chaotic system. Applied Mathematics and Computation, 346(??):272–286, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S009630031830866X.

Wang:2018:SUB

[WDB18] Guiyan Wang, Chunli Deng, and Changjiang Bu. Some upper bounds on Z_{t}-eigenvalues of tensors. Applied Mathematics and Computation, 329(??):266–277, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300936.

Wang:2019:ASF

[WDCJ19] Mei Wang, Feifei Du, Churong Chen, and Baoguo Jia. Asymptotic stability of (q, h)-fractional difference equations. Applied Mathematics and Computation, 349(??):158–167, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311007.

Wang:2019:FTSa

[WDG19a] Leimin Wang, Tiandu Dong, and Ming-Feng Ge. Finite-time synchronization of memristor chaotic systems and its application in image encryption. Applied Mathematics and Computation, 347(??):293–305, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309901.

Wang:2019:CLD

[WDG19b] Xiaofeng Wang, Weizhong Dai, and Shuangbing Guo. A conservative linear difference scheme for the 2D regularized longwave equation. Applied Mathematics and Computation, 342(??):55–70, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308014.

Wang:2015:NMM

[WDLL15] Xiao Wang, Jinqiao Duan, Xiaofan Li, and Yuanchao Luan. Numerical methods for the mean exit time and escape probability of twodimensional stochastic dynamical systems with non-Gaussian noises. Applied Mathematics and Computation, 258(??):282–295, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001496.

Wang:2018:NAM

[WDLS18] Xiao Wang, Jinqiao Duan, Xiaofan Li, and Renming Song. Numerical algorithms for mean exit time and escape probability of stochastic systems with asymmetric Lévy motion. Applied Mathematics and Computation, 337(??):607–617, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304478.

Wang:2019:GES

[WDLZ19] Xiaohong Wang, Haibo Du, Shihua Li, and Xingpeng Zhou. Globally exponential stabilization for a class of nonlinear systems with time delays both in nonlinearities and input. Applied Mathematics and Computation, 359(??):478–489, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303480.

Wang:2017:HSC

[WDZ17] Zhanshan Wang, Sanbo Ding, and Huaguang Zhang. Hierarchy of stability criterion for time-delay systems based on multiple integral approach. Applied Mathematics and Computation, 314(??):422–428, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304678.

Weijermars:2019:DFP

[WE19] Ruud Weijermars and Mahmood Ettehad. Displacement field potentials for deformation in elastic media: Theory and application to pressureloaded boreholes. Applied Mathematics and Computation, 340(??):276–295, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307434.

Wang:2018:TLA

[WEH18] Fei Wang, Joseph Eichholz, and Weimin Han. A two level algorithm for an obstacle problem. Applied Mathematics and Computation, 330(??):65–76, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318301322.

Wei:2017:ANF

[Wei17]  Leilei Wei.  Analy-

sis of a new finite difference/local discontinuous Galerkin method for the fractional diffusionwave equation. Applied Mathematics and Computation, 304(??):180–189, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300723.

Walicka:2015:RNE

[WF15a] A. Walicka and J. Falicki.

Reynolds number effects in the flow of an electrorheological fluid of a Casson type between fixed surfaces of revolution. Applied Mathematics and Computation, 250(??):636–649, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014878.

Wu:2015:SHB

[WF15b] Ranchao Wu and Tianbao Fang. Stability and Hopf bifurcation of a Lorenzlike system. Applied Mathematics and Computation, 262(??):335–343, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005342.

Wu:2015:NAS

[WF15c] Yong-Ping Wu and Guo-Lin Feng. A new algorithm for seasonal precipitation forecast based on global atmospheric hydrological water budget. Applied Mathematics and Computation, 268(??):478–488, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008401.

Wang:2017:CSM

[WFZ17] JinRong Wang, Michal Feckan, and Yong Zhou. Center stable manifold for planar fractional damped equations. Applied Mathematics and Computation, 296(??):257–269, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306129.

Gao:2015:ICI

[wG15a]  Pei wang Gao.  Im-

provement and its computer implementation of an artificial-free simplextype algorithm by Arsham. Applied Mathematics and Computation, 263(??):410–415, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005391.

Wei:2015:HNN

[WG15b] Jiao Wei and Xianguo Geng. A hierarchy of new nonlinear evolution equations and generalized bi-Hamiltonian structures. Applied Mathematics and Computation, 268(??):664–670, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008863.

Wozny:2015:EMM

[WGL15] Paweł Woźny, Przemysław Gospodarczyk, and Stanisław Lewanowicz. Efficient merging of multiple segments of Bézier curves. Applied Mathematics and Computation, 268(??):354–363, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008607.

Wang:2018:EHT

[WGL18] Chengyong Wang, Ziliang Guo, and Shuchao Li. Expected hitting times for random walks on the ktriangle graph and their applications. Applied Mathematics and Computation, 338(??):676–697, December 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318305459.

Wang:2016:TDH

[WGLZ16] Jinliang Wang, Min Guo, Xianning Liu, and Zhitao Zhao. Threshold dynamics of HIV-1 virus model with cell-to-cell transmission, cell-mediated immune responses and distributed delay. Applied Mathematics and Computation, 291(??):149–161, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304039.

Wang:2019:IAD

[WGSX19] Zhishuang Wang, Quantong Guo, Shiwen Sun, and Chengyi Xia. The impact of awareness diffusion on SIR-like epidemics in multiplex networks. Applied Mathematics and Computation, 349(??):134–147, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311068.

Welburn:2015:GTM

[WH15] Jonathan W. Welburn and Kjell Hausken. A game theoretic model of economic crises. Applied Mathematics and Computation, 266(??):738–762, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315007171.

Wang:2018:CSC

[WH18] Hui-Di Wang and Zheng-Da Huang. On convergence and semi-convergence of SSOR-like methods for augmented linear systems. Applied Mathematics and Computation, 326(??):87– 104, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300122.

Wang:2018:RDP

[WHC18] Qiang Wang, Nanrong He, and Xiaojie Chen. Replicator dynamics for public goods game with resource allocation in large populations. Applied Mathematics and Computation, 328(??):162–170, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300687.

Wang:2019:SLR

[WHM^{+}19] Jian-Li Wang, Ting-Zhu Huang, Tian-Hui Ma, Xi-Le Zhao, and Yong Chen. A sheared low-rank model for oblique stripe removal. Applied Mathematics and Computation, 360(??):167–180, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319302747.

Wang:2015:NMS

[WHSS15] Liqun Wang, Songming Hou, Liwei Shi, and James Solow. A numerical method for solving elasticity equations with interface involving multidomains and triple junction points. Applied Mathematics and Computation, 251(??):615–625, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401604X.

Wang:2017:OBP

[WHTW17] Qingzhi Wang, Yong He, Guanzheng Tan, and Min Wu. Observer-based periodically intermittent control for linear systems via piecewise Lyapunov function method. Applied Mathematics and Computation, 293(??):438–447, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305379.

Wang:2019:SDS

[WHWW19] Jing Wang, Xiaohui Hu, Yunliang Wei, and Zhen Wang. Sampled-data synchronization of semi-Markov jump complex dynamical networks subject to generalized dissipativity property. Applied Mathematics and Computation, 346(??):853–864, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318309688.

Wu:2015:CHH

[WHX15] Tingting Wu, Deren Han, and Yi Xu. Cross-Hill: a heuristic method for global optimization. Applied Mathematics and Computation, 266(??):959–967, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315007948.

Wang:2018:GES

[WHX18] Lingyu Wang, Tingwen Huang, and Qiang Xiao. Global exponential synchronization of nonautonomous recurrent neural networks with time delays on time scales. Applied Mathematics and Computation, 328(??):263–275, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830050X.

Wang:2015:NIF

[WIF15] JinRong Wang, Ahmed Gamal Ibrahim, and Michal Feckan. Nonlocal impulsive fractional differential inclusions with fractional sectorial operators on Banach spaces. Applied Mathematics and Computation, 257(??):103–118, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031400633X.

Williams:2018:SMI

[Wil18] Robert Williams. Strongly maximal intersection-complete neural codes on grids are convex. Applied Mathematics and Computation, 336(??):162–175, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303904.

Wang:2019:EDU

[WJ19] Qiuling Wang and Danyang Jia. Expectation driven by update willingness promotes cooperation in the spatial prisoner’s dilemma game. Applied Mathematics and Computation, 352(??):174–179, July 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300748.

Wang:2019:DSP

[WJGM19] Jianmin Wang, Ping Jiang, Yu Guo, and Jing Meng. Developable surface pencil pairs with special pairs as common asymptotes. Applied Mathematics and Computation, 362 (??):Article 124583, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305752.

Wang:2017:SHI

[WJHA17] Yan Wang, Daqing Jiang, Tasawar Hayat, and Bashir Ahmad. A stochastic HIV infection model with T-cell proliferation and CTL immune response. Applied Mathematics and Computation, 315(??):477–493, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305234.

Wang:2019:PMF

[WJL19] Tianjun Wang, Yujian Jiao, and Wenjie Liu. Pseudospectral method for Fisher equation in a disk. Applied Mathematics and Computation, 343(??):30–48, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830780X.

Wang:2015:UCB

[WJNW15] Xu-Wen Wang, Luo-Luo Jiang, Sen Nie, and Bing-Hong Wang. Uncovering cooperative behaviors with sparse historical behavior data in the spatial games. Applied Mathematics and Computation, 271(??):317–322, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012412.

Wang:2018:ECC

[WJWX18] Tingchun Wang, Jiaping Jiang, Hanquan Wang, and Weiwei Xu. An efficient and conservative compact finite difference scheme for the coupled Gross–Pitaevskii equations describing spin-1 Bose–Einstein condensate. Applied Mathematics and Computation, 323(??):164–181, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308007.

Wang:2019:ISR

[WJZF19] Ruihua Wang, Ticao Jiao, Tao Zhang, and Shumin Fei. Improved stability results for discrete-time switched systems: a multiple piecewise convex Lyapunov function approach. Applied Mathematics and Computation, 353(??):54– 65, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300827.

Wang:2016:CCI

[WK16a] Xiuhua Wang and Jisheng Kou. Convergence for a class of improved sixthorder Chebyshev–Halley type methods. Applied Mathematics and Computation, 273(??):513–524, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009777.

Wang:2016:SCA

[WK16b] Xiuhua Wang and Jisheng Kou. Semilocal convergence analysis on the modifications for Chebyshev– Halley methods under generalized condition. Applied Mathematics and Computation, 281(??):243–251, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300352.

Wang:2015:UHS

[WL15] JinRong Wang and Xuezhu Li. Ulam–Hyers stability of fractional Langevin equations. Applied Mathematics and Computation, 258(??):72–83, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315001435.

Wang:2016:NHW

[WL16a] P. Wang and D. Q.

Lu. Nonlinear hydroelastic waves traveling in a thin elastic plate floating on a two-layer fluid. Applied Mathematics and Computation, 274(??):700–710, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014265.

Wang:2016:AER

[WL16b] Weiping Wang and Hongmei Liu. Asymptotic expansions related to hyperfactorial function and Glaisher–Kinkelin constant. Applied Mathematics and Computation, 283(??):153–162, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301382.

Wang:2016:BKM

[WL16c] Yinxia Wang and Hengyan Li. Beale–Kato–Madja type criteria of smooth solutions to 3D Hall–MHD flows. Applied Mathematics and Computation, 286(??):41–48, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302478.

Wenqiong:2016:HDR

[WL16d] Liu Wenqiong and Shenghong Li. Hedging default risks of CDO tranches in nonhomogeneous Markovian contagion models. Applied Mathematics and Computation, 291(??):279–291, December 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316304295.

Wen:2017:TSI

[WL17] Rui-Ping Wen and Li-Xia Liu. The two-stage iteration algorithms based on the shortest distance for low-rank matrix completion. Applied Mathematics and Computation, 314(??):133–141, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304757.

Wang:2018:NO

[WL18a] Haijun Wang and Xianyi Li. A note on “Hopf bifurcation analysis and ultimate bound estimation of a new 4-D quadratic autonomous hyper-chaotic system” in [Appl. Math. Comput. 291 (2016) 323–339] by Amin Zarei and Saeed Tavakoli. Applied Mathematics and Computation, 329(??):1–4, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300481. See

[ZT16a].

Wu:2018:PSG

[WL18b] Tingzeng Wu and Hong-Jian Lai. On the permanental sum of graphs.

Applied Mathematics and Computation, 331(??): 334–340, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318301917.

Wang:2019:RBC

[WL19a] Renhai Wang and Yangrong Li. Regularity and backward compactness of attractors for nonautonomous lattice systems with random coefficients. Applied Mathematics and Computation, 354(??):86–102, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301407.

Wang:2019:RPS

[WL19b] Xianjia Wang and Shaojie Lv. The roles of particle swarm intelligence in the prisoner’s dilemma based on continuous and mixed strategy systems on scalefree networks. Applied Mathematics and Computation, 355(??):213–220, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300319301523.

Weng:2019:IMM

[WL19c] Lihui Weng and Wei Liu. Invariant measures of the Milstein method for stochastic differential equations with commutative noise. Applied Mathematics and Computation, 358(??):169–176, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303376.

Wu:2017:ETS

[WLFH17] Jilian Wu, Demin Liu, Xinlong Feng, and Pengzhan Huang. An efficient two-step algorithm for the stationary incompressible magnetohydrodynamic equations. Applied Mathematics and Computation, 302(??):21– 33, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300139.

Wu:2015:PFO

[WLFX15] Lifeng Wu, Sifeng Liu, Zhigeng Fang, and Haiyan Xu. Properties of the GM(1, 1) with fractional order accumulation. Applied Mathematics and Computation, 252(??): 287–293, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314016683.

Wang:2017:NPI

[WLG17] Xia Wang, Shengqiang Liu, and Hongjian Guo. A nosocomial-pathogens-infections model with impulsive antibiotics treatment on multiple bacteria. Applied Mathematics and Computation, 296(??):64– 87, March 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316306117.

Wang:2018:DFT

[WLH^{+}18] Jing Wang, Kun Liang, Xia Huang, Zhen Wang, and Hao Shen. Dissipative fault-tolerant control for nonlinear singular perturbed systems with Markov jumping parameters based on slow state feedback. Applied Mathematics and Computation, 328(??):247–262, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300729.

Wang:2015:NII

[WLJ15] Weizhe Wang, Yingzheng Liu, and Puning Jiang. Numerical investigation on influence of real gas properties on nonlinear behavior of labyrinth sealrotor system. Applied Mathematics and Computation, 263(??):12–24, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500452X.

Wu:2016:MLE

[WLL16] Jiancheng Wu, Xuan Li, and Xianbin Liu. The moment Lyapunov exponent of a co-dimension two bifurcation system driven by non-Gaussian colored noise. Applied Mathematics and Computation, 286(??):189–200, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302430.

Wu:2019:SPP

[WLL^{+}19] Zeyan Wu, Jianjuan Li, Shuying Liu, Liuting Zhou, and Yang Luo. A spatial predator–prey system with non-renewable resources. Applied Mathematics and Computation, 347(??):381–391, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310130.

Wang:2017:VUM

[WLLK17] Na-Na Wang, Ji-Cheng Li, Guo Li, and Xu Kong. Variants of the Uzawa method for three-order block saddle point problem. Applied Mathematics and Computation, 305(??):188–202, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300693.

Wen:2015:RPP

[WLLL15] Haijun Wen, Mingzhou Liu, Changyi Liu, and Conghu Liu. Remanufacturing production planning with compensation function approximation method. Applied Mathematics and Computation, 256(??):742–753, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001022.

Wang:2019:HOW

[WLQ^{+}19] Xiufang Wang, Gang Li, Shouguo Qian, Jiaojiao Li, and Zhen Wang. High order well-balanced finite difference WENO schemes for shallow water flows along channels with irregular geometry. Applied Mathematics and Computation, 363 (??):Article 124587, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930579X.

Wang:2018:ESI

[WLQZ18] Guangfu Wang, Shuchao Li, Dongchao Qi, and Huihui Zhang. On the edge-Szeged index of unicyclic graphs with given diameter. Applied Mathematics and Computation, 336(??):94–106, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830403X.

Wang:2015:WHV

[WLS15] Xia Wang, Shengqiang Liu, and Xinyu Song. A within-host virus model with multiple infected stages under time-varying environments. Applied Mathematics and Computation, 266(??):119–134, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315006505.

Wang:2018:SOS

[WLWA18] Liqing Wang, Yang Liu, Zhengguang Wu, and Fuad E. Alsaadi. Strategy optimization for static games based on STP method. Applied Mathematics and Computation, 316(??):390–399, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730574X.

Wu:2018:SCC

[WLWD18] Fengyan Wu, Dongfang Li, Jinming Wen, and Jinqiao Duan. Stability and convergence of compact finite difference method for parabolic problems with delay. Applied Mathematics and Computation, 322(??):129–139, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308147.

Wang:2018:ICI

[WLX18] Juan Wang, Chao Li, and Chengyi Xia. Improved centrality indicators to characterize the nodal spreading capability in complex networks. Applied Mathematics and Computation, 334(??):388–400, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303461.

Wei:2015:NBA

[WLZ15] Junqiang Wei, Gengyin Li, and Ming Zhou. Numerical bifurcation and its application in computation of available transfer capability. Applied Mathematics and Computation, 252(??):568–574, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016488.

Wang:2016:GDS

[WLZ16] Lianwen Wang, Zhijun Liu, and Xingan Zhang. Global dynamics of an SVEIR epidemic model with distributed delay and nonlinear incidence. Applied Mathematics and Computation, 284(??):47– 65, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301771.

Wang:2019:IBP

[WLZ19a] Chaojie Wang, Hongyi Li, and Di Zhao. Improved block preconditioners for linear systems arising from halfquadratic image restoration. Applied Mathematics and Computation, 363 (??):Article 124614, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930606X.

Wen:2019:TMC

[WLZ19b] Rui-Ping Wen, Shu-Zhen Li, and Fang Zhou. Toeplitz matrix completion via smoothing augmented Lagrange multiplier algorithm. Applied Mathematics and Computation, 355(??):299–310, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300319301316.

Wang:2015:PSA

[WLZW15] Ying Wang, Lishan Liu, Xinguang Zhang, and Yonghong Wu. Positive solutions of an abstract fractional semipositone differential system model for bioprocesses of HIV infection. Applied Mathematics and Computation, 258(??):312–324, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001125.

Wang:2017:RFT

[WLZY17] Guoliang Wang, Zhiqiang Li, Qingling Zhang, and Chunyu Yang. Robust finite-time stability and stabilization of uncertain Markovian jump systems with time-varying delay. Applied Mathematics and Computation, 293(??):377–393, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305392.

Wang:2015:SNR

[WM15] Rong-Nian Wang and Qing-Hua Ma. Some new results for multivalued fractional evolution equations. Applied Mathematics and Computation, 257(??):285–294, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314011291.

Wang:2018:ECDb

[WM18a] Shiying Wang and Xiaolei Ma. The gextra connectivity and diagnosability of crossed cubes. Applied Mathematics and Computation, 336(??):60–66, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303722.

Wang:2018:HCV

[WM18b] Wei Wang and Wanbiao Ma. Hepatitis C virus infection is blocked by HMGB1: A new nonlocal and time-delayed reaction–diffusion model. Applied Mathematics and Computation, 320(??): 633–653, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300317306744.

Wang:2017:ETS

[WMLG17] Hanquan Wang, Xiu Ma, Junliang Lu, and Wen Gao. An efficient time-splitting compact finite difference method for Gross–Pitaevskii equation. Applied Mathematics and Computation, 297(??):131–144, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306440.

Wang:2017:RNN

[WMS17] Xue-Zhong Wang, Haifeng Ma, and Predrag S. Stanimirović. Recurrent neural network for computing the W -weighted Drazin inverse. Applied Mathematics and Computation, 300(??):1–20, May 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307068.

Wongsaijai:2019:CSP

[WMSP19] B. Wongsaijai, T. Mouktonglang, N. Sukantamala, and K. Poochinapan. Compact structurepreserving approach to solitary wave in shallow water modeled by the Rosenau–RLW equation. Applied Mathematics and Computation, 340(??):84–100, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304971.

Wang:2016:SAS

[WMZ16] Jimin Wang, Shuping Ma, and Chenghui Zhang. Stability analysis and stabilization for nonlinear continuous-time descriptor semi-Markov jump systems. Applied Mathematics and Computation, 279(??):90–102, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300133.

Wu:2019:NNM

[WMZ19] Fuqiang Wu, Jun Ma, and Ge Zhang. A new neuron model under electromagnetic field. Applied Mathematics and Computation, 347(??):590–599, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318309676.

Withers:2015:SLR

[WN15] Christopher S. Withers and Saralees Nadarajah. Solutions of linear recurrence equations. Applied Mathematics and Computation, 271(??):768–776, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031501317X.

Watanabe:2016:NVM

[WN16] Yoshitaka Watanabe and Mitsuhiro T. Nakao. A numerical verification method for nonlinear functional equations based on infinitedimensional Newton-like iteration. Applied Mathematics and Computation, 276(??):239–251, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300096.

Wang:2019:DCC

[WNF19] Gang Wang, Min-Yao Niu, and Fang-Wei Fu. Deterministic constructions of compressed sensing matrices based on optimal codebooks and codes. Applied Mathematics and Computation, 343(??):128–136, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308208.

Wang:2015:ASM

[WNMC15] Xiao Wang, Yufu Ning, Tauqir A. Moughal, and Xiumei Chen. Adams– Simpson method for solving uncertain differential equation. Applied Mathematics and Computation, 271(??):209–219, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012278.

Waseem:2016:EMS

[WNN16] Muhammad Waseem, Muhammad Aslam Noor, and Khalida Inayat Noor. Efficient method for solving a system of nonlinear equations. Applied Mathematics and Computation, 275(??):134–146, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015635.

Wu:2019:FTS

[WNW^{+}19] Kai-Ning Wu, Ming-Ye Na, Liming Wang, Xiaohua Ding, and Boying Wu. Finite-time stability of impulsive reaction– diffusion systems with and without time delay. Applied Mathematics and Computation, 363 (??):Article 124591, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305831.

Wong:2015:EGC

[Won15] Patricia J. Y. Wong.

Eigenvalues of a general class of boundary value problem with derivative-dependent nonlinearity. Applied Mathematics and Computation, 259(??):908–930, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002878.

Wang:2015:SHB

[WPL15] Xuedi Wang, Miao Peng, and Xiuyu Liu. Stability and Hopf bifurcation analysis of a ratio-dependent predator–prey model with two time delays and Holling type III functional response. Applied Mathematics and Computation, 268(??):496–508, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008899.

Wang:2015:NSC

[WQ15] Hui Wang and Qing-Hua Qin. A new special coating/fiber element for analyzing effect of interface on thermal conductivity of composites. Applied Mathematics and Computation, 268(??):311–321, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008589.

Wang:2015:NSP

[WQK15] Dan Wang, Zhongfeng Qin, and Samarjit Kar. A novel single-period inventory problem with uncertain random demand and its application. Applied Mathematics and Computation, 269(??):133–145, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315008838.

Wang:2015:CDQ

[WqTjT15] Xing Wang, Chang qi Tao, and Guo ji Tang. A class of differential quadratic programming problems. Applied Mathematics and Computation, 270(??): 369–377, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315010966.

Waeytens:2015:CDP

[WR15] Julien Waeytens and Bojana Rosić. Comparison of deterministic and probabilistic approaches to identify the dynamic moving load and damages of a reinforced concrete beam. Applied Mathematics and Computation, 267(??):3–16, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011182.

Wang:2019:HOC

[WR19] Yuan-Ming Wang and Lei Ren. A high-order L_{2}- compact difference method for Caputo-type timefractional sub-diffusion equations with variable coefficients. Applied Mathematics and Computation, 342(??):71–93, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307793.

Wang:2016:BMO

[WRW16] Shaoli Wang, Libin Rong, and Jianhong Wu. Bistability and multistability in opinion dynamics models. Applied Mathematics and Computation, 289(??):388–395, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303393.

Wang:2019:FTSb

[WRX^{+}19] Jing Wang, Tingting Ru, Jianwei Xia, Yunliang Wei, and Zhen Wang. Finitetime synchronization for complex dynamic networks with semi-Markov switching topologies: an H_{∞} event-triggered control scheme. Applied Mathematics and Computation, 356(??):235–251, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300319302449.

Wang:2019:OIB

[WRZ19] Suxin Wang, Ximin Rong, and Hui Zhao. Optimal investment and benefit payment strategy under loss aversion for target benefit pension plans. Applied Mathematics and Computation, 346(??):205–218, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308968.

Wang:2016:PPM

[WS16] Guangbin Wang and Deyu Sun. Preconditioned parallel multisplitting USAOR method for H-matrices linear systems. Applied Mathematics and Computation, 275(??):156–164, February 15, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315015702.

Wang:2019:CHM

[WS19a] Xin Wang and Housheng Su. Consensus of hybrid multi-agent systems by event-triggered/self-triggered strategy. Applied Mathematics and Computation, 359(??):490–501, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319303819.

Wang:2019:IIT

[WS19b] Yu Ping Wang and Chung Tsun Shieh. The inverse interior transmission eigenvalue problem with mixed spectral data. Applied Mathematics and Computation, 343(??):285–298, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318307860.

Wu:2019:UGS

[WS19c] Tingzeng Wu and Wasin So. Unicyclic graphs with second largest and second smallest permanental sums. Applied Mathematics and Computation, 351(??):168–175, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300712.

Wang:2019:NEA

[WSC19] Chao Wang, Zhengjia Sun, and Jintao Cui. A new error analysis of a mixed finite element method for the quad–curl problem. Applied Mathematics and Computation, 349(??):23– 38, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310798.

Wang:2015:BBS

[WSdSY15] Liumei Wang, Wenyu Sun, Raimundo J. B. de Sampaio, and Jinyun Yuan. A Barzilai and Borwein scaling conjugate gradient method for unconstrained optimization problems. Applied Mathematics and Computation, 262(??):136–144, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005044.

Wang:2018:SSS

[WSH^{+}18] Jun Wang, Kaibo Shi, Qinzhen Huang, Shouming Zhong, and Dian Zhang. Stochastic switched sampleddata control for synchronization of delayed chaotic neural networks with packet dropout. Applied Mathematics and Computation, 335(??):211–230, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303564.

Wang:2017:FDL

[WSLC17] Huili Wang, Baochang Shi, Hong Liang, and Zhenhua Chai. Finitedifference lattice Boltzmann model for nonlinear convection-diffusion equations. Applied Mathematics and Computation, 309(??):334–349, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031730259X.

Wang:2019:OCC

[WSXR19] Xia Wang, Mingwang Shen, Yanni Xiao, and Libin Rong. Optimal control and costeffectiveness analysis of a Zika virus infection model with comprehensive interventions. Applied Mathematics and Computation, 359(??):165–185, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031930308X.

Wu:2018:FTB

[WSYL18] Kai-Ning Wu, Han-Xiao Sun, Baoqing Yang, and Cheng-Chew Lim. Finitetime boundary control for delay reaction-diffusion systems. Applied Mathematics and Computation, 329(??):52–63, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300717.

Wang:2016:SAE

[WSZ16] Haijin Wang, Chi-Wang Shu, and Qiang Zhang. Stability analysis and error estimates of local discontinuous Galerkin methods with implicit-explicit time-marching for nonlinear convection-diffusion problems. Applied Mathematics and Computation, 272 (part 2)(??):237–258, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002672.

Wu:2018:TIA

[WSZ18] Ai-Guo Wu, Hui-Jie Sun, and Ying Zhang. Two iterative algorithms for stochastic algebraic Riccati matrix equations. Applied Mathematics and Computation, 339(??): 398–409, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318305976.

Wan:2019:GEB

[WTD^{+}19] Pengfei Wan, Jianhua Tu, Matthias Dehmer, Shenggui Zhang, and Frank Emmert-Streib. Graph entropy based on the number of spanning forests of c-cyclic graphs. Applied Mathematics and Computation, 363(??): Article 124616, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306083.

Wang:2017:SFF

[WTJ17] Yu-Feng Wang, Bo Tian, and Yan Jiang. Soliton fusion and fission in a generalized variable-coefficient fifth-order Korteweg–de Vries equation in fluids. Applied Mathematics and Computation, 292(??):448–456, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304635.

Wang:2015:BDS

[WTSQ15] Pan Wang, Bo Tian, Kun Sun, and Feng-Hua Qi. Bright and dark soliton solutions and Bäcklund transformation for the Eckhaus–Kundu equation with the cubic-quintic nonlinearity. Applied Mathematics and Computation, 251(??):233–242, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015203.

Wu:2015:GOC

[WTUZ15] Zhiyou Wu, Jing Tian, Julien Ugon, and Liang Zhang. Global optimality conditions and optimization methods for constrained polynomial programming problems. Applied Mathematics and Computation, 262(??): 312–325, July 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315004981.

Wang:2016:IQP

[WTXZ16] Xiu-Bin Wang, Shou-Fu Tian, Mei-Juan Xua, and Tian-Tian Zhang. On integrability and quasiperiodic wave solutions to a (3 + 1)-dimensional generalized KdV-like model equation. Applied Mathematics and Computation, 283(??):216–233, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301394.

Wan:2018:CNI

[WTZL18] Pengfei Wan, Jianhua Tu, Shenggui Zhang, and Binlong Li. Computing the numbers of independent sets and matchings of all sizes for graphs with bounded treewidth. Applied Mathematics and Computation, 332(??):42–47, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301826.

Wu:2015:SDD

[Wu15] Zhaoyan Wu. Synchronization of discrete dynamical networks with non-delayed and delayed coupling. Applied Mathematics and Computation, 260(??):57– 62, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003586.

Wu:2017:EPA

[Wu17] Shulin Wu. An efficient parareal algorithm for a class of time-dependent problems with fractional Laplacian. Applied Mathematics and Computation, 307(??):329–341, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301169.

Wu:2018:NDD

[WV18] Ruijia Wu and Robert A.

Van Gorder. Nonlinear dynamics of discrete time multi-level leaderfollower games. Applied Mathematics and Computation, 320(??):240–250, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306628.

Wang:2015:SPP

[WW15] Yuanshi Wang and Hong Wu. Stability of plant– pollinator–ant co-mutualism. Applied Mathematics and Computation, 261(??):

231–241, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315003756.

Walawska:2016:IAV

[WW16a] Irmina Walawska and Daniel Wilczak. An implicit algorithm for validated enclosures of the solutions to variational equations for ODEs. Applied Mathematics and Computation, 291(??):303–322, December 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316304350.

Wan:2016:BSF

[WW16b] Youyan Wan and Zhengping Wang. Bound state for fractional Schrödinger equation with saturable nonlinearity. Applied Mathematics and Computation, 273(??):735–740, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013934.

Wang:2016:TIB

[WW16c] Shikun Wang and Yuanshi Wang. Transition of interactions between a cuckoo and its host: Fluctuating between parasitism and mutualism. Applied Mathematics and Computation, 273(??):664–677, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014186.

Wang:2017:GHP

[WW17] Weiping Wang and Hui Wang. Generalized Humbert polynomials via generalized Fibonacci polynomials. Applied Mathematics and Computation, 307(??):204–216, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301613.

Wang:2018:SEE

[WW18] Kun Wang and Hongyue Wang. Stability and error estimates of a new high-order compact ADI method for the unsteady 3D convection-diffusion equation. Applied Mathematics and Computation, 331(??):140–159, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301632.

Wei:2019:MRJ

[WW19] Zhaoying Wei and Guangsheng Wei. A method for recovering Jacobi matrices with mixed spectral data. Applied Mathematics and Computation, 359(??):426–432, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319303388.

Wu:2015:PDC

[WWC^{+}15] Jie Wu, Xiuling Wang, Jinjin Chen, Gang Shu, and Ya Li. The position of a door can significantly impact on pedestrians’ evacuation time in an emergency. Applied Mathematics and Computation, 258(??):29–35, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001447.

Wei:2018:PCM

[WWC^{+}18] Wei Wei, Jun Wang, Xuanyi Chen, Jing Yang, and Xiaowei Min. Psychological contract model for knowledge collaboration in virtual community of practice: An analysis based on the game theory. Applied Mathematics and Computation, 329(??):175–187, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300766.

Wang:2017:KTM

[WWDW17] Haiying Wang, Jun Wang, Liting Ding, and Wei Wei. Knowledge transmission model with consideration of self-learning mechanism in complex networks. Applied Mathematics and Computation, 304(??):83– 92, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300280.

Wang:2016:ITD

[WWH16] Ligong Wang, Qi Wang, and Bofeng Huo. Integral trees with diameter four. Applied Mathematics and Computation, 282(??):53– 64, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300868.

Wang:2017:SHE

[WWH17] Kun Wang, Yau Shu Wong, and Jizu Huang. Solving Helmholtz equation at high wave numbers in exterior domains. Applied Mathematics and Computation, 298(??): 221–235, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031630683X.

Wu:2018:DAF

[WWHZ18] Li-Bing Wu, Heng Wang, Xi-Qin He, and Da-Qing Zhang. Decentralized adaptive fuzzy tracking control for a class of uncertain large-scale systems with actuator nonlinearities. Applied Mathematics and Computation, 332(??):390–405, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302613.

Wang:2019:AAS

[WWK19] Jinliang Wang, Jing Wang, and Toshikazu Kuniya. Analysis of an age-structured multi-group heroin epidemic model. Applied Mathematics and Computation, 347(??):78–100, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309858.

Wang:2018:EMZ

[WWL18] Shaohui Wang, Chunxiang Wang, and Jia-Bao Liu. On extremal multiplicative Zagreb indices of trees with given domination number. Applied Mathematics and Computation, 332(??):338–350, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302315.

Wen:2016:MLE

[WWMX16] Jianghui Wen, Xiangjun Wang, Shuhua Mao, and Xinping Xiao. Maximum likelihood estimation of McKean–Vlasov stochastic differential equation and its application. Applied Mathematics and Computation, 274(??):237–246, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014812.

Wang:2015:IKS

[WWN^{+}15] Xu-Wen Wang, Zhen Wang, Sen Nie, Luo-Luo Jiang, and Bing-Hong Wang. Impact of keeping silence on spatial reciprocity in spatial games. Applied Mathematics and Computation, 250(??):848–853, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015458.

Wang:2018:ERD

[WwQqTW18] Xing Wang, Ya wei Qi, Chang qi Tao, and Qi Wu. Existence result for differential variational inequality with relaxing the convexity condition. Applied Mathematics and Computation, 331(??):297–306, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301693.

Weinbub:2015:VDM

[WWR^{+}15] Josef Weinbub, Matthias Wastl, Karl Rupp, Florian Rudolf, and Siegfried Selberherr. ViennaMaterials — a dedicated material library for computational science and engineering. Applied Mathematics and Computation, 267(??):282–293, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004087.

Wang:2019:RMP

[WWSM19] Haiying Wang, Jun Wang, Michael Small, and Jack Murdoch Moore. Review mechanism promotes knowledge transmission in complex networks. Applied Mathematics and Computation, 340(??):113–125, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306179.

Wang:2015:INR

[WWW15] Yuanshi Wang, Hong Wu, and Shikun Wang. Invasibility of nectar robbers in pollination-mutualisms. Applied Mathematics and Computation, 250(??): 908–919, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314015495.

Wang:2017:IRE

[WWW^{+}17] Chengjiang Wang, Li Wang, Juan Wang, Shiwen Sun, and Chengyi Xia. Inferring the reputation enhances the cooperation in the public goods game on interdependent lattices. Applied Mathematics and Computation, 293(??):18–29, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303976.

Wang:2018:EPS

[WWW18] Yiqiao Wang, Weifan Wang, and Ying Wang. Edge-partition and star chromatic index. Applied Mathematics and Computation, 333(??):480–489, September 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318302601.

Wang:2019:SEC

[WWW19] Ying Wang, Yiqiao Wang, and Weifan Wang. Star edge-coloring of graphs with maximum degree four. Applied Mathematics and Computation, 340(??):268–275, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307495.

Wu:2015:RRC

[WWZ15] Jinming Wu, Renhong Wang, and Xiaolei Zhang. Real root classification of parametric spline functions. Applied Mathematics and Computation, 250(??):212–220, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014696.

Wang:2019:OAA

[WWZ19] Qi Wang, Jiechang Wen, and Ping Zhang. Oscillation analysis of advertising capital model: Analytical and numerical studies. Applied Mathematics and Computation, 354(??):365–376, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930133X.

Wang:2016:MSS

[WWZD16] Liming Wang, Kai-Ning Wu, Ya-Nan Zhu, and Xiaohua Ding. Mean square H_{∞} synchronization of coupled stochastic partial differential systems. Applied Mathematics and Computation, 275(??):386–393, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015775.

Wei:2016:IEA

[WX16] Tao Wei and Mingtian Xu. An integral equation approach to the unsteady convectiondiffusion equations. Applied Mathematics and Computation, 274(??):55– 64, February 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315014356.

Wei:2018:GNC

[WX18] Yulong Wei and Min Xu. The 1, 2-good-neighbor conditional diagnosabilities of regular graphs. Applied Mathematics and Computation, 334(??):295–310, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300318303266.

Wang:2019:CFS

[WX19a] Junjie Wang and Aiguo Xiao. Conservative Fourier spectral method and numerical investigation of space fractional Klein– Gordon–Schrödinger equations. Applied Mathematics and Computation, 350(??):348–365, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031831107X.

Wang:2019:EEI

[WX19b] Wen-Huan Wang and Yi-Sai Xue. Extremal Estrada indices of the weighted trees with fixed total weight sum. Applied Mathematics and Computation, 354(??):32–41, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301171.

Wang:2019:SBD

[WXLL19] Zhen Wang, Yingkang Xie, Junwei Lu, and Yuxia Li. Stability and bifurcation of a delayed generalized fractional-order prey-predator model with interspecific competition. Applied Mathematics and Computation, 347(??): 360–369, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318309895.

Wei:2018:NSG

[WXM^{+}18] Na Wei, Chaoyang Xu, Yingfeng Meng, Gao Li, Xiao Ma, and Anqi Liu. Numerical simulation of gas-liquid two-phase flow in wellbore based on drift flux model. Applied Mathematics and Computation, 338(??):164–174, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302406.

Wang:2016:MSH

[WXP16] Jinyan Wang, Yanni Xiao, and Zhihang Peng. Modelling seasonal HFMD infections with the effects of contaminated environments in mainland China. Applied Mathematics and Computation, 274(??):615–627, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315015258.

Wang:2018:NAT

[WXW18] Jiao Wang, Tian-Zhou Xu, and Gang-Wei Wang. Numerical algorithm for timefractional Sawada–Kotera equation and Ito equation with Bernstein polynomials. Applied Mathematics and Computation, 338(??):ii, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304892.

Wang:2018:NSC

[WXWX18] Jiao Wang, Tian-Zhou Xu, Yan-Qiao Wei, and Jia-Quan Xie. Numerical simulation for coupled systems of nonlinear fractional order integrodifferential equations via wavelets method. Applied Mathematics and Computation, 324(??):36–50, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730869X.

Wang:2018:FTS

[WXX18] Ronghao Wang, Jianchun Xing, and Zhengrong Xiang. Finite-time stability and stabilization of switched nonlinear systems with asynchronous switching. Applied Mathematics and Computation, 316(??):229–244, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305684.

Wang:2019:TVF

[WXXY19] Jinhuan Wang, Yuling Xu, Yong Xu, and Dedong Yang. Timevarying formation for highorder multi-agent systems with external disturbances by event-triggered integral sliding mode control. Applied Mathematics and Computation, 359(??):333–343, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303546.

Wang:2015:MNE

[WXY15] Dongling Wang, Aiguo Xiao, and Wei Yang. Maximum-norm error analysis of a difference scheme for the space fractional CNLS. Applied Mathematics and Computation, 257(??):241–251, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015483.

Wang:2018:IBP

[WXZ18] Chengzhi Wang, Gang Xue, and Caidi Zhao. Invariant Borel probability measures for discrete longwave-short-wave resonance equations. Applied Mathematics and Computation, 339(??):846–852, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305484.

Wang:2015:PMM

[WY15a] Chuan-Long Wang and Xi-Hong Yan. Parallel multisplitting methods with optimal weighting matrices for linear systems. Applied Mathematics and Computation, 259(??):523–532, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315003306.

Wen:2015:NGP

[WY15b] Rui-Ping Wen and Xi-Hong Yan. A new gradient projection method for matrix completion. Applied Mathematics and Computation, 258(??):537–544, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002210.

Wen:2016:CRK

[WY16] Liping Wen and Yuexin Yu. Convergence of Runge–Kutta methods for neutral delay integrodifferential equations. Applied Mathematics and Computation, 282(??):84– 96, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300856.

Wang:2017:GNE

[WY17] Shiying Wang and Yuxing Yang. The 2-good-neighbor (2-extra) diagnosability of alternating group graph networks under the PMC model and MM* model. Applied Mathematics and Computation, 305(??):241–250, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301029.

Wang:2018:QSF

[WY18] Fei Wang and Yongqing Yang. Quasi-synchronization for fractional-order delayed dynamical networks with heterogeneous nodes. Applied Mathematics and Computation, 339(??):ii, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306064.

Wang:2019:FBS

[WY19] Bingjun Wang and Mingxia Yuan. Forward-backward stochastic differential equations driven by G-Brownian motion. Applied Mathematics and Computation, 349(??):39–47, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031831083X.

Wang:2017:NCS

[WYCZ17] Bo Wang, Juan Yan, Jun Cheng, and Shouming Zhong. New criteria of stability analysis for generalized neural networks subject to time-varying delayed signals. Applied Mathematics and Computation, 314(??):322–333, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304447.

Wei:2018:NTD

[WYD^{+}18] Yikun Wei, Hui Yang, Hua-Shu Dou, Zhe Lin, Zhengdao Wang, and Yuehong Qian. A novel two-dimensional coupled lattice Boltzmann model for thermal incompressible flows. Applied Mathematics and Computation, 339(??):546–555, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306131.

Wang:2015:EBF

[WYL15] Zhiqian Wang, Zhuangpeng Yi, and Yingshe Luo. Energy-based formulation for nonlinear normal modes in cable-stayed beam. Applied Mathematics and Computation, 265(??):176–186, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500627X.

Wei:2019:SAE

[WYL19a] Jia Wei, Zhifu You, and Hong-Jian Lai. Spectral analogues of Erdős’ theorem on Hamiltonconnected graphs. Applied Mathematics and Computation, 340(??):242–250, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306441.

Wei:2019:PAC

[WYL^{+}19b] Zongtian Wei, Chao Yue, Yinkui Li, Hongyun Yue, and Yong Liu. A polynomial algorithm for computing the weak rupture degree of trees. Applied Mathematics and Computation, 361(??):730–734, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319304849.

Wang:2019:HFV

[WYLG19] Xiufang Wang, Haiyan Yu, Gang Li, and Jinmei Gao. Hybrid finite volume weighted essentially non-oscillatory schemes with linear central reconstructions. Applied Mathematics and Computation, 359(??):132–147, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303078.

Wan:2016:POS

[WYW16] Long Wan, Jinjiang Yuan, and Lijun Wei. Pareto optimization scheduling with two competing agents to minimize the number of tardy jobs and the maximum cost. Applied Mathematics and Computation, 273(??):912–923, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014101.

Wu:2017:EWS

[WYZ^{+}17] Chufen Wu, Yong Yang, Qianyi Zhao, Yanling Tian, and Zhiting Xu. Epidemic waves of a spatial SIR model in combination with random dispersal and non-local dispersal. Applied Mathematics and Computation, 313(??):122–143, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303806.

Wang:2015:NIV

[WZ15a] JinRong Wang and Yuruo Zhang. Nonlocal initial value problems for differential equations with Hilfer fractional derivative. Applied Mathematics and Computation, 266(??):850–859, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007754.

Wang:2015:CFV

[WZ15b] Tongke Wang and Zhiyue Zhang. A compact finite volume method and its extrapolation for elliptic equations with third boundary conditions. Applied Mathematics and Computation, 264(??):258–271, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005494.

Wang:2016:IDF

[WZ16a] Peng Wang and Detong Zhu. An inexact derivativefree Levenberg–Marquardt method for linear inequality constrained nonlinear systems under local error bound conditions. Applied Mathematics and Computation, 282(??):32–52, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300728.

Wang:2016:SMS

[WZ16b] Yanxin Wang and Li Zhu. SCW method for solving the fractional integrodifferential equations with a weakly singular kernel. Applied Mathematics and Computation, 275(??):72–80, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015593.

Wang:2018:OGA

[WZ18a] Tao Wang and Xiaodan Zhao. Odd graph and its applications to the strong edge coloring. Applied Mathematics and Computation, 325(??):246–251, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308469.

Wu:2018:LQP

[WZ18b] Zhen Wu and Yi Zhuang.

Linear-quadratic partially observed forward-backward stochastic differential games and its application in finance. Applied Mathematics and Computation, 321(??):577–592, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730797X.

Wu:2019:SPA

[WZ19] Xiaoxia Wu and Lianzhu Zhang. On structural properties of ABC-minimal chemical trees. Applied Mathematics and Computation, 362(??): Article 124570, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305533.

Wang:2015:GPT

[WZAF15] Zhihui Wang, Junyong Zhai, Weiqing Ai, and Shumin Fei. Global practical tracking for a class of uncertain nonlinear systems via sampleddata control. Applied Mathematics and Computation, 260(??):257–268, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003884.

Wu:2017:STR

[WZC17] Xiaoxia Wu, Lianzhu Zhang, and Haiyan Chen. Spanning trees and recurrent configurations of a graph. Applied Mathematics and Computation, 314(??):25–30, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304435.

Wang:2018:RML

[WZC^{+}18] Qiao Wang, Wei Zhou, Yonggang Cheng, Gang Ma, Xiaolin Chang, Yu Miao, and E. Chen. Regularized moving leastsquare method and regularized improved interpolating moving least-square method with nonsingular moment matrices. Applied Mathematics and Computation, 325(??):120–145, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308767.

Wang:2019:AOI

[WZF^{+}19] Qiao Wang, Wei Zhou, Y. T. Feng, Gang Ma, Yonggang Cheng, and Xiaolin Chang. An adaptive orthogonal improved interpolating moving leastsquare method and a new boundary elementfree method. Applied Mathematics and Computation, 353(??):347–370, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301110.

Wu:2015:ISC

[WZK15] Xiangjun Wu, Changjiang Zhu, and Haibin Kan. An improved secure communication scheme based passive synchronization of hyperchaotic complex nonlinear system. Applied Mathematics and Computation, 252(??):201–214, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016816.

Wang:2015:FDB

[WZL15] Aizeng Wang, Gang Zhao, and Yong-Dong Li. Fairness degree based fairness criterion and fairing algorithm. Applied Mathematics and Computation, 253(??):184–192, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017354.

Wen:2018:RMB

[WZLP18] Baolian Wen, Hua Zheng, Wen Li, and Xiaofei Peng. The relaxation modulusbased matrix splitting iteration method for solving linear complementarity problems of positive definite matrices. Applied Mathematics and Computation, 321(??):349–357, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307737.

Wang:2015:FEM

[WZM15] Jue Wang, Lei Zhang, and FuMing Ma. Finite element method with nonlocal boundary condition for solving the nondestructive testing problem of wood moisture content. Applied Mathematics and Computation, 250(??):432–443, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015549.

Wang:2019:SCV

[WZS19] Pengfei Wang, Wenqing Zou, and Huan Su. Stability of complexvalued impulsive stochastic functional differential equations on networks with Markovian switching. Applied Mathematics and Computation, 348(??):338–354, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310476.

Wei:2015:CVM

[WZWD15] Dandan Wei, Weiwei Zhang, Linghui Wang, and Baodong Dai. The complex variable meshless local Petrov–Galerkin method for elasticity problems of functionally graded materials. Applied Mathematics and Computation, 268(??):1140–1151, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009327.

Wang:2018:ASW

[WZWJ18] Xiuli Wang, Qilong Zhai, Ruishu Wang, and Rabeea Jari. An absolutely stable weak Galerkin finite element method for the Darcy–Stokes problem. Applied Mathematics and Computation, 331(??):20–32, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830136X.

Wei:2019:BAT

[WZY^{+}19] Zhouchao Wei, Bin Zhu, Jing Yang, Matjaz Perc, and Mitja Slavinec. Bifurcation analysis of two disc dynamos with viscous friction and multiple time delays. Applied Mathematics and Computation, 347(??):265–281, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309706.

Wang:2015:SCT

[WZYS15] Guoliang Wang, Qingling Zhang, Chunyu Yang, and Chengli Su. Stabilization of continuous-time randomly switched systems via the LMI approach. Applied Mathematics and Computation, 266(??):527–538, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006852.

Wang:2018:CCB

[WZZ18] Huiru Wang, Chengjian Zhang, and Yongtao Zhou. A class of compact boundary value methods applied to semi-linear reactiondiffusion equations. Applied Mathematics and Computation, 325(??):69– 81, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308925.

Wang:2016:FTS

[WZZL16] Yijing Wang, Yanchao Zou, Zhiqiang Zuo, and Hongchao Li. Finite-time stabilization of switched nonlinear systems with partial unstable modes. Applied Mathematics and Computation, 291(??): 172–181, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300316303861.

Wei:2019:SCC

[WZZT19] Jingdong Wei, Jiangbo Zhou, Zaili Zhen, and Lixin Tian. Supercritical and critical traveling waves in a twocomponent lattice dynamical model with discrete delay. Applied Mathematics and Computation, 363 (??):Article 124621, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306137.

Xu:2019:IEP

[XBC19] Wei-Ru Xu, Natália Bebiano, and Guo-Liang Chen. An inverse eigenvalue problem for pseudo-Jacobi matrices. Applied Mathematics and Computation, 346(??):423–435, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309172.

Xu:2015:CFC

[XC15] Ai-Min Xu and Zhong-Di Cen. Closed formulas for computing higherorder derivatives of functions involving exponential functions. Applied Mathematics and Computation, 270(??):136–141, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315011066.

Xiong:2018:NIA

[XCCL18] Lianglin Xiong, Jun Cheng, Jinde Cao, and Zixin Liu. Novel inequality with application to improve the stability criterion for dynamical systems with two additive timevarying delays. Applied Mathematics and Computation, 321(??):672–688, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308020.

Xiao:2019:AFC

[XCDZ19] Wenbin Xiao, Liang Cao, Guowei Dong, and Qi Zhou. Adaptive fuzzy control for pure-feedback systems with full state constraints and unknown nonlinear dead zone. Applied Mathematics and Computation, 343(??):354–371, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307884.

Xiong:2016:MRM

[XCHW16] Xiangtuan Xiong, Xiaoxiao Cao, Shumei He, and Jin Wen. A modified regularization method for a Cauchy problem for heat equation on a two-layer sphere domain. Applied Mathematics and Computation, 290(??):240–249, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303757.

Xu:2018:QFC

[XCL18] Yao Xu, Chenyin Chu, and Wenxue Li. Quantized feedback control scheme on coupled systems with time delay and distributed delay: a finite-time inner synchronization analysis. Applied Mathematics and Computation, 337(??):302–314, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304314.

Xiong:2019:PCN

[XCW19] Shixun Xiong, Mou Chen, and Qianxian Wu. Predictive control for networked switch flight system with packet dropout. Applied Mathematics and Computation, 354(??):444–459, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930013X.

Xiong:2018:DRN

[XCY18] Jun Xiong, Xiao-Heng Chang, and Xiaojian Yi. Design of robust nonfragile fault detection filter for uncertain dynamic systems with quantization. Applied Mathematics and Computation, 338(??):758–773, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305101.

Xiao:2015:DIP

[XD15] Qizhen Xiao and Binxiang Dai. Dynamics of an impulsive predator–prey logistic population model with state-dependent. Applied Mathematics and Computation, 259(??): 220–230, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315002611.

Xi:2019:AFG

[XD19] Changjiang Xi and Jiuxiang Dong. Adaptive fuzzy guaranteed performance control for uncertain nonlinear systems with event-triggered input. Applied Mathematics and Computation, 363 (??):Article 124604, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930596X.

Xu:2019:PDS

[XDG^{+}19] Jiwei Xu, Zhenghong Deng, Bo Gao, Qun Song, Zhihong Tian, Qiuling Wang, Mingyu Gao, and Zhenxi Niu. Popularitydriven strategy updating rule promotes cooperation in the spatial prisoner’s dilemma game. Applied Mathematics and Computation, 353(??):82–87, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300773.

Xu:2019:ASM

[XDH19] Liguang Xu, Zhenlei Dai, and Hongxiao Hu. Almost sure and moment asymptotic boundedness of stochastic delay differential systems. Applied Mathematics and Computation, 361(??):157–168, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319304205.

Xu:2019:EIS

[XFTX19] Hedong Xu, Suohai Fan, Cunzhi Tian, and Xinrong Xiao. Evolutionary investor sharing game on networks. Applied Mathematics and Computation, 340(??):138–145, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307379.

Xu:2018:RBT

[XGG18] Kexiang Xu, Xiaoqian Gu, and Ivan Gutman. Relations between total irregularity and non-self-centrality of graphs. Applied Mathematics and Computation, 337(??): 452–460, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318304752.

Xu:2019:DOB

[XGQW19] Tianbo Xu, Xianwen Gao, Wenhai Qi, and Yunliang Wei. Disturbanceobserver-based control for semi-Markovian jump systems with generally uncertain transition rate and saturation nonlinearity. Applied Mathematics and Computation, 362 (??):Article 124569, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305521.

Xu:2017:GDA

[XGZ17] Jinhu Xu, Yan Geng, and Yicang Zhou. Global dynamics for an agestructured HIV virus infection model with cellular infection and antiretroviral therapy. Applied Mathematics and Computation, 305(??):62–83, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300905.

Xiao:2016:CMA

[XH16] Qiang Xiao and Zhenkun Huang. Consensus of multi-agent system with distributed control on time scales. Applied Mathematics and Computation, 277(??):54–71, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315300163.

Xiong:2019:CRQ

[XH19a] Yanqin Xiong and Jianqiang Hu. A class of reversible quadratic systems with piecewise polynomial perturbations. Applied Mathematics and Computation, 362(??): Article 124527, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305065.

Xu:2019:APE

[XH19b] Fei Xu and Qiumei Huang.

An accurate a posteriori error estimator for semilinear Neumann problem and its applications. Applied Mathematics and Computation, 362(??): Article 124540, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305193.

Xu:2015:TMA

[XHW15] Yan Xu, Zhimin He, and Peiguang Wang. p th moment asymptotic stability for neutral stochastic functional differential equations with Lévy processes. Applied Mathematics and Computation, 269(??):594–605, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009893.

Xie:2018:RSP

[XHW18] Liuliu Xie, Xiaotao Huang, and Lihe Wang. Radial symmetry for positive solutions of fractional p-Laplacian equations via constrained minimization method. Applied Mathematics and Computation, 337(??):41–53, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304375.

Xie:2017:EFG

[XJW17] Dan Xie, Kailin Jian, and Weibin Wen. An element-free Galerkin approach for rigid-flexible coupling dynamics in 2D state. Applied Mathematics and Computation, 310(??):149–168, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302242.

Xiao:2015:SSH

[XJZ^{+}15] Min Xiao, Guoping Jiang, Lindu Zhao, Wenying Xu, Youhong Wan, Chunxia Fan, and Zhengxin Wang. Stability switches and Hopf bifurcations of an isolated population model with delay-dependent parameters. Applied Mathematics and Computation, 264(??):99–115, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005330.

Xu:2019:SRU

[XJZ19] Pengfei Xu, Yanfei Jin, and Yanxia Zhang. Stochastic resonance in an underdamped triple-well potential system. Applied Mathematics and Computation, 346(??):352–362, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309263.

Xu:2015:ESD

[XKG15] Ruiping Xu, Yonggui Kao, and Cunchen Gao. Exponential synchronization of delayed Markovian jump complex networks with generally uncertain transition rates. Applied Mathematics and Computation, 271(??):682–693, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012680.

Xie:2015:DDR

[XKWG15] Jing Xie, Yonggui Kao, Changhong Wang, and Cunchen Gao. Delaydependent robust stability of uncertain neutraltype Itô stochastic systems with Markovian jumping parameters. Applied Mathematics and Computation, 251(??):576–585, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015823.

Xia:2015:MBM

[XL15] Zechen Xia and Chenliang Li. Modulus-based matrix splitting iteration methods for a class of nonlinear complementarity problem. Applied Mathematics and Computation, 271(??):34–42, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011649.

Xiao:2018:NPB

[XL18a] Chuanfu Xiao and Hanyu Li. New perturbation bounds of partitioned generalized Hermitian eigenvalue problem. Applied Mathematics and Computation, 321(??):422–430, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307725.

Xu:2018:SFE

[XL18b] Liping Xu and Zhi Li.

Stochastic fractional evolution equations with fractional Brownian motion and infinite delay. Applied Mathematics and Computation, 336(??):36–46, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303862.

Xu:2019:VSF

[XL19] Liping Xu and Jiaowan Luo. Viability for stochastic functional differential equations in Hilbert spaces driven by fractional Brownian motion. Applied Mathematics and Computation, 341:93–110, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307306.

Xu:2018:NSD

[XLG18] Changqing Xu, Jianguo Li, and Shan Ge. Neighbor sum distinguishing total chromatic number of planar graphs. Applied Mathematics and Computation, 332(??):189–196, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301784.

Xiao:2018:OSC

[XLJC18] Xiang Xiao, Xiao-Jian Li, Xiao-Zheng Jin, and Yan Cui. Output synchronization control for a class of complex dynamical networks with non-identical dynamics. Applied Mathematics and Computation, 335(??):38–49, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303473.

Xu:2019:IMT

[XLL^{+}19] Changjin Xu, Maoxin Liao, Peiluan Li, Ying Guo, Qimei Xiao, and Shuai Yuan. Influence of multiple time delays on bifurcation of fractional-order neural networks. Applied Mathematics and Computation, 361(??):565–582, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930459X.

Xu:2017:MFM

[XLR17] Wenwen Xu, Dong Liang, and Hongxing Rui. A multipoint flux mixed finite element method for the compressible Darcy– Forchheimer models. Applied Mathematics and Computation, 315(??): 259–277, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317305106.

Xue:2018:CPC

[XLS18] Jie Xue, Shuting Liu, and Jinlong Shu. The complements of path and cycle are determined by their distance (signless) Laplacian spectra. Applied Mathematics and Computation, 328(??):137–143, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300559.

Xue:2015:DTC

[XLW15] Bo Xue, Fang Li, and Hongyun Wang. Darboux transformation and conservation laws of a integrable evolution equations with 3 × 3 Lax pairs. Applied Mathematics and Computation, 269(??):326–331, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010048.

Xu:2016:NAN

[XLW16] M. Q. Xu, Y. Z. Lin, and Y. H. Wang. A new algorithm for nonlinear fourth order multi-point boundary value problems. Applied Mathematics and Computation, 274(??):163–168, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013922.

Xu:2019:SRE

[XLYZ19] Baogen Xu, Shuchao Li, Rong Yu, and Qin Zhao. On the spectral radius and energy of the weighted adjacency matrix of a graph. Applied Mathematics and Computation, 340(??):156–163, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307264.

Xie:2015:MMS

[XM15] Ya-Jun Xie and Chang-Feng Ma. The MGP-BiCG method for solving the generalized coupled Sylvester-conjugate matrix equations. Applied Mathematics and Computation, 265(??):68–78, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005408.

Xu:2015:ECH

[XMX15] Zhenhua Xu, Gradimir V.

Milovanović, and Shuhuang Xiang. Efficient computation of highly oscillatory integrals with Hankel kernel. Applied Mathematics and Computation, 261(??):312–322, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004579.

Xenos:2017:RET

[XP17] Michalis Xenos and Ioan Pop. Radiation effect on the turbulent compressible boundary layer flow with adverse pressure gradient. Applied Mathematics and Computation, 299(??):153–164, April 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306920.

Xu:2015:ESS

[XPG15] Yong Xu, Bin Pei, and Guobin Guo. Existence and stability of solutions to non-Lipschitz stochastic differential equations driven by Lévy noise. Applied Mathematics and Computation, 263(??): 398–409, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300315005329.

Xiao:2018:ETC

[XPZ18] Xiaoqing Xiao, Ju H.

Park, and Lei Zhou. Event-triggered control of discrete-time switched linear systems with packet losses. Applied Mathematics and Computation, 333(??):344–352, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303035.

Xia:2017:TSM

[XRB17] Liang Xia, Balaji Raghavan, and Piotr Breitkopf. Towards surrogate modeling of material microstructures through the processing variables. Applied Mathematics and Computation, 294(??):157–168, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305598.

Xiang:2015:EGF

[XS15] Xiuqiao Xiang and Baochang Shi. Evolving generation and fast algorithms of slantlet transform and slantlet-Walsh transform. Applied Mathematics and Computation, 269(??):731–743, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010164.

Xie:2019:CAR

[XS19a] Qichang Xie and Qiankun Sun. Computation and application of robust datadriven bandwidth selection for gradient function estimation. Applied Mathematics and Computation, 361(??):274–293, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304461.

Xu:2019:SAL

[XS19b] Chao Xu and Dongyang Shi. Superconvergence analysis of low order nonconforming finite element methods for variational inequality problem with displacement obstacle. Applied Mathematics and Computation, 348(??):1– 11, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S009630031830729X.

Xie:2017:PSS

[XSLW17] Falan Xie, Meijing Shan, Xinze Lian, and Weiming Wang. Periodic solution of a stochastic HBV infection model with logistic hepatocyte growth. Applied Mathematics and Computation, 293(??):630–641, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630399X.

Xing:2017:ARF

[XSS^{+}17] Yi Xing, Lipeng Song, Gui-Quan Sun, Zhen Jin, and Juan Zhang. Assessing reappearance factors of H7N9 avian influenza in China. Applied Mathematics and Computation, 309(??):192–204, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302424.

Xu:2019:AHC

[XSSW19] Zhaowen Xu, Hongye Su, Peng Shi, and Zheng-Guang Wu. Asynchronous H_{∞} control of semi-Markov jump linear systems. Applied Mathematics and Computation, 349(??):270–280, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310518.

Xu:2018:EIP

[XTXF18] Hedong Xu, Cunzhi Tian, Xinrong Xiao, and Suohai Fan. Evolutionary investors’ power-based game on networks. Applied Mathematics and Computation, 330(??):125–133, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301516.

Xiang:2015:IPC

[XTXW15] Zhongyi Xiang, Sanyi Tang, Changcheng Xiang, and Jianhong Wu. On impulsive pest control using integrated intervention strategies. Applied Mathematics and Computation, 269(??):930–946, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009959.

Xu:2015:AAS

[Xu15a]  Houbao Xu.  Avail-

ability analysis for software system with intrusion tolerance. Applied Mathematics and Computation, 252(??):64–76, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016439.

Xu:2015:IAC

[Xu15b] Jianhong Xu. An iterative algorithm for computing mean first passage times of Markov chains. Applied Mathematics and Computation, 250(??):372–389, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015069.

Xu:2015:DRS

[Xu15c]  Xi-Xiang Xu.  A de-

formed reduced semidiscrete Kaup–Newell equation, the related integrable family and Darboux transformation. Applied Mathematics and Computation, 251(??):275–283, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015951.

Xu:2017:NAS

[Xu17a] Da Xu. Numerical asymptotic stability for the integro-differential equations with the multi-term kernels. Applied Mathematics and Computation, 309(??):107–132, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302308.

Xu:2017:GSM

[Xu17b] Xuefeng Xu. Generalization of the Sherman– Morrison–Woodbury formula involving the Schur complement. Applied Mathematics and Computation, 309(??):183–191, September 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317302230.

Xu:2019:ENE

[Xu19a]  Ce Xu.  Evaluations

of nonlinear Euler sums of weight ten. Applied Mathematics and Computation, 346(??):594–611, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309081.

Xu:2019:NAV

[Xu19b] Da Xu. Numerical analysis of Volterra integrodifferential equations for viscoelastic rods and membranes. Applied Mathematics and Computation, 355(??):1–20, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301687.

Xu:2019:HOS

[Xu19c] Mingtian Xu. A high order scheme for unsteady heat conduction equations. Applied Mathematics and Computation, 348(??):565–574, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310713.

Xu:2019:LLA

[Xu19d] Renyu Xu. The list linear arboricity of planar graphs with 7-cycles containing at most two chords. Applied Mathematics and Computation, 363(??): Article 124565, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930548X.

Xu:2019:OPI

[Xu19e] Xiao Xu. OpenMP parallel implementation of stiffly stable time-stepping projection/GMRES(ILU(0)) implicit simulation of incompressible fluid flows on shared-memory, multicore architecture. Applied Mathematics and Computation, 355(??):238–252, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301572.

Xing:2019:NAF

[XW19] Zhiyong Xing and Liping Wen. Numerical analysis and fast implementation of a fourthorder difference scheme for two-dimensional spacefractional diffusion equations. Applied Mathematics and Computation, 346(??):155–166, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309238.

Xie:2017:DIP

[XWDW17] Youxiang Xie, Linjun Wang, Qicheng Deng, and Zhengjia Wu. The dynamics of an impulsive predator–prey model with communicable disease in the prey species only. Applied Mathematics and Computation, 292(??):320–335, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304805.

Xu:2015:GSH

[XWZ15] Rui Xu, Zhili Wang, and Fengqin Zhang. Global stability and Hopf bifurcations of an SEIR epidemiological model with logistic growth and time delay. Applied Mathematics and Computation, 269(??):332–342, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010061.

Xie:2019:MVE

[XX19a] Yan-Ting Xie and Shou-Jun Xu. On the maximum value of the eccentric distance sums of cubic transitive graphs. Applied Mathematics and Computation, 359(??):194–201, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303042.

Xiong:2019:FTR

[XX19b] Xiangtuan Xiong and Xuemin Xue. A fractional Tikhonov regularization method for identifying a space-dependent source in the time-fractional diffusion equation. Applied Mathematics and Computation, 349(??):292–303, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031831124X.

Xing:2019:AFN

[XXW^{+}19] Mengping Xing, Jianwei Xia, Jing Wang, Bo Meng, and Hao Shen. Asynchronous H_{∞} filtering for nonlinear persistent dwelltime switched singular systems with measurement quantization. Applied Mathematics and Computation, 362(??): Article 124578, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305703.

Xu:2018:NCP

[XXZ^{+}18] Quan Xu, Xiaohui Xu, Shengxian Zhuang, Jixue Xiao, Chunhua Song, and Chang Che. New complex projective synchronization strategies for drive-response networks with fractional complex-variable dynamics. Applied Mathematics and Computation, 338(??):544–551, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305447.

Xue:2019:RSI

[XXZY19] Huanbin Xue, Xiaohui Xu, Jiye Zhang, and Xiaopeng Yang. Robust stability of impulsive switched neural networks with multiple time delays. Applied Mathematics and Computation, 359(??):456–475, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303510.

Xiao:2015:NCM

[XY15a] Xiaoyong Xiao and Hongwei Yin. A new class of methods with higher order of convergence for solving systems of nonlinear equations. Applied Mathematics and Computation, 264(??):300–309, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005561.

Xingong:2015:SMS

[XY15b] Zhang Xingong and Wang Yong. Single-machine scheduling CON/SLK due window assignment problems with sum-of-processed times based learning effect. Applied Mathematics and Computation, 250(??):628–635, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015173.

Xiao-yong:2016:MLP

[Xy16] Zhang Xiao-yong. A multistep Legendre pseudospectral method for Volterra integral equations. Applied Mathematics and Computation, 274(??):480–494, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014733.

Xiao:2017:AHO

[XY17] Xiaoyong Xiao and Hongwei Yin. Achieving higher order of convergence for solving systems of nonlinear equations. Applied Mathematics and Computation, 311(??):251–261, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730334X.

Xiao:2018:ACS

[XY18] Xiao-Yong Xiao and Hong-Wei Yin. Accelerating the convergence speed of iterative methods for solving nonlinear systems. Applied Mathematics and Computation, 333(??):8–19, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302893.

Xia:2016:TWS

[XYDL16] Jing Xia, Zhixian Yu, Yucai Dong, and Hongyan Li. Traveling waves for n-species competitive system with nonlocal dispersals and delays. Applied Mathematics and Computation, 287– 288(??):201–213, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302752.

Xie:2018:TDC

[XYG^{+}18] Jiaquan Xie, Zhibin Yao, Hailian Gui, Fuqiang Zhao, and Dongyang Li. A two-dimensional Chebyshev wavelets approach for solving the Fokker–Planck equations of time and space fractional derivatives type with variable coefficients. Applied Mathematics and Computation, 332(??):197–208, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302133.

Xue:2018:NPP

[XYZ18] Lina Xue, Weihua Yang, and Shurong Zhang. Number of proper paths in edge-colored hypercubes. Applied Mathematics and Computation, 332(??):420–424, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302364.

Xiao:2015:NPF

[XYZK15] Yiyong Xiao, Yingying Yuan, Ren-Qian Zhang, and Abdullah Konak. Non-permutation flow shop scheduling with order acceptance and weighted tardiness. Applied Mathematics and Computation, 270(??):312–333, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010668.

Xu:2015:DDD

[XZ15] Zhiting Xu and Yingying Zhao. A diffusive dengue disease model with nonlocal delayed transmission. Applied Mathematics and Computation, 270(??):808–829, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011352.

Xiao:2018:EDC

[XZ18] Jianying Xiao and Shouming Zhong. Extended dissipative conditions for memristive neural networks with multiple time delays. Applied Mathematics and Computation, 323(??):145–163, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308354.

Xie:2019:AIC

[XZ19] Jianqiang Xie and Zhiyue Zhang. An analysis of implicit conservative difference solver for fractional Klein–Gordon–Zakharov system. Applied Mathematics and Computation, 348(??):153–166, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830897X.

Xiao:2018:ECK

[XZB^{+}18] Manyu Xiao, Guohua Zhang, Piotr Breitkopf, Pierre Villon, and Weihong Zhang. Extended co-kriging interpolation method based on multifidelity data. Applied Mathematics and Computation, 323(??):120–131, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307646.

Xie:2016:NLS

[XZC16a] Pinchen Xie, Zhongzhi Zhang, and Francesc Comellas. The normalized Laplacian spectrum of subdivisions of a graph. Applied Mathematics and Computation, 286(??):250–256, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302831.

Xie:2016:SNL

[XZC16b] Pinchen Xie, Zhongzhi Zhang, and Francesc Comellas. On the spectrum of the normalized Laplacian of iterated triangulations of graphs. Applied Mathematics and Computation, 273(??):1123–1129, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012953.

Xie:2015:APS

[XZCC15] Xiaoli Xie, Chunhua Zhang, Xiaoxing Chen, and Jiangyong Chen. Almost periodic sequence solution of a discrete Hassell–Varley predator– prey system with feedback control. Applied Mathematics and Computation, 268(??):35–51, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008152.

Xiao:2019:MHI

[XZG19] Yanni Xiao, Yunhu Zhang, and Min Gao. Modeling hantavirus infections in mainland China. Applied Mathematics and Computation, 360(??):28–41, November 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319303960.

Xu:2017:ETN

[XZHX17] Qiyi Xu, Yijun Zhang, Wangli He, and Shunyuan Xiao. Event-triggered networked H_{∞} control of discrete-time nonlinear singular systems. Applied Mathematics and Computation, 298(??):368–382, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300316306786.

Xiao:2015:FCF

[XZL15] Cuie Xiao, Biao Zeng, and Zhenhai Liu. Feedback control for fractional impulsive evolution systems. Applied Mathematics and Computation, 268(??):924–936, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008735.

Xue:2017:RAS

[XZLL17] Yanmei Xue, Bo-Chao Zheng, Tao Li, and Yuanlu Li. Robust adaptive state feedback sliding-mode control of memristor-based Chua’s systems with input nonlinearity. Applied Mathematics and Computation, 314(??):142–153, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304666.

Xi:2017:DAD

[XZLZ17] Changjiang Xi, Ding Zhai, Xiaojian Li, and Qingling Zhang. Decentralized adaptive delaydependent neural network control for a class of largescale interconnected nonlinear systems. Applied Mathematics and Computation, 311(??):148–163, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303247.

Xia:2016:NNC

[XZS16] Youshen Xia, Songchuan Zhang, and Predrag S. Stanimirović. Neural network for computing pseudoinverses and outer inverses of complex-valued matrices. Applied Mathematics and Computation, 273(??):1107–1121, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013971.

Xu:2015:EFP

[XZW^{+}15] Gang Xu, Yaguang Zhu, Guozhao Wang, André Galligo, Li Zhang, and Kin chuen Hui. Explicit form of parametric polynomial minimal surfaces with arbitrary degree. Applied Mathematics and Computation, 259(??):124–131, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002659.

Xing:2015:CGB

[XzXqT^{+}15] Yan Xing, Ren zheng Xu, Jie qing Tan, Wen Fan, and Ling Hong. A class of generalized B-spline quaternion curves. Applied Mathematics and Computation, 271(??):288–300, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315012436.

Xie:2018:EDE

[XZZ^{+}18] Wenqian Xie, Hong Zhu, Shouming Zhong, Dian Zhang, Kaibo Shi, and Jun Cheng. Extended dissipative estimator design for uncertain switched delayed neural networks via a novel triple integral inequality. Applied Mathematics and Computation, 335(??):82–102, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303552.

Yaseen:2017:CTB

[YAIN17] Muhammad Yaseen, Muhammad Abbas, Ahmad Izani Ismail, and Tahir Nazir. A cubic trigonometric B-spline collocation approach for the fractional sub-diffusion equations. Applied Mathematics and Computation, 293(??):311–319, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305239.

Yan:2017:EST

[Yan17] Weigen Yan. Enumeration of spanning trees of middle graphs. Applied Mathematics and Computation, 307(??):239–243, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301510.

Yan:2018:BIS

[Yan18a] Sherry H. F. Yan. Bijections for inversion sequences, ascent sequences and 3-nonnesting set partitions. Applied Mathematics and Computation, 325(??):24–30, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300317308809.

Yang:2018:SUH

[Yan18b] Xiangfeng Yang. Solving uncertain heat equation via numerical method. Applied Mathematics and Computation, 329(??):92– 104, July 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031830078X.

Yuan:2015:ASM

[YB15] Baoquan Yuan and Linna Bai. L^{2} asymptotic stability of the mild solution to the 3D MHD equation. Applied Mathematics and Computation, 269(??):443–455, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009170.

Yang:2018:ACR

[YBZ18] Yidu Yang, Hai Bi, and Yu Zhang. The adaptive Ciarlet–Raviart mixed method for biharmonic problems with simply supported boundary condition. Applied Mathematics and Computation, 339(??):199–205, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305721.

Yang:2015:AFG

[YC15] Zhen-Hang Yang and Yu-Ming Chu. Asymptotic formulas for gamma function with applications. Applied Mathematics and Computation, 270(??):665–680, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315011431.

Yang:2018:PCP

[YC18] Han-Xin Yang and Xiaojie Chen. Promoting cooperation by punishing minority. Applied Mathematics and Computation, 316(??):460–466, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306008.

Yu:2017:NNI

[YCD17] Zhaofei Yu, Feng Chen, and Jianwu Dong. Neural network implementation of inference on binary Markov random fields with probability coding. Applied Mathematics and Computation, 301(??):193–200, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307597.

Yang:2018:PEE

[YCH18] Jiming Yang, Yanping Chen, and Yunqing Huang. A priori error estimates of a combined mixed finite element and local discontinuous Galerkin method for an incompressible miscible displacement problem. Applied Mathematics and Computation, 334(??):141–151, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308810.

Yan:2019:RIS

[YCH19] Ke Yan, Ho-Lun Cheng, and Jing Huang. Representing implicit surfaces satisfying Lipschitz conditions by 4-dimensional point sets. Applied Mathematics and Computation, 354(??):42–57, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301298.

Yuan:2019:MMV

[YCLS19] Yirang Yuan, Ming Cui, Changfeng Li, and Tongjun Sun. The method of mixed volume elementcharacteristic mixed volume element and its numerical analysis for groundwater pollution in binary medium. Applied Mathematics and Computation, 362(??): Article 124536, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305156.

Yang:2019:RFP

[YCX^{+}19] Te Yang, Guoliang Chen, Jianwei Xia, Zhen Wang, and Qun Sun. Robust H_{∞} filtering for polytopic uncertain stochastic systems under quantized sampled outputs. Applied Mathematics and Computation, 347(??):688–701, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310099.

Yang:2018:SCN

[YCYL18] Yujun Yang, Yuliang Cao, Haiyuan Yao, and Jing Li. Solution to a conjecture on a Nordhaus–Gaddum type result for the Kirchhoff index. Applied Mathematics and Computation, 332(??):241–249, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302431.

Yang:2015:SBP

[YCZ15] Zhen-Hang Yang, Yu-Ming Chu, and Xiao-Hui Zhang. Sharp bounds for psi function. Applied Mathematics and Computation, 268(??):1055–1063, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009248.

Yang:2019:HAA

[YCZ19] Zhen-Hang Yang, Yu-Ming Chu, and Wen Zhang. High accuracy asymptotic bounds for the complete elliptic integral of the second kind. Applied Mathematics and Computation, 348(??):552–564, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310725.

Yang:2018:LCC

[YD18] Junmin Yang and Wei Ding. Limit cycles of a class of Liénard systems with restoring forces of seventh degree. Applied Mathematics and Computation, 316(??):422–437, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305593.

Yao:2015:ILR

[YDCS15] Guangming Yao, Jia Duo, C. S. Chen, and L. H. Shen. Implicit local radial basis function interpolations based on function values. Applied Mathematics and Computation, 265(??):91–102, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500569X.

Yin:2015:SPF

[YDLK15] J. Yin, D. Ding, Z. Liu, and S. Khoo. Some properties of finite-time stable stochastic nonlinear systems. Applied Mathematics and Computation, 259(??):686–697, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500288X.

Yan:2019:AGP

[YDWZ19] Yun Yan, Weizhong Dai, Longyuan Wu, and Shuying Zhai. Accurate gradient preserved method for solving heat conduction equations in double layers. Applied Mathematics and Computation, 354(??):58–85, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301420.

Yero:2017:CMD

[YEMRV17] Ismael G. Yero, Alejandro Estrada-Moreno, and Juan A. Rodríguez-Velázquez. Computing the k-metric dimension of graphs. Applied Mathematics and Computation, 300(??):60–69, May 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300316307317.

Ye:2017:ESG

[YF17] Wenxing Ye and Suohai Fan. Evolutionary snowdrift game with rational selection based on radical evaluation. Applied Mathematics and Computation, 294(??):310–317, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305677.

Yang:2019:ATP

[YFH19] Da-Wei Yang, Rongquan Feng, and Xiao-Hui Hua. On arc-transitive pentavalent graphs of order 2^{m}p^{n}. Applied Mathematics and Computation, 355(??):269–281, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301845.

Yang:2016:SPG

[YFL16] Xu Yang, Lihua Feng, and Weijun Liu. Some properties of graphs constructed from 2-designs. Applied Mathematics and Computation, 284(??):1– 11, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301564.

Ye:2017:MBP

[YFLF17] Wenxing Ye, Weiying Feng, Chen Lü, and Suohai Fan. Memory-based prisoner’s dilemma game with conditional selection on networks. Applied Mathematics and Computation, 307(??):31–37, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301467.

Yang:2018:ECE

[YFLZ18] Da-Wei Yang, Yan-Quan Feng, Jaeun Lee, and Jin-Xin Zhou. On extra connectivity and extra edge-connectivity of balanced hypercubes. Applied Mathematics and Computation, 320(??):464–473, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730694X.

Yi:2017:SDN

[YFW^{+}17] Chengbo Yi, Jianwen Feng, Jingyi Wang, Chen Xu, and Yi Zhao. Synchronization of delayed neural networks with hybrid coupling via partial mixed pinning impulsive control. Applied Mathematics and Computation, 312(??):78–90, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302825.

Yu:2017:LSL

[YFY17] Gui-Dong Yu, Yi-Zheng Fan, and Miao-Lin Ye. The least signless Laplacian eigenvalue of the complements of unicyclic graphs. Applied Mathematics and Computation, 306(??):13–21, August 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301224.

Yang:2015:PCB

[YG15a] Hyun Mo Yang and David Greenhalgh. Proof of conjecture in: the basic reproduction number obtained from Jacobian and next generation matrices — a case study of dengue transmission modelling. Applied Mathematics and Computation, 265(??):103–107, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005743.

Yin:2015:ECE

[YG15b] Zhengwei Yin and Siqing Gan. An error corrected Euler–Maruyama method for stiff stochastic differential equations. Applied Mathematics and Computation, 256(??):630–641, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000995.

Yao:2018:SDI

[YG18] Jinran Yao and Siqing Gan. Stability of the drift-implicit and doubleimplicit Milstein schemes for nonlinear SDEs. Applied Mathematics and Computation, 339(??): 286–293, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318305848.

Yang:2019:CHO

[YGZ19] Xiaojia Yang, Yongbin Ge, and Lin Zhang. A class of high-order compact difference schemes for solving the Burgers’ equations. Applied Mathematics and Computation, 358(??):394–417, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303054.

Yu:2018:LAL

[YHY^{+}18] Xiaowei Yu, Jie Hu, Donglei Yang, Jianliang Wu, and Guanghui Wang. Local antimagic labeling of graphs. Applied Mathematics and Computation, 322(??):30–39, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306975.

Yilmaz:2016:SDP

[Yil16] Fikriye Yilmaz. Semidiscrete a priori error analysis for the optimal control of the unsteady Navier–Stokes equations with variational multiscale stabilization. Applied Mathematics and Computation, 276(??):127–142, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315016112.

Yu:2019:MMT

[YJ19] Jui-Ling Yu and Sophia R.-J. Jang. A mathematical model of tumorimmune interactions with an immune checkpoint inhibitor. Applied Mathematics and Computation, 362(??):Article 124523, December 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319305028.

Yang:2019:AHI

[YJK19] Yun-Bo Yang, Yao-Lin Jiang, and Qiong-Xiang Kong. The Arrow– Hurwicz iterative finite element method for the stationary magnetohydrodynamics flow. Applied Mathematics and Computation, 356(??):347–361, September 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318309160.

Yu:2018:GCC

[YJMH18] Zhiyong Yu, Haijun Jiang, Xuehui Mei, and Cheng Hu. Guaranteed cost consensus for secondorder multi-agent systems with heterogeneous inertias. Applied Mathematics and Computation, 338(??):722–732, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305204.

Yu:2016:DRD

[YJSW16] Jianping Yu, Jian Jing, Yongli Sun, and Suping Wu. (n + 1)-dimensional reduced differential transform method for solving partial differential equations. Applied Mathematics and Computation, 273(??):697–705, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501348X.

Yu:2016:NAE

[YJW16] Bo Yu, Xiaoyun Jiang, and Chu Wang. Numerical algorithms to estimate relaxation parameters and Caputo fractional derivative for a fractional thermal wave model in spherical composite medium. Applied Mathematics and Computation, 274(??):106–118, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014320.

Yuan:2016:WRT

[YJY16] Xiaoguang Yuan, Quan Jiang, and Fengpeng Yang. Wave reflection and transmission in rotating and stressed pyroelectric half-planes. Applied Mathematics and Computation, 289(??):281–297, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630323X.

Yuan:2015:LST

[YJYH15] Yanbin Yuan, Bin Ji, Xiaohui Yuan, and Yuehua Huang. Lockage scheduling of Three Gorges– Gezhouba dams by hybrid of chaotic particle swarm optimization and heuristic-adjusted strategies. Applied Mathematics and Computation, 270(??):74–89, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010644.

Yang:2017:DDF

[YKJY17] Guowei Yang, Yonggui Kao, Baoping Jiang, and Jile Yin. Delay-dependent H_{∞} filtering for singular Markovian jump systems with general uncomplete transition probabilities. Applied Mathematics and Computation, 294(??):195–215, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304052.

Yu:2019:EIO

[YKLS19] Zhangqing Yu, Liying Kang, Lele Liu, and Erfang Shan. The extremal α-index of outerplanar and planar graphs. Applied Mathematics and Computation, 343(??):90–99, February 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S009630031830821X.

Yin:2015:AEM

[YKZ15] Jun-Feng Yin, Xiao-He Kong, and Ning Zheng. Aitken extrapolation method for computing the largest eigenvalue of nonnegative tensors. Applied Mathematics and Computation, 258(??):350–357, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000910.

Yan:2017:ACM

[YL17] Zuomao Yan and Fangxia Lu. Approximate controllability of a multivalued fractional impulsive stochastic partial integrodifferential equation with infinite delay. Applied Mathematics and Computation, 292(??):425–447, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304064.

Yang:2018:MCL

[YL18] Yuxuan Yang and Mei Lu. Matching of 5-γ-critical leafless graph with a cut edge. Applied Mathematics and Computation, 323(??):58–63, April 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308329.

Yang:2019:GSH

[YL19a] Bowen Yang and Ping Liu. Global stability and Hopf bifurcation of a three-component model for cell production systems. Applied Mathematics and Computation, 354(??):478–489, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301468.

Yu:2019:MRP

[YL19b] Xingchen Yu and Shiping Lu. A multiplicity result for periodic solutions of Liénard equations with an attractive singularity. Applied Mathematics and Computation, 346(??):183–192, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308798.

Yuan:2019:IHM

[YL19c] Baoquan Yuan and Chaoying Li. The 3D incompressible Hall magnetohydrodynamics equations with partial hyperdissipation. Applied Mathematics and Computation, 359(??):308–332, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303431.

Yang:2017:GGA

[YLCF17] Xu Yang, Weijun Liu, Jing Chen, and Lihua Feng. GCI-groups in the alternating groups. Applied Mathematics and Computation, 303(??):42–47, June 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300309.

Yang:2018:ROO

[YLCW18] Yuxia Yang, Chong Lin, Bing Chen, and Qing-Guo Wang. Reducedorder observer design for a class of generalized Lipschitz nonlinear systems with time-varying delay. Applied Mathematics and Computation, 337(??):234–266, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304144.

Yu:2019:RHS

[YLH^{+}19] Jianwu Yu, Hong Luo, Junzhi Hu, Thai Vinh Nguyen, and Yuetuo Lu.

Reconstruction of highspeed cam curve based on high-order differential interpolation and shape adjustment. Applied Mathematics and Computation, 356(??):272–281, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302565.

Yang:2017:MLS

[YLHS17] Xujun Yang, Chuandong Li, Tingwen Huang, and Qiankun Song. Mittag-Leffler stability analysis of nonlinear fractionalorder systems with impulses. Applied Mathematics and Computation, 293(??):416–422, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305343.

Yao:2015:HBA

[YLL15] Yong Yao, Zuxiong Li, and Zhijun Liu. Hopf bifurcation analysis of a turbidostat model with discrete delay. Applied Mathematics and Computation, 262(??):267–281, July 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005081.

Yuan:2016:ABC

[YLL^{+}16] Yirang Yuan, Yunxin Liu, Changfeng Li, Tongjun Sun, and Liqin Ma. Analysis on block-centered finite differences of numerical simulation of semiconductor device detector. Applied Mathematics and Computation, 279(??):1–15, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630011X.

Yang:2018:RUA

[YLL18] Xin-Guang Yang, Lu Li, and Yongjin Lu. Regularity of uniform attractor for 3D non-autonomous Navier–Stokes-Voigt equation. Applied Mathematics and Computation, 334(??):11–29, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302777.

Yan:2019:GTE

[YLLH19] Weijie Yan, Chen Ling, Liyun Ling, and Hongjin He. Generalized tensor equations with leading structured tensors. Applied Mathematics and Computation, 361(??): 311–324, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL https://www.sciencedirect.
com/science/article/
pii/S0096300319304448.

Yao:2015:COP

[YLNL15] Shengwei Yao, Xiwen Lu, Liangshuo Ning, and Feifei Li. A class of one parameter conjugate gradient methods. Applied Mathematics and Computation, 265(??):708–722, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007468.

Yao:2019:DOB

[YLP19] Xiuming Yao, Yue Lian, and Ju H. Park. Disturbanceobserver-based event-triggered control for semi-Markovian jump nonlinear systems. Applied Mathematics and Computation, 363(??): Article 124597, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.

com/science/article/
pii/S0096300319305892.

Yu:2017:SHQ

[YLQ17] Guihai Yu, Xin Liu, and Hui Qu. Singularity of Hermitian (quasi-)Laplacian matrix of mixed graphs. Applied Mathematics and Computation, 293(??):287–292, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305276.

Yang:2015:SCU

[YLW15] Ai-Li Yang, Xu Li, and Yu-Jiang Wu. On semiconvergence of the Uzawa-HSS method for singular saddle-point problems. Applied Mathematics and Computation, 252(??):88– 98, February 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314016336.

Yu:2016:ESU

[YLW16] Senlin Yu, Zaiming Liu, and Jinbiao Wu. Equilibrium strategies of the unobservable M/M/1 queue with balking and delayed repairs. Applied Mathematics and Computation, 290(??):56–65, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303654.

Yin:2019:QFT

[YLW19] Zeyang Yin, Jianjun Luo, and Caisheng Wei. Quasi fixed-time fault-tolerant control for nonlinear mechanical systems with enhanced performance. Applied Mathematics and Computation, 352(??): 157–173, July 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319300852.

Yang:2015:SSP

[YLWF15] Yu Yang, Hongbo Liu, Hua Wang, and Hongsun Fu. Subtrees of spiro and polyphenyl hexagonal chains. Applied Mathematics and Computation, 268(??):547–560, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008759. See

corrigendum [YLWF16].

Yang:2016:CSS

[YLWF16] Yu Yang, Hongbo Liu, Hua Wang, and Hongsun Fu. Corrigendum to “Subtrees of spiro and polyphenyl hexagonal chains” [Applied Mathematics and Computation 268 (2015) 547– 560]. Applied Mathematics and Computation, 287–288(??):232, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303551. See

[YLWF15].

Yang:2018:RJS

[YLX18] Xiaobo Yang, Anping Liao, and Jiaxin Xie. A remark on joint sparse recovery with OMP algorithm under restricted isometry property. Applied Mathematics and Computation, 316(??):18–24, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305428.

Yu:2017:SIR

[YLZT17] Jie-Ru Yu, Xue-Lu Liu, Xiu-Deng Zheng, and Yi Tao. Selection intensity and risk-dominant strategy: a two-strategy stochastic evolutionary game dynamics in finite population. Applied Mathematics and Computation, 297(??):1–7, March 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306464.

Yang:2018:CCH

[YM18] Hyunjin Yang and Mehrdad Massoudi. Conduction and convection heat transfer in a dense granular suspension. Applied Mathematics and Computation, 332(??):351–362, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302297.

Yan:2015:NCF

[YMD15] Yun Yan, Frederick Ira Moxley III, and Weizhong Dai. A new compact finite difference scheme for solving the complex Ginzburg– Landau equation. Applied Mathematics and Computation, 260(??):269–287, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300315003677.

Yu:2019:ANA

[YMeFG19] Yongyuan Yu, Min Meng, Jun e Feng, and Yan Gao. An adjoint network approach to design stabilizable switching signals of switched Boolean networks. Applied Mathematics and Computation, 357(??):12–22, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301778.

Yuan:2015:ATW

[YMHW15] Wenjun Yuan, Fanning Meng, Yong Huang, and Yonghong Wu. All traveling wave exact solutions of the variant Boussinesq equations. Applied Mathematics and Computation, 268(??):865–872, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008693.

You:2015:AES

[YMMH15] Surong You, Wei Mao, Xuerong Mao, and Liangjian Hu. Analysis on exponential stability of hybrid pantograph stochastic differential equations with highly nonlinear coefficients. Applied Mathematics and Computation, 263(??):73–83, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004737.

Yang:2016:NNT

[YMS16] Xiao-Jun Yang, J. A. Tenreiro Machado, and H. M. Srivastava. A new numerical technique for solving the local fractional diffusion equation: Twodimensional extended differential transform approach. Applied Mathematics and Computation, 274(??):143–151, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501423X.

Yin:2015:RSA

[YmZHC15] Chun Yin, Shou ming Zhong, Xuegang Huang, and Yuhua Cheng. Robust stability analysis of fractional-order uncertain singular nonlinear system with external disturbance. Applied Mathematics and Computation, 269(??):351–362, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009789.

Yuan:2016:SNT

[YNMW16] Xinpeng Yuan, Jianguo Ning, Tianbao Ma, and Cheng Wang. Stability of Newton TVD Runge– Kutta scheme for onedimensional Euler equations with adaptive mesh. Applied Mathematics and Computation, 282(??):1– 16, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630090X.

Yaman:2018:BIE

[YÖ18] Olha Ivanyshyn Yaman and Gazi Özdemir. Boundary integral equations for the exterior Robin problem in two dimensions. Applied Mathematics and Computation, 337(??):14–24, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303734.

Yordzhev:2015:CNA

[Yor15] Krasimir Yordzhev. Calculation of the number of all pairs of disjoint S-permutation matrices. Applied Mathematics and Computation, 268(??):1–11, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315008358.

Yoshikawa:2019:REM

[Yos19]  Shuji Yoshikawa.  Re-

marks on energy methods for structure-preserving finite difference schemes — small data global existence and unconditional error estimate. Applied Mathematics and Computation, 341:80–92, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318307446.

You:2015:SNQ

[You15] Xu You. Some new quicker convergences to Glaisher– Kinkelin’s and Bendersky– Adamchik’s constants. Applied Mathematics and Computation, 271(??): 123–130, November 15, 2015. CODEN AMHCBQ.

ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315011959.

Yu:2018:MPD

[YP18] Yongchao Yu and Jigen Peng. A modified primal-dual method with applications to some sparse recovery problems. Applied Mathematics and Computation, 333(??):76–94, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302704.

Yan:2018:UFA

[YPZ18] Zhiguo Yan, Ju H. Park, and Weihai Zhang. A unified framework for asymptotic and transient behavior of linear stochastic systems. Applied Mathematics and Computation, 325(??):31–40, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308822.

Yu:2015:HLM

[YQ15] Guihai Yu and Hui Qu.

Hermitian Laplacian matrix and positive of mixed graphs. Applied Mathematics and Computation, 269(??):70–76, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009649.

Yu:2018:CIP

[YQ18] Guihai Yu and Hui Qu. The coefficients of the immanantal polynomial. Applied Mathematics and Computation, 339(??):15– 37, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318305460.

Yu:2017:PMV

[YQD17] Guihai Yu, Hui Qu, and Matthias Dehmer. Principal minor version of matrix-tree theorem for mixed graphs. Applied Mathematics and Computation, 309(??):27–30, September 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317302187.

Yu:2015:ICU

[YQT15] Guihai Yu, Hui Qu, and Jianhua Tu. Inertia of complex unit gain graphs. Applied Mathematics and Computation, 265(??):619–629, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007365.

Yang:2015:AMS

[YR15] Xiangfeng Yang and Dan A.

Ralescu. Adams method for solving uncertain differential equations. Applied Mathematics and Computation, 270(??): 993–1003, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300315011650.

Younis:2015:ASS

[YRA15] Muhammad Younis, Syed Tahir Raza Rizvi, and Safdar Ali. Analytical and soliton solutions: Nonlinear model of nanobioelectronics transmission lines. Applied Mathematics and Computation, 265(??):994–1002, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007523.

Yin:2016:FTT

[YS16] Yi Yin and Pengjian Shang. Forecasting traffic time series with multivariate predicting method. Applied Mathematics and Computation, 291(??): 266–278, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316304477.

Ye:2017:RNL

[YSD^{+}17] Ya-Fen Ye, Yuan-Hai Shao, Nai-Yang Deng, Chun-Na Li, and Xiang-Yu Hua. Robust L_{p}-norm least squares support vector regression with feature selection. Applied Mathematics and Computation, 305(??):32–52, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317300887.

Yin:2016:MMC

[YSF16] Yi Yin, Pengjian Shang, and Guochen Feng. Modified multiscale crosssample entropy for complex time series. Applied Mathematics and Computation, 289(??):98–110, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303204.

Yan:2018:FTS

[YSL18] Zhiguo Yan, Yunxia Song, and Xiaoping Liu. Finitetime stability and stabilization for Itô-type stochastic Markovian jump systems with generally uncertain transition rates. Applied Mathematics and Computation, 321(??): 512–525, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317307580.

Yang:2019:SCE

[YSL19] Huizi Yang, Minghui Song, and Mingzhu Liu. Strong convergence and exponential stability of stochastic differential equations with piecewise continuous arguments for nonglobally Lipschitz continuous coefficients. Applied Mathematics and Computation, 341:111–127, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307513.

Yang:2017:GDH

[YSLS17] Wei Yang, Zhan Shu, James Lam, and Chengjun Sun. Global dynamics of an HIV model incorporating senior male clients. Applied Mathematics and Computation, 311(??):203–216, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317303284.

Yan:2017:GOF

[YSWZ17] Xuehua Yan, Xinmin Song, Zhonghua Wang, and Yun Zhang. Global output-feedback adaptive stabilization for planar nonlinear systems with unknown growth rate and output function. Applied Mathematics and Computation, 314(??):299–309, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304654.

Yin:2015:CST

[YSX15] Yi Yin, Pengjian Shang, and Jianan Xia. Compositional segmentation of time series in the financial markets. Applied Mathematics and Computation, 268(??):399–412, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008425.

Yang:2017:NSD

[YSY^{+}17] Donglei Yang, Lin Sun, Xiaowei Yu, Jianliang Wu, and Shan Zhou. Neighbor sum distinguishing total chromatic number of planar graphs with maximum degree 10. Applied Mathematics and Computation, 314(??):456–468, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304046.

Yu:2019:RMR

[YSZ19] Zhongbao Yu, Fangming Shao, and Zuyuan Zhang. Researches for more reliable arrangement graphs in multiprocessor computer system. Applied Mathematics and Computation, 363(??): Article 124611, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0096300319306034.

Yang:2019:NSF

[YTH19] Yin Yang, Zhuyan Tang, and Yunqing Huang. Numerical solutions for Fredholm integral equations of the second kind with weakly singular kernel using spectral collocation method. Applied Mathematics and Computation, 349(??):314–324, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318310877.

Yildirim:2015:NSH

[YU15] Ozgur Yildirim and Meltem Uzun. On the numerical solutions of high order stable difference schemes for the hyperbolic multipoint nonlocal boundary value problems. Applied Mathematics and Computation, 254(??):210–218, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017780.

Yu:2016:GEC

[Yu16] Yuehua Yu. Global exponential convergence for a class of HCNNs with neutral time-proportional delays. Applied Mathematics and Computation, 285(??):1–7, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302119.

Yuan:2015:OMP

[Yua15]  Quan Yuan.  Optimal

matrix pencil approximation problem in structural dynamic model updating. Applied Mathematics and Computation, 250(??):12–27, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401474X.

Yuan:2019:NTT

[Yua19] Long-Tu Yuan. A note on Turán’s theorem. Applied Mathematics and Computation, 361(??):13–14, November 15, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319304138.

Yue:2016:ASC

[Yue16] Jun Yue. Acyclic and star coloring of P_{4}-reducible and P_{4}-sparse graphs. Applied Mathematics and Computation, 273(??):68– 73, January 15, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315013223.

Yue:2019:SCC

[Yue19] Chao Yue. Strong convergence of compensated split-step theta methods for SDEs with jumps under monotone condition. Applied Mathematics and Computation, 340(??):72–83, January 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303060.

Yildiz:2019:SPR

[YUK19] Tugba Akman Yildiz, Murat Uzunca, and Bülent Karasözen. Structure preserving reduced order modeling for gradient systems. Applied Mathematics and Computation, 347(??):194–209, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309810.

Yun:2015:CRI

[Yun15] Jae Heon Yun. Convergence of relaxation iterative methods for saddle point problem. Applied Mathematics and Computation, 251(??):65–80, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015793.

Yuzbasi:2015:NSS

[Yüz15] Suayip Yüzbasi. Numerical solutions of system of linear Fredholm–Volterra integro-differential equations by the Bessel collocation method and error estimation. Applied Mathematics and Computation, 250(??):320–338, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014854.

Yuzbasi:2016:CMB

[Yüz16a] Suayip Yüzbasi. A collocation method based on Bernstein polynomials to solve nonlinear Fredholm–Volterra integro-differential equations. Applied Mathematics and Computation, 273(??):142–154, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013296.

Yuzbasi:2016:NSH

[Yüz16b] Suayip Yüzbasi. Numerical solutions of hyperbolic telegraph equation by using the Bessel functions of first kind and residual correction. Applied Mathematics and Computation, 287–288(??):83–93, September 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316302867.

Yuzbasi:2018:CAS

[Yüz18] Suayip Yüzbasi. A collocation approach for solving two-dimensional secondorder linear hyperbolic equations. Applied Mathematics and Computation, 338(??):87–100, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304703.

Yan:2015:FON

[YW15] Li Yan and Jiang Wei. Fractional order nonlinear systems with delay in iterative learning control. Applied Mathematics and Computation, 257(??):546–552, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000284.

Yang:2016:ACR

[YW16] Min Yang and Qiru Wang.

Approximate controllability of Riemann–Liouville fractional differential inclusions. Applied Mathematics and Computation, 274(??):267–281, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014794.

Yang:2018:OBD

[YW18a] Xiuwen Yang and Ligong Wang. On the ordering of bicyclic digraphs with respect to energy and iota energy. Applied Mathematics and Computation, 339(??):758–767, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306337.

Yang:2018:SEC

[YW18b] Zixuan Yang and Baoyindureng Wu. Strong edge chromatic index of the generalized Petersen graphs. Applied Mathematics and Computation, 321(??):431–441, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307464.

Yuan:2018:PSP

[YW18c] Hailong Yuan and Jianhua Wu. Positive solutions of a predator–prey model with cross-diffusion. Applied Mathematics and Computation, 331(??):232–250, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318302030.

Yang:2019:DAP

[YW19a] Junyuan Yang and Xiaoyan Wang. Dynamics and asymptotical profiles of an age-structured viral infection model with spatial diffusion. Applied Mathematics and Computation, 360(??):236–254, November 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0096300319303947.

Yang:2019:NTF

[YW19b] Tinggan Yang and Yihong Wang. A new tailored finite point method for strongly anisotropic diffusion equation on misaligned grids. Applied Mathematics and Computation, 355(??):85–95, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301870.

Yang:2019:MDG

[YwFW^{+}19] Yu Yang, Ai wan Fan, Hua Wang, Hailian Lv, and Xiao-Dong Zhang. Multi-distance granularity structural α-subtree index of generalized Bethe trees. Applied Mathematics and Computation, 359(??):107–120, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930325X.

Yun:2015:ALS

[YWH15] D. F. Yun, Z. H. Wen, and Y. C. Hon. Adaptive least squares finite integration method for higher-dimensional singular perturbation problems with multiple boundary layers. Applied Mathematics and Computation, 271(??):232–250, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011972.

Yang:2015:HTI

[YWL15] Rongjiang Yang, Bo Wu, and Yang Liu. A Halanaytype inequality approach to the stability analysis of discrete-time neural networks with delays. Applied Mathematics and Computation, 265(??):696–707, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007122.

Yan:2018:AQS

[YWL18] Zhi Yan, Wei Wang, and Xianbin Liu. Analysis of a quintic system with fractional damping in the presence of vibrational resonance. Applied Mathematics and Computation, 321(??):780–793, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730810X.

Yang:2019:FWT

[YWL19] Yong Yang, Chufen Wu, and Zunxian Li. Forced waves and their asymptotics in a Lotka–Volterra cooperative model under climate change. Applied Mathematics and Computation, 353(??):254–264, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930075X.

Yue:2018:DRD

[YWLL18] Jun Yue, Meiqin Wei, Min Li, and Guodong Liu. On the double Roman domination of graphs. Applied Mathematics and Computation, 338(??):661–668, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305228.

Yang:2018:CNN

[YWO18] Dan Yang, JinRong Wang, and D. O’Regan. A class of nonlinear noninstantaneous impulsive differential equations involving parameters and fractional order. Applied Mathematics and Computation, 321(??):654–671, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730807X.

Yimnet:2016:NIS

[YWRP16] S. Yimnet, B. Wongsaijai, T. Rojsiraphisal, and K. Poochinapan. Numerical implementation for solving the symmetric regularized long wave equation. Applied Mathematics and Computation, 273(??):809–825, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013077.

Yang:2018:NLE

[YWS18] Yanhong Yang, Yushun Wang, and Yongzhong Song. A new local energy-preserving algorithm for the BBM equation. Applied Mathematics and Computation, 324(??):119–130, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S009630031730872X.

Yang:2019:NSI

[YWS^{+}19] Huilan Yang, Xin Wang, Lan Shu, Guozhu Zhao, and Shouming Zhong. A new sampling interval fragmentation approach to synchronization of chaotic Luré systems. Applied Mathematics and Computation, 348(??):12–24, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309822.

Wang:2015:QPC

[yWyLyY^{+}15] Xiang yang Wang, Wei yi Li, Hong ying Yang, Pei Wang, and Yong wei Li. Quaternion polar complex exponential transform for invariant color image description. Applied Mathematics and Computation, 256(??):951–967, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001071.

Yang:2018:MAC

[YWYS18] Zhiqiang Yang, Ziqiang Wang, Zihao Yang, and Yi Sun. Multiscale analysis and computation for coupled conduction, convection and radiation heat transfer problem in porous materials. Applied Mathematics and Computation, 326(??):56–74, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317309062.

Yu:2018:GLS

[YWZ18] Hao Yu, Boying Wu, and Dazhi Zhang. A generalized Laguerre spectral Petrov–Galerkin method for the time-fractional subdiffusion equation on the semi-infinite domain. Applied Mathematics and Computation, 331(??):96– 111, August 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318301607.

Yang:2018:SNC

[YWZS18] Huilan Yang, Xin Wang, Shouming Zhong, and Lan Shu. Synchronization of nonlinear complex dynamical systems via delayed impulsive distributed control. Applied Mathematics and Computation, 320(??):75–85, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306409.

Yu:2015:IPT

[YX15] Yue Yu and Dinghua Xu. On the inverse problem of thermal conductivity determination in nonlinear heat and moisture transfer model within textiles. Applied Mathematics and Computation, 264(??):284–299, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005469.

Yang:2019:TGD

[YX19a] Jiming Yang and Xiaoqing Xing. A two-grid discontinuous Galerkin method for a kind of nonlinear parabolic problems. Applied Mathematics and Computation, 346(??):96– 108, April 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318308543.

Yang:2019:LSB

[YX19b] Shaojie Yang and Tianzhou Xu. Local-in-space blowup and symmetric waves for a generalized twocomponent Camassa–Holm system. Applied Mathematics and Computation, 347(??):514–521, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308981.

Yu:2015:GBF

[YY15] Xiu-Juan Yu and Chuan-Fu Yang. The gap between the first two eigenvalues of Schrödinger operators with single-well potential. Applied Mathematics and Computation, 268(??):275–283, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008590.

Yang:2017:NLC

[YY17] Junmin Yang and Pei Yu. Nine limit cycles around a singular point by perturbing a cubic Hamiltonian system with a nilpotent center. Applied Mathematics and Computation, 298(??):141–152, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300316306890.

Yin:2019:GSR

[YYL19] Jian-Hua Yin, Kun Ye, and Jia-Yun Li. Graphic sequences with a realization containing cycles C_{3}, …, C_{l}. Applied Mathematics and Computation, 353(??):88–94, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301018.

Yang:2019:TNA

[YYM19] Huizi Yang, Zhanwen Yang, and Shufang Ma. Theoretical and numerical analysis for Volterra integro-differential equations with Itô integral under polynomially growth conditions. Applied Mathematics and Computation, 360(??):70–82, November 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319302619.

Yu:2018:ELB

[YYP18] Zhuoxi Yu, Kai Yang, and Milan Parmar. Empirical likelihood based inference for generalized additive partial linear models. Applied Mathematics and Computation, 339(??):93–104, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305393.

Ye:2017:FTS

[YYS17] Dan Ye, Xiang Yang, and Lei Su. Faulttolerant synchronization control for complex dynamical networks with semi-Markov jump topology. Applied Mathematics and Computation, 312(??):36–48, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302965.

Yang:2019:VSV

[YYWH19] Ben-Zhang Yang, Jia Yue, Ming-Hui Wang, and Nan-Jing Huang. Volatility swaps valuation under stochastic volatility with jumps and stochastic intensity. Applied Mathematics and Computation, 355(??):73–84, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300319301675.

Yang:2019:EWA

[YYX^{+}19] Xiaoxia Yang, Xiaoli Yang, Shuqi Xue, Jihui Zhang, Fuquan Pan, Yuanlei Kang, and Qianling Wang. The effect of waiting area design at the metro platform on passengers’ alighting and boarding behaviors. Applied Mathematics and Computation, 358(??):177–193, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303005.

Yu:2019:SES

[YYZ19] Xingwang Yu, Sanling Yuan, and Tonghua Zhang. Survival and ergodicity of a stochastic phytoplankton– zooplankton model with toxin-producing phytoplankton in an impulsive polluted environment. Applied Mathematics and Computation, 347(??):249–264, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309780.

Yang:2015:RAB

[YZ15a] Zhihua Yang and Qianwei Zhang. Resource allocation based on DEA and modified Shapley value. Applied Mathematics and Computation, 263(??):280–286, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005251.

Yuan:2015:LSS

[YZ15b] Yongxin Yuan and Kezheng Zuo. Least-squares symmetric and skew-symmetric solutions of the generalized Sylvester matrix equation ∑_{i=1}^{s} A_{i}XB_{i} + ∑_{j=1}^{t} C_{j}Y D_{j} = E. Applied Mathematics and Computation, 265(??): 370–379, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315006529.

Yuan:2015:RND

[YZ15c] Yongxin Yuan and Kezheng Zuo. The Re-nonnegative definite and Re-positive definite solutions to the matrix equation AXB = D. Applied Mathematics and Computation, 256(??):905–912, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001307.

Yu:2016:SAD

[YZ16] Qiang Yu and Xudong Zhao. Stability analysis of discrete-time switched linear systems with unstable subsystems. Applied Mathematics and Computation, 273(??):718–725, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013909.

Ye:2017:CGC

[YZ17a] Xiaoli Ye and Zhizheng Zhang. A common generalization of convolved generalized Fibonacci and Lucas polynomials and its applications. Applied Mathematics and Computation, 306(??):31–37, August 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301200.

Ye:2017:RRC

[YZ17b] Zhuan Ye and Zujin Zhang. A remark on regularity criterion for the 3D Hall-MHD equations based on the vorticity. Applied Mathematics and Computation, 301(??):70– 77, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307378.

Yang:2018:SIG

[YZ18a] Jizhen Yang and Zhizheng Zhang. Some identities of the generalized Fibonacci and Lucas sequences. Applied Mathematics and Computation, 339(??):441–450, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306209.

Yang:2018:FEM

[YZ18b] Xu Yang and Weidong Zhao. Finite element methods and their error analysis for SPDEs driven by Gaussian and non-Gaussian noises. Applied Mathematics and Computation, 332(??):58–75, September 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0096300318302121.

Yang:2019:BMP

[YZ19] Zhanshan Yang and Bing Zheng. Block minimum perturbation algorithm based on block Arnoldi process for nonsymmetric linear systems with multiple righthand sides. Applied Mathematics and Computation, 347(??):741–766, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309986.

Yang:2015:EMI

[YZD15] Changhui Yang, Qiang Zhang, and Shuai Ding. An evaluation method for innovation capability based on uncertain linguistic variables. Applied Mathematics and Computation, 256(??):160–174, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000181.

Yao:2019:USA

[YZH^{+}19] Xueqi Yao, Shouming Zhong, Taotao Hu, Hong Cheng, and Dian Zhang.

Uniformly stable and attractive of fractional-order memristor-based neural networks with multiple delays. Applied Mathematics and Computation, 347(??):392–403, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310026.

Yao:2018:FEM

[YZJ18] Changhui Yao, Yuzhen Zhou, and Shanghui Jia. A finite element method for Maxwell polynomial chaos Debye model. Applied Mathematics and Computation, 325(??):59–68, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308780.

Yu:2015:VSS

[YZMM15] Shuhao Yu, Shenglong Zhu, Yan Ma, and Demei Mao. A variable step size firefly algorithm for numerical optimization. Applied Mathematics and Computation, 263(??):214–220, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005275.

Yu:2018:LEF

[YZYgG18] Guanglong Yu, Mingqing Zhai, Chao Yan, and Shu guang Guo. The least Q-eigenvalue with fixed domination number. Applied Mathematics and Computation, 339(??):459–476, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318306210.

Yue:2016:NPE

[YZZ16] Jun Yue, Shiliang Zhang, and Xia Zhang. Note on the perfect EICgraphs. Applied Mathematics and Computation, 289(??):481–485, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630340X.

Yuan:2018:EIU

[YZZ18] Yongxin Yuan, Kezheng Zuo, and Taizhong Zhang. An efficient iterative updating method for hysteretic damping models. Applied Mathematics and Computation, 320(??):86– 98, March 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317306434.

Zayed:2016:EAE

[ZA16] E. M. E. Zayed and K. A. E. Alurrfi. Extended auxiliary equation method and its applications for finding the exact solutions for a class of nonlinear Schrödingertype equations. Applied Mathematics and Computation, 289(??):111–131, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302648.

Zadehgol:2016:FIP

[Zad16] Ata Zadehgol. A frequencyindependent and parallel algorithm for computing the zeros of strictly proper rational transfer functions. Applied Mathematics and Computation, 274(??):229–236, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014848.

Zaman:2019:IMR

[Zam19] Tolga Zaman. Improvement of modified ratio estimators using robust regression methods. Applied Mathematics and Computation, 348(??):627–631, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310890.

Zada:2019:UTS

[ZAP19] Akbar Zada, Wajid Ali, and Choonkil Park. Ulam’s type stability of higher order nonlinear delay differential equations via integral inequality of Grönwall– Bellman–Bihari’s type. Applied Mathematics and Computation, 350(??):60– 65, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300220.

Zhang:2015:EIE

[ZAW15] Lihong Zhang, Bashir Ahmad, and Guotao Wang. Explicit iterations and extremal solutions for fractional differential equations with nonlinear integral boundary conditions. Applied Mathematics and Computation, 268(??):388–392, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008309.

Zhai:2018:CSV

[ZAY18] Shaohui Zhai, Dalal Alrowaili, and Dong Ye. Clar structures vs. Fries structures in hexagonal systems. Applied Mathematics and Computation, 329(??):384–394, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301164.

Zeng:2019:AEF

[ZAZ19] Weijun Zeng, Hongfeng Ai, and Man Zhao. Asymmetrical expectations of future interaction and cooperation in the iterated prisoner’s dilemma game. Applied Mathematics and Computation, 359(??):148–164, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303558.

Zapata:2017:HOI

[ZB17] Miguel Uh Zapata and Reymundo Itzá Balam. High-order implicit finite difference schemes for the two-dimensional Poisson equation. Applied Mathematics and Computation, 309(??):222–244, September 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302412.

Zeng:2017:FDE

[ZBBW17] Shengda Zeng, Dumitru Baleanu, Yunru Bai, and Guocheng Wu. Fractional differential equations of Caputo–Katugampola type and numerical solutions. Applied Mathematics and Computation, 315(??): 549–554, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S009630031730454X.

Zhao:2015:FEM

[ZBH^{+}15] Yanmin Zhao, Weiping Bu, Jianfei Huang, Da-Yan Liu, and Yifa Tang. Finite element method for two-dimensional spacefractional advection-dispersion equations. Applied Mathematics and Computation, 257(??):553–565, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000302.

Zaji:2016:FOA

[ZBKS16] Amir Hossein Zaji, Hossein Bonakdari, Saeed Reza Khodashenas, and Shahaboddin Shamshirband. Firefly optimization algorithm effect on support vector regression prediction improvement of a modified labyrinth side weir’s discharge coefficient. Applied Mathematics and Computation, 274(??):14– 19, February 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315014216.

Zhong:2015:FOP

[ZBL15] Yinhui Zhong, Qunfang Bao, and Shenghong Li. FX options pricing in logarithmic meanreversion jump-diffusion model with stochastic volatility. Applied Mathematics and Computation, 251(??):1–13, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031401563X.

Ziadi:2016:CGO

[ZBME16] Raouf Ziadi, Abdelatif Bencherif-Madani, and Rachid Ellaia. Continuous global optimization through the generation of parametric curves. Applied Mathematics and Computation, 282(??):65–83, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300765.

Zabadal:2015:ASW

[ZBR15] Jorge Zabadal, Bardo Bodmann, and Vinicius Ribeiro. Analytical solutions for water pollution problems using quasi-conformal mappings. Applied Mathematics and Computation, 264(??):218–222, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005445.

Zhou:2018:IRT

[ZC18] Mengmeng Zhou and Jianlong Chen. Integral representations of two generalized core inverses. Applied Mathematics and Computation, 333(??):187–193, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302662.

Zhang:2019:SGD

[ZC19a] Tian Zhang and Huabin Chen. The stability with a general decay of stochastic delay differential equations with Markovian switching. Applied Mathematics and Computation, 359(??):294–307, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303455.

Zhang:2019:AWG

[ZC19b] Tie Zhang and Yanli Chen. An analysis of the weak Galerkin finite element method for convection-diffusion equations. Applied Mathematics and Computation, 346(??):612–621, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309305.

Zhu:2019:LCI

[ZC19c] Bao-Xuan Zhu and Yun Chen. Log-concavity of independence polynomials of some kinds of trees. Applied Mathematics and Computation, 342(??):35– 44, February 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318308002.

Zhang:2019:FTA

[ZCAN19] Dian Zhang, Jun Cheng, Choon Ki Ahn, and Hongjie Ni. A flexible terminal approach to stochastic stability and stabilization of continuous-time semi-Markovian jump systems with time-varying delay. Applied Mathematics and Computation, 342(??):191–205, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308130.

Zhou:2015:MHI

[ZCC15] Duanmei Zhou, Guoliang Chen, and Qingyou Cai. On modified HSS iteration methods for continuous Sylvester equations. Applied Mathematics and Computation, 263(??):84– 93, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004713.

Zhou:2018:OEE

[ZCCL18] Yujie Zhou, Fanfan Chen, Jiaxiang Cai, and Hua Liang. Optimal error estimate for energy-preserving splitting schemes for Maxwell’s equations. Applied Mathematics and Computation, 333(??):32–41, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302649.

Zhang:2019:FTS

[ZCCZ19] Dian Zhang, Jun Cheng, Jinde Cao, and Dan Zhang. Finite-time synchronization control for semi-Markov jump neural networks with modedependent stochastic parametric uncertainties. Applied Mathematics and Computation, 344–345 (??):230–242, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318307859.

Zhu:2018:SEL

[ZCEW18] Mengkun Zhu, Yang Chen, Niall Emmart, and Charles Weems. The smallest eigenvalue of large Hankel matrices. Applied Mathematics and Computation, 334(??):375–387, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303242.

Zhang:2019:IFE

[ZCFW19] Huisen Zhang, Yongli Cai, Shengmao Fu, and Weiming Wang. Impact of the fear effect in a prey-predator model incorporating a prey refuge. Applied Mathematics and Computation, 356(??):328–337, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302358.

Zhao:2018:SEC

[ZCL18] Xingwang Zhao, Fuyuan Cao, and Jiye Liang. A sequential ensemble clusterings generation algorithm for mixed data. Applied Mathematics and Computation, 335(??):264–277, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303539.

Zhang:2018:ESS

[ZCLS18] Yanling Zhang, Xiaojie Chen, Aizhi Liu, and Changyin Sun. The effect of the stake size on the evolution of fairness. Applied Mathematics and Computation, 321(??):641–653, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307920.

Zhao:2019:BEC

[ZCML19] Bin Zhao, Wengu Chen, Jixiang Meng, and Fengxia Liu. Book embedding of complex network with community structure. Applied Mathematics and Computation, 361(??):747–751, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304850.

Zhu:2017:CAI

[ZCPM17] Huihui Zhu, Jianlong Chen, Pedro Patrício, and Xavier Mary. Centralizer’s applications to the inverse along an element. Applied Mathematics and Computation, 315(??):27–33, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305076.

Zhao:2015:MSC

[ZCQ15] Yulin Zhao, Haibo Chen, and Bin Qin. Multiple solutions for a coupled system of nonlinear fractional differential equations via variational methods. Applied Mathematics and Computation, 257(??):417–427, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017895.

Zhanlav:2015:HOA

[ZCU15] T. Zhanlav, O. Chuluunbaatar, and V. Ulziibayar. Higher-order accurate numerical solution of unsteady Burgers’ equation. Applied Mathematics and Computation, 250(??):701–707, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314015197.

Zhanlav:2017:GFM

[ZCU17] T. Zhanlav, O. Chuluunbaatar, and V. Ulziibayar. Generating function method for constructing new iterations. Applied Mathematics and Computation, 315(??):414–423, December 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317305398.

Zhang:2019:FTO

[ZCW19a] Jiancheng Zhang, Mohammed Chadli, and Yan Wang. A fixed-time observer for discrete-time singular systems with unknown inputs. Applied Mathematics and Computation, 363(??): Article 124586, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305788.

Zhang:2019:NHA

[ZCW^{+}19b] X. Zhang, M. Y. Chen, M. G. Wang, Y. E. Ge, and H. E. Stanley. A novel hybrid approach to Baltic Dry Index forecasting based on a combined dynamic fluctuation network and artificial intelligence method. Applied Mathematics and Computation, 361(??):499–516, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930445X.

Zhao:2017:NSI

[ZCX17] Yulin Zhao, Haibo Chen, and Chengjie Xu. Nontrivial solutions for impulsive fractional differential equations via Morse theory. Applied Mathematics and Computation, 307(??):170–179, August 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730156X.

Zhang:2019:NSC

[ZCX19] Hongmei Zhang, Jinde Cao, and Lianglin Xiong. Novel synchronization conditions for time-varying delayed Lur’e system with parametric uncertainty. Applied Mathematics and Computation, 350(??):224–236, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318311342.

Zhang:2018:CMG

[ZCXX18] Qifeng Zhang, Mengzhe Chen, Yinghong Xu, and Dinghua Xu. Compact θ-method for the generalized delay diffusion equation. Applied Mathematics and Computation, 316(??):357–369, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305842.

Zhao:2015:RPE

[ZCZM15] Jikun Zhao, Shaochun Chen, Bei Zhang, and Shipeng Mao. Robust a posteriori error estimates for conforming and nonconforming finite element methods for convection-diffusion problems. Applied Mathematics and Computation, 264(??):346–358, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315005536.

Zhang:2017:NIM

[ZD17] Jianhua Zhang and Hua Dai. A new iterative method for solving complex symmetric linear systems. Applied Mathematics and Computation, 302(??):9–20, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300024.

Zhou:2017:ETC

[ZDF17] Jiamu Zhou, Hailing Dong, and Jianwen Feng. Event-triggered communication for synchronization of Markovian jump delayed complex networks with partially unknown transition rates. Applied Mathematics and Computation, 293(??):617–629, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304106.

Zhou:2019:IDM

[ZDJ19] Hui Zhou, Qi Ding, and Ruiling Jia. Inverse of the distance matrix of a weighted cactoid digraph. Applied Mathematics and Computation, 362 (??):Article 124552, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305351.

Zhang:2015:ABF

[ZDK15] Yan Hui Zhang, Guan Tie Deng, and Kit Ian Kou. Asymptotic behavior of fractional Laplacians in the half space. Applied Mathematics and Computation, 254(??):125–132, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314017056.

Zhang:2015:ASC

[ZDPZ15] Chaolong Zhang, Feiqi Deng, Yunjian Peng, and Bo Zhang. Adaptive synchronization of Cohen– Grossberg neural network with mixed time-varying delays and stochastic perturbation. Applied Mathematics and Computation, 269(??):792–801, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009935.

Zhao:2015:BMA

[ZDZ15] Bin Zhao, Thirumalai Diraviyam, and Xiaoying Zhang. A biomathematical approach: Speculations to construct virtual placenta. Applied Mathematics and Computation, 256(??):344–351, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017846.

Zahra:2019:RSN

[ZEF19] W. K. Zahra, S. M.

Elkholy, and M. Fahmy. Rational spline-nonstandard finite difference scheme for the solution of timefractional Swift–Hohenberg equation. Applied Mathematics and Computation, 343(??):372–387, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307872.

Zeidan:2016:AMT

[Zei16] D. Zeidan. Assessment of mixture two-phase flow equations for volcanic flows using Godunov-type methods. Applied Mathematics and Computation, 272 (part 3)(??):707–719, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315012746.

Zheng:2019:SAV

[ZEW19] Xiangcheng Zheng, V. J. Ervin, and Hong Wang. Spectral approximation of a variable coefficient fractional diffusion equation in one space dimension. Applied Mathematics and Computation, 361(??):98–111, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304102.

Zhang:2018:VEM

[ZF18] Baiju Zhang and Minfu Feng. Virtual element method for two-dimensional linear elasticity problem in mixed weakly symmetric formulation. Applied Mathematics and Computation, 328(??):1– 25, July 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318300365.

Zhang:2018:GWP

[ZFL18] Rong Zhang, Mingshu Fan, and Shan Li. Global well-posedness of incompressible Bénard problem with zero dissipation or zero thermal diffusivity. Applied Mathematics and Computation, 321(??):442–449, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317307695.

Zhao:2015:FTS

[ZFZK15] Ping Zhao, Wei Feng, Yan Zhao, and Yu Kang. Finite-time stochastic inputto-state stability of switched stochastic nonlinear systems. Applied Mathematics and Computation, 268(??):1038–1054, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008565.

Zhao:2015:ASG

[ZG15a] Xiuyuan Zhao and Zuoxin Gan. The absolute stability for general discrete Lurie systems with superposition of multiple nonlinearities. Applied Mathematics and Computation, 259(??):490–496, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315002763.

Zheng:2015:STP

[ZG15b] Z. C. Zheng and A. S. Ghate. A solution of two-parameter asymptotic expansions for a two-dimensional unsteady boundary layer. Applied Mathematics and Computation, 270(??):90–104, November 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315010711.

Zhang:2018:GAA

[ZG18] Suxia Zhang and Hongbin Guo. Global analysis of age-structured multi-stage epidemic models for infectious diseases. Applied Mathematics and Computation, 337(??):190–213, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304235.

Zhan:2019:PTE

[ZGGY19] Weijun Zhan, Yan Gao, Qian Guo, and Xiaofeng Yao. The partially truncated Euler–Maruyama method for nonlinear pantograph stochastic differential equations. Applied Mathematics and Computation, 346(??):109–126, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309184.

Zhang:2016:MII

[ZGH16] Hong Zhang, Paul Georgescu, and Adamu Shitu Hassan. Mathematical insights and integrated strategies for the control of aedes aegypti mosquito. Applied Mathematics and Computation, 273(??):1059–1089, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014174.

Zhang:2017:HMS

[ZGH17] Yi Zhang, Jinwu Gao, and Zhiyong Huang. Hamming method for solving uncertain differential equations. Applied Mathematics and Computation, 313(??):331–341, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303983.

Zhang:2015:ORA

[ZGL15] Li-Tao Zhang, Tong-Xiang Gu, and Xing-Ping Liu. Overlapping restricted additive Schwarz method with damping factor for H-matrix linear complementarity problem. Applied Mathematics and Computation, 271(??):1–10, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501156X.

Zheng:2016:RIB

[ZGZ16] Baojing Zheng, Xiaowei Gao, and Ch. Zhang. Radial integration BEM for vibration analysis of twoand three-dimensional elasticity structures. Applied Mathematics and Computation, 277(??):111–126, March 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031501615X.

Zhang:2015:BTW

[ZH15a] Kelei Zhang and Junqiang Han. Bifurcations of traveling wave solutions for the (2 + 1)-dimensional generalized asymmetric Nizhnik–Novikov–Veselov equation. Applied Mathematics and Computation, 251(??):108–117, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015732.

Zhu:2015:QBB

[ZH15b] Yuanpeng Zhu and Xuli Han. Quasi-Bernstein– Bézier polynomials over triangular domain with multiple shape parameters. Applied Mathematics and Computation, 250(??):181–192, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014738.

Zhou:2016:NAN

[ZH16] Shaobo Zhou and Yangzi Hu. Numerical approximation for nonlinear stochastic pantograph equations with Markovian switching. Applied Mathematics and Computation, 286(??):126–138, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302417.

Zhang:2017:CPR

[ZH17] Zhimin Zhang and Xiao Han. The compound Poisson risk model under a mixed dividend strategy. Applied Mathematics and Computation, 315(??):1–12, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031730509X.

Zhang:2019:ELE

[ZH19a] Hai Zhang and Qing-Yi Hao. Eisenhart lift for Euler’s problem of two fixed centers. Applied Mathematics and Computation, 350(??):305–312, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300311.

Zhao:2019:GCE

[ZH19b] Shu-Li Zhao and Rong-Xia Hao. The generalized 4connectivity of exchanged hypercubes. Applied Mathematics and Computation, 347(??):342–353, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0096300318309974.

Zhang:2015:MDS

[Zha15a] Fuchen Zhang. On a model of the dynamical systems describing convective fluid motion in rotating cavity. Applied Mathematics and Computation, 268(??):873–882, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009017.

Zhang:2015:RNP

[Zha15b]  Jian-Jun Zhang.  The

relaxed nonlinear PHSSlike iteration method for absolute value equations. Applied Mathematics and Computation, 265(??):266–274, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315006347.

Zhang:2015:IDB

[Zha15c] Ren-Jiang Zhang. Improved derivative bounds of the rational quadratic Bézier curves. Applied Mathematics and Computation, 250(??):492–496, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014957.

Zhang:2015:NTA

[Zha15d] Renjiang Zhang. A note on the trigonometric approximation of Lip(ω(t), p)- class. Applied Mathematics and Computation, 269(??):129–132, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315010127.

Zhang:2015:PEN

[Zha15e] Tailei Zhang. Permanence and extinction in a nonautonomous discrete SIRVS epidemic model with vaccination. Applied Mathematics and Computation, 271(??):716–729, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013090.

Zhang:2015:CGS

[Zha15f] Xianmin Zhang. On the concept of general solution for impulsive differential equations of fractionalorder q ∈ (1, 2). Applied Mathematics and Computation, 268(??):103–120, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007547.

Zhao:2015:RON

[Zha15g] Dianli Zhao. A remark on one non-autonomous stochastic Gompertz model with delay. Applied Mathematics and Computation, 266(??):369–373, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500675X.

Zhao:2015:CWE

[Zha15h] Junjian Zhao. The convergence of wavelet expansion with divergence-free properties in vector-valued Besov spaces. Applied Mathematics and Computation, 251(??):143–153, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015756.

Zhao:2015:FTN

[Zha15i] Na Zhao. Finite termination of a Newton-type algorithm based on a new class of smoothing functions for the affine variational inequality problem. Applied Mathematics and Computation, 270(??):926–934, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011005.

Zhang:2016:CNP

[Zha16a] Jian Zhang. The critical Neumann problem of Kirchhoff type. Applied Mathematics and Computation, 274(??):519–530, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015064.

Zhang:2016:MMN

[Zha16b] Qian Zhang. Mathematical modeling and numerical study of carbonation in porous concrete materials. Applied Mathematics and Computation, 281(??):16–27, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300340.

Zhang:2016:GWP

[Zha16c]  Qin Zhang.  Global

well-posedness for the 21/2 D Bénard system with partial viscosity terms. Applied Mathematics and Computation, 283(??):282–289, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301540.

Zhang:2016:SGS

[Zha16d] Xiang Zhang. A system of generalized Sylvester quaternion matrix equations and its applications. Applied Mathematics and Computation, 273(??):74–81, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013120.

Zhang:2017:HOR

[Zha17a] Gui-Lai Zhang. High order Runge–Kutta methods for impulsive delay differential equations. Applied Mathematics and Computation, 313(??):12–23, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317303661.

Zhang:2017:OAL

[Zha17b] Ren-Jiang Zhang. Optimal asymptotic Lebesgue constant of Berrut’s rational interpolation operator for equidistant nodes. Applied Mathematics and Computation, 294(??):139–145, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630563X.

Zhang:2017:WRC

[Zha17c] Zujin Zhang. On weighted regularity criteria for the axisymmetric Navier– Stokes equations. Applied Mathematics and Computation, 296(??):18–22, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305999.

Zhang:2018:ERC

[Zha18a] Huamin Zhang. The eigenvalues range of a class of matrices and some applications in Cauchy– Schwarz inequality and iterative methods. Applied Mathematics and Computation, 321(??):37–48, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031730704X.

Zhang:2018:SES

[Zha18b]  Jingyuan Zhang.  A

stable explicitly solvable numerical method for the Riesz fractional advection-dispersion equations. Applied Mathematics and Computation, 332(??):209–227, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302339.

Zhang:2018:TPM

[Zha18c] Shougui Zhang. Two projection methods for the solution of Signorini problems. Applied Mathematics and Computation, 326 (??):75–86, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318300109.

Zhang:2018:RSU

[Zha18d] Zhehao Zhang. Renewal sums under mixtures of exponentials. Applied Mathematics and Computation, 337(??):267–280, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304405.

Zhang:2018:NSB

[Zha18e] Zhongxin Zhang. Normal solutions of a boundaryvalue problem arising in free convection boundarylayer flows in porous media. Applied Mathematics and Computation, 339(??):346–366, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305496.

Zhang:2019:UBS

[Zha19a] Xin Zhang. Upper bound on the sum of powers of the degrees of graphs with few crossings per edge. Applied Mathematics and Computation, 350(??):163–169, June 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300104.

Zhang:2019:CSC

[Zha19b] Yingnan Zhang. Controller synthesis for continuous T-S fuzzy systems with partly immeasurable premise variables. Applied Mathematics and Computation, 363 (??):Article 124592, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305843.

Zhang:2019:SEP

[Zha19c]  Zhehao Zhang.  On

the stochastic equation L(Z) = L[V (X +Z)] properties of Mittag-Leffler distributions. Applied Mathematics and Computation, 361(??):365–376, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930390X.

Zhang:2017:DDS

[ZHJ^{+}17] Chuan-Ke Zhang, Yong He, Lin Jiang, Wen-Juan Lin, and Min Wu. Delaydependent stability analysis of neural networks with time-varying delay: a generalized free-weighting-matrix approach. Applied Mathematics and Computation, 294(??):102–120, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305380.

Zhou:2015:GEB

[Zho15a] Jun Zhou. Global existence and blow-up of solutions for a Kirchhoff type plate equation with damping. Applied Mathematics and Computation, 265(??):807–818, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007298.

Zhou:2015:ESN

[Zho15b] Shaobo Zhou. Exponential stability of numerical solution to neutral stochastic functional differential equation. Applied Mathematics and Computation, 266(??):441–461, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500658X.

Zhou:2015:RAF

[Zho15c] Yong Zhou. Recent advances in fractional differential equations. Applied Mathematics and Computation, 257(??):1, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500315X.

Zhou:2015:FIE

[Zho15d] Zhengxin Zhou. On the first integral and equivalence of nonlinear differential equations. Applied Mathematics and Computation, 268(??):295–302, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008462.

Zhou:2015:SSH

[Zho15e] Zhengxin Zhou. Subharmonic solutions of Hill’s equation. Applied Mathematics and Computation, 253(??):17–22, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017688.

Zhou:2017:QPE

[Zho17] Jun Zhou. Quenching for a parabolic equation with variable coefficient modeling MEMS technology. Applied Mathematics and Computation, 314(??):7–11, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304393.

Zhao:2018:NSN

[ZHS18] Wei Zhao, Y. C. Hon, and Martin Stoll. Numerical simulations of nonlocal phase-field and hyperbolic nonlocal phase-field models via localized radial basis functions-based pseudo-spectral method (LRBF-PSM). Applied Mathematics and Computation, 337(??):504–513, November 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318304740.

Zhu:2015:CNL

[Zhu15]  Xiaojing Zhu.  Com-

puting the nearest lowrank correlation matrix by a simplified SQP algorithm. Applied Mathematics and Computation, 256(??):404–414, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000582.

Zhu:2016:ENG

[Zhu16] Xiangling Zhu. Essential norm of generalized weighted composition operators on Bloch-type spaces. Applied Mathematics and Computation, 274(??):133–142, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014125.

Zhu:2018:PPR

[Zhu18]  Yuanpeng Zhu.  C^{2}

positivity-preserving rational interpolation splines in one and two dimensions. Applied Mathematics and Computation, 316(??):186–204, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305775.

Zhang:2016:SBP

[ZHW16] Xueli Zhang, Yehui Huang, and Peixuan Weng. Stability and bifurcation of a predator–prey model with disease in the prey and temporal–spatial nonlocal effect. Applied Mathematics and Computation, 290(??):467–486, November 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304301.

Zhao:2018:RDE

[ZHW^{+}18] Xueqing Zhao, Keke Huang, Xiaoming Wang, Meihong Shi, Xinjuan Zhu, Quanli Gao, and Zhaofei Yu. Reaction–diffusion equation based image restoration. Applied Mathematics and Computation, 338(??):567–587, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305435.

Zhu:2015:MMM

[ZHWW15] L. H. Zhu, Y. Huang, R. H. Wang, and J. Y. Wang. A mathematical model of the motion of cutting particles in reverse circulation air drilling. Applied Mathematics and Computation, 256(??):192–202, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500017X.

Zhang:2017:ESCb

[ZHWW17] Zhi-Ming Zhang, Yong He, Min Wu, and Qing-Guo Wang. Exponential synchronization of chaotic neural networks with timevarying delay via intermittent output feedback approach. Applied Mathematics and Computation, 314(??):121–132, December 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304708.

Zhao:2017:CPA

[ZHX^{+}17] Fuqiang Zhao, Qingxue Huang, Jiaquan Xie, Yugui Li, Lifeng Ma, and Jianmei Wang. Chebyshev polynomials approach for numerically solving system of two-dimensional fractional PDEs and convergence analysis. Applied Mathematics and Computation, 313(??):321–330, November 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317303697.

Zhang:2016:IFE

[ZHY16] Gui-Qing Zhang, Tao-Ping Hu, and Zi Yu. An improved fitness evaluation mechanism with noise in prisoner’s dilemma game. Applied Mathematics and Computation, 276(??):31– 36, March 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315016197.

Zegeling:2018:NFD

[ZI18] Paul Andries Zegeling and Sehar Iqbal. Nonstandard finite differences for a truncated Bratu– Picard model. Applied Mathematics and Computation, 324(??):266–284, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308585.

Zilinskas:2015:VSA

[Zil15] Antanas Zilinskas. Visualization of a statistical approximation of the Pareto front. Applied Mathematics and Computation, 271(??):694–700, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315012904.

Zhang:2015:CSL

[ZJ15a] Qiumei Zhang and Daqing Jiang. The coexistence of a stochastic Lotka– Volterra model with two predators competing for one prey. Applied Mathematics and Computation, 269(??):288–300, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500973X.

Zhao:2015:FTC

[ZJ15b] Lin Zhao and Yingmin Jia. Finite-time consensus for second-order stochastic multi-agent systems with nonlinear dynamics. Applied Mathematics and Computation, 270(??):278–290, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011030.

Zhang:2019:CIS

[ZjA19] Liang Zhang and Ji jun Ao. On a class of inverse Sturm–Liouville problems with eigenparameter-dependent boundary conditions. Applied Mathematics and Computation, 362(??): Article 124553, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305363.

Zhang:2015:AML

[ZJJ15] Zhaozhong Zhang, Ziwu Jiang, and Tongsong Jiang. Algebraic methods for least squares problem in split quaternionic mechanics. Applied Mathematics and Computation, 269(??):618–625, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315009911.

Zhang:2018:DCN

[ZJJ18] Tong Zhang, JiaoJiao Jin, and Tao Jiang. The decoupled Crank–Nicolson/ Adams–Bashforth scheme for the Boussinesq equations with nonsmooth initial data. Applied Mathematics and Computation, 337(??):214–233, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303953.

Zu:2018:EPL

[ZJO^{+}18] Li Zu, Daqing Jiang, Donal O’Regan, Tasawar Hayat, and Bashir Ahmad. Ergodic property of a Lotka–Volterra predator– prey model with white noise higher order perturbation under regime switching. Applied Mathematics and Computation, 330(??):93–102, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301371.

Zhang:2019:PMN

[ZjW19] Chao Zhang and Tian jun Wang. Pseudospectral method for non-isotropic heat transfer using mixed Hermite-Legendre interpolation. Applied Mathematics and Computation, 343(??):328–341, February 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318307963.

Zhang:2018:TSS

[ZJY18] Hui Zhang, Xiaoyun Jiang, and Xiu Yang. A timespace spectral method for the time-space fractional Fokker–Planck equation and its inverse problem. Applied Mathematics and Computation, 320(??):302–318, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317306689.

Zhao:2017:SMB

[ZJYD17] Lin Zhao, Yingmin Jia, Jinpeng Yu, and Junping Du. H_{∞} sliding mode based scaled consensus control for linear multi-agent systems with disturbances. Applied Mathematics and Computation, 292(??):375–389, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304866.

Zhang:2015:BTE

[ZJZ15] Qiuyun Zhang, Zhaoxin Jiang, and Sining Zheng. Blow-up time estimate for a degenerate diffusion equation with gradient absorption. Applied Mathematics and Computation, 251(??):331–335, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314015902.

Zheng:2017:CCI

[ZK17] Xiying Zheng and Bo Kong. Cyclic codes and λ_{1}+λ_{2}u+ λ_{3}v + λ_{4}uv-constacyclic codes over F_{p} + uF_{p} + vF_{p} + uvF_{p}. Applied Mathematics and Computation, 306(??):86–91, August 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301212.

Zubik-Kowal:2018:APF

[ZK18] B. Zubik-Kowal. An algorithm for partial functional differential equations modeling tumor growth. Applied Mathematics and Computation, 321(??):85–92, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306562.

Zhang:2018:SMJ

[ZKKZ18] Caihong Zhang, Yonggui Kao, Binghua Kao, and Tiezhu Zhang. Stability of Markovian jump stochastic parabolic Itô equations with generally uncertain transition rates. Applied Mathematics and Computation, 337(??):386–398, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303680.

Zadin:2015:AMM

[ZKV^{+}15] V. Zadin, H. Kasemägi, V. Valdna, S. Vigonski, M. Veske, and A. Aabloo. Application of multiphysics and multiscale simulations to optimize industrial wood drying kilns. Applied Mathematics and Computation, 267(??):465–475, September 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001368.

Zhang:2015:AFV

[ZL15a] Tie Zhang and Zheng Li. An analysis of finite volume element method for solving the Signorini problem. Applied Mathematics and Computation, 270(??):830–841, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315011625.

Zhao:2015:CID

[ZL15b] Di Zhao and Hongyi Li. On the computation of inverses and determinants of a kind of special matrices. Applied Mathematics and Computation, 250(??):721–726, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314013484.

Zhu:2015:PFS

[ZL15c] Song-Ping Zhu and Guang-Hua Lian. Pricing forwardstart variance swaps with stochastic volatility. Applied Mathematics and Computation, 250(??): 920–933, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314014258.

Zhang:2016:CNS

[ZL16a] Jin Zhang and Xiaowei Liu. Convergence in L^{2} norm of the SDFEM on a Shishkin triangular mesh for problems with characteristic layers. Applied Mathematics and Computation, 287– 288(??):171–183, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302855.

Zhang:2016:ECG

[ZL16b] Minjie Zhang and Shuchao Li. Extremal cacti of given matching number with respect to the distance spectral radius. Applied Mathematics and Computation, 291(??):89–97, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316304027.

Zhang:2017:PEE

[ZL17a] Tong Zhang and ShiShun Li. A posteriori error estimates of finite element method for the timedependent Navier–Stokes equations. Applied Mathematics and Computation, 315(??):13–26, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317304563.

Zhao:2017:TDR

[ZL17b] Caidi Zhao and Bei Li. Time decay rate of weak solutions to the generalized MHD equations in R^{2}. Applied Mathematics and Computation, 292(??):1–8, January 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304660.

Zhang:2018:IIT

[ZL18] Wenpeng Zhang and Xin Lin. Identities involving trigonometric functions and Bernoulli numbers. Applied Mathematics and Computation, 334(??):288–294, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303278.

Zhao:2019:RAP

[ZL19a] Dazhi Zhao and Maokang Luo. Representations of acting processes and memory effects: General fractional derivative and its application to theory of heat conduction with finite wave speeds. Applied Mathematics and Computation, 346(??):531–544, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309032. See

supplementary remarks [LSZ^{+}19b].

Zhao:2019:SRR

[ZL19b] Dazhi Zhao and Maokang Luo. Supplementary remark to ‘Representations of acting processes and memory effects: General fractional derivative and its application to theory of heat conduction with finite wave speeds’ [Applied Mathematics and Computation 346 (2019) 531–544]. Applied Mathematics and Computation, 361(??):175–176, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319303777.

See [ZL19a].

Zhao:2019:EPT

[ZL19c] Jinqiu Zhao and Chao Luo. The effect of preferential teaching and memory on cooperation clusters in interdependent networks. Applied Mathematics and Computation, 363 (??):Article 124575, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305582.

Zheng:2015:BMO

[ZLC^{+}15] Xi Zheng, Xi Lu, Felix T. S. Chan, Yong Deng, and Zhen Wang. Bargaining models in opinion dynamics. Applied Mathematics and Computation, 251(??):162–168, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015859.

Zhang:2018:FNS

[ZLC^{+}18] H. Zhang, F. Liu, S. Chen, V. Anh, and J. Chen. Fast numerical simulation of a new time-space fractional option pricing model governing European call option. Applied Mathematics and Computation, 339(??):172–185, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305198.

Zhang:2018:HON

[ZLD18] Yuxin Zhang, Qian Li, and Hengfei Ding. High-order numerical approximation formulas for Riemann– Liouville (Riesz) tempered fractional derivatives: construction and application (I). Applied Mathematics and Computation, 329(??):432–443, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318301255.

Zhai:2016:AFD

[ZLDZ16] Ding Zhai, An-Yang Lu, Jiuxiang Dong, and Qing-Ling Zhang. Asynchronous H_{∞} filtering for 2D discrete Markovian jump systems with sensor failure. Applied Mathematics and Computation, 289(??):60–79, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316303216.

Zhai:2017:SAS

[ZLDZ17] Ding Zhai, An-Yang Lu, Jiuxiang Dong, and Qing-Ling Zhang. Stability analysis and state feedback control of continuoustime T -S fuzzy systems via anew switched fuzzy Lyapunov function approach. Applied Mathematics and Computation, 293(??):586–599, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305562.

Zhang:2016:RWR

[ZLH16] Yu Zhang, Chuanzhong Li, and Jingsong He. Rogue waves in a resonant erbium-doped fiber system with higher-order effects. Applied Mathematics and Computation, 273(??):826–841, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013478.

Zeng:2015:SAC

[ZLHH15] Xu Zeng, Chuandong Li, Tingwen Huang, and Xing He. Stability analysis of complex-valued impulsive systems with time delay. Applied Mathematics and Computation, 256(??):75– 82, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000144.

Zhang:2015:DAD

[ZLJ15] Ruixia Zhang, Deyu Li, and Zhen Jin. Dynamic analysis of a delayed model for vectorborne diseases on bipartite networks. Applied Mathematics and Computation, 263(??):342–352, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005366.

Zhong:2019:ECB

[ZLK^{+}19] Jie Zhong, Yang Liu, Kit Ian Kou, Liangjie Sun, and Jinde Cao. On the ensemble controllability of Boolean control networks using STP method. Applied Mathematics and Computation, 358(??):51–62, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031930267X.

Zhang:2015:PGS

[ZLL15a] Guo-Feng Zhang, Li-Dan Liao, and Zhao-Zheng Liang. On parameterized generalized skew-Hermitian triangular splitting iteration method for singular and nonsingular saddle point problems. Applied Mathematics and Computation, 254(??):340–359, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017810.

Zhang:2015:CVS

[ZLL15b] L. W. Zhang, Z. X. Lei, and K. M. Liew. Computation of vibration solution for functionally graded carbon nanotubereinforced composite thick plates resting on elastic foundations using the element-free IMLS– Ritz method. Applied Mathematics and Computation, 256(??):488–504, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000983.

Zhang:2015:EMP

[ZLL15c] Qi Zhang, Fuyi Li, and Zhanping Liang. Existence of multiple positive solutions to nonhomogeneous Schrödinger– Poisson system. Applied Mathematics and Computation, 259(??):353–363, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002325.

Zhang:2016:ASM

[ZLL16] Pengchong Zhang, Jun Liu, and Gao Lin. Axisymmetric solutions for the multi-layered transversely isotropic piezoelectric medium. Applied Mathematics and Computation, 290(??):355–375, November 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300316303848.

Zhu:2018:GSB

[ZLL18] Xiaoli Zhu, Fuyi Li, and Yuhua Li. Global solutions and blow up solutions to a class of pseudoparabolic equations with nonlocal term. Applied Mathematics and Computation, 329(??):38–51, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300973.

Zhang:2019:ANS

[ZLLA19] Jinghua Zhang, Fawang Liu, Zeng Lin, and Vo Anh. Analytical and numerical solutions of a multi-term timefractional Burgers’ fluid model. Applied Mathematics and Computation, 356(??):1–12, September 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300319301894.

Zhang:2018:SMD

[ZLLP18] Bo Zhang, Hui Li, Shengguo Li, and Jin Peng. Sustainable multi-depot emergency facilities locationrouting problem with uncertain information. Applied Mathematics and Computation, 333(??): 506–520, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300318302522.

Zhai:2016:SFD

[ZLLZ16] Ding Zhai, An-Yang Lu, Jing-Hao Li, and Qing-Ling Zhang. Simultaneous fault detection and control for switched linear systems with mode-dependent average dwell-time. Applied Mathematics and Computation, 273(??):767–792, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014137.

Zhang:2018:DAC

[ZLM^{+}18] Yanhui Zhang, Hongjing Liang, Hui Ma, Qi Zhou, and Zhandong Yu. Distributed adaptive consensus tracking control for nonlinear multi-agent systems with state constraints. Applied Mathematics and Computation, 326(??):16–32, June 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0096300317309050.

Zhang:2016:VPO

[ZLS16] Zhiqiang Zhang, Weiqi Liu, and Yuhong Sheng. Valuation of power option for uncertain financial market. Applied Mathematics and Computation, 286(??):257–264, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630282X.

Zhu:2018:ACM

[ZLSX18] Enqiang Zhu, Zepeng Li, Zehui Shao, and Jin Xu. On acyclically 4colorable maximal planar graphs. Applied Mathematics and Computation, 329(??):402–407, July 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301176.

Zhang:2015:EGE

[ZLW15] Xinhong Zhang, Wenxue Li, and Ke Wang. The existence and global exponential stability of periodic solution for a neutral coupled system on networks with delays. Applied Mathematics and Computation, 264(??):208–217, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005718.

Zhou:2018:NMC

[ZLW18] Zhongguo Zhou, Dong Liang, and Yaushu Wong. The new mass-conserving S-DDM scheme for twodimensional parabolic equations with variable coefficients. Applied Mathematics and Computation, 338(??):869–881, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305095.

Zeng:2019:GFM

[ZLW19] Hong-Bing Zeng, Xiao-Gui Liu, and Wei Wang. A generalized free-matrix-based integral inequality for stability analysis of timevarying delay systems.

Applied Mathematics and Computation, 354(??):1–8, August 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301079.

Zhang:2015:SUT

[ZLWS15] Ancai Zhang, Xuzhi Lai, Min Wu, and Jinhua She. Stabilization of underactuated two-link gymnast robot by using trajectory tracking strategy. Applied Mathematics and Computation, 253(??):193–204, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017196.

Zhang:2015:SAS

[ZLWW15] Xinguang Zhang, Lishan Liu, Yonghong Wu, and B. Wiwatanapataphee. The spectral analysis for a singular fractional differential equation with a signed measure. Applied Mathematics and Computation, 257(??):252–263, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0096300314017226.

Zhao:2019:EFO

[ZLX19] Jingjun Zhao, Yu Li, and Yang Xu. An explicit fourth-order energypreserving scheme for Riesz space fractional nonlinear wave equations. Applied Mathematics and Computation, 351(??):124–138, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319300499.

Zhang:2016:GBL

[ZLZ16] Fuchen Zhang, Xiaofeng Liao, and Guangyun Zhang. On the global boundedness of the Lü system. Applied Mathematics and Computation, 284(??):332–339, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302107.

Zhan:2018:CDE

[ZLZ^{+}18] Xiu-Xiu Zhan, Chuang Liu, Ge Zhou, Zi-Ke Zhang, Gui-Quan Sun, Jonathan J. H. Zhu, and Zhen Jin. Coupling dynamics of epidemic spreading and information diffusion on complex networks. Applied Mathematics and Computation, 332(??):437–448, September 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318302236.

Zhang:2019:IBL

[ZLZ^{+}19] Xueying Zhang, Ruixian Li, Bo Zhang, Yunxiang Yang, Jing Guo, and Xiang Ji. An instance-based learning recommendation algorithm of imbalance handling methods. Applied Mathematics and Computation, 351(??):204–218, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310671.

Zhi:2015:SCB

[ZLZW15] Yongfeng Zhi, Jieliang Li, Jun Zhang, and Zhen Wang. Statistical convergence behavior of affine projection algorithms. Applied Mathematics and Computation, 270(??):511–526, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011091.

Zhang:2015:RSK

[ZM15a] Yi Zhang and Wen-Xiu Ma. Rational solutions to a KdV-like equation. Applied Mathematics and Computation, 256(??):252–256, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000417.

Zhanlav:2015:LIQ

[ZM15b] T. Zhanlav and R. Mijiddorj. On local integro quartic splines. Applied Mathematics and Computation, 269(??):301–307, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009996.

Zheng:2015:FPI

[ZM15c] Qing-Qing Zheng and Chang-Feng Ma. Fast parameterized inexact Uzawa method for complex symmetric linear systems. Applied Mathematics and Computation, 256(??):11– 19, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000375.

Zheng:2016:PAP

[ZM16] Qing-Qing Zheng and Chang-Feng Ma. Preconditioned AHSS–PU alternating splitting iterative methods for saddle point problems. Applied Mathematics and Computation, 273(??):217–225, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315013119.

Zhanlav:2017:CMP

[ZM17] T. Zhanlav and R. Mijiddorj. Convexity and monotonicity properties of the local integro cubic spline. Applied Mathematics and Computation, 293(??):131–137, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305124.

Zhao:2018:SCH

[ZM18a] Shuang Zhao and Jixiang Meng. Sufficient conditions for hypergraphs to be maximally edgeconnected. Applied Mathematics and Computation, 333(??):362–368, September 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830290X.

Zheng:2018:CAP

[ZM18b] Qing-Qing Zheng and Chang-Feng Ma. A class of accelerated parameterized inexact Uzawa algorithms for complex symmetric linear systems. Applied Mathematics and Computation, 320(??):547–556, March 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317306963.

Zhang:2019:UCR

[ZM19] Chao Zhang and Igor Menshov. Using the composite Riemann problem solution for capturing interfaces in compressible two-phase flows. Applied Mathematics and Computation, 363 (??):Article 124610, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306022.

Zhang:2018:DBAa

[ZMA^{+}18] Ge Zhang, Jun Ma, Ahmed Alsaedi, Bashir Ahmad, and Faris Alzahrani. Dynamical behavior and application in Josephson junction coupled by memristor. Applied Mathematics and Computation, 321(??):290–299, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317307634.

Zafarghandi:2019:RBF

[ZMBJ19] Fahimeh Saberi Zafarghandi, Maryam Mohammadi, Esmail Babolian, and Shahnam Javadi. Radial basis functions method for solving the fractional diffusion equations. Applied Mathematics and Computation, 342(??):224–246, February 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318307574.

Zhan:2019:ISN

[ZMC19a] Tao Zhan, Shuping Ma, and Hao Chen. Impulsive stabilization of nonlinear singular switched systems with all unstablemode subsystems. Applied Mathematics and Computation, 344–345(??):57–67, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308695.

Zhao:2019:RSM

[ZMC^{+}19b] Wenying Zhao, Yuechao Ma, Aihong Chen, Lei Fu, and Yutong Zhang. Robust sliding mode control for Markovian jump singular systems with randomly changing structure. Applied Mathematics and Computation, 349(??):81– 96, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310610.

Zheng:2018:DCR

[ZMH18] Pan Zheng, Chunlai Mu, and Xuegang Hu. Deadcore rate for the fast diffusion equation with strong absorption in higher dimensional cases. Applied Mathematics and Computation, 335(??):292–304, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031830328X.

Zhang:2017:FHF

[ZML^{+}17] Jie Zhang, Lifeng Ma, Yurong Liu, Ming Lyu, Fuad E. Alsaadi, and Yuming Bo. H_{∞} and I_{2}– I_{∞} finite-horizon filtering with randomly occurring gain variations and quantization effects. Applied Mathematics and Computation, 298(??):171–187, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316306828.

Zhang:2015:PSP

[ZMMZ15] Tongqian Zhang, Wanbiao Ma, Xinzhu Meng, and Tonghua Zhang. Periodic solution of a preypredator model with nonlinear state feedback control. Applied Mathematics and Computation, 266(??):95–107, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315006323.

Meng:2015:ADA

[zMnZyZ15] Xin zhu Meng, Sheng nan Zhao, and Wen yan Zhang. Adaptive dynamics analysis of a predator–prey model with selective disturbance. Applied Mathematics and Computation, 266(??):946–958, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008012.

Zhang:2016:USL

[ZMQ16] Qing-Hua Zhang, Jian Ma, and Yuanyuan Qu. Unified solution for the Legendre equation in the interval [−1, 1] — an example of solving linear singular-ordinary differential equations. Applied Mathematics and Computation, 289(??):311–323, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630337X.

Zheng:2015:EBS

[ZMT15] Pan Zheng, Chunlai Mu, and Ya Tian. Extinction behavior of solutions for a quasilinear parabolic system with nonlocal sources. Applied Mathematics and Computation, 259(??):587–595, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300315002969.

Zhang:2018:SPE

[ZMZ18] Yufeng Zhang, Jianqin Mei, and Xiangzhi Zhang. Symmetry properties and explicit solutions of some nonlinear differential and fractional equations. Applied Mathematics and Computation, 337(??): 399–407, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318304399.

Zhang:2015:NRQ

[ZN15] Hao Zhang and Qin Ni. A new regularized quasi-Newton algorithm for unconstrained optimization. Applied Mathematics and Computation, 259(??):460–469, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500212X.

Zahra:2019:EFM

[ZNV19] W. K. Zahra, M. A. Nasr, and M. Van Daele. Exponentially fitted methods for solving time fractional nonlinear reaction– diffusion equation. Applied Mathematics and Computation, 358(??):468–490, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319303017.

Zarebnia:2016:NTA

[ZP16] M. Zarebnia and R. Parvaz. On the numerical treatment and analysis of Benjamin–Bona–Mahony– Burgers equation. Applied Mathematics and Computation, 284(??):79–88, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301485.

Zhou:2016:NFO

[ZPM16] Jianping Zhou, Ju H.

Park, and Qian Ma. Nonfragile observer-based H_{∞} control for stochastic timedelay systems. Applied Mathematics and Computation, 291(??):69–83, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303952.

Zanella:2018:SPA

[ZPRZ18] R. Zanella, F. Porta, V. Ruggiero, and M. Zanetti. Serial and parallel approaches for image segmentation by numerical minimization of a secondorder functional. Applied Mathematics and Computation, 318(??):153–175, February 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317304721.

Zhong:2015:ACM

[ZPT^{+}15] Xiaoyun Zhong, Minfang Peng, Chi K. Tse, Shangjiang Guo, and Mohammad Shahidehpour. Analysis and control of multiple chaotic attractors from a three-dimensional system. Applied Mathematics and Computation, 268(??):138–150, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008383.

Zeng:2019:SRS

[ZPZ^{+}19] Deqiang Zeng, Zhilin Pu, Ruimei Zhang, Shouming Zhong, Yajuan Liu, and Guo-Cheng Wu. Stochastic reliable synchronization for coupled Markovian reaction-diffusion neural networks with actuator failures and generalized switching policies. Applied Mathematics and Computation, 357(??):88–106, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302632.

Zhang:2018:CCM

[ZQ18] Jialing Zhang and Kun Qian. The computation of conformal map by harmonic diffeomorphisms between surfaces. Applied Mathematics and Computation, 333(??):536–546, September 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031830170X.

Zheng:2016:ACM

[ZQG16] Quan Zheng, Feng Qin, and Yue Gao. An adaptive coupling method for exterior anisotropic elliptic problems. Applied Mathematics and Computation, 273(??):410–424, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031501351X.

Zhang:2019:UCC

[ZRLX19] Qifeng Zhang, Yunzhu Ren, Xiaoman Lin, and Yinghong Xu. Uniform convergence of compact and BDF methods for the space fractional semilinear delay reaction-diffusion equations. Applied Mathematics and Computation, 358(??):91–110, October 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031930298X.

Zhou:2014:CRO

[ZS14] Xiaojian Zhou and Yongzhong Song. Convergence radius of Osada’s method under center–Hölder continuous condition. Applied Mathematics and Computation, 243(??):809–816, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314009114. See

note [HMGC18].

Zhang:2015:GDS

[ZS15a] Fuchen Zhang and Yonglu Shu. Global dynamics for the simplified Lorenz system model. Applied Mathematics and Computation, 259(??):53–60, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002076.

Zhang:2015:ASR

[ZS15b] G. L. Zhang and M. H.

Song. Asymptotical stability of Runge–Kutta methods for advanced linear impulsive differential equations with piecewise constant arguments. Applied Mathematics and Computation, 259(??):831–837, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002866.

Zhang:2015:PSN

[ZS15c] Guodong Zhang and Yi Shen.

Periodic solutions for a neutral delay Hassell– Varley type predator–prey system. Applied Mathematics and Computation, 264(??):443–452, August 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500572X.

Zhang:2015:ANP

[ZS15d] Shuyi Zhang and Xiaoguang Song. Another note on a paper “Convergence theorem for the common solution for a finite family of φstrongly accretive operator equations”. Applied Mathematics and Computation, 258(??):367–371, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315001563.

Zheng:2015:EDI

[ZS15e] Yanpeng Zheng and Sugoog Shon. Exact determinants and inverses of generalized Lucas skew circulant type matrices. Applied Mathematics and Computation, 270(??):105–113, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315010760.

Zamir:2016:LLS

[ZS16a] Z. Roshan Zamir and N. Sukhorukova. Linear least squares problems involving fixed knots polynomial splines and their singularity study. Applied Mathematics and Computation, 282(??):204–215, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630114X.

Zeidan:2016:PSI

[ZS16b] Dia Zeidan and Theodore E.

Simos. Preface to special issue of selected papers from the 9 ^{th} International Symposium on Numerical Analysis of Fluid Flow and Heat Transfer — Numerical Fluids 2014. Applied Mathematics and Computation, 272 (part 3)(??):581, January 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015106.

Zhang:2016:AST

[ZS16c] Le Zhang and Wolfgang M.

Schmidt. An approximation of small-time probability density functions in a general jump diffusion model. Applied Mathematics and Computation, 273(??):741–758, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315013351.

Zhukovsky:2017:ASH

[ZS17] K. V. Zhukovsky and H. M. Srivastava. Analytical solutions for heat diffusion beyond Fourier law. Applied Mathematics and Computation, 293(??):423–437, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305331.

Zhang:2019:ICR

[ZS19a] Gui-Lai Zhang and Ming-Hui Song. Impulsive continuous Runge–Kutta methods for impulsive delay differential equations. Applied Mathematics and Computation, 341:160–173, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318307331.

Zheng:2019:PIS

[ZS19b] Bo Zheng and Yueqiang Shang. Parallel iterative stabilized finite element algorithms based on the lowest equal-order elements for the stationary Navier–Stokes equations. Applied Mathematics and Computation, 357(??):35–56, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302504.

Zamir:2015:COB

[ZSA^{+}15] Z. Roshan Zamir, N. Sukhorukova, H. Amiel, A. Ugon, and C. Philippe. Convex optimisation-based methods for K-complex detection. Applied Mathematics and Computation, 268(??):947–956, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315009169.

Zhu:2017:SUT

[ZSF17] Tielei Zhu, Haiyan Su, and Xinlong Feng. Some Uzawa-type finite element iterative methods for the steady incompressible magnetohydrodynamic equations. Applied Mathematics and Computation, 302(??):34– 47, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300115.

Zeifman:2019:LCN

[ZSK^{+}19] A. Zeifman, Y. Satin, K. Kiseleva, V. Korolev, and T. Panfilova. On limiting characteristics for a non-stationary two-processor heterogeneous system. Applied Mathematics and Computation, 351(??):48–65, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319300414.

Zhang:2015:ASE

[ZSL15] G. L. Zhang, Minghui Song, and M. Z. Liu. Asymptotical stability of the exact solutions and the numerical solutions for a class of impulsive differential equations. Applied Mathematics and Computation, 258(??):12–21, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001472.

Zhang:2017:SUD

[ZSL17] Yifan Zhang, Gang Shu, and Ya Li. Strategyupdating depending on local environment enhances cooperation in prisoner’s dilemma game. Applied Mathematics and Computation, 301(??):224–232, May 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316307561.

Zhou:2018:HCN

[ZSL^{+}18] Jianping Zhou, Chengyan Sang, Xiao Li, Muyun Fang, and Zhen Wang. H_{∞} consensus for nonlinear stochastic multi-agent systems with time delay. Applied Mathematics and Computation, 325(??):41– 58, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308792.

Zada:2015:HUS

[ZSS15] Akbar Zada, Omar Shah, and Rahim Shah. Hyers– Ulam stability of nonautonomous systems in terms of boundedness of Cauchy problems. Applied Mathematics and Computation, 271(??):512–518, November 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S009630031501276X.

Zhang:2016:RNF

[ZSS16] Yingqi Zhang, Yan Shi, and Peng Shi. Robust and non-fragile finite-time H_{∞} control for uncertain Markovian jump nonlinear systems. Applied Mathematics and Computation, 279(??):125–138, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300121.

Zhang:2017:PIC

[ZSW^{+}17] Hai-Feng Zhang, Pan-Pan Shu, Zhen Wang, Ming Tang, and Michael Small. Preferential imitation can invalidate targeted subsidy policies on seasonalinfluenza diseases. Applied Mathematics and Computation, 294(??):332–342, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316305604.

Zhang:2015:RWK

[ZSX15] Yi Zhang, YanBo Sun, and Wen Xiang. The rogue waves of the KP equation with self-consistent sources. Applied Mathematics and Computation, 263(??):204–213, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500435X.

Zhao:2018:NDC

[ZSXW18] Guoyan Zhao, Mingbo Sun, Songbai Xie, and Hongbo Wang. Numerical dissipation control in an adaptive WCNS with a new smoothness indicator. Applied Mathematics and Computation, 330(??):239–253, August 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318300328.

Zhao:2015:MSS

[ZSY15] Guihua Zhao, Minghui Song, and Zhanwen Yang. Mean-square stability of analytic solution and Euler–Maruyama method for impulsive stochastic differential equations. Applied Mathematics and Computation, 251(??): 527–538, January 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300314016300.

Zhang:2017:MFV

[ZSY17] Qi Zhang, Zhiqiang Sheng, and Guangwei Yuan. A monotone finite volume scheme for diffusion equations on general nonconforming meshes. Applied Mathematics and Computation, 311(??): 300–313, October 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300317303429.

Zhou:2018:PPF

[ZSY18] Huifang Zhou, Zhiqiang Sheng, and Guangwei Yuan. Positivity preserving finite volume scheme for the Nagumo-type equations on distorted meshes. Applied Mathematics and Computation, 336(??):182–192, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303849.

Zhang:2017:SAF

[ZSZ17] Lei Zhang, Qiankun Song, and Zhenjiang Zhao. Stability analysis of fractionalorder complex-valued neural networks with both leakage and discrete delays. Applied Mathematics and Computation, 298(??):296–309, April 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316306956.

Zhi:2016:OSS

[ZSZW16] Yong-Feng Zhi, Fang-Fang Shang, Jun Zhang, and Zhen Wang. Optimal stepsize of pseudo affine projection algorithm. Applied Mathematics and Computation, 273(??):82–88, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315012977.

Zecova:2015:HCM

[ZT15] Monika Zecová and Ján Terpák. Heat conduction modeling by using fractional-order derivatives. Applied Mathematics and Computation, 257(??):365–373, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017974.

Zarei:2016:HBA

[ZT16a] Amin Zarei and Saeed Tavakoli. Hopf bifurcation analysis and ultimate bound estimation of a new 4-D quadratic autonomous hyper-chaotic system. Applied Mathematics and Computation, 291(??):323–339, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304611. See

note [WL18a].

Zhang:2016:WFE

[ZT16b] Tie Zhang and Lixin Tang. A weak finite element method for elliptic problems in one space dimension. Applied Mathematics and Computation, 280(??):1–10, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300182.

Zogheib:2016:NMM

[ZT16c] Bashar Zogheib and Emran Tohidi. A new matrix method for solving two-dimensional timedependent diffusion equations with Dirichlet boundary conditions. Applied Mathematics and Computation, 291(??):1–13, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316303940.

Zhou:2016:FDR

[ZTW16] Shengfan Zhou, Yongxiao Tian, and Zhaojuan Wang. Fractal dimension of random attractors for stochastic non-autonomous reactiondiffusion equations. Applied Mathematics and Computation, 276(??):80– 95, March 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300315016124.

Zhang:2019:MSI

[ZTW19] Wenjie Zhang, Jianhua Tu, and Lidong Wu. A multi-start iterated greedy algorithm for the minimum weight vertex cover P_{3} problem. Applied Mathematics and Computation, 349(??):359–366, May 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318311287.

Zahra:2017:DSM

[ZV17] W. K. Zahra and M. Van Daele. Discrete spline methods for solving two point fractional Bagley– Torvik equation. Applied Mathematics and Computation, 296(??):42–56, March 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316305768.

Zheng:2019:DPM

[ZVL19] Hua Zheng, Seakweng Vong, and Ling Liu. A direct preconditioned modulus-based iteration method for solving nonlinear complementarity problems of H-matrices. Applied Mathematics and Computation, 353(??): 396–405, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319301134.

Zhang:2015:DES

[ZW15a] Maozhu Zhang and Yicao Wang. Dependence of eigenvalues of Sturm– Liouville problems with interface conditions. Applied Mathematics and Computation, 265(??):31–39, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315005998.

Zhang:2015:TDE

[ZW15b] Yufeng Zhang and Lixin Wu. Two (2 + 1)- dimensional expanding dynamical systems associated to the mKP hierarchy. Applied Mathematics and Computation, 268(??):561–574, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008930.

Zhao:2015:PRS

[ZW15c] Jing Zhao and Lisha Wang. Pricing and retail service decisions in fuzzy uncertainty environments. Applied Mathematics and Computation, 250(??):580–592, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314015100.

Zhou:2015:PBR

[ZW15d] Weijun Zhou and Fei Wang. A PRP-based residual method for largescale monotone nonlinear equations. Applied Mathematics and Computation, 261(??):1–7, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003835.

Zhu:2015:NSV

[ZW15e] Li Zhu and Yanxin Wang. Numerical solutions of Volterra integral equation with weakly singular kernel using SCW method. Applied Mathematics and Computation, 260(??):63– 70, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315003793.

Zhang:2018:NAN

[ZW18] Jun Zhang and JinRong Wang. Numerical analysis for Navier–Stokes equations with time fractional derivatives. Applied Mathematics and Computation, 336(??):481–489, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303540.

Zhang:2019:SCS

[ZW19a] Guozhen Zhang and Dajin Wang. Structure connectivity and substructure connectivity of bubble-sort star graph networks. Applied Mathematics and Computation, 363 (??):Article 124632, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319306241.

Zhang:2019:ANC

[ZW19b] Yan Zhang and Fang Wang. Adaptive neural control of non-strict feedback system with actuator failures and time-varying delays. Applied Mathematics and Computation, 362 (??):Article 124512, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304916.

Zhang:2018:RGR

[ZWaY18] Zujin Zhang, Chupeng Wu, and Zheng an Yao. Remarks on global regularity for the 3D MHD system with damping. Applied Mathematics and Computation, 333(??):1–7, September 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300318302200.

Zhang:2018:SAF

[ZWBC18] Jianmei Zhang, Jianwei Wu, Haibo Bao, and Jinde Cao. Synchronization analysis of fractional-order three-neuron BAM neural networks with multiple time delays. Applied Mathematics and Computation, 339(??):431–440, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318305010.

Zhang:2015:TPM

[ZWD15] Xuelu Zhang, Jinting Wang, and Tien Van Do. Threshold properties of the M/M/1 queue under T -policy with applications. Applied Mathematics and Computation, 261(??):284–301, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315004233.

Zhai:2015:ALG

[ZWF15] Shuying Zhai, Zhifeng Weng, and Xinlong Feng. An adaptive local grid refinement method for 2D diffusion equation with variable coefficients based on block-centered finite differences. Applied Mathematics and Computation, 268(??):284–294, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315008644.

Zhang:2015:DDS

[ZWHC15] Lili Zhang, Yinhe Wang, Yuanyuan Huang, and Xuesong Chen. Delaydependent synchronization for non-diffusively coupled time-varying complex dynamical networks. Applied Mathematics and Computation, 259(??):510–522, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016889.

Zhang:2016:AOT

[ZWK16] Shao-Yan Zhang, Qi-Ru Wang, and Qingkai Kong. Asymptotics and oscillation of n th-order nonlinear dynamic equations on time scales. Applied Mathematics and Computation, 275(??):324–334, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015866.

Zhao:2016:MAM

[ZWL16] Guodong Zhao, Yuzhen Wang, and Haitao Li. A matrix approach to modeling and optimization for dynamic games with random entrance. Applied Mathematics and Computation, 290(??):9– 20, November 1, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316303198.

Zhou:2018:SSC

[ZWL18] Qiannan Zhou, Ligong Wang, and Yong Lu. Some sufficient conditions on kconnected graphs. Applied Mathematics and Computation, 325(??):332–339, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830002X.

Zhu:2019:IES

[ZWL^{+}19] Peican Zhu, Xinyu Wang, Shudong Li, Yangming Guo, and Zhen Wang. Investigation of epidemic spreading process on multiplex networks by incorporating fatal properties. Applied Mathematics and Computation, 359(??):512–524, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319301535.

Zhang:2015:SLS

[ZWLZ15] Fengxia Zhang, Musheng Wei, Ying Li, and Jianli Zhao. Special least squares solutions of the quaternion matrix equation AX = B with applications. Applied Mathematics and Computation, 270(??):425–433, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011017.

Zhou:2016:FDF

[ZWM16] Zhifu Zhou, Wei-Tao Wu, and Mehrdad Massoudi. Fully developed flow of a drilling fluid between two rotating cylinders. Applied Mathematics and Computation, 281(??):266–277, April 30, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300316300686.

Zhao:2019:FTA

[ZWO^{+}19] Nan-Nan Zhao, Li-Bing Wu, Xin-Yu Ouyang, Yan Yan, and Rui-Yan Zhang. Finite-time adaptive fuzzy tracking control for nonlinear systems with disturbances and dead-zone nonlinearities. Applied Mathematics and Computation, 362 (??):Article 124494, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304734.

Zhou:2015:GHS

[ZWT15] Rong Zhou, Xiang Wang, and Xiao-Bin Tang. A generalization of the Hermitian and skew-Hermitian splitting iteration method for solving Sylvester equations. Applied Mathematics and Computation, 271(??):609–617, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315012631.

Zhu:2016:IEP

[ZwX16] Lei Zhu and Wei wei Xu. The inverse eigenvalue problem of structured matrices from the design of Hopfield neural networks. Applied Mathematics and Computation, 273(??):1–7, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315011455.

Zhao:2015:FAY

[ZWXW15] Dawei Zhao, Lianhai Wang, Lijuan Xu, and Zhen Wang. Finding another yourself in multiplex networks. Applied Mathematics and Computation, 266(??):599–604, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007304.

Zhao:2015:DAR

[ZWY^{+}15a] Liqiang Zhao, Jianlin Wang, Tao Yu, Huan Jian, and Tangjiang Liu. Design of adaptive robust squareroot cubature Kalman filter with noise statistic estimator. Applied Mathematics and Computation, 256(??):352–367, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314016907.

Zhu:2015:FPP

[ZWY15b] Yun Zhu, Jian Wu, and Gaohang Yu. A fast proximal point algorithm for l_{1}-minimization problem in compressed sensing. Applied Mathematics and Computation, 270(??):777–784, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315011388.

Zhang:2019:EGA

[ZWY19] Suyong Zhang, Chuanxu Wang, and Chao Yu. The evolutionary game analysis and simulation with system dynamics of manufacturer’s emissions abatement behavior under cap-and-trade regulation. Applied Mathematics and Computation, 355(??):343–355, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300319301900.

Zhang:2019:DSS

[ZWZ19a] Jianhua Zhang, Zewen Wang, and Jing Zhao. Double-step scale splitting real-valued iteration method for a class of complex symmetric linear systems. Applied Mathematics and Computation, 353(??):338–346, July 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319301183.

Zhang:2019:SWI

[ZWZ19b] Jie Zhang, Hua Wang, and Xiao-Dong Zhang. The Steiner Wiener index of trees with a given segment sequence. Applied Mathematics and Computation, 344–345(??):20–29, March 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318308671.

Zhen:2018:WPN

[ZWZT18] Zaili Zhen, Jingdong Wei, Jiangbo Zhou, and Lixin Tian. Wave propagation in a nonlocal diffusion epidemic model with nonlocal delayed effects. Applied Mathematics and Computation, 339(??):1–14, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300318305575.

Zhu:2015:FBR

[ZX15] Chun-Gang Zhu and Bao-Yu Xia. A family of bivariate rational Bernstein operators. Applied Mathematics and Computation, 258(??):162–171, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001630.

Zhou:2016:TKC

[ZX16] Fengying Zhou and Xiaoyong Xu. The third kind Chebyshev wavelets collocation method for solving the time-fractional convection diffusion equations with variable coefficients. Applied Mathematics and Computation, 280(??):11–29, April 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300297.

Zhang:2019:FTI

[ZX19] H. M. Zhang and Y. F. Xing. A framework of time integration methods for nonsmooth systems with unilateral constraints. Applied Mathematics and Computation, 363(??): Article 124590, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S009630031930582X.

Zhu:2019:APD

[ZXC19] Min Zhu, Yong Xu, and Jinde Cao. The asymptotic profile of a dengue fever model on a periodically evolving domain. Applied Mathematics and Computation, 362 (??):Article 124531, December 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319305107.

Zhou:2015:MDR

[ZXD15] Zhongbao Zhou, Helu Xiao, and Yingchun Deng. Markov-dependent risk model with multi-layer dividend strategy. Applied Mathematics and Computation, 252(??):273–286, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016701.

Zhou:2018:RKG

[ZXDP18] Shengchao Zhou, Jianhui Xie, Ni Du, and Yan Pang. A random-keys genetic algorithm for scheduling unrelated parallel batch processing machines with different capacities and arbitrary job sizes. Applied Mathematics and Computation, 334(??):254–268, October 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318303424.

Zheng:2015:NSS

[ZXL15] Fu Zheng, Shuangshuang Xu, and Xin Li. Numerical solution of the steady-state probability and reliability of a repairable system with three unites. Applied Mathematics and Computation, 263(??):251–267, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031500466X.

Zhang:2018:FTT

[ZXS^{+}18] Jing Zhang, Jianwei Xia, Wei Sun, Guangming Zhuang, and Zhen Wang. Finite-time tracking control for stochastic nonlinear systems with full state constraints. Applied Mathematics and Computation, 338(??):192–206, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318304491.

Zeng:2016:HBN

[ZXW16] Xiaocai Zeng, Zuoliang Xiong, and Changjian Wang. Hopf bifurcation for neutral-type neural network model with two delays. Applied Mathematics and Computation, 282(??):17–31, May 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316300595.

Zhuang:2019:NFD

[ZXX^{+}19] Guangming Zhuang, Shengyuan Xu, Jianwei Xia, Qian Ma, and Zhengqiang Zhang. Non-fragile delay feedback control for neutral stochastic Markovian jump systems with time-varying delays. Applied Mathematics and Computation, 355(??):21–32, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300319301614.

Zhang:2019:EET

[ZXZB19] Bei Zhang, Yonghui Xia, Wenjing Zhu, and Yuzhen Bai. Explicit exact traveling wave solutions and bifurcations of the generalized combined double sinh-cosh-Gordon equation. Applied Mathematics and Computation, 363 (??):Article 124576, December 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319305685.

Zhang:2015:FSV

[ZY15a] Yinghan Zhang and Xiaoyuan Yang. Fractional stochastic Volterra equation perturbed by fractional Brownian motion. Applied Mathematics and Computation, 256(??):20– 36, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315000600.

Zhang:2015:HSN

[ZY15b] Ziheng Zhang and Rong Yuan. Homoclinic solutions for a nonperiodic fourth order differential equations without coercive conditions. Applied Mathematics and Computation, 250(??):280–286, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314014891.

Zhao:2015:HOC

[ZY15c] Guozhong Zhao and Xijun Yu. The high order control volume discontinuous Petrov–Galerkin finite element method for the hyperbolic conservation laws based on Lax– Wendroff time discretization. Applied Mathematics and Computation, 252(??):175–188, February 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314016786.

Zhou:2015:BDS

[ZY15d] Jun Zhou and Di Yang.

Blowup for a degenerate and singular parabolic equation with nonlocal source and nonlocal boundary. Applied Mathematics and Computation, 256(??):881–884, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001289.

Zhang:2016:RPC

[ZY16] Hong Zhang and Hang Ye. Role of perception cost in tag-mediated cooperation. Applied Mathematics and Computation, 279(??):76–89, April 10, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300091.

Zeng:2018:EPG

[ZY18a] De-Yan Zeng and Jian-Hua Yin. An extremal problem on graphic sequences with a realization containing every l-tree on k vertices. Applied Mathematics and Computation, 337(??):469–486, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304727.

Zhao:2018:SCE

[ZY18b] Dianli Zhao and Sanling Yuan. Sharp conditions for the existence of a stationary distribution in one classical stochastic chemostat. Applied Mathematics and Computation, 339(??):186–198, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305782.

Zhang:2019:GIE

[ZY19] Hairui Zhang and Yongxin Yuan. Generalized inverse eigenvalue problems for Hermitian and Jhamiltonian/skew-Hamiltonian matrices. Applied Mathematics and Computation, 361(??):609–616, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300319304692.

Zhang:2018:DBAb

[ZYC18] Yuting Zhang, Yongguang Yu, and Xueli Cui. Dynamical behaviors analysis of memristor-based fractional-order complexvalued neural networks with time delay. Applied Mathematics and Computation, 339(??):231–241, December 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318305319.

Zhang:2015:MRA

[ZYH15] Min Zhang, Lei Yang, and Zheng-Hai Huang. Minimum n-rank approximation via iterative hard thresholding. Applied Mathematics and Computation, 256(??):860–875, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001319.

Zhou:2015:ASL

[ZYJ15] Xian-Feng Zhou, Fuli Yang, and Wei Jiang. Analytic study on linear neutral fractional differential equations. Applied Mathematics and Computation, 257(??):295–307, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031401710X.

Zhao:2017:DAF

[ZYLY17] Lin Zhao, Jinpeng Yu, Chong Lin, and Haisheng Yu. Distributed adaptive fixed-time consensus tracking for second-order multi-agent systems using modified terminal sliding mode. Applied Mathematics and Computation, 312(??):23–35, November 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303612.

Zhang:2019:DNR

[ZYLY19] Ruigang Zhang, Liangui Yang, Quansheng Liu, and Xiaojun Yin. Dynamics of nonlinear Rossby waves in zonally varying flow with spatial-temporal varying topography. Applied Mathematics and Computation, 346(??):666–679, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309585.

Zhang:2019:CRT

[ZYSL19] Wei Zhang, Xunbo Yin, M. H. Song, and M. Z. Liu. Convergence rate of the truncated Milstein method of stochastic differential delay equations. Applied Mathematics and Computation, 357(??):263–281, September 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300319302772.

Zhou:2016:RCU

[ZYWW16] Qi Zhou, Deyin Yao, Jiahui Wang, and Chengwei Wu. Robust control of uncertain semi-Markovian jump systems using sliding mode control method. Applied Mathematics and Computation, 286(??):72–87, August 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316302065.

Zhao:2019:MSC

[ZYX19] Jingjun Zhao, Yulian Yi, and Yang Xu. Mean square convergence of explicit two-step methods for highly nonlinear stochastic differential equations. Applied Mathematics and Computation, 361(??):466–483, November 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319304394.

Zhang:2017:ESCa

[ZYXF17] Lan Zhang, Xinsong Yang, Chen Xu, and Jianwen Feng. Exponential synchronization of complexvalued complex networks with time-varying delays and stochastic perturbations via time-delayed impulsive control. Applied Mathematics and Computation, 306(??):22–30, August 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317301005.

Zhang:2017:MSA

[ZYY17] Shengliang Zhang, Yu Yang, and Hongqiang Yang. A meshless symplectic algorithm for nonlinear wave equation using highly accurate RBFs quasiinterpolation. Applied Mathematics and Computation, 314(??):110–120, December 1, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300317304617.

Zhao:2015:NSF

[ZYZ15] Yu Zhao, Sanling Yuan, and Qimin Zhang. Numerical solution of a fuzzy stochastic single-species age-structure model in a polluted environment. Applied Mathematics and Computation, 260(??): 385–396, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300315004117.

Zhi:2016:STB

[ZYZ^{+}16a] Yongfeng Zhi, Yunyi Yang, Xi Zheng, Jun Zhang, and Zhen Wang. Statistical tracking behavior of affine projection algorithm for unity step size. Applied Mathematics and Computation, 283(??):22– 28, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031630087X.

Zhou:2016:TBS

[ZYZ16b] Yanli Zhou, Sanling Yuan, and Dianli Zhao. Threshold behavior of a stochastic SIS model with Lévy jumps. Applied Mathematics and Computation, 275(??):255–267, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315015799.

Zhou:2016:TAB

[ZYZW16] Y. C. Zhou, Z. W. Yang, H. Y. Zhang, and Y. Wang. Theoretical analysis for blowup behaviors of differential equations with piecewise constant arguments. Applied Mathematics and Computation, 274(??):353–361, February 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315014319.

Zhang:2014:BME

[ZZ14] Songlin Zhang and Kun Zhang. On a basic multivariate EIV model with linear equality constraints. Applied Mathematics and Computation, 236(??):247–252, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314004081. See

corrigendum [ZZ15c].

Zeng:2015:POC

[ZZ15a] Min-Li Zeng and Guo-Feng Zhang. Preconditioning optimal control of the unsteady Burgers equations with H_{1} regularized term. Applied Mathematics and Computation, 254(??):133–147, March 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0096300314017664.

Zhang:2015:FCP

[ZZ15b] Lu Zhang and Yong Zhou. Fractional Cauchy problems with almost sectorial operators. Applied Mathematics and Computation, 257(??):145–157, April 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314009813.

Zhang:2015:CBM

[ZZ15c] Songlin Zhang and Kun Zhang. Corrigendum to ‘On a basic multivariate EIV model with linear equality constraints’ [Applied Mathematics and Computation 236 (2014) 247–252]. Applied Mathematics and Computation, 265(??):1155, August 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315007778. See

[ZZ14].

Zhao:2015:UBD

[ZZ15d] Caidi Zhao and Hongjin Zhu. Upper bound of decay rate for solutions to the Navier–Stokes–Voigt equations in R^{3}. Applied Mathematics and Computation, 256(??):183–191, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300314017925.

Zhao:2016:FPB

[ZZ16a] Shuang Zhao and Heping Zhang. Forcing polynomials of benzenoid parallelogram and its related benzenoids. Applied Mathematics and Computation, 284(??):209–218, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302004.

Zhao:2016:CAR

[ZZ16b] Wenqiang Zhao and Yijin Zhang. Compactness and attracting of random attractors for non-autonomous stochastic lattice dynamical systems in weighted space ℓ_{ρ}^{p}. Applied Mathematics and Computation, 291(??):226–243, December 1, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316304258.

Zou:2016:RHO

[ZZ16c] Yunlei Zou and Jiandong Zhu. Reachability of higher-order logical control networks via matrix method. Applied Mathematics and Computation, 287–288(??):50–59, September 5, 2016. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300316302636.

Zhang:2017:SME

[ZZ17a] Yuezhi Zhang and Jiansong Zhang. The splitting mixed element method for parabolic equation and its application in chemotaxis model. Applied Mathematics and Computation, 313(??):287–300, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031730423X.

Zheng:2017:NSC

[ZZ17b] Hongchan Zheng and Baoxing Zhang. A nonstationary combined subdivision scheme generating exponential polynomials. Applied Mathematics and Computation, 313(??):209–221, November 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317303788.

Zhang:2018:ESC

[ZZ18a] Guodong Zhang and Zhigang Zeng. Exponential stability for a class of memristive neural networks with mixed timevarying delays. Applied Mathematics and Computation, 321(??):544–554, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317308044.

Zhang:2018:STF

[ZZ18b] Hong Zhang and Paul Andries Zegeling. Simulation of thin film flows with a moving mesh mixed finite element method. Applied Mathematics and Computation, 338(??):260–273, December 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300318305058.

Zhang:2018:RHO

[ZZ18c] Xiaohua Zhang and Ping Zhang. A reduced high-order compact finite difference scheme based on proper orthogonal decomposition technique for KdV equation. Applied Mathematics and Computation, 339(??):516–534, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305757.

Zhang:2019:HMF

[ZZ19a] Jiansong Zhang and Yuezhi Zhang. A hybrid mixed finite element method for miscible displacement problem with MCC procedure. Applied Mathematics and Computation, 346(??):143–154, April 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318309111.

Zhang:2019:PPN

[ZZ19b] Mengqing Zhang and Qimin Zhang. A positivity preserving numerical method for stochastic R&D model. Applied Mathematics and Computation, 351(??):193–203, June 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310440.

Zhou:2019:OLM

[ZZ19c] Yongtao Zhou and Chengjian Zhang. One-leg methods for nonlinear stiff fractional differential equations with Caputo derivatives. Applied Mathematics and Computation, 348(??):594–608, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031831066X.

Zdravkovic:2016:LMW

[ZZBS16] Slobodan Zdravković, Slobodan Zeković, Aleksandr N. Bugay, and Miljko V. Satarić. Localized modulated waves and longitudinal model of microtubules. Applied Mathematics and Computation, 285(??):248–259, July 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316302120.

Zhang:2015:EFC

[ZZC15] Wei-Guo Zhang, Yan-Nan Zhao, and Ai-Hua Chen. The elastic-fusion-coupled interaction for the Boussinesq equation and new soliton solutions of the KP equation. Applied Mathematics and Computation, 259(??):251–257, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315002660.

Zhang:2016:AEG

[ZZC16] Junfeng Zhang, Xudong Zhao, and Xiushan Cai. Absolute exponential L_{1}- gain analysis and synthesis of switched nonlinear positive systems with timevarying delay. Applied Mathematics and Computation, 284(??):24–36, July 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300316301692.

Zheng:2019:DGF

[ZZC19a] Yunying Zheng, Zhengang Zhao, and Yanfen Cui. The discontinuous Galerkin finite element approximation of the multi-order fractional initial problems. Applied Mathematics and Computation, 348(??):257–269, May 1, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318310312.

Zhou:2019:BIC

[ZZC^{+}19b] Jie Zhou, Wei Zhou, Tong Chu, Ying xiang Chang, and Meng jia Huang. Bifurcation, intermittent chaos and multi-stability in a two-stage Cournot game with R&D spillover and product differentiation. Applied Mathematics and Computation, 341(??):358–378, January 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

//www.sciencedirect.
com/science/article/
pii/S0096300318307768.

Zhang:2016:WGM

[ZZCY16] Hongqin Zhang, Yongkui Zou, Shimin Chai, and Hua Yue. Weak Galerkin method with (r, r − 1, r − 1)-order finite elements for second order parabolic equations. Applied Mathematics and Computation, 275(??):24–40, February 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315015441.

Zhang:2017:NRP

[ZZG17] Ke Zhang, Ju-Li Zhang, and Chuan-Qing Gu. A new relaxed PSS preconditioner for nonsymmetric saddle point problems. Applied Mathematics and Computation, 308(??):115–129, September 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317301960.

Zhang:2018:NSD

[ZZG18] JiHong Zhang, JunSheng Zheng, and QinJiao Gao. Numerical solution of the Degasperis–Procesi equation by the cubic Bspline quasi-interpolation method. Applied Mathematics and Computation, 324(??):218–227, May 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317308470.

Zhao:2015:HBS

[ZZH15] Hongyong Zhao, Xuebing Zhang, and Xuanxuan Huang. Hopf bifurcation and spatial patterns of a delayed biological economic system with diffusion. Applied Mathematics and Computation, 266(??):462–480, September 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315007134.

Zhang:2017:TLN

[ZZHT17] Songlin Zhang, Kun Zhang, Jie Han, and Xiaohua Tong. Total least norm solution for linear structured EIV model. Applied Mathematics and Computation, 304(??):58–64, July 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317300140.

Zhao:2015:EIM

[ZZL15a] Guilan Zhao, Xiaoli Zhu, and Yuhua Li. Existence of infinitely many solutions to a class of Kirchhoff– Schrödinger–Poisson system. Applied Mathematics and Computation, 256(??):572–581, April 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315000521.

Zhu:2015:GLH

[ZZL15b] Mu-Zheng Zhu, Guo-Feng Zhang, and Zhao-Zheng Liang. On generalized local Hermitian and skew-Hermitian splitting iterative method for block two-by-two linear systems. Applied Mathematics and Computation, 250(??):463–478, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300314014866.

Zhang:2016:KSB

[ZZL16a] Chenghui Zhang, Huihong Zhao, and Tongxing Li. Krein space-based H_{∞} adaptive smoother design for a class of Lipschitz nonlinear discrete-time systems. Applied Mathematics and Computation, 287– 288(??):134–148, September 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316302727.

Zhou:2016:FBP

[ZZL16b] Ling Zhou, Shan Zhang, and Zuhan Liu. A free boundary problem of a predator–prey model with advection in heterogeneous environment. Applied Mathematics and Computation, 289(??):22–36, October 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630306X.

Zhai:2017:SCN

[ZZL17] Shidong Zhai, Yuan Zhou, and Qingdu Li. Synchronization for coupled nonlinear systems with disturbances in input and measured output. Applied Mathematics and Computation, 294(??):227–237, February 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031630580X.

Zeng:2018:PSS

[ZZL^{+}18a] Deqiang Zeng, Ruimei Zhang, Xinzhi Liu, Shouming Zhong, and Kaibo Shi. Pinning stochastic sampled-data control for exponential synchronization of directed complex dynamical networks with sampled-data communications. Applied Mathematics and Computation, 337(??):87–101, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S009630031830434X.

Zhao:2018:FSO

[ZZL18b] Yong-Liang Zhao, Pei-Yong Zhu, and Wei-Hua Luo. A fast secondorder implicit scheme for non-linear time-space fractional diffusion equation with time delay and drift term. Applied Mathematics and Computation, 336(??):231–248, November 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318304077.

Zhang:2015:TAS

[ZZLY15] Yang Zhang, L. W. Zhang, K. M. Liew, and J. L. Yu. Transient analysis of single-layered graphene sheet using the kp-Ritz method and nonlocal elasticity theory. Applied Mathematics and Computation, 258(??):489–501, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300315001769.

Zeng:2017:SDS

[ZZLZ17] Deqiang Zeng, Ruimei Zhang, Yajuan Liu, and Shouming Zhong. Sampleddata synchronization of chaotic Lur’e systems via input-delay-dependent-free-matrix zero equality approach. Applied Mathematics and Computation, 315(??):34–46, December 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317305003.

Zhang:2018:HAM

[ZZM18] Xueying Zhang, Yue Zhao, and Baofeng Min. High accuracy modeling of sharp wave fronts for hyperbolic problems. Applied Mathematics and Computation, 331(??):190–199, August 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318301838.

Zhang:2019:TCO

[ZZMS19] Kun Zhang, Huaguang Zhang, Yunfei Mu, and Shaoxin Sun. Tracking control optimization scheme for a class of partially unknown fuzzy systems by using integral reinforcement learning architecture. Applied Mathematics and Computation, 359(??):344–356, October 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL https:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0096300319303868.

Zhao:2018:SCS

[ZZQ18] Jing Zhao, Tong Zhang, and Yanxia Qian. Stability and convergence of second order time discrete projection method for the linearized Oldroyd model. Applied Mathematics and Computation, 316(??):342–356, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317305751.

Zhang:2019:ISS

[ZZS^{+}19] Baoxing Zhang, Hongchan Zheng, Weijie Song, Zengyao Lin, and Jie Zhou. Interpolatory subdivision schemes with the optimal approximation order. Applied Mathematics and Computation, 347(??):1–14, April 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300318309524.

Zhao:2017:GCN

[ZZW17] Xuan-Yi Zhao, Chun-Gang Zhu, and Han Wang. Geometric conditions of non-self-intersecting NURBS surfaces. Applied Mathematics and Computation, 310(??):89–96, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302606.

Zhao:2018:CCG

[ZZX18] Jingjun Zhao, Rui Zhan, and Yang Xu. Dconvergence and conditional GDN-stability of exponential Runge–Kutta methods for semilinear delay differential equations. Applied Mathematics and Computation, 339(??):38–44, December 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318305514.

Zhang:2019:TOH

[ZZX19] Huaibao Zhang, Fan Zhang, and Chunguang Xu. Towards optimal high-order compact schemes for simulating compressible flows. Applied Mathematics and Computation, 355(??):

221–237, August 15, 2019. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0096300319301948.

Zhang:2017:CSC

[ZZY^{+}17] Rongpei Zhang, Jiang Zhu, Xijun Yu, Mingjun Li, and Abimael F. D. Loula. A conservative spectral collocation method for the nonlinear Schrödinger equation in two dimensions. Applied Mathematics and Computation, 310(??):194–203, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300317302874.

Zhang:2018:AST

[ZZY18] Juan Zhang, Xindong Zhang, and Bohui Yang. An approximation scheme for the time fractional convection–diffusion equation. Applied Mathematics and Computation, 335(??):305–312, October 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300318303370.

Zang:2015:BAM

[ZZZ15a] Hong Zang, Tonghua Zhang, and Yanduo Zhang. Bifurcation analysis of a mathematical model for genetic regulatory network with time delays. Applied Mathematics and Computation, 260(??):204–226, June 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S009630031500346X.

Zhang:2015:MSA

[ZZZ15b] Wen Zhang, Jian Zhang, and Fukun Zhao. Multiple solutions for asymptotically quadratic and superquadratic elliptic system of Hamiltonian type. Applied Mathematics and Computation, 263(??):36– 46, July 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0096300315004221.

Zhang:2016:CSM

[ZZZ16] Jiansong Zhang, Jiang Zhu, and Rongpei Zhang. Characteristic splitting mixed finite element analysis of Keller–Segel chemotaxis models. Applied Mathematics and Computation, 278(??):33–44, March 31, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300316300212.

Zhang:2017:ETS

[ZZZY17] Ruimei Zhang, Deqiang Zeng, Shouming Zhong, and Yongbin Yu. Eventtriggered sampling control for stability and stabilization of memristive neural networks with communication delays. Applied Mathematics and Computation, 310(??):57–74, October 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0096300317302801.